A Bibliography of Publications about the Java Programming Language, 1995–1999

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org, beebe@ieee.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

22 August 2025 Version 3.250

Abstract

This bibliography records books about the Java Programming Language and related software.

Title word cross-reference

$1 [Gar97c]. $1-Million [GC97]. $19.95 [Ano96o]. 2 [Ano99b]. $25.00 [Ano98e]. $29.95 [Ano97l, Ano96n, Ano96p, Ano97h, Ano97g, Ano98i, Ano98h, Ano99b]. 3 [CH98a, Gro96b, Hor96a, Hor96b, MD99, MD98a, NGM^{+}97, Sun97c]. $32.95 [Ano97k, Ano98f, Ano98j, Ano99a, Ano99c]. $34.95 [Bay97a, Ano97j]. $39.95 [Ano97i, Ano98g]. $44.95 [Tee99]. $54.95 [Ano98k]. $54.99 [Ver98]. $64.95 [Ano97l]. light [NvO98]. k > 1 [PQ96]. Z_{n} [GDS98].

-D [CH98a].

.class [Kei95].

// [GS97b].

0 [vD99b]. 0-13-272295-X [Bay97a]. 0-471-18359-8 [Ver98]. 0-521-58388-8 [vD99b].

1 [Ano97-150, DK98]. 1-886801-65-7 [Tee99]. 1.0 [Ano96-105, Dar98a, Red97c, San98a, Til97c]. 1.1 [BD98b, Dav97c, Eck97d, HL99, JRE98, Knu97d, LO^{+}97b, MBH97, M^{+}97a, NS98a]. 1.1.6 [JDK98a]. 1.2 [And98a, Ano98u, Bro99a, DDJ98b, DDJ98c, D^{+}98b, For98d, GM98a, Gil98a, GMPS97, HR98, HL99, KG98a, KK97b, LO^{+}97b, LO^{+}99b, LC98, PP99, Zuk98]. 1.2/2.0 [Stexxa]. 1.5 [Oly96]. 1/ [Add99, Ano99d]. 100 [Ano97-50, Ano97-140]. 1200 [Chw98]. 12th [IEE96h]. 13th [BK96a]. 15th [Syd97]. 16-bit [Bro97c]. 166MHz [Ano96-135]. 18th [IEE96k, IDK98]. 1997 [Ano97-81, Cul97c]. 1998 [RD99]. 19th [Ano98a]. 1st [Abr96].

2 [Ano97-44, Ano98q, Bro99b, CLK99b, Cow99, DSM99, DiD97a, DiD97b, DiD97c, Fox99, Gon99, HR99, Jan99, Kro99d, LO99a, LG99, LC99, MD99, NRG^{+}99, Pis99, RHE99, Rom99, Sha96a, S^{+}96b, Tee99, WFC^{+}99, Zuk99]. 2-Card [Cha96e]. 2-D [Ano97-44]. 2.0 [Ano97-100, Ano97-111, Ano97-119, Ano97-147, Chw98, MD99, Rit96a, SUN97b]. 2.xx [PLS98]. 2/e [Tee99]. 2000 [Ano97r, Ano97-28, Bha98, Gar97b, NTS^{+}96]. 23rd [Ano98-34]. 24th [ACM97a]. 25th [ACM98b]. 26th [ACM99a, Isk96]. 29th [IEE96l]. 2D [Har99a, Loe99, Ano99f, Gea98, Knu99b, NNS97]. 2nd [Ano96-158, Cad98b, Mey97, Ass96a, USE96a, USE96b, USE99a, USE99b].

3 [Ano96-53, Ano96-133, Ano96-151, Anoxx-80, Bry98, CL98b, Dun97, Gre97h, JDK98b, Lau96, Mag95, Sch97j, Win95k]. 3-D [Ano96-53, Ano96-133, Ano96-151, Gre97h, Lau96, Mag95, Sch97j, Win95k]. 3.0 [Ano96l, Ano97-153, Bra96, Bra97a, Bra97b, How96]. 3.1 [Gau96g, Hay96y]. 300-nm [Gar97a]. 3000 [Wol97]. 32-bit [Ano96-126]. 3654-22 [Mau98]. 390 [Koo98]. 3Com [Gut95b]. 3D [SD99, And96a, Ano96-51, BV98, BP99, CJ98, DDJ99b, LMM96, Loe97, LCLL97, Mau98, MD98b, MCxx, Pel96, SML96, SRD98, Web97b]. 3rd [IEE96b, USE98a].

4 [BD99, Bra96, D^{+}98b, DiD97c, DC97, JC97, LO^{+}99b, S^{+}96b]. 4.0 [LO^{+}97b, LO98a, MP98, MP99]. 4.6 [Bal98a]. 4.6-Anwendungsentwicklung [H^{+}97]. 4.x [Knu97h]. 400 [Ano97-28, Den97, Oue96a]. 4Developers [Kro98a]. 4th [Cul97c, Ass96b].

5 [Ano96-28, TJ96]. 5.0 [Ano97r]. 500 [RC98]. 500-doru [IK96]. 5th [IEE96m, SMEN96, USE97a].

6.0 [Ano98y, Dav98b, Gil98b]. 60-fun [IK96]. 633MHz [Ano98-28]. 64 [Tho98]. 64-bit [KG97b]. 680x0 [Ano96-42]. 6th [Ano98o, DLM99].

7 [Ano97f, Dan96b, MD99].

7-Tage-Crashkurs [Dan96b]. 754 [Dar98a].

8 [vD99b]. 80 [Anoxx-32].

9 [Ano96-28]. 90 [DP99, MKS^{+}96]. ’90s [AL97]. ’95 [Boz95d, Ano97x, Bal97, Bro97b, Bro98d, Coo97c, Infxxa, Int96a, JC97, Sch97c, Sym95, Taf96, TJ96, Var97a, Wil98b]. 95/Java [Sha95]. 95/NT [Sym95, Wil98b]. ’96 [ACM96, Ano96-28, Ano96-147, Ano96-159, Bod96, BBJR96, BK96a, FGL96, IEE96c, IEE96d, IEE96s, IEE96m, PZ96, DF96, SLPM96, Ste96b, Str96, Ass96b, USE96a, Wae96, YN96]. ’97 [ACM97a, Ano97-139, USE97a, Ano97-107, San97a]. ’98 [ACM98b, Ano98a, IEE98b, IEE98c, Lew98]. ’99 [ACM99b, ACM99a, USE99a, USE99c, USE99b, USE99d]. 9th [IEE98c].

= [Ano96-133, PC97, Way96g, Way96f].

A&D [Ano97a]. A. [McC96e]. Aachen [SLPM96]. ABCs [PM97e]. Aboard [Des96]. Abstract [Hem96b, Jen98, MKS^{+}96]. abstraction [Nan95]. abstractions [AA98]. abstracts [ACM99d]. abuzz [Her96b]. academia [Bro98c]. Accelerated [Ano96-135]. accent [Cou96b]. Acceptance [Gil97a]. Access [Ano96-137, Gau97h, GFR98, Gre97d, HCF97a, Hamxxa, McC98a, PH99a, Sim97c, WCF98, ZMN98, Ano97-72, Ber97d, CDD97, CH98b, Dua96, Gau97e, Hay96b, Mey96a, MC97, Red97b, Reg96, Sau96, Way96g, WFC^{+}99, ZMN99, Hob98b]. Accessible [Sha97b]. Accessing [Bal99]. accord [Cha97c]. according [vSS96e]. account [Yel99]. Accounting [CvE98]. accuracy [Woe97]. AccuSoft [Ano98-29]. achievements [LMdS98]. ACM [ACM98b, ACM99b, IEE96l]. Acquires [GC97]. Acrobat [Ano95k, Kar96, Mat99]. Across [Ano97-34, Ano97-76, GS98c, NE98a, NE98b, GS98d, Ren98, Sha97b]. act [Chw98]. Active [BN96, FI96b, Lad97, Ladxxa, Ste97g, BR97, Ano96-73, Ano97-128, Gre96a, Hay96d, Hay96f, JC97, Kal95b, Lad98a]. ActiveX [Lad98a, Ladxxb, Ano96-61, Ano96-66, Ano96-87, Ano96-111, Ano97t, Ano97-92, Ano97-122, Ano97-159, Ano97-166, Anoxxm, Bra98, Bre96, CL97c, Cha97b, Chaxxd, Chu98a, Chw98, Cor97b, Cro97, GH96a, Har96b, Harxxa, Hay97a, Hop97, HS98a, Hus97, Wil97i, KSB97, KPF96, Kro98a, Kro98d, Kro98c, Kro99a, Kro99b, Lad97, Mac97c, Mac97b, Mey96b, Proxx, Raj96, SB97, SR97, VN96, Wil97a, Wil97j, WB^{+}96c, ZK98a, ZK98b]. ActiveXecutables [Ano97-28]. Activities [T^{+}98a, BT99]. Actor [IG99]. Actors [GK98, VA98]. Acu4GL [Ano96-137]. Acumen [Kro99a, Kro99b]. Ada [Ano97x, Bal97, Bro97b, Bro98b, Bro98d, Bro98c, BF99b, Carxx, EMN98, Fli98, Int96a, Kro98d, Kro98c, Lut98b, MH97b, Mun96, MH97c, PV99, Rie97a, Str96, Taf96, Wes98, Whe97, Woo99a]. Ada-Europe [Str96]. Ada95 [KFS97, Mun96, WN97].

Ada’97 [ACM97b]. Adaing [PV99]. Adaptable [KG98b, KG98c]. Adaptive [EM96, KB98a, KB98b, Led97, Sar99, DW97b, EH96, GRN99]. ADBC [Reg96]. Add [Cox95a, SBH97, Gau96a, GME99, Hay96i, Hay96q, Hib96, Pic96b]. add-on [Hay96i]. Add-Ons [SBH97]. Added [Ano98b]. Adding [AFM97, BOSW98, Nil98a, Dar98a]. Adds [Ano97y, Ano97-42, Web97b, Ano97-102, Ano97-136, Ano98p, DiD97a, Gau97l, Kar95b, O’C95d]. ADEPT [GSTV97, Ano97f]. administrators [Ano96-123, Gor96]. Adobe [Ano96-195, Com95a, Mat99]. Adoption [Hal98b, Che97]. Adopts [Knu99a]. Adra [Ano97b]. Advance [HC99c]. Advanced [Ano96-137, Ber98e, BF98, BF99a, BW99, CBT97a, CBT99, Cou98, DG96, Gei96, HC98b, HM98, Laf96, Lafxxa, Min97a, Nat98, RS97, RD99, SR97, Sri97a, SKN98, WB^{+}96b, Weh96, Amb96, SS97, TB96, Law99]. Advances [BK96a, MJA98, DLM99]. Advancing [Wor97]. advantageous [Bat96c]. Adventures [Mon96, Bal96c]. Advice [Jar97, Ano97-73, Kay97, You96a]. Advisor [Saa96, Wal96e, Wal96d, Wal96f, Int96b, Jav96a]. aepullit [Kim97d]. AES [RD99, Dra98, Fol99, Nat99]. AFC [Dav98a, SS98d, Wal98b, Wat98, Wat99, You98a]. Afraid [Bro98a, Bro99c, Hel97c]. after [Biexxa, Biexxb, VA98]. Again [Ano98u, Gar96b, Lew96b, Tym98, Urg96a]. Against [Ano96-163, Ano97-122, Ano96-196, Ano96-97]. Age [BCP^{+}99b, CBT98g, Gre97g, Atw96, Car96c]. Agency [Gar97a]. Agenda [Ano96-55]. Agent [Ano96-137, Ano96-154, EBR96a, Gla98, Ham97b, Wat95f, Ano97-69, EBR96b, NJC96, TSKA^{+}96, VI96]. Agent-Based [EBR96a, Ham97b, Wat95f, EBR96b]. agent-mediated [TSKA^{+}96]. Agent-Trudger [Ano96-137]. Agenten [KS97]. Agents [Ano96-154, BB97b, CR97, Doe97, Fue98, Fün98, Ham97b, KB98a, LOKK97, LO98b, Pet97, Dom96, KB98b, LO99c, Pog96, WPM99, dRPT98]. aging [Hay96k]. Aglets [LOKK97, LO98b, LO99c, TK99a]. Ago [Ano96b]. Agora [Dav97e, Davxxb]. Agree [Gar97a, Gar97b, Ano97-143]. Agreement [Gut95j, Kal95b]. Ahead [PTB^{+}97, Ano96g, Gau96h]. Ahora [Jam97c]. Ahpah [Kro97]. AI [Pog96]. aid [TGGP97]. aided [DF96]. Aim [Ano96-107, Hay96x]. aimed [Amb96]. Aims [Val96a, Kar95a, Pic96a]. AirJava [Mil99]. AIX [Ano96c]. AKA [TGxxb]. AKI [Anoxx-32]. AKI-80 [Anoxx-32]. Aktualisierung [BKP97b]. Al [Fra97, Holxxb, Holxxc, GGxxa, Sam97, Ver98]. Albuquerque [BBJR96]. alcance [Harxxc]. Alexandria [PSTZ99]. Alfonse [Har98c, Har99d]. Algebra [Sch99b, WKES98, WKE98]. Algebraic [Doo99, SS98a, SS98b]. Algor [Ano97c]. Algorithm [Boy98a, Boy98b, DW96b, Dra98, Nap97, Sha98a, BCLT95, Nap98, NAC99]. Algorithmen [NH98, NHxx]. Algorithms [BAC^{+}99, BCP99a, CH97g, GT97, JPK98a, JL96, Lad98b, Laf98a, Row97, Sed98, BV98, BCLT96, BA97, GME99, JPK98b, KH99]. Algortihmen [Laf98b]. Alike [Gar96e]. Alive [Aoy97, Aoyxxb, RBL^{+}97]. All-out [Win95c]. alla [Jaw97b, Jawxxa, Tis97]. aller [Sta96]. Alles [TGxxb, Ano96-56, DE96b, Sch97k, TG96a]. Alley [Boy98a, Boy98b, DW96b, Eck96b]. alliance [Ano97v]. Allocation [DH99]. allowing [Ano96-54, Stu98]. allows [MF97a, Pic96b]. Almanac [Cha98c, Cha99c]. Also [Ano96-53, Cou97d]. Alter [DS95a, Tub95]. Alternate [Joh97a]. Alternative [Gar97a, KF99a, KF99b, Ano96-54, WSK97]. Alto [Fox98b]. Always [BCP^{+}99b]. Amazing [GC97]. ambients [CG99].

Amelio [Ano96-195]. American [Chu98b]. AMIA [Cim96]. Amok [Lut98b]. among [Cli98a]. Amount [Jam96]. Amtec [Ano96-135]. Amzi [Ano96d]. Analogy [Kro98d, Kro98c]. Analysis [AFH98a, Ano97-28, Ano97-163, Ano98-30, Coc98, DK98, Dra98, Gil98a, HAT98a, HT98, HO99, HJZ^{+}98, Kro98d, Kro98c, MAS96, MST99, Sha98a, WH99, WO97, AFH98b, Ano96-128, Ano97-50, Bla99, CGS^{+}99, CGHS99, Cla97d, Com97a, Com99, Hin98, Kap99, Lan98, Lan99a, Ste96f, WR99b, Kro98d, Kro98c]. Analyzer [BD98a]. Analyzing [BAC^{+}99, RM99]. anarchy [Flo97]. Anatomy [CBT97b]. Andreessen [Par95h, Var95, Win95a]. Andrew [Ano96-137, Ano97l, Ano98k, vD99b, KW96]. anesthesia [IDK98]. Anfangen [Krü97a, Krüxxa]. anglais [BKP97a]. anglais-francais [BKP97a]. animal [Gol98b]. Animating [BCLT96, Lem97a]. Animation [Buc96a, Bur96b, CBT97e, Con97b, Fox97c, MD99, PWxx, Rao99b, tYpCxx, BCLT95, LCLL97, MN96b, SF96c]. Animations [TRMR97]. animierter [Amm97a, Ammxx]. Annasoft [Kro98d, Kro98c]. Annex [Bal97, TBA98]. Annotated [CL97b, Fre97b, HCF97a]. Annotating [HAKN97]. Announce [CRS98a, Gut95j, Inf95a, Net95a, Wat95b, Wat95d, Wat95c, CRS98b, Red97a]. Announce-Listen [CRS98a, CRS98b]. Announced [Ano96-167, Col95, MD99]. Announcement [Nat95]. Announces [Ano96d, Ano97-45, Ano97-124, Ano97-134, Bus95, Lot95, MD99, Ano96-72, Ano97b, Ano97-132, Ano97-157, Ano97-162, Com95g]. annoying [Gre97k]. Annual [ACM96, Ano96-160, Ano98a, Ano98-34, BBJR96, Chu98b, IEE96h, IEE96s, IEE98b, Isk96, Ass96b, USE97a, ACM99d, Cim96, IEE96l, USE96d, USE98c, USE98d, USE98e]. ANSI [Drixx]. Answer [Lew97b]. answered [Ano96-43]. Answering [Bra98].

Answers [Ano95t, WL97d]. ante [Mey98a]. Antennas [GM98b]. Antibes [Ano97w]. Antidotes [MF97b, Lea97c]. ANTLR [PQ95, PLS98, Sch99a]. Antonio [ACM99a, USE98f]. Antwerpen [IEE96b]. anwenden [Krü97a, Krüxxa]. Anwenderpraxis [KS96a]. Anwendungen [N^{+}97a, ER97]. Anwendungsentwicklung [H^{+}97]. Anwendungsprogrammierschnittstelle [G^{+}xxb]. Any [WL96, Ano96-59]. Anymore [CBS95, Sto96]. Anyone [Tho95]. Anyway [Sch95]. AOL [GC97]. Aonix [Ano97d, Kro98e, Kro99a, Kro99b]. Apache [Ano98-28]. API [Ano97-147, GYt96a, GYt96b, Gosxxa, Gosxxb, GSHP97, Groxxa, N^{+}xxa, SBH97, Ano96-62, Ano96-114, Ano96-181, Ano97s, Anoxx-35, Aoyxxa, FB96, FBxx, Gar98b, Gea98, Gon99, GYT97a, GYT97b, G^{+}xxd, G^{+}xxe, GHP^{+}96, Har99a, JZ98, Jon98, JRC97, Kro98a, Kro98b, Lea99, LS97b, Loe99, NGM^{+}97, NMS96, N^{+}96a, Rob99, Sim97c, SL99, SRD98, SD99, Van97m, WFC^{+}99, Wil97b]. APIs [Ano96-131, Fli98, Gon98, JMA98, JPA98, NMS96]. App [AW99a, DDJ99b, Lin96e, Ano96-124]. appeal [Ano96-153]. Appel [Ano97l, Ano98k, vD99b]. Apple [Ano96f, Ano96-195, Gar97a, Gar99, Swa97c, Swa97d]. Applet [Ano96-63, Ano96-71, Ano96-137, Ano96-169, Ano97-122, Ano97-125, Anoxx-35, BC98a, Ber96f, Ber97g, DR96, GS98a, Kie99, Kin97a, Mec97, Mee96, Moo98, N^{+}96a, Sav96, Ano96y, Ber97d, Ber97b, Cha96c, Gau97r, Hay96a, Hay96l, MF97a, WO97, pP98, Ano96-117, Ano97e, JM97c]. Applet-Browser [Kin97a]. applet-like [pP98]. Applet-Struktur [JM97c]. appleta [Gol96a]. Applets [Ano96-64, Ano96-107, Ano96-137, Ano96-172, Ano97-37, Anoxxb, Anoxxq, Anoxx-36, Anu96b, BC99a, BKKK98a, BM96a, BCE98, Bri99b, BLT98, Car96c, Chaxxb, Cho97a, CM97b, CM97c, CM97d, CM99, Cou97f, Couxx, Crexx, DMB96, EMe97e, EMe97d, EMe97g, EMe97f, ER97, Fra96a, Frö96, Fröxxb, G^{+}97, GYT96d, GYT97b, Gra96a, GRD96, GR96, GR97, Gul98, Gulxx, HI96a, HI96b, HIxxa, HIxxb, Hunxx, KFS97, KS96b, Küh97, Lea97c, LLxx, McF97, McF98a, MF97b, MPxx, Nag96, Per96c, Per97c, Pew97b, Rao99b, RMS97, Rit96a, Rod95, Rod96a, Rod96b, Rod97, Rodxxa, Rodxxb, Rodxxc, Sch97c, SBH97, Sim96e, Rapxx, SP97, VPM98a, VPM98b, Wal96j, WG97a, Wit97, ZK98a, ZK98b, AS96, Ano95w, Ano96-72, Ano96-102, Ano96-123, Ano97-70, Ano97-71, BKKK98b, Bru97, DiD97b, Ekxxa, GPO98a]. applets [GPO98b, Got98a, Hay96s, HL97b, HL97c, KW96, LR98, McC97b, NMS96, Raj96, Tur97, Way96g, Wes98, YLQ98, vSS96b, vSS96c, GYt96b, Gosxxb, vSS96d, vanxxd, Ek97a]. Appliance [Gar97b, Kal95a, Mil95d, Uim95]. Appliances [Ano96-53, DY99a, DY99b, Met95b]. Application [Ano96-154, Ano97-85, Ano98c, Ano98-31, Bar99, Ber97l, Ber97m, Bra99, Cob98, Dav98a, GYT96c, GYT96d, G^{+}xxf, G^{+}xxg, Gre98a, HNSS99, Kal98, Kro98e, Lad97, McG97, Smixxc, SKN98, Vac96, Ano97-47, BRS99, Bla99, FHKS98, FC96a, Ham96a, Han96, HW97, Joh96a, Mau98, Mur96a, Sli97, Stu98, SW99, TLSS99, You98a]. application-sharing [BRS99]. Applications [Abl96, Ano96-101, Ano99h, App95, AW99b, BN97, Ber98e, BB97b, Bin97, BLT98, BM98c, Cag96, CMH99, Dav97c, Desxx, Fis98a, Fri98, GM99, Ham97b, H^{+}xxb, IBM97, IEE96g, IEE96h, IEE96r, IEE96s, IEE98b, KT99, Laixx, LB97a, LBS98, Man96b, Man96d, McC98b, McF98a, Mit97b, Mor98a, Mor97a, MR97, MD97b, PM97b, PTB^{+}97, Rei97a, Sch97c, Sco96b, Sel99, Skl95, Ste98a, TPHB96, TJ96, T^{+}98a, Ude97c, Wat97b, Wat97a, Whi96d, Wil97a, Won98, ACM99d, Amb97, Ano95s, Ano96k, Ano96-54, Ano96-69, Ano96-125, Ano97t, Ano98b, Che96, CKTV96, DR98, DR99, DAB97, Dun97, Gol96a, Gol98b, Gre96c, GME99, HK99, IEE98a, Isa96, KW96, LQ99, MTU99, MLF97, Nan95, Sco96a, TTD96b, Ude97d, Val99, Way98c, Whi96b, Woo99b, Xu97, YLQ98, You96a, pP98, van98a]. applications-technology [Gol96a]. Applications/Applets [McF98a]. Applied [HH98, Hun97, Syd97, Ano98o, Ciz97]. Applikationen [G^{+}97, Her97, H^{+}97]. Applix [Ano96g]. Apply [SWJ97]. Applying [BF99b, EBB^{+}96, Ste96f]. appraisement [BCK99]. Apprentice [Cha98a]. Approach [AW98, Cul98d, Culxxa, Doe95, GM98b, Lor96b, MS99b, Sch97j, Wuxx, Ast98, Dav97a, Had99, Miz99, WSK97]. approaches [Ano96-193]. Approves [Gar98b]. Apps [And97b, Ano97-27, Ano97-35, Ano97-79, Ano97-155, CM97b, CM97c, CM97d, CM99, HB95, Inf95f, Kro97, LaM95g, Ren98, Win95n, Ano96-42, Ano96-72, Ano96-171, Ano97-69, Ano97-92, Ano97-165, Gau96e, Hay95a, LH96, Pic96b, Pic97, Red97b, Var96]. Aprend [Hob98a]. Aprenda [Cadxxa, Dav97e, Dav97f, Davxxc, Davxxb, Lem97b, Lemxxa, W^{+}xxa]. Aprendendo [Dam97]. Aprender [Walxxb]. Aprendiendo [Cad98a, Dan96e, Lem97c, LP98c, WL97c, WL98]. April [Ano96-154, Ano97w, IEE96d, Tau96, Ten99, Uni98a, USE98b]. Aproveche [Weh98a]. ArborText [Ano97f]. Architectural [VRG98a, VRG98b, PS98, SvK98]. Architecture [AT97, Ano96-93, CSY97, CSxx, Fre96b, Fre97c, Fre97d, GDS98, GMPS97, Hiu98, KG98a, KFL99, NK98, RNBV98, Sun99a, Tre99a, TSKA^{+}96, Ano96-153, CKTV96, FC98b, Gau96a, Gon99, JCS98b, KG96a, KG96b, MD98b, TYY96, Wal99, JCS98a, Pop97, JGKR96]. Architectures [Gon98, Lev99, ZMN98, FD96c, WBDF97, ZMN99]. Architektur [Dal97]. Archives [Fre97b]. Ardent [Gre98c]. ARDI [Ano96-135]. arena [Hay96i]. argumentation [KPG96]. Arguments [Whi96a]. Argus [Ano96h, Ano96i]. Aria [GSTV97]. Aria-Java [GSTV97]. Ariadne [JJG98]. Arizona [IEE99b]. ARM [Kuk98]. ARMI [RWB97]. Armidale [Ano98o]. armour [Col96c]. Arms [Ste96d, Ste96e]. Array [MMGL99, MMS98]. Arrays [GM98b, But99]. arrive [Ano96-69]. Arrives [Ano96-116, Ano96-43, Kar95a]. Art [FRT96, Pla96c, Rod96b, Eck97a, IDK98, W^{+}98]. Arthur [Gut95a]. article [Ano96y, Rob96]. articles [Far96e]. Artificial [IEE96i, BK96a]. Artlandia [Kro98d, Kro98c]. AS/400 [Ano97-28, Den97, Oue96a]. ASCII [Gop96]. ascp0.8.2 [Ano98-27]. Asia [Gut95d, Gut95c, Uni98a]. Asked [Kan96a, Kanxxc]. Asleep [BAC^{+}99]. aspects [EBB^{+}96, PL96]. assemble [Gre96c]. Assembling [Nau98, Ano96-153]. Assembly [Buc97, Red97c]. Assertions [PSS98]. Assessment [FW96, Zac97]. assist [Ano97-92]. Assistance [Gil97c]. Assisted [BBJR96, Gre97b]. assists [Red98]. Associates [Ano96n, Ano96o]. Association [Chu98b]. Assurance [Frü99, MRA98]. astronomy [Ref99]. Astuces [Ano96-120]. Asynchronous [Ver96, RWB97]. AT&T [Bus95]. atarashii [DM96a]. Atelier [Ladxxb]. Atlanta [ACM99b, Joy98]. Atlas [KSI^{+}96]. ATM [JZ98, Lut98b, Sau96]. attached [Mon97]. Attacks [Ano96-163, Kay97, McC98e]. Attitude [Esp96]. Audience [Bou96b]. Audience1 [SBB96]. audiences [Hay96v]. Audio [Moo98]. audit [Men96]. Auflage [Fla98c].

Augments [LaM95d]. August [IEE96e, IEE96n, IEE97, IEE98b, IEE99a, Nat98, RD99, Syd97, USE98a]. AUn [Red97a]. ausführliche [Her97]. Ausgabe [Fla98c, Fla98a, Fla98b, LJ96a]. Austin [USE99a]. Australia [Ano98o, MDM97]. Austria [IEE98b]. authentication [Ano97-38, Mey98b]. author [Sto97]. Authoring [Ano97-154, CBT97a, CBT99, Kro98e, MD99, TMP^{+}99, Wea97a]. Authors [AL97, Ano96-202]. Autobahn [Oly96]. automata [RHND99, RBL^{+}97]. Automate [Par99]. Automated [Lut99d, Lut99e, PPIG98, Sel99]. Automatic [JJBW98, JL96, Kro98d, Kro98c, MG97a, SLCM99, Spi99, CCCL97, vRvGS97]. Automatically [Ait96a, BG97b]. Automating [EMN98, MR99]. automation [Ban98]. Autonomous [BFD96]. AutoPlant [MD99]. Autorouter [Kro98e]. Available [Kro98d, Kro98c, MD99, Wol97]. Avalanche [SP98]. Aventail [Ano98-29]. Avilon [Kro99a, Kro99b]. Aviv [IEE98c]. Avoid [JN95b, Flo97]. avoided [JN95a]. Avoiding [Dic97a, Ano96-42]. await [Gau97a]. awake [Ano96-183, Cou97e]. Awards [Eri96a]. aware [App95]. Away [DDJ99a, Par95e, Hay96m]. Awesome [Hat97, Hat98b, Rod98]. AWT [Ano97i, GM97b, O’N97b, Cha96c, Anoxx-35, Chaxxb, Culxxb, Gat97, GM97c, GM97d, GM98a, GYT97b, Hol99, NMS96, N^{+}xxa, Nag96, OS96, Whi97b, Zuk97b]. AXP [Ano97-121].

B [Ano96-72]. Baan [Kro99a, Kro99b]. Baby [BDF96, Hay97d]. Babylon [Ano98-27]. Back [Bow96b, GC97, Mit96b, Mit97a, Mit96a]. Backbone [Lut98a]. Backend [RC97]. Backing [Cas96a]. Backstage [Ano96j, Nas95a, Ano96j]. backward [Mat96b]. Bad [All99a, Dar98b, Mar98c, Dar99a]. Bag [Anoxx-37]. Bahamas [Ano99h]. Bahasa [Wijxx]. Balance [Gut95c]. Balancing [Kro99a, Kro99b]. BALI [Lim95]. ball [Bab97a]. Ban [Gar97a]. Banco [Jep96]. Bancos [Jep96]. Band [GYt96a, GYt96b, Gosxxa, Gosxxb]. Bandwagon [Ano95c, Par95c, Uim95, Win96]. Bank [Cul97c]. Banking [KGG96]. banks [Hof97]. bar [Mee96]. bar-chart [Mee96]. Barcelona [Ano97-139, DLM99]. Barrier [Che97, O’B96a]. Barriers [GH96a, LM99]. Basas [FB98]. Base [Hob98a]. Based [Ano96-101, Ano97-114, Ano98-30, Buhxx, IG99, CKRW99, Cha98d, Cha98e, CCB96b, Div98, EBR96a, FFKO99, Fue98, Fün98, GFR98, GS98c, Ham97b, Hor96a, Hor96b, Isk96, KS98a, KS98b, KF99a, KF99b, KP99, Kro98b, LPR97, MK98, Mos98b, Mos98a, PSW^{+}98, PV98, PTTA99, SEH98, Sar98, Sar99, Sch99b, SN96, TBS97, TWP98, TMM^{+}98, TMMH99, UCD99, VPV99, VCCH99, WAES97a, WAES97b, ZMN98, AJ98a, AFH98a, AFH98b, Ale98, AISS97, Ano96t, Ano96-34, Ano96-47, Ano97b, Ano97c, Ano97t, Ano97-72, Ano97-135, Ano97-157, Ano98b, Ano98d, Ano98s, Asl98, Ast98, BN98b, BBC^{+}97, Bin97, BUP99, BK97b, CKRW97a, CKRW97b, CMM^{+}97, CT98, CCI^{+}97, CT96, CWS98, Cox95c, DF97, DAB97, Dug99, EBR96b, FT96, GH97, GRN99, GS98d, Gre98d, HW97, JADR96, JA98, JJG98]. based [KPG96, Kro98a, Lin96b, LQ99, LCLL97, MAS96, Mau98, Mey99, PSS97, Ref99, RHND99, RNBV98, SEH99a, SEH99b, SH99, SvG98, Sim96f, SW96, SHGD98a, SHGD98b, TB96, Tým99, Ude97b, Ung97, Val99, VO98, Wat95f, WOC99, WPM99, WCS99, YC98a, YC98b, ZMN99, vHO98, Don96, Ano96-72, Ano96-65, Ano96-192, Bar97, ERKE96, FF96b, JAS96a, JADR96, Jan96]. Bases [Lesxx]. Bashing [Swa97a]. Basic [Ano97l, App97, Ano96-127, Ano97-161, Buc97, Coo97c, GH96c, Gau97c, Gre97a, Gut95i, H^{+}97, JLW96, LaM95d, Met98, Nau98, Ubo96b, You97b]. Basics [AMPM96, AMP97, Bur96b, T^{+}98b, Ver98]. Basis [GPHL99, NB96]. Basiscursus [Amm97b, Kas97b]. Basispakete [GYt96a, Gosxxa]. Batch [Ano98-27]. Battle [Ano96r, Bal96a, Gar97a, Gar97c, Gar98b, Law99, Ort99, Fly95, Gau97n]. Battles [Gar99]. Bayanihan [Sar98, SHW98a, SHW98b, SH99]. B’blia [Hol98b]. BBSs [Ano96-151]. BC [Doo99, IEE98a, Tau96]. Bd.1 [CH98c]. Bd.2 [HC98c]. Be [Com95f, Als95a, Ano96-59, Dav97b, Fla96a, Gil97a, Gre97h, Gre97k, Kee96, LH96, McC97b, Pic96a, Sch96b, Sch97f, Sco96a, Zol96]. Beach [IEE96o]. beams [MC98]. Bean [Hou98]. BeanBox [Eng97b]. Beans [CBT98b, Joh98, JT97, Küh97, Mor97e, NK98, Swa97a, DiG95b, Gre97g, Lew96b, Ano96-66, Boy98c, Doh97, EMe97h, Eng97c, Feg97, Gau96d, Gig97, Pic96b, Rie97b, Sim96b, Van97j, W^{+}97b, Zif97, Ano97h]. beating [Nat96]. Beautiful [Wil99e]. became [Ban95a]. Becker [Anoxxh]. become [Ano96-54, Hal97a]. Becomes [Ano97-73, Ano95i, Gau97d, Ude97b]. Becoming [Tra95a, Lin98b]. been [Gan97]. before [MF97a]. Beginnen [WW96a]. Beginner [Cha98a, Ene97, M^{+}xxa, Wes98]. Beginners [McC97c, Amb96, BT98]. Beginning [Hor97b]. beginnt [Bat96d]. Begins [Gar98b, Law99, Win95c]. begs [Shu97b]. Behavior [DH99, Wol99]. beherrschen [GM97b]. behind [IEE96b, Mel96]. Being [Ano96-99]. Beispiele [Mor97e]. Beispielen [JM97c, TG96a]. Belgium [IEE96b]. Bell [Ano96-74, Gro96b]. bellatlantic.net [PJR^{+}98]. Benchmarks [DH99]. benefits [Ano96k]. Benioff [SW95]. Bentley [Ano98d]. Beowulf [DDJ98a]. Berechnung [Web98]. Berlin [IEE96k, IEE96j, Syd97]. Berlinda [Tol97b, Tol97a]. Beschreibung [Sta96]. Best [Ano96l, Bro96, Cou97f, Couxx, Hug97, TD97, WBC^{+}96]. Best-Selling [Bro96]. Beta [CL98b, Cox95b, DH97, DH98, Fla96e, Fla97h, JDK98b, Rit95b, Rit97a, Ritxxc, Ano96p, Ano96-195, Ano96-43, Ano97-164, DiD97c, Gau96d, Gol96b]. bets [Ano97-110, Gil97c, Wea97b]. Better [And96b, Ber96j, BCP^{+}99b, CM97b, CM97c, CM97d, CM99, DS95b, Orc98, SF96c, TMP^{+}99, Tur97, Way96b, Ano96-133, Ber96i, Cli98b, Gau96i, Hal98a, McC98c, PC97, Way96e]. BetterState [Ano97-128]. Between [Ano96r, JPT^{+}99, Tes96a, Hav98, Hop97, Kar98a, Kar98b, KMW99]. Beyond [Ano96-28, Cim96, Fra98a, GMPS97, Gre96c, Kar96, Nas95b, SHR^{+}99, SJRK97, van97b, Cou97c, DFW96, Gau97o, Jac96, KDH96, SLPM96, Tho98]. BHV [Kno96a]. Bible [Anoxx-78, Arm97, LA97a, LA97b, Wal97f, Goo96b]. Biblia [VPH97]. Bibliographic [Cob98]. Bibliography [Bee96, Bee98]. Biblioteca [LJ97, LJxxd, LJxxc, LJxxa]. Bibliothek [Ano97-31, Ano97-32]. Bibliotheken [Küh97]. bien [VB96]. Big [Als95c, Ano96-178, Ano97-73, Egr99, Chw98, Gil97c, Lew96a, SF96a]. Bigger [Bou96b]. Bill [Cha99e, Gib95, Ham96c, Ten95a]. Billion [Gar97c]. Binary [CL96, DWE98, Gar97c, Lew96a, Lew96b, Lew97b, Lew98, You98b]. bind [Ude97b]. Binding [EMN98, Kro98d, Kro98c, MG97a, MG97b, Ano97d, DHM99a, Gre98c]. bioinformatics [Stu98]. bis [CH97f]. bit [Ano96-126, Bro97c, KG97b, LZ97]. Bittco [Kuk98]. BitWizard [Ano97-118]. Black [Ano96-67, FRT96, Hoqxx, Knoxx, Medxx, Wal98b, Eck97a, Kro98d, Kro98c, Com95g]. BlackBox [War99]. BlackWidow [Ano96m]. BLAS [BG97c]. blend [Joh97b, Ano97-74]. Blending [BK99, Sta98b]. blends [Chw98]. Blood [Swa97b]. blows [Ano96-42]. Blue [Gar97c, Kro98e]. Blueprints [Jac96, KDH96]. Boîte [Anoxx-37]. Boards [MD99]. boast [Mes97]. bodies [RFM95]. Boek [Vanxxe, Anoxx-60]. BOF [Sha95]. Bongo [Afe97, Ano96s, BS97a, BS97b, BS98b, Bec97b, BP97, BPxx, CA97, Goo97a, Goo97b, Gooxxa]. Bonjour [Anoxx-51]. Book [Ano96n, Ano96o, Ano96p, Ano97h, Ano97g, Ano97n, Ano97m, Ano97j, Ano97i, Ano97k, Ano97l, Ano98f, Ano98j, Ano98i, Ano98h, Ano98g, Ano98e, Ano98k, Ano99a, Ano99c, Ano99b, BDF96, BS97c, BSxx, Bri99a, CCYxxb, Cla97c, Dun99, Heu97, Hoqxx, Hum98, Jar98, KK96a, Lea97c, Lea98b, Nic96a, NFP96, Nic96b, Nov98, O’N97c, O’N97a, O’N97b, O’N98a, Roo98, Sal96, SMC97, Str97, Van97m, Wal98b, Yal98, vD99b, Ano96-155, KK97b, TGGP97, Llo98]. Bookish [Ano97o]. Books [Amb96, Ano98z, Lut99a, BDF96, Cla97c, Lut98a, Lut98b, Lut99c, Lut99d, Lut99e, Lut99b, Mat96a, Swa97a, Urg96b, Ano97-109, Hay96v, Ano97n]. Bookshelf [Wil97e, Ano98l, Ano98m, Frü99, GF99, Wil97i, MRA98, Wil97f, Wil97j, Wil97g, Wil97h, WC99, Wil99d, Wil99e, You97c, McC96e]. Bookworm [Sal96]. Boost [Ano96-71, The98, Mor97b, Net95c, Par95g, Hay96k, Hay96x, Mes97]. Booster [MD99]. boosts [Gau97q]. Borland [Ano95e, Ano95h, JSA98, Kro97, Kro99d, Man97a, Mil95a, PM96a, Sch97c, TD97, TJ96, Vanxxf, Zuk97a]. Borneo [Dar98a]. Boston [USE97a, USE98a, Her96b]. Both [Hug97, Gre96a, Mac96]. Boulder [USE99b]. Boulevard [ACM99c]. Bouncing [Gre96c]. Boundaries [GS98c, GS98d]. Boundary [Gil98a]. Box [Ano96-132, BS98e, Kro99d, Way96c, Way96g]. Boxed [MPxx]. boxes [Ano96-53]. Boxing [Den97]. Brace [Ano96-28]. Brain [DiG95b, KSI^{+}96]. braintrust [DiG95a]. Branch [VR99a]. Branché [vSS96a]. brand [Ano96-195]. Brazil [BK96a, dN97]. Brazilian [BK96a, dN97]. Break [Gan96]. Breakaway [Sun97a, Sunxxc]. Breaking [Che97]. breaks [Hay96m]. breakthrough [Ano96-59]. breeding [Gol98b]. Breeze [Ano96-42]. Brew [D’S96c, Gar99, Ess95, Gor96, Wil96a]. brewed [Fla96a]. Brewing [Ano96q, Ano96-110, Fis98b, Ros96, Ano96-165, Fre95, Gil96a]. Brews [Gar97a, Gro96b, Gut95h, Lou95c, Tri95, WL95, Ano96-124, Ano96-174, Ano97-144, Hay95a]. Bridges [Gre98f, Swa97e, Ano96-133]. Bridging [Kar98a, Kar98b]. briefing [van99b]. Briefs [Ano95m, Ano96-139, Ano97p, Cha99e, Gar97a, Gar97b, Gar97c, GC97, Gar97d, Gar98a, Gar98b, Gar99, Law98, Law99, Lea99]. Bring [Gar97c, Cha97c, Cli98b, Eti97]. Bringing [LK99, Vac96, Ban95a, FC96a, Gil97a]. Brings [Ano95p, Tra95b, Ano97-145, Way98a]. Broad [Gut95f, Gut95j, MAS96]. Broadcasting [Mag95]. Broadway [Lav96, Mio96]. Broker [Kro98d, Kro98c, Pop97, CDF^{+}98, Com95g]. brokerage [Cha96a]. brokering [TB96]. Brook [Tee99]. Browse [Ber97e]. Browser [Ano96r, Ano96l, Dadxx, JOA98, Kin97a, Law98, Mes95a, Net95c, Par95e, Ric95c, Tic97, Wat95d, Way96g, Way96f, Ano95s, CS97a, LR96, Lev96c]. Browser-side [Tic97]. browsers [BGGR98, CCB96a, GG96a, Oue96b, Com95g]. Browsing [Koh95]. BSP [Sar99]. BT [BH96c]. BT/MCI [BH96c]. bubbling [Hay96u]. Buch [CH97f, HKK96, Kno96a, Wil98a]. Bucks [Ano96-68]. budget [Ano96-151]. Bug [Ano97q, The98, Hay96g, Nic99, Phi99]. bugs [Ano96-43, Ano97-75, PSS98]. Build [MAS96, SS96, Ano96-97, Atk99, Mee96, Tur97]. Builder [Ano97-78, D^{+}98b, Ano97r].

BuilderPro [MD99]. Builders [Pel95, Bal96a]. Building [Amb97, Ano97-85, Ano97-155, BN97, Cag96, Cel98, CM97b, CM97c, CM97d, CM99, Hob98b, Hob98c, Hob98d, Hor96a, Hor96b, MD98a, MCxx, Mor98a, Mor97a, Ren98, Rit96a, SHW98a, TRMR97, Tur97, Vog96, VR99b, Ano96k, Fri98, GG99, MD98b, Mur96a, SHW98b, SH99, Woo99b]. builds [Ano95o, Ano96-42, Ano96-49, Ano97t, DiD97b, Gau97b]. Built [Gar96e, Sch99c, S^{+}96b]. Built-In [S^{+}96b, Sch99c]. bullet [Gil96b]. Bundle [CBT98c, CBT98e, Pep97]. Bundles [Gut95b, Ano97-146]. Bunny [Egr99]. Business [And97a, Ano96-50, Ano97-73, Bow96a, GC97, Lut98a, MAS96, MF98, MF99, Piexx, Sco96a, Sco96b, Sun97a, Sunxxc, Ude96, Wil97a, van98a, Ano96-69, Ano97-43, Ano97-69, Ano97-83, Ano97-100, Ban95a, Car96c, DR98, DR99, Gau97k, GO96, Ham96a, Wat95f]. business-school [Ban95a]. businesses [FC96a]. Businessorientierte [Pie97]. Busting [Sho96a]. Button [Wag97a]. Button-down [Wag97a]. buttons [DiG96]. Buzz [San97a, Val95, Ban95a, Ven96]. Byte [Bal97, KF99a, KF99b, PV98]. Byte-Codes [KF99a, KF99b]. Bytecode [BFPV99, BG98b, BG98c, FM98, Gol98a, Int96a, MO98, MMBC97, PW97, Pus99, SA99, VP98, WdFD^{+}98, WDDF98, Ano96-200, BG98a, Cla97d, FM99a, FM99b, HGH96, LP96a, MJ99, MW98, O’C99, SA98b, Tow98]. Bytecode-to-Native [WdFD^{+}98, WDDF98]. Bytecodes [BM98a, BKR99, LZ97, CL97f, GTV98a, GTV98b, Gos95, HAKN97].

C [Ano96-42, Ano96-137, Wil97e, Biexxb, Bri99a, Her97, PM96a, Red96, Redxx, Sch97c, TJ96, Ver98, Wil97f, van97f, Ait96b, Ano96q, Ano96-54, Ano96-83, Ano96-197, Wil97e, Ano97r, Ano97-49, Ano97-103, Ano97-124, Ano97-165, Ast98, AC99, Ber96i, Ber96j, Biexxa, Biexxb, Bla96, Boo96c, Bor96, Buc97, CBT98e, CBT98d, Car96a, Carxx, Che98, CDL99, Cro99, DD96, D^{+}98a, D^{+}98b, DSM99, Dacxx, Dan98, Dav97d, Dem98b, Dem98a, Drixx, FLRT97, FC96a, Gar98b, Gau97g, Gau97j, Gre96b, Han97, Hav98, Hem96c, Hor97a, JS97, JLW96, Kal98, Kay96a, Kei95, Kro99e, Kro99c, Lafxxb, Lak96, Lin96f, MGL^{+}98, Mar97d, Mat96b, McK98, Met98, Mic97a, MMG98, Ols97, Par97, Pel96, Per96a, Phi99, Pla96b, Pla96a, Pla97, Pre99, Rag97, Roh96, Sad97a, Sam96b, Sch96a, Sch97h, Shu97b]. C [Smixxa, Spe97, Ste96d, Ste96e, Ste97e, Ste99, SW99, Syd96b, TMP^{+}99, Tuc98b, Van96f, VR99b, Whe97, Whi97a, WL97b, Wil97f, Xu97, Zei97, Zho97]. C/C [Ano96-197, DD96, Dacxx, Kro99e, Mic97a, Pla96b, Pla96a, Pre99, SW99, MMG98]. C/Dos/Us [Ano96-83]. C/Dos/Ww [Ano97-49]. C/S [FC96a, Lin96f]. C/W95/Ww [Per96a]. C/Ww95/Ww [Sam96b, Van96f]. C2J [Lafxxb]. CA [Ano96n, Ano96o, Ano96p, Ano97h, Ano97g, Ano97j, Ano97i, Ano97k, Ano98f, Ano98j, Ano98h, Ano98g, Ano99a, Ano99c, Ano99b, Fox98b, RD99, Ano98a, IEE96h, Uni96a, Ass96b, USE96b, Wae96, YN96]. Cabinet [Sch97i]. cable [Ano96-165, Ubo96b, Wat95f]. Cabletron [Dry97]. CACAO [KG97b]. Cache [CDL99, Kro98e, Lut98a, Yos97, San98a]. Cache-Conscious [CDL99]. caching [LM99, Tým99]. CAD [Kro99e, MD99, Web97b]. CAD-UL [Kro99e]. CAD/CAM [MD99]. Café [Anoxx-74, EM96, Ano96-35, Ano96-182, Ano97-154, Bar96, BLT98, Coh97e, CH97a, EMe97b, FKPR96, FKP97, F^{+}xxc, Frixx, FPxx, GPF97, Hem96a, JD97, JLP96, Kri96, Kro97, Krü97b, Krüxxb, Mes97, R^{+}xxa, Tit97, WG97b]. cafeteria [Vij97b]. caffeinated [Ano96-78]. Caffeinator [Met98]. Caffeine [HGH96]. cage [Ano97-70]. CAI [Kro98d, Kro98c]. Calculus [WJUH98, IPW99]. Caldera [Ano97-121]. Calendar [Kum98, Ude97b]. Calendars [Lam99]. CALICO [BBJR96]. California [ACM98b, BUP99, CEG96, IEE96c, IEE96e, IEE96a, IEE96m, IEE97, IEE99a, KJR98, Nat98, Spe98, Ten99, Uni96b, Uni97b, Uni98b, USE96d, USE97c, USE98d, USE99c, W^{+}98]. CALISCE [DF96]. call [Ano97-136, Isa96]. Callbacks [Ses98]. Called [vSS96e]. calls [Gol96a, Wal98a]. CAM [MD99]. Cambridge [Ano97l, Ano98i, Ano98e, Ano98k, USE99e, vD99b]. Campos [dN97]. camps [Cha97c]. Can [Ano96-45, Ber96h, Ber97c, EMN98, Hum96, Pit99, Ran99a, Sin95a, Sli97, Ano96-171, Ano97-30, Ber96b, Col96b, Eck96a, Gau97r, Gre97k, Hay96b, Hay96o, LP96a, Mey96a, Orc98, Stu98, Ude97d, Wal96h, Way96b]. Canada [Ano96-158, Doo99, FGL96, FW96, IEE98a, MJ98, Tau96, Ass96a, USE96c, Cou96b]. cancels [Ano96-195]. Candidate [Dra98, Nat98, Nat99, RD99]. cans [Ung97]. Capabilities [Ano97y, Lav96, Nil98a]. Capability [VCCH99, Ano96-188]. Capability-Based [VCCH99]. Capacities [The98]. Capacity [Gar97a, Gar98b]. Capitulation [JN95a]. capstone [Ada98]. Captioning [Fre96a]. Capturing [Fue98]. Car [The98]. Card [Ano97-147, BS97c, BSxx, Cha96d, Cha96e, Chaxxb, Coh97c, Gar97a, SUN97b, Mar98b, Mar99a, Bar99, Bro97c, HNSS99, MK99, OK99, JCV99]. Cards [Cha96e, DDJ99a, PV98, Wae96, Bro97c, Gau97q, GLV99]. care [BCF^{+}97, Chu98b, IDK98]. Career [Nau98, Saa96, Gof97]. Careers [Wil96a, Ano97-108]. Cargo [Wil97i, Wil97j]. Cargo-Cult [Wil97i, Wil97j]. CASCON’98 [MJ98]. Case [Ano97-49, Bin97, CJJ^{+}98b, Lut99c, TWP98, AGD97, Ban95a, CJJ^{+}98a]. Case-Based [TWP98]. Case-Driven [Lut99c]. Cases [ZJ96]. cash [Wil96c]. Castanet [Ano96s, BS97a, BS97b, BS98b, BP97, BPxx, Ber98a, Lem97d, Lem97e, Lemxxd]. Catalog [Gra99a]. Catch [PK97, S^{+}97]. Catching [San97a]. categories [Soo98b]. Cathedral [BUP99]. Causing [Lou95a]. CBT [BKxxa]. CC [Com95c]. CD [Buhxx, Tee99, Anoxx-53, BDF96, Dav96b, Fla99d, GG96a, Kay97, Mon99, NR96, Neo96a, Neo96b, Neo97, NR97, Neo98, Neo99, N^{+}96b, Sad96, Sad97b, Sad98, Ste96d, Ste96e, TMP^{+}99]. CD-ROM [Buhxx, Dav96b, Fla99d, GG96a, Kay97, Neo97, N^{+}96b, Ste96d, Ste96e]. CD-ROMs [BDF96]. CE [Kro98d, Kro98c]. Center [ACM99c, Ano98a, Ano98-29, IEE96d, IEE96r, Mor97b, Ano97-136, Law99]. Centers [Ano97-115]. Central [Ano97s]. Centre [IEE98a]. Centric [FC96a, Ham96b, Kro99e, WWR97a, WWR97b, Ham96a, MC97, SF96c, Wal99]. CEO [Ano97-52]. Certification [Boo97, Boo98, Jar98, RH97b, RH98, RHE99, Tuc98a, WL97b, Ano97-162]. Certified [Kro98b]. CEV [BRF98, RBF98]. CG [SL99]. CGI [Gar97e, Abexxb, Anoxxn, Bre96, Bro96, Chuxx, Cli98b, Coh96a, Coh97a, Con97b, CA98, GJ96d, Gol98b, Hal97b, JLW96, LO96, LOA^{+}96, Lad96, MAS96, Moh96, PWH97, P^{+}98b, S^{+}96a, SR97, SD96, Ste97d, Ste97e, SS96, T^{+}xxb, Ude97d]. CGI/Perl [Coh96a, Coh97a, PWH97]. Chagas [WBA^{+}96]. Chai [Hew99]. chairman [Ano97-52]. Challenge [Cou97a, DDJ98b, DDJ98c, Rym95]. challenges [Ano96-171]. Chance [Ada98]. Chance-It [Ada98]. change [Ano97-52, Fre95, LaM95a, Sco96a, SvK98]. changed [GS97b, Sco96a]. changes [Gau97n, HCD^{+}98, Mey98c]. Channel [Ber97c, Ber98a]. Channels [BC98a, RMS97]. Chapter [DHM99b]. character [Che97]. Charles [Tee99, Ano97t]. Charlotte [BKKW96]. Chart [Ano97-100, Ano97-118, Chaxxc, Mee96, Kro97]. Charting [Ano96-51]. Chases [VN96]. Chasing [Gar97c]. Chat [Ber96g]. Cheap [Ano98m, CTL98]. checked [Bok99]. Checker [Kro99a, Kro99b]. Checking [BFPV99, Kro99c, PV98, VP98, VN99, Ala99, McC97b, MMS98, Orc98]. checks [MF97a]. Chemistry [Lut99a, Lut99b, Mos98b, Mos98a]. Chemkin [Ano97-153, Bra97a, Bra97b]. CHI [Tau96]. Chicago [Ano96-160, USE99d, Ano97u]. Chief [Agh97]. chill [Lew96a]. Chinese [Lun99, Flaxxb, Kanxxb, Tri95]. chinks [Col96c]. CHIP [Ano96-70, GC97, Gro96b, Lea99, Wu97, Ano96-135, Ano96-188, Kan96b, Mon97, KS96a]. Chips [Ano96t, Ano96-71, IEE97, IEE99a, Way96e, Ano97-143, IEE96e]. Chisel [Bry98]. choice [Way98c]. choices [LS99]. Chores [Win95i]. chotsarang [Kim97d]. Chronicle [Sto97]. chuan [Nau96b]. CICS [Ano96-48, Ano97-72]. CIFEr [IEE96p]. Cimlogic [MD99]. Cisco [Ano97v]. City [IEE96p, Isk96, IDK98]. CJKV [Lun99]. Claims [Bet95]. Clara [IEE96c, Swa97e]. Clarifying [GF99]. ClarisWorks [Fei98]. Class [Add99, Ano96d, Ano96-137, Wil97e, Anoxx-38, Cha96b, CL97b, CL97a, CLK98, CL98a, CL98b, CL99, CLK99a, CLK99b, CLxxa, CLxxb, CLxxc, CDL99, Dav98a, Die97, EMe97i, Gar96f, Gra96b, HC98a, LB98b, Pug99, San97b, Tuc98b, Wal96d, Wil97f, You98b, CKH99, Cha96c, DV98, Nap98, SG97a, Whi96c, Won97, You98a]. Classes [Ano97-124, Ben99a, CBT97b, CBT98c, Com95d, DW99, FKF98, Gra96c, GK97b, Gri99, HHW^{+}99, JV98, LST99, MB98b, Nel98, Ran99a, Shu97b, Sta96, TWP98, Top98, Top99, Wal96f, Ano96y, Ast98, Cli98b, D’S97c, GHP^{+}96, JvdBH^{+}98, SS98a, SS98b, Wal96c, Fla99b, Gut98, Tur97]. classification [JJBW98, JJBW98]. Classroom [Ber96j, DD97b, DD97e, DD97f, DD97g]. cleans [Ant96]. clever [Den97]. Click [MO96]. Client [Al-97, Ano96u, Ano98n, Ano98-29, Bij96, Bla96, CH98b, Frü99, GDS98, Ham97b, Tou99, LaM95g, MC97, Mor97c, N^{+}xxc, Nov97, OH97c, OH98, OH99, SWJ97, TMM^{+}98, TMMH99, Uma97, Ale96, Ano96-72, Ano97b, Ano97r, Ano97-156, Col96a, Hay96q, HG96b, HL97b, HL97c, KW96, Sco96a, YB99, OHxxa]. Client-Server [GDS98, Ham97b, Nov97, TMM^{+}98, TMMH99, OHxxa]. Client-side [Al-97, HL97b, HL97c]. client/ [Sco96a]. Client/Server [Ano98n, Frü99, LaM95g, OH97c, OH98, OH99, SWJ97, Uma97, Ano96u, Bij96, Bla96, CH98b, Ale96, Ano96-72, Ano97r, KW96, YB99]. clients [Lev96c, Sco96a]. Clinical [ZJ96]. Clipper [Kan96b]. clips [Fre96a]. Clones [Gar96a]. close [Hac98]. closer [Gau97r]. closet [Ano97r]. Closures [Ben99a]. Cluster [KB98a, KB98b]. Clustering [DDJ98a, EKR97]. Clusters [GPHL99]. Clustor [Ano97-118]. Co [Die96, Diexx, Mag95]. Coarse [JLT98, BBC^{+}97]. Coarse-grained [JLT98]. Cobalt [Ano98-28]. COBOL [Met98, Ano96-168, Met98, HHW^{+}99]. CocoBase [Ano96-137]. Code [Ano97x, Ano97-128, Ano97-138, AC98, Bal97, Fra98a, Gre97h, Ham97b, Int96a, Kro98e, Kro98b, MF98, MF99, MMBC97, Pat99, PV98, TJ96, Wak98a, Zha97, ATCL^{+}98, Bru97, Che97, Dug99, EL98, FC98b, Gil97a, GA98, HGH96, Pla96a, Tho97a, Wil98b, Ano96z, Gre97h]. Codes [KF99a, KF99b]. Codetalk [Ano96v]. CodeWarrior [Met95d]. CodeWizard [Ano97-117]. coding [Hay97b, O’C97]. Coffee [Ano96-67, Beh95, CBS95, Knoxx, Medxx, Par95f, Lie96b]. CoffeeStrainer [Bok99]. Cold [Oly96, Sin95b]. Coldfire [Kro99d]. Collaborate [Ano96-48]. Collaborates [Met95d]. Collaboration [DW97a, BRS99, CC97, DAB97, Fuc96, Mar99b]. Collaborative [BRF98, IEE96m, JAS96a, KK97c, KK98a, MD97b, Ano96-128, BCF^{+}97, EHS^{+}99, LCLL97, RBF98, RMB^{+}98, BR97]. Collabra [Par95a]. Collecting [AR98]. Collection [AR98, ADM98, JL96, KT99, PB98, Spe97, SLC99, Eck96a, RBL^{+}97]. Collector [Roo97]. Collocation [Doe95]. Colloquium [Ano97-54]. Color [CLD99, MD99, Ano96-196, MD99]. Colorado [USE99b]. CoLoS [CT96]. Coloured [KMW99]. Column [Als95c, Als95a, Als95b, Met95a, Met95b, Met95c, Pet95, Rym95, Sin95b, Sin95a, DiG95c]. com [Ben96b, HI96a, HIxxa, Rit97a, Ritxxc, Lut99a, Ban98, Cha99f, FFHA98, FFKO99, Kro98a, Lut99b, PS98]. COM-CORBA [Lut99a, Lut99b]. COM/DCOM [PS98]. COM/MTS [Cha99f]. combination [Mac96]. combine [Ano96-117]. combined [Gre97h]. combines [Wag97b]. Comdex [Bus95, Gut95d, Gut95c]. come [Atw96, Lin96d, RBL^{+}97]. Comes [Gro96b, Ano97r, Bat96c]. coming [HCD^{+}98]. Comm [Ano98-28]. Command [Whi96a, BF99b]. Command-Line [Whi96a]. Comme [Anoxx-51]. comment [Ait96a]. Comments [Roo97]. Commerce [Bet95, Bus95, GS97a, Gol97, Ham97b, Jar97, Lit97, Mes95b, Mes95c, TW96, USE96b, USE98a, Ano95s, Gar97c, Her96b, LM98, NB96, VM96, Wil96c]. Commercial [Ano96-146, Boz95a, Kro98e, Woo99a]. Commitment [Sim97a]. Committment [Sim97b]. Commodity [FFHA98, FFKO99]. common [Ber99a, Pop97]. communicate [Ano96-59]. Communication [BC98a, CR97, GSTV97, GGS99, Gra96d, KMW99, Mec97, YC98a, YC98b, Bla96, FTT97, GH97, JS97, Neu97, Wel99a, Ano98-30]. Communications [Cli98a, Cou97g, GS98a, IEE96r, Wip96]. communicators [Ano97-54]. Community [IEE96b, Mun96, Car95, Jav96d]. Como [Anoxxc, Anoxxd, Ben96b, DDxx]. Compact [MD99, Cor98a]. Companies [And97a, Hes95, Urg96b, Ano96-53, Ano96-59, Ano96-183, Ano97-108, McC98d]. Companion [Fla97a, R^{+}xxa, SR97]. Company [GC97, Gau97c]. Compaq [DeV95]. comparative [Xu97]. compare [Shu97b]. Comparing [Cro99, Joh99, O’C98, Phi99, Pre99]. Comparison [Ano96-111, Bro97b, Bro98b, Bro98d, GFR98, Ano96l, Hop97, Lev96c, Mar97d, MMG98, Shu97b, Zho97]. Comparisons [Fol99, GGS99]. Compatibility [DWE98, Web97c, Web97d, Mat96b]. Compatible [CS98]. COMPCON [IEE96c]. compete [Ano96-74]. Competition [Ano96w, Wil96a]. Competitive [Ano97w]. competitor [Ano95i, Ano96-74, Ano96-73]. Compilation [Ano99g, Anoxx-53, BS98c, BS98d, Cla97a, Die98, CCCL97, CFM^{+}97a, CFM^{+}97b, vRvGS97]. Compiled [MMBC97, LP96a]. Compiler [App97, App98, BBGD98, BVG98, Kro98b, Kro99d, PTB^{+}97, RC97, Sch99a, SLC99, Wak98a, WdFD^{+}98, WDDF98, vD99b, ATCL^{+}98, Ano97l, Ano98k, BVG97, BGH99, DDG^{+}96, Gau96f, Gre96b, KG97b, RKCW98, Tow98, Wol97]. Compilers [Fli98, HCJ^{+}97, KFL99, Hal98a, WB99, Way96e, van97f]. Compiling [BKR99, Bot98a, CFM^{+}97a, CFM^{+}97b, DD97i, MJ99, Wak98b, Wak99, WJUH98, DDS99, HNS97a]. Completa [LP97b, LP97e]. Complete [Anu96b, Anu96a, Cobxxa, DD97b, DD97a, Dei97, Fla99d, LB97b, Man96c, NS97b, NS97a, NS98a, Nor96, RHE99, Sch97c, Sip98, Sta96, Vanxxe, WW96b, Wei97a, Ano96-75, FHKS98, McCxxa]. completely [Ano95r]. complex [Ano97-74, EM96]. complexity [Ano96-171, Rit97d]. Component [Ano98-30, Chu98a, Kro98a, Kro98d, Kro98e, Kro98c, Kro99a, Kro99b, Nicxx, NK98, SHGD98a, TBS97, UCD99, WKES98, Gre98d, LaM95a, SHGD98b, Val99, WKE98, CDF^{+}98]. Component-Based [UCD99, SHGD98a, SHGD98b, Val99]. Component-Savvy [Ano98-30]. Components [Ano96-76, Ano97-123, HS98a, Kro98e, Kro99e, Lad97, PH99b, PM97d, Ren98, Shi99, Wat97a, Ano96-181, CAZZ98, HC99a, PS98, Tra97]. Composants [Cha99f]. Composes [Way97b]. composition [AA98, CM97a]. Compositional [WR99b, Yel99]. Compound [BW98b]. Comprehensive [Bro96, JRC97, LV99, Ano98y, GHP^{+}96, JC97, NMS96, O’C99, SvJ96]. compress [NN97]. Compressing [Pug99]. Compression [HC98a]. COMPSAC [IEE96s, IEE98b]. CompuServe [GC97]. Computation [AT97, IG99, Gos98a, Num98, Syd97, BCK99, CCCL97, RHND99, Doo99]. Computational [ACM96, IEE96p, Mos98b, Mos98a, AGD97, Fox97a, Fox97b, FD96c, SZTR96]. Computations [AFH98a, GDS98, AFH98b, Fox98b]. Computer [AL97, Amm98, Ano95j, Ano97-137, Ano98-27, BAC^{+}99, BBJR96, BK97b, Buhxx, Gar97b, Gos98b, Hiu97c, Hiu98, IEE96c, IEE96h, IEE96s, IEE98b, JG96, Joy98, KMR98, Kuk98, MB96a, Par95b, DF96, SEH98, Sch99b, SPV99, Tra95a, Tra95b, WKES98, Web96, Wil96a, dN97, Ano97-52, Ano97-71, BV98, BNB^{+}98, BGGR98, CH98a, GJ96b, HCD^{+}98, Muk99, Neu97, SEH99a, SEH99b, SBB^{+}97, Sch97f, Sun97c, Tob99, WKE98, Mon97]. computer-communication [Neu97]. Computers [Kan96b, Lut98b, Gau97o, Mac97c, Way96c]. Computing [ACM98a, Ano98a, Ano98n, Anoxxk, BKKK98a, BS98a, Ber98h, Ber98i, BFD96, BC98b, Bue98, CBT98f, CCF99, Cha99a, Cha99e, Cul97c, DDJ99a, DBBE98, Far98a, Fer98b, FC96a, FFHA98, Gos97b, GS98c, Ham96b, Ham97b, Hor98a, HS98b, Hun97, IEE96n, IEE96q, IDK98, JG96, KB98a, Law98, Lun99, MO98, Mur99a, RA97, Raz98, Sar98, SHW98a, Sar99, Sch96b, Sun97a, Sunxxc, Tau96, Ten99, WW97, WR99a, WWR97a, WWR97b, WRB^{+}97, YH96, ZTLV98a, van97e, von98, AISS97, Ano96m, Ano96-192, Ano97-52, BKKK98b, BCF^{+}97, Ber98g, Ber99b, Ber99c, BBC^{+}97, BC99b, Car95, CKV98, Cha99b, CMM^{+}97, CCI^{+}97, Cul97a, Cul98g, Far96a, Fer98a, FF96b, FF97, Fox98a, Fox98b, GGMS99, Gon97, GS97c, GS98d, GME99, HVB^{+}99, Jav98, KJR98, KB98b, LG97, Liu96, Lu96, MMG98, NAC99, SHW98b, SH99, Sco96a]. computing [SF96c, Vil99, Wal99, Wal96i, YC97, ZTLV98b, Com95g, RSH97]. Con [DDJ99b, Anoxxd, Cha98b, Eng98, GR97, Hob97b, Hob98a, Lesxx, RVM97, Rit97b, Rit97c, Ritxxd, Weh98a]. Concentrated [Sau96]. Concepts [Hor98a, HS98c, Uzg96, Ano96u, BT99, DV98, G^{+}xxa]. Conceptual [PIPR97a, PIPR97b, PPIG98, Gre98a]. concerning [MKS^{+}96]. concerns [Nat96]. Concinity [Ham97b]. concorrente [Tis97]. Concurrency [Bro98d, CJJ^{+}98b, Lut98b, KCPC98, Ver96]. Concurrent [Agh97, CR99, Lea96a, Lea97b, LH97, Lea98a, Yal98, Cor98a, DIS99, Miz99, NAC99, Nev99]. Conducts [Law98]. Conference [ACM97a, ACM97b, ACM98b, ACM99b, Abr96, Ano96-142, Ano96-147, Ano96-158, Ano96-154, Ano96-160, Ano97-29, Ano98a, Ano99h, AOV^{+}99, BUP99, Cha99a, DDJ98a, FGL96, IEE96c, IEE96g, IEE96h, IEE96i, IEE96r, IEE96o, IEE96p, IEE96k, IEE96q, IEE96s, IEE98b, IEE98a, IEE98c, Isk96, MDM97, Nat98, Nat99, OW95, RD99, SMEN96, Sta97b, Str96, Tau96, Uni96b, Uni97a, Uni97b, Uni98a, Uni98b, Ass96a, USE96c, USE97b, USE98b, USE98d, USE99a, USE99c, Wae96, YH96, ACM99d, Ano96-73, Ano96-100, Ano97w, Ano98-34, Bod96, Cul97a, Cul98g, LM98, DF96, Uni96a, USE96d, USE98c, USE98e, USE98f, Cul97c, Ber98h, Ber98i, Her96b, Kal95b]. conferencing [ERKE96]. Configuration [HLRT97, Gol96a]. Confusion [GC97]. Congestion [Lea99]. Congress [Syd97, Ano98o]. Connect [Moh96]. Connecting [Hoq98]. Connection [Ano96-134, Hem96c, MAS96, Rei99, Ano96-125]. Connectivity [H^{+}xxa, Joc98, McD99b, San98b, San98c, Van97g]. connects [Com95g]. conquer [Lew96b]. Conquering [PF96]. Cons [von98]. Conscious [CDL99]. Consistency [HK98]. Consortium [BBJR96, WK95, Kal95b]. constraint [LQ99]. Constraints [Lut98a, Bok99]. Constructing [BB97b, Cor98a, WC98]. Construction [Mey97, Sch99a, Cul98c, Dav97a]. constructivist [GBS98, Had98a, Had99]. constructors [But99]. constructs [CTL98]. Consulta [Ano97-97]. Consumer [Boz95d, Gau97d]. consumer-friendly [Gau97d]. Contact [GM97a]. Contain [MB98b]. Content [IGO98, IAJR97, JPLI99, RK97a, SJRK97, Ste97g]. Contest [DDJ98a]. Context [JG96, For96]. Continued [Ano96-93, Fre97d, CWHt99]. continuous [CCCL97]. Contract [KHB99a, KHB99b, PJR^{+}98, Ran99b]. Contractor [KHB99a]. Contrasts [GGS99]. Control [All99a, CRS98a, CMSvE99, HHW^{+}99, JPLI99, Kro98d, Kro98c, NRG^{+}99, PH99a, TBS97, BF99b, CRS98b, Gau97n, Hay96b, Men96, MS96, Mes97, Mey96a, Mye99]. Controllable [SMY99]. Controlling [RK97a, Wil96f]. Controls [Inf95a, Kro98d, Kro98c, Kro99c, ZK98a, ZK98b, Cor97b]. Controversies [Ste96a]. Controversy [Law98]. Convention [ACM99c, Ano98a, IEE96d, IEE96r]. Convergence [Cha97c, CSY97, GC97]. Conversation [Woe97]. conversion [Hay96l]. Convert [Gre97a]. Converter [Met98, FLQZ97, Met98]. Converting [Ano97x, Del97, Dem98b, Dem98a, JS97]. Converts [Int96a]. Cookbook [Bra99, Coh97d, Sad96, Sad97b, Tee99, PWH97]. Cookies [Tic97, ZK98a, ZK98b]. Cooking [Anoxxe, JT97, JTxx]. Cool [Als95b, Per96c, Per97c, SS96, Tra95a, Wod96, Gil96c, Wag96]. Cooling [Gil96b]. coordinate [CKTV96]. Coordination [CRV^{+}96, CR97, HK98, Fri96, Tol97a, Tol97b]. COOTS [Ano96-158, Ass96a, USE96c, USE97b, USE98b, USE99c]. Coping [Gon97]. Corba [Cha99f, Ano97y, Ano98t, Lut99a, Anoxxf, Bak97a, Bakxx, BN98a, Ber97a, BF99b, Cha97b, Chaxxd, CSY97, Cor98b, Desxx, FFHA98, FFKO99, Frü99, Gar97b, Gau97e, Gau97h, G^{+}xxa, Hou98, Inf95f, JAS96a, JADR96, KS98a, KS98b, Kro98d, Kro98c, LPR97, LB97a, LBS98, Lut99b, MZ97, Mor98a, MR97, OH97a, OHE97a, OH97c, OHE97b, OH98, Orf98, OHxxa, OHxxb, P^{+}98a, PS98, Pop97, Com95g, Red96, Redxx, Ros98, SLPM96, Stu98, VD97, VD98, ZTLV98a, ZTLV98b, Chw98]. CORBA-based [JAS96a, JADR96, KS98a, KS98b, LPR97]. CORBA-Java [Stu98]. Corba/OTS [Cha99f]. CORBAplus [Ano97z]. Core [Anoxx-38, Atk99, CH96, CH97b, CH97c, CH97d, CH97e, C^{+}97b, CHxxa, CHxxb, CHxxc, EMe97i, Edw99, GYT96c, Hal97b, Hal99, HC97, HC98b, HC99b, HC99c, MT97, Nie97, PCxx, TK98, TK99b, Top98, Top99, deC99, IPW99, MN96b, CH98c, HC98c]. Corel [Ano98-27, Ano96x, Anoxxg, Cou96b, Hay97d, MK97, Pep97, Pic96a, Ung97]. Corner [Cha99a]. Corp [Ano98-27, Kuk98]. Corporate [Boo96b, Car96d, Hay96a, Sim97d, Wal96q, Ano96-72, Ano96-125, Dua96, GH96a]. Corporation [Ano98-29, Ano98-29]. Correct [BS98c, BS98d]. Correction [Ano96y, Ano00]. Correspondence [Han99b, Sak98]. COS [Ano96-135, Ano96-137]. COS/Print [Ano96-135]. COS/Relay [Ano96-137]. Cosmic [Gre97h]. Cosmo [Ano96z, Ano96-202, Ano97-138, Gre97h]. Cost [Kro98b, Ano96-53, Ano97-103, Gut96]. Costs [Ano96-28, Ano96-77]. Could [Als95a, Gar97b, Gar97c, WN97, Ano96-59, Car95, KW96, Kar95b]. Counting [LaM95a]. Countries [Gar97a]. counts [Gri97b]. County [Swa97e]. coupling [GRN99]. Course [Ano96-92, Anoxx-59, Bay97b, Dan97c, DD97b, DD97a, Dei97, LPMG97, LB97b, Sav99, Ano96-53, Ano96-145, Ano97-91, Asl98, Bös98, Car96c, CS99, KSW98, KR99, Sch97h]. Courses [Ano97-101, Inf95e, MHP98, Ast98]. Courseware [Div98]. Court [WBC^{+}96]. Covalent [Ano98-28]. Cover [Ano96-27, Van97n]. Coverage [SR97]. Covers [Cho97b, Kar96, Ano97d, Ano97-91, Way96c]. Covert [HS98b]. CoWeb [JGKR96]. CoWeb-architecture [JGKR96]. Cowell [Heu97]. Cozy [Hes97]. CPU [KH99]. CPUs [Gar96e, Way96e]. Cracked [Gar96b]. Cracking [GPO98a, GPO98b]. Craft [Wil97e, Wil97f]. Craftworks [Ano97-121]. Cram [Bro99a, Bro99b]. Crash [Bay97b, Ano96-53]. Crashkurs [KS96b, Dan96b]. crave [Gau97m]. Cray [Way96c]. CRC [BS97c, BSxx, Hen97b]. Creare [GGxxa]. Create [Ano97-27, BSS96, Ber96d, Ber96h, Ber97i, Ber97c, Ber97h, Ber97g, Ber97f, Ber98a, BLT98, CCB96b, Har96a, Moo98, S^{+}96a, Sch97c, Smixxc, Tes96b, Van97n, Ano95h, Ano95o, Ber96b, Bru97, Fet97, McC96d, Ubo96b]. Created [Law99]. Creates [Ano96x, Ano97-100, Ano98d, Not96]. Creating [Bri99b, Dav97c, DBK98, FSW98, GRD96, H^{+}xxb, IBM97, Lut99d, Man96d, McC98b, Per96c, Per97c, PWxx, Rod96b, SC96a, SS96, TPHB96, Wat97a, Wod96, tYpCxx, Anu96b, Hei97, KK97b, vSS96b, vSS96c]. Creation [Gul96a, Wae96, Coh96a, How96]. Creative [WW98]. Creator [Gut95a, McC96c]. CRED [Dra97]. Credit [Gro96b, Gau97q]. Cree [GR97]. Crie [McCxxb]. crime [Ano96b]. CRiSP [Ano96-137]. Critic [CL96, Lew96a, Lew96b, Lew97b, Lew98]. critical [DR98, DR99, Had98b, Mar97d]. critics [Mac96]. critique [Thi99]. Crop [UBF^{+}98, BDF^{+}98]. Cross [Ano96-48, Ano96-137, Che98, KGTF97, NRG^{+}99, Net95a, Ano96-90, Gau97c, Gau97e, Gau97f, Gon97, Gre97f, SG97b]. Cross-Platform [Ano96-137, KGTF97, NRG^{+}99, Net95a, Ano96-90, Gau97c, Gau97e, Gau97f, Gon97, Gre97f, SG97b]. Crossroads [Sch96b]. Crown [IEE96p, GH96c]. crunch [Stu98]. crusaded [Gan97]. Cryptographic [NK98, Way98c]. Cryptography [DDJ98a, Gon98, HH98, Knu98b, Ano98h]. CS [Ada98, Ciz97, RP98]. CS-1 [Ada98]. CS1 [KW99, War99]. CS1/CS2 [War99]. CS2 [CS99, GG99, War99]. CSE [Gro96c]. CSP [Nev99]. CSX [Meh97, Wil96b]. CT [Wal96a]. Cult [BCP^{+}99b, Wil97i, Wil97j]. Cultural [Lut99a, Lut99b]. Cup [Ano96-183, Com95f, Egr99, Sin96a, HFA^{+}98]. cupful [Ano95s]. Curitiba [BK96a]. Current [Soo98b, Fet97]. Curricula [Cha99a, Cha99b, CT98]. curricular [BN98b]. Curriculum [Ber98h, Ber98i, Cul96b, Cul97b, Cul97c, Cul97a, Cul98g, Had98a]. Curves [HHW^{+}99]. CurveSoft [Ano98-27]. Custom [Ano98-28, ZJ96]. Customer [Frü99]. Customers [And97a, Ano96-77, Ano96-59]. customizable [WOC99]. Customization [AA98, Vac96, LTC96]. customize [Fet97]. Cut [Win95j, Sli97]. cuts [Ano96-77]. Cutting [Ano96-28, Ano97-28, BSS96, Esp96, Wil97a]. Cutting-Edge [BSS96]. CVS [BAC^{+}99]. Cyber [DD97b, DD97e, DD97f, DD97g, Die97]. Cyberclassroom [Dei97]. cyberculture [BKP97a]. CyberCut [SW96]. Cyberdog [Ano96-133]. cyberpunks [LP96a]. Cyberspace [Lev96a]. cybertheatre [Sha98b]. Cyberwar [Kan96b]. cycles [Ano97-71]. Cyrix [Ano96-135, Ano96-188].

D [Ano98g, Ano99b, Kuk98, MD99, NGM^{+}97, Ano96-53, Ano96-72, Ano96-133, Ano96-151, Ano97-44, CH98a, Gre97h, Gro96b, Hor96a, Hor96b, Lau96, MD99, Mag95, MD98a, Sch97j, Sun97c, Win95k]. D&B [Boo96b]. Daconta [Ver98]. Dados [Jep96]. Daiwa [Sli97]. Dalang [WS99]. DAMSEL [PS96]. Dance [Tub95]. Dangerous [Mar98a, Pon98]. Dangerously [Sch97f]. Daniel [Ano98e]. Danny [Goo96a]. d’Applets [Gul96a]. DARP [AFH98a, AFH98b]. Data [AFH98a, Ano97-119, Ano98-28, Anoxxh, BAC^{+}99, Ber98d, Ber98c, Boy98a, BW99, CZF^{+}98b, Chw98, CJ98, DW96b, For96, GDS98, GPHL99, Gol99, GT97, Gre97d, IR97, Jen98, KK98a, Kro98d, Kro98c, Laf98a, LB98a, Li98, Lut98b, Mai98, MMGL99, Row97, Sak98, Sha98a, Sha97b, Sta98a, Tes96a, WR99a, Wei98, WCF98, AFH98b, Ano96-59, Ano96-153, Ano97w, Ano97-50, Ano97-142, BBG^{+}99, CZF^{+}98a, CH98b, Che99, Ham96a, Hav98, KK97c, Knu97f, Knu97e, MKS^{+}96, Reg96, Ste96f, Stu98, TGGP97, Var96, Wei97c, WFC^{+}99, Wol99, Ano97s, Ano97-133, Kro99d, LaM95c]. data-centric [Ham96a]. Data-Modeling [Lut98b]. Data-Parallel [IR97]. DataBase [Ano96-137, Ano96-185, Ano97-27, Ano97-118, Ano97-119, Anoxx-64, CW96b, Cog97, Col95, CMSvE99, The98, Dre98, FM97, GFR98, Gre98d, HCF97a, Hamxxa, H^{+}xxa, Hig98, Hob97c, Hor96a, Hor96b, Jep97, Joc98, KP99, Lei95, Les97, Lev96a, Lip99, LB97b, McC98b, McD99b, Minxxa, Moh96, Mor97f, Morxxe, Oga99, PM96b, PM97b, PM97a, Pat98b, Ree97, Ric95a, Rin97a, Sip98, Sul97, Thaxx, Tho99, Van97g, VPV99, WP95, Ano97-74, Ano97-146, Asl98, Ber97d, BLT98, BM98c, Dua96, Dun97, Gre98b, Gre98c, Gre98f, KT96, Ste97a, Tay97, Tay99, WM97b, Yan96, YLQ98, Ano97g]. Database-to-Web [Col95]. Databases [Ano96-28, Ano96e, Ano96-96, Ano97-30, AOV^{+}99, Gil97a, Key98, PM97b, ZMN98, Atw96, CSH^{+}98, Gre98f, PZ96, Ric96a, ZMN99]. DataBeam [Ano97-29]. DataCad [MD99]. DataCasting [Key98]. Datacomm [Ano96-135]. Dataflow [HT98]. DATAMATION [McC96c]. Datapro [Hes97]. Datawebs [Var96]. Date [BCP^{+}99b]. Datenbank [Dic97b, Dicxx, Pat98a]. Datenbank-Workshop [Pat98a]. Datenbankanwendungen [Klu97, Klu98, Kluxx]. Datenbankentwicklung [Til97a]. Datenbankprogrammierung [Hob97a, Hobxx]. Datenbankzugriff [HCF97b, Hamxxb, Her97]. Datos [FB98, Hob98a, Lesxx]. Dave [Lut99a, Lut99b]. DAVIC [CSY97]. David [Ano96o, Ano96p, Nic96a, Nov98, SMC97]. Day [Dec95a, Lie95b, Mic95, Lie95a]. Days [AM97, Ano96-185, Coh97e, The98, Doh97, FM97, Hob97c, LP96e, LPW96, LPM96, LPC96, LP97d, Lem98, LC98, LPxxb, Man97a, Man97b, Mor96b, Ros98, TA97, Tuc98a, W^{+}96, WPL97, Dow97, JLP96, KPF96, LC99]. DB [Gre98d]. dBase [Ric95b]. dbx [NE98a, NE98b]. DC [Cim96]. DCE [Lin96c]. DCOM [PS98]. dead [Nat96]. Deadlock [Tei98, Ver97, DIS99]. Deadlocking [KL97]. Deadlocks [SL97b]. Deal [Gut95k, Stu98, Ten95b, Com95e, Woo99a]. dealing [O’C95c]. Deals [Par95d, O’C95a]. Debate [Mil95d]. Debbuger [WCS99]. Debian [Ano96-135, Kuk98]. debts [Ano96-151]. Debug [Ano96-79]. Debugger [Sun99a]. Debugging [Ano98-32, Ano98-33, Kap97b, NE98a, NE98b, BS96, GHP^{+}96, Rit97d]. Debut [Gar98b]. débuter [VB96]. Debuts [Ano97-113, Mas95, Mes97]. DEC [Ano97-107]. Decaf [Ano96h, Ano96i]. December [IEE96h, IEE96l, IEE96o, MJ98, USE97c]. decent [Gre97k]. Decimal [JJBW98]. Decision [Sim96f]. declares [SG97a]. Decompilation [PW97]. Decompiler [Kro97, San98b, San98c, van96c]. Decompiling [Nol98]. decomposition [GRN99]. Decubierto [Ano97-68, Sam97]. dedicated [GLV99]. Deducing [BM98a]. deduction [LMdS98]. deductive [KT96]. Deep [Chaxxa, Proxx]. Deepens [Sim97a, Sim97b]. Default [Sch99c]. defeasible [KPG96]. defense [Ano96-97]. Defensive [Coh97b]. Defining [Ano97-30, BS98c, BS98d]. definitely [NvO98]. Definition [BS99, Gro96b, Bok99]. definitions [Hay96w]. Definitive [Fla96d, Fla96e, Fla97g, Fla97h, Fla98e, Nov98, PPJ96, SLO96, Zuk99, Ano96p, Ano98g]. degree [Mat96b]. Del [Anoxx-45, LJxxb, LJ97, LJxxc, Sal97]. delay [CKH99]. Delayed [Gar98b]. delegate} [Wil99b]. Delft [GV97a, GV97b]. Delft-Java [GV97a, GV97b]. Deliver [Ano95u, Bor95, LaM95f, Lot95, Ude97d].

Delivers

[Hor96a, Hor96b, MD99, Web97a, Gre97c]. Delivery [Kro98a]. Delphi [Ano95h, KSB97]. Delphi-like [Ano95h]. Deluxe [Ano96-103, F^{+}97b]. Demand [Hor96a, Hor96b, Wil96a, NJC96, Wip96]. demands [Ano97-150]. Demos [App95]. Demystified [CL97c]. Denotational [AFL99]. Deploying [BN97, LO98b, Gre97f, LO99c]. deployment [Ano97w]. Derby [The98]. derivatives [FD96c]. Desarrollo [Jawxxb]. Describing [DE99]. Description [Sta96, Fri96]. Desenvolvendo [HI96a, HIxxa]. Design [ACM99b, Ano97-118, BKCG98, Chaxxa, CSxx, CM97b, CM97c, CM97d, CM99, Coc98, Coo97b, DG96, Div98, Gar97c, GK97a, GM98c, Gra99a, JWS98, KHB99a, KHB99b, Kne96, Kro98e, Lea96a, Lea97a, LH97, Lut99e, MD99, PL96, PJR^{+}98, Ran97b, Ran99b, Rem97, R^{+}xxb, RRxx, RL98, Shixxb, Stu98, Sun99b, TA97, Tea97, Thaxx, TMM^{+}98, TMMH99, USE96a, Ver96, Yanxx, Zei97, CH98a, CM97a, D’S96b, Dys96, Gep98, Gon99, GSvD99, Gre98a, Had99, Lak96, LL98, MC97, O’C97, PSTZ99, RNBV98, Sch97h, SML96, SvG98, SW96, Ver97, VS96, WW98, Ano97-146, Ste97c]. design-to-fabrication [SW96]. designed [KCPC98, Sha98b]. Designer [Ano96j, Ano97-45, Ano97-168]. designers [Ano96-53, LS99]. Designing [BN97, Hei97, O’N98a, RP98, Whi96b, Whi96c, Arm99, RMB^{+}98, Rob99]. Designs [Cro95a, Ano96-145, Ano97-100]. Desk [Sin95b, Sin95a, PSTZ99]. Desktop [Ano96-28, Ano97-121, Fla95, FFCM99, Fla99c, Gil98c, HN96, MD99, Ano96o, Cha96a, Fla97d, Fla97e, Lea98b, PF97]. Desktops [Web97a]. Despite [Ano96-151]. Destabilizing [Ano97-108]. Detail [CH97f]. Detailed [DHM99b]. Details [Lut98a]. detection [DIS99]. determine [Ven96]. Deugo [Ano98i]. Deutsche [Fla98a, Fla98b, Fla98c, LJ96a]. deux [Eck97c]. dev [Ano96-124, Lin96e]. Develop [CRS98a, Mac95, Sch97c, TJ96, CRS98b, CWS98, Gau96a, Pog96, SW99, Nan95]. developed [Ano96-74]. Developer [Ait96c, And98a, Ano96-135, Ano97-160, Anoxxi, Anoxx-63, Bro97a, Broxxa, Broxxb, Cha98c, Cha99c, Coh96b, CMMN96, CC96, Wal97d, DT96, Dan96a, DiG95b, DiG95c, DiG96, Feg97, Fegxx, FF96a, Har96b, Har96c, HRSM97, HR98, HR99, Hor98b, Hus97, Jaw96, JSA98, Kas97a, Lei97, Lin96f, MS99a, McC98b, Mic97b, Min97b, Minxxc, MS97, Nas95c, Nic97, Pur97, SBH97, Sri97b, JAV96b, Tho99, Wal97b, Weh98b, WB98, Woo99b, You97a, You98a, Ano96-73, Ano97-148, HW97, HL97b, HL97c, HL97d, Mar97a, PM97d, Red97a, Tay97, Tay99, Wil96e, Ano97r, DiG95d]. Developers [Ano96-134, Ano97-51, Ano97-161, BF98, BF99a, Boo97, CBT98e, CBT98d, Gau97a, Gol96b, Hay96b, Hay96c, Hay96d, Jaw97c, Joc98, LaM95b, Mes96, Mey96a, Mey96b, Web97c, Web97d, Ano96-125, Ano96-191, Ano97-92, Gau97g, Gre97b, LS99, Ner97, Tic97, Tre98, Way98a]. Developing [Abl96, AW99b, AJ99, Bin97, BDP^{+}98a, BDP^{+}98b, CT98, Chu98a, Eng97c, Engxx, GM99, Goo99, HI96b, HIxxb, Kum98, LB97a, LBS98, Man96b, MBH97, McF97, McF98a, McF98b, Mit97b, PM97b, Pet97, SL99, SP97, Tra97, Wil97a, WR98b, WRxx, Wit97, Won98, YLQ98, Yeexxb, You96a, Amb97, Ciz97, EHS^{+}99, HK99, MTU99, Miz99, Val99, Whi96b, Ano97h]. Development [Ano95u, Ano96v, Ano96-40, Ano96-79, Ano96-197, Ano97-121, Ano98c, Ano98l, Ano98-29, Ano98-31, Ano98-32, Ano98-33, Anoxx-64, Bar99, BM98b, Ber98e, Dar99b, DBBE98, Det97, Fis98a, Fox99, GPF97, GMPS97, Gre97j, Gre98e, Ham97b, HNSS99, Hay95d, Hor97a, IEE96j, Infxxa, JDK98a, JDK98b, KSB97, KGTF97, Kro97, Kro98d, Kro98e, Kro98c, Kro98b, Kro99a, Kro99b, Kro99d, Kro99e, Kro99c, Lad97, Lea99, Man96c, MNO^{+}96, Mes95d, Per96b, Pit97, Rem97, Rin97a, Jav96c, Sut97, TJ96, Uni96b, Uni97b, Vac96, Van96f, Wal97c, Way98c, Wil97a, You96b, Ano95h, Ano96z, Ano96-97, Ano97r, Ano97-47, Ano97-46, Ano97-144, Ano97-157, Ano97-158, AA96a, BN98b, Bla96, Boo96a, BLT98, Buc97, DR98, DR99, Dic96, Dra97, Fie96, FI96a, FC96a, Gol96a, Gre96b, Gre97h, Ham96a, Hem96a, Hor96c, Joh96a, Klu98, LH96]. development [Lim95, Lin96b, Lin98b, McL96, Met96, Mey98c, Red98, RNBV98, Sem96, Sym95, Uni96a, Bee96, Kro99a, Kro99b]. Developments [Ano96-29, Ran97a]. développement [Salxx]. Développment [F^{+}xxc]. Device [Law99, WP95, Ano96-196]. Devices [BAC^{+}99, Won99, Gam96, YRE98]. Devilish [Gib95]. Dewey [JJBW98]. d’Experts [Morxxc, LO99a]. DGB [Anoxxj]. DHTML [Gil99]. Día [Fra97]. Diab [Kro98b, Kro99d]. Dial [Gil97c]. Dialect [Nil98b, YSP^{+}98a, YSP^{+}98b]. Dialog [Wel99b]. Dias [Hob98a, Lem97c, LP98c, Lemxxa, WL97c, W^{+}xxa, Lem97b]. Didattica [AG96c]. Diego [ACM98b, Ano98a, IEE96h, Ten99, USE96d, USE98d, USE99c, Wae96, W^{+}98]. Difference [Ano96-46, Nap97]. differences [Pre99]. Different [Gil98c, LaM95c, GLC99, Hay96v]. Differential [Loe97]. Digest [IEE96c]. Digital [Gar98b, KSI^{+}96, LR97, MD99, MM99, PSTZ99, PCWGM98, Tab96, Bry98]. Dijon [YH96]. Dimension [Ano96-30, Ano96-135]. dimensions [vD99a]. d’Internet [BKP97a]. Direct [MO98, Ano97-72, Wil98b]. directions [Ano95r]. directly [Ano96-97]. Director [Ano95k, MD99]. Directorio [T^{+}97b]. Directory [Gil97c, TSHH96, Ano97-102]. Dirty [Hay97a]. disabilities [Ano97-145]. Disabled [Laz97, Laz99]. disapproval [SG97b]. Disasters [Wil97e, Wil97f]. DISCIPLE [Mar99b]. Discounts [DDJ98b, DDJ98c]. Discover [Afe97, MBxx, PM97b, Per97d, TB97, WG97b, Gre97k]. Discovering [Eck97b, Eck97c]. Discovery [Gre97k]. Discrete [BSS98a, Cow98a, vHO98, BSS98b, Cow98b, VO98]. Discrete-Event [BSS98a, Cow98a, BSS98b, Cow98b]. Discrimination [BCP^{+}99b]. discussion [Ano96-151, Rob96]. Disease [WBA^{+}96]. Disk [Gar98b]. diskette [Ano97k]. Disks [Gar97a]. disparate [CL97d]. Display [Ano97-33, BA96, Ber96c]. Dispute [Gar98a]. Dissecting [Lut98b]. Distance [BBJR96, DDJ98b, DDJ98c]. Distributable [Ano97-30]. Distribute [Ano97-34, Sco96a]. Distributed [AR98, AFH98a, Als95c, Als95a, Als95b, Ano97-76, Lut99a, Anoxxk, Bak97a, Bakxx, Bal97, BA97, BFD96, BN96, IG99, CBT97d, CV98, CCF99, CRS98a, Cha98d, Cha98e, Desxx, DAB97, Dua96, Far98a, FFKO99, Gol99, GS98c, Ham97b, HJZ^{+}98, Hir97, HYI98a, Hiu98, HS98b, HLRT97, IEE96n, IEE96o, KS98a, KS98b, Kar98a, LP98b, Law98, Lut99b, Mah98, MCD99a, MZ97, Mor98a, Mor97a, MR97, PSW^{+}98, Rod96b, Rou98, Som97, SHGD98a, UCD99, VR99b, WW97, Wat97a, WRW96b, WWR97a, WWR97b, WRW96c, WAES97a, WAES97b, WCS99, YH96, Yos97, vHO98, AFH98b, Ano96-54, Ano96-60, Ano96-192, Ano96-205, AA98, BAS99, Ber99b, Ber99c, BBC^{+}97, CRS^{+}96, CRS98b, CC97, CMM^{+}97, CKTV96, Cli98a, ERKE96, Far96a, FF96b, GS97c, GS98d, HVB^{+}99, HYI98b, JA98, JCS98b, Kar98b, LM98, LH96, LCLL97, MHO96]. distributed [PS98, Pog96, RMB^{+}98, SHGD98b, SLPM96, Val99, VO98, Vck96, Vog96, WRW96a, AJ98a, JCS98a, TBA98]. distributed-object [RMB^{+}98]. Distribution [Spi99, Zha97, Ano96-77, pP98]. Dive [Wil96h]. Divide [Kin97a]. dividend [Ano96-195]. division [DiD97b]. DLLs [San98b, San98c]. DMI [Ano96-28]. DML [MD99]. DMS [Ano97-114]. DNA [Gau97h]. Do [AW99a, Bab97a, BA96, Ber96g, Ber96f, Ber96d, Ber96e, Ber97l, Ber97i, Ber97h, Ber97m, Ber97g, Ber97f, Ber97k, Ber97j, Ber98a, Ber98d, Cri98, DDJ98b, DDJ98c, Gof97, GS98a, Gri99, Moo98, Pit98a, Ses98, Wil99b, Ano97-30, Ber96a, Ber96b, Ber96c, Ber96h, Ber97a, Ber97d, Ber97b, Ber97e, Ber98b, Ber98c, LJxxd, LJxxa, dN97]. Dobb [Eri96a, Kim97b]. Doc [MD99, Ait96a]. Doc-to-Net [MD99]. Docs [Ano97-33]. document [Lu96]. Documentation [Ait96a]. Documenting [FHKS98]. Documents [Fei98, TMM^{+}98, TMMH99, Vad95b, BDKK96, Lu96]. dodging [Ude97d]. DOE [Boz95b, Inf95d]. doende [Holxxb, Holxxc]. Does [Eas99, Hay96e, MD99, MGL^{+}98, PC97, PW97, Pur98, Ran99b, Ano96-42, Ano96-133, Ano97r, Ber96c, Eti97, Rie97a]. dog [Ben98]. DOGMA [JCS98a, JCS98b]. Doing [Lut98a]. Dokumentation [GJS97]. Domain [Gar97a, Gar98b, USE99a, GRN99, LQ99]. Domain-Specific [USE99a]. Domains [HCC^{+}98, SKN98]. dominance [GH96a]. Dominando [Hol98b, Nauxxa]. Domine [Jep96]. Domino [H^{+}97, Bal98a, Cav97, Oue96c]. Domponents [Chu98a]. done [Ano97-30]. Données [Carxx, Staxxa]. Donnelley [Ano96-77]. Don’t [Cou96a, DDJ99b]. Doorbells [Law98]. DOOS [Cha98d, Cha98e]. d’Or [Lonxxa, Marxxb]. doru [IK96]. Dose [Ano96c]. Doubles [Ano95n]. DoubleTree [Nat98]. DOVE [AJ98a, JA98]. Down [GC97, Hay95c, MF98, MF99, Ols97, Pet95, van98a, Gil96b, Wag97a, Way96c]. Downing [Ano97k]. Downloadable [IGO98]. Downloaded [JPLI99, McC97b, MF97a].

Downloading [JRE98]. Downsizing [Ano96-28]. dozen [vSS96e]. DPLab [BAS99]. Dr [Eri96a]. Dr. [Kim97b]. Draft [Num98, Hay97d]. Drag [Ber96e, Pur98, Ber96a, Hay96y, Tur97]. Drag-and-Drop [Ber96e, Pur98, Ber96a]. DRAM [Ano96-153]. Draw [Bro95b]. drawer [Ano97-156]. Drawing [BH99a]. Draws [Gar97d]. Dreams [Swa98b]. Dreamweaver [MD99]. Dried [Wal96j]. Drive [Ano96-31, Ano96-40, The98, GC97, OH97a]. Driven [KSI^{+}96, Lut99c, MZ97]. Driver [Ano98-27, Ano97s]. Drivers [Soo98b, YLQ98]. Drives [DDJ98a]. driving [Ano95r]. Drop [Ber96e, Pur98, Ber96a, Tur97]. Drug [Stu98, Yat97]. DSL [The98, USE99a]. DSM [YC97]. DSP [Kro98d, Kro98c, RK97b]. dual [Ban98]. Dubbelboek [Anoxxl]. Due [Inf95c, Ano96-196, Gau96c]. Dukane [MD99]. Dummies [Anoxxg, Bur97, Coo97a, CA97, HB98, Hym97, KK96b, KK97d, KK97e, KK98c, KKxxc, LS97b, Mos96, Sha97a, Ste96d, Ste96e, Van96e, Van97j, Van97i, Van97h, Wal96l, Wal96o, Wal97h, Wal98c, Wal98d, Tit97, Van98b, Wal96n, Walxxf, Walxxg, KK97f, KWxx, Van97k, Wal97g, Walxxe]. Dundee [SMEN96]. duty [Sli97]. dv [MD99]. dv/MockUp [MD99]. DVD [GC97]. Dwarf [Ano96-112]. Dwight [Ano98i]. Dynamic [AFL99, Bot98a, JL96, KP99, LB98b, P^{+}98b, Rou98, S^{+}96a, TPHB96, Atk99, Fet97, Hei97, LTC96, Way96b, Kro98d, Kro98c]. Dynamically [Ber97e, FC98b]. dynamically-linked [FC98b]. Dynamics [CIF98, AGD97].

e-commerce [VM96, Ham97b, Gar97c]. E-Lab [Web98]. e-mail [Ber96f, Ano96-32, Gar97c]. e-mails [Bra96].

Eager [Oue95b]. early [Ast98, GJ97, Led97, PSS98]. Earnings [Ano95n]. Ease [Lea99, Hay96l]. eases [Ano97-74, LH96]. Easier [Hay95c, Hay97b, Zol96]. Easiest [Jam96]. Easing [Win95i]. East [Way98a]. Easy [Guaxx, MD99, Shixxc, Wat95e, Way98a, Cobxxb, Fet97, Pir97, Wil96a, Wil98b]. Eat [Ben98]. Eclipsing [Win95m]. ECMAScript [Dob98]. Economic [Ano98l, Wal96i]. Economics [Lut98b]. Ed [DDJ98b, DDJ98c, Jam97a, Jamxxa, MGL^{+}98, Mey97, New96d, Cad98b, Cor95]. Edge [Ano96-28, Ano97-28, BSS96, Esp96, Mur99a, Wil97a, Ban98, Kro98e]. Edicion [Cul98a, Lesxx]. Edinburgh [AOV^{+}99]. Edited [Ano98i]. Edith [Ano97-119]. Editing [Ano97-35, Kro99d, MD99]. Edition [Add99, Ano96p, Ano96-103, Anoxx-71, BM98b, CLK99b, CWHP96, DH98, Doh98, EMe97e, EMe97d, EMe97f, Fla96e, F^{+}97b, Fla97h, JM97d, Kro97, LO96, LOA^{+}96, Lad96, LO97a, LO^{+}99b, LPM96, M^{+}97d, N^{+}97a, Newxxa, Nor96, PS97, Que97, Rei97a, Rei97b, RWxx, Rom99, San97b, S^{+}96a, SA97, Sha98a, Sta97a, TJ96, Wal97c, WB^{+}96a, WBC^{+}97, Webxx, W^{+}97b, Wil97b, van97c, Ano98f, Ano98g, Ano98k, BLT98, DH97, LC99, N^{+}96b, WRTM98, Ano97-45, New98, Nov98, San98a]. Editions [Bro96]. Editor [Ano96-137, Bek97, BL97, CL96, Cra97, Fed97, Hen97b, KL97, McC97c, MF97d, Mon96, Nav97, Pla96b, Pla96a, Rag97, Raj96, Ste99, Urq97b, Wel97, WN97, WL97a, Wil96h, WL97d, Sto97, Agh97, MH97a, MH97b, Pla97]. Editor-in-Chief [Agh97]. Editorial [Cul98b, Fox97a, Fox98a, Lab95, Sch97f, Sch97e]. Editors [Kin99, SN96]. Edits [MD99]. Eds [BCP^{+}99b]. Education [Dal98, EBR96a, Isk96, Joy98, Ros96, BV98, BCF^{+}97, BGGR98, CH98a, CAZZ98, EBR96b, Owe96, SBB^{+}97, Sco96c, SZTR96, SPV99, WCL97, NTS^{+}96]. Educational [Ano98a, CAB^{+}98]. Educator [JOA98]. Educator/Researcher [JOA98]. educators [Gor96]. effect [Cla97d]. Effective [Coo97b, Lak96, TA97, ATCL^{+}98, Woo99b]. Efficace [Knaxxa, KG97a]. Efficage [Anoxxa]. Efficiency [ZMN98, ZMN99, Pre99]. Efficient [CIF98, CGHS99, Coo97b, HVB^{+}99, KWB^{+}98, LM99, McC98b, MMBC97, KG96a, KG96b, MvNV^{+}99, NAC99]. efficiently [KP98a, KP98b]. effizienten [Pie97]. effort [BH96c]. EG1A [Ano97-118]. Eiffel [Sch97h]. EiffelBase [Kro98d, Kro98c]. eigene [KS96b]. Eighth [IEE96i]. Einbinden [KS96b]. Einblicke [Prexx]. einem [Bat96d]. Einfuehrung [Koc97]. Einführung [BKP97b, Buhxx, DE96b, Min96c, Red96, Redxx, SST98, TG96a, Dec96b, Decxxb, Simxx, TGxxb, Wol98]. Einführungskurs [Kne98]. Einsatz [Piexx]. Einsatzmöglichkeiten [Mül96]. Einsteigerseminar [KSW97]. Einstieg [Sch97l, MG96]. Einzusetzen [Frö96, Fröxxb]. EJB [Cha99f]. EJO [Ano96-148]. Electro [IEE96d]. Electrochemical [DDJ99a]. electromagnetics [MLM98]. Electronic [Bus95, Dun99, FSW98, Gol97, Jar97, TW96, USE96b, USE98a, Wae96, Ano95s, LM98, Wil96c]. Electronics [Ano96-148, Kap99]. Electrotechnical [IEE98c]. Elektronik [Web98]. Elektronik-Labor [Web98]. elevating [Oua96]. Eleventh [Uni97b]. elicits [SG97b]. Eliminate [Moo95]. Elliotte [Ano97j]. Elliptic [Loe97]. Ellison [Fer95]. else [Gre97f]. Elumaze [Ano98-27]. Email [Ano96-163, Sim96e]. Embed [LQ99, Ano97-143]. Embedded [Ano96-171, Ano97-36, Ano97-137, Bar99, DDJ98a, Hew99, Ham97c, Kro98b, Kro99a, Kro99b, Kro99d, Kro99e, Mon97, USE99e, Ano96-100, Hol99, LK99, Lin98b, Wil96f, Wil96e]. Embedded-Centric [Kro99e]. Embedding [Skl95]. Embrace [KL97]. Embraces [LaM95e]. Emerge [Gar97b, RFM95]. Emerging [Ano96-33, AA96a, IEE96r, Urq97a]. empire [Wal96a]. Empirical [RR99]. employers [Pfe99]. empresa [Ben96b]. Empress [Ano98-27, Ano97-36]. Emulates [LaM95d]. Emulation [Ano97-37]. Emulator [San98b, San98c]. enable [Ano96-128, Gre98b]. Enabled [BD99, Won99, Col96a, MZ97]. enables [Cli98a]. Enabling [Ano96-34, Ano96-101, FT96, IEE96m, Sun97a, Sunxxc, BCF^{+}97, HC99a]. Encrypt [Ber98c]. Encrypted [AW99a]. Encryption [Ano97-38, Ano97-133, DDJ98b, DDJ98c, Gar97a, Gar97c, GC97, Gar98b, Law98, MGL^{+}98, Nat98, RD99, Ano97-132]. End [Dan98, IR99, Leo95, Lut98b, SHGD98a, Spi99, TBS97, Jav98, Mee96, SHGD98b]. End-User [IR99]. endorse [Dei96a]. Ends [Ano97-77]. engaged [Sch97f]. Engine [GV97a, GV97b, Kie99, MO98, Ang99, Ano97-43, Ano98p, YHW98]. Engineering [Ano96-59, Ano96-135, IEE96p, IEE96k, Isk96, Lev99, Mor97b, Sch97a, AC99, Com97a, Com99, Dei96b, DAB97, Dug99, Eti97, Fox97a, Fox97b, FF97, GS97b, GG99, Hin98, DF96, SZTR96]. Engineers [Cha99d, Dav99, Gar99, JG96, Ano96-59, Buc97]. Englander [Ano97h]. English [Beh95, Biexxb, Dre96, Ove96, Ove97]. Engst [Ano96-42]. Enhance [Ano96-148, Don96, Vad95a, Ano97-126]. enhanced [dRPT98]. Enhancements [MD99]. Enhances [Ano97-36, Moh98]. Enhancing [TPW96]. Enjoying [Lin96a]. Enough [Mac97d, Mac97e, New96a, Agn97, Mac97a]. Ensure [GS98a]. Enterprise [Ano96-118, Ano98-36, Anoxxe, App95, AW99b, BM98b, Ber98e, CCH97, Fis98a, JT97, JTxx, Kro98e, McC98a, Mit97b, Moo95, Net95a, RA97, Rom99, SWJ97, SA97, SAxx, SF96c, Sun97a, Sunxxc, Val99, WCF98, Hay96x, Lin98b, Mey99, Ano98c, BFGS98, CDF^{+}98, Dun97, FFCM99, Kro97, MH99, VR99b, Ano99a]. Enterprise-aware [App95]. Enterprises [IEE96m]. Enterprising [Ano98z]. Enters [Gar98a, Win95c]. Entertainment [Wit97]. enticing [Pro96]. entirely [FC96a]. Entrus [Ham97c]. entwickeln [Gulxx]. Entwickler [BBB^{+}96]. Entwicklung [Pie97, van97a]. Entwicklungssoftware [Klu98]. Entwicklungsumgebung [W^{+}97c]. Entwurfsprinzipien [Lea97b]. enum [Bri98]. Enumerated [Cai99]. enums [Eas99]. ENVI [Ano98-27]. Environment [AFH98a, AVN98, Ano96v, BRF98, CZFL98, DBBE98, GM99, Gol99, GM96, Ima97, JRE98, Kro99a, Kro99b, Kro99c, MMBC97, Per96b, AFH98b, Ano96z, Ano97-136, Ano97-144, Ano97-138, BAS99, Boy98c, DF97, DAB97, Dua96, GBS98, Gre97c, JA98, LG97, Lim95, Lin96b, LCLL97, MZ97, RBF98, Sym95, AJ98a, EM96, Tun98]. Environments [SWJ97, SN96, Uma97, Dic96, GGMS99, JAS96a, Mar99b, PS96, SZTR96]. envy [Woo99a]. Epic [Way97b]. Epidemic [Why96]. Epson [Ano96-196]. Equations [Loe97]. equivalent [Ude97d]. Equivalents [Mic97a]. ERA [Kro98e]. Erfolgreiche [Sch97l]. Eric [Par95a, Win95b]. Erica [Tee99]. Error [BN99a, Kro99c, MYD95, Mat96b, Orc98]. error-prone [Mat96b]. Erste [Gol96b, Golxx]. Erste-Hilfe-Kasten [Gol96b, Golxx]. erstellen [DE96b, vSS96d]. Ertweitungen [MS98b]. erweckt [Dre96]. erweitern [Bat96d]. ESA [Bue98]. eSafe [Mey98a]. Escape [Bla99, CGS^{+}99, WR99b]. Esempi [JM97b, JMxxb]. ESL [HNS97a]. eso [Kim97d]. Especial [Cul98a, Lesxx]. especially [GS97b]. Espresso [Hay95a, Hay95c, LaM95f]. esprit [BFGGxx]. Essence [Kin97c]. Essential [BLB99, Cow97, Cow99, GM98d, GT99, Heu97, Lan99b, Man96b, Min97c, Rob99]. Essentials [Boo96c, Hor98a, Hun98, JSA98, Man96d, PM97b, Per96b, Cul98c]. Estimation [Kro98b]. eSuite [Gil98c]. etaCOM [Ols96]. Ether [Met95a, Met95b, Met95c]. Ethernet [Lut98a]. EtherPage [Kuk98]. EUROMED [VPM98a, VPM98b]. EUROMED-JAVA [VPM98a, VPM98b]. Europe [Ano96-142, Ano97-139, Str96, Uni97a]. European [Ano97w, Bod96, DLM99]. EuroPVM [Bod96]. Evaluating [DDJ99b, BN99b]. Evaluation [BC98b, GLC99, HYI98a, Hos96b, RL98, BC99b, Had98b, Han96, HYI98b, Hos96a, KSW98, Ste96c]. Even [Ano96-84, The98, Ano96-171, KW96, Pro96]. Evening [WBC^{+}96]. Event [Ber96h, BSS98a, CKRW99, Cow98a, Gig97, vHO98, Ber96b, BSS98b, CKRW97a, CKRW97b, Cow98b, Lie95a, VO98]. Event-Based [CKRW99, CKRW97a, CKRW97b]. Events [CRS98a, CRS98b]. ever [Gil98b, NMS96, Wil96a]. Every [SLC99]. Everyone [Bri99c, KoCaB98, Kah98, KD98]. Everything [BSS96, Mee96, MRA98, M^{+}96b, M^{+}97e, M^{+}97c, Ano97r, GS97b, GHP^{+}96, NMS96, Way96c]. Everywhere [Ano96-28, Ano96-94, Ano96-95, KoCaB98, Kah98, KD98, Bab96, Sch96b]. Evolution [Gos97b, WS99, KNNL98, W^{+}98]. evolved [Ung97]. Evolves [Lin98b]. Evolving [Dar98b, Rey96, Sko96, Ano96-192, Dar99a, FF96b, Gon97]. eVote [Ano98-29]. Exactly [Ano96-141, CL97c]. exalted [Oua96]. Exam [Boo98, Bro99a, Bro99b, Jar98, Tuc98a]. examines [Ven96]. Examining [Ang99, Ano96-35, Gar96f, Hem96a, Soo98b, Wil98b, van97f]. Example [Anoxx-73, FCDW96, JM96, JM97a, JM97d, JM99, JMxxa, JT97, LH97, WR99a, Wal96k, WR98a, Wes97a, Wil99a, Ano96-60, SS98c, Bay97a]. Examples [Fla97a, Fla98a, Flaxxa, Por98, MAS96]. Excel [Ano96-133]. Excellence [Eri96a]. Excellent [Got98a, Gre97c]. Exception [Pit99, RM99]. Exceptional [All99a, Ano98m]. Exceptions [Gra96a, KP98a, Orc98, CGHS99, KP98b]. Exchanging [Hav98]. Exclusion [For98a]. exclusive [McC96c]. Executable [IGO98, JPLI99, RK97a, SH96, SJRK97]. Execute [YRE98, MF97a]. Executing [AW99a]. Execution [Hir97, MO98, MS99b, YK99, KCLZ98, MW98]. Executive [Wil97e, Wil97f]. Executor [Ano96-135]. Exemplos [Walxxb]. Existing [Vad95b]. expanded [Ano98k]. Expanding [MS98a]. Expands [Nas95b, Ano98-35]. Expedition [NP96b]. Experience [BVG98, Hol97e, Hol97i, KFS97, MHO96, Zuk97a, BP97, BPxx, Ciz97, Hol97d, TLSS99]. experienced [Mar97a]. Experiences [GFR98, JS97, McF97, Wei97c]. experiment [Led97]. Experiments [BDV^{+}98a, CCF^{+}97, BDV^{+}98b]. Expert [Eri96b, Han97, Jar97, Sch96a]. Experten [CH97f]. Expertenwissen [HC98c]. Explained [Bis97]. explanations [Kay97]. Exploiting [GFHM98, BG97b, Cim96, GHM98]. exploration [RR99, Ude97b]. exploratory [Ste96f]. Explorer [Ano96l, D^{+}98b, SBL96, TD97, Thexx, Ano96-43, BLT98]. Exploring [DW96a, Far96b, Nie95, NP96a, NP97b, NP97c, NP97a, NPxx, Shu97b, GPV97, SMC97, Ano96n, Ano98f]. Expo [Ano96-142, Kno95, Ano96-143]. Export [DDJ98b, DDJ98c, Gar97c, GC97]. EXPRESS [Sha97b]. Expressions [Ber99a, BCP^{+}99b, LS99]. Extend [Ano96-177, Sim97e, Bat96c, Mat96b]. Extended [Gri97b, GME99]. extended-precision [GME99]. Extending [But99, Hus96, MS98a, Fuc96]. Extensibility [SKN98]. Extensible [IR97, SHW98a, WBDF97, SHW98b]. Extension [DK98, DBK98, Par99, SMY99, WS99, vD99a]. Extensions [CZF^{+}98b, CMSvE99, Gos98a, Knu99a, Law98, S^{+}96a, SC96b, ZJ96, CZF^{+}98a, Eck96c, Her96b]. Extensive [SR97]. extract [NN97]. Extraction [Doe95]. extractor [TLSS99]. Extranet [Ano98-31]. Eye [Rei96]. Eyes [Mor97c, Ten95b, Uim95, GH96b, Ham97a].

F [Ano97-119]. F-Secure [Ano97-119]. Fabrication [DDJ99a, SW96]. FACADE [GLV99]. face [Ano96-196]. Fácil [Walxxb, Wei97b]. facile [Maixx]. Facility [Chw98]. fact [Wal96h]. Facto [Har96f]. Factor [RA97, SN98a, SN98b]. Factors [Tau96]. factory [Ath98, Kro99a, Kro99b]. Fad [Fra98b]. fails [Ban95a]. Failure [BN99a]. FairCom [Ano96-137]. Fall [Cim96]. falls [Sul95a]. family [Ano96-180, Ano97-151, Ano97-152]. Famous [MD99]. FAQ [Kan96a, Kanxxa, Kanxxb, Kanxxc]. FAQs [Ber98f]. Far [EMN98, Wil99d]. Fast [ATCL^{+}98, Cow97, Cow99, DDJ98b, DDJ98c, Gar97c, Hun98, McC98b, Red99, Syd96b, Heu97]. faster [Gau96j, Gau97p]. Fat [MC97]. Father [DDJ99a, Spe97]. Fault [Ano98m, Sar99]. Fault-Tolerant [Sar99]. favorite [Bös98]. FCC [Gar97c]. FCRC [ACM96]. FDBC [Sip98]. Fe [USE98b]. Feasibility [JPK98a, Mil95d, JPK98b]. Featherweight [BKMS98, IPW99]. Feature [Lei95]. Features [Bro97b, Bro98d, CBT98a, HC98b, HC99c, MS98a, Ano96-43, Ano97-92, KW99, Mat96b, Sim96f, Tur97, vSS96e]. February [Ano96-147, Cul97c, IEE96c, Joy98, Spe98, Ano96-196, Rob96]. Federated [Gre98d]. Feedback [Hen97b]. Feel [Gos97a, Gre97f, Lut99d, Sun99b, Gre97b]. feet [Hay96y]. Felleisen [Ano98e]. Fever [Car96b]. few [Ano98e, Col96c, FF98]. FFTs [GTV98a, GTV98b]. Fibel [Küh96, Küh97, Kühxx]. FIE’96 [Isk96]. Field [Ano96-36, Ano96-37, CDL99, DS97b, GDS98, HB95]. Fields [MB98b]. fifteen [Han96]. fifteen-puzzle [Han96]. Fifth [CI96, IEE96n, AOV^{+}99, USE99c]. Fight [Ano96-98]. File [AW99a, Kro98e, NN97, CKH99]. Files [HC98a, Pug99, Sch97i, Tuc98b, You98b, Kro98a, MLF97, NN97]. Filesystem [GS98c, GS98d]. Filling [JGKR96]. Filter [Ano98-27]. Filterfresh [BCH^{+}98]. Filters [NRG^{+}99]. finalizes [Ano97-147]. Financial [EM96, IEE96p, Sim96a, Ano96-54, Ano96-151]. Find [Pro96]. Finding [PSS98]. finds [Sli97]. Fine [LTC96, PH99a]. Fine-Grained [PH99a, LTC96]. finite [LQ99, RHND99]. Finland [JPT^{+}99]. Fire [Gar97d]. Firewall [Ano96-38, Kro98b, Hay96b, Mey96a]. Firing [Bur96a]. Firm [Bet95, Sim96a]. Firms [GC97, Gau96a, Hay96k, McK97, Pic96b, Ano96-54]. First [Ano95n, Ano95q, Ano96-28, Ano96-39, Ano96-40, Ano96-154, Ano96-160, Wil97e, Ano97-106, Ano97-130, Anoxxm, BDF96, CM96, Cul96a, Cul98d, Culxxa, GM97a, Hos96b, IEE96c, IEE96r, Kim97d, Law98, Nat98, Sav99, Wil97f, Ano96-165, Ano97-149, App95, Bös98, DiG95c, EHS^{+}99, Gau97n, Gos96, Had98b, Hay97d, Hos96a, KR99, Mee96, Pic96a, Sym95, WS98, RD99]. first-course [Bös98]. First-Year [Wil97e, Wil97f]. Fish [Als95c, DDJ99b]. Fissures [DDJ98b, DDJ98c]. fit [Soo98b]. Fits [Cou97b]. fitting [Rob98a]. Five [Ano97-124]. Fizz [Tub95]. FL [Chu98b]. Flanagan [Ano96o, Ano96p, Ano98g, Nic96a, Nov98, SMC97]. flank [Mey99]. flap [Ano96-179]. Flash [Com97b]. Flashcard [RC98]. Flat [MD99]. Flat-Panel [MD99]. Flatland [MD99]. flavor [Gau96a, Pic96b]. flavoring [Hay96h]. flavors [Ano96-174]. flaw [Sav97]. flexibility [Pic97]. Flexible [BRS99, IAJR97, JPLI99, MMBC97, Cli98a, pP98]. Flicker [Buc96a]. Flicker-free [Buc96a]. Floating [DK98, Dar98b, Gar97c, KoCaB98, Kah98, KD98, Dar98a, Dar99a, GME99]. Floating-Point [Gar97c, KoCaB98, Kah98, KD98, Dar99a]. flop [Woo99a]. FLORES [VPV99]. Florida [IEE96o]. flourishes [Ano96-151]. Flow [All99a, GO96, Mye99, RM99]. fluid [AGD97]. Fluke [Kro99c]. Fly [Sin95a]. flypaper [Chw98]. FMS [FFKO99]. Focus [Boz95c, Hay95d, Wal96i]. Focuses [Win95m]. foe [Ano97-28]. Folding [Lut99e]. follow [Lav96]. follow-on [Lav96]. FontScope [Ano98-27]. food [Ben98]. Footprint [Nil98b, MC97]. Forces [McK97, Ano97-108, Sil95]. forecast [Com97a, Com99, Hin98, Kap99, Lan98, Lan99a]. Forecasting [Ano98-32, Ano98-33]. Foreign [Gar97b]. Foresees [Kal95a]. Forge [Kuk98]. Forges [Kal95b]. Form [Coh97c]. Formal [AFF99, AF99, CR99, Die98, Fri96, JLMPH99, Qia99, FM99a, MKS^{+}96, MN96b, RBL^{+}97]. Formalizing [KG98a]. Format [BAC^{+}99, NN97]. Formation [Dav96a, Davxxa]. formats [Ber96c]. formerly [Cim96]. Forms [And96b, Ano96-81, Hor96a, Hor96b, JGKR96]. Fornax [VO98, vHO98]. Forrester [Ano97-73]. Fortgeschrittene [Hen97a]. Forth [BCP^{+}99b]. Fortran [Ano97-121, Ano97-153, Bra97a, Bra97b, DDS99, DP99, DD97i, FLQZ97, MKS^{+}96, MMG98]. Fortran-to-Java [FLQZ97]. Forty [IEE96c]. Forty-First [IEE96c]. Forum [Pla96b, Pla96a, Pla97, Coh97c, Fer95, Law99, Ros96]. found [Ano97q, Nic99]. Foundation [Bro98b, C^{+}xxc, Dav98a, DW99, Fla99b, Gut98, Jan97, JV98, Nel98, Top98, Top99, Tur97, You98a]. Foundations [Wal96m, Walxxc, LL98]. four [Soo98b].

Fourth [FW96, IEE96o, OW95, Her96b, USE98b]. Fox [Mag95]. FPGAs [MK98]. Fragment [OH97b]. Fragmentation [McC99d]. Frame [Tes96a, WH99, Sau96]. Frames [Tea97, Tes96a]. Framework [Anoxx-42, BD99, BS98e, CBT98f, CV98, Gol97, HHK98, Hou98, LaM95e, LP98a, Lin98a, Loe97, MD97b, OL98, SHW98a, Sem96, TMN^{+}98a, Win98, WAES97a, WAES97b, deC99, BKCG98, Ber99b, Ber99c, CMM^{+}97, DHM99a, DW97b, FM99a, GH97, Had98a, JS97, KPG96, Mar99b, Mey98c, SHW98b, Stu98, TMN^{+}98b, War99, DBK98, GT99, Kro98d, Kro98c]. frameworks [D’S96a, DW97a, Pol98]. Framing [Bat96a]. franca [Far96a, GG96a]. francais [BKP97a]. Francaise [Fla96c]. France [ACM97a, Ano97w, IEE96l, IEE96i, YH96]. Francisco [BUP99, CEG96, YN96, Pol98, RCB98]. Fraser [Doo99]. Fray [Gar97c, Gar98a, Law98, Win95c]. Free [Ano96-41, Ano96-146, Ano97-39, Ano97-41, Kro98d, Kro98c, Kro99d, Kro99c, O’B96b, Ric95c, Sri96, Tro97, Wal96i, Win95h, Buc96a, Gla97a, Gla97b]. Free-market [Wal96i]. FreeD [Hor96a, Hor96b]. FreeNIX [USE98c]. Freeze [Wal96j]. freight [Wil96b]. French [Anoxx-37]. Frequently [Kan96a, Kanxxc]. frets [Hay96a]. Friedman [Ano98e]. friend [Ano97-28]. Friendly [BS99, DeV95, Gau97d]. Friends [DDJ98b, DDJ98c]. Front [Ano97-77, Dan98, Gil97b, SHGD98a, Spi99, Joh97a, SHGD98b, Ano95w, Ano96-195, Ano96-196]. Front-End [SHGD98a, Spi99, SHGD98b]. front-page [Joh97a]. Frontiers [Isk96]. FrontPage [Ano97-154]. Frontrunner [KGTF97, A^{+}96, Aitxxb, BS97b, Dun96, FKPR96, Frixx, GPF97, GTFxx, Jarxx]. Fronts [Gar99]. frowns [Hay96a]. frustrating [Hay97d]. FSF [DDJ99b]. fueled [Ano96-82]. fuels [Dry96a]. Führer [Anoxxh]. Fuji [Ano96-196]. Fujitsu [Gut95f, Gut95e]. Full [TMP^{+}99, Ano96-54, Ano96-188, Ano98p, GM98b]. Full-JAVA [GM98b]. Full-Text [TMP^{+}99]. fun [Ano97r, IK96, Ram97]. Functional [Wak98b, MJ99, Wak99]. Functionality [Ano96-76, Ano96-130, MD99, JGKR96]. Functions [Hor96a, Hor96b, Dar98c]. Fundamental [GK97b, HCD^{+}98, TBxx]. Fundamentals [CMGR96, HC97, HC99b, Str99, TC95]. funktionsfahiger [Dav96b]. fused [GME99]. Fusion [Ano96f, Oly96]. Future [Bow96b, BOSW98, Fra98a, Gon97, Leo95, Liu96, Mes95d, Mit96b, Mit97a, Ric95b, Ste96a, Urq97b, Ano96-97, AA96a, BGGR98, HCD^{+}98, Mit96a, SFSS96, Ste96b, Tho98, Urg96a, Zol96]. Futures [Par95a]. FutureTense [Dys96]. Futurism [Tra95b]. Fuzzy [Lut99a, Cox96a, Lut99b, ZK98a, ZK98b, EKR97]. FX [Ano97-118].

ga [IK96]. Gage [Gut95c]. Gain [Gar96a]. Gains [Bou96b, Ric96b, WSL96]. Galaxy [Nan95]. Galorath [Kro98b]. galore [Ano96-43]. Gambles [Way96e]. Game [Ano96-83, BSS96, FRT96, Gut95e, HB98, Mor96b, Wil97c]. Gamelan [T^{+}97b, Ano97-125, TSHH96]. Games [BSS96, Wal97e]. Gantz [Ano96-102]. Gap [Kar98a, Kar98b]. Garbage [AR98, ADM98, JL96, KT99, PB98, Spe97, SLC99, Eck96a, McD98]. Garden [IEE96d]. gasoline [Ano96b]. Gates [Fer95, Gib95, Ten95a]. Gateway [Gol97]. GATEWAYS [Ano95a, Ano96-43, Ano96-42]. gather [Wol99]. Gcom [Ano98-28]. gear [Wil96b]. Gee [Tra95b]. Gee-whiz [Tra95b]. GeezerWare [The98]. Gegevensstructuren [Laf98b]. Geheimnis [Wie96, Wie97]. Gemplus [Coh97c]. Gems [Deu98, Due97a, Due97b, Due98, Ano98i].

Gen [Kro99d]. General [Bec98, Ber96h, Tun98, WC99, Ber96b, FF97, SPV99, Way96e, Gut95h, Win95d]. General-Purpose [Ber96h, Ber96b]. Generates [Ano97-128]. Generating [Ait96a, Fic96]. Generation [AC98, Doe95, Koh95, Kro98b, LP98b, Web96, ATCL^{+}98, And96a, EL98, O’C95e, You96a]. Generator [BD98a, HFA^{+}98, McC98a, PQ95, Pat99, Ano97-156]. Generators [CMH99]. Generic [Ano96-135, Ano96-146, Gla97d, CS99]. Genericity [BOSW98, CS98, Tho97b]. generics [TKH99]. Genetic [Lut99c]. Genetics [Ano98o]. gengo [IK96]. Gently [Bis97, Bis98, KW99]. GeoComputation [Abr96]. geometric [BCLT96]. Geometry [ACM96]. Geophysical [JPK98a, JPK98b]. geoprocessing [Vck96]. George [Ano97g]. Georgia [ACM99b, Joy98]. Geoscience [Ste96b]. Geospatial [AS96]. German [Sta96]. Germany [Ano96-147, Bod96, IEE96k, IEE96j, LM98, SLPM96, Uni97a]. Geschäftsanwendungen [Pie97]. gestalt [Zac97]. Gestion [FB98]. Get [Ano96c, Ano96-130, DDJ98a, Oue95b, Ano96-42, Ano96-188, DiD97d, Fri98, Gau96j, Gau97j, Hay96c, Mac97c, Mac97b, Sey96]. Gets [Ano95f, Ano96-84, Ano96-86, Ano96-85, Ano96-122, Ano97-98, Cas96a, The98, DeV95, Gar98b, Val95, Win95e, Ano96-133, Hay96h, McC98c, Nel95b, Pic96d, Ste97b, Vij97a]. Getting [Ano98m, BC97a, CBT98g, C^{+}xxa, CB98, Gra96b, Lut99d, MF98, MF99, S^{+}96b, Sun97c, van98a, Ano96-87]. GI [LM98, Wat95d]. giants [Ner97]. Gigapoche [Mar97b]. Give [Ano96-76, Mor97b, Par95e, Ano96-188, Ano97-75, Ban95b, Gau97c, Gof97, Gre97k, Kar95b, Pic96a]. Gives [Han97, McK95b, Web97c, Web97d, Col96a, GJ96a, Mes97]. GJ [IPW99]. Glasgow [Moh98]. glitches [GJ97]. glitz [Agn97]. Global [CRS98a, Dav97c, Jon98, Uni96b, Uni97a, Uni97b, Uni98a, Uni98b, Wae96, WRB^{+}97, AISS97, Ano98s, CRS98b, Che97, Uni96a, Wal96i, Ano98-32, Ano98-33]. globalization [Pir97]. globally [LH96]. Globbing [HHW^{+}99]. glory [Kin96]. glow [Hea95]. GNAT [CDG97, MB98c]. GNU [Kuk98]. GNU/Linux [Kuk98]. Go [Ano96-28, EMN98, Hay95c, Kru98, Ano96-136, CT96, The98, Gau97c, Saa96, Red97b]. GO-Joe [Red97b]. GoAhead [Kro99d]. Goes [Ano96-88, Gre97i, Oue96a, Wil96f]. Going [GMPS97, Uni96b, Uni97a, Uni97b, Uni98a, Uni98b, Sco96a, Uni96a]. gold [Kin96, BCC^{+}97, LS96, Wod96]. Good [All99a, Dar98b, Hin96, Kay97, Mar98c, Ban95a, Dar99a]. Goodies [SR97]. Gordian [Kro98d, Kro98c]. gorilla [Oua97]. Gosling [McC96b, McC96c]. gospel [Ham96c]. GPS [Wip96]. Grace [GC97]. grain [BBC^{+}97]. Grained [PH99a, JLT98, LTC96]. Grammar [AFF99]. Gran [F^{+}xxa, Kolxx]. Grand [Cadxxb, Hac96]. Grande [N^{+}xxb, CMH99]. Granularity [GSvD99]. GRAPh [BML97b, Tun98]. Graphic [Ano99d, Ano99e, Ano99f, CR99, Gea96, GM97c, GM97d, GM98a, Gea98, Gea99, GM97b]. Graphical [Ano97o, Dan98, Shi99, BDKK96]. Graphics [Har99a, Jan99, Knu99b, Kro98d, Kro98c, Kro99a, Kro99b, Kuk98, NRG^{+}99, Sch97j, Skl95, Ste97f, WG99, dN97, Amm98, Ano96-195, Ano96-151, Ano97-44, BV98, Muk99, RP98, SML96, SPV99, Sun97c, Tub95, WK99a, WK99b, Sil95, Ano99b]. Graphs [MB98a, Ano97r]. Grappa [BML97b]. great [Bru97]. Green [Kro98e]. greener [Sho96b]. GreenTree [Kro98d, Kro98c]. grenades [Knu98a]. Grep [Pit98b]. Grepping [HHW^{+}99]. grid [DW97b]. Grim [Ric95b]. groggy [Pic96c]. groom [Ano96-133]. große [Ano96-67, Ano97-31, Ano97-32, HKK96]. grosser [KS96b]. Ground [Law98, O’N98b].

Group [Ano96-160, Ano98-34, BN98b, DDJ99a, DLM99, JCS98a, Lea99, GH97, JCS98b, Ner97, Ude97b, WWR97a, WWR97b, Kro97, Kro98d, Kro98c, Kro99e, RNBV98]. Group’s [Pot99]. GroupWise [Col96a]. Grove [Det97]. Growing [Joc98]. Grows [Gro96b]. Growth [GC97, Ano96-153, Ano96-171]. Grundlagen [Min96c, CH98c]. guarantee [Ano96-102]. guard [Mey99]. Guarding [GS98b]. Guards [Ano97-122]. Guest [Urq97b]. GUI [Ano96-137, Ano97-78, Cul98e, Dar99b, Don96, Lut98a, Wal97i, Zol96]. Guia [Ano97-97, Car97a, Gooxxa, Jawxxb, KKxxa]. Guida [GGxxa, Jaw97b, Jawxxa, LP97b, LP97e, NS97a, N^{+}xxb, Lemxxd]. Guide [Ait96c, Ait97b, And98a, Wil97e, Ano97-160, Ano98x, Anu96a, Aoyxxa, AFHxx, BB97b, Boo98, Cha98a, Cobxxa, Con97b, Coo97b, Coo97c, Cor97a, Cue96, DT96, Dav98a, Dow98b, DS97b, Ene97, Eng97d, Eud97a, FR97, Feg97, Fegxx, FF96a, Fis96, Fla96d, Fla96e, Fla97g, Fla97h, Fla98e, Gai97, GPF97, GS96, GD98, Goo97a, Goo97b, Har96b, Har97a, Hel97a, Hel97b, Jar98, Jaw96, Jaw97c, JSA98, JRC97, Kas97a, KK96a, KGTF97, Kim97b, LG99, Lem97d, Lem97e, Lia99, MO96, Mic97b, M^{+}xxa, MS97, MP97b, Neo97, Nic97, Nor96, NS96, Pop97, Red99, RH97b, RH98, RHE99, SMC^{+}96, Sip98, Smi97, Tho99, TPHB96, Tit96b, TGxxa, Tuc98a, Urg96b, Wal97b, WC98, WW96b, Wei97a, WB98, Wil97f, Woo99b, You97a, Zuk99, Eud97b, van97e, Anu96b, HW97, JC97, Jur96, KK97b, Par97, TG95]. guide [Wes98, Nov98, Ano96p, Ano98g]. guided [JJG98]. Guidelines [Sun99b]. GUIs [WC98]. Gulp [CM96]. gulps [Dry97]. Gun [NJ95]. guns [Kin96]. Gurus [V^{+}96]. Gutenberg [DDJ99b].

Habanero [CGJ^{+}98, DAB97]. Hace [Anoxxd]. Hacker [Ano96-89]. Hackers [GC97, Ano97-75]. Hacking [WBB^{+}97]. Haefel [Ano99a]. Hahn [Par95a]. Hairy [Ste96d, Ste96e]. Hall [Gea96, Die97]. Hamamatsu [IDK98]. Hamburg [LM98]. hand [Knu98a, Renxx, Ano96-56]. Handboek [Gooxxb, Walxxd, McCxxa]. Handbook [Ano97-84, Anoxx-48, Che98, FP98, Goo96a, HRSM97, HR98, HR99, Jaw99, Mas97b, Nau96a, Nau96b, PF96, Woo97b, Yeexxb, Amb97, MP96, NFP96]. Handle [Gra96b, Pit99, Stu98]. Handling [Pit99]. Hands [AH97, Ano97-40, Anoxx-39, EMe97a, EMe97b, GPF97, KGTF97, Min96b, Ref99, RM97]. Hands-On [KGTF97, EMe97a, GPF97, Ref99]. hang [Sun97c]. Hard [MGL^{+}98]. hardback [Ano97l, vD99b]. Hardware [Ano96-153, Cla97a, Kno95, Kro98d, Kro98e, Kro98c, Kro98b, MD99, OT97, TO96]. Harissa [MMBC97, MS99b]. Harness [Win95n]. Harold [Ano97j]. Hashing [Uzg96]. Haskell [Rei99, Wak98a, Wak98b]. Haskell/Java [Rei99]. Hasn’t [Gar97a]. hassles [Ano96-42]. hastens [Ano96-195]. Hate [HHW^{+}99]. having [Stu98]. HDF [MLF97]. HDS [Way96g, Web96]. HDTV [Gro96b]. Head [Oue95a, Ric95a, Win95e]. Headaches [Pry96]. headed [MF96a]. headers [Ano96-153]. Heads [GC97, Gre96c]. Health [Ano98-30, Wae96, BCF^{+}97, Chu98b]. healthcare [W^{+}98]. Heap [HAT98a]. Heat [GC97, Ano96-87]. heating [Wil96a]. Heats [Gar98b, Gau96d]. heaven [PM97d]. heavy [Sli97]. heavy-duty [Sli97]. held [Cul97c, IEE98b, IDK98]. Help [Ano95b, Kro98e, Pel95, Win95n, Ano96-125, Ano96-188, Ano97-69, Flo97, Laz99, Mey99, PSTZ99, Way96b]. helped [Ubo96b]. Helping [Laz99]. Helps [Laz97, Ano96-90, Ano96-151]. Herbal [HC99a]. Herbal-T [HC99a]. Here [LP96a, LP96b, Lin96d, vSS96e, Har99d, Oua97, Urg96a]. heterogeneous [ERKE96, Mar99b, WWR97a, WWR97b, YC97]. Heures [BS97d]. hexagrams [Fic96]. Hidden [Tes96a]. Hidden-Frame [Tes96a]. hierarchy [Lak96]. High [ACM98a, AFH98a, Ano99g, Espxx, FFHA98, Gar97a, Gar98b, GFHM98, GHM98, Gro96b, HBG^{+}97, IEE96n, Lut98b, SEH98, SEH99a, SEH99b, Shi99, Ten99, WR99a, YSP^{+}98a, AFH98b, Ano96-192, Ano96b, Bar97, CL97e, FF96b, Fox98a, Fox98b, GGMS99, GME99, Jav98, LM99, Nan95, Wel99a, YSP^{+}98b]. High-Capacity [Gar97a, Gar98b]. High-End [Lut98b, Jav98]. high-level [Nan95]. high-octane [Ano96b]. High-Performance [ACM98a, GFHM98, HBG^{+}97, SEH98, Shi99, YSP^{+}98a, GHM98, SEH99a, SEH99b, CL97e, Fox98a, Fox98b, GME99, LM99, Wel99a, YSP^{+}98b]. high-score [Bar97]. Higher [The98]. Highlights [Gut95k, Ano98u]. Highway [Fer95, EBB^{+}96]. hijack [Cou96a]. Hike [Ano96-28]. Hilbert [HHW^{+}99]. Hilfe [Gol96b, Golxx]. Hill [BUP99]. Hills [Kro98e]. Hintergründe [van97a]. HiProf [Ham97b]. Hiroshima [IEE96q]. History [BAC^{+}99]. Hit [Met95a]. Hitachi [Ano97-118]. hits [Ano96-42, Gau96b, Gau97m]. HMTL [O’C97]. HMTL/SGML/JAVA/VRML [O’C97]. Hoff [McC96e, Gut95a]. Hogs [DDJ99b]. HOL [Pus99]. Holes [Lea97c, MF97b]. Hollywood [Ham96c]. Holy [Dob98, Lew98]. home [Gre97b]. Homepage [GK97a, PWxx, tYpCxx]. Homepage-Design [GK97a]. Homepages [Küh96]. Homer [Mag95]. homogeneous [WWR97a, WWR97b]. hones [Won97]. Hong [Tri95]. Honors [DDJ99b]. Hood [Hol97a, Ven97b]. Hooked [Ano96-44, Lea96b, McC96e, Nic96c, vSS96b, vSS96c, vSS97, v^{+}xxa, Nic96d]. hooks [Kar95b]. hoopla [Ano96-102]. Hopes [Hay95c, WP95]. Hora [Fra97]. Horas [Cad98a, Cadxxa]. HORB [Hir97]. hören [BKP97b]. horse [Nat96]. Horses [TW96]. Horseshoes [Knu98a]. Hostile [Ano97-122, Lea97c, MF97b, Sim96e]. hosts [Ano96-151]. Hot [Als95a, Als95b, Ami97, Ano96-183, Ano96-189, Ano96-190, BCH^{+}98, Boz95a, Cla97b, DeR96, Hay95b, Her96b, IEE96e, IEE97, IEE99a, IEE99b, LaM95f, Mac97d, Mac97e, Sin96a, SS96, Ven96, Wag96, Ano95v, Ano96-53, Ano96-59, Bab95, Ben99b, DM96a, Mac97a, SF96a, vSS96b, vSS96c, Ami96a, Ami96b, Ano96-148, Ban95b, Des96, Dre96, Gut95a, Kar95a, Kar95b, Lab95, Mag95, Nel95a]. Hotel [BUP99, Cim96, Nat98, Ten99]. HotGAMS [Bra96]. HotJava [Beh95, Dre96, Gol96b, Als95a, Ano95d, Ano95s, Ano97-104, Ano97-167, Anoxx-60, Beh95, DFW96, Dec95c, FLW96, Gon95a, Gon95b, Lou95c, Rod95, SvJ96, Tri95]. HoTMetaL [Ano97-154, How96]. HotOS [IEE99b]. HotOS-VII [IEE99b]. HotTEA [Leh96]. Hotter [Ano96-84, Ban95a, Wil96a]. hottest [Ban95a, Fla96f, WB^{+}96c]. Hour [Mee96, Ano96y]. Hours [Cad97, Cad98b, Mee96]. House [Beh95, DDJ99a, Sim96a]. How-To [PPJ96, SLO96, SL97a, SLxx, Sun97c, Pot99]. HP [Ano97p, CCF^{+}97, The98, GC97, Gar98a, GJ96a, Sto97, Ten95b, Wol97]. HPC [BDV^{+}98a, BDV^{+}98b, FFKO99]. HPC’99 [Ten99]. HPcc [FFHA98]. HPJava} [CZF^{+}98a, CZF^{+}98b]. HTML [Die96, Diexx, ER97, GK97a, KS96a, KS96b, Schxxe, Abexxb, And96b, Ano96-46, Ano96-202, Ano97-154, Anoxxa, Anoxxn, Anoxx-80, BT99, BKP97a, Bre96, Bro96, Bry98, Buc97, CC98, Chixx, Coh96a, Cro97, Gar97e, GFR98, Gol96a, Gol98b, Got98a, GG96a, Hal97b, Her96b, Hie97, HL97b, HL97c, JGKR96, Jac96, JLW96, Tou99, JOA98, Kar96, KG97a, KS96b, Knaxxb, Knaxxa, KDH96, Kro99a, Kro99b, Kro99d, LO96, LOA^{+}96, Lad96, LO97a, LO^{+}97b, LO98a, LO99a, LO^{+}99b, Lot95, McM96, Mee96, MHP98, Moh96, MPxx, MP98, MP99, NR96, Neo96a, Neo96b, Neo97, NR97, Neo98, Neo99, P^{+}98b, Rus98, S^{+}96a, SR97, Schxxc, Schxxf, Schxxd, Sha97b, Shixxb, SD96, Tea97, T^{+}xxb, VB96, Wor97, Wag97b, Wea97a, WW98, Wis97, Xu97, Yanxx, ZJ96, RNBV98]. HTML-Crashkurs [KS96b]. HTML-script [Gol96a]. HTML-Seiten [KS96b]. HTML/ [RNBV98]. HTMT [YK99]. http [GS97b, BCE98, JLW96, Lot95]. hub [Ano96-82, Ano96-133, Dry96a]. Human [Lut99c, Tau96, Ano97c]. Hunter [Ano99c]. Hunting [SC97a]. Hurry [New96b]. Hurt [Gar97a]. Hurts [KoCaB98, Kah98, KD98]. Hyatt [Ten99]. Hybrid [MS99b]. hyojun [IK96]. Hype [BBE^{+}99, Bek97, Oua96, McC98d, Sul95a, Ven96]. hyped [Ano96-54]. Hyper [ZKM99]. Hyper-Programming [ZKM99]. hyperbolic [LR96]. Hypermedia [Ano98-27, BV97, EH96, Gop96]. Hypertext [Div98]. Hyperwire [Jarxx].

I-Way [Ano96-53]. i.e [USE98d]. I/O [Ano96-137, Har99b, Har99c, McC98f, PL96, Wel99a]. I18N [Bec98]. IA [Tho98]. IA-64 [Tho98]. IA64 [KFL99]. IAFE [IEE96p]. IASTED [Ano99h]. Iband [Nas95a]. IBM [Ano95c, Ano95f, Ano96-47, Ano96-48, Ano97r, Ano97-41, Ano97-110, BFGS98, CBT98f, CBT98g, CLR98, Com95f, Cox95a, DiD96, DiD97a, DiD97b, DiD97c, DiD97d, Gar98b, Gar99, Gau96a, GJ96b, Gau97c, Gau97b, Got98b, Gre97g, Hay96j, Kar95b, LaM95b, McK95a, Oue96b, Par95c, RCB98, Sim97a, Sim97b, Ste97a, Wag97a, Wea97b, Win95e]. IC [Gep98, Kro98d, Kro98c]. ICE [IEE96m]. IceT [GS97c]. Ich [DE96a, LM97c]. ICL [Ano98-30]. Icon [PTTA99]. ICS [Ano97-42, The98]. ICSA [Kro98b]. ICSA-Certified [Kro98b]. IDBC [Joc98]. IDC [STV95]. IDE [Ano96-91, Ano97-120, Gre97h]. IDG [Jav96d]. Idiom [LG99]. Idioms [Laf96, Lafxxa]. Idiot [WW96b, Wei97a]. IDL [Küh97, LBS98]. IEEE [CEG96, Dar98a, Dar98c]. IEEE/ACM [IEE96l]. IEEE/IAFE [IEE96p]. IEEE/RSJ [IEE98a]. IETF [Lea99]. If [Lew97b, Zol96]. IFC [BCW97, Hol97g, PF97, Rao97]. IFIP [LM98]. IFIP/GI [LM98]. IGARSS [Ste96b]. Ignore [Wel97]. II [DY99b, Fox97b, Wal97a, Way98a, Wu97]. III [Ano96-43]. IIIMF [Hiu98]. IIOP [Sim97e]. I’ll [Bab97a]. Illinois [USE99d]. illustrate [Sim96f]. Illustrated [Gra99a]. illustrations [GSvD99]. ILOG [Ano97-43, Ano97-44]. im [BK96c, BK97b, BKxxb, G^{+}97, Pie97, Beh95]. IMACS [Syd97]. Image [Lut99a, BA96, Gol99, KP99, Lut99c, Lut99b, Lyo99, VPV99, dN97, Ref99]. Images [CBT97e, Hor96a, Hor96b, MM99, Vad95a, Ano96-151]. Imaging [Kro99a, Kro99b, Sko96]. Immagini [Dec97a, Decxxa]. Immedia [Bat96a]. Impact [SZTR96, FC96a, GGL^{+}97]. Imperative [ASG98]. Imperial [Ano96-49, Ano97-45, Ano97-46]. implantar [Ben96b]. Implement [Ber96e, Ses98, Wil99b, Ber96a, KP98a, KP98b, WB^{+}96c]. Implementation [ACM99b, App97, App98, Bot98b, BC97b, Cro99, Gut95f, HHK98, HBG^{+}97, KWB^{+}98, LS97a, Lan99b, PIPR97a, PIPR97b, PTTA99, Ran97a, Ran97b, RL98, Sar99, SMY99, TO96, TBA98, USE96a, vD99b, Ano97e, Ano97l, Ano98k, BKCG98, Gon99, Led98a, MvNV^{+}99, MKS^{+}96, PS96, SML96, Wal96b, Wal97a, Xu97, Zac97]. Implementations [CMH99, Gon98, Kro98d, Kro98c, Ste98a, Wal96e, Zei97]. implemented [DAB97, GBS98]. Implementing [AAC^{+}99a, AAC^{+}99b, Bow96a, GK98, HCC^{+}98, PSS98, Tai98, Tes96a, ZK98a, ZK98b, RMB^{+}98, Tol97a, Tol97b]. implications [NTS^{+}96]. implicit [BG97b]. import [All99a, All99b, All99c, All99d, All99e, All99f]. important [Gil97a, Lin98b, Nat96, Sco96a, Woe97, vSS96e]. Imprescindibles [Fra96b]. impressed [Ano96-102]. Impressions [Dec95a]. improve [CFM^{+}97a, CFM^{+}97b, Orc98, Way96b]. Improved [HCJ^{+}97, Lak96]. Improvement [Lut99d]. Improving [Num98]. IMSA [Ano99h]. in-house [Sim96a]. Inc. [Ano98-27, Hes95]. included [Ano97k, Her96b]. Includes [BS97c, BSxx, Lot95, Sch96a, Ano96-32, Ano97f, Ano97-151, Ano97-152]. Including [Desxx, Sta96, Ano95s, Hal98a, Key98, MAS96]. Incompatible [Rym95]. incorporate [Ano95l, GJ96b]. Incorporating [BGH99, Nap98]. increase [Eck96a]. Increased [Ano96-153, McK95b]. increases [Hay96u]. Independence [NRG^{+}99, BRS99, Cli98b]. Independent [BC99a, Gar97b, Hiu97a, Hiu97c, Hiu97b, BF99b, D’S97c, Gau97c, Lak96, Rob98a]. India [Ano98l]. Industrial [Ano98l, CDBM97, Kap99, Gre98a]. Industry [Ano98l, DB97, Dei96a, Kan96b, Ort99, Ano97-52, Liu96, Ner97, Sch97f, Kro98b]. Inferno [Ano96-28, Mac97a, Mac97d, Mac97e, YRE98]. infiltrated [Gan97]. Info [Fer95, Met95b]. InfoAccess [Kro99d]. InfoBus [Hoq98]. InfoDock [Ano97-121]. InfoMagic [Ano96-137]. Informant [Bra96]. Informatics [Chu98b, Cim96]. Informatik [Wol98]. Information [Ano97-30, Ano98-27, CEG96, DY99a, DY99b, Fis96, IEE96c, IEE96o, Lun99, Ano96-193, Ano96-186, BKP97a, BF99b, Car96c, CCB96a, Chu98b, Eck97b, Eck97c, EBB^{+}96, GO96, KT96, Mye99, NAC99, Not96, O’C97, TYY96, TB96]. Infoscape [Ano97-47]. Infrastructure [BKKK98a, IEE96m, BKKK98b, Wal96i]. infrastructures [JAS96a]. inheritance [Ber98g, CM97a, TKH99]. Initialization [FM98, Sch99c, FM99b]. initiative [Ano98-35]. initiatives [Got98b]. Injection [Ano98m]. inkl [BKP97b]. inklusive [BK97b, Sta96]. Inks [Win95g, Ano96-196]. Inktomi [Kro98e]. Inleiding [Anoxxp]. Inner [Ben99a, Shu97b]. Innotech [Ano98p]. innovatiions [IEE98a]. Innovation [Nas95a, Par95g]. Innovations [Par95h, Kro99d]. Innovative [Ano96-135]. Inprise [Ano98q]. Input [Hiu97a, Hiu97c, Hiu97b, Hiu98, Leo98, Lin98a, OL98, Rao99c, YHW98, GHP^{+}96]. input/output [GHP^{+}96]. Ins [SR97, You97a, CH97f, HL97b, HL97c, NP96b, Pew97b]. Insecure [Han99b, Han99a]. Insider [Mor97d, Morxxb]. Insignia [Web97a]. Inspector [Kro99c]. inspired [CM97a]. installation [Ano96-77, Stu98]. Installshield [San98a]. Instant [Ano97-49, DMB96, Gul98, HL97a, MPxx, OHE97b, Orf98, Pew96, Pew97a, PS97, PP99, Pewxxa, Syd96a, WG97a, Pew97b]. Instantáneas [Afe96a, Dar96]. instead [RFPG96]. Instruction [BBJR96, SLC99, BNB^{+}98, BK96b, GME99, KCPC98, DF96]. instructional [Ciz97]. Instructions [Qia99]. Instrumenting [LZ97]. Intanetto [IK96]. Intanneto [DM96a]. Integers [Gri97b]. Integra4 [Ano97-118]. Integrate [Ado95, Ora95, FFKO99]. Integrated [Ano97-55, The98, Div98, FFHA98, Lev99, HK99, Ren98, YLQ98]. Integrating [Ano98-29, Koo98, Owe96, Pog96, Sco96c, Sha97b, Yan96, Had98a]. Integration [Kro98e, Lot95, Lut99c, Por98, PJR^{+}98, AA97, Ano96-153, Boy98c, HC99a, JADR96, Wat95f, Yas96, Dic97b, Dicxx]. Intel [Ano97-107, Det97, Gar97c, GC97, Lut98b, Sch96b, Tho98]. Inteligente [Car97a]. intellectual [Ale98]. Intelligence [IEE96f, IEE96i, IEE96p, BK96a]. Intelligent [Ano96-154, BB97b, EBR96a, EBR96b, Flo97, FGL96, Gar98a, IEE98a, KP99, Pet97, Wat97b, Gam96, VI96]. Intel(R) [BGH99]. intensive [IDK98]. Inter [BC98a, MST99, Mec97]. Inter-Applet [BC98a, Mec97]. Inter-procedural [MST99]. Interacting [Ste96f]. Interaction [Doe97, Fet97]. Interactive [Ano95q, Ano96-51, Ano96-92, Anoxx-59, BSS96, Dan97c, GM98b, GJ96d, HNS97a, Hor96a, Hor96b, Kap97a, KK96a, LPMG97, Lor96b, Man96b, Man96d, MO96, Min96a, NE98a, NE98b, PS96, RK97b, S^{+}96a, SS96, TRMR97, Tes96b, Thaxx, TPHB96, TMM^{+}98, TMMH99, Tun98, Wea97a, Ano96-201, Ano97t, BKCG98, BB97a, Dav97a, GSvD99, HRW^{+}96, KB96, KK97b, Mag95, McC96d, PSTZ99, Rob98a, TGGP97, Wat95f, WK99a, WK99b]. interactive-oriented [Dav97a]. interactives [Coh96a]. Interactivity [Ano95p]. Interaktiv [BK96c, BKxxb, BK97b]. Interaktive [vSS96d]. interaktiver [DD97h, DD98b, Küh96]. Interativos [McCxxb]. Interattive [GGxxa]. InterCon [Net95c]. Interconnect [Kro98e]. Intercontinental [Ano97-139]. interdisciplinary [Ber98g]. Interent [Anoxx-49]. Interest [Gut95k, Wil96e]. Interface [Ano96-137, BCFK99, Bec98, Ber96e, Ber98b, CBT98j, CJ98, CvE98, GM98d, GYT96c, GYT96d, G^{+}xxf, G^{+}xxg, HLR97, Kin97a, KP99, LSxx, MSS98, Sun95a, Tea97, Tur97, Wal96b, Wal96e, Wal97a, Wil99c, Ano97c, Ber96a, Cul98c, DLM99, EBB^{+}96, EH96, FTT97, Gop96, WOC99, YB99, Lia99, Rao99e, Wil97d]. Interfaces [Ano97-30, CMH99, D’S96a, Gut98, Ban98]. Interfacing [NF96]. Interlanguage [BK99].

Intermediate

[HS96, LST99, Gos95, GLV99, Tým99]. Internal [KW96]. International [Abr96, Ano96-154, Ano96-160, Ano99h, AOV^{+}99, Ber97l, Ber97m, BUP99, CI96, Cha99e, Doo99, FGL96, FW96, IEE96c, IEE96f, IEE96b, IEE96l, IEE96i, IEE96n, IEE96o, IEE96k, IEE96q, IEE96j, IEE96s, IEE98b, IEE98a, IDK98, Kum98, MJA98, OW95, PZ96, PJR^{+}98, SLPM96, Ste96b, Str96, Uni96a, Uni96b, Uni97a, Uni97b, Uni98a, Uni98b, Wae96, YH96, Ano98-34, LM98, DF96, Kro98b, UCD99]. Internationalization [Ano96-93, Ano98-32, Ano98-33, Fre96b, Fre97c, Fre97d, Jon98, Pir97]. Internationalized [MBH97, McF97, McF98a, McF98b]. InTerNet [Ano96-137, Ano96-135, Ano97-118, Ano99h, Cim96, C^{+}97c, D^{+}98b, Dre96, Ene97, Gla97b, G^{+}97, H^{+}97, Hie97, Jim97, Klu98, Law99, Suaxx, Thoxxb, Uni96a, Uni98b, USE97c, USE99b, Vac97, Wat97b, AR98, Abexxa, Int96b, Ale96, And96a, Ano95d, Ano96l, Ano96-43, Ano96-53, Ano96-54, Ano96-55, Ano96-52, Ano96-74, Ano96-73, Ano96-97, Ano96-59, Ano96-106, Ano96-128, Ano96-166, Ano96-205, Ano97-53, Ano97-76, Ano97-122, Ano97-166, Anoxx-40, Anoxx-41, AL96, BSS96, BM96a, BN98b, Bet95, Bor95, BGGR98, Boz95b, Bri99b, Bro95b, CBT98h, CW96a, CW97, CW98, CWxxa, C^{+}xxb, CRS^{+}96, CSY97, CCI^{+}97, CP98, Com95b, Coo97a, Cow97, Cro95a, Dam97, DS95a, De 96, DeV95, Des96, Dre96, Fei98, Ferxx, Fla96a, FB98, Fra98a, Gar97a, Gar97c, GC97, GPV97, HMG^{+}95, Har96a, Hay95d, Hay96l, HC99a, Hiu97a]. Internet [Hiu97c, Hiu97b, Hiu98, Hor96a, Hor96b, IBM95, IF97, IAJR97, JCxx, JN95a, JN95b, JN95c, Kal95a, KP96, KSB97, Kie96, Klu97, Kluxx, Kro97, Kro98b, LaM95e, LO99a, Lan98, Lan99a, LP96a, Lie96a, LS98, LB96, LCxx, Mac95, MF96a, McC96a, McC97b, McK95b, Mes95b, MW95, Mes95c, Mic95, Mil95a, Mil95b, Mil95d, Mon97, Mor96a, Mor96b, Nel95a, NE98a, NE98b, Net95a, NB96, Not96, Oue95a, PM97b, PS96, Pie97, RA97, Rey96, Rod96b, Rus98, SBB96, Sch95, Sch96d, Sco96a, Sem96, SF96c, Shixxb, Sho96b, Sim96c, Sin95b, Sko96, Spi96, Sun96, Suš97, Taf96, Tay97, Tay98, Tay99, TPHB96, TPHBxxa, TPHBxxb, Thoxxa, T^{+}xxa, TG95, Tit96b, Tur97, T^{+}98b, Uni96b, Uni97a, Uni97b, Uni98a, Ubo96b, Uim95, Uma97, Vac96, VM96, Wal96a, Wat95a, Way98c, Wijxx, WSL96, Wil96c]. Internet [Wil97c, Win95a, Win95b, Win95c, Win95d, Win95n, Yanxx, You96a, ZK98a, ZK98b, ZTLV98a, ZTLV98b, dRPT98, van95, Hum98]. Internet-Applikationen [H^{+}97]. Internet-based [CCI^{+}97]. InTerNet-LINK [Ano96-137]. Internet-Programmierung [Vac97]. Internet-Software [Hie97]. internet.profi.praxis [MP97a]. Internet/Intranet [Ano97-118]. Internet/WWW [ZK98b, ZK98a]. InterNetivity [Ano97-50]. Internetsprache [Ano96-56]. Internett [Dan97b, Dan97a]. InternetWorking [Gon95a]. Interop [MW95]. Interoperability [BDV^{+}98a, PCWGM98, BDV^{+}98b, HC99a, MHO96]. interoperable [Vck96]. interpersonal [Ano97-54, Pre99]. interpretation [Fri96]. interpreter [BM96c, MB96c]. Interpreters [RLV^{+}96, WB99]. interprocess [Bla96]. Interval [NNS97]. Interview [Ano97-51, Ano97-52, Bow96b, Gut95a, Par95a, Par95b, SW95, Ten95a, Win95a, Win95b, McC96c]. Intonet [Kim97d]. Intra [Klu97, Kluxx]. Intra- [Klu97, Kluxx]. Intranet [Abl96, Ano96-135, Ano97-118, Anoxx-42, Klu98, Manxxb, PM97b, Pie97, Schxxf, Schxxd, Gil97c, Lin96e, RGT96, Ano96-72, Ano96-191, Ano98-31, G^{+}97, Schxxc, Schxxe].

Intranet/Extranet [Ano98-31].

Intranet/Web [Ano96-135]. Intranets [Ben96b, DC97, Flo97, KW96, Wat97b, Ano96-193, O’C97, Ano96-64, Vac96]. IntranetWare [Knu97h]. intrinsics [Reg96]. Intro [Bla98]. Introducão [dVGxx]. Introducción [Dec96a]. Introduce [Nas95c]. Introduced [MD99]. Introduces [Ano97-131]. Introduction [Ano96-137, Anoxxq, Anoxx-43, AW98, AC99, CBT97c, CCYxxa, Dec95c, DH98, DD97d, Die98, EMe97e, EMe97d, EMe97g, EMe97f, Hun97, Hun99, HLTxx, KMR98, Kon98, Liaxx, MRA98, Nie95, N^{+}97b, N^{+}xxd, Rie97b, Row97, Sad97a, Sha98a, SN98b, Rapxx, SP97, Urq97b, Wu99, Wuxx, DD97c, Ekxxc, Got98a, Gre97k, Wil97e, Wil97f]. Introductory [Dav99, MHP98, Ano96-155, Ast98, KSW98, WSK97]. introduktion [EExx]. Introduzione [BG97a, Tis97]. IntRprise [Ano97-53]. Intuit [Gut95d]. Intusoft [Kro98b]. invasive [CL97c]. inventing [Hal98a]. Invisible [Ano95j]. InvisibleWeb [Ano96-135]. Invocation [RWB97, VR99a, WW97, Ber97b, Cli98a, Led97, MvNV^{+}99, TTK98a, Dow98a, KWB^{+}98, Rao99d, TTK98b, Wal98a]. Invoke [Ber97e]. Involving [SR97]. IO [Cha96d, Eck97d]. Iomega [Ano96-40]. IONA [Kro98d, Kro98c, Kro98b]. IP [Gar98a, JLW96]. IQ [KT96]. Irix [Chw98]. Irregular [BCP99a]. Isabelle [Pus99]. Isabelle/HOL [Pus99]. ISAPI [KSB97]. ISBN [Bay97a, Tee99, Ver98, vD99b]. ISCA [YH96]. ISDN [Ano96-28]. ISE [Kro98d, Kro98c]. Islandia [Ten99]. Isn’t [CBS95, Ols97, Ano96-102, Gil96b, Gre97g]. ISO [Gar98b]. ISO10646 [Uni98a]. Israel [IEE98c]. ISSAC [Doo99]. issue [Gau97n]. Issues [Ano97-54, Ano99g, AJ99, EK98, Win95a, AISS97, DiD97c, Gau97f, Gon97, GA98, MKS^{+}96, Men96]. Italy [Nat99, PZ96]. Iterated [Ano96-195]. Itself [The98]. itsellesi [D^{+}xx, Kiu98, KNxx, NKxx]. ITU [Gro96b]. IUC12 [Uni98a]. IV [Her96a, Her96b]. IW [Ven96]. Iway [Met95a]. IX [IEE97].

J [A^{+}96, AFPW97, Aitxxa, Aitxxb, Ano96-124, Ano97-159, Ano97-160, Ano98x, Ano98y, Anoxxm, Anoxx-39, Anoxx-64, Anoxx-79, Anoxx-75, Anoxx-78, Anoxx-76, Anoxx-77, Bay97a, Bla98, Bre96, Cobxxb, CBJ97b, CBJ97a, Cohxx, Thexx, C^{+}97b, CS97b, CWHP96, Cul98a, Dav96b, Dav98b, D^{+}xx, DD97c, DD97d, EMe97a, EMe97d, F^{+}xxa, FC96b, FC98a, FM97, Garxx, Gil98b, Groxxb, Har96b, Harxxa, Hol97h, Hol97i, Hol98b, Holxxc, Holxxe, Holxxa, Hym97, JMM96, JC97, Lad97, Lad98a, Ladxxb, Ladxxa, LA97a, LA97b, Les97, Lesxx, Lin97, Lonxxb, Min96b, Manxxb, Marxxc, Mas97a, Mas97b, McH97, MRA98, Mic96, Mit97b, M^{+}97b, M^{+}97a, M^{+}xxc, Morxxe, MP96, MP97a, Pat98a, Pat98b, Per96b, Per97d, Per97a, Per97b, Perxxa, Perxxc, Perxxb, Rin97b, Ritxxi, Sch97d, Schxxa, Sha97a, Siy96, Siy97, Siy98, SP97]. J [Swa97f, Swaxx, Tee99, Wes97b, Wil98c, WOLP96, W^{+}96, WL97c, WPL97, W^{+}97c, WL98, W^{+}xxa, W^{+}xxb, Woo96, Woo97a, Woo97b, Wooxx, Wygxx, Ano96-39, AC98, BD99, JLW96, Rapxx, VCCH99, WDDF98, Wil98b, Ano97-99]. J-Kernel [VCCH99]. J/Direct [Wil98b]. J2C [Kei95]. jaba [Kim97d, Kim97d]. Jabs [SW95, Swa97b]. jackpot [Can96]. Jackson [Bay97a]. Jacl [LS97a]. Jada [CR97]. Jade [RL98]. Jakarta [Nor96]. jalapeño [AAC^{+}99a, AAC^{+}99b]. James [McC96c]. jangled [Joh97a]. January [ACM97a, ACM98b, ACM99a, Ano98o, BUP99, KJR98, USE96d, USE98f, W^{+}98]. Japan [IEE96q, IDK98, Uni98a, Gut95k]. Japanese [Lun99, BPxx, G^{+}xxd, G^{+}xxe, Gut95g, Lafxxa]. JAPE [Ima97]. Jason [Ano99c]. JAVA [Ano98-29, DE96b, ACM98a, All99b, Ammxx, Ano96o, Ano96-56, Ano96-135, Ano96-137, Ano96-142, Ano96-143, Ano96-161, Ano96-182, Ano97h, Ano97m, Ano97l, Ano97-31, Ano97-32, Ano97-45, Ano97-119, Ano97-120, Ano97-121, Ano97-134, Ano97-162, Ano98e, Ano98-27, Ano98-28, Anu96a, App95, BBB^{+}96, Bat96d, BM96a, Bay97a, Bee96, Ber98h, Ber98i, Biexxb, BH96a, BH96b, BD98b, Bri99a, CW96a, CW97, CWxxa, CH98c, DK98, Die96, Diexx, Dra98, DHM99a, Fox98b, Frö96, Fröxxb, GM97b, Gla97b, Gla98, Gol96b, Golxx, GYt96a, GYt96b, Gosxxa, Gosxxb, Gra96c, GHP^{+}96, Gulxx, HHK98, Her97, H^{+}97, Heu97, Hie97, HC98c, Hum98, Joh98, Klu98, KS96b, Küh96, Küh97, Kuk98, LOKK97, Lea97a, Lea97c, LPC96, Leo95, MB98a, MRA98, M^{+}97c, MJA98, NvO98, O’N97c, O’N97a, O’N97b, OHxxb, Pew97b, Pie97, Roo98, Ros96]. Java [San98a, SBH97, Sha96b, Str97, Til97c, TG96a, TGxxb, Tol97b, Ver98, Wil98a, W^{+}97c, Wol99, Wol98, Yal98, Ano95e, Ano96-46, Ano96-98, Ano96-174, Ano97-30, BV98, BM98c, CSY97, CH97f, Div98, ERKE96, Frü99, GM98b, GA98, Hac96, HJS98, Hub97, KG97a, KP99, O’C97, JAV96b, VPV99, VPM98a, VPM98b, VB96, VI96, YLQ98, Abexxa, Abexxb, Abl96, AM97, Add99, ATCL^{+}98, Ado95, Int96b, Jav96a, Afe96b, Afe96a, AFM97, ADM98, Agh97, Ait96a, Ait96b, AFPW97, Aitxxa, AFH98a, AFH98b, ALP99, ASG98, AVN98, Ale98, AISS97, All99c, AS96, AAC^{+}99a, AAC^{+}99b, Als95b, AFL99, AFF99, AF99, Amb96, Ami96a, Ami96b, Ami97, Amm97b, Amm98, And97a, AH99, And96b, And97b, And98b, AT97, AW99a, Ang99, Ano95w, Ano95a, Ano95b, Ano95c, Ano95f, Ano95g, Ano95h]. Java [Ano95o, Ano95j, Ano95k, Ano95l, Ano95u, Ano95t, Ano95q, Ano95r, Ano95v, Ano96d, Ano96h, Ano96i, Ano96g, Ano96f, Ano96c, Ano96q, Ano96m, Ano96t, Ano96s, Ano96x, Ano96z, Ano96-28, Ano96w, Ano96y, Ano96u, Ano96v, Ano96-195, Ano96-196, Ano96-41, Ano96-38, Ano96-40, Ano96-37, Ano96-43, Ano96-42, Ano96-44, Ano96-45, Ano96-53, Ano96-51, Ano96-54, Ano96-47, Ano96-49, Ano96-55, Ano96-58, Ano96-64, Ano96-62, Ano96-61, Ano96-63, Ano96-68, Ano96-69, Ano96-67, Ano96-66, Ano96-75, Ano96-74, Ano96-71, Ano96-76, Ano96-73, Ano96-70, Ano96-72, Ano96-77, Ano96-78, Ano96-79, Ano96-81, Ano96-84, Ano96-87, Ano96-88, Ano96-83, Ano96-86, Ano96-85, Ano96-89, Ano96-90, Ano96-82, Ano96-93, Ano96-92, Ano96-91, Ano96e, Ano96-96, Ano96-94, Ano96-95, Ano96-97, Ano96-109, Ano96-99, Ano96-59, Ano96-100, Ano96-101, Ano96-65, Ano96-80, Ano96-106, Ano96-102]. Java [Ano96-104, Ano96-105, Ano96-103, Ano96-107, Ano96-60, Ano96-117, Ano96-108, Ano96-110, Ano96-111, Ano96-112, Ano96-119, Ano96-124, Ano96-122, Ano96-120, Ano96-126, Ano96-32, Ano96-125, Ano96-123, Ano96-131, Ano96-132, Ano96-130, Ano96-133, Ano96-134, Ano96-136, Ano96-128, Ano96-145, Ano96-151, Ano96-148, Ano96-153, Ano96-146, Ano96-163, Ano96-166, Ano96-165, Ano96-156, Ano96-157, Ano96-162, Ano96-155, Ano96-167, Ano96-171, Ano96-170, Ano96-183, Ano96-175, Ano96-176, Ano96-177, Ano96-179, Ano96-180, Ano96-187, Ano96-184, Ano96-188, Ano96-189, Ano96-190, Ano96-186, Ano96-185, Ano96-194, Ano96-197, Ano96-198, Ano96-199, Ano96-202, Ano96-205, Ano96-200, Ano96-201, Ano96-204, Ano96-206, Ano96b, Ano97b, Ano97f, Ano97e, Ano97c, Ano97q, Ano97o, Ano97p, Wil97e, Ano97n, Ano97x, Ano97y, Ano97r, Ano97s, Ano97-27, Ano97v, Ano97u, Ano97z, Ano97-34, Ano97-29, Ano97-33, Ano97-38, Ano97-36, Ano97-35]. Java [Ano97-37, Ano97-39, Ano97-40, Ano97-42, Ano97-49, Ano97-48, Ano97-56, Ano97-55, Ano97-53, Ano97-43, Ano97-44, Ano97-47, Ano97-50, Ano97-46, Ano97-51, Ano97-52, Ano97-70, Ano97-71, Ano97-69, Ano97-73, Ano97-74, Ano97-75, Ano97-72, Ano97-91, Ano97-68, Ano97-76, Ano97-77, Ano97-78, Ano97-88, Ano97-66, Ano97-79, Ano97-80, Ano97-81, Ano97-84, Ano97-83, Ano97-82, Ano97-85, Ano97-86, Ano97-67, Ano97-87, Ano97-57, Ano97-58, Ano97-59, Ano97-60, Ano97-61, Ano97-62, Ano97-63, Ano97-64, Ano97-65, Ano97-103, Ano97-101, Ano97-102, Ano97-104, Ano97-113, Ano97-106, Ano97-105, Ano97-108, Ano97-109, Ano97-111, Ano97-112, Ano97-115, Ano97-114, Ano97-117, Ano97-122, Ano97-124, Ano97-125, Ano97-123, Ano97-127, Ano97-128, Ano97-132, Ano97-133, Ano97-130, Ano97-129, Ano97-135, Ano97-131, Ano97-136, Ano97-144, Ano97-146, Ano97-145, Ano97-137, Ano97-147, Ano97-151, Ano97-152, Ano97-148, Ano97-140, Ano97-138, Ano97-143, Ano97-142]. Java [Ano97-141, Ano97-155, Ano97-156, Ano97-153, Ano97-157, Ano97-161, Ano97-163, Ano97-164, Ano97-165, Ano97-158, Ano97-154, Ano97-166, Ano97-167, Ano97-168, Ano98c, Ano98b, Ano98l, Ano98i, Ano98d, Ano98m, Ano98n, Ano98p, Ano98s, Ano98w, Ano98t, Ano98r, Ano98-30, Ano98-31, Ano98-32, Ano98-33, Ano98-35, Ano98-36, Ano99d, Ano99e, Ano99f, Ano99g, Lut99a, Anoxxb, Anoxxf, Anoxxg, Anoxxc, Anoxxd, Anoxxh, Anoxxk, Anoxxj, Anoxxi, Anoxxa, Anoxxn, Anoxxp, Anoxxq, Anoxx-35, Anoxx-32, Anoxx-36, Anoxx-37, Anoxx-38, Anoxx-60, Anoxx-39, Anoxx-42, Anoxx-40, Anoxx-41, Anoxx-43, Anoxx-33, Anoxx-63, Anoxx-44, Anoxx-45, Anoxx-48, Anoxx-49, Anoxx-47, Anoxx-46, Anoxx-51, Anoxx-52, Anoxx-50, Anoxx-53, Anoxxr, Anoxxs, Anoxxt, Anoxx-34, Anoxxu, Anoxxv, Anoxxw, Anoxxx, Anoxxy, Anoxxz, Anoxx-27, Anoxx-28, Anoxx-29, Anoxx-30, Anoxx-31]. Java [Anoxx-68, Anoxx-67, Anoxx-65, Anoxx-66, Anoxx-70, Anoxx-69, Anoxx-72, Anoxx-73, Anoxx-74, Anoxx-81, Anoxx-80, Ant96, Anu96b, Anu96a, Anu97, Anuxx, Aoy97, Aoyxxa, Aoyxxb, App97, App98, Arm99, Arm97, AG96c, AG96d, AG96e, AG96a, AG96b, AG98, AGxxc, AGxxd, G^{+}xxc, AGxxb, AGxxe, AGxxa, AW98, AW99b, Ath98, ADJ^{+}96, AJ98b, AJ99, AGD97, AMPM96, AMP97, AC99, A^{+}99, AC98, Bab95, Bab96, Bab97b, BAC^{+}99, BH99a, Bacxxb, Bacxxa, BKMS98, BDF^{+}98, BM98a, BCFK99, Bak97b, Bal98a, Bal99, BDF96, Bal97, BC99a, Bam97, Ban98, Ban95a, BML97a, Ban95b, BCH^{+}98, BKKK98a, BKKK98b, BML97b, Bar97, Bar99, Bar96, Bar95b, BS98a, BB97a, BSS96, BW98a, BFPV99, Bat96c, BM96a, Bay97a, Bay97b, Bea96, BCF^{+}97, BD99, Bec97a, Bec98, Bek97, BP98, Bel98, BS97c, BSxx, BA97, Ben98, Ben99a]. Java [Ben99b, BKR99, BK99, Ber96a, BA96, Ber96g, Ber96f, Ber96d, Ber96c, Ber96h, Ber96e, Ber97l, Ber97a, Ber97d, Ber97b, Ber97i, Ber97c, Ber97h, Ber97m, Ber97g, Ber97f, Ber97k, Ber97j, Ber97e, Ber98e, BF98, Ber98f, Ber98a, Ber98d, Ber98b, Ber98c, BF99a, Ber99a, BCE98, BKP97a, Ber96i, Ber96j, BN98b, BNB^{+}98, BD98a, Ber99b, Ber99c, BCK99, BG97a, BDY99, BBC^{+}97, BT98, Biexxa, Biexxb, BB97b, BG97b, BVG97, BG97c, BBGD98, BVG98, BHxx, Bin97, Bin96, Bis97, Bis98, BSS98a, BSS98b, Bla96, Bla98, Bla99, BC98b, BC99b, BCP99a, BFxx, BS98c, BS99, BH99b, BDP^{+}98a, BDP^{+}98b, Bok99, BFGGxx, Boo96a, Boo96b, Boo96c, BM96b, Boo97, Boo98, Bor96, BS98d, Bor95, BGGR98, Bös98, Bot98a, Bot98b, Bou96a, Bou97, Bou96b, BS98e, BC98c, Bow96a, Boy98b, BC97b, Boy98c, BRF98, Boz95a, Boz95c]. Java [BOSW98, BFGS98, Bra98, Bra97a, Bra97b, BCP^{+}99b, BN99b, Bre96, BDV^{+}98a, BDV^{+}98b, BL97, BW99, Bri98, BH96c, Bri99b, Bri99c, BLT98, Bro99a, Bro99b, Bro98a, Bro99c, Bro97b, Bro98b, Bro98d, Bro98c, Bro96, Bro97c, BR97, BP99, Bru97, Bru98, BB99, Bry98, Buc97, Buc96a, BW98b, Bud98, BK97a, Bue98, BF99b, BK96c, BKP97b, BK97b, BKxxa, Buhxx, BKxxb, Bur96a, IG99, Bur96b, BS97d, BS97e, BCC^{+}97, CBS95, CBT97b, CBT97a, CBT97c, CBT97e, CBT97d, CBT97g, CBT97f, CBT98a, CBT98b, CBT98c, CBT98e, CBT98d, CBT98f, CBT98g, CBT98h, CBT98i, CBT98j, CBT99, Cad97, Cad98a, Cad98b, Cadxxa, Cadxxb, Cadxxc, Cai99, CKH99, CH98a, Cal99, CW96a, CW98, CWHt99, CWxxb, Can96, Car96a, Car96b, Car95, CV98, CKV98, CCF99, Car96c, CCF^{+}97, CZF^{+}98b, CZF^{+}98a, CCFL98, CZFL98, Car96d]. Java [CBA98, C^{+}xxa, Carxx, CS98, CAZZ98, CDD97, CD97, Cas96a, Cas96b, CDBM97, C^{+}xxb, Cav97, Cel98, Cel96a, CKRW97a, CKRW97b, CKRW99, CGJ^{+}98, Cha99a, Cha99b, CLR98, Cha96b, CL97b, CL97a, CGI97, CLK98, CL98a, CL98b, Cha98c, Cha98b, CL99, CLK99a, CLK99b, Cha99c, CLxxa, CLxxb, CLxxc, CRS^{+}96, CCCL97, CH98b, Cha96c, Cha96d, Cha96e, Cha99d, Chaxxb, Chaxxc, Cha98d, Cha98e, Cha97c, CL97d, Cha99f, C^{+}xxc, CC97, Che99, CIF98, Che97, Che98, CCYxxa, CCYxxb, CDL99, Chixx, Cho97a, CGS^{+}99, CGHS99, Cho97b, Chr99, CM96, CT98, CCI^{+}97, Chu98a, Chuxx, Chw98, CR97, CL97e, CL97f, Cil96, Cla97a, Cla97c, Cla97d, Claxx, CSxx, Cli98a, Cli98b, CSH^{+}98, CM97b, CM97c, CM97d, CM97a, CM99, CLD99, Cob98, Cobxxa, CP98, CMH99, Cof96, Cog97, CMGR96, Coh97b]. Java [Coh97c, Coh98, Coh96b, CMMN96, CJJ^{+}98b, CJJ^{+}98a, Col96a, CWS98, CS99, Col96c, CDG97, Com97b, Com95a, Com95d, Com95f, Con97b, Con97a, CC96, CCB96b, Coo97b, Coo97c, Coo98, CCH97, Cor98a, CH96, CH97b, CH97c, CH97d, CH97e, CH97g, CH97a, CA98, CHxxa, CHxxb, CHxxc, Ora95, C^{+}97c, CR99, Cou98, Cou96a, Cou96b, Cou97a, Cou97b, Cou97f, Couxx, Cou97g, Cow97, Cow99, Cow98a, Cow98b, Cox95b, Cox95c, Cox96a, Cox96b, CFM^{+}97a, CFM^{+}97b, Cra97, Wal97d, Wal97e, Crexx, Cri98, Cro97, Cro99, Cul96b, Cul96a, Cul97b, Cul97a, Cul98b, Cul98c, Cul98e, Cul98d, Cul98f, Cul98g, Culxxb, Culxxa, CvE98, CMSvE99, DDJ98b, DDJ98c, The98, DDJ98a, DDJ99a, DDJ99b, DW96b, DG96, D’S96a, D’S96b, D’S96c, DD96, D^{+}98a, D^{+}98b, DSM99, Dacxx, Dadxx, Dalxx, Dal97, Dal98, Dam97]. Java [Dan96a, Dan97b, Dan97a, Dan98, Dar98a, Dar98b, Dar99a, Dar99b, Dav97a, Dav99, DE96a, Dav96a, DMB96, Dav96b, Dav96c, Dav97e, Dav97f, Dav97b, Dav97d, DBBE98, Davxxc, Davxxb, Davxxa, D^{+}xx, DBK98, DS99, DFW96, Dec95a, Dec95b, Dec95c, Dec96a, Dec96b, DM96a, Dec97a, Dec97b, Decxxb, Decxxa, Decxxc, DD97b, DD97a, Dei97, DD97c, DD97d, DD97e, DD97f, DD97g, DD98a, DDxx, Dei96a, Dei96b, Del98, Del97, Dem98b, Dem98a, DIS99, DR98, DR99, DV98, Des98, Des96, Desxx, Det97, Deu98, DW96a, DiD96, DiD97a, DiD97b, DiD97c, DiD97d, DiG95a, DiG95b, DiG95c, DiG95d, DiG96, Dic97a, Dic97b, Dicxx, DH99, Die96, Die97, Die98, Diexx, DF97, DD97h, DD98b, Doe97, Doh97, Don96, DD97i, DDS99, Dow98a, Dow98b, Dow97, DS97b, DR96, Dra97, Dre96, Dre98, DAB97]. Java [Drixx, DE97, DWE98, DE99, Dry96a, Dry96b, Dry97, DW99, Dua96, Dub97, Due97a, Due97b, Due98, Dug99, DHM99b, Dun99, EMe97a, EMe97e, EMe97d, EMe97g, EMe97f, EMe97h, EMe97i, Eas99, Eck96b, Eck96d, Eck97a, Eck97d, Eck98, ELW98, EKR97, Egr99, EM99, Ek97a, Ek97b, Ekxxa, Ekxxb, Ekxxc, EExx, Ekxxd, EHS^{+}99, EMN98, EBB^{+}96, Ene97, Eng97a, Eng99, EK98, Eng97c, Eng97b, Ent97, Esp96, EBR96a, EBR96b, Ess95, EL98, Eti97, Eud97a, Fie96, Fah96, FRT96, Fan96, Far98a, Far96b, Far96e, Far96c, Far96d, Far97a, Far97b, Far97c, Far97d, Far98b, FR97, Fed97, Feg97, Fei98, FF98, Fer98b, Fer98a, Ferxx, Fic96, FI96a, FLRT97, FHKS98, Fla95, Fla96b, Fla96c, Fla97a, F^{+}97a, Fla97d, F^{+}97b, Fla97e, Fla97b, Fla97c, FGN^{+}97, Fla97f, Fla98a, Fla98b, Fla98c, FFCM99]. Java [Fla99b, Fla99c, Fla99d, Fla99a, Flaxxa, Flaxxb, F^{+}xxb, Fli98, FC96a, FC96b, FC98a, Fol99, FFHA98, For98a, For98b, For98c, FTT97, Fox97a, Fox97b, Fox97c, FF97, FLQZ97, Fox98a, Fox98b, FFKO99, Fox99, FLW96, Fra96a, FB96, FBxx, Fra96b, Fra97, Fra98b, Fra98a, FD96b, FD96a, FI96b, Fre95, FM98, FM99a, FM99b, Fre96b, Fre97c, Fre97d, FP96, FKP97, FP98, F^{+}xxc, FPxx, FT97, Froxxa, Fue98, Fün98, Gai97, GH97, Gan97, Gar97a, Gar97b, Gar97c, GC97, Gar97d, Gar98a, Gar98b, Gar99, Gar97e, Gar96a, Gar96b, Gar96c, Gar96d, GSTV97, GPF97, GM97a, Gat97, Gau96a, GJ96a, GJ96b, Gau96b, Gau96e, GH96a, GH96b, Gau96f, Gau96g, Gau96i, Gau96j, GH96c, Gau97a, Gau97c, Gau97b, Gau97d, Gau97e, Gau97f, Gau97g, GJ97, Gau97h, Gau97i, Gau97j, Gau97k, Gau97l, Gau97m]. Java [Gau97n, Gau97o, Gau97p, Gau97q, Gau97r, Gea96, GM97c, GM97d, GM98a, Gea98, Gea99, GRN99, GFHM98, GHM98, GGS99, GGMS99, GDS98, Gib98, GK97a, Gig97, GM97e, GM98c, Gil96a, Gil97a, Gil96b, Gil96c, Gil97b, Gil97c, Gil98b, GPHL99, GJ^{+}96c, GPO98a, GPO98b, Gla97a, Gla97c, G^{+}97, Gla97d, GV97a, GV97b, GTV98a, GTV98b, GG99, GFR98, Gof97, Gol96a, GD98, Gol98a, Gol98b, Gol97, GS98a, GMPS97, Gon97, Gon98, Gon99, Gon95a, Gon95b, GT97, Goo99, Gor96, GM98d, GT99, Gos95, Gv95, Gv96, GYT96c, GYT96d, GM96, GJS96, Gos96, GYT97a, GYT97b, Gos97b, Gos97a, GJS97, Gos98a, Gos98b, G^{+}xxb, G^{+}xxf, G^{+}xxg, G^{+}xxd, G^{+}xxe, G^{+}xxh, G^{+}xxi, Gosxxc, Got98a, Got98b, Gra96a, Gra96b, Gra96c, Gra96d, GK97b, Gra97a, Gra97b, Gra99a, Gra99b, GS97c, GBS98, Gre95]. Java [Gre96a, Gre98a, Gre96c, Gre97a, Gre97b, Gre97c, Gre97d, Gre97g, Gre97f, Gre97h, Gre97i, Gre97j, Gre97k, Gre98d, Gre98b, Gre98c, Gre98e, Gre98f, Gri97a, GCxx, Gri97b, Gri99, GHP^{+}96, Gro96a, GSHP97, Groxxa, Gro96b, GLC99, Guaxx, GGL^{+}97, GRD96, Gul96a, GR96, GR97, Gul98, GG96b, GG97, GGxxa, GGxxb, GME99, Gut95a, Gut95d, Gut95f, Gut95g, Gut95b, Gut95h, Gut95i, Gut95k, Gut96, Gut98, HKK96, Had98a, Had98b, HT98, Hal97a, Hal98a, Hal98b, HRW^{+}96, Hal97b, Ham97a, Ham96a, Ham96b, HCF97a, HCF97b, Ham97b, Ham97c, Hamxxa, H^{+}xxa, Hamxxb, Ham96c, Han97, Han96, Han99a, Han99b, HNSS99, HS96, HNS97a, Har99a, Har96a, HW97, HFR96, HFR98, Har96c, Har96d, Har96e, Har97a, Har97c, Har99b, Har99c, Harxxb, Harxxc, Har98b, Har96f, HHW^{+}99, HBL99, Har98c, Hav98, HCC^{+}98]. Java [Hay95a, Hay95b, Hay95c, HB95, Hay96a, Hay96b, Hay96c, Hay96e, HN96, Hay96d, Hay96f, Hay96g, Hay96h, Hay96i, Hay96j, Hay96k, Hay96m, Hay96l, Hay96n, Hay96q, Hay96r, HG96a, Hay96s, Hay96u, HG96b, Hay96t, Hay96v, Hay96w, Hay96x, Hay96y, Hay97a, Hay97b, Hay97c, Hay97d, HRSM97, Hel97a, Hel97b, Hel97c, HR98, HR99, Hem96a, Hem96b, Hem96c, Hen97a, Hen97b, Her96b, HJZ^{+}98, Herxx, Her97, Hes97, H^{+}xxb, Hib96, Hib97, HK98, Hin96, Hir97, HYI98a, HYI98b, HL99, Hiu97a, Hiu97c, Hiu97b, Hiu98, Hof97, H^{+}xxc, Hol96a, HB98, Hol98a, Hol99, Hol97a, HBG^{+}97, Hol96b, Hol96c, Hol97d, Hol97e, Hol97f, Hol97c, Hol97b, Holxxb, Holxxd, HI96a, HI96b, HIxxa, HIxxb, Hop97, Hor96a, Hor96b, HC98a, Hor96c, HC97, Hor97a, Hor98a, HC98b, HC99b, HC99c, Hor97b, HS98b, Hos96b, Hos96a, HGH96]. Java [HCJ^{+}97, HFA^{+}98, Hug96, HHSW97a, HHSW97b, HHW97, HH98, Hug97, HS98c, Hum96, HAKN97, HNS97b, Hunxx, H^{+}xxd, HLRT97, Hun97, HM98, Hun99, HC98d, Hus96, HL97b, HL97c, HLTxx, IBM95, IBM97, Jav96d, IPDxx, IR97, IPW99, IGO98, Ima97, IF97, Infxxa, Infxxb, Inf95a, Inf95c, Inf95e, Inf95f, Int96a, IK96, JDK98a, JDK98b, JMA98, JPA98, JRE98, JM96, JM97c, JM97a, JM97d, JM97b, JM99, JMxxa, JMxxb, JPK98a, JPK98b, JvdBH^{+}98, JLMPH99, JS97, Jam96, JLW96, Jam97a, Jam97c, Jam97b, Jamxxa, Jamxxb, Jan97, Jan99, Jar97, Jar98, JPT^{+}99, Jas96b, Jasxx, Jaw96, Jaw97c, Jaw97b, Jaw98, Jaw99, Jawxxa, Jawxxb, JJBW98, Jen98, JLT98, JSA98, JZ98, Jep96, Jep97, Jim97, Job96, Joc98, Joh96a, Joh97b, JN95a, Jon98, JLP96, JC97, Jos97a, Jos97b, JV97, JV98, Jos98, JJG98, Jun98]. Java [Jur96, KoCaB98, Kah98, KD98, KW96, Kal98, KMR98, Kan96a, Kanxxa, Kanxxb, Kanxxc, KK97a, Kap97b, KS98a, KS98b, KHB99a, KHB99b, Kar95a, Kar95b, Kar96, KG98a, KSB97, Kay96b, Kay96a, Kay97, Kei95, KB98a, KB98b, Kie99, Kie96, Kim97b, Kim97d, Kim97c, Kim97f, Kim97a, Kimxx, Kin96, Kin97c, Kin97b, KMS98, KF99a, KF99b, Kiu98, Kiuxx, KNxx, KG96a, KG96b, KSW98, KS96a, KS96b, Knaxxb, Knaxxa, Kne96, Kne98, Kno96a, KSW97, Kno96b, Knu97b, Knu97a, Knu97d, Knu97g, Knu98b, Knu99b, Knu99a, KS97, KW98, KW99, KR99, Koo98, KK96b, KK97d, KK97f, KK97e, KK98c, KWxx, KKxxb, KKxxc, KNNL98, KP98a, KP98b, KT99, Kra96, Kri96, KWB^{+}98, Kro97, Kro98a, Kro98b, Kro99a, Kro99b, Kro99d, Kro99e, Kro99c, Krü97a, Krü97b, Kru98, Krüxxa, Krüxxb, Kühxx, Kum98]. Java [KMW99, Lab95, LO96, LOA^{+}96, Lad96, Lad97, LO97a, LO^{+}97b, Lad98a, Lad98b, LO98a, LO99a, LO^{+}99b, Laf96, Lafxxa, Lafxxb, Laf98a, Laf98b, Laixx, LJ96b, LJ96a, LJ97, LJxxd, LJxxc, LJxxa, LJxxb, LS97a, LO98b, LO99c, Lan99b, LG99, LP98a, LP98b, Lau96, Lav96, Law98, Lea96a, LMM96, Lea97b, LH97, Lea98a, Lea96b, Lea96c, LST99, Lea99, LG97, Led97, Led98a, Led98b, LZ97, Lei97, LP96c, LP96d, LP96e, LPW96, LPM96, LP96a, LP96b, Lem97b, Lem97c, LP97b, LP97e, LP97c, LPMG97, Lem97f, LP97d, LP98c, Lem98, LC98, LC99, Lemxxa, Lemxxb, LPxxa, LPxxb, Leo95, Leo98, LPR97, Lev99, Lev96a, Lew97b, LB97a, Lew97a, Lew98, LL98, LBS98, LM99, LB98a, Li98, LB98b, LV99, Lia99, Liaxx, LSxx, LTxx, Lie95a, Lie95b, Lie95c, L^{+}xx, LLxx, Lin98a, LY97a, LYJ^{+}97, LY97b, LY99]. Java [LYxxa, LYxxb, LYxxc, Lin96a, Lin96b, Lin96d, Lin98b, Lip99, LS98, Lit97, LS97b, Loe99, Loe97, LQ99, Lon97, Lonxxa, Lor96a, Lot95, Lou95a, Lou95b, LM97c, LB96, LB97b, LCxx, Lut98a, Lut99c, Lut99d, Lut99e, Lut99b, LR97, Lyo99, Min96b, MvNV^{+}99, Mac98, MAS96, Macxx, MN96a, MN97, MD99, MGL^{+}98, MK98, Mac96, Mac97c, Mac97b, MST99, Mag95, Mah98, Mai98, Maixx, Man96a, Man96b, Man96c, Mar98a, MS99a, Marxxa, Mar98b, Mar99a, Mar97a, MK97, Mar97d, Mar97b, Mar97c, Mar98c, MLM98, MM99, Marxxb, MSS98, MTU99, Mas98a, Mas98b, Mas98c, Mas98d, Mas98e, Mas98f, Mat96b, Mat99, MNO^{+}96, MB96a, MB96b, MB97, Mau98, MF96a, McC96b, McC96c, McC96a, MD98a, MD98b, McC98a, MCxx, MG98, McC98b, MCD99a, McC95a, McC97b, McC98c, McC98d, McC98e, McC98g]. Java [McC98h, McC98i, McC98f, McC99d, McC99a, McC99b, McC99c, McD99b, McD98, MB98b, McF97, McF98a, McF98b, MO98, McG97, MLF97, MF96b, MF97b, MF97c, MF97a, MF98, MF99, McHxx, MO96, McK98, McL96, Mec97, MJ99, Mee96, Meh97, MS98a, Mel96, Men96, MRA98, Mes96, MS96, Mes97, Met98, Met95d, Met96, Mey96a, Mey96b, MD97a, Mey98a, Mey98b, Mey98c, Mey99, MDxx, Mic97a, MSS96, MS98b, MMS98, Mil98, MB98c, Mil95a, Min96a, Min97a, Min97b, Min97c, Minxxa, Minxxb, Minxxc, M^{+}xxa, Minxxd, Minxxe, MG97a, MG97b, MG96, Mit96b, Mit97a, MKT98, Miz99, Mof97, Moh96, MZ97, Mon97, MK99, MH97a, Moo98, MH97b, MMG98, MMGL99, Mor98a, Mor98b, Mor97a, Mor96a, MPxx, M^{+}96b, Mor96b, Mor97b, Mor97e, Mor97f, Mor97d, M^{+}97e, M^{+}97d, M^{+}97c, Morxxb]. Java [Morxxc, M^{+}xxb, M^{+}96c, Mos96, Mos98b, Mos98a, Mos98c, Muk99, Mül96, MMBC97, MS99b, Mun96, MH97c, MW98, MD97b, Mur99a, Mur99b, Mut97, Num98, NGM^{+}97, NMS96, N^{+}xxa, Nag96, Nap97, Nap98, NJ95, Nau96a, Nau96b, NS97b, NS97a, NS97c, Nau97, Nau98, NS98a, Nauxxa, Nauxxb, NAC99, Nav97, NRG^{+}99, Nel95a, Nel95b, NN97, Nel98, NH98, NHxx, NE98a, NE98b, Ner97, Net95b, Nev99, N^{+}96a, New96a, New96c, N^{+}96b, New96d, N^{+}97a, New97, New98, N^{+}xxb, Newxxa, Newxxb, Nghxx, Nic99, Nic96a, NFP96, Nic96c, Nic96d, N^{+}xxc, Nie95, NP96a, NP96b, NP96c, Nie97, NP97b, NP97c, NP97a, NPxx, NK98, Nil96, Nil98a, Nil98b, Nil97, N^{+}97b, N^{+}xxd, Nix98, Nol98, NF96, NS96, Not96, Nov98, Nov97, OH97a, OH97b, O’C99, O’C95b, O’C95a, O’C95c]. Java [O’C95d, O’C98, OT97, OS96, OW97a, Oak98, OW99, OWxx, OW97b, OK99, Oga99, OL98, Ols97, Orc98, OHE97a, OH97c, OH98, OHxxa, Ort99, Oua96, Oua97, Oue96a, Oue96b, Oue96c, Ove96, Ove97, Pur97, PH99a, PM96a, PK97, Par99, Par95c, Par95e, Par95f, Par95g, PSL^{+}98, PIPR97a, PIPR97b, PPIG98, PM96b, PM97b, PM97a, Pat99, PSS98, Pel96, Pel95, Pep97, PH99b, Per96a, Per96c, Per96b, Per97c, Pet97, PB98, PF97, Pew96, Pew97a, PS97, Pew97c, PP99, Pewxxa, Pewxxb, PSW^{+}98, PZ97, Phi96, PWxx, Phi99, Pic96a, Pic96b, Pic96c, Pic96d, Pic97, Piexx, Pir97, Pis99, Pit98a, Pit98b, Pit99, PS98, Pla96b, Pla96a, Pla97, PHM99, Pol98, Pon98, PK98, Por98, PV98, Com95g, PF96, Pot99, P^{+}98b, Pre99, Prexx, PJR^{+}98, PW97, PTB^{+}97, PTTA99]. Java [Pro96, Pry96, Pug99, Pur98, Pus99, Qia99, Que97, Rag97, Raj96, RWB97, RBF98, RMB^{+}98, RA97, Ram97, Ran97a, Ran97b, Ran99b, Ran99a, Rao98d, Rao98a, Rao98b, Rao98c, Rao99a, Rao99b, Rao99c, Rao99e, Rao99d, Raz98, Red99, Red97b, Red97c, Red98, Red96, Redxx, Ree97, Ree98, Ref99, Reg96, Rei97a, Rei97b, Rei99, Rem97, Renxx, RKCW98, Rey97, R^{+}xxb, RRxx, RS97, Ric96b, Ric96c, RM97, RMS97, Ric96d, Rie97b, RWWB96a, RWWB96b, RWWB96c, Rin97a, Rin97b, Rit95a, Rit95b, Rit96a, Rit96b, Rit96c, Rit97b, Rit97a, Ritxxb, Ritxxa, Ritxxd, Ritxxc, Ritxxf, Ritxxg, Ritxxe, Ritxxh, Ritxxi, Rit97d, Rob96, RH97b, RP98, RH98, RHE99, Rob99, RM99, Rob98b, RHND99, Rod96b, Rod97, Rodxxa, Rodxxb, Rodxxc, RCRBR97, Roh96, Rom99, Roo97, RGT96, Rou98, Row97, RCB98, Rus98, RSH97]. Java [RC98, Jav96c, Saa96, Sad97a, Sak98, Salxx, Sal97, Sam96a, Sam97, Sam96b, SEH98, SEH99a, SEH99b, San98b, San98c, San97b, Sar98, SHW98a, SHW98b, Sar99, SH99, SF96b, SF96a, SF97, SB97, Sav96, Sav97, SWJ97, Sav99, S^{+}96a, Sch97a, SST98, SB96, SBB^{+}97, Sch99a, Sch97b, SR97, Sch96a, Sch97c, SO97a, SO97b, Sch97d, Schxxa, Sch96b, Sch97e, Schxxb, Sch97g, Sch96d, SA97, SAxx, Schxxc, Schxxe, Schxxf, Schxxd, Sch97h, Sch98a, Sch99b, Sch97j, SLCM99, Sch97k, SH97, Sch98b, Schxxg, Sch97l, SS98a, SS98b, Sco96a, Sco96b, Sed98, SvG98, SMY99, Sem96, SF96c, Ses98, Sey96, Sha98a, Sha95, SvJ96, SS98c, Sha97b, SSxx, S^{+}97, SC97a, Shixxb, Shixxa, Shi99, Shixxc, Sho96a, SH96, Shu97a, Shu97b, SLO96, SL97a, SLOxx, SLxx, Sie98]. Java [Sim96a, Sim96b, Sim97a, Sim97b, Sim97c, Sim97e, Sim97d, SBL96, Simxx, Sim96f, SMC^{+}96, SC97b, Sin97a, Sin97b, Sinxx, SN98a, SZTR96, SN98b, Sin96a, Sip98, SMC97, SWM^{+}97, Sla97, Sli97, SG97a, SGB97, Smi97, Smixxa, Smixxc, Smixxb, SA98a, Rapxx, Som97, Soo98a, SHGD98a, SHGD98b, SL99, SRD98, SD99, Spi99, Spi96, Spy95, Sri97a, SJRK97, Sri96, Sta98a, Staxxa, SD96, SZ98, SZ99, Sta98b, Sta96, Staxxb, SA98b, SA99, Ste97a, Ste97b, Ste97c, Ste96a, Ste98a, SPV99, SKN98, Ste96c, Ste96d, Ste96e, Ste97f, Ste99, Ste97j, Ste97h, Ste97i, Ste98b, Stexxb, Stexxa, SLC99, SP97, S^{+}96b, Sto96, SL97b, Stu98, Suaxx, SBxx, Sul97, Sul95a, Sul95b, SS96, Sun95b, Sun95a, SUN97b, Sun97a, JCV99, Sun99b, Sun99a, Sunxxa, Sunxxb, Sunxxc, Sun97c, Suš97, Sut97]. Java [Swa97b, Swa97a, Swa97c, Swa97d, Swa98a, Swa97f, ST96, Syd96a, Syd96b, Sym95, Sym99, Tai98, T^{+}97a, TMN^{+}98a, TMN^{+}98b, Takxx, TRMR97, TSHH96, T^{+}97b, TBS97, TD97, TA97, TK98, TK99b, Tei98, Ten95b, TWP98, Thaxx, Thi99, TKH99, Tho99, Tho95, TPHB96, TPHBxxa, TPHBxxb, Thoxxa, T^{+}xxa, Thoxxc, Thoxxb, Tho98, Tho97b, Til97a, Til97b, TLSS99, Tis97, TG95, TG96b, Tit96a, Tit96b, Tit96c, TB97, Titxxa, Titxxb, TBxx, TGxxa, Tob99, Tol97a, Tol98, Tom96, Top98, Top99, TJ96, Tow98, Tra97, Tra95a, Tra95b, TO96, Tre98, Tre99a, Tre99b, TMP^{+}99, Tro97, TMH99, TMM^{+}98, TMMH99, Tub95, Tuc98a, Tuc96, Tuc98b, T^{+}xxb, Tun98, Tur97, T^{+}98a, T^{+}98b, TW96, TC95, TTD96b, TTD96a, Tym96, TTD97, Tym98, Tým99, TTDxx, T^{+}xxc, TBA98, Ubo96a, Ubo96b, Ude97d, Ude97b, Uim95]. Java [Ung97, UBF^{+}98, Urg96a, Urg96b, Urgxx, Urq97b, Urq97a, Uzg96, Val95, dVGxx, Val96b, Van97g, VO98, Van97j, Van96f, V^{+}96, V^{+}97, Van96g, VPH^{+}96, VPH97, Van97m, Van97n, VPHS97, Vanxxe, VA98, Var97b, VN96, Vck96, Ven96, Ven97a, Ven97b, Ven99, Ver96, Ver97, Ver98, Vij97a, VRG98a, VRG98b, VR99a, Vil99, Vog96, VD97, VD98, VR99b, VP98, VCCH99, VN99, WL96, Wag96, Wag97a, Wag97b, Wak98a, Wak98b, Wak99, Wal96b, Wal96e, Wal96d, Wal96f, Wal96c, Wal96g, Wal96h, WW97, Wal98a, WH99, WR99a, Wal96j, WG97a, WBDF97, Walxxa, Wal96k, WD^{+}96, Wal97b, Wal97c, Walxxb, Wal96m, Wal96l, Wal96n, Wal96o, Wal96p, Wal97f, Wal97h, Wal97g, Wal98c, Wal98d, Walxxc, Walxxe, Walxxd, Walxxf, Walxxg, Wal97i, WYxx, WCL97, Wat97b, Wat98, Wat99, WB99, Way96b]. Java [Way96e, Way96g, Way96f, Way97a, Way97b, Way98a, Way98c, WJ97, WJ98, WB^{+}96a, WBC^{+}97, WRTM98, Webxx, Web97a, Web97b, WB^{+}96b, Weh96, Weh98a, Wei97c, Wei98, WdFD^{+}98, WDDF98, WB98, WS99, Wel99a, WOC99, Wel99b, WM97a, Wel99c, WO97, WSK97, Wes98, WR99b, WBC^{+}96, Whe97, Whi96b, Whi96c, Whi96a, Whi97a, W^{+}97b, WCF98, WFC^{+}99, WN97, Why96, WL97a, Wie96, Wie97, WL97b, Wijxx, Wil96a, Wil96b, Wil97a, Wil96f, Wil96e, WM97b, Wil97b, Wil99a, Wil96g, Wil96h, Wil97f, Wil97c, Wil97d, Wil98b, Wil99b, Wil99c, WR98b, WRxx, WL95, Win95e, Win95l, WN96, Win96, WN98, Win98, Wis97, Wit97, Wol97, Wol99, WRW96a, WRW96b, WWR97a, WWR97b, WRW96c, Wolxxa, Wolxxb, WK99a, WK99b, Won97, Won98, WPM99, WRB^{+}97, Woo99a, Woo99b, WL97d, WJUH98, WAES97a, WAES97b, Wu99, WCS99]. Java [Wuxx, WBB^{+}97, WB^{+}96c, Xu97, YC98a, YC98b, Yan96, YLP^{+}99, YB99, Yanxx, Yat97, Yeexxa, YSP^{+}98a, YSP^{+}98b, Yel99, Yel95, YK99, tYpCxx, YHW98, Yos97, You97b, You97c, You96a, You96b, You97d, You98b, YC97, YRE98, ZK98a, ZK98b, Zan96, Zei97, Zho97, ZTLV98a, ZTLV98b, Ziexx, Zif96, Zif97, ZMN98, ZMN99, ZKM99, ZOxx, ZJ96, Zuk97b, Zuk98, Zuk99, deC99, pP98, Eud97b, vD99a, vHO98, van95, vSS96b, vSS96c, vSS96d, van96b, vSS96e, vSS97, van98a, vanxxd, v^{+}xxa, v^{+}xxb, van96a, van97d, van97b, van97c, van97a, van97e, van99a, van99b, vanxxa, vanxxb, vanxxc, von98, Knu97c, Cul97c, Ano97g, SMC97, Ano96n, Ano98f, Ano98k, McC96e, Nic96b, vD99b, vSS96a, Ano96o, Ano97j, Ano97i, Ano97k, Lea98b]. Java- [LPR97]. JAVA-1.1-Developer-Kits [SBH97]. Java-appleta [Gol96a]. Java-Applets [BM96a, vSS96d, Ek97a, Ekxxa]. JAVA-based [ERKE96, KP99, Ano96-72, Ano96-65, Ano96-192, Bar97, FF96b, Ano97-114, Ano98-30, IG99, Fue98, Fün98, Hor96a, Hor96b, Mos98b, Mos98a, PSW^{+}98, PTTA99, Sar99, WAES97a, WAES97b, ZMN98, AFH98a, AFH98b, Ale98, AISS97, Ano96-47, Ano97b, Ano97c, Ano97-72, Ano97-135, Ano97-157, Ano98d, Ano98s, Bin97, Cox95c, DAB97, GH97, GRN99, Gre98d, JJG98, Kro98a, Ref99, RHND99, Ung97, WOC99, WPM99, WCS99, YC98b, ZMN99]. Java-Centric [WWR97a, WWR97b]. Java-compiled [LP96a]. Java-Einführung [Dec96b, Decxxb]. Java-enable [Gre98b]. Java-enabled [Col96a]. Java-Fibel [Kühxx]. Java-fueled [Ano96-82]. Java-inspired [CM97a]. Java-Kompendium [Herxx]. Java-MPI [GGS99]. Java-Object [BCC^{+}97]. Java-ohjelmointi [Kiuxx, Ekxxb, Jamxxb, KNxx]. Java-Plattform [Ano97-81]. Java-Powered [Ano97-82]. Java-Processes [KMW99]. Java-programmering [Ekxxc, EExx, Ekxxd]. Java-Programmierhandbuch [FP96]. Java-Programmierung [BBB^{+}96, W^{+}97c, Bacxxb, FKP97, FPxx, KK97f, KWxx, Krü97b, Krüxxb, Sch97d, Schxxa]. Java-Referenz [Sta96]. Java-RMI [BRF98, RMB^{+}98]. Java-Sammlung [Ano96-67]. Java-Script [Nov98]. Java-SIG [Couxx]. Java-SIG’s [Cou97f]. Java-speak [Hay96l]. Java-supported [Reg96].

Java-Targeted [CR99]. Java-to-RDBMS [Ano96-109]. Java-to-Relational [Ano97-86]. Java-Tutorial [CWxxa, Her97]. Java-Zauberei [KS96b]. Java. [ER97]. java.* [All99a, All99b, All99c, All99d, All99e, All99f]. java.applet [FB96, FBxx, CL98a]. java.awt [FB96, FBxx, N^{+}96a, CL98a]. java.beans [CL98a]. java.io [CLK98, CLK99a]. java.lang [CLK98, CLK99a]. java.math [CLK98, CLK99a]. java.net [CLK98, CLK99a]. java.text [CLK98, CLK99a]. java.util [CLK98, CLK99a]. Java/AWT [Culxxb]. Java/C [Che98]. Java/DSM [YC97]. Java/EJB [Cha99f]. Java/Hot [Des96, Dre96]. Java/HotJava [Ano97-104, Ano97-167, Anoxx-60, FLW96]. Java/HotJava/und [Dre96]. Java/HPC [BDV^{+}98a, BDV^{+}98b]. Java/Internet [Ano97-53]. Java/intranet [RGT96]. Java/JavaScript [R^{+}xxb]. Java/mainframe [Hib97]. Java/RMI [Cha98d, Cha98e]. Java/Smart [Bro97c]. Java/VRML [Eng97a, L^{+}xx]. Java/Web [Ano97-106]. Java3D [DS97a]. javab [BG98b, BG98c]. JavaBean [Ber97f, Kro97, Kin99]. JavaBeans [AH97, Ano96-113, Ano96-181, Ano97-91, Ano97-92, Ano97-89, Ano97-90, Ano97-110, Ano98t, Ano98z, Anoxxe, Anoxx-54, Anoxx-55, Bro97a, Broxxa, Broxxb, Cha97b, Chaxxd, Chu98a, CDF^{+}98, C^{+}97a, Cox97, EMe97c, Eng97d, Eng98, Engxx, Fegxx, Gau96h, Gil98c, HGxx, Har98a, Hoq98, HS98a, Hoqxx, Hun98, JT97, JTxx, Kas97a, Kro98d, Kro98e, Kro98c, Lut98b, MS97, Moh98, MH99, Mor97g, Morxxd, Nic97, Nicxx, O’N98b, OH99, PCxx, Red97c, Ren98, Rod98, Rom99, Sri97b, Val99, Van97o, Vanxxg, VR99b, Wat97a, Way98b, Wea97b]. JavaBeans-Technik [Morxxd]. Javabeans^{T M} [Ano99a]. JavaCard [BBE^{+}99]. JavaCC [SW99]. JavaChip [Ano97-149]. JavaDev [DiG95c]. javadoc [Ait96a]. JavaEngine [Ano97-150, Mor97c]. JavaFit [Rob98a]. JavaHot [Ros96]. Javaing [PV99]. JavaLanche [SP98]. JavaMan [Ano97-28]. JavaManagement [Ano96-114]. JavaOne [Ano97-107, Kro99d, Sta97b, San97a]. Javaone-97 [San97a]. JavaOS [Ano96-101, Ano96-115, Ano97-126, MC97, Mit96a, SM99]. JavaParty [PZ97]. JavaPorts [GM99]. javar [BVG97, BVG98]. Java’s [Cla97b]. JavaScript [Ammxx, KK97b, Llo98, RNBV98, Sad98, SSxx, Tee99, Ver98, Ait96c, Ait97b, Ait97a, Al-97, And98a, Ano95i, Ano97-97, Ano97-66, Ano97-93, Ano97-94, Ano97-95, Ano97-96, Anoxxl, Anoxx-59, Anoxx-56, Anoxx-57, Anoxx-58, Anoxxo, AKM99, AL96, Bal96c, BT99, BLB99, Ben99c, Ben99b, BB96, BB97c, Bra99, Bre96, Buc97, Cag96, Car97a, Chaxxa, Chuxx, Coh96a, Coh97a, Cor97a, Cue96, D^{+}98b, DSM99, Dan96e, DT96, Dan96b, Dan96f, Dan96d, Dan97b, Dan97a, Dan97c, Danxx, DA96, Dar96, Die96, Diexx, DF97, Drexx, Dun97, Eis97, EE97, FCDW96, Fea97, FCD97, Fet97, FF96a, Fis98b, Fla96d, Fla96e, Fla97g, Fla97h, Fla98e, Fla98d, Fre97b, Fri98, Gau96c, GTW99, Gil99, GJ96d, Goo96a, Goo96b, Goo96c, Goo97c, Gooxxb, Hay96o, Hei96, Hei97, HL97a, Hus96, HL97b, HL97c, HL97d]. JavaScript [HK99, JLW96, Jaw97a, Joh96b, JOA98, Kan97, Kas97b, KK96a, KKxxa, Kim97e, Kolxx, KDH96, LO^{+}97b, LO98a, LL97, LM96, LS96, LM97a, LM97b, Lemxxc, Manxxa, McC95b, McC96d, McCxxa, McCxxb, MHP98, Min96c, Mon96, Mon99, Monxx, Morxxa, Mor96a, NS98b, NR96, Neo96a, Neo96b, Neo97, NR97, Neo98, Neo99, NKxx, Par95h, PM97c, PPJ96, PM97e, Purxx, RVM96, RVM97, RDD^{+}96, RWxx, R^{+}xxb, Rit97c, Rod96a, Sad96, Sad97b, S^{+}96a, SR97, Sha96a, Sha98b, SS97, SR96, Ste97e, Tea97, Tee99, Tes96a, Tes96b, Tic97, Tor96, TPW96, Ude97a, Vac97, Van96e, Van97i, Van97h, Van97k, Van97l, Van98b, Var95, WC96, WW96a, WW96b, WS98, Wat95b, Way97a, Wea97a, Wei97a, Wei97b, Wes97a, WF99, WMDH97, WB^{+}96c, Yeexxb, Anoxx-33, CC98, Coh97d, Col96b, D^{+}98a, Dan96c, Dew97, GS96]. Javascript [Hal97b, Hoq97, Jac96, LO97a, Man96d, MP98, MP99, Net95a, Pur96, Ritxxb, Vac96, W^{+}97a, WF97, Wod96, O’N98a, Ano96p, Ano98g]. JavaScript-programmering [EE97]. JavaScript. [Koc97]. JavaServer [Tre99b]. JavaSoccer [Bal98b]. JavaSoft [Ano96-173, Ano98u, Gau96d, GJS97, Kee96, Pic96b, Web97c, Web97d]. JavaSpace [AFHxx]. JavaSpaces [FHA99]. JavaStation [Ano97a, Ano97-98, Ano97-99, Car97b, Ano96-116]. JavaStations [Raz98]. JavaStudio [WR98a]. Javatalk [Gre97a, Har97b, Knu97b, Knu97c, Knu97d, Knu97f, Knu97e, Knu98a, Sto97, Gre97b, Gre97c, Gre97e, Gre97g, Gre97f, Gre97h]. Java^{T M} [BGH99, Dar98c, GS98b, Ano98j, Ano98i, Ano98h, Ano99c, Ano99b]. JavaVM [KG97b]. JavaWorld [Bee98]. Javelin [CCI^{+}97]. JavuNetwork [Kam99]. Jaz [Ano96-40]. jazyce [F^{+}xxb]. Jbed [TMH99]. JBuilder [Cha97a, Ano98q, Ano98v, Arm97, Bur97, Cog97, Doh98, Dun96, Espxx, GTFxx, Hol97f, JSA98, Joh98, KGTF97, Kiu98, Man96e, Man97a, Man97b, MBxx, M^{+}96a, MP97b, Rie96, TD97, Tayxx, Vanxxf, Zuk97a]. JCAT [BR97]. JClass [Ano96-137, Ano97-118, Ano96-117, Ano97-100, Kro97]. Jcon [PTTA99]. jContractor [KHB99b]. JDBC [Ano97g, Ano98-27, Hob97a, Hobxx, Joh98, Ano98w, Ano98t, CW96b, Cor98b, FB98, Ham96a, HCF97a, HCF97b, Hamxxa, Hamxxb, Hob97b, Hob97c, Hob98a, JRC97, Key98, Klu97, Klu98, Kluxx, Kon98, MT97, PM96b, PM97b, PM97a, P^{+}98a, Ree97, Ree98, Sip98, Soo98b, Tay97, Tay99, Van97g, Weh98b, WFC^{+}99, WM97b, YLQ98]. JDK [BD98b, Klu98, SBH97, Til97c, Ano97d, Ano98u, Anoxx-61, Anoxx-62, Bro99a, Bro99b, CWHt99, Cha98a, CL98b, Dav97c, For98d, Gil98a, HL99, KG98a, MBH97, Nghxx, Shu97b, VPH^{+}96]. JDK-Version [Til97c]. JDK1.1 [Anoxx-44]. JDKs [Gla97b]. JDuck [GG99]. Jeans [Ano96u]. Jeeves [Har97b]. JEM [Cha98d, Cha98e]. JEM-DOOS [Cha98d, Cha98e]. Jersey [IEE96d]. Jet [Gre97b, Red98, PSS97]. Jet-assisted [Gre97b]. JetAssist [Ano97-27]. jetzt [Dav96b, LM97c]. Jewel [Ano97-49]. Jewelbox [Golxx]. Jewels [Ano98i, Deu98, Due97b]. JFactory [Gre96c, Met96]. JFC [Ano99d, Ano99e, Ano99f, GM98a, Gea99, Joh99, JV97, JV98, Mic97b, Nel98, Red97a, Top98, WG99, WC98, WA98]. JFlow [Mye99]. JICC [Ber98h]. JIGGLE [Tun98]. Jini [AWO^{+}99, Edw99, Mat99, Swa98b, Wal99, Won99]. JIT [BGH99]. Jitters [Lou95a]. Jive [Ano96-187, BAC^{+}99, Ste96d, Ste96e, Ste99, Hin96, Mel96]. JLAPACK [DD97i, DDS99]. JLex [BD98a]. JMAS [IG99]. JMF [GT99]. JN [Mon97]. JNI [Coo98, GM98d]. Jockey [Fer95, Hin97]. Jockeys [DS95b]. Joe [Ano96-48, Ano96-118, LH96, LB97a, Lin96a, Red97b]. John [Ano97i, Heu97, Ver98, Ano96-102, Zuk99]. Join [Har96a, McK97, Mil95b, Sil95]. Joins [Com95a, Law98, Ano97v]. Joint [IEE96f]. Joke [Wil99d]. Jolt [Anoxx-63, Car95, DiD97d, GJ96a, Gof97, Gos96, Nel95b, Vij97a, van96b, Wal97d, Wal97e]. jolts [Hay96i]. Jon [Ano97k]. Jonathan [Ano98h, Ano99b]. Jordao [dN97]. Jos [Ano98-27]. Jose [KJR98, Uni96a, Uni96b, Uni97b, Uni98b]. Joshua [Ano96n, Ano98f, SMC97]. Journal [CC96, Pur97, JAV96b, Eri96a, OW95]. Journals [Ano96-148]. Journey [Meh97]. JPDA [Sun99a]. Jperl [Bal99]. JPT [BDY99]. JPVM [Fer98b, Fer98a]. JPython [Ang99]. Jr [ACM99c]. JRes [CvE98]. JSafe [Way98c]. JScape [Ano97-151, Ano97-152]. JScript [But99, RH97a]. JSDK [Klu98]. JSpace [Led98a]. JSPICE [SHGD98a, SHGD98b]. JSQL [McK97]. JTAPI [Rob99]. JTC [Gar97a]. JTED [Cow98a, Cow98b]. JTN [CR99]. JTS [VR99b]. Judge [GC97]. Juices [Ano96e, Ano96-96, Boo96b]. jukebox [Val96b]. July [Doo99, PZ96, SMEN96, DF96, Ass96b, USE97a]. Jump [Ano98m, Uim95, Gau96e, Win96]. Jump-Start [Ano98m]. jumpers [Gau97g]. Jumping [All99b, Syd96b, WF97, WF99]. Jumps [Ano95c, Gar97c, Par95c]. Jumpshot [ZLGS99]. Jumpstart [Ene97]. June [Ano96-158, Ano98a, BBJR96, FGL96, IEE96q, IEE96m, LM98, Str96, Ass96a, USE96c, USE97b, USE98c, USE98e]. Jury [Sch97b]. Just [And98b, Ano96-141, CBS95, CL97e, CFM^{+}97a, CFM^{+}97b, LP96a, Ste97j, Sto96, Str97, WC99, van96a, van97d, van97b, van97c, van99a, van99b, vanxxa, vanxxb, ATCL^{+}98, Ano96-102, Ben99b, Gre97g, KG97b, Mof97, PC97, Sem96, SF96c, Tre98, Ubo96a, Wal96h, Way96c, Nic96b]. Just-In-Time [Ano96-141, CL97e, ATCL^{+}98, KG97b]. JustEdit [MD99]. Justifying [Cul98b]. JVM [Rao98d]. JVMPI [Sun95a]. JVX [MK98]. JWarp [BSS98a, BSS98b]. JWAVE [Ano97-163].

Kaffeehaus [Beh95]. Kaffeemaschine [DE96a]. Kaffeemaschinen [KS97]. Kahan [Woe97]. KAI [Ano98-32, Ano98-33]. Kalakota [KW96]. Kanji [RC98, RC98]. Kannich [Jas96b]. kano [IK96]. Kasten [Gol96b, Golxx]. Kava [WS99]. Kawa [BM96c, Bot98a, MB96c]. Keep [Nan96, Hay96w, LP96a]. Keeping [Ano96-119, Ano96-183, Bay97a, Rei96]. Keeps [Ano96-38, Cou97e]. kernel [Dav97a, VCCH99]. kernels [GLC99]. Key [Fox97c, Gau97n, Gil97a, Hin96, HM98, Ano97-166, Ren98, You98b]. Keyboard [BAC^{+}99]. Kickass [Groxxb, Anoxx-64, Esp96]. Kicking [Ste97e]. Killer [Ano96a, DDJ99b, Ano96-97, Hal97a]. Kills [Ano96-172]. Kind [BCP^{+}99b, Wil99e]. King [ACM99c]. Kingdom [Ber98h]. kinks [Mes96]. Kiosk [Rey96]. Kit [Ano96-83, Ano96-104, Ano96-103, Ano96-135, Anoxx-52, Fox99, Gol96b, GMPS97, IR97, Infxxa, JDK98a, JDK98b, Kro99c, Man96c, Per97b, Sam96a, Sam96b, TJ96, Van96f, Ano96-105, Ano97-132, DiD97c, Gau96d, Hor96c, Klu98, McC98c, Mey98c, WBB^{+}97, Ano96-40, Dan96a, Gre97j]. Kits [SBH97]. KJPrice [PJR^{+}98]. KL [Kro97, Kro98d, Kro98c, Kro99e]. Klassen [Sta96]. klatch [Hin96]. Know [Shixxc, S^{+}96b, GHP^{+}96, Mee96, NMS96]. Knowing [Way96b]. Knowledge [Eri96b, Lut98a]. Knudsen [Ano98h, Ano99b]. Kodak [Vad95a]. Kompendium [Her97, Herxx, Ste98b, Stexxb]. Komplette [Sta96, Dav96b]. Kong [Tri95]. konkrete [Vac97]. Konzept [Dal97]. Konzepte [Mor97e, Mül96]. Korean [Lun99, Kanxxc, N^{+}xxa]. Krakatoa [Ano96-193, PW97]. kreativ [Hie97]. kurssi [Kiuxx]. kurze [DE96b].

L [Bay97a]. Lab [Ano96-204, Bar96, Cox97, DH97, Manxxb, Mor97f, Morxxe, Web98]. Labo [Lemxxc]. Labor [Web98]. Laboratorio [FLRT97]. Laboratory [DF97]. labs [CS99, Ano96-74, Gro96b, Ven96]. Labview [Chw98]. lack [Ano96-151]. lacks [HG96a]. Lagunita [WBC^{+}96]. Laid [Car96d]. Lake [Isk96]. Lambda [WJUH98]. Lambda-Calculus [WJUH98]. Lame [MF97d, Kay97]. LAN [S^{+}96b]. Landscape [Sch96c]. Landwehr [Par95b]. Lane [DDJ98b, DDJ98c]. Langage [AG96e, Drixx]. langages [Salxx]. Language [ACM99b, Ano96-63, AG96d, AG98, AGxxc, AGxxd, BBJR96, Bor95, BOSW98, BL97, Bro98b, Buc97, CBT97c, CW98, CCFL98, CS98, Cha96d, Che97, Com95d, Cul96a, Doe95, Eck96c, FM98, FP98, GH96a, GM96, GJS96, G^{+}xxh, G^{+}xxi, Gra97a, Gra97b, GS98c, GHP^{+}96, Gro96a, Ham97a, HS96, HHK98, Hos96b, LS99, Lea96c, LST99, Lor96a, McC97c, Minxxd, NRG^{+}99, Net95b, Net95a, Nghxx, NFP96, O’B96a, Par97, Par99, PF96, PJR^{+}98, Roo97, SMY99, Spe98, SUN97b, TPW96, Tra95a, Tra95b, Uzg96, Wal96d, WYxx, Yeexxa, Ziexx, ZJ96, AFM97, Ano95i, Ano95l, Ano95r, Ano96-54, Ano96-74, Ano96-59, Ano96-102, Ano96-188, Beh95, Bös98, Bot98b, BM98c, D’S97c, EM99, FF97, FM99a, FM99b, Fri98, GH96b, GGMS99, GJ96d, Gos96, GS98d, Gre98a, GLV99, Had98b, Hal97a, Hay96k, Hos96a]. language [Joh96a, Kar95b, KW99, Man96a, Mof97, Mur96b, NMS96, Orc98, Com95g, Pro96, Sem96, Sim96f, Ste96a, Tai98, TTD96b, Ubo96a, Ven96, Wal96h, WS98, WB99, Why96, van96b, vRvGS97, Bru98, Yal98]. language-independent [D’S97c]. Languages [ACM97a, ACM98b, Bot98a, Fre97a, MDM97, Sal97, Tol98, USE99a, ACM99a, Ano96-61, Ano96-171, Bla99, Bru96, DDG^{+}96, GA98, RBL^{+}97, RFM95, Tho97a, Tol97a, Tol97b, VS96]. LAPACK [DD97i, DDS99]. Large [Ano97-35, AOV^{+}99, GM98b, Gol99, JPK98a, JPK98b, KT99, Lak96, Fox98b, LR96, Mat96b, MC97]. Large-Scale [JPK98a, JPK98b, Lak96, MC97]. Larry [DDJ99b]. Laser [MD99]. last [Ano96-73, Joh97a]. Late [BCW97, JMM96, RCRBR97, TMP^{+}99, WB^{+}96b, Zif97, PC97]. Latest [Ano96-43, Gil97c]. Latté [Hol96b]. Latter [AL97]. Lättpocket [Ekxxd]. Launch [Cox95c, Mes95c]. Launched [Mes95b]. Launches [Ano95q, Ano96-173, Ano97-106, Par95d, Par95h, STV95, Ano97-148]. Laura [LM96, LS96, LPMG97]. Layered [Lut99d]. Layout [Bec98, Ber96d, BDKK96, Tun98, Bea96]. Lazy [Wak98b, McM96, MJ99, Wak99]. Lead [Win95j]. leaders [Dei96a]. leading [McL96]. leads [Chw98, GME99]. Leaks [DS99]. Learn [BM96b, Dav96c, Jam96, MO96, Ram97, SS97, WB^{+}96c, Wil96a]. Learning [BBJR96, Kim97a, Kimxx, Min96a, Pit97, Ver98, Ano96-201, GJ96d, GBS98, HRW^{+}96, JPT^{+}99, DF96, Ano97-101]. learnt [Roh96]. Least [Jam96, Rob98a]. least-squares [Rob98a]. Leave [Met95a, Ano96-78]. Leben [Dre96, Bat96d]. Lecture [Han97]. Lee [Ham96c]. Leeds [Abr96]. leert [Holxxb, Holxxc]. left [HN96]. leg [Ano96-133]. Legacy [Kro98e, McG97, Che97, Var96]. Leibniz [Gop96]. Leitfaden [BBB^{+}96]. Lemay [LM96, LM97b, LPMG97]. Lemays [LM97a]. Lemmens [Ano98-27]. Lenguaje [Abexxa, AGxxb, C^{+}97c, Ferxx, H^{+}xxc, LCxx, C^{+}xxb]. Lenguajes [Abexxb]. lens [KT96]. lerne [LM97c]. lernen [Krü97a, Krüxxa, Frö96, Fröxxb, Wes97b]. Lernpaket [Dav96b]. Less [Gar97b]. lessons [Ric96a, Roh96]. Let [DDJ98b, DDJ98c, Fuc96, Gil96c, Mac97c, Mac97b, Ano96-123, Cou96a]. lets [Ano96-181, Gau97r, MC97, Oue96c, Tur97]. Letter [Bek97, BL97, Cra97, Dav97b, Fed97, Gib95, Hen97b, KL97, McC97c, MF97d, Mon96, MH97b, Nav97, Rag97, Raj96, WN97, Wil96h, WL97d]. Letters [Ano96-121, BAC^{+}99, BCP^{+}99b, CL96, Coo98, HHW^{+}99, MGL^{+}98, Mar98a, MH97a, NRG^{+}99, Ols97, PJR^{+}98, SL97b, Tei98, TMP^{+}99, Wel97, WL97a]. Level [Mas98a, Mas98b, Mas98c, Mas98d, Mas98e, Mas98f, Yel95, BG97c, GGS99, KCPC98, Nan95, Rit97d, Sil95, SvK98]. Levenberg [Rob98a]. Lexer [McK98]. Lexical [BD98a, Doe95]. Lexical-semantic [Doe95]. Liao [Flaxxb]. Liberalisation [Ano98l]. Libero [iMa96]. Librarian [Sch96d]. Libraries [Add99, Ano96-137, Cha96b, CL97b, CL97a, CLK98, CL98a, CL98b, CL99, CLK99a, CLK99b, CLxxa, CLxxb, CLxxc, EMe97i, GFHM98, MG97a, San97b, BDP^{+}98a, BDP^{+}98b, CDD97, GHM98, HG96a, PCWGM98, SG97a, Whi96c, Won97]. Library [Ano96-41, Ano96-135, Ano96-146, Ano97-119, Ano97-125, Ano97-133, Ano98y, Ano98-32, Ano98-33, BSS98a, Bro96, CS99, FGN^{+}97, Gar96f, KHB99a, KHB99b, Kro98d, Kro98c, Kro98b, LJ96b, LJ96a, Mic97a, PSTZ99, SB96, ZOxx, Ano97-44, BSS98b, PL96, RP98, SGB97, Ste97d, Tab96, Ano97-164, Gla97d]. Libreria [Anoxx-45, LJxxb]. Libro [F^{+}xxa, Kolxx]. License [Ano96-48, Gre95, Gut95d, Gut95f, Gut95i, Mil95a, Net95b, Par95e, GH96b, JN95a]. Licensees [Com95a]. Licenses [Ano97-133, Spy95, Ano96-195, IBM95, Lie95c]. Licensing [Ten95b, O’C95a]. Lid [Par95f, PZ96]. Life [DDJ98b, DDJ98c, Lut99e, Bat96c, Gau97r, Sco96a]. Lifecycle [Kro99a, Kro99b]. lifestyle [Chw98]. Lifetimes [OK99]. Lift [Gar97a, Ban95b]. Light [Gar97c, KS98a, KS98b, Lav96, Lev96d]. Light-Weight [KS98a, KS98b]. lightens [Lin96c]. Lightweight [PH99b, YLP^{+}99, Ano95o]. like [Ano95h, Ano96-54, Hay96j, OH97b, WB^{+}96c, pP98]. likely [Cha97c, FC96a]. limitations [Fri98, Ude97d]. Limited [Kro98b]. Linda [Led98a]. Linden [Nic96b, Str97]. Line [Ano96-134, BH99a, Bur96a, Har96c, Whi96a, DS95b, Gau97q, SvK98, Vad95b].

Linear [Loe97]. lingua [Far96a, GG96a]. Linguaggio [GGxxa]. Linguistic [KMS98]. LINK [Ano96-137, Ano96-28, Var96, Ano96-125, Ano96b, Gau96a, Gau97e, Gau97i, Hof97, Ste97b]. Linked [Boy98b, FC98b, GJ96d]. Linking [PIPR97a, PIPR97b, FC98b, Way96b]. Links [Ano97-79, Hib97, Hor96a, Hor96b, Mes97, Oue96b]. LinkScan [Ano97-118]. Linux [Ano96-135, Ano96-137, Ano97-118, Ano97-121, Ano98-27, Ano98-28, Kuk98, Ano97-121, Cha98a, DDJ98a, Dal98, Kro99a, Kro99b, Oua97]. Linux/AXP [Ano97-121]. LinuxCAD [Kuk98]. Liquid [Ano96-30, Ano96-135]. Lisp [SFSS96, Ste96a]. List [Wel99b, Gau97m]. Listen [CRS98a, Hay96p, MO96, CRS98b]. Listening [Lut99a, Lut99b]. Lists [Boy98b]. Lite [Ano95o]. little [Ano98e, FF98, Hay97a, MMJ96]. Live [Lev96a, Min96a]. Liveness [ADM98]. Lives [Gre97e]. livestock [Ano98o]. LiveTable [Ano96-137]. LiveWire [CCB96b]. Living [Bri98, Sch97f, Rob96]. Livre [Lonxxa, Marxxb, Hac96]. lizenzierter [BKP97b]. LL [PQ95, PQ96]. LLC [Ano98-28, Ano98-28]. Load [Kro98b, Kro99a, Kro99b, Lin96c]. Load-Testing [Kro98b]. Loadable [Ora95]. Loading [Gol98a, LB98b]. lobbying [Ner97]. Local [ADM98]. Locale [O’C98]. Locks [BKMS98]. Logic [Ano96d, PZ96, PHM99, HVB^{+}99, Ano97-102]. London [Ano96-142, Ano96-154, Ano97-54, Cul97c]. Look [Ano96-39, Ano96-37, Ano98m, Sun99b, Ano96-188, Gau97r, Gre98b, Hay96j, McL96]. Lookahead [PQ96]. Looking [Mel96, Ano96-59]. Looks [OH97a, Win95d, Ano96-40]. Looms [Sim96e]. Loop [BVG98]. loosens [Wag97a]. Loses [Law98, Gau97n]. lot [Joh97a, Sem96, SF96c, vSS96e]. Lotus [Ano98-31, Cav97, DS95a, DS95b, Gil98c, Gut95d, H^{+}97, Oue95a, Oue96c]. LotusScript [H^{+}97]. Louis [ACM97b]. Love [Kim97d, Kim97f, Ham97a]. Low [Ano97-103, GGS99, Gut96, Kro98b, MD99, Yel95, Ano96-53]. Low-Cost [Kro98b, Ano97-103, Gut96, Ano96-53]. Low-level [GGS99]. lowcost [Way96g]. LR [PQ96]. Ltd [Ano98-27]. Lucent [Mac97a, Mac97d, Mac97e]. Lunch [DDJ98b, DDJ98c]. lure [Gau97k]. Lures [Win95l]. Luther [ACM99c].

M [Zol96]. MA [Ano98e, Ano98k, USE97a]. Mac [Gla97b, Ano95w, Ano96-195, Ano96-196, Ano96-42, Ano96-122, Ano97-104, Ess95, Kro99e, NR96]. Mach [MKT98]. Machine [Dalxx, Dal97, LY97b, LY99, LYxxb, LYxxc, MD97a, MDxx, Qia99, Sun95b, Sun95a, SUN97b, JCV99, Ven97a, VN99, WL96, Wak99, WH99, YK99, Ano96-192, Bod96, DLM99, FF96b, Ven97b, Won97, Yel99, Ano97q, Ano97-48, Ano97-87, AC98, BS98c, BS98d, Cla97a, Coh97b, CDG97, Eng99, Fra98b, HT98, LP96a, LM99, LB98b, LV99, LY97a, MB98c, Mon97, OT97, Tai98, Tol98, TO96, Ven99, Wak98a, Wak98b, Wol99, Ano97k]. Machine-Checking [VN99]. machines [BDF^{+}98, EL98, Hin97, RHND99, WWR97a, WWR97b, YC98b, ADM98, Cla97a, UBF^{+}98, YC98a]. MacIntosh [LPW96, Neo97, Ano96-104, BM96b, Dow97, Met95d, PM97c, SR96]. Macmillan [New96c]. macro [Cou97c, Gop96]. Macromedia [Ano96-177, Mil95b, Win95g]. Made [Wat95e, PM97d]. MagaPoche [Drexx]. Magazine [Bee98, Har96e, Weh96, Jav96d, Fra96a]. Magic [SC97b, Sin97a, Gil96b, Gut95h, Kro97, Sin97b, Sinxx, Win95d]. MagicFax [Ano97-119]. mail [Ano96-32, Ber96f, Gar97c, Com95c, Hor96a, Hor96b]. mails [Bra96]. main [Car96c]. Mainframe [WP95, Hib97, Oue96b]. Mainstream [Ano96-88, Gar97c, Gil97a]. maintain [Mat96b]. maintainability [Lak96]. Mainz [Uni97a]. Mais [Walxxb]. Major [Gar99, Law98, Mun96]. Make [And97a, DDJ99b, Gar97b, Hay95c, Hay97b, Sha97b, Way96e]. Maker [Ubo96b]. Makers [Ric95a]. Makes [Ano96h, Ano96i, FC96a, MD99, Ano96m, Ano96-153, Fet97, Hay96j, Nap97, Van97m]. Making [Ano96b, Lut99a, BOSW98, DDJ98b, DDJ98c, Jav98, Lut99b, Pir97, Ren98, SMC^{+}96, VM96, Way98c, Wil98b, BF99b, Lev96b, Pro96, RBL^{+}97]. Makkie [Jasxx]. Malice [Anoxx-37]. malicious [McC98e]. Mall [Mas95]. Manage [Ano97-105, CCB96b, Ano96-123, FC96a, McC97a]. Management [Ano96-135, Ano97e, Ano97-41, Ano97-82, Ano98-27, Cas96a, DW96b, HLRT97, JL96, JCS98a, Kro98d, Kro98c, Kro99e, Kro99c, LPR97, Lut98a, McG97, SEH98, Sul97, Tuc96, Ano96-54, Ano98s, AA98, BM98c, Che96, DR98, DR99, Dry96a, Dry96b, For96, Gre98c, Jan96, LP97a, SEH99a, SEH99b]. Manager [Ano97-119, Ber96d, Ber97h, Cue96, San98b, San98c, van97e, Bea96]. Managers [Ano97-56, SMC^{+}96, Ano96-82]. Managing [Ano96-123, Wil97e, Anoxx-50, Hor96a, Hor96b, Kro98a, MF97c, MRA98, Wil97f, Ver96]. mandulgi [Kim97d]. Maneira [Walxxb]. Manhattan [IEE96p, Dec95b]. manifesto [Sut96]. manipulation [Boy98c]. mannered [Dav97d]. Manual [BG98b, BG98c, DH97, FBxx, GG96b, GG97, Joh96b, LTxx, Mas97a, NS97c, Nau97, PLS98, TBxx, dVGxx]. Manuale [Hol97b, Holxxd, Manxxa, Nauxxb]. Manuales [Fra96b]. Manufacturing [CTL98, AA97]. many [Ano96-42, Ano97-91, Ude97d, Wal96h]. Map [Wat95a]. Maple [Bra96, Sch99b]. Mapping [Ano96-109]. Maps [Hor96a, Hor96b, Gre98f]. Marc [Win95a].

March [Ano96-160, Ano98-34, IEE96p, IEE96k, IEE96j, IEE99b, IDK98, Joy98, Nat99, USE99e, Ano96y, Uni97a]. Marimba [Ano97-106, Goo97a, Goo97b, Gooxxa, Lem97d, Lem97e, Lemxxd]. Market [Ano97-130, Boz95d, Ano96-153, Cox96b, Gro96b, Wal96i]. Marketplace [GC97]. Markets [FSW98, Ano97-140]. Marquardt [Rob98a]. Marries [Gre98c]. mars [Dys96, Gau97r]. Martin [ACM99c]. marvel [Ano96-166]. Maryland [IEE96f]. Maschine [LYJ^{+}97, LYxxa]. Mass [USE98a, Bat96c]. mass-media [Bat96c]. Massachusetts [USE99e]. Massenmedium [Bat96d]. Masses [Ben99a, Way96d]. massively [Ber99b, Ber99c]. Master [Gri97a, Per96b, Rod96b, TD97, Woo97b, M^{+}96b, M^{+}97e, M^{+}97c]. MasterConsole [Kuk98]. Mastering [Ano99d, Ano99e, Ano99f, GPF97, GM97c, GM97d, Gea99, Han97, Hol97g, Hol97h, Jaw97a, KGTF97, Pur96, Rom99, Swa97f, Swaxx, Van96g, VPH^{+}96, VPHS97, Van97o, Vanxxf, Zuk98, GM98a, Vanxxg]. Masters [Chu98a, CC96]. match [PM97d]. matches [Ano96q]. Matching [Ano96-65, MD99, CL97d]. Math [Gar97c, ZOxx, SGB97]. Mathematica [Bra96, Kro98d, Kro98c]. Mathematics [Syd97]. MATLAB [DP99, Ano97r, Ano97-153, AC99, Bra97a, Bra97b]. matrix [SGB97]. Matthias [Ano98e]. mature [Bur96a]. matures [Shu97b]. Maturing [Way98a]. Maximize [BLT98]. Maximum [V^{+}97]. May [ACM96, ACM99b, Als95b, Ano97-56, Ano97-54, Ano97-139, Bet95, BBJR96, FW96, IEE96d, IEE96a, IEE96b, IEE96r, IEE98c, USE99c, USE99d, Wae96, Ano96-54, GJ96b, Gil97a, Mac96]. Maybe [Ban95a, Gre96a, Hal97a]. Mayo [CCB96a]. McClellan [Bay97a]. MCI [BH96c]. McNealy [Ano96-112, Ano97-52]. Me [Macxx, MO96, Hay96k]. Measurement [MRA98, Ano97-135]. Measuring [Ano98l]. mechanics [MC98]. mechanism [Cli98a, MF97a, RWB97, pP98]. Mechanisms [KG98b, KG98c, For98a, SvK98]. Mecklermedia [TG95]. med [Dan97b, Dan97a]. Media [Bay97a, DBK98, GT99, JMA98, Tee99, Ver98, Win98, deC99, Bat96c, Ber96c, NJC96, Van97m]. Median [NRG^{+}99]. mediated [TSKA^{+}96]. Medical [CCB96a, Chu98b, HJS98, JOA98, VPM98a, VPM98b, Cim96]. Medicine [W^{+}98]. mediocre [Oua96]. Mediterranean [IEE98c]. MedJava [JWS98]. Meet [Ano95j, LH96, Oue95a]. Meeting [DLM99]. meets [Ano97-150, Hem96a, W^{+}98]. Megapoche [Marxxc, Rodxxb, Lon97]. Mehr [Mül96]. Meilleur [Ano96-120]. Meine [DE96a]. Melbourne [MDM97]. MELECON [IEE98c]. members [Ano96-78]. Memory [Ano96-153, DS99, HK98, JL96, Kar98a, KT99, McC99d, San98b, San98c, Kar98b, MC97]. men [Holxxb, Holxxc]. menace [Ano96-166]. Menu [Lou95a, Gau96a, Hay96r, Pic96b]. Merced [GC97, Tho98]. Mercury [Hor96a, Hor96b]. Mesh [GM98b]. Mesmerizing [Ano96b]. Message [Kar98a, Kro98d, Kro98c, MG97b, Ano96-73, DLM99, Kar98b, SZ99]. messaging [LH96]. meta [KG96a, KG96b]. MetaCard [Kuk98, PJR^{+}98]. metacomputing [BKKW96, CKV98, JCS98b, PSS97]. MetaCreation [Kro99a, Kro99b, Kro99c]. metadata [AS96]. Metafiles [Del97]. MetaJava [GK98, KG96a, KG98b, KG96b, KG98c]. Method [Dow98a, HT98, Hiu97a, Hiu97c, Hiu97b, Hiu98, KWB^{+}98, Lin98a, OL98, RWB97, Rao99d, TTK98b, VR99a, WW97, Wal98a, Wil97d, YHW98, Ber97b, Cli98a, Led97, MvNV^{+}99, PL97, Rob98a, TTK98a, Zac97]. Methode [G^{+}97, BFGGxx, Gla97c]. Methodology [Por98]. Methods [BC97b, Des98, Leo98, Qia99, WH99, Wel97]. Metrowerks [Ess95, Gre97k, Kro99a, Kro99b, Kro99e, Met95d]. Meyer [Ano97k]. MFC [Ano97-124, Proxx]. MFC/ActiveX [Proxx]. MHP [NAC99]. Miami [IEE96o]. Michael [Ver98]. Micro [Ano97-107, Ano97-108, Ano97-109, Ano97-110, Fla99a, Mat96b, Mat99, Wea97b, HB95]. Microarchitecture [IEE96l]. Micromanipulation [AVN98]. Microprocessor [Ano97-130, Lut98b, Zan96]. microprocessors [Ano97-126]. Microserver [Ano98-28]. Microsoft [Dav96b, Ladxxb, SP97, Ano96r, Ano96-39, Ano96-73, Ano96-97, Ano96-98, Ano96-124, Ano96-57, Ano98x, Ano98y, BM98b, Cou96a, CS97b, D^{+}98b, Dav98b, EMe97a, EMe97d, Gar97b, Gau96e, GH96b, Gau96f, Gau96g, Gau97g, Gil98b, Gre95, Gut95i, Hay96j, Hay96p, Hay97c, Hib96, JMM96, JN95a, JN95b, JN95c, Knu99a, Lie95c, Min96b, Mas97a, MW95, Mil98, NJ95, OH97b, Par95d, Par95e, Pat98a, Per96b, Sim96c, SW95, SG97a, SG97b, Ste97a, SS98d, Ten95a, Wat98, Wat99, Wat95a, WSL96, Wil98b, Wil99b, Win95c, WL95, Win95f, Won97, Wygxx, You97b]. Microsofts [W^{+}97c]. Microsystem [McC96c]. Microsystems [Ano95r, App95, Mag95, Ano96-54, Ano95o, Ano96-53, Ano96-174, Ano97-52, Ano97-129, IBM95, Mac95, McC95b, Met95d, Sil95, Spy95]. Microwave [Met95a]. Microway [Ano98-28]. Middleware [Ano96-125, Ano97-79, Ano97-86, BCC^{+}97, Kro98d, Kro98c, San98b, San98c, Ano96-191, AA98, DHM99a, HG96a]. Might [DDJ98b, DDJ98c]. Migratable [TMN^{+}98a, TMN^{+}98b]. Migrates [Ano97-168]. Migrating [Jos97b]. Migration [Ano98n, Chw98, Fue98, Fün98, SMY99]. Milk [Ano96-37]. milked [Fla96a].

Mimicking [For98a]. mind [DB97]. MindQ [Ano97-111]. Mindstorms [BAC^{+}99]. Miniato [PZ96]. Minidoc [Perxxa, GGxxb]. minimal [IPW99]. Mining [GPHL99, MMGL99, WR99a]. Minute [Gai97, Tit96b, TGxxa, TG95]. Mirror [Knu97b, Knu97c, Knu97a]. miss [Hay97b]. missing [Ano96b, Ste97b]. Mission [DR98, DR99]. Mission-critical [DR98, DR99]. Mistakes [Ano97-52]. Mita [MD99]. Mitchell [GM98c]. Mitsubishi [Ano96-126, MD99]. Mixed [Kro98d, Kro98c]. Mixed-Signal [Kro98d, Kro98c]. mixes [Ano96q]. Mixing [CL97d, MMBC97]. mixins [FKF98]. MKS [Ano96-90]. ML [BKR99]. MMX^{T M} [BGH99]. MO [ACM97b]. Mobile [AT97, IG99, BCC^{+}97, Dom96, Fra98a, Fue98, Fün98, Gar98a, LOKK97, LO98b, MF98, MF99, MD97b, VI96, Wak98b, CG99, FC98b, GA98, KCLZ98, LO99c, Lim95, Tho97a, WPM99]. Mocha [van96c]. MockUp [MD99]. Model [AKM99, BFPV99, BDV^{+}98a, For98d, GM98b, Gig97, Gol97, Gon98, KSI^{+}96, PV98, SKN98, VP98, WRW96b, WRW96c, BCLT95, BDV^{+}98b, Dav97a, FC96a, MN96b, RMB^{+}98, Ste96c, WRW96a, Zac97]. Model-Driven [KSI^{+}96]. Modeler [Kro98b]. Modeling [Bru98, BB99, BUP99, CIF98, CLD99, DDJ99b, DG96, Lut98b, Lut99e, PSL^{+}98, PPIG98, Spe98, CGHS99, D’S96b, Fox97a, Fox97b, FF97, Gre98a]. modélisation [Laixx]. Modelling [Lut99c, PIPR97a, PIPR97b, Syd97]. Models [Ano98l, Cha97c, CL97d, Cox96a, HJZ^{+}98, YK99, Cor98a, RHND99]. Modems [Gar97c, Ano96-165]. Modern [App97, App98, Ano97l, Ano98k, vD99b]. Modular [BS99, Dug99, FC98b]. Module [Ano98-28]. Modules [Div98, FFKO99]. Mojo [Dew97, Ano97-127, Bru97]. Momentum [Ric96b]. Monash [Kro99c]. Money [And97a, Kar95b]. Monitor [Ano97-112, Ano98-30, Wel97, YLP^{+}99]. monitoring [BS96, Red98]. Monitors [KP98a, KP98b, MD99]. Monroe [Tee99]. Monson [Ano99a]. Monson-Haefel [Ano99a]. Monstor [Gar98b]. Monterey [Spe98, Ass96b, USE97c]. month [Ano96-53]. Montréal [FGL96]. Montreux [Str96]. Morning [Sha96b]. Morphing [Ano96-127]. Mosaic [Des96, Spy95]. most [Mat96b, vSS96e]. mostly [Mye99]. mostly-static [Mye99]. Motif [Ano97-113, Ano97-168]. Motif-to-Java [Ano97-113]. Motion [Ano96-30, BAC^{+}99]. Motion-Tracking [BAC^{+}99]. Motorola [Gar97a]. Mountain [Kro98d, Kro98c]. Move [Cox95b, Gar98a, Vad95b, Ano97-69]. Movement [GF99]. movie [Sha98b]. movimento [Dec97a, Decxxa]. Moving [Ait96b, Ath98, HL99, WL97b, You97b]. MOWS [Yos97]. MP [Ano97-118]. MP-EG1A [Ano97-118]. MPEG [BD99]. MPEG-4 [BD99]. MPEG-J [BD99]. MPI [BCFK99, DLM99, GGS99, MG97b, NE98a, NE98b, WCS99]. mpiJava [BCFK99]. MQseries [Ano96-48]. Mr. [Egr99, O’B96a, O’B96b]. MS [Ano96-87]. MTS [Cha99f]. Much [Ano96-54, Agn97, Ven96]. Much-hyped [Ano96-54]. Multi [Ano96-154, Ano97-35, BC97b, CV98, CKRW99, Eng97a, GGMS99, Kro98d, Kro98c, RKCW98, CKRW97a, CKRW97b, PS96]. Multi-Agent [Ano96-154]. Multi-language [GGMS99]. Multi-Methods [BC97b]. Multi-ORB [Kro98d, Kro98c]. Multi-Threaded [CKRW99, RKCW98, CV98, CKRW97a, CKRW97b]. Multi-user [Eng97a, PS96]. MultiConference [BUP99]. Multilanguage [MHO96]. Multilateral [PSW^{+}98]. Multilevel [GM98b]. Multilingual [Kro98b, Won98]. Multimedia [Ano99h, Bat96b, Bry98, DBK98, DD97e, DD97f, DD97g, Gro96b, Hor96a, Hor96b, IEE96q, IEE96j, Kam99, KJR98, KGG96, Mes95a, Mil95b, TMM^{+}98, TMMH99, Bat96a, Die97, DHM99a, Kap97a, Kar95b, PS96, RB96, Rus98, TYY96, WS98]. multimediale [DE96b]. multimodal [Mar99b]. multiplataforma [C^{+}xxb]. Multiplatform [Ano96-137]. Multiple [HCC^{+}98, RC97, Ren98, TKH99]. Multiple-Backend [RC97]. multiply [GME99]. multiply-add [GME99]. Multithreaded [Ano98-32, Ano98-33, AG96a, Loe97, GLC99, Ver97]. Multiwavelength [Lut99d]. Munich [Ano96-147, Bod96]. muscles [Cox96b]. Muster [Lea97b]. Mutual [For98a]. MX4 [Kuk98]. My [Ber97h, Hol96a, Bab97a, Ber98b, Bös98, Har99d].

N [Ano96-133, HNS97a, RVM96]. Name [Gar97a, Gar98b, Ano96-195]. Names [Gre97e]. Nanomedicine [DDJ98b, DDJ98c]. Nashville [Ano98-34]. Nassau [Ano99h]. National [Ano98a, Cox95c]. Native [AC98, Des98, Fli98, GFHM98, GM98d, GS98c, GS98d, Gri97b, Lia99, LSxx, Lot95, MG97a, Rao99e, SZ99, WdFD^{+}98, WDDF98, Wil97d, BG97c, EL98, GHM98, HGH96, Oue96a, Wil98b]. Native-Language-Based [GS98c, GS98d]. Natural [Doe95]. Naughton [Ubo96b]. Navigates [Bar95a]. Navigating [Kin97a, Lor96b]. Navigator [Ano96l, Cag96, Ano95a, Ano95k, Ano95l, Ano96-133, Bal96b, Cro95b, DM96b, Gau96e, Hec95, Mes97, STV95, Win95h]. Navisoft [Net95c]. NC [The98, GJ97, Hiu97a, Hiu97c, Hiu97b, Lev96b, Que97, Rob96, Way96a, Way96c]. NCs [Lev96d, Mac97b]. NCSA [Ano96-128, Des96, MLF97]. NDP [Ano97-121]. NDRAM [Gar97c]. Near [Gon97]. nearly [Ban95a]. Nears [Gar97c]. NECC [Ano98a]. Need [BSS96, Jar97, PQ96, GHP^{+}96, Hay96e, Joc98, Mee96, M^{+}96b, M^{+}97e, M^{+}97c]. needed [NMS96]. Needs [Boz95a, Lou95b, Ude96, Bur96a, CAB^{+}98, Wal96h]. nelle [Dec97a, Decxxa]. NEO [Gar96d, LaM95e]. Nerds [Sto96]. nerves [Ano96j, Joh97a]. nested [Shu97b]. ’net [GH96c, Ano95s, Ano96-31, Ano96-43, Ano97r, Ano97-34, Ano98-31, Cha96d, Gar97a, Gar97c, Gro96b, Ham96b, Hor96a, Hor96b, Kro98e, Kro98c, Kro98b, Kro99a, Kro99b, Kro99d, KMW99, Lea99, Lut98a, MD99, Tuc96, Ude97c, Val95, Way96d, HCJ^{+}97, Ano96-150, Bar95b, Bat96c, Cro95b, Fly95, Gro96b, Kal95a, Mel96, NE98a, NE98b, Tub95, WSL96]. Net-Centric [Ham96b]. Net-dbx [NE98a, NE98b]. Net/SLE [WOC99]. NetAcquire [Ano97-134]. Netra [Ano97-99]. NetReXX [Cow96, Bue98, H^{+}xxb, IBM97, Tay98]. NetRight [Ano97-114]. Netscape [BK97b, Hol97g, Nic97, Wat95d, Als95c, And98a, Ano95k, Ano95l, Ano96r, Ano96l, Ano96-133, Ano96-129, Ban95a, BCW97, Ber96c, Cag96, CCB96b, DFW96, Des96, DM96b, Dun97, Gar97b, Gau96e, Gau96g, Gut95b, Har96f, Hec95, HL97b, HL97c, JV97, KK96a, KK97b, KKxxa, Law98, LS96, Llo98, Lut98a, McC95b, Mes95a, Mic97b, Mil95c, Nas95b, Nas95c, NJ95, OH97a, O’C95d, O’C95e, OBWR96, Par95a, Par95f, Par95g, PM97c, PF97, Rao97, Red97a, Rit96a, Sil95, Sha96a, Sim96d, SW95, STV95, Tat96, Ude97c, Val96a, Van98b, Var95, Wat95b, Win95g, Win95h, Win95i, Win95j, Wod96, You97a]. NetSolve [CD97]. Netspeak [Far96a]. NetTracker [Ano98-28]. NetVue [Ano98-29]. NetVue/JAVA [Ano98-29]. NetWare [Gil97c, Ano96-130, Smi97]. Network [ACM98a, Ano96-101, Ano97-80, Ano98l, Ano98m, Ano98n, Ano98-27, Ano98-30, BKKK98a, Bue98, CBT98f, Cou98, The98, Fer98b, FC96a, Gra96d, GS98c, Gro96c, Gut95c, Gut95e, Har97a, Harxxb, Hiu97c, Hiu98, HHSW97a, HHSW97b, HHW97, Kan96b, Kro98d, Kro98e, Kro98c, Kro99c, Law98, Law99, Lea99, LPR97, Lut98a, Lut99e, Mac98, Mac97c, Mon97, O’N97c, O’N97a, Pis99, Rei97a, Ren98, Roo98, SZ98, Web96, Wil98c, WP95, Won99, Ano97j, Ano98s, AA96a, BKKK98b, Bat96c, Che96, Fer98a, Fox98a, Fox98b, GJ96b, GS98d, Jan96, McC97a, Met96, MC97, Ols96, SF96c, Str99, Wal99, Way96c, WOC99, DC97, Ano98-30, Ora95, Kro99c]. Network-Based [Ano96-101]. Network-Centric [FC96a, MC97, SF96c, Wal99]. Network-Enabled [Won99]. Networked [Ano96-132, BK96b, Ano97-54, Wil96f]. Networking [Ano96-137, CBT98h, Cou97g, Gam96, IEE96b, Mil99, N^{+}xxa, Nie97, Sch96c, Sri97a, S^{+}96b, Ano96-42, Den97, KJR98, NMS96]. Networks [Cli98a, Lut99d, Net95a, Rei97b, Ano96-123, Nan96]. Networld [MW95]. Netz [Bat96d]. neuen [Küh97, SBH97]. neuer [W^{+}97c]. Neuron [LaM95c]. Never [TMP^{+}99, LH96, PC97]. News [Ano95m, Ano95n, Ano96-139, Ano96-138, Ano97-28, Ano97-107, Ano97-153, Bar95a, Bet95, Boz95a, Boz95c, Boz95b, Boz95d, Bra96, Bra97a, Bra97b, Bro95b, Bro95a, Bus95, Cha99e, Col95, Com95a, Com95b, Com95c, Com95d, Com95e, Com95f, Cox95a, Cox95b, Cox95c, Cox95d, Cro95a, Cro95b, DDJ98b, DDJ98c, The98, DDJ98a, DDJ99a, DDJ99b, DS95a, DS95b, DeV95, Fer95, FC95, Gar97a, Gar97b, Gar97c, GC97, Gar97d, Gar98a, Gar98b, Gar99, Gib95, Gro96b, Gut95d, Gut95c, Gut95f, Gut95e, Gut95g, Gut95b, Gut95h, Gut95i, Gut95j, Gut95k, Hay95b, Hay95c, HB95, Hay95d, Hay95e, Hes95, Inf95a, Inf95b, Inf95c, Inf95e, Inf95d, Inf95f, JN95b, JN95c, Kal95a, Kal95b, Kno95, LaM95a, LaM95b, LaM95c, LaM95d, LaM95f, LaM95e, LaM95g, Law98, Law99, Lea99, Mas95, McK95a, McK95b, Mes95b, MW95, Mes95a]. News [Mes95c, Mes95d, Mil95a, Mil95b, Mil95c, Mil95d, Moo95, Nas95a, Nas95b, Nas95c, NJ95, Net95c, Oue95a, Oue95b, Par95c, Par95d, Par95e, Par95f, Par95g, Par95h, Ric95a, Ric95b, Ric95c, Sin95b, Sin95a, Ten95b, Tri95, Uim95, Vad95a, Vad95b, Wat95a, Wat95b, Wat95d, Wat95c, Wat95e, Win95c, WK95, WP95, Win95d, WL95, Win95e, Win95f, Win95g, Win95h, Win95i, Win95j, Win95m, Win95l, Win95k, Win95n, TSKA^{+}96]. newspaper [KB96, Kam96]. NExS [Ano96-137]. Next [And97b, Ano96-140, Bec97a, Bec98, Koh95, Kro99d, OL98, SS96, And96a, Ano96s, HMG^{+}95, Hoy96, O’C95e, Sil95, Par95b]. Next-Gen [Kro99d]. Nexus [FTT97]. ni [IK96]. Niemeyer [Ano96n, Ano98f, SMC97]. Night [BCW97, JMM96, RCRBR97, WB^{+}96b, Zif97, Cou97e]. Nine [Way98a]. Ninflet [TMN^{+}98a, TMN^{+}98b]. Ninth [Uni96a, Uni96b, Joy98]. NIST [Bra96, Dra98]. nm [Gar97a, BBJR96, USE98b]. No [BP97, BPxx, DDJ98b, DDJ98c, Hol97e, Hol97i, McC97c, Zuk97a, Ano96-102, Che97, DM96a, Hol97d, IK96, Way96c]. Noches [Weh98a]. Nod [Ano95f, The98]. non [KCLZ98]. non-strict [KCLZ98]. nonlinear [Rob98a]. Nonprogrammer [KK96a, KK97b]. Norme [Drixx]. Northcon [IEE96g]. Northcon/96 [IEE96g]. Norton [Nor96, NS96]. Notation [CR99, Laixx]. Note [Bar96, BG97c, Pol98, Vil99]. Notebook [Mor96a]. notebooks [Ano96-40]. Notes [H^{+}97, Ano96-102, Hin96, Bar95a, Cav97, DS95a, DS95b, Lot95, Oue95b, Win95e]. Notes- [H^{+}97]. Notes/Domino [H^{+}97]. NOTIO [SL99]. Nov [MDM97]. novel [LS99]. Novell [Ano97r, Col96a, Gil97c, Knu97h, Smi97, WB98]. November [ACM99c, IEE96f, IEE96g, IEE96i, Isk96, MJ98, USE96b, YN96, ACM97b, Chu98b]. Novera [Ano98-31, Way97b]. novice [WK99a, WK99b]. noyau [GYT97a]. NT [DC97, JC97, Ano96-97, Ano98-32, Ano98-33, Coo97c, Infxxa, Mes97, Neo99, Sym95, TJ96, Wil98b]. NT/95 [Infxxa]. Nuls [KKxxb, Van97l]. Number [CMH99, Val96a]. Numbers [Bro95b, Knu98a]. Numeric [Ano97-119, Ano97-164]. Numerical [BC98b, Gos97b, Gos98a, Num98, BC99b, BDP^{+}98a, BDP^{+}98b, CDD97, MMG98, SS98a, SS98b, Woe97]. Numerics [Ano97-121, Ano97-162, Ano97-163, Ano97-164]. Nur [Frö96, Fröxxb, Wes97b]. Nutmeg [Gar96f]. Nutshell [Fla95, Fla96b, Fla96c, Fla97a, F^{+}97a, Fla97d, F^{+}97b, Fla97e, Fla97c, Fla98a, Fla98b, Fla98c, FFCM99, Fla99b, Fla99c, Flaxxa, Flaxxb, Ano96o, Lea98b, PF97, Fla97b, SMC97, Nic96a]. nutzbringenden [Bat96d]. nutzen [vSS96d, Hie97]. nutzliche [SBH97]. Nwanua [Ano98-27].

O [Ano96-137, McC96e, DM96a, Egr99, Hol98b, Nauxxa, Har99b, Har99c, McC98f, PL96, Wel99a]. OADD [Ham97c]. Oak [Mor95]. Oakland [IEE96a, USE96b]. Oaks [Ano98j]. Object [All99c, Ano96-28, Ano96-142, Ano96-144, Ano96-158, Ano96-159, Ano96-171, Ano97-124, AW98, Atw96, BCFK99, BS98a, Ber96j, Bla96, Boo96a, BDV^{+}98a, BLT98, Bro97b, Bud98, BCC^{+}97, CBT97f, CI96, CW98, Cha97c, CL97d, Cli98a, Coc98, Com95b, Con97a, Cow97, Cul98d, Culxxa, FC95, FFKO99, FM98, GM97e, GM98c, Gil97a, Gre98c, Hav98, HJZ^{+}98, HYI98a, Hun97, Hun99, JCS98a, Kro98a, LaM95e, LH96, Lor96a, Lor96b, Lut99c, MRA98, Mey97, MDM97, MK99, MJA98, MR97, Net95a, Nix98, OH97a, OK99, PIPR97a, PIPR97b, PPIG98, Pit97, PK98, Pop97, Rey97, Ric96a, Rym95, SHR^{+}99, Sch97h, Sch97j, SH96, Stu98, Ass96a, USE96c, USE97b, USE98b, USE99c, Uma97, VRG98a, VRG98b, Wel99b, WRW96b, WRW96c, Wu99, Wuxx, Ada98, Amb97, Ano96k, Ano96-73, Ano96-60, Ast98, Beh95, Bla99, BDV^{+}98b]. object [CGJ^{+}98, Coo97c, DDG^{+}96, DV98, FM99b, Gre98f, Had99, HYI98b, JCS98b, LM99, Nan95, RMB^{+}98, SML96, TB96, Tol97a, Tol97b, War99, WRW96a, Zol96, Ano96-143, Ano97-146, BN97, CW96a, Coo97b, FI96b, Hor97a, Kal95b, Kro98b, Kro99a, Kro99b, OHE97a, Com95g, Sad97a, Ste97c]. object-based [Ast98, TB96]. Object-Orientation [CI96]. Object-Oriented [All99c, Ano96-158, Ano97-124, AW98, BCFK99, BLT98, Bro97b, Bud98, CBT97f, CW98, Cli98a, Coc98, GM98c, Lor96a, MRA98, Mey97, MDM97, PIPR97a, PIPR97b, Pit97, PK98, Sch97j, Stu98, Ass96a, USE96c, USE97b, USE98b, USE99c, Uma97, VRG98a, VRG98b, Wu99, Wuxx, Ano96-171, Bla96, Boo96a, Sch97h, Ada98, Ano96k, Beh95, Bla99, Coo97c, DDG^{+}96, DV98, Gre98c, Had99, Nan95, SML96, Tol97a, Tol97b, War99, CW96a, Coo97b, FI96b, Hor97a, BN97, Sad97a]. object-relational [Gre98f]. object-sharing [CGJ^{+}98]. Object/Relational [Gil97a]. Objectgeorienteerd [MG98]. ObjecTime [Kro98b]. Objectivity [Gre98d]. Objectivity/DB [Gre98d]. Objects [AJ98a, All99d, All99e, Ano97-30, Ano97-33, Ano97-105, Lut99a, Bak97a, Bakxx, BCH^{+}98, BN97, Ber97i, Ber97e, BFD96, Bow96a, BN96, CBT97d, Cha98d, Cha98e, CJJ^{+}98b, CJJ^{+}98a, DHM99b, GM98b, GS98b, Lip99, Lor96b, Lut99b, Mai98, MCD99a, Qia99, Sad97a, Som97, TMN^{+}98a, WK95, WAES97a, WAES97b, Zei97, AA98, D’S96c, Dav97d, Gil97a, Gre98f, JA98, KR99, PZ97, PS98, Pog96, Com95g, TMN^{+}98b, VA98, Ora95]. ObjectShare [Ano97-123]. ObjectSpace [Ano96-135, Ano96-137, Ano96-135, Ano96-145, Ano96-146, Ano97-124, Gla98]. Objekt [Ano96-147]. Objekt-Orientiertes [Ano96-147]. Objektorientierte [CW97, CWxxa, Beh95, SST98]. objektorientiertes [Bou96a, Cul98f]. Objet [Drixx, Laixx]. Observations [Swa97c, Swa97d]. observed [Phi99]. obsolete [Way96e]. OCLC [Ano96-148]. octane [Ano96b]. October [Ano99h, Bod96, BK96a, CI96, CEG96, Cim96, IEE98a, SLPM96, USE96a, USE99a, USE99b, YN96, dN97]. ODBC [Ano96-137, Sip98]. Odd [WC99]. ODE [Ano97-153, Bra97a, Bra97b]. ODMG [Ala99]. OEMs [Ano97-150]. Off [DDJ98b, DDJ98c, Par95f, SB97, Ano96-196, Ban98, Cou97a, LP96a]. offer [Ano96-54, Ano96-151, Ano97-138, Cli98b, LS99, Ste97a]. Offered [Hor96a, Hor96b]. Offering [FC95]. Offers [Ano97-41, Ano97-55, Ano97-101, Ano97-114, Com95d, Gar97a, GC97, Inf95e, MD99, Ric95c, Vad95a, Ano95s, Ano96-196, Ano96-153, Ano97s, Ano97-73, Ste97a, Ubo96b, Vij97b]. Office [And97b, Hay97c, Hay97d, SC96a, Anoxxg, GH96b, MK97, Ung97]. Official [Ano97-125, BC97a, CB98, Dun97, Fis98a, Goo97a, Goo97b, JV97, KK96a, KK97b, Lem97d, Lem97e, Llo98, Mic97b, Nic97, PM97c, TSHH96, Tit96b, WL97b, Ano98y, TG95, Way96c]. offizielle [GJS97]. Offline [Ano96-135]. Oficial [Gooxxa, KKxxa, T^{+}97b]. ohjelmoinnin [Jam97b]. ohjelmointi [Ekxxb, Jamxxb, KNxx, Kiuxx]. ohne [Vac97, Wie96, Wie97]. Ojbects [Ano97-34]. OK [Gar98b]. older [Chw98]. OLE [SF96b, Inf95a]. Oltre [Var97b]. om [Ekxxd]. OMS [SKN98]. OMS/Java [SKN98]. On-Line [Ano96-134, Har96c, DS95b, SvK98, Vad95b]. One [Als95c, Mee96, Ano96y, App95, Gre97f, Val96a]. Online [Ano96-151, Ano96-148, Ano96-152, BCP^{+}99b, BKP97b, The98, Gau97h, KP99, MGL^{+}98, Ano96-166, Ano96-193, Bat96b, Cha96a, Fre95, Kam96, PSTZ99]. Online-Aktualisierung [BKP97b]. Only [Ano97-109, Ten95a, Che97, SF96a]. Ons [SBH97]. Ontario [Ano96-158, MJ98]. Onto [Ano96-106]. ontwikkelen [Rodxxa]. OODB [Yas96]. OODBMS [Gre97i, Gre98d, SKN98]. OOP [Ano96-147]. OOPSLA [Pfe99]. Op [BCP^{+}99b, MGL^{+}98]. Op-Ed [MGL^{+}98]. Op/Eds [BCP^{+}99b]. opaque [CTL98]. opas [Fla97f]. Open [Ano96-135, Fra98a, Gar97b, Jar97, MBxx, Net95a, Stu98, Tayxx, Urq97b, Ano96-66, Bab97b, Lea97a, FHKS98, Cha97a, GF99, Hac98, KGTF97, Mas95, NRG^{+}99, TD97]. OpenDoc [Ano96-133, Gau96a, Gau97i, LaM95b]. OpenDoc/Java [Gau97i]. OpenGL [Ano96-135, DDJ98b, DDJ98c]. openings [Ano97-75]. OpenMP [Ano97-153, Bra97a, Bra97b]. Opens [Law99, O’B96b, Ano96-143, Mio96]. Openscape [LaM95d]. Operating [Wil97e, CI96, Cha98d, Cha98e, IEE99b, KG98b, KG98c, Moh98, USE96a, VCCH99, Wil97f, Ano96-144, LTC96, Rey96, RR99, SM99]. Operating-System [KG98b, KG98c]. Operational [CKRW99, HBL99, CKRW97a, CKRW97b]. operations [Ber99a, Wat95f]. Opeta [D^{+}xx, Kiu98, KNxx, NKxx]. Opinion [Joh97a, Joh97b, Pre99]. opinions [Ubo96b]. opportunities [BN98b, D’S96b, Hay97b]. Opportunity [Mun96]. Optical [Lut99d]. Optima [BG96, Gre97b]. optimised [Zan96]. Optimization [Kro99a, Kro99b, HAKN97, RKCW98, SS98a, SS98b]. optimizations [CL97e]. optimizer [Cla97d]. Optimizing [BK97a, CL97f, HCJ^{+}97, MMS98, RC97, DDG^{+}96]. Options [Lei95]. Opts [JN95b]. OQL [Ala99]. Oracle [Ano96-53, Cel96a, Cro95a, Gau97j, Hay96q, Hob98c, Inf95a, Par95d, Ric95c, SW95, Ste97a, Win95j, Yan96].

Oracle8 [Won98]. ORB

[Ano97z, Gla98, Kro98d, Kro98c]. Orbix [Bak97a, Bakxx]. ORBs [Ano98-30]. ORDBMS [Gil97a]. order [LK99]. Oregon [ACM99c, IEE96r, USE97b]. O’Reilly [Ano96n, Ano96o, Ano96p, Ano97h, Ano97g, Ano97j, Ano97i, Ano97k, Ano98f, Ano98j, Ano98h, Ano98g, Ano99a, Ano99c, Ano99b]. Orientation [CI96, Hun97, Hun99, Ano96-144]. Oriented [All99c, Ano96-158, Ano97-124, AW98, BCFK99, BN97, BLT98, Bro97b, Bud98, CBT97f, CW98, Cli98a, Coc98, Con97a, Cow97, Frü99, GM97e, GM98c, Kro99a, Kro99b, Kro99c, Lor96a, MRA98, Mey97, MDM97, Nix98, PIPR97a, PIPR97b, PPIG98, Pit97, PK98, Rey97, Sad97a, Sch97j, Stu98, Ass96a, USE96c, USE97b, USE98b, USE99c, Uma97, VRG98a, VRG98b, Wu99, Wuxx, Ada98, Ano96k, Ano96-171, Beh95, Bla96, Bla99, Boo96a, Cli98b, Coo97c, Dav97a, DDG^{+}96, DV98, Gre98c, Had99, Lu96, Nan95, Ric96a, Sch97h, SML96, Tol97a, Tol97b, War99, Zol96, CW96a, Coo97b, FI96b, Hor97a]. Orientiertes [Ano96-147]. original [Ano96-78]. origins [Far96a]. Orlando [Chu98b]. Orleans [USE98c, USE98e]. Ornamentals [VPV99]. Orthogonal [ASG98, AJ98b]. Orthogonally [AJ99, ADJ^{+}96]. OS/2 [DiD97d, DiD97a, DiD97b, DiD97c, S^{+}96b]. OS/390 [Koo98]. OSDI [USE96a]. OSF [Mas95]. OT [Ano96-42, Cha97c]. Other [Mai98, Nan95, SR97, Swa98a, Ten95a, Ano96-53, McC97b, Nat96]. OTS [Cha99f, VR99b]. Our [Joh97a, GS97b, Soo98b]. Outdated [Wil97e, Wil97f]. Outils [F^{+}xxc]. Outline [Cro95a]. Outlines [Kro99d, Lot95]. Outlook [Ano96-153, Mil95c]. Outmuscle [WN97]. Output [Ano98-27, Gro96b, Rao99c, GHP^{+}96]. outshines [Chw98]. overcome [GJ97].

Overhauled [Gro96b]. Overlapping [KCLZ98]. Overnight [Con97b]. overriding [PL97]. Overview [Ano96-62, CBT97g, CBT98f, Car96a, Fet97, GMPS97, Sch96a, Bij96, Got98b, PS96, SvJ96]. Own [Ano96c, Ber97h, Hec95, Jar97, Ano96-74, Bea96, Ben98, Gre97h, Lav96, Van97n].

P [Ano98e, O’B96a]. PA [ACM96]. PAAM [Ano96-154]. pace [Che97]. Pack [Anoxx-53, MD99, Per96a]. Package [Ano96-161, MMGL99, WK99a, WK99b, BML97b]. Packages [FB96, N^{+}96a, GYT96c, Hay97b]. Packaging [All99d, Kro99a, Kro99b]. Pacotes [FBxx]. Pact [Bro97c, Win95g]. Page [Ano96-187, BKT96, LLxx, Tes96a, Ano96-53, Dys96, How96, Joh97a, KH99, Lem97a, Mee96, O’C97]. page/sites [O’C97]. Pages [Aoy97, Aoyxxb, Cro95b, DeR96, FLW96, Kan97, KK96a, S^{+}96a, Skl95, Tes96b, TPW96, Tre99b, Wod96, Ano96n, Ano96o, Ano96p, Ano97h, Ano97g, Ano97j, Ano97i, Ano97k, Ano97l, Ano98f, Ano98j, Ano98i, Ano98h, Ano98g, Ano98k, Ano99a, Ano99c, Ano99b, Coh96a, Fet97, Hei97, KK97b, vD99b]. PageSpace [CKTV96]. Pagine [GGxxa, Dec97a, Decxxa]. Pains [Joc98]. Painting [Kro99c]. Pair [Ano97-131]. PalmPilot [Ber98b, Chr99]. Palo [Fox98b]. Panel [MD99, UBF^{+}98, WBC^{+}96, BDF^{+}98, HCD^{+}98, NGM^{+}97, SBB^{+}97]. Pangaea [Spi99]. Paper [AL97, GM96, Ano97l]. Paperless [Fei98, SC96a, Wal96a]. Papers [Ber98h, Ber98i, IEE96c, ACM97a, ACM98b, Cul98g]. para [C^{+}xxb, Cha98b, Jim97, KKxxa, McCxxb, T^{+}xxa, Thoxxb, Walxxb]. Paradigm [IGO98, SF96a, SF97, BBC^{+}97, BGGR98, Chu98b, Had98a, Ubo96a, Wat95f]. Paradigms [Swa97e, Swa97b, Swa97a, Swa97c, Swa97d, Swa98a, Swa98b]. Paradise [Kro99d]. Paradox [Ric95b]. ParaGraph [Win95k]. Parallel [Bak97b, BSS98a, BCP99a, Bod96, IG99, CCF99, CZF^{+}98b, CCFL98, Cow98a, Fer98b, GM99, Gar98b, GFHM98, GDS98, GPHL99, Gol99, IEE96o, IR97, Ima97, JPK98a, KB98a, LP98a, LP98b, MMGL99, Sch99b, Sim97c, SZ98, TMN^{+}98a, WKES98, WKE98, YH96, Ano97s, BSS98b, CZF^{+}98a, CCCL97, CCI^{+}97, Cow98b, DF97, DLM99, DW97b, Fer98a, FF97, GRN99, GHM98, JPK98b, KB98b, TMN^{+}98b, YC98b, vRvGS97, YC98a]. Parallelism [Han99b, BG97b, Han99a, Led97]. Parallelization [BVG98, BG98b, BG98c, MST99, BDY99, BBGD98, BG98a]. Parameter [Sta96]. parameterization [AFM97]. Parameterized [BML97a, Dug99]. Parameters [Sta96]. Parametric [Ano98-32, Ano98-33, Kro98b, SA98a]. paranoid [Ano97-109]. Parasitic [BC97b]. ParaSoft [Kro99c]. Paris [ACM97a, IEE96l]. Parser [HFA^{+}98, PQ95]. Parsers [BN99a, MYD95]. Parsing [Ber99a, BCP^{+}99b, Par99]. Part [Ami97, Bec98, Coc98, DY99a, DY99b, EM99, Hal98a, JG96, OL98, Way98a, Ano96-97, Eck97c, SF96a, TC95, Ami96a, Ami96b]. Partial [Loe97, LMdS98]. partially [BN99b]. Particular [WC99]. Partner [Gut95e]. Parts [Gre97c, Sed98, Hug96, Gre97c]. Party [VPM98a, VPM98b, HG96a]. Pascal [Ano96-42, Buc97]. pasokon [IK96]. Passes [DDJ99a]. Passing [Fra98b, Kar98a, MG97b, DLM99, Kar98b, SZ99]. passphrase [Knu97f, Knu97e]. Passport [MGL^{+}98]. Password [Ber98c]. Past [BOSW98, AA96a]. patched [Nic99]. Patching [Com95c]. Patent [Bet95]. Patents [DDJ99b, Nat96]. Paths [LaM95c]. Patient [Wae96, JADR96]. Patrick [Ano96n, Ano98f, SMC97, Ubo96b]. Patriot [Ano97-126]. Pattern [Ano96-65, Pit98b, SvG98, Ver96].

Pattern-based [SvG98]. Patterns [DS99, Gra99a, Gra99b, Lan99b, Lea96a, LH97, Lut99e, Rem97, CH98a, FF98, FHA99, Ano98e]. paving [Wal96a]. Pay [Ano96-45, DDJ98b, DDJ98c, Hum96, Ano96-151]. PC [Ano97-119, Anoxx-65, Anoxx-66, Cro95a, Eud97a, Fra96a, GC97, GK97a, Har96e, MD99, MB96b, McC96a, Mil95c, Mon97, Way96g, Way96f, Weh96, Eud97b]. PCCTS [Par97]. PCs [Ano96-137, Ham97c]. PDA [Com95e, Pep97]. PDM [Ano97b]. peacemaker [Cha97c]. Peck [Ano96n, Ano98f, SMC97]. Pedagogy [Dal98]. peek [Ano96-188, Ano98u, Tho98]. Peer [S^{+}96b]. pegs [Ano97-73]. Pemograman [Wijxx]. Pencom [Ano96-155]. penetration [Ano97-73]. Pens [DDJ99b]. Pensée [Drixx]. Pentium [Ano96-40]. Pentium/133 [Ano96-40]. Penumbra [Ano97-127, Bru97]. People [Wil97e, Wil97f, Lav96, vSS96e]. Percolate [Boz95a]. Percolates [Ste97h, Tom96]. Perfect [WB^{+}96c]. perform [Kno96b]. Performance [ACM98a, AFH98a, Ano99g, BW98a, BDV^{+}98a, CP98, Espxx, FFHA98, GFHM98, GGS99, HO99, HYI98a, HYI98b, HBG^{+}97, HCJ^{+}97, Hum98, IEE96n, JWS98, LG99, Lut99e, McC98g, McC98h, McC98i, McC98f, McC99d, McC99a, McC99b, McC99c, Orc98, Red99, RLV^{+}96, SEH98, Shi99, Ten99, Tym96, WR99a, YC98a, YSP^{+}98a, ZLGS99, AFH98b, Ano96-153, Ano96-192, BDV^{+}98b, CL97e, CS97a, Cli98b, CFM^{+}97a, CFM^{+}97b, Eng97a, FF96b, Fox98a, Fox98b, GHM98, GGMS99, Gri97b, GLC99, GME99, LM99, SEH99a, SEH99b, Tho98, Way96b, Wel99a, YC98b, YSP^{+}98b]. performs [Chw98, MF97a, Ven97b]. Peripherals [MD99]. Perking [Ano96-119, Ano96-57, Ano96-156, Gau96j]. perks [Ano96-82, Ano96-100, Sul95b, Wil96e]. perl [Xu97, Bre96, Bal99, Coh96a, Coh97a, DDJ98a, PWH97, Roo97, Ste97d]. Permits [Gar97c]. Persistence [AJ98b, Gre98b, KS98a, KS98b, MJA98, SHR^{+}99, Tic97]. Persistent [AJ99, Ber97i, Bri99b, MJA98, Ude97b, ADJ^{+}96, LM99]. Personal [Gar97c, Kuk98, ZTLV98a, ZTLV98b]. personalised [TSKA^{+}96]. Personalized [BB97a, Kam96, WRB^{+}97, KB96]. PersonalJava [DY99a, DY99b]. Perspective [Kuk98, Har98b, MRA98]. Perspectives [Mos98b, Mos98a]. perusteet [Jam97b]. Petaflop [YK99]. Peter [Str97, Nic96b, Nor96, NS96]. Petri [KMW99]. petroleum [Liu96]. Phase [DD97i]. phenomena [Rob96]. Philadelphia [ACM96]. Phone [Gar97a]. Photo [MD99]. Photographs [Hor96a, Hor96b]. PHP [Atk99]. Physical [JG96, Ber98g, Lak96]. Physics [CT96, BK96b, CAZZ98, WCL97]. physlets [CT98]. Pick [Kuk98, Hay96k]. pick-me-up [Hay96k]. Pickin [Ano96-169]. Pickling [RWWB96a, RWWB96b, RWWB96c]. Picks [Ric96c]. PicoJava [MO98, TO96, HO99, OT97]. picoJava-I [OT97]. picoPERC [Nil98b]. Picture [NNS97, IEE96b, SF96a]. Pigs [Sin95a]. pillars [Dra97]. Pin [WP95, Ano96-153]. pioneer [DB97]. Pipeline [Inf95b, Inf95c]. Piping [DeR96]. pitched [Dry96b]. Pitches [Ung97]. Pitfalls [Dic97a, Laf96, Lafxxa, BT98]. Pixel [Ano97-118]. Pizza [OW97b]. PJama [AJ99]. PJAVA [MJA98]. Place [Kri96, Hay96r, SF96b]. Placement [KB98a, KB98b]. Plain [Ove96, Ove97, WC99]. Plan [Gar97a, Gar98b, Gau97b, Gau97e, Ano96-28]. Planar [GM98b]. planet [Gau97r]. Planetweb [Kro99d]. Planned [Ano96-198, Ano96-199, Cas96b]. Planning [Bro95a, KSI^{+}96]. Plans [Gar97d, Gut95d, Gut95f, Gut95b, Hal98b, JN95c, Lea99, Lot95, Wat95c, Ano95o, Ano96-97, Gau96e, Gau96f]. Plant [MD99]. plastic [Gau97q]. Platform [Ano96-101, Ano96-137, Ano97-39, AJ99, BS98c, BS98d, Bou96b, CLK99b, DM96b, Hiu97a, Hiu97c, Hiu97b, KGTF97, KG98b, KG98c, Kra96, NRG^{+}99, Net95a, PSW^{+}98, Rom99, Sun99a, WL96, WCF98, Ale98, Ano96-90, Ano96-59, Ano97-150, BF99b, Cal99, Gau97c, Gau97e, Gau97f, Gau97n, Gil98b, Gon97, Gon99, Gre97f, LC99, Rob98a, Sch97f, SF96c, SG97b, Tol97a, Tol97b, Way96c, Way98c, WFC^{+}99, YC97, MZ97, SF96c]. platform-independent [Gau97c]. Platforms [O’C98, Hay96f, Hay96k, Ren98]. Platinum [LO96, LOA^{+}96, Lad96, LO97a, LO^{+}99b, Mar97c, Ano96-136]. Plattform [Ano97-81]. Play [RVM96, Sim96c, Ano96-102, Ano96-188, Oue96c]. player [HG96a]. Plaza [IEE96p]. PLDI [ACM99b]. plenty [Sli97]. Plug [RVM96, SR97, You97a, Ano96-188, Gau96e, HL97b, HL97c]. plug-and-play [Ano96-188]. plug-in [Gau96e]. Plug-Ins [SR97, You97a, HL97b, HL97c]. Plug-N-Play [RVM96]. Plugging [Gul96b]. Plugs [Inf95f, Sim96d, Var95]. Plunge [Boz95d, Ano96-184, Gre97i]. Plus [Moh96, Tor96, TPW96, TTD96a, TTD97, TTDxx, T^{+}xxc, Ano97r, TTD96b, Ano96j, Ano96-135, Gre98d]. PNPACK [RSH97]. Poche [Anoxx-65, Anoxx-66, Cadxxb, Cadxxc, MB96b]. Pocket [Fla98d, SO97a]. Poet [Gre97i]. Point [DK98, Dar98b, Gar97c, KoCaB98, Kah98, KD98, Dar98a, Dar99a, GME99]. pointer [WR99b]. Pointers [Dem98b, Dem98a, MKS^{+}96]. Points [Gan96, Ano97-165]. poise [Ano96-153]. Poison [And98b, Ano96-157, Che99]. poky [Ham97a]. Policies [JLT98]. Policy [Ano98l, AA98]. Politics [Gil97b]. Polymorphism [SA98a]. Pop [Tub95]. POPL [ACM97a, ACM98b, ACM99a]. Popular [HYI98a, HYI98b]. por [GR97]. port [Ano97s, Hay96k]. Portability [Dic97a, Ano96-90, You98b]. Portable [MG97b, Ren98, SMY99, SC96b, TBA98, CS97a, Isa96]. Portables [Mil95c]. Porting [Kal98, MB98c, Dav97d, Hol99]. Portland [ACM99c, IEE96r, USE97b]. Ports [Ano96-47, Sim97c, Ano96-126, Chw98]. POS [MJA98]. poses [Ano96-171]. Position [Fer95, Hin97]. possibilities [LS99]. Possible [Ten95b, Ano96m]. possibly [Ano96-54]. Poster [CL98b, CL99, CLxxc]. Postgres95 [Bin96]. PostgreSQL [Klu98]. PostModern [Com95g]. Pot [Mat99, Lie96b]. Potential [RK97a, McC97b, Not96, Sul95a]. Pour [Gul96a, Anoxx-46, F^{+}xxc, KKxxb, Van97l]. Pouring [Wil96c]. Pourri [Mat99]. Pours [Ano96v, HG96b]. Power [BC97a, DDJ99b, Dow98b, Fla99d, Fla99a, GD98, Gut95c, Hat97, Hat98b, Lu96, NR96, Per96a, Rod98, Win95n, Ano97-145, Hat98b]. PowerBrowser [Ora95]. PowerBuilder [McK95b]. Powered [Ano95q, NE98a, NE98b, Ano97-82]. Powerful [Lot95, Cli98a, Dys96, Lav96]. PowerJ [CB98, Fis98a, Gre97b, Hat97, Hor98b]. Powerpack [DR96, Sav96]. Powersoft [Cox96b, Gre97b, Gre98d]. pp [Bay97a]. PR [And98b]. Practical [Ano96-154, Coo97b, Cue96, FR97, GFR98, Hoq97, Hor97a, Lut99e, MAS96, Sha98a, TLSS99, Cul98c, Mye99]. practice [BK97a, FHA99, IEE98a, OW97b]. Practices [Ano98l, BS98a, MO98]. Practitioners [Hun99, San98b, San98a, San98c]. Praktijkboek [Lonxxb]. Praktische [Red96, Redxx]. praktischen [TG96a]. Pratique [NP96c, G^{+}xxa]. Praxis [Klu97, Klu98, Kluxx, Wol98]. Praxisbuch [GK97a]. Praxislosungen [Vac97]. Pre [Uni96a, Ano96-195]. Pre-conference [Uni96a]. pre-K [Ano96-195]. preaches [Ham96c]. precise [CGHS99]. Precision [ADM98, Gri97b, GME99]. Predicated [PQ95]. Predicated-LL [PQ95]. predict [KW96]. Prediction [Hay97c]. predictive [GRN99]. predictive-adaptive [GRN99]. Predictor [SP98]. Predictors [VR99a]. prefetching [CKH99]. Preliminary [Lew97a, WJUH98, HGH96, JvdBH^{+}98]. Preload [McK95a]. Premia [Kro99d]. Premiere [Nor96]. Premium [Ano97-83]. PREMO [DHM99b, DHM99a]. Prentice [Gea96]. Preparation [Lut99a, Lut99b, Tuc98a]. Prepares [Law98, Oue95a]. Preprocessor [IR97, SW99]. Present [Gon97, AA96a]. presentation [RB96]. presented [ACM97a, ACM98b]. Presenting [Dec95c, Dec97b, Decxxc, Mor97g]. Press [Ano97l, Ano98i, Ano98e, Ano98k, Bay97a, DC97, Gea96, vD99b]. pressure [Ner97]. Prevent [Kay97]. Preventing [Wil97e, Wil97f]. Prevention [Kro99c]. Price [Ano96-28, TBS97, Ano96-153, Ano98-32, Ano98-33]. Price-Based [TBS97]. Priced [MD99]. Pricing [DS95a, FD96c]. prime [Gau97a, HG96a]. prime-time [Gau97a, HG96a]. Primer [Cul98e, Culxxb, For98b, Lut98b, Moh96, Tor96, TPW96, TTD96a, TTD97, TTDxx, T^{+}xxc, MAS96, TTD96b]. Primitive [All99f]. Princess [Ano97-139]. Princeton [Sav97]. Principled [Zei97]. Principles [ACM97a, ACM98b, ACM99a, Ano98l, Bis97, CBT97f, Coo97b, Lea96a, FHA99]. Print [Ano96-135, Ber97j, Gro96b, Van97m]. Printer [MD99]. prints [Ano96-196]. Privacy [IEE96a]. Privasuite [Ham97c]. Private [Law98]. Prize [GC97]. Pro [Ano96-137, Ano97-119, Ano97-121, How96, You97c, Kiuxx, Ano97-128, Ano97-154]. pro-kurssi [Kiuxx]. Pro/X11 [Ano97-119]. probabilities [RSH97]. Probably [DE97]. Probe [Ano98-30]. Problem [BS98a, KW98, PPJ96, SLO96, Wei98, Ano97-28, Zac97]. Problem-Solver [PPJ96, SLO96]. Problems [Gar97b, Ano96-90, Ano97-165, Hay96u, Stu98]. procedural [MST99]. procedure [Wal98a]. Proceeding [Isk96]. Proceedings [ACM96, ACM97b, ACM99b, ACM99a, ACM99d, Ano96-142, Ano96-147, Ano96-158, Ano96-159, Ano96-154, Ano96-160, Ano99h, AOV^{+}99, BBJR96, BUP99, CI96, FW96, IEE96h, IEE96l, IEE96i, IEE96r, IEE96n, IEE96o, IEE96p, IEE96k, IEE96q, IEE96j, IEE96s, IEE96m, Joy98, MJ98, MDM97, MJA98, Nat99, OW95, SMEN96, Spe98, Sta97b, Tau96, USE96c, USE96d, USE97c, USE98c, USE98b, USE98d, USE98e, USE98a, USE99a, USE99c, USE99b, USE99d, USE99e, Wae96, YN96, YH96, Ano97-139, Bod96, BK96a, Cim96, DLM99, Doo99, Fox98b, FGL96, IEE96d, IEE96b, IEE98b, IEE98a, IEE98c, IDK98, LM98, PZ96, DF96, SLPM96, Str96, Syd97, Uni96a, USE98f, dN97, Ten99]. Process [EMN98, HHW^{+}99, HK98, Kro99a, Kro99b, PJR^{+}98, Rob98b, AA97, MRA98]. Process-Oriented [Kro99a, Kro99b]. Processes [Wil97e, CJJ^{+}98a, Gar98b, KMW99, Wil97f]. Processing [Lut99a, Gar98b, GTV98a, Gol99, Lun99, Lut99c, Lut99b, LR97, Ran97a, Ran97b, Sak98, SZ98, WAES97a, WAES97b, dN97, Asl98, GTV98b, Lyo99, Ref99]. Processor [HO99, HBL99, VRG98a, VRG98b]. Processors [Ano96-175, Ano96-176, Ano96-153, Ano96-180, Ano97-149, WB99]. Produce [Kro99d, Cor98b]. Product [Ano96-161, Bru97, Cox95c, Hay97d, Hec95, Kri96, Met95d, Sha97b, Ano96-59, Ano97w, Ano97-142, Ban95a, Ung97]. Production [Kro99d, Ano98o, Kam99]. Productivity [Gil98c, Kuk98, MD99, Phi99]. Products [Ano96-135, Ano96-136, Ano96-137, Ano97-103, Ano97-106, Ano97-117, Ano97-118, Ano97-119, Ano97-120, Ano97-121, Ano97-116, Ano98-27, Ano98-28, Ano98-29, Chw98, Col95, Ham97b, Ham97c, Hor96a, Hor96b, JPA98, JN95c, Kro97, Kro98e, Kuk98, MD99, San98b, San98a, San98c, Urg96b, Ano97-153, Ano97-158, Bra97a, Bra97b, Flo97, IBM95, Mey99, Pic96c]. Professional [Ano96-137, A^{+}99, CMGR96, HI96b, HIxxb, Hor98b, Kas97a, Kro97, LPM96, LC99, M^{+}97d, Sam96a, Sta97a, Van96f, Wal97b, Coo97c, IEE96d, WBB^{+}97]. Professionals [Ano96-28, BLB99]. Professionelle [BM96a, BD98b, BBB^{+}96, Bacxxb]. professors [KW96]. Profiler [Ham97b, Wil99c]. Profiles [Net95c]. Profiling [LV99, Sun95a]. Profis [BB97c, Siy98]. profit [Ram97]. profound [FC96a]. Program [Bec97b, Ber96g, Cof96, C^{+}97a, Coo97b, DD97d, DD98a, Gre95, Hin96, Hob98a, MRA98, Pel95, Pit97, Tea97, Wygxx, Yeexxb, Zif96, ZJ96, Ano97-83, CFM^{+}97a, CFM^{+}97b, DD97c, IEE96d, LL98, Orc98, Rob98a, Way96b, Kro97]. Programacão [Dam97]. Programacion [AGxxb, Wal96p, BS98b, C^{+}xxb, C^{+}97c, Garxx, Goo97c, Jim97, LCxx, MN97, Morxxa, Thoxxb]. Programador [LJ97, LJxxd, LJxxc, LJxxa]. Programando [AGxxe, FC98a, Jep96, Rit97b, Rit97a, Ritxxd, Ritxxc, T^{+}xxa]. Programar [DDxx, Walxxb, Anoxxc]. Programas [McCxxb]. Programing [RRxx]. Programiranje [Suaxx, Suš97]. Programmable [IR99]. Programmar [Cha98b]. Programmare [Eng98, Hob97b, RVM97, Rit97c, WOLP96]. Programmatie [Anoxx-69]. Programmation [Anoxx-68, Anoxx-67, BS97d, CC98, Har97c, LO99a, LL97, MN96a]. Programmatore [Anoxx-45, LJxxb]. Programmazione [FLRT97, AG96c, Jaw97b, Jawxxa, Tis97]. Programme [SBH97]. Programmer [Ang99, Ano98x, Anoxx-37, Anuxx, AKM99, BB97b, BS99, Bro99a, Bro99b, CGI97, Cul98e, Culxxb, DDJ98a, Dav98a, Fre97b, Goo96c, GF99, HHW^{+}99, Wil97i, Joh99, Jos97a, Jos98, LJ96b, LJ96a, Lia99, McC96e, MP97b, Pel96, Rodxxb, Rou98, Sch99a, SO97a, SO97b, Soo98b, Sto97, WL97b, Wil97j, Wil97g, Wil97h, WC99, Wil99d, Wil99e, Wooxx, You97c, Gil96b, iMa96]. Programmeren [KK97e, KKxxc, Lem97f, Ritxxe, Thoxxc, MG98]. programmering [Ek97b, Ekxxc, EExx, Ekxxd, EE97]. Programmers [Anoxx-47, Boo96c, Boo97, Car96a, Coo97c, D^{+}98b, DSM99, Fla95, PM97c, Smixxa, Syd96b, Way98b, Amm98, DD96, D^{+}98a, Mar97a, PF97, Sch97h, Ver98, Bri99a, Ano96o]. Programmers/Microsoft [D^{+}98b]. Programmeur [Anoxx-37, Eng97d, LP96d, LPxxa, W^{+}xxb, Dan96f]. programmeurs [Dacxx]. Programmier [BS97e]. Programmierbuch [SBH97, BH96a]. Programmieren [Ano96-147, Bou96a, Cul98f, Frö96, Fröxxb, KS97, Rod96a, vanxxd, Job96, OHxxb, Staxxb, Woo97a]. Programmierhandbuch [MSS96, MS98b, FP96]. Programmierreich [NP96b]. Programmierrung [Bacxxa]. Programmiersprache [AG96b, AGxxa, Beh95, Mül96, Spi96]. Programmierung [AFPW97, Aitxxa, BBB^{+}96, BKP97b, CW97, CWxxa, DD97h, DD98b, Her97, Kie96, KS96b, Koc97, Küh96, Küh97, Lad98a, Pie97, Vac97, W^{+}97c, Amm97a, Ammxx, Bacxxb, Buhxx, CBJ97a, Cohxx, FKP97, FPxx, Kne98, KK97f, KWxx, Krü97b, Krüxxb, SST98, Sch97d, Schxxa]. Programming [ACM97a, ACM98b, ACM99b, A^{+}96, ALP99, All99c, Ano96-83, Ano96-137, Ano96-162, Ano96-185, Wil97e, Ano97-84, Ano98m, Anoxxg, Anoxx-40, Anoxx-41, Anoxx-43, Anoxx-48, Anoxx-49, Aoyxxa, AL96, AG96d, AG98, AGxxc, AGxxd, AW98, Atk99, AMPM96, AMP97, Bak97b, Bal98a, Bar96, BSS96, BS97b, Bis97, Bla98, Bor95, BOSW98, Bre96, BW99, Bri99b, Bri99c, BLT98, Bud98, CBT98j, Cad97, Cad98b, CW96a, CW98, CV98, CCFL98, CZFL98, CBA98, CS98, C^{+}xxc, Cho97b, Cog97, CBJ97b, Con97a, Coo97b, CCH97, Thexx, CH97g, CH97a, CA98, Cou98, Cox97, Dav98b, DH98, DF97, Dob98, Dub97, Dun96, EMe97a, EMe97e, EMe97d, EMe97g, EMe97f, EMe97i, Ene97, Eng99, Esp96, Espxx, FRT96, FR97, FC96b, FLW96, Fra96a, FD96b, FD96a, FI96b, FKPR96, FP98, Frixx, Fri98, FM97]. Programming [Frü99, GPF97, GFHM98, Gib98, GM97e, GJ^{+}96c, GTFxx, GYT96c, GYT96d, G^{+}xxf, G^{+}xxg, Gra96a, Gra96b, Gra96c, Gra96d, Gro96c, Groxxb, Hal97b, Hal99, Har96a, HFR96, HFR98, Har96e, Har97a, Harxxb, Hob97c, HB98, Hol98a, Hol96b, Hol96c, Hol97f, Hoq97, HS98a, Hug96, HHSW97a, HHSW97b, HHW97, HS98c, HC98d, Hus97, IF97, Inf95e, JLW96, JCxx, Jan97, Jarxx, JSA98, Jep97, JV97, KK97a, KFS97, KGTF97, Kim97b, Kim97c, KK96b, KK97d, KK98c, Laf96, Lafxxa, LOKK97, LO98b, LO99c, LP98a, Lea96a, Lea97b, LH97, Lea98a, Lea96c, Les97, Lev99, LB97a, LBS98, Li98, Liaxx, Lit97, LB96, Lut98a, Lut98b, Lut99c, Mah98, Man96e, Manxxb, McC98b, MT97, MO96, MHP98, Met95d, Minxxa, Minxxb, Mon99, Mor96b, Mor97f, Morxxe, Mos96, Net95b, NFP96, Nicxx, Nix98, NS96, OS96, O’N98b]. Programming [OH97c, OH98, OH99, OHxxa, PPIG98, PM96b, PM97b, PM97a, Pat98b, P^{+}98a, PHM99, PK98, PF96, Ree97, Rey97, R^{+}xxb, RS97, RMS97, Rin97b, Rit95b, Rit96b, Rit96c, Ritxxa, Ritxxf, Ritxxg, Rit97d, Rod95, Sad97a, SBB96, Sam96a, SR97, Shixxb, SBL96, Sip98, Rapxx, Ste97e, Ste97f, Ste99, SP97, Swa97e, Swa97b, Swa97a, Swa97c, Swa97d, Swa98a, Swa98b, SS98d, Syd96a, Taf96, TD97, Tay98, Thaxx, TPHB96, Thoxxa, Tho97a, Tol98, TPW96, Uzg96, V^{+}96, VD97, VD98, VR99b, Wal96g, Wal96h, WG99, Walxxa, Wal96m, Walxxc, WYxx, Wat98, WB99, Way96b, Way97a, Weh96, WA98, Wil97f, Wil98c, Win98, Wu99, Wuxx, Yal98, Yanxx, Yeexxa, Ziexx, Zif97, ZKM99, van95, ACM99a, Ano95l, Ano95r, Ano96-54, Ano96-61, Ano96-59, Ano96-102, Ano96-188, Ano97g, Ano97-74, AG96a]. programming [AC99, A^{+}99, BS96, Beh95, BAS99, BBC^{+}97, Bru96, BM98c, Cal99, CW96b, CS97a, Coo97c, Eck97a, EM99, EHS^{+}99, Gei96, GHM98, GGMS99, Gop96, Gre98a, Gre97c, GA98, GHP^{+}96, Had98b, Had99, Hal97a, Har98b, HCD^{+}98, Hem96a, Hor96c, HNS97b, HL97b, HL97c, HL97d, KPF96, KSW98, LQ99, Min96b, Man96a, MLM98, NMS96, Nev99, Pro96, Sch97h, SS98c, SS97, Sim96f, Ste96a, Ste96d, Ste96e, Tai98, Tay97, Tay99, T^{+}98b, TC95, TTD96b, WS98, Wat99, Way96d, Way98a, WSK97, Why96, WM97b, WK99a, WK99b, YB99, Zac97, vRvGS97, Eri96a, Gre97k, O’N97c, O’N97a, Ano97j, Ano99c]. programming-are [HCD^{+}98]. programming.java [DH97, DH98]. Programování [F^{+}xxb]. Programs [BC99a, Dan98, Fue98, Hir97, JLMPH99, LP98b, MST99, NE98a, NE98b, PH99a, Rod96b, Sch97c, SLCM99, Wak98b, Arm99, CH98a, CGHS99, CL97e, Cli98a, Cli98b, Cor98a, DP99, DIS99, GJ96d, HCD^{+}98, KCLZ98, McC96d, MJ99, MMS98, Miz99, Nap98, NAC99, RM99, Ude97b, Wak99, WR99b, Wol99, vRvGS97]. Progress [Bru96, Inf95a]. Progressive [MN96b]. Project [Wil97e, CD97, Cha98d, Cha98e, Kro98d, Kro98c, Kro99e, Ude97a, Ude97d, Ude97b, Wil97f, Ada98, Asl98, Ber98g, BM96c, DR98, DR99, Eng97a, GG99, MB96c, RCB98, TK99a, Ung97, DDJ99b]. Project-Processes [Wil97e, Wil97f]. Projection [NNS97]. Projector [MD99]. Projects [SD99, Rob99]. Projekte [FT97, Froxxa]. proliferate [Ano96-171]. proliferates [O’C95a]. Prolog [Ano98-29]. Promise [Law98, Moo95, Win95n, Far96a, RGT96].

Promised [Com95b]. promises [Joc98, Ste97a, Ude97d, Ung97]. Promising [Lou95b]. Promondia [GH97]. prompts [Ude97b]. prone [Mat96b]. Proof [FC98b]. Properly [Kim97a, Kimxx]. properties [LR98]. Property [Kin99, Ale98]. Proponents [Hal98a]. proportions [Why96]. Proposal [DK98, Lea99]. Propose [Gar98b]. Proposes [Gar97a]. proprietary [Ano96-54, Ano96-66]. Proprio [vanxxc]. Pros [von98, Gof97, Kin97b]. prospects [LMdS98]. Protect [Ano96-163, Knu97e]. Protected [Wal96f, Wal96c, Knu97f]. Protection [HCC^{+}98, Ale98]. Protocol [Hiu97a, Hiu97c, Hiu97b, O’B96b]. Protocols [Ano98-28, NK98, Fri96]. Prototype [BVG98, BG98b, BG98c, AGD97, BVG97, BG98a, FLQZ97, HGH96]. Prototyping [AFH98a, MK98, YK99, AFH98b]. ProtoVIew [Kro99a, Kro99b]. proud [Woo99a]. Provably [BS98c, BS98d]. Provide [Met95d, Ano97-72]. Provided [TMP^{+}99]. Provider [Tre99a]. Provides [Hor96a, Hor96b, Ano96z, Ano97-53, Ano97-83, Lin96b, Red97b, Wal96h]. Providing [AJ98b, PH99a, WWR97a, WWR97b]. Proview [MD99]. Proving [DE99, Pus99, Sym99, VN99]. Proxies [Lip99]. Proxy [Ano96-135, Ano98-28]. Pruning [BN99a]. PTR [Gea96]. PTR/Sunsoft [Gea96]. Public [Num98, Gau97r]. Publications [Bee96, Bee98]. Publishable [AL97]. publisher [McC96a]. Publishing [Ado95, Cro97, P^{+}98b, Sta97a, Ano98d, Bat96b, Fre95, Oga99, SD96, Tat96, DE96b]. Pulse [Ano96-164, Ano96-165]. punch [Gau97m]. pur [van97a]. Pure [Ano96-99, Ano97-50, Ano97-140, Ano98-27, Dar98c, GTW99, Pur97, Gre98f, Ano97-162, Gau97l, Wil99b]. purity [Ham97a].

Purpose [Ber96h, Ber96b, Way96e].

Purveyor [Var97a]. Push

[Ber97c, Vad95b, Hay96f]. Pushes

[DS95a, Ano96g]. pushing [Che97]. Put [Mil95b, Ano96-153, Ner97]. puts [Ang99, Cou96b, Sim96a]. Putting [Ano96-166, Lut98a]. puzzle [Han96]. PVM [BS96, Bod96, DLM99, Gei96, WCS99]. PVM/MPI [DLM99]. Python [Ang99, Hug97, Roo97, Tro97, Zho97].

Q [BA96, Ber96g, Ber96f, Ber96d, Ber96c, Ber96h, Ber96e, Ber97l, Ber97a, Ber97d, Ber97b, Ber97i, Ber97c, Ber97h, Ber97m, Ber97g, Ber97f, Ber97k, Ber97j, Ber97e, Ber98a, Ber98d, Ber98b, Ber98c, Cri98, Gri99, HL99, Pit98a, Pur98, Ran99a, Ses98, Soo98a, Wil99b, BC99a]. Q&A [AW99a, Eas99, GS98a, Moo98, Pit99, Ran99b, Tre99b, Wil99c]. QMASTER [Ano98-27]. QoS [TBS97, dRPT98]. qualitative [KPG96]. Quality [Frü99, MRA98, SMEN96]. Quantify [Ano97-165]. quantities [Ber98g]. Quantum [Mos98b, Mos98a]. Quark [Bat96a]. Quarter [Ano95n]. quasi [RB96]. quasi- [RB96]. Qube [Ano98-28]. queries [Ala99]. query [Ano98p, Asl98]. QUERYFLEX [Ano97-121]. Querying [MMM96, ZK98a, ZK98b, AS96]. Quest [Coh98, Gau97n]. Question [Lew97b]. questioned [Ano96-43]. Questions [Ano95t, Bra98, Kan96a, Kanxxc, WL97d]. Queue [Cri98]. Quick [Afe96b, Ano98m, DA96, Fla95, FFCM99, Fla99c, LS97b, San98b, San98c, Sha97a, Van97i, Ano96o, Fla97d, Fla97e, Lea98b, PF97]. quickly [Gre96c, Nan95]. Quickstart [GS96, Cor97a]. QuickTime [MS99a]. quite [Ano96-54, Hac98].

R [Ano96-77, Bay97a, LO99a, W^{+}98,

Ano97-128, Uni98b]. R-Active [Ano97-128]. R. [Ano96-77]. Race [DS95b, Nav97]. racy [Ano95r]. RAD [Ano96-86, Gau97k, Gre97d]. Radar [Ano96-165]. RADD [Kro97]. RAID [Ano96-28]. RAID-5 [Ano96-28]. Railroad [Meh97]. Raima [Ano97-119]. Raised [AJ99]. raises [Men96]. Raising [Sch97e]. Random [CMH99, Fic96, Knu98a]. rap [Hal98a]. Rapid [AFH98a, Ano97-85, Gre98e, MK98, AFH98b]. Rapida [Ano97-97]. rapidly [Gon97]. Raritan [Kuk98, Kuk98]. rates [Phi99]. rather [Bro98c, CM97a]. Rational [Ano96-167, Ano98-30, Kro98d, Kro98c, Kro98b]. rationale [DHM99a]. Raven [Ano98-28]. Ravi [KW96]. RC5 [GPO98a, GPO98b]. RDBMS [Ano97-118, Ano96-109, Ano97-36]. Re [Isk96]. Re-Engineering [Isk96]. Reaching [Bou96b, Wil96d]. Reaction [Ano97-98]. Reactive [BS98e, CR99, PSL^{+}98]. reader [Joh97a]. readers [Ano95v, Sun96]. Readies [Win95e]. Reading [Tuc98b, Hin96]. Ready [Bec97a, Bec98, BP99, OL98, Phi96, Uim95, Ven96, Gau96g, Har98c, Mac97c, Mac97b, van98a]. Ready-to-Run [BP99]. reaffirms [MS96]. Real [BW98a, Cla97a, DDJ99a, FC96a, GK98, Ham97c, HHW^{+}99, Kro98b, Lea99, Lut98b, MD99, MGL^{+}98, MKT98, Nil96, Ort99, PB98, ST96, TMH99, Way98b, Ano96-80, Ano96-171, BB97a, GH97, Kin97b, Nil98a, PV99, Sli97, van98a]. Real-Time [BW98a, DDJ99a, GK98, Ham97c, Kro98b, Lea99, MD99, MGL^{+}98, MKT98, Nil96, Ort99, PB98, TMH99, Ano96-80, Ano96-171, BB97a, GH97, Nil98a, PV99]. Real-World [Lut98b]. realities [Joh97a]. Reality [Ano96-135, BBE^{+}99, Bek97, Bru98, MD98b, SD99, Spe98, Gre98a, Ste96f, W^{+}98]. Really [Als95a, Cil96, DDJ98a, For98c, Hay96e]. Reasoning [JvdBH^{+}98, KPG96, KT96, RHND99]. reassures [DiD96]. Rebuilt [Gre97c]. receiver [Hav98]. Recipes [Way98a]. Recognition [S^{+}96b]. recommend [Sli97]. recommended [Dar98c]. Reconfigurable [DBBE98, LG97]. Record [IEE96g, Wae96, ACM97a, ACM98b, IEE96e, JADR96]. recording [Ano96-153]. recovery [SvG98]. Recreating [Red97c]. Recursive [Sak98]. Red [BH99a, Ano96-179, Kro99a, Kro99b, Swa97b]. redefine [Gau97q]. Redirector [Kro98d, Kro98c]. Redmond [Dav97b, Gau97g]. Reduce [Ano96-28, BN99a, MYD95, NF96]. Reducing [CKH99, McD98, MW98, Rit97d, Wip96]. reel [Gau97g]. Reels [Ano96-153]. Reengineering [McG97]. Reese [Ano97g]. Reestablishing [Fue98]. Refactoring [Lut99d]. Referência [FBxx]. Reference [Afe96b, Ano96-161, Ano98y, AFHxx, Bro97a, Broxxa, Broxxb, CL97b, Cha96d, Cha96e, Chaxxb, Che98, Coh96b, CMMN96, DA96, DS99, DW99, Fla95, FGN^{+}97, Fla98d, FFCM99, Fla99c, Fla99d, Fla99a, FB96, GK97b, Gra97a, Gra97b, Gri97a, HCF97a, Hor98b, Hun99, Jos97a, Jos98, LPM96, LS97b, MS99a, M^{+}97d, NS97b, NS98a, N^{+}96a, PLS98, PM97c, Pop97, San98b, San98c, Sch97c, SO97a, SO97b, Sha97a, Sta97a, Sta96, Van97i, Zuk97b, Ano96o, Ano97-150, Cha96c, Fla97d, Fla97e, GHP^{+}96, JC97, LC99, MMS98, NMS96, Par97, PF97, WFC^{+}99, O’N97b, Ano97i, Lea98b]. References [Bro96, Dem98b, Dem98a]. Referencia [NS97c, Car97a]. Referentiehandboek [BHxx]. Referenz [MS98b, N^{+}97a, SBH97, Her97, Hub97, Knu97h, Koc97, MSS96, Sta96]. Referenzbuch [BH96b]. Referenzteil [KS96b]. refills [Hay96e]. Reflecting [AH99]. Reflection [Ben98, KMS98, Par99, BN99b, Tre98]. Reflections [Ben98, Ben99a, Ben99b]. Reflective [KHB99a, KHB99b, SA98a, TBS97, TTK98a, WS99, TTK98b]. Refreshing [D’S96c]. refusal [SG97a]. Regular [LS99]. Rejects [Gar97a, GC97, Lea99]. Relational [Ano96-28, Ano97-86, Dre98, Gil97a, Gre98c, CSH^{+}98, Gre98f, Yan96]. Relax [Renxx]. Relay [Ano96-137, Sau96]. Release [Ano96-42, Ano96-91, Ano96-135, Ano96-137, Bra96, Inf95a, Inf95d, Gil97c, Red97a]. Released [Ano96-114, DDJ98a, MD99]. Releases [Ano96-146, Ano97-120, Ano97-129, Lea99, Sym95, Ano96-155, Ano97d, Ano97-135, Ano97-142, Ano97-163, Gau96e]. reliability [Wal96h]. Reliable [Str96]. reminds [Lav96]. Remote [Ber97e, Cli98a, KWB^{+}98, RWB97, San98b, San98c, Som97, Ste96b, TTK98b, WW97, Wal98a, ZMN98, Ber97b, Dry96a, Kam99, Led97, MvNV^{+}99, Mes97, PZ97, TTK98a, ZMN99, Dow98a, Rao99d]. Removing [BH99b]. renaming [SF96c]. Rendering [HJS98, Jan99, MD99]. Rent [Ano96-45, Hum96]. Reorganization [CDL99]. Repair [BN99a, Cha99e, MYD95]. Repeat [Ano97-52]. replace [KW96]. replacement [KH99]. replicated [BRS99]. Replication [BCH^{+}98, BCC^{+}97]. Report [Ano96-36, Ano96-37, Ano97-121, Bee96, Cul97c, Deu98, Dra98, Gof97, Hay97d, Hes97, IPDxx, McC95a, McC95b, RD99, Jav96c, SB97, Sch97b, STV95, WJUH98, JvdBH^{+}98, MHO96, Pfe99, Sha95, Ano98i, Due97b, Hig98]. Reported [Gar98a]. Reporting [Hig98, BB97a]. Repository [BW99]. Representation [NNS97, Tým99]. Representing [LST99]. Reproduce [Kan97]. Request [Pop97, Com95g]. Required [Hol97e, Hol97i, Zuk97a, BP97, BPxx, Hol97d]. requirements [MN96b]. rescue [Ano96-191]. rescue-middleware [Ano96-191]. Research [Ano98-27, AISS97, Ano96-188, Ciz97, Eng97a]. Researcher [JOA98]. Réseau [Har97c]. resilient [CTL98]. Resist [Ano97-56]. Resizable [Boy98a]. Resource [Bro96, CCYxxb, CvE98, CMSvE99, Fla99d, Har96c, Hus97, LP97a, SB96, Sri97b, Weh98b, Ano98y, HL97b, HL97c, HL97d, MP98, MP99, Tay97, Tay99, WBB^{+}97]. resources [Ano96-125, BNB^{+}98, EHS^{+}99, JJBW98]. Response [Cra97, JG96]. ressources [LO99a]. Rest [SMC^{+}96, CWHt99]. Restaurant [Tis97]. Restrictions [GC97]. Restructuring [BVG98, BVG97]. Results [The98, HGH96]. retirement [Ano96-151]. Retrieval [NNS97, TWP98, CCB96a]. return [Ano96-168]. reusability [GGL^{+}97, HC99a]. Reusable [Ano97-30, CAZZ98, Cel98, Gra99a, Ano96-181, Tra97]. Reuse [Ano98l, TWP98, Lak96, Pic96b]. Reveal [PW97]. Reveals [Hal98b, Win95j, Gre98b, KP96, Kin97b]. reverse [Dug99]. Review [Ano96n, Ano96o, Ano96p, Ano96-161, Ano97h, Ano97g, Ano97m, Ano97j, Ano97i, Ano97k, Ano97l, Ano97-108, Ano97-109, Ano98f, Ano98j, Ano98i, Ano98h, Ano98g, Ano98e, Ano98k, Ano99a, Ano99c, Ano99b, Bri99a, Bru97, Cla97c, Far96b, Fla97h, Fla99a, Fre97a, Gre97j, Hay97d, Hec95, Heu97, Hum98, Kay97, Lea96b, Lea96c, Lea97c, Lea98b, Mat96b, Mat99, MF97d, Nic96a, NFP96, Nic96b, Nic96d, Nov98, O’N97c, O’N97a, O’N97b, O’N98a, Roo98, SMC97, Str97, Van97m, Yal98, vD99b, McC96e]. Reviewing [RB96]. Reviews [Ano97n, Ano97-153, Bay97a, Bra96, Bra97a, Bra97b, Kri96, Sal96, Tee99, Ver98]. Revised [Ano98k]. Revision [DK98]. Revisited [Ude97a]. revives [Mag95]. Revolution [Phi96, RK97a, SBxx, Sut97, Cou97b, Jan96, Sko96]. revolutionize [Kar95a]. rewards [Ano96-195]. RF [Ano96-153]. ribbons [Ano96-196]. Richard [Ano99a]. Rico [IEE99b]. Ride [Win95d]. Rides [Ano98u]. Right [Kri96, Lav96]. Rio [IEE99b]. ripens [Gau96h]. rips [Mat96b]. RISC [Ano96-126]. Rise [Ano96-27, Ner97]. Rising [Car96b, Buc96b]. Risk [MAS96, Ano96-54]. Risks [Anoxx-50, Fel97, MF97c, McC98d, Neu97]. Rival [Gro96b]. Rivalry [OH97b]. Rivals [Win95b]. River [Tee99]. RMI [Küh97, Ano98t, BCH^{+}98, BRF98, BDV^{+}98a, BDV^{+}98b, Cha98d, Cha98e, Dow98a, Joh98, KS98a, KS98b, KWB^{+}98, PS98, RBF98, RMB^{+}98, Ses98, Sim97e]. RMI- [KS98a, KS98b]. Road [Wat95a, Ano98-35]. Roadmaps [Ano97c]. ROASTER [Nat95]. Robert [Ano97h, Bow96b]. Robotics [Fre97a]. Robots [IEE98a, Lim95]. robust [Amb97, Dar98c]. robustness [Shu97b]. Rockville [IEE96f]. Rockwell [Ano97-130]. Rogue [Ano96-172, Ano97-131, Kro98e]. roils [SG97b]. role [CC97]. Rolls [Cox95d, FC95]. ROM [Buhxx, Dav96b, Fla99d, GG96a, Kay97, Mon99, Neo97, N^{+}96b, Ste96d, Ste96e]. Rome [Nat99]. ROMs [BDF96]. Room [BCP^{+}99b, Rob96]. Roots [Nas95b]. Rose [Ano96-167]. Route [Gre97d, WBC^{+}96]. routers [Ano96-133]. routines [EM99]. Rover [San98a]. RS6000 [McK95a]. RSA [Ano97-132, Ano97-133, Way98c]. RSJ [IEE98a]. RST [Kro98e]. RTI [Ano97-134, Ano97-135]. RTOS [Kro98e]. RTOSs [Ano96-108]. rule [Sim96f]. rule-based [Sim96f]. Rules [Dun99, Sim96c, Ano97-43, Ber99a, BDKK96]. rumble [Mac97c, Mac97b]. Run [BP99, CS98, Gri99, Lut98b, Ano97-70, Eck97b, Eck97c, Kar95b, KG96a, KG96b, LP96a, MF97a]. Run-time [CS98, KG96a, KG96b, MF97a]. Running [Kim96, WB^{+}96c]. Runs [Gar97b, Ano96-59]. Runtime [Ano99g, DBBE98, JRE98, Kro98b, MB98c, S^{+}96b, CFM^{+}97a, CFM^{+}97b, Orc98]. Rush [BCC^{+}97]. Rusty [Ano97j]. Rx [McC97b].

S [McC96e, FC96a, Lin96f]. S.u.S.E [Ano98-28]. Sadun [Tee99]. Safe [BOSW98, DE97, Nan96, Nev99, NvO98, VM96, GA98]. Safe-Tcl [GA98]. Safety [DDJ99a, MF97a, VN99, Hay96d, Mey96b]. Saga [Ban95a]. said [SF96c]. Sales [Gar97a]. Salesman [Led98b]. Salt [Isk96]. Samba [Ano98-29]. same [LH96]. Sammlung [Ano96-67]. Samples [O’B96b]. SAMS [LC99]. San [ACM98b, ACM99a, Ano98a, BUP99, CEG96, IEE96h, KJR98, PZ96, RCB98, DF96, Ten99, Uni96a, Uni96b, Uni97b, Uni98b, USE96d, USE98d, USE98f, USE99c, Wae96, W^{+}98, YN96, Pol98]. Sandbox [GMPS97, Gos98b]. Sane [Ano98-28]. Santa [IEE96c, USE98b, Swa97e]. SAP [KW96, RGT96]. SAR [Gol99]. SAS [Ano96-160, Ano98-34, Hoy96]. Satellite [Gar97a, Wip96]. save [Ano96-151, OHE97a]. Savvy [Ano98-30, Kro99a, Kro99b]. say [Ham97a, Sun96]. Says [Ano96-112, Ano96-175, Ano96-176, Gut95c]. SBIA [BK96a]. SC’99 [ACM99c]. Scalable [KB98a, ZLGS99, Cli98a, KB98b, Woo99b]. Scale [JPK98a, Fox98b, JPK98b, Lak96, MC97]. SCAMC [Cim96]. scanner [Ano96-196]. scans [Bra96]. scarce [Ste97a]. Scene [Net95c]. scenes [Rob96]. Schaltungen [Web98]. schedule [Gau96h, San98a]. scheduler [KH99]. Scheifler [Bow96b]. Scheme [Gut95h, MB96c, AH99, BM96c]. Schildt [Sch96a]. Schlumberger [Coh97c]. schmava [Ano96-73, Ano96-102]. Schmidt [Kal95a, Win95b]. Schnell [Sch98b, Shu97a, SH97]. Schnellübersicht [Anoxx-79, Lin97, Schxxb, Sch97g]. Schnittstelle [Her97]. School [Ano96-28, The98, Ban95a]. Schools [DDJ98b, DDJ98c]. Schulung [Ano96-70]. Schwab [Ano97t]. Science [AL97, DDJ98b, DDJ98c, Wil97i, Joy98, KMR98, Wil97j, BNB^{+}98, BGGR98, CH98a, Fox97a, Fox97b, HCD^{+}98, DF96, SBB^{+}97, SZTR96, SPV99, W^{+}98]. Scientific [BW99, CCFL98, Dub97, KK98a, MG97a, Syd97, von98, AJ98a, DAB97, FF97, JA98, KK97c, MLM98, Vil99, Zac97, Ano97-153, Bra97a, Bra97b]. Scientist [Van96d]. Scientists [Cha99d, Dav99, Gau97h, JG96]. Scivis [KK98b]. SCM [Ano98m]. Scope [Gra96b]. Scopus [Ano97-136]. score [Bar97]. Scores [Ano96-115, Win95l]. Scotland [AOV^{+}99]. Scott [Ano98j, Ano96-112, Ano97-52]. SCREADER [Ano98-27]. Screamer [Ano98-28]. Script [NJ95, P^{+}98b, Sam96b, Nov98, Par95d, Pel95, T^{+}xxb, Gol96a]. ScriptEase [Chw98]. Scripting [BKT96, Cor97b, Hay95e, Net95a, RFM95, Ste97e, Bot98b, GJ96d, LS99]. Scripts [ER97, Lut98a, Ste97d]. ScriptX [Kar95b]. Scrolling [Ste99]. sculpting [And96a]. SDK [Kro99a, Kro99b]. se [Anoxxd]. Sea [GC97]. Seagate [Ano98-30]. Sealing [GS98b]. Seamless [CV98, CKV98]. Search [Chuxx, Kie99, Shixxa, Gau97r]. Searchable [Fla99d]. Searches [KP99]. Searching [Boy98b, DDJ98b, DDJ98c, TMP^{+}99]. Seattle [CI96, IEE96g, USE96a]. Sebastian [DF96]. Sebastopol [Ano96n, Ano96o, Ano96p, Ano97h, Ano97g, Ano97j, Ano97i, Ano97k, Ano98f, Ano98j, Ano98h, Ano98g, Ano99a, Ano99c, Ano99b]. Second [Add99, CLK99b, IEE98b, Kno95, Kno96b, Nat99, Nov98, S^{+}96a, USE96c, Web96, Ano98f, van96b, Ber98h, Ber98i]. Second-Generation [Web96]. Second-Tier [Kno95]. Second-wave [Kno96b]. secondary [Gop96]. secretos [Harxxc]. Secrets [Ano96-120, CC96, Har96d, Lut98b, Morxxc, Sha96b, Tit96a, Titxxa, Titxxb, Hay97a]. Section [Ber98i, Cul98g, Ano96-188]. Secure [Ano97-119, Ano97-137, BCE98, GS98a, HBL99, IGO98, LP96a, MAS96, MK99, Zha97, Ano97e, For98c, LP96b]. secured [Way98c]. Securely [Ber98d]. Securing [MF98, MF99, VPM98a, VPM98b].

Security [Ano97-80, Ano97-119, Ano98l, Anoxx-50, AKM99, Ber97h, Bou97, CBT98h, Fel97, For98d, Gar96b, GS97a, Gol97, GMPS97, Gon97, Gon98, Gos98b, GA98, IEE96a, IEE96h, IAJR97, Jaw99, JLT98, KG98a, Kro98b, Kro99c, Küh97, Lea97c, LR98, Lit97, Mac98, MAS96, McC98d, MF97b, MF97c, MF97a, MS98a, Neu97, Oak98, PSW^{+}98, Pis99, Rao98d, SH96, Ste97g, USE98f, Yel95, Ano97v, Ano98b, Ano98s, Ant96, Che99, DFW96, Gon99, Hay96a, Hop97, KNNL98, Mar98b, Mar99a, McC97b, McC98c, MF96b, Men96, Mey98a, Mey98c, Mey99, Nic99, Ste96c, Wal96h, WBDF97, Kro99c, Roo98, Ano98j]. See [Hay95e, Shu97b, Bar95b]. Seek [Gar97c, Kin96]. seeks [Gau96i, Mat96b]. Seems [Sin95a]. Sees [Gut95i]. Sega [Gut95e]. Segmentation [KSI^{+}96]. sehen [BKP97b]. Seiten [DE96b, KS96b, vSS96d, DD97h, DD98b]. sekai [DM96a]. selbst [DE96b]. selbstgemacht [vSS96d]. Selected [Ber98h, Ber98i, Cul98g]. Self [AW99a]. Self-Executing [AW99a]. Selling [Bro96]. Semana [Dan96e]. semantic [Doe95]. Semantics [AFL99, BS99, CKRW99, DK98, DE99, HBL99, AF99, CKRW97a, CKRW97b]. semi [vRvGS97]. semi-automatic [vRvGS97]. Semiconductor [Cox95c]. Seminar [Nev99, Ast98]. Send [Ber96f]. sender [Hav98]. Sense [SMC^{+}96]. sensible [PL97]. Sensing [Ste96b]. sent [Ano96-73]. sentence [Ano98p]. September [Abr96, Ano96-142, AOV^{+}99, DLM99, Uni96a, Uni96b, Uni97b, Uni98b, USE98d, USE98a, YH96]. Sequential [CV98, PHM99, CKRW97a, CKRW97b]. Serial [Ano96-137, Sim97c, Ano97s]. serial/parallel [Ano97s]. Serialization [Hav98, SHR^{+}99]. Series [CW96a, GM98c, Minxxb, Smixxa, Smixxc, Smixxb, Ano97-149, Far96e, Sto97, Ano97m].

Serious [Ano96-86, Ano97r, Hal97a, Sey96]. Serve [Ven96, Pic97, Wal96h]. Served [Ano96-99]. Server [Ano96f, Ano96-135, Ano96-137, Ano97-29, Ano98n, Ano98-30, Ano98-31, A^{+}99, BCH^{+}98, BFGS98, Dow98b, Frü99, GDS98, Ham97b, Hor96a, Hor96b, HK99, Kro99a, Kro99b, Kro99d, LaM95g, Lot95, Mey99, MKT98, Nov97, OH97c, OH98, OH99, OHxxa, Que97, San98b, San98c, SWJ97, TK98, TK99b, TMM^{+}98, TMMH99, Uma97, WM97a, Woo99b, Ale96, Ano96u, Ano96-42, Ano96-53, Ano96-72, Ano97r, Bar97, Bij96, Bla96, Cal99, CH98b, Cli98b, Dun97, Gre98d, HL97d, KW96, SBB96, Sco96a, Wil97c, YB99, Yos97, Ano96d, Ano96-137, Ano97-118, DiD97d, Hus97]. Server-based [Mey99]. server-oriented [Cli98b]. Server-Side [Dow98b, WM97a, HK99, Cal99, Dun97, HL97d, SBB96, Woo99b, Hus97]. Servers [Eri96b, Hes97, Mil95c, Hib96, KW96]. Serves [Sim97d, Ste97c]. Servet [Ber97k]. Serveur [N^{+}xxc]. Service [Ano96-28, MC97, TB96]. Services [BD99, BB99, Bus95, Gut95e, MF96a, Urg96b, VPM98a, VPM98b, MC97, NJC96, TB96, TSKA^{+}96]. Servidores [Gar97e]. Serving [Fis96]. Servlet [HC98d, Wil97b, Ano99c, Klu98, Ude97b]. servlet-based [Ude97b]. Servlets [Cli98b, Har97b, McD99b, Mos98c, Pit98a, Ude97d, Wel99c, Wil99a, Cal99, Goo99, WM97b]. Session [WBC^{+}96]. Set [Ano96-91, AC98, Cha96e, Cro95a, EL98, Inf95d, Inf95f, Kro99d, MPxx, Ude97c, Ano95g, Ano95h, Che97, Fox98b, Lav96, Way96c, Cox95d]. Set-Top [Kro99d, Way96c]. Sets [Gol99, Hec95, MW95, Win95h, Cou97d]. Setting [Coo97a]. Setups [Ste96d, Ste96e]. Seven [Cla97c, Dra97]. Seventh [IEE99b, USE98f]. Several [Gar99, Wat95c]. SGI [Ano96-177, Ano97-138, Gar99, Gre97h, Par95g, Win95k]. SGML [Ano97-30, Ano97-108, Ano97-139, O’C97, Sha97b, SD96]. Shaio [McC96e]. Shake [Ano96f, Car95]. Shared [Kar98a, EHS^{+}99, Kar98b]. Shareware [Anoxx-53, Ano96-151]. Sharing [Dal98, MK99, Tes96a, BRS99, CGJ^{+}98]. Shatters [Cro95b]. ShBoom [Ano97-126]. Shedding [Lav96, Lev96d]. Sheets [P^{+}98b, SR97]. shell [Gop96]. Shells [Lut99e]. Sheraton [Cim96]. Shiang [Kanxxb]. Shien [Flaxxb]. Shift [BN99a, Gut95c, Ham96b, MYD95, SF96a, SF97, BGGR98, Chu98b]. Shift-Reduce [BN99a, MYD95]. shine [Ano96-171]. shines [BH96c]. Shiny [LH96]. ship [SG97a]. Shipped [MD99]. shipping [Ano97-43]. Ships [Ano97-29, Ano97-123, Ano96-195, Ano97-44, Ano97-47, Ano97-50, Ano97-46]. shiru [DM96a]. Shockwave [Salxx, Ano96-129, Kar96, Nel95b]. shogeki [IK96]. Shoots [Ano96-178]. Shop [Ano97-118]. Shopper [Ano96-151]. shops [Ano97-28]. short [Hay96n, Sul95a]. Shortage [DDJ98a, Gar98a]. shortcomings [Nan95]. shortening [Rit97d]. Shots [Wu97]. shou [Nau96b]. Should [JG96, McC97b, Saa96, Fla96a, NGM^{+}97, Sch97f]. Show [Sha95, Nas95a]. showcases [Sul95a]. showgoers [Wag96]. Shows [Det97]. shrug [Hay96g]. siblings [Chw98]. sicher [SH97, Sch98b, Shu97a]. Sicherheits [SBH97]. Sicherheits- [SBH97]. Side [Dow98b, WM97a, Al-97, Cal99, Cla97d, Dun97, HL97b, HL97c, HL97d, HK99, SBB96, Sch97f, Tic97, Woo99b, Hus97]. side-effect [Cla97d]. sides [Ano97-91]. Siemens [Ano97-143, Bro97c]. SIG [Couxx]. SIGACT [ACM97a, ACM98b, ACM99a]. SIGCSE [Joy98]. SIGGRAPH [ACM99d, Sha95]. Sights [MW95]. Sign [Ano95e]. Signal [GTV98a, Kro98d, Kro98c, LR97, GTV98b, Har99d]. Signature [GM98c, JM97d, PS97, van97c]. Signed [Ber97g, Ber98a, Bri99b]. Signing [GS98b]. SIGPLAN [ACM97a, ACM98b, ACM99b, ACM99a]. SIGPLAN-SIGACT [ACM97a, ACM98b, ACM99a]. SIG’s [Cou97f]. silence [Mac96]. Silicon [WBC^{+}96]. SIM [Por98]. Similarities [BM98a]. Similarity [NNS97]. Simon [Doo99]. Simple [Anoxx-51, SMY99, O’C99, Sim96f]. simpleIO [WK99a, WK99b]. Simpler [Pel95]. simplifies [Ano97-111]. Simplify [Lea99]. Simpson [Mag95]. Simulated [DDJ99a, Str99]. Simulating [TKH99, RHND99]. Simulation [BUP99, Cow98a, Kro98b, MD99, Web98, vHO98, Cow98b, EKR97, Fox97a, Fox97b, FF97, VO98]. Simulation-Based [Kro98b]. Simulations [BSS98a, IR99, KMW99, AA97, BSS98b, CT96, DW97b, HNS97a, Tob99, WCL97]. Simulator [Kro98d, Kro98c, KCPC98]. simultaneously [JGKR96]. Sin [Ten95a]. since [Ban95a]. Single [Kro98e, Mon97]. single-chip [Mon97]. Single-Source [Kro98e]. sipping [Hay96t]. Site [Ano95q, Ano97-112, CCB96b, Coo97a, D^{+}98b, Fet97, HFR96, HFR98, Rei96, WB^{+}96c, Ano96-202, Bal96a, Mag95]. SiteMill [Ano96-195]. Sites [Hob98b, SS96, Win95k, Ano96-202, Atk99, Fet97, O’C97, vSS96b, vSS96c]. sixth [USE98d]. Sizes [Win95b]. Skeleton [Pat99]. Skills [BLT98, Wil96a]. Sky [Kro98e]. Skyline [MD99]. Slap [Zol96]. SLE [WOC99]. SlickEdit [Ano96-137]. Slim [Ano96-169]. Slow [Bet95, Ano96-153, Dys96, Gau96e, Hal98a]. Slowly [Cox95b, Hay96t]. Slows [Ano96-28]. slug [GH96a]. SM [MD99]. Small [Bet95, Chw98, Nil98b, Wil99e, Lim95, MC97, YRE98, NRG^{+}99]. Small-Footprint [Nil98b]. Smaller [DDJ98a, DDJ98a]. Smalltalk [EK98, Gau97c, Sut96]. Smart [DDJ99a, DDJ99b, Gar97a, Mil99, PV98, Gau97q, GLV99, Bro97c, HNSS99]. Smartcard [USE99d, Ano97-143, Bam97]. Smarter [BB97b, Pet97]. Smarties [Cel96b]. SmartStream [Ano96-72]. Smashing [DiG96]. Smell [Gut95g, Win95l]. Smells [Par95f]. Smith [Swa97b]. SML [BK99]. Smoking [Mac97d, Mac97e]. SNiFF [Ano96-170]. Snippets [Cue96]. snubbed [Kee96]. Soapbox [Pon98]. Soar [Gar97c]. Social [Sch96c]. Society [IEE96c]. socket [YB99]. sofa [Ano96-53]. Sofia [Ano97-139]. Soft [Kro98d, Kro98c]. SoftPlus [Ano97-140]. SoftQuest [Kro99e]. Software [Ano96-49, Ano96-135, Ano96-171, Ano96-172, Wil97e, Ano97-41, Ano97-119, Ano97-121, Ano97-122, Ano97-141, Ano98l, Ano98-27, Ano98-32, Ano98-33, Anoxx-69, Boz95b, Buc97, Buc96b, Com97a, Com99, Cow97, DDJ99b, Fox99, FW96, Frü99, Gar97a, Gil98c, Gre97j, Hew99, Ham97b, Ham97c, Hie97, Hin98, IEE96k, IEE96j, IEE96s, IEE98b, Jar97, Kro97, Kro98d, Kro98e, Kro98c, Kro98b, Kro99a, Kro99b, Kro99d, Kro99e, Kro99c, Kuk98, Lew97a, Lut99d, Lut99e, MD99, Mas95, MBH97, McF98b, McK95a, MRA98, Mes95c, Mey97, Nat96, Pit97, Ric95c, SMEN96, Sch97a, Sha96b, Stu98, SS96, Sut97, Tro97, Uni96a, Uni96b, Uni97a, Uni97b, Uni98a, Uni98b, Vad95a, WKES98, WBC^{+}96, Wil97f, WR98b, WRxx, You96b, ZMN98, Ano95s, Ano96-77, Ano97-47, Ano97-153, AA96a, BN98b, Bra97a, Bra97b, CAB^{+}98, Fla96a, GG99, Gol96a]. software [GGL^{+}97, Hoy96, JS97, Jon98, Lak96, Lew96a, LL98, Mes97, Oue96b, Pir97, SvG98, Str96, SvK98, Vck96, WKE98, ZMN99, pP98, Ano96-183, Ano96-72, Ano97-46, Ano98-31, Boo96b, Bru97, Kro97, Kro98d, Kro98e, Kro98c, Kro98b, Kro99a, Kro99b, Kro99d]. software- [Gol96a]. SoftwareTraining [Anoxx-70]. SoftWired [Kro98a]. Solaris [Cox95d, Inf95d, Vij97a, Vij97b]. SOLID [Ano96-137, Ano98-27, Lin98b]. Solstice [Anoxx-71]. Soluciones [Afe96a, Dar96]. Solution [Ano96-135, Ano97a, Ano98-32, Ano98-33, Hor96a, Hor96b, Kro98e, Law98, MD99, SA98a, Ano95w, Ano98s]. Solutions [Ano96-137, Ano98-27, Ano98-28, Ham97b, KS98a, KS98b, KSB97, Kro97, Kuk98, MD99, MRA98, Tro97, Ano97-53, LL98, Ren98, Way98a, MD99, San98a]. solve [Ano96-90]. Solver [PPJ96, SLO96]. Solving [BS98a, DS99, KW98, Loe97, Wei98]. Some [Ano95d, Ano96-204, Flo97, Fri98, Hay96s, Hay96t, LMdS98, O’B96b, Pry96, Swa97c, Swa97d, Ano95r, Lev96d, Lin96a, LQ99, Sim96f, Ubo96b, Ude97d]. Somerset [IEE96d]. something [Gre97f]. Sonic [Kro98b]. Sons [Ver98]. soon [Gre97h]. Soothsayers [Cel96a]. Sophisticated [How96, TJ96, MF97a]. sophomore [KCPC98]. sophomore-level [KCPC98]. sortieren [Til97c, Til97b]. Sorting [Boy98b, Lei95, RGT96, Ano96-61]. Sotwareentwicklung [BD98b]. Sound [MGL^{+}98, SB97]. Soundness [DE99, Pus99, Sym99]. Soup [Hal98a, Way98a]. Source [Bee96, Int96a, Kro98e, PF96, PW97, Jav96c, Urg96b, Urgxx, GF99, Hac98, NRG^{+}99]. SourceAgain [Kro97]. Sourcebook [Anu97, CCB96b, CS97b, Jar97, JD97, LB97a, McCxxb, Anu96b, Anu96a, McC96d]. SourceCraft [Ano96q, Wag97b]. Sources [BM98a]. South [Cul97c]. space [Chw98, LK99]. Spaces [Mil99]. Spain [Ano97-139, DLM99, DF96]. Spar [vRvGS97]. SPARC [Bab97b]. sparse [CCCL97]. speak [Ano95v, Ram97, Hay96l]. Speaks [MH97b, Cor95, DB97]. Spec [Cas96a, DDJ98a]. Special [Anoxx-71, BM98b, Ber98i, Bro96, CWHP96, Cul98g, Doh98, Gof97, Hay97d, KS96a, LTxx, N^{+}96b, New96d, N^{+}97a, New98, Que97, Rei97a, Rei97b, RWxx, San97b, SB97, Sch97b, SA97, TJ96, Wal97c, WB^{+}96a, WBC^{+}97, WRTM98, Webxx, W^{+}97b, Wil97b, Newxxa]. Special/Anwenderpraxis [KS96a]. Specialization [SLCM99]. Specific [USE99a]. Specification [Ano96-114, Coh97b, DS97a, Gol98a, GJS96, G^{+}xxh, G^{+}xxi, H^{+}xxa, Lea99, Lia99, LSxx, LY97b, LY99, Pus99, Qia99, SRD98, Sun95b, SUN97b, JCV99, VN99, Ano97-147, AWO^{+}99, DHM99a, MKS^{+}96]. Specifications [DHM99b, HGxx]. Specifying [Ber99a]. SPECjvm98 [DH99]. Specs [Ano96-113]. Speech [S^{+}96b]. Speed [Ano96-71, Ano99d, Gut98, Kim96, Eck96a, Eck96b, Gau97f, Gau97m, Hal98a, Ham97a, Ham97b]. Speeding [DDJ99b, Hib97]. Speeds [MD99]. Spell [Kro99c]. Spell-Checking [Kro99c]. SpellCaster [Ano98-27]. Spezifikation [LYJ^{+}97, LYxxa]. SPICE [SHGD98a, SHGD98b]. spill [Ant96]. spills [DiG95b]. spin [Ano96-153]. Spinning [Fis96]. Spire [Ano96-135]. splits [Ano95a]. splitting [CKH99]. SPMD [CZFL98, HNS97b]. Spoofing [HS98b]. sports [BB97a]. Spot [DS95b, SS96]. Spotlight [Wal98a]. Sprach [JM97c]. Sprach- [JM97c]. Sprache [Dalxx, Dal97]. Sprachen [Die96, Diexx]. Sprachkonzept [SBH97]. Sprachspezifikation [GJS97, G^{+}xxc, Gosxxc]. spread [Gau97o]. Spyglass [Ano95e, Mil95a, Par95d]. SQL [JRC97, Ber97d, Cel96b, EM99, Küh97, Lam99, McC98b, Sip98]. SQLJ [CSH^{+}98, EM99]. SQml [San98a]. squares [Rob98a]. squashes [Ano96-43]. SRAM [Ano96-153]. SSC [Ano98-29, Ano96-161]. SSL [Ano98-28]. St. [ACM97b]. Stab [Win95i]. Stack [WH99, MW98, Tým99]. stack-based [Tým99]. stances [SG97b]. Standard [CLK99b, Com95b, Dob98, EMe97e, Gar98b, Law98, Lea99, Mes95d, Nat98, RD99, Sch97e, Ude97c, ZMN98, Gau96c, Gau97n, GHP^{+}96, PL96, Ric96d, ZMN99, Ano97-153, BKR99, BCE98, Bra97a, Bra97b, CS99, DDJ98a, Gar98b]. Standardization [PJR^{+}98]. Standardize [WK95]. Standardizing [Mur96b]. Standards [FFHA98, Fra98a, Gar97a, Gar97d, Gar98b, Gro96b, Har96f, Hec95, JN95b, MR97, RFM95, Ano96-151, Isa96, JN95a, Lin98b, RB96]. Stanford [IEE96e, IEE96m, IEE97, IEE99a]. Stappen [Anoxxo]. star [MF96a]. Starbuch [McC96e]. Starfish [Mil95c]. Stars [Buc96b]. Start [Ano98m, Gau96e, Pic96c]. StarTech [Ano98-27]. Started [BC97a, CBT98g, C^{+}xxa, CB98]. Starter [Ano96-83, Ano96-104, Ano96-103, Anoxx-52, Anoxx-53, Per97b, Sam96b, Ano96-105]. Starting [Bay97b]. starts [Ano97-164, Cou97b, Mat96b]. Starwave [Ubo96b]. State [Ano98l, Fue98, IEE96d, IDK98, Pla96c, RWWB96b, RWWB96c, RWWB96a, Tub95]. State-of-the-art [IDK98]. states [MC98]. Static [Cro95b, MB98b, Mye99]. statically [Bok99]. statically-checked [Bok99]. Statistical [Doe95, WO97]. Statistics [Mur99a]. status [Oua96]. Stay [Gar97b, Oua97]. steal [Ano97-71]. stealthy [CTL98]. Steam [Ric96c, Win95e]. Steaming [Wil96a, Bab95]. Steelman [Whe97]. Step [Anoxxm, Bec98, Lut99e, OL98, Amb97, Ano96s, Ano97-142, Sha97b]. step-by-step [Amb97]. STEP/EXPRESS [Sha97b]. Stepping [GO96]. Steps [Bec97a, Hay97d, Cobxxb]. Steroids [HBG^{+}97]. Steuere [DE96a]. stick [Ano96-151]. Still [Ano96-110, Col96c, Cox95b, Sch97b, Gil96a, Ham97a, Hay96n, Hay96q]. stink [Hay96s]. Storage [Ano97-82, Ano97-103, BS98a, Ano96-153, Chw98, Kam99, Ude97b]. Store [AW99a, HL99, Jar97, Gre98b]. stories [IEE96b]. Storing [Dre98]. Storms [Win95f]. Story [Ano96-27, O’C95b, Van97n, Ban95a, Joh97a]. Strategies [Gol99, Sun97a, Sunxxc, DR98, DR99]. Strategy [Gar97a, Mic95, Win95j, Ano96-49, Car96c, EHS^{+}99, WSL96, Won97]. Stream [Key98]. Streaming [Moo98, Rus98, ST96]. Streams [Rao99c, Eck97d, Ano98-28]. Street [Cha96a, Gau96b, Law98, Tom96]. Strength [CDBM97]. stress [MC98]. strict [KCLZ98]. Strictly [Smixxa]. Strike [Com95e]. stripped [Way96c]. stripped-down [Way96c]. Strong [Gar97c, SG97b, Ubo96b]. Stronghold [Ano97-119]. Structural [CKRW99]. Structure [CDL99, RLV^{+}96, SN96]. Structure-Based [SN96]. Structure/Class [CDL99]. Structured [Gib98, Miz99]. Structures [Boy98a, Carxx, GT97, Laf98a, Li98, Mai98, Row97, Sha98a, Sta98a, Staxxa, Wei98, BBG^{+}99, TGGP97, Wei97c]. struggle [Hay96w]. Struktur [JM97c]. student [Ciz97, Nap98]. Studenten [Bel98]. Students [BP98, Gor96, Pfe99, Tob99, WS98]. Studio [Red97c, TA97, BM98b]. Study [Bin97, CJJ^{+}98b, DH99, JPK98a, RH97b, RH98, RHE99, WBA^{+}96, AGD97, Ban95a, CJJ^{+}98a, Die97, Gau97h, JPK98b, Xu97]. studying [Fox98b, SH99]. Stuff [Ham97b]. Stunden [Frö96, Fröxxb, Wes97b]. stunning [Ano96-151]. Stuns [SG97a]. Style [FLW96, GG96b, GG97, Joh96b, Lan99b, P^{+}98b, SR97, Wor97]. Styles [Laf96, Lafxxa]. sua [Ben96b]. Subito [Dadxx, Pew97c, Pewxxb, Jam97a, Jamxxa]. Subjects [Ten95a]. Submit [Law98]. Subroutines [HT98, Qia99, SA99, O’C99, SA98b]. Subscribers [GC97]. Subset [SUN97b]. substance [Wor97]. subtypes [D’S96a]. success [Hea95, Ude97b]. sucesso [Ben96b]. such [Tur97]. Suchen [Til97c, Til97b]. SugarCubes [BS98e]. SUGI [Ano96-160]. Suite [Ano96x, Bha98, Kro98e, Kro99a, Kro99b, Mes95a, Pep97, Ano96-32, Ano97-148, Ano97-158, Pic96a, Pic96d]. Suites [Kro99d]. Summer [Ano96-28, Gau96c]. Sun [EMe97g, Gau97l, Hum98, Sim97e, Vij97b, Ano95h, Ano95o, Ano95l, Ano95n, Ano95p, Ano95q, Ano95r, Ano95s, Ano96f, Ano96-27, Ano96-53, Ano96-54, Ano96-48, Ano96-74, Ano96-97, Ano96-112, Ano96-132, Ano96-181, Ano96-173, Ano96-174, Ano96-178, Ano96-175, Ano96-176, Ano96-177, Ano96-179, Ano96-180, Ano97d, Ano97-52, Ano97-129, Ano97-144, Ano97-146, Ano97-145, Ano97-147, Ano97-150, Ano97-148, Ano97-149, Ano98-35, Ant96, App95, Bab95, Ban95a, Ban95b, Bor95, Bou96b, Boz95a, Boz95c, Boz95b, Boz95d, BH96c, Bro97c, Car96c, Car97b, CP98, Com95d, Com95e, Cou97a, Cou97d, Cou97e, EMe97f, Ess95, Gar98b, Gau96h, Gau96i, Gau97m, Gau97n, Gau97o, Gau97p, Gre95, Gro96b, Gut95d, Gut95h, Gut95i, Gut95j, Gut95k, HB95, Hay96m, HG96a, HG96b, Hib96, Hof97, HBG^{+}97, Hor96c, IBM95, Inf95d, Jar98, Kal95a, KP96, Kar95a, Kar95b]. Sun [Kro99a, Kro99b, Kro99d, Law98, LH96, Lie96b, Lou95c, MD99, Mac95, Mag95, McC96c, McC95b, Mes95c, MS96, Met95d, Mey98c, Mil95a, Mil95b, Mor97c, NJ95, Net95b, Net95a, OH97b, Par95f, Par95g, Com95g, Red97c, Sil95, Sem96, SF96c, Sim96d, Sim97d, SG97a, Spy95, Sul95b, Swa98b, Ten95b, Tuc98a, Ubo96b, Ven96, Wat95b, Wat95d, Way96e, Win95b, Win95g, Win95m]. Sun/Microsoft [OH97b]. Sun’s [Gau97e]. SunScreen [Ano95t]. SunService [Inf95e]. SunSoft [Bay97a, Gea96, Ano95u, Ano96-182, Cox95d, FC95, Hes95, LaM95c, LaM95e, Tuc96, LPC96]. SunTest [Ano97-140]. SunWorld [Ano95v, Par95f]. superbible [NMS96, GHP^{+}96, Gro96a, Groxxa, N^{+}xxa, GSHP97]. Supercede [Wil97a, Gre97e]. supercharging [TTD96b]. supercomputer [Bam97, Van97n]. Supercomputers [Gar99]. supercomputing [Ano96-34, FT96, FTT97]. Superhighway [IEE96c, T^{+}98a, Ano96-186, Cim96]. SuperPack [Anoxx-80, GGA96]. SuperWeb [AISS97]. Supplement [CLK99b, Har96c]. suppliers [Ano96-59]. supply [Hay96n]. Support [Ano96-28, Ano97-42, BW99, CDF^{+}98, Cox95a, Dar98b, Eas99, Fah96, Hay96u, Inf95a, KHB99a, KHB99b, LV99, Lot95, McK95b, O’C98, Pir97, Ran99b, RNBV98, San98b, San98c, SLC99, S^{+}96b, VRG98a, VRG98b, VR99a, Win95e, Ano95k, Ano95s, Ano96-97, Ano97f, Ano97-83, Ber96c, BV97, BBGD98, CCCL97, Dar98a, Dar99a, DiD97a, EHS^{+}99, Fuc96, Hay96r, Hay96x, HAKN97, Sim96f, Ste97a, Wat95f, vD99a]. Supported [Ano97-77, Reg96]. Supporting [But99, Mon97, WCS99, ZTLV98a, ZTLV98b, BRS99, Ham96a, HG96a, Wel99a]. Supports [Ano97z]. Supremacy [Law99]. Surf [Ale96, Ano96-53]. Surges [Ano96-106]. Surgical [KSI^{+}96]. surpass [Ano97-166]. surprising [Ude97b]. surrounding [GA98]. suru [IK96]. Survey [AR98, Bro95a, Hal98b, Man96a]. survive [Ano97-109]. Sus [GR97]. sustainable [Ste96b]. Swaine [WL97a]. Swan [Swa97f]. Swing [DW99, ELW98, Gea99, Gut98, Ano99d, Ano99e, Soo98a, WC98, Zuk99]. Switch [Ano96-137]. SwitchBoard [Cro99]. Switzerland [Str96]. Sybase [Hob98d, Ste97a]. Symantec [Krü97b, Krüxxb, Ano96v, Ano96-35, Ano96-42, Ano96-182, Ano97-151, Ano97-152, Ano97-154, Bar96, CH97a, Hay95c, Hem96a, JD97, Kro97, LaM95f]. Symanticec [Mes97]. Symbolic [BCK99, Doo99, PJR^{+}98, Ber99b, Ber99c]. Symmetric [KMW99]. Symposia [IEE96f]. Symposium [ACM96, ACM97a, ACM98b, BBJR96, BK96a, CEG96, Cim96, Doo99, FW96, IEE96a, IEE96l, IEE96n, IDK98, Joy98, Spe98, Ste96b, Ten99, USE96a, USE97c, USE98f, USE99b, Wae96, dN97, Chu98b, IEE96e]. Sync [MD97b]. Synchronization [BKMS98, Chr99, For98a, BH99b, Ven97b]. syntax [AF99]. synthèse [Claxx]. Synthesis [Kro98b]. Syracuse [IEE96n]. System [Ano96-137, Lut99a, IG99, Cha98d, Cha98e, FM98, Gar97a, HLR97, IAJR97, KT96, KG98b, KG98c, Kro98d, Kro98c, Kro99a, Kro99b, Kro99d, Led98a, Lev99, Lut99b, MD99, MGL^{+}98, MK98, McK95a, MB98c, Ran97b, RWWB96b, RWWB96c, RC97, Sch99b, SA99, VCCH99, Wae96, WRW96a, WRW96b, WRW96c, Ano96-97, Ano96-205, Ano97-102, BF99b, CRS^{+}96, Cli98b, EH96, FTT97, FM99b, GJ96a, Gre97h, Gre98c, Gre98f, Hol99, JJG98, LVS^{+}96a, NJC96, O’C99, RMB^{+}98, Rey96, RFPG96, RWWB96a, Rit97d, SM99, Sim96f, SA98b, TYY96, Vck96, Wel99a, LVS^{+}96b, Ano96-43, Ano97e, Kro98d, Kro98c, RNBV98]. system-independence [Cli98b]. systematic [VS96]. Systems [Ano96-135, Ano96-158, Wil97e, Ano97-82, Ano98-27, Ano98-30, Ano99h, Bal97, Bar99, CI96, CRS98a, Cor98b, CR99, DDJ98a, DG96, Eri96b, Ham97c, IEE96f, IEE96o, IEE96q, IEE98a, IEE99b, Kro98b, Lew97a, MD99, MDM97, Moh98, MJA98, PB98, Ran97a, Rit97d, Rou98, Rym95, Sar99, Tau96, TMH99, TBA98, Ass96a, USE96c, USE96a, USE97b, USE97c, USE98b, USE99c, USE99b, USE99e, Wil97f, YH96, Zei97, Ala99, Ale96, Amb97, Ano96-54, Ano96-100, Ano96-144, BRS99, CRS98b, Chu98b, ERKE96, FGL96, Hay96k, LM98, Lea97a, LTC96, MHO96, Neu97, PSL^{+}98, PV99, Pog96, RR99, SH99, SLPM96, TB96, Vog96, VI96, Wat95f, Wil96f, Wil96e, Ano96-183, Ano98d, Ano98-32, Ano98-33, Kro98e, Kro98b, Kro99a, Kro99b, Kro99e, Kuk98, Por98, UCD99].

T [HC99a]. T3E [Chw98]. T3E-1200

[Chw98]. Table

[Ano96-76, Ano96-117, Ano96-117]. Tables [Hor96a, Hor96b]. tackle [Ano97-28]. Tackles [Win95a, Lin96e]. Tactician [Ano96-135]. Tactics [Mel96]. Tage [Dan96b]. Tagen [W^{+}97c, Cha97a, Hob97a, Hobxx, LP96c, LP97c, Lemxxb]. Tailored [HC98a]. Taiwan [Ano96-115, Uim95]. Take [Ano96-140, Gre97j, Hol96a, LaM95c, Bab97a, Gau97r, Sil95, Ano98q]. TakeFive [Kro98e]. Takes [Ano96-107, Ano96-118, Dec95b, Par95f, Tuc96, Win95i, Gre97i]. Taking [Ano96-184, Ban98]. Talent [DDJ98b, DDJ98c]. Taligent [Fah96]. Talk [Eng97b, Knu97a, Macxx, MO96, Sim96b, Fuc96, Her96b]. Talks [Par95a, Par95b, Ano95a, Ano97-52]. Taming [TMP^{+}99]. Tanner [Kro98e]. tantalizing [Hay97d]. Tanzania [JPT^{+}99]. tao [Nau96b, Fic96]. tap [Gut96, Ric96d]. Tape [Smixxa, Smixxc, Smixxb]. target [Hay96v]. Targeted [CR99, Gau97f]. Targeting [CDG97]. targets [Gau97e, Way98c]. Taschenbuch [MB96a, Schxxg, Stexxa]. Taschentabelle [Wolxxa, Wolxxb]. Tasking [MB98c]. taste [Ano96-188, Hay96c, Pic96a, Pic96d]. Tcl [Ass96b, USE97a, USE98d, Bek97, DBK98, GA98, LS97a, RC97, SBB96, Sim96d, Sta98b]. Tcl/Tk [Ass96b, USE97a, USE98d, Bek97, Sim96d]. Tclblend [Sta98b]. TclJava [SC96b]. TCP [JLW96]. TCP/IP [JLW96]. Tea [Com95f, TYY96]. Teach [AM97, Ano96-185, Cad97, Cad98b, Coh97e, Dan96d, Dan96c, Doh97, Dow97, FM97, Hob97c, JG96, JLP96, KPF96, LP96e, LPW96, LPM96, LPC96, LP97d, Lem98, LC98, LPxxb, Man97a, Man97b, Mor96b, Ros98, Tat96, Tuc98a, W^{+}96, WPL97, LC99]. Teaching [Ano96-186, Bro98b, KSI^{+}96, Muk99, Str99, Tob99, BBG^{+}99, DV98, EHS^{+}99, KR99, Ref99, TGGP97, Wei97c, YB99]. Team [Ano98-36, Lut99a, Lut99b, Ano96-74, Ano96-78, Hay96m, Sav97, Ano98v]. teams [Ano96-108]. teamwork [Fri96]. Tears [BK99]. Tech [Ano96-187, Ano96-188]. Technical [Ano96-152, Bra96, Got98b, Han99b, IEE96g, Joy98, Pol98, Pre99, Sak98, van97e, USE96d, USE98c, USE98e]. Technik [Mor97e, Morxxd]. Technique [BN99a, Tes96a, Gol96a]. Techniques [App97, BF98, BF99a, BKT96, CBT98i, CDL99, DP99, Esp96, HM98, JLMPH99, SR97, KR99, Mar97a, Ano97l]. Technologies [Ano96-135, Ano96-158, Ano98-28, FTT97, FFKO99, HYI98a, IEE96c, IEE96r, IEE96m, Lut98a, Moo95, Ass96a, USE96c, USE97b, USE97c, USE98b, USE99c, USE99b, Ano96-34, AA96a, BKP97a, Che96, Fla96f, FT96, HYI98b, Mar97a, Str96, Tub95, WB^{+}96c, Ano98-32, Ano98-33, Kro98d, Kro98c, Kro98b]. Technology [Ado95, Ano96-33, Ano96-47, Ano96-154, Ano96-198, Ano96-199, Ano97-153, Ano98-27, Ano98-29, Ano98-30, AFHxx, Ben96a, Ber96j, Bra97a, Bra97b, Bro95a, Cas96b, Cho97b, Ora95, EBR96a, Fol99, Gar98b, Hin98, Isk96, KGG96, Kro98a, Kro98d, Kro98e, Kro98c, Kro99a, Kro99b, Lan98, Lor96b, Lot95, Lut98b, Mac95, Mil95a, MDM97, USE99d, Urq97a, Amb97, And96a, Ano96-73, Ano96-60, Ano96-117, Ano96-133, Ano97w, Ano97-166, BGH99, CRV^{+}96, Ciz97, Com97a, Com99, EBR96b, Gau97j, Gol96a, IBM95, IDK98, Kap99, Lan99a, Mag95, Mar98b, Mar99a, NTS^{+}96, Owe96, Rus98, Sco96c, Ste96f, Sul95a, Sul95b, TYY96, W^{+}98, Yan96, Yas96, van99b, Ano96-189, Ano96-190, Ano96-49, Fah96]. Technology-Based [Isk96]. Technosoft [Kro99c]. Técnicas [Jep96]. Tecnologia [Gar97e]. TeCo3D [Mau98]. Tecplot [Ano96-135]. tehokóyttójón [Fla97f]. Tel [IEE98c]. Tel-Aviv [IEE98c]. telecom [Fla96f, Rob99, Boz95d]. Telecommunication [CJJ^{+}98b, CJJ^{+}98a]. Telecommunications [Ano98-27, Gro96b]. teleconferencing [Rus98]. telecooperation [Mau98]. telemedicine [LVS^{+}96a, LVS^{+}96b]. Teleoperated [AVN98]. Telephone [Gar97a]. Telephony [Gar97c, Marxxa, Rob99]. Teleporting [WRB^{+}97]. Telescript [Doe97, Dom96, Met95c]. television [Wat95f]. tell [Gof97]. Telling [Knu97d, Knu97g]. tells [DiG95a]. tempest [Lie96b]. Template [CS99]. Temporal [San98a]. Temptation [Kim97e]. Tenth [Uni97a]. Terminal [San98b, San98c, Way96f, Gut96]. terminals [Way96c]. terminologie [BKP97a]. terms [D’S96c]. TerraVista [MD99]. Test [Kro98e, Kro98b, Kro99a, Kro99b, Law98, Ran99a, Ang99, Ano96-188, Ano97-135, Ano97-164, Kin97b]. Testers [BBG^{+}99]. Testing [Wil97e, DDJ99b, HHW^{+}99, Kro98b, Lut99d, Lut99e, Sel99, UCD99, Wal97i, Wil97f, Ano97-148, Lak96]. Tests [Web97c, Web97d]. Tetherless [Law98]. Texas [USE98f, USE99a, Cou96b]. Text [Wil97e, Bec98, Gat97, Gil98a, TMP^{+}99, Wil97f]. textbooks [BR97]. TextSurgeon [Ano97-119]. textual [KT96]. Texture [Dys96]. Their [Ano96c, Wat95d, Ano96-59, MKS^{+}96, Sta96]. them [Sco96a, Way96e]. Themselves [Kan97]. theory [BK97a, IEE98a, OW97b, RBL^{+}97]. There [BCP^{+}99b, Ano96-102, Lie95a, Lie95b]. Thermal [Kro98d, Kro98c]. Thesaurus [Kro99c]. These [Cli98b]. Thi [Kanxxb]. Thin [BKMS98, Tou99, MC97, Mor97c]. Thing [BCP^{+}99b, Hol96a, vSS96e, Ban95a, Gre97f]. Things [Lut99e, Ano96-54, vSS96e]. Think [Sco96a]. Thinking [All99e, Als95c, Als95a, Als95b, D’S96c, Eck96d, Eck98, Hay98, GS97b, iMa96]. thinks [Ban95b]. Third [FGL96, Sch96d, USE97b, VPM98a, VPM98b, Ano98g, Bod96, Fox98b, HG96a, DF96, Spe98]. third-party [HG96a]. thirsting [HN96]. Thirteenth [Uni98b]. Those [Ano98z]. Thought [Ano96-135]. thread [GLC99, Ven97b]. Threaded [CKRW99, RKCW98, CV98, CKRW97a, CKRW97b]. threading [Eck97a]. Threads [Ber96h, Cri98, Gra96d, Küh97, OW97a, OW99, OWxx, Ber96b, Van97m]. Threat [Sim96e, SG97b, GH96b]. Three [Ano96-78, AJ99, Bro96, Cor98b, Hay96v, Kuk98, Ano96-174]. Three-Tier [Cor98b]. Thwarting [McC98e]. Tie [LaM95g]. Tier [Cor98b, Kno95]. Ties [Boz95b, Gil98b]. tightly [Gil98b]. till [Bar95b]. tillämpningar [Ek97a, Ekxxa]. Time [Ano96-141, Ano96-178, BW98a, CFM^{+}97a, CFM^{+}97b, DDJ99a, GK98, Ham97c, Jam96, Knu97d, Kro98b, Lea99, MD99, MGL^{+}98, MKT98, Nil96, Ort99, PB98, TMH99, ATCL^{+}98, Ano96-80, Ano96-171, BB97a, Bur96a, CS98, CL97e, Eck97b, Eck97c, GH97, Gau97a, HG96a, KG96a, KG96b, Knu97g, KG97b, MMJ96, MF97a, Nil98a, PV99, PC97, Rit97d, Sey96, ST96, Win96, PTB^{+}97]. timely [Ano96-60]. Tiny [IR97, Lew96b]. Tip [AL97]. Tips [Ano96-152, CBT98i, CGI97, Cha98b, GCxx, HM98, KS96b, Laf96, Lafxxa, Dav97d, Ek97a, Ekxxa, Flo97]. tired [Hay96k]. Titanium [YSP^{+}98a, YSP^{+}98b]. Tk [Ass96b, USE97a, USE98d, Bek97, Sim96d]. TLSI [Gop96]. TM [AJ99, Bro97b, Qia99]. TN [Ano98-34]. Toaster [Lie96a, KP96, Sun96]. Toba [PTB^{+}97]. Today [Hal97a, LP96a]. Todo [Siy97]. token [Mey98b]. Tokyo [Uni98a]. Tolerant [Sar99]. Tom [Swa97f]. tomb [Gre96b]. Tomorrow [Hal97a]. Tonga [Gar97a]. Tönne [FT97, Froxxa]. Too [Wil99d, Cli98b, Gau97r, Hal98a]. Tool [Ano96-59, Ano97-85, Ano97-113, Ano97-155, Ano97-163, Ano98-30, Ano98-32, Ano98-33, Bha98, BG98b, BG98c, BS98e, Car96d, Cox95d, Det97, Eng97b, GM98b, HAT98a, Hig98, Hor96a, Hor96b, Inf95f, Int96a, Kro98a, Kro98d, Kro98e, Kro98c, Kro98b, Kro99a, Kro99b, Kro99d, Kro99e, Kro99c, LaM95g, LZ97, MD99, Min96a, NE98a, NE98b, SF96c, Ano95h, Ano96-195, Ano96-124, Ano96-201, Ano97-50, Ano97-46, Ano97-132, Ano98d, Arm99, BDY99, BG98a, Dei96b, DIS99, Dys96, EKR97, Gau96d, GG99, HRW^{+}96, Mur96a, RHND99, SW96, Tre98, Wag97b, iMa96]. Toolbox [Ano97-153, Bra97a, Bra97b, MD99]. Toolbox/SM [MD99]. Toolchest [Ang99, Joh99, Rou98, Sch99a, Soo98b]. tooling [Gre98a]. Toolkit [Ano96-63, Ano96-122, Ano96-135, Ano96-137, Anoxx-63, Anoxx-47, Wal97d, D^{+}98b, GYt96b, GYT96d, Gosxxb, Kro99e, Lei97, Sch97j, Ano97-135, Coo97c, Hol99, Mar97a, SML96, Way98c, Ano97-134, Hem96b]. Toolmakers [Hay95d]. Tools [Ano95u, Ano96-37, Ano96-55, Ano96-79, Ano96-110, Ano96a, Ano96-132, Ano98-30, Ano98-31, Ano98-32, Ano98-33, Anoxxi, Anoxx-72, Ano00, AC98, BS96, Boo96b, Bor95, DDJ99b, De 96, FW96, Gau96j, Gla97b, HB95, Hay95e, Hay96l, Hes97, IEE96i, Klu98, KR99, Kro97, Kro98a, Kro98d, Kro98e, Kro98c, Kro98b, Kro99a, Kro99b, Kro99d, Kro99e, Kro99c, Kuk98, MD99, Mac95, Mes95b, Mes95a, Met95d, Nas95c, Ric96d, SL99, Vad95b, Wal97c, Ano95h, Ano96q, Ano96-171, Ano97-69, Ano97-72, Ano97-151, Ano97-152, Ano97-148, Ano97-140, Ano97-157, Dry96b, EL98, Gau97j, Gau97k, Gil96a, Ham96a, Hay96i, Hay96n, HG96a, Jur96, Kar95b, Kno96b, Lin96d, Lin98b, McL96, Mes97, Ref99, RBL^{+}97, Rus98, Sri96, You97d, MDM97, Ano97-142, MD99]. Toolset [Ano97-55, Ano97-121]. Top [Ano97-156, FFHA98, Kro99d, Kin96, Way96c, Won97].

Topics [IEE99b, Min97a, Her96b]. TopLink [Gre98f]. Toronto [Ano96-158, FW96, MJ98, Ass96a, USE96c]. Toshiba [Com95e, Gut95j, Gut95k]. Toshiba-Sun [Gut95k]. tot [BHxx]. Total [Lut98a]. totally [Sco96a]. Tough [Kin97b]. Toulouse [IEE96i]. Tour [Cla97c, KS98a, KS98b, JJG98]. Tout [Cadxxc]. touted [Gau97i]. Tower [Ano98-29]. TowerEiffel [Ano96-135]. TowerJ [Ano98-29, Tow98]. Toy [SF96c, Sem96]. toys [Ano96-195]. trace [Col96b, Hay96o]. Tracing [Wol99]. Track [Ano96-85, Syd96b, USE98c, Hay96w]. Tracking [BAC^{+}99]. tracks [Col96b, Hay96o]. Trade [Ano97u]. trading [Ano97t]. train [Hol97d]. Training [Ano97-40, Anoxx-69, Buhxx, BS97e, DD97b, DD97a, Dei97, Hel97a, Hel97b, LB97b, McH97, McHxx, Min97b, Min97c, Minxxc, WL97b, Ano96-145, BK97b]. Trainingsbuch [Hol97c]. tranlsation [Lafxxa]. Transaction [BCC^{+}97, Gre98d]. Transactions [BCE98, Tou99, LS98, VR99b, Cha99f, LP96a]. Transfer [Ber98d, CKH99, KCLZ98]. transform [KW96]. Transient [Tým99]. transistor [Ano96-153]. transit [Wip96]. transit-on-demand [Wip96]. transition [Ano96-153]. translating [EK98, OW97b]. Translation [N^{+}xxa, BPxx, DP99, Flaxxb, G^{+}xxd, G^{+}xxe, HGH96, Kanxxb, Kanxxc, Par97]. translator [Kei95, Lafxxb, Pla97]. Translators [PQ96]. Transliterating [Pla96a]. transparency [BRS99]. Transparent [Fue98, Fün98, SMY99, PZ97]. transporter [Wil96b]. Transvirtual [Knu99a]. Travel [Wat95e]. Traveling [Led98b]. Treaps [Nil97]. TREC’98 [LM98]. TreDS [SLPM96]. Tree [KF99a, KF99b, Ano97-101]. Tree-Based [KF99a, KF99b]. trees [LR96]. Trends [DDJ98b, DDJ98c, Kan96b, LM98, Ort99, SLPM96, Ste96a, And96a]. TRI [ACM97b]. TRI-Ada’97 [ACM97b]. Tricks [Anoxx-37, V^{+}96, Fri98]. tries [Gau97g]. Triveni [CJJ^{+}98b, CJJ^{+}98a]. Trojan [TW96]. Trouble [WL95, OHE97a]. troublesome [Mat96b]. Troy [Ano97k]. Trucs [Ano96-120]. Trudger [Ano96-137]. True [Ano95p, Har96a]. trust [DiG95b]. Trusted [VPM98a, VPM98b]. truth [Ano96-61]. tse [Nau96b]. tu [Harxxc]. tuned [Kno96b]. Tuning [CP98, Hum98, Red99, Tym96, VR99a, Cel96a]. Tunnel [dRPT98]. Turbo [WDDF98]. TurboImage [Reg96]. TurboJ [WdFD^{+}98]. Turing [RHND99]. Turkey [NRG^{+}99]. turn [MF96a]. Turnaround [Gar97a, Gar99]. Turns [Boz95c, Ano96-179, Bal96b, Gre98c]. Tutorial [AFHxx, CW96a, CW97, CW98, CWxxb, CWxxa, HCF97a, H^{+}xxa, HGxx, Har96c, Her97, KFL99, LSxx, Nie95, Sha96b, WC98, Ano97-111, BKCG98, CSH^{+}98, Gos96, Mee96, RK97b, Wea97a, WFC^{+}99, van96b]. tutorials [CT96, CWS98, Uni96a]. tutoring [FGL96]. Tutto [Var97b]. tutuksi [Ekxxb]. TV [Gar98b, Gro96b]. tweaks [Gau97p]. Twelfth [ACM96, Uni98a, Wae96]. Twentieth [IEE96s]. Twenty [Ano96-160, AOV^{+}99, IEE98b, Joy98]. Twenty-fifth [AOV^{+}99]. Twenty-First [Ano96-160]. Twenty-ninth [Joy98]. Twenty-Second [IEE98b]. twice [Sco96a]. Two [Ano96-193, Gre97j, Gre98f, GLC99, Mee96, RB96]. TX [ACM99a]. Tycho [HLR97]. Type [ADM98, Ala99, DE97, DE99, FM98, MF97a, SA99, Sym99, VN99, AFM97, Dug99, FM99b, NvO98, O’C99, SA98b]. type-based [Dug99]. Type-checking [Ala99]. Type-Precision [ADM98]. type-safe-definitely [NvO98]. Type-Safety [VN99]. Typed [LST99, WJUH98, GLV99, LR98]. Typen [Sta96]. Types [All99f, BW98b, CG99, CS98, D’S97c, Jen98, Sak98, Sch99c, Sta96, Tho97b, BML97a, Bok99, Cai99, Dug99, MKS^{+}96]. Typing [RC97]. Tze [Flaxxb].

U [WOC99]. U-Net [WOC99]. U-Net/SLE [WOC99]. U.K [Ano97l]. u.v.m [Klu98]. Überblick [Dec96b, Decxxb]. Übersetzung [MP97a]. ubersichtlicher [SBH97]. Ubiquitous [Cha99e, DDJ99a, WRB^{+}97, Ano96-34, CL97c, FT96, FTT97]. ufficiale [Lemxxd]. Ugly [Dar98b, Mar98c, Dar99a]. ui [Kim97d, Gat97]. UK [Ano98i, Ber98i, Abr96, Ano96-142, Ano96-159, Ano96-154, AOV^{+}99, SMEN96]. UL [Kro99e]. Ultimate [DD97b, Dei97, H^{+}xxd, PF96, JC97]. Ultimately [AL97]. Ultrafast [Gar99]. Ultrexx [Ano97-157]. UML [Ano96-136, Arm99, CLD99, Gra99a, HW97, Kro98b, Laixx, Lut99c]. Umstieg [BD98b]. Uncertain [Gre97k]. unclear [Gau96f]. Uncorked [Cha99e]. uncovers [Sav97]. undergraduate [BK96b, Had98a, Owe96, RR99, Sco96c]. underlying [MN96b]. Understanding [Bud98, HW97, Marxxa, Pop97, Smixxb, Whi97b]. Undocumented [Ano96-194, Tit96c]. Unfolds [Wat95a]. Unicode [Uni96a, Uni97b, Uni98a, Uni98b, Bin97, Che97, Cob98, Leo98, Ran97a, Ran97b]. Unicode/ISO10646 [Uni98a]. Unified [Sad97a]. unifiée [Laixx]. Uniscape [Ano98-32, Ano98-33, Kro98b]. United [Ber98h]. Universal [BL97, ZTLV98a, Gop96, LH96, WFC^{+}99, ZTLV98b, Ano97-102]. Universe [MRA98, Ref99]. University [Ano97l, Ano98i, Ano98k, Cul97c, Doo99, IEE96e, IEE97, IEE99a, vD99b, Kro99c, Tri95]. UNIX [Ano98-29, Gla97b, Lut99e, McG97, OH97b, Red97b, Ano96-74, Ano97p, Ano97-52, Ano98-32, Ano98-33, GJ96a, Kno95, Laz97, McK95a, PM97c]. Unix-to-NT [Ano98-32, Ano98-33]. Unleashed [Bre96, Cro97, DR^{+}97, Fan96, Jaw98, M^{+}96a, M^{+}97b, M^{+}97a, M^{+}97e, M^{+}97d, M^{+}xxb, M^{+}96c, OBWR96, P^{+}98b, Sta97a, WC96, W^{+}97a, Ano97-87, M^{+}96b, M^{+}97c, SD96]. Unlike [Mat96b]. Unlimited [McC98a]. Unloading [MB98b]. unnecessary [BH99b]. unplugs [Ano96-129]. unsafe [Raj96]. Untrusted [Gri99]. Unveil [GC97, Gar99, JN95c, Kno95, Mes95a]. Unveiled [Ano96-113, Ano97-99, Fed97, Ano96-73]. Unveils [Gar97a, Gar98b, Nas95a, Ano96-180, Ano97-149, Ano97-158, Red97c]. upcoming [Ano98u]. Update [Car95, MGL^{+}98, TMP^{+}99]. Updated [Anu96a]. Updates [Ano97-131, Ano96-133]. Upgrade [Oue95b, Gau96e, Ung97]. upgrades [Ano97-127, Stu98]. Uproar [Pla96b]. ups [Mey98a]. Urban [BB99, Lut99e]. URL [BAC^{+}99]. USA [ACM96, Ano96-160, BBJR96, IEE96g, IEE96h, Uni96b, Uni97b, USE96d, USE99a, USE99c, USE99b, USE99d, USE99e, YN96]. Usage [Kro98d, Kro98c, WH99, MW98]. Usando [New97, Newxxb, Per97a, Perxxb]. Usare [Dadxx, FCD97, Perxxc, Ritxxh]. USD [Ano96-53]. Use [Ano96-28, Ano96-146, Lut99a, CD97, Cob98, Gar97a, Gar97c, GC97, Gre96c, Lut99b, McC97a, OH97a, Pit98a, SR97, WR99a, Ano96-202, Ano97-73, AA96b, Ber97a, Ber97b, Bok99, CM97a, Fox98b, Gau97h, Hay96a, IBM95, MKS^{+}96, NGM^{+}97, Orc98, PL97, Vil99, Xu97]. used [Gau97q]. Useful [Cue96, Fri98, Tur97]. User [Ano96-137, Ano97-30, AKM99, Bec98, CBT98j, DLM99, Gut98, HLR97, IR99, Col96b, Cul98c, EBB^{+}96, Eng97a, Fet97, Gop96, Hay96o, LTC96, MP98, MP99, PS96, WOC99]. user-customizable [WOC99]. Users [Ano96c, Ano96-160, Ano98-34, Boz95a, DS97b, Hay95e, Hay96w, Oue95b, Amb96, Ano95s, Ano96-72, Ano96-181, Ano97-145, DiD96, GJ97, Gau97k, Gau97m, Ham97a, HN96, Hay96g, Hay96t, Red97b, SG97b, Ste97c]. Uses [Sco96a, Ano96-128, Ano97-126, Gil98c, MC97]. Usher [Win95k]. Using [AVN98, All99f, And97a, Ano97-30, Ano97-122, Anoxx-71, Anoxx-73, AW98, Ast98, BV98, Bak97a, Bakxx, BG96, BM98b, BCE98, Ber98g, BFD96, BC98c, BRF98, BLT98, Bro96, Buc96a, CRS98a, CRS98b, CSY97, CWS98, CJ98, Cor98b, Cox96a, CFM^{+}97a, CFM^{+}97b, CWHP96, DBK98, DF97, Doh98, EMe97a, EMe97g, EBR96a, Far96e, Far96c, Far96d, Far97a, Far97b, Far97c, Far97d, Far98b, Fli98, FFKO99, GM99, HNSS99, HJS98, H^{+}xxb, HS98b, IBM97, IAJR97, JOA98, KMR98, Knu97f, Knu97e, LO96, LOA^{+}96, Lad96, LO97a, LO^{+}97b, LO98a, LO^{+}99b, Lew97a, LB98a, Mai98, MLM98, McD99b, Mee96, MHP98, Moh96, Mor98b, NNS97, NN97, N^{+}96b, New96d, Par99, Per96b, Pet97, Pol98, Que97, Rao97, Rao98d, Rao98a, Rao98b, Rao98c, Rao99a, Rao99b, Rao99c, Rao99e, Rao99d, Rei97a, Rei97b, RDD^{+}96, RH97a, RWxx, RFPG96, Rie96]. Using [Sad97a, San97b, Sar98, SHW98a, Sav99, S^{+}96a, SB96, SBB^{+}97, SA97, SAxx, Sie98, Sla97, Smixxc, Rapxx, SP97, SS96, TBS97, Tay98, Tes96a, Tes96b, TJ96, VP98, WD^{+}96, Wal97c, Wat98, WB^{+}96a, WBC^{+}97, Webxx, Wei98, W^{+}97b, Wil97a, Wil97b, Wol99, WMDH97, YB99, YRE98, ZJ96, AS96, Ano96-205, Ano97t, Atk99, BB97a, Bot98b, BR97, CKH99, CRS^{+}96, CT98, CCI^{+}97, Cla97d, Cul98c, Dua96, EM99, EBR96b, FHKS98, HL97b, HL97c, JAS96a, JJBW98, KW99, KCLZ98, LR96, Min96b, MLF97, MMGL99, Muk99, Par97, Com95g, RBF98, RMB^{+}98, Rit97d, Rob98a, SHW98b, SH99, Sha98b, Str99, TMN^{+}98a, TMN^{+}98b, TRMR97, Tym98, Wat99, Wea97a, WRTM98, YLQ98]. USITS [USE99b]. Utah [Isk96]. util [Cha96c]. Utility [Ano96-63, Pit98b, Web97b]. Utilizing [KSB97]. Utils [Chaxxb].

V [Bra96, F^{+}xxb]. v.7.1 [Kuk98]. V1.0 [Gol96b]. V1.10 [MD99]. v1.2 [CLK99b]. V2.10 [Ano98-27]. v5 [Ano96-137]. v5-Multiplatform [Ano96-137]. Validation [Ano97-115, HO99]. Vancouver [Doo99, Tau96]. VanillaSearch [Ano96-135]. Vaporware [DDJ98b, DDJ98c]. Variable [ADM98, Tým99]. variables [HVB^{+}99]. VariCAD [Ano98-28]. variety [Way98a]. VB [Hay95e, Par95d, T^{+}xxb]. VBSCRIPT [Lam99, Ait96c, Ait97b, Ait97a, Bre96, Cor97b, JLW96, SR97, WB^{+}96c]. Velocis [Ano97-118]. Vendors [Ano96-37, Gar97a, Gar97c, Gut95g, Hay96x, Hay96y, Kno95, LaM95g, Mil95d, Vad95b, Wat95c, WP95, Win95l, Hay96r]. Ventura [Nat98, RD99]. VerCheck [MGL^{+}98]. Verification [BFPV99, Gol98a, PV98, VP98, FC98b]. Verifier [Pus99, Ano96-200, FM99a]. Verifying [RKCW98]. Verilog [Kro98b, Kro99a, Kro99b]. vers [Drixx]. Verschlusselungstechnologien [SBH97]. versie [Kas97b]. Version [ALP99, Ano96-137, Ano97f, Ano97-118, Anoxx-79, Dav96b, DS97a, Fla96c, Gar97c, HHW^{+}99, JRE98, Kuk98, Lew97a, Lin97, MD99, NRG^{+}99, Nghxx, Til97c, Tri95, Gau97p, Ano96-105]. versions [Ano96-97, Mes97]. verstehen [BKP97b, Krü97a, Krüxxa]. Versus [MGL^{+}98, Bek97, SZ99]. Very [AOV^{+}99]. vg500 [MD99]. VGJ [MB98a]. Via [Wat95e, Ale96, Ano96-123, Hof97, HLRT97, MN96b, Reg96, ZK98a, ZK98b]. VIBDaST [TGGP97]. Vibe [Ano97-120, Ano97-158, Gre97f]. Victoria [IEE98a]. Video [Ano98-27, Hor96a, Hor96b, Mas98a, Mas98b, Mas98c, Mas98d, Mas98e, Mas98f, Smixxa, Smixxc, Smixxb, Fre96a, Kam99]. Videoconferencing [Gro96b]. Videos [McH97, McHxx]. VideoSoft [Kro99c]. VideoTalk [MZ97]. Videotape [Bay97b]. vie [GH96c]. vielen [KS96b]. Vienna [IEE98b]. Vietnamese [Lun99]. View [Ano97-110, AKM99, Wal96q, Wea97b, D’S97c, PS98, WWR97a, WWR97b]. viewer [Beh95, Dys96, Beh95]. Viewing [Hor96a, Hor96b]. Views [DDJ98b, DDJ98c, The98, DDJ98a, DDJ99a, DDJ99b, Hor96a, Hor96b]. VII [IEE99b]. VIII [IEE96e]. village [Jon98]. violations [McC97b]. Virtual [ADM98, Ano96-204, Ano97q, Ano97k, Ano97-48, Ano97-87, AC98, BS98c, BS98d, Chw98, Cla97a, Coh97b, CDG97, DDJ99b, Dalxx, Dal97, Eng99, EBR96a, Fra98b, HT98, Hin97, Jar97, LP96a, LM99, LB98b, LV99, LY97a, LY97b, LY99, LYxxb, LYxxc, Lut99c, Lut99d, Lut99e, MD97a, MDxx, MB98c, OT97, Qia99, SD99, Spe98, Sun95b, Sun95a, SUN97b, JCV99, Tai98, Tho97b, Tol98, TO96, UBF^{+}98, Ven97a, Ven99, VR99a, WL96, Wak98a, Wak98b, Wak99, WH99, Wol99, Won97, YC98a, Ano96-192, BDF^{+}98, Bod96, DLM99, EBR96b, EL98, FF96b, GPV97, Gre98a, Sha98b, SZTR96, Ste96f, TGGP97, Ven97b, WOC99, W^{+}98, YC98b, Yel99, Bru98, DF97, Mon97]. Virtual-Human [Lut99c]. Virtually [WL96]. virtuelle [Web98]. Virtuellen [LYJ^{+}97, LYxxa]. Virtues [Car96d]. viruses [Cou97c, McC97b, Nan96]. Visa [Gau97q]. Visaj [Ano97-46]. Visas [Gar97b]. VisiBroker [P^{+}98a]. VisiCorpse [Swa97b]. Vision [Ano97r]. Visioneer [Ano96-196]. Visitor [Pit98b]. Visix [Ano97-120, Ano97-158]. Visual [AFPW97, Aitxxa, Ano96-188, Ano96-197, Ano97-152, Ano98x, Ano98y, Ano98-30, BM98b, CBT98g, Cho97b, CBJ97a, Cohxx, CS97b, Dav96b, Dav98b, D^{+}xx, EMe97a, EMe97d, GS96, Gre97g, Hem96a, H^{+}97, JD97, JMM96, KSW98, Kro99d, Krü97b, Krüxxb, Lad98a, Ladxxb, Llo98, Min96b, Mas97a, Mic96, MC98, MP97a, Pat98a, Per96b, Sch97d, Schxxa, Van98b, WSK97, Wes97b, Woo97a, Wygxx, You97b, Ano95h, BBC^{+}97, Boy98c, Gre97c, Hay96c, Red97c, A^{+}96, Aitxxb, Ano96-124, Ano96-127, Ano96-137, Ano97-151, Ano97-141, Ano97-159, Ano97-160, Ano97-161, Ano97-162, Ano97-163, Ano97-164, Ano97-165, Ano98-29, Anoxxm, Anoxx-39, Anoxx-64, Anoxx-74, Anoxx-79, Anoxx-75, Anoxx-78, Anoxx-76, Anoxx-77, Bla98, Bre96, BLT98, Buc97, Cobxxb, Coh97e, CBJ97b, Coo97c, Cor97a, Thexx, C^{+}97b, CWHP96, Cul98a, DD97c, DD97d, EMe97b, F^{+}xxa, FC96b]. Visual [FC98a, FM97, Garxx, GPF97, GH96c, Gil98b, Gre97a, Gre97g, Groxxb, Gut95i, Har96b, Harxxa, Hol97h, Hol97i, Hol98b, Holxxc, Holxxe, Holxxa, Hym97, JLW96, JC97, Kro97, Kro98e, LaM95d, Lad97, Ladxxa, LA97a, LA97b, Les97, Lesxx, Lin97, Lonxxb, Manxxb, Marxxc, Mas97b, McC98b, McH97, Mes97, Met98, Mit97b, M^{+}97b, M^{+}97a, M^{+}xxc, Morxxe, MP96, Pat98b, Per96b, Per97d, Per97a, Per97b, Perxxa, Perxxc, Perxxb, Rin97b, Ritxxi, Sha97a, Siy96, Siy97, Siy98, SP97, Swa97f, Swaxx, Tit97, Ubo96b, WG97b, Wil98c, WOLP96, W^{+}96, WL97c, WPL97, W^{+}97c, WL98, W^{+}xxa, W^{+}xxb, Woo96, Woo97b, Wooxx]. VisualAge [CLR98, ALP99, Ano96-198, Ano96-199, Ano98c, Ano98-36, CBA98, C^{+}xxa, Cas96b, Cox95a, McC98a, Pol98]. visualisation [CH98a]. visualisations [Nap98]. visualiser [Nap98]. Visualization [AJ98a, Ano97-121, BRF98, KK98a, LB98a, MD99, Sch97j, YN96, ZLGS99, AS96, DV98, JA98, KK97c, Nap97, RBF98, SML96]. Visualization’96 [CEG96]. visualize [BV98]. Visualizer [BW98a, Ano96-201, HRW^{+}96]. visualizers [BBG^{+}99]. Visualizing [DS99, KH99, LR96, MB98a]. Visuel [F^{+}xxc]. VM [Wil99c, Bot98a, Bue98, Mil98, YLP^{+}99]. VM/ESA [Bue98]. VMTs [Wip96]. Vocabulaire [BKP97a]. Voice [Ano97-51, Ano96-153]. Vol [DR96, GYT96c, GYT96d, G^{+}xxf, G^{+}xxg, G^{+}xxd, G^{+}xxe, Gra99b, Sav96]. voll [Dav96b]. Vollversion [BKP97b]. Volume [Add99, Ano99d, Ano99e, Ano99f, CLK98, CLK99b, Fla97a, GM98a, Gea98, Gea99, GYT97a, GYT97b, Gra99a, HJS98, HC97, HC98b, HC99b, HC99c, MD99]. Volunteer [Sar98, SHW98a, Sar99, SHW98b, SH99]. voor [Dacxx, KK97e, KKxxc, Walxxf, Walxxg]. Vortex [DDG^{+}96]. Vorwissen [Vac97]. Voyager [Gla98, Ham97b, Chw98]. VP [Ubo96b]. VPN [Ano98-29]. VR [Che96, SD99]. VRML [Bat96d, Die96, Diexx, AVN98, Ammxx, And96a, Ano96-202, Ano97-67, Anoxxn, BV98, Bat96c, BKP97a, Bot98b, BB99, CIF98, Cho97b, CJ98, DE96b, Dei96a, Div98, EBB^{+}96, Eng97a, Gar97b, Gar97c, GD98, GPV97, Gre98a, HJS98, Her96b, JLW96, Kar96, KSB97, LMM96, LB98a, L^{+}xx, Mac96, Mau98, MD98a, MD98b, MCxx, O’C97, Par95g, RCRBR97, Salxx, Sha98b, Spe98, SD96, SS96, Sun97c, TRMR97, Tub95, Ubo96b, Wat95c]. VRtuoso [Kuk98]. vs [Ano96-66, Ano96-111, Ano96-182, Ano97-107, GA98, PV99, Pre99, RGT96, SB97, VI96, Whe97]. VST [Ano96-196]. VXML [Law99].

W [Ano97l, Ano98k, vD99b]. W3 [NTS^{+}96]. W3Control [Ano98-27]. W95 [Per96a]. wa [Kim97d, IEE96g, USE96a]. Wabi [Ano97-121]. Wafers [Gar97a]. Wait [Bar95b, Hay95e, Har99d]. Waite [GM98c, Pot99]. wakaru [IK96]. Wake [Gut95g, Dry96b]. Wakes [Nel95a]. Waking [Gon95b]. Wall [Cou97a, Cha96a, DDJ99b, Gau96b, Law98, Tom96]. Walts [Mur96a]. WAN [Sau96]. Want [Ano96-53, Pfe99]. WAP [Lut99c]. War [Gar97b, Gar96c, JN95b, Law98, Lew98, Swa98a, Win95c, JN95a, Sch97f, SG97a, Sin95b]. Warehouse [CJ98, Ste96f]. Warp [DiD97c, DiD97d, S^{+}96b, DiD96, DiD97a, Kim96]. warring [Cha97c]. Wars [GC97]. Was [Lew97b, Her96b]. Washington [CI96, Cim96, DDJ98b, DDJ98c]. WAT [PTB^{+}97]. Watch [Ano96-29, Cla97b, Hes95, RA97, Ste97i]. Watchdog [Ano97-119]. Watered [Ols97]. Wave [Kro98e, Sch96d, Win95d, HMG^{+}95, Kno96b, VI96, Ano97-131, Ubo96b]. wavelets [Ano96-186]. Way [Cli98b, Ent97, Jam96, Pit97, TD97, Bab97a, Bru97, FC96a, Sco96a, Wal96a, Way96c, Ano96-53, PTB^{+}97]. ways [Fri98, Hal98a, Ude97b]. Weaknesses [Ano96-200]. Weapons [Swa98a]. Weather [Ano96-201, Whi96d, HRW^{+}96]. weatherman [Whi96d]. Weaves [Van96d]. weaving [Bat96b]. Web [Ano96-61, Ano96-135, Ano96-137, Ano96-206, Bat96d, BUP99, DE96b, Her96b, Hob98b, Lad98a, LM97a, MNO^{+}96, OW95, Salxx, San98a, Sin97b, Sinxx, vHO98, vSS96d, Bra96, Bat96b, O’C97, Ano97-154, Asl98, TTD96b, AA97, Ado95, Agn97, Ait96c, Ale98, Ano95g, Ano95p, Ano95r, Ano95s, Ano96l, Ano96-34, Ano96-195, Ano96-42, Ano96-53, Ano96-107, Ano96-118, Ano96a, Ano96-125, Ano96-123, Ano96-148, Ano96-192, Ano96-202, Ano96-204, Ano96-149, Ano97b, Ano97t, Ano97-36, Ano97-106, Ano97-112, Ano98b, Anu96b, Aoy97, Aoyxxb, App95, Atk99, AA96b, Bacxxa, BCLT96, Bal96a, Ban95a, BKKW96, Bar95a, BLB99, Bat96c, Beh95, BN97, BN98a, BBC^{+}97, BV97, BKT96, Boo96b, BDKK96, BGGR98, BC98c, Bra96, Bre96, BK96c, BKxxb, Bus95, CBT97a, CBT99, Cag96, Cha96a, CMM^{+}97, CT98, CKTV96, CBJ97b, CBJ97a, Cohxx, Col95].

Web [CT96, CWS98, CS99, CCB96b, Cor97a, Cox95a, Cox95c, Cro97, Cro95b, DR^{+}97, Dic96, Dic97b, Dicxx, DF97, Dom96, Don96, Eri96b, Fie96, Feg97, Fegxx, Fet97, FI96a, Fis96, FGN^{+}97, FFHA98, For96, FT96, FF96b, Fox97c, FFKO99, FLW96, Fre96a, FI96b, Fuc96, GH97, Gar97b, Gar98a, GS97a, Gau97e, Gep98, GS96, GJ^{+}96c, Gol98b, Gon95b, GG96a, GRD96, Gul96a, GR96, GR97, Gulxx, Gul96b, GGxxa, Hal97a, Hal97b, Hal99, Ham97b, HNS97a, Har96a, Har96b, HW97, HFR96, HFR98, Hay95e, Hay96s, Hei97, Her96b, Hib96, Hig98, HS98a, Hor96a, Hor96b, HS98b, How96, HLRT97, HK99, Isa96, JADR96, JLW96, JPT^{+}99, JJBW98, Jep96, JOA98, JJG98, KW96, Kal95b, Kan97, Kap97a, Kar95a, Kar95b, KK96a, KK97b, Key98, Kne96, Koh95, Kro98e, Kro98b, Kro99a, Kro99b, Kro99d]. Web [Kro99c, Küh96, LaM95c, LaM95g, Law98, Laz99, Lei95, LM96, LS96, Lem97a, LH96, Lev96c, LB97a, LBS98, LVS^{+}96a, LLxx, Lin96b, LQ99, Lot95, LCLL97, Lut98a, MAS96, MD99, Mag95, Man96b, Man96d, Man96e, MR99, Mar97a, MTU99, Mas95, MF96a, McC97a, McC98b, McC96d, McCxxb, MMJ96, McG97, Mel96, MMM96, Mes95a, Mes95d, Met95c, Moh96, MZ97, Mon99, Moo95, Nap97, Nas95a, Nas95c, NS98b, Nel95b, Net95c, OH97a, O’C97, OHE97a, Owe96, Par95g, Par95h, PSS97, Pel95, Per96c, Per97c, Com95g, PM97d, P^{+}98b, Pro96, RK97b, Rei96, Rem97, Rey96, Ric95a, RFPG96, Ric95c, Rod96a, RNBV98, Ros96, RC98, Sil95, Sar98, SH99, S^{+}96a, Sco96c, Sel99, SF96c, Sha97b, SH96, Sim96a, SC97b, Sin97a, SZTR96, Sin96b, Skl95, SW96, SD96, Sta97a, SPV99, Ste97g, Suaxx]. Web [SS96, Suš97, Tat96, TA97, TK98, TK99b, Tes96a, Tes96b, Tic97, TPW96, Ude96, Ude97a, Ude97d, Ude97b, Vad95a, Val96b, VO98, Van96d, Var96, VS96, Wag97b, Wal96h, Walxxb, Wal96m, Walxxc, Wat95d, Wat95e, Way96g, Way96f, Wea97a, WKES98, WKE98, WO97, Whi96d, Wil96d, Wil97a, WK95, Win95f, Win95i, Win95j, Win95k, Win95n, Wod96, Woo99b, WB^{+}96c, LVS^{+}96b, Xu97, YLQ98, Yos97, You96b, vSS96b, vSS96c]. Web-based [BUP99, vHO98, Asl98, Don96, Kro98b, Sar98, Ano96-34, Ano97t, Ano98b, CMM^{+}97, CT98, CT96, DF97, FT96, HW97, JADR96, Lin96b, LQ99, LCLL97, PSS97, RNBV98, SH99, VO98]. Web-Database [Lei95]. Web-Datenbank-Integration [Dic97b, Dicxx]. Web-Executable [SH96]. Web-Oriented [Kro99c]. Web-page [Ano96-53]. Web-Programmierrung [Bacxxa]. Web-Programmierung [Lad98a, CBJ97a, Cohxx]. Web-Savvy [Kro99a, Kro99b]. Web-Seiten [DE96b, vSS96d]. Web-to-Legacy [Kro98e]. Web-Viewer [Beh95]. Web/Commodity [FFKO99]. Web/Java [BBC^{+}97, FF96b, Ano96-192]. Web/Java-based [FF96b, Ano96-192]. Web Sight [Rei96]. WebAdmin [Wes97a]. Webbing [GJ96d]. WebBurst [Ano96-203]. WebCompiler [Kro99a, Kro99b]. WebEdge [Ano96-43]. Webfest [Her96a, Her96b]. WebFlow [BBC^{+}97]. WebHLA [FFKO99]. Webl [MR99]. Webmaster [Anoxx-80, Bro96, GGA96, SB96]. Webmastery [Fal96]. WebObjects [Par95b]. WebRunner [Mor95]. Webs [Die96, Diexx]. WebSite [MPxx]. Websites [Hob98c, Hob98d]. Webstat [WO97]. WebThreads [Ano97-121]. Webware [Fel97]. Week [Ano96-50, Dan96c, Ano96-73, Dan96d, Joh97a, Tat96]. Weg [Pie97]. wei [Nau96b]. Weight [KS98a, KS98b]. Welcome [DiG95d, Jan96]. Well [Dav97d]. Well-mannered [Dav97d]. Welten [Ammxx, Amm97a]. Werken [Holxxe, Holxxa, Purxx]. Werkzeuge [SBH97]. Western [BUP99]. WET [IEE96m]. WFC [Joh99, Wat98, Wat99]. wheeling [O’C95c]. Where [MF96a, SWJ97]. Which [Mes95d, Sch97f, Ano96-59, Han97, Mat96b]. While [Lav96, Oua96, Stu98, Ude97d]. Whinston [KW96]. Whirlwind [Cla97c]. White [GM96, Kro98d, Kro98c]. Whiteboard [Gut95b]. Whither [Gro96b, SF97]. whiz [Tra95b]. Who [Bro98a, Bro99c, CAB^{+}98, Gau97n, Hel97c, Ubo96b]. Whole [Hol96a, Nan95, SF96c, Ubo96a]. Whose [Sch95]. Wide [AA97, Ale98, Ano95g, Ano95p, Ano95r, Ano96-148, Ano96-206, AA96b, Bat96d, Bat96c, BV97, BK96c, BKxxb, Cha96a, Cor97a, DR^{+}97, Fis96, Fre96a, FI96b, Gep98, GS96, Gol98b, HLRT97, JJG98, Kap97a, Küh96, LVS^{+}96a, Lot95, MAS96, McC96d, McCxxb, MMM96, Moh96, Nap97, NS98b, Nel95b, OW95, Owe96, RC98, Sco96c, SZTR96, SW96, Suaxx, SS96, Suš97, Val96b, Wal96m, Walxxc, WO97, Whi96d, LVS^{+}96b, Xu97, Ano96-205, CRS^{+}96, MMJ96, SF96b, Way98a, Her96b, Man96b]. widen [Ano96-153]. widget [LH96]. Widgets [Cel98]. Widow [Com95g]. Wie [DE96a, Dre96]. Wiley [Ver98]. Will [Ano96-112, Ano97-166, Com95f, Gar97c, GC97, Hal97a, Hay95c, Lew96b, OH97b, Par95e, Sch96b, Ude97c, Ano95h, Ano95l, Ano96-102, Ano96-125, Ano97-52, Ban95b, Flo97, GH96a, GH96b, Gau97n, Gau97q, Gre97b, Gre97k, Hol97d, LH96, Mee96, Sco96a, Tub95, Way96e, Zol96]. William [Woe97]. Wilson [JG96]. Win [And97a, Ano96-188, Gau96g, Hay96y]. Win32 [For98a, O’C98, van97f]. WinCafé [Anoxx-81]. Window [Ano96-63, GYt96b, GYT96d, Gosxxb, Hem96b, Lut99d, Hol99, RFPG96]. windowing [NMS96]. Windows [Ano97-138, Ano98-29, Anoxx-44, DC97, Gla97b, Infxxa, Ano95a, Ano96-97, Ano96-105, Ano96-112, Ano96-140, Ano97-33, Ano97-55, Ano97-167, Anoxx-46, Coo97c, Del97, Gau96i, Gau97g, Gil98b, Hal97a, Hay96h, JCxx, JC97, Kro98d, Kro98c, LaM95b, NR97, Neo99, PM97c, Rie97a, San98b, San98c, Sch97c, Sch97i, Sym95, TJ96, Var97a, Wat98, Wat99, Way96g, Web97a, Wil97d, Wil98b, Win95m]. Winds [Ano97u]. Winner [Gv95]. Wins [Ano96-115]. WinWriters [Mat99]. wird [Dre96]. Wire [Pet95]. Wired [Sin96b]. wireless [Ano96-42]. Wiretap [Lea99]. Wissenswerte [DE96b]. within [AVN98, Ano96-59]. Without [BK99, RMS97, Syd96a, Ano96-196, Ano97-88, Bri98, Bru97, Nan95, Stu98, Whi96d]. Wizardry [Sha96a]. WMPI [MSS98]. Woes [Moo95, Ver97]. Wolfram [Bra96]. Womb [Gre96b]. Womb-to-tomb [Gre96b]. Wonder [Yat97]. Won’t [Ano96-98, Ano97-52, Kee96]. Woo [LaM95b]. Word [Ran97a, Ran97b]. Work [Ano96-52, Ano96-166, Lut99a, Ber96h, Gar98b, Kam99, KGG96, Kay96b, Lou95b, Lut99b, MGL^{+}98, Mes96, Pur98, Sto96, Amb97, Ber96b, Jav98, Sim96a, Wil96f]. Worker [Gar98a, BRS99]. Workers [Gar97b]. Workgroup [Ano96-137]. Working [Ano96-43, Ano96-204, BK99, DDJ99a, Lut99e, Ano96-53, Ano96-74, BN98b, Hay96q, LM98, Mee96]. WorkPlace [Ano97-153, Bra97a, Bra97b, Gil98c]. Works [Gos98b, Gau97p, Way96b, Ano97-102]. Workshop [ACM98a, Ano97-111, Ano98-28, Anoxx-71, Bro95b, CI96, CCF99, EMe97g, EMe97f, Hol96c, Hol97b, IEE96b, IEE96j, IEE99b, Kne98, LM96, MJA98, PZ96, SLPM96, Ass96b, USE96b, USE97a, USE98a, USE99d, USE99e, UCD99, WD^{+}96, You98a, LS96, Ano96-182, Cox95d, WJ98, Ano96-188, Ano97-146, Holxxd, Kanxxb, LPC96, LM97a, Lin96b, Pat98a, WJ97]. Workshops [IEE96m, Fox98b]. WorkSpace [WBA^{+}96]. Workstation [GPHL99, Way96g].

Workstations [Ano98-28, Gar98b]. World [Ano96-132, Ano96-159, Ano96-205, Ano98o, Bat96d, Bec97a, Bec98, BK96c, BKxxb, Hal97a, Küh96, Lut98b, Lut99d, Minxxe, OL98, Syd97, Tit96b, Van96d, And96a, CRS^{+}96, Cou97c, Jac96, KW96, KDH96, Lew96b, SF96b, Sli97, TG95, AA97, Ale98, Ano95g, Ano95p, Ano95r, Ano96-148, Ano96-206, AA96b, Bat96c, BV97, Cha96a, Cor97a, DR^{+}97, Fis96, Fre96a, FI96b, Gep98, GS96, Gol98b, Her96b, HLRT97, Jav96d, JJG98, Kal95a, Kap97a, LVS^{+}96a, Lot95, MAS96, Man96b, McC96d, McCxxb, MMM96, Moh96, Nap97, NS98b, Nel95b, OW95, Owe96, RC98, Sco96c, SZTR96, SW96, Suaxx, SS96, Suš97, Val96b, Wal96m, Walxxc, WO97, Whi96d, LVS^{+}96b, Xu97]. World-Ready [Bec97a]. World-wide [Ano96-205, CRS^{+}96, Her96b, Man96b]. Worlds [Gre97h, Gre98f, Hug97, LMM96, MCxx, GPV97, MD98a, MD98b]. Worldwide [MRA98, Urg96b, PCWGM98]. WorldWideWeb [BK97b]. worried [Hay96d, Mey96b]. worry [Stu98]. worst [Ude97d]. worthy [Gre97h]. Wrappers [All99f]. Wrapping [Gar97a]. Write [Ber96g, Ber97l, Ber97m, Ber97k, Cow97, Gau97g, Gre97h, TJ96, Ano96-181, Sco96a, Tob99]. Writer [Ano97-121]. Writers [DDJ99b]. Writing [Dar98c, Fri98, Har97b, Kin99, Rod96b, Rod97, Rodxxc, Shi99, Wes98, ZJ96, Ano95w, Bru97, Ste97d]. Written [Ano97-78, Ver96]. Wrong [CL96, Rob96]. Ww [Ano97-49, Per96a, Sam96b, Van96f]. Ww95 [Sam96b, Van96f]. WWW [AS96, Ano95q, Ano96-185, BCLT95, BCE98, Bou96a, C^{+}xxc, Che96, CSY97, CCB96a, CRV^{+}96, DE96b, Dec97a, Decxxa, DD97h, DD98b, EH96, Gar97e, HJS98, Hoy96, JAS96a, Jan96, Kal95b, KB96, MAS96, MH97c, Mur96a, OW95, Pew97b, Sul95b, TW96, Yas96, ZK98a, ZK98b]. WWW-based [Jan96]. WWW-OODB [Yas96]. WWW-Publishing [DE96b]. WWW-Seiten [DD97h, DD98b]. www.deliberate.com [Ano98-29]. www.security [MAS96]. Wyse [Ano96-115].

X [Ano96-135, Bay97a, vD99b, Ano96-30, Ano96-73, Ano97-45, Ano97-55, Ano97-168, Bow96b, Gre96a, Hay96d, Hay96f, JC97, Lav96, Mio96, Red97b, Way96f, dN97]. X-Designer [Ano97-45]. X11 [Ano97-119]. Xbase [Ano98-27]. XML [Ano97f, CH98b, Cha99f, FFKO99, LO99a, LO^{+}99b, MTU99, Wat98, Wat99]. XRT [Ano96-117]. XRT/table [Ano96-117]. XTerminal [Way96g]. Xynetix [Kro98e].

Y2K [Lut99a, BCP^{+}99b, Cha99e, Gar99, Law98, Lut99b, MGL^{+}98]. Ya [Dav97f, Davxxc]. Year [Wil97e, Wil97f, EHS^{+}99, Ano97-28, Bha98, Gar97b, Van96d]. Years [Ano96b, AJ99]. Yes [Hal97a, LP96a, LP96b, Sun96]. yin [Nau96b]. York [IEE96n, IEE96p]. Yourself [AM97, Ano96-28, Ano96-185, Cad97, Cad98b, Coh97e, Dan96c, Doh97, Dow97, FM97, Gof97, Hob97c, LP96e, LPW96, LPM96, LPC96, LP97d, Lem98, LC98, LPxxb, Man97a, Man97b, Mor96b, Ros98, Tuc98a, W^{+}96, WPL97, Dan96d, JLP96, KPF96, LC99, Tat96]. Yup [LP96a].

Z [BHxx, BC98c]. za [Suaxx, Suš97]. Zander [Cor95]. Zauberei [KS96b]. Ziel [SH97, Sch98b, Shu97a]. Zip [NN97]. ZOC [Ano96-40]. zone [Sha98b]. Zukai [IK96]. Zukowski [Ano97i, Zuk99]. zum [Dre96, KS96b, SH97, Sch98b, Shu97a]. zur [Pie97].

References

Aitken:1996:VJP

[A^{+}96] Peter Aitken et al. Visual J++ Programming FrontRunner. Coriolis Group Books, Scottsdale, AZ, USA, October 1996. ISBN 1-57610-064-2. ???? pp. LCCN ???? US$35.

URL http://www.coriolis.
com/site/msie/books/ind/jakprofr.
htm.

Ayers:1999:PJS

[A^{+}99] Danny Ayers et al. Professional Java Server programming. Wrox Press, Chicago, IL, USA, 1999. ISBN 1-86100-277-7. xxix + 1121 pp. LCCN QA76.73.J38 P76 1999.

Arano:1996:ETN

[AA96a] T. Arano and M. Aoyama.

Emerging technologies for network software development: past, present and future. In IEEE [IEE96s], pages 428–?? ISBN 0-8186-7579-9. LCCN QA76.6 .C6295 1996. IEEE Catalog number 96CB35986.

Ause:1996:HUW

[AA96b] Wayne Ause and Scott Arpajian. How to use the World Wide Web. Ziff-Davis Press, Emeryville, CA, USA, 1996. ISBN 1-56276-392-X (paperback). xiii + 223 pp. LCCN TK5105.888.A87 1996.

Adapalli:1997:WWW

[AA97] Sridhar Adapalli and Krishna Addepalli. World Wide Web integration of manufacturing process simulations. Concurrency: Practice and Experience, 9(11):1341–1350, November 1997. CODEN CPEXEI. ISSN 1040-3108. URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=13844;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=13844&PLACEBO=IE.pdf. Spe-

cial Issue: Java for computational science and engineering — simulation and modeling II.

Astley:1998:CCD

[AA98] Mark Astley and Gul A. Agha. Customization and composition of distributed objects: middleware abstractions for policy management. ACM SIGSOFT Software Engineering Notes, 23 (6):1–9, November 1998. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Alpern:1999:IJJ

[AAC^{+}99a] Bowen Alpern, C. R. Attanasio, Anthony Cocchi, Derek Lieber, Stephen Smith, Ton Ngo, John J. Barton, Susan Flynn Hummel, Janice C. Sheperd, and Mark Mergen. Implementing jalapeño in Java. ACM SIGPLAN Notices, 34 (10):314–324, October 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

URL http://www.acm.org/
pubs/citations/proceedings/
oops/320384/p314-alpern/.

Alpern:1999:IJN

[AAC^{+}99b] Bowen Alpern, C. R. Attanasio, Anthony Cocchi, Derek Lieber, Stephen Smith, Ton Ngo, John J. Barton, Susan Flynn Hummel, Janice C. Sheperd, and Mark Mergen. Implementing jalapeño in Java. ACM SIGPLAN Notices, 34 (10):314–324, October 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Abello:19xx:JLI

[Abexxa] Antonio M. Ferrer Abelló. Java el Lenguaje de Internet. Abeto Editorial, ????, 19xx. ISBN 84-89554-45-5. ???? pp. LCCN ???? 1995 Ptas.

Abello:19xx:LHJ

[Abexxb] Antonio M. Ferrer Abelló.

Lenguajes HTML, Java y CGI. Abeto Editorial, ????, 19xx. ISBN 84-89554-27-7. ???? pp. LCCN ???? 1995 Ptas.

Ablan:1996:DIA

[Abl96] Jerry Ablan. Developing Intranet Applications With Java. Sams.net, Indianapolis, IN 46268, USA, October 1996. ISBN 1-57521-166-1. xxiii + 492 pp. LCCN QA76.73.J38 A25 1996. US$45.00. URL http:

//www.mcp.com/cgi-bin/bag?
isbn=1-57521-166-1&last=/
bookstore;   http://www.mcp.
com/samsnet/. Includes CD-

ROM.

Abrahart:1996:GIC

[Abr96] R. J. Abrahart, editor. Geo-Computation 96. 1st International Conference on GeoComputation: Leeds, UK, 17–19 September 1996. University of Leeds, Leeds, UK, 1996. ISBN ???? LCCN ???? Two volumes.

AzanonEsteire:1998:JST

[AC98] Oscar Azañón Esteire and Juan Manual Cueva Lovelle. J — set of tools for native code generation for the Java Virtual Machine. ACM SIGPLAN Notices, 33(3):73–79, March 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Austin:1999:IEP

[AC99] Mark Austin and David Chancogne. Introduction to engineering programming: in C, Matlab and Java. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 1999. ISBN 0-471-00116-3. xv + 656 pp. LCCN QA76.73.C15 A98 1999. URL

ftp://uiarchive.cso.uiuc.
edu/pub/etext/gutenberg/;
 http://www.loc.gov/catdir/
description/wiley034/98008481.
html;   http://www.loc.gov/
catdir/toc/onix05/98008481.
html.

ACM:1996:PTA

[ACM96] ACM, editor. Proceedings of the Twelfth Annual Symposium on Computational Geometry, FCRC ’96. Philadelphia, PA, USA, May 24–26, 1996. ACM Press, New York, NY 10036, USA, 1996. CODEN PACGET. ISBN 0-89791-832-0. LCCN QA 448 D38 S96 1996.

ACM:1997:CRP

[ACM97a] ACM, editor. Conference record of POPL ’97, the 24th ACM SIGPLAN-SIGACT Symposium on Principles of Programming Languages: papers presented at the symposium, Paris, France, 15–17 January 1997. ACM Press, New York, NY 10036, USA, 1997. ISBN 0-89791-853-3. LCCN QA 76.7 A11 1997.

URL http://www.acm.org/
pubs/contents/proceedings/
plan/263699/index.html.

ACM:1997:PTA

[ACM97b] ACM, editor. Proceedings of the TRI-Ada’97 Conference, November 9–13, 1997, St. Louis, MO. ACM Press, New York, NY 10036, USA, 1997. ISBN 0-89791-981-5. LCCN QA 76.73 A35 T75 1997. Theme title: Ada; the right choice for reliable software. ACM order number: 825970.

ACM:1998:AWJ

[ACM98a] ACM, editor. ACM 1998 Workshop on Java for High-Performance Network Computing, Concurrency: Practice and Experience. ACM Press, New York, NY 10036, USA, 1998. CODEN CPEXEI. ISSN 1040-3108. LCCN ????

URL http://www.cs.ucsb.
edu/conferences/java98/program.

html. Also published as Concurrency: Practice and Experience, 10(11–13), September 1998, CODEN CPEXEI, ISSN 1040-3108.

ACM:1998:CRP

[ACM98b] ACM, editor. Conference record of POPL ’98: the 25th ACM SIGPLAN-SIGACT Symposium on Principles of Programming Languages: papers presented at the Symposium, San Diego, California, 19–21 January 1998. ACM Press, New York, NY 10036, USA, 1998. ISBN 0-89791-979-

3. LCCN QA76.7 .A15 1998.

URL http://www.acm.org/
pubs/contents/proceedings/
plan/268946/index.html. ACM

order number: 549981.

ACM:1999:PPA

[ACM99a] ACM, editor. POPL ’99. Proceedings of the 26th ACM SIGPLAN-SIGACT on Principles of programming languages, January 20–22, 1999, San Antonio, TX. ACM Press, New York, NY 10036, USA, 1999. ISBN 1-58113-095-3. LCCN ???? URL http://

www.acm.org/pubs/contents/
proceedings/plan/292540/index.
html.

ACM:1999:PASa

[ACM99b] ACM, editor. Proceedings of the ACM SIGPLAN ’99 Conference on Programming Language Design and Implementation (PLDI ’99), Atlanta, Georgia, 2–4 May 1999, volume 34(5) of ACM SIGPLAN Notices. ACM Press, New York, NY 10036, USA, 1999. ISBN 1-58113-094-5. LCCN ????

ACM:1999:SPO

[ACM99c] ACM, editor. SC’99: Oregon Convention Center 777 NE Martin Luther King Jr. Boulevard, Portland, Oregon, November 11–18, 1999. ACM Press and IEEE Computer Society Press, New York, NY 10036, USA and 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1999.

ACM:1999:SPCd

[ACM99d] ACM, editor. SIGGRAPH 99. Proceedings of the 1999 SIGGRAPH annual conference: Conference abstracts and applications, Computer Graphics. ACM Press, New York, NY 10036, USA, 1999. ISBN ???? LCCN ????

URL http://info.acm.org/
pubs/contents/proceedings/
graph/.

Adams:1998:CIO

[Ada98]  Joel C. Adams.  Chance-

it: an object-oriented capstone project for CS-1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(1):10–14, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

AddisonWesley:1999:JCL

[Add99] Addison Wesley. The Java Class Libraries, Second Edition, Volume 1/ The Java Class Libraries, Second Edition, Volume 2. Addison-Wesley, Reading, MA, USA, 1999. ISBN 1-199-45843-0 (??invalid checksum??). 3752 pp. LCCN ???? US$89.99.

Atkinson:1996:OPJ

[ADJ^{+}96] M. P. Atkinson, L. Daynes, M. J. Jordan, T. Printezis, and S. Spence. An orthogonally persistent Java. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(4): 68–75, December 1996. CO-DEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Agesen:1998:GCL

[ADM98] Ole Agesen, David Detlefs, and J. Eliot B. Moss. Garbage collection and local variable type-precision and liveness in Java Virtual Machines. ACM SIGPLAN Notices, 33(5):269– 279, May 1998. CODEN SIN-ODQ. ISBN 0-89791-987-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

                                        tronic). URL http://www.
                                        acm.org:80/pubs/citations/
                                        proceedings/pldi/277650/p269-
                                        agesen/;   http://www.cs.
                                        virginia.edu/pldi98/program.
                                        html.

Adobe:1995:AIJ

[Ado95] Adobe Systems Inc. Adobe to integrate Java technology for Web publishing. press release, December 7, 1995. URL

http://www.adobe.com/PDFs/
PR/9512/951206.java.pdf.

Alves-Foss:1999:FSS

[AF99] Jim Alves-Foss, editor. Formal syntax and semantics of Java, volume 1523 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. ISBN 3-540-66158-1. ISSN 0302-9743 (print), 1611-3349 (electronic). 404 pp. LCCN QA76.73.J38F67 1999. US$49.00.

Afergan:1996:JSI

[Afe96a] Michael Afergan. Java Soluciones Instantáneas. Prentice Hall HispanoAmericana, Mexico, DF, Mexico / Nueva York, NY, USA, 1996. ISBN 968-880-804-0. 432 pp. LCCN ???? 3500 Ptas. URL http:/

/prentice.com.mx/;   http:
//prentice.com.mx/libros/

javasi.html. Spanish translation of [Afe96b].

Afergan:1996:JQR

[Afe96b] Michael M. Afergan. Java Quick Reference. Que Corporation, Indianapolis, IN, USA, July 1, 1996. ISBN 0-7897-0868-X. xii + 419 pp. LCCN QA76.73.J38A34 1996. 40,00 DM, US$19.99. URL http://

www.mcp.com/2746397353063/
que/et/java_qr/;   http://
www.mcp.com/que/;   http:/
/www.mut.ch/profit.htx?&x_

                                        pk=kd84mf7ms9dm. Also avail-

able in Spanish translation [Afe96a].

Afergan:1997:DB

[Afe97]  Michael Afergan.  Discover

Bongo. IDG Books, San Mateo, CA, USA, 1997. ISBN 0-7645-3119-0. ???? pp. LCCN ???? US$25.

Alves-Foss:1999:FGJ

[AFF99] J. Alves-Foss and D. Frincke. Formal grammar for Java. Lecture Notes in Computer Science, 1523:1–??, 1999. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Akarsu:1998:DJBa

[AFH98a] Erol Akarsu, Geoffrey Fox, and Tomasz Haupt. DARP: Java-based data analysis and rapid prototyping environment for distributed high performance computations. In ACM [ACM98a], page ?? CODEN CPEXEI. ISSN 1040-3108. LCCN ????

                                        URL http://www.cs.ucsb.
                                        edu/conferences/java98/papers/
                                        darp.pdf;   http://www.cs.
                                        ucsb.edu/conferences/java98/
                                        papers/darp.ps. Also pub-

lished as Concurrency: Practice and Experience, 10(11– 13), September 1998, CODEN CPEXEI, ISSN 1040-3108.

Akarsu:1998:DJBb

[AFH98b] Erol Akarsu, Geoffrey Fox, and Tomasz Haupt. DARP: Java-based data analysis and rapid prototyping environment for distributed high performance computations. Concurrency: Practice and Experience, 10(11–13):1097–1105, September 1998. CODEN CPEXEI. ISSN 1040-3108. URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=10050424;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=10050424&PLACEBO=IE.pdf.

Special Issue: Java for Highperformance Network Computing.

Arnold:19xx:JTT

[AFHxx] Ken Arnold, Eric Freeman, and Susan Hupfer. JavaSpace Technology: a Tutorial and Reference Guide. Addison-Wesley, Reading, MA, USA, 19xx. ISBN 0-201-30955-6. ???? pp. LCCN

????  URL http://www2.
awl.com/cseng/javaseries/
index.html#Overview.

Alves-Foss:1999:DDS

[AFL99] J. Alves-Foss and F. S. Lam.

Dynamic denotational semantics of Java. Lecture Notes in Computer Science, 1523:201– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Agesen:1997:ATP

[AFM97] Ole Agesen, Stephen N. Freund, and John C. Mitchell. Adding type parameterization to the Java language. ACM SIGPLAN Notices, 32(10):49–65, October 1997. CODEN SINODQ.

ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Aitken:1997:JPV

[AFPW97] Peter Aitken, David H. Friedell, Jr., Anthony Potts, and Gerhard Wilhelms. Java Programmierung mit Visual J++. International Thomson Computer Press, 20 Park Plaza Suite 1001, Boston, MA 02116 USA, 1997. ISBN 3-8266-0318-4. ???? pp. LCCN ???? 69 DM. URL http:

                                        //www.itp.de/. Includes CD-

ROM.

Arnold:1996:MPJ

[AG96a] K. Arnold and J. Gosling. Multithreaded programming in Java. Web Techniques, 1(7):34– 40, 42–43, October 1996. CO-DEN WETEFA. ISSN 1086-556X.

Arnold:1996:PJ

[AG96b] Ken Arnold and James Gosling. Die Programmiersprache Java. Addison-Wesley, Reading, MA, USA, second edition, 1996. ISBN 3-8273-1034-2. xx + 366 pp. LCCN ???? 69,90 DM.

                                        URL http://www.addison-
                                        wesley.de/;   http://www.
                                        addison-wesley.de/katalog/
                                        item.ppml?textexpr=Java&id=
                                        00072;   http://www.javasoft.
                                        com/people/jag/.

Arnold:1996:JDP

[AG96c] Ken Arnold and James Gosling. Java — Didattica e programmazione. Addison Wesley Italia Editoriale, Via Deffenu, 7, 20133 Milano, Italy, 1996. ISBN 88-7192-058-9. 340 pp. LCCN ???? 39,000

Lire. URL http://www.doit.
it/Addison/;   http://www.
doit.it/Addison/cataloghi.
cgi?f=3D88-7192-058-9&pict=
3D1;   http://www.doit.it/
Addison/cataloghi.cgi?f=88-
7192-058-9&pict=1.

Arnold:1996:JPL

[AG96d] Ken Arnold and James Gosling.

The Java Programming Language. The Java Series. Addison-Wesley, Reading, MA, USA, May 1, 1996. ISBN 0-201-63455-4. xviii + 333 pp. LCCN QA76.73.J38A76 1996. US$34.50. URL http://www.

aw.com/cp/javaseries.html;
 http://www.aw.com/cseng/
titles/0-201-63455-4/. Also

available in French [AG96e] and Spanish [AGxxb] translations. See review in [Gar96c].

Arnold:1996:LJ

[AG96e] Ken Arnold and James Gosling.

Le Langage Java. International Thomson Publishing France, Paris, France, 1996. ISBN 2-84180-994-3. ???? pp. LCCN ???? 250 FF. URL http:

//www.thomson.com/intluk.

html. Translated by Serge Chaumette and Alain Miniussi. French translation of [AG96d]. Includes CD-ROM.

Arnold:1998:JPL

[AG98] Ken Arnold and James Gosling.

The Java Programming Language. Addison-Wesley, Reading, MA, USA, second edition, 1998. ISBN 0-201-31006-6. xix + 442 pp. LCCN QA76.73.J38A76 1998. US$39.

Arnold:19xx:PJb

[AGxxa] Ken Arnold and James Gosling. Die Programmiersprache Java. Addison-Wesley, Reading, MA, USA, 19xx. ISBN 3-8273-1034-2 (??invalid ISBN??). LCCN ???? 80 DM.

Arnold:19xx:LPJ

[AGxxb] Ken Arnold and James Gosling. El Lenguaje de Programacion Java. Addison-Wesley, Reading, MA, USA, 19xx. ISBN 84-7829-010-9. ???? pp. LCCN ???? 3950 Ptas. Spanish translation of [AG96d].

Arnold:19xx:JPLa

[AGxxc] Ken Arnold and James Gosling. The Java Programming Language. GOTOP Information Inc., 5F, No.7, Lane 50, Sec.3 Nan Kang Road Taipei, Taiwan; Unit 1905,Metro Plaza Tower 2, No.223 Hing Fong Road, Kwai Chung, N.T., Hong Kong, 19xx. ISBN 957-641-852-6. ???? pp. LCCN ???? 400 NT$.

                                        URL http://java.sun.com/
                                        people/arnold/;   http://
                                        www.gotop.com.hk/cl055.htm;
                                        http://www.gotop.com.tw. Chi-

nese translation by Khang Zien Hwa and others.

Arnold:19xx:JPLb

[AGxxd] Ken Arnold and James Gosling. The Java Programming Language. Piter Publishing Company, ????, 19xx. ISBN 5-88782-218-X. LCCN ???? ???? Russian translation.

Arnold:19xx:PJa

[AGxxe] Ken Arnold and James Gosling.

Programando en Java. Makron Books do Brasil Editora Ltda., Rua Tabapuã 1348, São Paulo SP - 04533-004 - Brasil, 19xx. ISBN 85-346-0772-9 (??invalid ISBN??). LCCN

???? ???? URL http://www.
awl.com/cseng/javaseries/
international/85-346-0772-
9.gif.

Atwood:1997:PCF

[AGD97] Christopher A. Atwood, Rajat P. Garg, and Dennis Deryke. A prototype computational fluid dynamics case study in Java. Concurrency: Practice and Experience, 9(11):1311–1318, November 1997. CODEN CPEXEI. ISSN 1040-3108. URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=13842;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=13842&PLACEBO=IE.pdf. Spe-

cial Issue: Java for computational science and engineering — simulation and modeling II.

Agha:1997:ECC

[Agh97] Gul Agha. From the Editor-in-Chief: Concurrent Java. IEEE Concurrency, 5(4):2–3, October/December 1997. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic).

                                        URL http://dlib.computer.
                                        org/pd/books/pd1997/pdf/p4002.
                                        pdf.

Agnvall:1997:WGH

[Agn97] E. Agnvall. Web glitz: how much is enough? Business Marketing, 82(1):M3–??, M21, January–February 1997. CO-DEN BUMAED. ISSN 0745-5933.

Allen:1997:HJ

[AH97] Mitch Allen and John B. Harvie. Hands On JavaBeans. Prima Publishing, Rocklin, CA, USA, 1997. ISBN 0-7615-1047-8. xxiv + 447 pp. LCCN QA76.73.J38A45 1997. US$40. URL http:

                                        //www.primapublishing.com/
                                        cpd/76151047.html.

Anderson:1999:RJS

[AH99] K. R. Anderson and T. J. Hickey. Reflecting Java into Scheme. Lecture Notes in Computer Science, 1616:154– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alexandrov:1997:SRI

[AISS97] Albert D. Alexandrov, Maximilian Ibel, Klaus E. Schauser, and Chris J. Scheiman. Super-Web: research issues in Javabased global computing. Concurrency: Practice and Experience, 9(6):535–553, June 1997. CODEN CPEXEI. ISSN 1040-3108. URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=13876;
http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=13876&PLACEBO=IE.pdf.

Aitken:1996:AGJ

[Ait96a] Gary Aitken. Automatically generating Java documentation: javadoc and the doc comment. Dr. Dobb’s Journal of Software Tools, 21(7):42, 44, 46–49, July 1996. CODEN DDJOEB. ISSN 1044-789X.

Aitken:1996:MCJ

[Ait96b] Gary Aitken. Moving from C++ to Java. Dr. Dobb’s Journal of Software Tools, 21(3):52, 54–56, March 1996. CODEN DDJOEB. ISSN 1044-789X.

Aitken:1996:WDG

[Ait96c] Peter G. Aitken. Web Developer’s Guide to JavaScript and VBScript. Coriolis Group Books, Scottsdale, AZ, USA, July 1, 1996. ISBN 1-883577-97-7. xxi + 293 pp. LCCN QA76.73.J39A48 1996. US$39.99; CDN$55.99.

URL http://165.247.200.
139/cgi-bin/java.cgi?page=
ibook&graphics=1#JPE;   http:
//www.coriolis.com/cgi-win/
KBsearch.exe. Includes CD-

ROM. Also available in French translation [Ait97b].

Aitken:1997:JV

[Ait97a] Peter Aitken. JavaScript und VBScript. International Thomson Computer Press, 20 Park Plaza Suite 1001, Boston, MA 02116 USA, 1997. ISBN 3-8266-0321-4 (??invalid ISBN??). 328 pp. LCCN ???? 50

DM. URL http://www.itp.
de/;   http://www.thomson.
com/intluk.html.  Includes

CD-ROM.

Aitken:1997:GJV

[Ait97b] Peter G. Aitken. Guide de JavaScript et VBScript. International Thomson Publishing France, Paris, France, 1997. ISBN 2-84180-145-4. xvii + 391 pp. LCCN ???? 298 FF.

URL http://www.thomson.
com/intluk.html.  French

translation of [Ait96c]. Includes CD-ROM.

Aitken:19xx:JPV

[Aitxxa] Peter Aitken. Java Programmierung mit Visual J++. International Thomson Computer Press, 20 Park Plaza Suite 1001, Boston, MA 02116 USA, 19xx. ISBN 3-8266-0318-4 (??invalid ISBN??). LCCN ???? 69 DM.

Aitken:19xx:VJF

[Aitxxb] Peter Aitken. Visual J++ FrontRunner. International Thomson Computer Press, 20 Park Plaza Suite 1001, Boston, MA 02116 USA, 19xx. ISBN 4-89369-513-4. ???? pp. LCCN ???? 4,500 yen. URL http:/

/www.itpj.co.jp/washo/513.
shtml.

Abbott:1998:DDO

[AJ98a] Mark Abbott and Lalit Kumar Jain. DOVE: Distributed Objects based scientific Visualization Environment. In ACM [ACM98a], page ?? CODEN CPEXEI. ISSN 1040-3108. LCCN ????

URL http://www.cs.ucsb.
edu/conferences/java98/papers/
dove.pdf;   http://www.cs.
ucsb.edu/conferences/java98/
papers/dove.ps. Also pub-

lished as Concurrency: Practice and Experience, 10(11– 13), September 1998, CODEN CPEXEI, ISSN 1040-3108.

Atkinson:1998:POP

[AJ98b] M. Atkinson and M. Jordan.

Providing orthogonal persistence for Java. Lecture Notes in Computer Science, 1445:383– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Atkinson:1999:IRT

[AJ99] M. Atkinson and M. Jordan.

Issues raised by three years of developing PJama: An orthogonally persistent, platform for Java[TM]. Lecture Notes in Computer Science, 1540:1– 30, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anupam:1999:UPV

[AKM99] Vinod Anupam, David M. Kristol, and Alain Mayer. A user’s and programmer’s view of the new JavaScript security model. In USENIX [USE99b], page ?? ISBN 1-880446-26-X. LCCN TK5105.875.I57

                                        U96.  URL http://www.
                                        usenix.org/publications/library/
                                        proceedings/usits99/anupam.
                                        html.

Al-VandeerVeer:1997:CJ

[Al-97] E. A. Al-Vandeer Veer. Clientside JavaScript. Web Techniques, 2(2):48–52, February 1997. CODEN WETEFA. ISSN 1086-556X.

Arnett:1996:IPJ

[AL96] Nick Arnett and David Land. Internet Programming with JavaScript. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, August 1996. ISBN 1-56592-205-0. 250 pp. LCCN ???? US$24.95. URL http:/

                                        /www.ora.com/.

Alpert:1997:UPC

[AL97] Sherman R. Alpert and Richard B. Lam. The ultimately publishable computer science paper for the latter ’90s: a tip for authors. Communications of the ACM, 40(1):94, January 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). This parody paper manages to use Java and JavaScript more than a dozen times.

Alagic:1999:TCO

[Ala99] Suad Alagíc. Type-checking OQL queries in the ODMG type systems. ACM Transactions on Database Systems, 24(3): 319–360, September 1999. CO-DEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

URL http://www.acm.org/
pubs/citations/journals/tods/
1999-24-3/p319-alagic/.

Alexander:1996:SCS

[Ale96] M. Alexander. Surf client/ server systems via the Internet. Datamation, 42(2):67–68, January 15, 1996. CODEN DTM-NAT. ISSN 0011-6963.

Alexandre:1998:JBP

[Ale98]  Thomas J. Alexandre.  A

Java-based platform for intellectual property protection on the World Wide Web. Computer Networks and ISDN Systems, 30(1–7):591–593, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.cl.cam.ac.uk/~fapp2/
steganography/bibliography/
072101.html;   http://www.
elsevier.com/cas/tree/store/
comnet/sub/1998/30/1-7/1863.
pdf.

Allison:1999:IJC

[All99a] Chuck Allison. import java.*:

Control flow — the bad, the good, the exceptional. C/C++ Users Journal, 17(5):??, May 1999. CODEN CCUJEX. ISSN 1075-2838.

Allison:1999:IJJ

[All99b] Chuck Allison. import java.*:

Jumping into java. C/C++ Users Journal, 17(1):??, January 1999. CODEN CCUJEX. ISSN 1075-2838.

Allison:1999:IJO

[All99c] Chuck Allison. import java.*: Object-oriented programming in Java. C/C++ Users Journal, 17(11):??, November 1999. CODEN CCUJEX. ISSN 1075-2838.

Allison:1999:IJP

[All99d] Chuck Allison. import java.*: Packaging your objects. C/C++ Users Journal, 17(9):42–??, September 1999. CODEN CCU-JEX. ISSN 1075-2838.

Allison:1999:IJT

[All99e] Chuck Allison. import java.*: Thinking in objects. C/C++ Users Journal, 17(7):79–??, July 1999. CODEN CCUJEX. ISSN 1075-2838.

Allison:1999:IJU

[All99f] Chuck Allison. import java.*: Using primitive types and wrappers. C/C++ Users Journal, 17(3):??, March 1999. CODEN CCUJEX. ISSN 1075-2838.

Akerley:1999:PVJ

[ALP99] John Akerley, Nina Li, and Antonello Parlavecchia. Programming with VisualAge for Java Version 2.0. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, second edition, 1999. ISBN 0-13-021298-9. xxx + 417 pp. LCCN ???? US$49.99.

Alsop:1995:CDTb

[Als95a] Stewart Alsop. Column: Distributed thinking — HotJava could be really hot. InfoWorld, page ??, May 30, 1995. CODEN INWODU. ISSN 0199-6649.

URL http://www.infoworld.
com/.

Alsop:1995:CDTc

[Als95b] Stewart Alsop. Column: Distributed thinking — Java’s hot but may cool. InfoWorld, page ??, December 18, 1995. CODEN INWODU. ISSN 0199-6649. URL http://www.

infoworld.com/.

Alsop:1995:CDTa

[Als95c] Stewart Alsop. Column: Distributed thinking — Netscape is one big fish. InfoWorld, page ??, April 11, 1995. CODEN IN-WODU. ISSN 0199-6649. URL

http://www.infoworld.com/.

Ablan:1997:TYM

[AM97] Jerry Ablan and Michael Morrison. Teach Yourself More Java 1.1 in 21 Days. Sams.net, Indianapolis, IN 46268, USA, November 1997. ISBN 1-57521-347-8. xxii + 479 pp. LCCN QA76.73.J38M6744 1997. US$40. URL http://

merchant.superlibrary.com:
8000/catalog/hg/PRODUCT/PAGE/
15752/bud/1575213478.html.

Ambrosio:1996:BJA

[Amb96] Johanna Ambrosio. Books on Java aimed at beginners, advanced users. Computerworld, 30(21):55, May 20, 1996. CO-DEN CMPWAB. ISSN 0010-

Ambler:1997:BOA

[Amb97]  Scott W. Ambler.  Build-

ing object applications that work: your step-by-step handbook for developing robust systems with object technology. SIGS Books, 71 W 23rd St., New York, NY 10010, USA, 1997. ISBN 0-13-789215-2, 0-521-64826-2. ???? pp. LCCN QA76.9.O35A45 1997. URL

http://www.ambysoft.com/;
 http://www.cup.org/sgs/0137892152.
html.

Amini:1996:WHAa

[Ami96a] Lisa Amini. What’s so HOT about Java? — Part 1. /AIXtra: IBM’s Magazine for AIX Professionals, 6(5):??, September/October 1996. URL http:

//pscc.dfw.ibm.com/aixtra/
aixtra11/tech/java.htm.

Amini:1996:WHAb

[Ami96b] Lisa Amini. What’s so HOT about Java? — Part 2. /AIXtra: IBM’s Magazine for AIX Professionals, 6(6):40– 51, November/December 1996.

URL http://pscc.dfw.ibm.

com/aixtra. Notes that IBM supplies the Java Development Toolkit (JDK) version 1.0.1 from the AIX V4.2 Bonus Pack, and version 1.0.2 is available from the IBM Centre for Java Technology home page at http://ncc.hursley.ibm.com/javainfo.

Amini:1997:WHA

[Ami97] Lisa Amini. What’s so hot about Java? — part 3. /AIXtra:

IBM’s Magazine for AIX Professionals, 7(1):45–47, 49–57, 59–60, January/February 1997.

Ammann:1997:PAW

[Amm97a] Eckhard Ammann. Programmierung animierter Welten. International Thomson Computer Press, 20 Park Plaza Suite 1001, Boston, MA 02116 USA, 1997. ISBN 3-8266-0329-X (??invalid ISBN??). 310 pp. LCCN ????

79 DM. URL http://www.itp.

de/. Includes CD-ROM.

Ammeraal:1997:BJ

[Amm97b] Leen Ammeraal. Basiscursus Java. Academic Service, Postbus 81, 2870 AB Schoonhoven, The Netherlands, 1997. ISBN 90-395-0718-X. 250 pp. LCCN ???? fl. 29,50. URL http:/

/acaserv.nl/bestel/5_0718.
htm;   http://www.acaserv.
nl; http://www.acaserv.nl/
5_0718.htm.

Ammeraal:1998:CGJ

[Amm98] L. (Leendert) Ammeraal. Computer graphics for Java programmers. Worldwide series in computer science. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 1998. ISBN 0-471-98142-7 (paperback). x + 271 pp. LCCN T385 .A488 1998. US$50.

URL http://www.wiley.com/
compbooks/catalog/98142-7.
htm.

Ammann:19xx:PAW

[Ammxx] Eckhard Ammann. Programmierung animierter Welten:

Java, JavaScript und VRML. International Thomson Computer Press, 20 Park Plaza Suite 1001, Boston, MA 02116 USA, 19xx. ISBN 3-8266-0329-X (??invalid ISBN??). LCCN ???? 79 DM.

Au:1997:JPB

[AMP97] Edith Au, Dave Makower, and Pencom Web Works. Java Programming Basics: For Java 1.1. MIS Press and M&T Books, P. O. Box 5277, Portland, OR 97208-5277, USA, Tel: (503) 282-5215 and M&T Publishing, Inc., 501 Galveston Drive, Redwood City, CA 94063, USA, second edition, April 1997. ISBN 1-55828-551-2 (softcover). ???? pp. LCCN ???? US$34.95. Includes CD-ROM.

Au:1996:JPB

[AMPM96] Edith Au, Dave Makower, Pencom Web Works, and Mark Mangan. Java Programming Basics. MIS Press and M&T Books, P. O. Box 5277, Portland, OR 97208-5277, USA, Tel: (503) 282-5215 and M&T Publishing, Inc., 501 Galveston Drive, Redwood City, CA 94063, USA, 1996. ISBN 1-55828-469-9. xix + 458 pp. LCCN QA76.73.J38 A9 1996. US$34.95. URL http://www.

mandt.com/javaprogbasics.
htm;   http://www.pencom.
com/javabasics/.  Includes

CD-ROM. See review in [Ano96-155].

Andoh:1996:VWS

[And96a] Y. Andoh. VRML’s 3D world: sculpting the next generation Internet — technology and trends in VRML. Joho-Shori (J. Information Processing Soc. Japan), 37(10):899–906, October 1996. CODEN JOSHA4. ISSN 0447-8053.

Andrews:1996:JFB

[And96b] David L. Andrews. Java forms:

Better than HTML. BYTE Magazine, 21(11):38, November 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic). Discuss the Jet-

Form (http://www.jetform.
com/) and Caere (http://www.

caere.com/) products for interactive forms applications.

Anderson:1997:JBH

[And97a] Thomas Anderson. Java For Business: How Companies Are Using Java to Win Customers and Make Money Now. Business Technology Ser. Van Nostrand Reinhold Co., New York, NY, USA, 1997. ISBN 0-442-02517-3. xiv + 354 pp. LCCN HF5548.5.J38J38 1997. US$24.95.

Andrews:1997:WNJ

[And97b] David L. Andrews. What’s next for Java office apps. BYTE Magazine, 22(11):24–??, November 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anderson:1998:NDG

[And98a] William F. Anderson. Netscape Developer’s Guide to JavaScript 1.2. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1998. ISBN 0-13-719279-7. 660 pp. LCCN QA76.73.J39A53 1998. US$39.95. URL http://

                                        www.prenhall.com/allbooks/
                                        ptr_0137192797.html.

Andrews:1998:JPJ

[And98b] Dave Andrews. Java poison or just PR? BYTE Magazine, 23 (1):22–23, January 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Angell:1999:PTE

[Ang99] Kirby W. Angell. Programmer’s toolchest: Examining JPython: a Java test engine puts Python to the test. Dr. Dobb’s Journal of Software Tools, 24(4):78, 81–83, April 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://

                                        www.ddj.com/ftp/1999/1999_
                                        04/jpython.txt;   http://
                                        www.ddj.com/ftp/1999/1999_
                                        04/jpython.zip.

Anonymous:1995:GNS

[Ano95a] Anonymous. GATEWAYS — Navigator 2.0 splits windows, talks Java. MacWEEK, 9(41): 18, ???? 1995. CODEN MWEEEI. ISSN 0892-8118.

Anonymous:1995:HJ

[Ano95b] Anonymous. Help from ‘Java’. The Chronicle of Higher Education, 42(6):A25, October 6, 1995. ISSN 0009-5982.

Anonymous:1995:IJJ

[Ano95c] Anonymous. IBM jumps on Java bandwagon. Electronic News, 41(2095):22, December 1995. CODEN ELNEAU. ISSN 0013-4937, 1061-6624.

Anonymous:1995:INS

[Ano95d] Anonymous. Internet: Now for some HotJava. Newsweek, 126 (20):87, 1995. ISSN 0028-9604.

Anonymous:1995:JBS

[Ano95e] Anonymous. JAVA: Borland and Spyglass sign up. Information Week, 554:20, November 20, 1995. CODEN INFWE4. ISSN 8750-6874.

Anonymous:1995:JGN

[Ano95f] Anonymous. Java gets the nod from IBM. Information Week, page 24, December 18, 1995. CODEN INFWE4. ISSN 8750-6874.

Anonymous:1995:JSW

[Ano95g]  Anonymous.  Java set for

World Wide Web. Graphic arts monthly: the magazine of the printing industry, 67(12):66, ???? 1995. CODEN GAMOE4. ISSN 1047-9325.

Anonymous:1995:JTB

[Ano95h] Anonymous. Java tools — Borland will create a Delphilike visual development tool set for Sun’s Java. Computerworld, 29(46):10, ???? 1995. CODEN CMPWAB. ISSN 0010-4841. URL http://www.

computerworld.com/.

Anonymous:1995:JLB

[Ano95i] Anonymous. JavaScript language becomes a competitor. Network Security, 1995 (12):3, December 1995. CO-DEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/1353485895901620.

Anonymous:1995:MJI

[Ano95j] Anonymous. Meet Java, the invisible computer. Business Week, 3453:82, December 04, 1995. CODEN BUWEA3. ISSN 0007-7135.

Anonymous:1995:NNA

[Ano95k] Anonymous. Netscape Navigator: Acrobat, Director, and Java support. MacUser, 11 (9):30, ???? 1995. CODEN MCUSEY. ISSN 0884-0997.

Anonymous:1995:NNW

[Ano95l] Anonymous. Netscape Navigator will incorporate Sun’s Java programming language. InfoWorld, 17(22):16, May 29, 1995. CODEN INWODU. ISSN 0199-6649. URL http://www.

infoworld.com/.

Anonymous:1995:NB

[Ano95m] Anonymous. News briefs. Dr. Dobb’s Developer Update, page 1, October 1995. ISSN 1079-8595. URL http://www.

ddj.com/dddu/issues/u510_

bri.htm. Discusses impact of an existing patent on executable content on Java.

Anonymous:1995:NSD

[Ano95n] Anonymous. News: Sun doubles first quarter earnings. Network World, page ??, October 20, 1995. ISSN 0887-7661 (print), 1944-7655 (elec-

tronic). URL http://www.
iquest.net/index.html.

Anonymous:1995:LSM

[Ano95o] Anonymous. OS Lite — Sun Microsystems plans to create a lightweight OS that builds on Java. LAN times, 12(26):37–??, 1995. ISSN 1040-5917.

Anonymous:1995:SBT

[Ano95p] Anonymous. Sun brings true interactivity to the World Wide Web. SunFlash, May 1995.

URL http://www.flashback.
com/www/may.1995/sunflash/
77.25.www.html.  Number

77.25.

Anonymous:1995:SLF

[Ano95q] Anonymous. Sun launches first interactive WWW site powered by Java. SunFlash, May 1995.

URL http://www.flashback.
com/www/may.1995/sunflash/
77.27.java.html.  Number

77.27.

Anonymous:1995:SMI

[Ano95r] Anonymous. Sun Microsystems Inc — Java, the racy new programming language for the World Wide Web, “is driving Sun in some completely new directions.”. Industry Week, 244(23):36–??, 1995. CODEN IWEEA4. ISSN 0039-0895.

Anonymous:1995:SOU

[Ano95s] Anonymous. Sun offers users a cupful of ’Net software to support electronic commerce applications, including its Hot-Java Web browser. Network World, 12(21):10, May 22, 1995. ISSN 0887-7661 (print), 1944-7655 (electronic). URL http://

                                        www.iquest.net/index.html.

Anonymous:1995:SJQ

[Ano95t] Anonymous. SunScreen and Java questions & answers. SunFlash, July 1995. URL

                                        http://www.flashback.com/
                                        www/july.1995/sunflash/79.
                                        05.sunscreen_qa.html. Num-

ber 79.05.

Anonymous:1995:SDJ

[Ano95u] Anonymous. SunSoft to deliver Java development tools. SunFlash, September 10, 1995.

                                        URL http://www.flashback.
                                        com/www/september.1995/sunflash/
                                        81.30.java_dev_tools.html.

Number 81.30.

Anonymous:1995:SRS

[Ano95v] Anonymous. SunWorld readers speak: Java is hot! Sunworld Online, 10:??, 1995. ISSN 1091-

                                        8914. URL http://www.sun.
                                        com/sunworldonline/swol-10-
                                        1995/swol-10-javadigest.html.

Anonymous:1995:FMS

[Ano95w] Anonymous. UP FRONT — Mac solution for writing Java applets. MacWEEK, 9(42):4, ???? 1995. CODEN MWEEEI. ISSN 0892-8118.

Anonymous:1996:KWT

[Ano96a] Anonymous. 200 killer Web tools. PC/Computing, 9(8): 131–??, August 1, 1996. CO-DEN PCMPEI. ISSN 0899-1847.

Anonymous:1996:YAM

[Ano96b]  Anonymous.  50, 100 and

150 years ago: 1946: Making high-octane gasoline 1896: The missing link in Java. 1846: Mesmerizing crime. Scientific American, 274(1):7–??, January 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.sciam.com/0196issue/
0196toc.html.

Anonymous:1996:AUG

[Ano96c] Anonymous. AIX users get their own dose of Java. RS/Magazine, 5(2):6–7, February 1996. ISSN 1061-0030.

Anonymous:1996:AAJ

[Ano96d] Anonymous. Amzi! announces Java class for Amzi! Logic Server. C/C++ Users Journal, 14(10):94, October 1996. CO-DEN CCUJEX. ISSN 1075-2838.

Anonymous:1996:JJDa

[Ano96e] Anonymous. And Java juices databases. SunExpert Magazine, 7(4):12, April 1996. ISSN 1053-9239.

Anonymous:1996:ASS

[Ano96f] Anonymous. Apple, Sun shake on server, Java fusion. RS/ Magazine, 5(11):11, November 1996. ISSN 1088-0844, 1061-0030.

Anonymous:1996:APA

[Ano96g] Anonymous. Applix pushes ahead with Java. Investment dealers’ digest, 62(13):11– ??, March 25, 1996. ISSN 0021-0080.

Anonymous:1996:AMDa

[Ano96h] Anonymous. Argus makes decaf for Java. SunExpert Magazine, 7(4):80, April 1996. ISSN 1053-9239.

Anonymous:1996:AMDb

[Ano96i] Anonymous. Argus makes decaf for Java. WebServer Magazine: For Managers of World Wide Web Sites, 1(1):78, March/June 1996. ISSN 1087-4232. URL

                                        http://www.cpg.com.

Anonymous:1996:BNB

[Ano96j] Anonymous. Backstage nerves [Backstage Designer Plus]. Publishing, ????(????):21–??, August 1996. CODEN PUBLFA.

Anonymous:1996:BBO

[Ano96k] Anonymous. The benefits of building object-oriented applications. Software Economics Letter, 5(8):1–3, August 1996. CODEN SECLE3. ISSN 1065-6146.

Anonymous:1996:BWB

[Ano96l] Anonymous. Best Web browser [Netscape Navigator 3.0 and Internet Explorer 3.0 comparison]. PC World, 14(8):136, 138–139, 142, 146, 148, August 1996. CO-DEN PCWDDV. ISSN 0737-8939 (print), 1944-9143 (electronic).

Anonymous:1996:BMP

[Ano96m] Anonymous. BlackWidow makes possible computing on Java. HP Chronicle, 13(9):26, August 1996. ISSN 0892-2829.

Anonymous:1996:BREn

[Ano96n]  Anonymous.  Book review:

Exploring Java: By Patrick Niemeyer and Joshua Peck. O’Reilly & Associates, Sebastopol, CA. (1996). 407 pages. $29.95. Computers and Mathematics with Applications, 32(6):135, September 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S089812219690259X.

Anonymous:1996:BRJa

[Ano96o] Anonymous. Book review: Java in a nutshell: a desktop quick reference for Java programmers: By David Flanagan. O’Reilly & Associates, Sebastopol, CA. (1996). 438 pages. $19.95. Computers and Mathematics with Applications, 31(11):139, June 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122196873469.

Anonymous:1996:BRJb

[Ano96p]  Anonymous.  Book review:

JavaScript: the definitive guide, (beta Edition): By David Flanagan. O’Reilly, Sebastopol, CA. (1996). 454 pages. $29.95. Computers and Mathematics with Applications, 32(10): 131, November 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic). URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0898122196900243.

Anonymous:1996:BJT

[Ano96q]  Anonymous.  Brewing Java

tools — SourceCraft mixes and matches C++ and Java. Computerworld, 30(7):48–??, ???? 1996. CODEN CMPWAB. ISSN 0010-4841.

Anonymous:1996:BBB

[Ano96r]  Anonymous.  The browser

battle between Microsoft and Netscape. PC Magazine, 15(10): 40–??, May 28, 1996. CODEN PCMGEP. ISSN 0888-8507.

Anonymous:1996:CBN

[Ano96s]  Anonymous.  Castanet and

Bongo: the next step for Java. Online & CDROM review: the international journal of online & optical information systems, 20(6):307–309, December 1996. CODEN ONCDEW. ISSN 0309-314X.

Anonymous:1996:CBJ

[Ano96t] Anonymous. Chips based on Java. Computer Design, 35(8):

14–??, July 1, 1996. CODEN CMPDAM. ISSN 0010-4566.

Anonymous:1996:CSC

[Ano96u] Anonymous. Client/server concepts: Jeans and Java. Manufacturing systems, 14(1):10–??, ???? 1996. CODEN MASYES. ISSN 0748-948X.

Anonymous:1996:CSP

[Ano96v] Anonymous. Codetalk: Symantec pours Java into its development environment. BYTE Magazine, 21(3):38–??, March 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:CJ

[Ano96w] Anonymous. Competition for Java. Information Week, 597: 13A–??, September 16, 1996. CODEN INFWE4. ISSN 8750-6874.

Anonymous:1996:CCJ

[Ano96x] Anonymous. Corel creates Java suite. Information Week, 595: 24–??, September 2, 1996. CO-DEN INFWE4. ISSN 8750-6874.

Anonymous:1996:CJC

[Ano96y] Anonymous. Correction in Java classes. (March 1, 1996 article, the one hour Java applet). Datamation, 42(????):21– ??, ???? 1996. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1996:CCP

[Ano96z]  Anonymous.  Cosmo Code

provides Java development environment. Silicon Graphics World, 6(6):12, June 1996. ISSN 1057-7041. URL http://

www.sgi.com/cosmo.

Anonymous:1996:CSS

[Ano96-27] Anonymous. Cover story: Sun’s rise. Business Week, ??(3459): 66, January 1996. CODEN BUWEA3. ISSN 0007-7135.

Anonymous:1996:CEU

[Ano96-28] Anonymous. Cutting edge — use DMI to reduce desktop support costs summer school for IS professionals. Java, Java everywhere. first Plan 9, now inferno. go beyond RAID-5. downsizing of IS slows in ’96. how to link relational and object databases. brace yourself for a price hike for ISDN service. Datamation, 42 (11):10–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1996:DW

[Ano96-29] Anonymous. Developments to watch. Business Week, ?? (3462):91–??, February 1996. CODEN BUWEA3. ISSN 0007-7135.

Anonymous:1996:DXL

[Ano96-30] Anonymous. Dimension X liquid motion. PC/Computing, 9 (10):111–??, October 1, 1996. CODEN PCMPEI. ISSN 0899-1847.

Anonymous:1996:DN

[Ano96-31] Anonymous. Drive for the net. Information Week, 592:37–??, August 12, 1996. CODEN IN-FWE4. ISSN 8750-6874.

Anonymous:1996:MSI

[Ano96-32] Anonymous. E-mail suite includes Java. Computerworld, 30 (52):65, December 1996. CO-DEN CMPWAB. ISSN 0010-4841.

Anonymous:1996:ET

[Ano96-33] Anonymous. Emerging technology. America’s network: technology for the information highway, 100(12):108–??, June 15, 1996. CODEN ANETE4. ISSN 0040-263X.

Anonymous:1996:ETW

[Ano96-34] Anonymous. Enabling technologies for Web-based ubiquitous supercomputing. IEEE International Symposium on High Performance Distributed Computing, Proceedings, pages 112–119, 1996. CODEN PIDCFB. ISSN 1082-8907. IEEE catalog number 96TB100069.

Anonymous:1996:ESC

[Ano96-35] Anonymous. Examining Symantec’s Café. Dr. Dobb’s Journal of Software Tools, 21(8):??, August 1996. CODEN DDJOEB. ISSN 1044-789X.

Anonymous:1996:FR

[Ano96-36] Anonymous. Field report.

Software magazine, 16(8):19– ??, August 1, 1996. CODEN SMWMEQ. ISSN 0279-9782 (or 0897-8085??).

Anonymous:1996:FRT

[Ano96-37] Anonymous. Field report: Tools vendors look to milk Java.

Software magazine, 16(3):23– ??, ???? 1996. CODEN SMWMEQ. ISSN 0897-8085.

Anonymous:1996:FKJ

[Ano96-38] Anonymous. Firewall keeps out Java. Information Week, 590: 30–??, July 29, 1996. CODEN INFWE4. ISSN 8750-6874.

Anonymous:1996:FLA

[Ano96-39] Anonymous. First look at Microsoft’s J++. Information Week, 595:77–??, September 2, 1996. CODEN INFWE4. ISSN 8750-6874.

Anonymous:1996:FLP

[Ano96-40] Anonymous. First looks — Pentium/133 notebooks. Java Development Kit 1.0. ZOC. Iomega Jaz drive. PC Magazine, 15(7):39–??, ???? 1996. CODEN PCMGEP. ISSN 0888-8507.

Anonymous:1996:FJL

[Ano96-41] Anonymous. Free Java library. WebServer Magazine: For Managers of World Wide Web Sites, 1(3):59, September/October 1996. ISSN 1087-4232. URL http://

www.objectspace.com/. Dis-

cusses ObjectSpaces Inc.’s Java Generic Library (JGL).

Anonymous:1996:GSC

[Ano96-42] Anonymous. GATEWAYS — Symantec C++ 8.0 release 5 builds Java, Pascal, 680x0 apps. breeze blows into Mac wireless networking. Engst on avoiding OT hassles. how many hits does your Web server get? MacWEEK, 10(14):16–??, ???? 1996. CODEN MWEEEI. ISSN 0892-8118.

Anonymous:1996:GSA

[Ano96-43] Anonymous. GATEWAYS — System 7.5.3 arrives. latest Internet Explorer beta squashes bugs. WebEdge III features Java galore. working the net: Java questioned and answered. MacWEEK, 10(10):14–??, ???? 1996. CODEN MWEEEI. ISSN 0892-8118.

Anonymous:1996:HJ

[Ano96-44] Anonymous. Hooked on Java.

Business Week, 3493:100–??, September 16, 1996. CODEN BUWEA3. ISSN 0007-7135.

Anonymous:1996:HJC

[Ano96-45] Anonymous. How Java can pay the rent. BYTE Magazine, 21(6):27–??, June 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:HJW

[Ano96-46] Anonymous. HTML and JAVA — what’s the difference? Library Systems, 16(2):15–??, ???? 1996. ISSN 0277-0288.

Anonymous:1996:IPJ

[Ano96-47] Anonymous. IBM ports Javabased technology. /AIXtra: IBM’s Magazine for AIX Professionals, 6(6):8, November/ December 1996. Notes that a new version of OS/2, known as Merlin, is Java-enabled, and includes a built-in developer’s kit for Java and run-time code. IBM’s OpenDoc-based Arabica provides support for the Java Beans initiative. Just-in-time Java compilation will be standard technology in AIX and OS/2. IBM has ported Java to the Windows 3.1 platform.

Anonymous:1996:ISC

[Ano96-48] Anonymous. IBM, Sun collaborate to cross license Joe, CICS, MQseries. /AIXtra: IBM’s Magazine for AIX Professionals, 6(6):8, November/ December 1996.

Anonymous:1996:IST

[Ano96-49] Anonymous. Imperial Software Technology builds Java strategy. Silicon Graphics World, 6 (7):17, July 1996. ISSN 1057-7041.

Anonymous:1996:BW

[Ano96-50] Anonymous. In business this week. Business Week, 3491:47– ??, September 2, 1996. CODEN BUWEA3. ISSN 0007-7135.

Anonymous:1996:IJC

[Ano96-51] Anonymous. Interactive 3D Java charting. WebServer Magazine: For Managers of World Wide Web Sites, 1(3):55–56, July/August 1996. ISSN 1087-4232. Discusses Infospace Inc.’s WebCharts product.

Anonymous:1996:IW

[Ano96-52] Anonymous. Internet @ work. LAN: the network solutions magazine, 5(1):75–??, June 1, 1996. CODEN LANNER. ISSN 1038-359X (or 1069-5621??).

Anonymous:1996:IAW

[Ano96-53] Anonymous. Internet appliances — want to surf the Web from your sofa? Oracle, Sun Microsystems, and other companies are working on USD 500 I-Way boxes. also hot this month a crash course in Java, 3-D Webpage designers, and a low-cost Web server. PC/Computing, 9 (3):291–??, ???? 1996. CODEN PCMPEI. ISSN 0899-1847.

Anonymous:1996:IMJ

[Ano96-54] Anonymous. Internet: Muchhyped Java, Sun Microsystems’ distributed applications programming language, may become a full alternative to C++, quite possibly allowing financial firms to offer things like proprietary risk management systems over the Internet. Wall Street and Technology, 14(3): 54–??, ???? 1996. CODEN WSTEE5. ISSN 1060-989X.

Anonymous:1996:ITJ

[Ano96-55] Anonymous. Internet tools:

The Java agenda. Information Week, ??(560):54–??, January 1996. CODEN INFWE4. ISSN 8750-6874.

Anonymous:1996:IJA

[Ano96-56] Anonymous. “Internetsprache”

Java — alles aus einer Hand. Elektronik, 45(7):12–??, ???? 1996. CODEN EKRKAR. ISSN 0013-5658.

Anonymous:1996:MP

[Ano96-57] Anonymous. Is Microsoft perking? Information Week, 584: 104–??, June 17, 1996. CODEN INFWE4. ISSN 8750-6874.

Anonymous:1996:J

[Ano96-58] Anonymous. Java. Ohmsha, ????, August 1996. ISBN 4-274-06164-7. LCCN ???? 1,900

                                        yen. URL http://www.ohmsha.
                                        co.jp/data/books/contents/
                                        4-274-06164-7.htm.

Anonymous:1996:JNT

[Ano96-59] Anonymous. Java: a new tool for engineering — this hot new Internet programming language, which runs on any platform, could be the breakthrough engineers are looking for to communicate product data within their companies and with suppliers and customers. Mechanical Engineering: the journal of the American Society of Mechanical Engineers, 118(4): 68–??, ???? 1996. CODEN MEENAH. ISSN 0025-6501.

Anonymous:1996:JTEa

[Ano96-60] Anonymous. Java: a timely example of distributed object technology. Client/Server Economics Letter, 3(6):7–8, June 1996. CODEN CSELF4. ISSN 1074-3138.

Anonymous:1996:JASa

[Ano96-61] Anonymous. Java and ActiveX: sorting out the truth [web programming languages]. Intranet and Networking Strategies Report, 4(9):5–7, September 1996. ISSN 1069-126X.

Anonymous:1996:JAO

[Ano96-62] Anonymous. Java API overview.

Sun Developer NEWS, 1(1):10, 11, Fall 1996.

Anonymous:1996:JAW

[Ano96-63] Anonymous. Java Applet: Window Toolkit & Utility / Java Language. Specialized Systems Consultants, P.O. Box 55549, Seattle, WA 98155, 1996. ISBN 0-916151-95-6 (set), 0-916151-93-X (vol. 1), 0-916151-94-8 (vol. 2). ???? pp. LCCN

????  URL http://www.
ssc.com/ssc/about.html. 2-

volume set.

Anonymous:1996:JAI

[Ano96-64] Anonymous. Java Applets for Intranets. New Riders Publishing, Carmel, IN, USA, 1996. ISBN 1-56205-627-1. ???? pp. LCCN ???? ???? URL http:/

/www.mcp.com/newriders/.

Anonymous:1996:JPM

[Ano96-65] Anonymous. Java-based pattern matching. SunExpert Magazine, 7(8):74, August 1996. ISSN 1053-9239. Describes VanillaSearch, a search class for Java developers from Thought, Inc..

Anonymous:1996:JBV

[Ano96-66] Anonymous. Java Beans vs.

ActiveX: open vs. proprietary. Computer Economics Report (International Edition), 18(12):

5–7, December 1996. ISSN 0739-0874.

Anonymous:1996:JBC

[Ano96-67] Anonymous. Java Black Coffee: Die große Java-Sammlung. Franzis-Verlag, Feldkirchen, Germany, 1996. ISBN 3-7723-8753-5. LCCN ???? DM 29,95; ATS 238,00; CHF 25,00.

                                        URL http://www.franzis-
                                        buch.de/katalog/blackcoffee/

. Includes CD-ROM.

Anonymous:1996:JB

[Ano96-68] Anonymous. Java bucks. Dr. Dobb’s Journal of Software Tools, 21(11):12, November 1996. CODEN DDJOEB. ISSN 1044-789X. Comments on the US$100,000,000 Kleiner, Perkins, Caulfield and Byers (KPCB) venture capital Java development fund.

Anonymous:1996:JBA

[Ano96-69] Anonymous. Java business applications arrive. BYTE Magazine, 21(8):30–??, August 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:JCSa

[Ano96-70] Anonymous. Java: CHIP Schulung. BHV, ??, Germany, October 1996. ISBN 3-89360-152-X. ???? pp. LCCN ???? 20

                                        DM. URL http://www.ctn-
                                        service.com/bhv.  Includes

CD-ROM.

Anonymous:1996:JCBb

[Ano96-71] Anonymous. Java chips boost applet speed. BYTE Magazine, 21(4):25–??, April 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:JCSb

[Ano96-72] Anonymous. Java client/server apps — D and B Software announces Java-based Smart-Stream applets for corporate Intranet users. Computerworld, 30 (14):45–??, ???? 1996. CODEN CMPWAB. ISSN 0010-4841.

Anonymous:1996:JCJ

[Ano96-73] Anonymous. Java competitor — “Java, schmava,” is the message Microsoft sent when it unveiled Active X, new object technology for the Internet, at a developer conference last week. Computerworld, 30(12): 14–??, ???? 1996. CODEN CM-PWAB. ISSN 0010-4841.

Anonymous:1996:JCBa

[Ano96-74] Anonymous. Java competitor — the Bell Labs team that developed Unix is now working on its own Internet language to compete with Sun’s Java. Computerworld, 30(8):6–??, ???? 1996. CODEN CMPWAB. ISSN 0010-4841. URL http://www.

computerworld.com/.

Anonymous:1996:JC

[Ano96-75] Anonymous. Java complete!

Datamation, 42(????):28–61, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1996:JCG

[Ano96-76] Anonymous. Java components give table functionality. Web-Server Magazine: For Managers of World Wide Web Sites, 1(1):74, 76, March/June 1996. ISSN 1087-4232. URL http:/

                                        /www.cpg.com.

Anonymous:1996:JCSc

[Ano96-77] Anonymous. Java cuts software distribution and installation costs for R. R. Donnelley’s customers. I/S Analyzer, 35(9): 12–16, September 1996. CO-DEN ISANEL. ISSN 0896-3231.

Anonymous:1996:JCT

[Ano96-78] Anonymous. Java de caffeinated? — three members of the original Java team leave. Computerworld, 30(6): 32–??, 1996. CODEN CMP-WAB. ISSN 0010-4841.

Anonymous:1996:JDT

[Ano96-79] Anonymous. Java development tools debug. WebServer Magazine: For Managers of World Wide Web Sites, 1(1):70, March/June 1996. ISSN 1087-

                                        4232. URL http://www.cpg.
                                        com.

Anonymous:1996:JR

[Ano96-80] Anonymous. Java for real-time. Real-Time Systems, 11(2):197– 205, September 1996. CODEN RESYE9. ISSN 0922-6443.

Anonymous:1996:JF

[Ano96-81] Anonymous. Java forms.

BYTE Magazine, 21(11):38–??, November 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1996:JHP

[Ano96-82] Anonymous. Java-fueled hub perks up managers. Computerworld, 30(27):90, July 1, 1996. CODEN CMPWAB. ISSN 0010-4841.

Anonymous:1996:JGP

[Ano96-83] Anonymous. Java Game Programming Starter Kit C/Dos/Us. Howard W. Sams, Indianapolis, IN 46268, USA, December 1996. ISBN 1-57521-201-3. ???? pp. LCCN ???? US$79.99.

Anonymous:1996:JGE

[Ano96-84] Anonymous. Java gets even hotter. Information Week, ??(562): 26–??, January 1996. CODEN INFWE4. ISSN 8750-6874.

Anonymous:1996:JGT

[Ano96-85] Anonymous. Java gets on track.

Information Week, 592:14–??, August 12, 1996. CODEN IN-FWE4. ISSN 8750-6874.

Anonymous:1996:JGS

[Ano96-86] Anonymous. Java gets serious with RAD. Software magazine, 16(10):77–??, October 1, 1996. CODEN SMWMEQ. ISSN 0279-9782 (or 0897-8085??).

Anonymous:1996:JGH

[Ano96-87] Anonymous. Java getting heat from MS ActiveX. Computer Design, 35(10):22–??, September 1, 1996. CODEN CMP-DAM. ISSN 0010-4566.

Anonymous:1996:JGM

[Ano96-88] Anonymous. Java goes mainstream. Information Week, ??(567):21–??, February 1996. CODEN INFWE4. ISSN 8750-6874.

Anonymous:1996:JH

[Ano96-89] Anonymous. Java Hacker. Franzis-Verlag, Feldkirchen, Germany, 1996. ISBN 3-7723-9753-0. ???? pp. LCCN ???? 49 DM.

Anonymous:1996:JHM

[Ano96-90] Anonymous. Java helps MKS solve its cross-platform portability problems. I/S Analyzer, 35(9):2–7, September 1996. CO-DEN ISANEL. ISSN 0896-3231.

Anonymous:1996:JIS

[Ano96-91] Anonymous. Java IDE set for release. Information Week, ??(568):83–??, February 1996. CODEN INFWE4. ISSN 8750-6874.

Anonymous:1996:JIC

[Ano96-92] Anonymous. Java Interactive Course. Waite Group Press, Corte Madera, CA, USA, 1996. ISBN ???? ???? pp. LCCN

???? URL http://www.waite.
com/waite/.

Anonymous:1996:JIA

[Ano96-93] Anonymous. Java internationalization architecture — continued. In Unicode Consortium [Uni96b], pages A4–?? ISBN ???? LCCN ???? URL http:

//www.unicode.org/unicode/
iuc9conf/index.html;   http:
//www.unicode.org/unicode/
uni2book/u2ord.html.

Anonymous:1996:JJEa

[Ano96-94] Anonymous. Java, Java everywhere. WebServer Magazine: For Managers of World Wide Web Sites, 1(1):10–11, March/ June 1996. ISSN 1087-4232.

URL http://www.cpg.com.

Anonymous:1996:JJEb

[Ano96-95] Anonymous. Java, Java everywhere. SunExpert Magazine, 7 (6):8, 10, June 1996. ISSN 1053-9239.

Anonymous:1996:JJDb

[Ano96-96] Anonymous. Java juices databases. WebServer Magazine: For Managers of World Wide Web Sites, 1(1):16, March/June 1996. ISSN 1087-4232. URL

http://www.cpg.com.

Anonymous:1996:JKM

[Ano96-97] Anonymous. Java killer? — Microsoft plans to build Internet support directly into future versions of Windows 95 and NT, as part of its defense against Sun’s Java development system. Computerworld, 30(8):46–??, ???? 1996. CODEN CMPWAB.

ISSN 0010-4841. URL http:/

                                        /www.computerworld.com/.

Anonymous:1996:JMW

[Ano96-98] Anonymous. JAVA: Microsoft won’t fight it. Information Week, ??(577):14–??, April 1996. CODEN INFWE4. ISSN 8750-6874.

Anonymous:1996:JNB

[Ano96-99] Anonymous. Java now being served “100% pure”. Electronic News, 42(2147):48, December 16, 1996. CODEN ELNEAU. ISSN 0013-4937, 1061-6624.

Anonymous:1996:JPE

[Ano96-100] Anonymous. Java perks up embedded systems conference. Electronic Design, 44(18):153– ??, September 3, 1996. CODEN ELODAW. ISSN 0013-4872.

Anonymous:1996:JPJ

[Ano96-101] Anonymous. The Java platform and JavaOS: Enabling networkbased applications. Sun Developer NEWS, 1(1):2, Fall 1996.

Anonymous:1996:JSJ

[Ano96-102] Anonymous. Java, schmava — John Gantz isn’t impressed by the Java hoopla. it’s just a programming language, he notes, and there’s no guarantee applets will play. Computerworld, 30 (13):37–??, ???? 1996. CODEN CMPWAB. ISSN 0010-4841.

Anonymous:1996:JSKc

[Ano96-103] Anonymous. Java Starter Kit, Deluxe Edition. Sams.net, Indianapolis, IN 46268, USA, August 14, 1996. ISBN 1-57521-190-4. ???? pp. LCCN ???? US$69.99.

Anonymous:1996:JSKa

[Ano96-104] Anonymous. Java Starter Kit for Macintosh. Hayden Books, 4300 West 62nd Street, Indianapolis, IN 46268, USA, 1996. ISBN 1-56830-293-2. ???? pp. LCCN ???? US$79.95. URL http://

merchant.superlibrary.com:
8000/catalog/hg/PRODUCT/PAGE/
15683/bud/1568302932.html.

Includes CD-ROM.

Anonymous:1996:JSKb

[Ano96-105] Anonymous. The Java starter kit. Version 1.0 for Windows. Sams.net, Indianapolis, IN 46268, USA, October 1996. ISBN 1-57521-077-0, 1-57521-097-5. ???? pp. LCCN ???? US$80. 1 CD-ROM and 1 book.

Anonymous:1996:JSI

[Ano96-106] Anonymous. Java surges onto the Internet. Library Systems, 16(1):3–??, 1996. ISSN 0277-0288.

Anonymous:1996:JTA

[Ano96-107] Anonymous. Java takes aim at Web applets. Graphic arts monthly: the magazine of the printing industry, 68(6):84– ??, June 1, 1996. CODEN GAMOE4. ISSN 1047-9325.

Anonymous:1996:JTR

[Ano96-108] Anonymous. Java teams up with RTOSs. Electronic Design, 44(7):153–??, ???? 1996. CO-DEN ELODAW. ISSN 0013-4872.

Anonymous:1996:JM

[Ano96-109] Anonymous. Java-to-RDBMS mapping. WebServer Magazine: For Managers of World Wide Web Sites, 3(4):50, November/ December 1996. ISSN 1087-4232. Discusses Thought Inc.’s CocoBase products for support of use of Java with relational database management systems.

Anonymous:1996:JTS

[Ano96-110] Anonymous. Java tools are still brewing. Information Week, 594:1A–??, August 26, 1996. CODEN INFWE4. ISSN 8750-6874.

Anonymous:1996:JVA

[Ano96-111] Anonymous. Java vs. ActiveX: a comparison. Sun Developer NEWS, 1(1):7, 11, Fall 1996.

Anonymous:1996:JWD

[Ano96-112] Anonymous. Java will dwarf Windows, says Sun’s Scott Mc-Nealy. Oracle Integrator, 7(5): 1, 19–21, September/October 1996. CODEN ORINEI. ISSN 1068-5537.

Anonymous:1996:JSU

[Ano96-113] Anonymous. JavaBeans specs unveiled. Information Week, 602:22–??, October 21, 1996. CODEN INFWE4. ISSN 8750-6874.

Anonymous:1996:JASb

[Ano96-114] Anonymous. JavaManagement API specification released. Sun Developer NEWS, 1(1):12, Fall 1996.

Anonymous:1996:JSW

[Ano96-115] Anonymous. JavaOS scores wins with Wyse, Taiwan. Sun-Expert Magazine, 7(8):6, August 1996. ISSN 1053-9239.

Anonymous:1996:JA

[Ano96-116] Anonymous. Javastation arrives. IEEE Micro, 16(6):80, November/December 1996. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1996:JTJ

[Ano96-117] Anonymous. JClass Table, JClass Table Applet combine XRT/table, Java technology. Silicon Graphics World, 6(8):25, August 1996. ISSN 1057-7041.

Anonymous:1996:JTEb

[Ano96-118] Anonymous. Joe takes enterprise to Web. Information Week, 594:11A–??, August 26, 1996. CODEN INFWE4. ISSN 8750-6874.

Anonymous:1996:KJP

[Ano96-119] Anonymous. Keeping Java perking. Business Week, ?? (3459):73, January 1996. CO-DEN BUWEA3. ISSN 0007-7135.

Anonymous:1996:MJT

[Ano96-120] Anonymous. Le Meilleur de Java: Trucs, Astuces et Secrets. Edition Micro Application, Paris, France, 1996. ISBN 2-7429-0774-2. 509 pp. LCCN ???? 149 FF. URL http:

//www.microapp.com/. French

translation by Hassina Abbasbhay. Based on the original edition by Data Becker in Dusseldorf. Includes CD-ROM.

Anonymous:1996:L

[Ano96-121] Anonymous. Letters. Forth Dimensions, 18(2):5–??, July 1, 1996. CODEN FODMD5. ISSN 0884-0822.

Anonymous:1996:MGJ

[Ano96-122] Anonymous. The Mac gets a Java toolkit. Information Week, ??(562):64–??, January 1996. CODEN INFWE4. ISSN 8750-6874.

Anonymous:1996:MWJ

[Ano96-123] Anonymous. Managing via the Web — Java applets let administrators manage networks over the Web. LAN times, 13(5):18– ??, ???? 1996. ISSN 1040-5917.

Anonymous:1996:MBJ

[Ano96-124] Anonymous. Microsoft brews Java with Visual J++ app dev tool. Datamation, 42(????):23– ??, ???? 1996. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1996:MWC

[Ano96-125] Anonymous. Middleware for the Web — Connection for Java will help developers link corporate resources to Web applications. LAN times, 13(5):45–??, ???? 1996. ISSN 1040-5917.

Anonymous:1996:MPJ

[Ano96-126] Anonymous. Mitsubishi ports Java to 32-bit RISC. Computer Design, 35(8):16–??, July 1, 1996. CODEN CMPDAM. ISSN 0010-4566.

Anonymous:1996:MVB

[Ano96-127] Anonymous. Morphing Visual Basic. Information Week, 597: 134–??, September 16, 1996. CODEN INFWE4. ISSN 8750-6874.

Anonymous:1996:NUJ

[Ano96-128] Anonymous. The NCSA uses Java to enable collaborative analysis over the Internet. I/S Analyzer, 35(9):8–11, September 1996. CODEN ISANEL. ISSN 0896-3231.

Anonymous:1996:NUS

[Ano96-129] Anonymous. Netscape unplugs Shockwave. Computerworld, 30 (37):68, September 9, 1996. CO-DEN CMPWAB. ISSN 0010-4841. Netscape has agreed to embed Macromedia’s Fireworks device, a Java-language-based animation API into Netscape Navigator.

Anonymous:1996:NGJ

[Ano96-130] Anonymous. Netware to get Java functionality. Information Week, 590:20–??, July 29, 1996. CODEN INFWE4. ISSN 8750-6874.

Anonymous:1996:NAJ

[Ano96-131] Anonymous. New APIs from Java. Electronic Design, 42 (2119):25, June 3, 1996. CO-DEN ELODAW. ISSN 0013-4872.

Anonymous:1996:NBT

[Ano96-132] Anonymous. A new box of tools for a networked world — Sun’s Java. Data communications, 25 (1):117–??, 1996. CODEN DA-CODM. ISSN 0363-6399.

Anonymous:1996:NNN

[Ano96-133] Anonymous. New Netscape Navigator does 3-D, better Java. N+I = bridges + routers + hub. OpenDoc gets a leg up; updates, Cyberdog groom technology. Excel does. MacWEEK, 10(13):14–??, ???? 1996. CO-DEN MWEEEI. ISSN 0892-8118.

Anonymous:1996:NOC

[Ano96-134] Anonymous. New on-line connection for Java developers. Sun Developer NEWS, 1(1):9, Fall 1996.

Anonymous:1996:NPO

[Ano96-135] Anonymous. New products: ObjectSpace Java Generic Library; ObjectSpace Web Toolkit; Innovative Software InvisibleWeb & Offline Proxy Server; Datacomm Internet and Intranet/Web Server with Cyrix 166MHz chip; ARDI Executor 2; TowerEiffel Release 2.0; Debian Linux 1.1; Open Systems Management COS/Print; Amtec Engineering Tecplot 7.0; Thought, Inc. VanillaSearch; Dimension X Liquid Reality Developer’s Kit; X Inside, Inc.

Accelerated OpenGL Solution for Linux; Spire Technologies Tactician Plus. Linux Journal, 30:??, October 1996. CO-DEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (elec-

tronic). URL http://www.
linuxjournal.com/issue30/
newprod.html.

Anonymous:1996:NPU

[Ano96-136] Anonymous. New products:

UML and Java go Platinum. Computer, 29(6):104–106, June 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:NPV

[Ano96-137] Anonymous. New products: Visual SlickEdit Version 2.0; C++ Version of Andrew User Interface; JClass LiveTable Pro & LiveTable Applet; FairCom Web Server; New Release of ObjectSpace C++ Class Libraries; CocoBase Java DataBase Access System; InTerNet-LINK; Linux Advanced GUI Toolkit; Introduction to Programming Java Applets; SOLID Server 2.1 for Linux; COS/Relay; Linux Serial I/O Solutions; CRiSP v5-Multiplatform Professional Editor; Cross-Platform Web Agent-Trudger; NExS version 1.2 for Linux; Networking Switch for Access 8 PCs; Acu4GL for ODBC; InfoMagic Workgroup Server. Linux Journal, 32:??, December 1996. CODEN LI-JOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Anonymous:1996:Nc

[Ano96-138] Anonymous. News. MacUser, 12(9):23–??, September 1, 1996. CODEN MCUSEY. ISSN 0884-0997.

Anonymous:1996:NB

[Ano96-139] Anonymous. News briefs. Information Week, 586:40–??, July 1, 1996. CODEN INFWE4. ISSN 8750-6874.

Anonymous:1996:NTW

[Ano96-140] Anonymous. The next take on Windows. Information Week, 588:18A–??, July 15, 1996. CO-DEN INFWE4. ISSN 8750-6874.

Anonymous:1996:EJ

[Ano96-141] Anonymous. Not exactly justin-time. Information Week, 597: 136–??, September 16, 1996. CODEN INFWE4. ISSN 8750-6874.

Anonymous:1996:OEEb

[Ano96-142] Anonymous, editor. Object Expo Europe — Java Expo Conference Proceedings. September 23–27, 1996, London, UK. SIGS Conferences, Newdigate, UK, September 1996. ISBN ???? LCCN ????

Anonymous:1996:OJE

[Ano96-143] Anonymous. Object, Java Expo opens. Computerworld, 30(32): 2, August 5, 1996. CODEN CM-PWAB. ISSN 0010-4841.

Anonymous:1996:OOO

[Ano96-144] Anonymous. Object orientation in operating systems. In Cabrera and Islam [CI96], page ?? CODEN PIWOF6. ISBN 0-8186-7692-2, 0-8186-7693-0, 0-8186-7694-9 (microfiche). ISSN 1063-5351. LCCN QA 76.76 O63 I59 1996.

Anonymous:1996:ODJ

[Ano96-145] Anonymous. ObjectSpace designs Java training course. Silicon Graphics World, 6(8):25, August 1996. ISSN 1057-7041.

Anonymous:1996:ORJ

[Ano96-146] Anonymous. ObjectSpace releases Java generic library for free commercial use. C/C++ Users Journal, 14(10):96, October 1996. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1996:OPO

[Ano96-147] Anonymous, editor. Objekt-Orientiertes Programmieren — OOP ’96 Conference Proceedings, February 12–16, 1996, Munich, Germany. SIGS Books, 71 W 23rd St., New York, NY 10010, USA, 1996. ISBN ???? LCCN ????

Anonymous:1996:OEEa

[Ano96-148] Anonymous. OCLC to enhance electronics journals online (EJO) with Hot Java on the World Wide Web. Information Intelligence, Online Libraries, and Microcomputers, 14 (2):9–??, February 1996. CO-DEN IIOMEI. ISSN 0737-7770.

Anonymous:1996:Wa

[Ano96-149] Anonymous. Of and for the Web. Information Week, 594:

35–??, August 26, 1996. CO-DEN INFWE4. ISSN 8750-6874.

Anonymous:1996:Na

[Ano96-150] Anonymous. On the Net. LAN: the network solutions magazine, 11(8):43–??, August 1, 1996. CODEN LANNER. ISSN 0898-0012.

Anonymous:1996:ODL

[Ano96-151] Anonymous. Online — despite a lack of standards, 3-D flourishes online; financial shareware helps you stick to a budget, save for retirement, and pay debts; graphics BBSs offer stunning images; shopper hosts an online discussion of Java. Computer Shopper, pages 615–??, February 1996. ISSN 0886-0556.

Anonymous:1996:OTT

[Ano96-152] Anonymous. Online: Technical tips. LAN: the network solutions magazine, 5(3):75–??, August 1, 1996. CODEN LAN-NER. ISSN 1038-359X (or 1069-5621??).

Anonymous:1996:OHP

[Ano96-153] Anonymous. Outlook — hardware processors poise to widen Java appeal. memory makes transition from voice recording to data storage. reels put new spin on assembling pin headers. memory architecture offers SRAM performance at DRAM price. increased integration to slow growth of RF transistor market. Electronic Products, 38 (11):17–??, ???? 1996. CODEN ELPOA2. ISSN 0013-4953.

Anonymous:1996:PPF

[Ano96-154] Anonymous, editor. PAAM 96. Proceedings of the First International Conference on the Practical Application of Intelligent Agents and Multi-Agent Technology, London, UK, 22–24 April, 1996. Practical Application Company, Blackpool, UK, 1996. ISBN ???? LCCN ????

Anonymous:1996:PRI

[Ano96-155] Anonymous. Pencom releases introductory Java book. Silicon Graphics World, 6(11): 7, November 1996. ISSN 1057-7041. Brief review of [AMPM96].

Anonymous:1996:PJa

[Ano96-156] Anonymous. Perking up Java.

Information Week, ??(564):70– ??, January 1996. CODEN IN-FWE4. ISSN 8750-6874.

Anonymous:1996:PJd

[Ano96-157] Anonymous. Poison Java. PC/ Computing, 9(10):332–??, October 1, 1996. CODEN PCMPEI. ISSN 0899-1847.

Anonymous:1996:PCO

[Ano96-158] Anonymous, editor. Proceedings of 2nd Conference on Object-Oriented Technologies and Systems (COOTS), Toronto, Ontario, Canada, 17–21 June 1996. Yggdrasil Computing, Inc., Berkeley, CA, USA, 1996. ISBN ???? LCCN ????

Anonymous:1996:POW

[Ano96-159] Anonymous, editor. Proceedings of Object World UK ’96. Object World, Hounslow, UK, 1996. ISBN ???? LCCN ????

Anonymous:1996:PTF

[Ano96-160] Anonymous, editor. Proceedings of the Twenty-First Annual SAS Users Group International Conference, SUGI 21, Chicago, IL, USA, 10–13 March 1996. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1996. ISBN ???? LCCN ???? Two volumes.

Anonymous:1996:PRJ

[Ano96-161] Anonymous. Product review: The Java Reference Package from SSC. Linux Journal, 30:??, October 1996. CO-DEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (elec-

tronic). URL http://www.
linuxjournal.com/issue30/
javarev.html.

Anonymous:1996:PJe

[Ano96-162] Anonymous. Programming with Java. UNIX review, 14(5):??, May 1996. CODEN UNRED5. ISSN 0742-3136.

Anonymous:1996:PAJ

[Ano96-163] Anonymous. Protect against Java email attacks. WebServer Magazine: For Managers of World Wide Web Sites, 1(4): 48, November/December 1996. ISSN 1087-4232. Discusses Finjan Software Ltd.’s SurfinBoard Java email firewall product.

Anonymous:1996:P

[Ano96-164] Anonymous. Pulse. PC/ Computing, 9(8):38–??, August 1, 1996. CODEN PCMPEI. ISSN 0899-1847.

Anonymous:1996:PJB

[Ano96-165] Anonymous. Pulse — Java brewing, the first cable modems, radar. PC/Computing, 9(3):48– ??, ???? 1996. CODEN PCM-PEI. ISSN 0899-1847.

Anonymous:1996:PIW

[Ano96-166] Anonymous. Putting the Internet to work: Is Java online marvel or menace? Communications news (Geneva, IL), 33(9): 44–??, September 1, 1996. CO-DEN CMUNA9. ISSN 0010-3632.

Anonymous:1996:RAR

[Ano96-167] Anonymous. Rational announced Rational Rose for Java. C/C++ Users Journal, 14(9): 94, September 1996. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1996:RC

[Ano96-168] Anonymous. The return of COBOL. Software Economics Letter, 5(6):4–7, June 1996. CODEN SECLE3. ISSN 1065-6146.

Anonymous:1996:SAP

[Ano96-169] Anonymous. Slim applet pickin’s. Information Week, 593:35–??, August 19, 1996. CODEN INFWE4. ISSN 8750-

Anonymous:1996:SJ

[Ano96-170] Anonymous. SNiFF+ for Java. SunExpert Magazine, 7(3):10, March 1996. ISSN 1053-9239.

Anonymous:1996:SES

[Ano96-171] Anonymous. Software: Embedded software poses challenges with growth, complexity; Java can shine in embedded apps — even real-time; object-oriented languages and tools proliferate. Computer Design, 35(10):91–??, September 1, 1996. CODEN CMPDAM. ISSN 0010-4566.

Anonymous:1996:SKR

[Ano96-172] Anonymous. Software kills rogue applets. RS/Magazine, 5(10):55, October 1996. ISSN 1088-0844, 1061-0030. Describes Surfinboard package from Finjan Software that monitors in-coming applets for hostile activity.

Anonymous:1996:SLJ

[Ano96-173] Anonymous. Sun launches Java-Soft. SunExpert Magazine, 7(3): 6, March 1996. ISSN 1053-9239.

Anonymous:1996:SMB

[Ano96-174] Anonymous. Sun Microsystems brews three flavors of JAVA. Computer Design, 35(5):34–??, ???? 1996. CODEN CMPDAM. ISSN 0010-4566.

Anonymous:1996:SSIa

[Ano96-175] Anonymous. Sun says it has Java processors. SunExpert Magazine, 7(4):12, April 1996. ISSN 1053-9239.

Anonymous:1996:SSIb

[Ano96-176] Anonymous. Sun says it has Java processors. WebServer Magazine: For Managers of World Wide Web Sites, 1(1): 16, 18, March/June 1996. ISSN 1087-4232. URL http://www.

cpg.com.

Anonymous:1996:SSM

[Ano96-177] Anonymous. Sun, SGI, and Macromedia to extend Java. Computer Graphics World, 19 (1):8, January 1996. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1996:SSB

[Ano96-178] Anonymous. Sun shoots for the big time. Information Week, ?? (576):14–??, April 1996. CO-DEN INFWE4. ISSN 8750-6874.

Anonymous:1996:STR

[Ano96-179] Anonymous. Sun turns red in Java flap. Computerworld, 30 (26):????, June 24, 1996. CO-DEN CMPWAB. ISSN 0010-4841.

Anonymous:1996:SUF

[Ano96-180] Anonymous. Sun unveils family of Java processors. Sun Observer, 10(5):4, May 1996. ISSN 1058-5400.

Anonymous:1996:SJA

[Ano96-181] Anonymous. Sun’s JavaBeans API lets users write reusable components. Silicon Graphics World, 6(12):14, December 1996. ISSN 1057-7041.

Anonymous:1996:SJW

[Ano96-182] Anonymous. SunSoft Java WorkShop 1.0 vs. Symantec Café 1.2. PC/Computing, 9(10): 108–??, October 1, 1996. CO-DEN PCMPEI. ISSN 0899-1847.

Anonymous:1996:SSH

[Ano96-183] Anonymous. SYSTEMS and SOFTWARE: Hot cup of Java: Keeping companies awake. Aerospace America, 34(4):12– ??, ???? 1996. CODEN AEAME2. ISSN 0740-722X.

Anonymous:1996:TPJb

[Ano96-184] Anonymous. Taking the plunge with Java. Irish Computer, 20 (6):18–21, August 1996. CO-DEN IRCODQ. ISSN 0332-0197.

Anonymous:1996:TYW

[Ano96-185] Anonymous. Teach Yourself WWW Database Programming with Java in 21 Days. Sams.net, Indianapolis, IN 46268, USA, 1996. ISBN ???? ???? pp. LCCN ???? US$39.95.

URL http://www.mcp.com/
samsnet/.

Anonymous:1996:TWJa

[Ano96-186] Anonymous. Teaching wavelets with Java on the information superhighway. IEEE Digital Signal Processing Workshop, pages 101–104, 1996. IEEE catalog number 96TH8225.

Anonymous:1996:TPJa

[Ano96-187] Anonymous. Tech page: Java jive. Electronic Musician, 12(3):

154–??, March 1996. ISSN 0884-4720.

Anonymous:1996:TSW

[Ano96-188] Anonymous. TECH SEC-TION — we look at the new Cyrix chip; give the new Java programming language a taste test; peek at the research of MIT’s Visual Language Workshop; and help you get full plugand-play capability from Win 95. Computer Shopper, pages 569–??, February 1996. ISSN 0886-0556.

Anonymous:1996:TWHa

[Ano96-189] Anonymous. TECHNOLOGY:

What’s so hot about Java? Time, 147(4):42–??, 1996. CO-DEN TYMEA9. ISSN 0040-781X.

Anonymous:1996:TWHb

[Ano96-190] Anonymous. TECHNOLOGY:

What’s so hot about Java? Time, 147(4):58–??, 1996. CO-DEN TYMEA9. ISSN 0040-781X.

Anonymous:1996:RID

[Ano96-191] Anonymous. To the rescuemiddleware for Intranet developers. Software Economics Letter, 5(8):4–6, August 1996. CO-DEN SECLE3. ISSN 1065-6146.

Anonymous:1996:TWJb

[Ano96-192] Anonymous. Towards Web/ Java-based high performance distributed computing — an evolving virtual machine. IEEE International Symposium on High Performance Distributed Computing, Proceedings, pages 308–317, 1996. CODEN PID-CFB. ISSN 1082-8907. IEEE catalog number 96TB100069.

Anonymous:1996:TNA

[Ano96-193] Anonymous. Two new approaches to online information [Krakatoa and intranets]. Online & CDROM review: the international journal of online & optical information systems, 20 (4):206–207, August 1996. CO-DEN ONCDEW. ISSN 0309-314X.

Anonymous:1996:UJ

[Ano96-194] Anonymous. Undocumented Java. Waite Group Press, Corte Madera, CA, USA, 1996. ISBN ???? ???? pp. LCCN ????

URL http://www.waite.com/
waite/.

Anonymous:1996:FAL

[Ano96-195] Anonymous. UP FRONT — Apple licenses brand name for pre-K toys. iterated tool hastens Web graphics. Adobe ships SiteMill. Mac beta of Java ships. Apple cancels dividend, rewards Amelio. MacWEEK, 10(7): 4–??, ???? 1996. CODEN MWEEEI. ISSN 0892-8118.

Anonymous:1996:FFD

[Ano96-196] Anonymous. UP FRONT — Fuji device prints color without inks or ribbons Epson scanner to face off against Visioneer. Java for Mac due in February — VST offers. MacWEEK, 10(4): 4–??, 1996. CODEN MWEEEI. ISSN 0892-8118.

Anonymous:1996:VDC

[Ano96-197] Anonymous. Visual development for C/C++, Java. SunExpert Magazine, 7(11):74, November 1996. ISSN 1053-9239.

Anonymous:1996:VTPa

[Ano96-198] Anonymous. VisualAge technology planned for Java. RS/ Magazine, 5(8):8, August 1996. ISSN 1088-0844, 1061-0030.

Anonymous:1996:VTPb

[Ano96-199] Anonymous. VisualAge technology planned for Java. SunExpert Magazine, 7(8):8, August 1996. ISSN 1053-9239.

Anonymous:1996:WJB

[Ano96-200] Anonymous. Weaknesses in Java bytecode verifier. Network Security, 1996(4):2, April 1996. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/S1353485896901737.

Anonymous:1996:WVJ

[Ano96-201] Anonymous. Weather visualizer: a Java tool for interactive learning. International Geoscience and Remote Sensing Symposium (IGARSS), 3: 1498–1500, 1996. CODEN IGRSE3. IEEE catalog number 96CB35875.

Anonymous:1996:WAU

[Ano96-202] Anonymous. Web authors use Cosmo site for HTML, Java and VRML sites. Silicon Graphics World, 6(11):25, November 1996. ISSN 1057-7041.

Anonymous:1996:Wb

[Ano96-203] Anonymous. WebBurst. MacUser, 12(11):69–??, November 1, 1996. CODEN MCUSEY. ISSN 0884-0997.

Anonymous:1996:WWV

[Ano96-204] Anonymous. Working the Web with a virtual lab and some Java. Science, 273(5275):591– ??, August 2, 1996. CO-DEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Anonymous:1996:WDS

[Ano96-205] Anonymous. World-wide distributed system using Java and the Internet. IEEE International Symposium on High Performance Distributed Computing, Proceedings, pages 11–18, 1996. CODEN PIDCFB. ISSN 1082-8907. IEEE catalog number 96TB100069.

Anonymous:1996:WWW

[Ano96-206] Anonymous. World Wide Web Java. Franzis-Verlag, Feldkirchen, Germany, 1996. ISBN 3-7723-8933-3. ???? pp. LCCN ???? 49 DM. Includes CD-ROM. Out of print.

Anonymous:1997:ASJ

[Ano97a] Anonymous. A&D solution for JavaStation. Object Magazine, 6 (11):77, January 1997. CODEN OBMAFO. ISSN 1055-3614.

Anonymous:1997:AAJ

[Ano97b] Anonymous. Adra announces Java-based PDM Web client. Silicon Graphics World, 7(5): 24, May 1997. ISSN 1057-7041.

URL http://www.adra.com/.

Anonymous:1997:ARJ

[Ano97c] Anonymous. Algor Roadmaps has Java-based human interface. Silicon Graphics World, 7(4):21, April 1997. ISSN 1057-7041.

Anonymous:1997:ARB

[Ano97d] Anonymous. Aonix releases binding covers for Sun’s JDK. SunServer, 11(10):23, October 1997. ISSN 1091-4986.

Anonymous:1997:AMS

[Ano97e] Anonymous. Applet Management System for secure Java implementation. Network Security, 1997(11):7, November 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/S1353485897900763.

Anonymous:1997:AIX

[Ano97f] Anonymous. ArborText includes XML, Java support in Adept Version 7. SunServer, 11 (12):21, December 1997. ISSN 1091-4986.

Anonymous:1997:BRDk

[Ano97g]  Anonymous.  Book review:

Database programming with JDBC and JAVA: By George Reese. O’Reilly, Sebastopol, CA. (1997). 224 pages. $29.95.

Computers and Mathematics with Applications, 34(10): 141, November 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0898122197902595.

Anonymous:1997:BRDi

[Ano97h]  Anonymous.  Book review:

Developing Java Beans: By Robert Englander. O’Reilly, Sebastopol, CA. (1997). 298 pages. $29.95. Computers and Mathematics with Applications, 34(10):140, November 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/S0898122197902571.

Anonymous:1997:BRJd

[Ano97i]  Anonymous.  Book review:

Java AWT reference: By John Zukowski. O’Reilly, Sebastopol, CA. (1997). 1045 pages. $39.95. Computers and Mathematics with Applications, 34(10):135, November 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/S0898122197901875.

Anonymous:1997:BRJc

[Ano97j] Anonymous. Book review: Java network programming: By Elliotte Rusty Harold. O’Reilly, Sebastopol, CA. (1997). 422 pages. $34.95. Computers and Mathematics with Applications, 33(9):147, May 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0898122197901413.

Anonymous:1997:BRJe

[Ano97k] Anonymous. Book review: Java Virtual Machine: By Jon Meyer and Troy Downing. O’Reilly, Sebastopol, CA. (1997). 426 pages. $32.95 (diskette included). Computers and Mathematics with Applications, 34 (10):135, November 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0898122197901899.

Anonymous:1997:BRMj

[Ano97l]  Anonymous.  Book review:

Modern compiler implementation in Java: Basic techniques: By Andrew W. Appel. Cambridge University Press, Cambridge, U.K. (1997). 398 pages. $64.95 (hardback), $29.95 (paper). Computers and Mathematics with Applications, 33 (7):139, April 1997. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0898122197846093.

Anonymous:1997:BRJb

[Ano97m] Anonymous. Book review:

The Java Series. Linux Journal, 40:??, August 1997. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http://

                                        www.linuxjournal.com/2061.
                                        html.

Anonymous:1997:BRJa

[Ano97n] Anonymous. Book reviews: Java Books. Linux Journal, 36:??, April 1997. CO-DEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (elec-

                                        tronic). URL http://www.
                                        linuxjournal.com/2038.html.

Anonymous:1997:BGJ

[Ano97o] Anonymous. Bookish on graphical Java. Object Magazine, 6 (11):77, January 1997. CODEN OBMAFO. ISSN 1055-3614.

Anonymous:1997:BJH

[Ano97p] Anonymous. Briefs: Java for HP Unix. Computerworld, 31 (11):51, March 17, 1997. CO-DEN CMPWAB. ISSN 0010-4841.

Anonymous:1997:BFJ

[Ano97q] Anonymous. Bug found in Java Virtual Machine. Network Security, 1997(3):2, March 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (elec-

                                        tronic). URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S1353485897830334.

Anonymous:1997:CBC

[Ano97r]  Anonymous.  C++ Builder

comes out of the closet, IBM does everything Java, a new vision for client/server development, Novell’s Developer Net 2000, plus Matlab 5.0: serious fun with graphs. .EXE: the software developers’ magazine, 11 (8):7–??, ???? 1997. CODEN EXEEE5. ISSN 0268-6872.

Anonymous:1997:CDO

[Ano97s] Anonymous. Central Data offers serial/parallel port API, driver for Java. Silicon Graphics World, 7(6):21, June 1997. ISSN 1057-7041.

Anonymous:1997:CSB

[Ano97t] Anonymous. Charles Schwab builds interactive, Web-based trading applications using ActiveX. I/S Analyzer, 36(1): 2–6, January 1997. CODEN ISANEL. ISSN 0896-3231.

Anonymous:1997:CJT

[Ano97u] Anonymous. Chicago’s Java trade winds. Object Magazine, 7(1):68, March 1997. CODEN OBMAFO. ISSN 1055-3614.

Anonymous:1997:CJJ

[Ano97v] Anonymous. Cisco joins Java security alliance. Network Security, 1997(8):5, August 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S1353485897832308.

Anonymous:1997:CDP

[Ano97w] Anonymous, editor. Competitive deployment of product data technology: European conference — April 1997, Antibes, France. Quality Marketing Services, ????, 1997. ISBN 1-901782-01-8. LCCN ????

Anonymous:1997:CAC

[Ano97x] Anonymous. Converting Ada 95 code to Java. Object Magazine, 7(1):68, March 1997. CODEN OBMAFO. ISSN 1055-3614. URL http://www.

                                        intermetrics.com/.

Anonymous:1997:CAJ

[Ano97y] Anonymous. CORBA adds Java capabilities. /AIXtra: IBM’s Magazine for AIX Professionals, 7(2):22, March/April 1997.

Anonymous:1997:CSJ

[Ano97z] Anonymous. CORBAplus supports Java ORB. SunExpert Magazine, 8(4):104, April 1997. ISSN 1053-9239.

Anonymous:1997:CJD

[Ano97-27] Anonymous. Create Java database apps with JetAssist. WebServer Magazine: For Managers of World Wide Web Sites, 2(2):49, March 1997. ISSN 1087-4232.

Anonymous:1997:CENa

[Ano97-28] Anonymous. Cutting edge: News analysis — JavaMan. ActiveXecutables. Are they friend or foe? AS/400 shops tackle Year 2000 problem. Datamation, 43(1):12–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1997:DSC

[Ano97-29] Anonymous. DataBeam ships conference server with Java. Cisco World: The Independent Journal for Internetworking Professionals, 3(8):27, August 1997. ISSN 1081-3187.

Anonymous:1997:DRD

[Ano97-30] Anonymous. Defining reusable, distributable information objects using or, how SGML can do for databases what JAVA has done for user interfaces. In SGML Europe ’97 proceedings 11–16 May 1997, Princess Sofia Intercontinental, Barcelona, Spain [Ano97-139], pages 307–310. ISBN ???? LCCN ???? One CD-ROM.

Anonymous:1997:GJBa

[Ano97-31] Anonymous. Die große Java Bibliothek. Carl Hanser, München, Germany, second edition, 1997. ISBN 3-446-18910-6. ???? pp. LCCN ???? 69 DM.

URL http://www.hanser.de/
computer/buecher/bi19019.

htm. Includes CD-ROM.

Anonymous:1997:GJBb

[Ano97-32] Anonymous. Die große Java Bibliothek. Carl Hanser, München, Germany, second edition, 1997. ISBN 3-446-19019-8. ???? pp. LCCN ???? 50 DM.

URL http://www.hanser.de/
computer/buecher/bi19019.

htm. Includes CD-ROM.

Anonymous:1997:DWD

[Ano97-33] Anonymous. Display Windows docs as Java objects. Web-Server Magazine: For Managers of World Wide Web Sites, 2(1):53, February 1997. ISSN 1087-4232.

Anonymous:1997:DJO

[Ano97-34] Anonymous. Distribute Java ojbects across the net. SunExpert Magazine, 8(4):108, April 1997. ISSN 1053-9239. Describes the Black & White Software Web/Enable product, which uses Java and the CORBA Internet Inter-ORB Protocol (IIOP).

Anonymous:1997:ELM

[Ano97-35] Anonymous. Editing large multi Java apps. Object Magazine, 6 (12):83, February 1997. CO-DEN OBMAFO. ISSN 1055-3614.

Anonymous:1997:EEE

[Ano97-36] Anonymous. Empress enhances embedded RDBMS with Web, Java. /AIXtra: IBM’s Magazine for AIX Professionals, 7 (4):13, July/August 1997.

Anonymous:1997:ETJ

[Ano97-37] Anonymous. Emulation through Java applets. /AIXtra: IBM’s Magazine for AIX Professionals, 7(3):20–21, May/June 1997.

Anonymous:1997:EAJ

[Ano97-38] Anonymous. Encryption and authentication for Java. Network Security, 1997(6):6, June 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/S1353485897895256.

Anonymous:1997:FJP

[Ano97-39] Anonymous. A free Java platform. BYTE Magazine, 22(7): 158, July 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:HJT

[Ano97-40] Anonymous. Hands On Java Training. Microsoft Press, Bellevue, WA, USA, November 1997. ISBN 1-57231-707-8. ???? pp. LCCN ???? ????

Anonymous:1997:IOF

[Ano97-41] Anonymous. IBM offers free management software. SunExpert Magazine, 8(4):10, April 1997. ISSN 1053-9239. Describes Java-based Network Printer Manager (NPM).

Anonymous:1997:IAJ

[Ano97-42] Anonymous. ICS adds Java support. UNIX Developer, 1(1):93, January/February 1997. ISSN 1090-2279.

Anonymous:1997:ISB

[Ano97-43] Anonymous. ILOG shipping business rules engine for Java. HP Chronicle, 14(7):18, June 1997. ISSN 0892-2829.

Anonymous:1997:ISD

[Ano97-44] Anonymous. ILOG ships 2-D Java graphics library. Sun-Server, 11(12):21, December 1997. ISSN 1091-4986.

Anonymous:1997:IAX

[Ano97-45] Anonymous. Imperial announces X-Designer: Java Edition. C/C++ Users Journal, 15 (2):88, February 1997. CODEN CCUJEX. ISSN 1075-2838. X-Designer is a GUI builder for Motif with Java, X/Motif, and Windows code generators.

Anonymous:1997:ISS

[Ano97-46] Anonymous. Imperial Software ships Visaj Java development tool. SunServer, 11(12):19, December 1997. ISSN 1091-4986.

Anonymous:1997:ISJ

[Ano97-47] Anonymous. Infoscape ships Java application development software. HP Chronicle, 14(8): 19, July 1997. ISSN 0892-2829.

Anonymous:1997:IJV

[Ano97-48] Anonymous. Inside the Java Virtual Machine. UNIX review, 15(1):31, 32, 34–36, 38–39, January 1997. CODEN UNRED5. ISSN 0742-3136. Adapted, with permission, from [LY97b].

Anonymous:1997:IJJ

[Ano97-49] Anonymous. Instant Java: Jewel Case C/Dos/Ww. Que Corporation, Indianapolis, IN, USA, March 1997. ISBN 0-7897-1008-0. ???? pp. LCCN ???? US$12.99.

Anonymous:1997:ISP

[Ano97-50] Anonymous. InterNetivity ships 100% Pure Java data analysis tool. SunServer, 11(11):20, November 1997. ISSN 1091-4986.

Anonymous:1997:IVJ

[Ano97-51] Anonymous. Interview: a voice for Java developers. BYTE Magazine, 22(12):44, December 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:IWJ

[Ano97-52] Anonymous. Interview: Why Java won’t repeat the mistakes of Unix. Scott McNealy, chairman and CEO of Sun Microsystems, talks about how Java will change computing and the computer industry. BYTE Magazine, 22(1):40, January 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic). Some McNealy quotes: “And I’m urging every CEO to make sure it [Windows] runs really slowly, so people stop using that stuff. The right way is to ban PowerPoint from your company.” “I banned PowerPoint from my company and we have had the best two quarters we’ve ever had… ” “I want to give everybody plastic Mylar sheets and all the pens they need to scribble on them. And I said use what I call the [Sun cofounder] Bill Joy font. You can see where he licked his thumb and erases; it’s so much faster.”.

Anonymous:1997:IPJ

[Ano97-53] Anonymous. IntRprise provides Java/Internet solutions.

SunServer, 11(10):23, October 1997. ISSN 1091-4986.

Anonymous:1997:INI

[Ano97-54] Anonymous, editor. Issues for networked interpersonal communicators: Colloquium — May 1997, London, number 139 in COLLOQUIUM DIGEST-IEE 1997. IEE, London, UK, 1997. ISSN 0963-3308. LCCN ????

Anonymous:1997:IOI

[Ano97-55] Anonymous. IST offers integrated X, Java, and Windows toolset. UNIX Developer, 1 (1):94, January/February 1997. ISSN 1090-2279.

Anonymous:1997:IMM

[Ano97-56] Anonymous. IT managers may resist Java. UNIX Developer, 1 (1):13, January/February 1997. ISSN 1090-2279.

Anonymous:1997:Jc

[Ano97-57] Anonymous. Java. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, October 1997. ISBN 4-89471-001-3. LCCN ???? 5,200 yen. URL http:

                                        //www.mmjp.or.jp/prentice/
                                        wa_int21-j.html.

Anonymous:1997:Je

[Ano97-58] Anonymous. Java. Softbank Publishing, ??, Japan, 1997. ISBN 4-89052-892-X. ???? pp. LCCN ???? 2700 yen.

                                        URL http://www.softbank.
                                        co.jp/softbank/publishing/
                                        ;   http://www.softbank.co.

jp/softbank/publishing/books/
kikan/kikan2/002.htm.

Anonymous:1997:Jf

[Ano97-59] Anonymous. Java. ????, ??, Japan, 1997. ISBN ???? ???? pp. LCCN ???? URL

http://www.im.noda.sut.ac.
jp/java/java-bon.html.

Anonymous:1997:Jg

[Ano97-60] Anonymous.  Java.  Soft-

bank Publishing, ??, Japan, 1997. ISBN ???? ???? pp. LCCN ???? 2400 yen.

URL http://www.softbank.
co.jp/softbank/publishing/
;   http://www.softbank.co.
jp/softbank/publishing/books/
editer/nw/appletbo.html.

Anonymous:1997:Jh

[Ano97-61] Anonymous. Java. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1997. ISBN

4-88735-011-2.  ????  pp.
LCCN ????  ????  URL

http://puma.asahi-net.or.
jp/prentice/index_washo-j.
html;   http://puma.asahi-
net.or.jp/prentice/wa_net05-
j.html;   http://www.mmjp.
or.jp/prentice/wa_net05-j.
html.

Anonymous:1997:Ji

[Ano97-62] Anonymous. Java. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1997. ISBN 4-88735-000-7. ???? pp. LCCN

????  3000 yen.  URL

http://puma.asahi-net.or.
jp/prentice/index_washo-j.

                                        html;   http://puma.asahi-
                                        net.or.jp/prentice/wa_net13-
                                        j.html.

Anonymous:1997:Jk

[Ano97-63] Anonymous. Java. ASCII Corporation, Tokyo, Japan, 1997. ISBN 4-7561-1650-7. ???? pp. LCCN ???? ???? URL http://

                                        www.ascii.co.jp/;   http:/
                                        /www.ascii.co.jp/pb/book1/
                                        shinkan/detail/1322183.html.

Anonymous:1997:Jm

[Ano97-64] Anonymous. Java. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1997. ISBN 4-88735-039-2. ???? pp. LCCN

????  2400 yen.  URL

                                        http://puma.asahi-net.or.
                                        jp/prentice/index_washo-j.
                                        html;   http://puma.asahi-
                                        net.or.jp/prentice/wa_net16-
                                        j.html.

Anonymous:1997:Jz

[Ano97-65] Anonymous. Java. Mizuki, ??, Japan, 1997. ISBN ???? ???? pp. LCCN ???? 800 yen.

                                        URL http://www.mizuki.co.
                                        jp/;   http://www.mizuki.
                                        co.jp/books/155-3b.html.

Anonymous:1997:JJb

[Ano97-66] Anonymous. Java & JavaScript. Softbank Publishing, ??, Japan, 1997. ISBN ???? ???? pp. LCCN ???? URL

                                        http://www.softbank.co.jp/
                                        softbank/publishing/;   http:
                                        //www.softbank.co.jp/softbank/
                                        publishing/books/editer/pg/
                                        javahomepage/index.html.

Anonymous:1997:JV

[Ano97-67] Anonymous. Java + VRML.

Prentice-Hall, Englewood Cliffs, NJ 07632, USA, June 1997. ISBN 4-88735-058-9. ???? pp. LCCN ???? 4000

yen. URL http://www.mmjp.
or.jp/prentice/wa_int11-j.
html.

Anonymous:1997:JD

[Ano97-68] Anonymous. Java 1.0 Decubierto. Prentice Hall HispanoAmericana, Mexico, DF, Mexico / Nueva York, NY, USA, 1997. ISBN 84-89660-59-X (??Invalid checksum??). ???? pp. LCCN ???? 6995 Ptas. URL http://

prentice.com.mx/;   http:
//prentice.com.mx/libros/
javadesc.html. Includes CD-

ROM.

Anonymous:1997:JAT

[Ano97-69] Anonymous. Java agent tools help move business apps. Sun-Server, 11(12):22, December 1997. ISSN 1091-4986.

Anonymous:1997:JAR

[Ano97-70] Anonymous. Java applets run from a ‘cage’. Network Security, 1997(10):5–6, October 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/S1353485897901033.

Anonymous:1997:JAS

[Ano97-71] Anonymous. Java applets that steal computer cycles. Network Security, 1997(5):3, May 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (elec-

                                        tronic). URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S1353485897856905.

Anonymous:1997:JBT

[Ano97-72] Anonymous. Java-based tools provide direct access to CICS. SunServer, 11(11):17, November 1997. ISSN 1091-4986.

Anonymous:1997:JBB

[Ano97-73] Anonymous. Java becomes big business: Forrester pegs Java’s penetration and offers advice on its use. Web Apps Magazine, 1 (1):11, January/February 1997. ISSN 1090-2287.

Anonymous:1997:JBE

[Ano97-74] Anonymous. Java Blend eases complex database programming. SunServer, 11(10):20, October 1997. ISSN 1091-4986.

Anonymous:1997:JBG

[Ano97-75] Anonymous. Java bugs give openings to hackers. Network Security, 1997(8):4, August 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/S135348589783228X.

Anonymous:1997:JDA

[Ano97-76] Anonymous. Java distributed across Internet. /AIXtra: IBM’s Magazine for AIX Professionals, 7(3):16, May/June 1997.

Anonymous:1997:JFE

[Ano97-77] Anonymous. Java front ends supported. Object Magazine, 6 (11):77, January 1997. CODEN OBMAFO. ISSN 1055-3614.

Anonymous:1997:JGB

[Ano97-78] Anonymous. Java GUI builder written in Java. WebServer Magazine: For Managers of World Wide Web Sites, 2(1):53, February 1997. ISSN 1087-4232.

Anonymous:1997:JML

[Ano97-79] Anonymous. Java middleware links apps. RS/Magazine, 6(2): 46, February 1997. ISSN 1088-0844, 1061-0030.

Anonymous:1997:JNS

[Ano97-80] Anonymous. Java Network Security. IBM Redbooks, 11400 Burnet Road, Austin, TX 78758-3493, USA, December 1997. ISBN ???? LCCN

???? ???? URL http://www.
redbooks.ibm.com/redpieces.
html.

Anonymous:1997:JP

[Ano97-81] Anonymous. Java-Plattform:

1997. BHV, ??, Germany, Winter 1997. ISBN 3-89360-302-6. ???? pp. LCCN ???? 60 DM.

Anonymous:1997:JPS

[Ano97-82] Anonymous. Java-Powered storage management systems. SunExpert Magazine, 8(3):82, March 1997. ISSN 1053-9239. Describes Legato Systems Inc.’s GEMS (Global Enterprise Management of Systems) front end to their Networker storage backup product.

Anonymous:1997:JPPa

[Ano97-83] Anonymous. Java Premium program provides business support. SunServer, 11(12):6, December 1997. ISSN 1091-4986.

Anonymous:1997:JPH

[Ano97-84] Anonymous. Java Programming Handbook. Softbank Publishing, ??, Japan, 1997. ISBN ???? ???? pp. LCCN ????

URL http://www.softbank.
co.jp/softbank/publishing/
;   http://www.softbank.co.
jp/softbank/publishing/books/
editer/nw/index.html.

Anonymous:1997:JRA

[Ano97-85] Anonymous. Java rapid application building tool. SunExpert Magazine, 8(9):104, September 1997. ISSN 1053-9239. Brief mention of Jyacc Inc.’s Prolifics 3 rapid application development tool.

Anonymous:1997:JRM

[Ano97-86] Anonymous. Java-to-relational middleware. WebServer Magazine: For Managers of World Wide Web Sites, 2(1):53, February 1997. ISSN 1087-4232.

Anonymous:1997:JVM

[Ano97-87] Anonymous. Java Virtual Machine unleashed. Howard W. Sams, Indianapolis, IN 46268, USA, 1997. ISBN 1-57521-247-1 (paperback). ???? pp. LCCN 9710 BOOK NOT YET IN LC.

Anonymous:1997:JJa

[Ano97-88] Anonymous. Java without Java. SunExpert Magazine, 8 (1):78, January 1997. ISSN 1053-9239. Discusses OpenConnect Systems’ integrated development environment, Open-Vista.

Anonymous:1997:Jx

[Ano97-89] Anonymous. JavaBeans.

Ohmsha, ????, November 1997. ISBN 4-274-06232-5. LCCN ???? 2,400 yen.

URL http://www.ohmsha.co.
jp/data/books/contents/4-
274-06232-5.htm.

Anonymous:1997:Jy

[Ano97-90] Anonymous. JavaBeans. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, November 1997. ISBN 4-89471-013-7. LCCN ???? 2,800 yen. URL http:

//www.mmjp.or.jp/prentice/
wa_int26-j.html.

Anonymous:1997:JCC

[Ano97-91] Anonymous. JavaBeans course covers the many sides of Java. SunServer, 11(9):32, September 1997. ISSN 1091-4986.

Anonymous:1997:JFA

[Ano97-92] Anonymous. JavaBeans features assist developers with ActiveX apps. SunServer, 11(9): 18, September 1997. ISSN 1091-4986.

Anonymous:1997:Ja

[Ano97-93] Anonymous. JavaScript. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1997. ISBN 4-88735-026-0. ???? pp. LCCN

????  URL http://puma.
asahi-net.or.jp/prentice/
wa_net11-j.html.

Anonymous:1997:Jb

[Ano97-94] Anonymous. JavaScript. Gijutsu Hyoron Co., Ltd., ??, Japan, 1997. ISBN 4-7741-0335-7. ???? pp. LCCN ????

1680 yen. URL http://www.
gihyo.co.jp/syoseki/promo/
js/index.html.

Anonymous:1997:Jj

[Ano97-95] Anonymous. JavaScript. Impress Corporation, ??, Japan, 1997. ISBN 4-8443-4778-0. ???? pp. LCCN ???? 1748 yen.

URL http://www.ips.co.jp/
;   http://www.ips.co.jp/
book/4778/4778.htm.

Anonymous:1997:Jl

[Ano97-96] Anonymous. JavaScript. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1997. ISBN 4-88735-040-6. ???? pp. LCCN

????  1600 yen.  URL

http://puma.asahi-net.or.
jp/prentice/wa_net17-j.html.

Anonymous:1997:JGC

[Ano97-97] Anonymous. JavaScript — Guia de Consulta Rapida. Novatec Editora Ltda., R Cons Moreira de Barros 1084 Conj. 01, 02018-012 São Paulo - SP, Brasil, 1997. ISBN ???? ???? pp. LCCN ???? URL http://

novatec1.com/.

Anonymous:1997:JGR

[Ano97-98] Anonymous. JavaStation gets a reaction. SunExpert Magazine, 8(1):6, January 1997. ISSN 1053-9239.

Anonymous:1997:JNJ

[Ano97-99] Anonymous. JavaStation, Netra j unveiled. SunExpert Magazine, 8(1):6, 8, January 1997. ISSN 1053-9239.

Anonymous:1997:JPPb

[Ano97-100] Anonymous. JClass Chart 2.0 creates business chart designs. SunServer, 11(12):20, December 1997. ISSN 1091-4986.

Anonymous:1997:LTO

[Ano97-101] Anonymous. Learning Tree offers Java courses. C/C++ Users Journal, 15(1):78, January 1997. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1997:LWA

[Ano97-102] Anonymous. Logic Works adds Java to Universal Directory system. SunServer, 11(10):28, October 1997. ISSN 1091-4986.

Anonymous:1997:LCC

[Ano97-103] Anonymous. Low-cost C++, Java storage products. Object Magazine, 6(11):76, January 1997. CODEN OBMAFO. ISSN 1055-3614.

Anonymous:1997:MJH

[Ano97-104] Anonymous. Mac Java/ HotJava. Cutt, ??, Japan, 1997. ISBN 4-906391-29-X. ???? pp. LCCN ???? 1,800

yen. URL http://www.cutt.
co.jp/;   http://www.cutt.
co.jp/book/macjava.html.

Anonymous:1997:MOJ

[Ano97-105] Anonymous. Manage objects in Java. BYTE Magazine, 22(7): 158, July 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:MLF

[Ano97-106] Anonymous. Marimba launches first Java/Web products. UNIX Developer, 1(1):92, January/ February 1997. ISSN 1090-2279.

Anonymous:1997:MNJ

[Ano97-107] Anonymous. Micro news: JavaOne 97. Intel vs. DEC. IEEE Micro, 17(3):5–??, May/ June 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//pascal.computer.org/mi/
books/mi1997/pdf/m3005.pdf.

Anonymous:1997:MRJ

[Ano97-108] Anonymous. Micro review: Java, SGML: Destabilizing forces for companies and careers? IEEE Micro, 17(3):3–5, May/June 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//pascal.computer.org/mi/
books/mi1997/pdf/m3003.pdf.

Anonymous:1997:MRO

[Ano97-109] Anonymous. Micro review: Only the paranoid survive; recent Java books. IEEE Micro, 17(2):4, 77, March/April 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:1997:MVI

[Ano97-110] Anonymous. Micro view: IBM bets on JavaBeans. IEEE Micro, 17(3):80–??, May/June 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//pascal.computer.org/mi/
books/mi1997/pdf/m3080.pdf.

Anonymous:1997:MTS

[Ano97-111] Anonymous. MindQ tutorial simplifies Java Workshop 2.0. SunServer, 11(11):4, November 1997. ISSN 1091-4986.

Anonymous:1997:MWS

[Ano97-112] Anonymous. Monitor a Web site with Java. WebServer Magazine: For Managers of World Wide Web Sites, 2(2):52–53, March 1997. ISSN 1087-4232.

Anonymous:1997:MJT

[Ano97-113] Anonymous. Motif-to-Java tool debuts. RS/Magazine, 6(1):53– 54, January 1997. ISSN 1088-0844, 1061-0030.

Anonymous:1997:NOJ

[Ano97-114] Anonymous. NetRight offers Java-based DMS. WebServer Magazine: For Managers of World Wide Web Sites, 2(1):52, February 1997. ISSN 1087-4232.

Anonymous:1997:NJV

[Ano97-115] Anonymous. New Java validation centers. /AIXtra: IBM’s Magazine for AIX Professionals, 7(1):6, January/February 1997.

Anonymous:1997:NPf

[Ano97-116] Anonymous. New products. C/C++ Users Journal, 15 (6):94–97, June 1997. CO-DEN CCUJEX. ISSN 1075-2838. DigitalThink Offers Online C++ and Java Courses

(http://www.digitalthink.

com/). Software Research Announces TCAT (coverage analyser) for Java (http://www.

soft.com/).

Anonymous:1997:NPC

[Ano97-117] Anonymous. New products: CodeWizard for Java. Computer, 30(6):106, June 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1997:NPI

[Ano97-118] Anonymous. New products: Integra4 RDBMS for Linux; Clustor 1.0; BitWizard; Hitachi MP-EG1A; LinkScan; Pixel!FX Version 5.1; JClass Chart; Internet/Intranet Design Shop; Velocis Database Server. Linux Journal, 38:??, June 1997. CO-DEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Anonymous:1997:NPR

[Ano97-119] Anonymous. New products: Raima Database Manager; PC Watchdog; TextSurgeon 2.0; Edith Pro/X11; Java Numeric Library; MagicFax; F-Secure Data Security Software; Stronghold 2.0. Linux Journal, 36:??, April 1997. CODEN LI-JOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Anonymous:1997:NPV

[Ano97-120] Anonymous. New products:

Visix releases Vibe Java IDE. C/C++ Users Journal, 15(7): 94, July 1997. CODEN CCU-JEX. ISSN 1075-2838.

Anonymous:1997:NPW

[Ano97-121] Anonymous. New products:

WebThreads 1.0.1; QUERYFLEX Report Writer; Linux Pro Desktop 1.0; NDP Fortran for Linux; Numerics and Visualization for Java; Craftworks Linux/AXP 2.2; InfoDock Linux Software Development Toolset; Caldera Wabi 2.2 for Linux. Linux Journal, 34:??, February 1997. CO-DEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Anonymous:1997:NSG

[Ano97-122] Anonymous. New software guards against ‘hostile applets’ using ActiveX and Java on Internet. Cisco World: The Independent Journal for Internetworking Professionals, 3(8):26, August 1997. ISSN 1081-3187.

Anonymous:1997:OSJ

[Ano97-123] Anonymous. ObjectShare ships Java components. WebServer Magazine: For Managers of World Wide Web Sites, 2(1):52– 53, February 1997. ISSN 1087-

Anonymous:1997:OAF

[Ano97-124] Anonymous. ObjectSpace announces five new objectoriented classes for C++, Java, and MFC. C/C++ Users Journal, 15(9):98, September 1997. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1997:OGJ

[Ano97-125] Anonymous. Official Gamelan Java Applet Library. Ziff-Davis Press, Emeryville, CA, USA, June 1997. ISBN 1-56276-461-6. ???? pp. LCCN ???? US$29.95. URL http://

                                        merchant.superlibrary.com:
                                        8000/catalog/hg/PRODUCT/PAGE/
                                        15627/bud/1562764616.html.

Anonymous:1997:PUJ

[Ano97-126] Anonymous. Patriot uses JavaOS to enhance ShBoom microprocessors. SunServer, 11(9): 23, September 1997. ISSN 1091-4986.

Anonymous:1997:PUM

[Ano97-127] Anonymous. Penumbra upgrades Mojo Java. UNIX Developer, 1(1):92, January/ February 1997. ISSN 1090-2279.

Anonymous:1997:RAB

[Ano97-128] Anonymous. R-Active’s Better-State PRO generates Java code. C/C++ Users Journal, 15(1): 79, January 1997. CODEN CCUJEX. ISSN 1075-2838.

Anonymous:1997:RRJ

[Ano97-129] Anonymous. Recent releases: Java: Sun Microsystems. UNIX review, 15(1):91, January 1997. CODEN UNRED5. ISSN 0742-3136.

Anonymous:1997:RFM

[Ano97-130] Anonymous. Rockwell first to market with Java microprocessor. Electronic News, 43(2187): 20, September 29, 1997. CO-DEN ELNEAU. ISSN 0013-4937, 1061-6624.

Anonymous:1997:RWI

[Ano97-131] Anonymous. Rogue Wave introduces a pair of Java updates. UNIX Developer, 1(1):92, January/February 1997. ISSN 1090-2279.

Anonymous:1997:RAJ

[Ano97-132] Anonymous. RSA announces Java encryption tool kit. Sun-Server, 11(11):19, November 1997. ISSN 1091-4986.

Anonymous:1997:RDL

[Ano97-133] Anonymous. RSA Data licenses Java encryption library. Cisco World: The Independent Journal for Internetworking Professionals, 3(8):21, August 1997. ISSN 1081-3187.

Anonymous:1997:RAN

[Ano97-134] Anonymous. RTI announces NetAcquire Java Toolkit. C/C++ Users Journal, 15(9):96, September 1997. CODEN CCU-JEX. ISSN 1075-2838.

Anonymous:1997:RRT

[Ano97-135] Anonymous. RTI releases test, measurement Java-based toolkit. HP Chronicle, 14(7):17, June 1997. ISSN 0892-2829.

Anonymous:1997:SAJ

[Ano97-136] Anonymous. Scopus adds Java to call center environment. Sun-Server, 11(9):18, September 1997. ISSN 1091-4986.

Anonymous:1997:SEJ

[Ano97-137] Anonymous. Secure embedded Java computer out. WebServer Magazine: For Managers of World Wide Web Sites, 2(2):52– 53, March 1997. ISSN 1087-4232.

Anonymous:1997:SOC

[Ano97-138] Anonymous. SGI to offer Cosmo Code Java environment on windows. Silicon Graphics World, 7 (6):3–4, June 1997. ISSN 1057-7041.

Anonymous:1997:SEP

[Ano97-139] Anonymous, editor. SGML Europe ’97 proceedings 11–16 May 1997, Princess Sofia Intercontinental, Barcelona, Spain. Inso Corporation, Boston, MA, USA, 1997. ISBN ???? LCCN ???? One CD-ROM.

Anonymous:1997:SMS

[Ano97-140] Anonymous. SoftPlus markets SunTest’s 100% Pure Java tools. SunServer, 11(10):20, October 1997. ISSN 1091-4986.

Anonymous:1997:SVJ

[Ano97-141] Anonymous. Software: Visual Java. Computer, 30(2):123, February 1997. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1997:STR

[Ano97-142] Anonymous. STEP Tools releases Java Tools for product data. Silicon Graphics World, 7(10):8, October 1997. ISSN 1057-7041.

Anonymous:1997:SSA

[Ano97-143] Anonymous. Su, Siemens agree to embed Java into smartcard chips. SunServer, 11(9):19, September 1997. ISSN 1091-4986.

Anonymous:1997:SBJ

[Ano97-144] Anonymous. Sun brews up Java development environment. SunServer, 11(11):4, November 1997. ISSN 1091-4986.

Anonymous:1997:SBP

[Ano97-145] Anonymous. Sun brings power of Java to users with disabilities. SunServer, 11(9):19, September 1997. ISSN 1091-4986.

Anonymous:1997:SBO

[Ano97-146] Anonymous. Sun bundles Object Design’s database with Java Workshop. SunServer, 11(11): 4, November 1997. ISSN 1091-4986.

Anonymous:1997:SFJ

[Ano97-147] Anonymous. Sun finalizes Java Card API 2.0 specification. Sun-Server, 11(12):20, December 1997. ISSN 1091-4986.

Anonymous:1997:SLS

[Ano97-148] Anonymous. Sun launches suite of Java testing developer tools. SunServer, 11(10):5, October 1997. ISSN 1091-4986.

Anonymous:1997:SUF

[Ano97-149] Anonymous. Sun unveils first JavaChip series of processors. SunServer, 11(12):5, December 1997. ISSN 1091-4986.

Anonymous:1997:SJR

[Ano97-150] Anonymous. Sun’s JavaEngine 1 reference platform meets demands of OEMs. SunServer, 11(10):20, October 1997. ISSN 1091-4986.

Anonymous:1997:SIJa

[Ano97-151] Anonymous. Symantec includes JScape’s Java tools in Visual family. SunServer, 11(9):32, September 1997. ISSN 1091-4986.

Anonymous:1997:SIJb

[Ano97-152] Anonymous. Symantec includes JScape’s Java tools in visual family. SunServer, 11(10):22, October 1997. ISSN 1091-4986.

Anonymous:1997:TNR

[Ano97-153] Anonymous. Technology news & reviews: Chemkin software; OpenMP Fortran Standard; ODE toolbox for Matlab; Java products; Scientific WorkPlace 3.0. IEEE Computational Science & Engineering, 4(4):75–??, October/ December 1997. CODEN IS-CEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL

http://dlib.computer.org/
cs/books/cs1997/pdf/c4075.
pdf.

Anonymous:1997:WAH

[Ano97-154] Anonymous. THE WEB authoring: HTML authoring, FrontPage 1.1, HoTMetaL PRO 3.0, Java, Symantec Café. Windows Magazine, ??(??):147–??, ???? 1997. CODEN WINMEV. ISSN 1060-1066.

Anonymous:1997:TBJ

[Ano97-155] Anonymous. Tool for building Java apps. SunExpert Magazine, 8(11):94, 96, November 1997. ISSN 1053-9239. Describes Tendril Software’s StructureBuilder 1.0.

Anonymous:1997:TDJ

[Ano97-156] Anonymous. Top drawer: Java client generator. IEEE Software, 14(2):125, March/April 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1997:UAJ

[Ano97-157] Anonymous. Ultrexx announces Java-based development tools. SunServer, 11(10):21, October 1997. ISSN 1091-4986.

Anonymous:1997:VUV

[Ano97-158] Anonymous. Visix unveils Vibe suite of Java development products. HP Chronicle, 14(7):18, June 1997. ISSN 0892-2829.

Anonymous:1997:VJA

[Ano97-159] Anonymous. Visual J++ & ActiveX. Softbank Publishing, ??, Japan, 1997. ISBN ???? ???? pp. LCCN ???? 2200 yen. URL

http://www.softbank.co.jp/
softbank/publishing/.

Anonymous:1997:VJD

[Ano97-160] Anonymous. Visual J++ Developer’s Guide. Sams.net, Indianapolis, IN 46268, USA, 1997. ISBN 1-57521-266-8. ???? pp. LCCN ???? US$55.

Anonymous:1997:VJV

[Ano97-161] Anonymous. Visual Java for Visual Basic developers. BYTE Magazine, 22(9):145–146, September 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:VNA

[Ano97-162] Anonymous. Visual Numerics announces Pure Java Certification. SunServer, 11(10):22, October 1997. ISSN 1091-4986.

Anonymous:1997:VNR

[Ano97-163] Anonymous. Visual Numerics releases JWAVE Java analysis tool. Silicon Graphics World, 7 (6):21, June 1997. ISSN 1057-7041.

Anonymous:1997:VNS

[Ano97-164] Anonymous. Visual Numerics starts Java Numeric Library beta test. Silicon Graphics World, 7(3):22, March 1997. ISSN 1057-7041.

Anonymous:1997:VQP

[Ano97-165] Anonymous. Visual Quantify points to problems in Java, C++ apps. SunServer, 11(9): 22, September 1997. ISSN 1091-4986.

Anonymous:1997:WAS

[Ano97-166] Anonymous. Will ActiveX surpass Java as the key Internet technology? Software Economics Letter, 6(2):7–8, February 1997. CODEN SECLE3. ISSN 1065-6146.

Anonymous:1997:WJH

[Ano97-167] Anonymous. Windows Java/ HotJava. Cutt, ??, Japan, 1997. ISBN 4-906391-27-3. ???? pp. LCCN ???? 1,800

yen. URL http://www.cutt.
co.jp/;   http://www.cutt.
co.jp/book/winjava.html.

Anonymous:1997:XDM

[Ano97-168] Anonymous. X Designer migrates Motif to Java. /AIXtra: IBM’s Magazine for AIX Professionals, 7(2):25, March/April 1997.

Anonymous:1998:ANE

[Ano98a] Anonymous, editor. The 19th Annual National Educational Computing Conference: San Diego Convention Center, San Diego, CA, June 22–24, 1998: NECC ’98. International Society for Technology in Education, Eugene, OR, USA, 1998. ISBN 1-56484-134-0. LCCN ????

Anonymous:1998:ASW

[Ano98b]  Anonymous.  Added secu-

rity for Web-based Java applications. Network Security, 1998(7):5, July 1998. CO-DEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S1353485898900393.

Anonymous:1998:ADV

[Ano98c] Anonymous. Application Development with VisualAge for Java Enterprise. IBM Redbooks, 11400 Burnet Road, Austin, TX 78758-3493, USA, January 1998. ISBN ???? LCCN

???? ???? URL http://www.
redbooks.ibm.com/SG245081/
sg245081.html.

Anonymous:1998:BSC

[Ano98d] Anonymous. Bentley Systems creates Java-based publishing tool. SunServer, 12(1):18, January 1998. ISSN 1091-4986.

Anonymous:1998:BRLd

[Ano98e]  Anonymous.  Book review:

A little Java, a few patterns: By Matthias Felleisen and Daniel P. Friedman. MIT Press, Cambridge, MA. (1998). $25.00. Computers and Mathematics with Applications, 35 (9):142, May 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0898122198907039.

Anonymous:1998:BREb

[Ano98f]  Anonymous.  Book review:

Exploring Java, (second edition): By Patrick Niemeyer and Joshua Peck. O’Reilly, Sebastopol, CA. (1997). 594 pages. $32.95. Computers and Mathematics with Applications, 35(9):144, May 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0898122198907179.

Anonymous:1998:BRJe

[Ano98g]  Anonymous.  Book review:

JavaScript: the definitive guide (third edition): By D. Flanagan. O’Reilly, Sebastopol, CA. (1998). 776 pages. $39.95. Computers and Mathematics with Applications, 36(8): 126, October 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0898122198911774.

Anonymous:1998:BRJd

[Ano98h]  Anonymous.  Book review:

Java^{T M} cryptography: By Jonathan Knudsen. O’Reilly, Sebastopol, CA. (1998). 344 pages. $29.95. Computers and Mathematics with Applications, 36(6):141, September 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0898122198901006.

Anonymous:1998:BRJc

[Ano98i]  Anonymous.  Book review:

Java^{T M} gems: Jewels from “Java Report”: Edited by Dwight Deugo. Cambridge University Press, Cambridge, UK.

(1998). 230 pages. $29.95. Computers and Mathematics with Applications, 36(3):126– 127, August 1998. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic). URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0898122198902097.

Anonymous:1998:BRJb

[Ano98j]  Anonymous.  Book review:

Java^{T M} security: By Scott Oaks. O’Reilly, Sebastopol, CA. (1998). 456 pages. $32.95. Computers and Mathematics with Applications, 36(3):126, August 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/S0898122198902061.

Anonymous:1998:BRMe

[Ano98k]  Anonymous.  Book review:

Modern compiler implementation in Java: Revised and expanded edition: By Andrew W. Appel. Cambridge University Press, Cambridge, MA. (1998). 548 pages. $54.95. Computers and Mathematics with Applications, 36(3):124, August 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0898122198901936.

Anonymous:1998:BJN

[Ano98l] Anonymous. Bookshelf: Java network security: Measuring software reuse: Principles, practices, and economic models: India’s software industry: State policy, liberalisation, and industrial development. IEEE Software, 15(6): 124–125, November/December 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://dlib.computer.org/
so/books/so1998/pdf/s6124.
pdf.

Anonymous:1998:BSQ

[Ano98m] Anonymous. Bookshelf: SCM:

The quick and the cheap; getting a jump-start in Java network programming; an exceptional look at fault injection. IEEE Software, 15(5): 115–118, September/October 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://dlib.computer.org/
so/books/so1998/pdf/s5115.
pdf.

Anonymous:1998:CSN

[Ano98n]  Anonymous.  From Client/

Server to Network Computing: a Migration to Java. IBM Redbooks, 11400 Burnet Road, Austin, TX 78758-3493, USA, March 1998. ISBN ???? LCCN

???? ???? URL http://www.
redbooks.ibm.com/SG242247/
sg242247.htm.

Anonymous:1998:GAL

[Ano98o] Anonymous, editor. Genetics applied to livestock production: World congress; 6th — January 1998, Armidale, Australia, volume 27 of PROCEEDINGS OF THE WORLD CONGRESS ON GENETICS APPLIED TO LIVESTOCK PRODUCTION 1998; 6th. ????, ????, 1998. ISBN 1-86389-454-3. LCCN ????

Anonymous:1998:IAF

[Ano98p] Anonymous. Innotech adds full sentence query to its Java engine. SunServer, 12(1):20, January 1998. ISSN 1091-4986.

Anonymous:1998:IJT

[Ano98q] Anonymous. Inprise’s JBuilder: Take 2. BYTE Magazine, 23 (7):85–??, July 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:Jc

[Ano98r] Anonymous. Java. Softbank Publishing, ??, Japan, April 1998. ISBN 4-7973-0550-9. LCCN ???? 3400 yen. URL

                                        http://www.softbank.co.jp/
                                        books/editor/2nd/java/.

Anonymous:1998:JBM

[Ano98s] Anonymous. Java-based management solution for global network security. Network Security, 1998(10):6, October 1998. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (elec-

                                        tronic). URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S135348589890191X.

Anonymous:1998:JRC

[Ano98t]  Anonymous.  Java: RMI,

CORBA, JavaBeans, JDBC. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, April 1998. ISBN 4-89471-044-7. LCCN ???? 4,300 yen. URL http:

//www.mmjp.or.jp/prentice/
washo/inet/wa_int36-j.html.

Anonymous:1998:JRA

[Ano98u] Anonymous. JavaSoft rides again — a peek at the highlights of the upcoming JDK 1.2. BYTE Magazine, 23(4):46–??, April 1998. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1998:JT

[Ano98v] Anonymous. JBuilder in Team.

Carl Hanser, München, Germany, 1998. ISBN 3-446-31121-

1. 900 pp. LCCN ???? 98 DM.

Anonymous:1998:JJ

[Ano98w] Anonymous. JDBC & Java.

Ohmsha, ????, May 1998. ISBN 4-274-06260-0. LCCN ???? 2,900 yen. URL http://www.

ohmsha.co.jp/data/books/contents/
4-274-06260-0.htm.

Anonymous:1998:MVJa

[Ano98x] Anonymous. Microsoft Visual J++ Programmer’s Guide. Microsoft Press, Bellevue, WA, USA, May 1998. ISBN 1-57231-869-4. xxvi + 534 pp. LCCN QA76.73.J38 M534 1998. US$28.

Anonymous:1998:MVJb

[Ano98y]  Anonymous.  Microsoft Vi-

sual J++ Reference Library: the comprehensive official resource for Microsoft Visual J++ 6.0. Microsoft Press, Bellevue, WA, USA, May 1998. ISBN 1-57231-872-4. ???? pp. LCCN QA76.73.J38M536 1998. US$100.

Anonymous:1998:NBb

[Ano98z] Anonymous. New books: All about those enterprising JavaBeans. Computer, 31(3):106, March 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://

pdf.computer.org/co/books/
co1998/pdf/r3106.pdf.

Anonymous:1998:NPE

[Ano98-27] Anonymous. New products: ENVI 3.0, Research Systems; Xbase for Linux, StarTech; Empress Hypermedia V2.10, Empress Software; SCREADER, Jos Lemmens; ascp0.8.2, Elumaze Nwanua; FontScope, CurveSoft; Pure Java JDBC Driver, Solid Information Technology Ltd.; QMASTER Batch & Output Management, QMASTER Software Solutions Inc.; Babylon, SpellCaster Telecommunications Inc.; Corel Video Network Computer, Corel Computer Corp.; W3Control Filter, W3Control Inc. Linux Journal, 46:??, February 1998. CO-DEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Anonymous:1998:NPS

[Ano98-28] Anonymous. New products:

STREAMS Data Comm Protocols, Gcom, Inc.; Raven SSL Module for Apache, Covalent Technologies; Java Workshop 2.0 for Linux, S.u.S.E.; VariCAD, VariCad LLC; Cobalt Qube Microserver, Cobalt Microserver, Inc.; Screamer 633MHz Custom Workstations, Microway; NetTracker Proxy 3.5, Sane Solutions LLC. Linux Journal, 51:??, July 1998. CO-DEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Anonymous:1998:NPV

[Ano98-29] Anonymous. New products:

Visual Prolog 5.0, Prolog Development Center; Samba: Integrating UNIX and Windows, SSC; VPN Client, Aventail Corporation; eVote 2.2, www.deliberate.com; TowerJ 2.0, Tower Technology Corporation; NetVue/JAVA, AccuSoft Corporation. Linux Journal, 49:??, May 1998. CODEN LI-JOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Anonymous:1998:NTC

[Ano98-30] Anonymous. New tools: Component technology: Rational’s component-savvy visual tool; ICL’s Java-based ORBs. systems analysis: Network Communication’s network probe; Seagate’s server health monitor. Computer, 31(6):101, June 1998. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

                                        http://dlib.computer.org/
                                        co/books/co1998/pdf/r6098.
                                        pdf.

Anonymous:1998:NTN

[Ano98-31] Anonymous. New tools: Net development: Lotus’ tools for Intranet/Extranet development; Novera Software’s Java application server. Computer, 31(6):100, 103, June 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

                                        http://dlib.computer.org/
                                        co/books/co1998/pdf/r6098.
                                        pdf.

Anonymous:1998:NTS

[Ano98-32] Anonymous. New tools: Software development: Uniscape’s internationalization library; Global Technologies’ Unix-to-NT solution; KAI’s multithreaded Java debugging tool; Price Systems’ parametric forecasting tool. Computer, 31(6):98, 102, June 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

                                        http://dlib.computer.org/
                                        co/books/co1998/pdf/r6098.
                                        pdf.

Anonymous:1998:NTSa

[Ano98-33] Anonymous. New tools: Software development: Uniscape’s internationalization library; Global Technologies’ Unix-to-NT solution; KAI’s multithreaded Java debugging tool; Price Systems’ parametric forecasting tool. Computer, 31(6):98, 102, June 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

http://dlib.computer.org/
co/books/co1998/pdf/r6098.
pdf.

Anonymous:1998:SUG

[Ano98-34] Anonymous, editor. SAS Users Group: Annual international conference; 23rd — March 1998, Nashville, TN, SUGI 1998; CONF 23. SAS, ????, 1998. ISBN 1-58025-149-8. LCCN ????

Anonymous:1998:SER

[Ano98-35] Anonymous. Sun expands Road to Java initiative. SunServer, 12(1):19, January 1998. ISSN 1091-4986.

Anonymous:1998:VJE

[Ano98-36] Anonymous. VisualAge for Java Enterprise Team. IBM Redbooks, 11400 Burnet Road, Austin, TX 78758-3493, USA, March 1998. ISBN ???? LCCN

???? ???? URL http://www.
redbooks.ibm.com/SG245245/
5245fm.htm.

Anonymous:1999:BREi

[Ano99a] Anonymous. Book review: Enterprise Javabeans^{T M}: By Richard Monson-Haefel. O’Reilly, Sebastopol, CA. (1999). 317 pages. $32.95. Computers and Mathematics with Applications, 38(11–12):284, December 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/S0898122199912130.

Anonymous:1999:BRJb

[Ano99b]  Anonymous.  Book review:

Java^{T M} 2D graphics: By Jonathan Knudsen. O’Reilly, Sebastopol, CA. (1999). 339 pages. $29.95. Computers and Mathematics with Applications, 38(11–12):282, December 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/S0898122199911914.

Anonymous:1999:BRJa

[Ano99c]  Anonymous.  Book review:

Java^{T M} servlet programming: By Jason Hunter. O’Reilly, Sebastopol, CA. (1998). 510 pages. $32.95. Computers and Mathematics with Applications, 37(8):145, April 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic). URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0898122199902079.

Anonymous:1999:GJM

[Ano99d] Anonymous. Graphic Java 1.2: Mastering the JFC, Volume 1/ Up to Speed with Swing. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1999. ISBN 1-199-45844-9 (??invalid checksum??). 1376 pp. LCCN ???? US$64.99.

Anonymous:1999:GJVb

[Ano99e] Anonymous. Graphic Java 1.2:

Volume 2. Mastering the JFC — Swing. SunSoft Press Java series. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, third edition, 1999. ISBN ???? ???? pp. LCCN ????

URL http://www.phptr.com/
bookseri/java.html.

Anonymous:1999:GJVc

[Ano99f] Anonymous. Graphic Java 1.2.

Volume 3. Mastering the JFC — 2D. SunSoft Press Java series. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, third edition, 1999. ISBN ???? ???? pp. LCCN ????

URL http://www.phptr.com/
bookseri/java.html.

Anonymous:1999:HPJ

[Ano99g] Anonymous. High performance Java compilation and runtime issues. Scientific Programming, 7(2):85, 1999. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Anonymous:1999:PII

[Ano99h] Anonymous, editor. Proceedings of the IASTED International Conference on Internet and Multimedia Systems and Applications (IMSA). Nassau, Bahamas, October 1999. Acta Press, Anaheim, CA, USA, 1999. ISBN ???? LCCN ????

Anonymous:19xx:EHJ

[Anoxxa] Anonymous. 100% Efficage HTML et Java. Easy Computing (Belgium), ????, 19xx. ISBN ???? LCCN ????

                                        920 BeF. URL http://www.
                                        easycomputing.com/francais/
                                        efficaceframe.htm.

Anonymous:19xx:AJ

[Anoxxb] Anonymous. Applets for Java. Universal CD-ROM, ????, 19xx. ISBN ???? LCCN ???? US$30.

                                        URL http://www.bigmall.
                                        com/program.html;   http:
                                        //www.bigmall.com/subucr.
                                        html.

Anonymous:19xx:CPJ

[Anoxxc] Anonymous. Como programar en Java. Prensa Técnica, ????, 19xx. ISBN 84-89245-24-X. ???? pp. LCCN ???? ????

Anonymous:19xx:CSH

[Anoxxd] Anonymous. Como se Hace con Java. Editorial Inforbooks, ????, 19xx. ISBN 84-89700-41-9 (??invalid ISBN??). ???? pp. LCCN ???? 4250 Ptas.

Anonymous:19xx:CJE

[Anoxxe]  Anonymous.  Cooking Jav-

aBeans in The Enterprise. IBM Redbooks, ????, 19xx. ISBN ???? ???? pp. LCCN ???? US$50. URL http://www.

                                        redbooks.ibm.com/SG247006/
                                        index.htm.

Anonymous:19xx:CJb

[Anoxxf] Anonymous. CORBA & Java. SRC, ????, 19xx. ISBN ???? ???? pp. LCCN ???? URL

                                        http://www.src-j.com/book/
                                        BOOK_084.html.

Anonymous:19xx:COJ

[Anoxxg] Anonymous. Corel Office for Java Programming for Dummies. IDG Books, San Mateo, CA, USA, 19xx. ISBN 0-7645-0203-4. ???? pp. LCCN ???? US$30.

Anonymous:19xx:DBF

[Anoxxh] Anonymous. Der Data Becker Führer Java. DATA-Becker, ????, 19xx. ISBN 3-8158-1561-4. ???? pp. LCCN ???? 30 DM.

Anonymous:19xx:DTJ

[Anoxxi] Anonymous. Developer Tools for Java. Universal CD-ROM, ????, 19xx. ISBN ???? LCCN ???? US$30. URL http://www.

bigmall.com/program.html.

Anonymous:19xx:DJb

[Anoxxj]  Anonymous.  DGB Java.

DATA-Becker, ????, 19xx. ISBN ???? ???? pp. LCCN ???? 69 DM.

Anonymous:19xx:DCJ

[Anoxxk] Anonymous. Distributed Computing for Java. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 19xx. ISBN ???? ???? pp. LCCN ???? ????

Anonymous:19xx:DJa

[Anoxxl] Anonymous. Dubbelboek JavaScript.

Sybex Nederland, ????, 19xx. ISBN 90-419-0162-0. ???? pp. LCCN ???? fl. 37,50; Bfr.

750. URL http://www.sybex.
nl/catalogus/index.html.

Anonymous:19xx:FSV

[Anoxxm] Anonymous. First Step of Visual J++ & ActiveX. Softbank Publishing, ??, Japan, 19xx. ISBN 4-7973-0168-6. ???? pp. LCCN ???? 2200 yen.

URL http://www.softbank.
co.jp/sbnet/books/editer/
pg/ActiveX/.

Anonymous:19xx:HCJ

[Anoxxn]  Anonymous.  HTML, CGI,

Java, VRML. Ohmsha, ????, 19xx. ISBN 4-274-13112-2. ???? pp. LCCN ???? 2,500

yen. URL http://www.ohmsha.
co.jp/data/books/contents/
4-274-13112-2.htm.

Anonymous:19xx:SJa

[Anoxxo]  Anonymous.  In 20 Stap-

pen JavaScript. Sybex Nederland, ????, 19xx. ISBN 90-419-0161-2. ???? pp. LCCN ???? fl. 19; Bfr.

380. URL http://www.sybex.
nl/catalogus/index.html.

Anonymous:19xx:IJ

[Anoxxp] Anonymous. Inleiding Java. Sybex Nederland, ????, 19xx. ISBN 90-5160-998-1. LCCN ???? fl. 49.

Anonymous:19xx:IJA

[Anoxxq] Anonymous. An Introduction to Java Applets. Softbank Publishing, ??, Japan, 19xx. ISBN 4-89052-968-3. ???? pp. LCCN ???? 2400 yen. URL

http://www.softbank.co.jp/
softbank/publishing/books/
kikan/kikan5/015.htm.

Anonymous:19xx:Jb

[Anoxxr] Anonymous. Java+. Universal CD-ROM, ????, 19xx. ISBN ???? ???? pp. LCCN ???? US$35.

Anonymous:19xx:Jca

[Anoxxs]  Anonymous.  Java.  Ai-

Publishing, ????, 19xx. ISBN 4-87193-439-X. LCCN ???? 2,524

yen. URL http://www.ai-
pub.co.jp/aipub/book/inet/
java.html.

Anonymous:19xx:Jea

[Anoxxt] Anonymous. Java. L-Gate Digital Group, ????, 19xx. ISBN 4-89621-155-3. LCCN ???? 2,800

yen. URL http://www.alles.
or.jp/~lgate/book.htm#g1.

Anonymous:19xx:Jib

[Anoxxu] Anonymous. Java. Piter Publishing Company, St. Petersburg, Russia, 19xx. ISBN ???? ???? pp. LCCN ???? URL

http://www.javaworld.com/
javaworld/books/jw-books-
alphabytitle.html;   http:
//www.piter-press.ru/;   http:
//www.piter-press.ru/koi/
ourbooks/comp/progr/jpj.html.

Anonymous:19xx:Jka

[Anoxxv] Anonymous. Java. CQ Publishing, ????, 19xx. ISBN 4-7898-3086-1. LCCN ???? 2,835

yen. URL http://www.cqpub.
co.jp/hanbai/sinkan/30861.
htm.

Anonymous:19xx:Jl

[Anoxxw] Anonymous. Java. Periplus Editions, ??, Italy, 19xx. ISBN ???? ???? pp. LCCN ???? 14000 Lire. URL ????.

Anonymous:19xx:Jla

[Anoxxx] Anonymous. Java. ASCII Corporation, Tokyo, Japan, 19xx. ISBN 4-7561-1694-9. LCCN

????  ????  URL http:/

                                        /www.ascii.co.jp/pb/book1/
                                        shinkan/detail/1190402.html.

Anonymous:19xx:Jn

[Anoxxy] Anonymous. Java. Softbank Publishing, ??, Japan, 19xx. ISBN 4-7973-0221-6. LCCN ???? 2500 yen. URL http:/

                                        /www.softbank.co.jp/sbnet/
                                        books/editer/os1/javapd/.

Anonymous:19xx:Jo

[Anoxxz] Anonymous. Java. ASCII Corporation, Tokyo, Japan, 19xx. ISBN 4-7561-1114-9. LCCN

????  ????  URL http:/

                                        /www.ascii.co.jp/pb/book1/
                                        shinkan/detail/1322406.html.

Anonymous:19xx:Jp

[Anoxx-27] Anonymous. Java. ASCII Corporation, Tokyo, Japan, 19xx. ISBN 4-7561-1721-X. LCCN

????  ????  URL http:/

                                        /www.ascii.co.jp/pb/book1/
                                        shinkan/detail/1322769.html.

Anonymous:19xx:Jq

[Anoxx-28] Anonymous.  Java.  GO-

TOP Information Inc., 5F, No.7, Lane 50, Sec.3 Nan Kang Road Taipei, Taiwan; Unit 1905,Metro Plaza Tower 2, No.223 Hing Fong Road, Kwai Chung, N.T., Hong Kong, 19xx. ISBN ???? LCCN ???? 480 NT$. URL http://

www.gotop.com/ch/book/new/
new9706cl062.html.

Anonymous:19xx:Js

[Anoxx-29] Anonymous. Java. Edition Micro Application, Paris, France, 19xx. ISBN ???? LCCN ???? 195 FF.

Anonymous:19xx:Ju

[Anoxx-30] Anonymous. Java. Chuen Hwa Books & Technology Co., Ltd., ????, 19xx. ISBN ???? LCCN ???? 550 NT$.

Anonymous:19xx:Jx

[Anoxx-31] Anonymous. Java. Flag Publishing, ??, Taiwan, 19xx. ISBN

????  LCCN ????  ????

URL http://www.flag.com.
tw/book_info/F8901.htm.

Anonymous:19xx:JAb

[Anoxx-32] Anonymous. Java & AKI-80. CQ Publishing, ????, 19xx. ISBN ???? ???? pp. LCCN ???? 1,800 yen. URL http://www.

cqpub.co.jp/hanbai/sinkan/
367080.HTM.

Anonymous:19xx:JJ

[Anoxx-33] Anonymous. Java & Javascript.

Softbank Publishing, ??, Japan, 19xx. ISBN 4-7973-0024-8. ???? pp. LCCN ???? 2700 yen. URL

http://www.softbank.co.jp/
softbank/publishing/books/
editer/pg/javahomepage/index.
html.

Anonymous:19xx:Jf

[Anoxx-34] Anonymous. Java 1.1. DaeRim Publishing, Seoul, Korea, 19xx. ISBN 89-7280-364-2 (??invalid ISBN??). ???? pp. LCCN ???? 17,000 won. URL http:

                                        //www.drbook.co.kr/daerim/
                                        tmjv11.htm.

Anonymous:19xx:JAA

[Anoxx-35] Anonymous. Java API (Applet & AWT). Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 19xx. ISBN 4-88735-036-8. ???? pp. LCCN ???? 1900

                                        yen. URL http://www.mmjp.
                                        or.jp/prentice/wa_net08-j.
                                        html.

Anonymous:19xx:JAc

[Anoxx-36] Anonymous. Java Applets. GOTOP Information Inc., 5F, No.7, Lane 50, Sec.3 Nan Kang Road Taipei, Taiwan; Unit 1905,Metro Plaza Tower 2, No.223 Hing Fong Road, Kwai Chung, N.T., Hong Kong, 19xx. ISBN ???? ???? pp. LCCN ???? 420 NT$.

Anonymous:19xx:JBM

[Anoxx-37] Anonymous. Java Boîte à Malice Le Programmeur (French) [Java Bag of Tricks: the Programmer]. Simon & Schuster Macmillan (France), Paris, France, 19xx. ISBN ???? ???? pp. LCCN ???? 99 FF. URL

                                        http://www.ssm.fr/.

Anonymous:19xx:JCC

[Anoxx-38] Anonymous. Java Core Class 1.1. Universal CD-ROM, ????, 19xx. ISBN ???? LCCN ???? US$50. URL http://www.

bigmall.com/program.html.

Anonymous:19xx:JHV

[Anoxx-39] Anonymous. Java Hands on Visual J++ 1.1. Universal CD-ROM, ????, 19xx. ISBN ???? LCCN ???? US$50. URL http:

//www.bigmall.com/program.
html.

Anonymous:19xx:JIPa

[Anoxx-40] Anonymous. Java Internet Programming. Universal CD-ROM, ????, 19xx. ISBN ???? ???? pp. LCCN ???? US$35. URL http://www.

bigmall.com/program.html.

Anonymous:19xx:JIPb

[Anoxx-41] Anonymous. Java Internet Programming. Frank Kasper, ????, 19xx. ISBN ???? ???? pp. LCCN ????

Anonymous:19xx:JIF

[Anoxx-42] Anonymous. Java Intranet Framework. Flag Publishing, ??, Taiwan, 19xx. ISBN ???? ???? pp. LCCN ???? ????

URL http://www.flag.com.
tw/book_info/F8903.htm.

Anonymous:19xx:JIPc

[Anoxx-43] Anonymous. Java Introduction to Programming 1.1. Universal CD-ROM, ????, 19xx. ISBN ???? LCCN ????

US$50.  URL http://www.
bigmall.com/program.html.

Anonymous:19xx:JJW

[Anoxx-44] Anonymous. Java JDK1.1 for Windows. Ai-Publishing, ????, 19xx. ISBN 4-87193-580-9. LCCN ???? 2,800

yen. URL http://www.ai-pub.
co.jp/aipub/book/inet/jdk.
html.

Anonymous:19xx:JLD

[Anoxx-45] Anonymous. Java Libreria Del Programmatore. Mondadori Informatica, ????, 19xx. ISBN ???? ???? pp. LCCN ???? 90000 L.

Anonymous:19xx:JPW

[Anoxx-46] Anonymous. Java pour Windows. Sybex France, Paris, France, 19xx. ISBN 2-7361-2225-9. ???? pp. LCCN ???? 128 FF. URL http://

www.sybex.com/. Includes CD-

ROM.

Anonymous:19xx:JPT

[Anoxx-47] Anonymous. Java Programmers Toolkit. Frank Kasper, ????, 19xx. ISBN ???? LCCN ???? ????

Anonymous:19xx:JPH

[Anoxx-48] Anonymous. Java Programming Handbook. Softbank Publishing, ??, Japan, 19xx. ISBN

4-89052-895-4.  ????  pp.
LCCN ????  ????  URL

http://www.softbank.co.jp/
softbank/publishing/books/
editer/nw/index.html.

Anonymous:19xx:JPI

[Anoxx-49] Anonymous. Java Programming on the Interent. Ventana Press, Chapel Hill, NC, USA, 19xx. ISBN 1-56604-442-1. ???? pp. LCCN ???? ????

Anonymous:19xx:JSM

[Anoxx-50] Anonymous. Java Security:

Managing Risks. Universal CD-ROM, ????, 19xx. ISBN ???? ???? pp. LCCN ???? US$60. URL http://www.

bigmall.com/program.html.

Anonymous:19xx:JSC

[Anoxx-51] Anonymous. Java Simple Comme Bonjour. Edition Micro Application, Paris, France, 19xx. ISBN 2-7429-0803-X. ???? pp. LCCN ???? 78 FF.

URL http://www.microapp.
com/.

Anonymous:19xx:JSKa

[Anoxx-52] Anonymous. Java Starter Kit 3.0. Sams.net, Indianapolis, IN 46268, USA, 19xx. ISBN 1-57521-311-7. ???? pp. LCCN ???? US$50.

Anonymous:19xx:JSP

[Anoxx-53] Anonymous. Java Starter Pack Shareware CD Compilation. Edition Micro Application, Paris, France, 19xx. ISBN 2-7429-0704-1. ???? pp. LCCN ???? 149 FF.

Anonymous:19xx:Ja

[Anoxx-54] Anonymous. JavaBeans.

Universal CD-ROM, ????, 19xx. ISBN ???? ???? pp. LCCN ???? US$60.

URL http://www.bigmall.
com/PJ0284.html;   http:/

/www.bigmall.com/program.
html.

Anonymous:19xx:Jia

[Anoxx-55] Anonymous. JavaBeans. Universal CD-ROM, ????, 19xx. ISBN ???? LCCN ????

US$60.  URL http://www.
bigmall.com/program.html.

Anonymous:19xx:Jc

[Anoxx-56] Anonymous. JavaScript. GO-TOP Information Inc., 5F, No.7, Lane 50, Sec.3 Nan Kang Road Taipei, Taiwan; Unit 1905,Metro Plaza Tower 2, No.223 Hing Fong Road, Kwai Chung, N.T., Hong Kong, 19xx. ISBN ???? ???? pp. LCCN ???? 340 NT$. URL http://

www.gotop.com.tw.

Anonymous:19xx:Je

[Anoxx-57] Anonymous. JavaScript. Unalis Corporation, ??, Taiwan, 19xx. ISBN 957-22-2390-9. ???? pp. LCCN ???? URL http:

//www.unalis.com.tw/index.
html;   http://www.unalis.
com.tw/~orderbk/chinese/g/
g51.htm.

Anonymous:19xx:Jg

[Anoxx-58] Anonymous. JavaScript. GO-TOP Information Inc., 5F, No.7, Lane 50, Sec.3 Nan Kang Road Taipei, Taiwan; Unit 1905,Metro Plaza Tower 2, No.223 Hing Fong Road, Kwai Chung, N.T., Hong Kong, 19xx. ISBN ???? ???? pp. LCCN ????

420 NT$. URL http://www.
gotop.com.hk/cl050.htm;
http://www.gotop.com.tw.

Anonymous:19xx:JIC

[Anoxx-59] Anonymous. JavaScript Interactive Course. La Nuova Italia, Firenze, Italy, 19xx. ISBN ???? ???? pp. LCCN ???? 118000 Lire. URL ????.

Anonymous:19xx:JHB

[Anoxx-60] Anonymous. Java/HotJava boek. Easy Computing SA-NV, Chaussée d’Alsemberg 610 - Alsembergsesteenweg 610, 1180 Brussels, Belgium; Korte Kleverlaan, 34Bf, 2061 EE Bloemendaal, The Netherlands, 19xx. ISBN 90-5167-117-2. ???? pp. LCCN ???? fl. 89. URL http:

//www.easycomputing.com/.

Anonymous:19xx:Jk

[Anoxx-61] Anonymous. JDK 1.1. Flag Publishing, ????, 19xx. ISBN ???? ???? pp. LCCN ???? 450 NT$.

Anonymous:19xx:Jw

[Anoxx-62] Anonymous. JDK 1.1. Flag Publishing, ??, Taiwan, 19xx. ISBN ???? LCCN ???? 450 NT$.

Anonymous:19xx:JJD

[Anoxx-63] Anonymous. Jolt Java Developer’s Toolkit. Universal CD-ROM, ????, 19xx. ISBN ???? LCCN ???? US$40. URL http:

//www.bigmall.com/program.
html.

Anonymous:19xx:KDD

[Anoxx-64] Anonymous. KickAss Database Development with Visual J++.

Coriolis Group Books, Scottsdale, AZ, USA, 19xx. ISBN 1-57610-068-5. ???? pp. LCCN ???? US$40.

Anonymous:19xx:PPJa

[Anoxx-65] Anonymous. PC Poche Java. Easy Computing (Belgium), ????, 19xx. ISBN ???? LCCN

????  400 BeF.  URL

http://www.easycomputing.
com/francais/pcpocheframe.
htm.

Anonymous:19xx:PPJb

[Anoxx-66] Anonymous. PC Poche Java. Edition Micro Application, Paris, France, 19xx. ISBN 2-7429-0752-1. 500 pp. LCCN ???? 59 FF. URL http://

www.microapp.com/. Includes

CD-ROM.

Anonymous:19xx:PJc

[Anoxx-67] Anonymous. Programmation Java. Easy Computing (Belgium), ????, 19xx. ISBN ???? LCCN ???? 945 BeF. URL

http://www.easycomputing.
com/francais/programmationframe.
htm.

Anonymous:19xx:PJa

[Anoxx-68] Anonymous. Programmation Java 100%. Edition Micro Application, Paris, France, 19xx. ISBN 2-7429-0779-3. ???? pp. LCCN ???? 99 FF. URL http:

//www.microapp.com/.  In-

cludes CD-ROM.

Anonymous:19xx:STP

[Anoxx-69] Anonymous. Software Training Programmatie Java. Easy Computing SA-NV, Chaussée d’Alsemberg 610 - Alsembergsesteenweg 610, 1180 Brussels, Belgium; Korte Kleverlaan, 34Bf, 2061 EE Bloemendaal, The Netherlands, 19xx. ISBN 90-5167-159-8. LCCN ???? fl. 50.

Anonymous:19xx:SJb

[Anoxx-70] Anonymous. SoftwareTraining Java. DATA-Becker, ????, 19xx. ISBN 3-8158-1303-X. LCCN ???? 49

DM. URL http://www.data-
becker.de/buecher/buecher.
htm#softwaretr.htmljava.

Anonymous:19xx:SEUb

[Anoxx-71] Anonymous. Special Edition Using Solstice Workshop. Que Corporation, Indianapolis, IN, USA, 19xx. ISBN 0-7897-0967-8. ???? pp. LCCN ????

Anonymous:19xx:TJ

[Anoxx-72] Anonymous. Tools For Java.

Universal CD-ROM, ????, 19xx. ISBN ???? LCCN ????

US$30.  URL http://www.
bigmall.com/program.html.

Anonymous:19xx:UJE

[Anoxx-73] Anonymous. Using Java By Example. Unalis Corporation, ??, Taiwan, 19xx. ISBN 957-99055-1-7 (??invalid ISBN??). ???? pp. LCCN ???? 430 NT$. URL

http://www.unalis.com.tw/

index.html. Includes CD-ROM.

Anonymous:19xx:VCJ

[Anoxx-74] Anonymous. Visual Café Java. Softbank Publishing, ??, Japan, 19xx. ISBN ???? ???? pp. LCCN ???? ???? URL http:/

                                        /www.softbank.co.jp/books/
                                        editor/4th/Cafe_Sample/index.
                                        htm.

Anonymous:19xx:VJa

[Anoxx-75] Anonymous. Visual J++. GOTOP Information Inc., 5F, No.7, Lane 50, Sec.3 Nan Kang Road Taipei, Taiwan; Unit 1905,Metro Plaza Tower 2, No.223 Hing Fong Road, Kwai Chung, N.T., Hong Kong, 19xx. ISBN ???? LCCN ???? 560 NTUS$.

Anonymous:19xx:VJc

[Anoxx-76] Anonymous. Visual J++. Edition Micro Application, ????, 19xx. ISBN ???? ???? pp. LCCN ???? 195 FF. URL

                                        http://sirius.imaginet.fr/
                                        cgi-bin/MicroApp/Webdriver?
                                        MIval=manouv01.htm&Action=
                                        Aff&ID_Categorie=1&ID_Collection=
                                        1#Produits.

Anonymous:19xx:VJd

[Anoxx-77] Anonymous. Visual J++. Unalis Corporation, ??, Taiwan, 19xx. ISBN 957-8426-03-8 (??invalid ISBN??). LCCN ????

                                        400 NT$. URL http://202.
                                        39.182.202/unalis/book/book.
                                        asp?sku=2101633.

Anonymous:19xx:VJb

[Anoxx-78] Anonymous. Visual J++ Bible. Edition Micro Application, Paris, France, 19xx. ISBN ???? ???? pp. LCCN ???? 195 FF. URL http:

//sirius.imaginet.fr/cgi-
bin/MicroApp/Webdriver?MIval=
manouv01.htm&Action=Aff&ID_
Categorie=1&ID_Collection=
1#Produits.

Anonymous:19xx:VJV

[Anoxx-79] Anonymous. Visual J++ Version 1.1 — Schnellübersicht. Markt & Technik, Haar bei München, Germany, 19xx. ISBN 382725257. LCCN ????

????  URL http://www.
mut.ch/html/buchtext.htx?
&x_pp=STb4XDV284xv2gpf&x_
pk=Direkt&x_pa=&x_pr=Fr&x_
artikelnr=25257.

Anonymous:19xx:WSJ

[Anoxx-80] Anonymous. Webmaster Super-Pack Java & HTML 3. Ziff-Davis Press, Emeryville, CA, USA, 19xx. ISBN 1-56276-425-X. ???? pp. LCCN ???? ????

Anonymous:19xx:WJ

[Anoxx-81] Anonymous. WinCafé Java.

BNN Corporation, ????, 19xx. ISBN 4-89369-435-9. ???? pp. LCCN ???? 3800 yen.

URL http://www.bnn.co.jp/
books/BOOK435-9.html.

Anonymous:2000:CNT

[Ano00] Anonymous. Correction: New tools. Computer, 33(2):7, February 2000. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

http://dlib.computer.org/

                                        co/books/co2000/pdf/r2004.

pdf. See [Kro99c].

Anthes:1996:SCJ

[Ant96] Gary H. Anthes. Sun cleans up Java security spill. Computerworld, 30(11):12, March 11, 1996. CODEN CMPWAB. ISSN 0010-4841.

Anuff:1996:JSCb

[Anu96a]  Ed Anuff.  The Java 1.1

sourcebook: a Complete Guide Updated for Java 1.1. Wiley Computer Publishers, New York, NY, USA, second edition, 1996. ISBN 0-471-17099-2 (paperback). ???? pp. LCCN ???? US$29.95, CDN$41.95.

                                        URL http://www.wiley.com/
                                        compbooks/.

Anuff:1996:JSCa

[Anu96b] Ed Anuff. The Java sourcebook: Complete guide to creating Java Applets for the Web. Wiley Computer Publishers, New York, NY, USA, 1996. ISBN 0-471-14859-8 (paperback). xiv + 498 pp. LCCN QA76.64.A58 1996. US$29.95. URL http:

                                        //www.wiley.com/compbooks/
                                        javasrbk/javahome.html.

Anuff:1997:JSb

[Anu97] Ed Anuff. The Java 1.1 Sourcebook. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, second edition, January 1997. ISBN 0-614-20340-6. ???? pp. LCCN ???? US$29.95.

Anuff:19xx:PJc

[Anuxx] Ed Anuff. Programmer en Java. Sybex France, Paris, France, 19xx. ISBN 2-7361-2259-3. ???? pp. LCCN ????

229 FF. URL http://www-
db.sybex.fr/cgi-bin/Sybex/
Webdriver?MIval=Fiche_Book.
html&Reference=2259.

Atkinson:1999:PTF

[AOV^{+}99] Malcolm P. Atkinson, Maria E.

Orlowska, Patrick Valduriez, Stanley B. Zdonik, and Michael L. Brodie, editors. Proceedings of the Twenty-fifth International Conference on Very Large Databases, Edinburgh, Scotland, UK, 7–10 September, 1999. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 1999. ISBN 1-55860-615-7. LCCN QA76.9.D3 I559 1999. Also known as VLDB’99.

Aoyagi:1997:JWP

[Aoy97] Tatsuya Aoyagi. Java: Web Pages Alive. Quality Books, ??, Japan, 1997. ISBN ????

????  pp.  LCCN ????

1700 yen. URL http://www.
qualitysoft.com/ (??).

Aoyagi:19xx:JAP

[Aoyxxa] Tatsuya Aoyagi. Java API Programming Guide. Quality Books, ????, 19xx. ISBN 4-7692-0372-1. ???? pp. LCCN ???? 3,689 yen.

URL http://www.quality.
co.jp/publishing/catalog/
catalog.html.

Aoyagi:19xx:JWP

[Aoyxxb] Tatsuya Aoyagi. Java: Web Pages Alive. Quality Books, ????, 19xx. ISBN 4-7692-0352-7. ???? pp. LCCN ???? 1,700

                                        yen. URL http://www.cup.
                                        com/java/.

Applix:1995:ADO

[App95] Applix. Applix demos one of the first enterprise-aware Sun Microsystems Java Web applications. press release, November 8, 1995. URL http://www.

                                        applix.com/pr/Java11-8-95.
                                        html.

Appel:1997:MCI

[App97]  Andrew W. Appel.  Mod-

ern Compiler Implementation in Java: Basic Techniques. Cambridge University Press, Cambridge, UK, February 1997. ISBN 0-521-58387-X (hardback), 0-521-58654-2 (paperback). x + 398 pp. LCCN QA76.73.J38A66 1997. US$74.95 (hardback), US$29.95 (paperback). URL http://

                                        www.cs.princeton.edu/~appel/

modern/java. See reviews in [Wil97i, Wil97j].

Appel:1998:MCI

[App98] Andrew W. Appel. Modern Compiler Implementation in Java. Cambridge University Press, Cambridge, UK, January 1998. ISBN 0-521-58388-8. x + 548 pp. LCCN QA76.73.J38A65 1998. US$55.

                                        URL http://193.60.94.193/

scripts/webbook.asp?isbn=
0521583888.

Abdullahi:1998:GCI

[AR98] Saleh E. Abdullahi and Graem A.

Ringwood. Garbage collecting the Internet: a survey of distributed garbage collection. ACM Computing Surveys, 30 (3):330–373, September 1998. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (elec-

tronic). URL http://www.
acm.org:80/pubs/citations/
journals/surveys/1998-30-
3/p330-abdullahi/.

Armstrong:1997:JJB

[Arm97] Eric Armstrong. JBuilder Java Bible. IDG Books, San Mateo, CA, USA, November 1997. ISBN 0-7645-3114-X. ???? pp. LCCN ???? US$50.

Armentrout:1999:TDJ

[Arm99] Anna Armentrout. A tool for designing Java programs with UML. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31 (3):180, September 1999. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Alper:1996:GMQ

[AS96] N. Alper and C. Stein. Geospatial metadata querying and visualization on the WWW using Java applets. In Card et al. [CEG96], page ?? ISBN 0-8186-7668-X. LCCN TA1634 .I325

Alagic:1998:OJI

[ASG98] S. Alagic, J. Solorzano, and D. Gitchell. Orthogonal to the Java imperative. Lecture Notes in Computer Science, 1445:212– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aslam:1998:WBQ

[Asl98]  Sohail Aslam.  WEB-based

query processing in a database course project. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(1):297–301, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

USENIX:1996:COO

[Ass96a] USENIX Association, editor. 2nd Conference on Object-Oriented Technologies & Systems (COOTS), June 17– 21, 1996. Toronto, Canada. USENIX, Berkeley, CA, USA, June 17–21, 1996. ISBN ???? LCCN ????

USENIX:1996:ATT

[Ass96b] USENIX Association, editor. 4th Annual Tcl/Tk Workshop ’96, July 10–13, 1996. Monterey, CA. USENIX, Berkeley, CA, USA, July 10–13, 1996. ISBN 1-880446-78-2. LCCN QA76.73.T44 T44 1996.

Astrachan:1998:UCE

[Ast98] Owen Astrachan. Using classes early, an object-based approach to using C++ in introductory courses (seminar). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(1):383, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ang:1997:AJM

[AT97] J.-P. Ang and Y.-T Tan. An architecture in Java for mobile computation. Lecture Notes in Computer Science, 1343:145– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Adl-Tabatabai:1998:FEC

[ATCL^{+}98] Ali-Reza Adl-Tabatabai, Michał Cierniak, Guei-Yuan Lueh, Vishesh M. Parikh, and James M. Stichnoth. Fast, effective code generation in a just-in-time Java compiler. ACM SIG-PLAN Notices, 33(5):280–290, May 1998. CODEN SIN-ODQ. ISBN 0-89791-987-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.
acm.org:80/pubs/citations/
proceedings/pldi/277650/p280-
adl-tabatabai/;   http://
www.cs.virginia.edu/pldi98/
program.html.

Atherton:1998:MJF

[Ath98] R. W. Atherton. Moving Java to the factory. IEEE Spectrum, 35(12):18–23, December 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Atkinson:1999:CPP

[Atk99] Leon Atkinson. Core PHP Programming: using PHP to build dynamic Web sites. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1999. ISBN 0-13-020787-X. xx + 568 pp. LCCN QA76.73.P224 A85 1999.

Atwood:1996:ODC

[Atw96] T. Atwood. Object databases come of age. Object Magazine, ????(????):60–63, 93, July 1996. CODEN OBMAFO. ISSN 1055-3614.

Alex:1998:TMW

[AVN98] J. Alex, B. Vikramaditya, and B. J. Nelson. Teleoperated micromanipulation within a VRML environment using Java. In IEEE [IEE98a], pages 1747– 1752. ISBN 0-7803-4465-0 (softbound), 0-7803-4466-9 (casebound), 0-7803-4467-7 (microfiche). LCCN TJ210.3.I443 1998. Three volumes. IEEE Catalog Number 98CH36190.

Arnow:1998:IPU

[AW98] David M. Arnow and Gerald Weiss. Introduction to Programming Using Java: An Object-Oriented Approach. Addison-Wesley, Reading, MA, USA, January 1998. ISBN 0-201-31184-4. xxxi + 783 pp. LCCN QA76.73.J38A77 1998. US$45. URL http:

                                        //cseng.aw.com/bookdetail.

qry?ISBN=0%2D201%2D31184%
2D4&ptype=0;   http://cseng.
aw.com/bookdetail.qry?ISBN=
0-201-31184-4&ptype=0.

Angel:1999:JQH

[AW99a] Dave Angel and Andy Wilson. Java Q&A: How do I store a Java app in a selfexecuting encrypted file? Dr. Dobb’s Journal of Software Tools, 24(2):115–116, 118, 120– 121, February 1999. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/1999/9902/9902toc.
htm;   http://www.ddj.com/
ftp/1999/1999_02/jqa299.txt;
 http://www.ddj.com/ftp/1999/
1999_02/jqa299.zip.

Asbury:1999:DJE

[AW99b] Stephen Asbury and Scott R.

Weiner. Developing Java Enterprise Applications. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 1999. ISBN 0-471-32756-5. xv + 780 pp. LCCN QA76.73.J38A83 1999. US$59.99.

Arnold:1999:JS

[AWO^{+}99] Ken Arnold, Ann Wollrath, Bryan O’Sullivan, Robert Scheifler, and Jim Waldo. The Jini specification. Addison-Wes-ley, Reading, MA, USA, 1999. ISBN 0-201-61634-3. xxii + 385 pp. LCCN QA76.9.D5 J56 1999. US$39.95.

Berg:1996:JQHa

[BA96] Cliff Berg and Steve Alexander.

Java Q and A: How do I display an image? Dr. Dobb’s Journal of Software Tools, 21 (5):103, 117, May 1996. CO-DEN DDJOEB. ISSN 1044-789X.

Ben-Ari:1997:DAJ

[BA97] Mordechai Ben-Ari. Distributed algorithms in Java. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(3):62–64, September 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Babcock:1995:SJS

[Bab95] Charles Babcock. Sun’s Java is steaming hot. Computerworld, 29(50):142, December 11, 1995. CODEN CMPWAB. ISSN 0010-4841. URL http://

                                        www.computerworld.com/.

Babcock:1996:JJE

[Bab96] Charles Babcock. Java, Java … everywhere? Computerworld, 30(42):137, October 14, 1996. CODEN CMPWAB. ISSN 0010-4841.

Babcock:1997:DIM

[Bab97a] Charles Babcock. “do it my way, or I’ll take my ball and … ”. Computerworld, 31(11): 51, March 17, 1997. CODEN CMPWAB. ISSN 0010-4841. Commentary on the Microsoft and Sun views of Java.

Babcock:1997:JMO

[Bab97b] Charles Babcock. Is Java more ‘open’ on SPARC? Computerworld, 31(10):105, March 10, 1997. CODEN CMPWAB. ISSN 0010-4841. Comments on IEEE 754 floating-point standard co-developer Jerome Coonan’s “A Note on Java Numerics” paper, which observes that Java’s requirement of identical floating-point results on all platforms imposes a run-time penalty on some platforms, notably IBM PowerPC and Intel x86, because it forces those systems to store intermediate results into memory and then immediately reload them, in order to prevent extra bits of precision in the PowerPC multiply-add instruction, or the 80-bit x86 floating-point registers, from affecting results. Babcock charges that Java is not platform neutral, as claimed by Sun, but has instead been ‘optimized for Sun SPARC and MIPS architectures.’, without understanding, or explaining, the reasoning behind the Java Virtual Machine floating-point design.

Babcock:1999:LCH

[BAC^{+}99] Dave Babcock, Peter T. Anderson, Norman S. Clerman, Ralph Hempel, Steve Grappone, Jim Cloughley, Kevin Johnson, John Bartley, and Cesar A. Gonzalez Perez. Letters: Computer history; motion-tracking devices; asleep at the keyboard?; mindstorms; Java jive; analyzing algorithms; the CVS data format URL. Dr. Dobb’s Journal of Software Tools, 24(7):12, 14, July 1999. CODEN DDJOEB.

ISSN 1044-789X. URL http://

                                        www.ddj.com/.

Back:19xx:WPJ

[Bacxxa] ?. Back. Web-Programmierrung mit Java. International Thomson Computer Press, 20 Park Plaza Suite 1001, Boston, MA 02116 USA, 19xx. ISBN ???? ???? pp. LCCN ????

                                        79 DM. URL http://www.
                                        thomson.com/itcp/default.

html. Includes CD-ROM.

Back:19xx:PJP

[Bacxxb] Svend Back. Professionelle Java-Programmierung. International Thomson Computer Press, 20 Park Plaza Suite 1001, Boston, MA 02116 USA, 19xx. ISBN 3-8266-0249-8 (??invalid ISBN??). LCCN ???? 79 DM.

Baker:1997:CDO

[Bak97a] Seán Baker. CORBA Distributed Objects Using Orbix. Addison-Wesley, Reading, MA, USA, 1997. ISBN 0-201-92475-7. xxiii + 518 pp. LCCN QA76.64.B28 1997. US$44. URL http:

                                        //cseng.aw.com/bookdetail.
                                        qry?ISBN=0%2D201%2D92475%
                                        2D7&ptype=0.

Bakkers:1997:PPJ

[Bak97b] A. Bakkers, editor. Parallel Programming and Java. IOS Press, Amsterdam, The Netherlands, 1997. ISBN 90-5199-336-

6. ???? pp. LCCN ???? US$73.

                                        URL http://www.iospress.
                                        nl/html/node98.html#SECTION00035200000000000000;

http://www.iospress.nl/html/
node99.html#SECTION00035200000000000000.

Baker:19xx:CDO

[Bakxx] Seán Baker. CORBA Distributed Objects Using Orbix. Addison-Wesley, Reading, MA, USA, 19xx. ISBN ???? LCCN ???? US$44.

URL http://cseng.aw.com/
bookdetail.qry?ISBN=0-201-
92475-7&ptype=0.

Baldazo:1996:BWS

[Bal96a] R. Baldazo. Battle of the Web site builders. BYTE Magazine, 21(7):157–158, July 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Baldazo:1996:NT

[Bal96b] R. Baldazo. Navigator turns 3.0. BYTE Magazine, 21(8):42– ??, August 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Baldazo:1996:JA

[Bal96c] Rex Baldazo. JavaScript adventures. BYTE Magazine, 21 (8):117–??, August 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Balfour:1997:AJB

[Bal97] B. Balfour. Ada 95, Java byte code, and the distributed systems annex. In ACM [ACM97b], pages 247–162. ISBN 0-89791-981-5. LCCN QA 76.73 A35 T75 1997. Theme title: Ada; the right choice for reliable software. ACM order number: 825970.

Balaban:1998:PDJ

[Bal98a] Bob Balaban. Programming Domino 4.6 with Java. MIS Press and M&T Books, P. O. Box 5277, Portland, OR 97208-5277, USA, Tel: (503) 282-5215 and M&T Publishing, Inc., 501 Galveston Drive, Redwood City, CA 94063, USA, January 1998. ISBN 1-55851-583-6. ?? pp. LCCN QA76.73.J38B35 1998. US$49.95. URL http:/

/www.idgbooks.com/cgi-bin/
db/fill_out_template.pl?idgbook:
1-5585-1583-6:book-idg::uidg1388.

Includes CD-ROM.

Balch:1998:J

[Bal98b] T. Balch. JavaSoccer. Lecture Notes in Computer Science, 1395:181–??, 1998. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Balamurugan:1999:JAP

[Bal99] S. Balamurugan. Jperl: Accessing Perl from Java. Dr. Dobb’s Journal of Software Tools, 24(2):103–105, February 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/1999/9902/9902toc.
htm;   http://www.ddj.com/
ftp/1999/1999_02/jperl.txt;
 http://www.ddj.com/ftp/1999/
1999_02/jperl.zip.

Bamforth:1997:JSS

[Bam97] R. Bamforth. Java — from smartcard to supercomputer. In Anonymous [Ano97-54], pages 1–?? ISSN 0963-3308. LCCN ????

Bank:1995:JSS

[Ban95a] David Bank. The Java saga — Sun’s Java is the hottest thing on the Web since Netscape. maybe hotter. but for all the buzz, Java nearly became a business-school case study in how a good product fails. the inside story of bringing Java to …. Wired, 3(12):166, ???? 1995. CODEN WREDEM. ISSN 1059-1028 (print), 1078-3148 (electronic).

Banks:1995:WST

[Ban95b]  David Banks.  Why Sun

thinks Hot Java will give you a lift. San Jose Mercury News, page ??, March 23, 1995. ISSN 0747-2099. URL

http://www.sjmercury.com/
archives/hotjava.htm.

Bandes:1998:JCA

[Ban98] Kenneth Bandes. Java and COM automation: Taking the edge off dual interfaces. Dr. Dobb’s Journal of Software Tools, 23(1):28, 30, 32, 34–35, 98–99, January 1998. CODEN DDJOEB. ISSN 1044-789X.

Barney:1995:NNN

[Bar95a] Doug Barney. News: Notes navigates Web. Computerworld, page ??, October 10, 1995. CO-DEN CMPWAB. ISSN 0010-4841.

Barr:1995:WTY

[Bar95b] C. Barr. Wait till you see the Net on Java. c∥net central, page ??, October 30, 1995.

URL http://www.cnet.com/
Central/Columns/Old/Barr/
10/.

Baron:1996:LNJ

[Bar96] Chris Baron. Lab note: Java programming with Symantec Café. Web Techniques, 6(1):59– 61, September 1996. CODEN WETEFA. ISSN 1086-556X.

Barlett:1997:JHS

[Bar97] N. Barlett. A Java-based highscore server. Web Techniques, 2 (3):51–54, March 1997. CODEN WETEFA. ISSN 1086-556X.

Barnes:1999:ESJ

[Bar99] Darryl Barnes. Embedded systems: Java Card application development. Dr. Dobb’s Journal of Software Tools, 24(2): 72, 74, 76–78, 80, February 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/1999/9902/9902toc.
htm;   http://www.ddj.com/
ftp/1999/1999_02/jcard.txt.

Ben-Ari:1999:DED

[BAS99] Mordechai Ben-Ari and Shawn Silverman. DPLab: an environment for distributed programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(3):

91–94, September 1999. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bath:1996:FMQ

[Bat96a] M. Bath. Framing multimedia [Quark Immedia]. Publishing, ????(????):29–??, August 1996. CODEN PUBLFA.

Bath:1996:MWW

[Bat96b] M. Bath. Multimedia web weaving [online publishing]. Publishing, ????(????):34–36, March 1996. CODEN PUBLFA.

Batt:1996:NCL

[Bat96c] P. Batt. The Net comes to life: Java and VRML extend the World Wide Web into an advantageous mass-media network. Bulletin des Schweizerischen Elektrotechnischen Vereins and des Verbandes Schweizerischer Elektrizitaetswerke, 87(9):11– 16, April 26, 1996. CODEN BEVEDP. ISSN 0036-1321.

Batt:1996:NBL

[Bat96d] Paul Batt. Das Netz beginnt zu leben — Java und VRML erweitern das World Wide Web zu einem nutzbringenden Massenmedium. Bulletin de l’Association Suisse des Electriciens, 87(9):11–??, ???? 1996. CODEN BUSEAH. ISSN 0004-587X.

Bayle:1997:MRK

[Bay97a] Aime J. Bayle. Media reviews: Keeping up with Java: Java by Example by J. R.

Jackson and A. L. McClellan (SunSoft Press, 1997, 426 pp., $34.95, ISBN 0-13-272295-X). IEEE MultiMedia, 4(3):91–92, July–September 1997. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://dlib.computer.org/
mu/books/mu1997/pdf/u3091.
pdf.

Bayomi:1997:SJC

[Bay97b] Donald Bayomi. Starting Java: Crash Course Videotape. Share-Views Video Tutor, ????, January 1997. LCCN ???? US$25.

URL http://www.mbnet.mb.
ca/shareviews/svvideos.html.

Bercik:1996:IJ

[BB96] Bill Bercik and Jill Bond. Inside JavaScript. New Riders Publishing, Carmel, IN, USA, October 1, 1996. ISBN 1-56205-593-3. xxii + 661 pp. LCCN QA76.73.J38 B47 1996. US$44.99.

Barstow:1997:PIR

[BB97a] D. R. Barstow and W. B. Brogden. Personalized interactive real-time sports reporting using Java. American Programmer, 10(1):32–37, January 1997. CO-DEN AMPRFD. ISSN 1048-5600.

Bigus:1997:CIA

[BB97b] Joseph Bigus and Jennifer Bigus. Constructing Intelligent Agents With Java: a Programmer’s Guide to Smarter Applications. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, December 1997. ISBN 0-471-19135-3. xxxv + 379 pp. LCCN QA76.76.I58B563 1997. US$45.

URL http://www.wiley.com/
compbooks/catalog/19135-3.
htm.

Bond:1997:JP

[BB97c] Jill Bond and Bill Berick.

JavaScript für Profis. Academic Service Verlag, ????, 1997. ISBN 3-932073-11-8 (??invalid ISBN??). ???? pp. LCCN ???? 85 DM. URL ????.

Bruzzone:1999:MUS

[BB99] A. Bruzzone and G. Berrino.

Modeling of urban services by VRML and Java. In Bruzzone et al. [BUP99], pages 34– 40. ISBN 1-56555-156-7. LCCN QA76.9.C65 I574 1999.

Back:1996:PJP

[BBB^{+}96] Svend Back, Klaus Bergius, Siegmar Beier, Peter Majorczyk, et al. Professionelle Java-Programmierung: Leitfaden für Entwickler. International Thomson Computer Press, 20 Park Plaza Suite 1001, Boston, MA 02116 USA, 1996. ISBN 3-8266-0249-8. 394 pp. LCCN ???? 79,00 DM. URL

http://www.itp.de/online/
0249/0249.html. Includes CD-

ROM.

Bhatia:1997:WVP

[BBC^{+}97] Dimple Bhatia, Vanco Burzevski, Maja Camuseva, Geoffrey Fox, Wojtek Furmanski, and Girish Premchandran. WebFlow — a visual programming paradigm for Web/Java based coarse grain distributed computing. Concurrency: Practice and Experience, 9(6):555–577, June 1997. CODEN CPEXEI. ISSN 1040-3108. URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=13877;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=13877&PLACEBO=IE.pdf.

Baentsch:1999:JHR

[BBE^{+}99] Michael Baentsch, Peter Buhler, Thomas Eirich, Frank Höring, and Marcus Oestreicher. JavaCard — from hype to reality. IEEE Concurrency, 7 (4):36–43, October/December 1999. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL

http://dlib.computer.org/
pd/books/pd1999/pdf/p4036.
pdf;   http://www.computer.
org/concurrency/pd1999/p4036abs.
htm.

Baker:1999:TVT

[BBG^{+}99] Ryan S. Baker, Michael Boilen, Michael T. Goodrich, Roberto Tamassia, and B. Aaron Stibel. Testers and visualizers for teaching data structures. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31 (1):261–265, March 1999. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bik:1998:CSJ

[BBGD98] Aart J. C. Bik, Fabian Breg, Dennis B. Gannon, and Paul F. Dubois. Compiler support for Java parallelization. Computers in Physics, 12(1):45–??, January 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:/

/aip.scitation.org/doi/10.
1063/1.168686.

Borchardt:1996:CAS

[BBJR96] F. L. Borchardt, C. L. Bradin, E. Johnson, and L. Rhodes, editors. CALICO ’96 Annual Symposium. Proceedings of the Computer Assisted Language Instruction Consortium 1996 Annual Symposium ‘Distance Learning’, Albuquerque, NM, USA, 27 May–1 June 1996. Duke University, Durham, NC, USA, 1996. ISBN ???? LCCN ????

Ball:1997:OPG

[BC97a] Derek Ball and David Cinderella. Official Power++: Getting Started. Ventana Press, Chapel Hill, NC, USA, November 1997. ISBN 1-85032-862-5. LCCN ???? US$45.

Boyland:1997:PMI

[BC97b] John Boyland and Giuseppe Castagna. Parasitic methods: An implementation of multimethods for Java. ACM SIG-PLAN Notices, 32(10):66–76, October 1997. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ball:1998:CIA

[BC98a] Steve Ball and John Miller Crawford. Channels for interapplet communication. Dr. Dobb’s Journal of Software Tools, 23(9):18–20, 22, 91, September 1998. CODEN DDJOEB. ISSN 1044-789X.

                                        URL http://www.ddj.com/
                                        ftp/1998/1998_09/channel.
                                        txt;   http://www.ddj.com/
                                        ftp/1998/1998_09/channel.
                                        zip.

Blount:1998:EJN

[BC98b] B. Blount and S. Chatterjee. An evaluation of Java for numerical computing. Lecture Notes in Computer Science, 1505:35– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bowen:1998:ZWU

[BC98c] J. P. Bowen and D. Chippington. Z on the Web using Java. Lecture Notes in Computer Science, 1493:66–??, 1998. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ball:1999:JQJ

[BC99a] Steve Ball and John Miller Crawford. Java Q and A: Are Java applets independent programs? Dr. Dobb’s Journal of Software Tools, 24(4): 101–105, April 1999. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.

ddj.com/ftp/1999/1999_04/
jqa499.txt;   http://www.
ddj.com/ftp/1999/1999_04/
webcad.zip.

Blount:1999:EJN

[BC99b] Brian Blount and Siddhartha Chatterjee. An evaluation of Java for numerical computing. Scientific Programming, 7(2):97–110, 1999. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL http://iospress.metapress.
com/app/home/contribution.
asp%3Fwasp=f277qlrwwjr5m4vxjywv%
26referrer=parent%26backto=
issue%2C3%2C8%3Bjournal%2C7%
2C9%3Blinkingpublicationresults%
2C1%2C1.

Butrico:1997:GRM

[BCC^{+}97] Maria A. Butrico, Henry Chang, Anthony Cocchi, Norman H. Cohen, Dennis G. Shea, and Stephen E. Smith. Gold Rush: Mobile transaction middleware with Java-object replication. In USENIX [USE97b], pages 91–101. ISBN 1-880446-86-3. LCCN QA76.64 .U845

1997. URL http://dxsting.
cern.ch/sting/COOTS97.html.

Bergadano:1998:SWT

[BCE98] F. Bergadano, B. Crispo, and M. Eccettuato. Secure WWW transactions using Standard HTTP and Java applets. In USENIX [USE98a], page ?? ISBN 1-880446-97-9. LCCN HF5004 .U74

1998.  URL http://www.

                                        usenix.org/publications/library/
                                        proceedings/ec98/crispo.html.

Beca:1997:JEC

[BCF^{+}97] Lukasz Beca, Gang Cheng, Geoffrey C. Fox, Tomasz Jurga, Konrad Olszewski, Marek Podgorny, Piotr Sokolowski, and Krzysztof Walczak. Java enabling collaborative education, health care, and computing. Concurrency: Practice and Experience, 9(6):521–533, June 1997. CODEN CPEXEI. ISSN 1040-3108. URL http:

                                        //www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=13875;
                                         http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=13875&PLACEBO=IE.pdf.

Baker:1999:MOO

[BCFK99] M. Baker, B. Carpenter, G. Fox, and Sung Hoon Koo. mpiJava: An object-oriented Java interface to MPI. Lecture Notes in Computer Science, 1586:748– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baratloo:1998:FHR

[BCH^{+}98] Arash Baratloo, P. Emerald Chung, Yennun Huang, Sampath Rangarajan, and Shalini Yajnik. Filterfresh: Hot replication of Java RMI server objects. In USENIX [USE98b], page ?? ISBN ???? LCCN

                                        ????  URL http://www.
                                        usenix.org/publications/library/
                                        proceedings/coots98/baratloo.
                                        html.

Bernardin:1999:SCJ

[BCK99] Laurent Bernardin, Bruce Char, and Erich Kaltofen. Symbolic computation in Java: an appraisement. In Dooley [Doo99], pages 237–244. ISBN 1-58113-073-2. LCCN QA76.95 .I57 1999. URL http://

www.acm.org/pubs/contents/
proceedings/issac/309831/.

Baker:1995:NMA

[BCLT95] J. E. Baker, I. F. Cruz, G. Liotta, and R. Tamassia. A new model for algorithm animation over the WWW. ACM Computing Surveys, 27(4):568– 572, December 1995. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic).

Baker:1996:AGA

[BCLT96] James E. Baker, Isabel F. Cruz, Giuseppe Liotta, and Roberto Tamassia. Animating geometric algorithms over the Web. In ACM [ACM96], pages C3–C4. CODEN PACGET. ISBN 0-89791-832-0. LCCN QA 448 D38 S96 1996.

Blount:1999:IPA

[BCP99a] B. Blount, S. Chatterjee, and M. Philippsen. Irregular parallel algorithms in Java. Lecture Notes in Computer Science, 1586:1026–??, 1999. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Braun:1999:LTA

[BCP^{+}99b] Jay Braun, David Cox, Christophe Ponsard, Reg Charney, Torken Danielsson, Christian Lanctot, John D. Kent, and Steve Furlong. Letters: There’s always room for Forth; online op/eds; parsing expressions in Java; age discrimination; A better date?; more Y2K; it’s a cult kind of thing. Dr. Dobb’s Journal of Software Tools, 24 (3):10, 12, March 1999. CO-DEN DDJOEB. ISSN 1044-

789X. URL http://www.ddj.
com/1999/9902/9902toc.htm.

Beaver:1997:LNN

[BCW97] Jason Beaver, Jamie Costa, and Jason Wehling. Late Night Netscape IFC. Ziff-Davis Press, Emeryville, CA, USA, 1997. ISBN 1-56276-540-X. xx + 575 pp. LCCN ???? US$50. URL http://

merchant.superlibrary.com:
8000/catalog/hg/PRODUCT/PAGE/
15627/bud/156276540X.html.

Berk:1998:JLA

[BD98a] Elliot Berk and Martin Dirichs. JLex: a lexical analyzer generator for Java, 1998. URL http:/

/www.cs.princeton.edu/~appel/
modern/java/JLex/. Lex is

the Java programming tool formerly known as Java-Lex. JLex is a lexical analyzer generator, written for Java, in Java. JLex was developed by Elliot Berk at Princeton University. It is now maintained with the assistance of Martin Dirichs, University of Oldenburg, Germany.

Borkner-Delcarlo:1998:UJP

[BD98b] Olaf Borkner-Delcarlo. Umstieg auf Java: Professionelle Sotwareentwicklung mit dem JDK 1.1. Carl Hanser, München, Germany, 1998. ISBN 3-446-19216-6. xv + 523 pp. LCCN ???? 89 DM. URL http:

//www.hanser.de/computer/
buecher/bo19216.htm.

Becchini:1999:JEM

[BD99] R. Becchini and G. De Petris.

Java enabled MPEG-4 services: The MPEG-J framework. Lecture Notes in Computer Science, 1597:373–384, 1999. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baldazo:1996:BCB

[BDF96] Rex Baldazo, Stanford Diehl, and Rich Friedman. Books and CD-ROMs: Baby’s first Java book. BYTE Magazine, 21(4):41–??, April 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Bak:1998:NCJ

[BDF^{+}98] Lars Bak, John Duimovich, Jesse Fang, Scott Meyer, and David Ungar. The new crop of Java virtual machines (panel). ACM SIGPLAN Notices, 33 (10):179–182, October 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Borchers:1996:LRG

[BDKK96] J. Borchers, O. Deussen, A. Klingert, and C. Knorzer. Layout rules for graphical Web documents. Computers and Graphics, 20(3):415–426, May– June 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Boisvert:1998:DNLa

[BDP^{+}98a] Ronald F. Boisvert, Jack J. Dongarra, Roldan Pozo, Karin A. Remington, and G. W. Stewart. Developing numerical libraries in Java. In ACM [ACM98a], page ?? CODEN CPEXEI. ISSN 1040-3108. LCCN ????

URL http://www.cs.ucsb.
edu/conferences/java98/papers/
jnt.pdf;   http://www.cs.
ucsb.edu/conferences/java98/

papers/jnt.ps. Also published as Concurrency: Practice and Experience, 10(11–13), September 1998, CODEN CPEXEI, ISSN 1040-3108.

Boisvert:1998:DNLb

[BDP^{+}98b] Ronald F. Boisvert, Jack J. Dongarra, Roldan Pozo, Karin A. Remington, and G. W. Stewart. Developing numerical libraries in Java. Concurrency: Practice and Experience, 10(11–13):1117–1129, September 1998. CODEN CPEXEI. ISSN 1040-3108. URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=10050395;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=10050395&PLACEBO=IE.pdf.

Special Issue: Java for Highperformance Network Computing.

Breg:1998:JRPa

[BDV^{+}98a] Fabian Breg, Shridhar Diwan, Juan Villacis, Jayashree Balasubramanian, Esra Akman, and Dennis Gannon. Java RMI performance and object model interoperability: Experiments with Java/ HPC++. In ACM [ACM98a], page ?? CODEN CPEXEI. ISSN 1040-3108. LCCN ????

URL http://www.cs.ucsb.
edu/conferences/java98/papers/
rmi.pdf;   http://www.cs.
ucsb.edu/conferences/java98/

papers/rmi.ps. Also published as Concurrency: Practice and Experience, 10(11–13), September 1998, CODEN CPEXEI, ISSN 1040-3108.

Breg:1998:JRPb

[BDV^{+}98b] Fabian Breg, Shridhar Diwan, Juan Villacis, Jayashree Balasubramanian, Esra Akman, and Dennis Gannon. Java RMI performance and object model interoperability: experiments with Java/HPC++. Concurrency: Practice and Experience, 10(11–13):941–955, September 1998. CODEN CPEXEI. ISSN 1040-3108. URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=10050400;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=10050400&PLACEBO=IE.pdf.

Special Issue: Java for Highperformance Network Computing.

Beyls:1999:JJP

[BDY99] K. Beyls, E. D’Hollander, and Y. Yu. JPT: a Java parallelization tool. In Dongarra et al. [DLM99], pages 173– 180. ISBN 3-540-66549-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 E973 1999.

Bear:1996:YOJ

[Bea96] C. Bear. Your own Java layout manager. Web Techniques, 1(7): 28–33, October 1996. CODEN WETEFA. ISSN 1086-556X.

Beck:1997:WJN

[Bec97a] B. Beck. World-ready Java: The next steps. In Unicode Consortium [Uni97b], pages A16–?? ISBN ???? LCCN ???? URL http:

                                        //www.unicode.org/unicode/
                                        iuc11/index.html;   http:
                                        //www.unicode.org/unicode/
                                        uni2book/u2ord.html.

Becker:1997:HPB

[Bec97b] Jocelyn Becker. How To Program Bongo. Ziff-Davis Press, Emeryville, CA, USA, 1997. ISBN 1-56276-534-5. xix + 326 pp. LCCN QA76.625 .B43 1997. US$40. URL http://

                                        merchant.superlibrary.com:
                                        8000/catalog/hg/PRODUCT/PAGE/
                                        15627/bud/1562765345.html.

Beck:1998:WRJ

[Bec98] Brian Beck. World ready Java — the next step — part 1: General I18N, text layout and user interface. In Unicode Consortium [Uni98a], page ?? ISBN ???? LCCN ???? URL http:

//www.unicode.org/unicode/
iuc12/program.html.

Beebe:1996:BPJd

[Bee96] Nelson H. F. Beebe. A bibliography of publications in Java Report: The Source for Java Development. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, September 6, 1996. 5 pp. URL

https://www.math.utah.edu/
pub/tex/bib/index-table-j.

html#javarep. This report is updated frequently.

Beebe:1998:BPJi

[Bee98] Nelson H. F. Beebe. A bibliography of publications in Java-World magazine. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, August 14, 1998. 101 pp. URL

https://www.math.utah.edu/
pub/tex/bib/index-table-j.
html#javaworld. This report

is updated frequently.

Behme:1995:IKH

[Beh95] Henning Behme. Im Kaffeehaus. HotJava: objektorientierte Programmiersprache und Web-Viewer (English: In the coffee house. HotJava: objectoriented programming language and Web viewer). WWW URL, July 1995. URL http:

                                        //www.ix.de/ix/raven/Web/
                                        9507/HotJava.html.

Beker:1997:LEJ

[Bek97] Harry Beker. Letter to the editor: Java hype versus Tcl/Tk reality. Dr. Dobb’s Journal of Software Tools, 22(1):8, 10, January 1997. CODEN DDJOEB. ISSN 1044-789X. Argues that Tcl/Tk provides a much more advanced windowing interface than Java currently does.

Bell:1998:JSb

[Bel98] Parr Bell. Java für Studenten. Markt & Technik, Haar bei München, Germany, September 1998. ISBN 3-8272-9563-7 (??invalid ISBN??). LCCN ???? 70 DM.

Bennahum:1996:T

[Ben96a] David S. Bennahum. Technology. New York, 29(1):18–??, January 1996. ISSN 0028-7369.

Bennet:1996:ICI

[Ben96b]  Gordon Bennet.  Intranets:

Como implantar com sucesso na sua empresa. Editore Campus Ltda., Rua Sete de Setembro, 111/16, Rio de Janeiro - RJ - 20050-002, Brazil, 1996. ISBN ???? 380 pp. LCCN ???? R$ 39,90. URL http:

                                        //www.campus.com.br/about-
                                        us.htm;   http://www.campus.
                                        com.br/catalogo/livro.cfm?
                                        cod_livro=20117.

Benson:1998:JRR

[Ben98] Brent W. Benson, Jr. Java reflections: Reflection: Eat your own dog food. ACM SIGPLAN Notices, 33(12):16–19, December 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Benson:1999:JRI

[Ben99a] Brent W. Benson, Jr. Java reflections: Inner classes: Closures for the masses. ACM SIGPLAN Notices, 34(2):32–35, February 1999. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://ftp.

informatik.rwth-aachen.de/
dblp/db/indices/a-tree/b/
Benson_Jr=:Brent_W=.html.

This paper describes recent extensions to the Java language languages to support inner classes and closures.

Benson:1999:JRJ

[Ben99b] Brent W. Benson, Jr. Java reflections: JavaScript — not Java (but just as hot). ACM SIGPLAN Notices, 34(4):25–27, April 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://ftp.
informatik.rwth-aachen.de/
dblp/db/indices/a-tree/b/
Benson_Jr=:Brent_W=.html.

Benson:1999:J

[Ben99c] Brent W. Benson, Jr. JavaScript.

ACM SIGPLAN Notices, 34 (4):25–27, April 1999. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Berg:1996:HDI

[Ber96a] C. Berg. How do I implement a drag-and-drop interface? [Java]. Dr. Dobb’s Journal of Software Tools, 21(12):110, 112, 116, December 1996. CODEN DDJOEB. ISSN 1044-789X.

Berg:1996:HDT

[Ber96b] C. Berg. How do threads work and how can I create a generalpurpose event? Dr. Dobb’s Journal of Software Tools, 21 (11):111–115, 126–127, November 1996. CODEN DDJOEB. ISSN 1044-789X.

Berg:1996:JQHe

[Ber96c] C. Berg. Java Q and A: How do I display media formats that Netscape does not support? Dr. Dobb’s Journal of Software Tools, 21(10):111, 114–115, 129, October 1996. CODEN DDJOEB. ISSN 1044-789X.

Berg:1996:JQHd

[Ber96d] Cliff Berg. Java Q and A: How do I create a layout manager? Dr. Dobb’s Journal of Software Tools, 21(9):101–??, September 1996. CODEN DDJOEB. ISSN 1044-789X.

Berg:1996:JQHg

[Ber96e] Cliff Berg. Java Q and A: How do I implement a drag-and-drop interface? Dr. Dobb’s Journal of Software Tools, 21(12):110– ??, December 1996. CODEN DDJOEB. ISSN 1044-789X.

Berg:1996:JQHc

[Ber96f] Cliff Berg. Java Q and A:

How do I send e-mail from a Java applet? Dr. Dobb’s Journal of Software Tools, 21(8): 111–??, August 1996. CODEN DDJOEB. ISSN 1044-789X.

Berg:1996:JQHb

[Ber96g] Cliff Berg. Java Q and A:

How do I write a chat program. Dr. Dobb’s Journal of Software Tools, 21(6):105–106, June 1996. CODEN DDJOEB. ISSN 1044-789X.

Berg:1996:JQHf

[Ber96h] Cliff Berg. Java Q and A: How do threads work and how can I create a general-purpose event? Dr. Dobb’s Journal of Software Tools, 21(11):111–??, November 1, 1996. CODEN DDJOEB. ISSN 1044-789X.

Bergin:1996:JBC

[Ber96i] Joseph Bergin. Java as a better C++. ACM SIGPLAN Notices, 31(11):21–27, November 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bergin:1996:OTC

[Ber96j] Joseph Bergin. Object technology in the classroom: Java as a better C++. ACM SIGPLAN Notices, 31(11):21–27, November 1996. CODEN SINODQ.

ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Berg:1997:JQHa

[Ber97a] C. Berg. Java Q and A: How do I use CORBA from Java? Dr. Dobb’s Journal of Software Tools, 22(1):100, 102, 103, 114, January 1997. CO-DEN DDJOEB. ISSN 1044-789X.

Berg:1997:JQHc

[Ber97b] C. Berg. Java Q and A: How do I use Java remote method invocation from an applet? Dr. Dobb’s Journal of Software Tools, 22(3):101– 103, March 1997. CODEN DDJOEB. ISSN 1044-789X.

Berg:1997:JQHe

[Ber97c] Cliff Berg. Java Q and A: How can I create a push Java channel? Dr. Dobb’s Journal of Software Tools, 22(5):99, 101– 103, 118, 122, 123, May 1997. CODEN DDJOEB. ISSN 1044-789X.

Berg:1997:JQHb

[Ber97d] Cliff Berg. Java Q and A: How do I access a SQL database from an applet? Dr. Dobb’s Journal of Software Tools, 22(2):103, 105, 113, 114, February 1997. CODEN DDJOEB. ISSN 1044-789X.

Berg:1997:JQHl

[Ber97e] Cliff Berg. Java Q and A: How do I browse and dynamically invoke remote objects. Dr. Dobb’s Journal of Software Tools, 22 (12):121–123, December 1997. CODEN DDJOEB. ISSN 1044-789X.

Berg:1997:JQHi

[Ber97f] Cliff Berg. Java Q and A:

How do I create a JavaBean? Dr. Dobb’s Journal of Software Tools, 22(9):102–104, 114, 115, September 1997. CODEN DDJOEB. ISSN 1044-789X.

Berg:1997:JQHh

[Ber97g] Cliff Berg. Java Q and A: How do I create a signed applet? Dr. Dobb’s Journal of Software Tools, 22(8):109–111, 122, August 1997. CODEN DDJOEB. ISSN 1044-789X.

Berg:1997:JQHf

[Ber97h] Cliff Berg. Java Q and A: How do I create my own security manager? Dr. Dobb’s Journal of Software Tools, 22(6):115, 117–120, 132, June 1997. CO-DEN DDJOEB. ISSN 1044-789X. See letter to the editor [Bou97] pointing out a flaw in the security manager.

Berg:1997:JQHd

[Ber97i] Cliff Berg. Java Q and A: How do I create persistent Java objects? Dr. Dobb’s Journal of Software Tools, 22(4):98–101, April 1997. CODEN DDJOEB. ISSN 1044-789X.

Berg:1997:JQHk

[Ber97j] Cliff Berg. Java Q and A: How do I print in Java? Dr. Dobb’s Journal of Software Tools, 22 (11):110–114, November 1997. CODEN DDJOEB. ISSN 1044-789X.

Berg:1997:JQHj

[Ber97k]  Cliff Berg.  Java Q and

A: How do I write a Java servet? Dr. Dobb’s Journal of Software Tools, 22(10):121–123, 130, 131, October 1997. CO-DEN DDJOEB. ISSN 1044-789X.

Berg:1997:JQH

[Ber97l] Cliff Berg. Java Q and A: How do I write an international application? Dr. Dobb’s Journal of Software Tools, 22(7):105– 106, 116, July 1997. CODEN DDJSDM. ISSN 1044-789X.

Berg:1997:JQHg

[Ber97m] Cliff Berg. Java Q and A: How do I write an international application? Dr. Dobb’s Journal of Software Tools, 22(7):105– 106, 116, July 1997. CODEN DDJSDM. ISSN 1044-789X.

Berg:1998:JQHa

[Ber98a] Cliff Berg. Java Q and A: How do I create a signed Castanet channel? Dr. Dobb’s Journal of Software Tools, 23(1):121–124, 129, January 1998. CODEN DDJOEB. ISSN 1044-789X.

Berg:1998:JQHc

[Ber98b] Cliff Berg. Java Q and A: How do I interface Java to my PalmPilot. Dr. Dobb’s Journal of Software Tools, 23 (4):118, 120, 122, 139, April 1998. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/ftp/1998/1998_04/
jqa498.txt.

Berg:1998:JQHd

[Ber98c] Cliff Berg. Java Q and A:

How do I password encrypt data? Dr. Dobb’s Journal of Software Tools, 23(8):107– 109, 117, August 1998. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/1998/1998_08/
jqa898.txt;   http://www.
ddj.com/ftp/1998/1998_08/
jqa898.zip.

Berg:1998:JQHb

[Ber98d] Cliff Berg. Java Q and A:

How do I transfer data securely? Dr. Dobb’s Journal of Software Tools, 23(2):119–121, February 1998. CODEN DDJOEB. ISSN 1044-789X.

Berg:1998:AJD

[Ber98e] Clifford J. Berg. Advanced Java Development for Enterprise Applications. SunSoft Press Java series. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, March 1998. ISBN 0-13-080461-4. 512 pp. LCCN QA76.73.J38B478 1998. US$49.95. URL http://www.

prenhall.com/ptrbooks/ptr_
0130804614.html.  Includes

CD-ROM.

Berg:1998:JF

[Ber98f] Clifford J. Berg. Java FAQs.

Prentice-Hall, Englewood Cliffs, NJ 07632, USA, February 1998.

ISBN 0-13-272980-6. ???? pp. LCCN ???? US$26.95.

Bergmann:1998:UIC

[Ber98g] Seth Bergmann. Using inheritance for computing with physical quantities: an interdisciplinary project. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(1):45–47, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Berman:1998:SPS

[Ber98h] A. Michael Berman. Selected papers from the Second United Kingdom: Java in the Computing Curriculum (JICC) Conference. ACM SIGPLAN Notices, 33(4):30, April 1998. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Berman:1998:SSS

[Ber98i] A. Michael Berman. Special section: Selected papers from the Second UK Java in the Computing Curriculum Conference. ACM SIGPLAN Notices, 33(4):30–??, April 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

                                        URL http://www.rowan.edu/
                                        sigplan/apr199~1.htm.

Berg:1999:PEJ

[Ber99a] Cliff Berg. Parsing expressions in Java: Specifying rules for common operations. Dr. Dobb’s Journal of Software Tools, 24 (1):50, 52–53, 56–58, January 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/ddj/ftp/1999/1999_
01/jparser.txt;   http://
www.ddj.com/ddj/ftp/1999/
1999_01/jparser.zip.

Bernardin:1999:JFMa

[Ber99b] Laurent Bernardin. A Java framework for massively distributed symbolic computing. Mathematics and Computers in Simulation, 49(3): 151–160, August 1999. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0378475499000488.

Bernardin:1999:JFMb

[Ber99c] Laurent Bernardin. A Java framework for massively distributed symbolic computing. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 33 (3):20–21, September 1999. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Betts:1995:NSF

[Bet95] Mitch Betts. News: Small firms’ patent claims may slow Internet commerce. Computerworld, page ??, September 01, 1995. CODEN CMPWAB. ISSN 0010-4841. URL http://

www.computerworld.com/.

Berg:1998:ATJ

[BF98] Daniel J. Berg and J. Stephen Fritzinger. Advanced Techniques for Java Developers. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 1998. ISBN 0-471-18208-7 (softcover). xvii + 426 pp. LCCN QA76.73.J38B48 1998. US$44.99.

                                        URL http://www.wiley.com/
                                        compbooks/catalog/18208-7.

htm. Includes CD-ROM.

Berg:1999:ATJ

[BF99a] Daniel J. Berg and J. Steven Fritzinger. Advanced Techniques for Java Developers. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, revised edition, 1999. ISBN 0-471-32718-2. xv + 511 pp. LCCN QA76.73.J38B48 1999. US$49.99.

Buhler:1999:AAJ

[BF99b] Gerhard Bühler and Heinz Faßbender. Applying Ada, Java and CORBA for making a command and control information system platform independent. ACM SIGADA Ada Letters, 19 (3):83–88, September 1999. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Boccato:19xx:J

[BFxx] Monica Boccato and Pierangelo Ferrari. Java. Jackson Libri, via XXV Aprile, 39 - 20091 Bresso (MI), Italy, 19xx. 80 pp. LCCN ???? 9500 L.

Bic:1996:DCU

[BFD96] Lubomir F. Bic, Munehiro Fukuda, and Michael B. Dillencourt. Distributed computing using autonomous objects. Computer, 29(8):55–61, August 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bonjour:19xx:JEM

[BFGGxx] Michel Bonjour, Gilles Falquet, Jacques Guyot, and André Le Grand. Java: De l’esprit à la méthode. International Thomson Computer Press, 20 Park Plaza Suite 1001, Boston, MA 02116 USA, 19xx. ISBN 2-84180-134-9. 442 pp. LCCN ???? 240 FF.

URL http://cuiwww.unige.
ch/java/livre/;   http://
cuiwww.unige.ch/~bonjour/
; http://cuiwww.unige.ch/~
falquet/;   http://cuiwww.
unige.ch/~guyot/; http://
cuiwww.unige.ch/~legrand/;
 http://www.thomson.com/intluk.

html. Includes CD-ROM.

Brackenbury:1998:IES

[BFGS98] I. F. Brackenbury, D. F. Ferguson, K. D. Gottschalk, and R. A. Storey. IBM’s Enterprise Server for Java. IBM Systems Journal, 37(3):323– ??, ???? 1998. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

URL http://www.almaden.
ibm.com/journal/sj/373/brackenbury.
html.

Basin:1999:JBV

[BFPV99] David Basin, Stefan Friedrich, Joachim Posegga, and Harald Vogt. Java bytecode verification by model checking. Lecture Notes in Computer Science, 1633:491–??, 1999. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://web.
                                        sec.uni-passau.de/papers/
                                        CAV99.pdf.

Bala:1996:UO

[BG96] Raghuram Bala and Kate Gregory. Using Optima++ 1.5. Que Corporation, Indianapolis, IN, USA, 1996. ISBN 0-7897-0894-9. xx + 634 pp. LCCN QA76.73.C153B335 1996. US$50. URL http:

                                        //www.mcp.com/cgi-bin/bag?
                                        isbn=0-7897-0894-9&last=/
                                        bookstore.

Bertacca-Guidi:1997:IJ

[BG97a] Bertacca-Guidi. Introduzione a Java. McGraw-Hill, New York, NY, USA, October 1997. ISBN 88-8386-076-4. LCCN ???? 35000 L.

Bik:1997:AEI

[BG97b] Aart J. C. Bik and Dennis B. Gannon. Automatically exploiting implicit parallelism in Java. Concurrency: Practice and Experience, 9(6):579–619, June 1997. CODEN CPEXEI. ISSN 1040-3108. URL http:

                                        //www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=13878;
                                         http://www3.interscience.

wiley.com/cgi-bin/fulltext?
ID=13878&PLACEBO=IE.pdf.

Bik:1997:NNL

[BG97c] Aart J. C. Bik and Dennis B.

Gannon. A note on native level 1 BLAS in Java. Concurrency: Practice and Experience, 9(11):1091–1099, November 1997. CODEN CPEXEI. ISSN 1040-3108. URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=13826;
http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=13826&PLACEBO=IE.pdf. Spe-

cial Issue: Java for computational science and engineering

— simulation and modeling II.

Bik:1998:PBPb

[BG98a] Aart J. C. Bik and Dennis B.

Gannon. A prototype bytecode parallelization tool. Concurrency: Practice and Experience, 10(11–13):879–885, September 1998. CODEN CPEXEI. ISSN 1040-3108. URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=10050417;
http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=10050417&PLACEBO=IE.pdf.

Special Issue: Java for Highperformance Network Computing.

Bik:1998:PBP

[BG98b] Aart J. C. Bik and Dennis B.

Gannon. A prototype bytecode parallelization tool (javab manual). In ACM [ACM98a], page ?? CODEN CPEXEI.

ISSN 1040-3108. LCCN ????

URL http://www.cs.ucsb.
edu/conferences/java98/papers/
javab.pdf;   http://www.cs.
ucsb.edu/conferences/java98/
papers/javab.ps. Also pub-

lished as Concurrency: Practice and Experience, 10(11– 13), September 1998, CODEN CPEXEI, ISSN 1040-3108.

Bik:1998:PBPa

[BG98c] Aart J. C. Bik and Dennis B. Gannon. A prototype bytecode parallelization tool (javab manual). In ACM [ACM98a], page ?? CODEN CPEXEI. ISSN 1040-3108. LCCN ????

URL http://www.cs.ucsb.
edu/conferences/java98/papers/
javab.pdf;   http://www.cs.
ucsb.edu/conferences/java98/
papers/javab.ps. Also pub-

lished as Concurrency: Practice and Experience, 10(11– 13), September 1998, CODEN CPEXEI, ISSN 1040-3108.

Boroni:1998:PSI

[BGGR98] Christopher M. Boroni, Frances W. Goosey, Michael T. Grinder, and Rockford J. Ross. A paradigm shift! The Internet, the Web, browsers, Java and the future of computer science education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30 (1):145–152, March 1998. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bik:1999:IIM

[BGH99] Aart J. C. Bik, Milind Girkar, and Mohammad R. Haghighat. Incorporating Intel(R) MMX^{T M} technology into a Java^{T M} JIT compiler. Scientific Programming, 7(1):167–184, 1999. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:

//iospress.metapress.com/
app/home/contribution.asp%
3Fwasp=64cr5a4mg33tuhcbdr02%
26referrer=parent%26backto=
issue%2C7%2C7%3Bjournal%2C8%
2C9%3Blinkingpublicationresults%
2C1%2C1.

Binas-Holz:1996:JP

[BH96a] Antje Binas-Holz. Das Java Programmierbuch. Sybex Verlag GmbH, Erkrather Straße 345-349, 40231 Düsseldorf, Germany, 1996. ISBN 3-8155-7225-8. 840 pp. LCCN ???? 98 DM. URL http://

www.abc.de/gutenb/verlage/
sybex/sybex.htm.  Includes

CD-ROM.

Binas-Holz:1996:JR

[BH96b] Antje Binas-Holz. Das Java Referenzbuch. Sybex Verlag GmbH, Erkrather Straße 345-349, 40231 Düsseldorf, Germany, October 1996. ISBN 3-8155-5009-2. ???? pp. LCCN ???? 69 DM. URL http://

www.abc.de/gutenb/verlage/
sybex/sybex.htm.  Includes

CD-ROM.

Briere:1996:SJE

[BH96c] Daniel Briere and Christine Heckart. Sun’s Java effort shines on BT/MCI. Network World, 13(51):22, December 16, 1996. ISSN 0887-7661 (print), 1944-7655 (electronic).

Back:1999:DRL

[BH99a] G. Back and W. Hsieh. Drawing the red line in Java. In IEEE [IEE99b], pages 116–

121. ISBN 0-7695-0237-7, 0-7695-0238-5 (case), 0-7695-0239-3 (microfiche). LCCN QA76.76.O63 W6666 1999. IEEE Computer Society Press order number PR00237.

Bogda:1999:RUS

[BH99b] Jeff Bogda and Urs Hölzle. Removing unnecessary synchronization in Java. ACM SIG-PLAN Notices, 34(10):35–46, October 1999. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://www.

                                        acm.org/pubs/citations/proceedings/
                                        oops/320384/p35-bogda/.

Binas-Holz:19xx:JRT

[BHxx] Antje Binas-Holz. Java Referentiehandboek — A tot Z. Sybex Nederland, ????, 19xx. ISBN 90-419-0070-5. ???? pp. LCCN ???? fl. 69. URL http:

                                        //www.sybex.nl/;   http://
                                        www.sybex.nl/catalogus/programmeren/
                                        azjava.htm;   http://www.
                                        sybex.nl/nieuw/azjava.htm.

Includes CD-ROM.

Bhattacharyya:1998:YTS

[Bha98] Dev Bhattacharyya. A Year 2000 tool suite. Dr. Dobb’s Journal of Software Tools, 23 (5):36, 38, May 1998. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/1998/1998_05/
javay2k.zip.

Bielecki:19xx:JAC

[Biexxa]  Jan Bielecki.  Java after

C++. Intersoftland, ????, 19xx. ISBN 83-86861-02-9 (??invalid ISBN??). ???? pp. LCCN ???? ????

Bielecki:19xx:JPC

[Biexxb] Jan Bielecki. Java po C++ [English: Java after C++]. Intersoftland, Warszawa, Poland, 19xx. ISBN ???? ???? pp. LCCN ???? URL http://

www.enter.pol.pl/ent96.10/
en5083.htm.

Bijl:1996:CSO

[Bij96]  D. Bijl.  Client/server, an

overview. Informatie, 38(????): 6–12, October 1996. CODEN INFTCR. ISSN 0019-9907.

Binko:1996:JP

[Bin96] Bill Binko. Java and Postgres95.

Linux Journal, 31:??, November 1996. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Bingler:1997:DJU

[Bin97] R. Bingler. Developing Javabased Unicode applications: a case study. In Unicode Consortium [Uni97b], pages A12–?? ISBN ???? LCCN ???? URL http:

//www.unicode.org/unicode/
iuc11/index.html;   http:
//www.unicode.org/unicode/
uni2book/u2ord.html.

Bishop:1997:JGP

[Bis97] Judy (Judith) M. Bishop. Java Gently: Programming Principles Explained. Addison-Wesley, Reading, MA, USA, 1997. ISBN 0-201-59399-8 (softcover). xxiii + 391 pp. LCCN QA76.73.J38B58 1997. US$36.95. URL http:

//cseng.aw.com/bookdetail.
qry?ISBN=0-201-59399-8&ptype=
0.

Bishop:1998:JG

[Bis98] Judy M. Bishop. Java Gently. Addison-Wesley, Reading, MA, USA, second edition, April 1998. ISBN 0-201-34297-9. xvii + 508 pp. LCCN QA76.73.J38 B55 1998. US$37.

URL http://jupiter.cs.up.
ac.za/javagently/.

Borges:1996:AAI

[BK96a] Dibio L. (Dibio Leander) Borges and Celso A. A. (Celso Antonio Alves) Kaestner, editors. Advances in artificial intelligence: 13th Brazilian Symposium on Artificial Intelligence, SBIA ’96, Curitiba, Brazil, October 23–25, 1996: proceedings, volume 1159 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 3-540-94900-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334 .S5434 1996 Bar.

Bothun:1996:NPU

[BK96b] G. D. Bothun and S. D. Kevan. Networked physics in undergraduate instruction. Computers in Physics, 10(4):318– 324, July–August 1996. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

Buhrmann:1996:JII

[BK96c] Peter Buhrmann and Thomas Kretzberg. Java: Interaktiv im World Wide Web. Addison-Wesley, Reading, MA, USA, second edition, 1996. ISBN 3-8273-1151-9. 96 pp. LCCN ????

79.90 DM. URL http://www.
addison-wesley.de/katalog/
item.ppml?textexpr=Java&id=
00090.

Budimlic:1997:OJT

[BK97a] Zoran Budimlic and Ken Kennedy. Optimizing Java: theory and practice. Concurrency: Practice and Experience, 9(6):445–463, June 1997. CODEN CPEXEI. ISSN 1040-3108. URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=13870;
http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=13870&PLACEBO=IE.pdf.

Buhrmann:1997:JII

[BK97b] Peter Buhrmann and Thomas Kretzberg. Java: interaktiv im WorldWideWeb; [Computer based training; inklusive Netscape 2.0]. Addison-Wes-ley, Reading, MA, USA, 1997. ISBN 3-8273-1027-X (??invalid ISBN??). 72 pp. LCCN ????

79,90 DM. URL http://www.
addison-wesley.de/. Includes

CD-ROM.

Benton:1999:IWT

[BK99] Nick Benton and Andrew Kennedy. Interlanguage working without tears: Blending SML with Java. ACM SIG-PLAN Notices, 34(9):126–137, September 1999. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://www.

acm.org/pubs/citations/proceedings/
fp/317636/p126-benton. Pro-

ceedings of the ACM SIGPLAN International Conference on Functional Programming (ICFP ’99).

Buhrmann:19xx:CJ

[BKxxa] Peter Buhrmann and Thomas Kretzberg. Das CBT zu Java. Addison-Wesley, Reading, MA, USA, 19xx. ISBN 3-8273-1027-X (??invalid ISBN??). LCCN ???? 80 DM.

Buhrmann:19xx:JII

[BKxxb] Peter Buhrmann and Thomas Kretzberg. Java: Interaktiv im World Wide Web. Addison-Wesley, Reading, MA, USA, 19xx. ISBN 3-8273-1151-9 (??invalid ISBN??). LCCN ???? 80 DM.

Barnett:1998:DII

[BKCG98] Lewis Barnett, Joseph F. Kent, Justin Casp, and David Green. Design and implementation of an interactive tutorial framework. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30 (1):87–91, March 1998. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Baratloo:1998:INCa

[BKKK98a] A. Baratloo, M. Karaul, H. Karl, and Z. M. Kedem. An infrastructure for network computing with Java applets. In ACM [ACM98a], page ?? CODEN CPEXEI. ISSN 1040-3108. LCCN ????

URL http://www.cs.ucsb.
edu/conferences/java98/papers/
knittingfactory.pdf;   http:
//www.cs.ucsb.edu/conferences/
java98/papers/knittingfactory.

ps. Also published as Concurrency: Practice and Experience, 10(11–13), September 1998, CODEN CPEXEI, ISSN 1040-3108.

Baratloo:1998:INCb

[BKKK98b] Arash Baratloo, Mehmet Karaul, Holger Karl, and Zvi M. Kedem. An infrastructure for network computing with Java applets. Concurrency: Practice and Experience, 10 (11–13):1029–1041, September 1998. CODEN CPEXEI.

ISSN 1040-3108. URL http:

                                        //www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=10050421;
                                         http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=10050421&PLACEBO=IE.pdf.

Special Issue: Java for Highperformance Network Computing.

Baratloo:1996:CMW

[BKKW96] A. Baratloo, M. Karaul, Z. Kedem, and P. Wyckoff. Charlotte: metacomputing on the Web. In Yetongnon and Hariri [YH96], pages 2–?? ISBN ???? LCCN ???? Two volumes.

Bacon:1998:TLF

[BKMS98] David F. Bacon, Ravi Konuru, Chet Murthy, and Mauricio Serrano. Thin locks: Featherweight synchronization for Java. ACM SIGPLAN Notices, 33(5): 258–268, May 1998. CODEN SINODQ. ISBN 0-89791-987-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

                                        tronic). URL http://www.
                                        acm.org:80/pubs/citations/
                                        proceedings/pldi/277650/p258-
                                        bacon/;   http://www.cs.virginia.
                                        edu/pldi98/program.html.

Bergeron:1997:VDH

[BKP97a] Marcel Bergeron, Corinne Kempa, and Yolande Perron. Vocabulaire d’Internet: HTML, Java, VRML, cyberculture: terminologie des technologies de l’information: vocabulaire anglais-français. Cahiers de l’Office de la langue française Cahiers de l’Office de la langue française (Unnumbered). Gouvernement du Québec, Office de la langue française, Québec, PQ, Canada, deuxième edition, 1997. ISBN 2-551-17286-1. 140 pp. LCCN ????

Buhrmann:1997:JCB

[BKP97b] Peter Buhrmann, Thomas Kretzberg, and Martin Pein. Java 1.1: ne Einführung in die Programmierung; mit Online-Aktualisierung; hören, sehen, verstehen; inkl. lizenzierter Vollversion. Addison-Wesley, Reading, MA, USA, third edition, September 1997. ISBN 3-8273-1236-1. 16 pp. LCCN ????

89,90 DM. URL http://www.
addison-wesley.de/katalog/
item.ppml?id=00207;   http:
//www.addison-wesley.de/katalog/
item.ppml?textexpr=Java&id=
00207.

Benton:1999:CSM

[BKR99] Nick Benton, Andrew Kennedy, and George Russell. Compiling Standard ML to Java bytecodes. ACM SIGPLAN Notices, 34(1): 129–140, January 1999. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bloomberg:1996:WPS

[BKT96] Jason Bloomberg, Jeff Kawski, and Paul Treffers. Web Page Scripting Techniques. Hayden Books, 4300 West 62nd Street, Indianapolis, IN 46268, USA, 1996. ISBN 1-56830-307-6. xviii + 510 pp. LCCN QA76.625.B556 1996. US$50.00, CDN$70.95, UK£46.95. URL http://www.

                                        mcp.com/hayden;   http://
                                        www.mcp.com/hayden/internet/
                                        scripting/index.html. Dis-

cusses JavaScript, VBScript, and advanced HTML scripting techniques.

Breuel:1997:LEJ

[BL97] Thomas M. Breuel and Ted Lewis. Letter to the editor: Java as universal language. Computer, 30(4):6–7, April 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Blaha:1996:OIC

[Bla96] Stephen Blaha. Object-oriented interprocess communication: Client/server development in C++ and Java. Dr. Dobb’s Journal of Software Tools, 21 (8):24–??, 26, 28, 30–31, 89– 92, August 1996. CODEN DDJOEB. ISSN 1044-789X.

Blaha:1998:JPI

[Bla98] Stephen Blaha. Java Programming with an Intro to Visual J++, volume 1. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, February 1, 1998. ISBN 0-13-262197-5. ???? pp. LCCN ???? US$35.95. URL http:

                                        //www.prenhall.com/books/
                                        ptr_0132621975.html.

Blanchet:1999:EAO

[Bla99] Bruno Blanchet. Escape analysis for object-oriented languages: application to Java. ACM SIGPLAN Notices, 34 (10):20–34, October 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

URL http://www.acm.org/
pubs/citations/proceedings/
oops/320384/p20-blanchet/.

Barrett:1999:EJW

[BLB99] Dan Barrett, Dan Livingston, and Micah Brown. Essential JavaScript for Web Professionals. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1999. ISBN 0-13-013056-7. xviii + 198 pp. LCCN QA76.76.J39 B37 1999. URL http:

//www.phptr.com/ptrbooks/
ptr_0130130567.html.

Brogden:1998:VCJ

[BLT98] William B. Brogden, Jeffrey A.

Louie, and Ed Tittel. Visual Café for Java Explorer: database development edition: Maximize Your Object-Oriented Programming Skills to Create Database Applets and Applications Using Java. Coriolis Group Books, Scottsdale, AZ, USA, 1998. ISBN 1-57610-219-X. xxii + 558 pp. LCCN QA76.73.J38 T59 1998. US$39.99; CDN$55.99.

URL https://www.coriolis.
com/cgi-win/tipsheet.exe/
tipsheet?name=vcproex. In-

cludes CD-ROM.

Baufeldt:1996:JPJ

[BM96a] Kai Baufeldt and Rolf Mäurers.

Java, Professionelle Java-Applets für das Internet. DATA-Becker, ????, 1996. ISBN 3-8158-4886-3 (??Invalid checksum??). ???? pp. LCCN ???? 20 DM.

Boone:1996:LJM

[BM96b] Barry Boone and Dave Mark. Learn Java on the Macintosh. Addison-Wesley Developers Press, Reading, MA, USA, 1996. ISBN 0-201-19157-1. x + 475 pp. LCCN QA76.8.M3B663 1996. US$34.95. URL http://

                                        www.aw.com/.

Bothner:1996:KSI

[BM96c] Per Bothner and R. Alexander Milowsk. The Kawa Scheme interpreter project. Web page, Winternet Corporation, 12 S 6th Street, Suite M110, Minneapolis, MN 55402-1507, USA, 1996. ???? pp. URL http:

                                        //www.winternet.com/~sgml/
                                        kawa (no   longer accessible)

. Implements a compiler of

Scheme into Java.

Baker:1998:DSJ

[BM98a] Brenda S. Baker and Udi Manber. Deducing similarities in Java sources from bytecodes. In USENIX [USE98e], page ?? ISBN 1-880446-94-4. LCCN QA76.76.O63

                                        U88 1998. URL http://www.
                                        usenix.org/publications/library/
                                        proceedings/usenix98/baker.
                                        html.

Benage:1998:SEU

[BM98b] Don Benage and Azam Mirza. Special Edition Using Microsoft Visual Studio for Enterprise Development. Que Corporation, Indianapolis, IN, USA, February 1998. ISBN 0-7897-1260-1. LCCN QA76.76.A65B46 1998. US$50. URL http://

merchant.superlibrary.com:
8000/catalog/mcp/PRODUCT/
PAGE/07897/bud/0789712601.
html.

Burton:1998:AJP

[BM98c] Bradley F. Burton and Victor W. Marek. Applications of JAVA programming language to database management. SIG-MOD Record (ACM Special Interest Group on Management of Data), 27(1):27–34, March 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bank:1997:PTJ

[BML97a] Joseph A. Bank, Andrew C.

Myers, and Barbara Liskov. Parameterized types for Java. In ACM [ACM97a], pages 132–145. ISBN 0-89791-853-3. LCCN QA 76.7 A11

1997.  URL http://www.
acm.org:80/pubs/citations/
proceedings/plan/263699/p132-
bank/.

Barghouti:1997:GGP

[BML97b] N. S. Barghouti, J. Mocenigo, and W. Lee. Grappa: A GRAPh PAckage in Java. Lecture Notes in Computer Science, 1353:336–??, 1997. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brown:1996:DAO

[BN96] Marc H. Brown and Marc A. Najork. Distributed active objects. Computer Networks and ISDN Systems, 28(7–11): 1037–1052, May 1, 1996. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/
                                        browse/browse.cgi?year=1996&
                                        volume=28&issue=7-11&aid=
                                        1586.

Ben-Natan:1997:WOD

[BN97] Ron Ben-Natan. Web Objects: Designing, Building, and Deploying Object-Oriented Applications for the Web. McGraw-Hill, New York, NY, USA, 1997. ISBN 0-07-006281-1. xix + 488 pp. LCCN QA76.64.B45 1997. US$40.

Ben-Natan:1998:CW

[BN98a] Ron Ben-Natan. CORBA on the Web. McGraw-Hill, New York, NY, USA, February 1998. ISBN 0-07-006724-4. xxi + 480 pp. LCCN QA76.625.B46 1998. US$45. Includes CD-ROM.

Bergin:1998:GWG

[BN98b] Joseph Bergin and Thomas L. Naps. Group 3 (working group): curricular opportunities of Java based Internet software development. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30 (3):270, September 1998. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bertsch:1999:FPT

[BN99a] Eberhard Bertsch and Mark-Jan Nederhof. On failure of the pruning technique in “Error Repair in Shift-Reduce Parsers”. ACM Transactions on Programming Languages and Systems, 21(1):1– 10, January 1999. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL http://www.acm.org/
pubs/citations/journals/toplas/
1999-21-1/p1-bertsch/. See

[MYD95].

Braux:1999:TPE

[BN99b] Mathias Braux and Jacques Noyé. Towards partially evaluating reflection in Java. ACM SIGPLAN Notices, 34 (11):2–11, November 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

URL http://www.acm.org/
pubs/citations/proceedings/
pepm/328690/p2-braux/.

Bergin:1998:JRC

[BNB^{+}98] Joseph Bergin, Thomas L.

Naps, Constance G. Bland, Stephen J. Hartley, Mark A. Holliday, Pamela B. Lawhead, John Lewis, Myles F. McNally, Christopher H. Nevison, Cheng Ng, George J. Pothering, and Tommi Teräsvirta. Java resources for computer science instruction. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(4):18–38, December 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bode:1996:PVM

[Bod96] Arndt Bode, editor. Parallel virtual machine, EuroPVM ’96: third European PVM conference, Munich, Germany, October 7–9, 1996: proceedings, volume 1156 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 3-540-61779-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .E975 1996 Bar.

Bokowski:1999:CSC

[Bok99] Boris Bokowski. CoffeeStrainer: statically-checked constraints on the definition and use of types in Java. ACM SIGSOFT Software Engineering Notes, 24 (6):355–374, November 1999. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Booch:1996:ODJ

[Boo96a] G. Booch. Object-oriented development with Java. Report on Object Analysis and Design, 2(6):2–4, March–April 1996. ISSN 1075-2528.

Booker:1996:JJD

[Boo96b] Ellis Booker. Java juices up D&B Software’s corporate Web tools. Web Week, 2(7):4, June 1996. CODEN WEWEFP. URL

http://www.webweek.com/.

Boone:1996:JEC

[Boo96c] Barry Boone. Java Essentials for C and C++ Programmers. Addison-Wesley Developers Press, Reading, MA, USA, May 1, 1996. ISBN 0-201-47946-X. xxi + 311 pp. LCCN QA76.64 .B675 1996. US$19.95.

URL http://cseng.aw.com/
bookdetail.qry?ISBN=0-201-
47946-X&ptype=0;   http://
www.aw.com/devpress/titles/
47946.html.

Boone:1997:JCP

[Boo97] Barry Boone. Java Certification for Programmers and Developers. McGraw-Hill, New York, NY, USA, July 1997. ISBN 0-07-913657-5 (hardcover). xxiv + 744 pp. LCCN QA76.3 .B66 1997. US$65. URL http:

//mcgraw-hill.inforonics.
com/cgi/getarec?mgh31641%

comp. Includes CD-ROM.

Boone:1998:JCE

[Boo98] Barry Boone. Java 1.2 Certification Exam Guide. McGraw-Hill, New York, NY, USA, February 1998. ISBN 0-07-913740-7. LCCN ???? US$55.

URL http://mcgraw-hill.
inforonics.com/cgi/getarec?
mgh50908%comp.

Borland:1995:BID

[Bor95] Borland International. Borland International to deliver tools for Sun’s Java Internet programming language. press release, November 8, 1995. URL http:

                                        //www.borland.com/Product/
                                        java/javapress.html.

Boreham:1996:JNC

[Bor96] B. Boreham. Java: the new C++. Object Expert, 1(4):58, 60–61, May–June 1996. CO-DEN OBEXFJ. ISSN 1360-3426.

Boszormenyi:1998:WJM

[Bös98]  László Böszörményi.  Why

Java is not my favorite firstcourse language. Software — Concepts and Tools, 19(3): 141–145, March 1998. CO-DEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic). URL https://

                                        link.springer.com/article/
                                        10.1007/s003780050017.

Bracha:1998:MFS

[BOSW98] Gilad Bracha, Martin Odersky, David Stoutamire, and Philip Wadler. Making the future safe for the past: Adding genericity to the Java programming language. ACM SIGPLAN Notices, 33(10):183–200, October 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bothner:1998:KCD

[Bot98a] Per Bothner. Kawa — compiling dynamic languages to the Java VM. In USENIX [USE98c], page ?? ISBN ????

LCCN ???? URL http://www.
usenix.org/publications/library/
proceedings/usenix98/freenix/
bothner.slides.ps.

Bottomley:1998:ISL

[Bot98b] Gareth Bottomley. Implementation of a scripting language for VRML using Java. University of Leeds, School of Computer Studies, Leeds, 1998.

Bouchard:1996:JOP

[Bou96a] Oliver Bouchard. Java: objektorientiertes Programmieren für das WWW. Beck/DTV, ????, 1996. ISBN 3-423-50185-5. ???? pp. LCCN ???? 20 DM.

Bournellis:1996:SJP

[Bou96b] Cynthia Bournellis. Sun’s Java platform gains in reaching bigger audience. Electronic News, 42(2146):6, December 9, 1996. CODEN ELNEAU. ISSN 0013-4937, 1061-6624.

Boukanov:1997:JS

[Bou97] Igor Boukanov. Java security. Dr. Dobb’s Journal of Software Tools, 22(9):10, September 1997. CODEN DDJOEB. ISSN 1044-789X. Points out a flaw in the security manager described in [Ber97h].

Bowman:1996:IBO

[Bow96a] Charles F. Bowman. Implementing Business Objects with Java. SIGS Books, 71 W 23rd St., New York, NY 10010, USA, 1996. ISBN 1-884842-56-9. 50 pp. LCCN QA76.9.O35A45

1997.  US$85.00.  URL

http://www.sigs.com/books/
wp_impjava.html.

Bowman:1996:XBF

[Bow96b] Charles F. Bowman. X: Back to the future: An interview with Robert Scheifler. The X Journal, 5(5):38–43, May 1996. CO-DEN XJOUEA. ISSN 1056-7003. Discusses relationship of Broadway to Java.

Boyer:1998:AAR

[Boy98a] John Boyer. Algorithm alley: Resizable data structures. Dr. Dobb’s Journal of Software Tools, 23(1):115–116, 118, 129, January 1998. CODEN DDJOEB. ISSN 1044-789X. Discusses some deficiencies of the Java library hash table support, and compares it with his algorithm and that used in the C++ Standard Template Library. Also compares the Jenkins hash function [Wil97i] with the one proposed in this paper.

Boyer:1998:AAS

[Boy98b] John Boyer. Algorithm alley: Sorting and searching linked lists in Java. Dr. Dobb’s Journal of Software Tools, 23(5):126– 129, 137–138, May 1998. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/1998/1998_05/
aa0598.txt;   http://www.
ddj.com/ftp/1998/1998_05/
aa0598.zip.

Boyle:1998:VEM

[Boy98c] John Boyle. A visual environment for the manipulation and integration of Java Beans. Bioinformatics, 14(8):739–748, September 1998. CODEN ???? ISSN 1367-4803 (print), 1367-4811 (electronic).

Bozman:1995:NSN

[Boz95a] Jean S. Bozman. News: Sun needs commercial users for hot Java to percolate. Computerworld, page ??, June 09, 1995. CODEN CMPWAB. ISSN 0010-4841.

Bozman:1995:NSTb

[Boz95b] Jean S. Bozman. News: Sun ties Internet software to DOE. Computerworld, page ??, September 22, 1995. CODEN CMPWAB. ISSN 0010-4841.

Bozman:1995:NSTa

[Boz95c] Jean S. Bozman. News: Sun turns focus on Java. Computerworld, page ??, September 08, 1995. CODEN CMPWAB. ISSN 0010-4841. URL http://

www.computerworld.com/.

Bozman:1995:NTS

[Boz95d] Jean S. Bozman. News: Telecom ’95: Sun to plunge into consumer market. Computerworld, page ??, October 06, 1995. CODEN CMPWAB. ISSN 0010-4841. URL http:/

/www.computerworld.com/.

Bercik:1997:CBN

[BP97] Bill Bercik and Sylvia Purcupile. Castanet & Bongo: No experience required. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 1997. ISBN 0-7821-2115-2. xix + 443 pp. LCCN TK5105.875.I57B467 1997. US$30. URL http://

www.sybex.com/cgi-bin/bookpg.
pl?2115back.html.

Bell:1998:JSa

[BP98] Doug Bell and Mike Parr. Java for Students. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1998. ISBN 0-13-858440-0. 448 pp. LCCN QA76.73.J38B45 1997. US$45.00.

Brown:1999:RRJ

[BP99] Kirk Brown and Daniel Petersen. Ready-to-Run Java 3D. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 1999. ISBN 0-471-31702-0. xii + 400 pp. LCCN QA76.73.J38B78 1999. US$44.99. With a foreword by James Gosling.

Bercik:19xx:CBNb

[BPxx] Bill Bercik and Sylvia Purcupile. Castanet & Bongo: No experience required (Japanese translation). Softbank Publishing, ??, Japan, 19xx. ISBN ???? ???? pp. LCCN ???? ???? URL

http://www.softbank.co.jp/
books/editor/os1/castanet/

. Japanese translation.

Brown:1997:JCA

[BR97] Marc H. Brown and Roope Raisamo. JCAT: Collaborative active textbooks using Java. Computer Networks and ISDN Systems, 29(14):1577– 1586, October 15, 1997. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

URL http://www.elsevier.
com/cas/tree/store/comnet/
sub/1997/29/14/1765.pdf.

Bramley:1996:TNR

[Bra96] Randall Bramley. Technical news & reviews: Maple V Release 4; Mathematica 3.0 from Wolfram; HotGAMS from NIST; Informant scans Web and e-mails you; more on the web. … IEEE Computational Science & Engineering, 3(3):80, Fall 1996. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). Discusses a Java-based front end for the Guide to Available Mathematical Software, with a pointer to

URL http://math.nist.gov/
HotGAMS/.

Bramley:1997:TNR

[Bra97a] Randall Bramley. Technology news & reviews: Chemkin software; OpenMP Fortran Standard; ODE toolbox for Matlab; Java products; Scientific WorkPlace 3.0. IEEE Computational Science & Engineering, 4(4):75–78, October/ December 1997. CODEN IS-CEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL

                                        http://dlib.computer.org/
                                        cs/books/cs1997/pdf/c4075.
                                        pdf.

Bramley:1997:TNRb

[Bra97b] Randall Bramley. Technology news & reviews: Chemkin software; OpenMP Fortran Standard; ODE toolbox for Matlab; Java products; Scientific WorkPlace 3.0. IEEE Computational Science & Engineering, 4(4):75–78, October/ December 1997. CODEN IS-CEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL

                                        http://dlib.computer.org/
                                        cs/books/cs1997/pdf/c4075.
                                        pdf.

Bradford:1998:AQJ

[Bra98] John Bradford. Answering questions on Java and ActiveX. Sun-Server, 12(1):13, January 1998. ISSN 1091-4986.

Bradenbaugh:1999:JAC

[Bra99] Jerry Bradenbaugh. JavaScript Application Cookbook. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1999. ISBN 1-56592-577-7. xiv + 462 pp. LCCN QA76.73.J39 B73 1999. US$34.95. URL http:

                                        //www.oreilly.com/catalog/
                                        jscook.

Breedlove:1996:WPU

[Bre96]  Bob Breedlove.  Web Pro-

gramming Unleashed: Java, HTML, CGI, PERL, Visual J++, JavaScript, VBScript, ActiveX. Sams.net, Indianapolis, IN 46268, USA, 1996. ISBN 1-57521-117-3 (softcover). xxvii + 1010 pp. LCCN QA76.625.B74 1996. US$44.99. Includes CD-ROM.

Boyles:1998:CCE

[BRF98] Michael Boyles, Rajeev Raje, and Shiaofen Fang. CEV: Collaborative environment for visualization using Java-RMI. In ACM [ACM98a], page ?? CODEN CPEXEI. ISSN 1040-3108. LCCN ????

URL http://www.cs.ucsb.
edu/conferences/java98/papers/
cev.pdf;   http://www.cs.
ucsb.edu/conferences/java98/

papers/cev.ps. Also published as Concurrency: Practice and Experience, 10(11–13), September 1998, CODEN CPEXEI, ISSN 1040-3108.

Bridges:1998:LEJ

[Bri98] Michael Bridges. Living without enum in Java. C/C++ Users Journal, 16(11):82–??, November 1998. CODEN CCUJEX. ISSN 1075-2838.

Briand:1999:BRC

[Bri99a] Marc Briand. Book review:

C++ for Java Programmers. C/C++ Users Journal, 17(12): ??, December 1999. CODEN CCUJEX. ISSN 1075-2838.

Brigner:1999:IPC

[Bri99b] Paul Brigner. Internet programming: Creating signed, persistent Java applets. Dr. Dobb’s Journal of Software Tools, 24(2):82, 84–88, February 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

                                        ddj.com/1999/9902/9902toc.
                                        htm;   http://www.ddj.com/
                                        ftp/1999/1999_02/persist.
                                        txt;   http://www.ddj.com/
                                        ftp/1999/1999_02/persist.
                                        zip.

BrinchHansen:1999:PEJ

[Bri99c] Per Brinch Hansen. Programming for Everyone in Java. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. ISBN 0-387-98683-9 (softcover), 1-4612-1514-5 (e-book). xviii + 238 pp. LCCN QA76.73.J38 B74 1999. US$34.95.

Brown:1995:NTP

[Bro95a] Jim Brown. News: 1996 technology planning survey. Network World, page ??, December 22, 1995. ISSN 0887-7661 (print), 1944-7655 (elec-

                                        tronic). URL http://www.
                                        iquest.net/index.html.

Brown:1995:NIW

[Bro95b] Jim Brown. News: Internet workshop draw numbers. Network World, page ??, November 22, 1995. ISSN 0887-7661 (print), 1944-7655 (elec-

                                        tronic). URL http://www.
                                        iquest.net/index.html.

Brown:1996:WRL

[Bro96] Mark R. Brown. Webmaster Resource Library: Three Best-Selling Comprehensive References: Special Editions of Using HTML, Using Java, and Using CGI. Que Corporation, Indianapolis, IN, USA, May 1996. ISBN 0-7897-0872-8. LCCN ???? US$150.

Brookshier:1997:JDR

[Bro97a] Dan Brookshier. JavaBeans Developer’s Reference. New Riders Publishing, Carmel, IN, USA, March 1997. ISBN 1-56205-716-2. xlvii + 733 pp. LCCN QA76.73.J38B76 1997. US$49.99. URL http://

merchant.superlibrary.com:
8000/catalog/hg/PRODUCT/PAGE/
15620/bud/1562057162.html.

Brosgol:1997:COF

[Bro97b] B. M. Brosgol. A comparison of the object-oriented features of Ada 95 and Java[TM]. In ACM [ACM97b], pages 213– 230. ISBN 0-89791-981-5. LCCN QA 76.73 A35 T75 1997. Theme title: Ada; the right choice for reliable software. ACM order number: 825970.

Brown:1997:JBC

[Bro97c] Peter Brown. Java 16-bit cards:

Sun, Siemens in Java/Smart Card pact. Electronic News, 43 (2178):14, July 28, 1997. CO-DEN ELNEAU. ISSN 0013-4937, 1061-6624.

Brookshier:1998:WAJ

[Bro98a] Dan Brookshier. Who’s Afraid Of Java? AP Professional, Boston, MA, USA, July 1998. ISBN 0-12-135455-5. LCCN ???? US$40.

Brosgol:1998:CAJ

[Bro98b] Ben Brosgol. A comparison of Ada and Java as a foundation teaching language. ACM SIGADA Ada Letters, 18(5):12– 38, September/October 1998. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Brosgol:1998:WAR

[Bro98c] Ben Brosgol. Why Ada rather than Java for academia? Ada User Journal, 19(2):135–??, ???? 1998. CODEN AUJOET. ISSN 1381-6551.

Brosgol:1998:CCF

[Bro98d] Benjamin M. Brosgol. A comparison of the concurrency features of Ada 95 and Java. ACM SIGADA Ada Letters, 18(6): 175–192, November/December 1998. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Brogden:1999:JPJa

[Bro99a] Bill Brogden. Java Programmer JDK 1.2 Exam Cram. Coriolis Group Books, Scottsdale, AZ, USA, 1999. ISBN ???? 350 pp. LCCN ???? US$30.

Brogden:1999:JPJb

[Bro99b] Bill Brogden. Java Programmer JDK 2 Exam Cram. Coriolis Group Books, Scottsdale, AZ, USA, 1999. ISBN ???? 1350 pp. LCCN ???? US$30.

Brookshier:1999:WAM

[Bro99c]  Daniel Brookshier.  Who’s

Afraid of More Java? Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 1999. ISBN 0-12-135455-5. 400 pp. US$39.95. URL http://

www.mkp.com/books_catalog/
0-12135-455-5.asp.

Brookshier:19xx:JDRa

[Broxxa] Dan Brookshier. JavaBeans Developer’s Reference. DaeRim Publishing, Seoul, Korea, 19xx. ISBN 89-7280-401-0 (??invalid ISBN??). LCCN ???? 25,000 won. URL http:/

/www.drbook.co.kr/daerim/
javabref.htm.

Brookshier:19xx:JDRb

[Broxxb] Dan Brookshier. JavaBeans Developer’s Reference. Softbank Publishing, ??, Japan, 19xx. ISBN 4-7973-0365-4. LCCN ???? 3400 yen. URL http:/

/www.softbank.co.jp/books/
editor/os1/JBean/.

Begole:1999:FCT

[BRS99] James Begole, Mary Beth Rosson, and Clifford A. Shaffer. Flexible collaboration transparency: supporting worker independence in replicated application-sharing systems. ACM Transactions on Computer-Human Interaction, 6(2):95– 132, June 1999. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

URL http://www.acm.org/

                                        pubs/citations/journals/tochi/
                                        1999-6-2/p95-begole/.

Bruce:1996:PPL

[Bru96] K. B. Bruce. Progress in programming languages. ACM Computing Surveys, 28(1):245– 247, March 1996. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic).

Bruckner:1997:PRP

[Bru97] Jared Bruckner. Product review: Penumbra Software’s Mojo: a great way to create Java applets without writing code. Web Apps Magazine, 1(1): 66–68, January/February 1997. ISSN 1090-2287.

Brutzman:1998:VRM

[Bru98] Don Brutzman. The Virtual Reality Modeling Language and Java. Communications of the ACM, 41(6):57–64, June 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL http://www.
                                        acm.org:80/pubs/citations/
                                        journals/cacm/1998-41-6/p57-
                                        brutzman/.

Bryan:1998:MJH

[Bry98] S. Bryan. Multimedia, Java, and HTML with Digital Chisel 3. In Anonymous [Ano98a], pages 48–?? ISBN 1-56484-134-0. LCCN ????

Beguelin:1996:TMD

[BS96] A. Beguelin and V. Sunderam. Tools for monitoring, debugging, and programming in PVM. In Bode [Bod96], page ?? ISBN 3-540-61779-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .E975 1996 Bar.

Bartlett:1997:CB

[BS97a] Neil Bartlett and Steve Simkin.

Castanet & Bongo. Ohmsha, ????, October 1997. ISBN 4-900718-44-0. ???? pp. LCCN ???? 4,000 yen.

URL http://www.ohmsha.co.
jp/data/books/contents/4-
900718-44-0.htm.

Bartlett:1997:CBP

[BS97b] Neil Bartlett and Steve Simkin.

Castanet and Bongo Programming FrontRunner. Coriolis Group Books, Scottsdale, AZ, USA, 1997. ISBN 1-57610-135-5. xix + 354 pp. LCCN TK5105.888.B37 1997. US$40.

URL http://www.coriolis.
com/Site/MSIE/Books/Ind/castbong.
htm.

Bellin:1997:CCB

[BS97c] David Bellin and Susan Suchman Simone. The CRC Card Book (Includes Java). Addison-Wesley, Reading, MA, USA, 1997. ISBN 0-201-89535-8. xxiii + 290 pp. LCCN QA76.64.B437

1997. URL http://cseng.aw.
com/bookdetail.qry?ISBN=0%
2D201%2D89535%2D8&ptype=0.

Buss:1997:PJH

[BS97d] Frank Buß and Stefan Schlöpke.

Programmation Java En 15 Heures. Edition Micro Application, Paris, France, 1997. ISBN 2-7429-0749-1. ???? pp. LCCN ???? 139 FF. URL http:/

                                        /www.microapp.com/. Includes

CD-ROM.

Buss:1997:PTJ

[BS97e] Frank Buß and Stefan Schlöpke. Programmier Training Java. DATA-Becker, ????, 1997. ISBN 3-8158-1303-4. 467 pp. LCCN ???? 49 DM. URL

                                        http://www.data-becker.de/
                                        buecher/buchinfos/1303.htm;
                                         http://www.data-becker.de/
                                        buecher/buecher.htm#softwaretr
                                        htmljava.

Barry:1998:CPS

[BS98a] Douglas Barry and Torsten Stanienda. Computing practices: Solving the Java object storage problem. Computer, 31(11):33–41, November 1998. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

                                        http://dlib.computer.org/
                                        co/books/co1998/pdf/ry033.
                                        pdf;   http://www.computer.
                                        org/computer/co1998/ry033abs.
                                        htm.

Bartlett:1998:PCB

[BS98b] Neil Bartlett and Steve Simkin. Programación en Castanet y Bongo. Anaya Multimedia, Madrid, Spain, February 1998. ISBN 84-415-0395-8. LCCN ???? 3495 Ptas. URL http:

                                        //www.anayamultimedia.es/
                                        cgi-multi/notas_prensa.pl?
                                        2321023&3&0&1.

Boerger:1998:DJV

[BS98c] E. Boerger and W. Schulte.

Defining the Java Virtual Machine as platform for provably correct Java compilation. Lecture Notes in Computer Science, 1450:17–??, 1998. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Borger:1998:DJV

[BS98d] Egon Börger and Wolfram Schulte. Defining the Java Virtual Machine as platform for provably correct Java compilation. Lecture Notes in Computer Science, 1450:17–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/
service/series/0558/bibs/
1450/14500017.htm;   http:
//link.springer-ny.com/link/
service/series/0558/papers/
1450/14500017.pdf.

Boussinot:1998:STB

[BS98e] Frédéric Boussinot and Jean-Ferdy Susini. The Sugar-Cubes tool box: a reactive Java framework. Software— Practice and Experience, 28 (14):1531–1550, December 10, 1998. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=5001669;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=5001669&PLACEBO=IE.pdf.

Boerger:1999:PFM

[BS99] E. Boerger and W. Schulte. A programmer friendly modular definition of the semantics of Java. Lecture Notes in Computer Science, 1523:353– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bellin:19xx:CCB

[BSxx] David Bellin and Susan Simone. The CRC Card Book (Includes Java). Addison-Wesley, Reading, MA, USA, 19xx. ISBN

????  LCCN ????  ????

                                        URL http://cseng.aw.com/
                                        bookdetail.qry?ISBN=0-201-
                                        89535-8&ptype=0.

Bartlett:1996:CEJ

[BSS96] Neil Bartlett, Steve Simkin, and Chris Stranc. Cutting-Edge Java Game Programming: Everything You Need to Create Interactive Internet Games with Java. Coriolis Group Books, Scottsdale, AZ, USA, November 1, 1996. ISBN 1-883577-98-5. xx + 523 pp. LCCN QA76.76.C672B37 1996. US$39.99; CDN$55.99.

                                        URL http://www.coriolis.
                                        com/site/msie/books/ind/jgdevw.

htm. Includes CD-ROM.

Bizarro:1998:JJLa

[BSS98a] Pedro Bizarro, Luís M. Silva, and João Gabriel Silva. JWarp: a Java library for parallel discrete-event simulations. In ACM [ACM98a], page ?? CODEN CPEXEI. ISSN 1040-3108. LCCN ????

URL http://www.cs.ucsb.
edu/conferences/java98/papers/
jwarp.pdf;   http://www.cs.
ucsb.edu/conferences/java98/
papers/jwarp.ps. Also pub-

lished as Concurrency: Practice and Experience, 10(11– 13), September 1998, CODEN CPEXEI, ISSN 1040-3108.

Bizarro:1998:JJLb

[BSS98b] Pedro Bizarro, Luís M. Silva, and João Gabriel Silva. JWarp: a Java library for parallel discrete-event simulations. Concurrency: Practice and Experience, 10(11–13):999–1005, September 1998. CODEN CPEXEI. ISSN 1040-3108.

URL http://www3.interscience.
wiley.com/cgi-bin/abstract?
ID=10050415;   http://www3.
interscience.wiley.com/cgi-
bin/fulltext?ID=10050415&
PLACEBO=IE.pdf. Special Issue:

Java for High-performance Network Computing.

Biddle:1998:JPB

[BT98] Robert Biddle and Ewan Tempero. Java pitfalls for beginners. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30 (2):48–52, June 1998. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Barksdale:1999:HJC

[BT99] Karl Barksdale and E. Turner.

HTML and JavaScript: concepts and activities. South-Western Educational Pub., Cincinnati, OH, USA, 1999. ISBN 0-538-68822-X. 192 pp. LCCN ????

Buchheit:1996:FFA

[Buc96a] Paul Buchheit. Flicker-free animation using Java. Linux Journal, 30:??, October 1996. CO-DEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (elec-

                                        tronic). URL http://www.
                                        linuxjournal.com/issue30/
                                        flicker.html.

Bucholtz:1996:SRS

[Buc96b] Chris Bucholtz. Software’s rising stars. Telephony, 231(8):50, 52, August 19, 1996. CODEN TLPNAS. ISSN 0040-2656.

Buchanan:1997:SDE

[Buc97] William Buchanan. Software development for engineers with C, Pascal, C++, Assembly Language, Visual Basic, HTML, JavaScript, and Java. Arnold, London, UK, 1997. ISBN 0-340-70014-9 (paperback). xiv + 674 pp. LCCN ????

Budd:1998:UOO

[Bud98] Timothy Budd. Understanding Object-Oriented Programming With Java. Addison-Wes-ley, Reading, MA, USA, February 1998. ISBN 0-201-30881-9. LCCN ???? ????

Buelens:1998:VEN

[Bue98] Kris Buelens. VM/ESA Network Computing with Java and NetRexx. IBM Redbooks, 11400 Burnet Road, Austin, TX 78758-3493, USA, April 1998. ISBN ???? LCCN

???? ???? URL http://www.
redbooks.ibm.com/SG245148/
jnr_bk.html.

Buhrmann:19xx:JCB

[Buhxx] Peter Buhrmann. Java 1.1 — Computer Based Training auf CD-ROM: Eine Einführung in die Programmierung. Addison-Wesley, Reading, MA, USA, 19xx. ISBN 3-8273-1236-1 (??invalid ISBN??). LCCN

???? ???? URL http://www.
addison-wesley.de/katalog/
item.ppml?id=00207.

Bruzzone:1999:PIC

[BUP99] Agostino G. Bruzzone, Adelinde Uhrmacher, and Ernest H. Page, editors. Proceedings of the 1999 International Conference on Web-based Modeling and Simulation: 1999 Western MultiConference, San Francisco, California, January 17– 20, 1999, Cathedral Hill Hotel, volume 31(3) of Simulation Series. Society for Computer Simulation International, San Diego, CA, USA, 1999. ISBN 1-56555-156-7. LCCN QA76.9.C65 I574 1999.

Burden:1996:FLJ

[Bur96a] Kevin Burden. Firing line: Java needs time to mature. Computerworld, 30(13):98, March 25, 1996. CODEN CMPWAB. ISSN 0010-4841. URL http://

www.computerworld.com/.

Burgess:1996:BJA

[Bur96b]  Jay Burgess.  The basics

of Java Animation. SIGS Books, 71 W 23rd St., New York, NY 10010, USA, 1996. ISBN 1-884842-64-X. 39 pp. LCCN QA76.73.J38B87

1996.  US$85.00.  URL

                                        http://www.sigs.com/books/
                                        wp_animation.html.

Burd:1997:JPD

[Bur97] Barry A. Burd. JBuilder For Dummies. IDG Books, San Mateo, CA, USA, October 1997. ISBN 0-7645-0078-3. 353 pp. LCCN QA76.73.J38B85 1997. US$24.99. URL http://www.

                                        dummies.com/cgi/fill_out_
                                        template.pl?book:0-7645-0078-
                                        3:book-Dummies+Press::u183717.

Includes CD-ROM.

Busse:1995:NCA

[Bus95] Torsten Busse. News: Comdex: AT&T announces electronic commerce, Web services. IDG International News Service, page ??, November 14, 1995.

Butcher:1999:EJS

[But99]  Paul Butcher.  Extending

JScript: Supporting constructors and arrays in JScript. Dr. Dobb’s Journal of Software Tools, 24(1):42, 44, 46, 48–49, January 1999. CO-DEN DDJOEB. ISSN 1044-

                                        789X. URL http://www.ddj.
                                        com/ddj/ftp/1999/1999_01/
                                        jstest.txt;   http://www.
                                        ddj.com/ddj/ftp/1999/1999_
                                        01/jstest.zip.

Bieber:1997:TSH

[BV97] M. Bieber and F. Vitali. Toward support for hypermedia on the World Wide Web. Computer, 30(1):62–70, January 1997. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Baerten:1998:UVJ

[BV98] Herbert Baerten and Frank Van Reeth. Using VRML and JAVA to visualize 3D algorithms in computer graphics education. Computer Networks and ISDN Systems, 30 (20–21):1833–1839, November 12, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/
store/comnet/sub/1998/30/
20-21/2013.pdf.

Bik:1997:JPJ

[BVG97] Aart J. C. Bik, Juan E. Villacis, and Dennis B. Gannon. javar: a prototype Java restructuring compiler. Concurrency: Practice and Experience, 9(11):1181–1191, November 1997. CODEN CPEXEI. ISSN 1040-3108. URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=13819;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=13819&PLACEBO=IE.pdf. Spe-

cial Issue: Java for computational science and engineering — simulation and modeling II.

Bik:1998:ELP

[BVG98] A. J. C. Bik, J. E. Villacis, and D. B. Gannon. Experience with loop parallelization in javar (A prototype restructuring compiler for Java). Lecture Notes in Computer Science, 1366:355–??, 1998. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barton:1998:RTP

[BW98a] John J. Barton and John Whaley. A real-time performance visualizer for Java. Dr. Dobb’s Journal of Software Tools, 23 (3):44, 46–48, 105, March 1998. CODEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/1998/1998_03/
                                        visualz.txt;   http://www.
                                        ddj.com/ftp/1998/1998_03/
                                        visualz.zip.

Buchi:1998:CTJ

[BW98b] Martin Büchi and Wolfgang Weck. Compound types for Java. ACM SIGPLAN Notices, 33(10):362–373, October 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Brezany:1999:ADR

[BW99] P. Brezany and M. Winslett. Advanced data repository support for Java scientific programming. Lecture Notes in Computer Science, 1593:1127– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Coffee:1997:HPJ

[C^{+}97a]  Peter Coffee et al.  How

to Program JavaBeans. Ziff-Davis Press, Emeryville, CA, USA, 1997. ISBN 1-56276-521-3. xiii + 367 pp. LCCN QA76.73.J38H69 1997. 79,00 DM; US$40. URL http://

merchant.superlibrary.com:
8000/catalog/hg/PRODUCT/PAGE/
15627/bud/1562765213.html;
http://www.mut.ch/profit.
htx?&x_pk=kd84mf7ms9dm. In-

cludes CD-ROM.

Cornell:1997:CVJ

[C^{+}97b] Gary Cornell et al. Core Visual J++. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, November 1997. ISBN 0-13-794520-5. ???? pp. LCCN ???? US$50. URL http://

www.prenhall.com/ptrbooks/
ptr_0137945205.html.

Cortes:1997:JLP

[C^{+}97c] José Luis Cortés et al. Java:

El Lenguaje de Programación de Internet. Editorial Marcombo, S.A., Barcelona, Spain, 1997. ISBN 84-267-1082-4. 168 pp. LCCN ???? 1900 Ptas.

Carrel:19xx:GSV

[C^{+}xxa] Marc Carrel et al. Getting Started with VisualAge for Java 1.0. IBM Redbooks, ????, 19xx. ISBN ???? ???? pp.

LCCN ???? URL http://www.
redbooks.ibm.com/SG242232/
javarpc.htm.

Castillo:19xx:JLP

[C^{+}xxb] Enrique Castillo et al. Java: Un lenguaje de Programación multiplataforma para Internet. Editorial Paraninfo, Madrid, Spain, 19xx. ISBN 84-283-2368-2. ???? pp. LCCN ???? 5200 Ptas. Includes CD-ROM.

Che:19xx:FJP

[C^{+}xxc] Chen Hung Che et al. Foundation of Java Programming for the WWW. Unalis Corporation, ??, Taiwan, 19xx. ISBN 957-22-2399-2. ???? pp. LCCN ???? 490 NT$.

Crawford:1997:BD

[CA97] Mike Crawford and Ted Alspach. Bongo For Dummies. IDG Books, San Mateo, CA, USA, June 1997. ISBN 0-7645-0176-

3. xxi + 359 pp. LCCN QA76.73.J38C73 1997. US$30.

URL http://www.idgbooks.
com/cgi-bin/db/fill_out_template.
pl?idgbook:0-7645-0176-3:
book-idg::uidg2862;   http:
//www.idgbooks.com/database/
book_result.msql?uid=uidg323&
isbn=0-7645-0176-3.

Cornell:1998:CPJ

[CA98] Gary Cornell and Kamal Abdali. CGI Programming With Java. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, April 1998. ISBN 0-13-287079-

7. 400 pp. LCCN ???? US$44.95. URL http://www.

prenhall.com/allbooks/ptr_
0132870797.html.

Carlisle:1998:WNE

[CAB^{+}98] Paul Carlisle, Mike Anderson, Ray Brown, James Mitchell, Anthony Ralston, Herb Grosch, Stewart A. Denenberg, and Rainer Schoenrank. Who needs educational software? Communications of the ACM, 41 (6):22–26, June 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:
80/pubs/articles/journals/
cacm/1998-41-6/p22-crawford/
p22-crawford.pdf;   http:/
/www.acm.org:80/pubs/citations/
journals/cacm/1998-41-6/p22-
carlisle/;   http://www.acm.
org:80/pubs/citations/journals/
cacm/1998-41-6/p22-crawford/

.

Cadenhead:1997:TYJ

[Cad97]  Rogers Cadenhead.  Teach

Yourself Java 1.1 Programming in 24 Hours. Sams.net, Indianapolis, IN 46268, USA, 1997. ISBN 1-57521-270-6. xviii + 384 pp. LCCN QA76.73.J38 C33 1997. US$25. URL http://

merchant.superlibrary.com:
8000/catalog/hg/PRODUCT/PAGE/
15752/bud/1575212706.html.

Includes CD-ROM.

Cadenhead:1998:AJH

[Cad98a] Rogers Cadenhead. Aprendiendo Java 1.1 en 24 Horas. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1998. ISBN 970-17-0085-6 (??invalid ISBN??). LCCN ????

???? URL http://www.mcl.

                                        com.mx/sotano/prentice.fm$
                                        Retrive?RecID=2752&html=HTMLRecord1.

Cadenhead:1998:TYJ

[Cad98b]  Rogers Cadenhead.  Teach

Yourself Java 1.2 Programming in 24 Hours 2nd ed. Sams.net, Indianapolis, IN 46268, USA, February 1998. ISBN 1-57521-391-5. LCCN ???? US$20.

Cadenhead:19xx:AEH

[Cadxxa] Rogers Cadenhead. Aprenda em 24 Horas Java 1.1. Editore Campus Ltda., Rua Sete de Setembro, 111/16, Rio de Janeiro - RJ - 20050-002, Brazil, 19xx. ISBN ???? ???? pp.

LCCN ????  ????  URL

                                        http://www.campus.com.br/
                                        catalogo/livro.cfm?cod_livro=
                                        20206.

Cadenhead:19xx:JGP

[Cadxxb] Rogers Cadenhead. Java 1.1 Le Grand Poche. Simon & Schuster Macmillan (France), Paris, France, 19xx. ISBN 2-7440-0302-6. ???? pp. LCCN ???? 99

                                        FF. URL http://www.ssm.fr/
                                        catalog/fiches/0302.htm.

Cadenhead:19xx:TPJ

[Cadxxc] Rogers Cadenhead. Le Tout en Poche Java. Simon & Schuster Macmillan (France), Paris, France, 19xx. ISBN 2-7440-0433-2. LCCN ???? 62

                                        FF. URL http://www.ssm.fr/
                                        catalog/fiches/0433.shtml.

Cagle:1996:BWA

[Cag96] Kurt Cagle. Building Web Applications With Netscape Navigator 2.1 and JavaScript. I D G Books Worldwide, Indianapolis, IN, USA, 1996. ISBN 0-7645-8008-6. 656 pp. LCCN ???? US$44.99.

Cairns:1999:ETJ

[Cai99]  Paul A. Cairns.  Enumer-

ated types in Java. Software—Practice and Experience, 29(3):291–297, March 1999. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=55000314;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=55000314&PLACEBO=IE.pdf.

Callaway:1999:ISS

[Cal99] Dustin R. Callaway. Inside servlets: server-side programming for the Java platform. Addison-Wesley, Reading, MA, USA, 1999. ISBN 0-201-37963-

5. xix + 434 pp. LCCN QA76.73.J38 C35 1999.

Canon:1996:JJ

[Can96] Maggie Canon. The Java jackpot. MacUser, 12(3):21–??, 1996. CODEN MCUSEY. ISSN 0884-0997.

Carlson:1995:UJJ

[Car95] Bob Carlson. Update: a jolt of Java could shake up the computing community. Computer, 28(11):81–82, November 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Cargill:1996:OJC

[Car96a] Tom Cargill. An overview of Java for C++ programmers. C++ Report, 8(2):46–49, February 1996. CODEN CRPTE7. ISSN 1040-6042.

Carl:1996:JFR

[Car96b] Jeremy Carl. Java fever rising. Web Week, 2(7):1, 58, June 1996. CODEN WEWEFP. URL

                                        http://www.webweek.com/.

Carpenter:1996:JAS

[Car96c] Joseph S. Carpenter. Java, applets, and Sun: the main course of business strategy for the information age. Senior thesis, Department of Economics, Colorado College, Colorado Springs, CO, USA, 1996. 140 pp.

Carr:1996:JVC

[Car96d] David F. Carr. Java’s virtues as corporate tool are laid out. Web Week, 2(7):25–26, June 1996. CODEN WEWEFP. URL

                                        http://www.webweek.com/.

Cardoso:1997:JGR

[Car97a] Carlos Cardoso. JavaScript: Guia de Referéncia Inteligente. Axcel Books do Brasil Editora, Rua da Glória, 344/10 And. - Ed. Glória Trade Center Glória - Rio de Janeiro/RJ, Brasil, 1997. ISBN 85-7323-016-9 (??Invalid checksum??).

????  pp.  LCCN ????

                                        R$ 35. URL http://www.
                                        axcelbooks.com.br/cgi-bin/

cgiwrap/axcel/pes-tit.pl?
livro=136;   http://www.visualnet.
com.br/shop/ciencia_moderna/
book23.htm.

Cartwright:1997:SJ

[Car97b] D. Cartwright. Sun JavaStation. LAN: the network solutions magazine, 5(2):27– 28, February 1997. CODEN LMAGEP. ISSN 0968-6320.

Carrez:19xx:SED

[Carxx] Christian Carrez. Structures et Données en Java, C++, et Ada. Masson Editeur, Masson, France, 19xx. ISBN 2-225-83007-X. ???? pp. LCCN ???? 325 FF. URL

http://www.masson.fr/cgi-
bin/bookf.pl?1:1:is=222583007X.

Casatelli:1996:JMS

[Cas96a] Christine Casatelli. Java management spec gets backing. RS/ Magazine, 5(6):10, June 1996. ISSN 1088-0844, 1061-0030.

Casatelli:1996:VTP

[Cas96b]  Christine Casatelli.  Visu-

alAge technology planned for Java. WebServer Magazine: For Managers of World Wide Web Sites, 1(3):10, September/ October 1996. ISSN 1087-4232. Discusses IBM’s VisualAge Java development tools.

Cavanaugh:1997:LDN

[Cav97]  Kevin Cavanaugh.  Lotus,

Domino, Notes and Java. In Unicode Consortium [Uni97a], page ?? ISBN ???? LCCN

????  URL http://www.
unicode.org/unicode/iuc10/
program.html.

Cary:1998:RJC

[CAZZ98] John R. Cary, David A. Alexander, Kevin Zollman, and Dean Zollman. Reusable Java components for physics education. Computers in Physics, 12(4):314–??, July 1998. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (elec-

tronic). URL https://aip.
scitation.org/doi/10.1063/
1.168714.

Cinderella:1998:OPG

[CB98] David Cinderella and Derek Ball. Official PowerJ: Getting Started. Ventana Press, Chapel Hill, NC, USA, January 1998. ISBN 1-85032-926-5. LCCN ???? US$45.

Carrel-Billiard:1998:PVJ

[CBA98] Marc Carrel-Billiard and John Akerley. Programming with VisualAge for Java. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, February 1998. ISBN 0-13-911371-1. xxix + 486 pp. LCCN QA76.65.C373 1998. US$49.95. URL http:

//www.prenhall.com/books/
ptr_0139113711.html.  In-

cludes CD-ROM.

Cohn:1997:WPVb

[CBJ97a] Mike Cohn, Jay Butten, and James Jory. Web-Programmierung mit Visual J++. Markt & Technik, Haar bei München, Germany, 1997. ISBN 3-8272-5211-3. 763 pp. LCCN ????

79,95 DM. URL http://www.
mut.ch/html/buchinfo.htx?
&x_pp=dJgxRqA9wlAYNijX&x_
az=10&x_pk=VpLvjnsx7b5V&x_
pa=B&x_pr=oS&x_bestell=25211&
x_stichwort=Cohn&x_sprache=
A&x_start=1&x_gefunden=5&
x_query=&x_lastpage=buchfund&
x_soft=&x_suchewo=A;   http:
//www.mut.ch/profit.htx?&
x_pk=kd84mf7ms9dm; http://
www.mut.com/;   http://www.
mut.com/listen/fmonatb.htm.

Cohn:1997:WPVa

[CBJ97b] Mike Cohn, Jay Butten, and James Jory. Web Programming with Visual J++. Sams.net, Indianapolis, IN 46268, USA, 1997. ISBN 1-57521-174-2. xxiii + 652 pp. LCCN QA76.625 .C64 1997. US$40. URL http://

merchant.superlibrary.com:
8000/catalog/hg/PRODUCT/PAGE/
15752/bud/1575211742.html.

CBS:1995:JIJ

[CBS95] CBS. Java isn’t just coffee anymore. news report, June

2, 1995. URL http://uttm.
com/drive/june_1995/june2.
html.

CBT:1997:AWA

[CBT97a] CBT Systems. Advanced Web Authoring: Java. CBT Systems, ????, June 1997. ISBN ???? LCCN ???? US$200. URL http://

www.clbooks.com/sqlnut/SP/
search/gtsumt?isbn=IT00215236.

CBT:1997:AJC

[CBT97b] CBT Systems. Anatomy of Java Classes. CBT Systems, ????, June 1997. ISBN ???? LCCN ???? US$200. URL

                                        http://www.cbtsys.com/cbt/
                                        curicula/courses/jav1105/
                                        jav1105.htm.

CBT:1997:IJL

[CBT97c] CBT Systems. Introduction to the Java Language. CBT Systems, ????, June 1997. ISBN ???? LCCN ???? US$200. URL

                                        http://www.cbtsys.com/cbt/
                                        curicula/courses/java03/java03.
                                        htm.

CBT:1997:JDO

[CBT97d] CBT Systems. Java and Distributed Objects. CBT Systems, ????, June 1997. ISBN ???? LCCN ???? US$200. URL

                                        http://www.cbtsys.com/cbt/
                                        curicula/courses/jav1111/
                                        jav1111.htm.

CBT:1997:JAI

[CBT97e] CBT Systems. Java Animation and Images. CBT Systems, ????, June 1997. ISBN ???? LCCN ???? US$200. URL

                                        http://www.cbtsys.com/cbt/
                                        curicula/courses/jav1112/
                                        jav1112.htm.

CBT:1997:OOP

[CBT97f] CBT Systems. Object-Oriented Principles and Java. CBT Systems, ????, June 1997. ISBN ???? LCCN ???? US$200. URL

                                        http://www.cbtsys.com/cbt/

curicula/courses/java04/java04.
htm.

CBT:1997:OJ

[CBT97g] CBT Systems. An Overview of Java. CBT Systems, ????, June 1997. ISBN ???? LCCN

????  US$200.  URL

http://www.cbtsys.com/cbt/
curicula/courses/java01/java01a.
htm.

CBT:1998:FJ

[CBT98a] CBT Systems. Features of Java. CBT Systems, ????, 1998. LCCN ???? US$200. URL

http://www.cbtsys.com/CBT/
curicula/courses/jav1107/
jav1107.htm.

CBT:1998:JB

[CBT98b] CBT Systems. Java Beans.

CBT Systems, ????, March 1998. ISBN ???? LCCN

????  US$200.  URL

http://www.cbtsys.com/cbt/
curicula/courses/jav1109/
jav1109.htm.

CBT:1998:JCB

[CBT98c] CBT Systems. Java Classes Bundle. Computer Literacy, ????, February 1998. ISBN 1-199-45619-5 (??invalid ISBN??). LCCN ???? US$200.

CBT:1998:JCDb

[CBT98d] CBT Systems. Java for C++ Developers. CBT Systems, ????, March 1998. ISBN ???? LCCN ???? US$200. URL

http://www.cbtsys.com/CBT/
curicula/courses/java02/java02.
htm.

CBT:1998:JCDa

[CBT98e] CBT Systems. Java for C++ Developers Bundle. Computer Literacy, ????, February 1998. ISBN 1-199-45628-4 (??invalid ISBN??). LCCN ???? US$200.

CBT:1998:JIN

[CBT98f] CBT Systems. Java: IBM Network Computing Framework Overview. CBT Systems, ????, 1998. LCCN

????  US$200.  URL

                                        http://www.cbtsys.com/CBT/
                                        curicula/courses/jav1115/
                                        jav1115.htm.

CBT:1998:JIV

[CBT98g] CBT Systems. Java IBM Visual Age For Java — Getting Started. CBT Systems, ????, 1998. LCCN

????  US$200.  URL

                                        http://www.cbtsys.com/CBT/
                                        curicula/courses/jav1114/
                                        jav1114.htm.

CBT:1998:JSN

[CBT98h] CBT Systems. Java Security, Networking, and the Internet. CBT Systems, ????, 1998. LCCN ???? US$200. URL

                                        http://www.cbtsys.com/CBT/
                                        curicula/courses/jav1108/
                                        jav1108.htm.

CBT:1998:JTT

[CBT98i]  CBT Systems.  Java Tips

and Techniques. CBT Systems, ????, 1998. LCCN

????  US$200.  URL

http://www.cbtsys.com/CBT/
curicula/courses/jav1113/
jav1113.htm.

CBT:1998:JUI

[CBT98j] CBT Systems. Java User Interface Programming. CBT Systems, ????, 1998. LCCN

????  US$200.  URL

http://www.cbtsys.com/cbt/
curicula/courses/jav1106/
jav1106.htm.

CBT:1999:AWA

[CBT99] CBT Systems. Advanced Web Authoring: Java. CBT Systems, ????, 1999. US$199.95. Interactive training course. Order number T10215236, Part number 655AW04.

Coombs:1996:JSJ

[CC96] Ted Coombs and Jason Coombs.

Java: Secrets of the Java Masters: Java Developer’s Journal. Java Training Ser.; Advanced. Sys-con publications, ????, June 1996. ISBN 1-886141-06-1. 480 pp. LCCN ???? US$38.95. URL http:

//www.sys-con.com///books/
index.html.

Chen:1997:JRD

[CC97] Marina Chen and James Cowie.

Java’s role in distributed collaboration. Concurrency: Practice and Experience, 9(6):509–519, June 1997. CODEN CPEXEI. ISSN 1040-3108. URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=13874;
http://www3.interscience.

wiley.com/cgi-bin/fulltext?
ID=13874&PLACEBO=IE.pdf.

Chaleat:1998:PHJ

[CC98] Philippe Chaleat and Daniel Charnay. Programmation HTML, Javascript. Eyrolles, Paris, France, 1998. ISBN 2-212-09024-2. 450 pp. LCCN ????

Chute:1996:MIR

[CCB96a] C. G. Chute, D. L. Crowson, and J. D. Buntrock. Medical information retrieval and WWW browsers at Mayo. In Waegemann [Wae96], pages 243–247. ISBN 0-9640667-7-7. LCCN R864.I68 1996. Two volumes.

Coombs:1996:NLS

[CCB96b] Ted Coombs, Jason Coombs, and Don Brewer. The Netscape LiveWire Sourcebook: Create and Manage a Java Based Web Site. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, June 1996. ISBN 0-471-15605-1. xii + 465 pp. LCCN TK5105.883.N48C66 1996. US$29.95. URL http:

//www.wiley.com/compbooks/
catalog/07/15605-1.html;
http://www.wiley.com/compbooks/
catalog/15605-1.htm.

Chang:1997:TAS

[CCCL97] Rong-Guey Chang, Cheng-Wei Chen, Tyng-Ruey Chuang, and Jenq Kuen Lee. Towards automatic support of parallel sparse computation in Java with continuous compilation. Concurrency: Practice and Experience, 9(11):1101–1111, November 1997. CODEN CPEXEI. ISSN 1040-3108. URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=13827;
http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=13827&PLACEBO=IE.pdf. Spe-

cial Issue: Java for computational science and engineering — simulation and modeling II.

Carpenter:1997:EHJ

[CCF^{+}97] Bryan Carpenter, Yuh-Jye Chang, Geoffrey Fox, Donald Leskiw, and Xiaoming Li. Experiments with HP Java. Concurrency: Practice and Experience, 9(6):633–648, June 1997. CODEN CPEXEI. ISSN 1040-3108. URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=13880;
http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=13880&PLACEBO=IE.pdf.

Caromel:1999:WJP

[CCF99] D. Caromel, S. Chaumette, and G. Fox. Workshop on Java for parallel and distributed computing. Lecture Notes in Computer Science, 1586:716– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carpenter:1998:JLS

[CCFL98] B. Carpenter, Y.-J. Chang, G. Fox, and X. Li. Java as a language for scientific parallel programming. Lecture Notes in Computer Science, 1366:340– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Copeland:1997:EPJ

[CCH97] G. S. Copeland, M. Conner, and G. Hambrick. Enterprise Programming with Java. ????, ????, October 1997. ISBN ???? ???? pp. LCCN ???? URL http:

//www.cs.tamu.edu/course-
info/cpsc681/fall97/welch/
copeland-ab.html.

Christiansen:1997:JIB

[CCI^{+}97] Bernd O. Christiansen, Peter Cappello, Mihai F. Ionescu, Michael O. Neary, Klaus E. Schauser, and Daniel Wu. Javelin: Internet-based parallel computing using Java. Concurrency: Practice and Experience, 9(11):1139–1160, November 1997. CODEN CPEXEI. ISSN 1040-3108. URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=13833;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=13833&PLACEBO=IE.pdf. Spe-

cial Issue: Java for computational science and engineering — simulation and modeling II.

Chiang:19xx:IJ

[CCYxxa] Frank Fu-Sun Chiang, H. S. Chen, and Jerry P. Yeh. Introduction to Java. GOTOP Information Inc., 5F, No.7, Lane 50, Sec.3 Nan Kang Road Taipei, Taiwan; Unit 1905,Metro Plaza Tower 2, No.223 Hing Fong Road, Kwai Chung, N.T., Hong Kong, 19xx. ISBN 957-641-756-

2. 452 pp. LCCN ???? 460 NT$.

URL http://oceaninfo.ife.
ntou.edu.tw/java/javaintr.

htm. Includes CD-ROM.

Chiang:19xx:JRB

[CCYxxb] Frank Fu-Sun Chiang, H. S.

Chen, and Jerry P. Yeh. Java Resource Book. GOTOP Information Inc., 5F, No.7, Lane 50, Sec.3 Nan Kang Road Taipei, Taiwan; Unit 1905,Metro Plaza Tower 2, No.223 Hing Fong Road, Kwai Chung, N.T., Hong Kong, 19xx. ISBN 957-641-757-

0. 493 pp. LCCN ???? 440 NT$.

URL http://oceaninfo.ife.
ntou.edu.tw/java/javareso.
htm.

Casanova:1997:UJN

[CD97] H. Casanova and J. Dongarra. The use of Java in the NetSolve project. In Sydow [Syd97], pages 791–796. ISBN 3-89685-550-6 (set), 3-89685-551-4 (vol. 1), 3-89685-552-2 (vol. 2), 3-89685-553-0 (vol. 3),3-89685-554-9 (vol. 4), 3-89685-555-7 (vol. 5), 3-89685-556-5 (vol. 6). LCCN Q183.9 .I46 1997. URL http://

www.netlib.org/utk/people/
JackDongarra/PAPERS/netsolve-
imacs897.ps;   http://www.
netlib.org/utk/people/JackDongarra/
pdf/netsolve-imacs897.pdf.

In cooperation with R.-P. Schafer, W. Rufeger, and H. Lehmann.

Cassady-Dorion:1997:ISJ

[CDBM97] Luke Cassady-Dorion, Matthew Brumbaugh, and Shirani Maheshwari. Industrial Strength Java. New Riders Publishing, Carmel, IN, USA, February 1, 1997. ISBN 1-56205-634-4. 842 pp. LCCN QA76.73.J38I53 1997. US$50. URL http://

                                        merchant.superlibrary.com:
                                        8000/catalog/hg/PRODUCT/PAGE/
                                        15620/bud/1562056344.html.

Casanova:1997:JAN

[CDD97] Henri Casanova, Jack Dongarra, and David M. Doolin. Java access to numerical libraries. Concurrency: Practice and Experience, 9(11): 1279–1291, November 1997. CODEN CPEXEI. ISSN 1040-3108. URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=13823;
                                         http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=13823&PLACEBO=IE.pdf. Spe-

cial Issue: Java for computational science and engineering — simulation and modeling II.

Codella:1998:SEJ

[CDF^{+}98] C. F. Codella, D. N. Dillenberger, D. F. Ferguson, R. D. Jackson, T. A. Mikalsen, and I. Silva-Lepe. Support for Enterprise JavaBeans in Component Broker. IBM Systems Journal, 37(4):502– ??, ???? 1998. CODEN IBMSA7. ISSN 0018-8670.

                                        URL http://www.almaden.

ibm.com/journal/sj/374/codella.
html.

Comar:1997:TGJ

[CDG97] C. Comar, G. Dismukes, and F. Gasperoni. Targeting GNAT to the Java Virtual Machine. In ACM [ACM97b], pages 149–

164. ISBN 0-89791-981-5. LCCN QA 76.73 A35 T75 1997. Theme title: Ada; the right choice for reliable software. ACM order number: 825970.

Chilimbi:1999:CCS

[CDL99] Trishul M. Chilimbi, Bob Davidson, and James R. Larus. Cache-conscious structure/class field reorganization techniques for C and Java. ACM SIG-PLAN Notices, 34(5):13–24, May 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.
acm.org:80/pubs/citations/
proceedings/pldi/301122/p13-

chilimbi/. See PLDI’99 proceedings [ACM99b].

Card:1996:ISI

[CEG96] Stuart Card, Stephen G. Eick, and Nahum Gershon, editors. IEEE Symposium on Information Visualization’96: October 28–29, 1996, San Francisco, California. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-8186-7668-X. LCCN TA1634 .I325 1996.

Celko:1996:SOJ

[Cel96a] J. Celko. Soothsayers for Oracle [Java and Oracle tuning]. DBMS, 9(8):20, 22, 24, 26, July 1996. CODEN DBMSEO. ISSN 1041-5173.

Celko:1996:SS

[Cel96b] Joe Celko. SQL for smarties. DBMS, 9(8):20–??, July 1, 1996. CODEN DBMSEO. ISSN 1041-5173.

Celestino:1998:BRJ

[Cel98]  R. J. Celestino.  Build-

ing reusable Java widgets. Linux Journal, 49:??, May 1998. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL

                                        ftp://ftp.ssc.com/pub/lj/
                                        listings/issue49/2627.tgz.

Cramer:1997:CJJ

[CFM^{+}97a] Timothy Cramer, Richard Friedman, Terrence Miller, David Seherger, Robert Wilson, and Mario Wolczko. Compiling Java just in time: Using runtime compilation to improve Java program performance. IEEE Micro, 17(3):36–43, May/ June 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //pascal.computer.org/mi/
                                        books/mi1997/pdf/m3036.pdf.

Cramer:1997:CLJ

[CFM^{+}97b] Timothy Cramer, Richard Friedman, Terrence Miller, David Seherger, Robert Wilson, and Mario Wolczko. Compiling Java just in time: Using runtime compilation to improve Java program performance. IEEE Micro, 17(3):36–43, May/ June 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//pascal.computer.org/mi/
books/mi1997/pdf/m3036.pdf.

Cardelli:1999:TMA

[CG99] Luca Cardelli and Andrew D.

Gordon. Types for mobile ambients. In ACM [ACM99a], pages 79–92. ISBN 1-58113-095-3.

LCCN ???? URL http://www.
acm.org:80/pubs/citations/
proceedings/plan/292540/p79-
cardelli/.

Choi:1999:EPM

[CGHS99] Jong-Deok Choi, David Grove, Michael Hind, and Vivek Sarkar. Efficient and precise modeling of exceptions for the analysis of Java programs. ACM SIGSOFT Software Engineering Notes, 24(5):21–31, September 1999. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Chan:1997:JPT

[CGI97] Mark C. Chan, Steven W.

Griffith, and Anthony F. Iasi. 1001 Java Programmer’s Tips. Jamsa Press, Las Vegas, NV, USA, January 1997. ISBN 1-884133-32-0. various pp. LCCN QA76.73.J38C46 1997. US$49.95. URL http://www.

jamsa.com/catalog/1001java/
1001java.htm.

Chabert:1998:JOH

[CGJ^{+}98] Annie Chabert, Ed Grossman, Larry S. Jackson, Stephen R. Pietrowiz, and Chris Seguin. Java object-sharing in Habanero. Communications of the ACM, 41(6):69–76, June 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.
acm.org:80/pubs/citations/
journals/cacm/1998-41-6/p69-
chabert/.

Choi:1999:EAJ

[CGS^{+}99] Jong-Deok Choi, Manish Gupta, Mauricio Serrano, Vugranam C. Sreedhar, and Sam Midkiff. Escape analysis for Java. ACM SIGPLAN Notices, 34(10):1–19, October 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

URL http://www.acm.org/
pubs/citations/proceedings/
oops/320384/p1-choi/.

Cornell:1996:CJa

[CH96] Gary Cornell and Cay S. Horstmann. Core Java. Sun-Soft Press, 2550 Garcia Avenue, Mountain View, CA 94043-1100, USA, April 1, 1996. ISBN 0-13-565755-5, 0-13-264516-5. xix + 622 pp. LCCN QA76.73.J38C67 1996. US$39.95. URL http://www.

prenhall.com/ptrbooks/ptr_
0135657555.html;   http://

www.sun.com/smi/ssoftpress/
books/Cornell/Cornell_TOC.

html. Includes CD-ROM. See review in [Gar96c, You97c]. Also available in Italian translation [CH97d].

Cornell:1997:JPS

[CH97a] Gary Cornell and Cay Horstmann. Java Programming With Symantec Café. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, December 1997. ISBN 0-13-270455-2. 750 pp. LCCN ???? US$44.95. URL http://www.

prenhall.com/ptrbooks/ptr_
0132704552.html. Check year:

1996 or 1997??

Cornell:1997:CJ

[CH97b] Gary Cornell and Cay S. Horstmann. Core Java. SunSoft Press Java series. SunSoft Press, 2550 Garcia Avenue, Mountain View, CA 94043-1100, USA, second edition, January 1997. ISBN 0-13-596891-7. xxxii + 766 pp. LCCN QA76.73.J38C67 1997. US$44.95. URL http:

//www.sun.com/books/books/
Cornell/Cornell.html. Avail-

able in German translation [CH97f]. See review in [You97c]. Includes CD-ROM.

Cornell:1997:CJa

[CH97c] Gary Cornell and Cay S. Horstmann. Core Java. SunSoft Press Java series. SunSoft Press, 2550 Garcia Avenue, Mountain View, CA 94043-1100, USA, second edition, January 1997. ISBN 0-13-596891-7. xxxii + 766 pp.

LCCN QA76.73.J38C67 1997.

                                        US$40.  URL http://www.
                                        prenhall.com/ptrbooks/ptr_
                                        0135968917.html;   http:/
                                        /www.sun.com/books/books/
                                        Cornell/Cornell.html. Avail-

able in German translation [CH97f]. Includes CD-ROM.

Cornell:1997:CJb

[CH97d] Gary Cornell and Cay S. Horstmann. Core Java. Mondadori Informatica, ??, Italy, 1997. ISBN ???? ???? pp. LCCN ???? 90000 Lire. URL

                                        http://www.horstmann.com/
                                        ;   http://www.mondadori.
                                        com/; mailto:gcornell@ix.

netcom.com. Italian translation of [CH97c].

Cornell:1997:CJc

[CH97e] Gary Cornell and Cay S. Horstmann. Core Java. ASCII Corporation, Tokyo, Japan, 1997. ISBN ???? ???? pp. LCCN ????

                                        URL http://75720.1524<at>
                                        compuserve.com;   http://
                                        www.ascii.co.jp/.

Cornell:1997:JBI

[CH97f] Gary Cornell and Cay S. Horstmann. JAVA bis ins Detail — Das Buch für Experten. Sun-Soft Press Java series. Verlag Heinz Heise, Helsdorfer Straße 7, D-30625, Hannover, Germany, 1997. ISBN 3-88229-087-0. 648 pp. LCCN ???? 99,80 DM. URL

                                        http://www.emedia.de/bin/
                                        bookshop?show=4553&id=;
                                        http://www.heise.de/buch/.

German translation of [CH97c]. Includes CD-ROM.

Cornell:1997:JPA

[CH97g] Gary Cornell and Cay S. Horstmann. Java Programming Algorithms. Microsoft Press, Bellevue, WA, USA, December 1997. ISBN 1-57231-690-X. ???? pp. LCCN ???? US$50.

Callaghan:1998:DVD

[CH98a] Michael Callaghan and Heiko Hirschmüller. 3-D visualisation of design patterns and Java programs in computer science education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3): 37–40, September 1998. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Chang:1998:CSD

[CH98b] Daniel Chang and Dan Harkey.

Client/server data access with Java and XML. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 1998. ISBN 0-471-24577-1 (paperback/CD-ROM). xxvii + 606 pp. LCCN QA76.9.C55 C45 1998. US$50. URL http:

//www.wiley.com/compbooks/
catalog/24577-1.htm.

Cornell:1998:CJBa

[CH98c] Gary Cornell and Cay S. Horstmann. Core Java, Bd.1 Grundlagen. Markt & Technik, Haar bei München, Germany, September 1998. ISBN 3-8272-9565-3 (??invalid ISBN??). LCCN ???? 90 DM.

Cornell:19xx:CJa

[CHxxa] Gary Cornell and Cay S. Horstmann. Core Java. Makron Books do Brasil Editora Ltda., Rua Tabapuã 1348, São Paulo SP - 04533-004 - Brasil, 19xx. ISBN 85-346-0839-3 (??invalid ISBN??). LCCN

????  ????  URL http:/

                                        /safe.tesla.com.br/makron/
                                        options/livro.asp?id=593.

Cornell:19xx:CJb

[CHxxb] Gary Cornell and Cay S. Horstmann. Core Java. Mondadori Informatica, ??, Italy, 19xx. ISBN 88-7131-791-2. LCCN ???? 90,000 L. URL http:

                                        //education.mondadori.com/
                                        books/SchBook.asp?Libro=123.

Cornell:19xx:CJc

[CHxxc] Gary Cornell and Cay S. Horstmann. Core Java. ASCII Corporation, Tokyo, Japan, second edition, 19xx. ISBN 4-7561-1719-8. LCCN

????  ????  URL http:/

                                        /www.ascii.co.jp/pb/book1/
                                        shinkan/detail/1322773.html.

Japanese translation.

Chabrow:1996:WSD

[Cha96a] E. R. Chabrow. Wall Street on the desktop [online brokerage on the World Wide Web]. Information Week, ????(570):61, 64, 66, March 11, 1996. CODEN IN-FWE4. ISSN 8750-6874.

Chan:1996:JCL

[Cha96b] Patrick Chan. Java Class Libraries. Addison-Wesley Longman, Reading, MA, USA, September 1996. ISBN 0-614-20264-7. ???? pp. LCCN ???? US$48.37. See review in [Gar96c].

Chapman:1996:JAA

[Cha96c] Randy Chapman. Java applet, awt and util class reference. Specialized Systems Consultants, P.O. Box 55549, Seattle, WA 98155, 1996. ISBN 0-916151-96-4. 20 pp. LCCN ???? US$7.00. URL http://

www.clbooks.com/sqlnut/SP/
search/gtsumt?isbn=0916151964.

Chapman:1996:JRCa

[Cha96d] Randy Chapman. Java Reference Card (Language, IO, Net). Specialized Systems Consultants, P.O. Box 55549, Seattle, WA 98155, December 1996. ISBN 0-916151-94-8 (softcover). ???? pp. LCCN ???? US$3.50.

Chapman:1996:JRCb

[Cha96e] Randy Chapman. Java Reference Cards (2-Card Set). Specialized Systems Consultants, P.O. Box 55549, Seattle, WA 98155, 1996. ISBN 0-916151-96-4. LCCN ???? US$7.

Charlesworth:1997:OJT

[Cha97a]  Susan Charlesworth.  Open

JBuilder in 21 Tagen. Markt & Technik, Haar bei München, Germany, November 1997. ISBN 3-87791-895-6. ???? pp. LCCN ???? 89,95

DM.  URL http://www.
mut.ch/html/buchinfo.htx?

                                        &x_pp=qvhiJEcdJrTbphZH&x_
                                        az=10&x_pk=Direkt&x_pa=&x_
                                        pr=Fr&x_bestell=91895&x_stichwort=
                                        jbuilder&x_sprache=D&x_start=
                                        1&x_gefunden=&x_query=&x_
                                        lastpage=buchfund&x_soft=&
                                        x_suchewo=S;   http://www.
                                        mut.ch/profit.htx?&x_pk=kd84mf7ms9dm;
                                        http://www.mut.com/.

Chauvet:1997:CAJ

[Cha97b] Jean-Marie Chauvet. CORBA, ActiveX et JavaBeans. Eyrolles, Paris, France, March 1997. ISBN 2-212-08950-

3.  LCCN ????  173

                                        FF. URL http://www.lmet.
                                        fr/cgi-bin/Nouveaute/en/n/
                                        2-212-08950-3.

Chauvet:1997:NCO

[Cha97c] Jean-Marie Chauvet. A new convergence of object models: Java is the likely peacemaker to bring the warring OT camps into accord. Object Magazine, 7 (2):46–49, April 1997. CODEN OBMAFO. ISSN 1055-3614.

Chamberlin:1998:LAB

[Cha98a] Gordon Chamberlin. Linux apprentice: Beginner’s guide to JDK. Linux Journal, 55:78–79, 82–83, November 1998. CO-DEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Chan:1998:TPP

[Cha98b] Mark Chan. 1001 Tips para Programmar con Java. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1998. ISBN 970-10-1762-5. LCCN ???? ????

Chan:1998:JDA

[Cha98c] Patrick Chan. The Java Developer’s Almanac. Addison-Wes-ley, Reading, MA, USA, March 1998. ISBN 0-201-37967-8. 976 pp. LCCN ???? US$18.96.

URL http://cseng.aw.com/
bookdetail.qry?ISBN=0-201-
37967-8&ptype=0.

Chaumette:1998:JDJ

[Cha98d] S. Chaumette. JEM-DOOS:

The Java/RMI based distributed objects operating system of the JEM project. Lecture Notes in Computer Science, 1505:135–??, 1998. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chaumette:1998:JJR

[Cha98e] S. Chaumette. JEM-DOOS:

The Java/RMI based distributed objects operating system of the JEM project. Lecture Notes in Computer Science, 1505:135–??, 1998. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chalk:1999:CCJ

[Cha99a] Peter Chalk. Conference corner: Java in the computing curricula. ACM SIGPLAN Notices, 34(12):9–11, December 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://ftp.
informatik.rwth-aachen.de/
dblp/db/indices/a-tree/c/
Chalk:Peter.html.

Chalk:1999:JCC

[Cha99b] Peter Chalk. Java in the computing curricula. ACM SIGPLAN Notices, 34(12):9–11, December 1999. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chan:1999:JDA

[Cha99c]  Patrick Chan.  The Java

Developer’s Almanac: 1999. The Java series. Addison-Wesley, Reading, MA, USA, March 1999. ISBN 0-201-43298-6. xi + 861 pp. LCCN QA76.73.J38C474 1999. US$19.95. URL http:/

/cseng.aw.com/bookdetail.
qry?ISBN=0-201-37967-8&ptype=
0.

Chapman:1999:JES

[Cha99d] Stephen J. Chapman. Java for Engineers and Scientists. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1999. ISBN 0-13-919523-8. ???? pp. LCCN QA76.73.J38C477 1999.

URL http://www.phptr.com/
ptrbooks/esm_0139195238.html.

Charles:1999:NUC

[Cha99e] John Charles. In the news: Ubiquitous computing uncorked: The international Y2K repair bill: News briefs. IEEE Software, 16(2):97–100, March/ April 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://dlib.computer.org/
so/books/so1999/pdf/s2097.

pdf. Discusses Sun’s plans for Jini.

Chauvet:1999:CTC

[Cha99f] Jean-Marie Chauvet. Composants et transactions: COM/ MTS, Corba/OTS, Java/EJB, XML. Collection dirigée par Guy Hervier. Eyrolles: Informatiques magazine, Paris, France, 1999. ISBN 2-212-09075-7. v + 274 pp. LCCN ????

Chang:19xx:DJD

[Chaxxa] Chen Ching Chang. In Deep With JavaScript Design. GO-TOP Information Inc., 5F, No.7, Lane 50, Sec.3 Nan Kang Road Taipei, Taiwan; Unit 1905,Metro Plaza Tower 2, No.223 Hing Fong Road, Kwai Chung, N.T., Hong Kong, 19xx. ISBN 957-641-820-8. ???? pp. LCCN ???? 460 NT$. URL

http://www.gotop.com/.

Chapman:19xx:JRCa

[Chaxxb] Randy Chapman. Java Reference Card (Applets, AWT, Utils). Specialized Systems Consultants, P.O. Box 55549, Seattle, WA 98155, 19xx. ISBN 0-916151-93-X. ???? pp. LCCN ???? US$4.

Charts:19xx:JBC

[Chaxxc] Bar Charts. Java Bar Chart.

IDG Books, San Mateo, CA, USA, 19xx. ISBN 1-57222-211-5. ???? pp. LCCN ????

Chauvet:19xx:CAJa

[Chaxxd] Jean-Marie Chauvet. CORBA, ActiveX y JavaBeans. Ediciones Gestión 2000, S.A., ????, 19xx. ISBN 84-8088-185-2. ???? pp. LCCN ???? 2985 Ptas.

URL http://www.intercom.
es/gestio2000/Novedad.html#
n_corba_activex.

Chen:1996:WVT

[Che96] Jiann-Liang Chen. WWW and VR technologies for network management applications. In IEEE [IEE96f], page ?? ISBN 0-8186-7728-7. LCCN QA76.9 .I4 1996. IEEE Computer Society Press Order Number PR07728. IEEE Order Plan Catalog Number 96TB100091.

Cherlin:1997:BLB

[Che97] Edward Cherlin. Breaking the language barrier: Java, with no legacy code, is pushing the pace of adoption of Unicode, the only global character set. UNIX Developer, 1(1):56–59, January/ February 1997. ISSN 1090-2279.

Chew:1998:JCC

[Che98] Frederick F. Chew. The Java/ C++ Cross Reference Handbook. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1998. ISBN 0-13-848318-3. xvii + 448 pp. LCCN QA76.73.J38C48 1998. US$40. URL http://

www.prenhall.com/ptrbooks/
ptr_0138483183.html.

Chen:1999:PJD

[Che99] E. Chen. Poison Java [data security]. IEEE Spectrum, 36 (8):38–43, August 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Chip:19xx:HJ

[Chixx] Chip. HTML a Java. Computer Press, ????, 19xx. ISBN ???? ???? pp. LCCN ???? 298 Kè.

URL http://vltava.cpress.
cz/cgi-bin/vltava/genbook.
cgi?book=prg402.

Choi:1997:JA

[Cho97a] Jeehwan Choi. Java Applets.

Power Book, ??, Korea, 1997. ISBN 89-8160-010-4 (??Invalid checksum??). ???? pp. LCCN ???? 10,000 won. URL ????.

Chorafas:1997:VPT

[Cho97b] Dimitris N. Chorafas. Visual Programming Technology: Covers VRML, Java. McGraw-Hill, New York, NY, USA, 1997. ISBN 0-07-011685-7 (softcover). xiv + 353 pp. LCCN QA76.65.C47 1997. US$49.00.

Christ:1999:JSP

[Chr99] Tilo Christ. Java, synchronization, and the Palmpilot. Dr. Dobb’s Journal of Software Tools, 24(7):58, 60–64, 66, July 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ftp/1999/1999_
07/syncbuild.txt;   http:/
/www.ddj.com/ftp/1999/1999_
07/syncbuild.zip.

Chung:1998:CJD

[Chu98a] David R. Chung. Component Java: Developing Domponents With JavaBeans and ActiveX (Java Masters). McGraw-Hill, New York, NY, USA, January 1998. ISBN 0-07-913690-7. ???? pp. LCCN ???? US$45.

Chute:1998:AMI

[Chu98b] Christopher G. Chute, editor. American Medical Informatics Association: a paradigm shift in health care information systems: Annual symposium — November 7–11, 1998, Orlando, FL, Journal of the American Medical Informatics Association: Symposium Supplement. Hanley and Belfus, Philadelphia, PA, USA, 1998. ISSN 1091-8280. LCCN ????

Chung:19xx:SCJ

[Chuxx] Whang You Chung. Search on CGI, Java and JavaScript. Ti Shan Bho Books, ??, Taiwan, 19xx. ISBN 957-23-0425-9. ???? pp. LCCN ???? 350 NT$. URL ????. Includes CD-ROM.

Chweh:1998:NPT

[Chw98] Crystal Chweh. New products: T3E-1200 outshines its older siblings; data migration facility ports to Irix; Voyager 2.0 blends Corba and Java; virtual flypaper; Labview leads ActiveX lifestyle; small space, big storage; Java performs ScriptEase act. IEEE Concurrency, 6(1):92–93, January/ March 1998. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL

http://dlib.computer.org/
pd/books/pd1998/pdf/p1092.
pdf.

Cabrera:1996:PFI

[CI96] L.-F. Cabrera and N. Islam, editors. Proceedings of the Fifth International Workshop on Object-Orientation in Operating Systems: October 27– 28, 1996, Seattle, Washington. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. CODEN PI-WOF6. ISBN 0-8186-7692-2, 0-8186-7693-0, 0-8186-7694-9 (microfiche). ISSN 1063-5351. LCCN QA 76.76 O63 I59 1996.

Chenney:1998:EDM

[CIF98] S. Chenney, J. Ichnowski, and D. Forsyth. Efficient dynamics modeling for VRML and Java. In Spencer [Spe98], pages 15– 24. ISBN 1-58113-022-8. LCCN ????

Cilibrasi:1996:WJR

[Cil96]  Rudi Cilibrasi.  What is

Java, really? Linux Journal, 30:??, October 1996. CO-DEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (elec-

tronic). URL http://www.
linuxjournal.com/issue30/
whatjava.html.

Cimino:1996:BSE

[Cim96] J. J. Cimino, editor. Beyond the superhighway: exploiting the Internet with medical informatics: proceedings, 1996 AMIA annual Fall Symposium, formerly SCAMC, October 26–30, 1996, Sheraton Washington Hotel, Washington, DC. Hanley and Belfus, Philadelphia, PA, USA, 1996. ISBN 1-56053-208-4. LCCN W 26.5 S978 1996.

Cizmar:1997:CSR

[Ciz97] Dawn Cizmar. CS student research experience applied to developing instructional technology. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(3): 120–126, September 1997. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Coors:1998:UVI

[CJ98] V. Coors and V. Jung. Using VRML as an interface to the 3D data warehouse. In Spencer [Spe98], pages 121–127. ISBN 1-58113-022-8. LCCN ????

Colby:1998:OPT

[CJJ^{+}98a] C. Colby, L. J. Jagadeesan, R. Jagadeesan, K. Läufer, and C. Puchol. Objects and processes in Triveni: a telecommunication case study in Java. In Proceedings of the 1998 Usenix Conference on Object Oriented Technologies and Systems, page ?? USENIX, Berkeley, CA, USA, 1998. To appear.

Colby:1998:OCT

[CJJ^{+}98b] Christopher Colby, Lalita Jategaonkar Jagadeesan, Radha Jagadeesan, Konstantin Laufer, and Carlos Puchol. Objects and concurrency in Triveni: a telecommunication case study in Java. In USENIX [USE98b], page ?? ISBN ???? LCCN

????  URL http://www.
usenix.org/publications/library/
proceedings/coots98/colby.
html.

Calder:1999:RTD

[CKH99] Brad Calder, Chandra Krintz, and Urs Hölzle. Reducing transfer delay using Java class file splitting and prefetching. ACM SIGPLAN Notices, 34 (10):276–291, October 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

URL http://www.acm.org/
pubs/citations/proceedings/
oops/320384/p276-calder/.

Cenciarelli:1997:SMJ

[CKRW97a] P. Cenciarelli, A. Knapp, B. Reus, and M. Wirsing. From sequential to multi-threaded Java: An event-based operational semantics. Lecture Notes in Computer Science, 1349:75– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cenciarelli:1997:SMT

[CKRW97b] P. Cenciarelli, A. Knapp, B. Reus, and M. Wirsing. From sequential to multi-threaded Java: An event-based operational semantics. Lecture Notes in Computer Science, 1349:75– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cenciarelli:1999:EBS

[CKRW99] P. Cenciarelli, A. Knapp, B. Reus, and M. Wirsing.

An event-based structural operational semantics of multithreaded Java. Lecture Notes in Computer Science, 1523:157– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ciancarini:1996:PAC

[CKTV96] Paolo Ciancarini, Andreas Knoche, Robert Tolksdorf, and Fabio Vitali. PageSpace: an architecture to coordinate distributed applications on the Web. Computer Networks and ISDN Systems, 28(7–11): 941–952, May 1, 1996. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/
browse/browse.cgi?year=1996&
volume=28&issue=7-11&aid=
1598.

Caromel:1998:TSC

[CKV98] Denis Caromel, Wilfried Klauser, and Julien Vayssière. Towards seamless computing and metacomputing in Java. Concurrency: Practice and Experience, 10(11–13):1043–1061, September 1998. CODEN CPEXEI. ISSN 1040-3108. URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=10050422;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=10050422&PLACEBO=IE.pdf.

Special Issue: Java for Highperformance Network Computing.

Cleaveland:1996:LEW

[CL96] Craig Cleaveland and Ted Lewis. Letters to the editor: What’s wrong with the binary critic? Computer, 29(8):10– 11, August 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). Debate on the merits of the Java programming language.

Chan:1997:JCLc

[CL97a] Patrick Chan and Rosanna Lee. The Java Class Libraries. GOTOP Information Inc., 5F, No.7, Lane 50, Sec.3 Nan Kang Road Taipei, Taiwan; Unit 1905,Metro Plaza Tower 2, No.223 Hing Fong Road, Kwai Chung, N.T., Hong Kong, July 1997. LCCN ???? ???? Chinese translation by Thi Shiang Workshop.

Chan:1997:JCLa

[CL97b] Patrick Chan and Rosanna Lee.

The Java Class Libraries: An Annotated Reference. Addison-Wesley, Reading, MA, USA, July 1997. ISBN 0-201-63458-9. xxvi + 1660 pp. LCCN QA76.73.J38C47 1997. US$52.67. URL http://www.

aw.com.

Chappell:1997:ADI

[CL97c] David Chappell and David S.

Linthicum. ActiveX demystified: It’s invasive. it’s ubiquitous. but what, exactly, is ActiveX? BYTE Magazine, 22(9): 56–62, 64, September 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic). Contains a comparison of Java and ActiveX.

Chauvet:1997:OMJ

[CL97d] Jean-Marie Chauvet and Marc Lerman. Object models and Java: Mixing and matching disparate object models. Dr. Dobb’s Journal of Software Tools, 22(12):20–22, 243, 99, 100, December 1997. CODEN DDJOEB. ISSN 1044-789X.

Cierniak:1997:JTO

[CL97e] Michał Cierniak and Wei Li. Just-in-time optimizations for high-performance Java programs. Concurrency: Practice and Experience, 9(11): 1063–1073, November 1997. CODEN CPEXEI. ISSN 1040-3108. URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=13828;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=13828&PLACEBO=IE.pdf. Spe-

cial Issue: Java for computational science and engineering — simulation and modeling II.

Cierniak:1997:OJB

[CL97f] Michał Cierniak and Wei Li. Optimizing Java bytecodes. Concurrency: Practice and Experience, 9(6):427–444, June 1997. CODEN CPEXEI. ISSN 1040-3108. URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=13869;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=13869&PLACEBO=IE.pdf.

Chan:1998:JCLb

[CL98a] Patrick Chan and Rosanna Lee. The Java Class Libraries:

java.applet, java.awt, java.beans,

volume 2. Addison-Wes-ley, Reading, MA, USA, second edition, 1998. ISBN 0-201-31003-1. xxix + 1682 pp. LCCN QA76.73.J38C47 1998. US$54.95. URL http:

//cseng.aw.com/bookdetail.
qry?ISBN=0-201-31003-1;
http://www2.awl.com/cgi-bin/
htsearch?restrict=&exclude=
&config=htdig&method=boolean&
format=builtin%2Dlong&words=
Chan%20AND%20Lee&page=2.

Chan:1998:JCLc

[CL98b] Patrick Chan and Rosanna Lee. The Java Class Libraries Poster, JDK 1.2 Beta 3. Addison-Wesley, Reading, MA, USA, 1998. ISBN 0-201-43292-

7. LCCN ???? US$4.95.

Chan:1999:JCLa

[CL99] Patrick Chan and Rosanna Lee. The Java Class Libraries Poster. Addison-Wesley, Reading, MA, USA, fourth edition, 1999. ISBN 0-201-43296-X. LCCN ???? US$9.95.

Chan:19xx:JCLa

[CLxxa] Patrick Chan and Rosanna Lee.

Java Class Libraries. Addison-Wesley, Reading, MA, USA, 19xx. ISBN 0-201-30011-7. LCCN ???? ???? Chinese translation.

Chan:19xx:JCLb

[CLxxb] Patrick Chan and Rosanna Lee. The Java Class Libraries. GOTOP Information Inc., 5F, No.7, Lane 50, Sec.3 Nan Kang Road Taipei, Taiwan; Unit 1905,Metro Plaza Tower 2, No.223 Hing Fong Road, Kwai Chung, N.T., Hong Kong, 19xx. ISBN ???? ???? pp. LCCN

                                        ???? URL http://www.gotop.

com.tw. Chinese translation by Thi Shiang Workshop.

Chan:19xx:JCLc

[CLxxc] Patrick Chan and Rosanna Lee. The Java Class Libraries Poster. Addison-Wesley, Reading, MA, USA, 19xx. ISBN 0-201-32575-6. LCCN ????

                                        US$5. URL http://cseng.aw.
                                        com/bookdetail.qry?ISBN=0%
                                        2D201%2D32575%2D6&ptype=0.

Cladingboel:1997:RJV

[Cla97a] Christopher Cladingboel. Real Java Virtual Machines: Hardware compilation and the Java Virtual Machine. Thesis (M.Sc.), Board of the Faculty of Mathematical Sciences, Oxford University, Oxford, UK, 1997. 107 pp.

Clark:1997:WJH

[Cla97b] Shelby Clark. Watch out, the Java’s hot. Contract Professional, 1(4):31, 32, 34, March/ April 1997. ISSN 1089-5825.

Clarke:1997:BRS

[Cla97c] G. Miller Clarke. Book review: Seven Java books — a whirlwind tour. C/C++ Users Journal, 15(5):37–39, May 1997. CODEN CCUJEX. ISSN 1075-2838.

Clausen:1997:JBO

[Cla97d] Lars R. Clausen. A Java bytecode optimizer using sideeffect analysis. Concurrency: Practice and Experience, 9(11):1031–1045, November 1997. CODEN CPEXEI. ISSN 1040-3108. URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=13838;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=13838&PLACEBO=IE.pdf. Spe-

cial Issue: Java for computational science and engineering

— simulation and modeling II.

Clavel:19xx:JS

[Claxx]  Gilles Clavel.  Java: La

synthèse. Masson Editeur, Masson, France, 19xx. ISBN 2-7296-0656-4. ???? pp. LCCN ???? 225 FF. URL

http://www.masson.fr/cgi-
bin/bookf.pl?1:1:is=2729606564.

Coad:1999:JMC

[CLD99] Peter Coad, Eric LeFebvre, and Jeff De Luca. Java Modeling in Color with UML. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1999. ISBN 0-13-011510-X. 221 pp. LCCN QA76.73.J38C652 1999. US$49.99. URL http:

//www.phptr.com/ptrbooks/
ptr_013011510X.html.

Clip:1998:NJO

[Cli98a] Paul Clip. Networks — Java’s object-oriented communications — remote method invocation is a powerful mechanism that enables flexible and scalable distributed communications among Java programs. BYTE Magazine, 23(2):53–54, February 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Clip:1998:SCJ

[Cli98b] Paul Clip. Servlets: CGI the Java way: These serveroriented Java classes offer better performance than CGI programs. they bring systemindependence, too. BYTE Magazine, 23(5):55–56, May 1998. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Chan:1998:JCLa

[CLK98] Patrick Chan, Rosanna Lee, and Doug Kramer. The Java Class Libraries, Volume 1: java.lang, java.io, java.math, java.net, java.text, java.util. Addison-Wesley, Reading, MA, USA, second edition, February 1998. ISBN 0-201-31002-3. 2080 pp. LCCN QA76.73.J38C47 1998. US$55.95. URL http://www.

                                        aw.com;   http://www2.awl.
                                        com/cseng/titles/0-201-31002-
                                        3/.

Chan:1999:JCLb

[CLK99a] Patrick Chan, Rosanna Lee, and Doug Kramer. The Java Class Libraries: java.io,

java.lang, java.math, java.net,
java.text, java.util, vol-

ume 1. Addison-Wesley, Reading, MA, USA, second edition, 1999. ISBN 0-201-31002-3. xxvi + 2050 pp. LCCN QA76.73.J38 C47 1998. US$59.99.

Chan:1999:JCLc

[CLK99b] Patrick Chan, Rosanna Lee, and Douglas Kramer. The Java Class Libraries, Second Edition, Volume 1: Supplement for Java 2 Platform, Standard Edition, v1.2. Addison-Wesley, Reading, MA, USA, 1999. ISBN 0-201-48552-4. xxix + 1157 pp. LCCN QA76.73.J38 C47 1998. US$34.95.

Chamberland:1998:IVJ

[CLR98] L. A. Chamberland, S. F. Lymer, and A. G. Ryman. IBM VisualAge for Java. IBM Systems Journal, 37(3):386– ??, ???? 1998. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

URL http://www.almaden.
ibm.com/journal/sj/373/chamberland.
html.

Christeson:1996:FGJ

[CM96] Brian Christeson and John D.

Mitchell. That first gulp of Java. Linux Journal, 30:??, October 1996. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http://www.

                                        linuxjournal.com/issue30/
                                        gulpjava.html.

Coad:1997:JDU

[CM97a] P. Coad and M. Mayfield. Javainspired design: use composition, rather than inheritance. American Programmer, 10(1): 22–31, January 1997. CODEN AMPRFD. ISSN 1048-5600.

Coad:1997:JDB

[CM97b] Peter Coad and Mark Mayfield. Java Design: Building Better Apps and Applets. Yourdon Press Computing Series. Yourdon Press, Upper Saddle River, NJ, USA, January 1, 1997. ISBN 0-13-271149-4. xiii + 238 pp. LCCN QA76.73.J38C65 1996. US$39.95. See [Wil97j].

Coad:1997:JDBa

[CM97c] Peter Coad and Mark Mayfield. Java Design: Building Better Apps and Applets. Yourdon Press Computing Series. Yourdon Press, Upper Saddle River, NJ, USA, 1997. ISBN 0-13-271149-4. xiii + 238 pp. LCCN QA76.73.J38C65 1996. US$39.95. URL http://www.

                                        oi.com/jd-book.htm. See re-

views in [Wil97i, Wil97j].

Coad:1997:JDBb

[CM97d] Peter Coad and Mark Mayfield. Java Design: Building Better Apps and Applets. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, November 1997. ISBN 4-89471-014-5. LCCN ???? 3,000 yen. URL http:/

/www.oi.com/jd-book.htm.

Coad:1999:JDB

[CM99] Peter Coad and Mark Mayfield. Java Design: Building Better Apps and Applets. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, second edition, 1999. ISBN 0-13-911181-6. 352 pp. LCCN QA76.73.J38C65 1999. US$44.99. URL http:

//www.phptr.com/ptrbooks/
ptr_0139111816.html.

Cohen:1996:PJF

[CMGR96] Shy Cohen, Tom Mitchell, Andres Gonzalez, and Larry Rodrigues. Professional Java Fundamentals. Wrox Press, Chicago, IL, USA, 1996. ISBN 1-86100-038-3. xix + 491 pp. LCCN QA76.73.J38P76 1996. US$35.00. URL http://www.

wrox.com/.

Coddington:1999:IIR

[CMH99] P. D. Coddington, J. A.

Mathew, and K. A. Hawick. Interfaces and implementations of random number generators for Java Grande applications. Lecture Notes in Computer Science, 1593:873–??, 1999. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1997:WBF

[CMM^{+}97] Zhikai Chen, Kurt Maly, Piyush Mehrotra, Praveen K. Vangala, and Mohammad Zubair. Web-based framework for distributed computing. Concurrency: Practice and Experience, 9(11):1175–1180, November 1997. CODEN CPEXEI. ISSN 1040-3108. URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=13839;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=13839&PLACEBO=IE.pdf. Spe-

cial Issue: Java for computational science and engineering

— simulation and modeling II.

Cohn:1996:JDRb

[CMMN96] Mike Cohn, Bryan Morgan, Michael Morrison, and Michael T. Nygard. Java Developer’s Reference. Howard W. Sams, Indianapolis, IN 46268, USA, November 1996. ISBN 1-57521-129-7. xxiv + 1258 pp. LCCN QA76.73.J38J365 1996. US$59.99. URL http:

//www.mcp.com/cgi-bin/bag?
isbn=1-57521-129-7&last=/
bookstore.

Czajkowski:1999:RCJ

[CMSvE99] Grzegorz Czajkowski, Tobias Mayr, Praveen Seshadri, and Thorsten von Eicken. Resource control for Java database extensions. In USENIX [USE99c], page ?? ISBN ???? LCCN QA76.64 .U84

1999.  URL http://www.
usenix.org/publications/library/
proceedings/coots99/czajkowski.
html.

Cobb:1998:UUJ

[Cob98] Joan Cobb. Use of Unicode and Java in a bibliographic application. In Unicode Consortium [Uni98b], page ?? ISBN ????

LCCN ???? URL http://www.
unicode.org/unicode/iuc13/
program.html.

Cobley:19xx:CGJ

[Cobxxa] Andy Cobley. The Complete Guide to Java. Computer Step, ????, 19xx. ISBN 1-874029-48-2. ???? pp. LCCN ????

UK£15. URL http://www.
computerstep.com/1874029482.
htm.

Cobley:19xx:VJE

[Cobxxb] Andy Cobley. Visual J++ in easy steps. Computer Step, ????, 19xx. ISBN 1-874029-75-X. ???? pp. LCCN

???? ???? URL http://www.
computerstep.com/187402975x.
htm.

Cockburn:1998:OOAb

[Coc98] Alistair Cockburn. Objectoriented analysis and design, part 2. C/C++ Users Journal, 16(6):??, June 1998. CODEN CCUJEX. ISSN 1075-2838.

Coffee:1996:HPJa

[Cof96] Peter Coffee. How to Program Java. Macmillan Computer Publishing, Indianapolis, IN, USA, October 1996. ISBN 0-614-20297-3. ???? pp. LCCN ???? US$39.99.

Cogswell:1997:JDP

[Cog97] Jeff Cogswell. Java Database Programming with JBuilder. MIS Press and M&T Books, P. O. Box 5277, Portland, OR 97208-5277, USA, Tel: (503)

282-5215 and M&T Publishing, Inc., 501 Galveston Drive, Redwood City, CA 94063, USA, 1997. ISBN 1-55851-512-7. 416 pp. LCCN ???? US$39.95.

Cohen:1996:CPJ

[Coh96a] Isaac Cohen. CGI/Perl et JavaScript: creation de pages HTML interactives. Eyrolles, Paris, France, 1996. ISBN 2-212-08918-X. vii + 294

pp.  LCCN ????  185

                                        FF. URL http://www.estp.

fr/eyrolles. Includes floppy disk.

Cohn:1996:JDRa

[Coh96b] Mike Cohn. Java Developer’s Reference. Macmillan Computer Publishing, Indianapolis, IN, USA, October 1996. ISBN 0-614-20293-0. ???? pp. LCCN ???? US$69.99.

Cohen:1997:CPJ

[Coh97a]  I. Cohen.  CGI/Perl y

JavaScript. Ediciones Gestión 2000, S.A., ????, 1997. ISBN 84-8088-143-7. ???? pp. LCCN ????

Cohen:1997:DJV

[Coh97b]  R. Cohen.  The defensive

Java Virtual Machine specification. Technical report, Computational Logic Inc., Austin, TX, USA, 1997.

Cohen:1997:GSF

[Coh97c]  Sarah Cohen.  Gemplus,

Schlumberger form Java Card Forum. Electronic News, 43 (2158):26, March 10, 1997. CO-DEN ELNEAU. ISSN 0013-4937, 1061-6624.

Cohen:1997:JC

[Coh97d] Yosef Cohen. Javascript Cookbook. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, April 1, 1997. ISBN 0-471-18145-5. 608 pp. LCCN QA76.73.J39C64 1997. US$39.95.

Cohn:1997:TYV

[Coh97e] Mike Cohn. Teach Yourself Visual Café 2 in 21 Days. Sams.net, Indianapolis, IN 46268, USA, 1997. ISBN 1-57521-303-6. xxx + 849 pp. LCCN QA76.73.J38C655 1997. US$40. URL http://

merchant.superlibrary.com:
8000/catalog/hg/PRODUCT/PAGE/
15752/bud/1575213036.html.

Cohen:1998:QJ

[Coh98] Sonya Freeman Cohen. Quest for Java. Communications of the ACM, 41(1):81–83, January 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org:80/pubs/citations/
journals/cacm/1998-41-1/p81-
cohen/.

Cohn:19xx:WPVa

[Cohxx] Mike Cohn. Web-Programmierung mit Visual J++. Markt & Technik, Haar bei München, Germany, 19xx. ISBN 3-8272-5211-3 (??invalid ISBN??). LCCN ???? 80 DM. URL http:/

                                        /www.mut.ch/html/buchinfo.
                                        htx?&x_pp=dJgxRqA9wlAYNijX&
                                        x_az=10&x_pk=VpLvjnsx7b5V&
                                        x_pa=B&x_pr=oS&x_bestell=
                                        25211&x_stichwort=Cohn&x_
                                        sprache=A&x_start=1&x_gefunden=
                                        5&x_query=&x_lastpage=buchfund&
                                        x_soft=&x_suchewo=A.

Cole:1995:NMD

[Col95]  Barb Cole.  News: More

database-to-Web products announced. Network World, page ??, December 08, 1995. ISSN 0887-7661 (print), 1944-7655 (electronic). URL http://

                                        www.iquest.net/index.html.

Cole:1996:NGG

[Col96a] Barb Cole. Novell gives Group-Wise Java-enabled client. Computerworld, 30(49):3, December 2, 1996. CODEN CMPWAB. ISSN 0010-4841.

Collinson:1996:JCT

[Col96b] Helen Collinson. Javascript can trace user tracks. Computers & Security, 15(2):126, ???? 1996. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

                                        tronic). URL https://www.
                                        sciencedirect.com/science/
                                        article/pii/0167404896893394.

Collinson:1996:SFC

[Col96c]  Helen Collinson.  Still a

few chinks in Java’s armour. Computers & Security, 15 (2):125, ???? 1996. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

                                        tronic). URL https://www.

sciencedirect.com/science/
article/pii/0167404896893333.

Computerworld:1995:NAJ

[Com95a] Computerworld staff. News:

Adobe joins Java licensees. Computerworld, page ??, December 08, 1995. CODEN CM-PWAB. ISSN 0010-4841.

Computerworld:1995:NIO

[Com95b] Computerworld staff. News: Internet object standard promised. Computerworld, page ??, December 15, 1995. CODEN CM-PWAB. ISSN 0010-4841. URL

http://www.computerworld.
com/.

Computerworld:1995:NPC

[Com95c] Computerworld staff. News:

Patching up CC:Mail. Computerworld, page ??, November 22, 1995. CODEN CMPWAB. ISSN 0010-4841.

Computerworld:1995:NSOa

[Com95d] Computerworld staff. News:

Sun offers classes in Java language. Computerworld, page ??, November 02, 1995. CO-DEN CMPWAB. ISSN 0010-

4841.  URL http://www.
computerworld.com/.

Computerworld:1995:NST

[Com95e] Computerworld staff. News:

Sun, Toshiba strike PDA deal. Computerworld, page ??, September 29, 1995. CODEN CM-PWAB. ISSN 0010-4841.

Computerworld:1995:NWJ

[Com95f] Computerworld staff. News: Will Java be IBM’s cup of tea? Computerworld, page ??, November 20, 1995. CO-DEN CMPWAB. ISSN 0010-

4841.  URL http://www.
computerworld.com/.

Postmodern:1995:PCA

[Com95g] Postmodern Computing. Post-Modern Computing announces Black Widow — an Object Request Broker that connects Web browsers to CORBA objects using Sun’s Java language. press release, December

11, 1995. URL http://www.
pomoco.com/bwdprs.html.

Comerford:1997:SET

[Com97a] R. Comerford. Software engineering [technology analysis and forecast]. IEEE Spectrum, 34 (1):65–69, January 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Comi:1997:JF

[Com97b] Luigi Comi. Java Flash! Apogeo srl, Viale Papiniano 38, 20123 Milano, Italy, 1997. ISBN 88-7303-197-8. 224 pp. LCCN ???? 16000 Lire. URL http:

//www.urra.it/285.html;
http://www.urra.it/apocat.
html.

Comerford:1999:SET

[Com99] Richard Comerford. Software engineering [technology 1999 analysis and forecast]. IEEE Spectrum, 36(1):84–87, January 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Conner:1997:OOP

[Con97a] Brook Conner. Object Oriented Programming in Java. Addison-Wesley, Reading, MA, USA, January 1997. ISBN 0-201-87013-4. ???? pp. LCCN ????

Conner:1997:OGJ

[Con97b] Nauman Conner. Overnight Guide to Java, CGI, Animation. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1997. ISBN 1-56592-283-2. ???? pp. LCCN ????

Coombs:1997:SIS

[Coo97a] Jason Coombs. Setting Up an Internet Site for Dummies. Softbank Publishing, ??, Japan, 1997. ISBN ???? ???? pp. LCCN ???? 1800 yen. URL http:/

/www.science.org/jasonc;
http://www.softbank.co.jp/
softbank/publishing/;   http:
//www.softbank.co.jp/softbank/
publishing/books/editer/os1/
internetsite/index.html.

Cooper:1997:POO

[Coo97b] James William Cooper. Principles of Object-Oriented Programming in Java 1.1: The Practical Guide to Effective, Efficient Program Design. Ventana Communications Group, Incorporated, Research Triangle Park, NC, USA, June 1997.

ISBN 1-56604-530-4. xx + 372 pp. LCCN QA76.64.C66 1997. US$39.99.

Cooper:1997:VBP

[Coo97c] James William Cooper. The Visual Basic Programmers Guide to Java for Windows 95 and Windows NT: your professional toolkit for object-oriented programming. Ventana Communications Group, Incorporated, Research Triangle Park, NC, USA, January 1, 1997. ISBN 1-56604-527-4. xxvi + 564 pp. LCCN QA76.73.J38C66 1997. US$39.99.

Cooper:1998:LJJ

[Coo98] Greg Cooper. Letters: Java JNI. Dr. Dobb’s Journal of Software Tools, 23(1):10, January 1998. CODEN DDJOEB. ISSN 1044-789X.

Corcoran:1995:EZS

[Cor95] Cate T. Corcoran. Ed Zander speaks out. WWW URL, November 1995. URL http://

                                        www.sun.com/sunworldonline/
                                        swol-11-1995/swol-11-fusion.
                                        zander.html.

Corcoran:1997:JWW

[Cor97a] Patrick Corcoran. JavaScript for the World Wide Web, (Visual QuickStart Guide). Peachpit Press, Inc., 1085 Keith Avenue, Berkeley, CA 94708, USA, second edition, September 1997. ISBN 0-201-69648-7 (softcover). xii + 195 pp. LCCN QA76.73.J39N44 1998. US$16.15.

Cornell:1997:SAC

[Cor97b]  G. Cornell.  Scripting Ac-

tiveX controls with VBScript. Web Techniques, 2(2):38–40, 42, 44, February 1997. CODEN WETEFA. ISSN 1086-556X.

Corbett:1998:CCM

[Cor98a] James C. Corbett. Constructing compact models of concurrent Java programs. ACM SIGSOFT Software Engineering Notes, 23 (2):1–10, March 1998. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Cornelius:1998:UCJ

[Cor98b] Barry Cornelius. Using CORBA and JDBC to produce three-tier systems. ACM SIGPLAN Notices, 33(4):44–??, April 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

URL http://www.rowan.edu/
sigplan/apr199~1.htm.

Coursey:1996:DLM

[Cou96a] David Coursey. Don’t let Microsoft hijack Java. Computerworld, 30(50):129, December 9, 1996. CODEN CMPWAB. ISSN 0010-4841.

Coursey:1996:TCC

[Cou96b]  David Coursey.  Texas?

Canada? Corel puts accent on Java. Computerworld, 30(41): 137, October 7, 1996. CODEN CMPWAB. ISSN 0010-4841. Notes that Corel will soon ship a Java-based WordPerfect suite.

Coursey:1997:CSW

[Cou97a] David Coursey. Challenge to Sun: Wall off Java. Computerworld, 31(17):113, April 28, 1997. CODEN CMPWAB. ISSN 0010-4841.

Coursey:1997:FSJ

[Cou97b] David Coursey. Fits and starts in the Java revolution. Computerworld, 31(10):105, March 10, 1997. CODEN CMPWAB. ISSN 0010-4841. Comments on IEEE 754 floating-point standard co-developer Jerome Coonan’s “A Note on Java Numerics” paper, which observes that Java’s requirement of identical floating-point results on all platforms imposes a run-time penalty on some platforms, notably IBM PowerPC and Intel x86, because it forces those systems to store intermediate results into memory and then immediately reload them, in order to prevent extra bits of precision in the PowerPC multiply-add instruction, or the 80-bit x86 floating-point registers, from affecting results. Babcock charges that Java is not platform neutral, as claimed by Sun, but has instead been ‘optimized for Sun SPARC and MIPS architectures.’, without understanding, or explaining, the reasoning behind the Java Virtual Machine floating-point design.

Coursey:1997:MWB

[Cou97c] David Coursey. The macro world beyond viruses. Computerworld, 31(22):28, June 2, 1997. CODEN CMPWAB. ISSN 0010-4841. Discusses the virus problem with reference to Java and ActiveX.

Coursey:1997:SS

[Cou97d] David Coursey. The Sun also sets. Computerworld, 31(23): 117, June 9, 1997. CODEN CM-PWAB. ISSN 0010-4841. Comments on Java’s commercial future.

Coursey:1997:WKS

[Cou97e] David Coursey. What keeps Sun awake at night? Computerworld, 31(16):129, April 21, 1997. CODEN CMPWAB. ISSN 0010-4841. Speculates about Java and the future of Sun Microsystems.

Courtney:1997:JSB

[Cou97f] Tom Courtney, editor. Java-SIG’s 100 Best Applets. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, July 1997. ISBN 0-471-18001-7 (softcover). viii + 231 pp. LCCN QA76.73.J38J383 1997. US$26.99.

URL http://www.wiley.com/
compbooks/catalog/18001-7.

htm. Includes CD-ROM.

Courtois:1997:JNC

[Cou97g] Todd Courtois. Java Networking and Communications. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, July 1997. ISBN 0-13-850454-7. 400 pp. LCCN QA76.73.J38C69 1997. US$39.95. URL http://

                                        www.prenhall.com/allbooks/
                                        ptr_0138504547.html.

Couch:1998:AJN

[Cou98] Justin Couch. Advanced Java Network Programming. Mc-Graw-Hill, New York, NY, USA, February 1998. ISBN 0-07-913741-5. LCCN ???? US$45.

Courtney:19xx:JSB

[Couxx] Tom Courtney, editor. Java-SIG’s 100 Best Applets. La Nuova Italia, Firenze, Italy, 19xx. ISBN 80-471-1800-1 (??Invalid checksum??). ???? pp. LCCN ???? 72000 L.

Cowlishaw:1996:N

[Cow96] Mike Cowlishaw. NetReXX. Web document, IBM Hursley, Hursley, UK, 1996. ???? pp.

                                        URL http://www.ibm.com/
                                        Technology/NetRexx/. Imple-

ments a Rexx-language compiler into Java.

Cowell:1997:EJF

[Cow97] John R. Cowell. Essential Java Fast: How to Write Object Oriented Software for the Internet in Java. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., May 1, 1997. ISBN 3-540-76052-0. x + 186 pp. LCCN QA76.64.C678 1997. US$24.95; US$39.95.

                                        URL http://www.springer-
                                        ny.com/catalog/np/jun97np/
                                        DATA/3-540-76052-0.html.

Cowie:1998:JPDa

[Cow98a] James Cowie. JTED: Parallel discrete-event simulation in Java. In ACM [ACM98a], page ?? CODEN CPEXEI. ISSN 1040-3108. LCCN ????

URL http://www.cs.ucsb.
edu/conferences/java98/papers/
jted.pdf;   http://www.cs.
ucsb.edu/conferences/java98/
papers/jted.ps. Also pub-

lished as Concurrency: Practice and Experience, 10(11– 13), September 1998, CODEN CPEXEI, ISSN 1040-3108.

Cowie:1998:JPDb

[Cow98b] James Cowie. JTED: parallel discrete-event simulation in Java. Concurrency: Practice and Experience, 10(11– 13):993–997, September 1998. CODEN CPEXEI. ISSN 1040-3108. URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=10050414;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=10050414&PLACEBO=IE.pdf.

Special Issue: Java for Highperformance Network Computing.

Cowell:1999:EJF

[Cow99] J. Cowell. Essential Java 2 Fast. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. ISBN 1-85233-071-6. 236 (est.) pp. US$29.95.

Cox:1995:NIA

[Cox95a] John Cox. News: IBM to add Web support to VisualAge. Network World, page ??, November 10, 1995. ISSN 0887-7661 (print), 1944-7655 (elec-

                                        tronic). URL http://www.
                                        iquest.net/index.html.

Cox:1995:NMS

[Cox95b] John Cox. News: Move slowly with Java, still in beta. Network World, page ??, December 11, 1995. ISSN 0887-7661 (print), 1944-7655 (elec-

                                        tronic). URL http://www.
                                        iquest.net/index.html.

Cox:1995:NNS

[Cox95c] John Cox. News: National Semiconductor to launch Javabased Web product. Network World, page ??, November 18, 1995. ISSN 0887-7661 (print), 1944-7655 (elec-

                                        tronic). URL http://www.
                                        iquest.net/index.html.

Cox:1995:NSR

[Cox95d] John Cox. News: SunSoft rolls out WorkShop Tool Set for Solaris 2.5. Network World, page ??, November 10, 1995. ISSN 0887-7661 (print), 1944-7655 (electronic). URL http://

                                        www.iquest.net/index.html.

Cox:1996:FMU

[Cox96a] Earl D. Cox. Fuzzy Models Using Java. Charles River Media, Inc., 403 VFW Drive, PO Box 417, Rockland, MA 02370, USA, November 1996. ISBN 1-886801-31-2. 650 pp. LCCN ???? US$49.95. URL http://www.

charlesriver.com/titles/fuzzyjava.
html.

Cox:1996:PMJ

[Cox96b] John Cox. Powersoft muscles into the Java market. Network World, 13(51):29, December 16, 1996. ISSN 0887-7661 (print), 1944-7655 (electronic).

Cox:1997:JPL

[Cox97] Ron Cox. JavaBeans Programming Lab. Prima Publishing, Rocklin, CA, USA, 1997. ISBN 0-7615-1047-8 (softcover). xxiv + 447 pp. LCCN QA76.73.J38 A45 1997. US$36.00. Includes CD-ROM.

Cockcroft:1998:SPT

[CP98] Adrian Cockcroft and Richard Pettit. Sun Performance and Tuning: Java and the Internet. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, second edition, February 1998. ISBN 0-13-095249-4. xxxvi + 587 pp. LCCN QA76.8.S86C63 1998. US$45. URL http:/

/www.fdds.com/books/books/
Cockcroft/Cockcroft.html;
 http://www.prenhall.com/
ptrbooks/ptr_0130952494.html.

Ciancarini:1997:JCC

[CR97] P. Ciancarini and D. Rossi. Jada — coordination and communication for Java agents. Lecture Notes in Computer Science, 1222:213–??, 1997. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Coscia:1999:JJT

[CR99] E. Coscia and G. Reggio. JTN: a Java-targeted graphic formal notation for reactive and concurrent systems. Lecture Notes in Computer Science, 1577:77– 97, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cravit:1997:LEJ

[Cra97] Matthew Cravit. Letter to the editor: Java response. Dr. Dobb’s Journal of Software Tools, 22(2):10, 12, February 1997. CODEN DDJOEB. ISSN 1044-789X.

Creek:19xx:JA

[Crexx] Walnut Creek. Java Applets. Walnut Creek, ????, 19xx. ISBN 1-57711-065-X. LCCN ???? US$30. URL http://

                                        www.cdrom.com/titles/prog/
                                        javaapp.htm.

Criscolo:1998:JQH

[Cri98] Mike Criscolo. Java Q and A: How do I queue Java threads? Dr. Dobb’s Journal of Software Tools, 23(10):127–129, October 1998. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

                                        ddj.com/ftp/1998/1998_10/
                                        jqa108.txt;   http://www.
                                        ddj.com/ftp/1998/1998_10/
                                        jqa108.zip.

Crothers:1995:NOS

[Cro95a] Brooke Crothers. News: Oracle set to outline Internet PC designs. InfoWorld, page ??, December 10, 1995. CODEN IN-WODU. ISSN 0199-6649. URL

http://www.infoworld.com/.

Crotty:1995:NNN

[Cro95b] Cameron Crotty. News: Net Navigator shatters static Web pages. MW (??), page ??, November 21, 1995.

Croft:1997:HJA

[Cro97] Brian A. Croft. HTML 4, Java 1.1, and ActiveX Web Publishing Unleashed. Sams.net, Indianapolis, IN 46268, USA, second edition, December 1997. ISBN 1-57521-363-X. ???? pp. LCCN ???? US$40.

Crowe:1999:CCJ

[Cro99] William L. Crowe. Comparing C++ and Java: a Java Switch-Board implementation. C/C++ Users Journal, 17(1):??, January 1999. CODEN CCUJEX. ISSN 1075-2838.

Chandy:1996:WDS

[CRS^{+}96] K. M. Chandy, A. Rifkin, P. A. G. Sivilotti, J. Mandelson, M. Richardson, W. Tanaka, and L. Weisman. A world-wide distributed system using Java and the Internet. In IEEE [IEE96n], page ?? ISBN 0-8186-7582-9. LCCN QA76.88.I52 1996. IEEE order plan catalog number 96TB100069. EEE Computer Society Press order number PR07582.

Chandy:1998:UALa

[CRS98a] K. Mani Chandy, Adam Rifkin, and Eve Schooler. Using announce-listen with global events to develop distributed control systems. In ACM [ACM98a], page ?? CODEN CPEXEI. ISSN 1040-3108. LCCN ????

URL http://www.cs.ucsb.
edu/conferences/java98/papers/
dcs.pdf;   http://www.cs.
ucsb.edu/conferences/java98/

papers/dcs.ps. Also published as Concurrency: Practice and Experience, 10(11–13), September 1998, CODEN CPEXEI, ISSN 1040-3108.

Chandy:1998:UALb

[CRS98b] K. Mani Chandy, Adam Rifkin, and Eve Schooler. Using announce-listen with global events to develop distributed control systems. Concurrency: Practice and Experience, 10(11–13):1021–1027, September 1998. CODEN CPEXEI. ISSN 1040-3108. URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=10050420;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=10050420&PLACEBO=IE.pdf.

Special Issue: Java for Highperformance Network Computing.

Ciancarini:1996:CTW

[CRV^{+}96] P. Ciancarini, D. Rossi, F. Vitali, A. Knoche, and R. Tolksdorf. Coordination technology for the WWW. In IEEE [IEE96m], page ?? ISBN 0-8186-7446-6. LCCN TS155.6 .W56 1996. IEEE Order Plan catalog number 96TB100058. IEEE Computer Society Press order number PR07445.

Cierniak:1997:PBP

[CS97a] Michał Cierniak and Suresh Srinivas. A portable browser for performance programming. Concurrency: Practice and Experience, 9(11):1243–1248, November 1997. CODEN CPEXEI. ISSN 1040-3108.

URL http://www3.interscience.
wiley.com/cgi-bin/abstract?
ID=13815;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=13815&PLACEBO=IE.pdf. Spe-

cial Issue: Java for computational science and engineering

— simulation and modeling II.

Cross:1997:MVJ

[CS97b] Jay Cross and Al Saganich. The Microsoft Visual J++ Sourcebook. Wiley Computer Publishers, New York, NY, USA, 1997. ISBN 0-471-17840-3. xiv + 539 pp. LCCN QA76.73.J38C76 1997. US$40. URL http:

//www.wiley.com/compbooks/
catalog/17840-3.htm.

Cartwright:1998:CGR

[CS98] Robert Cartwright and Guy L.

Steele, Jr. Compatible genericity with run-time types for the Java programming language. ACM SIGPLAN Notices, 33 (10):201–215, October 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Collins:1999:WLS

[CS99] William J. Collins and Yi Sun. Web labs for the Standard Template library and the Java generic library in a CS2 course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1): 174–178, March 1999. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Clements:19xx:JDA

[CSxx] Tom Clements and Tom Saulpaugh. Java OS Design and Architecture. Addison-Wesley, Reading, MA, USA, 19xx. ISBN 0-201-18393-5. ???? pp. LCCN ???? US$40.

Clossman:1998:JRD

[CSH^{+}98] Gray Clossman, Phil Shaw, Mark Hapner, Johannes Klein, Richard Pledereder, and Brian Becker. Java and relational databases (tutorial): SQLJ. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(2):500, June 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Choi:1997:DIC

[CSY97] T. Choi, E. Son, and K.-Y. Yu. DAVIC and the Internet convergence: a convergence architecture using WWW, JAVA, and CORBA. Lecture Notes in Computer Science, 1238:77– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cole:1996:PGW

[CT96] R. Cole and S. Tooker. Physics to go: Web-based tutorials for CoLoS physics simulations. In Iskander [Isk96], pages 3– ?? ISBN 0-7803-3348-9 (softbound), 0-7803-3349-7 (casebound), 0-7803-3350-0 (microfiche), 0-7803-3720-4 (CD-ROM). LCCN T 62 F76 1996. Three volumes. IEEE catalog number 96CH35946.

Christian:1998:DWB

[CT98] Wolfgang Christian and Aaron Titus. Developing Web-based curricula using Java physlets. Computers in Physics, 12(3): 227–??, May 1998. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (elec-

tronic). URL https://aip.
scitation.org/doi/10.1063/
1.168666.

Collberg:1998:MCR

[CTL98] Christian Collberg, Clark Thomborson, and Douglas Low. Manufacturing cheap, resilient, and stealthy opaque constructs. In ACM [ACM98b], pages 184–196. ISBN 0-89791-979-3. LCCN QA76.7 .A15

1998.  URL http://www.
acm.org:80/pubs/citations/
proceedings/plan/268946/p184-

collberg/. ACM order number: 549981.

Cuenca:1996:PMG

[Cue96] Mike Cuenca. The practical manager’s guide to useful JavaScript snippets. Computerworld, 30(34):74–75, August 1996. CODEN CMP-WAB. ISSN 0010-4841. Contains URLs for 7 sites with JavaScript code.

Culwin:1996:JFL

[Cul96a] Fintan Culwin. Java as a first language?! In Proceedings of 2nd All Ireland Computer on the teaching of computing, Dublin, 1996, page ?? ????, ????, 1996. ISBN ???? LCCN ????

                                        URL http://www.scism.sbu.
                                        ac.uk/~fintan/publications.
                                        html.

Culwin:1996:JC

[Cul96b] Fintan Culwin. Java in the curriculum. In 2nd SIGTOSE (Special Interest Group in the Teaching of Software Engineering), London, October 1996, page ?? ????, ????, 1996. ISBN ???? LCCN ????

                                        URL http://www.scism.sbu.
                                        ac.uk/~fintan/publications.
                                        html.

Culwin:1997:JCC

[Cul97a] Fintain Culwin. Java in the computing curriculum conference. ACM SIGPLAN Notices, 32(6):28–32, June 1997. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Culwin:1997:JC

[Cul97b] Fintan Culwin. Java in the HE curriculum. In ACM, editor, Workshop at ACM SIGCSE 1997, San Diego, CA, March 1997, page ?? ACM Press, New York, NY 10036, USA, 1997. ISBN ???? LCCN ????

URL http://www.scism.sbu.
ac.uk/~fintan/publications.
html.

Culwin:1997:RJC

[Cul97c] Fintan Culwin. A report on the ‘Java in the Computing Curriculum Conference’ held at South Bank University, London, on 4th February 1997. ACM SIGPLAN Notices, 32(6):28–32, June 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Culverhouse:1998:VJE

[Cul98a] Mark Culverhouse. Visual J++ Edicion Especial. Editorial Inforbooks, ????, 1998. ISBN 84-89660-88-3 (??invalid ISBN??). LCCN ???? ????

Culwin:1998:EJJ

[Cul98b] Fintan Culwin. Editorial — justifying Java? ACM SIG-PLAN Notices, 33(4):31–??, April 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.
rowan.edu/sigplan/apr199~
1.htm.

Culwin:1998:EPU

[Cul98c] Fintan Culwin. The essentials of practical user interface construction, using Java. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1998. ISBN ???? ???? pp. LCCN ????

                                        URL http://www.scism.sbu.
                                        ac.uk/~fintan/publications.
                                        html.

Culwin:1998:JOF

[Cul98d] Fintan Culwin. Java: An Object First Approach. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1998. ISBN 0-13-858457-5. 393 pp. LCCN QA76.73.J38 C86 1998. US$35. URL http://

                                        www.prenhall.com/ptrbooks/
                                        esm_0138584575.html;   http:
                                        //www.scism.sbu.ac.uk/~fintan/
                                        publications.html.

Culwin:1998:JGP

[Cul98e] Fintan Culwin. A Java GUI Programmers’ Primer. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, March 1998. ISBN 0-13-908849-0. 368 pp. LCCN QA76.73.J38C87 1998. US$35.00. URL http:

                                        //www.prenhall.com/books/
                                        esm_0139088490.html.

Culwin:1998:OPJ

[Cul98f] Fintan Culwin. Objektorientiertes Programmieren in Java. Markt & Technik, Haar bei München, Germany, May 1998. ISBN 3-8272-9564-5 (??invalid ISBN??). LCCN ???? 100 DM.

Culwin:1998:SSS

[Cul98g] Fintan Culwin. Special section:

selected papers from the Java in the computing curriculum conference. ACM SIGPLAN Notices, 33(4):31–33, April 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Culwin:19xx:JOF

[Culxxa] Fintan Culwin. Java an Object First Approach. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 19xx. ISBN 0-13-858457-X (??Invalid checksum??). ???? pp. LCCN ???? URL http://

www.scism.sbu.ac.uk/jfl/.

Culwin:19xx:JAP

[Culxxb] Fintan Culwin. A Java/AWT Programmer’s Primer. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 19xx. ISBN ???? ???? pp. LCCN

???? ???? URL http://www.
prenhall.com/allbooks/esm_
0139088490.html;   http://
www.scism.sbu.ac.uk/jfl/.

Caromel:1998:JFS

[CV98] Denis Caromel and Julien Vayssiere. A Java framework for seamless sequential, multithreaded, and distributed programming. In ACM [ACM98a], page ?? CODEN CPEXEI. ISSN 1040-3108. LCCN ????

URL http://www.cs.ucsb.
edu/conferences/java98/papers/
javapp.pdf;   http://www.
cs.ucsb.edu/conferences/java98/

                                        papers/javapp.ps. Also pub-

lished as Concurrency: Practice and Experience, 10(11– 13), September 1998, CODEN CPEXEI, ISSN 1040-3108.

Czajkowski:1998:JRA

[CvE98] Grzegorz Czajkowski and Thorsten von Eicken. JRes: a resource accounting interface for Java. ACM SIGPLAN Notices, 33 (10):21–35, October 1998. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Campione:1996:JTO

[CW96a] Mary Campione and Kathy Walrath. The Java Tutorial: Object-Oriented Programming for the Internet (Java Series). The Java Series. Addison-Wes-ley, Reading, MA, USA, 1996. ISBN 0-201-63454-6. xvii + 831 pp. LCCN QA76.64.C35 1996. US$39.76. URL http://www.

                                        aw.com/cp/javaseries.html;
                                         http://www.aw.com/cseng/
                                        titles/0-201-63454-6. Also

available in German translation [CW97]. See review in [Gar96c].

Carroll:1996:DPJ

[CW96b] E. Carroll and A. Wilson. Database programming with JDBC. Web Techniques, 1(7): 45–46, 48–50, October 1996. CODEN WETEFA. ISSN 1086-556X.

Campione:1997:JTO

[CW97] Mary Campione and Kathy Walrath. Das Java Tutorial: Objektorientierte Programmierung für das Internet. Addison-Wesley, Reading, MA, USA, 1997. ISBN 3-8273-1050-4. 784 pp. LCCN ????

99,90 DM. URL http://www.
addison-wesley.de/katalog/
item.ppml?textexpr=Java&id=

00125. German translation of [CW96a].

Campione:1998:JLT

[CW98] Mary Campione and Kathy Walrath. The Java Language Tutorial: Object-Oriented Programming for the Internet. Addison-Wesley, Reading, MA, USA, second edition, January 1998. ISBN 0-201-31007-4. xix + 964 pp. LCCN QA76.64.C35 1998. US$46. URL http://

www2.awl.com/cseng/titles/
0-201-31007-4/.

Campione:19xx:JTO

[CWxxa] Mary Campione and Kathy Walrath. Das Java-Tutorial: Objektorientierte Programmierung für das Internet. Addison-Wes-ley, Reading, MA, USA, 19xx. ISBN 3-8273-1050-4 (??invalid ISBN??). ???? pp. LCCN ????

URL http://java.sun.com/
people/kwalrath/;   http:/
/www.addison-wesley.de/.

Campione:19xx:JT

[CWxxb] Mary Campione and Kathy Walrath. The Java Tutorial. GOTOP Information Inc., 5F, No.7, Lane 50, Sec.3 Nan Kang Road Taipei, Taiwan; Unit 1905,Metro Plaza Tower 2, No.223 Hing Fong Road, Kwai Chung, N.T., Hong Kong, 19xx. ISBN ???? ???? pp. LCCN ????

                                        ???? URL http://www.gotop.

com.tw. Chinese translation by Thi Shiang Workshop.

Culverhouse:1996:SEU

[CWHP96] Mark Culverhouse, Clayton Walnum, Nelson Howell, and Greg Perry. Special Edition Using Visual J++. Que Corporation, Indianapolis, IN, USA, special edition, December 1996. ISBN 0-7897-0884-1. xxiii + 678 pp. LCCN QA76.73.J38C85 1996. US$50. URL http:

                                        //www.mcp.com/cgi-bin/bag?
                                        isbn=0-7897-0884-1&last=/
                                        bookstore.

Campione:1999:JTC

[CWHt99] Mary Campione, Kathy Walrath, Alison Huml, and the Tutorial Team. The Java tutorial continued: the rest of the JDK. Addison-Wesley, Reading, MA, USA, 1999. ISBN 0-201-48558-3. xii + 950 pp. LCCN TS1925.K5413 1999. US$39.95.

Cole:1998:UJD

[CWS98] David Cole, Roger Wainwright, and Dale Schoenefeld. Using Java to develop Web based tutorials. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30 (1):92–96, March 1998. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Carpenter:1998:HDPb

[CZF^{+}98a] Bryan Carpenter, Guansong Zhang, Geoffrey Fox, Xinying Li, and Yuhong Wen. HPJava: data parallel extensions to Java. Concurrency: Practice and Experience, 10(11–13):873–877, September 1998. CODEN CPEXEI. ISSN 1040-3108. URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=10050411;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=10050411&PLACEBO=IE.pdf.

Special Issue: Java for Highperformance Network Computing.

Carpenter:1998:HDPa

[CZF^{+}98b] Bryan Carpenter, Guansong Zhang, Geoffrey Fox, Xinying Li, and Yuhong Wen. HP-Java: Data parallel extensions to Java. In ACM [ACM98a], page ?? CODEN CPEXEI. ISSN 1040-3108. LCCN ????

URL http://www.cs.ucsb.
edu/conferences/java98/papers/
hpjava.pdf;   http://www.
cs.ucsb.edu/conferences/java98/
papers/hpjava.ps. Also pub-

lished as Concurrency: Practice and Experience, 10(11– 13), September 1998, CODEN CPEXEI, ISSN 1040-3108.

Carpenter:1998:TJE

[CZFL98] B. Carpenter, G. Zhang, G. Fox, and X. Li. Towards a Java environment for SPMD programming. Lecture Notes in Computer Science, 1470:659– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Daconta:1998:JJCa

[D^{+}98a] Michael C. Daconta et al. Java 1.2 and Javascript for C and C++ programmers. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, February 1998. ISBN 0-471-18359-8. ???? pp. LCCN QA76.73.J38 J3593 1997. US$55. URL http:

                                        //www.wiley.com/compbooks/
                                        catalog/18359-8.htm.

Daconta:1998:JJCb

[D^{+}98b] Michael C. Daconta et al. Java 1.2 and JavaScript for C and C++ Programmers/Microsoft Internet Explorer 4 Site Builder Toolkit. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 1998. ISBN ???? ???? pp. LCCN ???? US$94.98.

Davis:19xx:OIJ

[D^{+}xx] Stephen R. Davis et al. Opeta itsellesi Java Visual J++. Suomen Atk-kustannus Oy, PL 68, 02601 Espoo, Finland, 19xx. ISBN 951-762-480-8. 387 pp. LCCN ???? 295 Markka. URL http:/

                                        /www.satku.fi/;   http://
                                        www.satku.fi/4jvisual.htm.

Includes CD-ROM.

Darnell:1996:JQR

[DA96] Rick Darnell and Mike Afergan. JavaScript Quick Reference.

Que Corporation, Indianapolis, IN, USA, August 1, 1996. ISBN 0-7897-0869-8. xxvi + 182 pp. LCCN QA76.73.J39D37 1996. US$19.95. URL http://www. mcp.com/que. Also available in Spanish translation [Dar96].

Driggers:1997:DCE

[DAB97] Brent Driggers, Jay Alameda, and Ken Bishop. Distributed collaboration for engineering and scientific applications implemented in Habanero, a Javabased environment. Concurrency: Practice and Experience, 9(11):1269–1277, November 1997. CODEN CPEXEI. ISSN 1040-3108. URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=13841;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=13841&PLACEBO=IE.pdf. Spe-

cial Issue: Java for computational science and engineering — simulation and modeling II.

Daconta:19xx:JVC

[Dacxx] Mike Daconta. Java voor C/ C++ programmeurs. Addison-Wesley Nederland, ????, 19xx. ISBN 90-6789-759-0. ???? pp. LCCN ???? fl. 80. URL ????.

Dadda:19xx:USJ

[Dadxx] Roberto Dadda. Usare Subito Java E I Browser Java. Jackson Libri, via XXV Aprile, 39 - 20091 Bresso (MI), Italy, 19xx. ISBN ???? ???? pp. LCCN ???? 35000 L.

Dallmeier:1997:JVM

[Dal97] Matthias K. Dallmeier. Java Virtual Machine, Sprache, Konzept, Architektur. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1997. ISBN 3-930673-73-8. 209 pp. LCCN ???? 25 DM.

                                        URL http://www.ora.de/. In-

cludes CD-ROM.

Dalrymple:1998:LES

[Dal98] Robert A. Dalrymple. Linux in education: Sharing pedagogy with Java. Linux Journal, 56: 54–55, December 1998. CO-DEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Dalheimer:19xx:JVM

[Dalxx] Matthias K. Dalheimer. Java Virtual Machine, Sprache, Konzept. … O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 19xx. ISBN 3-930673-73-8 (??invalid ISBN??). LCCN ???? 29 DM.

                                        URL http://www.oreilly.
                                        de/german/essential/javavm/
                                        index.html.

Damasceno:1997:AJP

[Dam97] Americo Damasceno, Jr. Aprendendo Java — Programação Na Internet. Editora Erica, ??, Brazil, 1997. ISBN 85-7194-331-1. ???? pp. LCCN ???? R$ 33.

                                        URL ????.

Danesh:1996:JDK

[Dan96a] Arman Danesh. The Java Developer’s Kit. Linux Journal, 31:??, November 1996. CO-DEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Danesh:1996:JTC

[Dan96b] Arman Danesh. JavaScript 7-Tage-Crashkurs. Markt & Technik, Haar bei München, Germany, 1996. ISBN 3-8272-5176-1 (??invalid ISBN??). 611 pp. LCCN ???? 70 DM. URL

http://www.mut.ch/profit.
htx?&x_pk=kd84mf7ms9dm;
http://www.mut.com/. Ger-

man translation of [Dan96d].

Danesh:1996:TYJb

[Dan96c] Arman Danesh. Teach Yourself Javascript 1.1 in a Week. Howard W. Sams, Indianapolis, IN 46268, USA, second edition, November 1, 1996. ISBN 1-57521-195-5. xxiii + 587 pp. LCCN QA76.73.J39D356 1996. US$39.99.

Danesh:1996:TYJa

[Dan96d] Arman Danesh. Teach yourself JavaScript in a week. Howard W. Sams, Indianapolis, IN 46268, USA, April 1, 1996. ISBN 1-57521-073-8. xxiii + 543 + 9 pp. LCCN QA76.73.J39D36 1996. US$39.99. URL http://

merchant.superlibrary.com:
8000/catalog/hg/PRODUCT/PAGE/
15752/bud/1575210738.html;
 http://www.mcp.com/sams/
samspub/books/073-8/tjsfmfi.

htm#I1. Includes CD-ROM. Also available in German [Dan96b], Spanish [Dan96e], and Swedish [Dan97a] translations.

Danesh:1996:AJS

[Dan96e] Armen Danesh. Aprendiendo JavaScript en una Semana. Prentice Hall HispanoAmericana, Mexico, DF, Mexico / Nueva York, NY, USA, 1996. ISBN 968-880-640-4. 576 pp. LCCN ???? URL http:/

                                        /prentice.com.mx/;   http:
                                        //prentice.com.mx/libros/
                                        apren_java.html.  Spanish

translation of [Dan96d].

Danesh:1996:PJ

[Dan96f] Armen Danesh. Le programmeur JavaScript. Simon & Schuster Macmillan (France), Paris, France, 1996. ISBN 2-7440-0167-8. ix + 615 pp. LCCN ???? 199 FF. URL

                                        http://www.ssm.fr/.

Danesh:1997:IMJb

[Dan97a] Arman Danesh. Internett med Java och JavaScript. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1997. ISBN 0-13-286980-2. 188 pp. LCCN ????

                                        Skr 98. URL http://www.sn.
                                        no/prenhall/pren3000.htm;
                                         http://www.sn.no/prenhall/

pren3100.htm. Swedish translation of [Dan96d] by Magnus Gidlund.

Danesh:1997:IMJa

[Dan97b] Arman Danesh. Internett med Java og JavaScript. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1997. ISBN 0-13-271198-2. ???? pp. LCCN ????

NKr 149. URL http://www.
sn.no/prenhall/;   http:/
/www.sn.no/prenhall/oppg/
imjoj/index.htm;   http://
www.sn.no/prenhall/pren2000.
htm;   http://www.sn.no/prenhall/
pren2100.htm. Oversatt og

omarbeidet av K. Vilhelmsen [Translated and edited by K. Vilhelmsen]. Includes CD-ROM.

Danesh:1997:JIC

[Dan97c] Arman Danesh. JavaScript Interactive Course. Waite Group Press, Corte Madera, CA, USA, January 1, 1997. ISBN 1-57169-084-0. xx + 616 pp. LCCN QA76.73.J39D35 1997. US$39.99.

Danilko:1998:GJF

[Dan98] Brian Danilko. A graphical Java front end to C++ programs. C/C++ Users Journal, 16(1): ??, January 1998. CODEN CCUJEX. ISSN 1075-2838.

Danesh:19xx:J

[Danxx] Armen Danesh. JavaScript.

Simon & Schuster Macmillan (France), Paris, France, 19xx. ISBN ???? ???? pp. LCCN ???? 199 FF. URL http://

www.ssm.fr/.

Darnell:1996:JSI

[Dar96] Rick Darnell. JavaScript Soluciones Instantáneas. Prentice Hall HispanoAmericana, Mexico, DF, Mexico / Nueva York, NY, USA, 1996. ISBN 968-880-809-1. 224 pp. LCCN ???? URL

http://prentice.com.mx/;

                                        http://prentice.com.mx/libros/
                                        java_soluc.html.  Spanish

translation of [DA96].

Darcy:1998:BAI

[Dar98a] Joseph D. Darcy. Borneo 1.0: adding IEEE 754 floating point support to Java. Master of Science, Plan II, University of California, Berkeley, Dept. of Electrical Engineering and Computer Sciences, Berkeley, CA, USA, 1998. 131 pp. URL http:

                                        //www.sonic.net/~jddarcy/
                                        Borneo/borneo.pdf.

Darcy:1998:EJF

[Dar98b] Joseph D. Darcy. Evolving Java’s floating point support: The good, the bad, and the ugly. In MacKay and Johnson [MJ98], page ?? LCCN TK 5105.5 .C36 1998. URL http:

                                        //www.sonic.net/~jddarcy/
                                        Research/cascon.pdf.  31

slides.

Darcy:1998:WRI

[Dar98c] Joseph D. Darcy. Writing robust IEEE recommended functions in “100 pure Java^{T M}”. Report UCB/CSD-98-1009, University of California, Berkeley. Computer Science Division, Berkeley, CA, USA, August

                                        1998. URL http://sunsite.
                                        berkeley.edu/Dienst/UI/2.
                                        0/Describe/ncstrl.ucb/CSD-
                                        98-1009.

Darcy:1999:JEF

[Dar99a] J. D. Darcy. Java’s evolving floating-point support: The good, the bad, and the ugly. In Robert F. Enenkel, editor, CASCON 1998 Workshop Report, Numerical Computing: Compiler and Library Support, pages 2–10. IBM Center for Advanced Studies, Toronto, ON, Canada, 1999. URL https:

//www-927.ibm.com/ibm/cas/
publications/TR-74.165/n/

numcomp6.pdf. Technical report TR-74.165-n.

Darwin:1999:GDJ

[Dar99b] Ian Darwin. GUI development with Java. Linux Journal, 61:??, May 1999. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http://

www.linuxjournal.com/2673.
html.

Davis:1996:FJ

[Dav96a] Stephen R. Davis. Formation à Java. Microsoft Press France, ????, 1996. ISBN 2-84082-177-X. x + 413 pp. LCCN ????

198 FF. URL http://www.
estp.fr/eyrolles/ouvrages/

book003.html. French translation of [Dav96c]. Includes CD-ROM.

Davis:1996:JJ

[Dav96b] Stephen R. Davis. Java jetzt!:

das komplette Lernpaket: mit voll funktionsfahiger Version von Microsoft Visual J++ auf CD-ROM. Microsoft Press Germany, ????, 1996. ISBN 3-86063-021-0. 402 pp. LCCN ???? 79 DM. URL http://

194.120.227.99/prd.i/pgen/
press/FT16VFE47AS12P2E008MJ7W11XNLKLWL/

                                        product.html;   isbn=3-86063-
                                        021-0.

Davis:1996:LJN

[Dav96c] Stephen R. Davis. Learn Java Now. Microsoft Press, Bellevue, WA, USA, November 1, 1996. ISBN 1-57231-428-1. 448 pp. LCCN QA76.73.J38D38

1996.  US$39.95.  URL

                                        http://mspress.microsoft.
                                        com/order/bookpage.asp?title_
                                        id=284;   http://www.microsoft.

com/mspress/. Also available in Spanish translation [Davxxc].

David:1997:KAI

[Dav97a] Olaf David. A kernel approach for interactive-oriented model construction in Java. Concurrency: Practice and Experience, 9(11):1319–1326, November 1997. CODEN CPEXEI. ISSN 1040-3108. URL http:

                                        //www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=13836;
                                         http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=13836&PLACEBO=IE.pdf. Spe-

cial Issue: Java for computational science and engineering

— simulation and modeling II.

Davis:1997:LRJ

[Dav97b] Dwight B. Davis. Letter from Redmond: Java: To be or not to be. Datamation, 43(2):136–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Davis:1997:CGA

[Dav97c] M. Davis. Creating global applications with JDK 1.1. In Unicode Consortium [Uni97b], pages A11–?? ISBN ????

LCCN ???? URL http://www.
unicode.org/unicode/iuc11/
program.html.

Davis:1997:WMO

[Dav97d] M. Davis. Well-mannered objects: tips on porting C++ to Java. C++ Report, 9(4):14–21, April 1997. CODEN CRPTE7. ISSN 1040-6042.

Davis:1997:AJA

[Dav97e] Stephen R. Davis. Aprenda Java Agora. Editore Campus Ltda., Rua Sete de Setembro, 111/16, Rio de Janeiro - RJ - 20050-002, Brazil, 1997. ISBN ???? 432 pp. LCCN ???? R$ 52,20. URL http:

//www.campus.com.br/about-
us.htm;   http://www.campus.
com.br/catalogo/livro.cfm?
cod_livro=20147.  Includes

CD-ROM.

Davis:1997:AJY

[Dav97f] Stephen R. Davis. Aprenda Java Ya. McGraw-Hill Spain, ????, May 1997. ISBN 84-481-0976-7. ???? pp. LCCN ???? 5500 Ptas. URL http:/

/www.microsoft.com/mexico/
press/Program.htm.

Davis:1998:AFC

[Dav98a] Stephen R. Davis. Application Foundation Class (AFC) Programmer’s Guide. Microsoft Press, Bellevue, WA, USA, January 1998. ISBN 1-57231-732-9. xx + 487 pp. LCCN QA76.625 .D38 1998. US$40. URL

                                        http://mspress.microsoft.
                                        com/order/bookpage.asp?title_

id=1486. Includes CD-ROM.

Davis:1998:PMV

[Dav98b] Stephen R. Davis. Programming Microsoft Visual J++ 6.0. Microsoft Press, Bellevue, WA, USA, 1998. ISBN 1-57231-701-9. ???? pp. LCCN QA76.73.J38D39 1998. US$45.

Davies:1999:IJS

[Dav99] Richard J. Davies. Introductory Java for Scientists and Engineers. Addison-Wesley Longman, Reading, MA, USA, 1999. ISBN 0-201-39813-3. 312 pp. LCCN QA76.73.J38D35 1999.

                                        US$. URL http://www.awl-
                                        he.com/titles/11609.html;
                                         http://www.jscieng.co.uk/

.

Davis:19xx:FJa

[Davxxa] Stephen Davis. Formation 140 Java. Eyrolles, Paris, France, 19xx. ISBN ???? ???? pp.

                                        LCCN ???? URL http://www.
                                        estp.fr/eyrolles.

Davis:19xx:AJA

[Davxxb] Stephen R. Davis. Aprenda Java Agora. Editore Campus Ltda., Rua Sete de Setembro, 111/16, Rio de Janeiro - RJ - 20050-002, Brazil, 19xx. ISBN ???? LCCN ???? ???? URL

                                        http://www.campus.com.br/
                                        catalogo/livro.cfm?cod_livro=
                                        20147.

Davis:19xx:AJ

[Davxxc] Stephen R. Davis. Aprenda Java Ya. McGraw-Hill, New York, NY, USA, 19xx. ISBN ???? ???? pp. LCCN ???? 5500 Ptas. URL http:

//mexplaza.udg.mx/McGraw/.

Spanish translation of [Dav96c].

Deck:1997:IPS

[DB97] Stewart Deck and Dan Bricklin. Industry pioneer speaks his mind. Computerworld, 31(22): 28, June 2, 1997. CODEN CM-PWAB. ISSN 0010-4841. Visicalc inventor Bricklin calls Java a ‘write once, test many times’ language.

Davis:1998:JDR

[DBBE98] D. Davis, M. Bart, T. Bennett, and S. Edwards. A Java development and runtime environment for reconfigurable computing. Lecture Notes in Computer Science, 1388:43–??, 1998. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeJong:1998:CME

[DBK98] Moses DeJong, Brian Bailey, and Joseph A. Konstan. Creating a multimedia extension for Tcl using the Java Media Framework. In USENIX [USE98d], page ?? ISBN 1-880446-98-7. LCCN QA76.73.T44 T34

1998. URL http://db.usenix.
org/publications/library/
proceedings/tcl98/dejong.
html.

Dyson:1997:NPI

[DC97] Peter Dyson and Pat Coleman. NetWork Press: Intranets mit Windows NT 4. Sybex Verlag GmbH, Erkrather Straße 345-349, 40231 Düsseldorf, Germany, 1997. ISBN 3-8155-7260-6. xxx + 809 pp. LCCN ???? 129 DM. URL http:

//www.sybex.de/;   http:/
/www.sybex.de/bestell/tp/

tp07260.htm. Includes CD-ROM.

Daconta:1996:JCC

[DD96] Michael C. Daconta and Mike Daconta. Java for C/C++ programmers. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, April 1, 1996. ISBN 0-471-15324-9 (paperback). vii + 443 pp. LCCN QA76.73.J38D33 1996. US$39.95, CDN$48.95.

URL http://www.wiley.com/
compbooks/catalog/07/15324-
9.html.

Deitel:1997:CJTb

[DD97a] Harvey M. Deitel and Paul J. Deitel. Complete Java Training Course. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, January 1997. ISBN 0-13-842139-0 (softcover). ???? pp. LCCN ???? US$75.00; US$113.

Deitel:1997:CJTa

[DD97b] Harvey M. Deitel and Paul J. Deitel. A Complete Java Training Course: The Ultimate Cyber Classroom. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, May 1997. ISBN 0-13-841958-2 (softcover). 1056 pp. LCCN ???? US$113. URL http://

www.prenhall.com/ptrbooks/
ptr_0138419582.html.  In-

cludes CD-ROM.

Deitel:1997:JHPa

[DD97c] Harvey M. Deitel and Paul J.

Deitel. Java: how to program: with an introduction to Visual J++. How to program series. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1997. ISBN 0-13-263401-5, 0-614-20312-0. liii + 1056 pp. LCCN QA76.73.J38D45 1996. US$45. See review [Hay96v].

Deitel:1997:JHPb

[DD97d] Harvey M. Deitel and Paul J.

Deitel. Java: How to Program: With an Introduction to Visual J++. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, March 1997. ISBN 0-13-632589-0. liii + 1114 pp. LCCN QA 76.73 J38 D45 1997.

US$45.  URL http://www.
prenhall.com/ptrbooks/esm_
0136325890.html.

Deitel:1997:JMC

[DD97e] Harvey M. Deitel and Paul J.

Deitel. Java Multimedia Cyber Classroom. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, January 1997. ISBN 0-13-271974-6. ???? pp. LCCN ???? US$69.95. URL http://www.

deitel.com;   http://www.
prenhall.com/. See review

[Die97].

Deitel:1997:JMCa

[DD97f] Harvey M. Deitel and Paul J. Deitel. Java Multimedia Cyber Classroom. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, January 1997. ISBN 0-13-271974-6. ???? pp. LCCN ???? US$69.95. URL http://www.

                                        prenhall.com/ptrbooks/ptr_
                                        0132719746.html. See review

in [Die97].

Deitel:1997:JMCb

[DD97g] Harvey M. Deitel and Paul J. Deitel. Java Multimedia Cyber Classroom. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, second edition, November 1997. ISBN 0-13-905589-4. LCCN ???? ???? URL http:

                                        //www.prenhall.com/books/
                                        ptr_0139055894.html.

Doberenz:1997:JPI

[DD97h] Walter Doberenz and Uwe Druckenmüller. Java: Programmierung interaktiver WWW-Seiten. Carl Hanser, München, Germany, 1997. ISBN 3-446-18854-1. 494 pp. LCCN ???? 69 DM. URL http:

                                        //www.hanser.de/computer/
                                        buecher/do18854.htm.  In-

cludes CD-ROM.

Doolin:1997:JCL

[DD97i] David M. Doolin and Jack Dongarra. JLAPACK — compiling LAPACK Fortran to Java, phase 1. Technical report CS-97-367, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, 1997. URL http://

www.netlib.org/utk/people/
JackDongarra/PAPERS/f2jpaper.
ps;   http://www.netlib.org/
utk/people/JackDongarra/pdf/
f2jpaper.pdf.

Deitel:1998:JHP

[DD98a] Harvey M. Deitel and Paul J.

Deitel. Java How to Program. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, second edition, August 1998. ISBN 0-13-899394-7 (softcover). ???? pp. LCCN QA76.73.J38D45 1998. US$45.00. URL http://www.

prenhall.com/ptrbooks/esm_
0138993947.html.

Doberenz:1998:JPI

[DD98b] Walter Doberenz and Uwe Druckenmüller. Java: Programmierung interaktiver WWW-Seiten. Carl Hanser, München, Germany, second edition, 1998. ISBN 3-446-19047-3. 448 pp. LCCN ???? 59 DM; 54,80 SFr; 431,- ÖS. URL http://

www.buchhandel.de/cgi-bin/
bksuche.exe?aktion=full&opacdb=
VLB+Buchkatalog&lang=german&
uid=12573-02101997-16184600&
II=1268266&vcaller=12573&
aktionsort=on&sortby=KT;
http://www.hanser.de/computer/
buecher/do19047.htm.  In-

cludes CD-ROM.

Deitel:19xx:CPJ

[DDxx] Harvey M. Deitel and Paul J.

Deitel. Como Programar en Java. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 19xx. ISBN 970-17-0044-9 (??invalid ISBN??). LCCN ????

                                        ???? URL http://www.mcl.
                                        com.mx/sotano/prentice.fm$
                                        Retrive?RecID=2727&html=HTMLRecord1.

Dean:1996:VOC

[DDG^{+}96] Jeffrey Dean, Greg DeFouw, David Grove, Vassily Litvinov, and Craig Chambers. Vortex: an optimizing compiler for object-oriented languages. ACM SIGPLAN Notices, 31 (10):83–100, October 1996. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL

                                        http://www.cs.washington.
                                        edu/research/projects/cecil/

. Discusses performance of Vortex compiler for Cecil, C++, Java, and Modula-3.

DDJStaff:1998:NVS

[DDJ98a] DDJ Staff. News and views: a Standard Linux? cryptography contest; drives get smaller and Smaller; Perl conference; really embedded systems; programmer shortage?; Beowulf: Linux clustering; Java SPEC released. Dr. Dobb’s Journal of Software Tools, 23(11):16, November 1998. CODEN DDJOEB. ISSN 1044-789X. URL http://

                                        www.ddj.com/.

DDJ:1998:NVN

[DDJ98b] DDJ Staff. News and views: New trends in vaporware; distance ed might pay off; life in the fast lane; making friends in Washington; news on OpenGL 1.2; no discounts for schools; let’s do lunch; encryption export challenge?; Java fissures; searching for talent in science; nanomedicine. Dr. Dobb’s Journal of Software Tools, 23(6):18, June 1998. CODEN DDJOEB. ISSN 1044-789X.

DDJStaff:1998:NVN

[DDJ98c] DDJ Staff. News and views:

New trends in vaporware; distance ed might pay off; life in the fast lane; making friends in Washington; news on OpenGL 1.2; no discounts for schools; let’s do lunch; encryption export challenge?; Java fissures; searching for talent in science; nanomedicine. Dr. Dobb’s Journal of Software Tools, 23(6):18, June 1998. CODEN DDJOEB. ISSN 1044-789X.

DDJstaff:1999:NVR

[DDJ99a] DDJ staff. News and views:

Real-time Java working group; simulated safety; A house of smart cards; father of ubiquitous computing passes away; electrochemical fabrication. Dr. Dobb’s Journal of Software Tools, 24(7):18, July 1999. CO-DEN DDJOEB. ISSN 1044-

789X. URL http://www.ddj.
com/.

DDJstaff:1999:NVSa

[DDJ99b] DDJ Staff. News and views:

Speeding up 3D modeling; Project Gutenberg; FSF honors Larry Wall; smart pens don’t make smart writers; power hogs; virtual fish: Java’s killer app?; evaluating testing tools; software patents con. Dr. Dobb’s Journal of Software Tools, 24 (1):18, January 1999. CODEN DDJOEB. ISSN 1044-789X.

                                        URL http://www.ddj.com/.

Doolin:1999:JCL

[DDS99] David M. Doolin, Jack Dongarra, and Keith Seymour. JLAPACK — compiling LA-PACK FORTRAN to Java. Scientific Programming, 7 (2):111–138, ???? 1999. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:

                                        //iospress.metapress.com/
                                        app/home/contribution.asp%
                                        3Fwasp=f277qlrwwjr5m4vxjywv%
                                        26referrer=parent%26backto=
                                        issue%2C4%2C8%3Bjournal%2C7%
                                        2C9%3Blinkingpublicationresults%
                                        2C1%2C1;   http://www.netlib.
                                        org/utk/people/JackDongarra/
                                        PAPERS/f2jrep~1.pdf. The

software is available on the World-Wide Web at http://

                                        www.cs.utk.edu/f2j/.

DeCarmo:1996:IT

[De 96] Linden De Carmo. Internet tools. PC Magazine, 15(14): 297–??, August 1, 1996. CO-DEN PCMGEP. ISSN 0888-8507.

Davignon:1996:JWS

[DE96a] Bernhard Davignon and Georg Edelmann. Java: Wie Steuere Ich Meine Kaffeemaschine? tewi, München, Germany, 1996. ISBN 3-89362-459-7. 184 pp. LCCN ???? 20 DM.

URL http://www.tewi.de/
scripts/tewi/details_ISBN.
idc?ISBN=3-89362-459-7.

Davignon:1996:WPM

[DE96b] Bernhard Davignon and Georg Edelmann. WWW-Publishing: multimediale Web-Seiten selbst erstellen; alles Wissenswerte über JAVA; kurze Einführung in VRML. te-wi, München, Germany, 1996. ISBN 3-89362-442-2. ???? pp. LCCN ???? 59 DM.

URL http://www.tewi.de/
scripts/tewi/details_ISBN.
idc?ISBN=3-89362-442-2.

Drossopoulou:1997:JTS

[DE97] S. Drossopoulou and S. Eisenbach. Java is type safe — probably. Lecture Notes in Computer Science, 1241:389– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Drossopoulou:1999:DSJ

[DE99] S. Drossopoulou and S. Eisenbach. Describing the semantics of Java and proving type soundness. Lecture Notes in Computer Science, 1523:41–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

December:1995:IJD

[Dec95a]  John December.  Impres-

sions of Java day. Computer-Mediated Communication Magazine, page ??, October 1, 1995. ISSN 1076-027X. URL http:

                                        //sunsite.unc.edu/cmc/mag/
                                        1995/oct/java.html.

December:1995:JTM

[Dec95b] John December. Java takes Manhattan. Computer-Mediated Communication Magazine, page ??, October 1, 1995. ISSN 1076-

                                        027X. URL http://sunsite.
                                        unc.edu/cmc/mag/1995/oct/
                                        december.html.

December:1995:PJI

[Dec95c] John December. Presenting Java: An Introduction to Java and HotJava. Howard W. Sams, Indianapolis, IN 46268, USA, October 1, 1995. ISBN 1-57521-039-8. xii + 207 pp. LCCN QA76.75.J38D43 1995. US$25.00, CDN$34.95, UK£22.95. URL http://www.

                                        december.com/john//. Also

available in Hebrew translation [Dec97b].

December:1996:IJ

[Dec96a] John December. Introducción a Java. Prentice Hall HispanoAmericana, Mexico, DF, Mexico / Nueva York, NY, USA, 1996. ISBN 968-880-686-2. xv + 207 pp. LCCN ???? 2080 Ptas.; $18 Ar-

                                        gentine. URL http://www.pc-
                                        news.com/9608/27lect2.htm.

December:1996:JEU

[Dec96b] John December. Java-Einführung und Überblick. Markt & Technik, Haar bei München, Germany, 1996. ISBN 3-8272-5055-2 (??invalid ISBN??). 290 pp. LCCN ???? 49,00 DM.

URL http://www.december.
com/john/;   http://www.mut.
ch/html/buchinfo.htx?&x_pp=
dJgxRqA9wlAYNijX&x_az=10&
x_pk=VpLvjnsx7b5V&x_pa=B&
x_pr=oS&x_bestell=25055&x_
stichwort=December&x_sprache=
D&x_start=1&x_gefunden=2&
x_query=&x_lastpage=buchfund&
x_soft=&x_suchewo=A;   http:
//www.mut.ch/profit.htx?&
x_pk=kd84mf7ms9dm; http://
www.mut.com/.

December:1997:JIM

[Dec97a]  John December.  Java —

Immagini in movimento nelle pagine WWW. Tecniche Nuove, ??, Italy, 1997. ISBN ???? 240 pp. LCCN ???? 39.000 Lire.

URL http://www.tecnet.it/
;   http://www.tecnet.it/
libri/VisualizzaLibro.cfm?
ISBN=02530.

December:1997:PJ

[Dec97b] John December. Presenting Java. Focus-Computers Publishing, P.O. Box 873 Ramat-Gan 52108, Israel, 1997. ISBN ???? ???? pp. LCCN

???? URL http://www.focus.

co.il/. Hebrew translation of [Dec95c]. Includes floppy disk.

deCarmo:1999:CJM

[deC99] Linden deCarmo. The Core Java Media Framework. Addison-Wesley, Reading, MA, USA, 1999. ISBN 0-13-011519-3. ca. 500 pp. LCCN QA76.73.J38 D428 1999. URL

                                        http://www.phptr.com/ptrbooks/
                                        ptr_0130115193.html.

December:19xx:JIM

[Decxxa]  John December.  Java —

Immagini in movimento nelle pagine WWW. Tecniche Nuove, ??, Italy, 19xx. ISBN 88-8480-253-0 (??Invalid checksum??). ???? pp. LCCN

????  39000 L.  URL

                                        http://www.tecnet.it/cgi-
                                        shl/dbml.exe?Action=Query&
                                        Template=/LibriQueryDueLibro.
                                        dbm&ISBN=02530.

December:19xx:JEU

[Decxxb] John December. Java-Einführung und Überblick. Markt & Technik, Haar bei München, Germany, 19xx. ISBN 3-8272-5055-2 (??invalid ISBN??). LCCN ???? 49 DM. URL http:/

                                        /www.mut.ch/html/buchinfo.
                                        htx?&x_pp=dJgxRqA9wlAYNijX&
                                        x_az=10&x_pk=VpLvjnsx7b5V&
                                        x_pa=B&x_pr=oS&x_bestell=
                                        25055&x_stichwort=December&
                                        x_sprache=D&x_start=1&x_gefunden=
                                        2&x_query=&x_lastpage=buchfund&
                                        x_soft=&x_suchewo=A.

December:19xx:PJb

[Decxxc] John December. Presenting Java. Focus-Computers Publishing, ????, 19xx. LCCN ???? ???? Hebrew translation.

Deitz:1996:ILE

[Dei96a] Dan Deitz. Industry leaders endorse Java and VRML. Mechanical Engineering: the journal of the American Society of Mechanical Engineers, 118(????):

28–??, ???? 1996. CODEN MEENAH. ISSN 0025-6501.

Deitz:1996:JNT

[Dei96b] Dan Deitz. Java: a new tool for engineering. Mechanical Engineering: the journal of the American Society of Mechanical Engineers, 118(4):68–72, April 1996. CODEN MEENAH. ISSN 0025-6501.

Deitel:1997:CJTc

[Dei97] Deitel & Associates, Inc.

A Complete Java Training Course: The Ultimate Cyberclassroom. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, second edition, November 1997. ISBN 0-13-790569-6. ???? pp. LCCN ????

US$100. URL http://www.
prenhall.com/ptrbooks/ptr_
0137905696.html.

Delessio:1997:CWM

[Del97] Carmen Delessio. Converting Windows Metafiles to Java. Dr. Dobb’s Journal of Software Tools, 22(5):34, 36, 37, 74, 76, May 1997. CODEN DDJOEB. ISSN 1044-789X.

DelGrosso:1998:J

[Del98] Antonio Del Grosso. Java. vdf Hochschulverlag, ????, 1998. ISBN 3-7281-2573-3. LCCN ???? 48 DM. URL http://

vdf.ethz.ch/vdf/info/2573.
html.

Demaine:1998:CJC

[Dem98a] Erik D. Demaine. C to Java:

converting pointers into references. Concurrency: Practice and Experience, 10(11– 13):851–861, September 1998. CODEN CPEXEI. ISSN 1040-3108. URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=10050394;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=10050394&PLACEBO=IE.pdf.

Special Issue: Java for Highperformance Network Computing.

Demaine:1998:CCP

[Dem98b] Erik D. Demaine. Converting C pointers to Java references. In ACM [ACM98a], page ?? CODEN CPEXEI. ISSN 1040-3108. LCCN ????

URL http://www.cs.ucsb.
edu/conferences/java98/papers/
pointers.pdf;   http://www.
cs.ucsb.edu/conferences/java98/
papers/pointers.ps.  Also

published as Concurrency: Practice and Experience, 10(11– 13), September 1998, CODEN CPEXEI, ISSN 1040-3108.

Dennis:1997:BCN

[Den97] T. Dennis. Boxing clever [networking AS/400]. Network News, ????(????):27–29, ???? 5, 1997. CODEN NNEWFL.

DeRoest:1996:PHP

[DeR96] Jim DeRoest. Piping hot pages. RS/Magazine, 5(6):24–27, June 1996. ISSN 1088-0844, 1061-0030. Discusses ANDF, Java, and JavaScript, and provides pointers to a number of Java resources on the Internet.

Descy:1996:AAI

[Des96]  Don Descy.  All aboard

the Internet: NCSA Mosaic, Netscape, and Java/Hot Java!! TechTrends: for leaders in education and training, 41(1):6–??, 1996. CODEN TETREF. ISSN 8756-3894.

Descartes:1998:JNM

[Des98] Descartes. Java Native Methods. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, April 1998. ISBN 1-56592-345-6. LCCN ???? ????

Deshpande:19xx:CDA

[Desxx] Salil Deshpande. CORBA and Distributed Applications Including Java. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 19xx. ISBN 0-13-349960-X. ???? pp. LCCN ???? US$45.

Detar:1997:IGS

[Det97] Jim Detar. Intel’s Grove shows Java development tool. Electronic News, 43(2172):56, June 16, 1997. CODEN ELNEAU. ISSN 0013-4937, 1061-6624.

Deugo:1998:JGJ

[Deu98] Dwight Deugo, editor. Java Gems: Jewels from Java Report. Cambridge University Press, Cambridge, UK, January 1998. ISBN 0-521-64824-6. viii + 230 pp. LCCN A76.73.J38 J367 1998. UK£20.

URL http://193.60.94.193/
scripts/webbook.asp?isbn=
0521648246.

DeVoe:1995:NCG

[DeV95] Deborah DeVoe. News: Compaq gets friendly with Internet. InfoWorld, page ??, December 16, 1995. CODEN INWODU. ISSN 0199-6649. URL http://

www.infoworld.com/.

Dewolfe:1997:JM

[Dew97] Michele Dewolfe. Javascript Mojo. Ventana Communications Group, Incorporated, Research Triangle Park, NC, USA, April 1, 1997. ISBN 1-56604-649-1. ???? pp. LCCN ???? US$49.99.

Sanchez:1996:CAL

[DF96] Arantza Diaz de Ilarraza Sanchez and I. (Isabel) Fernandez de Castro, editors. Computer aided learning and instruction in science and engineering: third international conference, CALISCE ’96, San Sebastian, Spain, July 29– 31, 1996: proceedings, volume 1108 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 3-540-61491-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q181.A1I538 1996.

Dincer:1997:UJJ

[DF97] Kivanc Dincer and Geoffrey C. Fox. Using Java and JavaScript in the Virtual Programming Laboratory: a Web-based parallel programming environment.

Concurrency: Practice and Experience, 9(6):485–508, June 1997. CODEN CPEXEI. ISSN 1040-3108. URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=13873;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=13873&PLACEBO=IE.pdf.

Dean:1996:JSH

[DFW96] Drew Dean, Edward W. Felten, and Dan S. Wallach. Java security: from HotJava to Netscape and beyond. In IEEE [IEE96a], pages 190–200. CO-DEN PSSPEO. ISBN 0-8186-7417-2. ISSN 1063-7109. LCCN QA76.9.A25 S95 1996. IEEE catalog number 96CB35924.

DSouza:1996:AMD

[DG96] Desmond F. D’Souza and Petter Graff. Advanced modeling and design for Java systems. In USENIX [USE96c], page ?? ISBN 1-880446-77-4. LCCN QA76.64 .U85

1996.  URL http://www.
usenix.org/publications/library/
proceedings/coots96/.

Decker:1997:PJL

[DH97] Rick Decker and Stuart Hirshfield. programming.java: Lab Manual, Beta edition. PWS-Kent Publishing Company, Division of Wadsworth, Inc., 20 Park Plaza, Boston, MA 02116, USA, 1997. ISBN 0-534-95597-5. viii + 87 pp. LCCN ???? US$18.

URL http://www.thomson.

                                        com/cgi-bin/plweb/hitlist.
                                        cgi?isbn-text+0534955975.

Decker:1998:PJI

[DH98] Rick Decker and Stuart Hirshfield. programming.java: An Introduction to Programming, Beta Edition. PWS-Kent Publishing Company, Division of Wadsworth, Inc., 20 Park Plaza, Boston, MA 02116, USA, 1998. ISBN 0-534-95588-6. xviii + 574 pp. LCCN QA76.73.J38D44 1998.

                                        US$36. URL http://www.pws.
                                        com/compsci/javabook.html.

Dieckman:1999:SAB

[DH99] S. Dieckman and U. Hoelzle. A study of the allocation behavior of the SPECjvm98 Java benchmarks. Lecture Notes in Computer Science, 1628:92–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Duke:1999:PFM

[DHM99a] David J. Duke, I. Herman, and Scott Marshall. PREMO: a framework for multimedia middleware: specification, rationale, and Java binding, volume 1591 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. ISBN 3-540-66720-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). xii + 254 pp. LCCN QA76.575 .D85 1999.

Duke:1999:CDJ

[DHM99b] David J. Duke, Ivan Herman, and M. Scott Marshall. Chapter 8. detailed Java specifications of the PREMO objects. Lecture Notes in Computer Science, 1591:205–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/
service/series/0558/bibs/
1591/15910205.htm;   http:
//link.springer-ny.com/link/
service/series/0558/papers/
1591/15910205.pdf.

Dick:1996:WDE

[Dic96] K. Dick. Web development environments. Object Magazine, 6 (7):72–73, September 1996. CO-DEN OBMAFO. ISSN 1055-3614.

Dichter:1997:AJP

[Dic97a] Carl Dichter. Avoiding Java portability pitfalls. UNIX Developer, 1(1):26–28, January/ February 1997. ISSN 1090-2279.

Dicken:1997:WDI

[Dic97b] Hans Dicken. Web-Datenbank-Integration mit Java. International Thomson Computer Press, 20 Park Plaza Suite 1001, Boston, MA 02116 USA, 1997. ISBN 3-8266-0343-5. 377 pp. LCCN ???? 69 DM. URL http:

//www.itp.de/datenbanken/
0343/0343.html.

Dicken:19xx:WDI

[Dicxx] Hans Dicken. Web-Datenbank-Integration mit Java. International Thomson Computer Press, 20 Park Plaza Suite 1001, Boston, MA 02116 USA, 19xx. ISBN 3-8266-0343-5 (??invalid ISBN??). LCCN ???? 69

                                        DM. URL http://www.itp.
                                        de/datenbanken/0343/0343.
                                        html.

DiDio:1996:IRU

[DiD96] Laura DiDio. IBM reassures users with Java for Warp. Computerworld, 30(48):14, November 25, 1996. CODEN CMP-WAB. ISSN 0010-4841.

DiDio:1997:IAJ

[DiD97a] Laura DiDio. IBM adds Java support to OS/2 Warp. Computerworld, 31(8):67, February 24, 1997. CODEN CMPWAB. ISSN 0010-4841.

DiDio:1997:IDB

[DiD97b]  Laura DiDio.  IBM divi-

sion builds Java applets for OS/2. Computerworld, 31(8): 32, February 24, 1997. CODEN CMPWAB. ISSN 0010-4841.

DiDio:1997:IIB

[DiD97c]  Laura DiDio.  IBM issues

beta Java kit for OS/2 Warp 4. Computerworld, 31(26): 24, June 30, 1997. CODEN CMPWAB. ISSN 0010-4841.

                                        URL http://www.software.
                                        ibm.com/os2warp.

DiDio:1997:IWS

[DiD97d] Laura DiDio. IBM OS/2 Warp Server to get Java jolt. Computerworld, 31(20):24, May 19, 1997. CODEN CMPWAB. ISSN 0010-4841.

Diehl:1996:JCS

[Die96] Stephan Diehl. Java & Co: Die Sprachen des Webs: HTML, VRML, Java, JavaScript. Addison-Wesley, Reading, MA, USA, 1996. ISBN 3-8273-1124-1. 288 pp. LCCN ????

49,90 DM. URL http://www.
addison-wesley.de/katalog/
item.ppml?textexpr=Java&id=
00099.

Diehl:1997:JCC

[Die97] Stan Diehl. Java cyber class:

Java multimedia study hall. BYTE Magazine, 22(9):30, September 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic). Review of [DD97f].

Diehl:1998:FIC

[Die98] Stephan Diehl. A formal introduction to the compilation of Java. Software—Practice and Experience, 28(3):297–327, March 1998. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=1782;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=1782&PLACEBO=IE.pdf.

Diehl:19xx:JCS

[Diexx]  Stephan Diehl.  Java &

Co. — Die Sprachen des Webs: HTML, VRML, Java, JavaScript. Addison-Wesley Verlag, Bonn, Germany, 19xx. ISBN 3-8273-1124-1 (??invalid ISBN??). LCCN ???? 50 DM.

DiGiorgio:1995:JBT

[DiG95a]  Rinaldo DiGiorgio.  Java

braintrust tells all. Sun-World Online, December 1995.

                                        URL http://www.sun.com/
                                        sunworldonline/swol-12-1995/
                                        swol-12-java.html.

DiGiorgio:1995:JDB

[DiG95b] Rinaldo DiGiorgio. Java developer: Brain trust spills the beans. Sunworld Online, 12: ??, 1995. ISSN 1091-8914.

                                        URL http://www.sun.com/
                                        sunworldonline/swol-12-1995/
                                        swol-12-java.html.

DiGiorgio:1995:JDF

[DiG95c] Rinaldo DiGiorgio. Java developer: The first JavaDev column. Sunworld Online, 11: ??, 1995. ISSN 1091-8914.

                                        URL http://www.sun.com/
                                        sunworldonline/swol-11-1995/
                                        swol-11-java.html.

DiGiorgio:1995:WJD

[DiG95d] Rinaldo DiGiorgio. Welcome to Java Developer. Sun-World Online, November 1995.

                                        URL http://www.Sun.COM/
                                        sunworldonline/swol-11-1995/
                                        swol-11-java.html.

DiGiorgio:1996:JDS

[DiG96]  Rinaldo DiGiorgio.  Java

developer: Smashing buttons. Sunworld Online, 01: ??, 1996. ISSN 1091-8914.

URL http://www.sun.com/
sunworldonline/swol-01-1996/
swol-01-java.html.

Demartini:1999:DDT

[DIS99] Claudio Demartini, Radu Iosif, and Riccardo Sisto. A deadlock detection tool for concurrent Java programs. Software—Practice and Experience, 29(7):577–603, June 1999. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=62003577;
http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=62003577&PLACEBO=IE.pdf.

Divjak:1998:DHB

[Div98] S. Divjak. Design of hypertext based courseware with the integrated JAVA and VRML modules. In IEEE [IEE98c], pages 178–181. ISBN 0-7803-3879-0 (softbound), 0-7803-3880-4 (casebound), 0-7803-3881-2 (microfiche). LCCN TK5 .M43 1998; TK5101.A1 M38 1998. Two volumes. IEEE catalog number 98CH36056.

Darcy:1998:APE

[DK98] Joseph D. Darcy and William Kahan. Analysis of Proposal for Extension to Java Floating Point Semantics, Revision 1.

Technical report, Java Grande Numerics Working Group, August 7, 1998. 18 pp. URL http:

//www.sonic.net/~jddarcy/
Research/jgrande.pdf.

Dongarra:1999:RAP

[DLM99] J. J. Dongarra, E. Luque, and Tomas Margalef, editors. Recent advances in parallel virtual machine and message passing interface: 6th European PVM/MPI Users’ Group Meeting, Barcelona, Spain, September 26–29, 1999: proceedings, volume 1697 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. ISBN 3-540-66549-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 E973 1999.

December:1996:JHJ

[DM96a] John December and Koichi Matsuda. Java, hot Java o shiru: Intanneto no atarashii sekai. Kabushiki Kaisha Purentisuhoru, Tokyo, Japan, shohan edition, 1996. ISBN 4-931356-99-0. xv + 255 pp. LCCN

???? URL http://www.mmjp.
or.jp/prentice/wa_int06-j.
html.

Dragan:1996:INN

[DM96b] Richard V. Dragan and Thomas Mace. Inside the Netscape Navigator platform. PC Magazine, 15(17):283–??, October 8, 1996. CODEN PCMGEP. ISSN 0888-8507.

Davis:1996:IJA

[DMB96] Owen Davis, Tom McGinn, and Amit Bhatiani. Instant Java Applets. Ziff-Davis Press, Emeryville, CA, USA, June 1, 1996. ISBN 1-56276-386-5, 1-56276-415-2. xi + 220 + 6 pp. LCCN QA76.73.J38D36 1996. US$29.99, CDN$27.95.

URL http://www.mcp.com/
251356511370317/cgi-bin/bag?
isbn=1-56276-386-5&last=/
bookstore;   http://www.mcp.

com/zdpress. Includes CD-ROM.

deFigueiredo:1997:XBS

[dN97] Luiz Henrique de Figueiredo and Marcio Lobo Netto, editors. X Brazilian Symposium on Computer Graphics and Image Processing: proceedings, Campos do Jordao, Brazil, October 14–17, 1997. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1997. ISBN 0-8186-8102-0, 0-8186-8103-9, 0-8186-8104-7. LCCN T385 .S59 1997. IEEE Computer Society order number PR08102. IEEE order plan catalog number 97TB100164.

Dobson:1998:PEH

[Dob98] Rick Dobson. Programming ECMAScript: The holy standard? BYTE Magazine, 23 (7):47–??, July 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Doerr:1995:LSA

[Doe95]  Rita McCardell Doerr.  A

Lexical-semantic and Statistical Approach to Lexical Collocation Extraction for Natural Language Generation. Thesis (ph.d.), University of Maryland Baltimore County, Baltimore, MD, USA, 1995. 408 pp. See Dissertation Abstracts Index, DAI, Vol. 56-02B, Page 0915.

Doemel:1997:IJT

[Doe97] P. Doemel. Interaction of Java and Telescript agents. Lecture Notes in Computer Science, 1222:295–??, 1997. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Doherty:1997:TYJ

[Doh97] Donald Doherty. Teach Yourself Java Beans in 21 Days. Sams.net, Indianapolis, IN 46268, USA, June 1997. ISBN 1-57521-316-8 (softcover). xxii + 448 pp. LCCN QA76.73.J38D64 1997. US$35.99. URL http://

                                        merchant.superlibrary.com:
                                        8000/catalog/hg/PRODUCT/PAGE/
                                        15752/bud/1575213168.html.

Includes CD-ROM.

Doherty:1998:SEU

[Doh98] Don Doherty. Special Edition Using JBuilder. Que Corporation, Indianapolis, IN, USA, January 1998. ISBN 0-7897-0892-2. ???? pp. LCCN ???? US$50.

Domel:1996:MTA

[Dom96] Peter Domel. Mobile Telescript agents and the Web. In IEEE [IEE96c], pages 52–57. CODEN DCSIDU. ISBN 0-8186-7414-8. ISSN 1063-6390. LCCN QA76 .I49 1996. IEEE Cat. No. 96CB35911.

Donkers:1996:EYW

[Don96]  Arthur Donkers.  Enhance

your Web-Based GUI with Java. SysAdmin, 5(7):35, 36, 39, 40, 42, 43, 45, 47, 50, July 1996. CODEN SYADE7. ISSN 1061-2688.

Dooley:1999:IJS

[Doo99] Sam Dooley, editor. ISSAC 99: July 29–31, 1999, Simon Fraser University, Vancouver, BC, Canada: proceedings of the 1999 International Symposium on Symbolic and Algebraic Computation. ACM Press, New York, NY 10036, USA, 1999. ISBN 1-58113-073-2. LCCN QA76.95 .I57 1999.

Downs:1997:TYJ

[Dow97] Andrew Downs. Teach Yourself Java for Macintosh in 21 days. Hayden Books, 4300 West 62nd Street, Indianapolis, IN 46268, USA, second edition, November 1997. ISBN 1-56830-342-4. ???? pp. LCCN ???? 79,00 DM; US$40. URL http:/

/www.mut.ch/profit.htx?&x_
pk=kd84mf7ms9dm.

Downing:1998:JRR

[Dow98a] Troy Bryan Downing. Java RMI: Remote Method Invocation. IDG Books, San Mateo, CA, USA, January 1998. ISBN 0-7645-8043-4 (softcover). xxi + 298 pp. LCCN QA76.73.J38D69 1998. US$35.99. URL http:/

/www.idgbooks.com/cgi-bin/
db/fill_out_template.pl?idgbook:
0-7645-8043-4:book-idg::uidg323.

Downing:1998:SSJ

[Dow98b] Troy Bryan Downing. Server-Side Java Power Guide. IDG Books, San Mateo, CA, USA, 1998. ISBN 0-7645-8046-9. ???? pp. LCCN ???? US$40.

DeRose:1999:TTM

[DP99] Luiz De Rose and David Padua. Techniques for the translation of MATLAB programs into Fortran 90. ACM Transactions on Programming Languages and Systems, 21(2):286–323, March 1999. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). URL http://www.

acm.org:80/pubs/citations/
journals/toplas/1999-21-2/
p286-de_rose/.

Drake:1996:JAP

[DR96] Robin Drake and Tad Ringo. Java Applet Powerpack, Vol.

1. Prima Publishing, Rocklin, CA, USA, September 1, 1996. ISBN 0-7615-0678-0. xii + 344 pp. LCCN QA76.73.J38 J36 1996. US$30.00. URL http:

//www.primapublishing.com/
cpd/76150678.html.

December:1997:WWW

[DR^{+}97] John December, Neil Randall, et al. The World Wide Web 1997 Unleashed. Sams.net, Indianapolis, IN 46268, USA, fourth edition, February 1997. ISBN 1-57521-184-X. xxxv + 1226 pp. LCCN TK5105.888 .D43 1997. US$49.99. URL http://

merchant.superlibrary.com:
8000/catalog/hg/PRODUCT/PAGE/
15752/bud/157521184X.html.

Includes CD-ROM.

Dennis:1998:MCJ

[DR98] Gregory C. Dennis and James R.

Rubin. Mission-critical Java project management: business strategies, applications, and development. Addison-Wesley, Reading, MA, USA, August 1998. ISBN 0-201-32573-X. xviii + 245 pp. LCCN HF5548.5.J38 D46 1999. US$35.

Dennis:1999:MCJ

[DR99] Gregory C. Dennis and James R.

Rubin. Mission-critical Java project management: business strategies, applications, and development. Addison-Wesley, Reading, MA, USA, 1999. ISBN 0-201-32573-X. xviii + 245 pp. LCCN HF5548.5.J38D46 1999. US$34.95.

Drake:1997:CSP

[Dra97] R. Drake. CRED: seven pillars for Java development. American Programmer, 10(1):13–21, January 1997. CODEN AMPRFD. ISSN 1048-5600.

Dray:1998:RNJ

[Dra98] Jim Dray. Report on the NIST Java AES candidate algorithm analysis. In National Institute of Standards and Technology [Nat98], page 29. ISBN ???? LCCN ????

                                        URL http://csrc.nist.gov/
                                        encryption/aes/round1/r1-

java.pdf. See [RD99] for a conference overview. No formal proceedings were published, but the conference Web site contains pointers to slides and/or technical papers for most of the fifteen “complete and proper” candidates.

Drespling:1996:JHJ

[Dre96] W. Drespling. Java/HotJava/ und wie das Internet zum Leben erweckt wird [English: Java/ Hot Java and the Internet]. In Anonymous [Ano96-147], pages 118–122. ISBN ???? LCCN ????

Dreystadt:1998:SJR

[Dre98] John Dreystadt. Storing Java in a relational database. BYTE Magazine, 23(6):61–??, June 1998. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Dreyfus:19xx:JM

[Drexx] Michel Dreyfus. JavaScript MagaPoche. Sybex France, Paris, France, 19xx. ISBN 2-7361-2359-X. ???? pp. LCCN

                                        ???? URL http://www.sybex.

fr/. Includes CD-ROM.

Drix:19xx:LCN

[Drixx]  Drix.  Langage C Norme

ANSI vers une Pensée Objet en Java. Masson Editeur, Masson, France, third edition, 19xx. ISBN 2-225-82937-3. ???? pp. LCCN ???? 225 FF.

deMeer:1998:TAE

[dRPT98] Hermann de Meer, Jan-Peter Richter, Antonio Puliafito, and Orazio Tomarchio. Tunnel agents for enhanced Internet QoS. IEEE Concurrency, 6(2): 30–39, April–June 1998. CO-DEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic).

Dryden:1996:JFR

[Dry96a] Patrick Dryden. Java fuels remote hub management. Computerworld, 30(36):6, September 2, 1996. CODEN CMPWAB. ISSN 0010-4841.

Dryden:1996:JPW

[Dry96b] Patrick Dryden. Java pitched to wake up management tools. Computerworld, 30(24):24, June 10, 1996. CODEN CMPWAB. ISSN 0010-4841.

Dryden:1997:CGJ

[Dry97] Patrick Dryden. Cabletron gulps Java. Computerworld, 31 (14):14, April 7, 1997. CODEN CMPWAB. ISSN 0010-4841.

Davis:1995:NIP

[DS95a] Jessica Davis and Ed Scannell.

News: Internet pushes Lotus to alter Notes pricing. InfoWorld, page ??, December 10, 1995. CODEN INWODU. ISSN 0199-6649. URL http://www.

infoworld.com/.

Davis:1995:NLJ

[DS95b] Jessica Davis and Ed Scannell. News: Lotus jockeys Notes into better spot in on-line race. InfoWorld, page ??, December 16, 1995. CODEN INWODU. ISSN 0199-6649. URL http://www.

infoworld.com/.

DSouza:1996:ISF

[D’S96a]  D. D’Souza.  Interfaces,

subtypes, and frameworks [Java]. Journal of Object Oriented Programming, 9(7):19–22, November–December 1996. CO-DEN JOOPEC. ISSN 0896-8438.

DSouza:1996:JDM

[D’S96b] D. D’Souza. Java: design and modeling opportunities. Journal of Object Oriented Programming, 9(5):14–??, 16–18, September 1996. CODEN JOOPEC. ISSN 0896-8438.

DSouza:1996:JRB

[D’S96c] Desmond D’Souza. Java, a refreshing brew: Thinking in terms of objects. Web Techniques, 1(2):??, May 1996. CO-DEN WETEFA. ISSN 1086-

556X.  URL http://www.
webtechniques.com/features/
may96java.html.

Deering:1997:JSV

[DS97a] M. Deering and H. Sowizral. Java3D Specification, Version 1.0. Sun Microsystems, 2550 Garcia Avenue, Mountain View, CA 94043, USA, August 1997.

URL http://java.sun.com/
products/java-media/3D/.

Dragan:1997:JFG

[DS97b] Richard V. Dragan and Larry Seltzer. Java: a field guide for users. PC Magazine, 16(10): 100, 101, 103, 104, 108, 111– 113, 115, May 1997. CODEN PCMGEP. ISSN 0888-8507.

DSouza:1997:TCL

[D’S97c] D. D’Souza. Types and classes:

a language-independent view. Journal of Object Oriented Programming, 10(1):10–13, March– April 1997. CODEN JOOPEC. ISSN 0896-8438.

DePauw:1999:VRP

[DS99] W. De Pauw and C. Sevitski. Visualizing reference patterns for solving memory leaks in Java. Lecture Notes in Computer Science, 1628:116– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Daconta:1999:JJC

[DSM99] Michael C. Daconta, Al Saganich, and Eric Monk. Java 2 and JavaScript for C and C++ Programmers. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, revised edition, 1999. ISBN 0-471-32719-0. xxiv + 868 pp. LCCN QA76.73.J38D865 1999. US$59.99.

Danesh:1996:JDG

[DT96] Arman Danesh and Wes Tatters. JavaScript 1.1 Developer’s Guide. Howard W. Sams, Indianapolis, IN 46268, USA, December 1996. ISBN 1-57521-084-3. xxiv + 604 pp. LCCN QA76.73.J39D34 1996. US$49.99. URL http://www.

mcp.com/samsnet/.

Duan:1996:DDA

[Dua96] Nick N. Duan. Distributed database access in a corporate environment using Java. Computer Networks and ISDN Systems, 28(7–11):1149–1156, May 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/
browse/browse.cgi?year=1996&
volume=28&issue=7-11&aid=
1600.

Dubois:1997:JSP

[Dub97] Paul F. Dubois. Is Java for scientific programming? Computers in Physics, 11(6):611– ??, November 1997. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (elec-

tronic). URL https://aip.
scitation.org/doi/10.1063/
1.4822611.

Duego:1997:JG

[Due97a] Dwight Duego, editor. Java Gems. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, August 1997. ISBN 0-13-904863-4. ???? pp. LCCN ???? US$29. URL http://

www.prenhall.com/ptrbooks/
ptr_0139048634.html.

Duego:1997:JGJ

[Due97b] Dwight Duego, editor. Java Gems: Jewels from Java Report. SIGS Books, 71 W 23rd St., New York, NY 10010, USA, 1997. ISBN 1-884842-80-1. ???? pp. LCCN ????

US$. URL http://www.sigs.
com/books/deugo/.

Duego:1998:JG

[Due98] Dwight Duego, editor. Java Gems. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, second edition, 1998. ISBN 0-13-64824-6 (??Invalid checksum??). 248 pp. LCCN ???? US$29.95.

Duggan:1999:MTB

[Dug99] Dominic Duggan. Modular type-based reverse engineering of parameterized types in Java code. ACM SIGPLAN Notices, 34(10):97–113, October 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

URL http://www.acm.org/
pubs/citations/proceedings/
oops/320384/p97-duggan/.

Duntemann:1996:JPF

[Dun96] Jeff Duntemann. JBuilder Programming FrontRunner. Coriolis Group Books, Scottsdale, AZ, USA, 1996. ISBN 1-57610-004-9. ???? pp. LCCN ???? US$29.95. URL http://

                                        www.coriolis.com/. Duplicate

ISBN with [Dun96, GTFxx, KGTF97].

Duncan:1997:ONS

[Dun97] Luke Duncan. Official Netscape server-side JavaScript for database applications: for Enterprise Server 3. Ventana Press, Chapel Hill, NC, USA, 1997. ISBN 1-56604-745-5. ???? pp. LCCN QA76.73.J39 D86 1997.

Dunn:1999:JRE

[Dun99] Douglas Dunn. The Java Rules! Electronic Book. Integrated Warehouse Systems, Inc., PO Box 10643, Towson, Maryland 21285, USA; Tel: +1 800 828-5951, 1999. ISBN 1-893876-00-4. 1127 pp. US$39.50. URL

                                        http://www.javarules.com.

Dershem:1998:JCV

[DV98] Herbert L. Dershem and James Vanderhyde. Java class visualization for teaching objectoriented concepts. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(1):53–57, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Valle:19xx:JMI

[dVGxx] André Bittencourt do Valle and Claudia Guimarã Bittencourt do Valle. Java Manual de Introdução. Axcel Books do Brasil Editora, Rua da Glória, 344/10 And. - Ed. Glória Trade Center Glória - Rio de Janeiro/RJ, Brasil, 19xx. ISBN ???? ???? pp. LCCN ???? R$ 35. URL

http://www.axcelbooks.com.
br/cgi-bin/cgiwrap/axcel/
pes-tit.pl?livro=102.

Dhabolt:1996:EJ

[DW96a] J. Dhabolt and D. Weissman.

Exploring Java. SciTech Journal, 6(3):14–16, March 1996. CODEN SITJE8. ISSN 1072-0995.

DSouza:1996:AAD

[DW96b] Desmond D’Souza and Jun Wang. Algorithm alley: Data management with Java. Web Techniques, 1(6):30, 32, 33, September 1996. CODEN WETEFA. ISSN 1086-556X.

DSouza:1997:CF

[DW97a] D. D’Souza and A. Wills. Collaboration frameworks. Journal of Object Oriented Programming, 9(9):68–71, February 1997. CODEN JOOPEC. ISSN 0896-8438.

Dwyer:1997:FPA

[DW97b] Matthew B. Dwyer and Virgil Wallentine. A framework for parallel adaptive grid simulations. Concurrency: Practice and Experience, 9(11):1293–1310, November 1997. CODEN CPEXEI. ISSN 1040-3108. URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=13824;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?

                                        ID=13824&PLACEBO=IE.pdf. Spe-

cial Issue: Java for computational science and engineering — simulation and modeling II.

Drye:1999:JFC

[DW99] Stephen C. Drye and William C. Wake. Java Foundation Classes: Swing Reference. Manning Publications, Greenwich, CT, USA, July 1999. ISBN 1-884777-67-8. xxi + 1065 pp. LCCN QA76.73.J38 D79 1999.

                                        US$40.  URL http://www.
                                        manning.com/Drye/index.html.

Drossopoulou:1998:WJB

[DWE98] Sophia Drossopoulou, David Wragg, and Susan Eisenbach. What is Java binary compatibility? ACM SIGPLAN Notices, 33(10):341–358, October 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dolvane:1999:PIAa

[DY99a] Jaison Dolvane and Kumanan Yogaratnam. PersonalJava and information appliances, part I. Dr. Dobb’s Journal of Software Tools, 24(1):60, 62, 64–68, January 1999. CODEN DDJOEB. ISSN 1044-789X. URL http:/

                                        /www.ddj.com/ddj/ftp/1999/
                                        1999_01/pjia1.txt.

Dolvane:1999:PIAb

[DY99b] Jaison Dolvane and Kumanan Yogaratnam. PersonalJava and information appliances, part II. Dr. Dobb’s Journal of Software Tools, 24(2):64, 66–68, 70–71, February 1999. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/1999/9902/9902toc.
htm;   http://www.ddj.com/
ddj/1999/1999_02/9902f/9902f.
htm;   http://www.ddj.com/
ftp/1999/1999_02/pjphone.
txt;   http://www.ddj.com/
ftp/1999/1999_02/pjphone.
zip.

Dyson:1996:FTS

[Dys96] P. E. Dyson. FutureTense’s Texture: slow page viewer mars powerful design tool. Seybold Report on Internet Publishing, 1 (1):5–8, September 1996. ISSN 1090-4808.

Easton:1999:JQD

[Eas99] Evan Easton. Java Q&A: Does Java support enums? Dr. Dobb’s Journal of Software Tools, 24(10):113–117, October 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ftp/1999/1999_
10/jqa1099.txt;   http://
www.ddj.com/ftp/1999/1999_
10/jqa1099.zip.

Encarnacao:1996:NUI

[EBB^{+}96] L. Encarnação, Klaus M. Bauer, Klaus Böhm, Roland Holzapfel, Tanja Koop, and Ulrike Spierling. New user interface aspects to the information highway — applying Java and VRML. International Journal of Information Technology, 2(1):79–95, June 1996. ISSN 0218-7957.

Espinosa:1996:IAB

[EBR96a] E. Espinosa, A. Brito, and F. Ramos. Intelligent agentbased virtual education using the Java technology. Lecture Notes in Computer Science, 1086:270–278, 1996. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Espinosa:1996:IAV

[EBR96b] E. Espinosa, A. Brito, and F. Ramos. Intelligent agentbased virtual education using the Java technology. In Frasson et al. [FGL96], pages 270–278. ISBN 3-540-61327-7. LCCN LB1028.73 .I58 1996.

Eckel:1996:HGC

[Eck96a] B. Eckel. How garbage collection can increase speed. Web Techniques, 1(9):36–??, 38, December 1996. CODEN WETEFA. ISSN 1086-556X.

Eckel:1996:JAS

[Eck96b] B. Eckel. Java alley: speed. Web Techniques, 1(7):24–??, 26–27, October 1996. CODEN WETEFA. ISSN 1086-556X.

Eckel:1996:LE

[Eck96c] B. Eckel. Language extensions. Web Techniques, 1(8):36– ??, 38, November 1996. CODEN WETEFA. ISSN 1086-556X.

Eckel:1996:TJ

[Eck96d] Bruce Eckel. Thinking in Java. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1996. ISBN ???? ???? pp. LCCN

???? URL http://www.zone.
org/Eckel/books.htm.

Eckel:1997:BAT

[Eck97a]  B. Eckel.  The black art

of threading [Java programming]. Web Techniques, 2(1): 34–??, 36–37, January 1997. CODEN WETEFA. ISSN 1086-556X.

Eckel:1997:DIAa

[Eck97b] B. Eckel. Discovering information at run time. Web Techniques, 2(2):34–??, 36– 37, February 1997. CODEN WETEFA. ISSN 1086-556X.

Eckel:1997:DIAb

[Eck97c] B. Eckel. Discovering information at run time: part deux. Web Techniques, 2(3):28– ??, 30–31, March 1997. CODEN WETEFA. ISSN 1086-556X.

Eckel:1997:ISJ

[Eck97d] B. Eckel. IO streams in Java 1.1.

Web Techniques, 2(7):30, 32–36, July 1997. CODEN WETEFA. ISSN 1086-556X.

Eckel:1998:TJ

[Eck98]  Bruce Eckel.  Thinking in

Java. SunSoft Press Java series. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1998. ISBN 0-13-659723-8. 1098 pp. LCCN QA76.73.J38 E25 1998.

????  URL http://www.
eckelobjects.com/javabook.
html.

Edwards:1999:CJ

[Edw99] W. Keith Edwards. Core Jini. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1999. ISBN 0-13-014469-X. 772 pp. LCCN QA76.9.D5E38 1999. US$49.99. URL http:

//www.phptr.com/ptrbooks/
ptr_013014469X.html.

Ek:1997:JPb

[EE97] Jesper Ek and Rasmus Ekman. JavaScript-programmering. Pagina Förlags AB, Box 706, Travgatan 92, 194 27 Upplands-Väsby, Sweden, 1997. ISBN 91-636-0442-6 (??Invalid checksum??). ca. 250 pp. LCCN ????

Skr 312. URL http://www.
pagina.se/butik/product.html?
9,20442,-1,dvrobyaksz.

Ek:19xx:JPIb

[EExx] Jesper Ek and Rasmus Ekman. Java-programmering. En introduktion. Pagina Förlags AB, Box 706, Travgatan 92, 194 27 Upplands-Väsby, Sweden, second edition, 19xx. ISBN 91-636-0474-4 (??invalid ISBN??). ???? pp. LCCN ???? 317

krona.  URL http://www.
pagina.se/cgi-bin/ps_main_
show.pl?art=20474.

Egremont:1999:MBB

[Egr99] Carlton Egremont III. Mr. Bunny’s Big Cup o’ Java. Addison-Wesley, Reading, MA, USA, 1999. ISBN 0-201-61563-0. 115 pp. LCCN ???? US$14.95.

Espinoza:1996:WIA

[EH96] F. Espinoza and K. Hook. A WWW interface to an adaptive hypermedia system. In Anonymous [Ano96-154], pages 143– 167. ISBN ???? LCCN ????

Ellis:1999:CSD

[EHS^{+}99] Ainslie Ellis, Dianne Hagan, Judy Sheard, Jason Lowder, Wendy Doube, Angela Carbone, John Robinson, and Sylvia Tucker. A collaborative strategy for developing shared Java teaching resources to support first year programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(3): 84–87, September 1999. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Eisenmenger:1997:J

[Eis97] Richard Eisenmenger. JavaScript.

Markt & Technik, Haar bei München, Germany, 1997. ISBN 3-8272-5254-7. 400 pp. LCCN ???? 39,95 DM. URL

http://www.mut.ch/profit.
htx?&x_pk=kd84mf7ms9dm.

Ek:1997:JAT

[Ek97a] Jesper Ek. Java-applets tips och tillämpningar. Pagina Förlags AB, Box 706, Travgatan 92, 194 27 Upplands-Väsby, Sweden, 1997. ISBN 91-636-0427-2 (??Invalid checksum??). 250 pp. LCCN ???? Skr 471.

URL http://www.pagina.se/
butik/product.html?9,20427,
-1,cjnblpapne.

Ek:1997:JPa

[Ek97b]  Jesper Ek.  Java program-

mering. Pagina Förlags AB, Box 706, Travgatan 92, 194 27 Upplands-Väsby, Sweden, 1997. ISBN 91-636-0419-1 (??Invalid checksum??). 188 pp. LCCN ???? SKr 292.

                                        URL http://www.pagina.se/
                                        butik/product.html?9,20419,
                                        -1,cjnblpapne.

Engelbrecht:1998:ITS

[EK98] R. L. Engelbrecht and D. Kourie. Issues in translating Smalltalk to Java. Lecture Notes in Computer Science, 1383:249– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ek:19xx:JAT

[Ekxxa] Jesper Ek. Java-applets tips och tillämpningar. Pagina Förlags AB, Box 706, Travgatan 92, 194 27 Upplands-Väsby, Sweden, 19xx. ISBN 91-636-0427-2 (??invalid ISBN??). ???? pp. LCCN ???? 471 krona.

                                        URL http://www.pagina.se/
                                        butik/product.html?9,20427,
                                        -1,unjqpsbwrv.

Ek:19xx:JOT

[Ekxxb] Jesper Ek. Java-ohjelmointi tutuksi. Pagina Förlags AB, Box 706, Travgatan 92, 194 27 Upplands-Väsby, Sweden, 19xx. ISBN 951-644-089-4. ???? pp. LCCN ???? ????

                                        URL http://www.pagina.fi/
                                        kirjat/44089.html.

Ek:19xx:JPIa

[Ekxxc] Jesper Ek. Java-programmering:

En introduction. Pagina Förlags AB, Box 706, Travgatan 92, 194 27 Upplands-Väsby, Sweden, 19xx. ISBN 91-636-0419-1 (??invalid ISBN??). ???? pp. LCCN ???? 292 krona.

URL http://www.pagina.se/
butik/product.html?9,20419,
-1,cjnblpapne.

Ek:19xx:LOJ

[Ekxxd]  Jesper Ek.  Lättpocket om

Java-programmering. Pagina Förlags AB, Box 706, Travgatan 92, 194 27 Upplands-Väsby, Sweden, 19xx. ISBN 91-636-0477-9 (??invalid ISBN??). ???? pp. LCCN ???? 245

krona.  URL http://www.
pagina.se/cgi-bin/ps_main_
show.pl?art=20477.

Egan:1997:JST

[EKR97] M. A. Egan, M. Krishnamoorthy, and K. Rajan. A Java simulation tool for fuzzy clustering. Concurrency: Practice and Experience, 9(11): 1327–1332, November 1997. CODEN CPEXEI. ISSN 1040-3108. URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=13837;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=13837&PLACEBO=IE.pdf. Spe-

cial Issue: Java for computational science and engineering — simulation and modeling II.

Esteire:1998:STN

[EL98] Oscar Azañón Esteire and Juan Manuel Cueva Lovelle. Set of tools for native code generation for the Java virtual machines. ACM SIGPLAN Notices, 33(3):73–79, March 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Eckstein:1998:JS

[ELW98] Robert Eckstein, Marc Loy, and Dave Wood. Java Swing. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1998. ISBN 1-56592-455-X. xxviii + 1227 pp. LCCN QA76.73.J38 E27 1998. US$39.95.

Even:1996:CCA

[EM96] R. Even and B. Mishra. CAFE: a Complex Adaptive Financial Environment. In IEEE [IEE96p], pages 20– 25. ISBN 0-7803-3236-9 (softbound), 0-7803-3237-7 (microfiche). LCCN HG176.7.I33 1996. IEEE Cat. No. 96TH8177.

Eisenberg:1999:SPS

[EM99] Andrew Eisenberg and Jim Melton. SQLJ part 1: SQL routines using the Java programming language. SIG-MOD Record (ACM Special Interest Group on Management of Data), 28(4):58–63, December 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

EMedia:1997:HJP

[EMe97a] EMedia Professional, Inc. Handson Java Programming Using Microsoft Visual J++. MindQ Publishing, Inc., 450 Springpark Place Ste. 1200, Herndon, VA 20170, USA, Tel: +1 703 708-9380, 1997. ISBN 1-57559-005-0. ???? pp. LCCN ???? US$49.95. URL

http://www.mindq.com/java/
hovisjpp.html. CD ROM.

Emedia:1997:HVC

[EMe97b] EMedia Professional, Inc. Hands On Visual Café. MindQ Publishing, Inc., 450 Springpark Place Ste. 1200, Herndon, VA 20170, USA, Tel: +1 703 708-9380, 1997. ISBN 1-57559-017-4. ???? pp. LCCN

????  US$49.95.  URL

http://www.mindq.com/java/
viscafe.html. CD ROM.

EMedia:1997:IJ

[EMe97c] EMedia Professional, Inc. Inside JavaBeans. MindQ Publishing, Inc., 450 Springpark Place Ste. 1200, Herndon, VA 20170, USA, Tel: +1 703 708-9380, 1997. ISBN 1-57559-020-4. ???? pp. LCCN

????  US$49.95.  URL

http://www.mindq.com/java/
insbeans.html. CD ROM.

EMedia:1997:IPJb

[EMe97d] EMedia Professional, Inc. Introduction To Programming Java Applets: Microsoft Visual J++ Edition. MindQ Publishing, Inc., 450 Springpark Place Ste.

1200, Herndon, VA 20170, USA, Tel: +1 703 708-9380, 1997. ISBN 1-57559-003-4. ???? pp. LCCN ???? US$49.95. URL

http://www.mindq.com/java/
introjpp.html. CD ROM.

EMedia:1997:IPJa

[EMe97e] EMedia Professional, Inc. An Introduction to Programming Java Applets: Standard Edition. MindQ Publishing, Inc., 450 Springpark Place Ste. 1200, Herndon, VA 20170, USA, Tel: +1 703 708-9380, 1997. ISBN 1-57559-007-1 (??invalid ISBN??). ???? pp. LCCN ???? US$49.95.

URL http://www.mindq.com/
java/standard.html.  CD

ROM.

EMedia:1997:IPJd

[EMe97f] EMedia Professional, Inc. Introduction To Programming Java Applets: Sun Java Workshop Edition. MindQ Publishing, Inc., 450 Springpark Place Ste. 1200, Herndon, VA 20170, USA, Tel: +1 703 708-9380, 1997. ISBN 1-57559-007-7. ???? pp. LCCN ???? US$49.95. URL

http://www.mindq.com/java/
introsun.html. CD ROM.

EMedia:1997:IPJc

[EMe97g] EMedia Professional, Inc. An Introduction to Programming Java Applets Using Sun Java Workshop. MindQ Publishing, Inc., 450 Springpark Place Ste. 1200, Herndon, VA 20170, USA, Tel: +1 703 708-9380, 1997. ISBN 1-57559-003-4. LCCN ???? US$50.

EMedia:1997:JB

[EMe97h] EMedia Professional, Inc. Java Beans. MindQ Publishing, Inc., 450 Springpark Place Ste. 1200, Herndon, VA 20170, USA, Tel: +1 703 708-9380, 1997. ISBN ???? ???? pp. LCCN ???? US$49.95.

URL http://www.mindq.com/
java/javaintro.html.  CD

ROM.

EMedia:1997:JPC

[EMe97i] EMedia Professional, Inc. Java Programming and Core Class Libraries. MindQ Publishing, Inc., 450 Springpark Place Ste. 1200, Herndon, VA 20170, USA, Tel: +1 703 708-9380, 1997. ISBN 1-57559-008-5. ???? pp. LCCN ???? US$49.95.

URL http://www.mindq.com/
java/jpaccl.html. CD ROM.

Emery:1998:AAB

[EMN98] David E. Emery, Robert F.

Mathis, and Karl A. Nyberg. Automating the Ada binding process for Java — how far can we go? Lecture Notes in Computer Science, 1411:29– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/
service/series/0558/bibs/
1411/14110029.htm;   http:
//link.springer-ny.com/link/
service/series/0558/papers/
1411/14110029.pdf.

Enete:1997:JJB

[Ene97] Noel Enete. Java Jumpstart: a Beginner’s Guide to Internet Programming. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, March 1997. ISBN 0-13-565854-3. xv + 495 pp. LCCN QA76.73.J38E54 1997. US$29.95. URL http://www.

                                        prenhall.com/allbooks/ptr_
                                        0135658543.html.

Engberg:1997:MJV

[Eng97a] David A. Engberg. Multi-user Java/VRML performance: research project. Master of science, plan ii, University of California, Berkeley. Department of Electrical Engineering and Computer Sciences, Berkeley, CA, USA, 1997. 10 pp.

Englander:1997:JTB

[Eng97b] Rob Englander. Java talk: The BeanBox tool. Silicon Graphics World, 7(8):18–19, August 1997. ISSN 1057-7041. Excerpted from [Eng97c].

Englander:1997:DJB

[Eng97c] Robert Englander. Developing Java Beans. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1997. ISBN 1-56592-289-1. xvi + 298 pp. LCCN QA76.73.J38 E64 1997. US$29.95. URL http://www.

                                        ora.com/catalog/javabeans/

. Excerpted in [Eng97b].

Englander:1997:JGD

[Eng97d]  Robert Englander.  Jav-

aBeans Guide du Programmeur. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, December 1997. ISBN 2-84177-038-9. LCCN ???? 220 FF. URL http:

//www.editions-oreilly.fr/
international/france/prog/
java-beans.html.

Englander:1998:PCJ

[Eng98] Robert Englander. Programmare con JavaBeans. Jackson Libri, via XXV Aprile, 39 - 20091 Bresso (MI), Italy, February 1998. ISBN ???? LCCN ???? 45000 L.

Engel:1999:PJV

[Eng99] Joshua Engel. Programming for the Java Virtual Machine. Addison-Wesley, Reading, MA, USA, 1999. ISBN 0-201-30972-6 , 0-201-61654-8 (CD-ROM). xix + 488 pp. LCCN QA76.73.J38E543 1999. US$39.95.

Englander:19xx:DJa

[Engxx] Robert Englander. Developing JavaBeans. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 19xx. ISBN 4-900900-47-8. LCCN ???? 3,700

yen. URL http://www.ohmsha.
co.jp/data/books/contents/
4-900900-47-8.htm. Japanese

translation.

Entsminger:1997:WJ

[Ent97] Gary Entsminger. The Way of Java. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1997. ISBN 0-13-491978-5. xix + 392 pp. LCCN QA76.73.J38E58 1997. US$34.95. URL http://

                                        www.prenhall.com/allbooks/
                                        ptr_0134919785.html.

Erlenkotter:1997:JHS

[ER97] Helmut Erlenkötter and Volker Reher. Java. HTML, Scripts, Applets und Anwendungen. Rowohlt Taschenbuch, Reinbek bei Hamburg, Germany, 1997. ISBN 3-499-19872-X. 320 pp. LCCN ???? 20

                                        DM. URL http://www.abc.
                                        de/rowohlt.html;   http://
                                        www.rowohlt.de/rowohlt/rororo/
                                        klappen/2-97comp.htm;   http:
                                        //www.rowohlt.de/rowohlt/
                                        rororo/klappen/2-97comp.htm#
                                        Erlenk&#246;   tter2.

Erickson:1996:DDJ

[Eri96a] Jonathan Erickson. Dr. Dobb’s Journal Excellence in Programming awards. Dr. Dobb’s Journal of Software Tools, 21(3): 16–17, March 1996. CODEN DDJOEB. ISSN 1044-789X.

Eriksson:1996:ESW

[Eri96b] Henrik Eriksson. Expert systems and the Web: Expert systems as knowledge servers. IEEE expert: intelligent systems and their applications, 11 (3):14–19, June 1996. CODEN IEEXE7. ISSN 0885-9000. Discusses applications of Java.

El-Refai:1996:JDC

[ERKE96] M. Y. El-Refai, A. Kumar, and A. S. Elmaghraby. JAVAbased distributed conferencing for heterogeneous systems. In Yetongnon and Hariri [YH96], pages 2–?? ISBN ???? LCCN ???? Two volumes.

Espeset:1996:KJP

[Esp96] Tonny Espeset. KickAss Java Programming: Cutting Edge Java Techniques with an Attitude. Coriolis Group Books, Scottsdale, AZ, USA, August 1, 1996. ISBN 1-883577-99-3 (paperback). xii + 480 pp. LCCN QA76.73.J38E65 1996. US$39.99; CDN$55.99.

URL http://www.coriolis.
com/site/msie/books/ind/kajp.
htm;   http://www.sn.no/~
espeset; http://www.vmedia.
com/home.html. Includes CD-

ROM.

Espeset:19xx:HPJ

[Espxx] Tonny Espeset. High Performance JBuilder Programming. Coriolis Group Books, Scottsdale, AZ, USA, 19xx. ISBN 1-57610-067-7. 700 pp. LCCN ???? US$49.99, CDN$69.99.

URL http://www.coriolis.
com/.

Ess:1995:SMB

[Ess95] Robert Ess. Sun, Metrowerks brew Mac Java. MacWEEK, page ??, November 13, 1995. CODEN MWEEEI. ISSN 0892-

8118. URL http://www.ziff.
com/~macweek/mw_11-13-95/
sunmetro.html.

Etiel:1997:WDJ

[Eti97] Yoav Etiel. What does Java bring to engineering? Sun-Server, 11(10):22, October 1997. ISSN 1091-4986.

Eudida:1997:PGJ

[Eud97a] Eudida S.a.s. PC Guide Java. Jackson Libri, via XXV Aprile, 39 - 20091 Bresso (MI), Italy, November 1997. ISBN ???? ???? pp. LCCN ???? 49,000 L.

s:1997:PGJ

[Eud97b] Eudida S.a.s. PC Guide Java. Jackson Libri, via XXV Aprile, 39 - 20091 Bresso (MI), Italy, November 1997. ISBN ???? LCCN ???? 49,000 L.

Flanagan:1997:JN

[F^{+}97a] David Flanagan et al. Java in a Nutshell. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, deluxe edition, 1997. ISBN 1-56592-304-9. xvi + 610 pp. LCCN QA76.73.J38 J38 1997. US$69.95. URL http://www.

                                        ora.com/catalog/javadeluxe/
                                        ;   http://www.oreilly.com/
                                        catalog/9781565923041;   http:
                                        //www.oreilly.com/catalog/
                                        javadeluxe.

Flanagan:1997:JNDa

[F^{+}97b] David Flanagan et al. Java in a Nutshell, Deluxe Edition. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, deluxe edition, 1997. ISBN 1-56592-304-9. xvi + 610 pp. LCCN QA76.73.J38 J38 1997. US$69.95. URL http://www.

ora.com/catalog/javadeluxe/
;   http://www.ora.com/catalog/
{javadeluxe}/;   http://www.
oreilly.com/catalog/javadeluxe.

Feil:19xx:GLV

[F^{+}xxa] Peter Feil et al. El Gran Libro de Visual J++. Editorial Marcombo, S.A., Barcelona, Spain, 19xx. ISBN 84-267-1112-x. ???? pp. LCCN ???? 4200 Ptas.

Flanagan:19xx:PVJ

[F^{+}xxb] David Flanagan et al. Programování v jazyce Java. Computer Press, Brno-Bystrc, Czech Republic, 19xx. ISBN 80-85896-78-8 (??invalid ISBN??). ???? pp. LCCN ???? 270 Kè.

URL http://vltava.cpress.
cz/cgi-bin/vltava/genbook.
cgi?book=prg339;   http:/
/www.cpress.cz/;   http://
www.ora.com/catalog/jscriptbeta/

author.html. Czech translation of [Fla97b]. Includes CD-ROM.

Friedel:19xx:COD

[F^{+}xxc] David Friedel et al. Café:

Outils de Développment Visuel pour Java. International Thomson Computer Press, 20 Park Plaza Suite 1001, Boston, MA 02116 USA, 19xx. ISBN 2-84180-138-1. ???? pp. LCCN ???? 248 FF. URL http:

//www.thomson.com/intluk.

html. Includes CD-ROM.

Fahey:1996:TTS

[Fah96] Michael J. Fahey. Taligent Technology to support Java.

RS/Magazine, 5(12):12, December 1996. ISSN 1088-0844, 1061-0030. Discusses an agreement between Sun Microsystems Inc. and IBM’s Taligent Technology subsidiary to integrate the latter’s support for Java internationalization into the Java development environment.

Falla:1996:W

[Fal96]  Jane M. Falla.  Webmas-

tery. WebServer Magazine: For Managers of World Wide Web Sites, 1(1):38–40, 42, 44, 46, 48, 50, March/June 1996. ISSN 1087-4232. URL http://www.

                                        cpg.com.

Fanta:1996:JU

[Fan96] Rick Fanta. Java Unleashed. Howard W. Sams, Indianapolis, IN 46268, USA, April 1, 1996. ISBN 1-57521-049-5. 971 pp. LCCN QA76.73.J38J384 1996. US$49.99.

Farrell:1996:NJL

[Far96a] Tom Farrell. Netspeak: The origins and promise of the new lingua franca of distributed computing. IEE Review, 42(4):143– 146, July 18, 1996. CODEN IEREEF. ISSN 0953-5683 (or 0013-5127??).

Farrow:1996:REJ

[Far96b] Rik Farrow. Review: Exploring Java. ;login: the USENIX Association newsletter, 21(5):46– ??, October 1996. CODEN LOGNEM. ISSN 1044-6397.

Farrow:1996:UJa

[Far96c] Rik Farrow. Using Java. ;login: the USENIX Association newsletter, 21(5):39–42, October 1996. CODEN LOGNEM. ISSN 1044-6397.

Farrow:1996:UJb

[Far96d] Rik Farrow. Using Java. ;login: the USENIX Association newsletter, 21(6):24–26, December 1996. CODEN LOGNEM. ISSN 1044-6397.

Farrow:1996:UJN

[Far96e] Rik Farrow. Using Java: a new series of articles about Java. ;login: the USENIX Association newsletter, 21(5):39–??, October 1996. CODEN LOGNEM. ISSN 1044-6397.

Farrow:1997:UJa

[Far97a] Rik Farrow. Using Java. ;login: the USENIX Association newsletter, 22(2):19–23, April 1997. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

usenix.org/publications/java/
javaindex.html.

Farrow:1997:UJb

[Far97b] Rik Farrow. Using Java. ;login: the USENIX Association newsletter, 22(4):33–36, May 1997. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

usenix.org/publications/java/
javaindex.html.

Farrow:1997:UJc

[Far97c] Rik Farrow. Using Java. ;login: the USENIX Association newsletter, 22(5):32–35, June 1997. CODEN LOGNEM. ISSN 1044-6397.

Farrow:1997:UJd

[Far97d] Rik Farrow. Using Java. ;login: the USENIX Association newsletter, 22(4):33–36, August 1997. CODEN LOGNEM. ISSN 1044-6397.

Farley:1998:JDC

[Far98a] James Farley. Java Distributed Computing. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1998. ISBN 1-56592-206-9. xiv + 368 pp. LCCN QA76.9.D5F37 1998. US$30.

URL http://www.ora.com/
catalog/javadc/.

Farrow:1998:UJ

[Far98b] Rik Farrow. Using Java. ;login: the USENIX Association newsletter, 23(1):??, February 1998. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

usenix.org/publications/java/
javaindex.html;   http://
www.usenix.org/publications/
java/usingjava9.html.

Fraizer:1996:JAR

[FB96] Colin Fraizer and Jill Bond. Java API Reference: java.applet and java.awt API Packages. New Riders Publishing, Carmel, IN, USA, July 17, 1996. ISBN 1-56205-598-4. xxviii + 383 pp. LCCN QA76.73.J38 F73 1996. US$29.99. URL http://www.

mcp.com/295124552671766/cgi-

bin/bag?isbn=1-56205-598-
4&last=/bookstore;   http:
//www.mcp.com/nrp/.

FraminanTorres:1998:GBD

[FB98] José Manuel Framiñán Torres and José Miguel León Blanco. Gestion de Basas de Datos en Internet JDBC. Anaya Multimedia, Madrid, Spain, 1998. ISBN 84-415-0208-0. 544 pp. LCCN ???? 4,495

Ptas.  URL http://www2.
anaya.es/Catalogo/CGA?act=
L&art=2321018&ed=2300.

Fraizer:19xx:AJM

[FBxx] Colin Fraizer and Jill Bond.

API Java Manual De Referência: Pacotes de API java.applet e java.awt. Makron Books do Brasil Editora Ltda., Rua Tabapuã 1348, São Paulo SP - 04533-004 - Brasil, 19xx. ISBN 85-346-0758-3. 416 pp. LCCN ???? R$ 49,00. URL http://

safe.tesla.com.br/makron/
options/livro.asp?id=66;
http://www.makron.com/. Por-

tuguese translation of [FB96].

Fogarty:1995:NSR

[FC95] Kevin Fogarty and John Cox.

News: SunSoft rolls out object offering. Network World, page ??, September 22, 1995. ISSN 0887-7661 (print), 1944-7655 (electronic). URL http://

www.iquest.net/index.html.

Flynn:1996:HJM

[FC96a] Jim Flynn and Bill Clarke. How Java makes network-centric computing real — Java is likely to have a profound impact on the way businesses manage application development by bringing about an entirely new model — the network-centric model — of C/S computing. Datamation, 42(5):42–43, ???? 1996. CO-DEN DTMNAT. ISSN 0011-6963.

Flynn:1996:VJJ

[FC96b] Jim Flynn and Bill Clarke. Visual J++ Java Programming. New Riders Publishing, Carmel, IN, USA, December 1996. ISBN 1-56205-602-6. xvi + 666 pp. LCCN QA76.73.J38F56 1996. US$45.00. URL http:

//www.mcp.com/cgi-bin/bag?
isbn=1-56205-602-6&last=/
bookstore;   http://www.mcp.
com/nrp/.

Flynn:1998:VJP

[FC98a] Jim Flynn and Bill Clarke. Visual J++: Programando em Java. Makron Books do Brasil Editora Ltda., Rua Tabapuã 1348, São Paulo SP - 04533-004 - Brasil, April 1998. ISBN 85-346-0805-9 (??invalid ISBN??). LCCN ???? R$89. URL http:/

/safe.tesla.com.br/makron/
options/livro.asp?id=681.

Fong:1998:PLA

[FC98b] Philip W. L. Fong and Robert D. Cameron. Proof linking: an architecture for modular verification of dynamically-linked mobile code. ACM SIGSOFT Software Engineering Notes, 23(6):

222–230, November 1998. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Feather:1997:UJ

[FCD97] Stephen Feather and Luke Cassady-Dorion. Usare JavaScript. Jackson Libri, via XXV Aprile, 39 - 20091 Bresso (MI), Italy, 1997. ISBN ???? ???? pp. LCCN ???? 65000 Lire.

URL http://pallade.iol.
it/cgi-bin/cmsearch.pl;
http://pallade.iol.it/cgi-
bin/mercator2/mallurl/AMARCRUDSFAMBATTCORSa001EH/
bookstore/jackson/001886.
html;   http://www.jacksonlibri.

it/. Italian translation of [FCDW96]. Includes CD-ROM.

Feather:1996:JE

[FCDW96] Stephen Feather, Luke Cassady-Dorion, and Andrew Wooldridge. JavaScript by Example. Que Corporation, Indianapolis, IN, USA, July 1, 1996. ISBN 0-7897-0813-2. xviii + 410 pp. LCCN QA76.73.J39F43 1996. US$34.99. Also available in Italian translation [FCD97].

Freedman:1996:JPD

[FD96a] Edie (?) Freedman and ?. Di-Giorgio. Java Programming. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1996. ISBN 1-56592-185-2. ???? pp. LCCN ???? URL http:

//www.oreilly.com/catalog/
9781565921856.

Freedman:1996:JP

[FD96b] Edie (?) Freedman and Rinaldo DiGiorgio. Java Programming. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1996. ISBN 1-56592-185-2. ???? pp. LCCN

                                        ???? ???? URL http://www.
                                        ora.com/.

Freedman:1996:NCA

[FD96c] R. S. Freedman and R. Di Giorgio. New computational architectures for pricing derivatives. In IEEE [IEE96p], pages 14–19. ISBN 0-7803-3236-9 (softbound), 0-7803-3237-7 (microfiche). LCCN HG176.7.I33 1996. IEEE Cat. No. 96TH8177.

Feather:1997:J

[Fea97] Stephen Feather. JavaScript. Softbank Publishing, ??, Japan, 1997. ISBN ???? ???? pp. LCCN ???? 3300 yen.

                                        URL http://www.softbank.
                                        co.jp/sbnet/books/editer/
                                        4th/jsp/;   http://www.softbank.
                                        co.jp/softbank/publishing/

.

Fedyk:1997:LEJ

[Fed97] Roger Fedyk. Letter to the editor: Java unveiled. BYTE Magazine, 22(6):19, June 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Feghhi:1997:WDG

[Feg97]  Jalal Feghhi.  Web Devel-

oper’s Guide to Java Beans.

Coriolis Group Books, Scottsdale, AZ, USA, 1997. ISBN 1-57610-121-5. xvii + 398 pp. LCCN QA76.73.J38F44 1997. US$39.99; CDN$55.99.

URL http://www.coriolis.
com/cgi-win/KBSearch.exe;
http://www.coriolis.com/
Site/MSIE/Books/Ind/wdgjb.
htm.

Feghhi:19xx:WDGb

[Fegxx] Jalal Feghhi. Web Developer’s Guide to JavaBeans. La Nuova Italia, Firenze, Italy, 19xx. ISBN 81-576-1012-1 (??Invalid checksum??). ???? pp. LCCN ???? 107000 L.

Feiler:1998:CIJ

[Fei98]  Jesse Feiler.  ClarisWorks:

The Internet, Java and Paperless Documents. AP Professional, Boston, MA, USA, January 1998. ISBN 0-12-251332-0 (softcover). xxiii + 348 pp. LCCN QA76.76.I57F44 1997. US$31.45.

Felten:1997:IRW

[Fel97] Edward W. Felten. Inside risks: Webware security. Communications of the ACM, 40 (4):130, April 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org/
pubs/citations/journals/cacm/
1997-40-4/p130-felten/.

Ferranti:1995:NIF

[Fer95] Marc Ferranti. News: IT Forum: Ellison, Gates jockey for position on info highway.

IDG International News Service, page ??, September 04, 1995.

Ferrari:1998:JNPb

[Fer98a] Adam Ferrari. JPVM: network parallel computing in Java. Concurrency: Practice and Experience, 10(11– 13):985–992, September 1998. CODEN CPEXEI. ISSN 1040-3108. URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=10050413;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=10050413&PLACEBO=IE.pdf.

Special Issue: Java for Highperformance Network Computing.

Ferrari:1998:JNPa

[Fer98b] Adam J. Ferrari. JPVM: Network parallel computing in Java. In ACM [ACM98a], page ?? CODEN CPEXEI. ISSN 1040-3108. LCCN ????

URL http://www.cs.ucsb.
edu/conferences/java98/papers/
jpvm.pdf;   http://www.cs.
ucsb.edu/conferences/java98/
papers/jpvm.ps. Also pub-

lished as Concurrency: Practice and Experience, 10(11– 13), September 1998, CODEN CPEXEI, ISSN 1040-3108.

FerrerAbello:19xx:JLI

[Ferxx] Antonio M. Ferrer Abelló. Java el Lenguaje de Internet. Abeto Editorial, Marqués de Portugalete 10, 28017 Madrid, Spain, 19xx. ISBN ???? ???? pp.

LCCN ???? URL http://

www.virtualsw.es/abeto/ (?

?). Includes CD-ROM.

Fetik:1997:JWS

[Fet97] Richard Fetik. JavaScript Web site overview: JavaScript makes it easy to create dynamic sites that customize pages to the current user’s interaction with the site. Web Apps Magazine, 1(1): 35–43, January/February 1997. ISSN 1090-2287.

Fisco:1996:DGJ

[FF96a] David Fisco and Patrick Fitzgerald. A Developer’s Guide to JavaScript, volume 2 of Java Report Professional Series. SIGS Books and Multimedia, 71 W 23rd St., New York, NY 10010, USA, 1996. ISBN 0-13-579269-X. 200 pp. LCCN ???? US$24.95. URL http://www.

sigs.com/books/javascript.
html.

Fox:1996:TWJ

[FF96b] G. Fox and W. Furmanski.

Towards Web/Java-based high performance distributed computing — an evolving virtual machine. In IEEE [IEE96n], page ?? ISBN 0-8186-7582-9. LCCN QA76.88.I52 1996. IEEE order plan catalog number 96TB100069. EEE Computer Society Press order number PR07582.

Fox:1997:JPC

[FF97] Geoffrey C. Fox and Wojtek Furmanski. Java for parallel computing and as a general language for scientific and engineering simulation and modeling. Concurrency: Practice and Experience, 9(6):415–425, June 1997. CODEN CPEXEI. ISSN 1040-3108. URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=13868;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=13868&PLACEBO=IE.pdf.

Felleisen:1998:LJF

[FF98] Matthias Felleisen and Daniel P. Friedman. A little Java, a few patterns. MIT Press, Cambridge, MA, USA, 1998. ISBN 0-262-56115-8. xiv + 179 pp. LCCN QA76.73.J38F45 1998. Drawings by Duane Bibby, and foreword by Ralph E. Johnson.

Flanagan:1999:JEN

[FFCM99] David Flanagan, Jim Farley, William Crawford, and Kris Magnusson. Java Enterprise in a Nutshell: a Desktop Quick Reference. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1999. ISBN 1-56592-483-5. xvi + 604 pp. LCCN QA76.73.J38 J366 1999. US$29.95. URL http:

//www.oreilly.com/catalog/
jentnut.

For:1998:HHP

[FFHA98] G. C. For, W. Furmanski, T. Haupt, and E. Akarsu. HPcc as high performance commodity computing on top of integrated Java, CORBA, COM and Web standards. Lecture Notes in Computer Science, 1470:55–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fox:1999:UWI

[FFKO99] G. C. Fox, W. Furmanski, G. Krishnamurthy, and H. T. Ozdemir. Using WebHLA to integrate HPC FMS modules with Web/commodity based distributed object technologies of CORBA, Java, COM and XML. In Tentner [Ten99], pages 273–278. ISBN 1-56555-166-4. LCCN ????

Frasson:1996:ITS

[FGL96] Claude Frasson, Gilles Gauthier, and Alan Lesgold, editors. Intelligent tutoring systems: Third International Conference, ITS ’96 Montréal, Canada, June 12–14, 1996: proceedings, volume 1086 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 3-540-61327-7. LCCN LB1028.73 .I58 1996.

Flanagan:1997:JRL

[FGN^{+}97] David Flanagan, Mark Grand, Pat Niemeyer, Josh Peck, and John Zukowski. Java Reference Library on the Web. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, November 1997. ISBN 1-56592-359-6. 4000 (est.) pp. LCCN ????

US$59.95. URL http://www.

                                        ora.com/catalog/javarlw/.

Freeman:1999:JPP

[FHA99] Eric Freeman, Susanne Hupfer, and Ken Arnold. JavaSpaces principles, patterns, and practice. Addison-Wesley, Reading, MA, USA, 1999. ISBN 0-201-30955-6. 304 pp. LCCN QA76.73.J38F74 1999. US$39.95.

Firesmith:1998:DCJ

[FHKS98] Donald G. Firesmith, Greg Hendley, Scott Krutsch, and Marshall Stowe. Documenting a complete Java application using OPEN. Addison-Wes-ley, Reading, MA, USA, 1998. ISBN 0-201-34277-4. xxiv + 404 pp. LCCN QA76.64.D63 1998. US$41.35. Includes CD-ROM.

Field:1996:JDW

[FI96a] Field and Marketing Communications and Indiana University-Purdue University Indianapolis.Office of and Integrated Technologies. Java development on the Web, March 28, 1996. 1 videocassette (60 min.).

Freeman:1996:AJO

[FI96b] Adam Freeman and Darrel Ince. Active Java: Object-Oriented Programming for the World Wide Web. Addison-Wesley, Reading, MA, USA, March 1, 1996. ISBN 0-201-40370-6. xii + 235 pp. LCCN QA76.73.J38 F74 1996. US$25.77. URL http://

www.aw.com/cseng/authors/
freeman.a/actjava/.

Fichtelman:1996:TJG

[Fic96] M. Fichtelman. The Tao of Java: generating random hexagrams. Web Techniques, 1(7): 55–57, October 1996. CODEN WETEFA. ISSN 1086-556X.

FMC:1996:JDW

[Fie96] Field and Marketing Communications. Java development on the Web, 1996. 1 videocassette (60 min.).

Fisher:1996:SWG

[Fis96] Yuval Fisher. Spinning the Web:

a Guide to Serving Information on the World Wide Web. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 0-387-94539-3. xiv + 536 pp. LCCN TK5105.888.F58 1996. US$27.95. Discussions Java, JavaScript, and VRML.

Fish:1998:OPE

[Fis98a] David Fish. Official PowerJ:

Enterprise Applications Development. Ventana Press, Chapel Hill, NC, USA, February 1998. ISBN 1-85032-921-4. LCCN ???? US$45.

Fisher:1998:BJ

[Fis98b] Yuval Fisher. Brewing JavaScript.

Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., December 1, 1998. ISBN 0-387-94816-3. 350 pp. LCCN ???? US$29.95.

Flatt:1998:CM

[FKF98] Matthew Flatt, Shriram Krishnamurthi, and Matthias Felleisen. Classes and mixins. In ACM [ACM98b], pages 171–183. ISBN 0-89791-979-3. LCCN QA76.7 .A15 1998. URL http://www.

acm.org:80/pubs/citations/
proceedings/plan/268946/p171-

flatt/. ACM order number: 549981.

Friedel:1997:JPC

[FKP97] David H. Friedel, Joshua Kerievsky, and Anthony Potts. Java-Programmierung mit Café. International Thomson Computer Press, 20 Park Plaza Suite 1001, Boston, MA 02116 USA, 1997. ISBN 3-8266-0315-X (??invalid ISBN??). 461

pp.  LCCN ????  50

DM. URL http://www.itp.
de/;   http://www.thomson.
com/intluk.html.

Friedel:1996:CPF

[FKPR96] David H. Friedel, Jr., Joshua Kerievsky, Anthony Potts, and John Rodley. Café Programming FrontRunner. Coriolis Group Books, Scottsdale, AZ, USA, 1996. ISBN 1-57610-003-0. xvii + 430 pp. LCCN QA76.73.J38C34 1996. US$29.99. URL http:

//www.alexandria-home.com/
library/1-57610-003-0cover.
htm;   http://www.coriolis.
com/;   http://www.coriolis.
com/site/msie/books/ind/cafefrt.
htm.

Flanagan:1995:JND

[Fla95]  David Flanagan.  Java in

a Nutshell: a Desktop Quick Reference for Java Programmers. A Nutshell handbook. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1995. ISBN 1-56592-183-6. xix + 438 pp. LCCN QA76.73.J38F553 1996. US$19.95. URL http://www.

ora.com/info/java.

Flammia:1996:ISS

[Fla96a] G. Flammia. On the Internet, software should be milked, not brewed. IEEE expert: intelligent systems and their applications, 11(6):87–88, December 1996. CODEN IEEXE7. ISSN 0885-9000.

Flanagan:1996:JN

[Fla96b] David Flanagan. Java in a Nutshell. Specialized Systems Consultants, P.O. Box 55549, Seattle, WA 98155, December 1996. ISBN 1-55659-218-3 (softcover). ???? pp. LCCN ???? US$19.95.

Flanagan:1996:JNV

[Fla96c] David Flanagan. Java in a Nutshell, Version Française. Editions O’Reilly International Thomson, 981 Chestnut Street, Newton, MA 02164, USA, October 1996. ISBN 2-84177-009-5. xxii + 492 pp. LCCN ???? 220 FF.

URL http://www.editions-
oreilly.fr/international/
france/prog/javanut.html;
 http://www.ora.de/. French

translation by Serge Chaumette, Georges Eyrolles and Alain Miniussi. Includes CD-ROM.

Flanagan:1996:JDG

[Fla96d] David Flanagan. JavaScript: The Definitive Guide. A Nutshell handbook. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, beta edition, 1996. ISBN 1-56592-193-3. xv + 454 pp. LCCN QA76.73.J38 F555 1996; QA76.73.J39 F53 1996. US$29.95. URL http:

                                        //www.oreilly.com/catalog/
                                        9781565921931.

Flanagan:1996:JDGa

[Fla96e] David Flanagan. JavaScript: The Definitive Guide: Beta Edition. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, beta edition, August 1996. ISBN 1-56592-193-3. xv + 454 pp. LCCN QA76.73.J38F555 1996. US$29.95.

Flanagan:1996:HTT

[Fla96f] Patrick Flanagan. 10 hottest technologies in telecom. Telecommunications (Americas Edition), 30(5):??, May 1996. CO-DEN TLCMDV. ISSN 0278-4831.

Flanagan:1997:JEN

[Fla97a] David Flanagan. Java Examples in a Nutshell: a Companion Volume to Java in a Nutshell. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1997. ISBN 1-56592-371-5. xiv + 397 pp. LCCN QA76.73.J38F552 1997. US$19.95. URL http://www.

ora.com/catalog/jenut/.

Flanagan:1997:JNa

[Fla97b] David Flanagan. Java in a Nutshell. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, October 1997. ISBN 3-89721-100-9 (??invalid ISBN??). LCCN ???? 49 DM.

URL http://www.oreilly.
de/german/java/javanut2/.

Deutsche Ausgabe.

Flanagan:1997:JNb

[Fla97c] David Flanagan. Java in a Nutshell. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1997. ISBN 4-900900-08-7. ???? pp. LCCN ???? 3900 yen.

URL http://www.ora.com/
catalog/jscriptbeta/author.
html;   http://www.ora.com/
international/catalog/japanese/
inet/java.html;   http://
www.ora.com/international/
japan/;   http://www.oreilly.
co.jp/BOOK/java.htm. Japanese

translation.

Flanagan:1997:JND

[Fla97d] David Flanagan. Java in a Nutshell: a desktop quick reference. The Java series. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 1997. ISBN 1-56592-262-X. xvi + 610 pp. LCCN QA76.73.J38 F553 1997. US$24.95. URL http:

                                        //www.oreilly.com/catalog/
                                        9781565922624;   http://www.
                                        oreilly.com/catalog/doctor;
                                         http://www.oreilly.com/catalog/
                                        javanut2.

Flanagan:1997:JNDb

[Fla97e] David Flanagan. Java in a Nutshell: a desktop quick reference. Java Ser. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 1997. ISBN 1-56592-262-X, 1-56592-273-5. xvi + 610 pp. LCCN QA76.73.J38F553 1997. US$20.

                                        URL http://www.ora.com/
                                        catalog/javanut2/noframes.
                                        html.

Flanagan:1997:JTO

[Fla97f] David Flanagan. Java tehokóyttójón opas. Suomen Atk-kustannus Oy, ????, December 1997. ISBN 951-762-494-8. LCCN ???? ????

                                        URL http://www.satku.fi/
                                        kirjat/4javateho.htm.

Flanagan:1997:JDG

[Fla97g] David Flanagan. JavaScript: The Definitive Guide. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 1997. ISBN 1-56592-234-4. xiv + 647 pp. LCCN QA76.73.J39 F53 1997; QA76.73.V38 F53 1997.

                                        URL http://www.oreilly.
                                        com/catalog/9781565922341;
                                         http://www.oreilly.com/catalog/
                                        jscript2.

Flanagan:1997:RJD

[Fla97h]  David Flanagan.  Review:

JavaScript, the definitive guide (beta edition). ;login: the USENIX Association newsletter, 22(1):60–61, February 1997. CODEN LOGNEM. ISSN 1044-6397.

Flanagan:1998:JEN

[Fla98a] David Flanagan. Java Examples in a Nutshell — Deutsche Ausgabe. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, March 1998. ISBN 3-89721-112-2 (??invalid ISBN??). LCCN ???? 49 DM.

URL http://www.oreilly.
de/german/java/jenut/.

Flanagan:1998:JNDa

[Fla98b] David Flanagan. Java in a Nutshell, Deutsche Ausgabe. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1998. ISBN 3-930673-46-0 (??invalid ISBN??). ???? pp. LCCN ???? 39 DM. URL http:

//www.ora.com/gnn/bus/ora/
item/javanut.author.html;
http://www.ora.de/.

Flanagan:1998:JNDb

[Fla98c] David Flanagan. Java in a Nutshell (Deutsche Ausgabe), 2. Auflage. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1998. ISBN 3-89721-100-9. xvi + 638 pp. LCCN ???? 49 DM.

URL http://www.oreilly.
de/german/java/javanut2/.

Flanagan:1998:JPR

[Fla98d] David Flanagan. JavaScript Pocket Reference. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1998. ISBN 1-56592-521-1. v + 89 pp. LCCN QA76.73.J39 F52 1998. US$7.95. URL http:

//www.oreilly.com/catalog/
jscriptpr.

Flanagan:1998:JDG

[Fla98e] David Flanagan. JavaScript: The Definitive Guide. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, third edition, June 1998. ISBN 1-56592-392-8. xiv + 776 pp. LCCN A76.73.J39 F53 1998. US$39.95. URL http:

//www.oreilly.com/catalog/
jscript3.

Flanagan:1999:MRJ

[Fla99a] D. Flanagan. Micro review: Java power reference. IEEE Micro, 19(5):10–11, September/October 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Flanagan:1999:JFC

[Fla99b] David Flanagan. Java Foundation Classes in a Nutshell. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1999. ISBN 1-56592-488-6. xiii + 731 pp. LCCN QA76.73.J38 F53 1999. US$29.95. URL http:

//www.oreilly.com/catalog/
jfcnut.

Flanagan:1999:JNE

[Fla99c] David Flanagan. Java in a Nutshell: a Desktop Quick Reference. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, third edition, November 1999. ISBN 1-56592-487-8. xviii + 648 pp. LCCN QA76.73.J38 F553 1999. US$29.95. URL http:

//www.oreilly.com/catalog/
9781565924871;   http://www.
oreilly.com/catalog/javanut3.

Flanagan:1999:JPR

[Fla99d] David Flanagan. Java Power Reference: a Complete Searchable Resource on CD-ROM. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1999. ISBN 1-56592-589-0. xi + 51 pp. LCCN ???? US$19.95. URL http:

//www.oreilly.com/catalog/

jpower/. Includes CD-ROM.

Flanagan:19xx:JENa

[Flaxxa] David Flanagan. Java Examples in a Nutshell. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 19xx. ISBN 4-900900-09-5. LCCN ???? 4,300 yen. Japanese translation.

Flanagan:19xx:JNC

[Flaxxb] David Flanagan. Java in a Nutshell (Chinese translation by Liao Tze Shien). Image Art Publishing, ????, 19xx. ISBN 957-9641-16-1. ???? pp. LCCN ???? 520 NT$. Chinese translation by Liao Tze Shien.

Flint:1998:UJA

[Fli98]  Shayne Flint.  Using Java

APIs with native Ada compilers. ACM SIGADA Ada Letters, 18(6):193–203, November/December 1998. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Flohr:1997:IIS

[Flo97] Udo Flohr. Intelligent intranets — some tips and products that will help you avoid anarchy. BYTE Magazine, 22(8):69–??, August 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Fox:1997:PFJ

[FLQZ97] Geoffrey Fox, Xiaoming Li, Zheng Qiang, and Wu Zhigang. A prototype of Fortranto-Java converter. Concurrency: Practice and Experience, 9(11):1047–1061, November 1997. CODEN CPEXEI. ISSN 1040-3108. URL http:

                                        //www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=13832;
                                         http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=13832&PLACEBO=IE.pdf. Spe-

cial Issue: Java for computational science and engineering — simulation and modeling II.

Fiorentino:1997:CJL

[FLRT97] Giuseppe Fiorentino, Maria Rita Laganà, Francesco Romani, and Franco Turini. C E Java Laboratorio di Programmazione. McGraw-Hill, New York, NY, USA, September 1997. ISBN 88-386-0739-7. LCCN ???? 37000 L.

Fraize:1996:JSP

[FLW96] Scott Fraize, Chris Laurel, and Ryan Watkins. Java Style: Programming Web Pages with Java/HotJava. Ziff-Davis Press, Emeryville, CA, USA, 1996. ISBN 1-56276-368-7. 400 pp. LCCN ???? US$35. URL http:

//www.mcp.com/zdpress.

Flynn:1995:BN

[Fly95] Mary-Kathleen Flynn. The battle for the Net. U.S. news and world report, 119:55–56, December 18, 1995. CODEN XN-WRAV. ISSN 0041-5537.

Fronckowiak:1997:TYD

[FM97] John Fronckowiak and Gordon B. McMillan. Teach Yourself Database Programming With Visual J++ In 21 Days. Sams.net, Indianapolis, IN 46268, USA, 1997. ISBN 1-57521-262-5. xxiii + 774 pp. LCCN QA76.73.J38F76 1997. US$40. URL http://

merchant.superlibrary.com:
8000/catalog/hg/PRODUCT/PAGE/
15752/bud/1575212625.html.

Freund:1998:TSO

[FM98] Stephen N. Freund and John C. Mitchell. A type system for object initialization in the Java bytecode language. ACM SIG-PLAN Notices, 33(10):310–328, October 1998. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Freund:1999:FFJ

[FM99a] Stephen N. Freund and John C. Mitchell. A formal framework for the Java bytecode language and verifier. ACM SIGPLAN Notices, 34 (10):147–166, October 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

URL http://www.acm.org/
pubs/citations/proceedings/
oops/320384/p147-freund/.

Freund:1999:TSO

[FM99b] Stephen N. Freund and John C. Mitchell. The type system for object initialization in the Java bytecode language. ACM Transactions on Programming Languages and Systems, 21(6):1196–1250, November 1999. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL http://www.acm.org/
pubs/citations/journals/toplas/
1999-21-6/p1196-freund/.

Folmsbee:1999:AJT

[Fol99] Alan Folmsbee. AES Java technology comparisons. In National Institute of Standards and Technology [Nat99], page 28. ISBN ???? LCCN

???? URL http://csrc.nist.
gov/encryption/aes/round1/
conf2/Folmsbee.pdf. Only the

slides for the conference talk are available.

Forestier:1996:DMW

[For96] J.-P. Forestier. Data management in the Web context. Databases Journal, ????(4-5):8–15, July–October 1996. CODEN DAJOFP.

Ford:1998:MES

[For98a] Dan Ford. Mutual exclusion and synchronization in Java: Mimicking Win32 synchronization mechanisms. Dr. Dobb’s Journal of Software Tools, 23(1):62, 64, 66, 68–70, 73–75, January 1998. CODEN DDJOEB. ISSN 1044-789X.

Forster:1998:JP

[For98b]  David Forster.  The Java

Primer. Addison-Wesley, Reading, MA, USA, January 1998. ISBN 0-201-58685-1. ???? pp. LCCN ???? US$10.

Forte:1998:JRS

[For98c] Dario Forte. Is Java really secure? Network Security, 1998 (5):15–19, May 1998. CO-DEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S1353485898801220.

Forte:1998:SMJ

[For98d]  Dario Forte.  The security

model of JDK 1.2. ;login: the USENIX Association newsletter, 23(6):??, October 1998. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

usenix.org/publications/login/
1998-10/jdk.html.

Fox:1997:EJC

[Fox97a]  Geoffrey Fox.  Editorial:

Java for computational science and engineering — simulation and modeling. Concurrency: Practice and Experience, 9(6):413–414, June 1997. CODEN CPEXEI. ISSN 1040-3108. URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=13882;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=13882&PLACEBO=IE.pdf.

Fox:1997:JCS

[Fox97b] Geoffrey Fox. Java for computational science and engineering — simulation and modeling II. Concurrency: Practice and Experience, 9(11): 1001–1002, November 1997. CODEN CPEXEI. ISSN 1040-3108. URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=13845;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=13845&PLACEBO=IE.pdf. Spe-

cial Issue: Java for computational science and engineering — simulation and modeling II.

Fox:1997:JKW

[Fox97c] Jim Fox. Is Java the key to Web animation? SunExpert Magazine, 8(6):72, 75, 76, 78, 80, June 1997. ISSN 1053-9239.

Fox:1998:EJH

[Fox98a] Geoffrey Fox. Editorial: Java for high-performance network computing. Concurrency: Practice and Experience, 10(11– 13):821–824, September 1998. CODEN CPEXEI. ISSN 1040-3108. URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=10050427;
http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=10050427&PLACEBO=IE.pdf.

Special Issue: Java for Highperformance Network Computing.

Fox:1998:JHP

[Fox98b] Geoffrey C. Fox, editor. Java for high-performance network computing: [proceedings of the third of a set of workshops studying the use of Java in large scale computations, Palo Alto, CA], volume 10(11/13) of Concurrency: Practice and Experience. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 1998. CODEN CPEXEI. ISSN 1040-3108.

Foxwell:1999:JSD

[Fox99] Harry J. Foxwell. Java 2 software development kit. Linux Journal, 66:??, October 1999. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http://

noframes.linuxjournal.com/
lj-issues/issue66/3424.html.

Friedel:1996:JP

[FP96] David H. Friedel and Anthony Potts. Das Java-Programmierhandbuch. tewi, München, Germany, 1996. ISBN 3-89362-446-5. 430 pp. LCCN ???? 79 DM.

URL http://www.tewi.de/
scripts/tewi/details_ISBN.
idc?ISBN=3-89362-446-5. In-

cludes CD-ROM.

Friedel:1998:JPL

[FP98] David Friedel and Anthony Potts. Java Programming Language Handbook. Impress Corporation, ??, Japan, March 1998. ISBN 4-8443-4865-5. LCCN ???? 3,200 yen.

URL http://www.ips.co.jp/
book/4865/4865.htm.

Friedel:19xx:JPC

[FPxx] David Friedel and Anthony Potts. Java-Programmierung mit Café. International Thomson Computer Press, 20 Park Plaza Suite 1001, Boston, MA 02116 USA, 19xx. ISBN 3-8266-0315-X (??invalid ISBN??). LCCN ???? 50 DM. URL

http://www.itp.de/progra/
0315/0315.html.

Fawcett:1997:JPP

[FR97] Neil Fawcett and Terry Ridge. Java Programming: a Practical Guide. Butterworth-Heinemann, Oxford, England, July 1997. ISBN 0-7506-3344-1 (softcover). x + 434 pp. LCCN QA76.73.J38F39 1997. US$26.95.

Fraize:1996:PMP

[Fra96a] Scott Fraize. PC Magazine Programming Java Applets. Ziff-Davis Press, Emeryville, CA, USA, March 1996. ISBN ???? ???? pp. LCCN ???? US$34.99.

FraminanTorres:1996:JMI

[Fra96b] José Manuel Framiñán Torres. Java: Manuales Imprescindibles. Anaya Multimedia, Madrid, Spain, 1996. ISBN 84-415-0056-8. 336 pp. LCCN ???? 2995

Ptas. URL http://www.anaya.
es/Catalogo/CGA?act=L&art=
2311044&ed=2300;   http:/
/www.anaya.es/multimedia/;
http://www.anayamultimedia.
es/cgi-multi/notas_prensa.
pl?2311044&*2&torres&1.

FraminanTorres:1997:JAD

[Fra97] José Manuel Framiñán Torres.

Java: Al Día en una Hora. Anaya Multimedia, Madrid, Spain, 1997. ISBN 84-415-0061-4. 128 pp. LCCN ???? 495

Ptas. URL http://www.anaya.
es/Catalogo/CGA?act=L&art=
2316035&ed=2300;   http:/
/www.anaya.es/multimedia/;
http://www.anayamultimedia.
es/cgi-multi/notas_prensa.
pl?2316035&*2&torres&1.

Franz:1998:OSB

[Fra98a] M. Franz. Open standards beyond Java: On the future of mobile code for the Internet. J.UCS: Journal of Universal Computer Science, 4(5):522–??, May 28, 1998. CODEN ???? ISSN 0948-6968. URL http:

//medoc.springer.de:8000/
jucs/jucs_4_5/open_standards_
beyond_java;   internal&sk=
0C220489.

Franz:1998:JVM

[Fra98b] Michael Franz. The Java Virtual Machine: a passing fad? IEEE Software, 15(6):26–29, November/December 1998. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

URL http://computer.org/
software/so1998/s6026abs.
htm;   http://dlib.computer.
org/so/books/so1998/pdf/s6026.
pdf.

Frentzen:1995:JBO

[Fre95] Jeffrey Frentzen. Java brewing online publishing change. PC Week Online, October 23, 1995.

URL http://www.zdnet.com/
~pcweek/navigator/1023/nav1023.
html.

Freed:1996:CVC

[Fre96a] G. Freed. Captioning video clips in the World Wide Web. First Monday, 1(3):??, ???? 2, 1996. ISSN 1396-0466.

Freytag:1996:JIA

[Fre96b]  A. Freytag.  Java interna-

tionalization architecture. In Unicode Consortium [Uni96b], pages A3–?? ISBN ????

LCCN ???? URL http:

//www.unicode.org/unicode/
iuc9conf/index.html;   http:
//www.unicode.org/unicode/
uni2book/u2ord.html.

Frenger:1997:RRL

[Fre97a] Paul Frenger. A review of robotics languages. ACM SIG-PLAN Notices, 32(4):27–31, April 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Discusses PostScript, Ghostscript, Tps (transportable PostScript), Java, and others.

Frentzen:1997:JPA

[Fre97b] Jeff Frentzen. JavaScript Programmer’s Annotated Archives. Osborne/McGraw-Hill, Berkeley, CA, USA, November 1997. ISBN 0-07-882364-1 (softcover). ???? pp. LCCN ???? US$39.99. Includes CD-ROM.

Freytag:1997:JIAa

[Fre97c] Asmus Freytag. Java internationalization architecture. In Unicode Consortium [Uni97a], page ?? ISBN ???? LCCN

????  URL http://www.
unicode.org/unicode/iuc10/
program.html.

Freytag:1997:JIAb

[Fre97d] Asmus Freytag. Java internationalization architecture — continued. In Unicode Consortium [Uni97a], page ?? ISBN ???? LCCN ???? URL http:

//www.unicode.org/unicode/
iuc10/program.html.

Frick:1996:FDI

[Fri96] O. Frick. Formal description and interpretation of coordination protocols for teamwork. In Spaniol et al. [SLPM96], page ?? ISBN 3-540-61842-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D5 I63 1996.

Friesenhahn:1998:PWJ

[Fri98] Bob Friesenhahn. Programming — writing JavaScript applications — some tricks for building useful applications, and some ways to get around this language’s limitations. BYTE Magazine, 23(2):59–60, February 1998. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Friedel:19xx:CPFb

[Frixx] David Friedel. Café Programming FrontRunner. Editorial Paraninfo, Madrid, Spain, 19xx. ISBN 84-283-2413-1. ???? pp. LCCN ???? ????

Frohlich:1996:LNS

[Frö96] Marcus Fröhlich. Lernen Sie in Nur 14 Stunden Java Applets zu Programmieren und Einzusetzen. te-wi, München, Germany, 1996. ISBN 3-89362-465-1. 191 pp. LCCN ???? 39 DM.

                                        URL http://www.tewi.de/
                                        scripts/tewi/details_ISBN.
                                        idc?ISBN=3-89362-465-4. In-

cludes CD-ROM.

Froherz:19xx:TPJ

[Froxxa] Michael Froherz. Tönne: Projekte mit Java. International Thomson Computer Press, 20 Park Plaza Suite 1001, Boston, MA 02116 USA, 19xx. ISBN 3-8266-0316-8 (??invalid ISBN??). LCCN ???? 69 DM.

Frohlich:19xx:LNS

[Fröxxb] Marcus Fröhlich. Lernen Sie in Nur 14 Stunden Java Applets zu Programmieren und Einzusetzen. te-wi, München, Germany, 19xx. ISBN 3-89362-465-4. LCCN ???? 39 DM.

URL http://www.tewi.de/
scripts/tewi/details_ISBN.
idc?ISBN=3-89362-465-4.

Fan:1996:BAJ

[FRT96] Joel Fan, Eric Ries, and Calin Tenitchi. Black Art of Java Game Programming. Black Art Ser. Waite Group Press, Corte Madera, CA, USA, November 1, 1996. ISBN 1-57169-043-3. xxxi + 933 pp. LCCN QA76.73.J38F36 1996. US$49.99. URL http:

//www.mcp.com/cgi-bin/bag?
isbn=1-57169-043-3&last=/
bookstore;   http://www.waite.
com/waite/.

Fruhauf:1999:BCO

[Frü99] Karol Frühauf. Bookshelf: Customer oriented software quality assurance: Client/server programming with JAVA and CORBA. IEEE Software, 16(5): 131–133, September/October 1999. CODEN IESOEG.

ISSN 0740-7459 (print), 0740-7459 (electronic). URL

                                        http://dlib.computer.org/
                                        so/books/so1999/pdf/s5131.
                                        pdf.

Fan:1998:CEM

[FSW98] Ming Fan, Jan Stallaert, and Andrew B. Whinston. Creating electronic markets. Dr. Dobb’s Journal of Software Tools, 23 (11):52, 54–57, November 1998. CODEN DDJOEB. ISSN 1044-

                                        789X. URL http://www.ddj.
                                        com/ftp/1998/1998_11/fbts.
                                        txt.

Foster:1996:ETW

[FT96] I. Foster and S. Tuecke. Enabling technologies for Webbased ubiquitous supercomputing. In IEEE [IEE96n], page ?? ISBN 0-8186-7582-9. LCCN QA76.88.I52 1996. IEEE order plan catalog number 96TB100069. EEE Computer Society Press order number PR07582.

Froherz:1997:TPJ

[FT97] Michael Froherz and Peter Toenne. Tönne: Projekte mit Java. International Thomson Computer Press, 20 Park Plaza Suite 1001, Boston, MA 02116 USA, 1997. ISBN 3-8266-0316-8 (??invalid ISBN??). 450 pp. LCCN ???? 69 DM. URL http:

                                        //www.itp.de/. Includes CD-

ROM.

Foster:1997:TUS

[FTT97] Ian Foster, George K. Thiruvathukal, and Steven Tuecke.

Technologies for ubiquitous supercomputing: a Java interface to the Nexus communication system. Concurrency: Practice and Experience, 9(6):465–475, June 1997. CODEN CPEXEI. ISSN 1040-3108. URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=13871;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=13871&PLACEBO=IE.pdf.

Fuchs:1996:LTE

[Fuc96] M. Fuchs. Let’s talk: extending the Web to support collaboration. In IEEE [IEE96m], page ?? ISBN 0-8186-7446-6. LCCN TS155.6 .W56 1996. IEEE Order Plan catalog number 96TB100058. IEEE Computer Society Press order number PR07445.

Fuenfrocken:1998:TMJ

[Fue98] S. Fuenfrocken. Transparent migration of Java-based mobile agents: Capturing and reestablishing the state of Java programs. Lecture Notes in Computer Science, 1477:26–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Funfrocken:1998:TMJ

[Fün98] Stefan Fünfrocken. Transparent migration of Java-based mobile agents. Lecture Notes in Computer Science, 1477:26– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/
                                        service/series/0558/bibs/
                                        1477/14770026.htm;   http:
                                        //link.springer-ny.com/link/
                                        service/series/0558/papers/
                                        1477/14770026.pdf.

Frieder:1996:PFI

[FW96] O. Frieder and J. Wigglesworth, editors. Proceedings of the Fourth International Symposium on Assessment of Software Tools, May 22–24, 1996, Toronto, Canada. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-8186-7389-3, 0-8186-7390-7 (softbound), 0-8186-7391-5 (microfiche). LCCN QA76.76.D47I595 1996. IEEE order Plan catalog number 96TB100054.

Glahn:1997:JMA

[G^{+}97] Kay Glahn et al. Java 1.1 mit Methode: Applikationen und Applets im Internet und Intranet. Computer und Literaturverlag, ????, November 1997. ISBN 3-9823110-3-5 (??Invalid checksum??). ???? pp. LCCN ???? 98 DM.

Geib:19xx:CCP

[G^{+}xxa] Jean-Marc Geib et al. CORBA: Des concepts à la pratique. Masson Editeur, Masson, France, 19xx. ISBN 2-225-83046-0. ???? pp. LCCN ???? 165 FF. URL

                                        http://www.masson.fr/cgi-
                                        bin/bookf.pl?1:1:is=2225830460.

Gosling:19xx:JA

[G^{+}xxb] James Gosling et al. Java — Die Anwendungsprogrammierschnittstelle. Addison-Wesley, Reading, MA, USA, 19xx. ISBN ???? ???? pp. LCCN ????

URL http://java.sun.com/
people/jag/;   http://www.
addison-wesley.de/; http:
//www.javasoft.com/people/
jag/.

Arnold:19xx:JS

[G^{+}xxc] James Gosling et al. Java — Die Sprachspezifikation. Addison-Wesley, Reading, MA, USA, 19xx. ISBN ???? ???? pp. LCCN ???? URL http:

//www.addison-wesley.de/;
http://www.javasoft.com/
people/arnold/;   http://
www.javasoft.com/people/jag/

.

Gosling:19xx:JAVa

[G^{+}xxd] James Gosling et al. The Java API Vol. 1 (Japanese translation). Addison-Wesley, Reading, MA, USA, 19xx. ISBN

4-7952-9688-X.  ????  pp.
LCCN ????  ????  URL

http://www.awl.com/cseng/
javaseries/international/

japan.htm. Japanese translation.

Gosling:19xx:JAVb

[G^{+}xxe] James Gosling et al. The Java API Vol. 2 (Japanese translation). Addison-Wesley, Reading, MA, USA, 19xx. ISBN

4-7952-9689-8.  ????  pp.
LCCN ????  ????  URL

http://www.awl.com/cseng/
javaseries/international/

japan.htm. Japanese translation.

Gosling:19xx:JAPa

[G^{+}xxf] James Gosling et al. The Java Application Programming Interface, Vol. 1. GOTOP Information Inc., 5F, No.7, Lane 50, Sec.3 Nan Kang Road Taipei, Taiwan; Unit 1905,Metro Plaza Tower 2, No.223 Hing Fong Road, Kwai Chung, N.T., Hong Kong, 19xx. ISBN ???? ???? pp. LCCN ???? ????

URL http://java.sun.com/
people/jag/;   http://www.

gotop.com.tw. Chinese translation by Thi Shiang Workshop.

Gosling:19xx:JAPb

[G^{+}xxg] James Gosling et al. The Java Application Programming Interface, Vol. 2. GOTOP Information Inc., 5F, No.7, Lane 50, Sec.3 Nan Kang Road Taipei, Taiwan; Unit 1905,Metro Plaza Tower 2, No.223 Hing Fong Road, Kwai Chung, N.T., Hong Kong, 19xx. ISBN ???? ???? pp. LCCN ???? ????

URL http://java.sun.com/
people/jag/;   http://www.

gotop.com.tw. Chinese translation by Thi Shiang Workshop.

Gosling:19xx:JLS

[G^{+}xxh] James Gosling et al. The Java Language Specification. GOTOP Information Inc., 5F, No.7, Lane 50, Sec.3 Nan Kang Road Taipei, Taiwan; Unit 1905,Metro Plaza Tower 2, No.223 Hing Fong Road, Kwai Chung, N.T., Hong Kong, 19xx. ISBN ???? ???? pp. LCCN ????

URL http://java.sun.com/
people/jag/;   http://www.

gotop.com.tw. Chinese translation by Thi Shiang Workshop.

Gosling:19xx:JLSb

[G^{+}xxi] James Gosling et al. The Java Language Specification. GOTOP Information Inc., 5F, No.7, Lane 50, Sec.3 Nan Kang Road Taipei, Taiwan; Unit 1905,Metro Plaza Tower 2, No.223 Hing Fong Road, Kwai Chung, N.T., Hong Kong, 19xx. ISBN ???? ???? pp. LCCN

????  ????  URL http:

//java.sun.com/people/jag/
;   http://www.gotop.com.

tw. Chinese translation by Thi Shiang Workshop.

Gritzalis:1998:SIS

[GA98] Stefanos Gritzalis and George Aggelis. Security issues surrounding programming languages for mobile code: JAVA vs. Safe-Tcl. Operating Systems Review, 32(2):16–32, April 1998. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Gaither:1997:MGJ

[Gai97]  Mark Gaither.  60 Minute

Guide to Java. Softbank Publishing, ??, Japan, 1997. ISBN ???? ???? pp. LCCN ???? URL http:/

/www.lanw.com/mgbio.htm;

                                        http://www.softbank.co.jp/
                                        publishing/osbooks/java/java60.
                                        html;   http://www.softbank.
                                        co.jp/softbank/publishing/

.

Gameiro:1996:NID

[Gam96] G. Gameiro. Networking intelligent devices. Dr. Dobb’s Journal of Software Tools, 21(2):68, 70, 72, 74, 104–105, February 1996. CODEN DDJOEB. ISSN 1044-789X.

Ganssle:1996:BP

[Gan96] Jack G. Ganssle. Break points. Embedded Systems Programming, 9(8):89–??, August 1, 1996. CODEN EYPRE4. ISSN 1040-3272.

Gantz:1997:JCB

[Gan97] John Gantz. Has the Java crusaded been infiltrated? Computerworld, 31(24):33, June 16, 1997. CODEN CMP-WAB. ISSN 0010-4841. Argues that 100% Pure Java will be unattainable, as vendors promote proprietary changes to the language and environment.

Garfinkel:1996:JCG

[Gar96a] Simson L. Garfinkel. Java clones gain. SunExpert Magazine, 7(9):6, 8, September 1996. ISSN 1053-9239. Discusses Microsoft’s own implementation of a Java Virtual Machine, anda just-in-time (JIT) compiler now bundled into Microsoft Internet Explorer Version 3.0, and into a future version of the Windows operating system. Apple is collaborating with Natural Intelligence Inc. on a separate Java runtime system.

Garfinkel:1996:JSC

[Gar96b] Simson L. Garfinkel. Java security cracked: Again and again. WebServer Magazine: For Managers of World Wide Web Sites, 1(2):8, July/August 1996. ISSN 1087-4232. URL

http://www.cpg.com.

Garfinkel:1996:JW

[Gar96c] Simson L. Garfinkel. Java war.

SunExpert Magazine, 7(9):14– 15, September 1996. ISSN 1053-9239. Discusses the conflict and competition between Sun’s SunSoft and JavaSoft companies in the Java book market: the SunSoft Press Java Series [CH96, JM96, Pew96, van96a] is done with Prentice-Hall, and the JavaSoft Java series [AG96d, CW96a, Cha96b, GYT96c, GYT96d, GJS96, LY97b] from Addison-Wesley.

Garfinkel:1996:NJ

[Gar96d] Simson L. Garfinkel. NEO and Java. WebServer Magazine: For Managers of World Wide Web Sites, 1(2):6, 8, July/August 1996. ISSN 1087-4232. URL

http://www.cpg.com.

Garfinkel:1996:ACB

[Gar96e] Simson L. Garfinkel. Not all CPUs are built alike. Web-Server Magazine: For Managers of World Wide Web Sites, 1(2):12, 14, July/August 1996. ISSN 1087-4232. URL http:

//www.cpg.com. Discusses the

ShBoom CPU chip from Patriot Scientific Corporation; the chip is suited for Java and PostScript.

Garrett:1996:ENC

[Gar96f] Doug Garrett. Examining the Nutmeg class library. Dr. Dobb’s Journal of Software Tools, 21(10):80, 82, 84, 86, October 1, 1996. CODEN DDJOEB. ISSN 1044-789X.

Garber:1997:NBAa

[Gar97a]  Lee Garber.  News briefs:

Agency: Net use hasn’t hurt US phone system; Apple unveils turnaround strategy; JTC rejects Java standards plan; vendors plan 300-nm wafers; battle brews over smart card encryption; countries lift Internet telephone ban; Motorola proposes satellite system; vendors agree on highcapacity disks; “wrapping” software sales; Tonga offers domain name alternative. Computer, 30(9):19–22, September 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//neumann.computer.org/co/
books/co1997/pdf/r9019.pdf.

Garber:1997:NBAb

[Gar97b] Lee Garber. News briefs: Appliance war could make Web less open; Microsoft, Netscape agree of VRML; CORBA to stay independent of Java; US runs out of visas for foreign computer workers; new Year 2000 problems emerge. Computer, 30(10):20, 23–25, October 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://

pdf.computer.org/co/books/
co1997/pdf/rx020.pdf.

Garber:1997:NBB

[Gar97c] Lee Garber. News briefs: Binary version could bring VRML into the mainstream. FCC jumps into Internet fray. Java and floating-point math. Intel to design NDRAM. battle over net telephony. vendors seek fast modems. US permits export of strong encryption. E-commerce nears $1 billion. chasing the blue light. personal E-mail use will soar. Computer, 30(4):25–27, April 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Garber:1997:NBP

[Gar97d] Lee Garber. News briefs: Plans for Java standards draws fire. Computer, 30(7):17, July 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). Discusses the pending request by Sun to the Joint Technical Committee of the International Organization for Standardization and the International Electrotechnical Commission to standardize Java, and the opposition that has been raised by Sun’s competitors.

Garcia:1997:HCJ

[Gar97e] Alonso Alvarez Garcia. HTML CGI Java Servidores Tecnologia WWW. Anaya Multimedia, Madrid, Spain, 1997. ISBN 84-7614-959-X. ???? pp. LCCN ???? 3250

                                        Ptas. URL http://www.anaya.
                                        es/Catalogo/CGA?act=L&art=
                                        2321005&ed=2300;   http://
                                        www.anaya.es/multimedia/.

Garber:1998:NBHa

[Gar98a] Lee Garber. News briefs: HP enters the Java fray: Toward a more intelligent Web: Mobile IP on the move: Dispute over reported IT worker shortage. Computer, 31(6):19– 21, June 1998. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

                                        http://dlib.computer.org/
                                        co/books/co1998/pdf/r6019.
                                        pdf.

Garber:1998:NBS

[Gar98b] Lee Garber. News briefs: Sun gets OK to propose Java standards; domain name plan delayed; encryption battle heats up; monstor processes to debut in 1998; work begins on API for digital TV; ISO approves C++ Standard; new standard for parallel processing workstations; IBM unveils high-capacity disk technology. Computer, 31(1):21–24, January 1998. CODEN CPTRB4.

ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://

pdf.computer.org/co/books/
co1998/pdf/r1021.pdf.

Garber:1999:NBA

[Gar99] Lee Garber. News briefs: Apple engineers a major turnaround; Java battles brew on several fronts; IBM and SGI unveil ultrafast supercomputers; Y2K briefs. Computer, 32(1):23–25, January 1999. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

http://dlib.computer.org/
co/books/co1999/pdf/r1023.
pdf.

Garcia:19xx:PVJ

[Garxx] Alvaro Ojeda García. Programación en Visual J++. Anaya Multimedia, Madrid, Spain, 19xx. ISBN 84-415-0171-8. ???? pp. LCCN ???? 4,595 Ptas. URL http:

//www.anayamultimedia.es/
cgi-multi/notas_prensa.pl?
2318006&3&0&1.

Gathman:1997:TUJ

[Gat97] Stuart D. Gathman. A text UI for the Java AWT. Dr. Dobb’s Journal of Software Tools, 22 (9):38, 40–42, 45, 84, 86, September 1997. CODEN DDJOEB. ISSN 1044-789X.

Gaudin:1996:FAF

[Gau96a] Sharon Gaudin. Firms add flavor to Java menu: IBM to develop architecture to link Open-Doc and Java. Computerworld, 30(23):14, June 3, 1996. CO-DEN CMPWAB. ISSN 0010-4841.

Gaudin:1996:JHW

[Gau96b] Sharon Gaudin. Java hits Wall Street. Computerworld, 30(42): 51, 60, October 14, 1996. CO-DEN CMPWAB. ISSN 0010-4841.

Gaudin:1996:JSD

[Gau96c] Sharon Gaudin. JavaScript standard due by summer. Computerworld, 30(49):139, December 2, 1996. CODEN CMP-WAB. ISSN 0010-4841.

Gaudin:1996:JHB

[Gau96d] Sharon Gaudin. JavaSoft heats up Beans tool kit beta. Computerworld, 30(50):3, December 9, 1996. CODEN CMPWAB. ISSN 0010-4841. Notes that Sun will release a beta version of the Beans Development Kit (BDK) at Internet World ’96.

Gaudin:1996:JSS

[Gau96e]  Sharon Gaudin.  A jump

start for slow Java apps: Microsoft releases Java plug-in for Navigator; Netscape plans upgrade. Computerworld, 30(48): 2, November 25, 1996. CODEN CMPWAB. ISSN 0010-4841.

Gaudin:1996:MJC

[Gau96f] Sharon Gaudin. Microsoft’s Java compiler plans unclear. Computerworld, 30(50):53, 58, December 9, 1996. CODEN CMPWAB. ISSN 0010-4841.

Gaudin:1996:NMR

[Gau96g] Sharon Gaudin. Netscape, Microsoft ready Java for Win 3.1. Computerworld, 30(46):6, November 11, 1996. CODEN CMPWAB. ISSN 0010-4841.

Gaudin:1996:SRJ

[Gau96h] Sharon Gaudin. Sun ripens JavaBeans ahead of schedule. Computerworld, 30(43):8, October 21, 1996. CODEN CMPWAB. ISSN 0010-4841.

Gaudin:1996:SSB

[Gau96i] Sharon Gaudin. Sun seeks a better Java for Windows. Computerworld, 30(49):3, December 2, 1996. CODEN CMPWAB. ISSN 0010-4841.

Gaudin:1996:TGJ

[Gau96j] Sharon Gaudin. Tools to get Java perking faster. Computerworld, 30(46):54, 56, November 11, 1996. CODEN CMPWAB. ISSN 0010-4841. Notes that Sun expects to ship a just-in-time compiler for Java in the first half of 1997, and a native compiler in the second half of 1997. Project Studio with a drag-and-drop Java application builder, and Project Ice-T for building and porting client/server applications, are expected in the second quarter of 1997.

Gaudin:1997:DAP

[Gau97a] Sharon Gaudin. Developers await prime-time Java. Computerworld, 31(16):47–48, April 21, 1997. CODEN CMPWAB. ISSN 0010-4841.

Gaudin:1997:IPB

[Gau97b] Sharon Gaudin. IBM plan builds on Java. Computerworld, 31(25):61–62, June 23, 1997. CODEN CMPWAB. ISSN 0010-4841. IBM’s Universal Virtual Machine (UVM) will support Java, Smalltalk, Basic, and possibly Cobol and C++.

Gaudin:1997:IGJ

[Gau97c] Sharon Gaudin. IBM to give Java cross-platform company: Smalltalk, Basic go platformindependent. Computerworld, 31(24):1, 12, June 16, 1997. CO-DEN CMPWAB. ISSN 0010-4841.

Gaudin:1997:JBM

[Gau97d] Sharon Gaudin. Java becomes more consumer-friendly. Computerworld, 31(27):64, July 7, 1997. CODEN CMPWAB. ISSN 0010-4841.

Gaudin:1997:JCL

[Gau97e] Sharon Gaudin. Java, CORBA to link up: Sun’s crossplatform plan targets Web access. Computerworld, 31(12):1, 121, March 24, 1997. CODEN CMPWAB. ISSN 0010-4841.

Gaudin:1997:JSC

[Gau97f] Sharon Gaudin. Java’s speed, cross-platform issues targeted. Computerworld, 31(14):1, 113, April 7, 1997. CODEN CMP-WAB. ISSN 0010-4841.

Gaudin:1997:MTR

[Gau97g]  Sharon Gaudin.  Microsoft

tries to reel in Java jumpers: Redmond to C++ developers: Write Java for Windows. Computerworld, 31(27):20, July 7, 1997. CODEN CMPWAB. ISSN 0010-4841.

Gaudin:1997:OAS

[Gau97h] Sharon Gaudin. Online access:

Scientists use Java, CORBA to study DNA. Computerworld, 31(31):55, 57, August 4, 1997. CODEN CMPWAB. ISSN 0010-4841.

Gaudin:1997:OJL

[Gau97i] Sharon Gaudin. OpenDoc/Java link touted. Computerworld, 31 (4):49, January 27, 1997. CO-DEN CMPWAB. ISSN 0010-4841.

Gaudin:1997:OTG

[Gau97j] Sharon Gaudin. Oracle tools to get Java, C++ technology. Computerworld, 31(6):16, February 10, 1997. CODEN CMPWAB. ISSN 0010-4841.

Gaudin:1997:RJT

[Gau97k] Sharon Gaudin. RAD Java tools lure business users. Computerworld, 31(24):57, 60, June 16, 1997. CODEN CMPWAB. ISSN 0010-4841.

Gaudin:1997:SAP

[Gau97l] Sharon Gaudin. Sun adds on to 100% Pure Java. Computerworld, 31(23):12, June 9, 1997. CODEN CMPWAB. ISSN 0010-4841.

Gaudin:1997:SHJ

[Gau97m] Sharon Gaudin. Sun hits Java punch list, but users crave speed. Computerworld, 31(29): 10, July 21, 1997. CODEN CM-PWAB. ISSN 0010-4841.

Gaudin:1997:SLF

[Gau97n] Sharon Gaudin. Sun loses first battle in Java standard quest: Key issue: Who will control changes to platform? Computerworld, 31(23):3, June 9, 1997. CODEN CMPWAB. ISSN 0010-4841.

Gaudin:1997:SSJ

[Gau97o] Sharon Gaudin. Sun to spread Java beyond computers. Computerworld, 31(13):16, March 31, 1997. CODEN CMPWAB. ISSN 0010-4841.

Gaudin:1997:STJ

[Gau97p] Sharon Gaudin. Sun tweaks Java, but faster version in works. Computerworld, 31(8): 14, February 24, 1997. CODEN CMPWAB. ISSN 0010-4841.

Gaudin:1997:VBJ

[Gau97q] Sharon Gaudin. Visa boosts Java’s credit line: smart cards will redefine how plastic is used. Computerworld, 31(31):1, 16, August 4, 1997. CODEN CM-PWAB. ISSN 0010-4841.

Gaudin:1997:YTC

[Gau97r] Sharon Gaudin. You, too, can search for life on Mars: Java applet lets public take a closer look at planet. Computerworld, 31(28):53–54, July 14, 1997. CODEN CMPWAB. ISSN 0010-4841. URL http://

mars.graham.com/wits.

Gray:1998:CLE

[GBS98] John Gray, Tom Boyle, and Colin Smith. A constructivist learning environment implemented in Java. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):94–97, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Garber:1997:NBJ

[GC97] Lee Garber and David Clark.

News briefs: Judge rejects US restrictions on export of encryption; Intel, HP unveil Merced chip; Java wars heat up; AOL acquires CompuServe subscribers; confusion in the DVD marketplace; ‘amazing grace’ heads to sea; PC firms back down on convergence; business use will drive Internet growth; company offers $1million prize for hackers. Computer, 30(11):22–25, November 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://

pdf.computer.org/co/books/
co1997/pdf/ry022.pdf.

Griffith:19xx:JT

[GCxx] Steven Griffith and Mark Chan.

Java: 1001 Tips. Ohmsha, ????, 19xx. ISBN 4-274-06223-6. ???? pp. LCCN ???? 5,000

yen. URL http://www.ohmsha.

                                        co.jp/data/books/contents/
                                        4-274-06223-6.htm.

Goldberg:1998:JVP

[GD98] Athomas Goldberg and Troy Downing. Java VRML Power Guide. IDG Books, San Mateo, CA, USA, February 1998. ISBN 0-7645-8051-5. ???? pp. LCCN ???? US$50.

Giavitto:1998:DPJ

[GDS98] J.-L. Giavitto, D. De Vito, and J.-P. Sansonnet. A data parallel Java client-server architecture for data field computations over Z_{n}. Lecture Notes in Computer Science, 1470:742– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Geary:1996:PHP

[Gea96] David Geary. Prentice Hall PTR/Sunsoft Press Graphic Java. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, September 1996. ISBN 0-614-20319-8. ???? pp. LCCN ???? US$39.95.

Geary:1998:GJVc

[Gea98] David Geary. Graphic Java Volume 3: 2D API. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, third edition, June 1998. ISBN 0-13-079668-9. LCCN

????  ????  URL http:

                                        //www.prenhall.com/books/
                                        ptr_0130796689.html.

Geary:1999:GJM

[Gea99] David M. Geary. Graphic Java 2: Mastering the JFC, Volume 2: Swing. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, third edition, 1999. ISBN 0-13-079667-0. 1664 pp. LCCN ???? US$54.99. URL http:

//www.prenhall.com/books/
ptr_0130796670.html.

Geist:1996:APP

[Gei96] G. A. Geist. Advanced programming in PVM. In Bode [Bod96], pages 1–?? ISBN 3-540-61779-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .E975 1996 Bar.

Geppert:1998:IDW

[Gep98] L. Geppert. IC design on the World Wide Web. IEEE Spectrum, 35(6):45–50, June 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Grinzo:1999:PBC

[GF99] Lou Grinzo and Laryn Fernandes. Programmer’s bookshelf: Clarifying the Open Source movement. Dr. Dobb’s Journal of Software Tools, 24(9):119– 120, September 1999. CODEN DDJOEB. ISSN 1044-789X.

Getov:1998:HPPa

[GFHM98] Vladimir Getov, Susan Flynn-Hummel, and Sava Mintchev. High-performance parallel programming in Java: Exploiting native libraries. In ACM [ACM98a], page ?? CODEN CPEXEI. ISSN 1040-3108. LCCN ????

URL http://www.cs.ucsb.
edu/conferences/java98/papers/
hpjavampi.pdf;   http://www.
cs.ucsb.edu/conferences/java98/
papers/hpjavampi.ps. Also

published as Concurrency: Practice and Experience, 10(11– 13), September 1998, CODEN CPEXEI, ISSN 1040-3108.

Goeschka:1998:DAH

[GFR98] K. M. Goeschka, J. Falb, and W. Radinger. Database access with HTML and Java — a comparison based on practical experiences. In IEEE [IEE98b], pages 588–595. ISBN 0-8186-8585-9 (softbound), 0-7803-5168-1 (casebound), 0-8186-8587-5 (microfiche). ISSN 0730-3157. LCCN QA76.6 .I546 1998.

Guenette:1996:CWB

[GG96a] D. R. Guenette and R. Gustavson. CD-ROM and Web browsers: ¡HTML¿ as the Lingua Franca. CD-ROM Professional, 9(8):26–30, 32–34, 36, 38–42, 44, 46–48, 50–51, August 1996. CODEN CRPFEX. ISSN 1049-0833.

Gurewich:1996:JMS

[GG96b] Nathan Gurewich and Ori Gurewich. Java Manual of Style. Ziff-Davis Press, Emeryville, CA, USA, April 1, 1996. ISBN 1-56276-408-X. xvii + 170 pp. LCCN QA76.73.J38G87 1996. US$24.99. URL http://www.

mcp.com/zdpress.

Gurewich:1997:JMS

[GG97] Nathan Gurewich and Ori Gurewich. Java Manual of Style. Impress Corporation, ??, Japan, 1997. ISBN 4-8443-4763-X (??Invalid checksum??). ???? pp. LCCN ???? 1631 yen. URL http:

//www.ips.co.jp/;   http:
//www.ips.co.jp/book/4763/

4763.htm. Japanese translation.

Godfrey:1999:JBS

[GG99] Michael Godfrey and Dan Grossman. JDuck: building a software engineering tool in Java as a CS2 project. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1):48–52, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gurewich:19xx:JGA

[GGxxa] Nathan Gurewich and Ori Gurewich. Java Guida al Linguaggio per Creare Pagine Web Interattive. Mondadori Informatica, ??, Italy, 19xx. ISBN 88-7131-743-2. ???? pp. LCCN ???? 25000 Lire. URL http:/

/www.mondadori.com/. Italian

translation of [GG96b].

Gurewich:19xx:JM

[GGxxb] Nathan Gurewich and Ori Gurewich. Java MiniDoc. Simon & Schuster Macmillan (France), Paris, France, 19xx. ISBN 2-7440-0143-0. ???? pp. LCCN ???? 69 FF. Includes CD-ROM.

Gurewich:1996:WS

[GGA96] Nathan Gurewich, Ori Gurewich, and Larry Aronson. Webmaster SuperPack. Ziff-Davis Press, Emeryville, CA, USA, May 1996. ISBN 1-56276-395-4 (softcover). 520 pp. LCCN ???? US$40.49. Includes Both Java Manual of Style and HTML 3 Manual of Style.

Guerrieri:1997:IJS

[GGL^{+}97] Ernesto Guerrieri, Martin Griss, Doug Lea, Prashant Sridharan, and Will Tracz. The impact of Java on software reusability. ACM SIGSOFT Software Engineering Notes, 22(3):210–211, May 1997. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Getov:1999:MLP

[GGMS99] Vladimir Getov, Paul Gray, Sava Mintchev, and Vaidy Sunderam. Multi-language programming environments for high performance Java computing. Scientific Programming, 7(2):139–146, ???? 1999. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:

                                        //iospress.metapress.com/
                                        app/home/contribution.asp%
                                        3Fwasp=f277qlrwwjr5m4vxjywv%
                                        26referrer=parent%26backto=
                                        issue%2C5%2C8%3Bjournal%2C7%
                                        2C9%3Blinkingpublicationresults%
                                        2C1%2C1.

Getov:1999:MJM

[GGS99] Vladimir Getov, Paul Gray, and Vaidy Sunderam. MPI and Java-MPI: Contrasts and comparisons of low-level communication performance. In ACM [ACM99c], page ??

Gaudin:1996:LBJ

[GH96a] Sharon Gaudin and Frank Hayes. Language barriers: In 1997, Java and ActiveX will slug it out for corporate dominance. Computerworld, 30(52):38, 39, December 23, 1996. CODEN CMPWAB. ISSN 0010-4841.

Gaudin:1996:MEJ

[GH96b] Sharon Gaudin and Frank Hayes. Microsoft eyes Java threat: will license Office language. Computerworld, 30(22): 1, 14, May 27, 1996. CODEN CMPWAB. ISSN 0010-4841.

Gaudin:1996:VBJ

[GH96c] Sharon Gaudin and Frank Hayes. Visual Basic, Java vie for ’net crown. Computerworld, 30 (24):1, 109, June 10, 1996. CO-DEN CMPWAB. ISSN 0010-4841.

Gall:1997:PJB

[GH97] Ulrich Gall and Franz J.

Hauck. Promondia: a Javabased framework for real-time group communication in the Web. Computer Networks and ISDN Systems, 29(8–13):917– 926, September 30, 1997. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/
browse/browse.cgi?year=1997&
volume=29&issue=08-13&aid=
1682.

Getov:1998:HPPb

[GHM98] Vladimir Getov, Susan Flynn Hummel, and Sava Mintchev. High-performance parallel programming in Java: exploiting native libraries. Concurrency: Practice and Experience, 10(11–13):863–872, September 1998. CODEN CPEXEI. ISSN 1040-3108. URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=10050396;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=10050396&PLACEBO=IE.pdf.

Special Issue: Java for Highperformance Network Computing.

Groner:1996:JLA

[GHP^{+}96] Dan Groner, K. C. Hopson, Harish Prabandham, Todd Sundsted, and Daniel Groner. Java Language API Super-Bible: the comprehensive reference to the Java programming language: everything you need to know about the standard language, input/output, and debugging classes. Waite Group Press, Corte Madera, CA, USA, October 1, 1996. ISBN 1-57169-038-7. xxiii + 851 pp. LCCN QA76.73.J38J37 1996. US$59.99. URL http:

//www.mcp.com/cgi-bin/bag?
isbn=1-57169-038-7&last=/

bookstore;   http://www.waite.
com/waite/.

Gibbs:1995:NDL

[Gib95] Mark Gibbs. News: a devilish letter to Bill Gates. Network World, page ??, December 19, 1995. ISSN 0887-7661 (print), 1944-7655 (elec-

tronic). URL http://www.
iquest.net/index.html.

Gibbons:1998:SPJ

[Gib98] Jeremy Gibbons. Structured programming in Java. ACM SIGPLAN Notices, 33(4):40–43, April 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.
rowan.edu/sigplan/apr199~
1.htm.

Giguere:1997:JBN

[Gig97]  Eric Giguère.  Java Beans

and the new event model. Dr. Dobb’s Journal of Software Tools, 22(4):50, 52, 53, 79, 80, April 1997. CODEN DDJOEB. ISSN 1044-789X.

Gill:1996:JTS

[Gil96a] P. Gill. Java tools are still brewing. Information Week, ???? (594):1A, 3A–4A, 8A, ???? 26, 1996. CODEN INFWE4. ISSN 8750-6874.

Gillin:1996:CJJ

[Gil96b] Paul Gillin. Cooling down Java — Java isn’t the programmer’s magic bullet. Computerworld, 30(14):36–??, ???? 1996. CO-DEN CMPWAB. ISSN 0010-4841.

Gillin:1996:LJC

[Gil96c] Paul Gillin. Let Java cool. Computerworld, 30(14):36, April 1, 1996. CODEN CMPWAB. ISSN 0010-4841.

Gill:1997:ORD

[Gil97a] Philip Gill. Object/relational databases: The key to mainstream acceptance? ORDBMS may be as important as Java in bringing objects in from the code. Object Magazine, 7(2):14– 15, April 1997. CODEN OB-MAFO. ISSN 1055-3614.

Gillin:1997:FJP

[Gil97b] Paul Gillin. Up front: Java politics. Computerworld, 31(19):2, May 12, 1997. CODEN CMP-WAB. ISSN 0010-4841.

Gillmor:1997:DDA

[Gil97c] Steve Gillmor. Dial 411 for directory assistance — Novell’s latest NetWare release bets big on Java and the intranet. BYTE Magazine, 22(1):137–??, January 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Gillam:1998:TBA

[Gil98a] Richard Gillam. Text boundary analysis in JDK 1.2. In Unicode Consortium [Uni98b], page ?? ISBN ???? LCCN

                                        ????  URL http://www.
                                        unicode.org/unicode/iuc13/
                                        program.html.

Gillmor:1998:JWM

[Gil98b] Steve Gillmor. Java for Windows: Microsoft’s Visual J++ 6.0 ties Java ever more tightly into the Windows platform. BYTE Magazine, 23(5):41, May 1998. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Gillmor:1998:PSD

[Gil98c] Steve Gillmor. Productivity software — a different desktop for the workplace — Lotus’ eSuite uses JavaBeans. BYTE Magazine, 23(2):41, February 1998. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Gilorien:1999:DJ

[Gil99] Gilorien. DHTML and JavaScript.

P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1999. ISBN 0-13-086334-3. 1168 pp. LCCN QA76.76.H94 G55 1999. US$42.99. URL http:

//www.phptr.com/ptrbooks/
ptr_0130863343.html.

Gaudin:1996:HGU

[GJ96a] Sharon Gaudin and April Jacobs. HP gives its Unix system a Java jolt. Computerworld, 30(47):53, 57, November 18, 1996. CODEN CM-PWAB. ISSN 0010-4841. Notes that HP is readying the HP-UX Java Developer Kit, the HP-UX Java Virtual Machine, and an HP-UX Just-In-Time compiler for Java. The first two are available now at http://www.hp.com/go/JAVA.

Gaudin:1996:IMI

[GJ96b] Sharon Gaudin and April Jacobs. IBM may incorporate Java into its network computer. Computerworld, 30(46): 14, November 11, 1996. CO-DEN CMPWAB. ISSN 0010-4841.

Girdley:1996:WPJ

[GJ^{+}96c] Michael Girdley, Kathryn A. Jones, et al. Web Programming With Java. Sams.net, Indianapolis, IN 46268, USA, September 1996. ISBN 1-57521-113-0. xxii + 481 pp. LCCN TK5105.888.G57 1996. US$39.99, CDN$56.95, UK£37.50. URL http://

                                        merchant.superlibrary.com:
                                        8000/catalog/hg/PRODUCT/PAGE/
                                        15752/bud/1575211130.html;
                                         http://www.mcp.com/cgi-bin/
                                        bag?isbn=1-57521-113-0&last=
                                        /bookstore;   http://www.
                                        mcp.com/samsnet/.

Godwin-Jones:1996:IWC

[GJ96d]  R. Godwin-Jones.  Interac-

tive Webbing: CGI scripting, JavaScript and linked programs for language learning. In Borchardt et al. [BBJR96], page ?? ISBN ???? LCCN ????

Gaudin:1997:NUO

[GJ97] Sharon Gaudin and April Jacobs. NC users overcome early Java glitches. Computerworld, 31(27):1, 107, July 7, 1997. CO-DEN CMPWAB. ISSN 0010-4841.

Gosling:1996:JLS

[GJS96] James Gosling, Bill Joy, and Guy L. Steele Jr. The Java Language Specification. The Java Series. Addison-Wesley, Reading, MA, USA, 1996. ISBN 0-201-63451-1. xxv + 825 pp. LCCN QA76.73.J38G68 1996. US$36.75. URL http://www.

aw.com/cp/javaseries.html;
 http://www.aw.com/cseng/
titles/0-201-63451-1/.

Gosling:1997:JS

[GJS97] James Gosling, Bill Joy, and Guy L. Steele, Jr. Java: Die Sprachspezifikation: Die offizielle Dokumentation von JavaSoft. Addison-Wesley, Reading, MA, USA, 1997. ISBN 3-8273-1038-5. xxi + 758 pp. LCCN ???? 89,90

DM.  URL http://www.
addison-wesley.de/katalog/
item.ppml?textexpr=Java&id=

00113. German translation of [GJS96].

Gieseke:1997:PPH

[GK97a] Wolfram Gieseke and Michael Knapp. PC Praxisbuch Homepage-Design mit HTML & Java. DATA-Becker, ????, 1997. ISBN 3-8158-1309-3. 357 pp. LCCN ???? 49 DM. URL

http://www.data-becker.de/
buecher/buchinfos/1309.htm;
 http://www.data-becker.de/
buecher/buecher.htm#pcp homepage.

Grand:1997:JFC

[GK97b] Mark Grand and Jonathan Knudsen. Java Fundamental Classes Reference. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, May 1997. ISBN 1-56592-241-7. xxi + 1090 pp. LCCN QA76.73.J38 G728 1997. US$34.95. URL http://www.

                                        ora.com/catalog/javafund/;
                                         http://www.oreilly.com/catalog/
                                        javafund.

Golm:1998:IRT

[GK98] Michael Golm and Jürgen Kleinöder. Implementing realtime actors with MetaJava. Lecture Notes in Computer Science, 1357:68–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/
                                        service/series/0558/bibs/
                                        1357/13570068.htm;   http:
                                        //link.springer-ny.com/link/
                                        service/series/0558/papers/
                                        1357/13570068.pdf.

Glahn:1997:FJ

[Gla97a] Kay Glahn. Free 4 Java. Carl Hanser, München, Germany, 1997. ISBN 3-446-31130-

0. ???? pp. LCCN ???? 40 DM.

Glahn:1997:FJT

[Gla97b] Kay Glahn. Free 4 Java: Tools für das Internet. JDKs für Windows, Mac und UNIX. Computer und Literaturverlag, ????, July 1997. ISBN 3-932311-30-2 (??invalid ISBN??). ???? pp. LCCN ???? 39 DM.

Glahn:1997:JM

[Gla97c] Kay Glahn. Java mit Methode.

Carl Hanser, München, Germany, 1997. ISBN 3-446-31103-3. 1000 pp. LCCN ???? 98 DM.

URL http://www.hanser.de/.

Includes CD-ROM.

Glass:1997:JGL

[Gla97d] G. Glass. The Java Generic Library. C++ Report, 9(1): 70–74, January 1997. CODEN CRPTE7. ISSN 1040-6042.

Glass:1998:OVA

[Gla98] G. Glass. ObjectSpace voyager — the agent ORB for java. Lecture Notes in Computer Science, 1368:38–??, 1998. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gu:1999:EJT

[GLC99] Yan Gu, B. S. Lee, and Wentong Cai. Evaluation of Java thread performance on two different multithreaded kernels. Operating Systems Review, 33(1):34– 46, January 1999. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Grimaud:1999:FTI

[GLV99] Gilles Grimaud, Jean-Louis Lanet, and Jean-Jacques Vandewalle. FACADE: a typed intermediate language dedicated to smart cards. ACM SIGSOFT Software Engineering Notes, 24 (6):476–493, November 1999.

CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gosling:1996:JLE

[GM96] James Gosling and Henry McGilton. The Java language environment: a white paper. Technical report, Sun Microsystems, 2550 Garcia Avenue, Mountain View, CA 94043, USA, October 1996. ???? pp.

Garside:1997:JFC

[GM97a] Roger Garside and John A. Mariani. Java First Contact. International Thomson Computer Press, 20 Park Plaza Suite 1001, Boston, MA 02116 USA, 1997. ISBN 1-85032-316-X. xxvi + 660 pp. LCCN QA76.73.J38G37 1997. US$40.

Geary:1997:GJA

[GM97b] David M. Geary and Alan L. McClellan. Graphic Java. Das AWT beherrschen. Verlag Heinz Heise, Helsdorfer Straße 7, D-30625, Hannover, Germany, 1997. ISBN 3-88229-089-7. 607 pp. LCCN ???? 88,00 DM.

URL http://www.emedia.de/
bin/bookshop?show=4864&id=

. German translation of [GM97c].

Geary:1997:GJMa

[GM97c] David M. Geary and Alan L. McClellan. Graphic Java: Mastering the AWT. SunSoft Press Java series Java series (Mountain View, CA). Sun-Soft Press, 2550 Garcia Avenue, Mountain View, CA 94043-1100, USA, 1997. ISBN 0-13-565847-0. xl + 600 pp. LCCN QA76.73.J38G43 1997. US$39.95. URL http://www.

sun.com/smi/ssoftpress/books/
Geary/Geary.html. Available

in German translation [GM97b].

Geary:1997:GJMb

[GM97d] David M. Geary and Alan L.

McClellan. Graphic Java: Mastering the AWT. SunSoft Press Java series Java series (Mountain View, CA). Sun-Soft Press, 2550 Garcia Avenue, Mountain View, CA 94043-1100, USA, second edition, July 1997. ISBN 0-13-863077-1. 900 pp. LCCN QA76.73.J38G43 1997. US$49.95. URL http://

merchant.superlibrary.com:
8000/catalog/hg/PRODUCT/PAGE/
01386/bud/0138630771.html;
 http://www.sun.com/smi/ssoftpress/
books/Geary/Geary.html.

Gilbert:1997:OOP

[GM97e] Stephen Gilbert and Bill Mc-Carty. Object Oriented Programming In Java. Mitchell Waite Signature Series. Waite Group Press, Corte Madera, CA, USA, 1997. ISBN 1-57169-086-7. xxii + 953 pp. LCCN QA76.64.G55 1997. 118,00 DM; US$60. URL http://

merchant.superlibrary.com:
8000/catalog/hg/PRODUCT/PAGE/
15716/bud/1571690867.html;
 http://www.mut.ch/profit.
htx?&x_pk=kd84mf7ms9dm.

Geary:1998:GJVa

[GM98a] David M. Geary and Alan L. McClellan. Graphic Java 1.2: mastering the JFC: Volume 1

— AWT. SunSoft Press Java series. SunSoft Press, 2550 Garcia Avenue, Mountain View, CA 94043-1100, USA, third edition, August 1998. ISBN 0-13-079666-2. 800 pp. LCCN ???? US$50. URL http:

                                        //www.phptr.com/ptrbooks/
                                        ptr_0130796662.html;   http:
                                        //www.prenhall.com/books/
                                        ptr_0130796662.html;   http:
                                        //www.sun.com/books/catalog/
                                        books_comingsoon.html.

Gheri:1998:MIO

[GM98b] G. Gheri and J. R. Mosig. The multilevel interactive objects approach to model and mesh large planar antennas and arrays: a Full-JAVA tool. Lecture Notes in Computer Science, 1401:969–??, 1998. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gilbert:1998:OOD

[GM98c] Stephen Gilbert and Bill Mc-Carthy. Object-Oriented Design in Java (Mitchell Waite Signature Series). Waite Group Press, Corte Madera, CA, USA, May 1998. ISBN 1-57169-134-0. LCCN QA76.73.J38G55 1998.

                                        US$50. URL http://iq-mcp3.
                                        iquest.net/catalog/mcp/PRODUCT/
                                        PAGE/15716/bud/1571691340.
                                        html.

Gordon:1998:EJJ

[GM98d] Robert Gordon and Alan Mc-Clellan. Essential JNI: Java Native Interface. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1998. ISBN 0-13-679895-0. xxvii + 499 pp. LCCN QA76.73.J38 G665 1998. US$34.95. URL http://www.

prenhall.com/ptrbooks/ptr_
0136798950.html.

Galatopoullos:1999:DPA

[GM99] D. G. Galatopoullos and E. S.

Manolakos. Developing parallel applications using the Java-Ports environment. Lecture Notes in Computer Science, 1586:813–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gustavson:1999:FMA

[GME99] Fred G. Gustavson, José E.

Moreira, and Robert F. Enenkel. The fused multiply-add instruction leads to algorithms for extended-precision floating point: applications to Java and high-performance computing. In Stephen A. MacKay and J. Howard Johnson, editors, CASCON ’99: Proceedings of the 1999 Conference of the Centre for Advanced Studies on Collaborative Research. November 8–11, 1999, Mississauga, Ontario, Canada, page [4]. IBM Corporation, San Jose, CA, USA, 1999. ISBN ???? LCCN ???? Dedicated to Cleve Moler on his 60th birthday.

Gong:1997:GBS

[GMPS97] Li Gong, Marianne Mueller, Hemma Prafullchandra, and Roland Schemers. Going beyond the sandbox: An overview of the new security architecture in the Java Development Kit 1.2. In USENIX [USE97c], pages 103–112. ISBN 1-880446-91-X. LCCN TK5105.875.I57

                                        U96 1997. URL http://www.
                                        usenix.org/publications/library/
                                        proceedings/usits97/gong.
                                        html.

Guadagno:1996:SFB

[GO96] L. Guadagno and O. B. Okon, III. Stepping up the flow of business information. Object Magazine, ????(????):41–45, July 1996. CODEN OBMAFO. ISSN 1055-3614.

Goff:1997:SRJ

[Gof97] Leslie Goff. Special report: Java: Do it yourself: Java pros tell how to give your career the Java jolt. Computerworld, 31 (19):88–89, May 12, 1997. CO-DEN CMPWAB. ISSN 0010-4841.

Goldammer:1996:HCJ

[Gol96a] G. Goldammer. HTML-script calls Java-applet… a new development technique under a software- and applicationstechnology configuration. IM Information Management, 11 (3):6–14, ???? 2, 1996. ISSN 0930-5181.

Golla:1996:EHK

[Gol96b] Andreas F. Golla. Erste-Hilfe-Kasten Java: HotJava Developers Kit V1.0 Beta. Sybex Verlag GmbH, Erkrather Straße 345-349, 40231 Düsseldorf, Germany, 1996. ISBN 3-8155-9926-1. 7 pp. LCCN ???? 40 DM. Includes CD-ROM.

Goldstein:1997:GSM

[Gol97] T. Goldstein. The gateway security model in the Java electronic commerce framework. Lecture Notes in Computer Science, 1318:340–??, 1997. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goldberg:1998:SJL

[Gol98a]  A. Goldberg.  A specifica-

tion of Java loading and bytecode verification. In ACM, editor, 5th ACM Conference on Computer and Communications Security: November 3–5, 1998, San Francisco, California, pages 49–58. ACM Press, New York, NY 10036, USA, 1998. ISBN 1-58113-007-4. LCCN QA76.9.A25 A33 1998.

Golden:1998:WWW

[Gol98b] B. L. Golden. The World Wide Web: HTML, CGI and Java applications in animal breeding. In Anonymous [Ano98o], pages 409–414. ISBN 1-86389-454-3. LCCN ????

Goller:1999:PPS

[Gol99] A. Goller. Parallel processing strategies for large SAR image data sets in a distributed environment. Computing, 62(4):277–291, 1999. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http://

link.springer-ny.com/link/
service/journals/00607/bibs/
9062004/90620277.htm;   http:
//link.springer-ny.com/link/
service/journals/00607/papers/
9062004/90620277.pdf.

Golla:19xx:EHKb

[Golxx]  Andreas F. Golla.  Erste-

Hilfe-Kasten Java mit Jewelbox. Sybex Verlag GmbH, Erkrather Straße 345-349, 40231 Düsseldorf, Germany, 19xx. ISBN 3-8155-9942-3. ???? pp. LCCN ???? 40 DM. URL http://

www.abc.de/gutenb/verlage/
sybex/sybex.htm;   http://
www.sybex.de/. Includes CD-

ROM.

Gonzalez:1995:IJH

[Gon95a] Sean González. InternetWorking — Java and HotJava. PC Magazine, 14(18):265, 1995. CODEN PCMGEP. ISSN 0888-8507.

Gonzalez:1995:JHW

[Gon95b] Sean González. Java & Hot-Java: Waking up the Web. PC Magazine, page ??, October 24, 1995. CODEN PCMGEP. ISSN 0888-8507.

URL http://www.zdnet.com/
~pcmag/issues/1418/pcm00085.
htm.

Gong:1997:JSP

[Gon97]  Li Gong.  Java security:

Present and near future: Coping with the rapidly evolving security issues of crossplatform computing. IEEE Micro, 17(3):14–19, May/June 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//pascal.computer.org/mi/
books/mi1997/pdf/m3014.pdf.

Gong:1998:JSM

[Gon98] Li Gong. The Java Security Model: Cryptography, Architectures, APIs, and Implementations. Addison-Wesley, Reading, MA, USA, 1998. ISBN 0-201-31000-7 (softcover). xiv + 262 pp. LCCN QA76.73.J38 G65 1999. US$36.53. URL

http://www2.awl.com/cseng/
javaseries/security.html.

Gong:1999:IJP

[Gon99] Li Gong. Inside Java 2 platform security: architecture, API design, and implementation. Addison-Wesley, Reading, MA, USA, 1999. ISBN 0-201-31000-7. ???? pp. LCCN QA76.73.J38 G65 1999. US$34.95.

Goodman:1996:DGJ

[Goo96a] Danny Goodman. Danny Goodman’s JavaScript Handbook. I D G Books Worldwide, Indianapolis, IN, USA, 1996. ISBN 0-7645-3003-8. xxi + 520 pp. LCCN QA76.73.J39G66 1996. US$34.99. URL http://db.

www.idgbooks.com/database/

                                        book/isbn/generic-book.tmpl?
                                        query=0-7645-3003-8.

Goodman:1996:JB

[Goo96b] Danny Goodman. JavaScript Bible. I D G Books Worldwide, Indianapolis, IN, USA, second edition, 1996. ISBN 0-7645-3022-4. xxvii + 607 pp. LCCN QA76.73.J39G66 1996. US$39.99.

Goodman:1996:PAJ

[Goo96c] Danny Goodman. Programmer avec JavaScript. Sybex Verlag GmbH, Erkrather Straße 345-349, 40231 Düsseldorf, Germany, 1996. ISBN 2-7361-2212-7. xix + 546 pp. LCCN ???? URL ????. Includes CD-ROM. French translation of [Goo96a].

Goodman:1997:OMGa

[Goo97a]  Danny Goodman.  Official

Marimba Guide to Bongo. Sams.net, Indianapolis, IN 46268, USA, 1997. ISBN 1-57521-254-4. xxxi + 511 pp. LCCN QA76.9.U83G66 1997. US$40. URL http://

                                        merchant.superlibrary.com:
                                        8000/catalog/hg/PRODUCT/PAGE/
                                        15752/bud/1575212544.html.

Includes CD-ROM.

Goodman:1997:OMGb

[Goo97b]  Danny Goodman.  Official

Marimba Guide to Bongo. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, August 1997. ISBN 4-88735-072-4. ???? pp. LCCN ???? 4,300

                                        yen. URL http://www.mmjp.

or.jp/prentice/wa_int15-j.
html.

Goodman:1997:PJ

[Goo97c]  Danny Goodman.  Progra-

mación en JavaScript. Anaya Multimedia, Madrid, Spain, 1997. ISBN 84-415-0080-0. ???? pp. LCCN ???? 4495 Ptas. URL http:

//www.anayamultimedia.es/
cgi-multi/notas_prensa.pl?
2321010&*2&goodman&1.

Goodwill:1999:DJS

[Goo99] James Goodwill. Developing Java servlets. Howard W. Sams, Indianapolis, IN 46268, USA, 1999. ISBN 0-672-31600-5. vii + 515 pp. LCCN QA76.73.J38 G6625 1999. US$29.99.

Goodman:19xx:BMG

[Gooxxa]  Danny Goodman.  Bongo

Marimba Guia Oficial. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 19xx. ISBN 84-8322-033-4. ???? pp. LCCN ???? ????

Goodman:19xx:JH

[Gooxxb] Danny Goodman. Het JavaScript Handboek. Addison-Wesley, Reading, MA, USA, 19xx. ISBN 90-6789-816-3. ???? pp. LCCN ???? fl. 90. URL http:/

/www.dannyg.com/. Includes

CD-ROM.

Goppold:1996:LTS

[Gop96]  A. Goppold.  The Leibniz

TLSI: a secondary macro programming interface and universal ASCII user interface shell for hypermedia. In Diaz de Ilarraza Sanchez and Fernandez de Castro [DF96], page ?? ISBN 3-540-61491-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q181.A1I538 1996.

Gordon:1996:JNB

[Gor96] B. Gordon. Java — a new brew for educators, administrators and students. Educom review, 31(2):44–46, March–April 1996. CODEN EDREEW. ISSN 1045-9146.

Gosling:1995:JIB

[Gos95] James Gosling. Java intermediate bytecodes. ACM SIG-PLAN Notices, 30(3):111–118, March 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

                                        tronic). URL http://www.
                                        acm.org:80/pubs/citations/
                                        proceedings/plan/202529/p111-
                                        gosling/.

Gosling:1996:JLT

[Gos96] James Gosling. Java language tutorial: your first jolt, 1996. 1 videocassette (82 min.).

Gosling:1997:FJ

[Gos97a] James Gosling. The feel of Java. Computer, 30(6):53–57, June 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gosling:1997:ENC

[Gos97b] James A. Gosling. The evolution of numerical computing in Java. World Wide Web document., 1997. URL http:

//java.sun.com/people/jag/
FP.html.

Gosling:1998:EJN

[Gos98a]  James Gosling.  Extensions

to Java for numerical computation. In ACM [ACM98a], page ?? CODEN CPEXEI. ISSN 1040-3108. LCCN ????

URL http://java.sun.com/
people/jag/FP.html.  Also

published as Concurrency: Practice and Experience, 10(11– 13), September 1998, CODEN CPEXEI, ISSN 1040-3108.

Gosling:1998:HCS

[Gos98b] James Gosling. How computer security works: The Java sandbox. Scientific American, 279(4):109–109 (Intl. ed. 78–??), October 1998. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (elec-

tronic). URL http://www.
sciam.com/1998/1098issue/
1098currentissue.html.

Gosling:19xx:JABa

[Gosxxa]  James Gosling.  Das Java

API, Band 1: Die Basispakete. Addison-Wesley, Reading, MA, USA, 19xx. ISBN 3-8273-1040-7 (??invalid ISBN??). LCCN ???? ????

Gosling:19xx:JABb

[Gosxxb] James Gosling. Das Java API, Band 2: Das Window Toolkit und Applets. Addison-Wes-ley, Reading, MA, USA, 19xx.

ISBN 3-8273-1084-9 (??invalid ISBN??). LCCN ???? ????

Gosling:19xx:JS

[Gosxxc] James Gosling. Java — Die Sprachspezifikation. Addison-Wesley, Reading, MA, USA, 19xx. ISBN 3-8273-1038-5 (??invalid ISBN??). LCCN ???? ????

Gottleber:1998:EHI

[Got98a] Timothy T. Gottleber. Excellent HTML with an introduction to Java applets. Irwin/McGraw Hill, Boston, MA, USA, 1998. ISBN 0-07-024772-2, 0-07-913675-3. xx + 428 pp. LCCN QA76.76.H94G68 1998.

Gottschalk:1998:TOI

[Got98b] K. D. Gottschalk. Technical overview of IBM’s Java initiatives. IBM Systems Journal, 37(3):308–??, ???? 1998. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

                                        URL http://www.almaden.
                                        ibm.com/journal/sj/373/gottschalk.
                                        html.

Garrett:1997:VCP

[GPF97] Doug Garrett, Anthony Potts, and David H. Friedel, Jr. Visual Café Programming FrontRunner: The Hands-on Guide to Mastering Java Development with Visual Café. Coriolis Group Books, Scottsdale, AZ, USA, 1997. ISBN 1-57610-160-6. xx + 417 pp. LCCN QA76.73.J38 G36 1997. US$29.99, CDN$41.99. URL

http://www.coriolis.com/;
 http://www.coriolis.com/
Site/MSIE/Books/Ind/cprfrmac.
htm;   http://www.coriolis.
com/site/msie/books/ind/cprfrmac.
htm.

Gimbel:1999:JBP

[GPHL99] M. Gimbel, M. Philippsen, B. Haumacher, and P. C. Lockemann. Java as a basis for parallel data mining in workstation clusters. Lecture Notes in Computer Science, 1593:884– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gladychev:1998:CRJa

[GPO98a] Pavel Gladychev, Ahmed Patel, and Donal O’Mahony. Cracking RC5 with Java applets. In ACM [ACM98a], page ?? CODEN CPEXEI. ISSN 1040-3108. LCCN ????

URL http://www.cs.ucsb.
edu/conferences/java98/papers/
rc5.pdf;   http://www.cs.
ucsb.edu/conferences/java98/

papers/rc5.ps. Also published as Concurrency: Practice and Experience, 10(11–13), September 1998, CODEN CPEXEI, ISSN 1040-3108.

Gladychev:1998:CRJb

[GPO98b] Pavel Gladychev, Ahmed Patel, and Donal O’Mahony. Cracking RC5 with Java applets. Concurrency: Practice and Experience, 10(11–13):1165–1171, September 1998. CODEN CPEXEI. ISSN 1040-3108. URL http:

//www3.interscience.wiley.

                                        com/cgi-bin/abstract?ID=10050408;
                                         http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=10050408&PLACEBO=IE.pdf.

Special Issue: Java for Highperformance Network Computing.

Goralski:1997:VEV

[GPV97] Walter Goralski, Matthew Poli, and Peter Vogel. VRML: exploring virtual worlds on the Internet. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1997. ISBN 0-13-486960-5, 0-614-14500-7. xi + 481 pp. LCCN QA76.76.H94G67 1997. US$39.95. Includes some discussion of Java.

Gulbransen:1996:WAJ

[GR96] David Gulbransen and Kenrick Rawlings. Web Applets mit Java. Markt & Technik, Haar bei München, Germany, October 1996. ISBN 3-8272-5182-

6. ???? pp. LCCN ????

                                        50 DM. URL http://www.
                                        mut.ch/html/buchinfo.htx?
                                        &x_pp=dJgxRqA9wlAYNijX&x_
                                        az=10&x_pk=VpLvjnsx7b5V&x_
                                        pa=B&x_pr=oS&x_bestell=25182&
                                        x_stichwort=Gulbransen&x_
                                        sprache=A&x_start=1&x_gefunden=
                                        4&x_query=&x_lastpage=buchfund&
                                        x_soft=&x_suchewo=A;   http:
                                        //www.mut.ch/profit.htx?&
                                        x_pk=kd84mf7ms9dm; http://
                                        www.mut.com/. Includes CD-

ROM.

Gulbransen:1997:CSA

[GR97] David Gulbransen and Kenrick Rawlings. Cree Sus Applets por Web con Java. Prentice Hall HispanoAmericana, Mexico, DF, Mexico / Nueva York, NY, USA, 1997. ISBN 968-880-782-6. xi + 307 (or 328??) pp. LCCN ???? 3120 Ptas. URL http://

prentice.com.mx/;   http:
//prentice.com.mx/libros/
cree_applets.html;   http:
//www.mcl.com.mx/sotano/prentice.
fm$Retrive?RecID=2641&html=

HTMLRecord1. Spanish translation of [GRD96]. Includes CD-ROM.

Graham:1996:JPE

[Gra96a] John R. Graham. Java programming: Exceptions and applets. WebServer Magazine: For Managers of World Wide Web Sites, 1(3):38–43, September/October 1996. ISSN 1087-

4232. URL http://www.cpg.
com.

Graham:1996:JPG

[Gra96b] John R. Graham. Java programming: Getting a handle on class and scope. WebServer Magazine: For Managers of World Wide Web Sites, 1(2):44– 49, July/August 1996. ISSN 1087-4232. URL http://www.

cpg.com.

Graham:1996:JPJ

[Gra96c] John R. Graham. Java programming: Java classes. Web-Server Magazine: For Managers of World Wide Web Sites, 1(1):52–54, 56, 58, 60, March/ June 1996. ISSN 1087-4232.

URL http://www.cpg.com.

Graham:1996:JPT

[Gra96d] John R. Graham. Java programming: Threads and network communication. Web-Server Magazine: For Managers of World Wide Web Sites, 1(4):34–36, 38, 39, November/ December 1996. ISSN 1087-

                                        4232. URL http://www.cpg.
                                        com.

Grand:1997:JLRa

[Gra97a] Mark Grand. Java Language Reference. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, January 1997. ISBN 1-56592-204-2. xiv + 448 pp. LCCN QA76.73.J38 G73 1997. US$24.95. URL http://www.

                                        ora.com/catalog/javalang/
                                        noframes.html.

Grand:1997:JLRb

[Gra97b]  Mark Grand.  Java Lan-

guage Reference. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 1997. ISBN 1-56592-326-X. xiv + 475 pp. LCCN QA76.73.J38 G73 1997b. US$32.95. URL http://www.

                                        ora.com/catalog/javalang2/
                                        ;   http://www.ora.com/catalog/
                                        {javalang}2/;   http://www.
                                        oreilly.com/catalog/javalang2.

Grand:1999:PJC

[Gra99a] Mark Grand. Patterns In Java: a Catalog of Reusable Design Patterns Illustrated With UML, Volume 1. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 1999. ISBN 0-471-25839-3. 640 pp. LCCN QA76.76.P37G73 1998. US$49.99.

Grand:1999:PJV

[Gra99b] Mark Grand. Patterns in Java, Vol. 2. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 1999. ISBN 0-471-25841-5. 368 pp. LCCN ???? US$49.99. Includes CD-ROM.

Gulbransen:1996:CWA

[GRD96] David Gulbransen, Kenrick Rawlings, and John December. Creating Web Applets With Java. Sams.net, Indianapolis, IN 46268, USA, 1996. ISBN 1-57521-070-3. 307 pp. LCCN QA76.73.J38G85 1996. US$39.99. URL http://www.

mcp.com/210167322173490/cgi-
bin/bag?isbn=1-57521-070-
3&last=/bookstore;   http:
//www.mcp.com/samsnet/.

Green:1995:MLS

[Gre95]  Heather Green.  Microsoft

to license Sun Java program. New York Times Syndicate, page ??, December 7,

1995. URL http://nytsyn.
com/live/News3/341_120795_
134125_6450.html.

Greenberg:1996:JAX

[Gre96a] Ross M. Greenberg. Java or Active X? maybe both. Computerworld, 30(39):98, September 23, 1996. CODEN CMP-WAB. ISSN 0010-4841.

Grehan:1996:WCD

[Gre96b] R. Grehan. Womb-to-tomb C++ development [C++ compiler]. BYTE Magazine, 21 (7):159–160, July 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1996:BBH

[Gre96c]  Rick Grehan.  Beyond the

bouncing heads: Use JFactory to quickly assemble Java applications. BYTE Magazine, 21 (10):216, October 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic). Discusses Rogue Wave Software’s JFactory, a GUI tool for rapid prototyping of Java screens.

Grehan:1997:JCV

[Gre97a] Rich Grehan. Javatalk: Convert Visual Basic to Java. BYTE Magazine, 22(3):188, March 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1997:JJA

[Gre97b] Rich Grehan. Javatalk: Jetassisted Java: Optima++ developers will feel at home in Powersoft’s PowerJ for Java. BYTE Magazine, 22(5):103– 104, May 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1997:JRP

[Gre97c] Rich Grehan. Javatalk: Rebuilt parts: Parts for Java delivers an excellent visual programming environment for Java. BYTE Magazine, 22(10):117–118, October 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1997:JRR

[Gre97d] Rick Grehan. Java’s RAD route to data access. BYTE Magazine, 22(2):192, February 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1997:JSL

[Gre97e] Rick Grehan. Javatalk: SuperCede lives up to its names. BYTE Magazine, 22(4):125– 126, April 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1997:JVJ

[Gre97f] Rick Grehan. Javatalk: Vibe Java: Feel that cross-platform Vibe is one thing; deploying it is something else. BYTE Magazine, 22(9):89–90, September 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic). Praises Visix’s [Nan95] Vibe springs-and-strut geometry management, but is severely critical of its nonstandardness and lack of security.

Grehan:1997:JVA

[Gre97g] Rick Grehan. Javatalk: Visual Age for Java: IBM’s Visual Age for Java isn’t just beans. BYTE Magazine, 22 (7):101–102, July 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1997:JWC

[Gre97h] Rick Grehan. Javatalk: Write cosmic code: Cosmo Code, soon to be combined with SGI’s Cosmo Worlds 3-D development system, is a worthy Java IDE on its own. BYTE Magazine, 22 (8):113–114, August 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1997:PGJ

[Gre97i] Rick Grehan. Poet goes Java — an OODBMS takes the plunge into Java. BYTE Magazine, 22(12):113–??, December 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1997:RDS

[Gre97j] Rick Grehan. Review: Development software: Java Development Kit, take two. BYTE Magazine, 22(4):145–146, April 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1997:UDM

[Gre97k] Rick Grehan. An uncertain discovery — Metrowerks’ Discover Programming will give you a decent introduction to Java, but it can be annoying. BYTE Magazine, 22(11):101–??, November 1997. CODEN BYTEDJ.

ISSN 0360-5280 (print), 1082-7838 (electronic).

Gregoire:1998:AVR

[Gre98a] David Remi Gregoire. Application of virtual reality modeling language (VRML), and the Java programming language to conceptual design of industrial tooling. ????, ????, ????, 1998. x + 97 pp.

Grehan:1998:MJP

[Gre98b]  Rich Grehan.  More Java

persistence: a look at another Java-enable database reveals what’s in store for you. BYTE Magazine, 23(1):93–94, January 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1998:OMR

[Gre98c] Rich Grehan. Object marries relational — Ardent’s Java Relational Binding turns a relational database into a Java objectoriented database management system. BYTE Magazine, 23 (3):101–102, March 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1998:FDO

[Gre98d] Rick Grehan. The federated database — Objectivity/DB, a Java-based OODBMS Plus, Powersoft’s component transaction server. BYTE Magazine, 23(2):101–102, February 1998. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1998:RJD

[Gre98e] Rick Grehan. Rapid Java development. BYTE Magazine, 23(6):109–??, June 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Grehan:1998:TBT

[Gre98f] Rick Grehan. TopLink bridges two worlds: a pure Java object-relational database system maps Java objects to relational databases. BYTE Magazine, 23(5):107–108, May 1998. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Griffith:1997:JMR

[Gri97a] Arthur Griffith. The Java Master Reference. I D G Books Worldwide, Indianapolis, IN, USA, May 1997. ISBN 0-7645-3084-4 (softcover). xv + 1615 pp. LCCN QA76.73.J38G75 1998. US$59.99. URL http://

                                        www.idgbooks.com/database/
                                        book_result.msql?uid=uidg81&
                                        isbn=0-7645-3084-4.

Grinzo:1997:EPN

[Gri97b] Lou Grinzo. Extended precision native integers for Java: When performance counts. Dr. Dobb’s Journal of Software Tools, 22 (11):42, 44–45, 91–92, November 1997. CODEN DDJOEB. ISSN 1044-789X.

Grinzo:1999:JQH

[Gri99] Lou Grinzo. Java Q and A: How do you run untrusted classes? Dr. Dobb’s Journal of Software Tools, 24(5):121–123, May 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/ftp/1999/1999_05/
jqa599.txt;   http://www.
ddj.com/ftp/1999/1999_05/
jqa599.zip.

Germain-Renaud:1999:JBC

[GRN99] Cécile Germain-Renaud and Vincent Néri. Java-based coupling for parallel predictiveadaptive domain decomposition. Scientific Programming, 7(2):185–189, ???? 1999. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:

//iospress.metapress.com/
app/home/contribution.asp%
3Fwasp=f277qlrwwjr5m4vxjywv%
26referrer=parent%26backto=
issue%2C8%2C8%3Bjournal%2C7%
2C9%3Blinkingpublicationresults%
2C1%2C1.

Groner:1996:JLS

[Gro96a] Daniel Groner. Java Language SuperBible. Macmillan Computer Publishing, Indianapolis, IN, USA, September 1996. ISBN 0-614-20294-9. ???? pp. LCCN ???? US$59.99.

Grosky:1996:NWH

[Gro96b]  William Grosky.  In the

news: Whither the high definition in HDTV? TV on the Net; telecommunications overhauled; videoconferencing market grows; ITU videoconferencing standards; credit on the net; Bell Labs brews a rival to Sun’s Java; 3D multimedia comes on a chip; new print output. IEEE MultiMedia, 3(1):6–9, Spring 1996. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://dlib.computer.org/
mu/books/mu1996/pdf/u1006.

pdf. Discusses the Inferno programming language.

Grosse:1996:NPC

[Gro96c] Eric Grosse. Network programming and CSE. IEEE Computational Science & Engineering, 3(2):40–41, Summer 1996. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). Discusses the emergence of systems like Java and Inferno for scientific visualization.

Groner:19xx:JAS

[Groxxa] Daniel Groner. Java API SuperBible. Markt & Technik, Haar bei München, Germany, 19xx. ISBN 3-8272-4507-9 (??invalid ISBN??). LCCN ????

119 DM. URL http://www.
mut.ch/html/buchinfo.htx?
&x_pp=dJgxRqA9wlAYNijX&x_
az=10&x_pk=VpLvjnsx7b5V&x_
pa=B&x_pr=oS&x_bestell=24507&
x_stichwort=Groner&x_sprache=
A&x_start=1&x_gefunden=2&
x_query=&x_lastpage=buchfund&
x_soft=&x_suchewo=A.

Group:19xx:KVJ

[Groxxb] Coriolis Group. Kickass Visual J++ Programming. Coriolis Group Books, Scottsdale, AZ, USA, 19xx. ISBN 1-57610-093-6. ???? pp. LCCN ???? US$50.

Gesing:1996:JWW

[GS96] Ted Gesing and Jeremy Schneider. Javascript for the World Wide Web: Visual Quickstart Guide. Visual quickstart guide. Peachpit Press, Inc., 1085 Keith Avenue, Berkeley, CA 94708, USA, December 1, 1996. ISBN 0-201-68814-X. ix + 182 pp. LCCN QA76.73.J38G47 1997. US$17.95.

Garfinkel:1997:WSC

[GS97a] Simson Garfinkel and Gene Spafford. Web Security & Commerce. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1997. ISBN 1-56592-269-7. xx + 483 pp. LCCN TK5105.59 .G385 1997. US$32.95. URL http://www.

ora.com/catalog/websec/. Cov-

ers Java, JavaScript, and ActiveX.

Geppert:1997:HIC

[GS97b] L. Geppert and W. Sweet.

http://: it has changed everything, especially our engineering thinking. IEEE Spectrum, 34 (1):23–26, January 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gray:1997:IDC

[GS97c] Paul A. Gray and Vaidy S. Sunderam. IceT: distributed computing and Java. Concurrency: Practice and Experience, 9(11):1161–1167, November 1997. CODEN CPEXEI. ISSN 1040-3108. URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=13820;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=13820&PLACEBO=IE.pdf. Spe-

cial Issue: Java for computational science and engineering — simulation and modeling II.

Golomb:1998:JQH

[GS98a] Kenneth Golomb and Thomas Sorgie. Java Q&A: How do I ensure secure communications from a Java applet? Dr. Dobb’s Journal of Software Tools, 23 (6):107–109, 124, June 1998. CODEN DDJOEB. ISSN 1044-789X.

Gong:1998:SSG

[GS98b] Li Gong and Roland Schemers. Signing, sealing, and guarding Java^{T M} objects. Lecture Notes in Computer Science, 1419:206– 216, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/
service/series/0558/bibs/
1419/14190206.htm;   http:
//link.springer-ny.com/link/
service/series/0558/papers/
1419/14190206.pdf.

Gray:1998:NLBa

[GS98c] Paul A. Gray and Vaidy S. Sunderam. Native-language-based distributed computing across network and filesystem boundaries. In ACM [ACM98a], page ?? CODEN CPEXEI. ISSN 1040-3108. LCCN ????

URL http://www.cs.ucsb.
edu/conferences/java98/papers/
icet.pdf;   http://www.cs.
ucsb.edu/conferences/java98/
papers/icet.ps. Also pub-

lished as Concurrency: Practice and Experience, 10(11– 13), September 1998, CODEN CPEXEI, ISSN 1040-3108.

Gray:1998:NLBb

[GS98d] Paul A. Gray and Vaidy S.

Sunderam. Native-language-based distributed computing across network and filesystem boundaries. Concurrency: Practice and Experience, 10(11–13):901–909, September 1998. CODEN CPEXEI. ISSN 1040-3108. URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=10050397;
http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=10050397&PLACEBO=IE.pdf.

Special Issue: Java for Highperformance Network Computing.

Groner:1997:JAS

[GSHP97] Daniel Groner, Todd Sundsted, K. C. Hopson, and Harish Prabandham. Java API Superbible. Markt & Technik, Haar bei München, Germany, 1997.

ISBN 3-8272-4507-9 (??invalid ISBN??). ???? pp. LCCN ????

119 DM. URL http://www.
mut.ch/html/buchinfo.htx?
&x_pp=dJgxRqA9wlAYNijX&x_
az=10&x_pk=VpLvjnsx7b5V&x_
pa=B&x_pr=oS&x_bestell=24507&
x_stichwort=Groner&x_sprache=
A&x_start=1&x_gefunden=2&
x_query=&x_lastpage=buchfund&
x_soft=&x_suchewo=A;   http:
//www.mut.ch/profit.htx?&
x_pk=kd84mf7ms9dm; http://

www.mut.com/. German translation of [Gro96a]. Includes CD-ROM.

Gargaro:1997:ACA

[GSTV97] A. Gargaro, G. Smith, R. J. Theriault, and R. A. Volz. Aria-Java communication in ADEPT. In ACM [ACM97b], pages 231–246. ISBN 0-89791-981-5. LCCN QA 76.73 A35 T75 1997. Theme title: Ada; the right choice for reliable software. ACM order number: 825970.

Gould:1999:GDI

[GSvD99] Daniel L. Gould, Rosemary M. Simpson, and Andries van Dam. Granularity in the design of interactive illustrations. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1):306–310, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Goodrich:1997:DSA

[GT97] Michael T. Goodrich and Roberto Tamassia. Data Structures and Algorithms in Java. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, January 1997. ISBN 0-471-19308-9 (paperback). xviii + 738 pp. LCCN QA76.73.J38 G66 1997. US$22. URL

http://www.loc.gov/catdir/
enhancements/fy0706/97039794-
d.html;   http://www.loc.
gov/catdir/toc/onix01/97039794.
html;   http://www.wiley.
com/college/cs2java/.

Gordon:1999:EJJ

[GT99] Rob Gordon and Stephen Talley. Essential JMF: Java Media Framework. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1999. ISBN 0-13-080104-6. xxxvii + 624 pp. LCCN QA76.73.J38G663 1998.

URL http://www.phptr.com/
ptrbooks/ptr_0130801046.html.

Goggin:19xx:JPF

[GTFxx] Terence Goggin, Matt Telles, and Steven Fraser. JBuilder Programming FrontRunner. Coriolis Group Books, Scottsdale, AZ, USA, 19xx. ISBN 1-57610-004-9. 450 pp. LCCN ???? US$29.99, CDN$41.99. URL

http://www.coriolis.com/.

Glossner:1998:JSPa

[GTV98a] John Glossner, Jesse Thilo, and Stamatis Vassiliadis. Java signal processing: FFTs with bytecodes. In ACM [ACM98a], page ?? CODEN CPEXEI. ISSN 1040-3108. LCCN ????

                                        URL http://www.cs.ucsb.
                                        edu/conferences/java98/papers/
                                        jfft.pdf;   http://www.cs.
                                        ucsb.edu/conferences/java98/
                                        papers/jfft.ps. Also pub-

lished as Concurrency: Practice and Experience, 10(11– 13), September 1998, CODEN CPEXEI, ISSN 1040-3108.

Glossner:1998:JSPb

[GTV98b] John Glossner, Jesse Thilo, and Stamatis Vassiliadis. Java signal processing: FFTs with bytecodes. Concurrency: Practice and Experience, 10(11–13): 1173–1178, September 1998. CODEN CPEXEI. ISSN 1040-3108. URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=10050412;
                                         http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=10050412&PLACEBO=IE.pdf.

Special Issue: Java for Highperformance Network Computing.

Gilliam:1999:PJ

[GTW99] Jason Gilliam, Charlton Ting, and R. Allen Wyke. Pure JavaScript. Howard W. Sams, Indianapolis, IN 46268, USA, 1999. ISBN 0-672-31547-5. xxxix + 1448 pp. LCCN QA76.73.J39 G54 1999. US$34.99.

Guang:19xx:EJ

[Guaxx] Ting Tzung Guang. Easy Java. Chuen Hwa Books, ??, Taiwan, 19xx. ISBN 957-21-1653-X (??invalid ISBN??). ???? pp.

LCCN ???? 430 NT$. URL

http://www.chwa.com.tw. In-

cludes CD-ROM.

Gulbransen:1996:JCA

[Gul96a] David Gulbransen. Java Creation d’Applets Pour le Web. Simon & Schuster Macmillan (France), Paris, France, August 1996. ISBN 2-7440-0163-5. ???? pp. LCCN ???? 129 FF. URL

http://www.ssm.fr/. Includes

CD-ROM.

Gulesian:1996:PW

[Gul96b] M. Gulesian. Plugging into the Web. DBMS, 9(13):69–70, 72, December 1996. CODEN DBM-SEO. ISSN 1041-5173.

Gulbransen:1998:IJA

[Gul98] David Gulbransen. 101 Instant Java Applets. Sams.net, Indianapolis, IN 46268, USA, March 1998. ISBN 1-57521-091-6. ???? pp. LCCN ???? US$40.

Gulbransen:19xx:WAJ

[Gulxx] David Gulbransen. Web Applets mit Java entwickeln. Markt & Technik, Haar bei München, Germany, 19xx. ISBN 3-8272-5182-6 (??invalid ISBN??). LCCN ???? 50

DM.  URL http://www.
mut.ch/html/buchinfo.htx?
&x_pp=dJgxRqA9wlAYNijX&x_
az=10&x_pk=VpLvjnsx7b5V&x_
pa=B&x_pr=oS&x_bestell=25182&
x_stichwort=Gulbransen&x_
sprache=A&x_start=1&x_gefunden=
4&x_query=&x_lastpage=buchfund&
x_soft=&x_suchewo=A.

Guth:1995:IAV

[Gut95a] Rob Guth. Interview: Arthur van Hoff, creator of Hot Java. IDG International News Service, page ??, December 14, 1995.

Guth:1995:NPB

[Gut95b] Rob Guth. News: 3Com plans bundles with Netscape, Whiteboard, Java. IDG International News Service, page ??, December 04, 1995.

Guth:1995:NCAb

[Gut95c] Rob Guth. News: Comdex Asia: Network to shift power balance, Gage says. IDG International News Service, page ??, October 31, 1995.

Guth:1995:NCAa

[Gut95d] Rob Guth. News: Comdex Asia: Sun plans to license Java to Lotus, Intuit. IDG International News Service, page ??, October 26, 1995.

Guth:1995:NFS

[Gut95e] Rob Guth. News: Fujitsu, Sega partner over network game services. IDG International News Service, page ??, October 31, 1995.

Guth:1995:NFL

[Gut95f] Rob Guth. News: Fujitsu to license Java, plans broad implementation. IDG International News Service, page ??, December 19, 1995.

Guth:1995:NJV

[Gut95g] Rob Guth. News: Japanese vendors wake up to smell of Java. IDG International News Service, page ??, December 12, 1995.

Guth:1995:NSB

[Gut95h] Rob Guth. News: Sun brews Java scheme for General Magic. IDG International News Service, page ??, December 08, 1995.

Guth:1995:NSS

[Gut95i] Rob Guth. News: Sun sees Visual Basic in Microsoft’s Java license. IDG International News Service, page ??, December 11, 1995.

Guth:1995:NST

[Gut95j] Rob Guth. News: Sun, Toshiba to announce broad agreement. IDG International News Service, page ??, September 21, 1995.

Guth:1995:NTS

[Gut95k] Rob Guth. News: Toshiba-Sun deal highlights Japan’s interest in Java. IDG International News Service, page ??, September 25, 1995.

Guth:1996:LJT

[Gut96] Rob Guth. Low-cost Java terminal on tap. Computerworld, 30 (18):42, April 29, 1996. CODEN CMPWAB. ISSN 0010-4841.

Gutz:1998:SSU

[Gut98] Steven Gutz. Up to Speed with Swing: User Interfaces with Java Foundation Classes. Manning Publications, Greenwich, CT, USA, May 1998. ISBN 1-884777-64-3. xxxviii + 498 pp. LCCN QA76.9.U83 G88 1998. US$43. URL http://

www.browsebooks.com/Gutz2/
index.html.

Gosling:1995:WJ

[Gv95] James Gosling and Arthur van Hoff. Winner: Java. PC Magazine, page ??, December 19, 1995. CODEN PCMGEP. ISSN 0888-8507.

URL http://www.zdnet.com/
~pcmag/issues/1422/pcm00073.
htm.

Gosling:1996:J

[Gv96] James Gosling and Arthur van Hoff. Java, 1996. 2 videocassettes (ca. 167 min.).

Glossner:1997:DE

[GV97a] C. J. Glossner and S. Vassiliadis. The Delft-Java engine. Lecture Notes in Computer Science, 1300:766–??, 1997. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Glossner:1997:DJE

[GV97b] C. J. Glossner and S. Vassiliadis. The Delft-Java engine. Lecture Notes in Computer Science, 1300:766–??, 1997. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gosling:1996:JABa

[GYt96a] James Gosling, Frank Yellin, and the Java-Team. Das Java API, Band 1: Die Basispakete. Addison-Wesley, Reading, MA, USA, 1996. ISBN 3-8273-1040-7. xxxvii + 441 pp. LCCN ????

89,90 DM. URL http://www.
addison-wesley.de/katalog/
item.ppml?textexpr=Java&id=

00075. German translation of [GYT96c].

Gosling:1996:JABb

[GYt96b] James Gosling, Frank Yellin, and the Java-Team. Das Java API: Band 2: Das Window Toolkit und Applets. Addison-Wesley, Reading, MA, USA, 1996. ISBN 3-8273-1084-9. xxxvii + 367 pp. LCCN ????

79,90 DM. URL http://www.
addison-wesley.de/katalog/
item.ppml?textexpr=Java&id=

00076. German translation of [GYT96d].

Gosling:1996:JAPa

[GYT96c] James Gosling, Frank Yellin, and The Java Team. The Java Application Programming Interface. Vol. 1: Core packages. The Java Series. Addison-Wesley, Reading, MA, USA, 1996. ISBN 0-201-63453-8. 494 (or 600??) pp. LCCN QA76.73.J38G67 1996. US$35.95. URL http:/

/www.aw.com/cp/javaseries.
html;   http://www.aw.com/
cseng/titles/0-201-63453-
8/.

Gosling:1996:JAPb

[GYT96d] James Gosling, Frank Yellin, and The Java Team. The Java Application Programming Interface. Vol 2: Window Toolkit and Applets. The Java Series. Addison-Wesley, Reading, MA, USA, 1996. ISBN 0-201-63459-7. 406 pp. LCCN QA76.73.J38G67 1996. US$35.95. URL http://www.

aw.com/cp/javaseries.html;
 http://www.aw.com/cseng/
titles/0-201-63459-7/.

Gosling:1997:AJVa

[GYT97a] James Gosling, Frank Yellin, and The Java Team. Les API de Java, Volume 1: Le noyau. International Thomson Publishing France, Paris, France, 1997. ISBN 2-84180-998-6. xxxi + 438 pp. LCCN ???? 328 FF. URL

http://www.awl.com/cseng/
javaseries/international/
france.htm;   http://www.
thomson.com/intluk.html. French

translation of [GYT96c]. Includes CD-ROM.

Gosling:1997:AJVb

[GYT97b] James Gosling, Frank Yellin, and The Java Team. Les API de Java, Volume 2: Applets et l’AWT. International Thomson Computer Press, 20 Park Plaza Suite 1001, Boston, MA 02116 USA, 1997. ISBN 2-84180-999-4. xxix + 349 pp. LCCN ???? 318 FF. URL

http://www.awl.com/cseng/
javaseries/international/
france.htm;   http://www.
thomson.com/intluk.html. French

translation of [GYT96d]. Includes CD-ROM.

Hertel:1997:LND

[H^{+}97] Matthias Hertel et al. Lotus Notes/Domino 4.6-Anwendungsentwicklung: Notes- und Internet-Applikationen mit LotusScript, Java und Visual Basic. Addison-Wesley Verlag, Bonn, Germany, December 1997. ISBN 3-8273-1320-1. ???? pp. LCCN

????  URL http://www.
addison-wesley.de/katalog/
item.ppml?id=00272.

Hamilton:19xx:JDC

[H^{+}xxa] Graham Hamilton et al. The Java Database Connectivity Specification and Tutorial. Addison-Wesley, Reading, MA, USA, 19xx. ISBN 0-201-92454-

4. ???? pp. LCCN ???? US$37.

Heuchert:19xx:CJA

[H^{+}xxb] Peter Heuchert et al. Creating Java Applications Using NetRexx. IBM Redbooks, 11400 Burnet Road, Austin, TX 78758-3493, USA, 19xx. ISBN ???? ???? pp. LCCN

???? ???? URL http://www.
redbooks.ibm.com/SG242216/
2216ht.htm.

Hojas:19xx:LJ

[H^{+}xxc] Luis Ignacio Hojas Hojas et al.

Lenguaje Java. Html Editores, ????, 19xx. ISBN 84-89807-02-7 (??invalid ISBN??). ???? pp. LCCN ???? ????

Hung:19xx:UJ

[H^{+}xxd] Kao Wen Hung et al. Ultimate Java. Unalis Corporation, ??, Taiwan, 19xx. ISBN 957-22-2378-4. LCCN ???? 320 NT$.

Hackl:1996:GLJ

[Hac96] Yann Hackl. Le grand livre de JAVA. Edition Micro Application, Paris, France, 1996. ISBN 2-7429-0764-5. 605 pp. LCCN ???? 195 FF. URL http:/

/www.microapp.com/. Includes

CD-ROM.

Hackvan:1998:QOS

[Hac98] Stig Hackvän. Not quite Open Source, but close. Linux Today, December 10, 1998. URL http:

//linuxtoday.com/news_story.
php3?ltsn=1998-12-10-001-
05-NW-CY.

Hadjerrouit:1998:CFI

[Had98a] Said Hadjerrouit. A constructivist framework for integrating the Java paradigm into the undergraduate curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3): 105–107, September 1998. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hadjerrouit:1998:JFP

[Had98b]  Said Hadjerrouit.  Java as

first programming language: a critical evaluation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(2):43–47, June 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hadjerrouit:1999:CAO

[Had99]  Said Hadjerrouit.  A con-

structivist approach to objectoriented design and programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(3): 171–174, September 1999. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hummel:1997:AJB

[HAKN97] Joseph Hummel, Ana Azevedo, David Kolson, and Alexandru Nicolau. Annotating the Java bytecodes in support of optimization. Concurrency: Practice and Experience, 9(11):1003–1016, November 1997. CODEN CPEXEI. ISSN 1040-3108. URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=13830;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=13830&PLACEBO=IE.pdf. Spe-

cial Issue: Java for computational science and engineering — simulation and modeling II.

Halfhill:1997:TWT

[Hal97a] Tom H. Halfhill. Today the Web, tomorrow the world. is Java a serious programming language? yes. will it become a Windows killer? maybe. BYTE Magazine, 22(1):68–71, 74, 76, 78, 80, January 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hall:1997:CWP

[Hal97b] Marty Hall. Core Web Programming: HTML, Java, CGI, & Javascript. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, December 1997. ISBN 0-13-625666-X. ???? pp. LCCN QA76.625.H35 1997.

                                        US$50.  URL http://www.
                                        prenhall.com/ptrbooks/ptr_
                                        013625666x.html.

Halfhill:1998:HSJ

[Hal98a] Tom R. Halfhill. How to soup up Java: Part 1: The rap on Java is that it’s too slow. proponents are inventing ways to speed it up, including better compilers. BYTE Magazine, 23(5):60–64, 66, 68, 70, 72, 74, May 1998. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Halfhill:1998:SRJ

[Hal98b] Tom R. Halfhill. Survey reveals Java adoption plans. BYTE Magazine, 23(3):26, 30, March 1998. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hall:1999:CWP

[Hal99] Marty Hall. Core Web Programming. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1999. ISBN 0-13-625666-X. xl + 1279 pp. LCCN QA76.625.H35 1998.

Hamilton:1996:JJT

[Ham96a] James Hamilton. Java and JDBC: tools supporting datacentric business application development. In Frieder and Wigglesworth [FW96], pages 121–138. ISBN 0-8186-7389-3, 0-8186-7390-7 (softbound), 0-8186-7391-5 (microfiche). LCCN QA76.76.D47I595 1996. IEEE catalog number 96TB100054.

Hamilton:1996:JSN

[Ham96b] Marc A. Hamilton. Java and the shift to net-centric computing. Computer, 29(8):31– 39, August 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hamit:1996:BLP

[Ham96c]  Francis Hamit.  Bill Lee

preaches Java gospel to Hollywood. Sun Observer, 10(5):4, May 1996. ISSN 1058-5400.

Hamblin:1997:AEJ

[Ham97a] Matt Hamblin. All eyes on Java’s purity, speed: Language poky, but users say they still love it. Computerworld, 31(23): 12, June 9, 1997. CODEN CM-PWAB. ISSN 0010-4841.

Hamilton:1997:NPA

[Ham97b] Scott Hamilton. New products: Agent-based distributed computing in Java: Voyager agents. E-commerce solutions for the Web: Concinity and Web Speed. software development: HiProf code profiler, distributed applications, clientserver to the Web. stuff. Computer, 30(5):97–98, May 1997.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hamilton:1997:NPE

[Ham97c] Scott Hamilton. New products: Embedded systems software: Real-time Java. embedded OADD. entrus for PCs. Privasuite. Computer, 30(4):107, April 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hamilton:19xx:JDAb

[Hamxxa] Graham Hamilton. JDBC Database Access with Java. DaeRim Publishing, Seoul, Korea, 19xx. ISBN 89-7280-395-2 (??invalid ISBN??). LCCN ???? 18,000 won. URL http:

                                        //www.drbook.co.kr/daerim/
                                        jdbc.htm.

Hamilton:19xx:JDJ

[Hamxxb] Graham Hamilton. JDBC Datenbankzugriff mit Java. Addison-Wesley, Reading, MA, USA, 19xx. ISBN 3-8273-1306-6 (??invalid ISBN??). LCCN ????

                                        80 DM. URL http://www.
                                        addison-wesley.de/katalog/
                                        item.ppml?id=00269.

Hang:1996:EJA

[Han96] Raksas Hang. An evaluation of Java and its application to the fifteen-puzzle. Thesis (m.s.), University of California, Santa Cruz, Santa Cruz, CA, USA, 1996. viii + 82 pp.

Han:1997:JMW

[Han97] Sangbeom Han. Java Mastering on which C Expert Gives a Lecture. Cyber Publishing, ??, Korea, 1997. ISBN 89-7636-027-3 (??Invalid checksum??). ???? pp. LCCN ???? 10,000 won. URL http://www.

chomdan.com/sister.html.

Hansen:1999:JIP

[Han99a] Per Brinch Hansen. Java’s insecure parallelism. ACM SIGPLAN Notices, 34(4):38–45, April 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hansen:1999:TCJ

[Han99b] Per Brinch Hansen. Technical correspondence: Java’s insecure parallelism. ACM SIG-PLAN Notices, 34(4):38–45, April 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://ftp.
informatik.rwth-aachen.de/
dblp/db/indices/a-tree/h/
Hansen:Per_Brinch.html.

Hare:1996:JWJ

[Har96a] Dwight Hare. Java and the Web join to create true Internet programming. Sun Developer NEWS, 1(1):1, 2, Fall 1996.

Harmon:1996:WDG

[Har96b] Trevor Harmon. Web Developer’s Guide to Visual J++ & ActiveX. Coriolis Group Books, Scottsdale, AZ, USA, 1996. ISBN 1-57610-062-6. xviii + 409 pp. LCCN QA76.73.J38H36 1996. US$39.99, CDN$55.99.

URL http://www.coriolis.
com/cgi-win/KBsearch.exe;
 http://www.course.com/templates/
catalog/detail.cfm?isbn=1-

57610-062-6. Includes CD-ROM.

Harold:1996:JDR

[Har96c] Elliotte Rusty Harold. The Java Developer’s Resource: a Tutorial and On-Line Supplement. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1996. ISBN 0-13-570789-7. vii + 608 pp. LCCN QA76.73.J38H37 1996. US$26.95. URL http://www.

prenhall.com/ptrbooks/ptr_
0135707897.html.

Harold:1996:JS

[Har96d] Elliotte Rusty Harold. Java Secrets. IDG Books, San Mateo, CA, USA, 1996. ISBN 0-7645-8007-8. xxvi + 885 pp. LCCN QA76.73.J38 H373 1997. US$60. URL http://

www.idgbooks.com/database/
book_result.msql?uid=uidg323&
isbn=0-7645-8007-8.

Harold:1996:PMJ

[Har96e] Elliotte Rusty Harold. PC Magazine Java Programming. Ziff-Davis Press, Emeryville, CA, USA, May 1996. ISBN 1-56276-407-1. ???? pp. LCCN ???? US$34.95. URL http://www.

mcp.com/zdpress.

Harrison:1996:NJF

[Har96f] Brad Harrison. Are Netscape and Java de facto standards. Internetwork, 7(5):8, May 1996. ISSN 1079-0373.

Harold:1997:JNP

[Har97a] Elliotte Rusty Harold. Java Network Programming Guide. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1997. ISBN 1-56592-227-1. xvii + 422 pp. LCCN QA76.73.J38 H37 1997. US$29.95. URL http://www.

ora.com/catalog/javanetwk/
noframes.html;   http://www.
oreilly.com/catalog/javanetwk.

Harold:1997:JWS

[Har97b] Elliotte Rusty Harold. JavaTalk: Writing servlets with Jeeves. HP Chronicle, 14(11): 14, 16, October 1997. ISSN 0892-2829.

Harold:1997:PRA

[Har97c] Elliotte Rusty Harold. Programmation Réseau avec Java. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, October 1997. ISBN 2-84177-034-6. LCCN ???? 220 FF. URL http:

//www.editions-oreilly.fr/
international/france/prog/
javanet.html.

Harold:1998:J

[Har98a] Elliotte Rusty Harold. JavaBeans. I D G Books Worldwide, Indianapolis, IN, USA, 1998. ISBN 0-7645-8052-

3. xxiii + 355 pp. LCCN QA76.73.J38H3733 1998. US$39.99.

URL http://www.idgbooks.
com/cgi-bin/db/fill_out_template.
pl?idgbook:0-7645-8052-3:
book-idg::uidg781.

Harrington:1998:JPP

[Har98b] Jan L. Harrington. Java programming: an IS perspective. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 1998. ISBN 0-471-19665-7 (paper). xx + 460 pp. LCCN QA76.73.J38 H374 1998. US$59.

Hartley:1998:AYJ

[Har98c] Stephen J. Hartley. “Alfonse, your Java is ready!”. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(1):247–251, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hardy:1999:JAG

[Har99a] Vincent J. Hardy. Java 2D API Graphics. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1999. ISBN 0-13-014266-2. xli + 509 pp. LCCN 385 .H334 2000. US$49.99.

Harold:1999:J

[Har99b] Elliotte Rusty Harold. Java I/O. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1999. ISBN 1-56592-485-1. xxvi + 568 pp. LCCN QA76.73.J38 H372 1999. US$32.95. URL http:

//metalab.unc.edu/javafaq/
books/javaio/;   http://www.
oreilly.com/catalog/javaio/

.

Harold:1999:JIO

[Har99c] Elliotte Rusty Harold. Java I/O. The Java series. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1999. ISBN 1-56592-485-1. xxvi + 568 pp. LCCN QA76.73.J38 H372 1999. US$32.95. URL http:

//www.oreilly.com/catalog/
9781565924857;   http://www.
oreilly.com/catalog/javaio/

.

Hartley:1999:AWH

[Har99d]  Stephen J. Hartley.  “Al-

fonse, wait here for my signal!”. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31 (1):58–62, March 1999. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Harmon:19xx:VJA

[Harxxa] Trevor Harmon. Visual J++ y ActiveX. Editorial Paraninfo, Madrid, Spain, 19xx. ISBN 84-283-2416-6. ???? pp. LCCN ???? ????

Harold:19xx:JNPb

[Harxxb] Elliotte Rusty Harold. Java Network Programming. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 19xx. ISBN 4-900900-56-7. LCCN ???? 4,300

                                        yen. URL http://www.ohmsha.
                                        co.jp/data/books/contents/
                                        4-900900-56-7.htm. Japanese

translation.

Harold:19xx:SJT

[Harxxc] Elliotte Rusty Harold. Los secretos de Java a tu alcance. Anaya Multimedia, Madrid, Spain, 19xx. ISBN 84-415-0342-7. LCCN ???? 5995

                                        Ptas. URL http://www.anaya.
                                        es/Catalogo/CGA?act=L&art=
                                        2317019&ed=2300.

Hatton:1997:APP

[Hat97] Tim Hatton. The Awesome Power of PowerJ. Manning Publications, Greenwich, CT, USA, November 1997. ISBN 1-884777-53-8. ???? pp. LCCN QA76.73.J38H377 1997.

                                        US$44.  URL http://www.
                                        manning.com/Hatton2/index.

html. Includes CD-ROM.

HAT:1998:HAT

[HAT98a] Heap analysis tool. Sun Microsystems, 1998. URL http://

                                        java.sun.com/people/billf/
                                        heap/.

Hatton:1998:APP

[Hat98b] Tim Hatton. The Awesome Power of Power++. Manning Publications, Greenwich, CT, USA, May 1998. ISBN 1-884777-54-6. ???? pp. LCCN QA76.73.C153H38 1998.

                                        US$44.  URL http://www.
                                        manning.com/Hatton3/index.
                                        html.

Haverlock:1998:OSC

[Hav98] Kevin Haverlock. Object serialization, C++, and Java: Exchanging data between a sender and a receiver. Dr. Dobb’s Journal of Software Tools, 23(8):32, 34, 36–37, August 1998. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ftp/1998/1998_
08/serial.zip.

Hayes:1995:EBJ

[Hay95a] Frank Hayes. Espresso brews up Java apps. Computerworld, 29(51):8, December 1995. CO-DEN CMPWAB. ISSN 0010-4841.

Hayes:1995:NJH

[Hay95b] Frank Hayes. News: Java is hot.

Computerworld, page ??, December 08, 1995. CODEN CM-PWAB. ISSN 0010-4841. URL

http://www.computerworld.
com/.

Hayes:1995:NSH

[Hay95c] Frank Hayes. News: Symantec hopes Espresso will make Java go down easier. Computerworld, page ??, December 15, 1995. CODEN CMPWAB. ISSN 0010-4841.

Hayes:1995:NTF

[Hay95d] Frank Hayes. News: Toolmakers focus on Internet development. Computerworld, page ??, December 22, 1995. CODEN CMPWAB. ISSN 0010-4841.

Hayes:1995:NVU

[Hay95e]  Frank Hayes.  News: VB

users wait and see about Web scripting tools. Computerworld, page ??, December 14, 1995. CODEN CMPWAB. ISSN 0010-4841. URL http://www.

computerworld.com/.

Hayes:1996:CIF

[Hay96a] Frank Hayes. Corporate IS frets about Java security, frowns on applet use. Computerworld, 30 (37):68, September 9, 1996. CO-DEN CMPWAB. ISSN 0010-4841. Notes that Boeing Co., the U.S. Defense Intelligence Agency, and Telstra Corporation have all banned, or otherwise discouraged, the use of Java because of security concerns.

Hayes:1996:DCC

[Hay96b] Frank Hayes. Developers can control Java access with firewall. Computerworld, 30(43):60, October 21, 1996. CODEN CM-PWAB. ISSN 0010-4841.

Hayes:1996:DGT

[Hay96c] Frank Hayes. Developers get taste of visual Java. Computerworld, 30(20):50, May 13, 1996. CODEN CMPWAB. ISSN 0010-4841.

Hayes:1996:DWA

[Hay96d] Frank Hayes. Developers worried about safety of Java and Active X. Computerworld, 30 (41):55, October 7, 1996. CO-DEN CMPWAB. ISSN 0010-4841.

Hayes:1996:DJR

[Hay96e] Frank Hayes. Does Java really need the refills? Computerworld, 30(40):49, 52, September 30, 1996. CODEN CMPWAB. ISSN 0010-4841.

Hayes:1996:JAX

[Hay96f] Frank Hayes. Java, Active X push for platforms. Computerworld, 30(39):6, September 23, 1996. CODEN CMPWAB. ISSN 0010-4841.

Hayes:1996:JBU

[Hay96g] Frank Hayes. Java bug; users shrug. Computerworld, 30(14): 16, April 1, 1996. CODEN CM-PWAB. ISSN 0010-4841.

Hayes:1996:JGW

[Hay96h] Frank Hayes. Java gets Windows flavoring. Computerworld, 30(43):24, October 21, 1996. CODEN CMPWAB. ISSN 0010-4841.

Hayes:1996:JJN

[Hay96i] Frank Hayes. Java jolts new add-on tools arena. Computerworld, 30(23):47, 50, June 3, 1996. CODEN CMPWAB. ISSN 0010-4841.

Hayes:1996:JMM

[Hay96j] Frank Hayes. Java makes Microsoft look like IBM. Computerworld, 30(50):58, December 9, 1996. CODEN CMP-WAB. ISSN 0010-4841.

Hayes:1996:JPT

[Hay96k] Frank Hayes. A Java pick-me-up for tired systems: Firms port language to boost aging platforms. Computerworld, 30(40): 12, September 30, 1996. CO-DEN CMPWAB. ISSN 0010-4841.

Hayes:1996:JTE

[Hay96l] Frank Hayes. Java-speak: Tools ease Internet applet conversion. Computerworld, 30(17):47, April 22, 1996. CODEN CM-PWAB. ISSN 0010-4841. Describes tools from a dozen companies that assist in converting C/C++ GUI code to Java.

Hayes:1996:JTB

[Hay96m] Frank Hayes. Java team breaks away from Sun. Computerworld, 30(6):32, February 5, 1996. CODEN CMPWAB. ISSN 0010-4841. URL http:/

/www.computerworld.com/.

Hayes:1996:JTS

[Hay96n] Frank Hayes. Java tools still in short supply. Computerworld, 30(34):50, August 19, 1996. CO-DEN CMPWAB. ISSN 0010-4841.

Hayes:1996:JCT

[Hay96o] Frank Hayes. JavaScript can trace user tracks. Computerworld, 30(12):1, 113, March 18, 1996. CODEN CMPWAB. ISSN 0010-4841.

Hayes:1996:LM

[Hay96p] Frank Hayes. Listen up, Microsoft. Computerworld, 30(47): 61, November 18, 1996. CO-DEN CMPWAB. ISSN 0010-4841. Chides Microsoft for not offering Java support on Windows 3.1.

Hayes:1996:OSW

[Hay96q] Frank Hayes. Oracle still working to add Java client. Computerworld, 30(41):55, October 7, 1996. CODEN CMPWAB. ISSN 0010-4841.

Hayes:1996:OVP

[Hay96r] Frank Hayes. OS vendors place Java support on 1996 menu. Computerworld, 30(18):6, April 29, 1996. CODEN CMPWAB. ISSN 0010-4841.

Hayes:1996:SJA

[Hay96s] Frank Hayes. Some Java applets stink up Web. Computerworld, 30(21):57, May 20, 1996. CO-DEN CMPWAB. ISSN 0010-4841.

Hayes:1996:SUS

[Hay96t] Frank Hayes. Some users are sipping Java slowly. Computerworld, 30(38):6, September 16, 1996. CODEN CMPWAB. ISSN 0010-4841. Discusses Java GUI builders from Symantec.

Hayes:1996:SJI

[Hay96u] Frank Hayes. Support for Java increases, but problems bubbling up. Computerworld, 30 (23):14, June 3, 1996. CODEN CMPWAB. ISSN 0010-4841.

Hayes:1996:TBJ

[Hay96v] Frank Hayes. Three books on Java target different audiences. Computerworld, 30(45): 59, November 4, 1996. CO-DEN CMPWAB. ISSN 0010-4841. Brief reviews of [DD97c, Lev96a, Laf96].

Hayes:1996:USK

[Hay96w] Frank Hayes. Users struggle to keep track of Java definitions. Computerworld, 30(29):53, July 15, 1996. CODEN CMPWAB. ISSN 0010-4841.

Hayes:1996:VBJ

[Hay96x] Frank Hayes. Vendors boost Java support, aim for enterprise. Computerworld, 30(36): 48, September 2, 1996. CODEN CMPWAB. ISSN 0010-4841.

Hayes:1996:VDF

[Hay96y] Frank Hayes. Vendors drag feet on Java for Win 3.1. Computerworld, 30(48):53, November 25, 1996. CODEN CMP-WAB. ISSN 0010-4841.

Hayes:1997:DLS

[Hay97a] Frank Hayes. Dirty little secrets of Java, ActiveX. Computerworld, 31(24):127, June 16, 1997. CODEN CMPWAB. ISSN 0010-4841. Criticizes Sun’s Java Development Kit for not yet having a just-in-time compiler, and browsers for not caching applets, and Microsoft for not addressing the severe security problems of ActiveX.

Hayes:1997:JPM

[Hay97b] Frank Hayes. Java packages make coding easier but miss opportunities. Computerworld, 31 (11):51, 56, March 17, 1997. CODEN CMPWAB. ISSN 0010-4841.

Hayes:1997:PMO

[Hay97c] Frank Hayes. Prediction: Microsoft Office for Java. Computerworld, 31(18):127, May 5, 1997. CODEN CMPWAB. ISSN 0010-4841. Has Microsoft given up on ActiveX? It sure looks like it from the way Microsoft’s top people have clammed up about it. . . . .

Hayes:1997:SRJ

[Hay97d] Frank Hayes. Special report:

Java: Product review: Corel Office for Java is a tantalizing but frustrating first draft: Baby steps. Computerworld, 31(19): 85, 87, May 12, 1997. CODEN CMPWAB. ISSN 0010-4841.

Hayashi:1998:TIT

[Hay98] Alden M. Hayashi. Thinking it through. Datamation, 44(1):46– ??, ???? 1998. CODEN DTM-NAT. ISSN 0011-6963.

Hayes:1995:NSM

[HB95] Frank Hayes and Jean S. Bozman. News: Sun Micro to field apps, tools with Java. Computerworld, page ??, November 10, 1995. CODEN CMPWAB. ISSN 0010-4841. URL http://

www.computerworld.com/.

Holder:1998:JGP

[HB98] Wayne Holder and Doug Bell. Java Game Programming for Dummies. IDG Books, San Mateo, CA, USA, 1998. ISBN 0-7645-0168-2 (softcover). xviii + 359 pp. LCCN QA76.76.C672H65 1998. US$22.49. URL http:/

/www.idgbooks.com/cgi-bin/
db/fill_out_template.pl?idgbook:
0-7645-0168-2:book-idg::uidg1388.

Includes CD-ROM.

Holzle:1997:JSS

[HBG^{+}97] Urs Hölzle, Lars Bak, Steffen Grarup, Robert Griesemer, and Srdjan Mitrovic. Java on steroids: Sun’s highperformance Java implementation. In IEEE [IEE97], page ?? ISBN ???? LCCN ????

Hartel:1999:OSJ

[HBL99] P. H. Hartel, M. J. Butler, and M. Levy. The operational semantics of a Java secure processor. Lecture Notes in Computer Science, 1523:313– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Horstmann:1997:CJV

[HC97] Cay S. Horstmann and Gary Cornell. Core Java 1.1: Volume 1: Fundamentals. SunSoft Press Java series. SunSoft Press, 2550 Garcia Avenue, Mountain View, CA 94043-1100, USA, third edition, 1997. ISBN 0-13-766957-7. ???? pp. LCCN QA76.73.J38C67 1997. US$40.

URL http://www.horstmann.
com/corejava.html;   http:
//www.sun.com/books/books/
Horstmann/Horstmann.html.

Includes CD-ROM.

Horspool:1998:TCJ

[HC98a] R. Nigel Horspool and Jason Corless. Tailored compression of Java class files. Software—Practice and Experience, 28(12):1253–1268, October 1998. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=1762;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=1762&PLACEBO=IE.pdf.

Horstmann:1998:CJV

[HC98b] Cay S. Horstmann and Gary Cornell. Core Java 1.1: Volume 2: Advanced Features. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1998. ISBN 0-13-766965-8 (softcover). 661 pp. LCCN QA76.73.J38 C67 1998. US$31.45. URL http:/

/www.fdds.com/books/books/
Horstmann2/Horstmann2.html;
 http://www.prenhall.com/
ptrbooks/ptr_0137669658.html.

Includes CD-ROM.

Horstmann:1998:CJB

[HC98c] Cay S. Horstmann and Gary Cornell. Core Java, Bd.2 Expertenwissen. Markt & Technik, Haar bei München, Germany, September 1998. ISBN 3-8272-9566-1 (??invalid ISBN??). LCCN ???? 90 DM.

Hunter:1998:JSP

[HC98d] Jason Hunter and William Crawford. Java Servlet Programming. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1998. ISBN 1-56592-391-X. 528 pp. LCCN ???? US$32.95. URL http:

                                        //www.oreilly.com/catalog/
                                        jservlet.

Hilerio:1999:HEI

[HC99a] Israel Hilerio and Weidong Chen. Herbal-T, enabling integration, interoperability, and reusability of Internet components. ACM SIGSOFT Software Engineering Notes, 24(2):49–58, March 1999. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Horstmann:1999:CJVa

[HC99b] Cay S. Horstmann and Gary Cornell. Core Java 1.2: Volume 1: Fundamentals. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, fourth edition, 1999. ISBN 0-13-081933-6. 742 pp. LCCN QA76.73.J38 C67 1999. US$42.99.

Horstmann:1999:CJVb

[HC99c] Cay S. Horstmann and Gary Cornell. Core Java 1.2: Volume 2: Advance Features. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, fourth edition, 1999. ISBN 0-13-081934-4. ca. 750 pp. LCCN ???? US$49.99.

Hawblitzel:1998:IMP

[HCC^{+}98] Chris Hawblitzel, Chi-Chao Chang, Grzegorz Czajkowski, Deyu Hu, and Thorsten von Eicken. Implementing multiple protection domains in Java. In USENIX [USE98e], page ?? ISBN 1-880446-94-4. LCCN QA76.76.O63

U88 1998. URL http://www.
usenix.org/publications/library/
proceedings/usenix98/hawblitzel.
html.

Hart:1998:FPF

[HCD^{+}98] Hal Hart, Jim Caristi, Robert Dewar, Mark Gerhardt, Drew Hamilton, Christopher Haynes, and Sam Rebelsky. The future of programming-are fundamental changes in computer science programs coming? (panel). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(1): 370–371, March 1998. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hamilton:1997:JDA

[HCF97a] Graham Hamilton, Rick Cattell, and Maydene Fisher. JDBC Database Access With Java: a Tutorial and Annotated Reference. Addison-Wesley, Reading, MA, USA, 1997. ISBN 0-201-30995-5. xvi + 462 pp. LCCN QA76.73.J38H35 1997. US$37.61. URL http:

//cseng.aw.com/bookdetail.
qry?ISBN=0-201-30995-5&ptype=

Hamilton:1997:JDJ

[HCF97b] Graham Hamilton, Rick Cattell, and Maydene Fisher. JDBC Datenbankzugriff mit Java. Addison-Wesley, Reading, MA, USA, November 1997. ISBN 3-8273-1306-

6. 500 pp. LCCN ????

                                        79,90 DM. URL http://www.
                                        addison-wesley.de/katalog/
                                        item.ppml?id=00269;   http:
                                        //www.addison-wesley.de/katalog/
                                        item.ppml?textexpr=Java&id=

00265. German translation of

[HCF97a].

Hsieh:1997:ONC

[HCJ^{+}97] Cheng-Hsueh A. Hsieh, Marie T. Conte, Teresa L. Johnson, John C. Gyllenhaal, and Wen mei W. Hwu. Optimizing NET compilers for improved Java performance. Computer, 30(6): 67–75, June 1997. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Headden:1995:GS

[Hea95] Susan Headden. The glow of success. U.S. news and world report, 119:63–64, December 4, 1995. CODEN XN-WRAV. ISSN 0041-5537.

Heck:1995:PRN

[Hec95] Mike Heck. Product review: Netscape Navigator sets own standards. InfoWorld, page ??, November 29, 1995. CODEN INWODU. ISSN 0199-6649.

                                        URL http://www.infoworld.
                                        com/.

Heib:1996:J

[Hei96] David Heib. JavaScript. Prima Publishing, Rocklin, CA, USA, 1996. ISBN 0-7615-0685-3. xii + 337 pp. LCCN QA76.73.J39K66 1996. US$34.95. URL http://

www.webcom.com/~prima.

Heinle:1997:DJC

[Hei97] Nick Heinle. Designing with JavaScript: creating dynamic Web pages. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1997. ISBN 1-56592-300-6. xii + 241 pp. LCCN QA76.73.J39 H45 1997. US$29.95. URL http://www.

ora.com/catalog/designjs/.

Includes CD-ROM.

Heller:1997:JTGa

[Hel97a]  Steve Heller.  Java Train-

ing Guide. AP Professional, Boston, MA, USA, August 1997. ISBN 0-12-339106-7. 350 pp. LCCN QA76.73.J38H45 1997. US$49.95. URL http://

www.apcatalog.com/cgi-bin/
AP?ISBN=0123391067&LOCATION=
US&FORM=FORM2. Includes CD-

ROM.

Heller:1997:JTGb

[Hel97b]  Steve Heller.  Java Train-

ing Guide. AP Professional, Boston, MA, USA, September 1997. ISBN 0-12-339107-5. ???? pp. LCCN QA76.73.J38H45 1997. US$10.

Heller:1997:WAJ

[Hel97c] Steve Heller. Who’s Afraid Of Java? AP Professional, Boston, MA, USA, 1997. ISBN 0-12-339101-6. xxx + 372 pp. LCCN QA76.73.J38H454 1997. US$39.95. URL http://www.

                                        apcatalog.com/cgi-bin/AP?
                                        ISBN=0123391016&LOCATION=
                                        US&FORM=FORM2. Includes CD-

ROM.

Hemrajani:1996:ESC

[Hem96a] Anil Hemrajani. Examining Symantec’s Café: Visual development meets Java programming. Dr. Dobb’s Journal of Software Tools, 21(8):78, 80, 82, August 1996. CODEN DDJOEB. ISSN 1044-789X.

Hemrajani:1996:JAW

[Hem96b] Anil Hemrajani. The Java Abstract Window Toolkit. Dr. Dobb’s Journal of Software Tools, 21(9):40, 42, 44, September 1996. CODEN DDJOEB. ISSN 1044-789X.

Hemrajani:1996:JCC

[Hem96c] Anil Hemrajani. The Java and C connection. C/C++ Users Journal, 14(9):43–46, 48– 50, September 1996. CODEN CCUJEX. ISSN 1075-2838.

Hendrich:1997:JF

[Hen97a] Norman Hendrich. Java für Fortgeschrittene. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. ISBN 3-540-61531-8. viii + 596 pp. LCCN ???? 48 DM. URL http:/

/www.springer.de/. Includes

CD-ROM.

Henson:1997:LEJ

[Hen97b] Scott G. Henson. Letter to the editor: Java and CRC feedback. Dr. Dobb’s Journal of Software Tools, 22(6):10, June 1997. CO-DEN DDJOEB. ISSN 1044-789X.

Herbst:1996:WI

[Her96a] Kris Herbst. Webfest IV. Internet World, 7(3):22–??, March 1, 1996. CODEN IERNE8. ISSN 1064-3923.

Herbst:1996:WIF

[Her96b] Kris Herbst. Webfest IV — the fourth World-Wide Web Conference in Boston was abuzz with Web talk. hot topics included Web commerce, Java, VRML, and HTML extensions. Internet World, 7(3): 22–??, ???? 1996. CODEN IERNE8. ISSN 1064-3923.

Hermelink:1997:JKJ

[Her97] Jan Hermelink. Java — Das Kompendium: Java-Tutorial, Programmierung von Applikationen, die Schnittstelle zu C und C++, Datenbankzugriff, ausführliche Referenz. Markt & Technik, Haar bei München, Germany, 1997. ISBN 3-8272-5195-8. ???? pp. LCCN ????

90 DM. URL http://www.mut.

com/. Includes CD-ROM.

Hermeling:19xx:JK

[Herxx]  Wegen Hermeling.  Java-

Kompendium. Markt & Technik, Haar bei München, Germany, 19xx. ISBN 3-8272-5195-8 (??invalid ISBN??). LCCN ???? 90 DM.

Heskett:1995:NCW

[Hes95] Ben Heskett. News: Companies to watch in 1996 — SunSoft Inc. Network World, page ??, December 22, 1995. ISSN 0887-7661 (print), 1944-7655 (electronic). URL http://

                                        www.iquest.net/index.html.

Hess:1997:DRN

[Hes97] Deborah Hess. Datapro report: New Java tools cozy up to servers. BYTE Magazine, 22(7): 28, July 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Heuring:1997:BRE

[Heu97] Vincent P. Heuring. Book review: Essential Java Fast, John Cowell. Operating Systems Review, 31(4):2, October 1997. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

HP:1999:CES

[Hew99] Hewlett–Packard Corp. Chai embedded software. World-Wide Web documents, 1999.

                                        URL http://www.chai.hp.

com/. The TurboChai compiler converts Java byte code to ANSI/ISO Standard C, which can then be compiled with native compilers. The Chai system is a clean-room implementation of Java intended for embedded applications. It is available for HP-UX 10.20 and Microsoft Windows NT 4.0 platforms.

Hudson:1998:CPG

[HFA^{+}98] Scott E. Hudson, Frank Flannery, C. Scott Ananian, Dan Wang, and Andrew W. Appel. CUP parser generator for Java, March 1998. URL http://www.

cs.princeton.edu/~appel/modern/
java/CUP/;   http://www.cs.
princeton.edu/~appel/modern/
java/CUP/manual.html. From

the manual: “This manual describes the basic operation and use of the Java(tm) Based Constructor of Useful Parsers (CUP for short). CUP is a system for generating LALR parsers from simple specifications. It serves the same role as the widely used program YACC [1] and in fact offers most of the features of YACC. However, CUP is written in Java, uses specifications including embedded Java code, and produces parsers which are implemented in Java.”.

Harms:1996:WSP

[HFR96] David Harms, Barton C.

Fiske, and Jeffrey C. Rice. Web Site Programming With Java. McGraw-Hill, New York, NY, USA, 1996. ISBN 0-07-912986-2. xxix + 578 pp. LCCN TK5105.888 .H37

1996.  US$39.95.  URL

                                        http://www.incunabula.com/
                                        websitejava/index.html. In-

cludes CD-ROM. Covers Java 1.0.

Harms:1998:WSP

[HFR98] David Harms, Barton C. Fiske, and Jeffrey C. Rice. Web Site Programming With Java 1.1. McGraw-Hill, New York, NY, USA, January 1, 1998. ISBN 0-07-913178-6. 604 pp. LCCN ???? US$39.95.

Hayes:1996:PPS

[HG96a] Frank Hayes and Sharon Gaudin. Not a prime-time player: Sun’s Java lacks supporting tools, middleware, third-party libraries. Computerworld, 30(44):45, October 28, 1996. CODEN CMPWAB. ISSN 0010-4841.

Hayes:1996:SPJ

[HG96b] Frank Hayes and Sharon Gaudin. Sun pours Java client. Computerworld, 30(41):1, 15, October 7, 1996. CODEN CM-PWAB. ISSN 0010-4841.

Hamilton:19xx:JST

[HGxx] Graham Hamilton and David Geary. JavaBeans Specifications & Tutorial. Addison-Wesley, Reading, MA, USA, 19xx. ISBN 0-201-31001-5. ???? pp. LCCN ???? ????

Hsieh:1996:JBN

[HGH96] C.-H. A. Hsieh, J. C. Gyllenhaal, and W. W. Hwu. Java bytecode to native code translation: the Caffeine prototype and preliminary results. In IEEE [IEE96l], page ?? ISBN 0-8186-7641-8. LCCN QA76.9 .A73 I585 1996.

Hughes:1998:AJC

[HH98] Merlin Hughes and Conrad Hughes. Applied Java Cryptography. Manning Publications, Greenwich, CT, USA, May 1998. ISBN 1-884777-63-5. LCCN ???? US$37.

Harsu:1998:LIF

[HHK98] M. Harsu, J. Hautamaeki, and K. Koskimies. A language implementation framework in java. Lecture Notes in Computer Science, 1357:141–??, 1998. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hughes:1997:JNP

[HHSW97a] Merlin Hughes, Conrad Hughes, Michael Shoffner, and Maria Winslow, editors. Java Network Programming. Prentice-Hall and Manning Publications, Englewood Cliffs, NJ 07632, USA and Greenwich, CT, USA, January 1997. ISBN 1-884777-34-1 (Manning), 0-13-841206-5 (Prentice-Hall). xxiv + 519 pp. LCCN QA76.73.J38J377 1997. US$44.95. URL http://www.

prenhall.com/ptrbooks/ptr_
0138412065.html.  Includes

CD-ROM.

Hughes:1997:JNPa

[HHSW97b] Merlin Hughes, Conrad Hughes, Michael Shoffner, and Maria Winslow, editors. Java Network Programming. Prentice-Hall and Manning Publications, Englewood Cliffs, NJ 07632, USA and Greenwich, CT, USA, January 1997. ISBN 1-884777-34-1 (Manning), 0-13-841206-5 (Prentice-Hall). xxiv + 519 pp. LCCN QA76.73.J38J377 1997. US$44.95. URL http://www.

prenhall.com/ptrbooks/ptr_
0138412065.html.  Includes

CD-ROM.

Hughes:1997:JNPb

[HHW97] Merlin Hughes, Conrad Hughes, and Maria Winslow. Java Network Programming. Manning Publications, Greenwich, CT, USA, second edition, September 1997. ISBN 1-884777-49-X (softcover). ???? pp. LCCN QA76.73.J38J377 1997. US$38.65.

Harrison:1999:LRP

[HHW^{+}99] Brian Harrison, Philip Hendrickson, Murel Warren, Jr., Lee Kamentsky, Ron Gutman, Brenton Hoff, Martin Handwerker, Tom Culliton, and Aspi Havewala. Letters: Real programmer’s hate Cobol; 1984; Hilbert curves; grepping and globbing; testing Java classes; the version control process. Dr. Dobb’s Journal of Software Tools, 24(9):10, 12, September 1999. CODEN DDJOEB. ISSN 1044-789X.

Hopson:1996:DAC

[HI96a] K. C. Hopson and Stephen E. Ingram. Desenvolvendo Applets com Java. Editore Campus Ltda., Rua Sete de Setembro, 111/16, Rio de Janeiro - RJ - 20050-002, Brazil, 1996. ISBN ???? 668 pp. LCCN ???? R$ 83,00. URL http:

//www.campus.com.br/about-
us.htm;   http://www.campus.
com.br/catalogo/resposta.

cfm. Portuguese translation of [HI96b].

Hopson:1996:DPJ

[HI96b] K. C. (Casey) Hopson and Stephen E. Ingram. Developing Professional Java Applets. Sams.net, Indianapolis, IN 46268, USA, July 1, 1996. ISBN 1-57521-083-5. xv + 528 pp. LCCN QA76.73.J38H66 1996. US$49.99. URL http://www.

mcp.com/251356511370317/cgi-
bin/bag?isbn=1-57521-083-
5&last=/bookstore;   http:
//www.mcp.com/samsnet/.

Hopson:19xx:DAC

[HIxxa] K. C. Hopson and Stephen E.

Ingram. Desenvolvendo Applets com Java. Editore Campus Ltda., Rua Sete de Setembro, 111/16, Rio de Janeiro - RJ - 20050-002, Brazil, 19xx. ISBN ???? LCCN ???? R$83.

Hopson:19xx:DPJ

[HIxxb] K. C. Hopson and Stephen E.

Ingram. Developing Professional Java Applets. DaeRim Publishing, Seoul, Korea, 19xx. LCCN ???? 18,000 won.

URL http://www.drbook.co.
kr/daerim/jvpapp.htm. Ko-

rean translation.

Hibbard:1996:MSA

[Hib96] Justin Hibbard. Microsoft, Sun add Java to Web servers. Computerworld, 30(51):12, December 16, 1996. CODEN CMP-WAB. ISSN 0010-4841.

Hibbard:1997:SJM

[Hib97] Justin Hibbard. Speeding up Java/mainframe links. Computerworld, 31(9):28, March 3, 1997. CODEN CMPWAB. ISSN 0010-4841.

Hierle:1997:HJI

[Hie97]  Alexander Hierle.  HTML

und Java: Internet-Software kreativ nutzen. Sybex Verlag GmbH, Erkrather Straße 345-349, 40231 Düsseldorf, Germany, 1997. ISBN 3-8155-7509-5. ???? pp. LCCN ???? 20

                                        DM. URL http://www.sybex.

de/. Includes two CD-ROMs.

Hightower:1998:WRD

[Hig98] Lauren Hightower. The Web Report database reporting tool. Dr. Dobb’s Journal of Software Tools, 23(10):90, 92–93, 95, 113–114, October 1998. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/1998/1998_10/
                                        webrept.txt;   http://www.
                                        ddj.com/ftp/1998/1998_10/
                                        webrept.zip.

Hines:1996:PNJ

[Hin96] J. R. Hines. Program notes: Java: jive?; key klatch; good reading. IEEE Spectrum, 33 (3):15, March 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hines:1997:VMJ

[Hin97] J. R. Hines. Virtual machines jockey for position. IEEE Spectrum, 34(7):16, July 1997. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hines:1998:SET

[Hin98] J. Hines. Software engineering [technology 1998 analysis and forecast]. IEEE Spectrum, 35 (1):48–51, January 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hirano:1997:HDE

[Hir97] S. Hirano. HORB: Distributed execution of Java programs. Lecture Notes in Computer Science, 1274:29–??, 1997. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hiura:1997:PII

[Hiu97a] H. Hiura. Platform independent input method protocol for Internet, NC, and Java. In Unicode Consortium [Uni97b], pages A14–?? ISBN ???? LCCN ???? URL http:

//www.unicode.org/unicode/
iuc11/index.html;   http:
//www.unicode.org/unicode/
uni2book/u2ord.html.

Hiura:1997:PIIb

[Hiu97b] H. Hiura. Platform independent input method protocol for Internet, NC, and Java. In Unicode Consortium [Uni97b], pages A14–?? ISBN ????

LCCN ???? URL http://www.
unicode.org/unicode/iuc11/
program.html.

Hiura:1997:PIIa

[Hiu97c] Hideki Hiura. Platform independent input method protocol for the Internet, network computer (NC) and Java. In Unicode Consortium [Uni97a], page ?? ISBN ???? LCCN

????  URL http://www.
unicode.org/unicode/iuc10/
program.html.

Hiura:1998:IDI

[Hiu98] Hideki Hiura. IIIMF — distributed input method architecture for the Internet, network computer, and Java. In Unicode Consortium [Uni98a], page ?? ISBN ???? LCCN

????  URL http://www.
unicode.org/unicode/iuc12/
program.html.

Hendin:1998:MVR

[HJS98] O. Hendin, N. W. John, and O. Shochet. Medical volume rendering over the WWW using VRML and JAVA. In Westwood et al. [W^{+}98], pages 34– 40. ISBN 90-5199-386-2, 4-274-90210-2. ISSN 0926-9630. LCCN W26.5 M45 1998.

Hericko:1998:JDO

[HJZ^{+}98] Marjan Hericko, Matjaz B. Juric, Ales Zivkovic, Ivan Rozman, Tomaz Domajnko, and Marjan Krisper. Java and distributed object models: An analysis. ACM SIGPLAN Notices, 33(12):57–65, December 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Higham:1998:JMC

[HK98] L. Higham and J. Kawash. Java:

Memory consistency and process coordination. Lecture Notes in Computer Science, 1499:201– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Husted:1999:SSJ

[HK99] Robert Husted and J. J. Kuslich. Server-side JavaScript: developing integrated Web applications. Addison-Wesley, Reading, MA, USA, 1999. ISBN 0-201-43329-X, 0-201-61624-6 (CD-ROM). xiv + 316 pp. LCCN QA76.73.J39H877 1999. US$39.95.

Hackl:1996:GBJ

[HKK96] Yan Hackl, Michael Knapp, and Bernd Kretschmer. Das große Buch Java. DATA-Becker, ????, 1996. ISBN 3-8158-1259-3. 606 pp. LCCN ???? 69

DM. URL http://www.data-
becker.de/.

Holman:1997:IJ

[HL97a] Brian Holman and William Lund. Instant JavaScript. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1997. ISBN 0-13-268434-9, 0-614-20316-3. xx + 399 pp. LCCN QA76.73.J39 H64 1996. US$29.95.

Husain:1997:JDR

[HL97b] Kamran Husain and Jason Levitt. JavaScript developer’s resource: client-side programming using HTML, Netscape plug-ins and Java applets. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1997. ISBN 0-13-267923-X. xxiv + 568 pp. LCCN QA76.73.J39 H87 1997. US$39.95.

Husain:1997:JDRa

[HL97c] Kamran Husain and Jason Levitt. JavaScript developer’s resource: client-side programming using HTML, Netscape plug-ins and Java applets. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1997. ISBN 0-13-267923-X. xxiv + 568 pp. LCCN QA76.73.J39 H87 1997. US$39.95.

Husain:1997:JDRb

[HL97d] Kamran Husain and Jason Levitt. JavaScript developer’s resource: server-side programming. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1997. ISBN 0-13-270786-1. 300 pp. LCCN QA76.73.J39 H87 1997. US$39.95.

Hittleman:1999:JQW

[HL99] Kenneth Hittleman and Ted Leung. Java Q and A: What’s in store when moving from JDK 1.1 to JDK 1.2? Dr. Dobb’s Journal of Software Tools, 24 (1):112–115, January 1999. CO-DEN DDJOEB. ISSN 1044-

                                        789X. URL http://www.ddj.

com/ddj/ftp/1999/1999_01/
jqa199.txt.

Hylands:1997:TUI

[HLR97] Christopher Hylands, Edward A. Lee, and H. John Reekie. The Tycho user interface system. In USENIX [USE97a], pages 149–?? ISBN 1-880446-87-1. LCCN QA76.73.T44 T44 1997.

Hunt:1997:DCM

[HLRT97] J. J. Hunt, F. Lamers, J. Reuter, and W. F. Tichy. Distributed configuration management via Java and the World Wide Web. Lecture Notes in Computer Science, 1235:161–??, 1997. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hwang:19xx:JI

[HLTxx] C. K. Hwang, C. M. Lee, and S. Y. Tzeng. Java Introduction. GOTOP Information Inc., 5F, No.7, Lane 50, Sec.3 Nan Kang Road Taipei, Taiwan; Unit 1905,Metro Plaza Tower 2, No.223 Hing Fong Road, Kwai Chung, N.T., Hong Kong, 19xx. ISBN 957-641-768-9 (??invalid ISBN??). 500 pp. LCCN ???? 460 NT$. Includes CD-ROM.

Hunt:1998:KJA

[HM98] John Hunt and Alexander Mc-Manus. Key Java: Advanced Tips and Techniques. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. ISBN 3-540-76259-0. xvii + 331 pp. LCCN QA76.73.J38 H87 1998. ????

Hacking:1995:INW

[HMG^{+}95] Edwin Hacking, Denise Marcucci, John Gage, Arthur van Hoff, Bill Joy, Geoffrey Baehr, and Karl Jacob. Internet: the next wave, July 11, 1995. 1 videocassette (60 min.).

Hayes:1996:DUL

[HN96] Frank Hayes and Kim S. Nash. Desktop users left thirsting for Java. Computerworld, 30(17):1, April 22, 1996. CODEN CMP-WAB. ISSN 0010-4841.

Hardwick:1997:ISW

[HNS97a] Jonathan C. Hardwick, Girija Narlikar, and Jay Sipelstein. Interactive simulations on the Web: compiling N ESL into Java. Concurrency: Practice and Experience, 9(11):1075–1089, November 1997. CODEN CPEXEI. ISSN 1040-3108. URL http:

                                        //www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=13829;
                                         http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=13829&PLACEBO=IE.pdf. Spe-

cial Issue: Java for computational science and engineering — simulation and modeling II.

Hummel:1997:SPJ

[HNS97b] Susan Flynn Hummel, Ton Ngo, and Harini Srinivasan. SPMD programming in Java. Concurrency: Practice and Experience, 9(6):621–631, June 1997. CODEN CPEXEI. ISSN 1040-3108. URL http:

                                        //www3.interscience.wiley.

com/cgi-bin/abstract?ID=13879;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=13879&PLACEBO=IE.pdf.

Hansmann:1999:SCA

[HNSS99] U. Hansmann, M. S. Nicklous, T. Schäck, and F. Seliger. Smart Card Application Development Using Java. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. ISBN 3-540-65829-7. 304 pp. US$49.95.

Hangal:1999:PAV

[HO99] Sudheendra Hangal and Mike O’Connor. Performance analysis and validation of the picoJava processor. IEEE Micro, 19(3):66–72, May/June 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL

http://dlib.computer.org/
mi/books/mi1999/pdf/m3066.
pdf;   http://www.computer.
org/micro/mi1999/m3066abs.
htm.

Hobbs:1997:DJT

[Hob97a] Ashton Hobbs. Datenbankprogrammierung mit JDBC in 21 Tagen. Markt & Technik, Haar bei München, Germany, 1997. ISBN 3-8272-2005-X (??invalid ISBN??). 579 pp. LCCN ????

99,95 DM. URL http://www.
mut.ch/html/buchtext.htx?
&x_pp=STb4XDV284xv2gpf&x_
pk=Direkt&x_pa=&x_pr=Fr&x_
artikelnr=22005;   http:/
/www.mut.ch/profit.htx?&x_

                                        pk=kd84mf7ms9dm; http://www.

mut.com/. Includes CD-ROM.

Hobbs:1997:PCJ

[Hob97b] Ashton Hobbs. Programmare con JDBC. Jackson Libri, via XXV Aprile, 39 - 20091 Bresso (MI), Italy, November 1997. ISBN ???? LCCN ???? 65,000 L.

Hobbs:1997:TYD

[Hob97c] Ashton Hobbs. Teach Yourself Database Programming with JDBC in 21 Days. Sams.net, Indianapolis, IN 46268, USA, April 1, 1997. ISBN 1-57521-123-8. xxvii + 524 pp. LCCN QA76.625.H63 1997. US$39.99. URL http://

                                        merchant.superlibrary.com:
                                        8000/catalog/hg/PRODUCT/PAGE/
                                        15752/bud/1575211238.html.

Hobbs:1998:APB

[Hob98a] Ashton Hobbs. Aprend Program Base de Datos con JDBC en 21 Dias. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1998. ISBN 970-17-0084-8 (??invalid ISBN??). LCCN ???? ????

Hobuss:1998:BAW

[Hob98b]  James J. Hobuss.  Build-

ing Access Web Sites. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1998. ISBN 0-13-079830-4. 512 pp. LCCN QA76.9.D3H589 1998. US$44.95. Discusses Java, JavaScript, and HTML.

Hobuss:1998:BOW

[Hob98c] James J. Hobuss. Building Oracle Websites. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1998. ISBN 0-13-079841-X. xii + 488 pp. LCCN QA76.625.H64 1998. US$44.95. Discusses Java, JavaScript, and HTML.

Hobuss:1998:BSW

[Hob98d] James J. Hobuss. Building Sybase Websites. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1998. ISBN 0-13-798357-3. xxii + 437 pp. LCCN QA76.625.H642 1998. US$44.95. Discusses Java, JavaScript, and HTML.

Hobbs:19xx:DJT

[Hobxx] Ashtom Hobbs. Datenbankprogrammierung mit JDBC in 21 Tagen. Markt & Technik, Haar bei München, Germany, 19xx. ISBN 3-8272-2005-X (??invalid ISBN??). LCCN ????

100 DM. URL http://www.
mut.ch/html/buchtext.htx?
&x_pp=STb4XDV284xv2gpf&x_
pk=Direkt&x_pa=&x_pr=Fr&x_
artikelnr=22005.

Hoffman:1997:SLB

[Hof97] Thomas Hoffman. Sun to link banks via Java. Computerworld, 31(25):1, 16, June 23, 1997. CO-DEN CMPWAB. ISSN 0010-4841.

Holden:1996:MTW

[Hol96a] Julie Holden. My take on the whole Java thing. Internetwork, 7(3):18, March 1996. ISSN 1079-0373.

Holzner:1996:JPLa

[Hol96b] Steve Holzner. Java Programming with Latté. Henry Holt and Company, Incorporated, New York, NY, USA, November 1996. ISBN 0-614-20301-5. ???? pp. LCCN ???? US$39.95.

Holzner:1996:JWP

[Hol96c] Steven Holzner. Java Workshop Programming. MIS Press and M&T Books, P. O. Box 5277, Portland, OR 97208-5277, USA, Tel: (503) 282-5215 and M&T Publishing, Inc., 501 Galveston Drive, Redwood City, CA 94063, USA, 1996. ISBN 1-55851-491-0. xv + 464 pp. LCCN QA76.73.J38 H65 1996. US$34.95. URL http://www.

                                        mandt.com/javaworkshopprog.
                                        htm;   www.mispress.com.

Holub:1997:JUH

[Hol97a] Allen I. Holub. Java Under the Hood. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 1997. ISBN 1-55860-457-X. 450 (est.) pp. LCCN ???? US$40.

Holzner:1997:MJW

[Hol97b] Steven Holzner. Il Manuale di Java Workshop. Jackson Libri, via XXV Aprile, 39 - 20091 Bresso (MI), Italy, May 1997. ISBN ???? LCCN ???? 55,000 L.

Holzner:1997:JT

[Hol97c] Steven Holzner. Java 1.1, Das Trainingsbuch. Sybex Verlag GmbH, Erkrather Straße 345-349, 40231 Düsseldorf, Germany, 1997. ISBN 3-8155-2061-4. x + 559 pp. LCCN ???? 50

DM. URL http://www.sybex.
de/bestell/tp/tp02061.htm.

Includes CD-ROM.

Holzner:1997:JNEa

[Hol97d] Steven Holzner. Java 1.1: no experience required, will train. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, March 1997. ISBN 0-7821-2083-0. 521 pp. LCCN QA76.73.J38 H64 1997. US$34.99. URL http://www.

sybex.com/cgi-bin/bookpg.
pl?2083back.html.

Holzner:1997:JNEb

[Hol97e] Steven Holzner. Java 1.2: No Experience Required. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, second edition, December 1997. ISBN 0-7821-2171-3 (softcover). ???? pp. LCCN ???? US$19.99. Includes CD-ROM.

Holzner:1997:JPJ

[Hol97f] Steven Holzner. Java Programming With JBuilder. MIS Press and M&T Books, P. O. Box 5277, Portland, OR 97208-5277, USA, Tel: (503) 282-5215 and M&T Publishing, Inc., 501 Galveston Drive, Redwood City, CA 94063, USA, 1997. ISBN 1-55851-507-0. xvii + 477 pp.

LCCN QA76.73.J39H65 1997. US$39.95. URL http://www.

                                        mandt.com/javaprogjbuilder.

htm. Includes CD-ROM.

Holzner:1997:MNI

[Hol97g]  Steven Holzner.  Mastering

Netscape IFC. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, May 1997. ISBN 0-7821-2116-0. xviii + 816 pp. LCCN QA76.73.J38H653 1997.

                                        US$40.  URL http://www.
                                        sybex.com/cgi-bin/bookpg.
                                        pl?2116back.html.

Holzner:1997:MVJ

[Hol97h] Steven Holzner. Mastering Visual J++. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, February 1997. ISBN 0-7821-2041-5. xxi + 681 pp. LCCN QA76.73.J38 H654 1997 Interactive Learning Center.

                                        US$40.  URL http://www.
                                        sybex.com/cgi-bin/bookpg.
                                        pl?2041back.html.

Holzner:1997:VJN

[Hol97i] Steven Holzner. Visual J++: No Experience Required. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, March 1997. ISBN 0-7821-2078-4. xvii + 524 pp. LCCN QA76.73.J38H654 1997.

                                        US$30.  URL http://www.
                                        sybex.com/cgi-bin/bookpg.
                                        pl?2078back.html.

Holmes:1998:PJ

[Hol98a]  Barry Holmes.  Program-

ming with Java. Jones and Bartlett Publishers, ????, March 1998. ISBN 0-7637-0707-4. LCCN ???? US$49.

URL http://www.jbpub.com/
catalog/Detail.CFM?titles_
_ISBN=0763707074.

Holzner:1998:DVJ

[Hol98b] Steve Holzner. Dominando o Visual J++ “A B’blia”. Makron Books do Brasil Editora Ltda., Rua Tabapuã 1348, São Paulo SP - 04533-004 - Brasil, April 1998. ISBN 85-346-0849-0 (??invalid ISBN??). LCCN ???? R$89. URL http:/

/safe.tesla.com.br/makron/
options/livro.asp?id=655.

Holt:1999:EWT

[Hol99] William Holt. The embedded window toolkit: porting Java’s AWT to an embedded system. Thesis (m.s.), University of California, Santa Cruz, Santa Cruz, CA, USA, 1999.

Holzner:19xx:WMVb

[Holxxa] Steve Holzner. Werken met Visual J++. Sybex Nederland, ????, 19xx. ISBN 90-419-0119-1. ???? pp. LCCN ???? fl.

89. URL http://www.sybex.
nl/catalogus/programmeren/
visualj.htm.

Holzner:19xx:ADLa

[Holxxb] Steven Holzner. Al doende leert men: Java 1.1. Sybex Nederland, ????, 19xx. ISBN 90-419-0148-5. ???? pp. LCCN ???? fl.

69. URL http://www.sybex.
nl/catalogus/programmeren/
aldjava.htm.

Holzner:19xx:ADLb

[Holxxc] Steven Holzner. Al doende leert men: Visual J++ 1.1. Sybex Nederland, ????, 19xx. ISBN 90-419-0158-2. ???? pp. LCCN ???? fl. 49. URL http:

//www.sybex.nl/catalogus/
programmeren/aldvj.htm.

Holzner:19xx:MJW

[Holxxd] Steven Holzner. Il Manuale di Java Workshop. Jackson Libri, via XXV Aprile, 39 - 20091 Bresso (MI), Italy, 19xx. ISBN ???? ???? pp. LCCN ???? 55,000 L. URL http:/

/www.jacksonlibri.it/. Ital-

ian translation of [Hol96c].

Holzner:19xx:WMVa

[Holxxe] Steven Holzner. Werken met Visual J++. Sybex Nederland, ????, 19xx. ISBN 90-419-0019-1. LCCN ???? fl. 89.

Hopwood:1997:CBJ

[Hop97] David Hopwood. A comparison between Java and ActiveX security. Network Security, 1997(12):15–20, December 1997. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/S1353485897885522.

Hoque:1997:PJP

[Hoq97] Reaz Hoque. Practical Javascript Programming. M&T Books, M&T Publishing, Inc., 501 Galveston Drive, Redwood City, CA 94063, USA, 1997. ISBN 1-55851-513-5. xiii + 496 pp. LCCN QA76.73.J38H67 1997. US$29.95.

Hoque:1998:CJI

[Hoq98] Reaz Hoque. Connecting JavaBeans with InfoBus. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 1998. ISBN 0-471-29652-X. 462 pp. LCCN QA76.76.A65H67 1998. US$49.99.

Hoque:19xx:JBB

[Hoqxx] Reaz Hoque. JavaBeans Black Book. Coriolis Group Books, Scottsdale, AZ, USA, 19xx. ISBN 1-57610-163-0. ???? pp. LCCN ???? US$50.

Horowitz:1996:NP

[Hor96a] Bradley Horowitz. New products: FreeD links photographs and 3D; server views images on the Web; Mercury Mail delivers; Web tool provides database functions; building Java-based tables and forms; viewing interactive maps on the net; managing multimedia on demand; Internet video solution offered. IEEE MultiMedia, 3(4):89–90, Winter 1996. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://dlib.computer.org/
mu/books/mu1996/pdf/u4089.
pdf.

Horowitz:1996:NPd

[Hor96b] Bradley Horowitz. New products: FreeD links photographs and 3D; server views images on the Web; Mercury Mail delivers; Web tool provides database functions; building Java-based tables and forms; viewing interactive maps on the net; managing multimedia on demand; Internet video solution offered. IEEE MultiMedia, 3(4):89–90, Winter 1996. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://dlib.computer.org/
mu/books/mu1996/pdf/u4089.
pdf.

Horstmann:1996:JPS

[Hor96c] C. S. Horstmann. Java programming with Sun’s Java development kit. C++ Report, 8 (3):63–70, March 1996. CODEN CRPTE7. ISSN 1040-6042.

Horstmann:1997:POO

[Hor97a] Cay S. Horstmann. Practical Object-Oriented Development in C++ and Java. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, April 1997. ISBN 0-471-14767-2. xiii + 562 pp. LCCN QA76.64.H675 1997. US$29.95.

URL http://www.wiley.com/
compbooks/catalog/14767-2.
htm.

Horton:1997:BJ

[Hor97b] Ivor Horton. Beginning Java.

Wrox Press, Chicago, IL, USA, 1997. ISBN 1-86100-027-8. 600 pp. LCCN ???? US$35.00.

URL http://www.wrox.com/
scripts/bookdetail.idc?Code=
0278.

Horstmann:1998:CCJ

[Hor98a] Cay S. Horstmann. Computing Concepts With Java Essentials. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 1998. ISBN 0-471-17223-5 (softcover). ???? pp. LCCN QA76.H665 1998. US$65.95.

URL http://www.wiley.com/
college/horstmannjava/.

Horwood:1998:PDP

[Hor98b] Peter J. Horwood. PowerJ Developer’s Professional Reference. McGraw-Hill, New York, NY, USA, May 1998. ISBN 0-07-913660-5. LCCN ???? US$50. URL http:

//mcgraw-hill.inforonics.
com/cgi/getarec?mgh41686%
new.

Hosch:1996:JST

[Hos96a] F. Hosch. Java as a first language: an evaluation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(3):45–50, September 1996. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hosch:1996:JFL

[Hos96b] Frederick Hosch. Java as a first language: an evaluation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28(3): 45–??, September 1, 1996. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Houlding:1998:CBF

[Hou98] David Houlding. A CORBA bean framework. Dr. Dobb’s Journal of Software Tools, 23 (11):34, 36, 38–40, November 1998. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

                                        ddj.com/ftp/1998/1998_11/
                                        CORBABeans.exe;   http://
                                        www.ddj.com/ftp/1998/1998_
                                        11/CORBEAN.TXT.

Howlett:1996:SWP

[How96] D. Howlett. Sophisticated Web page creation [HoTMetaL Pro 3.0]. PC User, ????(287):47–??, ???? 1996. CODEN PCUSDW. ISSN 0263-5720.

Hoyle:1996:SSW

[Hoy96] L. Hoyle. SAS software and the WWW — what next? In Anonymous [Ano96-160], pages 2–?? ISBN ???? LCCN ???? Two volumes.

Heller:1998:JDH

[HR98] Philip Heller and Simon Roberts. Java 1.2 Developer’s Handbook. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 1998. ISBN 0-7821-2179-9. ???? pp. LCCN ???? US$60.

Heller:1999:JDH

[HR99] Philip Heller and Simon Roberts.

Java 2 Developer’s Handbook. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 1999. ISBN ???? 1010 pp. LCCN ???? US$60.

Heller:1997:JDHa

[HRSM97] Philip Heller, Simon Roberts, Peter Seymour, and Tom McGinn. Java 1.1 Developer’s Handbook. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, May 1997. ISBN 0-7821-1919-0. xxvii + 1172 pp. LCCN QA76.73.J38 J359 1997. US$59.99. URL http://www.

sybex.com/cgi-bin/bookpg.
pl?1919back.html.

Hall:1996:WVJ

[HRW^{+}96] S. E. Hall, M. K. Ramamurthy, R. B. Wilhelmson, J. Plutchak, D. Wojtowicz, and M. Sridhar. The Weather Visualizer: a Java tool for interactive learning. In Stein [Ste96b], pages 4– ?? ISBN 0-7803-3068-4. LCCN QE33.2.R4 I57 1996. Four volumes. IEEE catalog number 96CH35875.

Hardwick:1996:JIL

[HS96] Jonathan C. Hardwick and Jay Sipelstein. Java as an intermediate language. Technical Report CMU-CS-96-161, School of Computer Science, Carnegie Mellon University, Pittsburgh, PA, USA, August 12, 1996. 17

                                        pp. URL http://www.cs.cmu.
                                        edu/~scandal/html-papers/

javanesl/. Describes the use of Java as the output language of a compiler for NESL, a high-level parallel programming language.

Hoque:1998:PWC

[HS98a] Reaz Hoque and Tarun Sharma. Programming Web Components: JavaBeans & ActiveX. McGraw-Hill, New York, NY, USA, November 1998. ISBN 0-07-912316-3. ???? pp. LCCN QA76.625.H674 1998. US$50. URL http:

                                        //mcgraw-hill.inforonics.
                                        com/cgi/getarec?mgh41710%
                                        comp.

Horton:1998:CDC

[HS98b] J. Horton and J. Seberry. Covert distributed computing using Java through Web spoofing. Lecture Notes in Computer Science, 1438:48–??, 1998. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hume:1998:PCJ

[HS98c] J. N. Patterson Hume and Christine Stephenson. Programming Concepts in Java. Holt Software Assoc., ????, January 1998. ISBN 0-921598-28-9. LCCN ???? US$50. URL http:

                                        //www.holtsoft.com/books/
                                        java_concepts.html.

Hagiya:1998:NMD

[HT98] M. Hagiya and A. Tozawa. On a new method for dataflow analysis of Java Virtual Machine subroutines. Lecture Notes in Computer Science, 1503:17–32, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Huber:1997:JR

[Hub97] Gerald Huber. JAVA — Die Referenz. Verlag Heinz Heise, Helsdorfer Straße 7, D-30625, Hannover, Germany, 1997. ISBN 3-88229-078-1. 490 pp. LCCN ???? 89,90 DM. URL

http://www.emedia.de/bin/
bookshop?show=4551&id=;
http://www.heise.de/buch/.

Includes CD-ROM.

Hughes:1996:PPJ

[Hug96] J. Hughes. Programming with PARTS for Java. Web Techniques, 1(8):71–73, November 1996. CODEN WETEFA. ISSN 1086-556X.

Hugunin:1997:PJB

[Hug97] Jim Hugunin. Python and Java: The best of both worlds. Discusses an implementation of the Python scripting language in Java. The article and the code are available electronically., 1997. URL http://www.

python.org/jpython/.

Hummel:1996:HJC

[Hum96] Robert L. Hummel. How Java can pay the rent. BYTE Magazine, 21(6):40, 42, June 1996.

CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hume:1998:BRS

[Hum98] Andrew Hume. Book review: Sun Performance and Tuning: Java and the Internet. ;login: the USENIX Association newsletter, 23(6):??, October 1998. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

usenix.org/publications/login/
1998-10/sun.html.

Hunt:1997:JOO

[Hun97] John Hunt. Java and Object Orientation: An Introduction (Applied Computing). Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. ISBN 3-540-76201-9. xxii + 473 pp. LCCN QA76.73.J38H86 1998. US$29.

Hunt:1998:EJF

[Hun98] John Hunt. Essentials JavaBeans Fast. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. ISBN 1-85233-032-5. xiv + 174 pp. LCCN QA76.73.J38 H85 1998. US$29.95.

Hunt:1999:JPI

[Hun99] J. Hunt. Java for Practitioners: An Introduction and Reference to Java and Object Orientation. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. ISBN 1-85233-093-7. xxv + 668 pp. LCCN QA76.73.J38 H865 1999. US$64.00.

Hung:19xx:JA

[Hunxx] Whang Yue Hung. Java Applets. DrMaster Consulting Co., Ltd., ??, Taiwan, 19xx. ISBN 957-9625-08-5. ???? pp. LCCN ???? 450 NT$. URL http://

www.drmaster.com/.

Husain:1996:EJJ

[Hus96] K. Husain. Extending JavaScript with Java. Web Techniques, 1 (7):58–??, 60–62, October 1996. CODEN WETEFA. ISSN 1086-556X.

Husain:1997:ADR

[Hus97] Kamran Husain. ActiveX Developer’s Resource: Server-Side Programming. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, July 1997. ISBN 0-13-270786-1. xxi + 457 pp. LCCN QA76.76.D47H88 1997. US$49.95.

Haridi:1999:ELV

[HVB^{+}99] Seif Haridi, Peter Van Roy, Per Brand, Michael Mehl, Ralf Scheidhauer, and Gert Smolka. Efficient logic variables for distributed computing. ACM Transactions on Programming Languages and Systems, 21(3):569–626, May 1999. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). URL http://

www.acm.org/pubs/articles/
journals/toplas/1999-21-3/
p569-haridi/p569-haridi.pdf;
http://www.acm.org/pubs/
citations/journals/toplas/
1999-21-3/p569-haridi/.

Harmon:1997:UUD

[HW97] Paul Harmon and Mark Watson. Understanding UML: the developer’s guide: with a Webbased application in Java. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, November 1997. ISBN 1-55860-465-

0. xvi + 367 pp. LCCN QA76.9.O35H36 1997. US$30.

Hirano:1998:PEPa

[HYI98a] Satoshi Hirano, Yoshiji Yasu, and Hirotaka Igarashi. Performance evaluation of popular distributed object technologies for Java. In ACM [ACM98a], page ?? CODEN CPEXEI. ISSN 1040-3108. LCCN ????

URL http://www.cs.ucsb.
edu/conferences/java98/papers/
dots.pdf;   http://www.cs.
ucsb.edu/conferences/java98/
papers/dots.ps. Also pub-

lished as Concurrency: Practice and Experience, 10(11– 13), September 1998, CODEN CPEXEI, ISSN 1040-3108.

Hirano:1998:PEPb

[HYI98b] Satoshi Hirano, Yoshiji Yasu, and Hirotaka Igarashi. Performance evaluation of popular distributed object technologies for Java. Concurrency: Practice and Experience, 10(11–13):927–940, September 1998. CODEN CPEXEI. ISSN 1040-3108. URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=10050399;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?

ID=10050399&PLACEBO=IE.pdf.

Special Issue: Java for Highperformance Network Computing.

Hyman:1997:VJD

[Hym97] Michael I. Hyman. Visual J++ For Dummies. –For dummies. IDG Books, San Mateo, CA, USA, 1997. ISBN 0-7645-0079-1 (paperback). xx + 330 pp. LCCN QA 76.73 J38 H96 1997. US$24.99.

URL http://www.idgbooks.
com/database/book_result.
msql?uid=uidg81&isbn=0-7645-
0079-1.

Islam:1997:FSS

[IAJR97] Nayeem Islam, Rangachari Anand, Trent Jaeger, and Josyula R. Rao. A flexible security system for using Internet content. IEEE Software, 14 (5):52–59, September/October 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/neumann.computer.org/so/
books/so1997/pdf/s5052.pdf;
 http://www.computer.org/
software/so1997/s5052abs.

htm. Discusses the FlexxGuard security system built using the Java Development Kit 1.1.

IBM:1995:ILJ

[IBM95] IBM Corporation. IBM licenses Java technology from Sun Microsystems for use in Internet products. press release, December 6, 1995. URL http://www.

ibm.com/News/javapr.html.

IBM:1997:CJA

[IBM97] IBM. Creating Java Applications Using NetRexx. IBM Technical Support Organization, San Jose, CA, USA, September 1997. xx + 287 pp. URL http://

                                        www.redbooks.ibm.com/pubs/
                                        pdfs/redbooks/sg242216.pdf.

SG24-2216-00.

Ikeda:1998:SAT

[IDK98] Kazuyuki Ikeda, Matsuyuki Doi, and Tomiei Kazama, editors. State-of-the-art technology in anesthesia and intensive care: proceedings of the 18th International Symposium on Computing in Anesthesia and Intensive Care held in Hamamatsu city, Japan on 18– 21 March 1998, volume 1168 of International congress series. Elsevier Science Publishers B.V., Amsterdam, The Netherlands, 1998. ISBN 0-444-50008-1. LCCN WO 200 I61-2 1998; RD80.95 .I68 1998.

IEEE:1996:ISS

[IEE96a] IEEE, editor. 1996 IEEE Symposium on Security and Privacy: May 6–8, 1996, Oakland, California. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. CODEN PSSPEO. ISBN 0-8186-7417-2. ISSN 1063-7109. LCCN QA76.9.A25 S95 1996. IEEE catalog number 96CB35924.

IEEE:1996:IWC

[IEE96b] IEEE, editor. 3rd International Workshop on Community Networking: proceedings: stories behind the picture, May 23–24, 1996, Antwerpen, Belgium. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-7803-3304-7. LCCN TK5103.4.I57 1996. IEEE catalog number 96TH8187.

IEEE:1996:DPC

[IEE96c] IEEE, editor. Digest of Papers. COMPCON ’96. Technologies for the Information Superhighway. Forty-First IEEE Computer Society International Conference: February 25–28, 1996 Santa Clara, California. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. CODEN DCSIDU. ISBN 0-8186-7414-8. ISSN 1063-6390. LCCN QA76 .I49 1996. IEEE Cat. No. 96CB35911.

IEEE:1996:EPP

[IEE96d] IEEE, editor. Electro ’96: professional program proceedings, April 30–May 2, 1996, Garden State Convention Center, Somerset, New Jersey. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-7803-3272-5 (casebound), 0-7803-3271-7, 0-7803-3273-3. LCCN TK7801.E45 1996. IEEE catalog number 96CB35926.

IEEE:1996:HCV

[IEE96e] IEEE, editor. Hot chips VIII: symposium record: Stanford University, Stanford, California, August 18–20, 1996. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN ???? LCCN ????

IEEE:1996:IIJ

[IEE96f] IEEE, editor. IEEE International Joint Symposia on Intelligence and Systems: November 4–5, 1996, Rockville, Maryland. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-8186-7728-7. LCCN QA76.9 .I4 1996. IEEE Computer Society Press Order Number PR07728. IEEE Order Plan Catalog Number 96TB100091.

IEEE:1996:ITA

[IEE96g] IEEE, editor. IEEE Technical Applications Conference. Northcon/96. Conference Record. Seattle, WA, USA, 4–6 November 1996. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-7803-3277-6 (softbound), 0-7803-3278-4 (casebound). LCCN TK7868.D5 I34 1996. IEEE catalog number 96CH35928.

IEEE:1996:PAC

[IEE96h]  IEEE, editor.  Proceedings.

12th Annual Computer Security Applications Conference.

San Diego, CA, USA. 9–13 December 1996. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-8186-7606-X. LCCN QA76.9.A25 C6375 1996. IEEE Computer Society Press Order Number PRO7606.

IEEE:1996:PEI

[IEE96i]  IEEE, editor.  Proceedings.

Eighth IEEE International Conference on Tools with Artificial Intelligence. November 16–19, 1996, Toulouse, France. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. CODEN PCTIFX. ISBN 0-8186-7686-8. ISSN 1063-6730. LCCN Q 334 I564 1996.

IEEE:1996:PIW

[IEE96j] IEEE, editor. Proceedings International Workshop on Multimedia Software Development: March 25–26, 1996, Berlin, Germany. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, March 1996. ISBN 0-8186-7511-X. LCCN QA76.575 .M84 1996.

IEEE:1996:PICb

[IEE96k] IEEE, editor. Proceedings of the 18th International Conference on Software Engineering: March 25–29, 1996, Berlin, Germany. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-8186-7247-1. LCCN QA 76.758 I5725 1996. IEEE Cat. No.96CB35918.

IEEE:1996:PAI

[IEE96l] IEEE, editor. Proceedings of the 29th annual IEEE/ACM International Symposium on Microarchitecture, December 2– 4, 1996, Paris, France. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-8186-7641-8. LCCN QA76.9 .A73 I585 1996.

IEEE:1996:PWE

[IEE96m] IEEE, editor. Proceedings of the 5th Workshops on Enabling Technologies: Infrastructure for Collaborative Enterprises (WET ICE ’96): June 19–21, 1996, Stanford, California. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-8186-7446-6. LCCN TS155.6 .W56 1996. IEEE Order Plan catalog number 96TB100058. IEEE Computer Society Press order number PR07445.

IEEE:1996:PFIa

[IEE96n] IEEE, editor. Proceedings of the Fifth IEEE International Symposium on High Performance Distributed Computing, August 6–9, 1996, Syracuse, New York. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-8186-7582-9. LCCN QA76.88.I52 1996.

IEEE order plan catalog number 96TB100069. EEE Computer Society Press order number PR07582.

IEEE:1996:PFIb

[IEE96o] IEEE, editor. Proceedings of the Fourth International Conference on Parallel and Distributed Information Systems: December 18–20, 1996, Miami Beach, Florida. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-8186-7475-X. LCCN QA76.58 .I544 1996. IEEE Computer Society Press order number PR07475. IEEE catalog number 96TB100085.

IEEE:1996:PICa

[IEE96p] IEEE, editor. Proceedings of the IEEE/IAFE 1996 Conference on Computational Intelligence for Financial Engineering (CIFEr): March 24–26, 1996, Crown Plaza Manhattan, New York City. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-7803-3236-9 (softbound), 0-7803-3237-7 (microfiche). LCCN HG176.7.I33 1996. IEEE Cat. No. 96TH8177.

IEEE:1996:PICc

[IEE96q] IEEE, editor. Proceedings of the International Conference on Multimedia Computing and Systems: June 17–23, 1996, Hiroshima, Japan. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-8186-7436-9. LCCN QA76.575 .I623 1996. IEEE Computer Society Press order number PR07436. IEEE Catalog No. 96TB100057.

IEEE:1996:PFC

[IEE96r]  IEEE, editor.  Proceedings:

the First Conference on Emerging Technologies and Applications in Communications, Oregon Convention Center, May 7–10, 1996, Portland, Oregon. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-8186-7585-3. LCCN TK5101.A1C677 1996. IEEE catalog number 96TB100035.

IEEE:1996:PTA

[IEE96s] IEEE, editor. Proceedings. The Twentieth Annual International Computer Software and Applications Conference (COMPSAC ’96). IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-8186-7579-9. LCCN QA76.6 .C6295 1996. IEEE Catalog number 96CB35986.

IEEE:1997:HCI

[IEE97] IEEE, editor. Hot Chips IX: Stanford University, Stanford, California, August 24–26, 1997. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1997. ISBN ???? LCCN ????

IEEE:1998:IRI

[IEE98a] IEEE, editor. 1998 IEEE/ RSJ International Conference on Intelligent Robots and Systems: proceedings: innovatiions in theory, practice, and applications: October 13–17, 1998, Victoria Conference Centre, Victoria, BC, Canada, volume 3. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1998. ISBN 0-7803-4465-0 (softbound), 0-7803-4466-9 (casebound), 0-7803-4467-7 (microfiche). LCCN TJ210.3.I443 1998. Three volumes. IEEE Catalog Number 98CH36190.

IEEE:1998:CPT

[IEE98b] IEEE, editor. COMPSAC ’98:

proceedings [of the] Twenty-Second Annual International Computer Software & Applications Conference: held August 19–21, 1998, Vienna, Austria, COMPSAC -NEW YORK-1998. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1998. ISBN 0-8186-8585-9 (softbound), 0-7803-5168-1 (casebound), 0-8186-8587-5 (microfiche). ISSN 0730-3157. LCCN QA76.6 .I546 1998.

IEEE:1998:MME

[IEE98c]  IEEE, editor.  MELECON

’98, 9th Mediterranean Electrotechnical Conference: proceedings, May 18–20. 1998, Tel-Aviv, Israel, volume 1. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1998. ISBN 0-7803-3879-0 (softbound), 0-7803-3880-4 (casebound), 0-7803-3881-2 (microfiche). LCCN TK5 .M43 1998; TK5101.A1 M38 1998. Two volumes. IEEE catalog number 98CH36056.

IEEE:1999:HCS

[IEE99a] IEEE, editor. Hot Chips 11: Stanford University, Stanford, California, August 15–17, 1999. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1999. ISBN ???? LCCN ????

                                        URL http://www.hotchips.
                                        org/hotc11_index.html.

IEEE:1999:SWH

[IEE99b] IEEE, editor. The Seventh Workshop on Hot Topics in Operating Systems: [HotOS-VII]: 29–30 March 1999, Rio Rico, Arizona. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1999. ISBN 0-7695-0237-7, 0-7695-0238-5 (case), 0-7695-0239-3 (microfiche). LCCN QA76.76.O63 W6666 1999. IEEE Computer Society Press order number PR00237.

Ince:1997:PIJ

[IF97] Darrel Ince and Adam Freeman. Programming the Internet With Java. Addison-Wesley, Reading, MA, USA, 1997. ISBN 0-201-17549-5 (softcover). xiv + 440 pp. LCCN QA76.73.J38I52 1997. US$37.95. URL http:

//cseng.aw.com/bookdetail.
qry?ISBN=0-201-17549-5.

Burge:1999:JJB

[IG99] Legand L. Burge III and K. M.

George. JMAS: a Java-based mobile actor system for distributed parallel computation. In USENIX [USE99c], page ?? ISBN ???? LCCN QA76.64 .U84 1999. URL http://www.

usenix.org/publications/library/
proceedings/coots99/burge.
html.

Iliadis:1998:TSD

[IGO98] J. Iliadis, S. Gritzalis, and V. Oikonomou. Towards secure downloadable executable content: The Java paradigm. Lecture Notes in Computer Science, 1516:117–??, 1998. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ito:1996:ZWJ

[IK96] Tomohachiro Ito and Mitsuo Kawano. Zukai 60-fun de wakaru Java: Intanetto hyojun gengo ga kano ni suru 500doru pasokon no shogeki. PHP Kenkyujo, Tokyo, Japan, dai 1han edition, 1996. ISBN 4-569-55133-5. 206 pp. LCCN ????

iMatrix:1996:LTP

[iMa96]  iMatrix, ????  Libero

— the thinking programmer’s tool, ???? edition, ????

1996.  URL http://www.
imatrix.com/html/libero/ (

                                        no   longer accessible). Com-

piles a finite-state machine into one of several languages, including Java.

Imasaki:1997:JJP

[Ima97] K. Imasaki. JAPE: The Java parallel environment. Lecture Notes in Computer Science, 1343:137–??, 1997. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

InfoWorld:1995:NOA

[Inf95a] InfoWorld staff. News: Oracle to announce Java support, progress to release Ole controls. InfoWorld, page ??, October 21, 1995. CODEN INWODU. ISSN 0199-6649. URL http://www.

                                        infoworld.com/.

InfoWorld:1995:NP

[Inf95b] InfoWorld staff. News: Pipeline. InfoWorld, page ??, November 22, 1995. CODEN INWODU. ISSN 0199-6649. URL http://

                                        www.infoworld.com/.

InfoWorld:1995:NPJ

[Inf95c] InfoWorld staff. News: Pipeline — Java due in 1996. InfoWorld, page ??, October 03, 1995. CODEN INWODU. ISSN 0199-6649. URL http://www.

                                        infoworld.com/.

InfoWorld:1995:NSS

[Inf95d] InfoWorld staff. News: Sun set to release DOE on Solaris. InfoWorld, page ??, September 16, 1995. CODEN INWODU. ISSN 0199-6649. URL http://

                                        www.infoworld.com/.

InfoWorld:1995:NSOb

[Inf95e] InfoWorld staff. News: SunService offers Java programming courses. InfoWorld, page ??, December 20, 1995. CODEN INWODU. ISSN 0199-6649.

URL http://www.infoworld.
com/.

InfoWorld:1995:NTS

[Inf95f] InfoWorld staff. News: Tool set plugs Java apps into CORBA. InfoWorld, page ??, December 12, 1995. CODEN INWODU. ISSN 0199-6649. URL http://

www.infoworld.com/.

InfoMagic:19xx:JDK

[Infxxa] InfoMagic. Java Development Kit For Windows NT/95. Specialized Systems Consultants, P.O. Box 55549, Seattle, WA 98155, 19xx. ISBN ???? ???? pp. LCCN ???? US$25.

InfoMagic:19xx:JI

[Infxxb] InfoMagic. Java from Info-Magic. Universal CD-ROM, ????, 19xx. ISBN ???? ???? pp. LCCN ???? US$25. URL http:

//www.bigmall.com/PJ0279.
htm;   http://www.bigmall.
com/program.html.

Intermetrics:1996:TCA

[Int96a] Intermetrics. Tool converts Ada 95 source code to Java bytecode. Java Report Online,

1996. URL http://www.sigs.
com/publications/docs/jro/
twij96/twij961216.html#TOOL.

Intermetrics Inc. debuted AppletMagic, a tool that converts Ada 95 source code to Java bytecode for execution by any Java-capable Web browser. AppletMagic simplifies the development of complex, highreliability applets and can be used as a supplement or an alternative to the Java language. Ada provides compile-time advantages such as enumeration types and generic templates, as well as in, in-out, and out parameter modes. The Java execution technology contributes runtime flexibility through automatic garbage collection, dynamic linking, and platform independence.

Advisor:1996:IJA

[Int96b] Internet and Java advisor, 1996. ISSN 1084-158X. Advisor Publications, Inc., San Diego, CA, USA.

IPD:19xx:JR

[IPDxx] IPD. Java Report. IPD, ????, 19xx. ISBN ???? ???? pp. LCCN ???? 50 FF.

Igarashi:1999:FJM

[IPW99] Atshushi Igarashi, Benjamin Pierce, and Philip Wadler. Featherweight Java: a minimal core calculus for Java and GJ. ACM SIGPLAN Notices, 34(10):132–146, October 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

URL http://www.acm.org/
pubs/citations/proceedings/

oops/320384/p132-igarashi/

.

Ichisugi:1997:EJP

[IR97] Y. Ichisugi and Y. Roudier. The extensible Java preprocessor kit and a tiny data-parallel Java. Lecture Notes in Computer Science, 1343:153–??, 1997. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ioannidou:1999:EUP

[IR99] Andri Ioannidou and Alexander Repenning. End-user programmable simulations. Dr. Dobb’s Journal of Software Tools, 24(8):40, 42–48, August 1999. CODEN DDJOEB. ISSN 1044-789X.

Isaak:1996:CSP

[Isa96] Jim Isaak. A call for standards for portable Web applications. Computer, 29(6): 89–90, June 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). Recommends standardization of Java, JavaScript, VBScript, and Java’s byte codes.

Iskander:1996:TBR

[Isk96] Magdy F. Iskander, editor.

Technology-Based Re-Engineering. Engineering Education. Proceeding of Frontiers in Education FIE’96 26th Annual Conference, November 6–9, 1996, Salt Lake City, Utah. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-7803-3348-9 (softbound), 0-7803-3349-7 (casebound), 0-7803-3350-0 (microfiche), 0-7803-3720-4 (CD-ROM). LCCN T 62 F76 1996. Three volumes. IEEE catalog number 96CH35946.

Jain:1998:DDO

[JA98] Lalit Kumar Jain and Mark Abbott. DOVE: distributed objects based scientific visualization environment. Concurrency: Practice and Experience, 10(11–13):1087–1095, September 1998. CODEN CPEXEI. ISSN 1040-3108. URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=10050410;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=10050410&PLACEBO=IE.pdf.

Special Issue: Java for Highperformance Network Computing.

Jacobs:1996:JBW

[Jac96] Russell Jacobs. Javascript: Blueprints to a world beyond HTML. Prima Publishing, Rocklin, CA, USA, July 1996. ISBN 0-7615-0685-3. xii + 337 pp. LCCN QA76.73.J39K66 1996. US$35.00.

Jagannathan:1996:CWP

[JADR96] V. Jagannathan, G. Almasi, T. Davis, and R. Raman. CORBA-based and Web-based patient record integration. In Waegemann [Wae96], pages 243–247 (vol. 2). ISBN 0-9640667-7-7. LCCN R864.I68 1996. Two volumes.

Jamsa:1996:JNE

[Jam96] Kris A. Jamsa. Java Now:

The Easiest Way to Learn Java in the Least Amount of Time. Jamsa Press, Las Vegas, NV, USA, June 1996. ISBN 1-884133-30-4. 218 pp. LCCN QA76.73.J38J355 1996. US$16.95. URL http://www.

jamsa.com/catalog/javanow/
javanow.htm.

Jamsa:1997:ESJ

[Jam97a] Kris Jamsa. Ed è subito Java!

Tecniche Nuove, ??, Italy, 1997. ISBN ???? 242 pp. LCCN ???? 39000 Lire. URL

http://www.tecnet.it/cgi-
shl/dbml.exe?Action=3DQuery&
Template=3D/LibriQueryDueLibro.
dbm&ISBN=3D03863;   http:/
/www.tecnet.it/libri/VisualizzaLibro.
cfm?ISBN=03863.

Jamsa:1997:JOP

[Jam97b] Kris Jamsa. Java — ohjelmoinnin perusteet. Teknolit, Helsinki, Finland, September 1997. ISBN 952-9823-77-0. 233 pp. LCCN ???? 230 Markka. URL http://

www.jamsa.com/authors/bio.
htm#kjamsa;   http://www.
teknolit.co.jyu.fi/books/
muut/java.htm;   http://www.

teknolit.fi/. Finnish translation of [Jam96].

Jamsa:1997:JA

[Jam97c] Kris Jamsa. Java Ahora! Mc-Graw-Hill Spain, ????, 1997. ISBN 958-600-624-7. ???? pp. LCCN ???? 3365

                                        Ptas.  URL http://www.
                                        mcgraw-hill.es/McGrawHill/
                                        catalogo.nsf/b013777771f7e01bc125647a00607b0e/
                                        94973e13fcf9c5e6c12564d60040555c?
                                        OpenDocument.

Jamsa:19xx:ESJ

[Jamxxa]  Kris Jamsa.  Ed è subito

Java! Tecniche Nuove, ??, Italy, 19xx. ISBN 88-8481-038-6 (??Invalid checksum??). ???? pp. LCCN ???? 39000 L. URL

                                        http://www.tecnet.it/cgi-
                                        shl/dbml.exe?Action=Query&
                                        Template=/LibriQueryDueLibro.
                                        dbm&ISBN=03863.

Jamsa:19xx:JO

[Jamxxb] Kris Jamsa. Java-ohjelmointi. Teknolit, ????, 19xx. ISBN 952-9823-77-0 (??invalid ISBN??). ???? pp. LCCN ???? ????

Jander:1996:WRW

[Jan96] M. Jander. Welcome to the revolution [WWW-based network management]. Data Communications International, 25(16): 38–42, 44, 46, 48, 50, 52– 53, ???? 21, 1996. CODEN DCINEU. ISSN 0363-6399.

Jang:1997:FJP

[Jan97] Kyuo Jang. The Foundation of the Java Programming. Seongandang, ??, Korea, 1997. ISBN 89-315-4381-6 (??Invalid checksum??). ???? pp. LCCN ???? 3,500 won. URL ????.

Jannak:1999:JGR

[Jan99] Torpum Jannak. Java 2 graphics rendering. Dr. Dobb’s Journal of Software Tools, 24(9):

19–20, 22, 24, 26–28, September 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ftp/1999/1999_
09/java2gra.txt;   http://
www.ddj.com/ftp/1999/1999_
09/java2gra.zip.

Jardin:1997:JEC

[Jar97] Cary A. Jardin. Java Electronic Commerce Sourcebook: All the Software and Expert Advice You Need to Open Your Own Virtual Store. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, February 1, 1997. ISBN 0-471-17611-7. 473 pp. LCCN 96-049117. US$39.95. URL http:

//www.wiley.com/compbooks/
catalog/17611-7.htm.

Jardin:1998:SJC

[Jar98] Cary A. Jardin. Sun Java 1.1 Certification Exam Guide Book. Que Corporation, Indianapolis, IN, USA, March 1998. ISBN 0-7897-1390-X. xiii + 605 pp. LCCN QA76.3.J37 1998. 138,00 DM, US$70.00. URL http://

merchant.superlibrary.com:
8000/catalog/mcp/PRODUCT/
PAGE/07897/bud/078971390X.
html;   http://www.mut.ch/
profit.htx?&x_pk=kd84mf7ms9dm.

Includes CD-ROM.

Jarol:19xx:HPF

[Jarxx] Scott Jarol. Hyperwire Programming FrontRunner. Coriolis Group Books, Scottsdale, AZ, USA, 19xx. ISBN 1-57610-066-9. LCCN ???? US$30.

Jagannathan:1996:CIU

[JAS96a] V. Jagannathan, G. Almasi, and A. Suvaiala. Collaborative infrastructures using the WWW and CORBA-based environments. In IEEE [IEE96m], page ?? ISBN 0-8186-7446-6. LCCN TS155.6 .W56 1996. IEEE Order Plan catalog number 96TB100058. IEEE Computer Society Press order number PR07445.

Jasper:1996:JK

[Jas96b]  Dirk Jasper.  Java Kan-

nich. DATA-Becker, ????, 1996. ISBN 3-8158-1300-X. 254 pp. LCCN ???? 30 DM. URL http:

                                        //www.data-becker.de/.

Jasper:19xx:JM

[Jasxx] Dirk Jasper. Java Makkie. Easy Computing SA-NV, Chaussée d’Alsemberg 610 - Alsembergsesteenweg 610, 1180 Brussels, Belgium; Korte Kleverlaan, 34Bf, 2061 EE Bloemendaal, The Netherlands, 19xx. ISBN 90-5167-154-7. ???? pp. LCCN ???? fl. 30. URL http://www.

                                        easycomputing.com/.

Advisor:1996:JA

[Jav96a] Java advisor, 1996. ISSN 1087-0776. Advisor Publications, San Diego, CA, USA.

Sys-Con:1996:JDJ

[JAV96b] JAVA Developer’s Journal, 1996. ISSN 1087-6944. Sys-Con Publications, Jersey City, NJ, USA.

SIGS:1996:JRS

[Jav96c] Java Report: The Source for Java Development, 1996. ISSN 1086-4660. SIGS Books, 71 W 23rd St., New York, NY 10010, USA.

IDG:1996:JWI

[Jav96d] Java World: IDG’s magazine for the Java community, 1996. ISSN ???? IDG Books, San Mateo, CA, USA.

JGF:1998:MJW

[Jav98] Java Grande Forum. Making Java work for high-end computing. Technical report TR-1, ????, ????, 1998. URL

http://www.javagrande.org/
reports.htm.

Jaworski:1996:JDG

[Jaw96] Jamie Jaworski. Java Developer’s Guide. Sams.net, Indianapolis, IN 46268, USA, July 1, 1996. ISBN 1-57521-069-X. xxvi + 729 pp. LCCN QA76.73.J38J39 1996. US$49.99. URL http://www.

mcp.com/288190891680543/cgi-
bin/bag?isbn=1-57521-069-
X&last=/bookstore.

Jaworski:1997:MJ

[Jaw97a]  James Jaworski.  Master-

ing JavaScript. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 1997. ISBN 0-7821-2014-8. xxxv + 1085 pp. LCCN QA76.73.J38J393 1997. US$39.99.

Jaworski:1997:JGA

[Jaw97b] Jamie Jaworski. Java — Guida alla programmazione. Tecniche Nuove, ??, Italy, 1997. ISBN ???? 704 pp. LCCN ???? 89000 Lire. URL

http://www.tecnet.it/cgi-
shl/dbml.exe?Action=3DQuery&
Template=3D/LibriQueryDueLibro.
dbm&ISBN=3D03804;   http:/
/www.tecnet.it/libri/VisualizzaLibro.
cfm?ISBN=03804. Includes CD-

ROM.

Jaworski:1997:JDG

[Jaw97c] Jamie Jaworski. Java 1.1 Developers Guide. Sams.net, Indianapolis, IN 46268, USA, second edition, 1997. ISBN 1-57521-283-8. xxxvii + 1100 pp. LCCN QA76.73.J38J389 1997. 99,00 DM; US$50. URL http:/

/merchant.superlibrary.com:
8000/catalog/hg/PRODUCT/PAGE/
15752/bud/1575212838.html;
 http://www.mut.ch/profit.
htx?&x_pk=kd84mf7ms9dm.

Jaworski:1998:JU

[Jaw98] Jamie Jaworski. Java 1.2 Unleashed. Sams.net, Indianapolis, IN 46268, USA, fourth edition, May 1998. ISBN 1-57521-389-3. LCCN ???? US$50.

Jaworski:1999:JSH

[Jaw99] Jamie Jaworski. Java Security Handbook. Howard W. Sams, Indianapolis, IN 46268, USA, 1999. ISBN 0-672-31602-1. ???? pp. LCCN ???? US$50. Edited by Tim Ryan et al.

Jaworski:19xx:JGA

[Jawxxa] Jamie Jaworski. Java — Guida alla programmazione. Tecniche Nuove, ??, Italy, 19xx. ISBN ???? LCCN ???? 89000 L. URL

http://www.tecnet.it/cgi-
shl/dbml.exe?Action=Query&
Template=/LibriQueryDueLibro.
dbm&ISBN=03804.

Jaworski:19xx:JGD

[Jawxxb] Jamie Jaworski. Java Guia Desarrollo. Prentice Hall HispanoAmericana, Mexico, DF, Mexico / Nueva York, NY, USA, 19xx. ISBN 84-89660-61-1 (??invalid ISBN??). ???? pp. LCCN ???? 6990

Ptas. URL http://prentice.
com.mx/;   http://www.mcl.
com.mx/sotano/prentice.fm$
Retrive?RecID=2702&html=HTMLRecord1.

Spanish translation of [Jaw97a].

Joshi:1997:CGV

[JC97] Daniel Joshi and Ramesh Chandak. The comprehensive guide to Visual J++: the ultimate reference for Java, Visual J++ & Active X: Windows 95 & Windows NT 4. Ventana Press, Chapel Hill, NC, USA, 1997. ISBN 1-56604-533-9. xxiv + 750 pp. LCCN QA76.73.J38 J67 1997. US$50. URL http:

//www.joshipublishing.com/
visj.htm.

Jamsa:19xx:IPU

[JCxx] Kris Jamsa and Ken Cope.

Internet Programming Under Windows. Piter Publishing Company, St. Petersburg, Russia, 19xx. ISBN ???? ???? pp. LCCN ???? URL http://

www.jamsa.com/authors/bio.
htm#kjamsa;   http://www.
piter-press.ru/;   http://
www.piter-press.ru/koi/ourbooks/
comp/progr/pdisw_i.html.

Judd:1998:DDOa

[JCS98a] Glenn Judd, Mark Clement, and Quinn Snell. DOGMA: Distributed Object Group Management Architecture. In ACM [ACM98a], page ?? CODEN CPEXEI. ISSN 1040-3108. LCCN ????

URL http://www.cs.ucsb.
edu/conferences/java98/papers/
dogma.pdf;   http://www.cs.
ucsb.edu/conferences/java98/
papers/dogma.ps. Also pub-

lished as Concurrency: Practice and Experience, 10(11– 13), September 1998, CODEN CPEXEI, ISSN 1040-3108.

Judd:1998:DDOb

[JCS98b] Glenn Judd, Mark Clement, and Quinn Snell. DOGMA: distributed object group metacomputing architecture. Concurrency: Practice and Experience, 10(11–13):977–983, September 1998. CODEN CPEXEI. ISSN 1040-3108. URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=10050409;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=10050409&PLACEBO=IE.pdf.

Special Issue: Java for Highperformance Network Computing.

Sun:1999:JCV

[JCV99] SUN Microsystems, Inc. Java Card 2.1 Virtual Machine Specification, March 3, 1999. Final Revision 1.0.

Jardin:1997:SVC

[JD97] Cary A. Jardin and Pam Dixon.

Symantec Visual Café Sourcebook. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, June 1997. ISBN 0-471-17804-7. xv + 395 pp. LCCN QA76.76.D47J366 1997. US$39.99. URL http:

//www.wiley.com/compbooks/
catalog/17804-7.htm.

JDK:1998:JDKa

[JDK98a] Java Development Kit 1.1.6.

Sun Microsystems, 1998. URL

http://java.sun.com/products/
jdk/1.1/.

JDK:1998:JDKb

[JDK98b] Java Development Kit 1.2 Beta 3. Sun Microsystems, 1998.

URL http://java.sun.com/
products/jdk/1.2/.

Jenkins:1998:ADT

[Jen98] Michael S. Jenkins. Abstract Data Types in Java. McGraw-Hill, New York, NY, USA, 1998. ISBN 0-07-913270-7. 388 pp. LCCN QA76.73.J38J45 1998. US$39.95. URL http:

//mcgraw-hill.inforonics.
com/cgi/getarec?mgh31409%
comp.

Jepson:1996:PBD

[Jep96] Brian Jepson. Programando Banco de Dados em Java: Domine as Técnicas dos Bancos de Dados da Web. Makron Books do Brasil Editora Ltda., Rua Tabapuã 1348, São Paulo SP - 04533-004 - Brasil, 1996. ISBN 85-346-0759-1. 504 pp. LCCN ???? R$ 65,00. URL http://

safe.tesla.com.br/makron/;
 http://safe.tesla.com.br/
makron/options/livro.asp?
id=7;   http://www.makron.

com/. Portuguese translation of [Jep97].

Jepson:1997:JDP

[Jep97] Brian Jepson. Java Database Programming. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 1997. ISBN 0-471-16518-2. x + 485 pp. LCCN QA76.73.J38J47 1997. US$29.95. URL http:

//www.wiley.com/compbooks/
catalog/16518-2.htm.

Jessup:1996:WSC

[JG96] Elizabeth R. Jessup and Roscoe C. Giles. What should computer scientists teach to physical scientists and engineers? part 2. response to Wilson: Teach computing in context. IEEE Computational Science & Engineering, 3(3):54–62, Fall 1996. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

URL http://www.computer.
org/cse/cs1998/c3054abs.htm.

Proposes that Java be included in a week-long course of “computer science for physical scientists.”.

Jacobs:1996:FHF

[JGKR96] Stephan Jacobs, Michael Gebhardt, Stefanie Kethers, and Wojtek Rzasa. Filling HTML forms simultaneously: CoWebarchitecture and functionality. Computer Networks and ISDN Systems, 28(7–11):1385– 1395, May 1, 1996. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/
browse/browse.cgi?year=1996&
volume=28&issue=7-11&aid=
1607.

Jimenez:1997:PJP

[Jim97] Pedro Manuel Cuenca Jiménez.

Programación en Java para Internet. Guias Practicas. Anaya Multimedia, Madrid, Spain, May 1997. ISBN 84-415-0174-2, 84-7614-958-1. 352 pp. LCCN ???? 1,695 Ptas. (or 3750 Pesatas??). URL http://

www.anaya.es/Catalogo/CGA?
act=3DL&art=3D2321006&ed=
3D2300;   http://www.anaya.
es/Catalogo/CGA?act=L&art=
2321006&ed=2300;   http://
www.anayamultimedia.es/;
http://www.anayamultimedia.
es/cgi-multi/notas_prensa.
pl?2335127&3&0&1.

Jenkins:1998:ACW

[JJBW98] Charlotte Jenkins, Mike Jackson, Peter Burden, and Jon Wallis. Automatic classification of Web resources using Java and Dewey Decimal Classification. Computer Networks and ISDN Systems, 30(1–7): 646–648, April 1, 1998. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

URL http://www.elsevier.
com/cas/tree/store/comnet/
sub/1998/30/1-7/1846.pdf.

Juhne:1998:AJB

[JJG98] Jesper Jühne, Anders T. Jensen, and Kaj Grønbæk. Ariadne: a Java-based guided tour system for the World Wide Web. Computer Networks and ISDN Systems, 30(1–7):131– 139, April 1, 1998. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

URL http://www.elsevier.
com/cas/tree/store/comnet/
sub/1998/30/1-7/1935.pdf.

Jones:1996:GCA

[JL96] Richard Jones and Rafael Lins. Garbage Collection: Algorithms for Automatic Dynamic Memory Management. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 1996. ISBN 0-471-94148-4. xxvi + 377 pp. LCCN QA76.9.G37J66 1996. US$60.00. URL http:

//www.ukc.ac.uk/computer_
science/Html/Jones/gc.html.

Reprinted in 1999 with improved index, and corrected errata.

Jacobs:1999:FTJ

[JLMPH99] Bart Jacobs, Gary T. Leavens, Peter Müller, and Arnd Poetzsch-Heffter. Formal techniques for Java programs. Lecture Notes in Computer Science, 1743:97–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/
service/series/0558/bibs/
1743/17430097.htm;   http:
//link.springer-ny.com/link/
service/series/0558/papers/
1743/17430097.pdf.

Joshi:1996:TYJ

[JLP96] Daniel I. Joshi, Laura Lemay, and Charles L. Perkins. Teach yourself Java in Café in 21 days. Sams.net, Indianapolis, IN 46268, USA, 1996. ISBN 1-57521-157-2. xxiv + 614 pp. LCCN QA76.73.J38J676 1996. US$40. URL http://

merchant.superlibrary.com:
8000/catalog/hg/PRODUCT/PAGE/
15752/bud/1575211572.html;
http://www.joshipublishing.
com/cafe.htm.

Jensen:1998:CGJ

[JLT98] T. Jensen, D. Le Metayer, and T. Thorn. Coarse-grained Java security policies. Lecture Notes in Computer Science, 1543:296– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jamsa:1996:WPV

[JLW96] Kris Jamsa, Suleiman ‘Sam’ Lalani, and Steve Weakley. Web Programming: Visual Basic, TCP/IP, JavaScript, HTML, HTTP, VRML, VB-Script, Java, CGI, J++, C++. Jamsa Press, Las Vegas, NV, USA, 1996. ISBN 1-884133-27-4 (softcover). xxii + 506 pp. LCCN QA76.625 .J36 1996. US$44.95. Includes CD-ROM.

Jackson:1996:JE

[JM96] Jerry R. Jackson and Alan L. McClellan. Java by Example. Java series. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1996. ISBN 0-13-565763-6. xxx + 345 pp. LCCN QA76.76.J38J33 1996. US$34.95. URL http://www.

sun.com/smi/ssoftpress/books/
Jackson/Jackson.html.

Jackson:1997:JE

[JM97a] Jerry Jackson and Alan McClellan. Java by Example. ASCII Corporation, Tokyo, Japan, second edition, 1997. ISBN ???? ???? pp. LCCN

???? URL http://www.ascii.
co.jp/;   http://www.sun.
com/smi/ssoftpress/books/
Jackson/Jackson.html.

Jackson:1997:JPE

[JM97b] Jerry Jackson and Alan McClellan. Java Per Esempi. Mondadori Informatica, ??, Italy, 1997. ISBN ???? ???? pp.

LCCN ???? URL http://www.
mondadori.com/;   http://
www.sun.com/smi/ssoftpress/
books/Jackson/Jackson.html.

Italian translation of [JM97a]. Includes CD-ROM.

Jackson:1997:JBS

[JM97c] Jerry R. Jackson and Alan L.

McClellan. Java an Beispielen — Sprach- und Applet-Struktur. Verlag Heinz Heise, Helsdorfer Straße 7, D-30625, Hannover, Germany, 1997. ISBN 3-88229-085-4. 336 pp. LCCN ???? 84,80 DM. URL

http://www.emedia.de/bin/
bookshop?show=4562&id=;
http://www.heise.de/buch/.

German translation of [JM97a]. Includes CD-ROM.

Jackson:1997:JES

[JM97d] Jerry R. Jackson and Alan L.

McClellan. Java by Example: Signature Edition. Java series. SunSoft Press and Prentice-Hall, 2550 Garcia Avenue, Mountain View, CA 94043-1100, USA and Englewood Cliffs, NJ 07632, USA, second edition, January 1997. ISBN 0-13-272295-X. xl + 386 pp. LCCN QA76.73.J38J33 1997. US$34.95. URL http://www.

prenhall.com/ptrbooks/ptr_
013272295x.html;   http:/
/www.sun.com/books/books/
Jackson/Jackson.html. Avail-

able in German translation [JM97c]. Includes CD-ROM.

Jackson:1999:JE

[JM99] Jerry R. Jackson and Alan L.

McClellan. Java By Example 1.2. SunSoft Press Java series. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1999. ISBN 0-13-079669-7. 500 pp. LCCN ???? US$39.99. URL http:

                                        //www.phptr.com/ptrbooks/
                                        ptr_0130796697.html;   http:
                                        //www.sun.com/books/catalog/
                                        jackson2/index.html.

Jackson:19xx:JEc

[JMxxa] Jerry Jackson and Alan Mc-Clellan. Java by Example. ASCII Corporation, Tokyo, Japan, second edition, 19xx. ISBN 4-7561-1709-0. LCCN

????  ????  URL http:/

                                        /www.ascii.co.jp/pb/book1/
                                        shinkan/detail/1322352.html.

Jackson:19xx:JPE

[JMxxb] Jerry Jackson and Alan McClellan. Java Per Esempi. Mondadori Informatica, ??, Italy, 19xx. LCCN ???? 70000 L.

JMA:1998:JMA

[JMA98] Java Media APIs. Sun Microsystems, 1998. URL http://

                                        java.sun.com/products/java-
                                        media/.

Johnson:1996:LNM

[JMM96] Marc Johnson, Robert Mc-Daniel, and Michael Morrison. Late Night Microsoft Visual J++. Ziff-Davis Press, Emeryville, CA, USA, 1996. ISBN 1-56276-452-7. xxi + 598 pp. LCCN ???? US$50.

                                        URL http://www.mcp.com/
                                        cgi-bin/bag?isbn=1-56276-
                                        452-7&last=/bookstore.

Johnston:1995:CML

[JN95a] Stuart J. Johnston and Kim S.

Nash. Capitulation! Microsoft to license Java; Internet standards war avoided. Computerworld, 29(50):1, 141, December 11, 1995. CODEN CMPWAB. ISSN 0010-4841.

Johnston:1995:NMO

[JN95b] Stuart J. Johnston and Kim S.

Nash. News: Microsoft opts to avoid Internet standards war. Computerworld, page ??, December 08, 1995. CODEN CM-PWAB. ISSN 0010-4841. URL

http://www.computerworld.
com/.

Johnston:1995:NMP

[JN95c] Stuart J. Johnston and Kim S.

Nash. News: Microsoft plans to unveil Internet products. Computerworld, page ??, December 01, 1995. CODEN CMPWAB. ISSN 0010-4841.

Johnson:1998:WBM

[JOA98] C. W. Johnson, G. Oser, and A. J. Abedor. Web browser as medical educator/researcher using HTML and JavaScript. In Chute [Chu98b], pages 1023–?? ISSN 1091-8280. LCCN ????

Jobst:1996:PJ

[Job96] Fritz Jobst. Programmieren in Java. Carl Hanser, München, Germany, September 1996. ISBN 3-446-18789-8. ???? pp. LCCN ???? 42 DM. URL http:

//www.hanser.de/computer/
buecher/jo18789.htm.  In-

cludes CD-ROM.

Joch:1998:IGP

[Joc98] Alan Joch. IDBC’s growing pains: Java database connectivity: Developers need more than promises. BYTE Magazine, 23(5):112M, 112N, 112P, May 1998. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Johnson:1996:JAD

[Joh96a] J. Johnson. Java as an application development language. Object Magazine, 6(4):59–61, June 1996. CODEN OBMAFO. ISSN 1055-3614.

Johnson:1996:JMS

[Joh96b] Marcus Johnson. JavaScript 2.1 Manual of Style. Ziff-Davis Press, Emeryville, CA, USA, 1996. ISBN 1-56276-423-3. xiv + 277 pp. LCCN QA76.73.J39J65 1996. US$24.99. URL http://www.

                                        mcp.com/zdpress.

Johnson:1997:OAR

[Joh97a] Maryfran Johnson. Opinion: Alternate realities: Our frontpage story last week jangled a lot of readers’ nerves. Computerworld, 31(31):32, August 4, 1997. CODEN CMPWAB. ISSN 0010-4841. See [SG97a] for original story, and [SG97b] for further comments.

Johnson:1997:OJB

[Joh97b] Maryfran Johnson. Opinion: Java blend. Computerworld, 31 (15):34, April 14, 1997. CODEN CMPWAB. ISSN 0010-4841.

Johann:1998:JJB

[Joh98] Michael Johann. JBuilder: Java Beans, JDBC, RMI. Addison-Wesley Verlag, Bonn, Germany, March 1998. ISBN 3-8372-1319-8 (??invalid ISBN??). LCCN ???? 80 DM. URL

http://www.addison-wesley.
de/katalog/item.ppml?id=00264.

Johnson:1999:PTC

[Joh99] David M. Johnson. Programmer’s toolchest: Comparing WFC and JFC. Dr. Dobb’s Journal of Software Tools, 24 (2):90, 92–95, February 1999. CODEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/1999/9902/9902toc.
htm;   http://www.ddj.com/
ftp/1999/1999_02/relay.txt;
 http://www.ddj.com/ftp/1999/
1999_02/relay.zip.

Jones:1998:JIA

[Jon98] Carol A. Jones. The Java internationalization API: Global software for the global village. Dr. Dobb’s Journal of Software Tools, 23(1):54, 56–69, 103– 104, January 1998. CODEN DDJOEB. ISSN 1044-789X.

Joshi:1997:JPR

[Jos97a] Daniel Joshi. The Java 1.1 Programmer’s Reference. Ventana Press, Chapel Hill, NC, USA, 1997. ISBN 1-56604-687-4. ???? pp. LCCN QA76.73.J38 J672 1997. US$60. URL http:

//www.joshipublishing.com/
jpref.htm.

Joshi:1997:MJJ

[Jos97b] Daniel I. Joshi. Migrating from Java 1.0 to Java 1.1. Ventana Press, Chapel Hill, NC, USA, June 1997. ISBN 1-56604-686-6 (softcover). ???? pp. LCCN QA76.73.J674 1997. US$35.99. URL http://www.

                                        vmedia.com/bookshop/book.
                                        html?isbn=1566046866.

Joshi:1998:JPR

[Jos98] Daniel Joshi. The Java Programmer’s Reference. Ventana Press, Chapel Hill, NC, USA, February 1998. ISBN 1-56604-857-5. LCCN ???? US$50.

Joyce:1998:PTN

[Joy98] Daniel T. Joyce, editor. Proceedings of the Twenty-ninth SIGCSE Technical Symposium on Computer Science Education: February 25–March 1, 1998, Atlanta, Georgia, volume 30(1) of SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education). ACM Press, New York, NY 10036, USA, 1998. ISBN 0-89791-994-7. ISSN 0097-8418 (print), 2331-3927 (electronic). LCCN QA1 .A86 v.30, no.1.

JPA:1998:JPA

[JPA98] Java products & APIs. Sun Microsystems, 1998. URL http:/

                                        /java.sun.com/products/.

Jacob:1998:LSPa

[JPK98a] Matthias Jacob, Michael Philippsen, and Martin Karrenbach. Largescale parallel geophysical algorithms in Java: a feasibility study. In ACM [ACM98a], page ?? CODEN CPEXEI. ISSN 1040-3108. LCCN ????

URL http://www.cs.ucsb.
edu/conferences/java98/papers/
veltran.pdf;   http://www.
cs.ucsb.edu/conferences/java98/
papers/veltran.ps. Also pub-

lished as Concurrency: Practice and Experience, 10(11– 13), September 1998, CODEN CPEXEI, ISSN 1040-3108.

Jacob:1998:LSPb

[JPK98b] Matthias Jacob, Michael Philippsen, and Martin Karrenbach. Largescale parallel geophysical algorithms in Java: a feasibility study. Concurrency: Practice and Experience, 10 (11–13):1143–1153, September 1998. CODEN CPEXEI. ISSN 1040-3108. URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=10050403;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=10050403&PLACEBO=IE.pdf;
 http://wwwipd.ira.uka.de/
~phlipp/veltran.ps.gz. Spe-

cial Issue: Java for Highperformance Network Computing.

Jaeger:1999:FCD

[JPLI99] Trent Jaeger, Atul Prakash, Jochen Liedtke, and Nayeem Islam. Flexible control of downloaded executable content. ACM Transactions on Information and System Security, 2 (2):177–228, May 1999. CO-DEN ATISBQ. ISSN 1094-9224 (print), 1557-7406 (electronic).

                                        URL http://www.acm.org/
                                        pubs/citations/journals/tissec/
                                        1999-2-2/p177-jaeger/.

Jarvinen:1999:BTF

[JPT^{+}99] Kimmo Järvinen, Tuukka Pienimäki, Tommi Teräsvirta, John Joel Kyaruzi, and Erkki Sutinen. Between Tanzania and Finland: learning Java over the Web. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31 (1):217–221, March 1999. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Joshi:1997:CGJ

[JRC97] Daniel I. Joshi, Rodney Runolfson, and Ramesh Chandak. The Comprehensive Guide to the JDBC SQL API. Ventana Press, Chapel Hill, NC, USA, 1997. ISBN 1-56604-637-8. xxiv + 422 pp. LCCN QA76.625.J67 1997. US$50. URL http:

                                        //www.vmedia.com/bookshop/
                                        book.html?isbn=1566046378.

Includes CD-ROM.

JRE:1998:DJR

[JRE98] Downloading the Java Runtime Environment Version 1.1. Sun Microsystems,

                                        1998. URL http://java.sun.
                                        com/products/jdk/1.1/jre/
                                        index.html.

Jain:1997:ECC

[JS97] P. Jain and D. C. Schmidt. Experiences converting a C++ communication software framework to Java. C++ Report, 9 (1):50–66, January 1997. CO-DEN CRPTE7. ISSN 1040-6042.

Jensen:1998:JED

[JSA98] Cary Jensen, Blake Stone, and Loy Anderson. JBuilder Essentials: The Developer’s Guide to Java Programming with Borland’s JBuilder. Osborne/McGraw-Hill, Berkeley, CA, USA, 1998. ISBN 0-07-882223-8. xxii + 528 pp. LCCN ???? US$35.

Jubin:1997:JEC

[JT97] Henry Jubin and The Jalapeño Team. JavaBeans by Example: Cooking Beans in the Enterprise. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, December 1997. ISBN 0-13-790338-3. ???? pp. LCCN QA76.73.J38J83 1997.

US$45.  URL http://www.
prenhall.com/ptrbooks/ptr_
0137903383.html.

Jubin:19xx:CJE

[JTxx] Henri Jubin and The Jalapeño Team. Cooking JavaBeans in The Enterprise. IBM Redbooks, 11400 Burnet Road, Austin, TX 78758-3493, USA, 19xx. ISBN ???? LCCN ???? US$50. URL http://www.

redbooks.ibm.com/SG247006/
index.htm.

Jun:1998:J

[Jun98] Yoo Hong Jun. Java. Gijutsu Hyoron Co., Ltd., ????, May 1998. ISBN 4-7741-0591-0. LCCN ???? 1880 yen.

URL http://www.gihyo.co.
jp/syoseki/may/bn-5d.htm.

Jurvis:1996:GJT

[Jur96] J. Jurvis. A guide to Java tools. Information Week, ???? (599):55–56, 58, 60, 64, ???? 30, 1996. CODEN INFWE4. ISSN 8750-6874.

Joshi:1997:ONP

[JV97] Daniel J. Joshi and Pavel A. Vorobiev. Official Netscape Programming Java 1.1 With JFC. Ventana Press, Chapel Hill, NC, USA, October 1997. ISBN 1-56604-766-8 (softcover). ???? pp. LCCN QA76.73.J38J675 1997. US$49.99. URL http:

//www.vmedia.com/bookshop/
book.html?isbn=1566047668.

Joshi:1998:JJF

[JV98] Daniel Joshi and Pavel A. Vorobiev. JFC, Java Foundation Classes. IDG Books, San Mateo, CA, USA, May 1998. ISBN 0-7645-8041-8. LCCN ???? US$50.

Jacobs:1998:RAJ

[JvdBH^{+}98] Bart Jacobs, Joachim van den Berg, Marieke Huisman, Martijn van Berkum, Ulrich Hensel, and Hendrik Tews. Reasoning about Java classes: preliminary report. ACM SIGPLAN Notices, 33(10):329–340, October 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jain:1998:DPM

[JWS98] Prashant Jain, Seth Widoff, and Douglas C. Schmidt. The design and performance of Med-Java. In USENIX [USE98b], page ?? ISBN ???? LCCN

????  URL http://www.
usenix.org/publications/library/
proceedings/coots98/jain.
html.

Jepsen:1998:AAJ

[JZ98] T. C. Jepsen and Z. Zhang.

An ATM API for Java. Lecture Notes in Computer Science, 1430:167–??, 1998. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kahan:1998:HJFa

[Kah98] William Kahan. How Java’s floating-point hurts everyone everywhere. In ACM [ACM98a], page ?? CODEN CPEXEI. ISSN 1040-3108. LCCN ????

URL http://www.cs.ucsb.
edu/conferences/java98/papers/

javahurt.pdf. Also published as Concurrency: Practice and Experience, 10(11–13), September 1998, CODEN CPEXEI, ISSN 1040-3108.

Kalin:1995:NIW

[Kal95a] Sari Kalin. News: Internet World 95: Sun’s Schmidt foresees Net appliance. IDG International News Service, page ??, October 30, 1995.

Kalin:1995:NWC

[Kal95b] Sari Kalin. News: WWW Conference: Web Consortium forges Active Object agreement. IDG International News Service, page ??, December 12, 1995.

Kalev:1998:PCA

[Kal98] Danny Kalev. Porting a C++ application to Java. C/C++ Users Journal, 16(2):??, February 1998. CODEN CCUJEX. ISSN 1075-2838.

Kamba:1996:PON

[Kam96] T. Kamba. Personalized online newspaper. NEC Technical Journal = NEC giho, 49 (7):11–16, July 1996. CODEN NECGEZ. ISSN 0285-4139.

Kamara:1999:MWJ

[Kam99] Sheila Kamara. Multimedia at work: JavuNetwork: remote video production and storage. IEEE MultiMedia, 6(3):78–80, July–September 1999. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

                                        http://dlib.computer.org/
                                        mu/books/mu1999/pdf/u3078.
                                        pdf.

Kanerva:1996:JFFa

[Kan96a]  Jonni Kanerva.  The Java

FAQ: Frequently Asked Questions. Addison-Wesley, Reading, MA, USA, July 1996. ISBN 0-201-63456-2. 352 pp. LCCN ???? US$28. URL http://

                                        www.aw.com/cseng/titles/0-
                                        201-63456-2.

Kaneshige:1996:ITC

[Kan96b] Thomas Kaneshige. Industry trends: Cyberwar; Clipper chip; network computers. Computer, 29(7):20–23, July 1996. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). Contains news brief: Lots of Java brewing at conference.

Kanada:1997:WPR

[Kan97] Yasusi Kanada. Web pages that reproduce themselves by JavaScript. ACM SIGPLAN Notices, 32(11):49–56, November 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kanerva:19xx:JF

[Kanxxa]  Jonni Kanerva.  The Java

FAQ. GOTOP Information Inc., 5F, No.7, Lane 50, Sec.3 Nan Kang Road Taipei, Taiwan; Unit 1905,Metro Plaza Tower 2, No.223 Hing Fong Road, Kwai Chung, N.T., Hong Kong, 19xx. ISBN ???? ???? pp. LCCN ????

???? URL http://www.gotop.

com.tw. Chinese translation by Thi Shiang Workshop.

Kanerva:19xx:JFC

[Kanxxb] Jonni Kanerva. The Java FAQ (Chinese translation by Thi Shiang Workshop). GOTOP Information Inc., ????, 19xx. ISBN ???? ???? pp. LCCN ????

Kanerva:19xx:JFFb

[Kanxxc] Jonni Kanerva. The Java FAQ:

Frequently Asked Questions (Korean translation). DaeRim Publishing, Seoul, Korea, 19xx. ISBN 89-7280-392-8 (??invalid ISBN??). ???? pp. LCCN ???? 13,000 won. URL http:

//www.drbook.co.kr/daerim/

javafaq.htm. Korean translation.

Kaplan:1997:IMW

[Kap97a] H. Kaplan. Interactive multimedia and the World Wide Web. Educom review, 32(1): 48–51, January–February 1997. CODEN EDREEW. ISSN 1045-9146.

Kaplan:1997:DJ

[Kap97b]  Philippe Kaplan.  Debug-

ging in Java. ;login: the USENIX Association newsletter, 22(7):??, December 1997. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

usenix.org/publications/java/
javaindex.html;   http://
www.usenix.org/publications/
java/usingjava8.html.

Kaplan:1999:IET

[Kap99] Gadi Kaplan. Industrial electronics [technology 1999 analysis and forecast]. IEEE Spectrum, 36(1):68–72, January 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Karpinski:1995:HJA

[Kar95a] Richard Karpinski. Hot Java arrives: Sun aims to revolutionize the Web. InteractiveAge, page ??, May 22, 1995. URL

http://techweb.cmp.com/ia/
15issue/15hotjava.html.

Karpinski:1995:SLC

[Kar95b] Richard Karpinski. ScriptX language could give Sun’s Hot Java a run for its money— IBM adds Web hooks to multimedia tools. InteractiveAge, page ??, July 17, 1995. URL

http://techweb.cmp.com/ia/
19issue/19ibm.html.

Karpinski:1996:BHC

[Kar96] Richard Karpinski. Beyond HTML: Covers Acrobat, Java, Shockwave, VRML, and More. Osborne/McGraw-Hill, Berkeley, CA, USA, 1996. ISBN 0-07-882198-3. xxiii + 472 pp. LCCN QA76.76.H94K37 1996. US$27.95.

Karl:1998:BGBa

[Kar98a] Holger Karl. Bridging the gap between distributed shared memory and message passing. In ACM [ACM98a], page ?? CODEN CPEXEI. ISSN 1040-3108. LCCN ????

URL http://www.cs.ucsb.
edu/conferences/java98/papers/
bridging.pdf;   http://www.
cs.ucsb.edu/conferences/java98/
papers/bridging.ps.  Also

published as Concurrency: Practice and Experience, 10(11– 13), September 1998, CODEN CPEXEI, ISSN 1040-3108.

Karl:1998:BGBb

[Kar98b] Holger Karl. Bridging the gap between distributed shared memory and message passing. Concurrency: Practice and Experience, 10(11– 13):887–900, September 1998. CODEN CPEXEI. ISSN 1040-3108. URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=10050425;
                                         http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=10050425&PLACEBO=IE.pdf.

Special Issue: Java for Highperformance Network Computing.

Kassabgi:1997:PDG

[Kas97a] George Kassabgi. The Professional Developer’s Guide to JavaBeans. Que Corporation, Indianapolis, IN, USA, 1997. ISBN 0-7897-1168-0. LCCN ???? ????

Kassenaar:1997:BJV

[Kas97b] Peter Kassenaar. Basiscursus JavaScript: versie 1.1. Academic Service, Postbus 81, 2870 AB Schoonhoven, The Netherlands, October 1997. ISBN 90-395-0781-3. 250 pp. LCCN ???? fl. 29,50. URL http:/

                                        /acaserv.nl/bestel/5_0781.
                                        htm.

Kay:1996:JC

[Kay96a] Jeffrey Kay. Java and C++. Dr. Dobb’s Journal of Software Tools, 21(11):8, 10, November 1996. CODEN DDJOEB. ISSN 1044-789X. Comments on [Ste96d].

Kay:1996:JAW

[Kay96b] Roger L. Kay. Java at work. Computerworld, 30(20):85, 88, May 13, 1996. CODEN CMP-WAB. ISSN 0010-4841.

Kay:1997:CRR

[Kay97] Russell Kay. CD-ROM review: Prevent Java attacks: Good explanations but lame advice. BYTE Magazine, 22(4): 38, April 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic). Brief review of [MF97c]. See rebuttal in [MF97d].

Kamba:1996:IPN

[KB96] T. Kamba and K. Bharat. An interactive, personalized, newspaper on the WWW. Proceedings of the SPIE — The International Society for Optical Engineering, 2667:290–301, ???? 1996. CODEN PSISDG. ISSN 0277-786X (print), 1996-756X (electronic).

Keren:1998:APPa

[KB98a] Arie Keren and Amnon Barak.

Adaptive placement of parallel Java agents in a scalable computing cluster. In ACM [ACM98a], page ?? CODEN CPEXEI. ISSN 1040-3108. LCCN ????

URL http://www.cs.ucsb.
edu/conferences/java98/papers/
agents.pdf;   http://www.
cs.ucsb.edu/conferences/java98/
papers/agents.ps. Also pub-

lished as Concurrency: Practice and Experience, 10(11– 13), September 1998, CODEN CPEXEI, ISSN 1040-3108.

Keren:1998:APPb

[KB98b] Arie Keren and Amnon Barak. Adaptive placement of parallel Java agents in a scalable computing cluster. Concurrency: Practice and Experience, 10(11–13):971–976, September 1998. CODEN CPEXEI. ISSN 1040-3108. URL http:

                                        //www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=10050404;
                                         http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=10050404&PLACEBO=IE.pdf.

Special Issue: Java for Highperformance Network Computing.

Krintz:1998:OET

[KCLZ98] Chandra Krintz, Brad Calder, Han Bok Lee, and Benjamin G. Zorn. Overlapping execution with transfer using non-strict execution for mobile programs. ACM SIG-PLAN Notices, 33(11):159– 169, 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

                                        tronic). URL http://www.
                                        acm.org:80/pubs/citations/
                                        proceedings/asplos/291069/
                                        p159-krintz/.

Kurtz:1998:CSD

[KCPC98] Barry L. Kurtz, Hong Cai, Chris Plock, and Xijia Chen. A concurrency simulator designed for sophomore-level instruction. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(1):

237–241, March 1998. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kahan:1998:HJFb

[KD98] W. Kahan and Joseph D. Darcy.

How Java’s floating-point hurts everyone everywhere. Technical report, Department of Mathematics and Department of Electrical Engineering and Computer Science, University of California, Berkeley, Berkeley, CA, USA, June 18, 1998.

80 pp. URL http://www.cs.
berkeley.edu/~wkahan/JAVAhurt.
pdf;   http://www.cs.berkeley.
edu/~wkahan/JAVAhurt.ps.

Kooros:1996:JBW

[KDH96] Paul Kooros, Michele DeWolfe, and David Heib. JavaScript: Blueprints to a world beyond HTML. Prima Publishing, Rocklin, CA, USA, August 1, 1996. ISBN 0-7615-0685-3. xii + 337 pp. LCCN A76.73.J39K66 1996. US$35.00.

Keefe:1996:JWB

[Kee96] Patricia Keefe. JavaSoft won’t be snubbed. Computerworld, 30 (44):130, October 28, 1996. CO-DEN CMPWAB. ISSN 0010-4841. ????

Keishiro:1995:JJC

[Kei95] T. Keishiro. J2C Java .class to C translator. J2C is compared with Harissa in [MMBC97].,

1995.  URL http://www.
webity.co.jp/info/andoh/java/
j2c.html.

Keyes:1998:DHS

[Key98]  Jessica Keyes.  DataCast-

ing: How to Stream Databases over the Web (including the JDBC). McGraw-Hill, New York, NY, USA, 1998. ISBN 0-07-034678-X. xvii + 555 pp. LCCN QA76.9.D3K35898 1998. US$40. URL http://

                                        www.betabooks.mcgraw-hill.
                                        com/keyes2/index.htm.

Kistler:1999:TAJ

[KF99a] Thomas Kistler and Michael Franz. A tree-based alternative to Java byte-codes. International Journal of Parallel Programming, 27(1):21–??, ???? 1999. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Kistler:1999:TBA

[KF99b] Thomas Kistler and Michael Franz. A tree-based alternative to Java byte-codes. International Journal of Parallel Programming, 27(1):21– 33, February 1999. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (elec-

                                        tronic). URL http://www.
                                        springerlink.com/openurl.
                                        asp?genre=article&issn=0885-
                                        7458&volume=27&issue=1&spage=
                                        21.

Knies:1999:TIA

[KFL99] Allan Knies, Jesse Fang, and Wei Li. Tutorial: IA64 architecture and compilers. In IEEE [IEE99a], page ?? ISBN ????

LCCN ???? URL http://www.
hotchips.org/hotc11_index.
html.

Kann:1997:EPA

[KFS97] Charles W. Kann, Michael B.

Feldman, and John Sibert. Experience programming applets with Ada95. ACM SIGADA Ada Letters, 17(3):17–29, May/ June 1997. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). Uses an early beta release of AppletMagic for compiling Ada95 programs into code for the Java Virtual Machine.

Kleinoder:1996:MER

[KG96a] J. Kleinoder and M. Golm.

MetaJava: an efficient runtime meta architecture for Java. In Cabrera and Islam [CI96], page ?? CODEN PIWOF6. ISBN 0-8186-7692-2, 0-8186-7693-0, 0-8186-7694-9 (microfiche). ISSN 1063-5351. LCCN QA 76.76 O63 I59 1996.

Kleinoeder:1996:MER

[KG96b] Jürgen Kleinoeder and Michael Golm. MetaJava: An efficient run-time meta architecture for Java. In Cabrera and Islam [CI96], pages 54–61. CODEN PIWOF6. ISBN 0-8186-7692-2, 0-8186-7693-0, 0-8186-7694-9 (microfiche). ISSN 1063-5351. LCCN QA 76.76 O63 I59 1996.

Knapp:1997:HJE

[KG97a] Michael Knapp and Wolfram Gieseke. HTML and JAVA 100% efficace. Ed. Micro application, Paris, France, 1997. ISBN 2-7429-0924-9. 364 pp. LCCN ????

Krall:1997:CBJ

[KG97b] Andreas Krall and Reinhard Grafl. CACAO — a 64bit JavaVM just-in-time compiler. Concurrency: Practice and Experience, 9(11): 1017–1030, November 1997. CODEN CPEXEI. ISSN 1040-3108. URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=13831;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=13831&PLACEBO=IE.pdf. Spe-

cial Issue: Java for computational science and engineering — simulation and modeling II.

Kassab:1998:TFJ

[KG98a] L. L. Kassab and S. J. Greenwald. Towards formalizing the Java security architecture of JDK 1.2. Lecture Notes in Computer Science, 1485:191– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kleinoder:1998:MPA

[KG98b] Jürgen Kleinöder and Michael Golm. MetaJava — a platform for adaptable operating-system mechanisms. Lecture Notes in Computer Science, 1357:507– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/

service/series/0558/bibs/
1357/13570507.htm;   http:
//link.springer-ny.com/link/
service/series/0558/papers/
1357/13570507.pdf.

Kleinoeder:1998:MPA

[KG98c] J. Kleinoeder and M. Golm.

MetaJava — a platform for adaptable operating-system mechanisms. Lecture Notes in Computer Science, 1357:507– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Karduck:1996:MWM

[KGG96] Achim P. Karduck, André Geiser, and Thomas Gutekunst. Multimedia at work: Multimedia technology in banking. IEEE MultiMedia, 3(4): 82–86, Winter 1996. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

URL http://dlib.computer.
org/mu/books/mu1996/pdf/u4082.

pdf. Discusses the Portfolio Analysis System (PAS) using Java.

Kerievsky:1997:OJP

[KGTF97] Joshua Kerievsky, Terence Goggin, Matt Telles, and Steven Fraser. Open JBuilder Programming Frontrunner: The Hands-On Guide to Mastering Cross-Platform Development With Open JBuilder. Coriolis Group Books, Scottsdale, AZ, USA, August 1997. ISBN 1-57610-004-9. ???? pp. LCCN ???? US$29.99. Duplicate ISBN with [Dun96, GTFxx, KGTF97].

Khuri:1999:VCS

[KH99] Sami Khuri and Hsiu-Chin Hsu. Visualizing the CPU scheduler and page replacement algorithms. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1): 227–231, March 1999. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Karaorman:1999:CRJ

[KHB99a] M. Karaorman, U. Hoelzle, and J. Bruno. Contractor: a reflective Java library to support design by contract. Lecture Notes in Computer Science, 1616:175– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Karaorman:1999:JRJ

[KHB99b] Murat Karaorman, Urs Hölzle, and John Bruno. jContractor: a reflective Java library to support design by contract. Lecture Notes in Computer Science, 1616:175–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/
                                        service/series/0558/bibs/
                                        1616/16160175.htm;   http:
                                        //link.springer-ny.com/link/
                                        service/series/0558/papers/
                                        1616/16160175.pdf.

Kierdorf:1996:IPJ

[Kie96] Wolfgang H.-D. Kierdorf. Internet Programmierung mit Java. Addison-Wesley, Reading, MA, USA, 1996. ISBN 3-612-28134-8. 308 pp. LCCN ????

40 DM. URL http://www.
addison-wesley.de/. Includes

CD-ROM.

Kientzle:1999:JAS

[Kie99] Tim Kientzle. A Java applet search engine. Dr. Dobb’s Journal of Software Tools, 24 (2):32, 36, 38–40, February 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/1999/9902/9902toc.
htm;   http://www.ddj.com/
ftp/1999/1999_02/jsearch.
txt;   http://www.ddj.com/
ftp/1999/1999_02/jsearch.
zip.

Kim:1996:RAW

[Kim96] Eugene Eric Kim. Running at Warp speed. Dr. Dobb’s Journal of Software Tools, 21(12): 12, December 1996. CODEN DDJOEB. ISSN 1044-789X. Brief discussion of Java products under IBM OS/2.

Kim:1997:LJP

[Kim97a] Chungho Kim. Learning Java Properly. Moonyeimadang, ??, Korea, 1997. ISBN ???? ???? pp. LCCN ???? 15,000 won.

URL ????.

Kim:1997:DDG

[Kim97b] Eugene Eric Kim, editor. Dr.

Dobb’s Guide to Java Programming. R&D Publications, ????, 1997. ISBN 0-87930-499-5 (softcover). ???? pp. LCCN ???? US$19.95. URL http://www.

                                        rdbooks.com/scripts/store/
                                        vsc/store/products/rd2830.
                                        htm?L+/htdocs/rdbooks/config/
                                        store+aotp4118.

Kim:1997:JP

[Kim97c] Hyongchol Kim. The Java Programming. Game Pub., ??, Korea, 1997. ISBN 89-8078-035-4 (??Invalid checksum??). ???? pp. LCCN ???? 12,000 won.

                                        URL ????.

Kim:1997:FLJ

[Kim97d] Seokjoo Kim. First Love with Java [Jaba wa ui chotsarang: Intonet eso jaba aepullit mandulgi]. Kanam Press, Seoul, Korea, 1997. ISBN 89-7354-144-7 (??Invalid checksum??). 311 pp. LCCN ???? 11,000 won. URL

                                        ????.

Kim:1997:TJ

[Kim97e] Seokjoo Kim. Temptation to JavaScript. Kanam Press, Seoul, Korea, 1997. ISBN 89-7354-157-9 (??Invalid checksum??). ???? pp. LCCN ???? 13,000 won. URL ????.

Kim:1997:LJ

[Kim97f] Taehyon Kim. I Love Java! Crown Publishing, ??, Korea, 1997. ISBN 89-406-3014-X (??Invalid checksum??). ???? pp. LCCN ???? 12,000 won.

                                        URL ????.

Kim:19xx:LJP

[Kimxx] Chungho Kim. Learning Java Properly. Moonyeimadang, ????, 19xx. LCCN ???? 15,000 won.

King:1996:JTG

[Kin96] Julia King. Java top guns seek glory, not gold. Computerworld, 30(42):2, October 14, 1996. CO-DEN CMPWAB. ISSN 0010-4841. ????

Kindlund:1997:INA

[Kin97a]  Erika Kindlund.  Interface:

Navigating the applet-browser divide. IEEE Software, 14 (5):22–25, September/October 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/neumann.computer.org/so/
books/so1997/pdf/s5022.pdf.

Discusses design issues for Java applets.

King:1997:TTR

[Kin97b] Julia King. Tough test reveals real Java pros. Computerworld, 31(5):61–62, February 3, 1997. CODEN CMPWAB. ISSN 0010-4841.

King:1997:EJ

[Kin97c] K. N. King. Essence of Java. W.

W. Norton & Co., New York, NY, USA, June 1997. ISBN 0-393-31661-0. 300 pp. LCCN ???? US$24.95. URL http://

norton3.wwnorton.com/catnos/
tl031661.htm.

Kinne:1999:WJP

[Kin99] Morgan Kinne. Writing Javabean property editors. Dr. Dobb’s Journal of Software Tools, 24(2):52, 54–56, 58, 60, 62, February 1999. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/1999/9902/9902toc.
htm;   http://www.ddj.com/
ftp/1999/1999_02/propedit.
txt;   http://www.ddj.com/
ftp/1999/1999_02/propedit.
zip.

Kiuttu:1998:OIJ

[Kiu98] Petri Kiuttu. Opeta itsellesi Java JBuilder. Suomen Atkkustannus Oy, ????, May 1998. ISBN 951-762-642-8. LCCN

???? ???? URL http://www.
satku.fi/kirjat/4jbuilder.
htm.

Kiuttu:19xx:JOP

[Kiuxx] Petri Kiuttu. Java-ohjelmointi pro-kurssi. Suomen Atkkustannus Oy, ????, 19xx. ISBN 951-762-548-0. ???? pp. LCCN

???? ???? URL http://www.
satku.fi/4javapro.htm.

Kandlur:1998:MCN

[KJR98] Dilip D. Kandlur, Kevin Jeffay, and Timothy Roscoe, editors. Multimedia computing and networking 1999: 25–27 January, 1999, San Jose, California, number 3654 in SPIE Proceedings. SPIE Optical Engineering Press, Bellingham, WA, USA, 1998. ISBN 0-8194-3125-7. ISSN 0277-786X (print), 1996-756X (electronic). LCCN TS510.S63 v.3654.

Kent:1996:ONJ

[KK96a] Peter Kent and John Kent. Official Netscape JavaScript Book: The Nonprogrammer’s Guide to Interactive Web Pages. Ventana Communications Group, Incorporated, Research Triangle Park, NC, USA, July 1, 1996. ISBN 1-56604-465-0. xx + 480 pp. LCCN QA76.73.J39K46 1996. US$29.99.

Koosis:1996:JPD

[KK96b] Donald J. Koosis and David S.

Koosis. Java Programming for Dummies. I D G Books Worldwide, Indianapolis, IN, USA, June 1, 1996. ISBN 1-56884-995-8. xxvi + 382 (or 408??) pp. LCCN ???? US$29.99. URL http://db.

www.idgbooks.com/database/
book/isbn/generic-book.tmpl?
query=1-56884-995-8;   http:
//www.idgbooks.com/database/
book_result.msql?uid=uidg323&
isbn=1-56884-995-8.

Kang:1997:PJ

[KK97a] Hee-Chun Kang and Bok-Sik Kim. Programming with Java! CA Books c/o DRT International, ??, Korea, 1997. ISBN 89-86604-08-6. ???? pp. LCCN ???? 15,000 Won. URL ????.

Kent:1997:ONJ

[KK97b] Peter Kent and John Kent.

Official Netscape JavaScript 1.2 book: the nonprogrammer’s guide to creating interactive Web pages. Ventana Press, Chapel Hill, NC, USA, second edition, 1997. ISBN 1-56604-675-0. xxx + 550 pp. LCCN QA76.73.J39 K459 1997. US$29.99.

Ki:1997:CSD

[KK97c] Byeongseob Ki and Scott Klasky. Collaborative scientific data visualization. Concurrency: Practice and Experience, 9(11):1249–1259, November 1997. CODEN CPEXEI. ISSN 1040-3108. URL http:

                                        //www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=13822;
                                         http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=13822&PLACEBO=IE.pdf. Spe-

cial Issue: Java for computational science and engineering — simulation and modeling II.

Koosis:1997:JPDa

[KK97d] Donald Koosis and David Koosis. Java Programming for Dummies. IDG Books, San Mateo, CA, USA, second edition, 1997. ISBN 0-7645-0141-0 (softcover). xxvi + 382 pp. LCCN QA76.73.J38K66 1997. US$25.

                                        URL http://www.idgbooks.
                                        com/database/book_result.
                                        msql?uid=uidg323&isbn=0-7645-

0141-0. Includes CD-ROM.

Koosis:1997:PJV

[KK97e] Donald Koosis and David Koosis. Programmeren in Java voor Dummies. Addison-Wesley Nederland, ????, second edition, December 1997. ISBN 90-6789-898-8. ???? pp. LCCN ???? fl. 70.

Koosis:1997:JPDb

[KK97f] Donald J. Koosis and David Koosis. Java-Programmierung für Dummies. International Thomson Computer Press, 20 Park Plaza Suite 1001, Boston, MA 02116 USA, 1997. ISBN 3-8266-2732-6 (??invalid ISBN??). 372 pp. LCCN ???? 50 DM.

URL http://www.itp.de/;
http://www.itp.de/dummies/
2732/2732.html. Includes CD-

ROM.

Ki:1998:CSD

[KK98a] Byeongseob Ki and Scott Klasky. Collaborative scientific data visualization. In ACM [ACM98a], page ?? CODEN CPEXEI. ISSN 1040-3108. LCCN ????

URL http://www.cs.ucsb.
edu/conferences/java98/papers/
scivis.pdf;   http://www.
cs.ucsb.edu/conferences/java98/
papers/scivis.ps. Also pub-

lished as Concurrency: Practice and Experience, 10(11– 13), September 1998, CODEN CPEXEI, ISSN 1040-3108.

Ki:1998:S

[KK98b] Byeongseob Ki and Scott Klasky. Scivis. Concurrency: Practice and Experience, 10(11–13):1107–1115, September 1998. CODEN CPEXEI. ISSN 1040-3108. URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=10050426;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=10050426&PLACEBO=IE.pdf.

Special Issue: Java for Highperformance Network Computing.

Koosis:1998:JPD

[KK98c] Donald Koosis and David Koosis. Java Programming for Dummies. IDG Books, San Mateo, CA, USA, third edition, June 1998. ISBN 0-7645-0388-X. xxiv + 375 pp. LCCN QA76.73.J38 K66 1998. US$30.

Kent:19xx:JPN

[KKxxa] Peter Kent and John Kent. JavaScript para Netscape: Guia Oficial. Makron Books do Brasil Editora Ltda., Rua Tabapuã 1348, São Paulo SP - 04533-004 - Brasil, 19xx. ISBN 85-346-0722-2. 510 pp. LCCN ???? R$ 59,00. URL http:/

                                        /safe.tesla.com.br/makron/
                                        options/livro.asp?id=24;
                                        http://www.makron.com/; http:
                                        //www.mcp.com/authors/pkent/

. Portuguese translation of [KK96a]. Includes floppy disk.

Koosis:19xx:JPN

[KKxxb] Donald Koosis and David Koosis. Java pour Les Nuls. Sybex France, Paris, France, 19xx. ISBN 2-7361-2349-2. ???? pp. LCCN ????

                                        140 FF. URL http://www-
                                        db.sybex.fr/cgi-bin/Sybex/
                                        Webdriver?MIval=Fiche_Book.
                                        html&Reference=2349;   http:
                                        //www.sybex.fr/.  Includes

CD-ROM.

Koosis:19xx:JPV

[KKxxc] Donald Koosis and David Koosis. Java Programmeren voor Dummies. Addison-Wesley Nederland, ????, 19xx. ISBN 90-6789-754-X. ???? pp. LCCN ???? fl. 50.

Kim:1997:LED

[KL97] Jihong Kim and Ted Lewis. Letter to the editor: Deadlocking embrace. Computer, 30(4):7, April 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Klute:1997:JPD

[Klu97] Rainer Klute. JDBC in der Praxis: Datenbankanwendungen in Intra- und Internet. Addison-Wesley Verlag, Bonn, Germany, 1997. ISBN 3-8273-1301-5. 300 pp. LCCN ????

60 DM. URL http://www.
addison-wesley.de/katalog/
item.ppml?id=00264.

Klute:1998:JPD

[Klu98] Rainer Klute. JDBC in der Praxis: Datenbankanwendungen in Intranet und Internet; [Entwicklungssoftware und Tools: Java development kit (JDK), Java servlet development kit (JSDK), PostgreSQL u.v.m.]. Addison-Wesley Longman, Reading, MA, USA, 1998. ISBN 3-8273-1301-5. 287 (est.) pp. LCCN ???? Includes CD-ROM.

Klute:19xx:JPD

[Kluxx]  Rainer Klute.  JDBC in

der Praxis: Datenbankanwendungen in Intra- und Internet. Addison-Wesley Verlag, Bonn, Germany, 19xx. ISBN 3-8273-1301-5 (??invalid ISBN??). LCCN ???? 60

                                        DM.  URL http://www.
                                        addison-wesley.de/katalog/
                                        item.ppml?id=00264.

Kamin:1998:ICS

[KMR98] Samuel N. Kamin, M. Dennis Mickunas, and Edward M. Reingold. An Introduction to Computer Science Using Java. McGraw-Hill, New York, NY, USA, 1998. ISBN 0-07-034224-

5. ???? pp. LCCN QA76.73.J38K36 1998. US$55.

Kirby:1998:LRJ

[KMS98] Graham Kirby, Ron Morrison, and David Stemple. Linguistic reflection in Java. Software—Practice and Experience, 28(10):1045–1077, August 1998. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http:

                                        //www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=10007353;
                                         http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=10007353&PLACEBO=IE.pdf.

Kummer:1999:SCB

[KMW99] O. Kummer, D. Moldt, and F. Wienberg. Symmetric communication between coloured Petri net simulations and Javaprocesses. Lecture Notes in Computer Science, 1639:86–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kiuttu:19xx:OIJ

[KNxx] Petri Kiuttu and Juha Niemi.

Opeta itsellesi Java-ohjelmointi. Suomen Atk-kustannus Oy, ????, 19xx. ISBN 951-762-468-9. ???? pp. LCCN ???? ????

Knappe:19xx:HJE

[Knaxxa] Heiko Knappe. HTML et Java 100% Efficace. Edition Micro Application, Paris, France, 19xx. ISBN 2-7429-0924-9. ???? pp. LCCN ???? 135 FF.

Knappe:19xx:HJ

[Knaxxb] Heiko Knappe. HTML und Java. Vogel, ????, 19xx. ISBN 3-8259-1382-1 (??invalid ISBN??). LCCN ???? 34 DM.

Knecht:1996:WDJ

[Kne96] Joachim Knecht. Web Design mit Java. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., October 1996. ISBN 3-540-61182-7. ???? pp. LCCN ???? 40 DM. URL http:/

/www.springer.de/. Includes

CD-ROM.

Knecht:1998:JWE

[Kne98] Joachim Knecht. Java Workshop 2.0: ein Einführungskurs in die Programmierung. dpunkt-Verlag, Heidelberg, Germany, 1998. ISBN 3-920993-65-9. x + 267 pp. LCCN ???? 68 DM.

Koved:1998:EJS

[KNNL98] L. Koved, A. J. Nadalin, D. Neal, and T. Lawson. The evolution of Java security. IBM Systems Journal, 37(3):349– ??, ???? 1998. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

URL http://www.almaden.
ibm.com/journal/sj/373/koved.
html.

Knowles:1995:NUE

[Kno95] Anne Knowles. News: Unix expo: Second-tier vendors unveil hardware. IDG International News Service, page ??, September 21, 1995.

Knobloch:1996:JBB

[Kno96a] Frank Knobloch. Java: (Das BHV Buch). BHV, ??, Germany, September 1996. ISBN 3-89360-335-2. ???? pp. LCCN ???? 79 DM. URL ????. Includes CD-ROM.

Knowles:1996:SWJ

[Kno96b] Anne Knowles. Second-wave Java tools tuned to perform. Datamation, 42(18):82–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

KnowledgeMedia:19xx:BC

[Knoxx] Knowledge Media. Black Coffee. Knowledge Media, ????, 19xx. ISBN ???? ???? pp. LCCN ????

Knudsen:1997:JTM

[Knu97a] Jonathan Knudsen. Java talk: Mirror, mirror, on the Java …. Silicon Graphics World, 7 (11):12, 17, 27, November 1997. ISSN 1057-7041.

Knudsen:1997:JMMa

[Knu97b] Jonathan Knudsen. JavaTalk:

Mirror, mirror on the Java. Sun-Server, 11(10):19, 25, 27–28, October 1997. ISSN 1091-4986.

Knudsen:1997:JMMb

[Knu97c] Jonathan Knudsen. JavaTalk:

Mirror, mirror, on the Java. … HP Chronicle, 14(10):10, 22, September 1997. ISSN 0892-2829. Discusses the Java Reflection API, a set of classes that allows finding out information about Java classes at runtime.

Knudsen:1997:JTT

[Knu97d] Jonathan Knudsen. JavaTalk:

Telling time in Java 1.1. Silicon Graphics World, 7(9):12, 18, September 1997. ISSN 1057-7041.

Knudsen:1997:JUPb

[Knu97e] Jonathan Knudsen. JavaTalk; using a passphrase to protect data. SunServer, 11(11):18, 20, November 1997. ISSN 1091-4986.

Knudsen:1997:JUPa

[Knu97f] Jonathan Knudsen. JavaTalk:

Using a passphrase to protected data. HP Chronicle, 14(12):20– 21, 23, November 1997. ISSN 0892-2829. Contains brief comparison of Java and Limbo.

Knudsen:1997:TTJ

[Knu97g] Jonathan Knudsen. Telling time in Java 1.1. HP Chronicle, 14 (9):20, 25, August 1997. ISSN 0892-2829.

Knut:1997:NIX

[Knu97h] Detlef Knut. Novell Intranet-Ware 4.x Referenz. Sybex Verlag GmbH, Erkrather Straße 345-349, 40231 Düsseldorf, Germany, 1997. ISBN 3-8155-7262-2. xviii + 842 pp. LCCN ???? 89 DM.

Knudsen:1998:JHH

[Knu98a] Jonathan Knudsen. JavaTalk: Horseshoes, hand grenades and random numbers. SunServer, 12 (1):16–17, January 1998. ISSN 1091-4986.

Knudsen:1998:JC

[Knu98b] Jonathan B. Knudsen. Java Cryptography. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, May 1998. ISBN 1-56592-402-9. xvi + 344 pp. LCCN QA76.73.J38 K59 1999. US$29.95. URL http://www.

ora.com/catalog/javacrypt/

.

Knudsen:1999:TAM

[Knu99a] Craig Knudsen. Transvirtual adopts Microsoft Java extensions. Linux Journal, 66:??, October 1999. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http://

noframes.linuxjournal.com/
lj-issues/issue66/3482.html.

Knudsen:1999:JG

[Knu99b]  Jonathan Knudsen.  Java

2D Graphics. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1999. ISBN 1-56592-484-3. xvi + 339 pp. LCCN QA76.73.J38 K58 1999. US$29.95. URL http:

//www.oreilly.com/catalog/
java2d/.

Koch:1997:JEP

[Koc97] Stefan Koch. JavaScript. Einfuehrung, Programmierung und Referenz. dpunkt-Verlag, Heidelberg, Germany, February 15, 1997. ISBN 3-920993-64-0. 494 pp. LCCN ???? 78,00 DM.

URL http://www.dpunkt.de/
javascript;   http://www.
dpunkt.de/welcome.html;
http://www.dpunkt.de:80/produkte/
JavaScript.html;   http://
www.emedia.de/bin/bookshop?
show=4873&id=. Includes CD-

ROM.

Kahan:1998:HJF

[KoCaB98] William Kahan and University of California at Berkeley. How Java’s floatingpoint hurts everyone everywhere. In ACM [ACM98a], page ?? CODEN CPEXEI. ISSN 1040-3108. LCCN ????

URL http://www.cs.ucsb.
edu/conferences/java98/papers/

javahurt.pdf. Also published as Concurrency: Practice and Experience, 10(11–13), September 1998, CODEN CPEXEI, ISSN 1040-3108.

Kohlhepp:1995:NGW

[Koh95] Robert J. Kohlhepp. Next generation Web browsing. Network Computing, 6(9):48–??, August 1, 1995. CODEN NCOMEV.

ISSN 1046-4468. URL http:

                                        //techweb.cmp.com/techweb/
                                        nc/609/609sneak1.html.

Kolbeck:19xx:GLJ

[Kolxx] Rainer Kolbeck. El Gran Libro De JavaScript. Editorial Marcombo, S.A., Barcelona, Spain, 19xx. ISBN ???? ???? pp. LCCN ????

Konchady:1998:IJ

[Kon98] Manu Konchady. An introduction to JDBC. Linux Journal, 55:??, November 1998. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL

                                        ftp://ftp.ssc.com/pub/lj/
                                        listings/issue55/2846.tgz.

Kooijmans:1998:IJ

[Koo98] Alex Louwe Kooijmans. Integrating Java with OS/390. IBM Redbooks, 11400 Burnet Road, Austin, TX 78758-3493, USA, May 1998. ISBN ???? LCCN

                                        ???? ???? URL http://www.
                                        redbooks.ibm.com/SG245142/
                                        sg245142.html.

Kalin:1996:SRI

[KP96] Sari Kalin and Michael Parsons. Sun reveals its Internet Toaster. Sunworld Online, 02:??, 1996. ISSN 1091-8914.

                                        URL http://www.sun.com/
                                        sunworldonline/swol-02-1996/
                                        swol-02-appliance.html.

Krall:1998:MEHa

[KP98a] Andreas Krall and Mark Probst. Monitors and exceptions: How to implement Java efficiently. In ACM [ACM98a], page ?? CODEN CPEXEI. ISSN 1040-3108. LCCN ????

URL http://www.cs.ucsb.
edu/conferences/java98/papers/
cacao.pdf;   http://www.cs.
ucsb.edu/conferences/java98/
papers/cacao.ps. Also pub-

lished as Concurrency: Practice and Experience, 10(11– 13), September 1998, CODEN CPEXEI, ISSN 1040-3108.

Krall:1998:MEHb

[KP98b] Andreas Krall and Mark Probst. Monitors and exceptions: how to implement Java efficiently. Concurrency: Practice and Experience, 10 (11–13):837–850, September 1998. CODEN CPEXEI. ISSN 1040-3108. URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=10050393;
http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=10050393&PLACEBO=IE.pdf.

Special Issue: Java for Highperformance Network Computing.

Konstantinou:1999:DJB

[KP99] V. Konstantinou and A. Psarrou. A dynamic JAVA-based intelligent interface for online image database searches. Lecture Notes in Computer Science, 1614:211–??, 1999. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaufman:1996:TYA

[KPF96] Sanders Kaufman, Jr., Jeff Perkins, and Dina Fleet. Teach yourself ActiveX programming in 21 days. Sams.net, Indianapolis, IN 46268, USA, 1996. ISBN 1-57521-163-7. xvi + 575 pp. LCCN QA76.625 .K38 1996. URL http://

merchant.superlibrary.com:
8000/catalog/hg/PRODUCT/PAGE/
15752/bud/1575211637.html.

Also covers JavaScript.

Karacapilidis:1996:ABF

[KPG96] N. I. Karacapilidis, D. Papadias, and T. F. Gordon. An argumentation based framework for defeasible and qualitative reasoning. In Borges and Kaestner [BK96a], page ?? ISBN 3-540-94900-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Q334 .S5434 1996 Bar.

Kolling:1999:TTT

[KR99] Michael Kölling and John Rosenberg. Tools and techniques for teaching objects first in a Java course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1):368, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kramer:1996:JP

[Kra96]  Doug Kramer.  The Java

platform. Report, Sun Microsystems, 2550 Garcia Avenue, Mountain View, CA 94043, USA, 1996. ???? pp.

URL http://java.sun.com/

doc/whitePaper.Platform/JavaPlatform.
doc.html.

Krishna:1996:PRC

[Kri96] Dileep Krishna. Product reviews: Café: the right place for Java. Computer, 29(7):104–108, July 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kroeker:1997:NPJ

[Kro97] Kirk L. Kroeker. New products: Java and Internet development tools: Borland’s Internet Solutions Program, Magic Software Enterprise’s RADD for Enterprise and Internet apps, Symantec’s Visual Café for Java 2.0 professional development edition, Ahpah Software’s SourceAgain Java decompiler, KL Group’s JClass Chart 2.0 JavaBean. Computer, 30(11):128–129, November 1997. CODEN CMPWAB. ISSN 0010-4841. URL http://

pdf.computer.org/co/books/
co1997/pdf/ry125.pdf.

Kroeker:1998:NTC

[Kro98a]  Kirk L. Kroeker.  New

tools: Component technology: 4Developers’ tool for managing ActiveX and COM files; SoftWired’s Java-based object delivery API. Computer, 31(12):121, December 1998. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

http://dlib.computer.org/
co/books/co1998/pdf/rz118.
pdf.

Kroeker:1998:NTSf

[Kro98b] Kirk L. Kroeker. New tools: Software development: Galorath’s parametric estimation tool; Uniscape’s multilingual API and runtime library; Object International’s UML modeler; ObjecTime Limited’s realtime code generation tool. hardware development: Diab Industry’s Java compiler for embedded systems; Verilog’s embedded systems tool; Intusoft’s simulation-based test synthesis tool. net development: Rational Software’s Webbased load-testing tool; IONA Technologies’ Internet security tool; Sonic Systems’ ICSAcertified low-cost firewall. Computer, 31(11):113–116, November 1998. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

http://dlib.computer.org/
co/books/co1998/pdf/ry113.
pdf.

Kroeker:1998:NTSe

[Kro98c] Kirk L. Kroeker. New tools: Software development: Rational Software’s development tools for Ada, Artlandia’s Mathematica graphics tool, ISE’s EiffelBase library available for free, Annasoft System’s Windows CE development tool. hardware development: White Mountain’s DSP tools, Analogy’s mixed-signal IC simulator, Dynamic Soft Analysis’ thermal analysis tool; net development: Framework Technologies’ project management system, CAI’s message broker, Gordian’s network redirector software; component technology: Black & White’s usage control software for multi-ORB CORBA implementations, IONA Technologies’ middleware development tools, Green-Tree Technologies’ ActiveX controls, KL Group’s JavaBeans with automatic data binding. Computer, 31(10):110–114, October 1998. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

http://dlib.computer.org/
co/books/co1998/pdf/rx110.
pdf.

Kroeker:1998:NTM

[Kro98d] Kirk L. Kroeker. New tools:

Software development: Rational Software’s development tools for Ada, Artlandia’s Mathematica graphics tool, ISE’s EiffelBase library available for free, Annasoft System’s Windows CE development tool; hardware development: White Mountain’s DSP tools, Analogy’s mixed-signal IC simulator, Dynamic Soft Analysis’ thermal analysis tool; new development: Framework Technologies’ project management system, CAI’s message broker, Gordian’s network redirector software; component technology: Black & White’s usage control software for multi-ORB CORBA implementations, IONA Technologies’ middleware development tools, Green-Tree Technologies’ ActiveX controls, KL Group’s JavaBeans with automatic data binding. Computer, 31(10):110–114, October 1998. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

                                        http://dlib.computer.org/
                                        co/books/co1998/pdf/rx110.
                                        pdf.

Kroeker:1998:NTSc

[Kro98e] Kirk L. Kroeker. New tools: Software development: Take-Five Software’s source code tool; RST’s test tools; Blue Sky’s help file authoring solution. hardware development: Green Hills Software’s singlesource solution for tools and RTOS; Tanner ERA’s interconnect design tools; Xynetix Design Systems’ autorouter. net development: Aonix’s Web-to-legacy tool; Inktomi’s commercial network cache. component technology: Visual Edge’s enterprise application integration products; Rogue Wave Software’s suite of JavaBeans components. Computer, 31(7):90– 93, July 1998. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

                                        http://dlib.computer.org/
                                        co/books/co1998/pdf/r7090.
                                        pdf.

Kroeker:1999:NTN

[Kro99a] Kirk L. Kroeker. New tools: Net development: Sun’s Java embedded server; MetaCreation’s Web-savvy graphics tool; WebCompiler’s HTML packaging tool. component technology: ProtoVIew Development’s ActiveX tools; Avilon Software’s load balancing component system. software development: Verilog’s test checker; Red Hat and Metrowerks’ development tools for Linux; the Object Factory’s optimization tool; Acumen Systems’s SDK for imaging; Aonix’s process-oriented lifecycle environment; Baan’s embedded software development suite. Computer, 32(5):103– 107, May 1999. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

http://dlib.computer.org/
co/books/co1999/pdf/r5103.
pdf.

Kroeker:1999:NTNa

[Kro99b] Kirk L. Kroeker. New tools: Net development: Sun’s Java embedded server; MetaCreation’s Web-savvy graphics tool; WebCompiler’s HTML packaging tool. component technology: ProtoVIew Development’s ActiveX tools; Avilon Software’s load balancing component system. software development: Verilog’s test checker; Red Hat and Metrowerks’ development tools for Linux; the Object Factory’s optimization tool; Acumen Systems’s SDK for imaging; Aonix’s process-oriented lifecycle environment; Baan’s embedded software development suite. Computer, 32(5):103– 107, May 1999. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

                                        http://dlib.computer.org/
                                        co/books/co1999/pdf/r5103.
                                        pdf.

Kroeker:1999:NTSf

[Kro99c] Kirk L. Kroeker. New tools: Software development: Monash University’s free Java development environment; Para-Soft’s error prevention tool; Technosoft’s C++ development kit; VideoSoft’s spell-checking and thesaurus controls. network development: Network Security’s security management tool; Fluke’s network inspector; MetaCreation’s Weboriented painting tool. Computer, 32(11):120–122, November 1999. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

                                        http://dlib.computer.org/
                                        co/books/co1999/pdf/ry120.

pdf. See correction [Ano00].

Kroeker:1999:NTSd

[Kro99d] Kirk L. Kroeker. New tools: Software development: Premia’s editing system; Diab Data’s compiler suites for Coldfire; Borland’s visual development tools for Java 2; Sun outlines Java 2 at JavaOne. net development: InfoAccess’ HTML production tool; GoAhead Software’s free embedded Web server; Planetweb and Paradise Innovations produce next-gen set-top box. Computer, 32(8):91–93, August 1999. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

http://dlib.computer.org/
co/books/co1999/pdf/r8091.
pdf.

Kroeker:1999:NTSe

[Kro99e] Kirk L. Kroeker. New tools:

Software development: Soft-Quest Systems’ project management software; Metrowerks’ Java development tool for Mac OS; CAD-UL’s embeddedcentric C/C++ toolkit; KL Group’s Java components. Computer, 32(9):106–107, September 1999. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

http://dlib.computer.org/
co/books/co1999/pdf/r9106.
pdf.

Kruger:1997:JLA

[Krü97a]  Guido Krüger.  Java 1.1

lernen: Anfangen, anwenden, verstehen. Addison-Wes-ley Verlag, Bonn, Germany, 1997. ISBN 3-8273-1299-X. 595 pp. LCCN ????

60 DM. URL http://www.
addison-wesley.de/katalog/
item.ppml?id=00251.

Kruger:1997:JPS

[Krü97b] Guido Krüger. Java-Programmierung mit Symantec Visual Café. Addison-Wesley, Reading, MA, USA, 1997. ISBN 3-8273-1141-1 (??invalid ISBN??). 556 pp. LCCN ???? 70 DM. URL

                                        http://www.addison-wesley.

de/. Includes CD-ROM.

Kruger:1998:GJ

[Kru98] Guido Kruger. Go to Java 1.2. Addison-Wesley Verlag, Bonn, Germany, August 1998. ISBN 3-8273-1370-8 (??invalid ISBN??). LCCN ???? 80 DM.

Kruger:19xx:JLA

[Krüxxa] Guido Krüger. Java 1.1 lernen: Anfangen, anwenden, verstehen. Addison-Wesley Verlag, Bonn, Germany, 19xx. ISBN 3-8273-1299-X (??invalid ISBN??). LCCN ????

                                        60 DM. URL http://www.
                                        addison-wesley.de/katalog/
                                        item.ppml?id=00251.

Kruger:19xx:JPS

[Krüxxb] Guido Krüger. Java-Programmierung mit Symantec Visual Café. Addison-Wesley, Reading, MA, USA, 19xx. ISBN 3-8273-1141-1 (??invalid ISBN??). LCCN ???? 70 DM.

Knappe:1996:CSA

[KS96a] Heiko Knappe and Till S. Schwalm. Chip Special/ Anwenderpraxis: HTML und Java. Vogel, Würzburg, Germany, 1996. ISBN 3-8259-1382-1 (??invalid checksum??). 138 pp. LCCN ???? 34

                                        DM. URL http://medweb.
                                        uni-muenster.de/zbm/liti/
                                        design.html#473.  Includes

CD-ROM.

Knappe:1996:HJH

[KS96b] Heiko Knappe and Till S.

Schwalm. HTML und Java: HTML-Crashkurs mit vielen Tips, Applets zum Einbinden in eigene HTML-Seiten, Programmierung mit Java, grosser Referenzteil, Java-Zauberei. Vogel, Würzburg, Germany, 1996. ISBN 3-8259-1382-1. 138 pp. LCCN ???? 34 DM. URL http:

//medweb.uni-muenster.de/
zbm/liti/design.html#473.

Includes CD-ROM.

Koch-Steinheimer:1997:JPA

[KS97] Peter Koch-Steinheimer. Java:

Das Programmieren von Agenten und Kaffeemaschinen. Harri Deutsch Verlag, Thun, Switzerland, 1997. ISBN 3-8171-1500-8. xix + 397 pp. LCCN ???? 68 DM. URL http:

//www.germany.eu.net/shop/
HD/verlag/index.htm.  In-

cludes CD-ROM.

Kappel:1998:DLP

[KS98a] G. Kappel and B. Schroeder.

Distributed light-weight persistence in Java — a tour on RMIand CORBA-based solutions. Lecture Notes in Computer Science, 1460:411–??, 1998. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kappel:1998:DLW

[KS98b] G. Kappel and B. Schroeder.

Distributed light-weight persistence in Java — a tour on RMIand CORBA-based solutions.

Lecture Notes in Computer Science, 1460:411–??, 1998. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaster:1997:DID

[KSB97] John Kaster, Loren Scott, and Ted Blue. Delphi Internet Development: Utilizing ActiveX, ISAPI, Java and VRML for Internet Solutions. McGraw-Hill, New York, NY, USA, July 1997. ISBN 0-07-913254-5 (softcover). ???? pp. LCCN ???? US$50; US$44.95. Includes CD-ROM.

Kikinis:1996:DBA

[KSI^{+}96] Ron Kikinis, Martha E. Shenton, Dan V. Iosifescu, Robert W. McCarley, Pairash Saiviroonporn, Hiroto H. Hokama, Andre Robatino, David Metcalf, Cynthia G. Wible, Chiara M. Portas, Robert M. Donnino, and Ferenc A. Jolesz. A digital brain atlas for surgical planning, model-driven segmentation, and teaching. IEEE transactions on visualization and computer graphics, 2(3):232–241, September 1996. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1996/pdf/v0232.
pdf;   http://www.computer.
org/tvcg/tg1996/v0232abs.

htm. Illustrates the use of Java for display and visualization of medical images.

Knobloch:1997:EJ

[KSW97] Frank Knobloch and M. Seeboerger-Weichselbaum. Das Einsteigerseminar Java. BHV, ??, Germany, 1997. ISBN 3-89360-925-3. 350 pp. LCCN ???? 20

DM. URL http://www.ctn-
service.com/b-treff/bookshop/
bhv/buch/deinst.htm#Java.

Kluit:1998:VPJ

[KSW98] Peter G. Kluit, Marleen Sint, and Frank Wester. Visual programming with Java: evaluation of an introductory programming course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):143–147, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kero:1996:ISD

[KT96] B. Kero and S. Tsur. The IQ system: a deductive database information lens for reasoning about textual information. In Pedreschi and Zaniolo [PZ96], page ?? ISBN 3-540-61814-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 I5857 1996.

Krall:1999:GCL

[KT99] A. Krall and P. Tomsich.

Garbage collection for large memory Java applications. Lecture Notes in Computer Science, 1593:895–??, 1999. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kuhnel:1996:JF

[Küh96] Ralf Kühnel. Die Java Fibel: Programmierung interaktiver Homepages für das World Wide Web. Addison-Wesley, Reading, MA, USA, 1996. ISBN 3-8273-1024-X (??Invalid checksum??). 224 pp. LCCN ???? 59,90 DM. URL http:/

/www.addison-wesley.de/;
http://www.addison-wesley.
de/KD/NE/1024.html. Includes

CD-ROM.

Kuhnel:1997:JFP

[Küh97] Ralf Kühnel. Die Java 1.1 Fibel: Programmierung von Threads und Applets, mit den neuen Bibliotheken: Beans, Security, RMI, IDL, SQL. Addison-Wes-ley, Reading, MA, USA, second edition, 1997. ISBN 3-8273-1233-7. 352 (or 333??) pp. LCCN ???? 49,90 DM.

URL http://www.addison-
wesley.de/;   http://www.
addison-wesley.de/katalog/
item.ppml?textexpr=Java&id=
00212;   http://www.addison-
wesley.de/KD/NE/1024.html.

Includes CD-ROM.

Kuhnel:19xx:JF

[Kühxx] Ralf Kühnel. Die Java-Fibel. Addison-Wesley, Reading, MA, USA, 19xx. ISBN 3-8273-1024-X (??invalid ISBN??). LCCN ???? 60 DM.

Kukuk:1998:NPR

[Kuk98] Amy Kukuk. New products: Raritan MasterConsole MX4, Raritan Computer, Inc.; VRtuoso, Bittco Solutions; Debian GNU/Linux ARM, Debian GNU/Linux; LinuxCAD, Software Forge, Inc.; Perspective for Java, Three D Graphics; MetaCard 2.2, MetaCard Corp.; EtherPage Version 3.0, Personal Productivity Tools, Inc.; D 3 Linux v.7.1, Pick Systems. Linux Journal, 56:80, 95, December 1998. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Kumhyr:1998:DIC

[Kum98] David B. Kumhyr. Developing an international calendar in Java. In Unicode Consortium [Uni98b], page ?? ISBN ????

LCCN ???? URL http://www.
unicode.org/unicode/iuc13/
program.html.

Kalakota:1996:IIW

[KW96] Ravi Kalakota and Andrew Whinston. Intranets — internal Web servers that have Java applets could transform the client/server world–and even replace SAP applications, professors Ravi Kalakota and Andrew Whinston predict. Computerworld, 30(11):37–??, ???? 1996. CODEN CMPWAB. ISSN 0010-4841.

Koffman:1998:PSJ

[KW98] Elliot B. Koffman and Ursula Wolz. Problem Solving with Java. Addison-Wesley, Reading, MA, USA, August 1998. ISBN 0-201-35743-7. LCCN ???? ????

URL http://cseng.aw.com/

                                        bookdetail.qry?ISBN=0-201-
                                        35743-7&ptype=0.

Koffman:1999:CUJ

[KW99] Elliot Koffman and Ursula Wolz. CS1 using Java language features gently. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(3):40–43, September 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Koosis:19xx:JPDc

[KWxx] Donald Koosis and Wimmershoff. Java-Programmierung für Dummies. International Thomson Computer Press, 20 Park Plaza Suite 1001, Boston, MA 02116 USA, 19xx. ISBN 3-8266-2732-6 (??invalid ISBN??). LCCN ???? 50 DM. URL http:

                                        //www.itp.de/dummies/2732/
                                        2732.html.

Krishnaswamy:1998:EIJ

[KWB^{+}98] Vijaykumar Krishnaswamy, Dan Walther, Sumeer Bhola, Ethendranath Bommaiah, George Riley, Brad Topol, and Mustaque Ahamad. Efficient implementation of Java Remote Method Invocation (RMI). In USENIX [USE98b], page ?? ISBN ???? LCCN ???? URL

                                        http://www.usenix.org/publications/
                                        library/proceedings/coots98/
                                        krishnaswamy.html.

Lien:19xx:JV

[L^{+}xx] Shock Lien et al. Java/VRML. Acer TWP Corp., ????, 19xx.

ISBN 957-23-0429-1. ???? pp. LCCN ???? 450 NT$.

Leinecker:1997:VJBa

[LA97a] Richard C. Leinecker and Tom Archer. Visual J++ Bible. IDG Books, San Mateo, CA, USA, 1997. ISBN 0-7645-8021-3 (paperback). xxix + 896 pp. LCCN ???? US$50.

URL http://www.idgbooks.
com/database/book_result.
msql?uid=uidg323&isbn=0-7645-
8021-3.

Leinecker:1997:VJBb

[LA97b] Richard C. Leinecker and Tom Archer. Visual J++ Bible. Flag Publishing, ??, Taiwan, February 1997. ISBN 0-7645-8021-3. LCCN QA76.73.J38 L434 1997. URL http://www.

flag.com.tw/new_book/book/

F8908.htm. With Michael Farmer and Jerry Muelver. Includes CD-ROM.

Laberis:1995:EHJ

[Lab95] Bill Laberis. Editorial: Hot Java! Computerworld, 29 (50):36, December 11, 1995. CODEN CMPWAB. ISSN 0010-4841. URL http://www.

computerworld.com/.

Ladd:1996:UHJb

[Lad96]  Eric Ladd.  Using HTML,

Java and CGI, Platinum Edition. Macmillan Computer Publishing, Indianapolis, IN, USA, November 1996. ISBN 0-614-20289-2. ???? pp. LCCN ???? US$69.99.

Ladd:1997:AVJ

[Lad97] Scott Robert Ladd. Active Visual J++ — Application Development with Java, ActiveX & Components. Microsoft Press, Bellevue, WA, USA, August 1997. ISBN 1-57231-609-8. vii + 331 pp. LCCN QA76.76.J38L33 1997.

                                        US$40. URL http://mspress.
                                        microsoft.com/mspress/Books/
                                        Des/1368.HTM.

Ladd:1998:AVJ

[Lad98a]  Scott Robert Ladd.  Ac-

tive Visual J++ — Web-Programmierung mit ActiveX und Java. Microsoft Press Germany, ????, March 1998. ISBN 3-86063-393-7. LCCN ???? 80 DM. URL http://

                                        194.120.227.99/prd.i/pgen/
                                        press/FT16VFE47AS12P2E008MJ7W11XNLKLWL/
                                        product.html;   isbn=3-86063-
                                        393-7.

Ladd:1998:JA

[Lad98b] Scott Robert Ladd. Java Algorithms. Osborne/McGraw-Hill, Berkeley, CA, USA, January 1998. ISBN 0-07-913696-6. ???? pp. LCCN QA76.73.J38L33 1997. US$45. URL http:

                                        //mcgraw-hill.inforonics.
                                        com/cgi/getarec?mgh42092%
                                        comp.

Ladd:19xx:AVJ

[Ladxxa] Scott Robert Ladd. Active Visual J++. Mondadori Informatica, ??, Italy, 19xx. ISBN 88-7131-846-3. LCCN ???? 80,000 L. URL http:

//education.mondadori.com/
books/SchBook.asp?Libro=203.

Ladd:19xx:AMA

[Ladxxb]  Scott Robert Ladd.  Ate-

lier Microsoft ActiveX — Visual J++. Microsoft Press France, ????, 19xx. ISBN 2-84082-282-2. LCCN ???? 299 FF. URL http://

www.microsoft.com/france/
mspress/ouvrage.asp?OuvrageID=
171.

Laffra:1996:AJI

[Laf96] Chris Laffra. Advanced Java:

Idioms, Pitfalls, Styles and Programming Tips. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, September 12, 1996. ISBN 0-13-534348-8. xv + 270 pp. LCCN QA76.73.J38L34 1997. US$35.95. URL http:

//members.aol.com/laffra/
book.html.

Lafore:1998:DSA

[Laf98a] Robert Lafore. Data Structures and Algorithms in Java. Mitch Waite Signature Series. Waite Group Press, Corte Madera, CA, USA, March 1998. ISBN 1-57169-095-6. xxi + 625 pp. LCCN QA76.9.D35L34 1998. US$50. URL http://

merchant.superlibrary.com:
8000/catalog/mcp/PRODUCT/
PAGE/15716/bud/1571690956.
html.

Lafore:1998:GAJ

[Laf98b] Robert Lafore. Gegevensstructuren, Algortihmen, Java. Academic Service, Postbus 81, 2870 AB Schoonhoven, The Netherlands, April 1998. ISBN 90-395-0927-1. LCCN ???? fl.

                                        85. URL http://acaserv.nl/
                                        bestel/5_0927.htm.

Laffra:19xx:AJIb

[Lafxxa] Chris Laffra. Advanced Java: Idioms, Styles, Programming Tips, and Pitfalls (Japanese tranlsation). ASCII Corporation, Tokyo, Japan, 19xx. ISBN 4-7561-1399-0. ???? pp. LCCN ???? 2800 yen. URL http:/

                                        /www.ascii.co.jp/pb/book1/
                                        shinkan/detail/1322412.html.

Laffra:19xx:CCJ

[Lafxxb] Chris Laffra. C2J, a C++ to Java translator. World-wide web document, IBM T.J. Watson Research Center, P.O. Box 704, Yorktown Heights, NY 10598. USA, 19xx. URL http:

                                        //members.aol.com/laffra/
                                        c2j.html.

Lai:19xx:UNU

[Laixx] Michel Lai. UML. La notation unifiée de modélisation objet. Applications en Java. Masson Editeur, Masson, France, 19xx. ISBN 2-7296-0659-9. ???? pp. LCCN ???? 295 FF.

                                        URL http://www.masson.fr/
                                        cgi-bin/bookf.pl?2:2:'tx=(
                                        java*)'.

Lakos:1996:LCS

[Lak96] J. Lakos. Large-scale C++ software design: physical hierarchy. 2 effective testing, improved maintainability, independent reuse. C++ Report, 8 (8):55–64, September 1996. CO-DEN CRPTE7. ISSN 1040-6042.

LaMonica:1995:NCC

[LaM95a]  Martin LaMonica.  News:

Counting the component change. InfoWorld, page ??, November 14, 1995. CODEN INWODU. ISSN 0199-6649. URL http://

www.infoworld.com/.

LaMonica:1995:NIW

[LaM95b] Martin LaMonica. News: IBM to woo Windows developers for OpenDoc. InfoWorld, page ??, November 04, 1995. CODEN INWODU. ISSN 0199-6649.

URL http://www.infoworld.
com/.

LaMonica:1995:NND

[LaM95c] Martin LaMonica. News: Neuron Data, SunSoft take different paths to Web. InfoWorld, page ??, December 16, 1995. CODEN INWODU. ISSN 0199-6649. URL http://www.

infoworld.com/.

LaMonica:1995:NOE

[LaM95d] Martin LaMonica. News: Openscape emulates, augments Visual Basic. InfoWorld, page ??, December 03, 1995. CODEN INWODU. ISSN 0199-6649.

URL http://www.infoworld.
com/.

LaMonica:1995:NSN

[LaM95e] Martin LaMonica. News: Sun-Soft’s Neo object framework embraces Internet. InfoWorld, page ??, September 23, 1995. CODEN INWODU. ISSN 0199-6649. URL http://www.

                                        infoworld.com/.

LaMonica:1995:NSH

[LaM95f]  Martin LaMonica.  News:

Symantec hot to deliver Espresso. InfoWorld, page ??, December 16, 1995. CODEN INWODU. ISSN 0199-6649. URL http://

                                        www.infoworld.com/.

LaMonica:1995:NTV

[LaM95g] Martin LaMonica. News: Tool vendors tie Web to client/server apps. InfoWorld, page ??, December 10, 1995. CODEN IN-WODU. ISSN 0199-6649. URL

                                        http://www.infoworld.com/.

Lambert:1999:VSC

[Lam99] John Donovan Lambert. VB-SCRIPT and SQL calendars. Dr. Dobb’s Journal of Software Tools, 24(5):40, 42, 46–48, May 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

                                        ddj.com/ftp/1999/1999_05/
                                        vbsql.txt;   http://www.ddj.
                                        com/ftp/1999/1999_05/vbsql.
                                        zip.

Lange:1998:ITA

[Lan98] L. Lange. The Internet [technology 1998 analysis and forecast]. IEEE Spectrum, 35(1): 37–42, January 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lange:1999:ITA

[Lan99a] L. Lange. The Internet [technology 1999 analysis and forecast]. IEEE Spectrum, 36(1): 35–40, January 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Langr:1999:EJS

[Lan99b] Jeff Langr. Essential Java Style:

Patterns for Implementation. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1999. ISBN 0-13-085086-1. ???? pp. LCCN QA76.13.J38L357 1999.

URL http://www.phptr.com/
ptrbooks/ptr_0130850861.html.

Laurel:1996:J

[Lau96] Chris Laurel. 3-D Java. New Riders Publishing, Carmel, IN, USA, July 1, 1996. ISBN 1-56205-597-6. xxviii + 383 pp. LCCN QA76.73.J38F73 1996.

US$40. URL http://www.mcp.
com/nrp/.

Lavallee:1996:SLB

[Lav96] Paul C. Lavallee. Shedding light on Broadway: While it reminds people of Java, the X follow-on has a set of powerful capabilities in its own right. The X Journal, 5(5):32–34, 36, May 1996. CODEN XJOUEA. ISSN 1056-7003.

Lawton:1998:NBS

[Law98] George Lawton. News briefs:

Sun joins tetherless network fray; private doorbells: Solution to encryption controversy; Web extensions promise distributed computing; Wall Street conducts major Y2K test; Sun prepares to submit first Java standard; Netscape loses more ground in browser war. Computer, 31(10):17–19, October 1998. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

http://dlib.computer.org/
co/books/co1998/pdf/rx017.
pdf.

Lawton:1999:NBB

[Law99] George Lawton. News briefs: Battle begins for device network supremacy; VXML Forum is created; Advanced Internet Center opens. Computer, 32(6): 20–21, June 1999. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

http://dlib.computer.org/
co/books/co1999/pdf/r6020.

pdf. Discusses Sun/Sony/Philips Jini/HAVi vs. Microsoft Universal Plug and Play (UPnP) vs. Lucent’s InfernoSpaces.

Lazarro:1997:UHD

[Laz97] Joseph J. Lazarro. Unix helps the disabled. BYTE Magazine, 22(4):51–52, April 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic). Discusses speech output, Braille output, screen readers, and voice recognition tools for UNIX, some of them freely available.

Lazzaro:1999:HWH

[Laz99] J. J. Lazzaro. Helping the Web help the disabled. IEEE Spectrum, 36(3):54–59, March 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Loy:1996:JPI

[LB96] Marc Loy and Thomas J. Berry.

Java Programming for the Internet. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1996. ISBN 0-13-270778-0. LCCN QA76.73.J38 1997. US$69.95. URL http://www.

prenhall.com/allbooks/ptr_
0132707780.html. CD-ROM

only; no printed book.

Lewis:1997:JPS

[LB97a] Geoffrey Lewis and Steven Barber. Joe Programming Sourcebook: Developing Web Applications With Java and CORBA. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, July 1997. ISBN 0-471-18147-1 (softcover). ???? pp. LCCN QA76.625.L48 1997. US$26.99. Publication cancelled.

Loy:1997:CJD

[LB97b] Marc Loy and Thomas J.

Berry. A Complete Java Database Training Course. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1997. ISBN 0-13-759507-7. 32 pp. LCCN QA76.73.J38¡1997 00439. US$70. URL http://

www.prenhall.com/allbooks/
ptr_0137595077.html;   http:
//www.prenhall.com/ptrbooks/
ptr_0137595077.html.  CD

ROM only; no printed book.

Li:1998:DVU

[LB98a] L. Li and A. Barnes. Data visualization using Java and VRML. In Anonymous [Ano98-34], pages 911–916. ISBN 1-58025-149-8. LCCN ????

Liang:1998:DCL

[LB98b] Sheng Liang and Gilad Bracha. Dynamic class loading in the Java Virtual Machine. ACM SIGPLAN Notices, 33(10):36– 44, October 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lewis:1998:PJI

[LBS98] Geoffrey Lewis, Steven Barber, and Ellen Siegel. Programming with Java IDL: Developing Web Applications with Java and CORBA. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 1998. ISBN 0-471-24797-9. xiv + 322 pp. LCCN QA76.625.L483 1998. US$35. URL http:

                                        //www.wiley.com/compbooks/
                                        catalog/24797-9.htm.

Lemay:1998:TYJb

[LC98] Laura Lemay and Rogers Cadenhead. Teach Yourself Java 1.2 in 21 Days. Howard W. Sams, Indianapolis, IN 46268, USA, third edition, 1998. ISBN 1-57521-390-7. LCCN QA76.73.J38 L448 1998. Revised edition, with same ISBN, of Teach Yourself Java 1.1 in 21 Days.

Lemay:1999:STY

[LC99] Laura Lemay and Rogers Cadenhead. SAMS teach yourself Java 2 platform in 21 days, Professional reference edition. Howard W. Sams, Indianapolis, IN 46268, USA, 1999. ISBN 0-672-31438-X. xix + 948 pp. LCCN QA76.73.J38 L4493 1999. US$49.99.

Luis:19xx:JLP

[LCxx] José Luis and Otros Cortés.

Java: El Lenguaje de Programación de Internet. Editorial Marcombo, S.A., Barcelona, Spain, 19xx. ISBN ???? ???? pp. LCCN ????

Lu:1997:WBD

[LCLL97] Tain-Chi Lu, Chung-Wen Chiang, Chungnan Lee, and Tong-Yee Lee. A Web-based distributed and collaborative 3D animation environment. Concurrency: Practice and Experience, 9(11):1261–1268, November 1997. CODEN CPEXEI. ISSN 1040-3108. URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=13840;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=13840&PLACEBO=IE.pdf. Spe-

cial Issue: Java for computational science and engineering — simulation and modeling II.

Lea:1996:CPJ

[Lea96a] Douglas Lea. Concurrent Programming in Java: Design Principles and Patterns. Addison/Wesley Java series Java series. Addison-Wesley, Reading, MA, USA, November 1, 1996. ISBN 0-201-69581-2 (paperback). xii + 339 pp. LCCN QA76.73.J38L4 1996. US$39.76. URL http://www.

                                        aw.com/cseng/titles/0-201-
                                        69581-2/.

Leach:1996:RHJ

[Lea96b] George W. Leach. Review: Hooked on Java. ;login: the USENIX Association newsletter, 21(3):55–??, June 1996. CODEN LOGNEM. ISSN 1044-6397.

Leach:1996:RJP

[Lea96c] George W. Leach. Review: The Java programming language. ;login: the USENIX Association newsletter, 21(5):46– ??, October 1996. CODEN LOGNEM. ISSN 1044-6397.

Lea:1997:DOS

[Lea97a] D. Lea. Design for open systems in java. Lecture Notes in Computer Science, 1282:32–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lea:1997:CPJb

[Lea97b] Doug Lea. Concurrent Programming in Java. Entwurfsprinzipien und Muster. Addison-Wesley Verlag, Bonn, Germany, July 1997. ISBN 3-8273-1243-4 (??invalid ISBN??). LCCN ???? 70 DM. URL

                                        http://www.addison-wesley.
                                        de/katalog/item.ppml?id=00219.

Leach:1997:BRJ

[Lea97c] George W. Leach. Book review: Java Security: Hostile Applets, Holes, and Antidotes. ;login: the USENIX Association newsletter, 22(5):??, October 1997. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

usenix.org/publications/login/
1997-10/mcgraw.html.

Lea:1998:CPJ

[Lea98a] Doug Lea. Concurrent Programming in Java. Addison-Wesley, Reading, MA, USA, second edition, December 1998. ISBN 0-201-31009-0 (softcover). 448 pp. LCCN ????

Leach:1998:BRB

[Lea98b] Michael R. Leach. Book review:

Java in a nutshell: a desktop quick reference. Journal of the American Society for Information Science, 49(14):1329–1330, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lear:1999:NBI

[Lea99] Anne C. Lear. News briefs:

IETF rejects net wiretap proposal; new plans ease net congestion; standard API to simplify network chip development; group releases realtime Java specification. Computer, 32(12):20–21, December 1999. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

http://dlib.computer.org/

                                        co/books/co1999/pdf/rz020.
                                        pdf.

Ledru:1997:APE

[Led97] Pascal Ledru. Adaptive parallelism: an early experiment with Java remote method invocation. Operating Systems Review, 31(4):24–29, October 1997. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ledru:1998:JIL

[Led98a] Pascal Ledru. JSpace: implementation of a Linda system in Java. ACM SIGPLAN Notices, 33(8):48–50, August 1998. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ledru:1998:TSJ

[Led98b] Pascal Ledru. A traveling salesman in Java. ACM SIGPLAN Notices, 33(8):51–56, August 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lehman:1996:H

[Leh96] Mike Lehman. HotTEA. Web page, Cereus Design Corp., 1199 Forest Ave., Suite 264, Pacific Grove, CA 93950, USA, Voice & Fax: 408-626-3016, 1996.

                                        ???? pp. URL http://www.
                                        mbay.net/~cereus7/HotTEA.

html. Implements a simple Basic interpreter in Java.

Leinfuss:1995:FSY

[Lei95] Emily Leinfuss. Feature: Sorting out your Web-database options. InfoWorld, page ??, December 05, 1995. CODEN IN-WODU. ISSN 0199-6649. URL

http://www.infoworld.com/.

Leininger:1997:JDT

[Lei97] Kevin E. Leininger. Java Developer’s Toolkit. McGraw-Hill, New York, NY, USA, 1997. ISBN 0-07-913106-9. xv + 201 pp. LCCN QA76.73.J38L43 1997. US$40. URL http:

//mcgraw-hill.inforonics.
com/cgi/getarec?mgh30378%
comp;   http://www.mcgraw-
hill.com/.

Lemay:1997:AYW

[Lem97a] L. Lemay. Animating your Web page. Web Techniques, 2(2):18– ??, 20–22, February 1997. CO-DEN WETEFA. ISSN 1086-556X.

Lemay:1997:AED

[Lem97b] Laura Lemay. Aprenda em 21 dias Java. Editore Campus Ltda., Rua Sete de Setembro, 111/16, Rio de Janeiro - RJ - 20050-002, Brazil, 1997. ISBN ???? 676 pp. LCCN ???? R$ 79. URL http:

//www.campus.com.br/about-
us.htm;   http://www.campus.
com.br/catalogo/livro.cfm?
cod_livro=20104.

Lemay:1997:AJD

[Lem97c] Laura Lemay. Aprendiendo Java En 21 Dias. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1997. ISBN 968-880-741-9. ???? pp. LCCN ???? $199.00 pesos nuevos; 5616 Ptas.

Lemay:1997:OMGa

[Lem97d] Laura Lemay. Official Marimba Guide to Castanet. Sams.net, Indianapolis, IN 46268, USA, 1997. ISBN 1-57521-255-2. xx + 354 pp. LCCN TK5105.887.L46 1997. US$40. URL http://

merchant.superlibrary.com:
8000/catalog/hg/PRODUCT/PAGE/
15752/bud/1575212552.html.

Lemay:1997:OMGb

[Lem97e] Laura Lemay. Official Marimba Guide to Castanet. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, August 1997. ISBN 4-88735-073-2. ???? pp. LCCN ???? 3,200

yen. URL http://www.mmjp.
or.jp/prentice/wa_int14-j.
html.

Lemay:1997:PJ

[Lem97f] Laura Lemay. Programmeren in Java 1.1. Academic Service, Postbus 81, 2870 AB Schoonhoven, The Netherlands, October 1997. ISBN 90-395-0837-2. ca. 500 pp. LCCN ???? fl. 68,00. URL http:/

/acaserv.nl/bestel/5_0837.

htm. Dutch translation of [LPMG97].

Lemay:1998:TYJa

[Lem98] Laura Lemay. Teach Yourself Java 1.1 in 21 Days. Sams.net, Indianapolis, IN 46268, USA, third edition, June 1998. ISBN 1-57521-390-7. LCCN ???? US$30. URL

http://iq-mcp3.iquest.net/
catalog/mcp/PRODUCT/PAGE/
15752/bud/1575213907.html.

Lemay:19xx:AED

[Lemxxa] Laura Lemay. Aprenda em 21 Dias Java. Editore Campus Ltda., Rua Sete de Setembro, 111/16, Rio de Janeiro - RJ - 20050-002, Brazil, 19xx. ISBN ???? ???? pp. LCCN ???? R$79.

Lemay:19xx:JTb

[Lemxxb] Laura Lemay. Java 1.1 in 21 Tagen. Markt & Technik, Haar bei München, Germany, 19xx. ISBN 3-8272-2009-2 (??invalid ISBN??). LCCN ????

80 DM. URL http://www.
mut.ch/html/buchtext.htx?
&x_pp=STb4XDV284xv2gpf&x_
pk=Direkt&x_pa=&x_pr=Fr&x_
artikelnr=22009.

Lemay:19xx:JL

[Lemxxc] Laura Lemay. JavaScript Labo.

Simon & Schuster Macmillan (France), Paris, France, 19xx. ISBN ???? ???? pp. LCCN ????

URL http://www.ssm.fr/.

Lemay:19xx:MCG

[Lemxxd]  Laura Lemay.  Marimba

Castanet La guida ufficiale. Apogeo srl, Viale Papiniano 38, 20123 Milano, Italy, 19xx. ISBN 88-7303-352-0. ???? pp. LCCN ????

46,000 L. URL http://www.
apogeonline.com/catalogo/

                                        352.html;   http://www.urra.
                                        it/352.html.

Leonard:1995:JJJ

[Leo95] Andrew Leonard. Java Java Java: Is it the future or the end? Web Review, page ??, November 22, 1995. URL http:/

                                        /www.gnn.com/gnn/wr/nov22/
                                        features/java/index.html.

Leong:1998:UJI

[Leo98] Kok Yong Leong. Unicode, Java and input methods. In Unicode Consortium [Uni98a], page ?? ISBN ???? LCCN

                                        ????  URL http://www.
                                        unicode.org/unicode/iuc12/
                                        program.html.

Lester:1997:DPV

[Les97] Christopher S. Lester. Database Programming with Visual J++. Que Corporation, Indianapolis, IN, USA, 1997. ISBN 0-7897-1015-3. xxii + 688 pp. LCCN ???? US$50. URL http:

                                        //dev.mcp.com:8001/about/
                                        coinfo/imprints/que/books/
                                        descriptions/10153d.html.

Lester:19xx:BDC

[Lesxx] Christopher S. Lester. Bases de Datos con Visual J++ Edicion Especial. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 19xx. ISBN 84-8322-058-X. LCCN ???? ????

Levine:1996:LJD

[Lev96a]  David Levine.  Live Java:

Database to Cyberspace. AP Professional, Boston, MA, USA, August 1996. ISBN 0-12-445485-2. xxi + 323 pp. LCCN QA76.73.J38L48 1996. US$24.95. URL http://www.

apcatalog.com/cgi-bin/AP?
ISBN=0124454852&LOCATION=
US&FORM=FORM2.

Levitt:1996:MN

[Lev96b] J. Levitt. The making of an NC. Information Week, ???? (606):64, 68, 72, ???? 18, 1996. CODEN INFWE4. ISSN 8750-6874.

Levitt:1996:NWC

[Lev96c] J. Levitt. The new Web clients [browser comparison]. Information Week, 576(576):73–74, 76, 78, April 22, 1996. CODEN IN-FWE4. ISSN 8750-6874.

Levitt:1996:SSL

[Lev96d] J. Levitt. Shedding some light on NCs. Information Week, ????(606):57–58, 62, 64, ???? 18, 1996. CODEN INFWE4. ISSN 8750-6874.

Levia:1999:IEP

[Lev99] Oz Levia. Integrated engineering: Programming system architectures with Java. Computer, 32(8):96–98, August 1999. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

http://dlib.computer.org/
co/books/co1999/pdf/r8096.
pdf.

Lewis:1996:BCB

[Lew96a] Ted Lewis. Binary critic: The big software chill. Computer, 29(3):12–14, March 1996. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lewis:1996:BCW

[Lew96b] Ted Lewis. Binary critic: Will tiny beans conquer the world again? Computer, 29(9):13–14, September 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). Discusses Java and its component technology, Java Beans.

Lewis:1997:SSUa

[Lew97a] John Lewis. Software Systems Using Java Preliminary Version. Addison-Wesley Longman, Reading, MA, USA, January 1997. ISBN 0-201-69588-X. ???? pp. LCCN ????

Lewis:1997:BCI

[Lew97b] Ted Lewis. Binary critic: If Java is the answer, what was the question? Computer, 30(3): 136, 133–135, March 1997. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). See responses [Ols97, Wel97].

Lewis:1998:BCJa

[Lew98] Ted Lewis. Binary critic: Java Holy War ’98. Computer, 31 (3):128, 126–127, March 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://

                                        pdf.computer.org/co/books/
                                        co1998/pdf/r3128.pdf.

Lechner:1997:JER

[LG97] E. Lechner and S. A. Guccione.

The Java environment for reconfigurable computing. Lecture Notes in Computer Science, 1304:284–??, 1997. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Larman:1999:JPI

[LG99] Craig Larman and Rhett Guthrie. Java 2 Performance and Idiom Guide. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1999. ISBN 0-13-014260-3. xvii + 299 pp. LCCN QA76.73.J38 L359 2000. URL http:

//www.phptr.com/ptrbooks/
ptr_0130142603.html.

Levin:1996:SSN

[LH96] Rich Levin and Mary Hayes.

Sun’s shiny new object: Meet Joe, a universal messaging widget that eases development of globally distributed apps. the Web will never be the same. Information Week, 572:14–15, March 25, 1996. CODEN IN-FWE4. ISSN 8750-6874. From the article: “Sun’s product, named Joe, will be a universal messaging widget that will make developing globally distributed apps as easy as adding three lines of Java code.”.

Lea:1997:DPC

[LH97] Doug Lea and David Holmes.

Design patterns for concurrent programming: a Java example. Tutorial session at [USE97b].,

                                        1997. URL http://dxsting.
                                        cern.ch/sting/COOTS97.html.

Li:1998:JDS

[Li98] Liwu Li. Java, Data Structures and Programming. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., March 1998. ISBN 3-540-63763-X. 440 pp. LCCN ???? US$49.95.

Liang:1999:JNI

[Lia99] Sheng Liang. Java Native Interface: Programmer’s Guide and Specification. Addison-Wes-ley, Reading, MA, USA, 1999. ISBN 0-201-32577-2. xiv + 303 pp. LCCN QA76.38 .L53 1999. US$39.95.

Liang:19xx:IJP

[Liaxx] Daniel Liang. Introduction to Java Programming. Que E&T, ????, 19xx. ISBN 1-57576-548-9, 1-57576-549-7. ???? pp. LCCN ???? ????

Liederman:1995:JDE

[Lie95a] Erica Liederman. ‘Java Day’ event: You are there. Sunworld Online, 11:??, 1995. ISSN 1091-

                                        8914. URL http://www.sun.
                                        com/sunworldonline/swol-11-
                                        1995/swol-11-java.event.html.

Liederman:1995:JDY

[Lie95b] Erica Liederman. Java day: You are there. SunWorld Online, November 1995. URL http://

                                        www.Sun.COM/sunworldonline/
                                        swol-11-1995/swol-11-java.
                                        event.html.

Liederman:1995:MLJ

[Lie95c] Erica Liederman. Microsoft licenses Java. Sunworld Online, 12:??, December 1995. ISSN 1091-8914. URL http://www.

sun.com/sunworldonline/swol-
12-1995/swol-12-microsoft.
html.

Liederman:1996:IT

[Lie96a] Erica Liederman. The Internet toaster. Sunworld Online, 01:??, 1996. ISSN 1091-8914.

URL http://www.sun.com/
sunworldonline/swol-01-1996/
swol-01-appliance.html.

Liederman:1996:STC

[Lie96b] Erica Liederman. Sun’s tempest in the coffee pot. Sunworld Online, 02:??, 1996. ISSN 1091-

8914. URL http://www.sun.
com/sunworldonline/swol-02-
1996/swol-02-javacup.html.

Lim:1995:BDE

[Lim95] W. Lim. The BALI development environment for small mobile robots. Proceedings of the SPIE — The International Society for Optical Engineering, 2591:55–65, ???? 1995. CO-DEN PSISDG. ISSN 0277-786X (print), 1996-756X (electronic).

Linkins:1996:EJH

[Lin96a] Kim Fulcher Linkins. Enjoying Java? how about some Joe? Sun Observer, 10(5):1, 7, May 1996. ISSN 1058-5400.

Linkins:1996:JWP

[Lin96b] Kim Fulcher Linkins. Java Workshop provides Web-based development environment. Sun Observer, 10(5):1, 6, May 1996. ISSN 1058-5400.

Linthicum:1996:DLL

[Lin96c] D. S. Linthicum. DCE lightens its load. DBMS, 9(7):24, 26, 28, June 1996. CODEN DBMSEO. ISSN 1041-5173.

Linthicum:1996:HCJ

[Lin96d] D. S. Linthicum. Here come the Java tools. DBMS, 9(9):24, 26, 28, August 1996. CODEN DBMSEO. ISSN 1041-5173.

Linthicum:1996:ADT

[Lin96e] David S. Linthicum. App dev tackles the intranet. Datamation, 42(15):113–??, September 1, 1996. CODEN DTMNAT. ISSN 0011-6963.

Linthicum:1996:CD

[Lin96f] David S. Linthicum. C/S developer. DBMS, 9(9):24–??, August 1, 1996. CODEN DBM-SEO. ISSN 1041-5173.

Linke:1997:VJV

[Lin97] Thomas Linke. Visual J++ Version 1.1 Schnellübersicht. Markt&Technik, ????, May 1997. ISBN 3-8272-5257-1. ???? pp. LCCN ????

                                        40 DM. URL http://www.
                                        mut.ch/html/buchinfo.htx?
                                        &x_pp=dJgxRqA9wlAYNijX&x_
                                        az=10&x_pk=VpLvjnsx7b5V&x_
                                        pa=B&x_pr=oS&x_bestell=25257&

x_stichwort=Linke&x_sprache=
A&x_start=1&x_gefunden=2&
x_query=&x_lastpage=buchfund&
x_soft=&x_suchewo=A.

Lindenberg:1998:JIM

[Lin98a] Norbert Lindenberg. Java input method framework. In Unicode Consortium [Uni98b], page ?? ISBN ???? LCCN

????  URL http://www.
unicode.org/unicode/iuc13/
program.html.

Linthicum:1998:JES

[Lin98b] David S. Linthicum. Java evolves: With solid standards and tools, Java is becoming important in enterprise and embedded development. BYTE Magazine, 23(1):60, January 1998. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Lipton:1999:JPD

[Lip99] Paul Lipton. Java proxies for database objects. Dr. Dobb’s Journal of Software Tools, 24 (5):34, 36–39, May 1999. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/1999/1999_05/
jproxy.txt.

Littleford:1997:JCS

[Lit97] Alan Littleford. Java Commerce and Security Programming. Coriolis Group Books, Scottsdale, AZ, USA, February 1997. ISBN 1-57610-101-0. 500 pp. LCCN ???? US$39.99, CDN$55.99. URL http://www.

coriolis.com/.

Liu:1996:FCP

[Liu96] Dianbin Liu. Future of computing in the petroleum industry. Proceedings — Petroleum Computer Conference, pages 245– 254, 1996. CODEN PPCMEG.

Lalani:1996:JPLb

[LJ96a] Suleiman Lalani and Kris Jamsa. Java Programmer’s Library, Deutsche Ausgabe. Franzis-Verlag, Feldkirchen, Germany, September 1996. ISBN 3-7723-4573-5. ???? pp. LCCN ???? 89 DM.

URL http://www.franzis-
buch.de/katalog/javalibrary/

. German translation of [LJ96b]. Includes CD-ROM.

Lalani:1996:JPLa

[LJ96b] Suleiman ‘Sam’ Lalani and Kris A. Jamsa. Java Programmer’s Library. Jamsa Press, Las Vegas, NV, USA, 1996. ISBN 1-884133-26-6. iv + 556 pp. LCCN QA76.73.J38L35 1996. US$49.95. URL http://www.

jamsa.com/catalog/javalib/
javalib.htm.

Lalani:1997:JBD

[LJ97] Suleiman Lalani and Kris Jamsa. Java Biblioteca del Programador. McGraw-Hill Spain, ????, November 1997. ISBN 970-10-1350-6. ???? pp. LCCN ???? 6500

Ptas.  URL http://www.
mcgraw-hill.es/McGrawHill/

catalogo.nsf/b013777771f7e01bc125647a00607b0e/
2ed8c207cc71905dc125651c00399d6b?
OpenDocument.

Lalani:19xx:JBDc

[LJxxa] Suleiman Lalani and Kris Jamsa. Java — Biblioteca do Programador. Makron Books do Brasil Editora Ltda., Rua Tabapuã 1348, São Paulo SP - 04533-004 - Brasil, 19xx. ISBN 85-346-0517-3 (??invalid ISBN??). LCCN

????  ????  URL http:/

/safe.tesla.com.br/makron/
options/livro.asp?id=260.

Lalani:19xx:JLD

[LJxxb] Suleiman Lalani and Kris Jamsa. Java Libreria Del Programmatore. Mondadori Informatica, ??, Italy, 19xx. LCCN ???? 90000 L.

Lalani:19xx:JBDb

[LJxxc] Suleiman ‘Sam’ Lalani and Kris A. Jamsa. Java — Biblioteca del Programador. Mc-Graw-Hill, New York, NY, USA, 19xx. ISBN ???? ???? pp. LCCN ???? URL http:

//mexplaza.udg.mx/McGraw/
;   http://www.jamsa.com/
authors/bio.htm#kjamsa;
http://www.wizard.com/~slalani.

Spanish translation of [LJ96b].

Lalani:19xx:JBDa

[LJxxd] Suleiman ‘Sam’ Lalani and Kris A. Jamsa. Java — Biblioteca do Programador. Makron Books do Brasil Editora Ltda., Rua Tabapuã 1348, São Paulo SP - 04533-004 - Brasil, 19xx. ISBN ???? ???? pp. LCCN

????  URL http://www.
jamsa.com/authors/bio.htm#
kjamsa; http://www.makron.
com/;   http://www.wizard.
com/~slalani.  Portuguese

translation of [LJ96b].

Lazarus-Karaoglan:1999:BOE

[LK99] T. Lazarus-Karaoglan. Bringing order to the embedded space. IEEE Spectrum, 36(4):65–72, April 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lecomte:1997:PJ

[LL97] Cyrille Lecomte and Thomas Leduc. Programmation JavaScript. Eyrolles, Paris, France, 1997. ISBN 2-212-08937-6. viii + 318 pp. LCCN ???? URL http:

//www.estp.fr/eyrolles. In-

cludes CD-ROM.

Lewis:1998:JSS

[LL98] John Lewis and William Loftus. Java software solutions: foundations of program design. Addison-Wesley, Reading, MA, USA, 1998. ISBN 0-201-57164-1. xxii + 857 pp. LCCN QA76.73.J38L49 1998. US$48.36. URL http:/

/www.awl.com/cseng/titles/
0-201-57164-1.

Lin:19xx:JAY

[LLxx] Fu Yan Lin and Mitch Lai. Java Applets For Your Web Page. Unalis Corporation, ??, Taiwan, 19xx. ISBN 957-22-2171-X. ???? pp. LCCN ????

390 NT$. URL http://www.
unalis.com.tw/index.html.

Lloyd:1998:ONV

[Llo98] Doug Lloyd. Official Netscape Visual JavaScript Book. Ventana Press, Chapel Hill, NC, USA, 1998. ISBN 1-56604-761-7 (softcover). xxiv + 436 pp. LCCN QA76.73.J39 L56 1997. US$49.99, CDN$69.99.

Lemay:1996:LLW

[LM96] Laura Lemay and Michael G.

Moncur. Laura Lemay’s Web Workshop JavaScript. Laura Lemay’s Web Workshop Series. Sams.net, Indianapolis, IN 46268, USA, September 1, 1996. ISBN 1-57521-141-6. xxix + 393 pp. LCCN QA76.73.J39L47 1996. US$39.99, CDN$56.95, CDN$37.50. URL http://

merchant.superlibrary.com:
8000/catalog/hg/PRODUCT/PAGE/
15752/bud/1575211416.html.

Lemay:1997:JLL

[LM97a] Laura Lemay and Michael Moncur. JavaScript: Laura Lemays Web Workshop. Markt & Technik, Haar bei München, Germany, 1997. ISBN 3-8272-5204-0 (??invalid ISBN??). 470 pp. LCCN ???? 59,95 DM. URL

http://www.emedia.de/bin/
bookshop?show=4938&id=;
http://www.mut.ch/profit.
htx?&x_pk=kd84mf7ms9dm; http:
//www.mut.com/;   http://
www.mut.com/listen/fmonatb.

htm. German translation of [LM96].

Lemay:1997:LJ

[LM97b] Laura Lemay and Michael Moncur. Lemay JavaScript. Academic Service, Postbus 81, 2870 AB Schoonhoven, The Netherlands, 1997. ISBN 90-395-0695-7. 387 pp. LCCN ???? fl. 58.

                                        URL http://slack.lne.com/
                                        lemay/index.html;   http:
                                        //www.acaserv.nl; http://
                                        www.acaserv.nl/5_0695.htm.

Dutch translation of [LM96]. Includes CD-ROM.

Louis:1997:JLI

[LM97c] Dirk Louis and Peter Müller. Jetzt lerne ich Java. Markt & Technik, Haar bei München, Germany, 1997. ISBN 3-8272-5311-X. ???? pp. LCCN ???? 39,95 DM. URL http:/

                                        /www.mut.ch/profit.htx?&x_
                                        pk=kd84mf7ms9dm.

Lamersdorf:1998:TDS

[LM98] Winfried Lamersdorf and Michael Merz, editors. Trends in distributed systems for electronic commerce: international IFIP/GI working conference, TREC’98, Hamburg, Germany, June 3–5, 1998, proceedings, volume 1402 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. ISBN 3-540-64564-0 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1402.

Lewis:1999:EBP

[LM99] Brian T. Lewis and Bernd Mathiske. Efficient barriers for persistent object caching in a high-performance Java Virtual Machine. Technical report, Sun Microsystems, Palo Alto, CA, USA, December 1999. 9 pp.

Leuschel:1998:SAP

[LMdS98] Michael Leuschel, B. Martens, and D. de Schreye. Some achievements and prospects in partial deduction. ACM Computing Surveys, 30(3es):4:1– 4:??, September 1998. CO-DEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (elec-

tronic). URL http://www.
acm.org:80/pubs/citations/
journals/surveys/1998-30-
3es/a4-leuschel/. Article 4.

Lea:1996:JVW

[LMM96] Rodger Lea, Kouichi Matsuda, and Ken Miyashita. Java for 3D and VRML Worlds. New Riders Publishing, Carmel, IN, USA, December 1, 1996. ISBN 1-56205-689-1. xiii + 399 pp. LCCN QA76.73.J38L43 1996. US$45.00. URL http:

//www.mcp.com/cgi-bin/bag?
isbn=1-56205-689-1&last=/
bookstore.

Ladd:1996:PEU

[LO96] Eric Ladd and Jim O’Donnell.

Platinum Edition Using HTML, Java and CGI. Que Corporation, Indianapolis, IN, USA, November 1996. ISBN 0-7897-0932-5. xlvi + 1471 pp.

LCCN QA76.76.H94L33 1996. US$69.99.

Ladd:1997:PEU

[LO97a] Eric Ladd and Jim O’Donnell. Platinum Edition Using HTML 3.2, Java 1.2 and Javascript. Que Corporation, Indianapolis, IN, USA, second edition, December 1997. ISBN 0-7897-1477-9. ???? pp. LCCN ???? US$60; 138,00

                                        DM. URL http://merchant.
                                        superlibrary.com:8000/catalog/
                                        mcp/PRODUCT/PAGE/07897/bud/
                                        0789714779.html;   http:/
                                        /www.mut.ch/profit.htx?&x_
                                        pk=kd84mf7ms9dm.

Ladd:1997:UHJ

[LO^{+}97b] Eric Ladd, Jim O’Donnell, et al., editors. Using HTML 4.0, Java 1.1, and JavaScript 1.2. Que Corporation, Indianapolis, IN, USA, 1997. ISBN 0-7897-1477-9. xxix + 1395 pp. LCCN QA76.76.H94L333 1998. 138,00 DM. URL http:/

                                        /www.mut.ch/profit.htx?&x_
                                        pk=kd84mf7ms9dm.

Ladd:1998:UHJ

[LO98a] Eric Ladd and Jim O’Donnell. Using HTML 4.0, Java 1.1, and JavaScript 1.2. Que Corporation, Indianapolis, IN, USA, second edition, 1998. ISBN 0-7897-1477-9. xxix + 1395 pp. LCCN QA76.76.H94L333 1998.

Lange:1998:PDM

[LO98b] Danny B. Lange and Mitsuru Oshima. Programming and Deploying Mobile Agents with Java Aglets. Addison-Wesley, Reading, MA, USA, September 1998. ISBN 0-201-32582-9. xxiv + 225 pp. LCCN ???? US$33.

Ladd:1999:PIH

[LO99a] Eric Ladd and Jim O’Donnell.

Programmation Internet, HTML 4, XML et Java R 2: ressources d’experts. CampusPress France, Paris, France, 1999. ISBN 2-7440-0628-9. xvii + 1175 pp. LCCN ????

Ladd:1999:UHX

[LO^{+}99b] Eric Ladd, Jim O’Donnell, et al.

Using HTML 4, XML, and Java 1.2: Platinum Edition. Que Corporation, Indianapolis, IN, USA, 1999. ISBN 0-7897-1759-X. xxiii + 1282 pp. LCCN QA76.76.H94L332. US$49.99.

Lange:1999:PDJ

[LO99c] Danny B. Lange and Mitsuru Oshima. Programming and deploying Java mobile agents with Aglets. Addison-Wesley, Reading, MA, USA, 1999. ISBN 0-201-32582-9. xxiv + 225 pp. LCCN QA76.76.I58L36 1998. US$37.95.

Ladd:1996:UHJa

[LOA^{+}96] Eric Ladd, Jim O’Donnell, Jerry Ablan, et al. Using HTML 3.2, Java 1.1, and CGI: Platinum Edition. Que Corporation, Indianapolis, IN, USA, November 1, 1996. ISBN 0-7897-0932-5. xlvi + 1471 pp. LCCN QA76.76.H94L33 1996.

US$69.99. Includes two CD-ROMs.

Loeffler:1997:MJF

[Loe97] G. Loeffler. A multithreaded Java framework for solving linear elliptic partial differential equations in 3D. Lecture Notes in Computer Science, 1343:121– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Loeb:1999:JA

[Loe99] Bill Loeb. The Java 2D API. Dr. Dobb’s Journal of Software Tools, 24(2):44, 46–49, February 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/1999/9902/9902toc.
htm;   http://www.ddj.com/
ftp/1999/1999_02/j2d.txt.

Lange:1997:APM

[LOKK97] D. B. Lange, M. Oshima, G. Kargoth, and K. Kosaka. Aglets: Programming mobile agents in java. Lecture Notes in Computer Science, 1274:253– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Longuet:1997:JM

[Lon97] Patrick Longuet. Java MegaPoche. Sybex France, Paris, France, 1997. ISBN 2-7361-2406-5. 453 pp. LCCN ????

????  URL http://www-
db.sybex.fr/cgi-bin/Sybex/
Webdriver?MIval=Fiche_Book.
html&Reference=2406;   http:
//www.sybex.fr/.  Includes

CD-ROM.

Longuet:19xx:LDJ

[Lonxxa] Patrick Longuet. Le Livre d’Or Java. Sybex France, Paris, France, 19xx. ISBN 2-7361-2274-7. ???? pp. LCCN ????

279 FF. URL http://www-
db.sybex.fr/cgi-bin/Sybex/
Webdriver?MIval=Fiche_Book.
html&Reference=2274.

Longuet:19xx:PVJ

[Lonxxb] Patrick Longuet. Praktijkboek Visual J++. Sybex Nederland, ????, 19xx. ISBN 90-419-0076-4. ???? pp. LCCN ???? fl.

49. URL http://www.sybex.
nl/catalogus/programmeren/
pbvisj.htm.

Lorenz:1996:JOO

[Lor96a] Mark Lorenz. Java As an Object-Oriented Language. Java Report Professional Ser.; Vol.

1. SIGS Books and Multimedia, 71 W 23rd St., New York, NY 10010, USA, June 1996. ISBN 1-884842-40-2. 50 pp. LCCN QA76.73.J38L67

1996.  US$85.00.  URL

http://www.sigs.com/books/
wp_java.html;   http://www.
sigs.com/books/wp_javaoo.
html.

Lorenz:1996:NOI

[Lor96b] Mark Lorenz. Navigating Objects: The Interactive Approach to Object Technology. SIGS Books, 71 W 23rd St., New York, NY 10010, USA, March 1996. ISBN 1-884842-32-1. ???? pp. LCCN ???? US$249.00. Includes CD-ROM.

Lotus:1995:LOP

[Lot95] Lotus Development Corporation. Lotus outlines plans to deliver powerful integration of Notes and World Wide Web— announces Lotus Notes server that includes native Notes support for HTTP, HTML and Java technology. press release, December 13, 1995. URL http:

                                        //www.Lotus.com/corpcomm/
                                        3406.htm.

Louis:1995:JMC

[Lou95a] Tristan Louis. Java on the menu: Causing jitters? Web Week, page ??, June 1995. CODEN WEWEFP. URL

                                        http://www.mecklerweb.com/
                                        ww/news/june-95/paper/df/
                                        1-2java.html.

Louis:1995:JPN

[Lou95b] Tristan Louis. Java: Promising, but needs work. Web Week, page ??, August 1995. CODEN WEWEFP. URL

                                        http://www.mecklerweb.com/
                                        ww/news/aug-95/1-4java.html.

Louis:1995:SBH

[Lou95c] Tristan Louis. Sun brews Hot-Java. Internet World, page ??, September 1995. CODEN IERNE8. ISSN 1064-3923. URL

                                        http://www.mecklerweb.com/
                                        iw/v6n9/news.htm#sun.

Lemay:1996:YJI

[LP96a] Laura Lemay and Charles Perkins. Yes, Java is secure. here’s why — secure transactions on the Internet? yup. here, today, now. just run Javacompiled bytecode on a Java Virtual Machine, and you can keep the cyberpunks off the Internet. Datamation, 42(5):47– ??, ???? 1996. CODEN DTM-NAT. ISSN 0011-6963.

Lemay:1996:YJS

[LP96b] Laura Lemay and Charles Perkins. Yes, Java’s secure. here’s why. Datamation, 42 (????):47–49, ???? 1996. CO-DEN DTMNAT. ISSN 0011-6963.

Lemay:1996:JT

[LP96c] Laura Lemay and Charles L.

Perkins. Java in 21 Tagen. Markt & Technik, Haar bei München, Germany, 1996. ISBN 3-87791-865-4. 571 pp. LCCN ???? 79 DM. URL http:

//rendezvous.com/java/book.
html;   http://slack.lne.
com/lemay/index.html; http:
//www.mut.com/. Includes CD-

ROM.

Lemay:1996:PJ

[LP96d] Laura Lemay and Charles L.

Perkins. Le Programmeur Java. Le Programmeur. Simon & Schuster Macmillan (France), Paris, France, 1996. ISBN 2-7440-0104-X. xvii + 498 pp. LCCN ???? 220 FF.

URL http://slack.lne.com/
lemay/index.html.  French

translation of [LP96e]. Includes CD-ROM.

Lemay:1996:TYJa

[LP96e] Laura Lemay and Charles L. Perkins. Teach Yourself Java in 21 Days. Sams.net, Indianapolis, IN 46268, USA, 1996. ISBN 1-57521-030-4. xxiv + 527 pp. LCCN QA76.73.J38L46 1996. US$39.99. URL http://

                                        merchant.superlibrary.com:
                                        8000/catalog/hg/PRODUCT/PAGE/
                                        15752/bud/1575210304.html;
                                         http://www.mcp.com/13113543955138/
                                        cgi-bin/bag?isbn=1-57521-
                                        030-4&last=/bookstore. In-

cludes CD-ROM. 250,000 copies of this book were sold in its first eight weeks.

LeJacq:1997:RM

[LP97a] J. P. LeJacq and D. M. Papurt. Resource management. Journal of Object Oriented Programming, 9(9):13–17, February 1997. CODEN JOOPEC. ISSN 0896-8438.

Lemay:1997:JGCa

[LP97b] Laura Lemay and Charles L. Perkins. Java — Guida Completa. Apogeo srl, Viale Papiniano 38, 20123 Milano, Italy, 1997. ISBN 88-7303-189-7. 544 pp. LCCN ???? 64.000 Lire. URL http://www.

                                        apogeonline.com/catalogo/
                                        278.html;   http://www.urra.
                                        it/278.html; http://www.urra.
                                        it/apocat.html.

Lemay:1997:JT

[LP97c] Laura Lemay and Charles L. Perkins. Java 1.1 in 21 Tagen. Markt & Technik, Haar bei München, Germany, 1997. ISBN 3-8272-2009-2. 840 pp. LCCN ???? 79,95 DM. URL http://

rendezvous.com/java/book.
html;   http://slack.lne.
com/lemay/index.html;   http:
//www.mut.ch/html/buchtext.
htx?&x_pp=STb4XDV284xv2gpf&
x_pk=Direkt&x_pa=&x_pr=Fr&
x_artikelnr=22009;   http:
//www.mut.com/.  German

translation of [LP96e]. Includes CD-ROM.

Lemay:1997:TYJ

[LP97d] Laura Lemay and Charles L.

Perkins. Teach Yourself Java 1.1 in 21 Days. Sams.net, Indianapolis, IN 46268, USA, second edition, April 1, 1997. ISBN 1-57521-142-4. xxxi + 775 pp. LCCN QA76.73.J38 L449 1997. US$39.99. URL http://

merchant.superlibrary.com:
8000/catalog/hg/PRODUCT/PAGE/
15752/bud/1575211424.html;
 http://www.mcp.com/samsnet/

.

Lemay:1997:JGCb

[LP97e] Laura Lemay and Charles Perlins. Java 1.1 Guida Completa. Apogeo srl, Viale Papiniano 38, 20123 Milano, Italy, 1997. ISBN 88-7303-351-2. ???? pp. LCCN ????

69,000 L. URL http://www.
apogeonline.com/catalogo/
351.html;   http://www.urra.
it/351.html.

Launay:1998:FPP

[LP98a] P. Launay and J.-L Pazat. A framework for parallel programming in Java. Lecture Notes in Computer Science, 1401:628– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Launay:1998:GDP

[LP98b] P. Launay and J.-L Pazat. Generation of distributed parallel Java programs. Lecture Notes in Computer Science, 1470:729– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lemay:1998:AJD

[LP98c] Laura Lemay and Charles L. Perkins. Aprendiendo Java 1.1 en 21 Dias. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, second edition, 1998. ISBN 970-17-0054-6 (??invalid ISBN??). LCCN ????

                                        ???? URL http://www.mcl.
                                        com.mx/sotano/prentice.fm$
                                        Retrive?RecID=2743&html=HTMLRecord1.

Lemay:19xx:PJb

[LPxxa] Laura Lemay and Charles L. Perkins. Le Programmeur Java 1.1. Simon & Schuster Macmillan (France), Paris, France, 19xx. ISBN 2-7440-0297-6. ???? pp. LCCN ???? 199

                                        FF. URL http://www.ssm.fr/
                                        catalog/fiches/0297.htm.

Lemay:19xx:TYJd

[LPxxb] Laura Lemay and Charles L. Perkins. Teach Yourself Java 1.1 in 21 Days. Sams.net, Indianapolis, IN 46268, USA, 19xx.

ISBN 1-57521-142-2. LCCN ???? US$40.

Lemay:1996:TYS

[LPC96] Laura Lemay, Charles L.

Perkins, and Rogers Cadenhead. Teach Yourself Sunsoft Java Workshop in 21 Days. Sams.net, Indianapolis, IN 46268, USA, 1996. ISBN 1-57521-159-9. xxvi + 622 pp. LCCN QA76.73.J39L46 1996. US$39.99. URL http://

merchant.superlibrary.com:
8000/catalog/hg/PRODUCT/PAGE/
15752/bud/1575211599.html;
http://www.mcp.com/cgi-bin/
bag?isbn=1-57521-159-9&last=
/bookstore.

Lemay:1996:TYJd

[LPM96] Laura Lemay, Charles L.

Perkins, and Michael Morrison. Teach Yourself Java in 21 Days: Professional Reference Edition. Sams.net, Indianapolis, IN 46268, USA, 1996. ISBN 1-57521-183-1. xxxvi + 1247 pp. LCCN QA76.73.J38 L46 1996b. US$59.99. URL http://

merchant.superlibrary.com:
8000/catalog/hg/PRODUCT/PAGE/
15752/bud/1575211831.html;
http://www.mcp.com/cgi-bin/
bag?isbn=1-57521-183-1&last=
/bookstore.

Lemay:1997:LLJ

[LPMG97] Laura Lemay, Charles Perkins, Michael Morrison, and Daniel Groner. Laura Lemay’s Java 1.1 Interactive Course. Waite Group Press, Corte Madera, CA, USA, 1997. ISBN 1-57169-083-2. xl + 1194 pp. LCCN QA76.73.J38 L447 1997. US$45. URL http://

merchant.superlibrary.com:
8000/catalog/hg/PRODUCT/PAGE/
15716/bud/1571690832.html.

Leppinen:1997:JCB

[LPR97] Mika Leppinen, Pekka Pulkkinen, and Aapo Rautiainen. Java- and CORBA-based network management. Computer, 30(6):83–87, June 1997. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lemay:1996:TYJb

[LPW96] Laura Lemay, Charles L. Perkins, and Tim Webster. Teach Yourself Java for Mac-Intosh in 21 Days. Hayden Books, 4300 West 62nd Street, Indianapolis, IN 46268, USA, April 1996. ISBN 1-56830-280-0 (paperback). xviii + 567 pp. LCCN QA76.73.J38L45 1996. US$40.00. URL http://

merchant.superlibrary.com:
8000/catalog/hg/PRODUCT/PAGE/
15683/bud/1568302800.html;
 http://www.mcp.com/hayden/
internet/tyjava/index.html.

Loia:1999:EFD

[LQ99] Vincenzo Loia and Michel Quaggetto. Embed finite domain constraint programming into Java and some Webbased applications. Software —Practice and Experience, 29 (4):311–339, April 10, 1999. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=55001842;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=55001842&PLACEBO=IE.pdf.

Lamping:1996:VLT

[LR96] John Lamping and Ramana Rao. Visualizing large trees using the hyperbolic browser. In Tauber [Tau96], pages 388–389. ISBN 0-201-94687-4 (paperback), 0-89791-777-4 (ACM). LCCN ????

Lyon:1997:JDS

[LR97] Douglas A. Lyon and Hayagriva V. Rao. Java Digital Signal Processing. MIS Press and M&T Books, P. O. Box 5277, Portland, OR 97208-5277, USA, Tel: (503) 282-5215 and M&T Publishing, Inc., 501 Galveston Drive, Redwood City, CA 94063, USA, November 1997. ISBN 1-55851-568-

2. LCCN ???? US$45.

URL http://www.mandt.com/
javadigsignproc.htm.

Leroy:1998:SPT

[LR98] Xavier Leroy and François Rouaix. Security properties of typed applets. In ACM [ACM98b], pages 391–

403. ISBN 0-89791-979-

3. LCCN QA76.7 .A15

1998.  URL http://www.
acm.org:80/pubs/citations/
proceedings/plan/268946/p391-

leroy/. ACM order number: 549981.

Lemay:1996:WWN

[LS96] Laura Lemay and Ned Snell. Laura Lemay’s Web workshop: Netscape Gold and JavaScript. Sams.net, Indianapolis, IN 46268, USA, 1996. ISBN 1-57521-128-9. xxii + 368 pp. LCCN TK5105.883.N48L466

                                        1996. URL http://merchant.
                                        superlibrary.com:8000/catalog/
                                        hg/PRODUCT/PAGE/15752/bud/
                                        1575211289.html;   http://
                                        www.mcp.com/samsnet/.

Lam:1997:JTI

[LS97a] Ioi K. Lam and Brian C. Smith. Jacl: a Tcl implementation in Java. In USENIX [USE97a], pages 31–?? ISBN 1-880446-87-1. LCCN QA76.73.T44 T44

                                        1997. URL http://www.cs.
                                        cornell.edu/home/ioi/Jacl.

Lockwood:1997:JAD

[LS97b] Stephen D. Lockwood and Madhu Siddalingaiah. Java API for Dummies: Quick Reference. I D G Books Worldwide, Indianapolis, IN, USA, May 1997. ISBN 0-7645-0118-6. xvi + 224 pp. LCCN QA76.73.J38 L63 1997. US$14.99. URL http:/

                                        /www.idgbooks.com/cgi-bin/
                                        db/fill_out_template.pl?idgbook:
                                        0-7645-0112-7:book-idg::uidg323;
                                         http://www.idgbooks.com/
                                        database/book_result.msql?
                                        uid=uidg323&isbn=0-7645-0118-
                                        6.

Little:1998:JTI

[LS98] M. C. Little and S. K. Shrivastava. Java transactions for the Internet. In USENIX [USE98b], page ?? ISBN ????

LCCN ???? URL http://www.
usenix.org/publications/library/
proceedings/coots98/little.
html.

Laird:1999:REN

[LS99] Cameron Laird and Kathryn Soraiz. Regular expressions: New choices for scripting: Language designers offer novel possibilities for scripting developers. Sunworld Online, 5(2):??, February 1999. ISSN 1091-8914.

URL http://www.sunworld.
com/swol-02-1999/swol-02-
regex.html?0202a. Discusses

new scripting languages: Ficl, FIJI, Rebol, Ruby and WebL. Ficl is a descendant of Forth. FIJI is real Forth, but with objects that are Java objects, and with full access to Java.

Liang:19xx:JNI

[LSxx] Sheng Liang and Beth Sterns.

The Java Native Interface: a Tutorial and Specification. Addison-Wesley, Reading, MA, USA, 19xx. ISBN 0-201-18393-

5. LCCN ???? US$40. URL

http://www2.awl.com/cseng/
javaseries/jni.html.

League:1999:RJC

[LST99] Christopher League, Zhong Shao, and Valery Trifonov. Representing Java classes in a typed intermediate language. ACM SIGPLAN Notices, 34 (9):183–196, September 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

                                        URL http://www.acm.org/
                                        pubs/citations/proceedings/
                                        fp/317636/p183-league. Pro-

ceedings of the ACM SIGPLAN International Conference on Functional Programming (ICFP ’99).

Lie:19xx:JSM

[LTxx] Chung Whei Lie and Yeung E Tai. Java Special Manual. Flag Publishing, ??, Taiwan, 19xx. ISBN 957-717-197-4 (??invalid ISBN??). ???? pp. LCCN ????

                                        500 NT$. URL http://www.
                                        seed.net.tw/~flaginfo/.

Liao:1996:FDU

[LTC96] W. S. Liao, See-Mong Tan, and R. H. Campbell. Fine-grained, dynamic user customization of operating systems. In Cabrera and Islam [CI96], page ?? CO-DEN PIWOF6. ISBN 0-8186-7692-2, 0-8186-7693-0, 0-8186-7694-9 (microfiche). ISSN 1063-5351. LCCN QA 76.76 O63 I59 1996.

Lu:1996:PDD

[Lu96] C. Lu. Power documents [document oriented computing]. Inc, 18(9):35–??, ???? 1996. CO-DEN INCCDU. ISSN 0162-8968.

Lunde:1999:CIP

[Lun99] Ken Lunde. CJKV Information Processing: Chinese, Japanese, Korean & Vietnamese Computing. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1999.

ISBN 1-56592-224-7. 1174 pp. LCCN PL1074.5 .L86 1999. US$64.95. URL http:

//www.oreilly.com/catalog/
cjkvinfo/.

Lutz:1998:NBN

[Lut98a]  Michael J. Lutz.  New

books: Network management with scripts; Ethernet backbone; Web technologies; total cache knowledge; putting on the constraints; Java GUI details; doing business on the net; programming Netscape. Computer, 31(7): 89, July 1998. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

http://dlib.computer.org/
co/books/co1998/pdf/r7089.
pdf.

Lutz:1998:NBPb

[Lut98b] Michael J. Lutz. New books:

Programming for high-end JavaBeans; Intel secrets; Ada and concurrency; data-modeling primer; technology and economics; dissecting a microprocessor; real-world ATM; when computers run amok. Computer, 31(10):109, October 1998. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

http://dlib.computer.org/
co/books/co1998/pdf/rx109.
pdf.

Anonymous:1999:NB

[Lut99a] Michael J. Lutz. New books:

Are you listening, Dave?; cultural team chemistry; Java image processing; Y2K preparation; making distributed objects work for COM-CORBA; fuzzy system use. Computer, 32(3): 103, March 1999. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

                                        http://dlib.computer.org/
                                        co/books/co1999/pdf/r3103.
                                        pdf.

Lutz:1999:NBY

[Lut99b] Michael J. Lutz. New books: Are you listening, Dave?; cultural team chemistry; Java image processing; Y2K preparation; making distributed objects work for COM-CORBA; fuzzy system use. Computer, 32(3): 103, March 1999. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

                                        http://dlib.computer.org/
                                        co/books/co1999/pdf/r3103.
                                        pdf.

Lutz:1999:NBCa

[Lut99c] Michael J. Lutz. New books: Case-driven object modelling with UML; virtual-human integration; genetic programming; Java image processing; WAP programming. Computer, 32(7): 99, July 1999. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

                                        http://dlib.computer.org/
                                        co/books/co1999/pdf/r7099.

pdf. Includes review of [Lyo99].

Lutz:1999:NBCb

[Lut99d] Michael J. Lutz. New books:

Creating a virtual window to the world; optical networks in layered multiwavelength; automated software testing; software improvement through refactoring; getting a feel for Java. Computer, 32(10):116, October 1999. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

http://dlib.computer.org/
co/books/co1999/pdf/rx116.

pdf. Includes review of [Sun99b].

Lutz:1999:NBF

[Lut99e] Michael J. Lutz. New books:

Folding all things virtual into urban life; performance modeling in network design; practical automated software testing; Java patterns; testing software step by step; working with UNIX shells. Computer, 32(11):119, November 1999. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

http://dlib.computer.org/
co/books/co1999/pdf/ry119.

pdf. Includes review of [Lan99b].

Liang:1999:CPS

[LV99] Sheng Liang and Deepa Viswanathan.

Comprehensive profiling support in the Java Virtual Machine. In USENIX [USE99c], page ?? ISBN ???? LCCN QA76.64 .U84 1999. URL http:

//www.usenix.org/publications/
library/proceedings/coots99/
liang.html.

Li:1996:WWW

[LVS^{+}96a] Xin Li, D. J. Valentino, G. J. So, R. Lufkin, and R. K. Taira. A World Wide Web telemedicine system. Proceedings of the SPIE — The International Society for Optical Engineering, 2711 (????):427–439, ???? 1996. CO-DEN PSISDG. ISSN 0277-786X (print), 1996-756X (electronic).

XinLi:1996:WWW

[LVS^{+}96b] Xin Li, D. J. Valentino, G. J. So, R. Lufkin, and R. K. Taira. A World Wide Web telemedicine system. Proceedings of the SPIE — The International Society for Optical Engineering, 2711:427–439, ???? 1996. CO-DEN PSISDG. ISSN 0277-786X (print), 1996-756X (electronic).

Lindholm:1997:IJV

[LY97a] Tim Lindholm and Frank Yellin. Inside the Java Virtual Machine. UNIX review, 15(1):31, 32, 34–36, 38, 39, January 1997. CODEN UNRED5. ISSN 0742-3136.

Lindholm:1997:JVM

[LY97b] Tim Lindholm and Frank Yellin. The Java Virtual Machine Specification. The Java Series. Addison-Wesley, Reading, MA, USA, January 1997. ISBN 0-201-63452-X. xvi + 475 pp. LCCN QA76.73.J38L56 1997. US$36.53. URL http://www.

                                        aw.com/cp/javaseries.html;
                                         http://www.aw.com/cseng/
                                        titles/0-201-63452-X.

Lindholm:1999:JVM

[LY99] Tim Lindholm and Frank Yellin.

The Java Virtual Machine Specification. Addison-Wes-ley, Reading, MA, USA, second edition, 1999. ISBN 0-201-43294-3. xv + 473 pp. LCCN QA76.73.J38L56 1999. US$42.95.

Lindholm:19xx:JSV

[LYxxa] Tim Lindholm and Frank Yellin.

Java — Die Spezifikation der Virtuellen Maschine. Addison-Wesley Verlag, Bonn, Germany, 19xx. ISBN 3-8273-1045-8 (??invalid ISBN??). LCCN ???? ????

Lindholm:19xx:JVMa

[LYxxb] Tim Lindholm and Frank Yellin.

The Java Virtual Machine. GO-TOP Information Inc., 5F, No.7, Lane 50, Sec.3 Nan Kang Road Taipei, Taiwan; Unit 1905,Metro Plaza Tower 2, No.223 Hing Fong Road, Kwai Chung, N.T., Hong Kong, 19xx. ISBN ???? ???? pp. LCCN

???? URL http://www.gotop.

com.tw. Chinese translation by Thi Shiang Workshop.

Lindholm:19xx:JVMb

[LYxxc] Tim Lindholm and Frank Yellin.

The Java Virtual Machine. GO-TOP Information Inc., 5F, No.7, Lane 50, Sec.3 Nan Kang Road Taipei, Taiwan; Unit 1905,Metro Plaza Tower 2, No.223 Hing Fong Road, Kwai Chung, N.T., Hong Kong, 19xx. ISBN ???? LCCN ???? ????

Chinese translation by Thi Shi-

ang Workshop.

Lindholm:1997:JSV

[LYJ^{+}97] Tim Lindholm, Frank Yellin, Bill Joy, et al. Java — Die Spezifikation der Virtuellen Maschine. Addison-Wesley, Reading, MA, USA, 1997. ISBN 3-8273-1045-8. 464 pp. LCCN ???? 79,90

                                        DM.  URL http://www.
                                        addison-wesley.de/katalog/
                                        item.ppml?textexpr=Java&id=

00142. German translation of

[LY97b].

Lyon:1999:IPJ

[Lyo99]  Douglas A. Lyon.  Image

processing in Java. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1999. ISBN 0-13-974577-7. 450 pp. LCCN QA76.73.J38L94 1999. US$49.99. URL http:

                                        //www.phptr.com/ptrbooks/
                                        ptr_0139745777.html.

Lee:1997:BTI

[LZ97] Han Bok Lee and Benjamin G. Zorn. BIT: a tool for instrumenting Java bytecodes. In USENIX [USE97c], pages 73–82. ISBN 1-880446-91-X. LCCN TK5105.875.I57

                                        U96 1997. URL http://www.
                                        usenix.org/publications/library/
                                        proceedings/usits97/lee.html.

Mendinets:1996:JU

[M^{+}96a] Dave Mendinets et al. JBuilder Unleashed. Sams.net, Indianapolis, IN 46268, USA, 1996.

ISBN 1-57521-103-3. ???? pp. LCCN ???? US$50.

Morrison:1996:JU

[M^{+}96b] Michael Morrison et al. Java unleashed: Everything you need to master Java. Sams.net, Indianapolis, IN 46268, USA, 1996. ISBN 1-57521-049-5 (paperback). xxxii + 971 pp. LCCN QA76.73.J38J384 1996. US$49.99, CDN$69.99, UK£46.95. URL http://

merchant.superlibrary.com:
8000/catalog/hg/PRODUCT/PAGE/
15752/bud/1575210495.html;
 http://www.mcp.com/194152836633/
cgi-bin/bag?isbn=1-57521-
049-5&last=/bookstore;   http:
//www.mcp.com/cgi-bin/bag?
isbn=1--57521-049-5&last=/

. Includes CD-ROM. CD-ROM includes Symantec Café Lite, Sun’s Java 1.01 Developers Kit, plug-ins and helper apps, and HTML editors and utilities for Windows and Macintosh. Covers Version 1.0.

Morrow:1996:JU

[M^{+}96c] Cindy Morrow et al., editors.

Java Unleashed. Sams.net, Indianapolis, IN 46268, USA, 1996. ISBN 1-57521-049-5. xxxii + 971 pp. LCCN QA76.73.J38J384 1996. US$49.99.

URL http://merchant.superlibrary.
com:8000/catalog/hg/PRODUCT/
PAGE/15752/bud/1575210495.

html. Includes CD-ROM with Symantec Café Lite, Sun’s Java 1.01 Developers Kit, plug-ins and helper apps, and HTML editors and utilities for Windows and Macintosh.

Morgan:1997:VJUb

[M^{+}97a] Bryan Morgan et al. Visual J++ 1.1 Unleashed. Sams.net, Indianapolis, IN 46268, USA, second edition, 1997. ISBN 1-57521-356-7. xxix + 845 pp. LCCN QA76.73.J38M66 1997. US$50. URL http://

                                        merchant.superlibrary.com:
                                        8000/catalog/hg/PRODUCT/PAGE/
                                        15752/bud/1575213567.html.

Morgan:1997:VJUa

[M^{+}97b] Bryan Morgan et al. Visual J++ Unleashed. Sams.net, Indianapolis, IN 46268, USA, 1997. ISBN 1-57521-161-0. xxix + 834 pp. LCCN QA76.73.J38M67 1997. US$50. URL http://

                                        merchant.superlibrary.com:
                                        8000/catalog/hg/PRODUCT/PAGE/
                                        15752/bud/1575211610.html;
                                         http://www.superlibrary.
                                        com/cgi-bin/bag?isbn=1-57521-
                                        161-0&last=/bookstore.

Morrison:1997:JUPb

[M^{+}97c] Michael Morrison et al. Java 1.1 unleashed: Everything you need to master Java. Sams.net, Indianapolis, IN 46268, USA, third edition, 1997. ISBN 1-57521-298-6. xlix + 1469 pp. LCCN QA76.73.J38M6738 1997. 118,00 DM; US$60. URL

                                        http://merchant.superlibrary.
                                        com:8000/catalog/hg/PRODUCT/
                                        PAGE/15752/bud/1575212986.
                                        html;   http://www.mut.ch/
                                        profit.htx?&x_pk=kd84mf7ms9dm.

Includes CD-ROM.

Morrison:1997:JUPa

[M^{+}97d] Michael Morrison et al. Java 1.1 Unleashed, Professional Reference Edition. Sams.net, Indianapolis, IN 46268, USA, November 1997. ISBN 1-57521-361-3. xliv + 1164 pp. LCCN QA76.73.J38M674 1997. US$59.99. URL http://

merchant.superlibrary.com:
8000/catalog/hg/PRODUCT/PAGE/
15752/bud/1575211971.html.

Morrison:1997:JUE

[M^{+}97e] Michael Morrison et al. Java Unleashed: Everything you need to master Java. Sams.net, Indianapolis, IN 46268, USA, second edition, 1997. ISBN 1-57521-197-1. xliv + 1164 pp. LCCN QA76.73.J38M674 1997. US$49.99. URL http://

merchant.superlibrary.com:
8000/catalog/hg/PRODUCT/PAGE/
15752/bud/1575211971.html;
 http://www.superlibrary.
com/cgi-bin/bag?isbn=1-57521-
197-1&last=/bookstore.

Ming:19xx:BGJ

[M^{+}xxa] Li Cia Ming et al. Beginner’s Guide To Java. GOTOP Information Inc., 5F, No.7, Lane 50, Sec.3 Nan Kang Road Taipei, Taiwan; Unit 1905,Metro Plaza Tower 2, No.223 Hing Fong Road, Kwai Chung, N.T., Hong Kong, 19xx. ISBN 957-641-768-6. ???? pp. LCCN ???? 460 NT$.

Morrison:19xx:JUb

[M^{+}xxb] Michael Morrison et al. Java Unleashed. DaeRim Publishing, Seoul, Korea, 19xx. ISBN 89-7280-303-0 (??invalid ISBN??). ???? pp. LCCN ???? ????

                                        URL http://www.drbook.co.
                                        kr/daerim/ulsjv.htm. Korean

translation.

Morrison:19xx:VJ

[M^{+}xxc] Michael Morrison et al. Visual J++ 1.1. DaeRim Publishing, Seoul, Korea, 19xx. ISBN 89-7280-375-8 (??invalid ISBN??). ???? pp. LCCN ???? ????

                                        URL http://www.drbook.co.
                                        kr/daerim/ltvj.htm. Korean

translation.

Macromedia:1995:MSM

[Mac95] Macromedia. Macromedia & Sun Microsystems to develop Internet tools and technology. press release, October

                                        30, 1995. URL http://www.
                                        macromedia.com/Industry/Macro/
                                        Ucon/News/sun.html.

Mackie:1996:JVC

[Mac96] Jamieson Mackie. Java and VRML combination may silence critics of both. Silicon Graphics World, 6(11):27, 29, November 1996. ISSN 1057-7041.

Mackie:1997:LIH

[Mac97a] Jamieson Mackie. Is Lucent’s Inferno hot enough for you? HP Chronicle, 14(12):17, 22–23, November 1997. ISSN 0892-2829. Contains brief comparison of Java and Limbo.

Mackie:1997:NJA

[Mac97b] Jamieson Mackie. NCs, Java and ActiveX: Let’s get ready to rumble? HP Chronicle, 14(9): 18–19, August 1997. ISSN 0892-2829.

Mackie:1997:NCJ

[Mac97c] Jamieson Mackie. Network computers, Java and ActiveX; let’s get ready to rumble? Sun-Server, 11(9):1, 20, September 1997. ISSN 1091-4986.

Mackie:1997:SLIa

[Mac97d] Jamieson Mackie. A smoking OS: Is Lucent’s Inferno hot enough for you? Cisco World: The Independent Journal for Internetworking Professionals, 3(9):14–15, September 1997. ISSN 1081-3187. Comparison of Limbo and Java.

Mackie:1997:SLIb

[Mac97e] Jamieson Mackie. A smoking OS: Is Lucent’s Inferno hot enough for you? Silicon Graphics World, 7(9):16, 18, September 1997. ISSN 1057-7041. Comparison of Limbo and Java.

MacGreggor:1998:JNS

[Mac98]  Robert MacGreggor.  Java

Network Security. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1998. ISBN 0-13-761529-9. xii + 232 pp. LCCN QA76.73.J38J378 1998.

US$30.  URL http://www.
prenhall.com/ptrbooks/ptr_
0137615299.html.

MacIntosh:19xx:TJM

[Macxx] Harry MacIntosh. Talk Java to Me. GOTOP Information Inc., 5F, No.7, Lane 50, Sec.3 Nan Kang Road Taipei, Taiwan; Unit 1905,Metro Plaza Tower 2, No.223 Hing Fong Road, Kwai Chung, N.T., Hong Kong, 19xx. ISBN ???? ???? pp. LCCN ???? 300 NT$. URL http://

www.gotop.com.tw.

Maguire:1995:HJH

[Mag95] John G. Maguire. Hot Java for Homer Simpson: Sun Microsystems Inc revives its interactive technology for Fox Broadcasting Co’s 3-D Web site. Open Computing, 12(11):72–??, November 1995. CODEN OPCOEB. ISSN 1078-2370. URL http:/

/www.wcmh.com/oc/products/
reviews/java9511.html.

Mahmoud:1998:DPJ

[Mah98] Qusay H. Mahmoud. Distributed Programming with Java. Manning Publications, Greenwich, CT, USA, October 1998. ISBN 1-884777-65-1. LCCN ???? US$44. URL http:

//www.manning.com/Mahmoud/
index.html.

Main:1998:DSO

[Mai98] Michael Main. Data Structures and Other Objects Using Java. Addison-Wesley, Reading, MA, USA, 1998. ISBN 0-201-35744-5. xxxii + 768 pp. LCCN QA76.73.J38M33

1999. URL http://www.awl-
he.com/titles/26621.html.

Mairet:19xx:JF

[Maixx] Alexandre Mairet. Java facile.

Les Editions Marabout, Alleur, Belgique, 19xx. ISBN 2-501-02911-7 (??Invalid checksum??). ???? pp. LCCN

????  46 FF.  URL

http://www.marabout.com/;
 http://www.marabout.com/
c-javfac.htm;   http://www.
marabout.com/framecat.htm.

Mane:1996:SJP

[Man96a] I. Mane. Survey of the Java programming language. Elektronik, 45(17):84–87, ???? 20, 1996. CODEN EKRKAR. ISSN 0013-5658.

Manger:1996:EJD

[Man96b]  Jason J. Manger.  Essen-

tial Java: Developing Interactive Applications for the World-Wide Web. McGraw-Hill, New York, NY, USA, May 1996. ISBN 0-07-709292-9 (paperback). xi + 364 pp. LCCN QA76.73.J38M36

1996.  US$39.95.  URL

http://www.gold.net/users/
ag17/index.htm;   http://
www.infor.com:53311/cgi/getarec?

mgh30062. Includes CD-ROM.

Manger:1996:JCD

[Man96c]  Jason J. Manger.  Java:

The Complete Development Kit. McGraw-Hill, New York, NY, USA, 1996. ISBN 0-07-709332-1, 0-07-709337-

2. ???? pp. LCCN ???? US$39.95. URL http:

//mcgraw-hill.inforonics.

                                        com/cgi/getarec?mgh30929%
                                        comp.

Manger:1996:JEC

[Man96d] Jason J. Manger. Javascript Essentials: Creating Interactive Web Applications. Osborne/McGraw-Hill, Berkeley, CA, USA, 1996. ISBN 0-07-882234-3. xxxvii + 541 pp. LCCN QA76.73.J39M37 1996. US$34.95.

                                        URL http://www.osborne.
                                        com/.

Manning:1996:WPJ

[Man96e] Michael M. Manning. Web Programming with JBuilder. Sams.net, Indianapolis, IN 46268, USA, 1996. ISBN 1-57521-152-1. ???? pp. LCCN ???? US$40.

Manning:1997:TYBa

[Man97a] Michelle Manning. Teach Yourself Borland JBuilder 2 in 21 Days. Sams.net, Indianapolis, IN 46268, USA, November 1997. ISBN 0-672-31318-9. LCCN ???? US$40.

Manning:1997:TYJ

[Man97b] Michelle M. Manning. Teach Yourself JBuilder in 21 Days. Sams.net, Indianapolis, IN 46268, USA, 1997. ISBN 1-57521-104-1. xxix + 752 pp. LCCN QA76.73.J38M346 1997. US$40. URL http://

                                        merchant.superlibrary.com:
                                        8000/catalog/hg/PRODUCT/PAGE/
                                        15752/bud/1575211041.html.

Manger:19xx:MJ

[Manxxa]  Jason Manger.  Manuale

JavaScript. McGraw-Hill, New York, NY, USA, 19xx. ISBN ???? ???? pp. LCCN ???? 62000 Lire. URL http:

//pallade.iol.it/cgi-bin/
mercator2/mallurl/CESSECCOCEDRBILEDOMAa000eC/
mcghill/index.html;   http:
//www.gold.net/users/ag17/

index.html. Includes floppy disk.

Manning:19xx:VJI

[Manxxb] Michelle M. Manning. Visual J++ Intranet Programming Lab. Prima Publishing, Rocklin, CA, USA, 19xx. ISBN 0-7615-1043-5. ???? pp. LCCN ???? ????

Marketos:1997:JDT

[Mar97a] Joshua Marketos. The Java developer’s toolkit: techniques and technologies for experienced Web programmers. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 1997. ISBN 0-471-16519-0 (paperback). xv + 383 pp. LCCN QA76.73.J38M35 1997. US$29.95. URL http:

//www.wiley.com/compbooks/
catalog/16519-0.htm.

Martin:1997:JG

[Mar97b] Michel Martin. Java 1.1 Gigapoche. Sybex France, Paris, France, December 1997. ISBN 2-7361-2792-7. LCCN ????

98 FF. URL http://www-
db.sybex.fr/cgi-bin/Sybex/

                                        Webdriver?MIval=Fiche_Book.
                                        html&Reference=2792.

Martin:1997:JP

[Mar97c] Michel Martin. Java 1.1 Le Platinum. Sybex France, Paris, France, December 1997. ISBN 2-7361-2758-7. LCCN ????

                                        229 FF. URL http://www-
                                        db.sybex.fr/cgi-bin/Sybex/
                                        Webdriver?MIval=Fiche_Book.
                                        html&Reference=2758.

Martin:1997:JCC

[Mar97d] R. C. Martin. Java and C++: a critical comparison. C++ Report, 9(1):42–49, January 1997. CODEN CRPTE7. ISSN 1040-6042.

Marceau:1998:LWJ

[Mar98a] Carla Marceau. Letters: Why is Java dangerous? IEEE Software, 15(3):10, May/June 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

                                        http://dlib.computer.org/
                                        so/books/so1998/pdf/s3010.
                                        pdf.

Markantonakis:1998:JCT

[Mar98b] Constantinos Markantonakis. Java card technology and security. Information Security Technical Report, 3(2):82–89, ???? 1998. CODEN ISTRFR. ISSN 1363-4127 (print), 1873-605X (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/S0167404898800083.

Martin:1998:JGB

[Mar98c] Peter Martin. Java, the good, the bad and the ugly. ACM SIGPLAN Notices, 33(4):34–39, April 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.
rowan.edu/sigplan/apr199~
1.htm.

Markantonakis:1999:JCT

[Mar99a] Constantinos Markantonakis.

Java card technology and security. Information Security Technical Report, 4(S1):26–27, ???? 1999. CODEN ISTRFR. ISSN 1363-4127 (print), 1873-605X (electronic). URL http://www.

sciencedirect.com/science/
article/pii/S1363412799800508.

Marsic:1999:DFM

[Mar99b] Ivan Marsic. DISCIPLE: a framework for multimodal collaboration in heterogeneous environments. ACM Computing Surveys, 31(2es):??, June 1999. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic). URL http://

www.acm.org/pubs/articles/
journals/surveys/1999-31-
2es/a4-marsic/a4-marsic.pdf;
 http://www.acm.org/pubs/
citations/journals/surveys/
1999-31-2es/a4-marsic/;
http://www.acm.org/pubs/citations/
journals/surveys/1999-31-
2es/a4-marsic/#abstract;
http://www.acm.org/pubs/citations/
journals/surveys/1999-31-
2es/a4-marsic/#indterms.

Margulies:19xx:UJT

[Marxxa] Edwin K. Margulies. Understanding Java Telephony. Flatiron Publishing, ????, 19xx. ISBN 1-57820-003-2. ???? pp. LCCN ???? US$35. URL http:

                                        //www.ctexpo.com/catalog/
                                        demo.html.

Martin:19xx:JL

[Marxxb] Michel Martin. Java Livre d’Or. Sybex France, Paris, France, 19xx. ISBN 2-7361-2274-7. LCCN ????

                                        279 FF. URL http://www-
                                        db.sybex.fr/cgi-bin/Sybex/
                                        Webdriver?MIval=Fiche_Book.
                                        html&Reference=2274.

Martin:19xx:VJM

[Marxxc] Michel Martin. Visual J++ Megapoche. Sybex France, Paris, France, 19xx. ISBN 2-7361-2319-0. ???? pp. LCCN ???? 98 FF. URL http://www-

                                        db.sybex.fr/cgi-bin/Sybex/
                                        Webdriver?MIval=Fiche_Book.
                                        html&Reference=2319.

Mason:1995:NOD

[Mas95] Sara Mason. News: OSF debuts Open Software “Mall” on Web. IDG International News Service, page ??, August 01, 1995.

MacGregor:1996:WSH

[MAS96] Robert S. MacGregor, Alberto Aresi, and Andreas Siegert. www.security: How to Build a Secure World Wide Web Connection: Broad based primer on WWW Security. Practical examples including CGI and Java. Business on the Web Risk Analysis. ITSO networking series. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1996. ISBN 0-13-612409-7 (softcover). xi + 211 pp. LCCN TK5105.888.M236 1996. US$35.00. URL http://www.

prenhall.com/ptrbooks/ptr_
0136124097.html.

Maso:1997:MMV

[Mas97a] Brian Maso. Manual de Microsoft Visual J++. McGraw-Hill Spain, ????, 1997. ISBN 84-481-1054-4. LCCN ???? 4200 Ptas. URL http://www.

mcgraw-hill.es/McGrawHill/
catalogo.nsf/b013777771f7e01bc125647a00607b0e/
d61dcf49c3b59661c12564d60040531d?
OpenDocument.

Maso:1997:VJH

[Mas97b] Brian Maso. The Visual J++ Handbook. Osborne/McGraw-Hill, Berkeley, CA, USA, 1997. ISBN 0-07-882266-1 (paperback). xxiii + 440 pp. LCCN QA76.73.J38M353 1997.

US$30.  URL http://www.
osborne.com/int/jplus.htm.

Maso:1998:JLVa

[Mas98a] Brian Maso. Java 1.1, Level 1 Video. KeyStone Learning Systems, ????, 1998. US$71.95. Video Training: Order number VT10018356.

Maso:1998:JLVb

[Mas98b] Brian Maso. Java 1.1, Level 2 Video. KeyStone Learning Systems, ????, 1998. US$71.95.

Video Training: Order number VT10018357.

Maso:1998:JLVc

[Mas98c] Brian Maso. Java 1.1, Level 3 Video. KeyStone Learning Systems, ????, 1998. US$89.95. Video Training: Order number VT10018358.

Maso:1998:JLVd

[Mas98d] Brian Maso. Java 1.1, Level 4 Video. KeyStone Learning Systems, ????, 1998. US$89.95. Video Training: Order number VT10018359.

Maso:1998:JLVe

[Mas98e] Brian Maso. Java 1.1, Level 5 Video. KeyStone Learning Systems, ????, 1998. US$89.95. Video Training: Order number VT10018360.

Maso:1998:JLVf

[Mas98f] Brian Maso. Java 1.1, Level 6 Video. KeyStone Learning Systems, ????, 1998. US$89.95. Video Training: Order number VT10018361.

Mateosian:1996:BBB

[Mat96a] Richard Mateosian. Books, books, books. IEEE Micro, 16(4):77–79, July/August 1996. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). Brief reviews of several Java books.

Mateosian:1996:MRU

[Mat96b] Richard Mateosian. Micro review — unlike C++, which seeks to extend C and maintain a large degree of backward compatibility, Java starts from C and rips out its most troublesome and error-prone features. IEEE Micro, 16(3):3–5, May/ June 1996. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mateosian:1999:MRP

[Mat99]  Richard Mateosian.  Micro

review: Pot pourri: Jini, Java, WinWriters, Adobe Acrobat. IEEE Micro, 19 (5):10–12, September/October 1999. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL

http://dlib.computer.org/
mi/books/mi1999/pdf/m5010.
pdf.

Mauve:1998:TTA

[Mau98] M. Mauve. TeCo3D: a 3D telecooperation application based on VRML and Java [3654-22]. In Kandlur et al. [KJR98], pages 240–251. ISBN 0-8194-3125-7. ISSN 0277-786X (print), 1996-756X (electronic). LCCN TS510.S63 v.3654.

Maurers:1996:CTJ

[MB96a] Rolf Mäurers and Kai Baufeldt.

Das Computer Taschenbuch Java. DATA-Becker, ????, 1996. ISBN 3-8158-1553-3. 370 pp. LCCN ???? 20 DM. URL

http://www.data-becker.de/
javatest/projekt.htm.

Maurers:1996:PPJ

[MB96b] Rolf Mäurers and Kai Baufeldt. PC Poche Java. Edition Micro Application, Paris, France, 1996. ISBN 2-7429-0752-1. 384 pp. LCCN ???? 59 FF.

Milowski:1996:KSI

[MB96c] R. A. Milowski and P. Bothner. The Kawa scheme interpreter project. Kawa translates the scheme programming language (a simplified Lisp) to Java bytecodes. Kawa is compared with Harissa in [MMBC97].,

                                        1996.  URL http://www.
                                        copsol.com/kawa/index.html.

Maurers:1997:J

[MB97] Rolf Mäurers and Kai Baufeldt. Java. Micro application, Paris, France, second edition, 1997. ISBN 2-7429-0752-1. 369 pp. LCCN ????

McCreary:1998:VVG

[MB98a] C. McCreary and L. Barowski. VGJ: Visualizing graphs through java. Lecture Notes in Computer Science, 1547:454–455, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

McDowell:1998:UJC

[MB98b] C. E. McDowell and E. A. Baldwin. Unloading Java classes that contain static fields. ACM SIGPLAN Notices, 33(1):56–60, January 1998. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Millet:1998:PGT

[MB98c] Laurent Millet and Ted Baker.

Porting the GNAT tasking runtime system to the Java Virtual Machine. Lecture Notes in Computer Science, 1411:19– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/
service/series/0558/bibs/
1411/14110019.htm;   http:
//link.springer-ny.com/link/
service/series/0558/papers/
1411/14110019.pdf.

McIntosh:19xx:DOJ

[MBxx] Harry McIntosh and Mike Britton. Discover Open JBuilder. IDG Books, San Mateo, CA, USA, 19xx. ISBN 0-7645-3061-5. ???? pp. LCCN ???? US$25.

McFarland:1997:DIS

[MBH97] T. McFarland, B. Beck, and H. Hiura. Developing internationalized software with JDK 1.1. In Unicode Consortium [Uni97b], pages TA2–?? ISBN ???? LCCN ???? URL http:

//www.unicode.org/unicode/
iuc11/program.html.

Mirho:1997:JTC

[MC97] Charles Mirho and Tom Clements.

JavaOS: Thin client, fat service — JavaOS uses a small memory footprint, yet its networkcentric design lets it access large-scale services. BYTE Magazine, 22(7):53–54, July 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Miller:1998:VMB

[MC98] Gregory R. Miller and Stephen C. Cooper. Visual mechanics: beams and stress states. PWS Publishing Company, Boston, MA, USA, 1998. ISBN 0-534-95587-8. x + 149 pp. LCCN TA660.B4 M55 1998.

McCarthy:19xx:BWJ

[MCxx] Martin McCarthy and Alistair Carty. Building 3D Worlds in Java and VRML. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 19xx. ISBN 0-13-748625-1. LCCN ???? US$55.

McChesney:1995:RJ

[McC95a] John McChesney. Report on Java. National Public Radio Morning Edition, October

                                        27, 1995. URL http://www.
                                        realaudio.com/rafiles/npr/
                                        password/nb-102701-7.ram.

McChesney:1995:RNS

[McC95b] John McChesney. Report on Netscape and Sun Microsystems’ JavaScript. National Public Radio All Things Considered, December 4, 1995.

                                        URL http://www.realaudio.
                                        com/rafiles/npr/password/
                                        120401-9.ram.

McCarthy:1996:PJI

[McC96a] Michael McCarthy. From the publisher: Java and the Internet PC. Sunworld Online, 01:??, 1996. ISSN 1091-8914.

URL http://www.sun.com/
sunworldonline/swol-01-1996/
swol-01-editorial.html.

McCarthy:1996:GJ

[McC96b] Vance McCarthy. Gosling on Java. Datamation, 42(????):30– 37, ???? 1996. CODEN DTM-NAT. ISSN 0011-6963.

McCarthy:1996:GJD

[McC96c] Vance McCarthy. Gosling on Java — a DATAMATION exclusive interview with the creator of Java, Sun Microsystems’ James Gosling. Datamation, 42 (5):30–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

McComb:1996:JSC

[McC96d] Gordon McComb. The JavaScript sourcebook: create interactive JavaScript programs for the World Wide Web. Wiley Computer Publishers, New York, NY, USA, August 1, 1996. ISBN 0-614-20338-4, 0-471-16185-3 (paperback). xxv + 726 pp. LCCN QA76.73.J38M353 1996. US$44.95 (hardback), US$39.95 (paperback). URL

http://www.wiley.com/.

McCoy:1996:PBH

[McC96e] John H. McCoy. Programmer’s bookshelf: Hooked on Java, by A. van Hoff, S. Shaio, and O. Starbuch [review]. Dr. Dobb’s Journal of Software Tools, 21 (4):128–??, April 1996. CODEN DDJOEB. ISSN 1044-789X.

McCarthy:1997:UWM

[McC97a] V. McCarthy. Use the Web to manage your network. Datamation, 43(1):128–133, January 1997. CODEN DTMNAT. ISSN 0011-6963.

McClure:1997:IRS

[McC97b] Ben McClure. Internet Rx: Should I be checking downloaded Java applets for viruses or other potential security violations? Computers & Security, 16(8):677, ???? 1997. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.
sciencedirect.com/science/
article/pii/S016740489787592X.

McCormick:1997:LEN

[McC97c] John W. M. McCormick. Letter to the editor: No language for beginners. Computer, 30(4):7, April 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). Decries adoption of Java as a dominant language in beginning Computer Science courses.

McCarthy:1998:UEA

[McC98a] Pat McCarthy. Unlimited Enterprise Access: Java and VisualAge Generator. IBM Redbooks, 11400 Burnet Road, Austin, TX 78758-3493, USA, June 1998. ISBN ???? LCCN

???? ???? URL http://www.
redbooks.ibm.com/SG245246/
sg245246.html.

McCarty:1998:VDS

[McC98b] Bill McCarty. Visual Developer SQL Database Programming with Java: Creating Fast, Efficient Database Applications for the Web. Coriolis Group Books, Scottsdale, AZ, USA, 1998. ISBN 1-57610-176-2. xvii + 493 pp. LCCN QA76.9.D3M39367 1998. US$39.99, CDN$55.99.

URL http://www.coriolis.
com/cgi-win/KBSearch.exe;
 http://www.coriolis.com/
site/msie/books/ind/hpjsqudp.
htm.

McClure:1998:JKG

[McC98c] Ben McClure. Java kit gets better security. Computers & Security, 17(2):158, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.
sciencedirect.com/science/
article/pii/S0167404897820081.

McClure:1998:SCH

[McC98d] Ben McClure. Security companies hype up Java risks. Computers & Security, 17 (3):226, ???? 1998. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.
sciencedirect.com/science/
article/pii/S0167404898803313.

McClure:1998:TMJ

[McC98e] Ben McClure. Thwarting malicious Java attacks. Computers & Security, 17(4): 326, ???? 1998. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

                                        tronic). URL https://www.
                                        sciencedirect.com/science/
                                        article/pii/S0167404898800187.

McCluskey:1998:JPd

[McC98f] Glen McCluskey. Java I/O performance. ;login: the USENIX Association newsletter, 23(7):??, December 1998. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

                                        usenix.org/publications/login/
                                        1998-12/javaio.html.

McCluskey:1998:JPa

[McC98g] Glen McCluskey. Java performance. ;login: the USENIX Association newsletter, 23 (4), June 1998. CODEN LOGNEM. ISSN 1044-

                                        6397.  URL http://www.
                                        usenix.org/publications/login/
                                        1998-6/jperf.html.

McCluskey:1998:JPb

[McC98h] Glen McCluskey. Java performance. ;login: the USENIX Association newsletter, 23(5): ??, August 1998. CO-DEN LOGNEM. ISSN 1044-

                                        6397.  URL http://www.
                                        usenix.org/publications/login/
                                        1998-8/java_perf.html.

McCluskey:1998:JPc

[McC98i] Glen McCluskey. Java performance. ;login: the USENIX Association newsletter, 23(6): ??, October 1998. CO-DEN LOGNEM. ISSN 1044-

                                        6397.  URL http://www.
                                        usenix.org/publications/login/

1998-10/java_performance.
html.

McCluskey:1999:JPa

[McC99a] Glen McCluskey. Java performance. ;login: the USENIX Association newsletter, 24(2): ??, April 1999. CODEN LOGNEM. ISSN 1044-

6397.  URL http://www.
usenix.org/publications/login/
1999-4/java_perf.html.

McCluskey:1999:JPb

[McC99b] Glen McCluskey. Java performance. ;login: the USENIX Association newsletter, 24(4): ??, August 1999. CO-DEN LOGNEM. ISSN 1044-

6397.  URL http://www.
usenix.org/publications/login/
1999-8/features/performance.
html.

McCluskey:1999:JPc

[McC99c] Glen McCluskey. Java performance. ;login: the USENIX Association newsletter, 24(6): ??, December 1999. CO-DEN LOGNEM. ISSN 1044-

6397.  URL http://www.
usenix.org/publications/login/
1999-12/features/performance.
html.

McCluskey:1999:JPM

[McC99d] Glen McCluskey. Java performance: Memory fragmentation. ;login: the USENIX Association newsletter, 24(1):??, February 1999. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

usenix.org/publications/login/
1999-2/java_perf.html.

McComb:19xx:CHJ

[McCxxa] Gordon McComb. Het complete handboek JavaScript. Academic Service, Postbus 81, 2870 AB Schoonhoven, The Netherlands, 19xx. ISBN ???? ???? pp. LCCN ???? fl.

                                        109,00.  URL http://www.
                                        acaserv.nl;   http://www.
                                        acaserv.nl/5_0508.htm. In-

cludes CD-ROM.

McComb:19xx:JSC

[McCxxb] Gordon McComb. JavaScript Sourcebook: Crie Programas Interativos JavaScript para a World Wide Web. Makron Books do Brasil Editora Ltda., Rua Tabapuã 1348, São Paulo SP - 04533-004 - Brasil, 19xx. ISBN 85-346-0728-1. 760 pp. LCCN ???? R$ 85.00. URL

                                        http://safe.tesla.com.br/
                                        makron/options/livro.asp?
                                        id=9;   http://www.makron.

com/. Portuguese translation of [McC96d]. Includes CD-ROM.

McDowell:1998:RGJ

[McD98]  C. E. McDowell.  Reduc-

ing garbage in Java. ACM SIGPLAN Notices, 33(9):84–86, September 1998. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

McCarty:1999:JDO

[MCD99a] Bill McCarty and Luke Cassady-Dorion. Java Distributed Objects. Macmillan Computer Publishing, Indianapolis, IN, USA, 1999. ISBN 0-672-31537-8. xvii + 936 pp. LCCN QA76.73.J38M3528 1999. US$49.99.

McDonald:1999:UJS

[McD99b] Bruce McDonald. Using Java servlets with database connectivity. Linux Journal, 67:??, November 1999. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http://

noframes.linuxjournal.com/
lj-issues/issue67/3243.html.

McFarland:1997:EDI

[McF97] T. McFarland. Experiences developing internationalized applets with Java. In Unicode Consortium [Uni97b], pages A15–?? ISBN ???? LCCN ???? URL http:

//www.unicode.org/unicode/
iuc11/index.html;   http:
//www.unicode.org/unicode/
uni2book/u2ord.html.

McFarland:1998:DIA

[McF98a]  Tom McFarland.  Develop-

ing internationalized applications/applets with Java. In Unicode Consortium [Uni98a], page ?? ISBN ???? LCCN

????  URL http://www.
unicode.org/unicode/iuc12/
program.html.

McFarland:1998:DIS

[McF98b] Tom McFarland. Developing internationalized software with Java. In Unicode Consortium [Uni98b], page ?? ISBN ????

LCCN ???? URL http://www.

                                        unicode.org/unicode/iuc13/
                                        program.html.

McGibbon:1997:UAM

[McG97] Barry McGibbon. UNIX application management: Legacy reengineering with Java and the Web. UNIX Developer, 1(1): 31–34, January/February 1997. ISSN 1090-2279.

McHale:1997:VJT

[McH97]  Robert McHale.  Visual

J++ Training Videos. Key-Stone Learning Systems, ????, May 1997. LCCN ????

                                        US$90.  URL http://www.
                                        keylearnsys.com/outlines/
                                        j++.html.

McHale:19xx:JTV

[McHxx] Robert McHale. Java Training Videos. KeyStone Learning Systems, ????, 19xx. LCCN ???? US$90. URL http://www.

                                        keylearnsys.com/outlines/
                                        java1.html.

McKay:1995:NIP

[McK95a] Niall McKay. News: IBM to preload Unix system software on RS6000. IDG International News Service, page ??, December 18, 1995.

McKay:1995:NPG

[McK95b] Niall McKay. News: Power-Builder 5.0 gives increased Internet support. IDG International News Service, page ??, December 14, 1995.

McKeon:1997:FJF

[McK97] Maureen McKeon. Firms join forces for JSQL. SunExpert Magazine, 8(6):22–23, June 1997. ISSN 1053-9239. Discusses a proposed Java interface to SQL (Standard Query Language), and JDBC (Java-to-database connectivity).

McKeeman:1998:LJC

[McK98]  Bill McKeeman.  A lexer

for Java in C++. C/C++ Users Journal, 16(1):??, January 1998. CODEN CCUJEX. ISSN 1075-2838.

McLaughlin:1996:LAL

[McL96] J. McLaughlin. A look at the leading Java development tools. Software Economics Letter, 5 (11):4–5, November 1996. CO-DEN SECLE3. ISSN 1065-6146.

McMurdo:1996:HL

[McM96] G. McMurdo. HTML for the lazy. Journal of Information Science, Principles and Practice, 22(3):198–212, ???? 1996. CODEN JIOSED. ISSN 0165-5515 (print), 1741-6485 (electronic).

Meyer:1997:JVM

[MD97a] Jon Meyer and Troy Downing. The Java Virtual Machine. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1997. ISBN 1-56592-194-1. xxiv + 426 pp. LCCN QA76.73.J38M49

1997.  US$32.95.  URL

http://www.loc.gov/catdir/
enhancements/fy0915/97146808-
d.html;   http://www.ora.
com/www/item/javavm.html.

Munson:1997:SJF

[MD97b] Jonathan P. Munson and Prasun Dewan. Sync: a Java framework for mobile collaborative applications. Computer, 30(6): 59–66, June 1997. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

McCarthy:1998:BWJ

[MD98a] Martin McCarthy and Alligator Descartes. Building 3D worlds in Java and VRML. Prentice Hall Europe, New York, NY, USA, 1998. ISBN 0-13-748625-1 (paperback). 382 pp. LCCN QA76.73.J38M3527 1998.

McCarthy:1998:RAB

[MD98b] Martin McCarthy and Alligator Descartes. Reality architecture: building 3D worlds with Java and VRML. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1998. ISBN 0-13-748625-1. xiv + 382 pp. LCCN QA76.73.J38M3527 1998.

Machover:1999:NPHa

[MD99] Carl Machover and John Dill. New products: Hardware: Peripherals: Mita offers desktop color laser printer; Proview announces low priced monitors; digital flat-panel solution pack shipped; compact projector from Dukane. systems:

PC real-time 3D simulation introduced. boards: Mitsubishi vg500 does volume rendering. software: Web tools: 3DML released by Flatland; Sun delivers Java 2; doc-to-net introduced by Skyline Tools; software makes Web photo editing easy; JustEdit speeds up Web edits. authoring tools: Director 7 and Dreamweaver introduced; Famous offers animation solutions. CAD/CAM: 3D BuilderPro shipped; Toolbox/ SM available from Cimlogic; AutoPlant 97 V1.10 plant design tool; productivity booster for DataCad; enhancements to dv/MockUp functionality. visualization tools: TerraVista Version 2.0 announced. color: Color matching system by Color Solutions. IEEE Computer Graphics and Applications, 19(2):96–99, March/April 1999. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

http://dlib.computer.org/
cg/books/cg1999/pdf/g2096.
pdf.

Meyer:19xx:JVMb

[MDxx] Jon Meyer and Troy Downing. The Java Virtual Machine. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 19xx. ISBN 4-900900-63-X. LCCN ???? 4,500 yen. Japanese translation.

Mingins:1997:TPC

[MDM97] Christine Mingins, Roger Duke, and Bertrand Meyer, editors.

TOOLS 25: Proceedings of the 1997 Conference on Technology of Object-Oriented Languages and Systems Nov 24– 28 1997 1997 Melbourne, Australia. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1997. ISBN 0-8186-8485-2, 0-8186-8487-9 (microfiche). LCCN QA76.64.T66 1997b.

Meckler:1997:JIA

[Mec97] Andrew Meckler. Java and inter-applet communication. Dr. Dobb’s Journal of Software Tools, 22(10):46, 48, 50– 53, 103, October 1997. CODEN DDJOEB. ISSN 1044-789X.

Media:19xx:BC

[Medxx] Knowledge Media. Black Coffee. Knowledge Media, ????, 19xx. LCCN ???? ????

Meese:1996:OHJ

[Mee96] Philip David Meese. The one hour Java applet — everything you need to know to build your first Java applet. using this tutorial, at the end of one or two hours you will have a working Java bar-chart applet and HTML page. Datamation, 42 (5):51–61, March 1996. CODEN DTMNAT. ISSN 0011-6963.

Mehler:1997:CRJ

[Meh97] Mark Mehler. CSX: a railroad’s Java journey. Object Magazine, 7(1):63, March 1997. CODEN OBMAFO. ISSN 1055-3614.

Melford:1996:WNT

[Mel96] Bob Melford. Web and Net tactics: Looking behind the Java jive. Digital News and Review, 13(4):17–??, ???? 1996. ISSN 1065-7452.

Menkus:1996:JRC

[Men96] B. Menkus. Java raises control, audit, and security issues. EDPACS, 24(6):11–16, December 1996. CODEN EDPCDF. ISSN 0736-6981 (print), 1936-1009 (electronic).

Messmer:1995:NNU

[Mes95a] Ellen Messmer. News: Netscape to unveil multimedia Web browser, tools suite. Network World, page ??, September 15, 1995. ISSN 0887-7661 (print), 1944-7655 (elec-

tronic). URL http://www.
iquest.net/index.html.

Messmer:1995:NNIb

[Mes95b] Ellen Messmer. News: New Internet commerce tools launched. Network World, page ??, October 06, 1995. ISSN 0887-7661 (print), 1944-7655 (elec-

tronic). URL http://www.
iquest.net/index.html.

Messmer:1995:NSL

[Mes95c] Ellen Messmer. News: Sun to launch Internet software for commerce. Network World, page ??, May 19, 1995. ISSN 0887-7661 (print), 1944-7655 (electronic). URL http://www.

iquest.net/index.html.

Messmer:1995:NWS

[Mes95d] Ellen Messmer. News: Which standard for future Web development? Network World, page ??, December 11, 1995. ISSN 0887-7661 (print), 1944-7655 (electronic). URL http://

www.iquest.net/index.html.

Messmer:1996:DWJ

[Mes96] Ellen Messmer. Developers work out Java kinks. Network World, 13(34):35, August 19, 1996. ISSN 0887-7661 (print), 1944-7655 (electronic).

URL http://www.nwfusion.

com/. Describes problems in getting the same look-and-feel on different platforms from AWT.

Messmer:1997:SGR

[Mes97] Ellen Messmer. Symanticec gives remote control software NT boost, debuts Java tools: New versions of Visual Café for Java boast Navigator links. Network World, 14(33):15, August 18, 1997. ISSN 0887-7661 (print), 1944-7655 (electronic).

Metcalfe:1995:CEH

[Met95a] Bob Metcalfe. Column: From the ether — hit the iway, but leave the microwave. InfoWorld, page ??, December 12, 1995. CODEN INWODU. ISSN 0199-6649. URL http://www.

infoworld.com/.

Metcalfe:1995:CEI

[Met95b] Bob Metcalfe. Column: From the ether — info appliances. InfoWorld, page ??, October 24, 1995. CODEN INWODU. ISSN 0199-6649. URL http://www.

infoworld.com/.

Metcalfe:1995:CET

[Met95c] Bob Metcalfe. Column: From the ether — Telescript on the Web. InfoWorld, page ??, November 06, 1995. CODEN INWODU. ISSN 0199-6649.

URL http://www.infoworld.
com/.

Metrowerks:1995:MCS

[Met95d]  Inc. Metrowerks.  Metrow-

erks collaborates with Sun Microsystems to provide Java programming tools in CodeWarrior product for Macintosh. press release, November 10, 1995. URL

http://www.metrowerks.com/
news/press/java.html.

Metz:1996:JJN

[Met96]  Eldon Metz.  Java, JFac-

tory, and network development. Dr. Dobb’s Journal of Software Tools, 21(9):78, 80–82, 85– 87, September 1996. CODEN DDJOEB. ISSN 1044-789X.

Metamorphic:1998:MCC

[Met98] Metamorphic. Metamorphic COBOL Converter: COBOL to Visual Basic, COBOL to C++, COBOL Caffeinator (a COBOL to Java converter),

1998.  URL http://www.
metamorphic.com/default.htm;
 http://www.metamorphic.com/
html/cobol.html. This is an

expensive commercial translator, available as a service charged by lines converted, or as a complete product to run locally.

Meyer:1996:DCC

[Mey96a] Helen Meyer. Developers can control Java access with firewall. Computers & Security, 15(8):688, ???? 1996. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

                                        tronic). URL https://www.
                                        sciencedirect.com/science/
                                        article/pii/S0167404897897143.

Meyer:1996:DWA

[Mey96b] Helen Meyer. Developers worried about safety of Java and ActiveX. Computers & Security, 15(8):687, ???? 1996. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

                                        tronic). URL https://www.
                                        sciencedirect.com/science/
                                        article/pii/S0167404897897106.

Meyer:1997:OOS

[Mey97]  Bertrand Meyer.  Object-

Oriented Software Construction, 2nd Ed. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, second edition, 1997. ISBN 0-13-629155-4. xxvii + 1254 pp. LCCN QA76.64 .M493 1997. US$70. URL http://

                                        www.prenhall.com/allbooks/
                                        ptr_0136291554.html.

Meyer:1998:EUA

[Mey98a] Helen Meyer. eSafe ups ante for Java, security. Computers & Security, 17(1):57, ???? 1998.

CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.
sciencedirect.com/science/
article/pii/S0167404897802532.

Meyer:1998:JAT

[Mey98b] Helen Meyer. Java authentication token. Computers & Security, 17(1):59, ???? 1998. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.
sciencedirect.com/science/
article/pii/S0167404897802623.

Meyer:1998:SCSa

[Mey98c] Helen Meyer. Sun changes security framework for Java development kit. Computers & Security, 17(1):58–59, ???? 1998. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.
sciencedirect.com/science/
article/pii/S0167404897802593.

Meyer:1999:SBJ

[Mey99]  Helen Meyer.  Server-based

Java security products help guard your enterprise flank. Computers & Security, 18(2): 153–154, ???? 1999. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.
sciencedirect.com/science/
article/pii/S0167404899800479.

McCandless:1996:ISW

[MF96a] Michael McCandless and Giovanni Flammia. Internet services: Where the Web is headed: Java’s star turn. IEEE expert: intelligent systems and their applications, 11(2):82–83, April 1996. CODEN IEEXE7. ISSN 0885-9000.

McGraw:1996:JS

[MF96b] Gary McGraw and Edward Felten. Java security. SIGS Books and Multimedia, 71 W 23rd St., New York, NY 10010, USA, 1996. ISBN 1-884842-72-0. 46 pp. LCCN ???? US$85. URL

http://www.sigs.com/books/
wp_javasec.html.

McGraw:1997:JST

[MF97a] Gary McGraw and Edward Felten. Java security and type safety — Java’s sophisticated security mechanism performs run-time safety checks before it allows a downloaded applet to execute. BYTE Magazine, 22 (1):63–64, January 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

McGraw:1997:JSH

[MF97b] Gary McGraw and Edward Felten. Java Security: Hostile Applets, Holes, and Antidotes. Wiley Computer Publishing. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 1997. ISBN 0-471-17842-X (paperback). xiv + 192 pp. LCCN QA76.73.J38M354

1997.  US$19.95.  URL

http://www.mindq.com/java/
javasecu.html;   http://www.
wiley.com/compbooks/catalog/
17842-X.htm. A companion

CD-ROM “Java Security: Managing the Risks” is also available from MindQ Publishing.

McGraw:1997:JSM

[MF97c] Gary McGraw and Edward Felten. Java Security: Managing the Risks. MindQ Publishing, Inc., 450 Springpark Place Ste. 1200, Herndon, VA 20170, USA, Tel: +1 703 708-9380, 1997. ISBN 1-57559-012-3. ???? pp. LCCN ???? US$59.95. URL

http://www.mindq.com/java/
javasecu.html. CD ROM.

McGraw:1997:LEL

[MF97d] Gary McGraw and Edward Felten. Letter to the editor: Lame review? BYTE Magazine, 22(6):20, June 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic). Authors’ response to [Kay97].

McGraw:1998:SJG

[MF98] Gary McGraw and Edward Felten. Securing Java: Getting Down to Business with Mobile Code. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 1998. ISBN 0-471-31952-X. xii + 324 pp. LCCN QA76.73.J38 M354 1999.

McGraw:1999:SJG

[MF99] Gary McGraw and Edward W.

Felten. Securing Java: Getting Down to Business with Mobile Code. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 1999. ISBN 0-471-31952-X. xii + 324 pp. LCCN QA76.73.J38M354 1999. US$34.99.

Misgeld:1996:EJ

[MG96] Wolfgang Misgeld and Jürgen Gruber. Einstieg in Java. Carl Hanser, München, Germany, 1996. ISBN 3-446-18836-3. 187 pp. LCCN ???? 56 DM.

URL http://www.hanser.de/
computer/buecher/mi18836.

htm. Includes CD-ROM.

Mintchev:1997:ABN

[MG97a] S. Mintchev and V. Getov. Automatic binding of native scientific libraries to Java. Lecture Notes in Computer Science, 1343:129–??, 1997. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mintchev:1997:TPM

[MG97b] S. Mintchev and V. Getov. Towards portable message passing in Java: Binding MPI. Lecture Notes in Computer Science, 1332:135–??, 1997. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

McCarty:1998:OPJ

[MG98] Bill McCarty and Stephan Gilbert. Objectgeorienteerd programmeren in Java 1.1. Academic Service, Postbus 81, 2870 AB Schoonhoven, The Netherlands, January 1998. ISBN 90-395-0856-9. ca. 1000 pp. LCCN ???? fl. 119,00. URL http:/

/acaserv.nl/bestel/5_0856.
htm.

Mackenzie:1998:LPS

[MGL^{+}98] Don Mackenzie, Andrew J.

Gryc, Graziano Lo Russo, Gary Clouse, C. J. Hinke, Bruce E. Hogman, Thomas Fleischer, and John Graham-Cumming. Letters: The passport system does work; real-time sound; C++ versus Java; online op-ed; hard encryption; Y2K; vercheck update. Dr. Dobb’s Journal of Software Tools, 23(12):12, 16– 17, December 1998. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/.

Moore:1997:LEJ

[MH97a] David L. Moore and Tom R.

Halfhill. Letters to the Editor: On Java. BYTE Magazine, 22(4):19–20, April 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Moran:1997:LEA

[MH97b] Tom Moran and Tom R.

Halfhill. Letter to the Editor: Ada speaks Java. BYTE Magazine, 22(4):20, April 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Munck:1997:AJW

[MH97c] Robert G. Munck and Richard F. Hilliard II. Ada and Java on the WWW. ACM SIGADA Ada Letters, 17(3):3–16, May/ June 1997. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Monson-Haefel:1999:EJ

[MH99] Richard Monson-Haefel. Enterprise JavaBeans. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1999. ISBN 1-56592-605-6. xvi + 317 pp. LCCN QA76.76.J38 M66 1999. US$32.95. URL http:

                                        //www.oreilly.com/catalog/
                                        entjbeans.

Maybee:1996:MID

[MHO96] M. J. Maybee, D. M. Heimbigner, and L. J. Osterweil. Multilanguage interoperability in distributed systems experience report. In IEEE [IEE96k], pages 451–463. ISBN 0-8186-7247-1. LCCN QA 76.758 I5725 1996. IEEE Cat. No.96CB35918.

Mercuri:1998:UHJ

[MHP98] R. Mercuri, N. Herrmann, and J. Popyack. Using HTML and JavaScript in introductory programming courses. In Joyce [Joy98], pages 176–180. ISBN 0-89791-994-7. ISSN 0097-8418 (print), 2331-3927 (electronic). LCCN QA1 .A86 v.30, no.1.

Microsoft:1995:ISD

[Mic95] Microsoft Corporation. Internet strategy day. press release, December 7, 1995. URL

                                        http://www.microsoft.com/
                                        internet/press.htm.

Microsoft:1996:MVJ

[Mic96] Microsoft Corporation. Microsoft Visual J++, 1996. Includes one CD ROM, getting started card, and poster.

Michaels:1997:CCJ

[Mic97a]  Laura Michaels.  C/C++

and Java library equivalents. C/C++ Users Journal, 15(6): 53–55, June 1997. CODEN CCUJEX. ISSN 1075-2838.

Michaels:1997:ONJ

[Mic97b] Sean Michaels. Official Netscape JFC Developer’s Guide. Ventana Press, Chapel Hill, NC, USA, 1997. ISBN 1-56604-756-0. ???? pp. LCCN ???? US$50. URL http:

//www.vmedia.com/bookshop/
book.html?isbn=1566047560.

Mills:1995:NBS

[Mil95a] Elinor Mills. News: Borland, Spyglass license Sun’s Java Internet technology. IDG International News Service, page ??, November 08, 1995.

Mills:1995:NMS

[Mil95b] Elinor Mills. News: Macromedia, Sun join to put multimedia on the Internet. IDG International News Service, page ??, October 31, 1995.

Mills:1995:NPO

[Mil95c] Elinor Mills. News: PC outlook:

Servers, portables, Starfish and Netscape. IDG International News Service, page ??, December 05, 1995.

Mills:1995:NVD

[Mil95d] Elinor Mills. News: Vendors debate Internet appliance feasibility. IDG International News Service, page ??, December 08, 1995.

Miller:1998:MVJ

[Mil98] Kevin Miller. Microsoft VM for Java. In ACM [ACM98a], page ?? CODEN CPEXEI. ISSN 1040-3108. LCCN ????

URL http://www.cs.ucsb.
edu/conferences/java98/program.

html. Invited talk.

Mills:1999:ANS

[Mil99] Kevin L. Mills. AirJava: Networking for smart spaces. In USENIX [USE99e], page ?? ISBN 1-880446-31-6. LCCN TK7895.E42 W67 1999. URL

http://www.usenix.org/publications/
library/proceedings/es99/
mills.html.

MindQ:1996:JLI

[Min96a] MindQ. Java Live: An Interactive Learning Tool. Ziff-Davis Press, Emeryville, CA, USA, 1996. ISBN 1-56276-482-9. ???? pp. LCCN ???? US$49.99.

MPI:1996:HJP

[Min96b] MindQ Publishing, Inc. Hands on Java programming using Microsoft Visual J++, 1996. Includes CD-ROM.

Mintert:1996:JGE

[Min96c] Stefan Mintert. JavaScript: Grundlagen und Einführung. Addison-Wesley, Reading, MA, USA, 1996. ISBN 3-8273-1087-3 (??invalid ISBN??). 300 pp. LCCN ???? 60 DM. URL

http://www.addison-wesley.
de/;   http://www.addison-
wesley.de/KD/PR/1087.html.

MindQ:1997:AJT

[Min97a] MindQ. Advanced Java Topics. MindQ Publishing, Inc., 450 Springpark Place Ste. 1200, Herndon, VA 20170, USA, Tel: +1 703 708-9380, 1997. US$995.00. Interactive traing system. Order number T10216128, Part number 901715.

MindQ:1997:DTJ

[Min97b] MindQ. Developer Training For Java. MindQ Publishing, Inc., 450 Springpark Place Ste. 1200, Herndon, VA 20170, USA, Tel: +1 703 708-9380, 1997. US$1595.00. Interactive traing system. Order number IT10216129, Part number 901704.

MindQ:1997:EJT

[Min97c] MindQ. Essential Java Training. MindQ Publishing, Inc., 450 Springpark Place Ste. 1200, Herndon, VA 20170, USA, Tel: +1 703 708-9380, 1997. US$795.00. Interactive traing system. Order number IT10216127, Part number 901746.

MindQ:19xx:JDP

[Minxxa] MindQ. Java Database Programming. MindQ Publishing, ????, 19xx. ISBN ???? ???? pp. LCCN ????

MindQ:19xx:JPS

[Minxxb] MindQ. Java Programming Series. MindQ Publishing, Inc., 450 Springpark Place Ste. 1200, Herndon, VA 20170, USA, Tel: +1 703 708-9380, 19xx. ISBN JAVPROSER-. LCCN ???? ????

MindQ:19xx:MDT

[Minxxc] MindQ. MindQ’s Developer Training for Java. MindQ Publishing, Inc., 450 Springpark Place Ste. 1200, Herndon, VA 20170, USA, Tel: +1 703 708-9380, 19xx. ISBN ???? ???? pp. LCCN ???? US$995.

                                        URL http://www.mindq.com/
                                        java/javaover.html.

Ming:19xx:JL

[Minxxd] Chang Shin Ming. The Java Language. Rulin Books, ????, 19xx. ISBN 957-652-900-X. ???? pp. LCCN ???? 250 NT$.

Ming:19xx:WJ

[Minxxe] Chen Tjien Ming. World Of Java. LIWIL Publishing Co Ltd., ????, 19xx. ISBN 957-8982-64-X. ???? pp. LCCN ???? 320 NT$.

Mione:1996:XOB

[Mio96] A. N. Mione. X opens on Broadway. Digital Systems Report, 18 (4):8–10, Fall 1996. CODEN DSREFP. ISSN 1086-9638.

Mitchell:1996:JBFb

[Mit96a] James G. Mitchell. JavaOS: back to the future. Operating Systems Review, 30(SI):1, October 1996. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Mitchell:1996:JBFa

[Mit96b] Jim Mitchell. Java OS: Back to the future. In USENIX [USE96a], pages 1–?? ISBN ???? LCCN ????

Mitchell:1997:JBF

[Mit97a] Jim Mitchell. Java OS: Back to the future, 1997. URL http://

www.usenix.org/publications/
library/proceedings/osdi96/
invited_talks/mitchell/abstract.

html. Unpublished invited talk at the USENIX 2nd Symposium on OS Design and Implementation (OSDI ’96), October 28–31, 1996, Seattle, Washington.

Mitchell:1997:DEA

[Mit97b]  Michael A. Mitchell.  De-

veloping Enterprise Applications with Visual J++. Waite Group Press, Corte Madera, CA, USA, 1997. ISBN 1-57169-085-9. xxvii + 522 pp. LCCN QA76.73.J38M58 1997. US$50. URL http://

merchant.superlibrary.com:
8000/catalog/hg/PRODUCT/PAGE/
15716/bud/1571690859.html.

Mizuno:1999:SAD

[Miz99] Masaaki Mizuno. A structured approach for developing concurrent programs in Java. Information Processing Letters, 69(5): 233–238, March 12, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

MacKay:1998:PCT

[MJ98] Stephen A. MacKay and J. Howard Johnson, editors. Proceedings of CASCON’98: Toronto, Ontario, Canada, 30 November–3 December 1998. IBM Toronto Laboratory, Centre for Advanced Studies, Toronto, ON, Canada, 1998. LCCN TK 5105.5 .C36 1998.

Meehan:1999:CLF

[MJ99] Gary Meehan and Mike Joy. Compiling lazy functional programs to Java bytecode. Software—Practice and Experience, 29(7):617–645, June 1999. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=62003575;
                                         http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=62003575&PLACEBO=IE.pdf.

Morrison:1998:APO

[MJA98] Ron Morrison, Mick Jordan, and Malcolm Atkinson, editors. Advances in Persistent Object Systems: Proceedings of the International Workshop on Persistent Object Systems (POS) and the International Workshop on Persistence & Java (PJAVA). Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 1998. ISBN 1-55860-585-1. LCCN QA76.9.D3 I59 1998. US$54.95.

Marmel:1997:COJ

[MK97] Elaine Marmel and Sherry Kinkoph. Corel Office for Java 6 in 1. Que Corporation, Indianapolis, IN, USA, June 1997. ISBN 0-7897-1314-4 (softcover). ???? pp. LCCN ???? US$26.99.

Macketanz:1998:JRP

[MK98] R. Macketanz and W. Karl.

JVX — a rapid prototyping system based on Java and FPGAs. Lecture Notes in Computer Science, 1482:99–??, 1998. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Montgomery:1999:SOS

[MK99] Michael Montgomery and Ksheerabdhi Krishna. Secure object sharing in Java Card. In USENIX [USE99d], page ?? ISBN 1-880446-34-0. LCCN TK7895.S62 U84 1999. URL

http://www.usenix.org/publications/
library/proceedings/smartcard99/
montgomery.html.

Maley:1996:FSA

[MKS^{+}96] Maley, L. Kilpatrick, E. W.

Schreiner, N. S. Scott, and G. H. F. Diercksen. The formal specification of abstract data types and their implementation in Fortran 90: implementation issues concerning the use of pointers. Computer Physics Communications, 98(1-2):167–180, October 1996. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Miyoshi:1998:RTJ

[MKT98] Akihiko Miyoshi, Takuro Kitayama, and Hideyuki Tokuda. A real-time Java server for real-time Mach. Parallel and Distributed Computing Practices, 1(2):??, ???? 1998. CODEN ???? ISSN 1097-

2803.  URL http://www.
cs.okstate.edu/~pdcp/vols/
vol01/vol01no2abs.html#miyoshi.

McGrath:1997:JTA

[MLF97] Robert E. McGrath, Xinjian Lu, and Michael Folk. Java applications using NCSA HDF files. Concurrency: Practice and Experience, 9(11): 1113–1125, November 1997. CODEN CPEXEI. ISSN 1040-3108. URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=13825;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=13825&PLACEBO=IE.pdf. Spe-

cial Issue: Java for computational science and engineering

— simulation and modeling II.

Martin:1998:UJS

[MLM98] Jared Martin, Lyle N. Long, and Kevin M. Morooney. Using Java for scientific programming and electromagnetics. Computers in Physics, 12(6):601– ??, November 1998. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (elec-

tronic). URL https://aip.
scitation.org/doi/10.1063/
1.168742.

Martin:1999:JDI

[MM99] David H. Martin and Johnny Martin. Java and digital images. Dr. Dobb’s Journal of Software Tools, 24(5):72, 74–76, 78–79, May 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ftp/1999/1999_
05/snapshot.txt;   http://
www.ddj.com/ftp/1999/1999_
05/snapshot.zip.

Muller:1997:HFE

[MMBC97] Gilles Muller, Bárbara Moura, Fabrice Bellard, and Charles Consel. Harissa: a flexible and efficient Java environment mixing bytecode and compiled code. In USENIX [USE97b], pages 1–20. ISBN 1-880446-86-3. LCCN QA76.64 .U845 1997.

URL http://www.irisa.fr/
compose/harissa/coots_web.
ps.gz;   http://www.irisa.
fr/compose/harissa/harissa.
html;   http://www.oasis.
leo.org/java/development/
java-to-c/Harissa.dsc.html.

Moreira:1998:CFC

[MMG98] J. E. Moreira, S. P. Midikiff, and M. Gupta. A comparison of Fortran, C/C++, and Java for numerical computing. IEEE Antennas and Propagation Magazine, 40(5):102–105, October 1998. CODEN IAP-MEZ. ISSN 1045-9243.

Moreira:1999:PDM

[MMGL99] Jose Moreira, Sam Midkiff, Manish Gupta, and Rick Lawrence. Parallel data mining using the array package for Java. In ACM [ACM99c], page ??

McConville:1996:WWL

[MMJ96] D. McConville, J. Magid, and P. Jones. So wide a Web, so little time. Educom review, 31 (3):44–48, May–June 1996. CO-DEN EDREEW. ISSN 1045-9146.

Mendelzon:1996:QWW

[MMM96] A. O. Mendelzon, G. A. Mihaila, and T. Milo. Querying the World Wide Web. In IEEE [IEE96o], page ?? ISBN 0-8186-7475-X. LCCN QA76.58 .I544 1996. IEEE Computer Society Press order number PR07475. IEEE catalog number 96TB100085.

Midkiff:1998:OAR

[MMS98] S. P. Midkiff, J. E. Moreira, and M. Snir. Optimizing array reference checking in Java programs. IBM Systems Journal, 37(3):409–??, ???? 1998. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

URL http://www.almaden.
ibm.com/journal/sj/373/midkiff.
html.

Macary:1996:PJ

[MN96a] Jean-François Macary and Cédric Nicolas. Programmation Java. Eyrolles, Paris, France, 1996. ISBN 2-212-08916-3. 236 pp. LCCN ???? 168

FF. URL http://www.estp.
fr/eyrolles;   http://www.

fisystem.fr/java/javabook.
htm#A3;   http://www.fisystem.
fr/java/{javabook}.htm. In-

cludes CD-ROM.

Moulding:1996:PAC

[MN96b] M. R. Moulding and M. J. Newton. Progressive animation of CORE requirements via an underlying formal model. In Samson et al. [SMEN96], page ?? ISBN ???? LCCN ????

Macary:1997:PJ

[MN97] Jean-François Macary and Cédric Nicolas. Programación Java. Ediciones Gestión 2000, S.A., ????, 1997. ISBN 84-8088-128-3. ???? pp. LCCN ???? 3995 Ptas. URL http://www.

fisystem.fr/java/javabook.
htm.

Matsumura:1996:JDW

[MNO^{+}96] Miko Matsumura, Steven Nathan, Steve Outtrim, Tengku Mohd Azzman Shariffadeen, and Francis Yeoh. Java development on the web, March 21, 1996. 1 videocassette (60 min.). Order number Sun-20.

McIntosh:1996:TJM

[MO96] Harry McIntosh and Lynnzy Orr. Talk Java to Me: The Interactive Click, Listen, and Learn Guide to Java Programming. Waite Group Press, Corte Madera, CA, USA, 1996. ISBN 1-57169-044-1. 200 pp. LCCN QA76.73.J38M355 1996. US$24.99. URL http:

//www.mcp.com/cgi-bin/bag?

isbn=1-57169-044-1&last=/
bookstore.

McGhan:1998:CPP

[MO98] Harlan McGhan and Mike O’Connor. Computing practices: PicoJava: a direct execution engine for Java bytecode. Computer, 31(10):22–30, October 1998. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

http://dlib.computer.org/
co/books/co1998/pdf/rx022.
pdf;   http://www.computer.
org/computer/co1998/rx022abs.
htm.

Moffat:1997:JJA

[Mof97] David Moffat. Java is not just another language. SunServer, 11(9):21, September 1997. ISSN 1091-4986.

Mohseni:1996:WDP

[Moh96] Piroz Mohseni. Web Database Primer Plus: Connect Your Database to the World Wide Web Using HTML, CGI, and Java. Waite Group Press, Corte Madera, CA, USA, 1996. ISBN 1-57169-070-0 (softcover). xxvii + 482 pp. LCCN QA76.9.D26M66 1996. US$44.99. Includes CD-ROM.

Mohseni:1998:OSG

[Moh98] Piroz Mohseni. Operating systems — Glasgow enhances JavaBeans. BYTE Magazine, 23 (2):51–52, February 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Moncur:1996:LEA

[Mon96] Michael Moncur. Letter to the editor: Adventures in JavaScript. BYTE Magazine, 21(12):20, December 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Montague:1997:JEJ

[Mon97] Bruce R. Montague. JN: OS for an embedded Java Network Computer — supporting Java’s Virtual Machine on a single-chip embedded PC attached to the Internet. IEEE Micro, 17(3):54–60, May/ June 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//pascal.computer.org/mi/
books/mi1997/pdf/m3054.pdf.

Moncur:1999:WPJ

[Mon99] Michael Moncur. Web Programming with JavaScript, with CD ROM. Sams.net, Indianapolis, IN 46268, USA, January 1999. ISBN 1-57521-122-X (softcover). ???? pp. LCCN ???? US$39.99.

Moncur:19xx:J

[Monxx] Michael Moncur. JavaScript.

Mondadori Informatica, ??, Italy, 19xx. ISBN ????

????  pp.  LCCN ????

URL http://slack.lne.com/
lemay/index.html;   http:/
/www.mondadori.com/;   http:
//www.starlingtech.com/books/
javascript/author.html.

Moore:1995:NWT

[Moo95] Steve Moore. News: Web technologies promise to eliminate enterprise woes. Computerworld, page ??, August 11, 1995. CODEN CMPWAB. ISSN 0010-4841.

Moore:1998:JQH

[Moo98] L. Richard Moore. Java Q&A: How do I create a streaming audio Java applet? Dr. Dobb’s Journal of Software Tools, 23 (5):122–124, May 1998. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/1998/1998_05/
                                        jqa598.txt;   http://www.
                                        ddj.com/ftp/1998/1998_05/
                                        jqa598.zip.

Morin:1995:OW

[Mor95] Richard Morin. Oak and WebRunner. SunExpert Magazine, 6(2):??, February 1995. ISSN 1053-9239. URL http://java.

                                        sun.com/./press/940200sunexpert/

index.html. Oak was the internal name at Sun Microsystems for this project, but was changed to Java for trademark reasons.

Morin:1996:INJ

[Mor96a] Richard Morin. The Internet Notebook Java and JavaScript. UNIX review, 14(8):87–??, July 1, 1996. CODEN UNRED5. ISSN 0742-3136.

Morrison:1996:TYI

[Mor96b] Michael Morrison. Teach Yourself Internet Game Programming With Java in 21 Days. Sams.net, Indianapolis, IN 46268, USA, September 1996. ISBN 1-57521-148-3. xvi + 416 pp. LCCN QA76.76.C672M677 1996. US$39.99. URL http://

merchant.superlibrary.com:
8000/catalog/hg/PRODUCT/PAGE/
15752/bud/1575211483.html;
http://www.mcp.com/30549809800454/
cgi-bin/bag?isbn=1-57521-
148-3&last=/bookstore;   http:
//www.mcp.com/samsnet/.

Morgenthal:1997:BDJ

[Mor97a] Jeffrey Morgenthal. Building Distributed Java Applications. McGraw-Hill, New York, NY, USA, September 1997. ISBN 0-07-913679-6 (softcover). ???? pp. LCCN ???? US$44.95. Includes CD-ROM.

Morrison:1997:ECG

[Mor97b] Gale B. Morrison. Engineering center to give Java a boost. Electronic News, 43(2183):10, 60, September 1, 1997. CO-DEN ELNEAU. ISSN 0013-4937, 1061-6624.

Morrison:1997:SJE

[Mor97c]  Gale B. Morrison.  Sun’s

JavaEngine eyes ‘thin clients’. Electronic News, 43(2182):16, August 25, 1997. CODEN EL-NEAU. ISSN 0013-4937, 1061-6624.

Morrison:1997:JI

[Mor97d] Michael Morrison. Java 1.1 für Insider. Markt & Technik, Haar bei München, Germany, August 1997. ISBN 3-8272-2007-6. 1112 pp. LCCN ????

99,95 DM. URL http://www.
mut.ch/html/buchtext.htx?
&x_pp=STb4XDV284xv2gpf&x_
pk=VpLvjnsx7b5V&x_pa=B&x_
pr=oS&x_artikelnr=22007;
http://www.mut.ch/profit.
htx?&x_pk=kd84mf7ms9dm.

Morrison:1997:JBT

[Mor97e]  Michael Morrison.  Java

Beans: Technik, Konzepte, Beispiele. Markt & Technik, Haar bei München, Germany, 1997. ISBN 3-8272-5284-9. ???? pp. LCCN ????

59,95 DM. URL http://www.
mut.ch/html/buchtext.htx?
&x_pp=STb4XDV284xv2gpf&x_
pk=Direkt&x_pa=&x_pr=Fr&x_
artikelnr=25284;   http:/
/www.mut.ch/profit.htx?&x_
pk=kd84mf7ms9dm.

Morrison:1997:JDP

[Mor97f]  Michael Morrison.  Java

Database Programming Lab. Prima Publishing, Rocklin, CA, USA, 1997. ISBN 0-7615-1040-0. 500 pp. LCCN ???? US$40.00. Includes CD-ROM.

Morrison:1997:PJ

[Mor97g] Michael Morrison. Presenting JavaBeans. Sams.net, Indianapolis, IN 46268, USA, 1997. ISBN 1-57521-287-0. xix + 352 pp. LCCN QA76.76.J38 M6743 1997. US$35; 70,00

DM. URL http://merchant.
superlibrary.com:8000/catalog/
hg/PRODUCT/PAGE/15752/bud/

1575212870.html;   http:/
/www.mut.ch/profit.htx?&x_
pk=kd84mf7ms9dm.  Includes

CD-ROM.

Morgan:1998:BDA

[Mor98a] Bryan Morgan. Building distributed applications with Java and CORBA. Dr. Dobb’s Journal of Software Tools, 23 (4):94, 96–99, 104–105, April 1998. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ftp/1998/1998_
04/corbajav.txt;   http://
www.ddj.com/ftp/1998/1998_
04/corbajav.zip.

Morgan:1998:UJ

[Mor98b] Michael Morgan. Using Java 1.2. Que Corporation, Indianapolis, IN, USA, June 1998. ISBN 0-7897-1627-5. xi + 843 pp. LCCN QA76.73.J38 M6725 1998. US$30. URL

http://iq-mcp3.iquest.net/
catalog/mcp/PRODUCT/PAGE/
07897/bud/0789716275.html.

Moreno:19xx:PJ

[Morxxa] Oscar González Moreno. Programación en JavaScript. Anaya Multimedia, Madrid, Spain, 19xx. ISBN ???? ???? pp. LCCN ???? Ptas

1695. URL http://www.anaya.
es/Catalogo/CGA?act=L&art=
2335130&ed=2300;   http://
www.anaya.es/Catalogo/CGA?
act=Q&autor_id=09032&max_
rows=all;   http://www.anayamultimedia.
es/.

Morrison:19xx:JI

[Morxxb] Michael Morrison. Java 1.1 für Insider. Markt & Technik, Haar bei München, Germany, 19xx. ISBN 3-8272-2007-6 (??invalid ISBN??). LCCN ????

                                        100 DM. URL http://www.
                                        mut.ch/html/buchtext.htx?
                                        &x_pp=STb4XDV284xv2gpf&x_
                                        pk=VpLvjnsx7b5V&x_pa=B&x_
                                        pr=oS&x_artikelnr=22007.

Morrison:19xx:JSE

[Morxxc] Michael Morrison. Java 1.1 Secrets d’Experts. Simon & Schuster Macmillan (France), Paris, France, 19xx. ISBN 2-7440-0194-5. ???? pp. LCCN ???? ???? URL http://

                                        www.ssm.fr/catalog/fiches/
                                        0194.htm.

Morrison:19xx:JT

[Morxxd] Michael Morrison. JavaBeans-Technik. Markt & Technik, Haar bei München, Germany, 19xx. ISBN 3-8272-5284-9 (??invalid ISBN??). LCCN ????

                                        60 DM. URL http://www.
                                        mut.ch/html/buchtext.htx?
                                        &x_pp=STb4XDV284xv2gpf&x_
                                        pk=Direkt&x_pa=&x_pr=Fr&x_
                                        artikelnr=25284.

Morrison:19xx:VJD

[Morxxe] Mike Morrison. Visual J++ Database Programming Lab. Prima Publishing, Rocklin, CA, USA, 19xx. ISBN 0-7615-1045-1. ???? pp. LCCN ???? ????

Moseley:1996:DJP

[Mos96] Tom Moseley. Dummies 101:

Java Programming. I D G Books Worldwide, Indianapolis, IN, USA, September 1996. ISBN 0-7645-0068-6. ???? pp. LCCN ???? US$24.99.

Mosley:1998:PJC

[Mos98a] D. H. Mosley. Perspectives for Java-based computational quantum chemistry. International Journal of Quantum Chemistry, 70(1):159–??, ???? 1998. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mosley:1998:PJB

[Mos98b] David H. Mosley. Perspectives for Java-based computational quantum chemistry. International Journal of Quantum Chemistry, 70(1):159–165, ???? 1998. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=74360;
http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=74360&PLACEBO=IE.pdf. Spe-

cial Issue: The Ninth International Congress of Quantum Chemistry (Part I of II). Issue Edited by Keiji Morokuma, Ernest R. Davidson, Henry F. Schaefer III.

Moss:1998:JS

[Mos98c]  Karl Moss.  Java Servlets.

McGraw-Hill, New York, NY, USA, July 1998. ISBN 0-07-913779-2. LCCN ????

US$45. URL http://www.pbg.
mcgraw-hill.com/betabooks/
moss/moss-home.html.

Murray:1996:VJH

[MP96] William H. Murray and Chris H. Pappas. The Visual J++ handbook. AP Professional, Boston, MA, USA, October 1996. ISBN 0-12-511915-1. xv + 574 pp. LCCN A76.73.J38 M87 1996.

US$40.  URL http://www.
apcatalog.com/cgi-bin/AP?
ISBN=0125119151&LOCATION=
US&FORM=FORM2.

Murray:1997:IPP

[MP97a] William H. Murray and Chris H. Pappas. internet.profi.praxis Visual J++ (Übersetzung). Sybex Verlag GmbH, Erkrather Straße 345-349, 40231 Düsseldorf, Germany, 1997. ISBN 3-8155-2016-9. 555 pp. LCCN ???? 59

DM. URL http://www.sybex.
de/bestell/tp/tp02016.htm.

Murray:1997:JPG

[MP97b] William H. Murray and Chris H. Pappas. JBuilder Programmer’s Guide. AP Professional, Boston, MA, USA, August 1997. ISBN 0-12-511920-8 (softcover). ???? pp. LCCN ???? US$35.95. URL http://www.

apcatalog.com/cgi-bin/AP?
ISBN=0125119208&LOCATION=
US&FORM=FORM2. Publication

cancelled.

Murray:1998:JHU

[MP98] William H. Murray and Chris H.

Pappas. Javascript and HTML 4.0 user’s resource. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1998. ISBN 0-13-977422-X. ???? pp. LCCN QA76.73.J39M87 1998.

Murray:1999:JHU

[MP99] William H. Murray and Chris H.

Pappas. Javascript and HTML 4.0 user’s resource. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1999. ISBN 0-13-977422-X. xxi + 521 pp. LCCN QA76.73.J39M87 1998.

URL http://www.phptr.com/
ptrbooks/ptr_013977422x.html.

Morris:19xx:IWH

[MPxx] Mary E. S. Morris and John A.

Pew. Instant WebSite with HTML and Java Applets Boxed Set. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 19xx. ISBN 0-13-270067-0. ???? pp. LCCN ???? US$70.

Mowbray:1997:ICD

[MR97] Thomas J. Mowbray and William A. Ruh. Inside CORBA: Distributed Object Standards and Applications. The Addison-Wesley object technology series. Addison-Wes-ley, Reading, MA, USA, 1997. ISBN 0-201-89540-4 (paperback). xxiii + 376 pp. LCCN QA76.64 .R88 1997. US$47.

URL http://cseng.aw.com/
bookdetail.qry?ISBN=0-201-
89540-4&ptype=0.

Marais:1999:AWW

[MR99] Hannes Marais and Tom Rodeheffer. Automating the Web with Webl. Dr. Dobb’s Journal of Software Tools, 24(1):20– 23, 26–27, January 1999. CO-DEN DDJOEB. ISSN 1044-

                                        789X. URL http://www.ddj.
                                        com/ddj/ftp/1999/1999_01/
                                        webl.txt.

Messer:1998:BJU

[MRA98] Robert Messer, Steven R. Rakitin, and Charles Ashbacher. Bookshelf: Java, the universe, and everything: Java How to Program with an Introduction to J++; Software Quality Assurance and Measurement: A Worldwide Perspective; Object Solutions: Managing the Object-Oriented Process. IEEE Software, 15(3):115–118, May/ June 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

                                        http://dlib.computer.org/
                                        so/books/so1998/pdf/s3115.
                                        pdf.

Messmer:1996:SRJ

[MS96] Ellen Messmer and Carol Sliwa. Sun reaffirms Java control. Network World, 13(51):1, 13, December 16, 1996. ISSN 0887-7661 (print), 1944-7655 (electronic).

Mohseni:1997:JDG

[MS97] Piroz Mohseni and Tom Stewart. JavaBeans Developer’s Guide. MIS Press and M&T Books, P. O. Box 5277, Portland, OR 97208-5277, USA, Tel: (503) 282-5215 and M&T Publishing, Inc., 501 Galveston Drive, Redwood City, CA 94063, USA, September 1997. ISBN 1-55851-573-9 (softcover). xix + 311 pp. LCCN QA76.73.J38M64 1997. US$35.95. URL http:/

/www.idgbooks.com/cgi-bin/
db/fill_out_template.pl?idgbook:
1-55851-573-9:book-idg::uidg781;
 http://www.mandt.com/{javabeansdevgde}

.htm. Includes CD-ROM.

Mehta:1998:EES

[MS98a] Nimisha V. Mehta and Karen R.

Sollins. Expanding and extending the security features of Java. In USENIX [USE98f], page ?? ISBN 1-880446-92-8. LCCN QA76.9.A25

U83 1998. URL http://www.
usenix.org/publications/library/
proceedings/sec98/mehta.html.

Middendorf:1998:JPR

[MS98b] Stefan Middendorf and Reiner Singer. Java Programmierhandbuch und Referenz Ertweitungen. dpunkt-Verlag, Heidelberg, Germany, April 1998. ISBN 3-920993-87-X. LCCN ???? 78 DM.

Maremaa:1999:QJD

[MS99a] Tom Maremaa and William Stewart. QuickTime for Java: a Developer Reference. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 1999. ISBN 0-12-305440-0. xviii + 655 pp. LCCN QA76.73.J38 M348 1999; TR899 .M37 1999. US$49.95. URL http://

                                        www.mkp.com/books_catalog/
                                        catalog.asp?ISBN=0-12305-
                                        440-0.

Muller:1999:HHA

[MS99b] Gilles Muller and Ulrik Pagh Schultz. Harissa: a hybrid approach to Java execution. IEEE Software, 16(2):44–51, March/April 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

                                        URL http://computer.org/
                                        software/so1999/s2044abs.
                                        htm;   http://dlib.computer.
                                        org/so/books/so1999/pdf/s2044.
                                        pdf.

Middendorf:1996:JPR

[MSS96] Stefan Middendorf, Reiner Singer, and Stefan Strobel. Java: Programmierhandbuch und Referenz. dpunkt-Verlag, Heidelberg, Germany, June 1996. ISBN 3-920993-38-

1. 811 pp. LCCN ???? 88,00 DM. URL http:

                                        //www.dpunkt.de/produkte/
                                        java.html;   http://www.dpunkt.
                                        de/welcome.html;   http://
                                        www.dpunkt.de:80/produkte/
                                        Java.html;   http://www.emedia.
                                        de/bin/bookshop?show=4605&
                                        id=.

Martins:1998:JIW

[MSS98] P. Martins, L. M. Silva, and J. Silva. A Java interface for WMPI. Lecture Notes in Computer Science, 1497:121– ??, 1998. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic).

Magnaghi:1999:IPA

[MST99] A. Magnaghi, S. Sakai, and H. Tanaka. Inter-procedural analysis for parallelization of Java programs. Lecture Notes in Computer Science, 1557:594– 595, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

McElroy:1997:CJP

[MT97] David McElroy and Janet L.

Traub. Core JDBC Programming. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, November 1997. ISBN 0-13-764994-0. ???? pp. LCCN ???? US$50. URL http://

www.prenhall.com/ptrbooks/
ptr_0137649940.html.

Maruyama:1999:XJD

[MTU99] Hiroshi Maruyama, Kento Tamura, and Naohiko Uramoto. XML and Java: developing Web applications. Addison-Wesley, Reading, MA, USA, 1999. ISBN 0-201-48543-5, 0-201-61611-4 (CD-ROM). 368 pp. LCCN QA76.76.H94M28 1999. US$39.95.

Mukundan:1999:TCG

[Muk99] R. Mukundan. Teaching computer graphics using Java. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(4): 66–69, December 1999. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Muller:1996:JMP

[Mül96]  Claus M. Müller.  Java:

Mehr als eine Programmiersprache: Konzepte und Einsatzmöglichkeiten. dpunkt-Verlag, Heidelberg, Germany, 1996. ISBN 3-920993-44-6. 98 pp. LCCN ???? 49,00 DM.

URL http://www.dpunkt.de/
produkte/javamehr.html;
http://www.dpunkt.de:80/produkte/
Javamehr.html. Includes CD-

ROM.

Munck:1996:AJM

[Mun96] Bob Munck. Ada95 and Java: a major opportunity for the Ada community. ACM SIGADA Ada Letters, 16(1):18–20, January/February 1996. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). New mailing list web_ada@acm. org created for discussion of Ada-Java issues. Send subscription requests to mailserv@acm. org with no subject line and a body consisting of the lines

 subscribe
  web_ada help

.

Muramoto:1996:WAB

[Mur96a] E. Muramoto. WWW application building tool Walts. NEC Technical Journal = NEC giho, 49(7):152–156, July 1996. CO-DEN NECGEZ. ISSN 0285-4139.

Murphy:1996:SL

[Mur96b] P. Murphy. Standardizing on a language? Object Magazine, 6 (4):34–37, June 1996. CODEN OBMAFO. ISSN 1055-3614.

Murdoch:1999:ESCa

[Mur99a] Duncan J. Murdoch. On the edge: Statistics & computing: What’s Java? Chance, 12(1):44–47, 1999. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Murphy:1999:JT

[Mur99b] Timothy Murphy. Java and TEX. TUGboat, 20(3):307–312, September 1999. ISSN 0896-3207.

Muto:1997:J

[Mut97] Kenshi Muto. Java. Gijutsu Hyoron Co., Ltd., ??, Japan, 1997. ISBN 4-7741-0277-6. ???? pp. LCCN ????

1950 yen. URL http://www.
gihyo.co.jp/syoseki/promo/
java/java.html.

Maassen:1999:EIJ

[MvNV^{+}99] Jason Maassen, Rob van Nieuwpoort, Ronald Veldema, Henri E. Bal, and Aske Plaat. An efficient implementation of Java’s remote method invocation. ACM SIGPLAN Notices, 34(8):173–182, August 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

URL http://www.acm.org/
pubs/citations/proceedings/

ppopp/301104/p173-maassen/

.

Messmer:1995:NNMa

[MW95] Ellen Messmer and Peggy Watt. News: Networld+Interop: Microsoft sets sights on Internet. Network World, page ??, September 29, 1995. ISSN 0887-7661 (print), 1944-7655 (elec-

tronic). URL http://www.
iquest.net/index.html.

Munsil:1998:RSU

[MW98] Wes Munsil and Chia-Jiu Wang. Reducing stack usage in Java bytecode execution. ACM SIGARCH Computer Architecture News, 26(1):7–11, March 1998. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

McKenzie:1995:ERS

[MYD95] Bruce J. McKenzie, Corey Yeatman, and Lorraine De Vere. Error repair in shiftreduce parsers. ACM Transactions on Programming Languages and Systems, 17(4): 672–689, July 1995. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL http://www.acm.org/
pubs/toc/Abstracts/0164-0925/

210193.html. See failure report [BN99a].

Myers:1999:JPM

[Mye99] Andrew C. Myers. JFlow: practical mostly-static information flow control. In ACM [ACM99a], pages 228– 241. ISBN 1-58113-095-3.

LCCN ???? URL http://www.
acm.org:80/pubs/citations/
proceedings/plan/292540/p228-
myers/.

Mojsa:1997:WEC

[MZ97] Tomasz Mojsa and Krzysztof Zieliński. Web enabled, CORBA driven, distributed VideoTalk environment on the Java Platform. Computer Networks and ISDN Systems, 29 (8–13):865–873, September 30, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/
browse/browse.cgi?year=1997&
volume=29&issue=08-13&aid=
1681.

NewRiders:1996:JAR

[N^{+}96a] New Riders Publishing Staff et al. Java API Reference: Java, Applet and Java.awt API Packages. New Riders Publishing, Carmel, IN, USA, July 1996. ISBN 1-56205-598-4. 416 pp. LCCN 96-028598. US$29.99.

Newman:1996:UJC

[N^{+}96b] Alexander Newman et al. Using Java with CD-ROM: Special edition. Que Corporation, Indianapolis, IN, USA, special edition, April 1, 1996. ISBN 0-7897-0604-0. xliv + 869 pp. LCCN QA76.73.J38U85 1996. US$49.99. URL http://www.

mcp.com/21855341933474/cgi-
bin/bag?isbn=0-7897-0604-

                                        0&last=/bookstore. Includes

CD-ROM.

Newman:1997:JRA

[N^{+}97a] Alexander Newman et al. Java: Referenz und Anwendungen. (Special Edition). Markt & Technik, Haar bei München, Germany, 1997. ISBN 3-8272-1001-1. 1091 pp. LCCN ????

                                        79,95 DM. URL http://www.
                                        mut.ch/html/buchinfo.htx?
                                        &x_pp=dJgxRqA9wlAYNijX&x_
                                        az=10&x_pk=VpLvjnsx7b5V&x_
                                        pa=B&x_pr=oS&x_bestell=21001&
                                        x_stichwort=Newman&x_sprache=
                                        A&x_start=1&x_gefunden=2&
                                        x_query=&x_lastpage=buchfund&
                                        x_soft=&x_suchewo=A;   http:
                                        //www.mut.ch/profit.htx?&
                                        x_pk=kd84mf7ms9dm.

Nishimura:1997:IJ

[N^{+}97b] Toshihiro Nishimura et al. Introduction to Java. Shoeisha Co., Ltd., ??, Japan, 1997. ISBN ???? ???? pp. LCCN ???? URL ????.

Nagarathnam:19xx:JNAb

[N^{+}xxa] Nataraj Nagarathnam et al. Java Networking & AWT API SuperBible (Korean Translation). DaeRim Publishing, Seoul, Korea, 19xx. ISBN 89-7280-332-4 (??invalid ISBN??). ???? pp. LCCN ???? ????

                                        URL http://www.drbook.co.
                                        kr/daerim/jvnaapi.htm. Ko-

rean translation.

Newman:19xx:GGJ

[N^{+}xxb] Alex Newman et al. Grande Guida a Java. Jackson Libri, via XXV Aprile, 39 - 20091 Bresso (MI), Italy, 19xx. ISBN ???? ???? pp. LCCN ???? 95000 Lire. URL http:/

/www.jacksonlibri.it/. Ital-

ian translation of [N^{+}96b].

Nicolas:19xx:JCS

[N^{+}xxc] C. Nicolas et al. Java Client Serveur. Eyrolles, Paris, France, 19xx. ISBN 2-212-08938-X (??Invalid checksum??). ????

pp.  LCCN ????  285

FF. URL http://www.estp.
fr/eyrolles;   http://www.
estp.fr/eyrolles/ouvrages/
book159.html.

Nishimura:19xx:IJ

[N^{+}xxd] Toshihiro Nishimura et al. Introduction to Java. Shoeisha Co., Ltd., ????, 19xx. ISBN 4-88135-351-9. ???? pp. LCCN ???? 2800 yen. URL http://

www.coara.or.jp/coara/eto/
jbook/.

Naumovich:1999:EAC

[NAC99] Gleb Naumovich, George S.

Avrunin, and Lori A. Clarke. An efficient algorithm for computing MHP information for concurrent Java programs. ACM SIGSOFT Software Engineering Notes, 24(6):338–354, November 1999. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Nagaratnam:1996:JAA

[Nag96] Nataraj Nagaratnam. Java applets and the AWT. In USENIX [USE96c], page ?? ISBN 1-880446-77-4. LCCN QA76.64 .U85 1996. URL http://www.

usenix.org/publications/library/
proceedings/coots96/.

Nance:1995:WOG

[Nan95] Barry Nance. A whole other galaxy: quickly develop objectoriented applications without the shortcomings of high-level abstraction. BYTE Magazine, 20(7):155–??, July 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nance:1996:KNS

[Nan96] B. Nance. Keep networks safe from viruses. BYTE Magazine, 21(11):167, 169, 171, 173, 175, November 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Naps:1997:AVW

[Nap97] Thomas L. Naps. Algorithm visualization on the World Wide Web — the difference Java makes! SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(3): 59–61, September 1997. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Naps:1998:JVC

[Nap98] Thomas L. Naps. A Java visualiser class: incorporating algorithm visualisations into students’ programs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):181–184, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Nash:1995:NIU

[Nas95a] Kim S. Nash. News: Iband unveils backstage at Web Innovation Show. Computerworld, page ??, December 08, 1995. CODEN CMPWAB. ISSN 0010-4841. URL http://www.

computerworld.com/.

Nash:1995:NNE

[Nas95b] Kim S. Nash. News: Netscape expands beyond roots. Computerworld, page ??, September 29, 1995. CODEN CMPWAB. ISSN 0010-4841.

Nash:1995:NNI

[Nas95c] Kim S. Nash. News: Netscape to introduce Web developer tools. Computerworld, page ??, September 15, 1995. CO-DEN CMPWAB. ISSN 0010-

4841.  URL http://www.
computerworld.com/.

NaturalIntelligence:1995:RA

[Nat95] Natural Intelligence, Inc. ‘ROASTER’ announcement. press release, October 18, 1995.

URL http://www.natural.
com/pages/products/roaster/
roasterpr.html.

Natoli:1996:SPB

[Nat96] Anthony J. Natoli. Software patents: beating a dead horse and other important concerns.

In IEEE [IEE96d], pages 269– 276. ISBN 0-7803-3272-5 (casebound), 0-7803-3271-7, 0-7803-3273-3. LCCN TK7801.E45 1996. IEEE catalog number 96CB35926.

NIST:1998:FAE

[Nat98] National Institute of Standards and Technology, editor. The First Advanced Encryption Standard Candidate Conference, August 20– 22, 1998, DoubleTree Hotel, Ventura, California. National Institute for Standards and Technology, Gaithersburg, MD, USA, 1998. ISBN ???? LCCN ???? URL http:/

/csrc.nist.gov/encryption/
aes/round1/conf1/aes1conf.
htm;   http://www.nist.gov/

aes. See [RD99] for a conference overview. No formal proceedings were published, but the conference Web site contains pointers to slides and/or technical papers for most of the fifteen “complete and proper” candidates.

NIST:1999:SAC

[Nat99] National Institute of Standards and Technology, editor. Second AES Candidate Conference Proceedings, March 22– 23, 1999, Rome, Italy. National Institute for Standards and Technology, Gaithersburg, MD, USA, March 1999. ISBN ???? LCCN ???? URL http:/

/csrc.nist.gov/encryption/
aes/round1/conf2/aes2conf.
htm;   http://csrc.nist.gov/

encryption/aes/round1/conf2/
agenda-final.pdf;   http:/
/www.nist.gov/aes. No for-

mal proceedings were published, but the conference Web site contains pointers to slides and/or technical papers for most of the fifteen “complete and proper” candidates.

Naughton:1996:JH

[Nau96a]  Patrick Naughton.  The

Java Handbook. Osborne/ McGraw-Hill, Berkeley, CA, USA, 1996. ISBN 0-07-882199-1. xxiv + 424 pp. LCCN QA76.73.J38 N38 1996. US$27.95. URL http://db.

www.idgbooks.com/database/
book/isbn/generic-book.tmpl?
query=0-7645-3003-8;   http:
//www.osborne.com/int/{javah}
.htm.

Naughton:1996:JHJ

[Nau96b] Patrick Naughton. The Java Handbook [Java chuan wei tao yin shou tse]. McGraw-Hill Inc. [Hudson Technology], Taipei, Taiwan, 1996. ISBN 957-8496-07-9 (??invalid ISBN??). 20 + 594 pp. LCCN

????  500 NT$.  URL

http://www.hudson.com.tw;
 http://www.hudson.com.tw/
cbookn.html#N20C;   http:
//www.starwave.com/people/

naughton/. Chinese translation of [Nau96a] by Chen Phi Rhong.

Naughton:1997:MJ

[Nau97] Patrick Naughton. Manual de Java. McGraw-Hill Spain, ????, 1997. ISBN 84-481-0693-8. ???? pp. LCCN ???? 4000 Ptas.

                                        URL http://mexplaza.udg.
                                        mx/McGraw/;   http://www.
                                        mcgraw-hill.es/McGrawHill/
                                        catalogo.nsf/b013777771f7e01bc125647a00607b0e/
                                        19a221cd0649f88cc12564d600405519?
                                        OpenDocument.

Naughton:1998:BJA

[Nau98]  Patrick Naughton.  Basic

to Java: Assembling a career. IEEE Software, 15 (1):14–17, January/February 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

                                        http://dlib.computer.org/
                                        so/books/so1998/pdf/s1014.
                                        pdf.

Naughton:19xx:DJ

[Nauxxa] Patrick Naughton. Dominando o Java. Makron Books do Brasil Editora Ltda., Rua Tabapuã 1348, São Paulo SP - 04533-004 - Brasil, 19xx. ISBN 85-346-0566-1 (??invalid ISBN??). LCCN ???? ???? URL http:/

                                        /safe.tesla.com.br/makron/
                                        options/livro.asp?id=262;
                                         http://www.makron.com/; http:
                                        //www.makron.com/G57.htm.

Naughton:19xx:JM

[Nauxxb]  Patrick Naughton.  Java

— Il Manuale. McGraw-Hill, New York, NY, USA, 19xx. ISBN ???? ???? pp. LCCN ???? 52000

                                        L.  URL http://pallade.
                                        iol.it/cgi-bin/mercator2/
                                        mallurl/CESSECCOCEDRBILEDOMAa000eC/
                                        mcghill/001074.html.

Navas:1997:LEJ

[Nav97] John Navas. Letter to the editor: The Java race. BYTE Magazine, 22(2):20, February 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Nieuwenhuis:1996:IBC

[NB96] L. Nieuwenhuis and F. Bergsma.

Internet as a basis for commerce. Informatie, 38(????):26– 30, November 1996. CODEN INFTCR. ISSN 0019-9907.

Neophytou:1998:NDJ

[NE98a] N. Neophytou and P. Evripidou. Net-dbx: a Java powered tool for interactive debugging of MPI programs across the Internet. Lecture Notes in Computer Science, 1470:181– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Neophytou:1998:NJP

[NE98b] N. Neophytou and P. Evripidou. Net-dbx: a Java powered tool for interactive debugging of MPI programs across the Internet. Lecture Notes in Computer Science, 1470:181– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nelson:1995:HJW

[Nel95a] Brian Nelson. Hot Java wakes up Internet. Cable News Network Computer Connection, December 9, 1995. URL http:

                                        //www.cnn.com/TRANSCRIPTS/
                                        conn/conn120995+1.html.

Nelson:1995:WWW

[Nel95b] Brian Nelson. World Wide Web gets jolt from Java and Shockwave. Cable News Network, December 8, 1995.

                                        URL http://www.cnn.com/
                                        TECH/9512/java/index.html.

Nelson:1998:JFC

[Nel98] Matthew Nelson. Java Foundation Classes (JFC). McGraw-Hill, New York, NY, USA, March 1998. ISBN 0-07-913758-X. LCCN ???? US$45.

                                        URL http://mcgraw-hill.
                                        inforonics.com/cgi/getarec?
                                        mgh50918%comp.

Neou:1996:HCJb

[Neo96a] Vivian Neou. HTML 3.0 CD with JavaScript. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, September 1996. ISBN 0-13-243353-2. ???? pp. LCCN ???? US$44.95.

Neou:1996:HCJc

[Neo96b] Vivian Neou. HTML 3.0 CD with JavaScript. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, September 1996. ISBN 0-13-243353-2. ???? pp. LCCN ???? US$44.95.

Neou:1997:HCJa

[Neo97] Vivian Neou. HTML 3.0 CD JavaScript for MacIntosh: CD-ROM With Guide. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, August 1997.

ISBN 0-13-243353-2. ???? pp. LCCN ???? US$44.95.

Neou:1998:HCJ

[Neo98] Vivian Neou. HTML 4.0 CD with JavaScript. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1998. ISBN 0-13-095783-6. 496 pp. LCCN ???? US$49.95. URL http:

//www.phptr.com/ptrbooks/
ptr_0130957836.html.  In-

cludes CD-ROM.

Neou:1999:HCJ

[Neo99] Vivian Neou. HTML 4.0 CD with JavaScript for Windows 95, 98 and NT. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1999. ISBN 0-13-095783-6. xi + 619 pp. LCCN QA76.76.H94N435 1999.

Nerny:1997:JDR

[Ner97] Chris Nerny. Java developers rise up: New lobbying group to put pressure on industry giants. Network World, 14(33):1, 80, August 18, 1997. ISSN 0887-7661 (print), 1944-7655 (electronic).

Netscape:1995:NSA

[Net95a] Netscape Communications Corporation. Netscape and Sun announce Javascript, the open, cross-platform object scripting language for enterprise networks and the Internet. press release, December 4, 1995. URL

http://home.netscape.com/
newsref/pr/newsrelease67.
html.

Netscape:1995:NLS

[Net95b] Netscape Communications Corporation. Netscape to license Sun’s Java programming language. press release, May 23, 1995. URL http://

                                        home.netscape.com/newsref/
                                        pr/newsrelease25.html.

NetworkWorld:1995:NIN

[Net95c] Network World staff. News: InterCon, Navisoft boost profiles on Web browser scene. Network World, page ??, September 19, 1995. ISSN 0887-7661 (print), 1944-7655 (elec-

                                        tronic). URL http://www.
                                        iquest.net/index.html.

Neumann:1997:SRC

[Neu97] P. Neumann. Security risks in computer-communication systems. SIG security, audit & control review, 15(1):15–18, January 1997. CODEN SSARE7. ISSN 0277-920X (print), 1558-0261 (electronic).

Nevison:1999:SSC

[Nev99] Christopher H. Nevison. Seminar: safe concurrent programming in Java with CSP. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1):367, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Newman:1996:EJ

[New96a] Alex Newman. Had enough Java? SunExpert Magazine, 7(4):49–51, April 1996. ISSN 1053-9239.

Newman:1996:HHH

[New96b] Alex Newman. Hurry, hurry, hurry, …. SunExpert Magazine, 7(6):42, 44, June 1996. ISSN 1053-9239. Discusses Military Java and JUGMAP (Java User Group for Military Appli-

cations). See http://www.sug.
org/jugmap.html.

Newman:1996:MJ

[New96c] Alex Newman. Le Macmillan Java. Simon & Schuster Macmillan (France), Paris, France, 1996. ISBN 2-7440-0147-3. xxv + 771 pp. LCCN ???? 249 FF. URL http://

www.ssm.fr/catalog/fiches/

0147.htm. French translation of [N^{+}96b]. Includes CD-ROM.

Newman:1996:UJS

[New96d] Alex Newman. Using Java, Special Ed. MacMillan Publishing Company, New York, NY, USA, March 1996. ISBN 0-614-14450-7. ???? pp. LCCN ???? US$49.99.

Newman:1997:UJ

[New97] Alexander Newman. Usando Java. Editore Campus Ltda., Rua Sete de Setembro, 111/16, Rio de Janeiro - RJ - 20050-002, Brazil, 1997. ISBN ???? 904 pp. LCCN ???? R$ 89,00. URL http:

//www.campus.com.br/about-
us.htm;   http://www.campus.
com.br/cgi-bin/catalog.cgi?
livro=20098.

Newman:1998:JSE

[New98] Alex Newman. Java 1.2 Special Edition. Markt & Technik, Haar bei München, Germany, March 1998. ISBN 3-8272-1032-1 (??invalid ISBN??). LCCN ???? 90 DM. URL

                                        http://www.mut.ch/html/v2/
                                        Produkt-Information.htx?&
                                        x_pp=204%2e94%2e76%2e206VanW4xQbKi&
                                        x_pk=Direkt-Fr&x_stichwort=
                                        java&x_suchsprache=A&x_search=
                                        T&x_artikelnr=21032.

Newman:19xx:SEJ

[Newxxa] Alex Newman. Special Edition Java. Markt & Technik, Haar bei München, Germany, 19xx. ISBN 3-8272-1001-1 (??invalid ISBN??). LCCN ????

                                        80 DM. URL http://www.
                                        mut.ch/html/buchinfo.htx?
                                        &x_pp=dJgxRqA9wlAYNijX&x_
                                        az=10&x_pk=VpLvjnsx7b5V&x_
                                        pa=B&x_pr=oS&x_bestell=21001&
                                        x_stichwort=Newman&x_sprache=
                                        A&x_start=1&x_gefunden=2&
                                        x_query=&x_lastpage=buchfund&
                                        x_soft=&x_suchewo=A.

Newman:19xx:UJ

[Newxxb] Alex Newman. Usando Java. Editore Campus Ltda., Rua Sete de Setembro, 111/16, Rio de Janeiro - RJ - 20050-002, Brazil, 19xx. ISBN ???? LCCN ???? R$89.

Norman:1996:IRJ

[NF96] A. Norman and J. Fitch. Interfacing REDUCE to Java. Lecture Notes in Computer Science, 1128:271–276, 1996. CO-DEN LNCSD9. ISBN 3-540-61697-7. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nickerson:1996:BRJc

[NFP96] Doug Nickerson, David H.

Friedel Jr., and Anthony Potts. Book review: Java Programming Language Handbook. Dr. Dobb’s Journal of Software Tools, 21(12):134–??, December 1996. CODEN DDJOEB. ISSN 1044-789X.

Nghiem:19xx:LJV

[Nghxx] Nghiem. Language Java Version JDK 1.1. Infoprax, ????, 19xx. ISBN 2-901683-15-0. ???? pp. LCCN ???? 195 FF.

Nadeau:1997:WAJ

[NGM^{+}97] Dave Nadeau, Brad Grantham, Colin McCartney, Mitra, and Henry Sowizral. What 3D API for Java should I use and why? (panel). Computer Graphics, 31(Annual Conference Series): 452–453, August 1997. CODEN CGRADI, CPGPBZ. ISSN 0097-8930 (print), 1558-4569 (electronic). URL http://www.

acm.org:80/pubs/citations/
proceedings/graph/258734/
p452-nadeau/.

Nemirovski:1998:AJ

[NH98] German Nemirovski and Michael Heise. Algorithmen in Java. Addison-Wesley, Reading, MA, USA, 1998. ISBN 3-8273-1234-5 (??invalid ISBN??). ???? pp. LCCN ???? 40 DM. URL

http://www.addison-wesley.

de/. Includes CD-ROM.

Nemirovski:19xx:AJ

[NHxx] German Nemirovski and Michael Heise. Algorithmen in Java. Addison-Wesley Verlag, Bonn, Germany, 19xx. ISBN 3-8273-1234-5 (??invalid ISBN??). LCCN ???? 40 DM. URL

                                        http://www.addison-wesley.
                                        de/katalog/item.ppml?id=00201.

Nickerson:1996:BRD

[Nic96a] Doug Nickerson. Book review: David Flanagan’s Java in a Nutshell. Dr. Dobb’s Journal of Software Tools, 21(12):134– ??, December 1996. CODEN DDJOEB. ISSN 1044-789X.

Nickerson:1996:BRP

[Nic96b] Doug Nickerson. Book review: Peter van der Linden, Just Java. Dr. Dobb’s Journal of Software Tools, 21(12):134–??, December 1996. CODEN DDJOEB. ISSN 1044-789X.

Nickerson:1996:HJ

[Nic96c] Douglas A. Nickerson. Hooked on Java. C/C++ Users Journal, 14(11):79–80, November 1996. CODEN CCUJEX. ISSN 1075-2838. Extensive review of [vSS96b].

Nickerson:1996:RHJ

[Nic96d] Douglas A. Nickerson. Review: Hooked on Java. C/C++ Users Journal, 14(11):79–80, November 1996. CODEN CCUJEX. ISSN 1075-2838. Extensive review of [vSS96c].

Nickerson:1997:ONJ

[Nic97]  Doug Nickerson.  Official

Netscape JavaBeans Developer’s Guide. Ventana Press, Chapel Hill, NC, USA, November 1997. ISBN 1-56604-825-7. ???? pp. LCCN QA76.73.J38N53 1997. US$40.

Nichol:1999:JSB

[Nic99] Sandy Nichol. Java security bug found, patched. Computers & Security, 18(5):434, ???? 1999. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.
sciencedirect.com/science/
article/pii/S0167404899800960.

Nickerson:19xx:CPJ

[Nicxx] Doug Nickerson. Component Programming With JavaBeans. Ventana Press, Chapel Hill, NC, USA, 19xx. ISBN 1-56604-753-6. ???? pp. LCCN ???? US$40.

Niemeyer:1995:EJA

[Nie95] Pat Niemeyer. Exploring Java:

a tutorial introduction. The X Resource, 16(1):77–102, December 1995. CODEN XRE-SEA. ISBN 0-937175-79-X. ISSN 1058-5591.

Niemeyer:1997:CJN

[Nie97] Pat Niemeyer. Core Java Networking. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, November 1997. ISBN 0-13-761537-X. ???? pp. LCCN ???? US$50.

Nilsen:1996:JR

[Nil96] K. Nilsen. Java for real-time. Real-Time Systems, 11(2):197– 205, September 1996. CODEN RESYE9. ISSN 0922-6443.

Nilsson:1997:TJ

[Nil97] Stefan Nilsson. Treaps in Java. Dr. Dobb’s Journal of Software Tools, 22(7):40, 42–44, 84, 86, July 1997. CODEN DDJOEB. ISSN 1044-789X.

Nilsen:1998:ARC

[Nil98a] Kelvin Nilsen. Adding realtime capabilities to Java. Communications of the ACM, 41 (6):49–56, June 1998. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL http://www.
                                        acm.org:80/pubs/citations/
                                        journals/cacm/1998-41-6/p49-
                                        nilsen/.

Nilsen:1998:PSF

[Nil98b] Kelvin Nilsen. picoPERC: a small-footprint dialect of Java. Dr. Dobb’s Journal of Software Tools, 23(3):50, 52–54, March 1998. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

                                        ddj.com/ftp/1998/1998_03/
                                        javabj.zip.

Nixon:1998:OOP

[Nix98] Nixon. Object Oriented Programming With Java. Ventana Press, Chapel Hill, NC, USA, June 1998. ISBN 1-85032-336-4. LCCN ???? US$40.

Nash:1995:NSN

[NJ95] Kim S. Nash and Stuart J.

Johnston. News: Sun, Netscape gun for Microsoft with Java script. Computerworld, page ??, December 01, 1995. CODEN CMPWAB. ISSN 0010-4841.

Nygren:1996:ASM

[NJC96] K. Nygren, I.-M. Jonsson, and O. Carlvik. An agent system for media on demand services. In Anonymous [Ano96-154], pages 437–454. ISBN ???? LCCN ????

Nikander:1998:JBC

[NK98] Pekka Nikander and Arto Karila. A Java beans component architecture for cryptographic protocols. In USENIX [USE98f], page ?? ISBN 1-880446-92-8. LCCN QA76.9.A25

U83 1998. URL http://www.
usenix.org/publications/library/
proceedings/sec98/nikander.
html.

Niemi:19xx:OIJ

[NKxx] Juha Niemi and Petri Kiuttu. Opeta itsellesi JavaScript. Suomen Atk-kustannus Oy, PL 68, 02601 Espoo, Finland, 19xx. ISBN 951-762-447-6. 290 pp. LCCN ???? 265 Markka. URL http:/

/www.satku.fi/;   http://
www.satku.fi/4javascr.htm.

Includes CD-ROM.

Nagarathnam:1996:JNA

[NMS96] Nataraj Nagarathnam, Brian Maso, and Arvind Srinivasan.

Java networking and AWT API superbible: the comprehensive reference to the Java programming language: everything you ever needed to know about Java applets and its networking and windowing APIs. Waite Group Press, Corte Madera, CA, USA, July 1, 1996. ISBN 1-57169-031-X. xvii + 930 pp. LCCN QA76.73.J38N34 1996. 119,00 DM; US$59.95. URL http:

//www.mcp.com/cgi-bin/bag?
isbn=1-57169-031-X&last=/
bookstore;   http://www.mut.
ch/profit.htx?&x_pk=kd84mf7ms9dm;
 http://www.waite.com/waite/

. Includes CD-ROM.

Nelson:1997:JZF

[NN97] Mark R. Nelson and Mark Nelson. Java and the zip file format: Using Java to compress and extract files. Dr. Dobb’s Journal of Software Tools, 22 (12):50, 54–54, 102, December 1997. CODEN DDJOEB. ISSN 1044-789X.

Nabil:1997:PSR

[NNS97] Mohammad Nabil, Anne H. H. Ngu, and John Shepherd. Picture similarity retrieval using the 2D projection interval representation. IEEE Transactions on Knowledge and Data Engineering, 8(4):533–539, August 1997. CODEN ITKEEH. ISSN 1041-4347. Uses Java to construct a prototype image database system.

Nolan:1998:DJ

[Nol98] Godfrey Nolan. Decompiling Java. McGraw-Hill, New York, NY, USA, June 1998. ISBN 0-07-913767-9. LCCN ????

US$45. URL http://www.pbg.
mcgraw-hill.com/betabooks/
nolan/index.html.

Norton:1996:PNC

[Nor96] Peter Norton. Peter Norton’s Complete Guide To Jakarta, Premiere Edition. Sams.net, Indianapolis, IN 46268, USA, 1996. ISBN 1-57521-176-9. ???? pp. LCCN ???? US$50.

Notess:1996:JCN

[Not96] Greg R. Notess. Java creates new potential for Internet information. Online, 20(2):82–84, 86, March–April 1996. CODEN ONLIDN. ISSN 0146-5422.

Novosel:1997:JCS

[Nov97]  Joe Novosel.  Java and

client-server. Linux Journal, 33:??, January 1997. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http://

www.linuxjournal.com/0155.
html.

Novak:1998:BRD

[Nov98] Gregor M. Novak. Book review:

David Flanagan, Java-Script: The Definitive Guide, Second Edition. Computers in Physics, 12(1):41–??, January 1998. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (elec-

tronic). URL https://aip.
scitation.org/doi/10.1063/
1.168647.

Niemeyer:1996:EJ

[NP96a] Pat Niemeyer and Josh Peck. Exploring Java. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, June 1, 1996. ISBN 1-56592-184-4. xv + 407 pp. LCCN QA76.73.J38N54 1996. US$24.95. URL http://www.

                                        ora.com/www/item/expjava.
                                        html.

Niemeyer:1996:JEI

[NP96b] Patrick Niemeyer and Joshua Peck. Java — Expedition ins Programmierreich. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, December 1996. ISBN 3-930673-52-5 (??invalid ISBN??). ???? pp. LCCN ???? 69 DM.

                                        URL http://www.ora.de/. In-

cludes CD-ROM.

Niemeyer:1996:JPP

[NP96c] Patrick Niemeyer and Joshua Peck. Java par La Pratique. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, October 1996. ISBN 2-84177-022-2. xxi + 413 pp. LCCN ???? 220 FF.

                                        URL http://www.editions-
                                        oreilly.fr/international/
                                        france/prog/expjava.html;
                                         http://www.ora.de/. French

translation of [NP96a]. Includes CD-ROM.

Niemeyer:1997:EJb

[NP97a] Pat Niemeyer and Josh Peck. Exploring Java. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1997. ISBN 4-900900-17-

6. ???? pp. LCCN ???? 4300

yen. URL http://www.ora.
com/international/catalog/
japanese/inet/exjava.html;
http://www.ora.com/international/
japan/;   http://www.oreilly.
co.jp/BOOK/exjava.htm. Japanese

translation.

Niemeyer:1997:EJ

[NP97b] Patrick Niemeyer and Joshua Peck. Exploring Java. The Java series. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 1997. ISBN 1-56592-271-9. xviii + 594 pp. LCCN QA76.73.J38 N54 1997. US$32.95. URL http://www.

ora.com/catalog/expjava2/;
http://www.oreilly.com/catalog/
9781565922716.

Niemeyer:1997:EJa

[NP97c] Patrick Niemeyer and Joshua Peck. Exploring Java. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 1997. ISBN 1-56592-271-9. xviii + 594 pp. LCCN QA76.73.J38N54 1997. US$32.95. URL http://www.

ora.com/catalog/expjava2/.

Niemeyer:19xx:EJd

[NPxx] Pat Niemeyer and Josh Peck.

Exploring Java. Image Art Publishing, ????, 19xx. LCCN ???? ???? Chinese translation.

Neou:1996:HCJa

[NR96] Vivian Neou and Mimi Recker. HTML 3.0 CD With JavaScript for the Mac and Power Mac. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, June 1, 1996. ISBN 0-13-439399-6. xix + 441 pp. LCCN QA76.76.H94 N43 1996. US$44.95.

Neou:1997:HCJb

[NR97] Vivian Neou and Mimi Recker. HTML 3.2 CD With JavaScript for Windows 95. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1997. ISBN 0-13-270125-1. xxi + 439 pp. LCCN QA76.76.H94N434 1997. US$44.95.

Neal:1999:LSL

[NRG^{+}99] John Neal, Thiadmer Riemersma, Jeff Genender, Torpum Jannak, Richard A. Clarke, William C. Brown, Jeffrey Simmers, and Kemal Gencay. Letters: The Small language; Java 2 graphics; cross-platform independence; version control; median filters; Open Source in Turkey. Dr. Dobb’s Journal of Software Tools, 24(12):10, 12, December 1999. CODEN DDJOEB. ISSN 1044-789X.

Norton:1996:PNG

[NS96] Peter Norton and William R. Stanek. Peter Norton’s Guide to Java Programming. Sams.net, Indianapolis, IN 46268, USA, 1996. ISBN 1-57521-088-6. xxiv + 875 pp. LCCN QA76.73.J38N67 1996. US$39.99, CDN$56.95, UK£37.50. URL

http://merchant.superlibrary.
com:8000/catalog/hg/PRODUCT/
PAGE/15752/bud/1575210886.
html;   http://www.mcp.com/
samsnet/.

Naughton:1997:JGC

[NS97a] Patrick Naughton and Herbert Schildt. Java Guida Complete. McGraw-Hill, New York, NY, USA, March 1997. ISBN 88-8386-041-6. LCCN ???? 79000 L.

Naughton:1997:JCRa

[NS97b] Patrick Naughton and Herbert Schildt. Java: The Complete Reference. Osborne/McGraw-Hill, Berkeley, CA, USA, November 1997. ISBN 0-07-882231-9. xxiv + 886 pp. LCCN QA76.73.J38N378 1997. US$39.95. URL http:/

/www.osborne.com/;   http:
//www.starwave.com/people/
naughton/jtcr/.

Naughton:1997:JMR

[NS97c] Patrick Naughton and Herbert Shildt. Java: Manual de Referencia. McGraw-Hill Spain, ????, 1997. ISBN 84-481-1131-

1. ???? pp. LCCN ???? 6900 Ptas. URL http://www.

mcgraw-hill.es/McGrawHill/
catalogo.nsf/b013777771f7e01bc125647a00607b0e/
82287940fc60553ec125650c00352778?
OpenDocument.

Naughton:1998:JCR

[NS98a] Patrick Naughton and Herbert Schildt. Java 1.1: The Complete Reference. Osborne/McGraw-Hill, Berkeley, CA, USA, second edition, 1998. ISBN 0-07-882436-2. xxiv + 1028 pp. LCCN QA76.73.J38N378 1998. US$39.99, CDN$57.95.

                                        URL http://www.osborne.
                                        com/program/java11cr.htm.

Negrino:1998:JWW

[NS98b] Tom Negrino and Dori Smith. JavaScript for the World Wide Web. Peachpit Press, Inc., 1085 Keith Avenue, Berkeley, CA 94708, USA, second edition, 1998. ISBN 0-201-69648-7. xii + 195 pp. LCCN QA76.73.J39 N44 1998. US$17.95.

Neilson:1996:EIW

[NTS^{+}96] I. Neilson, R. Thomas, C. Smeaton, A. Slater, and G. Chand. Education 2000: implications of W3 technology. Computers and Education, 26(1-3):113–122, April 1996. CODEN COMEDR. ISSN 0360-1315 (print), 1873-782X (electronic).

NWGJGF:1998:PDI

[Num98] Numerics Working Group, Java Grande Forum. Public draft: Improving Java for numerical computation. World-Wide Web document., October 30,

                                        1998. URL http://math.nist.
                                        gov/javanumerics/reports/
                                        jgfnwg-01.html.

Nipkow:1998:JTS

[NvO98] Tobias Nipkow and David von Oheimb. Java_{light} is type-safe-definitely. In ACM [ACM98b], pages 161–170. ISBN 0-89791-979-3. LCCN QA76.7 .A15 1998. URL http://www.

acm.org:80/pubs/citations/
proceedings/plan/268946/p161-

nipkow/. ACM order number: 549981.

Oaks:1998:JS

[Oak98] Scott Oaks. Java Security.

O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, May 1998. ISBN 1-56592-403-7. xv + 454 pp. LCCN A76.73.J38 O247 1998. US$32.95. URL http:

//www.oreilly.com/catalog/
javasec.

OBrien:1996:MPL

[O’B96a] Michael O’Brien. Mr. P and the language barrier. SunExpert Magazine, 7(2):20, 22, 24– 25, February 1996. ISSN 1053-9239.

OBrien:1996:MPO

[O’B96b] Michael O’Brien. Mr. Protocol opens some free samples. Sun-Expert Magazine, 7(8):14, 16, 18, 20, August 1996. ISSN 1053-9239. Extensive commentary on Java.

Oliver:1996:NU

[OBWR96] Dick Oliver, Selby Bateman, David Wade, and Neil Randall. Netscape Unleashed. Howard W. Sams, Indianapolis, IN 46268, USA, February 1, 1996. ISBN 1-57521-007-X. xxxv + 942 pp. LCCN TK5105.883.N48N48 1996. US$49.99. URL http://

merchant.superlibrary.com:

                                        8000/catalog/hg/PRODUCT/PAGE/
                                        15752/bud/157521007X.html.

OConnell:1995:JPT

[O’C95a] Michael O’Connell. Java proliferates through licensing deals. SunWorld Online, December

                                        1995. URL http://www.sun.
                                        com/sunworldonline/swol-12-
                                        1995/swol-12-javadeal.html.

OConnell:1995:JIS

[O’C95b] Michael O’Connell. Java: The inside story. Sunworld Online, 07:??, July 1995. ISSN 1091-

                                        8914. URL http://www.sun.
                                        com/sunworldonline/swol-07-
                                        1995/swol-07-java.html.

OConnell:1995:JWD

[O’C95c] Michael O’Connell. Java wheeling and dealing. Sunworld Online, 12:??, 1995. ISSN 1091-

                                        8914. URL http://www.sun.
                                        com/sunworldonline/swol-12-
                                        1995/swol-12-javadeal.html.

OConnell:1995:NAJ

[O’C95d] Michael O’Connell. Netscape adds Java and more. Sunworld Online, 10:??, 1995. ISSN 1091-

                                        8914. URL http://www.sun.
                                        com/sunworldonline/swol-10-
                                        1995/swol-10-netscape.html.

OConnell:1995:NNG

[O’C95e] Michael O’Connell. Netscape’s next generation. SunWorld Online, October 1995. URL http:

                                        //www.sun.com/sunworldonline/
                                        swol-10-1995/swol-10-netscape.
                                        html.

OConnor:1997:WSI

[O’C97] Hugh O’Connor. Web sites for information on intranets, HMTL/SGML/JAVA/VRML coding and web page/sites design. American Association of Retired Persons, Washington, DC, USA, version 2 edition, February 20, 1997. 10 pp. LCCN ????

OConner:1998:CLS

[O’C98] John O’Conner. Comparing locale support in the Java and Win32 platforms. In Unicode Consortium [Uni98b], page ?? ISBN ???? LCCN

????  URL http://www.
unicode.org/unicode/iuc13/
program.html.

OCallahan:1999:SCT

[O’C99] Robert O’Callahan. A simple, comprehensive type system for Java bytecode subroutines. In ACM [ACM99a], pages 70–78. ISBN 1-58113-095-3.

LCCN ???? URL http://www.
acm.org:80/pubs/citations/
proceedings/plan/292540/p70-
o_callahn/.

Ogawa:1999:DPJ

[Oga99] Arthur Ogawa. Database publishing with Java and TEX. TUGboat, 20(3):231, September 1999. ISSN 0896-3207.

OBrien:1997:NLJ

[OH97a] Timothy O’Brien and Douglas Heise. Netscape looks to Java and CORBA to drive Web object use. Object Magazine, 6 (12):66–67, February 1997. CO-DEN OBMAFO. ISSN 1055-3614.

OBrien:1997:SMR

[OH97b] Timothy O’Brien and Douglas Heise. Sun/Microsoft rivalry: Will Java fragment like UNIX? Object Magazine, 6(12):14–16, February 1997. CODEN OB-MAFO. ISSN 1055-3614.

Orfali:1997:CSP

[OH97c] Robert Orfali and Dan Harkey. Client/Server Programming with Java and CORBA. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, January 1997. ISBN 0-471-16351-1. 657 pp. LCCN QA76.9.C55 O73 1997. US$44.95. URL http:

//www.wiley.com/compbooks/
catalog/16351-1.htm.

Orfali:1998:CSP

[OH98] Robert Orfali and Dan Harkey. Client/Server Programming with Java and CORBA. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, second edition, January 1998. ISBN 0-471-24578-X. 832 pp. LCCN QA76.9.C55 O73 1997. US$49.99. URL http:/

/www.wiley.com/compbooks/
catalog/24578-X.htm.

Orfali:1999:CSP

[OH99] Robert Orfali and Dan Harkey. Client/Server Programming with JavaBeans. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, January 1999. ISBN 0-471-18931-6. ???? pp. LCCN ???? US$45.

Orfali:19xx:CSPb

[OHxxa] Robert Orfali and Dan Harkey.

Client-Server Programming with Java and CORBA. La Nuova Italia, Firenze, Italy, 19xx. ISBN 80-471-1635-1 (??Invalid checksum??). ???? pp. LCCN ???? 108000 L.

Orfali:19xx:PJC

[OHxxb] Robert Orfali and Dan Harkey.

Programmieren mit Java und CORBA. Addison-Wesley Verlag, Bonn, Germany, 19xx. ISBN 3-8273-1324-4 (??invalid ISBN??). LCCN ???? 90 DM.

Orfali:1997:CJO

[OHE97a] Robert Orfali, Dan Harkey, and Jeri Edwards. CORBA, Java, and the Object Web: The Web is in trouble, CORBA and Java are out to save it. BYTE Magazine, 22(10):95–98, 100, October 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Orfali:1997:IC

[OHE97b] Robert Orfali, Dan Harkey, and Jeri Edwards. Instant CORBA. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, March 1997. ISBN 0-471-18333-4. xviii + 313 pp. LCCN QA76.9.C55O764 1997. US$20. URL http:

                                        //www.wiley.com/compbooks/
                                        catalog/18333-4.htm.

Oestreicher:1999:OLJ

[OK99] Marcus Oestreicher and Ksheerabdhi Krishna. Object lifetimes in Java Card. In USENIX [USE99d], page ?? ISBN 1-880446-34-0. LCCN TK7895.S62 U84 1999. URL

                                        http://www.usenix.org/publications/
                                        library/proceedings/smartcard99/
                                        oestreicher.html.

Okutsu:1998:WRJ

[OL98] Masayoshi Okutsu and Norbert Lindenberg. World ready Java

— the next step — part 2: Input method framework. In Unicode Consortium [Uni98a], page ?? ISBN ???? LCCN

                                        ????  URL http://www.
                                        unicode.org/unicode/iuc12/
                                        program.html.

Olsen:1996:EN

[Ols96] L. Olsen. The etaCOM network. In IEEE [IEE96r], page ?? ISBN 0-8186-7585-3. LCCN TK5101.A1C677 1996. IEEE catalog number 96TB100035.

Olsen:1997:LWJ

[Ols97] Kurt Olsen. Letters: Why Java isn’t watered down C++. Computer, 30(5):9, March 1997. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). Response to [Lew97b]. See also [Wel97].

Olympia:1996:CFA

[Oly96] P. L. Olympia. Cold Fusion 1.5, Autobahn. DBMS, 9(8):33– 34, 36, 38, July 1996. CODEN DBMSEO. ISSN 1041-5173.

ONeel:1997:BRJa

[O’N97a] Bruce O’Neel. Book review:

Java Network Programming. ;login: the USENIX Association newsletter, 22(4):33–36, August 1997. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

usenix.org/publications/java/
javaindex.html. Reviews of

[Har97a, HHSW97a].

ONeel:1997:BRJb

[O’N97b] Bruce O’Neel. Book review:

Java AWT Reference. ;login: the USENIX Association newsletter, 22(5):??, October 1997. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

usenix.org/publications/login/
1997-10/zukowski.html.

ONeel:1997:BRJ

[O’N97c] Bruce O’Neel. Book review:

Java Network Programming. ;login: the USENIX Association newsletter, 22(4):33–36, August 1997. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

usenix.org/publications/java/
javaindex.html. Reviews of

[Har97a, HHSW97b].

ONeel:1998:BRD

[O’N98a] Bruce O’Neel. Book review: Designing with JavaScript. ;login: the USENIX Association newsletter, 23(4):??, June 1998. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

usenix.org/publications/login/
1998-6/heinle.html.

ONeil:1998:JPG

[O’N98b] Joseph O’Neil. JavaBeans Programming from the Ground Up. Osborne/McGraw-Hill, Berkeley, CA, USA, March 1998. ISBN 0-07-882477-X. xxiv, 487 pp. LCCN QA76.73.J38 O817 1998. US$35. URL http:

                                        //www.osborne.com/program/
                                        javabeans.htm. Edited by

Herb Schildt.

Corporation:1995:OPI

[Ora95]  Oracle Corporation.  Ora-

cle PowerBrowser to integrate Java technology with Oracle’s Network Loadable Objects. press release, October

                                        30, 1995. URL http://www.
                                        oracle.com/info/news/java.
                                        html.

Orchard:1998:BPE

[Orc98] David Orchard. Better performance with exceptions in Java — you can use the language’s runtime error checking to improve a program’s performance. BYTE Magazine, 23(3):53–54, March 1998. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Orfali:1998:IC

[Orf98] Robert Orfali. Instant CORBA. Addison-Wesley Verlag, Bonn, Germany, January 1998. ISBN 3-8273-1325-2 (??invalid ISBN??). LCCN ???? 80 DM. URL

                                        http://www.addison-wesley.
                                        de/katalog/item.ppml?id=00270.

Ortiz:1999:ITB

[Ort99]  Sixto Ortiz Jr.  Industry

trends: The battle over realtime Java. Computer, 32(6): 13–15, June 1999. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

http://dlib.computer.org/
co/books/co1999/pdf/r6013.
pdf.

ONeal:1996:APJ

[OS96] Miles O’Neal and Tom Stewart. AWT Programming for Java. MIS Press and M&T Books, P. O. Box 5277, Portland, OR 97208-5277, USA, Tel: (503) 282-5215 and M&T Publishing, Inc., 501 Galveston Drive, Redwood City, CA 94063, USA, November 1996. ISBN 1-55851-494-5. xx + 444 pp. LCCN QA76.73.J38O82 1997. US$39.95. URL http:/

/www.idgbooks.com/cgi-bin/
db/fill_out_template.pl?idgbook:
1-55851-494-5:book-idg::uidg1388.

OConnor:1997:PJV

[OT97] J. Michael O’Connor and Marc Tremblay. picoJava-I: the Java Virtual Machine in hardware. IEEE Micro, 17(2):45– 53, March/April 1997. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Oualline:1996:HEM

[Oua96] Steve Oualline. Hype elevating mediocre Java to exalted status

— for a while. HP Chronicle, 13 (9):20, August 1996. ISSN 0892-

Oualline:1997:LGH

[Oua97] Steve Oualline. Linux and the gorilla are here to stay; not so Java. HP Chronicle, 14(2):15, January 1997. ISSN 0892-2829.

Ouellette:1995:NLP

[Oue95a] Tim Ouellette. News: Lotus prepares to meet Internet head on. Computerworld, page ??, December 15, 1995. CODEN CMPWAB. ISSN 0010-4841. URL http://www.

                                        computerworld.com/.

Ouellette:1995:NNU

[Oue95b] Tim Ouellette. News: Notes users eager to get upgrade. Computerworld, page ??, December 22, 1995. CODEN CM-PWAB. ISSN 0010-4841.

Ouellette:1996:AGN

[Oue96a] Tim Ouellette. AS/400 goes native with Java. Computerworld, 30(47):39, November 18, 1996. CODEN CMPWAB. ISSN 0010-4841. URL http://

                                        www.computerworld.com/.

Ouellette:1996:ISL

[Oue96b] Tim Ouellette. IBM software links Java browsers to mainframe. Computerworld, 30 (47):12, November 18, 1996. CODEN CMPWAB. ISSN 0010-4841. URL http://www.

                                        computerworld.com/.

Ouellette:1996:LLJ

[Oue96c] Tim Ouellette. Lotus lets Java play with Domino. Computerworld, 30(46):28, November 11, 1996. CODEN CMPWAB. ISSN 0010-4841.

Overland:1996:JPE

[Ove96] Brian R. Overland. Java in Plain English. MIS Press and M&T Books, P. O. Box 5277, Portland, OR 97208-5277, USA, Tel: (503) 282-5215 and M&T Publishing, Inc., 501 Galveston Drive, Redwood City, CA 94063, USA, October 1, 1996. ISBN 1-55828-503-2. xxiii + 552 pp. LCCN QA76.73.J38O83 1996. US$16.95. URL http://www.

mispress.com/catagories/internetprog.
html.

Overland:1997:JPE

[Ove97] Brian R. Overland. Java in Plain English. MIS Press and M&T Books, P. O. Box 5277, Portland, OR 97208-5277, USA, Tel: (503) 282-5215 and M&T Publishing, Inc., 501 Galveston Drive, Redwood City, CA 94063, USA, second edition, September 1997. ISBN 1-55828-563-6 (softcover). xxviii + 628 pp. LCCN QA76.73.J38O83 1997. US$17.95. URL http:/

/www.idgbooks.com/cgi-bin/
db/fill_out_template.pl?idgbook:
1-55828-563-6:book-idg::uidg781;
http://www.mandt.com/javaplaineng2e.
htm.

OReilly:1995:WWW

[OW95] O’Reilly and Associates and Web Consortium (W3C), editors. World Wide Web Journal: The Fourth International WWW Conference Proceedings.

O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1995. ISBN 1-56592-169-0. ISSN 1085-2301. LCCN TK5105.888 .I68 1995. US$39.95. URL http://www.

ora.com/gnn/bus/ora/item/

wj1.html. The World Wide Web Journal is a quarterly publication that provides timely, indepth coverage of the issues, techniques, and research developments in the World Wide Web. The December issue contains the Conference Proceeding papers that were chosen for the 4th International World Wide Web conference in Boston, MA.

Oaks:1997:JT

[OW97a] Scott Oaks and Henry Wong. Java Threads. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1997. ISBN 1-56592-216-6. xiii + 252 pp. LCCN QA76.73.J38O25 1997. US$29.95. URL http://www.

ora.com/catalog/jthreads/
noframes.html.

Odersky:1997:PJT

[OW97b] Martin Odersky and Philip Wadler. Pizza into Java: translating theory into practice. In ACM [ACM97a], pages 146–159. ISBN 0-89791-853-3. LCCN QA 76.7

A11 1997. URL http://www.
acm.org:80/pubs/citations/
proceedings/plan/263699/p146-
odersky/.

Oaks:1999:JT

[OW99] Scott Oaks and Henry Wong.

Java Threads. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 1999. ISBN 1-56592-418-5. xiii + 319 pp. LCCN QA76.73.J38 O25 1999. US$32.95.

Oaks:19xx:JTb

[OWxx] Scott Oaks and Henry Wong.

Java Threads. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 19xx. ISBN 4-900900-54-0. ???? pp. LCCN ???? 3,400 yen. URL http:

//www.oreilly.co.jp/BOOK/

jthread.htm. Japanese translation.

Owen:1996:IWW

[Owe96] G. Scott Owen. Integrating World Wide Web technology into undergraduate education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28 (??):101–103, ???? 1996. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Pedrick:1998:PVC

[P^{+}98a] Doug Pedrick et al. Programming with VisiBroker (CORBA & JDBC). John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, February 1998. ISBN 0-471-23901-1. xvi + 435 pp. LCCN QA76.73.J38 P79 1998. US$45.

                                        URL http://www.wiley.com/
                                        compbooks/catalog/23901-1.
                                        htm.

Powers:1998:DWP

[P^{+}98b] Shelley Powers et al. Dynamic Web Publishing Unleashed (HTML, Java Script, CGI, Java, Style Sheets). Sams.net, Indianapolis, IN 46268, USA, second edition, 1998. ISBN 1-57521-363-X. xxviii + 844 pp. LCCN TK5105.888.D96 1998.

                                        US$40. URL http://iq-mcp3.
                                        iquest.net/catalog/mcp/PRODUCT/
                                        PAGE/15752/bud/157521363X.
                                        html.

Parsons:1995:IEH

[Par95a] Michael Parsons. Interview: Eric Hahn of Collabra talks Netscape futures. IDG International News Service, page ??, October 06, 1995.

Parsons:1995:INC

[Par95b] Michael Parsons. Interview: Next Computer’s Landwehr talks WebObjects. IDG International News Service, page ??, August 15, 1995.

Parsons:1995:NIJ

[Par95c] Michael Parsons. News: IBM jumps on Java bandwagon. IDG International News Service, page ??, December 06, 1995.

Parsons:1995:NML

[Par95d] Michael Parsons. News: Microsoft launches VB Script, deals with Oracle, Spyglass.

IDG International News Service, page ??, December 07, 1995.

Parsons:1995:NMW

[Par95e] Michael Parsons. News: Microsoft will license Java and give away browser. IDG International News Service, page ??, December 07, 1995.

Parsons:1995:NSS

[Par95f] Michael Parsons. News: Sun-World: Sun takes lid off Java, Netscape smells coffee. IDG International News Service, page ??, May 24, 1995.

Parsons:1995:NWIa

[Par95g] Michael Parsons. News: Web innovation: SGI, Sun and Netscape boost VRML, Java. IDG International News Service, page ??, December 05, 1995.

Parsons:1995:NWIb

[Par95h]  Michael Parsons.  News:

Web innovations: Andreessen launches JavaScript. IDG International News Service, page ??, December 05, 1995.

Parr:1997:LTU

[Par97] Terence John Parr. Language translation using PCCTS and C++: a reference guide. Automata Publishing Company, San Jose, CA, USA, January 1997. ISBN 0-9627488-5-4. v + 310 pp. LCCN ???? US$34.95. URL http://www.

parr-research.com/~parrt/

buybook.html. PCCTS is a precursor to the ANTLR system [PLS98].

Parson:1999:UJR

[Par99] Dale Parson. Using Java reflection to automate extension language parsing. In USENIX [USE99a], page ?? ISBN 1-880446-27-8, 1-58113-255-7. LCCN QA76.7 .C663

                                        1999.  URL http://www.
                                        usenix.org/publications/library/
                                        proceedings/dsl99/parson.
                                        html.

Pattison:1998:MVJ

[Pat98a] Ted Pattison. Microsoft Visual J++ Datenbank-Workshop. Microsoft Press Germany, ????, March 1998. ISBN 3-86063-438-

0. LCCN ???? 98 DM.

Pattison:1998:VJD

[Pat98b] Ted Pattison. Visual J++ Database Programming. Microsoft Press, Bellevue, WA, USA, February 1998. ISBN 1-57231-689-6. LCCN ???? US$40.

Paternostro:1999:JSC

[Pat99] Pat Paternostro. A Java skeleton code generator. C/C++ Users Journal, 17(7):34, 36–38, July 1999. CODEN CCUJEX. ISSN 1075-2838.

Petit-Bianco:1998:JGC

[PB98] Alexandre Petit-Bianco. Java garbage collection for real-time systems. Dr. Dobb’s Journal of Software Tools, 23(10):20–22, 24, 26–29, October 1998. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/1998/1998_10/
kaffe.zip.

Plezbert:1997:DQT

[PC97] Michael P. Plezbert and Ron K.

Cytron. Does ‘just in time’ = ‘better late than never’? In ACM [ACM97a], pages 120–131. ISBN 0-89791-853-3. LCCN QA 76.7 A11

1997.  URL http://www.
acm.org:80/pubs/citations/
proceedings/plan/263699/p120-
plezbert/.

Pew:19xx:CJ

[PCxx] John A. Pew and John Cable. Core JavaBeans. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 19xx. ISBN 0-13-786294-6. LCCN ???? ????

Paepcke:1998:IDL

[PCWGM98] Andreas Paepcke, Chen-Chuan K.

Chang, Terry Winograd, and Héctor García-Molina. Interoperability for digital libraries worldwide. Communications of the ACM, 41(4):33–42, April 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org:80/pubs/citations/
journals/cacm/1998-41-4/p33-
paepcke/.

Pelline:1995:NSJ

[Pel95] Jeff Pelline. New simpler Java Script program to help Web builders. New York Times Syndicate, page ??, December 2,

                                        1995. URL http://nytsyn.
                                        com/live/News3/336_120295_
                                        185637_6719.html.

Pelletier:1996:PAJ

[Pel96] Guillaume Pelletier. Programmer la 3D avec Java et C++. Sybex France, Paris, France, 1996. ISBN 2-7361-2275-5. xxi + 439 pp. LCCN ????

                                        229 FF. URL http://www-
                                        db.sybex.fr/cgi-bin/Sybex/
                                        Webdriver?MIval=Fiche_Book.
                                        html&Reference=2275.

Pepper:1997:CPB

[Pep97] Jon Pepper. Corel PDA to bundle Java suite. BYTE Magazine, 22(1):38, January 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Perkins:1996:JPP

[Per96a] Charles Perkins. Java Power Pack C/W95/Ww. Sams.net, Indianapolis, IN 46268, USA, 1996. ISBN 1-57521-074-6. ???? pp. LCCN ???? US$79.95.

Perry:1996:UVJ

[Per96b]  Greg Perry.  Using Visual

J++: Master the Essentials of Microsoft’s New Visual Java Development Environment. Que Corporation, Indianapolis, IN, USA, 1996. ISBN 0-7897-0899-X. xix + 384 pp. LCCN QA76.73.J38P47 1996. US$34.99, CDN$49.95, UK£32.49. URL http://

                                        merchant.superlibrary.com:
                                        8000/catalog/hg/PRODUCT/PAGE/
                                        07897/bud/078970899X.html.

Perry:1996:CCW

[Per96c] Paul J. Perry. Creating Cool Web Applets With Java. I D G Books Worldwide, Indianapolis, IN, USA, May 1, 1996. ISBN 1-56884-881-1. xxiv + 355 pp. LCCN TK5105.888.P4227 1996. US$29.99. URL http://db.

www.idgbooks.com/database/
book/isbn/generic-book.tmpl?
query=1-56884-881-1;   http:
//www.idgbooks.com/database/
book_result.msql?uid=uidg323&
isbn=1-56884-881-1. Includes

CD-ROM.

Perry:1997:UVJ

[Per97a]  Greg Perry.  Usando Vi-

sual J++. Editore Campus Ltda., Rua Sete de Setembro, 111/16, Rio de Janeiro - RJ - 20050-002, Brazil, 1997. ISBN ???? 428 pp. LCCN ???? R$ 53,91. URL http:

//www.campus.com.br/about-
us.htm;   http://www.campus.
com.br/catalogo/livro.cfm?
cod_livro=20148.

Perry:1997:VJS

[Per97b]  Greg Perry.  Visual J++

Starter Kit. Que Corporation, Indianapolis, IN, USA, 1997. ISBN 0-7897-1282-2. xix + 395 pp. LCCN ???? US$35. URL http://

merchant.superlibrary.com:
8000/catalog/hg/PRODUCT/PAGE/
07897/bud/0789712822.html.

Perry:1997:CCW

[Per97c] Paul J. Perry. Creating Cool Web Applets with Java. Softbank Publishing, ??, Japan, 1997. ISBN 4-7973-0146-5. ???? pp. LCCN ???? 2500 yen.

                                        URL http://www.softbank.
                                        co.jp/softbank/publishing/
                                        ;   http://www.softbank.co.
                                        jp/softbank/publishing/books/
                                        editer/os1/JavaSS/index.html;
                                         http://www.softbank.co.jp/
                                        softbank/publishing/books/
                                        editer/os1/javass/index.html.

Perry:1997:DVJ

[Per97d] Paul J. Perry. Discover Visual J++. IDG Books, San Mateo, CA, USA, 1997. ISBN 0-7645-3077-1. xxiii + 390 pp. LCCN QA76.73.J38P474 1997¡. US$25. URL http://

                                        www.idgbooks.com/database/
                                        book_result.msql?isbn=0-7645-
                                        3077-1.

Perry:19xx:MVJ

[Perxxa]  Greg Perry.  Minidoc Vi-

sual J++. Simon & Schuster Macmillan (France), Paris, France, 19xx. ISBN 2-7440-0230-5. ???? pp. LCCN ???? 79

                                        FF. URL http://www.ssm.fr/
                                        catalog/fiches/0230.htm.

Perry:19xx:UVJs

[Perxxb] Greg Perry. Usando Visual J++. Editore Campus Ltda., Rua Sete de Setembro, 111/16, Rio de Janeiro - RJ - 20050-002, Brazil, 19xx. ISBN ????

LCCN ????  ????  URL

                                        http://www.campus.com.br/
                                        catalogo/livro.cfm?cod_livro=
                                        20148.

Perry:19xx:UVJb

[Perxxc] Greg Perry. Usare Visual J++.

Jackson Libri, via XXV Aprile, 39 - 20091 Bresso (MI), Italy, 19xx. ISBN ???? LCCN ???? 45000 L.

Petreley:1995:CW

[Pet95] Nicholas Petreley. Column:

Down to the wire. InfoWorld, page ??, December 18, 1995. CODEN INWODU. ISSN 0199-6649. URL http://www.

infoworld.com/.

Peterson:1997:DSI

[Pet97] David Peterson. Developing Smarter Intelligent Agents Using Java. McGraw-Hill, New York, NY, USA, March 1997. ISBN 0-07-913643-5. 400 pp. LCCN ???? US$44.95.

Pew:1996:IJ

[Pew96] John A. Pew. Instant Java.

Java series. SunSoft Press, 2550 Garcia Avenue, Mountain View, CA 94043-1100, USA, March 1996. ISBN 0-13-565821-7. xi + 340 pp. LCCN QA76.73.J38 P48 1996. US$29.95. URL http://www.

sun.com/smi/ssoftpress/books/
Pew2/Pew2.html.

Pew:1997:IJ

[Pew97a] John A. Pew. Instant Java.

ASCII Corporation, Tokyo, Japan, 1997. ISBN ????

????  pp.  LCCN ????

URL http://www.ascii.co.
jp/;   http://www.vivids.

com/java/author.html. Japanese

translation of [Pew97a].

Pew:1997:IJA

[Pew97b] John A. Pew. Instant Java — Mit Applets ins WWW. Verlag Heinz Heise, Helsdorfer Straße 7, D-30625, Hannover, Germany, 1997. ISBN 3-88229-084-6. 246 (or 245??) pp. LCCN ???? 79,90 DM (or 79,80 DM??). URL

http://www.emedia.de/bin/
bookshop?show=4552&id=;
http://www.heise.de/buch/

. German translation of [Pew97a]. Includes CD-ROM.

Pew:1997:SJ

[Pew97c] John A. Pew. Subito Java. Mondadori Informatica, ??, Italy, 1997. ISBN ???? ???? pp. LCCN ???? 70000 L.

URL http://www.mondadori.
com/;   http://www.vivids.
com/java/author.html. Ital-

ian translation of [Pew97a].

Pew:19xx:IJ

[Pewxxa] John A. Pew. Instant Java. ASCII Corporation, Tokyo, Japan, 19xx. ISBN 4-7561-1684-1. ???? pp. LCCN

????  ????  URL http:/

/www.ascii.co.jp/pb/book1/
shinkan/detail/1322237.html.

Japanese translation.

Pew:19xx:SJ

[Pewxxb] John A. Pew. Subito Java. Mondadori Informatica, ??, Italy, 19xx. LCCN ???? 70000 L.

Potts:1996:JPL

[PF96] Anthony Potts and David H.

Friedel, Jr. Java Programming Language Handbook: The Ultimate Source for Conquering the Java Programming Language. Internet programming series. Coriolis Group Books, Scottsdale, AZ, USA, May 1, 1996. ISBN 1-883577-77-2. xvii + 414 pp. LCCN QA76.73.J38P68 1996. US$24.99; CDN$34.99.

URL http://165.247.200.
139/cgi-bin/books.cgi?page=
books&subject=jprogl&graphics=
0;   http://www.coriolis.
com/site/msie/books/ind/jprogl.
htm.

Petrich:1997:NIN

[PF97] Dean Petrich and David Flanagan. Netscape IFC in a nutshell: a desktop quick reference for Java programmers. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1997. ISBN 1-56592-343-X. xiii + 354 pp. LCCN QA76.73.J38 P38 1997. US$19.95. URL http://www.

ora.com/catalog/ifcnut/. Cov-

ers IFC 1.1.

Pfeiffer:1999:WEW

[Pfe99] Phil Pfeiffer. What employers want from students: a report from OOPSLA. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(2):69–70, June 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Pandey:1999:PFG

[PH99a] R. Pandey and B. Hashii. Providing fine-grained access control for Java programs. Lecture Notes in Computer Science, 1628:449–??, 1999. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Perelman-Hall:1999:JLC

[PH99b] David K. Perelman-Hall. Java and lightweight components. Dr. Dobb’s Journal of Software Tools, 24(2):22–24, 26– 28, February 1999. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/1999/9902/9902toc.
                                        htm;   http://www.ddj.com/
                                        ftp/1999/1999_02/lightwei.
                                        txt;   http://www.ddj.com/
                                        ftp/1999/1999_02/lightwei.
                                        zip.

Phillips:1996:JYR

[Phi96] Paul Phillips. Java 1.0 — are you ready for the Java revolution? Database Directory: the buyer’s guide to database products and services, 14(5):42–??, ???? 1996. CODEN DBADEL. ISSN 0740-5200.

Phipps:1999:COB

[Phi99] Geoffrey Phipps. Comparing observed bug and productivity rates for Java and C++. Software—Practice and Experience, 29(4):345–358, April 10, 1999. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http:

                                        //www3.interscience.wiley.

com/cgi-bin/abstract?ID=55001844;
http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=55001844&PLACEBO=IE.pdf.

Poetzsch-Heffter:1999:PLS

[PHM99] A. Poetzsch-Heffter and P. Mueller.

A programming logic for sequential Java. Lecture Notes in Computer Science, 1576:162– 176, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Picarille:1996:CAB

[Pic96a] Lisa Picarille. Corel aims to be first to give Java a suite taste. Computerworld, 30(22):8, May 27, 1996. CODEN CMPWAB. ISSN 0010-4841.

Picarille:1996:FAF

[Pic96b] Lisa Picarille. Firms add flavor to Java menu: JavaSoft’s Beans allows reuse of apps. Computerworld, 30(23):14, June 3, 1996. CODEN CMPWAB. ISSN 0010-4841.

Picarille:1996:GSJ

[Pic96c] Lisa Picarille. A groggy start for Java products. Computerworld, 30(24):49–52, June 10, 1996. CODEN CMPWAB. ISSN 0010-4841.

Picarille:1996:JGS

[Pic96d] Lisa Picarille. Java gets a suite taste. Computerworld, 30(48): 47, 51, November 25, 1996. CO-DEN CMPWAB. ISSN 0010-

Picarille:1997:JAS

[Pic97] Lisa Picarille. Java apps serve up flexibility. Computerworld, 31(6):43–44, February 10, 1997. CODEN CMPWAB. ISSN 0010-4841.

Piemont:1997:BPJ

[Pie97] Claudia Piemont. Businessorientierte Programmierung mit Java: Der Weg zur effizienten Entwicklung von Geschäftsanwendungen im Intranet und Internet. Vieweg & Son, Braunschweig, Germany, 1997. ISBN 3-528-05556-1. 440 pp. LCCN ???? 88

DM. URL http://members.
aol.com/ClaudiaPi/jbook1.
htm.

Piemont:19xx:JBE

[Piexx] Claudia Piemont. Java für den Business Einsatz. Vieweg & Son, Braunschweig, Germany, 19xx. ISBN ???? ???? pp. LCCN ???? 88 DM. Includes CD-ROM.

Pastor:1997:LOC

[PIPR97a] O. Pastor, E. Insfran, V. Pelechano, and S. Ramirez. Linking objectoriented conceptual modelling with object-oriented implementation in Java. Lecture Notes in Computer Science, 1308:132– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pastor:1997:LOO

[PIPR97b] O. Pastor, E. Insfran, V. Pelechano, and S. Ramirez. Linking objectoriented conceptual modelling with object-oriented implementation in Java. Lecture Notes in Computer Science, 1308:132– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Piroumian:1997:ISJ

[Pir97] Vartan Piroumian. Internationalization support in Java: Making software globalization easy. IEEE Micro, 17(3):20–29, May/ June 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//pascal.computer.org/mi/
books/mi1997/pdf/m3020.pdf.

Pistoia:1999:JNS

[Pis99] Marco Pistoia. Java 2 Network Security. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1999. ISBN 0-13-015592-6. xxi + 713 pp. LCCN QA76.9.A25 J38 1999. US$49.99. URL http:

//www.phptr.com/ptrbooks/
ptr_0130155926.html.

Pitts:1997:SDL

[Pit97] Robert I. Pitts. Software development: Learning to program the object-oriented way. SunExpert Magazine, 8(1):56– 62, 64, 66, 68, 70, 71, January 1997. ISSN 1053-9239. Discusses C++, Objective C, and Java.

Pitt:1998:JQH

[Pit98a] W. David Pitt. Java Q and A: How do I use servlets?

Dr. Dobb’s Journal of Software Tools, 23(12):122, 124– 125, December 1998. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ddj/1998/1998_12/
../../../ftp/1998/1998_12/
jqa128.txt;   http://www.
ddj.com/ddj/1998/1998_12/
../../../ftp/1998/1998_12/
jqa128.zip.

Pitt:1998:VPJ

[Pit98b] W. David Pitt. The visitor pattern and a Java grep utility. Dr. Dobb’s Journal of Software Tools, 23(6):30, 32, 93, June 1998. CODEN DDJOEB. ISSN 1044-789X.

Pitt:1999:JQC

[Pit99] W. David Pitt. Java Q&A: Can Java handle exception handling? Dr. Dobb’s Journal of Software Tools, 24(8):103, 106–108, August 1999. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/1999/1999_08/
jqa899.txt;   http://www.
ddj.com/ftp/1999/1999_08/
jqa899.zip.

Prigozhin:1998:LDC

[PJR^{+}98] Alexander Prigozhin, Steve Johnson, Jonathan Rosenne, Sy Wong, Scott Raney, Andreas Eder, and Arthur Nunes. Letters: Design by contract; the standardization process; international Java; to:

<KJPrice@bellatlantic.net>;

the Metacard language; symbolic integration. Dr. Dobb’s Journal of Software Tools, 23 (4):10, 12, April 1998. CO-DEN DDJOEB. ISSN 1044-

789X. URL http://www.ddj.
com/.

Park:1997:CJ

[PK97] Kibae Park and Hodong Kim.

Catch Java. HanCom Press, ??, Korea, 1997. ISBN 89-85770-50-0 (??Invalid checksum??). ???? pp. LCCN ???? 10,000

won. URL http://magazine.
hnc.net/ (??).

Poo:1998:OOP

[PK98] Danny C. C. Poo and Derek B. K. (Derek Beng Kee) Kiong. Object-Oriented Programming in Java. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. ISBN 981-3083-96-4. xvii + 316 pp. LCCN QA76.64.P65 1998. US$24.95. Guest contributor: Dennis Lee.

Papurt:1996:DAS

[PL96] D. M. Papurt and J. P. LeJacq.

Design aspects of the standard I/O library. Journal of Object Oriented Programming, 9(6):6– ??, 8, 10, 12, October 1996. CODEN JOOPEC. ISSN 0896-8438.

Papurt:1997:SUM

[PL97] D. M. Papurt and J. P. LeJacq.

The sensible use of method overriding. Journal of Object Oriented Programming, 10(1):62–5, 71, March–April 1997. CODEN JOOPEC. ISSN 0896-8438.

Plauger:1996:CCEb

[Pla96a] P. J. Plauger. C/C++ editor’s forum: Transliterating C and C++ code to Java. C/C++ Users Journal, 14(9):6, September 1996. CODEN CCUJEX. ISSN 1075-2838.

Plauger:1996:CCEa

[Pla96b] P. J. Plauger. C/C++ editor’s forum: Uproar over Java. C/C++ Users Journal, 14(5):6, May 1996. CODEN CCUJEX. ISSN 1075-2838.

Plauger:1996:SA

[Pla96c] P. J. Plauger. State of the art. Embedded Systems Programming, 9(8):93–??, August 1, 1996. CODEN EYPRE4. ISSN 1040-3272.

Plauger:1997:EFJ

[Pla97] P. J. Plauger. Editor’s forum: a Java to C translator. C/C++ Users Journal, 15(6):6, June 1997. CODEN CCUJEX. ISSN 1075-2838. URL http://www.

                                        dinkumware.com/;   http://
                                        www.tiac.net/users/pjp.

Parr:1998:AXR

[PLS98] Terence John Parr, John Lilley, and Scott Stanchfield. ANTLR 2.xx Reference Manual. ????, ????, 1998. ISBN ???? v + 310 pp. LCCN ???? US$34.95. URL

                                        http://www.parr-research.
                                        com/antlr/doc.html;   http:
                                        //www.parr-research.com/antlr/
                                        doc/antlr2story.html;   http:
                                        //www.parr-research.com/antlr/
                                        parr.phd.thesis.ps.zip;

http://www.parr-research.
com/antlr/performance.html.

ANTLR [PQ95] is a redesign and rewrite of PCCTS [PQ96] in Java. ANTLR generates parsers in C++ and Java. This book is still in preparation, and the title has not been settled.

Pappas:1996:JBC

[PM96a] Chris H. Pappas and William H.

Murray. Java With Borland C++. AP Professional, Boston, MA, USA, August 1996. ISBN 0-12-511960-7. xi + 428 pp. LCCN QA76.73.J38P36 1996. US$34.95. URL http://www.

apcatalog.com/cgi-bin/AP?
ISBN=0125119607&LOCATION=
US&FORM=FORM2;   http://www.
apnet.com/.

Patel:1996:JDP

[PM96b] Pratik Patel and Karl Moss.

Java Database Programming With JDBC. Coriolis Group Books, Scottsdale, AZ, USA, 1996. ISBN 1-57610-056-1. xiii + 480 pp. LCCN QA76.73.J38P38 1996. US$39.99. Includes CD-ROM.

Patel:1997:JDPb

[PM97a] Pratik Patel and Karl Moss.

Java Database Programming with JDBC. Ohmsha, ????, October 1997. ISBN 4-900718-55-6. LCCN ???? 2400

yen. URL http://www.ohmsha.
co.jp/data/books/contents/
4-900718-55-6.htm.

Patel:1997:JDPa

[PM97b] Pratik Patel and Karl Moss. Java Database Programming With JDBC: Discover the Essentials for Developing Databases for Internet and Intranet Applications. Coriolis Group Books, Scottsdale, AZ, USA, second edition, 1997. ISBN 1-57610-159-2 (softcover). xviii + 491 pp. LCCN QA76.73.J38P38 1997. US$35.99. URL http://www.

coriolis.com/Site/mise/books/
ind/javjdbc2.htm.

Peter:1997:ONJ

[PM97c] Kent Peter and Kent Multer. Official Netscape JavaScript 1.2 Programmers Reference: Windows, Macintosh, and Unix. Ventana Press, Chapel Hill, NC, USA, eleventh edition, October 1997. ISBN 1-56604-757-9 (softcover). xxxii + 490 pp. LCCN QA76.73.J39 K458 1997. US$35.99. Includes CD-ROM.

Pountain:1997:WCC

[PM97d] Dick Pountain and John Montgomery. Web components: Components and the Web are a match made in developer heaven. BYTE Magazine, 22 (8):56–60, 62, 64, 66, 68, August 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic). Discusses Sun’s Java Beans, Microsoft’s ActiveX, IBM’s San Francisco Project, and OMG’s CORBA.

Purcell:1997:AJ

[PM97e] Lee Purcell and Mary Jane Mara. The ABCs of JavaScript. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, January 1, 1997. ISBN 0-7821-1937-9. xxi + 456 pp. LCCN QA76.73.J38P87 1997. US$29.99.

Poggi:1996:IAO

[Pog96] A. Poggi. Integrating agents and objects to develop distributed AI systems. In IEEE [IEE96i], pages 443–444. CO-DEN PCTIFX. ISBN 0-8186-7686-8. ISSN 1063-6730. LCCN Q 334 I564 1996.

Polan:1998:TNU

[Pol98] M. G. Polan. Technical note:

Using the San Francisco frameworks with VisualAge for Java. IBM Systems Journal, 37(2): 215–??, ???? 1998. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

URL http://www.almaden.
ibm.com/journal/sj/372/polan.
html.

Pont:1998:SWJ

[Pon98]  Michael J. Pont.  Soap-

box: Why Java is dangerous. IEEE Software, 15 (1):20–22, January/February 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

http://dlib.computer.org/
so/books/so1998/pdf/s1020.
pdf.

Pope:1997:CRG

[Pop97] Alan Pope. The CORBA Reference Guide: Understanding the Common Object Request Broker Architecture. Addison-Wesley, Reading, MA, USA, December 1997. ISBN 0-201-63386-8. ???? pp. LCCN QA76.64 .P66 1997. US$34.

Port:1998:SSI

[Por98]  Dan Port.  SIM — Sys-

tems Integration Methodology with Java Examples. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., February 1998. ISBN 0-387-98359-7. ???? pp. LCCN ???? US$55.

Potts:1999:WGJ

[Pot99] Steve Potts. The Waite Group’s Java 1.2 how-to. Waite Group Press, Corte Madera, CA, USA, 1999. ISBN 1-57169-157-X. LCCN QA76.73.J38 P69 1999. US$40.

pratimPal:1998:FAL

[pP98] Partha pratim Pal. A flexible, applet-like software distribution mechanism for Java applications. ACM SIGSOFT Software Engineering Notes, 23(4):56–60, July 1998. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Pew:1999:IJ

[PP99] John Pew and Stephen Pew. Instant Java 1.2. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1999. ISBN 0-13-010533-3. xxiv + 275 pp. LCCN QA76.73.J38 P48 1999. US$34.99.

Pastor:1998:OOC

[PPIG98] O. Pastor, V. Pelechano, E. Insfran, and J. Gomez. From object oriented conceptual modeling to automated programming in Java. Lecture Notes in Computer Science, 1507:183– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pickering:1996:JHD

[PPJ96] George Pickering, Shelley Powers, and Ron Johnson. JavaScript How-To: The Definitive JavaScript Problem-Solver. Waite Group Press, Corte Madera, CA, USA, 1996. ISBN 1-57169-047-6. xvi + 598 pp. LCCN QA76.73.J39 P53 1996. US$44.99.

Parr:1995:APL

[PQ95] Terence J. Parr and Russell W. Quong. ANTLR: a predicated-LL(k) parser generator. Software—Practice and Experience, 25(7):789–810, July 1995. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL ftp://

ftp.parr-research.com/pub/
pccts/papers/antlr.ps;   ftp:
//ftp.parr-research.com/pub/
pccts/papers/needlook.ps;
 ftp://ftp.parr-research.
com/pub/pccts/papers/predicates.
ps.Z;   ftp://ftp.parr-research.
com/pub/pccts/papers/sorcerer.

                                        ps.Z;   http://java.magelang.
                                        com/antlr/entry.html;   http:
                                        //www.parr-research.com/antlr/
                                        parr.phd.thesis.ps.zip.

Parr:1996:LLT

[PQ96] Terence J. Parr and Russell W. Quong. LL and LR translators need k > 1 lookahead. ACM SIGPLAN Notices, 31(2): 27–34, February 1996. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Prechelt:1999:TOC

[Pre99] Lutz Prechelt. Technical opinion: comparing Java vs. C/ C++ efficiency differences to interpersonal differences. Communications of the ACM, 42 (10):109–112, October 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL http://www.
                                        acm.org:80/pubs/citations/
                                        journals/cacm/1999-42-10/
                                        p109-prechelt/.

Preis:19xx:EJ

[Prexx] Rotter Preis. Einblicke in Java. Carl Hanser, München, Germany, 19xx. ISBN 3-446-18908-4. ???? pp. LCCN ???? 49 DM.

                                        URL http://www.hanser.de/.

Includes CD-ROM.

Prosise:1996:JPL

[Pro96] Jeff Prosise. The Java programming language is making the Web even more enticing: Find out what it’s all about. PC Magazine, 15(14):279–??, August 1, 1996. CODEN PCMGEP. ISSN 0888-8507.

Prosise:19xx:DMA

[Proxx] Jeff Prosise. In Deep with MFC/ ActiveX. Image Art Publishing, ????, 19xx. ISBN 957-9641-27-7. ???? pp. LCCN ???? 1000 NT$. Chinese translation by Liu The Ming.

Pryor:1996:JSH

[Pry96] Doug Pryor. Java and some headaches. SunExpert Magazine, 7(2):4, February 1996. ISSN 1053-9239.

Pazandak:1996:IMM

[PS96] P. Pazandak and J. Srivastava. Interactive multi-user multimedia environments on the Internet: an overview of DAMSEL and its implementation. In IEEE [IEE96p], pages 287–290. ISBN 0-7803-3236-9 (softbound), 0-7803-3237-7 (microfiche). LCCN HG176.7.I33 1996. IEEE Cat. No. 96TH8177.

Pew:1997:IJS

[PS97] John A. Pew and Sunsoft Books Staff. Instant Java: Signature Edition. SunSoft Press Java series. Prentice-Hall and Sun-Soft Press, Englewood Cliffs, NJ 07632, USA and 2550 Garcia Avenue, Mountain View, CA 94043-1100, USA, second edition, January 1997. ISBN 0-13-272287-9. xxviii + 423 pp. LCCN QA76.73.J38P48 1997.

US$30.  URL http://www.
prenhall.com/ptrbooks/ptr_
0132722879.html;   http:/

/www.sun.com/books/books/
Pew2/Pew2.html. Available in

German translation [Pew97b]. Includes CD-ROM.

Plasil:1998:AVD

[PS98] František Plášil and Michael Stal. An architectural view of distributed objects and components in CORBA, Java RMI and COM/DCOM. Software

— Concepts and Tools, 19 (1):14–28, March 1998. CO-DEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (electronic). URL https://

link.springer.com/article/
10.1007/s003780050004.

Passerone:1998:MRS

[PSL^{+}98] C. Passerone, C. Sansoe, L. Lavagno, R. McGeer, J. Martin, R. Passerone, and A. Sangiovanni-Vincentelli. Modeling reactive systems in Java. ACM Transactions on Design Automation of Electronic Systems, 3(4):515–523, October 1998. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL http://

www.acm.org/pubs/articles/
journals/todaes/1998-3-4/
p515-passerone/p515-passerone.
pdf;   http://www.acm.org/
pubs/citations/journals/todaes/
1998-3-4/p515-passerone/.

Pedroso:1997:WBM

[PSS97] Hernâni Pedroso, Luis M. Silva, and João Gabriel Silva. Webbased metacomputing with JET. Concurrency: Practice and Experience, 9(11):

1169–1173, November 1997. CODEN CPEXEI. ISSN 1040-3108. URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=13834;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=13834&PLACEBO=IE.pdf. Spe-

cial Issue: Java for computational science and engineering — simulation and modeling II.

Payne:1998:IAJ

[PSS98] Jeffery E. Payne, Michael A.

Schatz, and Matthew Schmid. Implementing assertions for Java: Finding bugs early. Dr. Dobb’s Journal of Software Tools, 23(1):40, 42–44, January 1998. CODEN DDJOEB. ISSN 1044-789X.

Prince:1999:DIO

[PSTZ99] Robert Prince, Jianwen Su, Hong Tang, and Yonggang Zhao. The design of an interactive online help desk in the Alexandria digital library. ACM SIGSOFT Software Engineering Notes, 24(2):217–226, March 1999. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Pfitzmann:1998:JBD

[PSW^{+}98] A. Pfitzmann, A. Schill, A. Westfeld, G. Wicke, G. Wolf, and J. Zöllner. A Java-based distributed platform for multilateral security. Lecture Notes in Computer Science, 1402:52– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/
                                        service/series/0558/bibs/
                                        1402/14020052.htm;   http:
                                        //link.springer-ny.com/link/
                                        service/series/0558/papers/
                                        1402/14020052.pdf.

Proebsting:1997:TJA

[PTB^{+}97] Todd A. Proebsting, Gregg Townsend, Patrick Bridges, John H. Hartman, Tim Newsham, and Scott A. Watterson. Toba: Java for applications — a Way Ahead of Time (WAT) compiler. In USENIX [USE97b], pages 41– 53. ISBN 1-880446-86-3. LCCN QA76.64 .U845 1997. URL

                                        ftp://ftp.cs.arizona.edu/
                                        reports/1997/TR97-01.pdf;
                                         ftp://ftp.cs.arizona.edu/
                                        reports/1997/TR97-01.ps.

Proebsting:1999:JJB

[PTTA99] Todd Proebsting, Gregg Townsend, Denise Todd, and Bob Alexander. Jcon: a Java-based Icon implementation. Version 2 of Jcon is an essentially complete implementation of Icon, including graphics and large integers. A few minor features are missing — mostly things like chdir() that cannot be done in Java.,

                                        1999. URL http://www.cs.
                                        arizona.edu/icon/jcon/.

Pugh:1999:CJC

[Pug99] William Pugh. Compressing Java class files. ACM SIG-PLAN Notices, 34(5):247–258, May 1999. CODEN SINODQ.

ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.
acm.org:80/pubs/citations/
proceedings/pldi/301122/p247-

pugh/. See PLDI’99 proceedings [ACM99b].

Purcell:1996:MJ

[Pur96] Lee Purcell. Mastering Javascript.

Prima Publishing, Rocklin, CA, USA, November 1996. ISBN ???? 550 pp. LCCN ???? US$40.00.

PJDJ:1997:PJD

[Pur97] Pure Java Developer’s Journal, 1997. ISSN 1093-7471. Cobb Group, Louisville, KY, USA. Publication began with volume 1, number 1, in July 1997.

Purdy:1998:JQ

[Pur98] Jason W. Purdy. Java Q and A: How does Java drag-and-drop work? Dr. Dobb’s Journal of Software Tools, 23(11): 127–130, November 1998. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/1998/1998_11/
JQA118.TXT;   http://www.
ddj.com/ftp/1998/1998_11/
JQA118.ZIP.

Purcell:19xx:WMJ

[Purxx]  Lee Purcell.  Werken met

JavaScript. Sybex Nederland, ????, 19xx. ISBN 90-5160-997-3. ???? pp. LCCN

????  fl. 49,00.  URL

http://www.sybex.nl/nieuw/
wmjavas.htm.

Pusch:1999:PSJ

[Pus99] C. Pusch. Proving the soundness of a Java bytecode verifier specification in Isabelle/HOL. Lecture Notes in Computer Science, 1579:89–103, 1999. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Posegga:1998:BCV

[PV98] J. Posegga and H. Vogt. Byte code verification for Java smart cards based on model checking. Lecture Notes in Computer Science, 1485:175–??, 1998. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pinho:1999:AAA

[PV99] Luís Miguel Pinho and Francisco Vasques. To Ada or not to Ada: Adaing vs. Javaing in real-time systems. ACM SIGADA Ada Letters, 19(4): 37–43, December 1999. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Proebsting:1997:KDJ

[PW97] Todd A. Proebsting and Scott A. Watterson. Krakatoa: Decompilation in Java (does bytecode reveal source?). In USENIX [USE97b], pages 185– 197. ISBN 1-880446-86-3. LCCN QA76.64 .U845 1997.

URL http://dxsting.cern.
ch/sting/COOTS97.html. The

authors demonstrate that, because of the strongly-type nature of the Java Virtual Machine, it is possible to reverseengineer Java byte codes back to quite readable correct Java source code, making it difficult to conceal Java source code by distributing only compiled byte codes. This has significant ramifications for commercial software development in Java.

Phing:19xx:JCA

[PWxx] Chen Ing Phing and E Wen Wei. Java: Creating Animation Homepage. Flag Publishing, ??, Taiwan, 19xx. ISBN ???? ???? pp. LCCN ???? 522

NT$. URL http://www.seed.
net.tw/~flaginfo/.

Patchett:1997:CPC

[PWH97] Craig Patchett, Matthew Wright, and Peter Holfelder. CGI/Perl cookbook. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, October 1997. ISBN 0-471-16896-3. 656 pp. LCCN QA76.73.P22P37 1997. US$44.95.

Pedreschi:1996:LDI

[PZ96] Dino Pedreschi and Carlo Zaniolo, editors. Logic in databases: International Workshop LID ’96, San Miniato, Italy, July 1–2, 1996: proceedings, volume 1154 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 3-540-61814-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 I5857 1996.

Philippsen:1997:JTR

[PZ97] Michael Philippsen and Matthias Zenger. JavaParty — transparent remote objects in Java. Concurrency: Practice and Experience, 9(11):1225–1242, November 1997. CODEN CPEXEI. ISSN 1040-3108. URL http:

                                        //www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=13816;
                                         http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=13816&PLACEBO=IE.pdf;
                                        http://wwwipd.ira.uka.de/~
                                        phlipp/partyd.ps.gz. Special

Issue: Java for computational science and engineering — simulation and modeling II.

Qian:1999:FSJ

[Qia99] Z. Qian. A formal specification of Java[TM] virtual machine instructions for objects, methods and subroutines. Lecture Notes in Computer Science, 1523:271– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Que:1997:SEUa

[Que97] Que Development Group. Special Edition Using Java NC Server. Que Corporation, Indianapolis, IN, USA, 1997. ISBN 0-7897-1279-2. ???? pp. LCCN ???? 101,00 DM; US$50. URL

                                        http://www.mut.ch/profit.
                                        htx?&x_pk=kd84mf7ms9dm.

Reed:19xx:CC

[R^{+}xxa] Jim Reed et al. Café Companion. Symantec, ????, 19xx.

ISBN ???? ???? pp. LCCN ???? US$30.

Rhung:19xx:JJP

[R^{+}xxb] Chen Jien Rhung et al. Java/ JavaScript Programming Design. Ru Lin Books Co., ??, Taiwan, 19xx. ISBN ???? ???? pp. LCCN ???? 300 NT$.

Ram:1997:IWE

[RA97] Prabhu Ram and Robert Abarbanel. Internet watch: Enterprise computing: The Java factor. Computer, 30(6):115–119, June 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Raghunathan:1997:LEJ

[Rag97] Balaji Raghunathan. Letter to the editor: Java for C++. BYTE Magazine, 22(5):19–20, May 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Rajan:1996:LEA

[Raj96] Satish Rajan. Letter to the editor: ActiveX, not Java, applets are unsafe. Computerworld, 30 (43):36, October 21, 1996. CO-DEN CMPWAB. ISSN 0010-4841. ????

Ramel:1997:LSJ

[Ram97] David Ramel. Learn to speak Java for fun and profit. Computerworld, 31(9):46, March 3, 1997. CODEN CMPWAB. ISSN 0010-4841.

Raner:1997:NDI

[Ran97a] M. Raner. New developments for the implementation of Unicode word processing systems in Java. In Unicode Consortium [Uni97b], pages A13–?? ISBN ???? LCCN ???? URL http:

//www.unicode.org/unicode/
iuc11/index.html;   http:
//www.unicode.org/unicode/
uni2book/u2ord.html.

Raner:1997:UJD

[Ran97b] Mirko Raner. Unicode and Java — design and implementation of a Unicode word processing system in Java. In Unicode Consortium [Uni97a], page ?? ISBN ???? LCCN ???? URL http:

//www.unicode.org/unicode/
iuc10/program.html.

Rangaraajan:1999:JQH

[Ran99a] Krishnan Rangaraajan. Java Q and A: How can I test Java classes? Dr. Dobb’s Journal of Software Tools, 24 (7):107–109, July 1999. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/1999/1999_07/
jqa799.txt;   http://www.
ddj.com/ftp/1999/1999_07/
jqa799.zip.

Rangaraajan:1999:JQD

[Ran99b] Krishnan Rangaraajan. Java Q&A: Does Java support design by contract? Dr. Dobb’s Journal of Software Tools, 24 (11):113–114, 116, November 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/ftp/1999/1999_11/
jqa1199.txt.

Rao:1997:UNI

[Rao97] Arun Rao. Using Netscape IFC. Que Corporation, Indianapolis, IN, USA, 1997. ISBN 0-7897-1251-2. xii + 294 pp. LCCN TK5105.883.N48R36 1997. US$35. URL http:

//dev.mcp.com:8001/about/
coinfo/imprints/que/books/
descriptions/12512d.html.

Rao:1998:UJa

[Rao98a] Prithvi Rao. Using Java. ;login: the USENIX Association newsletter, 23(4), June 1998. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

usenix.org/publications/java/
usingjava11.html.

Rao:1998:UJb

[Rao98b] Prithvi Rao. Using Java. ;login: the USENIX Association newsletter, 23(5):??, August 1998. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

usenix.org/publications/java/
usingjava12.html.

Rao:1998:UJc

[Rao98c] Prithvi Rao. Using Java. ;login: the USENIX Association newsletter, 23(6):??, October 1998. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

usenix.org/publications/java/
usingjava13.html.

Rao:1998:UJS

[Rao98d] Prithvi Rao. Using Java: Security outside the JVM. ;login: the USENIX Association newsletter, 23(7):??, December 1998. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

                                        usenix.org/publications/java/
                                        usingjava14.html.

Rao:1999:UJ

[Rao99a] Prithvi Rao. Using Java. ;login: the USENIX Association newsletter, 24(6):??, December 1999. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

                                        usenix.org/publications/java/
                                        usingjava19.html.

Rao:1999:UJA

[Rao99b]  Prithvi Rao.  Using Java:

Applets and animation. ;login: the USENIX Association newsletter, 24(2):??, April 1999. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

                                        usenix.org/publications/java/
                                        usingjava16.html.

Rao:1999:UJI

[Rao99c] Prithvi Rao. Using Java: Input and output streams. ;login: the USENIX Association newsletter, 24(4):??, August 1999. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

                                        usenix.org/publications/java/
                                        usingjava18.html.

Rao:1999:UJR

[Rao99d] Prithvi Rao. Using Java: Remote Method Invocation. ;login: the USENIX Association newsletter, 24(3):??, June 1999. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

usenix.org/publications/java/
usingjava17.html.

Rao:1999:UJJ

[Rao99e] Prithvi Rao. Using Java: The Java Native Interface. ;login: the USENIX Association newsletter, 24(1):??, February 1999. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

usenix.org/publications/java/
usingjava15.html.

Solutions:19xx:IPJc

[Rapxx] Rapid Systems Solutions. An Introduction to Programming Java Applets Using J++. MindQ Publishing, Inc., 450 Springpark Place Ste. 1200, Herndon, VA 20170, USA, Tel: +1 703 708-9380, 19xx. ISBN 1-57559-009-3. LCCN ???? US$50.

Razavi:1998:JJC

[Raz98] Behfar Razavi. JavaStations and Java Computing. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, May 1998. ISBN 0-13-887423-9. LCCN ????

US$40.  URL http://www.
prenhall.com/ptrbooks/ptr_
0138874239.html.

Rosch:1996:RTM

[RB96] P. Rosch and M. Baentsch. Reviewing two multimedia presentation (quasi-) standards. In IEEE [IEE96j], page ?? ISBN 0-8186-7511-X. LCCN QA76.575 .M84 1996.

Raje:1998:CCE

[RBF98] Rajeev R. Raje, Michael Boyles, and Shiaofen Fang. CEV: collaborative environment for visualization using Java RMI. Concurrency: Practice and Experience, 10 (11–13):1079–1085, September 1998. CODEN CPEXEI. ISSN 1040-3108. URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=10050406;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=10050406&PLACEBO=IE.pdf.

Special Issue: Java for Highperformance Network Computing.

Rodger:1997:CTM

[RBL^{+}97] Susan H. Rodger, Anna O. Bilska, Kenneth H. Leider, Magdalena Procopiuc, Octavian Procopiuc, Jason R. Salemme, and Edwin Tsang. A collection of tools for making automata theory and formal languages come alive. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(1):15–19, March 1997. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rouse:1997:TSO

[RC97] Forest Rouse and Wayne Christopher. A typing system for an optimizing multiplebackend Tcl compiler. In USENIX [USE97a], pages 37– ?? ISBN 1-880446-87-1. LCCN QA76.73.T44 T44 1997.

URL ftp.dnai.com/users/i/
icemcfd/tcl;   http://www.
icemcfd.com/tcl/ice.html.

The backends include C and the Java Virtual Machine. Demonstration licenses for the Tcl Lint system are also available.

Ryner:1998:JKF

[RC98] Stephen Wood Ryner, Jr. and Nobuko Chikamatsu. Java Kanji Flashcard 500: Kanji, Java and the World Wide Web. In Unicode Consortium [Uni98a], page ?? ISBN ????

LCCN ???? URL http://www.
unicode.org/unicode/iuc12/
program.html.

Rubin:1998:JIS

[RCB98] B. S. Rubin, A. R. Christ, and K. A. Bohrer. Java and the IBM San Francisco project. IBM Systems Journal, 37(3): 365–??, ???? 1998. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

URL http://www.almaden.
ibm.com/journal/sj/373/rubin.
html.

Roehl:1997:LNV

[RCRBR97] Bernie Roehl, Justin Couch, Cindy Reed-Ballreich, and Tim Rohaly. Late Night VRML 2.0 With Java. Ziff-Davis Press, Emeryville, CA, USA, 1997. ISBN 1-56276-504-3. xxxvi + 710 pp. LCCN ???? US$45. URL http://

merchant.superlibrary.com:
8000/catalog/hg/PRODUCT/PAGE/
15627/bud/1562765043.html.

Roback:1999:CRF

[RD99] Edward Roback and Morris Dworkin. Conference report: First Advanced Encryption Standard (AES) Candidate Conference, Ventura, CA, August 20–22, 1998. Journal of Research of the National Institute of Standards and Technology, 104(1):97– 105, January/February 1999. CODEN JRITEF. ISSN 1044-677X (print), 2165-7254 (electronic). URL http://

                                        csrc.nist.gov/encryption/
                                        aes/round1/conf1/j41ce-rob.
                                        pdf.

Reynolds:1996:UJ

[RDD^{+}96] Mark C. Reynolds, Ray Daly, Rick Darnell, Bill Dortch, Mo Everett, and Gordon Mc-Comb. Using JavaScript. Que Corporation, Indianapolis, IN, USA, special edition, May 1, 1996. ISBN 0-7897-0789-6, 0-7897-1079-X. xxvi + 831 pp. LCCN QA76.73.J39R49 1996. US$49.99. URL http://www.

                                        mcp.com/que.

Redlich:1996:CPE

[Red96] Jens-Peter Redlich. CORBA 2.0: Praktische Einführung für C++ und Java. Addison-Wes-ley, Reading, MA, USA, 1996. ISBN 3-8273-1060-1 (??invalid ISBN??). xiv + 314 pp. LCCN ???? 60 DM. URL http://

                                        www.addison-wesley.de/.

Redfern:1997:ANA

[Red97a]  Heather Redfern.  AUn,

Netscape announce developer release of JFC. SunServer, 11 (9):4, September 1997. ISSN 1091-4986.

Redfern:1997:GJP

[Red97b] Heather Redfern. GO-Joe provides Java users access to UNIX X apps. SunServer, 11(9):6, September 1997. ISSN 1091-4986.

Redfern:1997:RJT

[Red97c] Heather Redfern. Recreating JavaBeans through visual assembly: Sun unveils Java Studio 1.0. SunServer, 11(12):1, 16, December 1997. ISSN 1091-4986.

Redfern:1998:JAM

[Red98] Heather Redfern. Java assists in monitoring of jet development. SunServer, 12(1):15, January 1998. ISSN 1091-4986.

Reddy:1999:FJP

[Red99] Achut Reddy. Fast Java: Performance Tuning Guide. Sun-Soft Press Java series. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1999. ISBN ???? ???? pp. LCCN ????

URL http://www.phptr.com/
bookseri/java.html.

Redlich:19xx:CPE

[Redxx] Jens-Peter Redlich. CORBA 2.0: Praktische Einführung für C++ und Java. Addison-Wes-ley, Reading, MA, USA, 19xx.

ISBN 3-8273-1060-1 (??invalid ISBN??). LCCN ???? 60 DM.

Reese:1997:DPJ

[Ree97] George Reese. Database Programming With JDBC and Java. Java series. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, July 1, 1997. ISBN 1-56592-270-0. 300 (or 224??) pp. LCCN QA76.625 .R44 1997 Sci-Eng. US$29.95.

URL http://www.ora.com/
catalog/javadata/. Java 1.1.

Reese:1998:JJ

[Ree98] George Reese. JDBC et Java. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, February 1998. ISBN 2-84177-042-7. LCCN ???? 220 FF. URL http:

//www.editions-oreilly.fr/
international/france/prog/
java-jdbc.html.

Refling:1999:HUT

[Ref99] John Refling. Hands-on universe: teaching astronomy with Java-based image processing tools. In ACM [ACM99d], page 58. ISBN ???? LCCN ????

URL http://www.acm.org/
pubs/citations/proceedings/
graph/311625/p58-refling/.

Rego:1996:AJA

[Reg96] F. Alfredo Rego. ADBC: Javasupported access via intrinsics in TurboImage data. HP Chronicle, 13(6):18, May 1996. ISSN 0892-2829.

Reisman:1996:WKE

[Rei96] Sorel Reisman. Web Sight:

Keeping an eye on the Web; what is Web Sight Web site? IEEE MultiMedia, 3(1):5, Spring 1996. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://dlib.computer.org/
mu/books/mu1996/pdf/u1005.
pdf.

Reichert:1997:SEUa

[Rei97a] Jake Reichert. Special Edition Using Java Network Applications. Que Corporation, Indianapolis, IN, USA, October 1997. ISBN 0-7897-1465-5. ???? pp. LCCN ???? 89,00 DM, US$45.00. URL http:/

/www.mut.ch/profit.htx?&x_
pk=kd84mf7ms9dm.  Includes

CD-ROM.

Reichert:1997:SEUb

[Rei97b] Jake Reichert. Special Edition Using Java Networks. Ziff-Davis Press, Emeryville, CA, USA, special edition, October 1997. ISBN 1-56276-569-8 (softcover). ???? pp. LCCN ???? US$40.49. Includes CD-ROM.

Reinke:1999:THJ

[Rei99] C. Reinke. Towards a Haskell/ Java connection. Lecture Notes in Computer Science, 1595:200– 215, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Remy:1997:DPJ

[Rem97] Martin Remy. Design patterns, Java and Web development. Dr. Dobb’s Journal of Software Tools, 22(6):28, 30, 32, 34, 98–100, June 1997. CODEN DDJOEB. ISSN 1044-789X.

Renshaw:1998:BNA

[Ren98] David S. Renshaw. Building network apps — making components portable with JavaBeans — a key to integrated solutions across multiple platforms. BYTE Magazine, 23(2):81–84, 86–88, February 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Ren:19xx:RHJ

[Renxx] Huang Jung Ren. Relax on hand with Java. Acer TWP Corp., ????, 19xx. ISBN 957-23-0435-6. ???? pp. LCCN ???? 350 NT$.

Reynolds:1996:IKE

[Rey96] Franklin D. Reynolds. Internet kiosk: Evolving an operating system for the Web. Computer, 29(9):90–92, September 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Reynolds:1997:OOP

[Rey97] Mark C. Reynolds. Object Oriented Programming In Java. McGraw-Hill, New York, NY, USA, March 1997. ISBN 0-07-913250-2. 336 pp. LCCN ???? US$39.95.

Rodriguez:1995:SSL

[RFM95] Arturo A. Rodriguez, Martin Fisher, and Brian Markey. Standards: Scripting languages emerge in standards bodies. IEEE MultiMedia, 2(4):88– 92, Winter 1995. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

URL http://dlib.computer.
org/mu/books/mu1995/pdf/u40088.

pdf. Discusses the proposed Standard Multimedia/ Hypermedia Scripting Language (SMSL), a superset of Java.

Rice:1996:UWI

[RFPG96] James Rice, Adam Farquhar, Philippe Piernot, and Thomas Gruber. Using the Web instead of a window system. In Tauber [Tau96], pages 103–110. ISBN 0-201-94687-4 (paperback), 0-89791-777-4 (ACM). LCCN ????

Rosenblatt:1996:SJI

[RGT96] Max Rosenblatt, Joshua Greenbaum, and Anne Thomas. Sorting out the Java/intranet vs. SAP promise. Computerworld, 30(15):40, April 8, 1996. CO-DEN CMPWAB. ISSN 0010-4841.

Reynolds:1997:UJ

[RH97a] Mark Reynolds and Jerry Honeycutt. Using JScript. Que Corporation, Indianapolis, IN, USA, special edition, 1997. ISBN 0-7897-1079-X. xiv + 586 pp. LCCN QA76.73.J39 R49 1997.

Roberts:1997:JCS

[RH97b] Simon Roberts and Philip Heller. Java 1.1 Certification Study Guide. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, September 1997. ISBN 0-7821-2069-5 (softcover). xxvi + 720 pp. LCCN QA76.3 .R63 1997. US$35.99. URL http://www.

sybex.com/cgi-bin/bookpg.
pl?2069back.html. Includes

CD-ROM.

Roberts:1998:JCS

[RH98] Simon Roberts and Philip Heller. Java 1.1 Certification Study Guide. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, August 1998. ISBN 0-7821-2330-9. LCCN ???? US$55.

Roberts:1999:CJC

[RHE99] Simon Roberts, Philip Heller, and Michael Ernest. The Complete Java 2 Certification Study Guide. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 1999. ISBN 0-7821-2463-1. xxvi + 885 pp. LCCN QA76.3 .R627 1999. US$45.

Robinson:1999:JBT

[RHND99] Matthew B. Robinson, Jason A. Hamshar, Jorge E. Novillo, and Andrew T. Duchowski. A Javabased tool for reasoning about models of computation through simulating finite automata and Turing machines. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1):105–109, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ricciuti:1995:NDM

[Ric95a] Mike Ricciuti. News: Database makers head for the Web. InfoWorld, page ??, December 10, 1995. CODEN INWODU. ISSN 0199-6649. URL http://www.

infoworld.com/.

Ricciuti:1995:NPD

[Ric95b] Mike Ricciuti. News: Paradox, dBase future grim. InfoWorld, page ??, December 03, 1995. CODEN INWODU. ISSN 0199-6649. URL http://www.

infoworld.com/.

Richman:1995:NOO

[Ric95c] Dan Richman. News: Oracle offers free Web browser and software. Computerworld, page ??, October 27, 1995. CODEN CMPWAB. ISSN 0010-4841. URL http://www.

computerworld.com/.

Richards:1996:OLO

[Ric96a] J. Richards. Object lessons [object oriented databases]. Conspectus, ????(????):28–29, July 1996. CODEN CONSF8. ISSN 1351-0908.

Richardson:1996:JGM

[Ric96b] Robert Richardson. Java gains momentum. LAN: the network solutions magazine, 4(11):

69–??, May 1, 1996. CODEN LANNER. ISSN 1038-359X (or 1069-5621??).

Richardson:1996:JPS

[Ric96c] Robert Richardson. Java picks up steam. LAN: the network solutions magazine, 11(4):79–??, ???? 1996. CODEN LANNER. ISSN 1069-5621.

Richman:1996:TTJ

[Ric96d] Dan Richman. Tools on tap for Java standard. Computerworld, 30(20):24, May 13, 1996. CO-DEN CMPWAB. ISSN 0010-4841.

Riesebeck:1996:UJ

[Rie96] Jim Riesebeck, editor. Using JBuilder. Que Corporation, Indianapolis, IN, USA, 1996. ISBN 0-7897-0880-9. ???? pp. LCCN ???? US$30.

Riehle:1997:ADW

[Rie97a] R. Riehle. Ada does Windows. Journal of Object Oriented Programming, 10(1):74–7, March– April 1997. CODEN JOOPEC. ISSN 0896-8438.

Rieken:1997:IJB

[Rie97b]  Bill Rieken.  An introduc-

tion to Java Beans. Tutorial session at [USE97b]., 1997.

                                        URL http://dxsting.cern.
                                        ch/sting/COOTS97.html.

Rinehart:1997:JDD

[Rin97a]  Martin L. Rinehart.  Java

Database Development. Osborne/McGraw-Hill, Berkeley, CA, USA, 1997. ISBN 0-07-882356-0 (softcover). xxiii + 917 pp. LCCN QA76.73.J38R558 1998. US$39.99. URL http:

//www.osborne.com/program/
javadbdev.htm. Includes CD-

ROM.

Rinehart:1997:JPV

[Rin97b] Martin L. Rinehart. Java Programming With Visual J++. MIS Press and M&T Books, P. O. Box 5277, Portland, OR 97208-5277, USA, Tel:

(503) 282-5215 and M&T Publishing, Inc., 501 Galveston Drive, Redwood City, CA 94063, USA, 1997. ISBN 1-55851-506-2. xiii + 462 pp. LCCN QA76.73.J38R56 1996. US$39.95. URL http://www.

mandt.com/javaprogvisualj+
+.htm.

Ritchey:1995:J

[Rit95a] Tim Ritchey. Java! New Riders Publishing, Carmel, IN, USA, 1995. ISBN 1-56205-533-X. x + 366 pp. LCCN TK5105.888.R58

1995. URL http://java.sun.

com/. Includes CD-ROM.

Ritchey:1995:PJB

[Rit95b] Tim Ritchey. Programming with Java. Beta 2.0 Java. New Riders Publishing, Carmel, IN, USA, November 1, 1995. ISBN 1-56205-533-X. xiii + 389 pp. LCCN QA 76.73 J3 R482 1995. US$35.00. URL http:

//java.sun.com/;   http://
www.mcp.com/113779062800198/
cgi-bin/bag?isbn=1-56205-
533-X&last=/bookstore.

Ritchey:1996:BJA

[Rit96a] Tim Ritchey. Building Java Applets for Netscape 2.0. New Riders Publishing, Carmel, IN, USA, April 1996. ISBN 1-56205-585-2. xiii + 469 pp. LCCN QA76.73.J38R58 1996. US$35.00. URL http://www.

                                        mcp.com/175067741767563/cgi-
                                        bin/bag?isbn=1-56205-585-
                                        2&last=/bookstore. Includes

CD-ROM.

Ritchey:1996:JP

[Rit96b] Tim Ritchey. Java Programming. DrMaster Consulting Co, Ltd, Taipei, Taiwan, 1996. ISBN 957-9625-09-3. viii + 372 pp. LCCN ???? 450 NT$. URL ????. Chinese translation by Cheng Chuin Wen and Liao Shiu Yee. Includes CD-ROM.

Ritchey:1996:PJ

[Rit96c] Tim Ritchey. Programming with Java! New Riders Publishing, Carmel, IN, USA, 1996. ISBN 1-56205-533-X. xiii + 389 pp. LCCN QA 76.73 J3 R482 1995. US$35.00, 70,00 DM. URL http://

                                        www.mut.ch/profit.htx?&x_
                                        pk=kd84mf7ms9dm. CD-ROM

includes: Java developers kit for Windows 95, Windows NT, and Sun Solaries 2.x, Java applet tools, example applets from the text.

Ritchey:1997:PCJb

[Rit97a] Tim Ritchey. Programando com Java! Beta 2.0. Editore Campus Ltda., Rua Sete de Setembro, 111/16, Rio de Janeiro - RJ - 20050-002, Brazil, 1997. ISBN ???? 460 pp. LCCN ???? R$ 69,90. URL http:

//www.campus.com.br/about-
us.htm;   http://www.campus.
com.br/catalogo/livro.cfm?
cod_livro=20060.

Ritchey:1997:PCJa

[Rit97b] Tim Ritchey. Programando con Java. Prentice Hall HispanoAmericana, Mexico, DF, Mexico / Nueva York, NY, USA, 1997. ISBN 84-89660-60-3. 290 pp. LCCN ????

URL http://prentice.com.
mx/libros/progjava.html. Span-

ish translation of [Rit96c].

Ritchey:1997:PCJc

[Rit97c] Tim Ritchey. Programmare con JavaScript. Tecniche Nuove, ??, Italy, 1997. ISBN ???? ???? pp. LCCN ???? 55000 Lire. URL

http://www.tecnet.it/cgi-
shl/dbml.exe?Action=Query&
Template=/LibriQueryDueLibro.
dbm&ISBN=03456.

Ritchie:1997:SPJ

[Rit97d] Stuart Ritchie. Systems programming in Java: Reducing complexity and shortening debugging time by using Java at the system level. IEEE Micro, 17(3):30–35, May/ June 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//pascal.computer.org/mi/
books/mi1997/pdf/m3030.pdf.

Ritchey:19xx:JPb

[Ritxxa] Tim Ritchey. Java Programming. CA Books c/o DRT Int’l, ????, 19xx. ISBN 89-86604-08-6 (??invalid ISBN??). ???? pp. LCCN ???? 15,000 Won. Korean translation by Hee-Chun Kang and Bok-Sik Kim.

Ritchey:19xx:JJ

[Ritxxb] Tim Ritchey. Javascript & Java. Impress Corporation, ??, Japan, 19xx. ISBN 4-8443-4918-X. ???? pp. LCCN ???? 2,980

                                        yen. URL http://www.ips.co.
                                        jp/book/4918/4918.htm.

Ritchey:19xx:PCJb

[Ritxxc] Tim Ritchey. Programando com Java! Beta 2.0. Editore Campus Ltda., Rua Sete de Setembro, 111/16, Rio de Janeiro - RJ - 20050-002, Brazil, 19xx. ISBN ???? LCCN ???? R$60.

Ritchey:19xx:PCJa

[Ritxxd] Tim Ritchey. Programando con Java. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 19xx. ISBN 84-89660-60-3 (??invalid ISBN??). ???? pp. LCCN ???? 4495 Ptas.

Ritchey:19xx:PMJ

[Ritxxe] Tim Ritchey. Programmeren met Java! Academic Service, Postbus 81, 2870 AB Schoonhoven, The Netherlands, 19xx. ISBN 90-395-0390-7. 377 pp. LCCN ???? fl. 69; Bfr 1388. URL http:/

                                        /acaserv.nl/bestel/5_0390.
                                        htm;   http://www.acaserv.

nl; http://www.acaserv.nl/
5_0390.html.

Ritchey:19xx:PJa

[Ritxxf] Tim Ritchey. Programming with Java. New Riders, ????, second edition, 19xx. ISBN ???? ???? pp. LCCN ???? US$35.

Ritchey:19xx:PJb

[Ritxxg] Tim Ritchey. Programming with Java. New Riders Publishing, Carmel, IN, USA, second edition, 19xx. LCCN ???? US$35.

Ritchey:19xx:UJ

[Ritxxh] Tim Ritchey. Usare Java. Jackson Libri, via XXV Aprile, 39 - 20091 Bresso (MI), Italy, 19xx. ISBN ???? 352 pp. LCCN ???? 59000 Lire. URL http://

www.jacksonlibri.it/Libri/
libri-aut.htm.

Ritchey:19xx:VJJ

[Ritxxi] Tim Ritchey. Visual J++ & Java. Impress Corporation, ??, Japan, 19xx. ISBN 4-8443-4809-4. ???? pp. LCCN ???? 2,800

yen. URL http://www.ips.co.
jp/book/4809/4809.htm.

Radermacher:1997:ECC

[RK97a] Todd Radermacher and Peter Kocks. Executable content: Controlling the potential revolution. SysAdmin, 6(11):61–63, 65–66, November 1997. CO-DEN SYADE7. ISSN 1061-2688.

Rahkila:1997:IDT

[RK97b] Martti Rahkila and Matti Karjalainen. Interactive DSP tutorial on the Web. In ICASSP, IEEE International Conference on Acoustics, Speech and Signal Processing — Proceedings, volume 3, pages 2253–2256. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1997. CODEN IPRODJ. ISBN ???? ISSN 0736-7791. LCCN ????

Reus:1998:VCO

[RKCW98] B. Reus, A. Knapp, P. Cenciarelli, and M. Wirsing. Verifying a compiler optimization for multi-threaded Java. Lecture Notes in Computer Science, 1376:402–??, 1998. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rinard:1998:DIE

[RL98] Martin C. Rinard and Monica S. Lam. The design, implementation, and evaluation of Jade. ACM Transactions on Programming Languages and Systems, 20(3):483–??, May 1998. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Romer:1996:SPI

[RLV^{+}96] Theodore H. Romer, Dennis Lee, Geoffrey M. Voelker, Alec Wolman, Wayne A. Wong, Jean-Loup Baer, Brian N. Bershad, and Henry M. Levy. The structure and performance of interpreters. ACM SIGPLAN Notices, 31(9):150–159, September 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Compares the MIPSI, Java, Perl, and Tcl interpreters on a DEC Alpha platform.

Richardson:1997:HJ

[RM97] Eric C. Richardson and Michael Morrison. Hands On Java. Prima Publishing, Rocklin, CA, USA, 1997. ISBN 0-7615-1013-3. 500 pp. LCCN ???? US$40.00.

Robillard:1999:AEF

[RM99] Martin P. Robillard and Gail C.

Murphy. Analyzing exception flow in Java programs. ACM SIGSOFT Software Engineering Notes, 24(6):322–337, November 1999. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Raje:1998:DIC

[RMB^{+}98] R. R. Raje, S. Mukhopadhyay, M. Boyles, N. Patel, and A. Papiez. On designing and implementing a collaborative system using the distributed-object model of Java-RMI. Parallel and Distributed Computing Practices, 1(4):??, ???? 1998. CODEN ???? ISSN 1097-2803. URL http://www.

cs.okstate.edu/~pdcp/vols/
vol01/vol01no4abs.html#raje.

Richardson:1997:JAC

[RMS97] Ronny Richardson, Bruce Murray, and Marc Singer. Java Applets and Channels Without Programming. Manning Publications, Greenwich, CT, USA, August 1997. ISBN 1-884777-39-2 (softcover). ???? pp. LCCN QA76.73.J38J363 1997. US$34.15. URL http:/

/www.manning.com/Shoffner/

index.html. Includes CD-ROM.

Romano:1998:ADD

[RNBV98] Nicholas C. Romano, Jr., Jay F. Nunamaker, Jr., Robert O. Briggs, and Douglas R. Vogel. Architecture, design, and development of an HTML/ JavaScript Web-based Group Support System. Journal of the American Society for Information Science, 49(7):649–667, May 15, 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Roberts:1996:WWJ

[Rob96] Ken Roberts. What’s wrong with Java? (discussion of the February 1996 article, the NC phenomena: scenes from your living room). Computer, 29(6): 8–9, June 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Robinson:1998:JPI

[Rob98a] A. W. Robinson. JavaFit: a platform independent program for interactive nonlinear least-squares fitting using the Levenberg–Marquardt method. Computer Physics Communications, 112(2–3): 183–190, August 1998. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0010465598000526.

Robinson:1998:JP

[Rob98b] G. S. Robinson. Java and the PAS process. IEEE Micro, 18(1):4–5, January/February 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Roberts:1999:EJJ

[Rob99] Spencer Roberts. Essential JTAPI: Java telephony API: designing telecom projects with Java. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1999. ISBN 0-13-080360-X. xxi + 565 pp. LCCN TK6421.R63 1999. US$89.99. URL http:

//www.phptr.com/ptrbooks/
ptr_013080360X.html.

Rodley:1995:PHA

[Rod95] John Rodley. Programming HotJava applets. Dr. Dobb’s Journal of Software Tools, 20 (??):7–??, November/December 1995. CODEN DDJOEB. ISSN 1044-789X.

Rodley:1996:WAP

[Rod96a] John Rodley. Web Applets programmieren mit JavaScript. tewi, München, Germany, 1996. ISBN 3-89362-472-4. 389 pp. LCCN ???? 79 DM. URL http:

//www.channel1.com/users/
ajrodley/;   http://www.tewi.
de/;   http://www.tewi.de/
scripts/search/details_ISBN.
idc?ISBN=3-89362-472-4.

Rodley:1996:WJA

[Rod96b] John Rodley. Writing Java Applets: Master the Art of Creating Distributed Internet Programs. Coriolis Group Books, Scottsdale, AZ, USA, April 1, 1996. ISBN 1-883577-78-0. xvi + 430 pp. LCCN QA76.73.J38R63 1996. US$39.99. URL http://165.

247.200.139/cgi-bin/books.
cgi?page=books&subject=wjavap&
graphics=0;   http://www.
coriolis.com/cgi-win/KBSearch.

exe. Includes CD-ROM.

Rodley:1997:WJA

[Rod97] John Rodley. Writing Java Applets. Softbank Publishing, ??, Japan, 1997. ISBN 4-7973-0110-4. ???? pp. LCCN ???? 3900 yen. URL http:

//www.channel1.com/users/
ajrodley/;   http://www.softbank.
co.jp/softbank/publishing/
;   http://www.softbank.co.
jp/softbank/publishing/books/
kikan/kikan10/008.htm.

Rodrigues:1998:APJ

[Rod98] Lawrence H. Rodrigues. Awesome Power of JavaBeans. Manning Publications, Greenwich, CT, USA, May 1998. ISBN 1-884777-56-2. xxiii + 545 pp. LCCN QA76.73.J38R64 1998. US$45. URL http://

www.manning.com/Rodrigues/
index.html.

Rodley:19xx:OVJ

[Rodxxa] John Rodley. Het ontwikkelen van Java Applets. Addison-Wesley Nederland, ????, 19xx. ISBN 90-6789-734-5. ???? pp. LCCN ???? fl. 80. URL http:

//www.channel1.com/users/

ajrodley/. Includes CD-ROM.

Rodley:19xx:PAJ

[Rodxxb]  John Rodley.  Programmer

des Applets Java Megapoche. Sybex France, Paris, France, 19xx. ISBN 2-7361-2142-2. ???? pp. LCCN ????

98 FF. URL http://www-
db.sybex.fr/cgi-bin/Sybex/
Webdriver?MIval=Fiche_Book.
html&Reference=2142.

Rodley:19xx:WJAc

[Rodxxc] John Rodley. Writing Java Applets. Image Art Publishing, ????, 19xx. ISBN ???? ???? pp. LCCN ???? Chinese translation by Chiang Yhung Shiang.

Rohnert:1996:JLL

[Roh96] H. Rohnert. Java: lessons learnt from C++. In Anonymous [Ano96-142], pages 348–??, 283. ISBN ???? LCCN ????

Roman:1999:MEJ

[Rom99] Ed Roman. Mastering Enterprise JavaBeans and the Java 2 Platform, Enterprise Edition. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 1999. ISBN 0-471-33229-1. xxviii + 722 pp. LCCN QA76.73.J38 R66 1999. US$49.99. URL http://www.

wiley.com/compbooks/. In-

cludes CD-ROM.

Rook:1997:LCC

[Roo97] David Rook. A language collector comments on: Java, Perl & Python, October 1997. URL

                                        http://www.chips.navy.mil/
                                        chips/archives/97_oct/file12.
                                        htm.

Rooker:1998:BRJ

[Roo98] Terry Rooker. Book review: Java Network Security. ;login: the USENIX Association newsletter, 23(6):??, October 1998. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

                                        usenix.org/publications/login/
                                        1998-10/java_sec.html.

Ross:1996:EFJ

[Ros96] Rocky Ross. Education Forum: Java… Hot Java!: What is that brewing on the Web? ACM SIGACT News, 27(1): 24–27, March 1996. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Rosenberger:1998:TYC

[Ros98] Jeremy Rosenberger. Teach Yourself CORBA in 21 Days. Howard W. Sams, Indianapolis, IN 46268, USA, January 1998. ISBN 0-672-31208-5. xiv + 452 pp. LCCN QA76.9.O35R67 1998. US$30. URL http://

                                        merchant.superlibrary.com:
                                        8000/catalog/hg/PRODUCT/PAGE/
                                        06723/bud/0672312085.html.

Rousselle:1998:PTD

[Rou98] Philip Rousselle. Programmer’s toolchest: Dynamic distributed systems in Java. Dr. Dobb’s Journal of Software Tools, 23 (4):88, 90–92, 102–104, April 1998. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ftp/1998/1998_
04/sojournr.txt;   http://
www.ddj.com/ftp/1998/1998_
04/sojournr.zip.

Rowe:1997:IDS

[Row97] Glenn Rowe. An Introduction to Data Structures and Algorithms with Java. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, November 1997. ISBN 0-13-857749-8. ???? pp. LCCN QA76.73.J38R69 1997.

US$44.  URL http://www.
prenhall.com/ptrbooks/esm_
0138577498.html.

Roberts:1998:DJG

[RP98] Eric Roberts and Antoine Picard. Designing a Java graphics library for CS 1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):213–218, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Robbins:1999:EEU

[RR99] Steven Robbins and Kay A.

Robbins. Empirical exploration in undergraduate operating systems. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1): 311–315, March 1999. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rhung:19xx:JPD

[RRxx] Chen Jien Rhung and Wang Tze Rhung. Java Programing and Design. Ru Lin Books Co., ????, 19xx. ISBN 957-652-848-8. ???? pp. LCCN ???? 270 NT$.

Rice:1997:AJP

[RS97] Jeffrey C. Rice and Irving Salisbury, III. Advanced Java 1.1 Programming. McGraw-Hill, New York, NY, USA, August 1997. ISBN 0-07-913089-5 (set), 0-07-052426-2 (book), 0-07-852998-0 (CD-ROM). xv + 476 pp. LCCN QA76.73.J38R53 1997. US$39.95. URL http:

                                        //mcgraw-hill.inforonics.
                                        com/cgi/getarec?mgh30379%
                                        comp;   http://www.mcgraw-
                                        hill.com/.

Russell:1997:PCP

[RSH97] Stuart Russell, Lewis Stiller, and Othar Hansson. PN-PACK: Computing with probabilities in Java. Concurrency: Practice and Experience, 9(11):1333–1339, November 1997. CODEN CPEXEI. ISSN 1040-3108. URL http:

                                        //www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=13843;
                                         http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=13843&PLACEBO=IE.pdf. Spe-

cial Issue: Java for computational science and engineering — simulation and modeling II.

Ruskin:1998:NIT

[Rus98] K. J. Ruskin. New Internet technology: HTML tools, Java, teleconferencing, and streaming multimedia. In Ikeda et al. [IDK98], pages 229–234. ISBN 0-444-50008-1. LCCN WO 200 I61-2 1998; RD80.95 .I68 1998.

Ready:1996:PPJ

[RVM96] Kevin Ready, Paul Vachier, and Benoit Marsot. Plug-N-Play JavaScript. New Riders Publishing, Carmel, IN, USA, October 1, 1996. ISBN 1-56205-674-3. xii + 340 pp. LCCN QA76.73.J39 R43 1996. US$34.99.

Ready:1997:PCJ

[RVM97] Kevin Ready, Paul Vachier, and Benoit Marsot. Programmare con JavaScript. Apogeo srl, Viale Papiniano 38, 20123 Milano, Italy, 1997. ISBN 88-7303-305-9. 352 pp. LCCN ???? 42000 Lire. URL http://www.

urra.it/305.html;   http:
//www.urra.it/apocat.html.

Italian translation of [RVM96]. Includes CD-ROM.

Reynolds:19xx:SEU

[RWxx] Mark Reynolds and Andrew Wooldridge. Special Edition Using JavaScript. Simon and Schuster (Asia), Singapore, 19xx. ISBN 957-9625-56-5. ???? pp. LCCN ???? 460 NT$. URL http:

//www.sbpa.org.sg/members/

4764688.htm. Chinese translation by Yhou Chi Fhu of [RDD^{+}96]. Includes CD-ROM.

Raje:1997:ARM

[RWB97] Rajeev R. Raje, Joseph I. Williams, and Michael Boyles. Asynchronous Remote Method Invocation (ARMI) mechanism for Java. Concurrency: Practice and Experience, 9(11):1207–1211, November 1997. CODEN CPEXEI. ISSN 1040-3108. URL http:

                                        //www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=13835;
                                         http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=13835&PLACEBO=IE.pdf. Spe-

cial Issue: Java for computational science and engineering — simulation and modeling II.

Riggs:1996:PSJa

[RWWB96a] Roger Riggs, Jim Waldo, Ann Wollrath, and Krishna Bharat. Pickling state in the Java system. In Association [Ass96a], pages 261–??, 241–250. ISBN ???? LCCN ????

Riggs:1996:PSJb

[RWWB96b] Roger Riggs, Jim Waldo, Ann Wollrath, and Krishna Bharat. Pickling state in the Java system. Computing Systems, 9(4): 291–312, Fall 1996. CODEN CMSYE2. ISSN 0895-6340.

Riggs:COMPSYS-9-4-291

[RWWB96c] Roger Riggs, Jim Waldo, Ann Wollrath, and Krishna Bharat. Pickling state in the Java system. Computing Systems, 9(4): 291–312, Fall 1996. CODEN CMSYE2. ISSN 0895-6340.

Rymer:1995:CCI

[Rym95] John Rymer. Column: The challenge of incompatible object systems. Network World, page ??, May 24, 1995. ISSN 0887-7661 (print), 1944-7655 (electronic). URL http://www.

iquest.net/index.html.

Savola:1996:UHS

[S^{+}96a]  Tom Savola et al.  Us-

ing HTML Second Edition: Create Dynamic and Interactive Pages Web with Java, JavaScript, CGI, and new HTML 3 Extensions. Que Corporation, Indianapolis, IN, USA, 1996. ISBN 0-7897-0758-6 (softcover). xxxiii + 921 pp. LCCN QA76.76.H94U85 1996¡. US$44.99. Includes CD-ROM.

Stokes:1996:GKW

[S^{+}96b] Neil Stokes et al. Getting to Know OS/2 Warp 4: Speech Recognition, Built-In LAN Support, Java Runtime Support, Peer to Peer Networking. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, November 1996. ISBN 0-13-842147-1 (softcover). xx + 621 pp. LCCN ???? US$35.95.

Shin:1997:CJ

[S^{+}97] Jeongho Shin et al. Catch Java!

Miraejungbosa, ??, Korea, 1997. ISBN 89-7836-024-6 (??Invalid checksum??). ???? pp. LCCN ???? 16,000 won. URL ????.

Schneider:1997:UEJ

[SA97] Jeff Schneider and Rajeev Arora. Using Enterprise Java, Special Edition. Que Corporation, Indianapolis, IN, USA, 1997. ISBN 0-7897-0887-6 (softcover). 496 pp. LCCN QA76.73.J38S36 1997. US$53.99. URL http://

merchant.superlibrary.com:
8000/catalog/hg/PRODUCT/PAGE/
07897/bud/0789708876.html.

Solorzano:1998:PPJ

[SA98a] Jose H. Solorzano and Suad Alagić. Parametric polymorphism for Java: a reflective solution. ACM SIGPLAN Notices, 33(10):216–225, October 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Stata:1998:TSJ

[SA98b] Raymie Stata and Martín Abadi. A type system for Java bytecode subroutines. In ACM [ACM98b], pages 149–160. ISBN 0-89791-979-3. LCCN QA76.7 .A15

1998.  URL http://www.
acm.org:80/pubs/citations/
proceedings/plan/268946/p149-

stata/. ACM order number: 549981.

Stata:1999:TSJ

[SA99] Raymie Stata and Martin Abadi. A type system for Java bytecode subroutines. ACM Transactions on Programming Languages and Systems, 21(1): 90–137, January 1999. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL http://www.acm.org/
pubs/citations/journals/toplas/
1999-21-1/p90-stata/.

Schneider:19xx:UEJ

[SAxx] Jeff Schneider and Rajeev Arora. Using Enterprise Java. La Nuova Italia, Firenze, Italy, 19xx. ISBN 80-7897-088-7 (??Invalid checksum??). ???? pp. LCCN ???? 142000 L.

Saadi:1996:CAS

[Saa96]  Edgar Saadi.  Career ad-

visor: Should I go with Java? Sunworld Online, 02: ??, 1996. ISSN 1091-8914.

URL http://www.sun.com/
sunworldonline/swol-02-1996/
swol-02-career.html.

Sadun:1996:JCC

[Sad96] Erica Sadun. JavaScript CD Cookbook. Charles River Media, Inc., 403 VFW Drive, PO Box 417, Rockland, MA 02370, USA, 1996. ISBN 1-886801-35-5. ???? pp. LCCN ???? US$39.95. URL http://www.

amazon.com/exec/obidos/ISBN%
3D1886801355/wholesaleproductA/
1624-2272339-016690;   http:
//www.charlesriver.com/titles/
java.html.

Sadr:1997:UOO

[Sad97a]  Babak Sadr.  Unified Ob-

jects Object-Oriented Programming Using C++ (with an Introduction to Java). IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1997. ISBN 0-8186-7733-3. xxvii + 429 pp. LCCN QA76.64.S25 1997. US$40. URL http:/

                                        /www.computer.org/cspress/
                                        catalog/bp07733.htm.

Sadun:1997:JCC

[Sad97b] Erica Sadun. JavaScript CD Cookbook. Charles River Media, Inc., 403 VFW Drive, PO Box 417, Rockland, MA 02370, USA, second edition, July 1, 1997. ISBN 1-886801-65-7. ???? pp. LCCN ???? US$44.95. URL http://www.

                                        amazon.com/exec/obidos/ISBN=
                                        1886801657/1624-2272339-016690.

Sadun:1998:JC

[Sad98] Erica Sadun. Die JavaScript CD. Midas Verlag, Zürich, Switzerland, 1998. ISBN 3-907020-40-5. ???? pp. LCCN ???? 79 DM. Includes CD-ROM.

Sakharov:1998:TCP

[Sak98] Alex Sakharov. Technical correspondence: Processing recursive data types in Java. ACM SIG-PLAN Notices, 33(12):33, December 1998. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Salus:1996:BRB

[Sal96] Peter H. Salus. Book reviews: The bookworm. ;login: the USENIX Association newsletter, 21(5):45–48, October 1996.

CODEN LOGNEM. ISSN 1044-6397. Review of eight books on Java and JavaScript.

Salus:1997:LDJ

[Sal97] Peter H. Salus. Languages:

From DEL to Java. UNIX Developer, 1(1):29–30, January/ February 1997. ISSN 1090-2279.

Saleh:19xx:LDW

[Salxx]  Imad Saleh.  Les langages

de développement sur Web: Java, Shockwave, VRML. Editions Hermès, ????, 19xx. ISBN 2-86601-542-8. LCCN ???? 145 FF. URL http:

//www.editions-hermes.fr/
principal/html/saleh.htm.

Sams:1996:PJP

[Sam96a] Sams Development Team Staff.

Professional Java Programming Kit. Sams.net, Indianapolis, IN 46268, USA, 1996. ISBN 1-57521-133-5. ???? pp. LCCN ???? US$129.99.

Samsnet:1996:JSS

[Sam96b] Sams.net. Java Script Starter Kit C/Ww95/Ww. Sams.net, Indianapolis, IN 46268, USA, August 1, 1996. ISBN 1-57521-170-X. ???? pp. LCCN ???? US$69.99.

Sams:1997:JAD

[Sam97] Sams. Java al Decubierto.

Prentice Hall HispanoAmericana, Mexico, DF, Mexico / Nueva York, NY, USA, 1997. ISBN 84-89660-59-X. 898 pp.

LCCN ???? URL http:/

/prentice.com.mx/;   http:
//prentice.com.mx/libros/
javadesc.html. Includes CD-

ROM.

Sanders:1997:JCB

[San97a] B. Sanders. Javaone-97 — catching the buzz. IEEE Micro, 17(3):5–6, May/June 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//pascal.computer.org/mi/
books/mi1997/pdf/m3005.pdf.

Sankar:1997:SEU

[San97b] Krishna Sankar. Special Edition Using Java Class Libraries. Que Corporation, Indianapolis, IN, USA, 1997. ISBN 0-7897-1292-X. ???? pp. LCCN ???? 118,00 DM; US$60. URL http:/

/www.mut.ch/profit.htx?&x_
pk=kd84mf7ms9dm.

Sanders:1998:PPC

[San98a] James Sanders. Products for practitioners: Caché; SQml Web 1.0; Temporal Rover; Solutions: Schedule; Installshield Java Edition. IEEE Software, 15(1):103, 102, January/February 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

URL http://dlib.computer.
org/so/books/so1998/pdf/s1103.
pdf.

Sanders:1998:PP

[San98b]  James Sanders.  Products

for practitioners: Windows DLLs; Windows memory manager; middleware support; remote server; Java decompiler; Java connectivity; Java quick reference; terminal emulator. IEEE Software, 15(4):83–86, July/August 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

URL http://dlib.computer.
org/so/books/so1998/pdf/s4083.
pdf.

Sanders:1998:PPW

[San98c]  James Sanders.  Products

for practitioners: Windows DLLs; Windows memory manager; middleware support; remote server; Java decompiler; Java connectivity; Java quick reference; terminal emulator. IEEE Software, 15(4):83–86, July/August 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

URL http://dlib.computer.
org/so/books/so1998/pdf/s4083.
pdf.

Sarmenta:1998:BWB

[Sar98] L. F. G. Sarmenta. Bayanihan:

Web-based volunteer computing using Java. Lecture Notes in Computer Science, 1368:444– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sarmenta:1999:AFT

[Sar99] L. F. G. Sarmenta. An adaptive, fault-tolerant implementation of BSP for Java-based volunteer computing systems. Lecture Notes in Computer Science, 1586:763–??, 1999. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Saunders:1996:CWA

[Sau96] Stephen Saunders. Concentrated WAN access to ATM and frame relay. Data communications, 25(10):??, August 1996. CODEN DACODM. ISSN 0363-6399.

Savail:1996:JAP

[Sav96] L. Savail. Java Applet Powerpack, Vol. 1. Prima Publishing, Rocklin, CA, USA, 1996. ISBN 0-7615-0678-0. xii + 344 pp. LCCN QA76.73.J38J36 1996. US$35.00. URL http://

                                        www.webcom.com/~prima/. The

book contains no author credit, although library catalogs credit L. Savail.

Savarnejad:1997:PTU

[Sav97] Atoosa Savarnejad. Princeton team uncovers flaw in Java. Network Security, 1997 (6):6–7, June 1997. CO-DEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (elec-

                                        tronic). URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S1353485897895268.

Savitch:1999:FCU

[Sav99] Walter Savitch. A First Course Using Java. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1999. ISBN 0-13-287426-1. xxiv + 726 pp. LCCN A76.73.J38 S27 1998. ???? URL http:

//www.prenhall.com/books/
esm_0132874261.html.

Schafer:1996:WRL

[SB96] Steve Schafer and Elizabeth Bruns. Webmaster Resource Library: Using Java. Que Corporation, Indianapolis, IN, USA, October 1996. ISBN 0-7897-0872-8. ???? pp. LCCN ???? US$150.00.

Savage:1997:SRJ

[SB97] Colin Savage and Steven R.

Bothe. Special report: Java: Sound off: Java vs. ActiveX. Computerworld, 31(19):80–82, May 12, 1997. CODEN CMP-WAB. ISSN 0010-4841.

Suematsu:19xx:JR

[SBxx] Chihiro Suematsu and Keith Bennett. Java Revolution. Diamond, Inc., ????, 19xx. ISBN 4-483-71970-0. ???? pp. LCCN ???? 1,900 yen.

Sah:1996:PIS

[SBB96] A. Sah, K. Brown, and E. Brewer. Programming the Internet from the server-side with Tcl and Audience1. In Association [Ass96b], pages 235– ??, 183–188. ISBN 1-880446-78-2. LCCN QA76.73.T44 T44 1996.

Schaller:1997:UJC

[SBB^{+}97] Nan C. Schaller, Michael Berman, Judith Bishop, Paddy Nixon, Evelyn Rozanski, and Peter Welch. Using Java in computer science education (panel).

SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(3): 140–142, September 1997. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Schumann:1997:PJS

[SBH97] Michael Schumann and Antje Binas-Holz. Das Programmierbuch Java 1.1: das Sprachkonzept: Programme und Applets; die Werkzeuge des JAVA-1.1-Developer-Kits (JDK) und nutzliche Add-Ons; Sicherheits- und Verschlusselungstechnologien unter JAVA 1.1; mit ubersichtlicher Referenz des neuen JAVA 1.1 API . Sybex Verlag GmbH, Erkrather Straße 345-349, 40231 Düsseldorf, Germany, 1997. ISBN 3-8155-5420-9. xxxii + 1096 pp. LCCN ???? 89 DM. URL http:

//www.sybex.de/bestell/tp/
tp05420.htm.

Simkin:1996:JPE

[SBL96] Steve Simkin, Neil Bartlett, and Alex Leslie. The Java Programming Explorer. Coriolis Group Books, Scottsdale, AZ, USA, March 1, 1996. ISBN 1-883577-81-0. xxiv + 820 pp. LCCN QA76.73.J38S55 1996. US$39.99, CDN$55.99.

URL http://165.247.200.
139/cgi-bin/java.cgi?page=
ibook&graphics=1#JPE;   http:
//www.coriolis.com/site/msie/
books/ind/javapex.htm. In-

cludes CD-ROM.

Shetterly:1996:CPO

[SC96a] Rick Shetterly and Alan Chmura.

Creating the paperless office. IEEE Software, 13(2):124–126, March 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). Discusses the Java development environment.

Stanton:1996:TTP

[SC96b] Scott Stanton and Ken Corey.

TclJava: Toward portable extensions. In Association [Ass96b], pages 25–29. ISBN 1-880446-78-2. LCCN QA76.73.T44 T44 1996.

Shin:1997:JH

[SC97a] Myungki Shin and Heechang Choi. Java Hunting. Information Age Co., ??, Korea, 1997. ISBN 89-85346-50-4 (??Invalid checksum??). 12,000 won pp. LCCN ???? URL http://

www.infoage.co.kr/publish/
index.html.

Sinclair:1997:JWMa

[SC97b] Joseph T. Sinclair and Lee Callister. Java Web Magic. Hayden Books, 4300 West 62nd Street, Indianapolis, IN 46268, USA, March 1, 1997. ISBN 1-56830-341-6. x + 224 pp. LCCN QA76.73.J38S555 1997. US$39.99. URL http://

merchant.superlibrary.com:
8000/catalog/hg/PRODUCT/PAGE/
15683/bud/1568303416.html.

Schlender:1995:WII

[Sch95] Brent Schlender. Whose Internet is it, anyway? Fortune, page ??, December 11, 1995. CODEN FORTAP. ISSN 0015-8259. URL http:

                                        //pathfinder.com/fortune/
                                        magazine/1995/951211/infotech.
                                        internet.html.

Schildt:1996:SEC

[Sch96a]  Herbert Schildt.  Schildt’s

Expert C++: Includes Java Overview. Osborne/McGraw-Hill, Berkeley, CA, USA, 1996. ISBN 0-07-882209-2 (softcover). xviii + 402 pp. LCCN QA76.73.C153S337 1996. US$31.45.

Schlack:1996:CCW

[Sch96b]  Mark Schlack.  Computing

crossroads: Will it be Intel inside or Java everywhere? BYTE Magazine, 21(11):14, November 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Schmidt:1996:NNS

[Sch96c] Eric Schmidt. Networking and the new social landscape, 1996. 1 videocassette (57 min.). Order number 14-Schmidt.

Schneider:1996:ILJ

[Sch96d] Karen G. Schneider. Internet librarian: “Java: The third wave of the Internet”. American Libraries, 27(2):60–??, 1996. ISSN 0002-9769.

Schach:1997:SEJ

[Sch97a] Stephen R. Schach. Software Engineering With Java. Richard D. Irwin, Chicago, IL, February 1997. ISBN 0-256-24167-8. xxii + 618 pp. LCCN QA76.73.J38S3 1997. US$42.

URL http://www.irwin.com/
catalogs/924167.html.

Scheir:1997:SRJ

[Sch97b] Robert L. Scheir. Special report: Java: Jury’s still out. Computerworld, 31(19):74–76, May 12, 1997. CODEN CMP-WAB. ISSN 0010-4841.

Schildt:1997:BCC

[Sch97c] Herbert Schildt. Borland C++:

The Complete Reference: Create Java Applets and Applications, Develop Windows 95 Programs. Osborne/McGraw-Hill, Berkeley, CA, USA, 1997. ISBN 0-07-882230-0 (softcover). xvi + 1114 pp. LCCN QA76.73.C153S318 1997. US$35.95.

Schlabach:1997:JPV

[Sch97d]  Torsten Schlabach.  Java-

Programmierung mit Visual J++. Addison-Wesley, Reading, MA, USA, 1997. ISBN 3-8273-1191-8. 400 pp. LCCN ???? 60 DM.

Schlack:1997:ERJ

[Sch97e] Mark Schlack. Editorial — raising the Java standard. BYTE Magazine, 22(11):10, November 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Schlack:1997:ELD

[Sch97f] Mark Schlack. Editorial: Living dangerously: The computer industry is engaged in a platform war. which side should you be on? BYTE Magazine, 22(1):14, January 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic). Discusses Java, Network Computers, and NetPCs.

Schmidt:1997:JS

[Sch97g]  Volker Schmidt.  Java 1.1

Schnellübersicht. Markt & Technik, Haar bei München, Germany, July 1997. ISBN 3-8272-5219-9. 543 pp. LCCN ????

39,95 DM. URL http://www.
mut.ch/html/buchinfo.htx?
&x_pp=dJgxRqA9wlAYNijX&x_
az=10&x_pk=VpLvjnsx7b5V&x_
pa=B&x_pr=oS&x_bestell=25219&
x_stichwort=Java&x_sprache=
A&x_start=1&x_gefunden=92&
x_query=&x_lastpage=buchfund&
x_soft=&x_suchewo=;   http:
//www.mut.ch/profit.htx?&
x_pk=kd84mf7ms9dm.

Schoenefeld:1997:OOD

[Sch97h] Dale A. Schoenefeld. Objectoriented design and programming: an Eiffel, C++, and Java course for C programmers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(1): 135–139, March 1997. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Schreiber:1997:IWC

[Sch97i] Sven B. Schreiber. Inside Windows Cabinet files. Dr. Dobb’s Journal of Software Tools, 22 (5):20–??, May 1997. CODEN DDJOEB. ISSN 1044-789X.

Schroeder:1997:VTJ

[Sch97j] Will Schroeder. The Visualization Toolkit (in Java): An Object-Oriented Approach to 3-D Graphics. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, second edition, December 1997. ISBN 0-13-954694-4. LCCN

????  ????  URL http:

//www.prenhall.com/books/
ptr_0139546944.html.

Schulz:1997:AJ

[Sch97k]  Kay Schulz.  Alles über

Java. Managementwissen Zukunft, ????, October 1997. ISBN 3-932506-08-1 (??invalid ISBN??). LCCN ???? 79 DM. URL http://

home.t-online.de/home/GS.
VerlagMWZ/javaabs.htm.

Schumann:1997:EEJ

[Sch97l] Michael Schumann. Der Erfolgreiche Einstieg in Java. Ullstein, Frankfurt/Main und Berlin, Germany, 1997. ISBN 3-548-41090-1. 256 pp. LCCN ???? 20 DM.

Scholten:1998:J

[Sch98a] Scholten. Java. Addison-Wesley Nederland, ????, January 1998. ISBN 90-6789-938-0. LCCN ???? fl. 50.

Schulz:1998:JSS

[Sch98b] Paul Schulz. Java 1.2. Schnell und sicher zum Ziel. Markt & Technik, Haar bei München, Germany, May 1998. ISBN 3-453-14251-9. LCCN ???? 20 DM.

Schaps:1999:PTC

[Sch99a] Gary L. Schaps. Programmer’s toolchest: Compiler construction with Antlr and Java. Dr. Dobb’s Journal of Software Tools, 24(3):84, 86–89, March 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/1999/9902/9902toc.
htm;   http://www.ddj.com/
ftp/1999/1999_03/antlr.txt;
 http://www.ddj.com/ftp/1999/
1999_03/antlr.zip.

Schreiner:1999:PCA

[Sch99b] Wolfgang Schreiner. A parallel computer algebra system based on Maple and Java. Technical report 99-08, RISC-Linz, Johannes Kepler University, Linz, Austria, March 1999.

URL http://www.geocities.
com/%7Etechnionges/Pages/
Sympos.html;   http://www.
risc.uni-linz.ac.at/people/
schreine/papers/swcomm99/
distmaple.rtf;   http://www.
risc.uni-linz.ac.at/people/
schreine/papers/swcomm99/
swcomm99.ps.gz. Presented

at Software for Communication Technologies: Third International Austrian-Israeli Technion Symposium with Industrial Forum, April 26–27, 1999, Hagenberg, Austria.

Schwartz:1999:DIB

[Sch99c] Robert Allan Schwartz. Default initialization of built-in types.

C/C++ Users Journal, 17(6): ??, June 1999. CODEN CCU-JEX. ISSN 1075-2838.

Schlabach:19xx:JPV

[Schxxa]  Torsten Schlabach.  Java-

Programmierung mit Visual J++. Addison-Wesley, Reading, MA, USA, 19xx. ISBN 3-8273-1191-8 (??invalid ISBN??). LCCN ???? 60 DM.

Schmid:19xx:JS

[Schxxb]  Volker Schmid.  Java 1.2

Schnellübersicht. Markt & Technik, Haar bei München, Germany, 19xx. ISBN 3-8272-9040-6 (??invalid ISBN??). LCCN ???? 46 SF. URL

http://www.mut.ch/html/v2/
Produkt-Information.htx?&
x_pp=204%2e94%2e76%2e206VanW4xQbKi&
x_pk=Direkt-Fr&x_stichwort=
java&x_suchsprache=A&x_search=
T&x_artikelnr=25435.

Schnoor:19xx:IHJa

[Schxxc] Peter Schnoor. Intranet, HTML og Java. KnowWare, Ordrupvej 63 C, 2920 Charlottenlund, Denmark, 19xx. ISBN ???? ???? pp. LCCN

????  URL http://www.
knowware.dk/dk/dk06.htm.

Schnoor:19xx:IHO

[Schxxd] Peter Schnoor. Intranet, HTML og Java. KnowWare, ????, 19xx. LCCN ???? ????

Schnoor:19xx:IHJb

[Schxxe] Peter Schnoor. Intranet, HTML und Java. KnowWare, Ordrupvej 63 C, 2920 Charlottenlund, Denmark, 19xx. ISBN ???? 48 pp. LCCN ????

URL http://www.knowware.
dk;   http://www.knowware.
dk/de/de06.htm. Nr. 133.

Schnoor:19xx:IHJc

[Schxxf] Peter Schnoor. Intranet, HTML und Java. KnowWare, ????, 19xx. LCCN ???? ???? URL

http://www.knowware.dk/de/
internet.htm.

Schulz:19xx:JT

[Schxxg] Paul Schulz. Java 1.1 Taschenbuch. Markt & Technik, Haar bei München, Germany, 19xx. ISBN 3-8272-5227-X (??invalid ISBN??). LCCN ????

20 DM. URL http://www.
mut.ch/html/buchinfo.htx?
&x_pp=dJgxRqA9wlAYNijX&x_
az=10&x_pk=VpLvjnsx7b5V&x_
pa=B&x_pr=oS&x_bestell=25227&
x_stichwort=Schulz&x_sprache=
A&x_start=1&x_gefunden=2&
x_query=&x_lastpage=buchfund&
x_soft=&x_suchewo=A.

Scotkin:1996:BUJ

[Sco96a] Joel Scotkin. Business uses for Java — think client/ server computing and the Internet have changed your life? Java’s going to be twice as important — it will totally change the way you write applications and distribute them to clients. Datamation, 42(5):40–41, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Scott:1996:BAJ

[Sco96b] Adrian Scott. Business Applications of Java. SIGS Books, 71 W 23rd St., New York, NY 10010, USA, 1996. ISBN 1-884842-65-8. 50 pp. LCCN QA76.9.O35O34 1997. US$85.

URL http://www.sigs.com/
books/wp_javaapps.html.

ScottOwen:1996:IWW

[Sco96c] G. Scott Owen. Integrating World Wide Web technology into undergraduate education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28 (????):101–103, ???? 1996. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Stanek:1996:WPU

[SD96] William Robert Stanek and Steven J. DeRose. Web publishing unleashed: HTML, Java, CGI, VRML, SGML. Howard W. Sams, Indianapolis, IN 46268, USA, March 1, 1996. ISBN 1-57521-051-7. xxii + 916 pp. LCCN QA76.76.H94S73 1996. US$49.99. URL http://

merchant.superlibrary.com:
8000/catalog/hg/PRODUCT/PAGE/
15752/bud/1575210517.html.

Sowizral:1999:PVJ

[SD99] Henry A. Sowizral and Michael F.

Deering. Projects in VR: The Java 3D API and virtual reality. IEEE Computer Graphics and Applications, 19(3):12–15, May/June 1999. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL

                                        http://dlib.computer.org/
                                        cg/books/cg1999/pdf/g3012.
                                        pdf.

Sedgewick:1998:AJ

[Sed98] Robert Sedgewick. Algorithms in Java, Parts 1–4. Addison-Wesley, Reading, MA, USA, third edition, 1998. ISBN 0-201-36120-5. 720 (est) pp. LCCN QA76.73.J38S4

                                        1998. URL http://www.awl-
                                        he.com/titles/26809.html.

Sander:1998:HPC

[SEH98] V. Sander, D. Erwin, and V. Huber. High-performance computer management based on Java. Lecture Notes in Computer Science, 1401:526– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sander:1999:HCM

[SEH99a] Volker Sander, Dietmar Erwin, and Valentina Huber. Highperformance computer management based on Java. Future Generation Computer Systems, 15(3):425–??, ???? 1999. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Sander:1999:HPC

[SEH99b] Volker Sander, Dietmar Erwin, and Valentina Huber. Highperformance computer management based on Java. Future Generation Computer Systems, 15(3):425–432, April 1, 1999.

CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/19/
19/30/19/27/abstract.html.

Selvakumar:1999:ATW

[Sel99] M. Selvakumar. Automated testing for Web applications. Dr. Dobb’s Journal of Software Tools, 24(5):88, 90, 92, 95–96, May 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ftp/1999/1999_
05/webtest.txt.

Semich:1996:FSJ

[Sem96] J. William Semich. Framework — Sun’s Java development language is a lot more than just another new toy for the Internet. Datamation, 42(5):5–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Seshadri:1998:JQH

[Ses98] Govind Seshadri. Java Q and A:

How do I implement callbacks with Java’s RMI? Dr. Dobb’s Journal of Software Tools, 23 (3):123–124, 137–138, March 1998. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/ftp/1998/1998_03/
jqa398.txt.

Seybold:1996:ITG

[Sey96] Patricia B. Seybold. It’s time to get serious about Java. Computerworld, 30(32):2, August 5, 1996. CODEN CMPWAB. ISSN 0010-4841. Describes a 200,000-line Java project for travel bureau business.

Sarna:1996:PSJ

[SF96a] D. Sarna and G. Febish. Paradigm shift — Java’s hot, but only part of the big picture. Datamation, 42(3):25–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Sarna:1996:JPW

[SF96b] David E. Y. Sarna and George J. Febish. Java’s place in the wide OLE world. Datamation, 42 (????):25–27, ???? 1996. CO-DEN DTMNAT. ISSN 0011-6963.

Semich:1996:JIT

[SF96c] J. William Semich and David Fisco. Java: Internet toy or enterprise tool — Java’s a whole lot more than just an animation tool for the Web. Sun’s renaming it the “Java Enterprise Platform”. better said, it’s the Java network-centric computing platform. Datamation, 42(5):28–29, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Sarna:1997:PSWb

[SF97] David E. Y. Sarna and George J. Febish. Paradigm shift: Whither Java. Datamation, 43 (7):23–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Shrobe:1996:FL

[SFSS96] H. Shrobe, S. E. Fahlman, B. Stroustrup, and G. L. Steele, Jr. The future of Lisp. IEEE expert: intelligent systems and their applications, 11(1):10–16, February 1996. CODEN IE-EXE7. ISSN 0885-9000.

Sliwa:1997:MDW

[SG97a] Carol Sliwa and Sharon Gaudin.

Microsoft declares war: Stuns Sun with refusal to ship Java class libraries. Computerworld, 31(30):1, 17, July 28, 1997. CO-DEN CMPWAB. ISSN 0010-4841. See [SG97a, Joh97a] for backslash.

Sliwa:1997:MSR

[SG97b] Carol Sliwa and Sharon Gaudin.

Microsoft stances roils users: Threat to cross-platform elicits strong disapproval. Computerworld, 31(31):1, 100, August 4, 1997. CODEN CMPWAB. ISSN 0010-4841. See [SG97a] for original story, and [Joh97a] for further comments.

Smith:1997:MML

[SGB97] Taber H. Smith, Aaron E.

Gower, and Duane S. Boning. A matrix math library for Java. Concurrency: Practice and Experience, 9(11):1127–1137, November 1997. CODEN CPEXEI. ISSN 1040-3108. URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=13821;
http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=13821&PLACEBO=IE.pdf. Spe-

cial Issue: Java for computational science and engineering — simulation and modeling II.

Shoffner:1996:JWO

[SH96] M. Shoffner and M. Hughes.

Java and Web-executable object security. Dr. Dobb’s Journal of Software Tools, 21(11):38, 40, 42, 44, 46, 48, 49, November 1996. CODEN DDJOEB. ISSN 1044-789X.

Schulz:1997:JSS

[SH97] Paul Schulz and Jan Hermelink. Java 1.1: schnell und sicher zum Ziel. Markt & Technik, Haar bei München, Germany, 1997. ISBN 3-8272-5227-X. 361 pp. LCCN ????

20 DM. URL http://www.
mut.ch/html/buchinfo.htx?
&x_pp=dJgxRqA9wlAYNijX&x_
az=10&x_pk=VpLvjnsx7b5V&x_
pa=B&x_pr=oS&x_bestell=25227&
x_stichwort=Schulz&x_sprache=
A&x_start=1&x_gefunden=2&
x_query=&x_lastpage=buchfund&
x_soft=&x_suchewo=A;   http:
//www.mut.ch/zbin/webdbc.
dll/support/leer/htx/&/html/
buchinfo.htx?&x_merkenbestell=
25227&x_stichwort=java&x_
sprache=D&x_start=1&x_pp=
UBqkUYviWj&x_lang=D&x_az=
75&x_hg=FFFFFF&x_gefunden=
9&x_query=&x_vonwo=buchfund&
x_soft=1;   http://www.mut.
com/.

Sarmenta:1999:BBS

[SH99] Luis F. G. Sarmenta and Satoshi Hirano. Bayanihan: building and studying Webbased volunteer computing systems using Java. Future Generation Computer Systems, 15 (5–6):675–686, October 1, 1999. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.

elsevier.com/gej-ng/10/19/
19/30/21/28/abstract.html.

Shah:1995:SRS

[Sha95]  Rawn Shah.  Show re-

port: SIGGRAPH 95/Java BOF. Sunworld Online, 09: ??, 1995. ISSN 1091-8914.

URL http://www.sun.com/
sunworldonline/swol-09-1995/
swol-09-siggraph.html.

Shafer:1996:JNW

[Sha96a] Dan Shafer. JavaScript and Netscape 2 Wizardry. Coriolis Group Books, Scottsdale, AZ, USA, May 1, 1996. ISBN 1-883577-86-1. xv + 365 pp. LCCN QA76.J39S48 1996. US$34.99. URL http://165.

247.200.139/cgi-bin/java.
cgi?page=ibook&graphics=1#
JPE;   http://www.coriolis.
com/site/msie/books/ind/nscp2wiz.

htm. Includes CD-ROM.

Shao:1996:MTJ

[Sha96b] Sami Shao. Morning tutorial:

Java software secrets. In IEEE [IEE96e], page ?? ISBN ???? LCCN ????

Shammas:1997:VJD

[Sha97a] Namir Clement Shammas. Visual J++ for Dummies Quick Reference. IDG Books, San Mateo, CA, USA, August 1997. ISBN 0-7645-0253-

0. xii + 212 pp. LCCN QA76.73.J38S45 1997. US$15.

URL http://www.dummies.
com/cgi/fill_out_template.
pl?book:0-7645-0253-0:book-
Dummies+Press::u151219.

Shaw:1997:ISS

[Sha97b] N. Shaw. Integrating SGML, STEP/EXPRESS, HTML, and Java to make product data accessible across the Web. In Anonymous [Ano97w], pages 115–118. ISBN 1-901782-01-8. LCCN ????

Shaffer:1998:PID

[Sha98a] Clifford A. Shaffer. A Practical Introduction to Data Structures and Algorithm Analysis: Java Edition. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1998. ISBN 0-13-660911-2. xvii + 488 pp. LCCN QA76.9.D35S45 1998. US$54.95. URL http:

                                        //www.prenhall.com/books/
                                        esm_0136609112.html.

Shahi:1998:CVM

[Sha98b] Avantika Shahi. A cybertheatre a virtual movie zone designed using VRML and JavaScript. ????, ????, 1998. 44 pp.

Souder:1998:JCBa

[SHGD98a] Dan Souder, Morgan Herrington, Rajat P. Garg, and Dennis DeRyke. JSPICE: a component-based distributed Java front-end for SPICE. In ACM [ACM98a], page ?? CODEN CPEXEI. ISSN 1040-3108. LCCN ????

                                        URL http://www.cs.ucsb.
                                        edu/conferences/java98/papers/
                                        jspice.pdf;   http://www.
                                        cs.ucsb.edu/conferences/java98/
                                        papers/jspice.ps. Also pub-

lished as Concurrency: Practice and Experience, 10(11– 13), September 1998, CODEN CPEXEI, ISSN 1040-3108.

Souder:1998:JCBb

[SHGD98b] Dan Souder, Morgan Herrington, Rajat P. Garg, and Dennis Deryke. JSPICE: a componentbased distributed Java frontend for SPICE. Concurrency: Practice and Experience, 10(11–13):1131–1141, September 1998. CODEN CPEXEI. ISSN 1040-3108. URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=10050401;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=10050401&PLACEBO=IE.pdf.

Special Issue: Java for Highperformance Network Computing.

Shinsato:1999:WHP

[Shi99] Harold Shinsato. Writing highperformance graphical Java components. Dr. Dobb’s Journal of Software Tools, 24(9):50, 52–54, September 1999. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/1999/1999_09/
javacomp.txt.

Shing:19xx:SJ

[Shixxa] Chou Chung Shing. Search on Java. Unalis Corporation, ??, Taiwan, 19xx. ISBN 957-22-2084-5. ???? pp. LCCN ????

220 NT$. URL http://www.
unalis.com.tw/index.html.

Shing:19xx:IPD

[Shixxb] Yang We Shing. Internet Programming and Design (Java and HTML). Chuen Hwa Books & Technology Co., Ltd., ????, 19xx. ISBN ???? LCCN ???? 250 NT$.

Shiu:19xx:EKJ

[Shixxc] Chi Shiu. Easy To Know Java. Infowide Book Co., ????, 19xx. ISBN 957-726-335-6 (??invalid ISBN??). ???? pp. LCCN ???? 130 NT$.

Shoffner:1996:JBA

[Sho96a] Michael Shoffner. Java’s busting out all over. BYTE Magazine, 21(5):32, May 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Shotsberger:1996:TGI

[Sho96b] P. G. Shotsberger. Toward a greener Internet. Computer, 29 (12):113–114, December 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Salo:1999:OPB

[SHR^{+}99] Timo Salo, Justin Hill, Scott Rich, Chuck Bridgham, and Daniel Berg. Object persistence: Beyond serialization. Dr. Dobb’s Journal of Software Tools, 24(5):19, 22, 24– 27, 30–33, May 1999. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/1999/1999_05/
                                        persist.txt.

Shulz:1997:JSS

[Shu97a] Paul Shulz. Java 1.1. Schnell und sicher zum Ziel. Markt & Technik, Haar bei München, Germany, 1997. ISBN 3-453-13559-8. 391 pp. LCCN ???? 20 DM.

Shur:1997:EIC

[Shu97b] Jim Shur. Exploring inner classes in Java 1.1: As Java matures, it begs comparison to the robustness of C++. see how JDK 1.1’s inner classes compare to nested classes. C++ Report, 9(5):18–29, 68, May 1997. CODEN CRPTE7. ISSN 1040-6042.

Sarmenta:1998:TBBa

[SHW98a] Luis F. G. Sarmenta, Satoshi Hirano, and Stephen A. Ward. Towards Bayanihan: Building an extensible framework for volunteer computing using Java. In ACM [ACM98a], page ?? CODEN CPEXEI. ISSN 1040-3108. LCCN ????

URL http://www.cs.ucsb.
edu/conferences/java98/papers/
bayanihan.pdf;   http://www.
cs.ucsb.edu/conferences/java98/
papers/bayanihan.ps. Also

published as Concurrency: Practice and Experience, 10(11– 13), September 1998, CODEN CPEXEI, ISSN 1040-3108.

Sarmenta:1998:TBBb

[SHW98b] Luis F. G. Sarmenta, Satoshi Hirano, and Stephen A. Ward. Towards Bayanihan: building an extensible framework for volunteer computing using Java. Concurrency: Practice and Experience, 10(11–13): 1015–1019, September 1998. CODEN CPEXEI. ISSN 1040-3108. URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=10050419;
                                         http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=10050419&PLACEBO=IE.pdf.

Special Issue: Java for Highperformance Network Computing.

Sienkiewicz:1998:UJ

[Sie98] Mark Sienkiewicz. Using Java. ;login: the USENIX Association newsletter, 23(2):??, April 1998. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

                                        usenix.org/publications/java/
                                        usingjava10.html.

SGI:1995:SGS

[Sil95] Silicon Graphics, Inc. Silicon Graphics, Sun Microsystems and Netscape join forces to take the Web to the next level. press release, December 4,

                                        1995. URL http://www.sgi.
                                        com/Products/cosmo/sgisun.
                                        htm.

Sim:1996:FFP

[Sim96a] Philip Sim. Financial firm puts Java to work on Web, in-house. Computerworld, 30(36):48, September 2, 1996. CODEN CMP-WAB. ISSN 0010-4841.

Simeonides:1996:JBT

[Sim96b] Alex Simeonides. Java and the Beans talk. SunExpert Magazine, 7(12):6, December 1996. ISSN 1053-9239. URL http://

splash.javasoft.com/. Notes

the imminent release of the JavaBeans Developer’s Kit, and that this will be integrated into Sun’s Java Developer’s Kit version 1.1; parts are already available in JDK version 1.02. The JavaBeans specification is available on the World Wide

Web at URL http://splash.
javasoft.com/.

Simeonides:1996:MPI

[Sim96c] Alex Simeonides. Microsoft to play by Internet rules. WebServer Magazine: For Managers of World Wide Web Sites, 1(3):6, September/ October 1996. ISSN 1087-4232. Discusses ActiveX and Java.

Simeonides:1996:SPT

[Sim96d] Alex Simeonides. Sun plugs Tcl/Tk into Netscape. Sun-Expert Magazine, 7(9):6, September 1996. ISSN 1053-9239. Discusses SpecTcl and Spec-Java GUI builders, and the Safe-Tcl feature for giving Javalike safety to scripts written in Tcl/Tk.

Simeonides:1996:THE

[Sim96e] Alex Simeonides. Threat of hostile email applets looms. Web-Server Magazine: For Managers of World Wide Web Sites, 1(4):10, 12, November/ December 1996. ISSN 1087-4232. Discusses Java applets that forge email and gather email addresses, and the possibility of applets embedded in email that are executed when the email is read. Mentions a related book [MF97b].

Simonsen:1996:DSJ

[Sim96f] P. Simonsen. Decision support in Java: a simple rule-based decision support system to illustrate some of the features of the Java programming language. .EXE: the software developers’ magazine, 10(8):30–??, 1996. CODEN EXEEE5. ISSN 0268-6872.

Simeonides:1997:IDCa

[Sim97a] Alex Simeonides. IBM deepens commitment to Java. Web-Server Magazine: For Managers of World Wide Web Sites, 2(1):10–11, February 1997. ISSN 1087-4232.

Simeonides:1997:IDCb

[Sim97b] Alex Simeonides. IBM deepens committment to Java. RS/ Magazine, 6(2):9–10, February 1997. ISSN 1088-0844, 1061-0030.

Simeonides:1997:JAA

[Sim97c] Alex Simeonides. Java API to access serial, parallel ports. SunExpert Magazine, 8(6):20– 22, June 1997. ISSN 1053-9239.

Simeonides:1997:SSC

[Sim97d] Alex Simeonides. Sun serves up corporate Java. SunExpert Magazine, 8(6):6–7, June 1997. ISSN 1053-9239. URL http://

                                        www.cpg.com.

Simeonides:1997:SEJ

[Sim97e] Alex Simeonides. Sun to extend Java RMI with IIOP. SunExpert Magazine, 8(8):14, August 1997. ISSN 1053-9239.

Simon:19xx:JE

[Simxx]  Hanno Simon.  Java, eine

Einführung. Sybex Verlag GmbH, Erkrather Straße 345-349, 40231 Düsseldorf, Germany, 19xx. ISBN 3-8155-7219-3. ???? pp. LCCN ???? 49 DM.

Singh:1995:CNDb

[Sin95a] Jai Singh. Column: From the news desk — it seems pigs can fly. InfoWorld, page ??, December 12, 1995. CODEN IN-WODU. ISSN 0199-6649. URL

http://www.infoworld.com/.

Singh:1995:CNDa

[Sin95b] Jai Singh. Column: From the news desk — the Cold War of the Internet. InfoWorld, page ??, December 04, 1995. CODEN INWODU. ISSN 0199-6649. URL http://www.

infoworld.com/.

Singleton:1996:HCJ

[Sin96a] Andrew Singleton. A hot cup of Java. BYTE Magazine, 21 (4):129–??, April 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Singleton:1996:WW

[Sin96b] Andrew Singleton. Wired on the Web. BYTE Magazine, 21(1):

77–78, 80, January 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic). Discusses Java and Hot-Java.

Sinclair:1997:JWMb

[Sin97a] Joseph Sinclair. Java Web Magic. Academic Service, Postbus 81, 2870 AB Schoonhoven, The Netherlands, 1997. ISBN 90-395-0710-4. ca. 225 pp. LCCN ???? fl. 45. URL http:/

/acaserv.nl/bestel/5_0710.
htm;   http://www.acaserv.
nl; http://www.acaserv.nl/

5_0710.htm. Dutch translation of [SC97b]. Includes CD-ROM.

Sinclair:1997:JWMc

[Sin97b] Joseph Sinclair. Java Web Magic. Academic Service Verlag, ????, 1997. ISBN 3-932073-25-8 (??invalid ISBN??). ???? pp. LCCN ???? 45 DM.

URL http://www.sybex.de/
bestell/tp/tp02016.htm. In-

cludes CD-ROM.

Sinclair:19xx:JWMb

[Sinxx] Joseph Sinclair. Java Web Magic. Academic Service Verlag, ????, 19xx. ISBN 3-932073-25-8 (??invalid ISBN??). LCCN ???? 45 DM.

Siple:1998:CGJ

[Sip98] Matthew D. Siple. The Complete Guide to Java Database Programming with FDBC: JDBC, ODBC & SQL. Java Masters Series. McGraw-Hill, New York, NY, USA, January 1998. ISBN 0-07-913286-3 (softcover). LCCN QA76.625 .S557 1998. US$44.95. URL http:

//mcgraw-hill.inforonics.
com/cgi/getarec?mgh31395%
comp.

Siyan:1996:IVJ

[Siy96] Karanjit S. Siyan. Inside Visual J++. New Riders Publishing, Carmel, IN, USA, 1996. ISBN 1-56205-682-4. xxxiii + 899 pp. LCCN QA76.73.J38 S59 1996. US$45. URL http:

//www.mcp.com/cgi-bin/bag?
isbn=1-56205-682-4&last=/
bookstore.

Siyan:1997:TVJ

[Siy97] Karanjit Siyan. Todo Visual J++ 1.0 y 1.1. Editorial Inforbooks, ????, 1997. ISBN 84-89700-42-7 (??invalid ISBN??). LCCN ???? ????

Siyan:1998:VJP

[Siy98] Karanjit Siyan. Visual J++ für Profis. Academic Service Verlag, ????, 1998. ISBN 3-932073-12-6 (??invalid ISBN??). LCCN ???? 98 DM.

Srinivas:1997:JBE

[SJRK97] Kankanahalli Srinivas, V. “Juggy”

Jagannathan, Y. V. Ramana Reddy, and Raghu Karinthi. Java and beyond: Executable content. Computer, 30(6):52, June 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sklar:1995:EGA

[Skl95]  David F. Sklar.  Embed-

ding graphics applications inside Web pages. The X Journal, 5(5):68–69, May 1995. CO-DEN XJOUEA. ISSN 1056-7003. Discusses Java use on the Web.

Steiner:1998:OJM

[SKN98] Andreas Steiner, Adrian Kobler, and Moira C. Norrie. OMS/ Java : Model extensibility of OODBMS for advanced application domains. Lecture Notes in Computer Science, 1413:115– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/
service/series/0558/bibs/
1413/14130115.htm;   http:
//link.springer-ny.com/link/
service/series/0558/papers/
1413/14130115.pdf.

Skok:1996:EII

[Sko96] D. Skok. Evolving imaging into the Internet revolution. Imaging Magazine, 5(1):104–106, 108, 110–112, 114, 118, 120–122, January 1996. CODEN IM-NYE9. ISSN 1063-4320.

Siddalingaiah:1997:JH

[SL97a] Madhu Siddalingaiah and Stephen D. Lockwood. Java How-To. Impress Corporation, ??, Japan, 1997. ISBN 4-8443-4798-5. ???? pp. LCCN ???? 4900 yen.

URL http://www.ips.co.jp/
;   http://www.ips.co.jp/
book/4798/4798.htm. Includes

CD-ROM. Japanese translation.

Struble:1997:LDJ

[SL97b] Craig A. Struble and Ted Lewis.

Letters: Deadlocks, and Java. Computer, 30(5):10, May 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). Response to [Lew97b].

Southey:1999:NJA

[SL99] F. Southey and J. G. Linders.

NOTIO — a Java API for developing CG tools. Lecture Notes in Computer Science, 1640:262– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Siddalingaiah:19xx:JHb

[SLxx] Madhu Siddalingaiah and Stephen D.

Lockwood. Java How-To. DaeRim Publishing, Seoul, Korea, 19xx. ISBN 89-7280-338-3 (??invalid ISBN??). ???? pp. LCCN ???? ???? URL http:

//www.drbook.co.kr/daerim/

jvht.htm. Korean translation.

Slattery:1997:UJ

[Sla97] Terry Slattery. Using Java. ;login: the USENIX Association newsletter, 22(5):??, October 1997. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

usenix.org/publications/java/
javaindex.html;   http://
www.usenix.org/publications/
java/usingjava7.html.

Stichnoth:1999:SGC

[SLC99] James M. Stichnoth, Guei-Yuan Lueh, and Michał Cierniak. Support for garbage collection at every instruction in a Java compiler. ACM SIG-PLAN Notices, 34(5):118–127, May 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.
acm.org:80/pubs/citations/
proceedings/pldi/301122/p118-

stichnoth/. See PLDI’99 proceedings [ACM99b].

Schultz:1999:TAS

[SLCM99] U. P. Schultz, J. L. Lawall, C. Consel, and G. Muller. Towards automatic specialization of Java programs. Lecture Notes in Computer Science, 1628:367– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sliwa:1997:CJC

[Sli97] Carol Sliwa. Can Java cut in in the real world with a heavy-duty application? Daiwa finds plenty to recommend Java. Computerworld, 31(34):1, 16, August 25, 1997. CODEN CMPWAB. ISSN 0010-4841.

Siddalingaiah:1996:JHD

[SLO96] Madhu Siddalingaiah, Stephen D. Lockwood, and Chris Ott. Java How-To: The Definitive Java Problem-Solver. Waite Group Press, Corte Madera, CA, USA, 1996. ISBN 1-57169-035-2. xviii + 644 pp. LCCN QA76.73.J38S53 1996. US$39.99. URL http:/

/www.kaizen.net/javabook/;
 http://www.mcp.com/cgi-bin/
bag?isbn=1-57169-035-2&last=
/bookstore.

Siddalingaiah:19xx:JHa

[SLOxx] Madhu Siddalingaiah, Stephen Lockwood, and Chris Ott. Java How To. Image Art Publishing, ????, 19xx. ISBN 957-9641-51-X. ???? pp. LCCN ???? 480 NT$. Chinese translation by Chai Bao Zing of [SLO96]. Includes CD-ROM.

Spaniol:1996:TDS

[SLPM96] Otto Spaniol, Claudia Linnhoff-Popien, and Bernd E. Meyer, editors. Trends in distributed systems: CORBA and beyond: International Workshop TreDS ’96, Aachen, Germany, October 1–2, 1996: proceedings, volume 1161 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 3-540-61842-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D5 I63 1996.

Saulpaugh:1999:IJ

[SM99] Tom Saulpaugh and Charles Mirho. Inside the JavaOS operating system. Addison-Wes-ley, Reading, MA, USA, 1999. ISBN 0-201-18393-5. xix + 184 pp. LCCN QA76.76.O63S3563 1999. US$29.95.

Simpson:1996:MSJ

[SMC^{+}96] Bruce Simpson, John Mitchell, Brian Christeson, A. Rehan Zaidi, and Jonathan Levine. Making Sense of Java: a Guide for Managers and the Rest of Us. Manning Publications, Greenwich, CT, USA, June 1, 1996. ISBN 1-884777-24-4 (softcover), 0-13-263294-2. xiii + 159 pp. LCCN QA76.73.J38S56 1996. US$29.95. URL http://www.

                                        browsebooks.com/Simpson/;
                                         http://www.manning.com/Simpson/
                                        index.html.

Sipson:1997:BRD

[SMC97] Roger Sipson, Susan McKay, and Wolfgang Christian. Book review: David Flanagan, Java in a Nutshell; Patrick Niemeyer and Joshua Peck, Exploring JAVA. Computers in Physics, 11(1):46–??, January 1997. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (elec-

                                        tronic). URL https://aip.
                                        scitation.org/doi/10.1063/
                                        1.4822513.

Samson:1996:PSQ

[SMEN96] W. B. Samson, I. M. Marshall, and D. G. Edgar-Nevill, editors. Proceedings of the 5th Software Quality Conference, Dundee, UK, 9–10 July 1996. University Abertay Dundee, Dundee, UK, 1996. ISBN ???? LCCN ????

Smith:1997:NGN

[Smi97] J. Douglas Smith. Novell’s Guide to Netware Java. I D G Books Worldwide, Indianapolis, IN, USA, January 1997. ISBN 0-7645-4524-8. ???? pp. LCCN ???? US$39.99.

Smith:19xx:JSC

[Smixxa] Randy J. Smith. Java: Strictly for C Programmers Video Tape Series. RJS Int’l Research, ????, 19xx. LCCN ????

US$100. URL http://www.
rjsinternational.com/java/
index.html.

Smith:19xx:UJV

[Smixxb] Randy J. Smith. Understanding Java Video Tape Series. RJS Int’l Research, ????, 19xx. ISBN ???? ???? pp. LCCN ???? US$250. URL http://

www.javatrain.com;   http:
//www.rjsinternational.com/
java/index.html.

Smith:19xx:UJC

[Smixxc] Randy J. Smith. Using Java to Create an Application Video Tape Series. RJS Int’l Research, ????, 19xx. LCCN ????

US$200. URL http://www.
rjsinternational.com/java/
index.html.

Schroeder:1996:DIO

[SML96] W. J. Schroeder, K. M. Martin, and W. E. Lorensen. The design and implementation of an object-oriented toolkit for 3D graphics and visualization. In Yagel and Nielson [YN96], pages 516–??, 93–100, 472. ISBN 0-89791-864-9. LCCN T385 .V59 1996. IEEE catalog number 96CB36006. IEEE Computer Society Press order number RS00118.

Sekiguchi:1999:SEJ

[SMY99] T. Sekiguchi, H. Masuhara, and A. Yonezawa. A simple extension of Java language for controllable transparent migration and its portable implementation. Lecture Notes in Computer Science, 1594:211–226, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Szwillus:1996:SBE

[SN96] Gerd Szwillus and Lisa Neal. Structure-Based Editors and Environments. Academic Press, New York, NY, USA, 1996. ISBN 0-12-681890-8. viii + 374 pp. LCCN QA76.9.H85 S89 1996. US$75.00. Discusses editors for Java, HTML, Visual Basic, Visual C, ….

Singal:1998:JF

[SN98a] Sandeep Singal and Binh Nguyen. The Java factor. Communications of the ACM, 41 (6):34–37, June 1998. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL http://www.
                                        acm.org:80/pubs/citations/
                                        journals/cacm/1998-41-6/p34-
                                        singal/.

Singhal:1998:JFI

[SN98b] Sandeep Singhal and Binh Nguyen. The Java factor: Introduction. Communications of the ACM, 41(6):34–37, June 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org:80/pubs/citations/
journals/cacm/1998-41-6/p34-
singal/.

Schildt:1997:JPP

[SO97a] Herbert Schildt and Joe O’Neil.

Java Programmer’s Pocket Reference. Osborne/McGraw-Hill, Berkeley, CA, USA, June 1997. ISBN 0-07-882368-4 (softcover). xi + 462 pp. LCCN QA76.73.J38S33 1997. US$13.49. URL http:

//www.osborne.com/program/
javaref.htm.

Schildt:1997:JPR

[SO97b] Herbert Schildt and Joe O’Neil.

Java Programmer’s Reference. Osborne/McGraw-Hill, Berkeley, CA, USA, 1997. ISBN 0-07-882364-4 (??Invalid checksum??). xi + 462 pp. LCCN QA76.73.J38 S33 1997. US$16.99, CDN$24.95.

Sommers:1997:DJR

[Som97]  Bret Sommers.  Distributed

Java: Remote Objects for Java. McGraw-Hill, New York, NY, USA, March 1997. ISBN 0-07-913249-9. 304 pp. LCCN ???? US$39.95.

Sood:1998:JQW

[Soo98a] Mukul Sood. Java Q and A:

What is Swing? Dr. Dobb’s Journal of Software Tools, 23 (9):111–114, 125, September 1998. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

                                        ddj.com/ftp/1998/1998_09/
                                        jqa998.txt;   http://www.
                                        ddj.com/ftp/1998/1998_09/
                                        jqa998.zip.

Sood:1998:PTE

[Soo98b] Mukul Sood. Programmer’s toolchest: Examining JDBC drivers: Current drivers fit into our of four categories. Dr. Dobb’s Journal of Software Tools, 23(1):82, 84–87, January 1998. CODEN DDJOEB. ISSN 1044-789X.

Stoker:1997:IJP

[SP97] Carol Stoker and Thomas Plew. An Introduction to Java Programming: Developing Applets Using Microsoft Visual J++. Course Technology, ????, December 1997. ISBN 0-7600-5043-0. ???? pp. LCCN

????  ????  URL http:/

                                        /www.course.com/templates/
                                        catalog/detail.cfm?isbn=0-
                                        7600-5043-0.

Sevenich:1998:JAP

[SP98] Richard Sevenich and Rick Price. JavaLanche: An avalanche predictor. Linux Journal, 51:??, July 1998. CO-DEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Spertus:1997:CGC

[Spe97]  Mike Spertus.  C++ and

garbage collection: Not your father’s garbage collection. Dr. Dobb’s Journal of Software Tools, 22(12):36, 38, 40–41, December 1997. CODEN DDJOEB. ISSN 1044-789X.

Spencer:1998:PVT

[Spe98] S. N. Spencer, editor. Proceedings VRML 98: third Symposium on the Virtual Reality Modeling Language, Monterey, California, February 16– 19, 1998. ACM Press, New York, NY 10036, USA, 1998. ISBN 1-58113-022-8. LCCN ????

Spiller:1996:JPI

[Spi96] Falko Spiller. Java: Die Programmiersprache für das Internet. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 3-540-61116-9. ca. 200 pp. LCCN ????

40 DM. URL http://www.
springer.de/. Includes CD-

ROM.

Spiegel:1999:PAD

[Spi99] A. Spiegel. Pangaea: An automatic distribution front-end for Java. Lecture Notes in Computer Science, 1586:93–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stegink:1999:CSG

[SPV99] Gordon Stegink, Jon Pater, and Daron Vroon. Computer science and general education: Java, graphics, and the Web. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1):146–149, March 1999. CODEN SIGSD3.

ISSN 0097-8418 (print), 2331-3927 (electronic).

Spyglass:1995:SLS

[Spy95] Spyglass, Inc. Spyglass licenses Sun Microsystems’ Java for Spyglass Mosaic. press release, November 8, 1995. URL http:/

/www.spyglass.com/current/
nov895.html.

Shobe:1996:JM

[SR96] Matt Shobe and Tim Ritchey. JavaScript for the Macintosh. Hayden Books, 4300 West 62nd Street, Indianapolis, IN 46268, USA, May 1996. ISBN 1-56830-278-9. viii + 326 pp. LCCN QA76.73.J38S48 1996. US$40.00, CDN$54.95, UK£37.50. URL http://

merchant.superlibrary.com:
8000/catalog/hg/PRODUCT/PAGE/
15683/bud/1568302789.html;
 http://www.mcp.com/221684522211009/
cgi-bin/bag?isbn=1-56830-
278-9&last=/hayden. Includes

CD-ROM.

Schengili-Roberts:1997:AHC

[SR97] Keith Schengili-Roberts. The Advanced HTML Companion: Extensive Coverage of Programming Techniques Involving CGI, JavaScript and VB-Script, How to Use HTML With Style Sheets, Java, ActiveX, Plug-Ins and Other Goodies. AP Professional, Boston, MA, USA, 1997. ISBN 0-12-623540-6 (softcover). xvii + 387 pp. LCCN QA76.76.H94S257 1997. US$26.95.

Sowizral:1998:JAS

[SRD98] Henry Sowizral, Kevin Rushforth, and Michael Deering. The Java 3D API Specification. Addison-Wesley, Reading, MA, USA, 1998. ISBN 0-201-32576-4. xx + 482 pp. LCCN QA76.73.J38S69 1997. US$35.

Sriram:1996:FTJ

[Sri96] K. B. Sriram. Free tools for Java. This important archive contains a number of Java tools, including Obsfuscator, Installer, Assembler, Jack (LL(k) parser generator), JavaCUP (LALR(1) parser generator), jb (convert flex/bison C output to Java), JavaLex (lex rewritten in Java), Pat (perl-style pattern matching), Guavac (Java compiler in C), lex/yacc grammars for Java, ANTLR specification, BNF grammar for Java, Sumatra Java-bytecode-to-C translator, j2c Java-bytecode-to-C translator, Jasmin assembler, Javaa assembler, Jobe (Java obfuscator), Mocha (Java decompiler), Kawa Scheme compiler to Java bytecodes, c2j (C++ to Java). Jas (Scheme to bytecode translator), Jax (regular-expression tokenizer), Jel (LL(1) parser generator), and more., 1996. URL http:

//www.blackdown.org/~kbs/.

Sridharan:1997:AJN

[Sri97a] Prashant Sridharan. Advanced Java Networking. SunSoft Press Java series. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1997. ISBN 0-13-749136-0 (softcover). xxiv + 368 pp. LCCN QA76.73.J38S75 1997. US$44.95. URL http://www.

                                        prenhall.com/allbooks/ptr_
                                        0137491360.html.  Includes

CD-ROM.

Sridharan:1997:JDR

[Sri97b]  Prashant Sridharan.  Jav-

aBeans: Developer’s Resource. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, August 1997. ISBN 0-13-887308-9 (softcover). xxi + 353 pp. LCCN QA76.73.J38S754 1997. US$44.95.

                                        URL http://www.prenhall.
                                        com/allbooks/ptr_0138873089.

html. Includes CD-ROM.

Summitt:1996:CCI

[SS96] Paul M. Summitt and Mary J. Summitt. Creating Cool Interactive Web Sites: Build The Next Hot Spot On The World Wide Web Using Cool Software — Java, CGI, VRML, And More! I D G Books Worldwide, Indianapolis, IN, USA, 1996. ISBN 1-56884-842-0 (softcover). xx + 412 pp. LCCN TK5105.888.S85 1996. US$26.99. Includes CD-ROM.

Shiran:1997:LAJ

[SS97] Yehuda Shiran and Tomer Shiran. Learn advanced JavaScript programming. Wordware Pub., Plano, TX, USA, 1997. ISBN 1-55622-552-0. ???? pp. LCCN QA76.73.J38 S46 1997.

Schwab:1998:AJCa

[SS98a] Matthias Schwab and Joel Schroeder. Algebraic Java classes for numerical optimization. In ACM [ACM98a], page ?? CODEN CPEXEI. ISSN 1040-3108. LCCN ????

URL http://www.cs.ucsb.
edu/conferences/java98/papers/
jest.pdf;   http://www.cs.
ucsb.edu/conferences/java98/
papers/jest.ps. Also pub-

lished as Concurrency: Practice and Experience, 10(11– 13), September 1998, CODEN CPEXEI, ISSN 1040-3108.

Schwab:1998:AJCb

[SS98b] Matthias Schwab and Joel Schroeder. Algebraic Java classes for numerical optimization. Concurrency: Practice and Experience, 10(11–13): 1155–1164, September 1998. CODEN CPEXEI. ISSN 1040-3108. URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=10050405;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=10050405&PLACEBO=IE.pdf.

Special Issue: Java for Highperformance Network Computing.

Sharma:1998:JPE

[SS98c] Rajiv Sharma and Vivek Sharma. Java programming by example. Cambridge University Press, Cambridge, UK, 1998. ISBN 0-521-64442-9. xi + 430 pp. LCCN QA76.73.J38 S455

Swildens:1998:PMA

[SS98d] Eric Swildens and Selena Sol. Programming With Microsoft AFC. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, January 1998. ISBN 0-471-24891-6. xviii + 573 pp. LCCN QA76.73.J38 S94 1998. US$50. URL http:

                                        //www.wiley.com/compbooks/
                                        catalog/24891-6.htm.

Shen:19xx:JJ

[SSxx] Chuang Thung Shen and Syu Sheng Shiung. Java (JavaScript). Computer and Information Publishing, ??, Taiwan, 19xx. ISBN ???? ???? pp. LCCN ???? 380 NT$.

Schader:1998:JEO

[SST98] M. Schader and L. Schmidt-Thieme. Java: Einführung in die objektorientierte Programmierung. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., June 1998. ISBN 3-540-63770-2. LCCN ???? 45 DM.

Swanson:1996:RTS

[ST96] S. Swanson and G. Travis. Real time streaming and Java. Object Magazine, ????(????):36– ??, 38–39, July 1996. CODEN OBMAFO. ISSN 1055-3614.

Stark:1996:JRK

[Sta96] Brian Stark. Java-Referenz: Komplette Beschreibung aller Klassen inklusive aller Parameter und Typen. (German)

[Java Reference: Complete Description of all Classes Including their Parameters and Types]. Franzis-Verlag, Feldkirchen, Germany, 1996. ISBN 3-7723-4422-4. 494 pp. LCCN ???? 49 DM.

Stanek:1997:WPU

[Sta97a] William R. Stanek. Web Publishing Unleashed: Professional Reference Edition. Howard W. Sams, Indianapolis, IN 46268, USA, January 1, 1997. ISBN 1-57521-198-X. xliii + 1501 pp. LCCN TK5105.888.S732 1997. US$69.99. URL http://

merchant.superlibrary.com:
8000/catalog/hg/PRODUCT/PAGE/
15752/bud/157521198X.html.

Starbuck:1997:JCP

[Sta97b] Orca Starbuck. JavaOne Conference Proceedings. Addison-Wesley Longman, Reading, MA, USA, 1997. ISBN 0-201-69586-3. 848 pp. LCCN ???? US$41.91.

Standish:1998:DSJ

[Sta98a] Thomas A. Standish. Data Structures in Java. Addison-Wesley, Reading, MA, USA, 1998. ISBN 0-201-30564-X (hardcover). xvii + 555 pp. LCCN QA76.73.J38 S78 1998. US$49.95. URL http:

//cseng.aw.com/bookdetail.
qry?ISBN=0%2D201%2D30564%
2DX&ptype=0;   http://www.
awl.com/cseng/titles/0-201-
30564-X.

Stanton:1998:TBT

[Sta98b] Scott Stanton. Tclblend: Blending Tcl and Java. Dr. Dobb’s Journal of Software Tools, 23 (2):50, 52–54, 100–101, February 1998. CODEN DDJOEB. ISSN 1044-789X.

Standish:19xx:SDJ

[Staxxa] Thomas A. Standish. Structures de Données en Java. International Thomson Computer Press, 20 Park Plaza Suite 1001, Boston, MA 02116 USA, 19xx. ISBN 2-84180-194-2. ???? pp. LCCN ???? 250 FF.

Stark:19xx:PJ

[Staxxb] Brian Stark. Programmieren in Java. Franzis-Verlag, Feldkirchen, Germany, 19xx. ISBN 3-7723-4432-1. LCCN ???? 49 DM.

Steele:1996:TCF

[Ste96a] Guy L. Steele, Jr. Trends & controversies: The future of Lisp: The Java programming language. IEEE expert: intelligent systems and their applications, 11(1):10–16, February 1996. CODEN IEEXE7. ISSN 0885-9000.

Stein:1996:IIG

[Ste96b] T. I. Stein, editor. IGARSS ’96: 1996 International Geoscience and Remote Sensing Symposium: remote sensing for a sustainable future. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-7803-3068-4. LCCN QE33.2.R4 I57 1996. Four volumes. IEEE catalog number 96CH35875.

Sterbenz:1996:EJS

[Ste96c] A. Sterbenz. An evaluation of the Java security model. In IEEE [IEE96h], page ?? ISBN 0-8186-7606-X. LCCN QA76.9.A25 C6375 1996. IEEE Computer Society Press Order Number PRO7606.

Stevens:1996:CPH

[Ste96d] Al Stevens. C programming:

Hairy arms for dummies, CD-ROM setups, and Java jive. Dr. Dobb’s Journal of Software Tools, 21(7):91–95, 98–99, 115–116, July 1996. CODEN DDJOEB. ISSN 1044-789X. Compares Java with C++. See rebuttal in [Kay96a].

Stevens:1996:CPHb

[Ste96e] Al Stevens. C programming:

Hairy arms for dummies, CD-ROM setups, and Java jive. Dr. Dobb’s Journal of Software Tools, 21(7):91–95, 98–99, 115–116, July 1996. CODEN DDJOEB. ISSN 1044-789X. Compares Java with C++. See comments [Kay96a].

Stewart:1996:IDW

[Ste96f] V. Stewart. Interacting with the data warehouse: applying virtual reality technology for exploratory data analysis. In Abrahart [Abr96], pages 2–?? ISBN ???? LCCN ???? Two volumes.

Stedman:1997:JDS

[Ste97a] Craig Stedman. Java database support scarce: IBM offers support, Oracle Microsoft, Sybase offer promises. Computerworld, 31(26):53, 56, June 30, 1997. CODEN CMPWAB. ISSN 0010-4841.

Stedman:1997:JGM

[Ste97b] Craig Stedman. Java gets missing link. Computerworld, 31(3): 43–44, January 20, 1997. CO-DEN CMPWAB. ISSN 0010-4841. Discusses Java links to databases, and replacement of PCs with Java-based network computers.

Stedman:1997:ODS

[Ste97c] Craig Stedman. Object Design serves Java to users. Computerworld, 31(14):3, April 7, 1997. CODEN CMPWAB. ISSN 0010-4841.

Stein:1997:PLW

[Ste97d] L. Stein. A Perl library for writing CGI scripts. Web Techniques, 2(2):61–65, February 1997. CODEN WETEFA. ISSN 1086-556X.

Stevens:1997:CPK

[Ste97e]  Al Stevens.  C program-

ming: Kicking and scripting: JavaScript and CGI. Dr. Dobb’s Journal of Software Tools, 22 (4):92–95, 97, 114, April 1997. CODEN DDJOEB. ISSN 1044-789X.

Stevens:1997:GPJ

[Ste97f] Roger T. Stevens. Graphics Programming With Java. Charles River Media, Inc., 403 VFW Drive, PO Box 417, Rockland, MA 02370, USA, 1997. ISBN 1-886801-62-2. 400 pp. LCCN ???? US$44.95. URL http://www.

charlesriver.com/titles/graphprog.
html.

Stewart:1997:WSA

[Ste97g] John N. Stewart. Web security: Active content. WebServer Magazine: For Managers of World Wide Web Sites, 2(1):20– 22, February 1997. ISSN 1087-4232. Discusses and compares Java, JavaScript, ActiveX, and VBScript.

Stewart:1997:JP

[Ste97h] Michael Kei Stewart. The Java percolates. Object Magazine, 7 (1):14, March 1997. CODEN OBMAFO. ISSN 1055-3614.

Stewart:1997:JW

[Ste97i] Michael Kei Stewart. The Java watch. Object Magazine, 7(2): 12, April 1997. CODEN OB-MAFO. ISSN 1055-3614.

Stewart:1997:JJ

[Ste97j] Michael Kei Stewart. Just Java.

Object Magazine, 1(11):12, January 1997. CODEN OBMAFO. ISSN 1055-3614.

Steele:1998:JAI

[Ste98a] G. L. Steele. Java applications and implementations. Lecture Notes in Computer Science, 1368:18–??, 1998. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Steyer:1998:JK

[Ste98b] Ralph Steyer. Java 1.2 Kompendium. Markt & Technik, Haar bei München, Germany, 1998. ISBN 3-8272-5317-9. 700 pp. LCCN ????

89,95 DM. URL http://www.
mut.ch/html/buchinfo.htx?
&x_pp=dJgxRqA9wlAYNijX&x_
az=10&x_pk=VpLvjnsx7b5V&x_
pa=B&x_pr=oS&x_bestell=25317&
x_stichwort=Steyer&x_sprache=
A&x_start=1&x_gefunden=1&
x_query=&x_lastpage=buchfund&
x_soft=&x_suchewo=A;   http:
//www.mut.ch/profit.htx?&
x_pk=kd84mf7ms9dm.

Stevens:1999:CPJ

[Ste99] Al Stevens. C programming: Java jive and scrolling the editor. Dr. Dobb’s Journal of Software Tools, 24(4):96, 98, April 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/.

Steyer:19xx:JT

[Stexxa]  R. Steyer.  Java 1.2/2.0

Taschenbuch. Markt & Technik, Haar bei München, Germany, 19xx. ISBN 3-8272-5409-4 (??invalid ISBN??). LCCN

????  20 SF.  URL

http://www.mut.ch/html/v2/
Produkt-Information.htx?&
x_pp=204%2e94%2e76%2e206VanW4xQbKi&
x_pk=Direkt-Fr&x_stichwort=

java&x_suchsprache=A&x_search=
T&x_artikelnr=25409.

Steyer:19xx:JK

[Stexxb] Ralph Steyer. Java 1.2 Kompendium. Markt & Technik, Haar bei München, Germany, 19xx. ISBN 3-8272-5317-9 (??invalid ISBN??). LCCN ????

90 DM. URL http://www.
mut.ch/html/buchinfo.htx?
&x_pp=dJgxRqA9wlAYNijX&x_
az=10&x_pk=VpLvjnsx7b5V&x_
pa=B&x_pr=oS&x_bestell=25317&
x_stichwort=Steyer&x_sprache=
A&x_start=1&x_gefunden=1&
x_query=&x_lastpage=buchfund&
x_soft=&x_suchewo=A.

Storiguard:1996:JAW

[Sto96] Victor Storiguard. Java at work: Not just for nerds anymore. WebServer Magazine: For Managers of World Wide Web Sites, 1(2):50–54, July/ August 1996. ISSN 1087-4232.

URL http://www.cpg.com.

Storer:1997:PEA

[Sto97] Larry Storer. Programmer, editor to author HP Chronicle’s JavaTalk series. HP Chronicle, 14 (7):14, June 1997. ISSN 0892-2829.

Strohmeier:1996:RST

[Str96] Alfred Strohmeier, editor. Reliable software technologies, Ada-Europe ’96: 1996 Ada-Europe International Conference on Reliable Software Technologies, Montreux, Switzerland, June 10–14, 1996: proceedings, volume 1088 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 3-540-61317-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.73.A16 A23 1996.

Strauss:1997:BRJ

[Str97] Fred J. Strauss. Book review: “Just Java”, by Peter van der Linden. UNIX Developer, 1(1): 79–80, January/February 1997. ISSN 1090-2279.

Stratton:1999:TNF

[Str99] David Stratton. Teaching network fundamentals using a simulated network. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(2):62–64, June 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Studt:1998:OOO

[Stu98] Tim Studt. Open, objectoriented software deal with drug design: CORBA-Java application framework can handle the data crunch in bioinformatics, while allowing for upgrades without having to worry about installation problems. Research & Development, 40(9):67–68, August 1998. CODEN RE-DEEA. ISSN 0746-9179.

Sullivan-Trainor:1995:IRN

[STV95] Michael Sullivan-Trainor and Richard Villars. IDC report: Netscape launches Navigator 2.0. International Data Corp. (research), page ??, September

21, 1995. URL http://www.
idcresearch.com/.

Suanj:19xx:JPZ

[Suaxx] Dario Suanj. Java: Programiranje za Internet i World Wide Web. Znak, ????, 19xx. ISBN

????  LCCN ????  199

Kn. URL http://www.znak.
tel.hr/znak/java.

Sullivan:1995:JFS

[Sul95a] Eamonn Sullivan. Java falls short of hype, but showcases technology’s potential. PC Week, page ??, November 9, 1995. ISSN 0740-

1604. URL http://www.pcmag.
ziff.com/~pcweek/reviews/
1106/tr45java.html.

Sullivan:1995:SJT

[Sul95b] Eamonn Sullivan. Sun’s Java technology perks up WWW. PC Week Online, June 5, 1995.

URL http://www.ziff.com/
~pcweek/reviews/june_1995/
java.html.

Sulger:1997:DMJ

[Sul97] Art Sulger. Database management and Java. Dr. Dobb’s Journal of Software Tools, 22 (5):28, 30–32, 73, 74, May 1997. CODEN DDJOEB. ISSN 1044-789X.

Sun:1995:JVMb

[Sun95a] Sun Microsystems. Java Virtual Machine Profiling Interface (JVMPI). Sun Microsystems, 1995. URL

                                        http://java.sun.com/j2se/

1.3/docs/guide/jvmpi/. http://java.sun.com/j2se/1.3/docs/guide/jvmpi/.

Sun:1995:JVMa

[Sun95b] Sun Microsystems. The Java Virtual Machine Specification, 1.0 beta edition, August 1995.

                                        URL http://java.sun.com/
                                        doc/vmspec/VMSpec.ps.

SunWorld:1996:ITR

[Sun96] SunWorld Online staff. Internet Toaster: And the readers say: Yes! Sunworld Online, 02:??, 1996. ISSN 1091-8914.

                                        URL http://www.sun.com/
                                        sunworldonline/swol-02-1996/
                                        swol-02-appliance.results.
                                        html.

Sun:1997:JEC

[Sun97a] Sun Microsystems. Java Enterprise Computing: Enabling Breakaway Business Strategies. Sun Microsystems, 2550 Garcia Avenue, Mountain View, CA 94043, USA, 1997. xvii + 202 pp.

Sun:1997:JCL

[SUN97b] SUN Microsystems, Inc., Palo Alto/CA. Java Card 2.0 Language Subset and Virtual Machine Specification, revision 1.0 final edition, October 13, 1997.

                                        URL ftp://ftp.javasoft.
                                        com/docs/javacard/JC20-Language.
                                        pdf.

Sundsted:1997:HJC

[Sun97c] Todd Sundsted. How-to Java: 3D computer graphics: Getting the hang of VRML. JavaWorld:

IDG’s magazine for the Java community, 2(8), August 1997. CODEN ???? ISSN 1091-8906.

URL http://www.javaworld.
com/javaworld/jw-08-1997/
jw-08-howto.html.

Sun:1999:JPD

[Sun99a]  Sun Microsystems.  Java

Platform Debugger Architecture (JPDA). Sun Microsystems, Mountain View, CA, USA, 1999. URL

http://java.sun.com/j2se/
1.3/docs/guide/jpda/.

Sun:1999:JLF

[Sun99b] Sun Microsystems Inc. Java Look and Feel Design Guidelines. Addison-Wesley Longman, Reading, MA, USA, 1999. ISBN 0-201-61585-1. xxviii + 230 pp. LCCN QA76.73.J38 S86 1999. US$39.95.

Sun:19xx:Ja

[Sunxxa] Sun Microsystems. Java. ASCII Corporation, Tokyo, Japan, 19xx. ISBN 4-7561-1664-7. LCCN ???? ???? URL http:/

/www.ascii.co.jp/pb/book1/
shinkan/detail/1322200.html.

Sun:19xx:Jb

[Sunxxb] Sun Microsystems. Java. ASCII Corporation, Tokyo, Japan, 19xx. ISBN 4-7561-1688-4. LCCN ???? ???? URL http:/

/www.ascii.co.jp/pb/book1/
shinkan/detail/1322245.html.

SunSoft:19xx:JEC

[Sunxxc] SunSoft. Java Enterprise Computing: Enabling Breakaway Business Strategies. Sun Microsystems, 2550 Garcia Avenue, Mountain View, CA 94043, USA, 19xx. ISBN ???? ???? pp. LCCN ???? ????

Susanj:1997:JPZ

[Suš97] Dario Sušanj. Java: Programiranje za Internet i World Wide Web. Znak, Zagreb, Croatia, 1997. ISBN 953-189-019-6. 750 pp. LCCN ???? 199 kuna.

                                        URL http://www.znak.tel.
                                        hr/znak/java/info.html. In-

cludes CD-ROM.

Sutherland:1996:SM

[Sut96] J. Sutherland. The Smalltalk manifesto. Object Magazine, 6 (7):56–59, September 1996. CO-DEN OBMAFO. ISSN 1055-3614.

Sutherland:1997:SDJ

[Sut97] Jeff Sutherland. Software development: The Java revolution. SunExpert Magazine, 8(1): 42–44, 46, 48, 50, 52, 54, January 1997. ISSN 1053-9239.

Seemann:1998:PBD

[SvG98] Jochen Seemann and Jürgen Wolff von Gudenberg. Pattern-based design recovery of Java software. ACM SIGSOFT Software Engineering Notes, 23(6): 10–16, November 1998. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Shaio:1996:JHC

[SvJ96] Sami Shaio, Arthur van Hoff, and Herb Jellinek. Java and HotJava: a comprehensive overview. In IEEE [IEE96c], pages 424–429. CODEN DC-SIDU. ISBN 0-8186-7414-8. ISSN 1063-6390. LCCN QA76 .I49 1996. IEEE catalog number 96CB35911.

Stuurman:1998:LCM

[SvK98] Sylvia Stuurman and Jan van Katwijk. On-line change mechanisms: the software architectural level. ACM SIGSOFT Software Engineering Notes, 23 (6):80–86, November 1998. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Singh:1995:IOB

[SW95] Jai Singh and Nick Wingfield.

Interview: Oracle’s Benioff jabs at Netscape, Microsoft. InfoWorld, page ??, November 14, 1995. CODEN INWODU. ISSN 0199-6649. URL http://www.

infoworld.com/.

Smith:1996:CWW

[SW96] C. S. Smith and P. K. Wright.

CyberCut: a World Wide Web based design-to-fabrication tool. Journal of Manufacturing Systems, 15(6):432–442, ???? 1996. CODEN JMSYEB. ISSN 0278-6125.

Succi:1999:AJD

[SW99] Giancarlo Succi and Raymond W. Wong. The application of JavaCC to develop a C/C++ preprocessor. ACM SIGAPP Applied Computing Review, 7(3):11– 18, September 1999. CO-DEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/333630.333633.

Swaine:1997:PPJb

[Swa97a]  Michael Swaine.  Program-

ming paradigms: Java bashing, books, and beans. Dr. Dobb’s Journal of Software Tools, 22 (7):93–95, July 1997. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1997:PPJa

[Swa97b] Michael Swaine. Programming paradigms: Java jabs, Visi-Corpse, and the blood of Red Smith. Dr. Dobb’s Journal of Software Tools, 22(5):83, 84, 86, 89, May 1997. CODEN DDJOEB. ISSN 1044-789X.

Swaine:1997:PPS

[Swa97c] Michael Swaine. Programming paradigms: Some observations on Apple and Java. Dr. Dobb’s Journal of Software Tools, 22 (2):91–93, February 1997. CO-DEN DDJOEB. ISSN 1044-789X.

Swaine:1997:PPSa

[Swa97d] Michael Swaine. Programming paradigms: Some observations on Apple and Java. Dr. Dobb’s Journal of Software Tools, 22 (2):91–93, February 1997. CO-DEN DDJOEB. ISSN 1044-789X.

Swaine:1997:PPBa

[Swa97e]  Michael Swaine.  Program-

ming paradigms: The bridges of Santa Clara County. Dr. Dobb’s Journal of Software Tools, 22 (9):93–95, September 1997. CO-DEN DDJOEB. ISSN 1044-789X. Comments on the role of Java in the Apple’s survival struggle.

Swan:1997:TSM

[Swa97f]  Tom Swan.  Tom Swan’s

Mastering Java with Visual J++. Sams.net, Indianapolis, IN 46268, USA, 1997. ISBN 1-57521-210-2 (softcover). xxi + 658 pp. LCCN QA76.73.J38S93 1997. US$39.99. Includes CD-ROM.

Swaine:1998:PPJ

[Swa98a] Michael Swaine. Programming paradigms: Java and other weapons of war. Dr. Dobb’s Journal of Software Tools, 23 (1):107–109, January 1998. CO-DEN DDJOEB. ISSN 1044-789X.

Swaine:1998:PPSb

[Swa98b]  Michael Swaine.  Program-

ming paradigms: Sun dreams of Jini. Dr. Dobb’s Journal of Software Tools, 23(11):113– 114, 117, November 1998. CO-DEN DDJOEB. ISSN 1044-

789X. URL http://www.ddj.
com/.

Swan:19xx:MVJ

[Swaxx] Tom Swan. Mastering Visual J++. Dialectica Computer Pub., ????, 19xx. ISBN ???? ???? pp. LCCN ????

???? URL http://www.dcp.
kiev.ua/koi8/books/master/
m_vjpp.html.

Savit:1997:EJW

[SWJ97] Jeffrey Savit, Sean Wilcox, and Bhuvana Jayaraman. Enterprise Java: Where, How, When, and When Not, to Apply Java in Client/Server Environments. McGraw-Hill, New York, NY, USA, October 1997. ISBN 0-07-057991-1 (softcover). ???? pp. LCCN HF5548.5.J38S28 1997. US$44.95. URL http:

                                        //mcgraw-hill.inforonics.
                                        com/cgi/getarec?mgh41735%
                                        comp.

Siyan:1997:IJ

[SWM^{+}97] Karanjit Siyan, James L. Weaver, Jim Mathis, Luke Cassady-Dorion, and Tim Ritchey. Inside Java. New Riders Publishing, Carmel, IN, USA, March 19, 1997. ISBN 1-56205-664-6. xiv + 912 pp. LCCN QA76.73.J38S588 1997. US$35.00. URL http://

                                        merchant.superlibrary.com:
                                        8000/catalog/hg/PRODUCT/PAGE/
                                        15620/bud/1562056646.html.

Sydow:1996:IJP

[Syd96a] Dan P. Sydow. Instant Java Without Programming. I D G Books Worldwide, Indianapolis, IN, USA, September 1996. ISBN 0-7645-4008-4. ???? pp. LCCN ???? US$34.99.

Sydow:1996:JJF

[Syd96b] Dan Parks Sydow. Jumping To Java: Fast Track For C And C++ Programmers. I D G Books Worldwide, Indianapolis, IN, USA, 1996.

ISBN 0-7645-4007-6. xi + 473 pp. LCCN QA76.73.J38S95 1996. US$39.99. URL http://

www.idgbooks.com/database/
book_result.msql?uid=uidg323&
isbn=0-7645-4007-6.

Sydow:1997:IWC

[Syd97] Achim Sydow, editor. 15th IMACS World Congress on Scientific Computation, Modelling and Applied Mathematics: Berlin, August 1997: proceedings, IMACS World Congress. Wissenschaft and Technik, Berlin, Germany, 1997. ISBN 3-89685-550-6 (set), 3-89685-551-4 (vol. 1), 3-89685-552-2 (vol. 2), 3-89685-553-0 (vol. 3),3-89685-554-9 (vol. 4), 3-89685-555-7 (vol. 5), 3-89685-556-5 (vol. 6). LCCN Q183.9 .I46 1997. In cooperation with R.- P. Schafer, W. Rufeger, and H. Lehmann.

Symantec:1995:SRF

[Sym95] Symantec Corporation. Symantec releases first Java development environment for Windows 95/NT. press release, December 13, 1995. URL http:/

/www.Symantec.com/lit/dev/
javapress.html.

Syme:1999:PJT

[Sym99] D. Syme. Proving Java type soundness. Lecture Notes in Computer Science, 1523:83–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stankovic:1998:JNP

[SZ98] N. Stankovic and K. Zhang. Java and network parallel processing. Lecture Notes in Computer Science, 1497:239– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stankovic:1999:NVJ

[SZ99] N. Stankovic and K. Zhang. Native versus Java message passing. In Dongarra et al. [DLM99], pages 165–172. ISBN 3-540-66549-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 E973 1999.

Singh:1996:IWW

[SZTR96] T. Singh, M. Zhu, U. Thakkar, and U. Ravaioli. Impact of World Wide Web, Java, and virtual environments on education in computational science and engineering. In Iskander [Isk96], pages 3–?? ISBN 0-7803-3348-9 (softbound), 0-7803-3349-7 (casebound), 0-7803-3350-0 (microfiche), 0-7803-3720-4 (CD-ROM). LCCN T 62 F76 1996. Three volumes. IEEE catalog number 96CH35946.

Takagi:1997:J

[T^{+}97a] K. Takagi et al. Java. RESET Publishing, ??, Japan, 1997. ISBN ???? ???? pp. LCCN

                                        ???? URL http://www.reset.
                                        co.jp/home.html.

Tapley:1997:DOG

[T^{+}97b] Rebecca Tapley et al. Directorio Oficial De Gamelan Java. Anaya Multimedia, Madrid, Spain, 1997. ISBN 84-415-0099-1. xiv + 443 pp. LCCN ???? 3995

Ptas. URL http://www.anaya.
es/Catalogo/CGA?act=L&art=
2321008&ed=2300;   http:/
/www.anaya.es/multimedia/.

Includes CD-ROM.

Turner:1998:JAA

[T^{+}98a] E. Shane Turner et al. Java:

Activities and Applications on the Superhighway. DF-Computer Applications Ser. South-Western Educational Publishing, Cincinnati, OH, USA, 1998. ISBN 0-538-68012-1. viii + 181 pp. LCCN QA76.73.J38J3598 1997. US$19.95.

Turner:1998:JPB

[T^{+}98b] E. Shane Turner et al. Java programming basics for the Internet. South-Western Educational Publishing, Cincinnati, OH, 1998. ISBN 0-538-68012-1. ???? pp. LCCN QA76.73.J38 J379 1998.

Thomas:19xx:PEJ

[T^{+}xxa] Michael D. Thomas et al. Programando em Java para a Internet. Makron Books do Brasil Editora Ltda., Rua Tabapuã 1348, São Paulo SP - 04533-004 - Brasil, 19xx. ISBN 85-346-0689-7 (??invalid ISBN??). ???? pp. LCCN ???? ???? URL

                                        http://safe.tesla.com.br/
                                        makron/options/livro.asp?
                                        id=62.

Tung:19xx:HCJ

[T^{+}xxb] Wu Shiu Tung et al. HTML, CGI, Java, and VB Script. Run PC, ??, Taiwan, 19xx. ISBN 957-99138-0-3 (??invalid ISBN??). ???? pp. LCCN ???? 450 NT$. Includes CD-ROM.

Tyma:19xx:JPPc

[T^{+}xxc] Paul Tyma et al. Java Primer Plus. Softbank Publishing, ??, Japan, 19xx. ISBN 4-7973-0119-8. LCCN ???? 5900 yen. URL

                                        http://www.softbank.co.jp/
                                        softbank/publishing/books/
                                        editer/os1/JavaPower/index.

html. Japanese translation.

Taylor:1997:JSE

[TA97] Lee Taylor and Jennifer Atkinson. Java Studio: Effective Web Design in 3 Days. International Thomson Computer Press, 20 Park Plaza Suite 1001, Boston, MA 02116 USA, November 1997. ISBN 1-85032-923-0. ???? pp. LCCN ???? US$40.

Tabata:1996:WDL

[Tab96] K. Tabata. What’s a digital library? Joho-Shori (J. Information Processing Soc. Japan), 37(9):814–819, September 1996. CODEN JOSHA4. ISSN 0447-8053.

Taft:1996:PIA

[Taf96] Tucker Taft. Programming the Internet in Ada 95. In Strohmeier [Str96], pages 1– 16. ISBN 3-540-61317-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.73.A16 A23 1996. Describes the Intermetrics Ada 95 compiler, which generates Java Virtual Machine bytecodes directly.

Taivalsaari:1998:IJV

[Tai98] Antero Taivalsaari. Implementing a Java Virtual Machine in the Java programming language. Technical report, Sun Microsystems, Palo Alto, CA, USA, March 1998. 23 pp.

Takagi:19xx:J

[Takxx] K. Takagi. Java. RESET Publishing, ????, 19xx. ISBN 4-8103-9304-6. LCCN ????

???? URL http://www.reset.
co.jp/jb/home.html.

Tatters:1996:TYN

[Tat96] Wes Tatters. Teach yourself Netscape Web publishing in a week. Sams.net, Indianapolis, IN 46268, USA, 1996. ISBN 1-57521-068-1. xxvi + 634 pp. LCCN TK5105.883.N48 T38 1996. URL http://

merchant.superlibrary.com:
8000/catalog/hg/PRODUCT/PAGE/
15752/bud/1575210681.html.

Tauber:1996:PCH

[Tau96] M. J. (Michael J.) Tauber, editor. Proceedings of the 1996 Conference on Human Factors in Computing Systems, CHI 96: April 13–18, 1996, Vancouver, BC, Canada. ACM Press, New York, NY 10036, USA, April 1996. ISBN 0-201-94687-4 (paperback), 0-89791-777-4 (ACM). LCCN ????

Taylor:1997:JDR

[Tay97] Art Taylor. JDBC developer’s resource: database programming on the Internet. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1997. ISBN 0-13-842352-0. xx + 730 pp. LCCN QA76.625 .T39 1997.

US$50.  URL http://www.
prenhall.com/allbooks/ptr_
0138423520.html.

Taylor:1998:IPU

[Tay98] Pamela J. Taylor. Internet programming: Using NetRexx. Dr. Dobb’s Journal of Software Tools, 23(4):84–87, 101, April 1998. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ftp/1998/1998_
04/netrexx.txt.

Taylor:1999:JDR

[Tay99]  Art Taylor.  JDBC devel-

oper’s resource: database programming on the Internet. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1999. ISBN 0-13-901661-9. xx + 730 pp. LCCN QA76.625 .T39 1999. URL http:

//www.phptr.com/ptrbooks/
ptr_0139016619.html.

Taylor:19xx:OJ

[Tayxx] Taylor. Open JBuilder. International Thomson Computer Press, 20 Park Plaza Suite 1001, Boston, MA 02116 USA, 19xx. ISBN 1-85032-891-9. ???? pp. LCCN ???? US$40.

Tokmakoff:1996:SBO

[TB96] A. Tokmakoff and J. Billington. Service brokering in objectbased systems: advanced information services. In IEEE [IEE96b], page ?? ISBN 0-7803-3304-7. LCCN TK5103.4.I57 1996. IEEE catalog number 96TH8187.

Tittel:1997:DJ

[TB97] Ed Tittel and Bill Brogden.

Discover Java. IDG Books, San Mateo, CA, USA, March 1997. ISBN 0-7645-8024-8 (softcover). xx + 312 pp. LCCN QA76.73.J38 T58 1997. US$22.49. URL http://

www.idgbooks.com/database/
book_result.msql?isbn=0-7645-

8024-8. Includes CD-ROM.

Tittel:19xx:MFJ

[TBxx] Ed Tittel and Bill Brogden.

Manual fundamental de Java. Anaya Multimedia, Madrid, Spain, 19xx. ISBN 84-415-0274-9. LCCN ???? 3750

Ptas. URL http://www.anaya.
es/Catalogo/CGA?act=L&art=
2310138&ed=2300.

Tzruya:1998:PID

[TBA98] Yoav Tzruya and Mordecai Ben-Ari. A portable implementation of the Distributed Systems Annex in Java. ACM SIGADA Ada Letters, 18(6): 204–211, November/December 1998. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Tassel:1997:EEP

[TBS97] J. Tassel, B. Briscoe, and A. Smith. An end to end price-based QoS control component using reflective Java. Lecture Notes in Computer Science, 1356:18–??, 1997. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tyma:1995:JPF

[TC95] Paul Tyma and Computer Channel, Inc. Java programming fundamentals: part 1, 1995. 1 videocassette.

Taylor:1997:JPE

[TD97] Don Taylor and Jeff Duntemann. JBuilder Programming Explorer: The Best Way to Master Java Programming With Borland’s Open JBuilder. Coriolis Group Books, Scottsdale, AZ, USA, June 1997. ISBN 1-57610-005-7. 600 pp. LCCN ???? US$39.99, CDN$55.99. URL http://www.

                                        coriolis.com/.

Teague:1997:HPH

[Tea97] Jason Cranford Teague. How to Program HTML Frames: Interface Design and JavaScript. Ziff-Davis Press, Emeryville, CA, USA, 1997. ISBN 1-56276-495-0. ???? pp. LCCN ???? US$39.99.

Teeter:1999:MRC

[Tee99] Brian R. Teeter. Media reviews:

a cookbook for JavaScript: The JavaScript CD Cookbook 2/e by Erica Sadun and J. Brook Monroe (Charles River Media, 1997, $44.95, ISBN 1-886801-65-7). IEEE MultiMedia, 6(1): 90, January–March 1999. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Teig:1998:LJD

[Tei98] Oyvind Teig. Letters: Java deadlock. Dr. Dobb’s Journal of Software Tools, 23(1):10, January 1998. CODEN DDJOEB. ISSN 1044-789X.

Tennant:1995:IBG

[Ten95a] Don Tennant. Interview: Bill Gates on Microsoft’s “only sin” and other subjects. Computerworld Hong Kong, page ??, December 19, 1995.

Tennant:1995:NHE

[Ten95b] Don Tennant. News: HP eyes Sun’s Java for possible licensing deal. Computerworld Hong Kong, page ??, November 30, 1995.

Tentner:1999:PHP

[Ten99] Adrian Tentner, editor. Proceedings of the High Performance Computing Symposium — HPC’99: San Diego, California, April 11–15, 1999, Hyatt Islandia Hotel. Society for Computer Simulation, San Diego, CA, USA, 1999. ISBN 1-56555-166-4. LCCN ????

Tessier:1996:SDB

[Tes96a] Tom Tessier. Sharing data between Web page frames using JavaScript: Implementing a hidden-frame technique. Dr. Dobb’s Journal of Software Tools, 21(5):72, 74–75, 86–87, May 1996. CODEN DDJOEB. ISSN 1044-789X.

Tessier:1996:UJC

[Tes96b] Tom Tessier. Using JavaScript to create interactive Web pages. Dr. Dobb’s Journal of Software Tools, 21(3):84, 86–89, 96–97, March 1996. CODEN DDJOEB. ISSN 1044-789X.

Tittel:1995:MOI

[TG95] Ed Tittel and Mark Gaither. Mecklermedia’s official Internet world: 60 minute guide to Java. I D G Books Worldwide, Indianapolis, IN, USA, October 1, 1995. ISBN 1-56884-711-4. xxiii + 253 pp. LCCN TK 5105.888 T58 1995. US$19.99. URL http://db.

www.idgbooks.com/database/
book/isbn/generic-book.tmpl?
query=1-56884-711-4.

Tittel:1996:AJ

[TG96a] Ed Tittel and Mark Gaither. Alles Java! Einführung mit praktischen Beispielen. International Thomson Computer Press, 20 Park Plaza Suite 1001, Boston, MA 02116 USA, 1996. ISBN 3-8266-0259-5 (??invalid ISBN??). 413 pp. LCCN ???? 59 DM. URL http://

www.itp.de/;   http://www.
lanw.com/etbio.htm; http:
//www.thomson.com/intluk.

html. Includes CD-ROM.

Tittel:1996:J

[TG96b] Ed Tittel and Mark Gaither.

Java. Apogeo srl, Viale Papiniano 38, 20123 Milano, Italy, February 1996. ISBN 88-7303-165-0 (??Invalid checksum??). ???? pp. LCCN ???? 32000 Lire. URL http://www.

apogeonline.com/catalogo/
244.html;   http://www.urra.
it/244.html; http://www.urra.
it/apocat.html.

Tittel:19xx:MGJ

[TGxxa] Ed Tittel and Mark Gaither. 60 Minute Guide to Java. Softbank Publishing, ??, Japan, 19xx. ISBN 4-89052-894-6. ???? pp.

LCCN ????  ????  URL

http://www.softbank.co.jp/
softbank/publishing/books/
kikan/kikan6/003.htm.

Tittel:19xx:EJA

[TGxxb] Ed Tittel and Mark Gaither.

Einführung in Java (AKA Alles Java!). International Thomson Computer Press, 20 Park Plaza Suite 1001, Boston, MA 02116 USA, 19xx. ISBN 3-8266-0259-5 (??invalid ISBN??). LCCN ????

59 DM. URL http://www.itp.
de/progra/0259/0259.html.

Theotokis:1997:VVI

[TGGP97] Dimitrios Theotokis, George Gyftodimos, Panagiotis Georgiadis, and George Philokyprou.

VIBDaST: a virtual interactive book as a data structures teaching aid. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(3): 111–113, September 1997. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Thai:19xx:JDI

[Thaxx] Liu Pang Thai. Java Database and Interactive Programming Design. Unalis Corporation, ??, Taiwan, 19xx. ISBN 957-22-2208-2. ???? pp. LCCN ????

                                        220 NT$. URL http://www.
                                        unalis.com.tw/index.html.

DDJStaff:1998:NVO

[The98] The DDJ staff. News & views: Online database derby results; integrated ICs; Java gets the nod; HP to go; drive capacities go even higher; DSL gets a boost; school days; NC: The network car?; the bug itself; GeezerWare. Dr. Dobb’s Journal of Software Tools, 23(2): 18, February 1998. CODEN DDJOEB. ISSN 1044-789X.

Coriolis:19xx:VJP

[Thexx]  The Coriolis Group.  Vi-

sual J++ Programming EXplorer. Coriolis Group Books, Scottsdale, AZ, USA, 19xx. ISBN 1-57610-122-3. 700 pp. LCCN ???? US$39.99, CDN$55.99. URL http://www.

                                        coriolis.com/.

Thimbleby:1999:CJ

[Thi99]  Harold Thimbleby.  A cri-

tique of Java. Software— Practice and Experience, 29 (5):457–478, April 25, 1999. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=55003857;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=55003857&PLACEBO=IE.pdf.

Thomas:1995:JA

[Tho95] Wes Thomas. Java, anyone?

Web Review, page ??, September 29, 1995. URL http://

www.gnn.com/gnn/wr/sept29/
tech/safari/index.html.

Thorn:1997:PLM

[Tho97a] Tommy Thorn. Programming languages for mobile code. ACM Computing Surveys, 29(3):213– 239, September 1997. CO-DEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (elec-

tronic). URL http://www.
acm.org:80/pubs/citations/
journals/surveys/1997-29-
3/p213-thorn/.

Thorup:1997:GJV

[Tho97b] Kresten Krab Thorup. Genericity in Java with virtual types. Lecture Notes in Computer Science, 1241:444–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/
service/series/0558/bibs/
1241/12410444.htm;   http:
//link.springer-ny.com/link/

                                        service/series/0558/papers/
                                        1241/12410444.pdf.

Thompson:1998:JIP

[Tho98] Carol Thompson. Java on IA-64 — a peek into the future of Java performance on Intel’s Merced and beyond. UNIX review, 16 (11):41–48, ???? 1998. CODEN UNRED5. ISSN 0742-3136.

Thodler:1999:JDD

[Tho99] Dora Thodler. Java Database Developer’s Guide. Prima Publishing, Rocklin, CA, USA, January 1999. ISBN 0-7615-0763-9. 700 pp. LCCN ???? US$50.00.

Thomas:19xx:JPIb

[Thoxxa] Michael Thomas. Java 1.1 Programming for the Internet. Ventana Press, Chapel Hill, NC, USA, second edition, 19xx. ISBN 1-56604-746-3. ???? pp. LCCN ???? ????

Thomas:19xx:PJP

[Thoxxb] Michael Thomas. Programación en Java para Internet. Editorial Paraninfo, Madrid, Spain, 19xx. ISBN 84-283-2394-1. LCCN ???? 6726 Ptas.

Thomas:19xx:PJ

[Thoxxc] Michael Thomas. Programmeren in Java. Sybex Nederland, ????, 19xx. ISBN 90-419-0018-7. LCCN ???? fl.

                                        89. URL http://www.sybex.
                                        nl/catalogus/programmeren/
                                        progjava.htm.

Tichenor:1997:JCB

[Tic97] Charles B. Tichenor. JavaScript cookies: Browser-side persistence for Web developers. Dr. Dobb’s Journal of Software Tools, 22(5):42–45, May 1997. CODEN DDJOEB. ISSN 1044-789X.

Tilli:1997:DJ

[Til97a] Thomas Tilli. Datenbankentwicklung mit Java. Franzis-Verlag, Feldkirchen, Germany, September 1997. ISBN 3-7723-6033-5. ???? pp. LCCN ???? 49 DM.

Tilli:1997:SSJb

[Til97b] Thomas Tilli. Suchen und sortieren in Java. Franzis-Verlag, Feldkirchen, Germany, 1997. ISBN 3-7723-7604-

1. LCCN ???? 49 DM.

URL http://www.franzis-
buch.de/katalog/java-suchensortieren/

.

Tilli:1997:SSJa

[Til97c] Thomas Tilli. Suchen und sortieren in Java ab JDK-Version 1.0. Franzis-Verlag, Feldkirchen, Germany, 1997. ISBN 3-7723-6004-1. 192 pp. LCCN ???? 49 DM.

URL http://www.franzis-
buch.de/katalog/java-suchensortieren/

.

Tisato:1997:JRI

[Tis97] Nigro Tisato. Java Restaurant: Introduzione alla programmazione concorrente in Java. Apogeo srl, Viale Papiniano 38, 20123 Milano, Italy, 1997. ISBN 88-7303-350-4. 224 pp. LCCN ???? 28,000 Lire. URL http://www.

                                        apogeonline.com/catalogo/
                                        350.html;   http://www.urra.
                                        it/350.html; http://www.urra.
                                        it/apocat.html;   http://
                                        www.urra.it/cataapo.html.

Includes floppy disk.

Tittel:1996:JS

[Tit96a] Ed Tittel. Java Secrets. I D G Books Worldwide, Indianapolis, IN, USA, October 1996. ISBN 0-614-20273-6. ???? pp. LCCN ???? US$44.99.

Tittel:1996:OIW

[Tit96b] Ed Tittel. The Official Internet World 60 Minute Guide to Java. I D G Books Worldwide, Indianapolis, IN, USA, second edition, May 1996. ISBN 0-7645-8006-X. ???? pp. LCCN ???? US$24.99.

Tittel:1996:UJ

[Tit96c]  Ed Tittel.  Undocumented

Java. I D G Books Worldwide, Indianapolis, IN, USA, October 1996. ISBN 0-7645-8007-8. xxvi + 885 pp. LCCN QA76.73.J38H373 1997. US$44.99.

Tittel:1997:VCD

[Tit97] Ed Tittel. Visual Café for dummies. IDG Books, San Mateo, CA, USA, 1997. ISBN 0-7645-0077-5. ???? pp. LCCN ???? US$25.

Tittel:19xx:JSa

[Titxxa] Ed Tittel. Java Secrets. Dialectica Computer Pub., ????, 19xx. ISBN ???? ???? pp. LCCN ????

URL http://www.dcp.kiev.
ua/koi8/books/secrets/s_java.
html.

Tittel:19xx:JSb

[Titxxb] Ed Tittel. Java Secrets. Dialectica Computer Pub., ????, 19xx. LCCN ???? ????

URL http://www.dcp.kiev.
ua/koi8/books/secrets/s_java.
html.

Toupin:1996:UBC

[TJ96] Ed Toupin and Russ Jacobs.

Using Borland C++ 5, Special Edition: Develop Sophisticated Applications For Windows 95 And Windows NT, Write Java Code With The Java Development Kit. Que Corporation, Indianapolis, IN, USA, 1996. ISBN 0-7897-0284-3 (softcover). xxvii + 870 pp. LCCN QA76.73.C153K56 1996. US$44.99. Includes CD-ROM.

Taylor:1998:CJW

[TK98] Christopher Taylor and Tim Kimmet. Core Java Web Server. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1998. ISBN 0-13-080559-9. 608 pp. LCCN ????

URL http://www.phptr.com/
ptrbooks/ptr_0130805599.html.

Includes CD-ROM.

Tai:1999:AP

[TK99a] Hideki Tai and Kazuya Kosaka.

The Aglets project. Communications of the ACM, 42(3): 100–101, March 1999. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.
acm.org:80/pubs/citations/
journals/cacm/1999-42-3/p100-
tai/.

Taylor:1999:CJW

[TK99b] Christopher Taylor and Tim Kimmett. Core Java Web Server. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1999. ISBN 0-13-080559-9. xxii + 593 pp. LCCN TK5105.888.T39 1999. US$49.99.

Thirunarayan:1999:SMI

[TKH99] Krishnaprasad Thirunarayan, Günter Kniesel, and Haripriyan Hampapuram. Simulating multiple inheritance and generics in Java. Computer Languages, 25(4):189–210, December 1999. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/15/
18/28/30/24/abstract.html.

Tip:1999:PEA

[TLSS99] Frank Tip, Chris Laffra, Peter F. Sweeney, and David Streeter. Practical experience with an application extractor for Java. ACM SIGPLAN Notices, 34(10):292–305, October 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

URL http://www.acm.org/

pubs/citations/proceedings/
oops/320384/p292-tip/.

Tryggvesson:1999:JJR

[TMH99] Jörgen Tryggvesson, Torbjörn Mattsson, and Hansruedi Heeb. Jbed: Java for real-time systems. Dr. Dobb’s Journal of Software Tools, 24(11):78, 80, 82–84, 86, November 1999. CO-DEN DDJOEB. ISSN 1044-

789X. URL http://www.ddj.
com/ftp/1999/1999_11/jbed.
txt;   http://www.ddj.com/
ftp/1999/1999_11/jbed.zip.

Tsirikos:1998:CSD

[TMM^{+}98] D. Tsirikos, T. Markousis, I. Mouroulis, M. Hatzopoulos, M. Vazirgiannis, and Y. Stavrakas. A client-server design for interactive multimedia documents based on Java. Lecture Notes in Computer Science, 1483:248–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/
service/series/0558/bibs/
1483/14830248.htm;   http:
//link.springer-ny.com/link/
service/series/0558/papers/
1483/14830248.pdf.

Tsirikos:1999:CDI

[TMMH99] D. Tsirikos, T. Markousis, Y. Mouroulis, and M. Hatzopoulos. A client-server design for interactive multimedia documents based on Java. Lecture Notes in Computer Science, 1483:248–??, 1999. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Takagi:1998:NMPa

[TMN^{+}98a] Hiromitsu Takagi, Satoshi Matsuoka, Hidemoto Nakada, Satoshi Sekiguchi, Mitsuhisa Satoh, and Umpei Nagashima. Ninflet: a migratable parallel objects framework using Java. In ACM [ACM98a], page ?? CODEN CPEXEI. ISSN 1040-3108. LCCN ????

URL http://www.cs.ucsb.
edu/conferences/java98/papers/
ninflet.pdf;   http://www.
cs.ucsb.edu/conferences/java98/
papers/ninflet.ps. Also pub-

lished as Concurrency: Practice and Experience, 10(11– 13), September 1998, CODEN CPEXEI, ISSN 1040-3108.

Takagi:1998:NMPb

[TMN^{+}98b] Hiromitsu Takagi, Satoshi Matsuoka, Hidemoto Nakada, Satoshi Sekiguchi, Mitsuhisa Satoh, and Umpei Nagashima. Ninflet: a migratable parallel objects framework using Java. Concurrency: Practice and Experience, 10(11–13): 1063–1078, September 1998. CODEN CPEXEI. ISSN 1040-3108. URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=10050423;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=10050423&PLACEBO=IE.pdf.

Special Issue: Java for Highperformance Network Computing.

Tremblett:1999:LJP

[TMP^{+}99] Paull Tremblett, A. P. Madden, Serguei Patchkovskii, Edd Dumbill, Michael Olson, Tim Kientzle, and Andrew Bowley. Letters: Java provided update; better late than never; CD authoring; full-text searching; taming C++. Dr. Dobb’s Journal of Software Tools, 24(5): 10, 12, May 1999. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/.

Tremblay:1996:PHI

[TO96] Marc Tremblay and Michael O’Connor. PicoJava: a hardware implementation of the Java Virtual Machine. In IEEE [IEE96e], pages 131–144. ISBN ???? LCCN ????

Tobochnik:1999:TSW

[Tob99] Jan Tobochnik. Teaching students to write computer simulations in Java. Computer Physics Communications, 121–122:562– 568, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0010465599004075.

Tolksdorf:1997:BOO

[Tol97a] R. Tolksdorf. Berlinda: An object-oriented platform for implementing coordination languages in Java. Lecture Notes in Computer Science, 1282:430– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tolksdorf:1997:BOP

[Tol97b] R. Tolksdorf. Berlinda: An object-oriented platform for implementing coordination languages in java. Lecture Notes in Computer Science, 1282:430– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tolksdorf:1998:PLJ

[Tol98] Robert Tolksdorf. Programming languages for the Java Virtual Machine. Technical report, Technische Universität Berlin, Fachbereich 13, Informatik, Formale Methoden, Logik und Programmierung (FLP), Sekr. FR 6-10, Franklinstraße 28/29, D-10587 Berlin, Germany, 1998. URL http:

                                        //grunge.cs.tu-berlin.de/~
                                        tolk/vmlanguages.html. World-

Wide Web document with pointers to more than 60 compilers and translators between various programming languages and Java.

Tomasula:1996:JPT

[Tom96] D. Tomasula. Java percolates through Wall Street. Wall Street and Technology, 14(3): 54–??, 56, March 1996. CODEN WSTEE5. ISSN 1060-989X.

Topley:1998:CJF

[Top98] Kim Topley. Core Java Foundation Classes (JFC). P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, May 1998. ISBN 0-13-080301-4. xxxv + 1168 pp. LCCN QA76.73.J38 T67 1998.

US$49.95. URL http://www.

prenhall.com/ptrbooks/ptr_
0130803014.html.  Includes

CD-ROM.

Topley:1999:CJF

[Top99] Kim Topley. Core Java Foundation Classes. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1999. ISBN 0-13-080301-4. xxxv + 1168 pp. LCCN QA76.73.J38T67 1998. US$49.99.

Torok:1996:JPPa

[Tor96]  Gabriel Torok.  JavaScript

Primer Plus. Macmillan Computer Publishing, Indianapolis, IN, USA, September 1996. ISBN 0-614-20295-7. ???? pp. LCCN ???? US$39.99.

Jean-Fran:1999:HTC

[Tou99] Jean-François Touchette. HTML thin client and transactions. Dr. Dobb’s Journal of Software Tools, 24(10):80, 82, 84–86, October 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ftp/1999/1999_
10/htmlthin.txt;   http://
www.ddj.com/ftp/1999/1999_
10/htmlthin.zip.

TowerTechnology:1998:TJB

[Tow98] Tower Technology. TowerJ 2.0 Java bytecode compiler, 1998.

URL http://www.twr.com/
java/java-products.html. The

(commercial) TowerJ compiler translates Java source and Java bytecodes to Java bytecodes, native executables, or native shared libraries, for Windows NT/95, SPARC Solaris and SunOS, IBM AIX, HP/UX on PA/RISC, SGI IRIX and Linux.

Thomas:1996:JPI

[TPHB96] Michael D. Thomas, Pratik R. Patel, Alan D. Hudson, and Donald A. Ball, Jr. Java Programming for the Internet: a Guide to Creating Dynamic, Interactive Internet Applications. Ventana Communications Group, Incorporated, Research Triangle Park, NC, USA, 1996. ISBN 1-56604-355-7. xxiv + 780 pp. LCCN QA76.73.J38 J38 1996. US$49.95. URL http:/

/www.course.com/templates/
catalog/detail.cfm?isbn=1-
56604-355-7;   http://www.
vmedia.com/cat/press/store/
java/.

Thomas:19xx:JIa

[TPHBxxa] Michael D. Thomas, Pratik R. Patel, Alan D. Hudson, and Donald A. Ball, Jr. Java For The Internet. GOTOP Information Inc., 5F, No.7, Lane 50, Sec.3 Nan Kang Road Taipei, Taiwan; Unit 1905,Metro Plaza Tower 2, No.223 Hing Fong Road, Kwai Chung, N.T., Hong Kong, 19xx. ISBN 957-641-874-7. ???? pp. LCCN ???? 580

NT$. URL http://www.gotop.

com.tw. Chinese translation by Infocom of [TPHB96]. Includes CD-ROM.

Thomas:19xx:JIb

[TPHBxxb] Michael D. Thomas, Pratik R. Patel, Alan D. Hudson, and Donald A. Ball, Jr. Java For The Internet. Piter Publishing Company, St. Petersburg, Russia, 19xx. ISBN 5-88782-119-1. ???? pp. LCCN

????  URL http://www.
piter-press.ru/;   http://
www.piter-press.ru/koi/ourbooks/
comp/progr/er_pdij.html. Rus-

sian translation of [TPHB96]. Includes CD-ROM.

Torok:1996:JPPb

[TPW96] Gabriel Torok, Jeffrey Payne, and Matt Weisfeld. JavaScript Primer Plus: Enhancing Web Pages With the JavaScript Programming Language. Waite Group Press, Corte Madera, CA, USA, December 1996. ISBN 1-57169-041-7. xxviii + 588 pp. LCCN QA76.73.J39J39 1996. US$39.99.

Trager:1995:JBC

[Tra95a] Louis Trager. Java becoming cool new computer language. New York Times Syndicate, page ??, November 1, 1995. URL http://

nytsyn.com/live/Features1/
305_110195_061353_32246.html.

Trager:1995:JCL

[Tra95b] Louis Trager. Java computer language brings gee-whiz futurism. New York Times Syndicate, page ??, October 31, 1995. URL http://

nytsyn.com/live/Features1/
304_103195_083533_21012.html.

Tracz:1997:DRJ

[Tra97] Will Tracz. Developing reusable Java components. ACM SIG-SOFT Software Engineering Notes, 22(3):100–103, May 1997. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tremblett:1998:JRJ

[Tre98] Paul Tremblett. Java Reflection: Not just for tool developers. Dr. Dobb’s Journal of Software Tools, 23(1):36, 38–39, 99–101, January 1998. CODEN DDJOEB. ISSN 1044-789X.

Tremblett:1999:JPA

[Tre99a] Paul Tremblett. The Java provider architecture. Dr. Dobb’s Journal of Software Tools, 24(3):40, 42, 44–47, 49, March 1999. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/1999/9902/9902toc.
                                        htm;   http://www.ddj.com/
                                        ftp/1999/1999_03/provider.
                                        txt;   http://www.ddj.com/
                                        ftp/1999/1999_03/provider.
                                        zip.

Tremblett:1999:JQW

[Tre99b] Paul Tremblett. Java Q&A: What are JavaServer pages? Dr. Dobb’s Journal of Software Tools, 24(12):109–113, 115, December 1999. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/1999/1999_12/
                                        jqa1299.txt;   http://www.

ddj.com/ftp/1999/1999_12/
jqa1299.zip.

Trinkle:1995:NHK

[Tri95] Veronika Trinkle. News: Hong Kong University brews Chinese version of HotJava. Computerworld Hong Kong, page ??, October 04, 1995.

Tamiosso:1997:BIA

[TRMR97] F. Tamiosso, A. Raposo, L. Magalhaes, and I. Ricarte. Building interactive animations using VRML and Java. In de Figueiredo and Netto [dN97], pages 42–48. ISBN 0-8186-8102-0, 0-8186-8103-9, 0-8186-8104-7. LCCN T385 .S59 1997. IEEE Computer Society order number PR08102. IEEE order plan catalog number 97TB100164.

Troan:1997:FSS

[Tro97] Eric Troan. Free software solutions: From Python to Java. UNIX Developer, 1(1):77–78, January/February 1997. ISSN 1090-2279.

Tapley:1996:OGJ

[TSHH96] Rebecca Tapley, Nova Spivak, Jack D. Hidary, and Murray Hidary. The Official Gamelan Java Directory. Ziff-Davis Press, Emeryville, CA, USA, October 1996. ISBN 1-56276-449-7. xv + 244 pp. LCCN ???? US$29.99, CDN$42.95, UK£27.95. URL http:/

/www.mcp.com/cgi-bin/bag?
isbn=1-56276-449-7&last=/
bookstore.

Turpeinen:1996:AAP

[TSKA^{+}96] M. Turpeinen, J. Saarela, M. Korkea-Aho, T. Puskala, and R. Sulonen. Architecture for agent-mediated personalised news services. In Anonymous [Ano96-154], pages 615–

628. ISBN ???? LCCN ????

Tyma:1996:JPPb

[TTD96a] Paul Tyma, Gabriel Torok, and Troy Downing. Java Primer Plus. DaeRim Publishing, Seoul, Korea, May 1996. ISBN 89-7280-301-4 (??Invalid checksum??). ???? pp. LCCN ???? 23,000 Won. URL

http://www.drbook.co.kr/;
 http://www.interpia.net/~

yoonforh/. Korean translation by Yoon Kyung Koo of [TTD96b].

Tyma:1996:JPPa

[TTD96b] Paul Tyma, Gabriel Torok, and Troy Downing. Java primer plus: supercharging WEB applications with the Java programming language. Waite Group Press, Corte Madera, CA, USA, 1996. ISBN 1-57169-062-X. xxv + 612 pp. LCCN QA76.64 .T96 1996. US$39.99; 80,00 DM. URL http:/

/www.mcp.com/waite/books.
new/Java_Primer_Plus/html/
jppcov.html;   http://www.
mut.ch/profit.htx?&x_pk=kd84mf7ms9dm;
 http://www.preemptive.com/
javap.html;   http://www.
preemptive.com/{javap}.html.

Includes CD-ROM.

Tyma:1997:JPPa

[TTD97] Paul Tyma, Gabriel Torok, and Troy Downing. Java Primer Plus. Softbank Publishing, ??, Japan, 1997. ISBN ???? ???? pp. LCCN ???? 5900 yen. URL

http://www.preemptive.com/
aboutauth.html;   http://
www.softbank.co.jp/softbank/
publishing/;   http://www.
softbank.co.jp/softbank/publishing/
books/editer/os1/JavaPower/
index.html.

Tyma:19xx:JPPb

[TTDxx] Paul Tyma, Gabriel Torok, and Troy Downing. Java Primer Plus. Waite Group Press, Corte Madera, CA, USA, 19xx. ISBN 957-99224-9-7 (??invalid ISBN??). ???? pp. LCCN ????

580 NT$. URL http://www.
preemptive.com/aboutauth.
html;   http://www.waite.

com/waite/. Chinese translation by Chiang Yung Hsiang of [TTD96b]. Includes CD-ROM.

Thiruvathukal:1998:RRMb

[TTK98a] G. K. Thiruvathukal, L. S.

Thomas, and A. T. Korczynski. Reflective remote method invocation. Concurrency: Practice and Experience, 10(11–13):911–925, September 1998. CODEN CPEXEI. ISSN 1040-3108. URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=10050398;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=10050398&PLACEBO=IE.pdf.

Special Issue: Java for Highperformance Network Computing.

Thiruvathukal:1998:RRMa

[TTK98b] George K. Thiruvathukal, Lovely S. Thomas, and Andy T. Korczynski. Reflective Remote Method Invocation. In ACM [ACM98a], page ?? CODEN CPEXEI. ISSN 1040-3108. LCCN ????

                                        URL http://www.cs.ucsb.
                                        edu/conferences/java98/papers/
                                        reflective.pdf;   http://
                                        www.cs.ucsb.edu/conferences/
                                        java98/papers/reflective.

ps. Also published as Concurrency: Practice and Experience, 10(11–13), September 1998, CODEN CPEXEI, ISSN 1040-3108.

Tubbs:1995:PFD

[Tub95] Dave Tubbs. Pop, fizz, and dance — Java and VRML are new graphics technologies that will alter the state of the Net. Internet World, 6(11):72, ???? 1995. CODEN IERNE8. ISSN 1064-3923.

Tucker:1996:STJ

[Tuc96] Michael J. Tucker. SunSoft takes Java to net management. WebServer Magazine: For Managers of World Wide Web Sites, 1(2):18, 20, July/ August 1996. ISSN 1087-4232.

                                        URL http://www.cpg.com.

Tuck:1998:TYJ

[Tuc98a] Devon Tuck. Teach Yourself Java 1.1 in 21 Days, Sun Certification Exam Preparation Guide. Sams.net, Indianapolis, IN 46268, USA, January 1998. ISBN 1-57521-365-6 (softcover). LCCN ???? US$50.

Tucker:1998:RJC

[Tuc98b] Andrew Tucker. Reading Java class files in C++. C/C++ Users Journal, 16(4):??, April 1998. CODEN CCUJEX. ISSN 1075-2838.

Tunkelang:1998:JJI

[Tun98] D. Tunkelang. JIGGLE: Java Interactive General Graph Layout Environment. Lecture Notes in Computer Science, 1547:413– 422, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Turk:1997:BBI

[Tur97] Andy Turk. Building a better interface with Java: Internet Foundation Classes lets you build Java applets with useful features such as drag and drop. BYTE Magazine, 22(8):45–46, August 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Tygar:1996:WEC

[TW96] J. D. Tygar and Alma Whitten. WWW electronic commerce and Java Trojan horses. In USENIX [USE96b], pages 243–250. ISBN ???? LCCN ????

Tessem:1998:RJC

[TWP98] B. Tessem, R. A. Whitehurst, and C. L. Powell. Retrieval of Java classes for case-based reuse. Lecture Notes in Computer Science, 1488:148–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tyma:1996:TJP

[Tym96] Paul Tyma. Tuning Java performance. Dr. Dobb’s Journal of Software Tools, 21(4):52, 55– 56, 58, 90, April 1996. CODEN DDJOEB. ISSN 1044-789X.

Tyma:1998:WWU

[Tym98] Paul Tyma. Why are we using Java again? Communications of the ACM, 41(6):38–42, June 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org:80/pubs/citations/
journals/cacm/1998-41-6/p38-
tyma/.

Tyma:1999:TVC

[Tým99] Paul Týma. Transient variable caching in Java’s stack-based intermediate representation. Scientific Programming, 7(2):157– 166, 1999. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:

//iospress.metapress.com/
app/home/contribution.asp%
3Fwasp=f277qlrwwjr5m4vxjywv%
26referrer=parent%26backto=
issue%2C7%2C8%3Bjournal%2C7%
2C9%3Blinkingpublicationresults%
2C1%2C1.

Yih:19xx:JCA

[tYpCxx] Wen tau Yih and Ying ping Chen. Java: Creating Animation Homepage. Flag Publishing, ??, Taiwan, 19xx. ISBN 957-717-192-3 (??invalid ISBN??). ???? pp. LCCN ???? 580 NT$.

Takata:1996:TMI

[TYY96] S. Takata, C. Yamashita, and S. Yamamoto. TEA: multimedia information system architecture technology. NTT R and D, 45(8):745–752, ???? 1996. CODEN NTTDEC. ISSN 0915-2326.

Ungar:1998:PNC

[UBF^{+}98] David Ungar, Lars Bak, Jesse Fang, John Duimovich, and Scott Meyer. Panel 2: The new crop of Java Virtual Machines. ACM SIGPLAN Notices, 33(10):179–182, October 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ubois:1996:JJA

[Ubo96a] J. Ubois. Java: just another language, or a whole new paradigm? Digital Media, 6 (2):3–6, July–August 1996. CO-DEN DMEDEG. ISSN 1056-7038.

Ubois:1996:WMS

[Ubo96b] Jeff Ubois. Wave Maker Starwave VP Patrick Naughton — who helped create Java at Sun — offers some strong opinions on Java, Visual Basic, VRML, Internet over cable, and more. Internet World, 7(3): 96–??, ???? 1996. CODEN IERNE8. ISSN 1064-3923.

Ulrich:1999:IWT

[UCD99] Andreas Ulrich and Gunther Chrobok-Diening. International Workshop on Testing Distributed Component-Based Systems. ACM SIGSOFT Software Engineering Notes, 24(4):43–46, July 1999. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Udell:1996:YBN

[Ude96] Jon Udell. Your business needs the Web. BYTE Magazine, 21 (8):68–70, 72, 74, 76, 78, 80, August 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic). Discusses NASA’s distributed data visualization using Java.

Udell:1997:WPJa

[Ude97a]  John Udell.  Web project:

JavaScript revisited. BYTE Magazine, 22(5):99, 100, 102, May 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1997:WPP

[Ude97b] John Udell. Web project: Persistent Java: a servlet-based group calendar becomes a surprising success and prompts an exploration of ways to bind Java programs to persistent storage. BYTE Magazine, 22(8): 109, 110, 112, August 1997. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1997:NAW

[Ude97c] Jon Udell. Net applications:

Will Netscape set the standard? BYTE Magazine, 22 (3):66–69, 72, 74, 76, 80, 82, 84, March 1997. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic). Discusses Java, Javascript, and the Java Runtime Environment.

Udell:1997:WPJb

[Ude97d] Jon Udell. Web project: Java servlets: Servlets, the Java equivalent of CGI applications, can deliver on many of Java’s promises while dodging some of its worst limitations. BYTE Magazine, 22(6):115, 116, 118, June 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Uimonen:1995:NTE

[Uim95] Terho Uimonen. News: Taiwan eyes Java, ready to jump on “Internet appliance” bandwagon. IDG International News Service, page ??, December 14, 1995.

Umar:1997:OOC

[Uma97] Amjad Umar. Object-Oriented Client/Server Internet Environments. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1997. ISBN 0-13-375544-4. xxi + 527 pp. LCCN QA76.9.C55U53 1997. US$49.00. Includes some discussion of Java.

Ung:1997:CCJ

[Ung97] Gordon Mah Ung. Corel cans Java Office project: Pitches an ‘evolved’ product, promises a Java-based upgrade. Computerworld, 31(34):14, August 25, 1997. CODEN CMPWAB. ISSN 0010-4841.

UC:1996:PCT

[Uni96a] Unicode Consortium, editor. Pre-conference tutorials proceedings: Software development + the Internet: going global with Unicode: Ninth International Unicode Conference, San Jose, CA, September 4–6 1996. The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-

5082, E-mail: unicode-inc@

unicode.org, 1996. ISBN ???? LCCN ???? A1. Input method design / by Mark Leisher – A2. Weaving the multilingual web : standards and their implementations / by Martin Durst … [et al.] – B1. National language and Unicode support in relational databases and SQL2/3 / by Stefan Buchta – B2. The Unicode Standard : version 2 / by Asmus Freytag – C1/C2. Non-Latin writing systems : characteristics and impacts on multinational product design / by Richard Ishida.

UC:1996:SDI

[Uni96b] Unicode Consortium, editor. Software Development + the Internet: Going Global with Unicode: Ninth International Unicode Conference, 4–6 September, San Jose, California, USA. The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-5082, E-mail: unicode-

inc@unicode.org, 1996. ISBN

???? LCCN ???? URL http:

//www.unicode.org/unicode/
iuc9conf/index.html;   http:
//www.unicode.org/unicode/
uni2book/u2ord.html.

UC:1997:ESI

[Uni97a] Unicode Consortium, editor.

Europe, Software + the Internet: Going Global with Unicode: Tenth International Unicode Conference, March 10–12, 1997, Mainz, Germany. The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-

5082, E-mail: unicode-inc@

unicode.org, 1997. ISBN ????

LCCN ???? URL http://www.
unicode.org/unicode/iuc10/
;   http://www.unicode.org/
unicode/uni2book/u2ord.html.

UC:1997:SDI

[Uni97b] Unicode Consortium, editor.

Software Development + the Internet: Going Global with Unicode: The Eleventh International Unicode Conference, September 2–5, 1997, San Jose, California, USA. The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-

                                        5082, E-mail: unicode-inc@

unicode.org, 1997. ISBN ???? LCCN ???? URL http:

                                        //www.unicode.org/unicode/
                                        iuc11/index.html;   http:
                                        //www.unicode.org/unicode/
                                        uni2book/u2ord.html.

UC:1998:ASI

[Uni98a] Unicode Consortium, editor. Asia, Software + the Internet: Going Global with Unicode: The Twelfth International Unicode/ ISO10646 Conference (IUC12) April 8–10, 1998, Tokyo, Japan. The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-

                                        5082, E-mail: unicode-inc@

unicode.org, 1998. ISBN ???? LCCN ???? URL http:

                                        //www.unicode.org/unicode/
                                        iuc12/index.html;   http:
                                        //www.unicode.org/unicode/
                                        uni2book/u2ord.html.

UC:1998:TIU

[Uni98b] Unicode Consortium, editor. Thirteenth International Unicode Conference: Software + the Internet: Going Global with Unicode (R), September 8–11, 1998, San Jose, California. The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-

                                        5082, E-mail: unicode-inc@

unicode.org, 1998. ISBN ????

                                        LCCN ???? URL http://www.
                                        reuters.com/unicode/iuc13;
                                         http://www.unicode.org.

Urgo:1996:JFH

[Urg96a] M. Urgo. Java — the future is here again. Business Information Review, 13(2):107–108, June 1996. CODEN BIREEY. ISSN 0266-3821.

Urgo:1996:JSW

[Urg96b] Marisa Urgo. The Java Source:

The Worldwide Guide to Java Companies, Products, Services, and Books. SIGS Books, 71 W 23rd St., New York, NY 10010, USA, 1996. ISBN ???? ???? pp. LCCN

???? URL http://www.sigs.
com/books/worldjava.html.

Urgo:19xx:JS

[Urgxx] Marissa Urgo. The Java Source.

SIGS Books, 71 W 23rd St., New York, NY 10010, USA, 19xx. LCCN ???? ????

Urquhart:1997:JET

[Urq97a] K. Urquhart. Java and the emerging technology. Lecture Notes in Computer Science, 1274:1–??, 1997. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Urquhart:1997:GEI

[Urq97b]  Ken Urquhart.  Guest ed-

itor’s introduction: Java’s open future. IEEE Micro, 17(3):10–13, May/June 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//pascal.computer.org/mi/
books/mi1997/pdf/m3010.pdf.

USENIX:1996:SOS

[USE96a] USENIX, editor. 2nd Symposium on Operating Systems Design and Implementation (OSDI ’96), October 28–31, 1996. Seattle, WA. USENIX, Berkeley, CA, USA, October 28–31, 1996. ISBN ???? LCCN ????

USENIX:1996:WEC

[USE96b] USENIX, editor. 2nd Workshop on Electronic Commerce, November 18–21, 1996. Oakland, CA. USENIX, Berkeley, CA, USA, November 18–21, 1996. ISBN ???? LCCN ????

USENIX:1996:PSUb

[USE96c] USENIX, editor. Proceedings of the Second USENIX Conference on Object-Oriented Technologies and Systems (COOTS), June 17–21, 1996, Toronto, Canada. USENIX, Berkeley, CA, USA, 1996. ISBN 1-880446-77-4. LCCN QA76.64 .U85 1996. URL http://www.

usenix.org/publications/library/
proceedings/coots96/.

USENIX:1996:PUA

[USE96d] USENIX Association, editor. Proceedings of the USENIX 1996 annual technical conference: January 22–26, 1996, San Diego, California, USA, USENIX Conference Proceedings 1996. USENIX, Berkeley, CA, USA, 1996. ISBN 1-880446-76-6. LCCN QA 76.76 O63 U88 1996.

USENIX:1997:ATT

[USE97a]  USENIX, editor.  5th An-

nual Tcl/Tk Workshop ’97, July 14–17, 1997. Boston, MA. USENIX, Berkeley, CA, USA, 1997. ISBN 1-880446-87-1. LCCN QA76.73.T44 T44 1997.

USENIX:1997:TUC

[USE97b] USENIX, editor. The Third USENIX Conference on Object-Oriented Technologies and Systems (COOTS), June 16– 19, 1997. Portland, Oregon. USENIX, Berkeley, CA, USA, 1997. ISBN 1-880446-86-3. LCCN QA76.64 .U845 1997.

URL http://dxsting.cern.
ch/sting/COOTS97.html.

USENIX:1997:USI

[USE97c] USENIX, editor. USENIX Symposium on Internet Technologies and Systems Proceedings, Monterey, California, December 8–11, 1997. USENIX, Berkeley, CA, USA, 1997. ISBN 1-880446-91-X. LCCN TK5105.875.I57

U96 1997. URL http://www.
usenix.org/publications/library/
proceedings/usits97/.

USENIX:1998:PUWb

[USE98a] USENIX, editor. Proceedings of the 3rd USENIX Workshop on Electronic Commerce: August 31–September 3, 1998, Boston, Mass. USENIX, Berkeley, CA, USA, 1998. ISBN 1-880446-97-9. LCCN HF5004 .U74

1998. URL http://db.usenix.
org/publications/library/
proceedings/ec98.

USENIX:1998:PFU

[USE98b] USENIX, editor. Proceedings of the fourth USENIX Conference on Object-Oriented Technologies and Systems (COOTS): April 27–30, 1998, Santa Fe, NM. USENIX, Berkeley, CA, USA, 1998. ISBN ???? LCCN

                                        ???? URL http://db.usenix.
                                        org/publications/library/
                                        proceedings/coots98.

USENIX:1998:PFT

[USE98c] USENIX, editor. Proceedings of the FreeNIX Track: USENIX 1998 annual technical conference: June 15–19, 1998, New Orleans, LA. USENIX, Berkeley, CA, USA, 1998. ISBN ????

                                        LCCN ???? URL http://www.
                                        usenix.org/publications/library/
                                        proceedings/usenix98/freenix/

.

USENIX:1998:PSA

[USE98d] USENIX, editor. Proceedings of the sixth annual Tcl/Tk Conference, September 18–24 [i.e. 14–18], 1998, San Diego, California. USENIX, Berkeley, CA, USA, 1998. ISBN 1-880446-98-7. LCCN QA76.73.T44 T34

                                        1998. URL http://db.usenix.
                                        org/publications/library/
                                        proceedings/tcl98/.

USENIX:1998:PUA

[USE98e]  USENIX, editor.  Proceed-

ings of the USENIX 1998 annual technical conference: June 15–19, 1998, New Orleans, LA. USENIX, Berkeley, CA, USA, 1998. ISBN 1-880446-94-4. LCCN QA76.76.O63 U88

1998. URL http://db.usenix.
org/publications/library/
proceedings/usenix98.

USENIX:1998:SUS

[USE98f]  USENIX, editor.  Seventh

USENIX Security Symposium proceedings: conference proceedings: San Antonio, Texas, January 26–29, 1998. USENIX, Berkeley, CA, USA, 1998. ISBN 1-880446-92-8. LCCN QA76.9.A25 U83 1998. URL

http://db.usenix.org/publications/
library/proceedings/sec98.

USENIX:1999:PCD

[USE99a]  USENIX, editor.  Proceed-

ings of the 2nd Conference on Domain-Specific Languages (DSL ’99), October 3–5, 1999, Austin, Texas, USA. USENIX, Berkeley, CA, USA, 1999. ISBN 1-880446-27-8, 1-58113-255-7. LCCN QA76.7 .C663

1999.  URL http://www.
usenix.org/publications/library/
proceedings/dsl99/.

USENIX:1999:PUS

[USE99b] USENIX, editor. Proceedings of the 2nd USENIX Symposium on Internet Technologies and Systems (USITS ’99), October 11– 14, 1999, Boulder, Colorado, USA. USENIX, Berkeley, CA, USA, 1999. ISBN 1-880446-26-X. LCCN TK5105.875.I57

U96. URL http://db.usenix.
org/publications/library/
proceedings/usits99.

USENIX:1999:PFU

[USE99c] USENIX, editor. Proceedings of the fifth USENIX Conference on Object-Oriented Technologies and Systems (COOTS ’99): May 3–7, 1999, San Diego, California, USA. USENIX, Berkeley, CA, USA, 1999. ISBN ???? LCCN QA76.64 .U84

                                        1999.  URL http://www.
                                        usenix.org/publications/library/
                                        proceedings/coots99/.

USENIX:1999:PUWa

[USE99d]  USENIX, editor.  Proceed-

ings of the USENIX Workshop on Smartcard Technology (Smartcard ’99): May 10– 11, 1999, Chicago, Illinois, USA. USENIX, Berkeley, CA, USA, 1999. ISBN 1-880446-34-0. LCCN TK7895.S62

                                        U84 1999. URL http://www.
                                        usenix.org/publications/library/
                                        proceedings/smartcard99/.

USENIX:1999:PWE

[USE99e] USENIX, editor. Proceedings of the Workshop on Embedded Systems, March 29–31, 1999, Cambridge, Massachusetts, USA. USENIX, Berkeley, CA, USA, 1999. ISBN 1-880446-31-

6. LCCN TK7895.E42 W67

                                        1999.  URL http://www.
                                        usenix.org/publications/library/
                                        proceedings/es99/.

Uzgalis:1996:HCJ

[Uzg96] Robert Uzgalis. Hashing concepts and the Java programming language. Web site,

1996. URL http://www.serve.
net/buz/hash.adt/java.000.
html;   https://bit.ly/4gjxUuA.

Vanderburg:1996:TJP

[V^{+}96] Glenn L. Vanderburg et al.

Tricks of the Java Programming Gurus. Sams.net, Indianapolis, IN 46268, USA, July 1, 1996. ISBN 1-57521-102-5. xxviii + 847 pp. LCCN QA76.73.J38V37 1996. US$39.99, CDN$56.95, UK£37.50. URL http://

merchant.superlibrary.com:
8000/catalog/hg/PRODUCT/PAGE/
15752/bud/1575211025.html;
 http://www.mcp.com/288190891680543/
cgi-bin/bag?isbn=1-57521-
102-5&last=/bookstore;   http:
//www.mcp.com/samsnet/.

Vanderburg:1997:MJ

[V^{+}97] Glenn Vanderburg et al. Maximum Java 1.1. Sams.net, Indianapolis, IN 46268, USA, 1997. ISBN 1-57521-290-

0. xxiv + 892 pp. LCCN QA76.73.J38V369 1997. 99,00 DM; US$50. URL http://

merchant.superlibrary.com:
8000/catalog/hg/PRODUCT/PAGE/
15752/bud/1575212900.html;
 http://www.mut.ch/profit.
htx?&x_pk=kd84mf7ms9dm. In-

cludes CD-ROM.

vanHoff:19xx:HJb

[v^{+}xxa] Arthur van Hoff et al. Hooked On Java. Unalis Corporation, ??, Taiwan, 19xx. ISBN 957-22-2172-8. ???? pp. LCCN ???? 400 NT$. URL http:

//www.unalis.com.tw/index.

html. Chinese translation by Cheng Tze Wen of [vSS96c]. Includes CD-ROM.

vanHoff:19xx:J

[v^{+}xxb] Arthur van Hoff et al. Java. Wydawnictwa Helion, ????, 19xx. ISBN 83-86718-30-7 (??invalid ISBN??). ???? pp. LCCN ???? 200 stron. URL

                                        http://www.helion.com.pl/
                                        ksiazki/java.htm.  Polish

translation by Tomasz Widuch.

Varela:1998:WAJ

[VA98] Carlos A. Varela and Gul A. Agha. What after Java? From objects to actors. Computer Networks and ISDN Systems, 30(1–7):573–577, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/
                                        store/comnet/sub/1998/30/
                                        1-7/1890.pdf.

Vacca:1996:JDB

[Vac96] John R. Vacca. Javascript Development: Bringing Application Development and Customization for Intranets and the Internet. AP Professional, Boston, MA, USA, 1996. ISBN 0-12-710005-9 (paperback), 0-12-710006-7 (CD-ROM). xxxi + 559 pp. LCCN QA76.73.J38V33 1997. US$34.95.

Vacca:1997:JIP

[Vac97] John R. Vacca. JavaScript: Internet-Programmierung ohne Vorwissen; konkrete Praxislosungen. Sybex Verlag GmbH, Erkrather Straße 345-349, 40231 Düsseldorf, Germany, 1997. ISBN 3-8155-2020-7. xxviii + 602 pp. LCCN ???? 59 DM, 48,- SFr, 431,- ÖS. URL http:

//www.sybex.de/;   http:/
/www.sybex.de/bestell/tp/

tp02020.htm. Translated by Dieter Winkler.

Vadlamudi:1995:NKO

[Vad95a] Pardhu Vadlamudi. News: Kodak offers software to enhance images on the Web. InfoWorld, page ??, December 16, 1995. CODEN INWODU. ISSN 0199-6649. URL http://www.

infoworld.com/.

Vadlamudi:1995:NVP

[Vad95b] Pardhu Vadlamudi. News: Vendors push tools to move existing documents on-line. InfoWorld, page ??, October 05, 1995. CODEN INWODU. ISSN 0199-6649. URL http://www.

infoworld.com/.

Valdes:1995:NGJ

[Val95] Ray Valdés. Net gets a Java buzz. Dr. Dobb’s Developer Update, 2(8):1, 3–4, 6, August 1995. CODEN ???? ISSN 1079-8595.

Valentine:1996:NAN

[Val96a] Jon L. Valentine. Netscape aims for Number ONE. Web-Server Magazine: For Managers of World Wide Web Sites, 1(3):6, 8, 9, September/October 1996. ISSN 1087-4232. Discusses JavaScript, LiveConnect, and CORBA.

Valles:1996:JJW

[Val96b] David J. Valles. A Java jukebox for the World Wide Web. Thesis (b.s.), California Polytechnic State University, San Luis Obispo, CA, USA, 1996. iv + 26 + 18 pp.

Valesky:1999:EJD

[Val99]  Thomas C. Valesky.  En-

terprise JavaBeans: developing component-based distributed applications. Addison-Wes-ley, Reading, MA, USA, 1999. ISBN 0-201-60446-9. 352 pp. LCCN QA76.73.J38V35 1999. US$39.95.

vanHoff:1995:JIP

[van95] Arthur van Hoff. Java and Internet programming. Dr. Dobb’s Journal of Software Tools, 20(8):56–??, 58, 60–61, 101–102, August 1995. CO-DEN DDJOEB. ISSN 1044-

789X. URL http://www.ddj.
com/ddj/issues/j508a.htm.

vanderLinden:1996:JJ

[van96a] Peter van der Linden. Just Java. Java series. SunSoft Press, 2550 Garcia Avenue, Mountain View, CA 94043-1100, USA, April 1996. ISBN 0-13-565839-X, 0-13-493982-4 (disc). xxiii + 354 pp. LCCN QA76.73.J38V36 1996. US$34.95. URL http://www.

sun.com/smi/ssoftpress/books/
vanderLinden2/vanderLinden2.
html.

vanHoff:1996:JLT

[van96b] Arthur van Hoff. Java language tutorial: your second jolt, 1996. 1 videocassette (83 min.).

vanVliet:1996:MD

[van96c] Hanpeter van Vliet. Mocha decompiler. Mocha turns Java bytecodes back into Java source code. The author died 31 December 1996; see

http://www.blackdown.org/
~kbs/hanpeter.html., 1996.
URL http://www.brouhaha.
com/~eric/computers/mocha-
bl.zip.

Vandendorpe:1996:SYW

[Van96d] Laura Vandendorpe. Scientist of the Year weaves Web over the world. Research & Development, 38(12):14–16, 18, November 1996. CODEN REDEEA. ISSN 0746-9179. Profile of WWW inventor Tim Berners-Lee. Also includes discussion of the Jigsaw HTTP server, written entirely in Java.

VanderVeer:1996:JD

[Van96e] Emily A. Vander Veer. JavaScript for Dummies. For Dummies Ser. I D G Books Worldwide, Indianapolis, IN, USA, November 1, 1996. ISBN 0-7645-0071-6. xx + 378 pp. LCCN QA76.73.J39V36 1996. US$24.99.

Vanderburg:1996:PJD

[Van96f] Glenn Vanderburg. Professional Java Development Kit C/Ww95/Ww. Sams.net, Indianapolis, IN 46268, USA, October 1, 1996. ISBN 1-57521-200-5. ???? pp. LCCN ???? US$89.95. URL http://

merchant.superlibrary.com:
8000/catalog/hg/PRODUCT/PAGE/
15752/bud/1575212005.html.

Vanhelsuwe:1996:MJ

[Van96g] Laurence Vanhelsuwé. Mastering Java. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, September 1996. ISBN 0-614-20326-0. ???? pp. LCCN ???? US$49.99.

vanderLinden:1997:JPH

[van97a] Peter van der Linden. Java pur: Hintergründe und Entwicklung. Verlag Heinz Heise, Helsdorfer Straße 7, D-30625, Hannover, Germany, 1997. ISBN 3-88229-086-2. 326 (or 336??) pp. LCCN ???? 89,80 DM. URL

http://www.emedia.de/bin/
bookshop?show=4564&id=;
http://www.heise.de/buch/.

German translation of [van97c]. Includes CD-ROM.

vanderLinden:1997:JJB

[van97b] Peter van der Linden. Just Java 1.1 and Beyond. SunSoft Press Java series. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, third edition, September 1997. ISBN 0-13-784174-4 (softcover). 654 pp. LCCN ???? US$40.45.

URL http://www.fdds.com/
books/books/vanderLinden2/
vanderLinden2.html;   http:
//www.prenhall.com/ptrbooks/
ptr_0137841744.html.

vanderLinden:1997:JJS

[van97c] Peter van der Linden. Just Java: Signature Edition. Java series. Prentice-Hall and Sun-Soft Press, Englewood Cliffs, NJ 07632, USA and 2550 Garcia Avenue, Mountain View, CA 94043-1100, USA, second edition, November 1997. ISBN 0-13-272303-4. xxxv + 538 pp. LCCN QA76.73.J38V36 1997. US$34.95. URL http://www.

prenhall.com/ptrbooks/ptr_
0132723034.html;   http:/
/www.sun.com/books/books/
vanderLinden2/vanderLinden2.

html. Includes CD-ROM. Available in German [van97a] and Italian [vanxxc] translations.

vanderLinden:1997:JJ

[van97d] Peter van der Linden. Not Just Java. SunSoft Press Java series. SunSoft Press, 2550 Garcia Avenue, Mountain View, CA 94043-1100, USA, 1997. ISBN 0-13-864638-4 (softcover). xii + 313 pp. LCCN QA76.73.J38V363 1997. US$31.45. URL http://www.

prenhall.com/allbooks/ptr_
0138646384.html.

vanderLinden:1997:TMG

[van97e] Peter van der Linden. A Technical Manager’s Guide to Java Computing. McGraw-Hill, New York, NY, USA, April 1997. ISBN 0-07-066941-4. 384 pp. LCCN ???? US$29.95.

vanderWal:1997:EWC

[van97f]  R. van der Wal.  Exam-

ining Win32 C++ compilers. Dr. Dobb’s Journal of Software Tools, 22(3):64–??, 66–68, 70, 72–73, March 1997. CODEN DDJOEB. ISSN 1044-789X.

VanHaecke:1997:JJD

[Van97g] Bernard Van Haecke. JDBC: Java Database Connectivity. IDG Books, San Mateo, CA, USA, 1997. ISBN 0-7645-3144-1 (softcover). xix + 385 pp. LCCN QA76.73.J38V364

1997.  URL http://www.
cbtsys.com/cbt/curicula/courses/
jav1110/jav1110.htm;   http:
//www.idgbooks.com/cgi-bin/
db/fill_out_template.pl?idgbook:
0-7645-3144-1:book-idg::uidg1985.

Includes CD-ROM.

VanderVeer:1997:JDa

[Van97h] Emily Vander Veer. JavaScript for Dummies. IDG Books, San Mateo, CA, USA, second edition, July 1997. ISBN 0-7645-0223-9 (softcover). xxiv + 375 pp. LCCN QA76.73.J39 V36 1997. US$26.99. URL http:/

/www.idgbooks.com/cgi-bin/
db/fill_out_template.pl?idgbook:
0-7645-0223-9:book-idg::uidg323.

Includes CD-ROM.

VanderVeer:1997:JDQ

[Van97i] Emily Vander Veer. JavaScript for Dummies: Quick Reference. I D G Books Worldwide, Indianapolis, IN, USA, March 1, 1997. ISBN 0-7645-0112-7. xiv + 210 pp. LCCN QA76.73.J38V368 1997. US$12.99.

VanderVeer:1997:JBD

[Van97j] Emily A. Vander Veer. Java Beans For Dummies. IDG Books, San Mateo, CA, USA, 1997. ISBN 0-7645-0153-4 (softcover). xxii + 385 pp. LCCN QA76.73.J38V367 1997. US$22.49. URL http://

www.idgbooks.com/database/
book_result.msql?uid=uidg81&
isbn=0-7645-0153-4. Forward

by Joseph H. McIntyre. Includes CD-ROM.

VanderVeer:1997:JDb

[Van97k] Emily A. Vander Veer. JavaScript für Dummies. International Thomson Computer Press, 20 Park Plaza Suite 1001, Boston, MA 02116 USA, 1997. ISBN 3-8266-2742-3 (??invalid ISBN??). 400 pp. LCCN ???? 50 DM.

URL http://www.itp.de/;
http://www.itp.de/dummies/
2742/2742.html. Includes CD-

ROM.

VanderVeer:1997:JPN

[Van97l] Emily A. Vander Veer. JavaScript pour Les Nuls. Sybex France, Paris, France, 1997. ISBN 2-7361-2375-1. ???? pp. LCCN

???? URL http://www.sybex.

fr/. French translation of [Van97i]. Includes CD-ROM.

Vanhelsuwe:1997:BRJ

[Van97m] Laurence Vanhelsuwé. Book review: The Java Threads API makes it to print media. JavaWorld: IDG’s magazine for the Java community, 2(7):??, July 1997. CO-DEN ???? ISSN 1091-8906.

                                        URL http://www.javaworld.
                                        com/javaworld/jw-07-1997/
                                        jw-07-threads.htm.

Vanhelsuwe:1997:CSC

[Van97n] Laurence Vanhelsuwé. Cover story: Create your own supercomputer with Java. Java-World: IDG’s magazine for the Java community, 2(1), January 1997. CODEN ???? ISSN 1091-8906.

                                        URL http://www.javaworld.
                                        com/javaworld/jw-01-1997/
                                        jw-01-dampp.htm.

Vanhelsuwe:1997:MJb

[Van97o] Laurence Vanhelsuwé. Mastering JavaBeans. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, May 1997. ISBN 0-7821-2097-0. xxx + 799 pp. LCCN QA76.73.J38V374 1997. US$49.99. URL http://www.

                                        sybex.com/cgi-bin/bookpg.
                                        pl?2097back.html. Includes

CD-ROM.

vanHoff:1998:JGB

[van98a] Arthur van Hoff. Java: Getting down to business: Is IS ready for real applications? Dr. Dobb’s Journal of Software Tools, 23 (1):20–22, 24, January 1998. CODEN DDJOEB. ISSN 1044-789X.

VanderVeer:1998:NVJ

[Van98b] Emily A. Vander Veer. Netscape Visual JavaScript for dummies. I D G Books Worldwide, Indianapolis, IN, USA, 1998. ISBN 0-7645-0286-7. xxvi + 380 pp. LCCN QA76.73.J39V3624 1998. US$29.99. URL http:/

/www.idgbooks.com/cgi-bin/
db/fill_out_template.pl?idgbook:
0-7645-0286-7:book-idg::uidg323.

vanderLinden:1999:JJ

[van99a] Peter van der Linden. Just Java 1.2. SunSoft Press Java series. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, fourth edition, March 1999. ISBN 0-13-010534-1. xxi + 777 pp. LCCN QA76.73.J38 V36 1999. US$44.99. URL http:

//www.javaworld.com/books/
jw-clbooks.html?030899txt;
 http://www.phptr.com/bookseri/
java.html;   http://www.phptr.
com/ptrbooks/ptr_0130105341.
html.

vanderLinden:1999:JJT

[van99b] Peter van der Linden. Not Just Java: a technology briefing. SunSoft Press Java series. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, second edition, 1999. ISBN 0-13-079660-3. x + 329 pp. LCCN QA76.73.J38V363 1999.

URL http://www.phptr.com/
ptrbooks/ptr_0130796603.html;
 http://www.sun.com/books/
catalog/vanderLinden3/vanderLinden3.
html;   http://www1.clbooks.
com/asp/BookInfo/BookInfo.
asp?theisbn=0130796603.

vanderLinden:19xx:JJa

[vanxxa] Peter van der Linden. Just Java.

Makron Books do Brasil Editora Ltda., Rua Tabapuã 1348, São Paulo SP - 04533-004 - Brasil, 19xx. ISBN 85-346-0843-1 (??invalid ISBN??). ???? pp. LCCN ???? R$75. URL http:/

                                        /safe.tesla.com.br/makron/
                                        options/livro.asp?id=592.

vanderLinden:19xx:JJe

[vanxxb] Peter van der Linden. Just Java. ASCII Corporation, Tokyo, Japan, 19xx. ISBN 4-7561-0799-0. ???? pp. LCCN ???? 4,800 yen. URL http:/

                                        /www.ascii.co.jp/pb/book1/
                                        shinkan/detail/1322396.html.

Japanese translation.

vanderLinden:19xx:PJ

[vanxxc] Peter van der Linden. Proprio Java. Mondadori Informatica, ??, Italy, 19xx. ISBN ???? ???? pp. LCCN ???? 70000 Lire. URL http://www.

                                        mondadori.com/;   http://
                                        www.sun.com/smi/ssoftpress/
                                        books/vanderLinden2/vanderLinden2.

html. Italian translation of

[van97c].

vanHoff:19xx:APJ

[vanxxd] Arthur van Hoff. Applets programmieren mit Java. Addison-Wesley, Reading, MA, USA, 19xx. ISBN 3-8273-1032-6 (??invalid ISBN??). LCCN ???? 50 DM.

Vanhelsuwe:19xx:CJB

[Vanxxe]  Laurence Vanhelsuwé.  Hét

Complete Java Boek. Sybex Nederland, ????, 19xx. ISBN 90-5160-951-5. LCCN ???? fl.

                                        129. URL http://www.sybex.

nl/catalogus/internet/java.
htm.

Vanhelsuwe:19xx:MBJ

[Vanxxf] Laurence Vanhelsuwé. Mastering Borland JBuilder. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 19xx. ISBN 0-7821-1989-1. ???? pp. LCCN ???? US$50.

Vanhelsuwe:19xx:MJ

[Vanxxg] Laurence Vanhelsuwé. Mastering JavaBeans. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 19xx. ISBN 3-8155-2021-5. ???? pp. LCCN ???? 59 DM.

Varney:1995:NAP

[Var95] Sarah Varney. Netscape’s Andreessen plugs JavaScript. c∥net central, page ??, December 4,

1995. URL http://www.cnet.
com/Central/News/webinno.
html.

Varney:1996:DLW

[Var96] Sarah E. Varney. Datawebs!

link the Web to your legacy data and apps. Datamation, 42 (????):38–41+, ???? 1996. CO-DEN DTMNAT. ISSN 0011-6963.

Varhol:1997:PW

[Var97a] P. Varhol. Purveyor for Windows 95. Web Techniques, 2(1): 71–73, January 1997. CODEN WETEFA. ISSN 1086-556X.

Various:1997:JTO

[Var97b] Various. Java 1.1 Tutto & Oltre. Apogeo srl, Viale Papiniano 38, 20123 Milano, Italy, November 1997. ISBN 88-7303-375-X. LCCN ????

88,000 L. URL http://www.
apogeonline.com/catalogo/
375.html.

Viegnes:1996:HJB

[VB96] Laurent Viegnes and Stéphane Boix. HTML & JAVA: bien débuter. Edition Micro Application, Paris, France, 1996. ISBN 2-7429-0733-5. 358 pp. LCCN ???? 59 FF. URL http://www.

microapp.com/. Includes CD-

ROM.

VonEicken:1999:JKC

[VCCH99] T. Von Eicken, C.-C. Chang, G. Czajkowski, and C. Hawblitzel. J-Kernel: a capabilitybased operating system for Java. Lecture Notes in Computer Science, 1603:369–394, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vckovski:1996:JSS

[Vck96] A. Vckovski. Java as a software system for distributed and interoperable geoprocessing. In Yetongnon and Hariri [YH96], pages 2–?? ISBN ???? LCCN ???? Two volumes.

Vogel:1997:JPC

[VD97] Andreas Vogel and Keith Duddy. Java Programming With CORBA. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, March 1, 1997. ISBN 0-471-17986-8. xvi + 425 pp. LCCN QA76.73.J38V64 1997. US$29.99. URL http:

//www.wiley.com/compbooks/
catalog/17986-8.htm.

Vogel:1998:JPC

[VD98] Andreas Vogel and Keith Duddy. Java Programming With CORBA. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, second edition, February 1998. ISBN 0-471-24765-0. 496 pp. LCCN QA76.73.J38V64 1998. US$34.99. URL http:

//www.wiley.com/compbooks/
catalog/24765-0.htm.

vanDelft:1999:JES

[vD99a] André van Delft. A Java extension with support for dimensions. Software—Practice and Experience, 29(7):605–616, June 1999. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=62003574;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=62003574&PLACEBO=IE.pdf.

vanDeursen:1999:BRB

[vD99b] Arie van Deursen. Book review: Modern Compiler Implementation in Java, Andrew W. Appel, Cambridge University Press, 1998, x + 548 pages, hardback, ISBN 0-521-58388-

8. Science of Computer Programming, 34(2):159–161, June 1999. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/S0167642399000039.

Venditto:1996:JIH

[Ven96] Gus Venditto. Java: It’s hot, but is it ready to serve? — IW Labs examines Sun’s Java language to determine how much of the buzz is hype. Internet World, 7(2):76–??, 1996. CO-DEN IERNE8. ISSN 1064-3923.

Venners:1997:IJV

[Ven97a] Bill Venners. Inside the Java Virtual Machine. McGraw-Hill, New York, NY, USA, November 1997. ISBN 0-07-913248-0. 384 pp. LCCN QA76.73.J38 V46 1998. US$39.95. URL http:

                                        //mcgraw-hill.inforonics.
                                        com/cgi/getarec?mgh31406%
                                        comp.

Venners:1997:UHH

[Ven97b] Bill Venners. Under the hood: How the Java virtual machine performs thread synchronization. JavaWorld: IDG’s magazine for the Java community, 2(7):??, July 1997. CO-DEN ???? ISSN 1091-8906.

                                        URL http://www.javaworld.
                                        com/javaworld/jw-07-1997/
                                        jw-07-hood.htm.

Venners:1999:IJV

[Ven99]  Bill Venners.  Inside the

Java Virtual Machine. Mc-Graw-Hill, New York, NY, USA, second edition, 1999. ISBN 0-07-135093-4. xxxi + 703 pp. LCCN QA76.73.J38 V46 1999. URL http://

www.gbv.de/dms/bowker/toc/
9780071350938.pdf;   http:
//www.loc.gov/catdir/description/
mh024/00269375.html.

Vermeulen:1996:ADP

[Ver96] Allan Vermeulen. An asynchronous design pattern: a pattern for managing concurrency— written in Java! Dr. Dobb’s Journal of Software Tools, 21 (6):42–??, 44, 84–86, June 1996. CODEN DDJOEB. ISSN 1044-789X.

Vermeulen:1997:JDW

[Ver97] Alaln Vermeulen. Java deadlock: The woes of multithreaded design. Dr. Dobb’s Journal of Software Tools, 22(9):52, 54–56, 88, 89, September 1997. CO-DEN DDJOEB. ISSN 1044-789X.

Vermillion:1998:MRL

[Ver98]  Colleen Vermillion.  Media

reviews: Learning Java basics: Java 1.2 and JavaScript for C and C++ Programmers by Michael C. Daconta et al. (John Wiley & Sons,1998, $54.99, ISBN 0-471-18359-8). IEEE MultiMedia, 5(4):83, October–December 1998. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://dlib.computer.org/
mu/books/mu1998/pdf/u4083.
pdf.

vanHalderen:1998:FWBa

[vHO98] Berry A. W. van Halderen and Benno J. Overeinder. Fornax: Web-based distributed discrete event simulation in Java. In ACM [ACM98a], page ?? CODEN CPEXEI. ISSN 1040-3108. LCCN ????

                                        URL http://www.cs.ucsb.
                                        edu/conferences/java98/papers/
                                        fornax.pdf;   http://www.
                                        cs.ucsb.edu/conferences/java98/
                                        papers/fornax.ps. Also pub-

lished as Concurrency: Practice and Experience, 10(11– 13), September 1998, CODEN CPEXEI, ISSN 1040-3108.

Vuong:1996:MIA

[VI96] Son Vuong and Ivailo Ivanov. Mobile intelligent agent systems: WAVE vs. JAVA. In IEEE [IEE96r], pages 196–199. ISBN 0-8186-7585-3. LCCN TK5101.A1C677 1996. IEEE catalog number 96TH100035.

Vijayan:1997:SGJ

[Vij97a] Jaikumar Vijayan. Solaris gets Java jolt. Computerworld, 31 (23):3, June 9, 1997. CODEN CMPWAB. ISSN 0010-4841.

Vijayan:1997:SOS

[Vij97b] Jaikumar Vijayan. Sun offers Solaris cafeteria. Computerworld, 31(25):61–62, June 23, 1997. CODEN CMPWAB. ISSN 0010-4841.

Villacis:1999:NUJ

[Vil99] Juan Villacis. A note on the use of Java in scientific computing. ACM SIGAPP Applied Computing Review, 7 (1):14–17, April 1999. CO-DEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/570150.570155.

Vigil:1996:MIS

[VM96] H. P. Vigil and M. Mueller.

Making the Internet safe for ecommerce. Datamation, 42(16): 64–65, October 1996. CODEN DTMNAT. ISSN 0011-6963.

Vaughan-Nichols:1996:ACJ

[VN96] Steven J. Vaughan-Nichols. ActiveX chases Java. BYTE Magazine, 21(6):27, June 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

VonOheimb:1999:MCJ

[VN99] D. Von Oheimb and T. Nipkow. Machine-checking the Java specification: Proving type-safety. Lecture Notes in Computer Science, 1523:119– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanHalderen:1998:FWBb

[VO98] Berry A. W. Van Halderen and Benno J. Overeinder. Fornax: Web-based distributed discrete event simulation in Java. Concurrency: Practice and Experience, 10(11– 13):957–970, September 1998. CODEN CPEXEI. ISSN 1040-3108. URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=10050402;
                                         http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=10050402&PLACEBO=IE.pdf.

Special Issue: Java for Highperformance Network Computing.

Vogel:1996:BDS

[Vog96] A. Vogel. Building distributed systems in Java. Object Expert, 1(6):24–27, September–October 1996. CODEN OBEXFJ. ISSN 1360-3426.

vonGudenberg:1998:JSC

[von98] J. W. von Gudenberg. Java for scientific computing, pros and cons. J.UCS: Journal of Universal Computer Science, 4(1):11–??, January 28, 1998. CODEN ???? ISSN 0948-6968. URL http:

                                        //medoc.springer.de:8000/
                                        jucs/jucs_4_1/java_for_scientific_
                                        computing.

Vogt:1998:JBV

[VP98] H. Vogt and J. Possegga. Java bytecode verification using model checking. In Proc. OOPSLA’98 Workshop on Formal Underpinnings of the Java Paradigm, Vancouver, October 18–22, 1998, page ?? ACM Press, New York, NY 10036, USA, 1998. ISBN ???? LCCN ????

Vanhelsuwe:1996:MJJ

[VPH^{+}96] Laurence Vanhelsuwé, Ivan Phillips, Goang-Tay Hsu, Krishna Sankar, Eric Ries, Tim Rohaly, and John Zukowski. Mastering Java 2 JDK. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, September 1, 1996. ISBN 0-7821-1935-2. xxxvi + 967 pp. LCCN QA76.73.J38M355 1996. US$49.99. URL http://www.

sybex.com/cgi-bin/bookpg.
pl?1935back.html.

Vanhelsuwe:1997:BJ

[VPH97] Laurence Vanhelsuwé, Ivan Philips, and Goang-Tay Hsu. La Biblia de Java. Anaya Multimedia, Madrid, Spain, May 1997. ISBN 84-415-0092-4. 816 pp. LCCN ???? 5,995 Ptas. URL http://

www.anayamultimedia.es/;
http://www.anayamultimedia.
es/cgi-multi/notas_prensa.
pl?2310120&3&0&1. Spanish

translation of [VPHS97].

Vanhelsuwe:1997:MJa

[VPHS97] Laurence Vanhelsuwé, Ivan Phillips, Goang-Tay Hsu, and Krishna Sankar. Mastering Java 1.1. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, second edition, April 1997. ISBN 0-7821-2070-9. xxxv + 1049 pp. LCCN QA76.73.J38M37 1997. US$49.99. URL http://www.

sybex.com/cgi-bin/bookpg.
pl?2070back.html.

Varvitsiotis:1998:EJT

[VPM98a] A. Varvitsiotis, D. Polemi, and A. Marsh. EUROMED-JAVA: Trusted third party services for securing medical Java applets. Lecture Notes in Computer Science, 1485:209–??, 1998. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Varvitsiotis:1998:ETT

[VPM98b] A. Varvitsiotis, D. Polemi, and A. Marsh. EUROMED-JAVA: Trusted third party services for securing medical Java applets. Lecture Notes in Computer Science, 1485:209–??, 1998. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanderHeijden:1999:FJB

[VPV99] G. Van der Heijden, G. Polder, and J. W. Van Eck. FLORES: a JAVA based image database for ornamentals. Lecture Notes in Computer Science, 1614:641– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vijaykrishnan:1999:TBP

[VR99a] N. Vijaykrishnan and N. Ranganathan. Tuning branch predictors to support virtual method invocation in Java. In USENIX [USE99c], page ?? ISBN ???? LCCN QA76.64 .U84 1999. URL http://www.

                                        usenix.org/publications/library/
                                        proceedings/coots99/vijaykrishnan.
                                        html.

Vogel:1999:PEJ

[VR99b] Andreas Vogel and Madhavan Rangarao. Programming with Enterprise JavaBeans, JTS, and OTS: Building Distributed Transactions with Java and C++. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 1999. ISBN 0-471-31972-4. 368 pp. LCCN QA76.73.J38P767 1999. US$39.99.

Vijaykrishnan:1998:OAS

[VRG98a] N. Vijaykrishnan, N. Ranganathan, and R. Gadekarla. Object-oriented architectural support for a Java processor. Lecture Notes in Computer Science, 1445:330–??, 1998. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vijaykrishnan:1998:OOA

[VRG98b] N. Vijaykrishnan, N. Ranganathan, and R. Gadekarla. Object-oriented architectural support for a Java processor. Lecture Notes in Computer Science, 1445:330–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/
service/series/0558/bibs/
1445/14450330.htm;   http:
//link.springer-ny.com/link/
service/series/0558/papers/
1445/14450330.pdf.

vanReeuwijk:1997:SPL

[vRvGS97] Kees van Reeuwijk, Arjan J. C.

van Gemund, and Henk J. Sips. Spar: a programming language for semi-automatic compilation of parallel programs. Concurrency: Practice and Experience, 9(11):1193–1205, November 1997. CODEN CPEXEI. ISSN 1040-3108. URL http:

                                        //www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=13818;
                                         http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=13818&PLACEBO=IE.pdf. Spe-

cial Issue: Java for computational science and engineering — simulation and modeling II.

Volpano:1996:SDW

[VS96] D. Volpano and G. Smith. On the systematic design of Web languages. ACM Computing Surveys, 28(2):315–317, June 1996. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic).

vanHoff:1996:BJ

[vSS96a] Arthur van Hoff, Sami Shaio, and Orca Starbuck. Branché Java. International Thomson Publishing France, Paris, France, 1996. ISBN 2-84180-078-4. xviii + 204 pp. LCCN ???? 220 FF. URL http:

                                        //www.thomson.com/intluk.

html. French translation by Serge Chaumette and Alain Miniussi of [vSS96c]. Includes CD-ROM.

vanHoff:1996:HJ

[vSS96b] Arthur van Hoff, Sami Shaio, and Orca Starbuck. Hooked on Java: creating hot Web sites with Java applets. The Java Series. Addison-Wesley, Reading, MA, USA, 1996. ISBN 0-201-48837-X (paperback), 0-201-85274-8 (computer laser optical disk). xviii + 181 pp. LCCN QA76.64.V36 1996. US$29.95. URL http:/

/www.aw.com/devpress/java/

index.html. Includes CD-ROM. See [Nic96c].

vanHoff:1996:HJC

[vSS96c] Arthur van Hoff, Sami Shaio, and Orca Starbuck. Hooked on Java: creating hot Web sites with Java applets. The Java Series. Addison-Wesley, Reading, MA, USA, January 1996. ISBN 0-201-48837-X (paperback), 0-201-85274-8 (computer laser optical disk). xviii + 181 pp. LCCN QA76.64.V36 1996. US$29.95. URL http:/

/www.aw.com/devpress/java/

index.html. Includes CD-ROM.

vanHoff:1996:JAE

[vSS96d] Arthur van Hoff, Sami Shaio, and Orca Starbuck. Java-Applets erstellen und nutzen: Interaktive Web-Seiten selbstgemacht. Addison-Wesley, Reading, MA, USA, 1996. ISBN 3-8273-1032-6. 221 (or

220??)  pp.  LCCN ????

49,90 DM. URL http://www.
addison-wesley.de/katalog/
item.ppml?textexpr=Java&id=

00004. German translation of [vSS96c]. Includes CD-ROM.

vanHoff:1996:WTC

[vSS96e] Arthur van Hoff, Sami Shaio, and Orca Starbuck. What is this thing called Java — Java is a lot of things to a lot of people. here are a dozen of its most important features, according to the people. Datamation, 42(5):45– 46, ???? 1996. CODEN DTM-NAT. ISSN 0011-6963.

vanHoff:1997:HJa

[vSS97] Arthur van Hoff, Sami Shaio, and Orca Starbuck. Hooked on Java. Addison-Wesley, Reading, MA, USA, 1997. ISBN 4-7952-9677-4. ???? pp. LCCN

????  ????  URL http:

                                        //www.javasoft.com/people/

avh/. Japanese translation.

Winters:1996:TYV

[W^{+}96] Patrick Winters et al. Teach Yourself Visual J++ in 21 Days. Sams.net, Indianapolis, IN 46268, USA, 1996. ISBN 1-57521-158-0. xvii + 560 pp. LCCN QA76.73.J38 T43 1996. US$39.99. URL http://

                                        merchant.superlibrary.com:
                                        8000/catalog/hg/PRODUCT/PAGE/
                                        15752/bud/1575211580.html;
                                         http://www.mcp.com/cgi-bin/
                                        bag?isbn=1-57521-158-0&last=
                                        /bookstore.

Wagner:1997:JU

[W^{+}97a] Richard Wagner et al. Javascript Unleashed. Sams.net, Indianapolis, IN 46268, USA, second edition, 1997. ISBN 1-57521-306-0. xxxii + 1014 pp. LCCN QA76.73.J39W33 1997. 99,00 DM. URL http://

                                        merchant.superlibrary.com:
                                        8000/catalog/hg/PRODUCT/PAGE/
                                        15752/bud/1575213060.html;
                                         http://www.mut.ch/profit.

htx?&x_pk=kd84mf7ms9dm. In-

cludes CD-ROM.

White:1997:SEU

[W^{+}97b] Barbara White et al. Special Edition Using Java Beans. Que Corporation, Indianapolis, IN, USA, 1997. ISBN 0-7897-1460-

4. ???? pp. LCCN ???? 99,00 DM; US$50. URL http://

merchant.superlibrary.com:
8000/catalog/hg/PRODUCT/PAGE/
07897/bud/0789714604.html;
 http://www.mut.ch/profit.
htx?&x_pk=kd84mf7ms9dm. In-

cludes CD-ROM.

Winters:1997:VJT

[W^{+}97c] Patrick Winters et al. Visual J++ in 21 Tagen: Java-Programmierung in Microsofts neuer Entwicklungsumgebung. Markt & Technik, Haar bei München, Germany, 1997. ISBN 3-87791-896-4. LCCN ???? 90 DM. URL http:/

/www.mut.ch/html/buchinfo.
htx?&x_pp=dJgxRqA9wlAYNijX&
x_az=10&x_pk=VpLvjnsx7b5V&
x_pa=B&x_pr=oS&x_bestell=
21158SAN&x_stichwort=Winters&
x_sprache=A&x_start=1&x_gefunden=
1&x_query=&x_lastpage=buchfund&
x_soft=&x_suchewo=A.

Westwood:1998:MMV

[W^{+}98] James D. Westwood et al., editors. Medicine meets virtual reality: art, science, technology: healthcare (r)evolution. San Diego, California, January 28–31, 1998, volume 50 of Studies in health technology and informatics. IOS Press, Amsterdam, The Netherlands, 1998. ISBN 90-5199-386-2, 4-274-90210-2. ISSN 0926-9630. LCCN W26.5 M45 1998.

Winters:19xx:AED

[W^{+}xxa] Patrick Winters et al. Aprenda em 21 Dias Visual J++. Editore Campus Ltda., Rua Sete de Setembro, 111/16, Rio de Janeiro - RJ - 20050-002, Brazil, 19xx. ISBN ???? ???? pp.

LCCN ????  ????  URL

                                        http://www.campus.com.br/
                                        catalogo/livro.cfm?cod_livro=
                                        20170.

Winters:19xx:VJP

[W^{+}xxb] Patrick Winters et al. Visual J++ Le Programmeur. Simon & Schuster Macmillan (France), Paris, France, 19xx. ISBN 2-7440-0229-1. ????

pp.  LCCN ????  199

                                        FF. URL http://www.ssm.fr/
                                        catalog/fiches/0229.htm.

Weiner:1998:PJ

[WA98] Scott Weiner and Stephen Asbury. Programming with JFC. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, April 1998. ISBN 0-471-24731-6. x + 564 pp. LCCN QA76.73.J38 W44 1998. US$50. URL http:

                                        //www.wiley.com/compbooks/
                                        catalog/24731-6.htm.

Waegemann:1996:PTE

[Wae96] C. Peter Waegemann, editor. Proceedings. Toward an Electronic Patient Record ’96.

Twelfth International Symposium on the Creation of Electronic Health Record System and Global Conference on Patient Cards, May 1996, San Diego, CA, volume 1. Medical Records Inst, Newton, MA, USA, 1996. ISBN 0-9640667-7-7. LCCN R864.I68 1996. Two volumes.

Wu:1997:JBF

[WAES97a] D. Wu, D. Agrawai, A. El Abbadi, and A. Singh. A Javabased framework for processing distributed objects. Lecture Notes in Computer Science, 1331:333–??, 1997. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wu:1997:JFP

[WAES97b] D. Wu, D. Agrawai, A. El Abbadi, and A. Singh. A Javabased framework for processing distributed objects. Lecture Notes in Computer Science, 1331:333–??, 1997. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wagner:1996:HSC

[Wag96] Mitch Wagner. Hot showgoers cool toward Java. Computerworld, 30(19):14, May 6, 1996. CODEN CMPWAB. ISSN 0010-4841.

Wagner:1997:BIL

[Wag97a] Mitch Wagner. Button-down IBM loosens up for Java. Computerworld, 31(19):12, May 12, 1997. CODEN CMPWAB. ISSN 0010-4841.

Wagner:1997:SWT

[Wag97b] Mitch Wagner. SourceCraft Web tool combines Java, HTML. Computerworld, 31(6): 51, 53, February 10, 1997. CO-DEN CMPWAB. ISSN 0010-4841.

Wakeling:1998:HJV

[Wak98a] D. Wakeling. A Haskell to Java Virtual Machine code compiler. Lecture Notes in Computer Science, 1467:39–??, 1998. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wakeling:1998:MHC

[Wak98b] D. Wakeling. Mobile Haskell: Compiling lazy functional programs for the Java Virtual Machine. Lecture Notes in Computer Science, 1490:335– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wakeling:1999:CLF

[Wak99] David Wakeling. Compiling lazy functional programs for the Java virtual machine. Journal of Functional Programming, 9(6):579–603, November 1999. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

                                        //www.cambridge.org/core/
                                        product/AC59FAFC68063D9EC71E4FCF283E7EFC.

Walder:1996:CIP

[Wal96a] B. Walder. CT and the Internet: paving the way for a paperless empire. Network Week, 1(42):

29–31, ???? 10, 1996. CODEN NEWEFS.

Waldo:1996:IIJ

[Wal96b] J. Waldo. Interface and implementation [Java]. UNIX review, 14(13):81–82, 85–86, December 1996. CODEN UNRED5. ISSN 0742-3136.

Waldo:1996:PCJ

[Wal96c] J. Waldo. Protected classes [in Java]. UNIX review, 14(11): 97–98, 100–101, October 1996. CODEN UNRED5. ISSN 0742-3136.

Waldo:1996:JAL

[Wal96d] Jim Waldo. Java advisor: a language with class. UNIX review, 14(9):83–92, August 1, 1996. CODEN UNRED5. ISSN 0742-3136.

Waldo:1996:JAI

[Wal96e] Jim Waldo. Java advisor: Interface and implementations. UNIX review, 14(??):81–84, December 1996. CODEN UNRED5. ISSN 0742-3136.

Waldo:1996:JAP

[Wal96f] Jim Waldo. Java advisor: Protected classes. UNIX review, 14 (??):97–101, October 1996. CO-DEN UNRED5. ISSN 0742-3136.

Waldo:1996:PJ

[Wal96g] Jim Waldo. Programming with Java. UNIX review, 14(5):31– 32, 34, 36–37, May 1, 1996. CO-DEN UNRED5. ISSN 0742-

Waldo:1996:PJJ

[Wal96h] Jim Waldo. Programming with Java — Java is more than just a language for Web programming. in fact, it provides reliability and security that can serve many programming needs. UNIX review, 14(5):31–??, ???? 1996. CODEN UNRED5. ISSN 0742-3136.

Walker:1996:FFM

[Wal96i] D. W. Walker. Focus: Freemarket computing and the global economic infrastructure. IEEE parallel and distributed technology: systems and applications, 4(3):60–62, Fall 1996. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). Discusses the role of Java as a standard for the Web, and how payment for computing resources might be triggered by author bank account information stored in Java applets.

Wall:1996:IJP

[Wal96j] David A. Wall. 101 Freeze Dried Java Applets. IDG Books, San Mateo, CA, USA, 1996. ISBN 0-7645-4008-4. LCCN ???? US$35.

Walnum:1996:JE

[Wal96k] Clayton Walnum. Java By Example. Que Corporation, Indianapolis, IN, USA, June 1, 1996. ISBN 0-7897-0814-0. xxv + 549 pp. LCCN QA76.73.J38W35 1996. US$34.99. URL http://www.

mcp.com/251356511370317/cgi-

bin/bag?isbn=0-7897-0814-
0&last=/bookstore;   http:
//www.mcp.com/que.

Walsh:1996:JDa

[Wal96l] Aaron Walsh. Java for Dummies. I D G Books Worldwide, Indianapolis, IN, USA, May 1996. ISBN 0-614-14433-7. ???? pp. LCCN ???? US$24.99.

Walsh:1996:FJP

[Wal96m] Aaron E. Walsh. Foundations of Java Programming for the World Wide Web. I D G Books Worldwide, Indianapolis, IN, USA, 1996. ISBN 1-56884-811-0. xxviii + 906 pp. LCCN QA76.73.J38W357 1996. US$44.99. URL http://db.

www.idgbooks.com/database/
book/isbn/generic-book.tmpl?
query=1-56884-811-0;   http:
//www.idgbooks.com/database/
book_result.msql?uid=uidg323&
isbn=1-56884-811-0.

Walsh:1996:JDb

[Wal96n] Aaron E. Walsh. Java for dummies. I D G Books Worldwide, Indianapolis, IN, USA, Macintosh and Windows 95 edition, 1996. ISBN ???? 1 laser optical disc pp. LCCN ????

Walsh:1996:JDc

[Wal96o] Aaron E. Walsh. Java for Dummies. I D G Books Worldwide, Indianapolis, IN, USA, August 1996. ISBN 1-56884-641-X. xviii + 365 pp. LCCN QA76.73.J38 W36 1996. US$19.99. URL http://db.

                                        www.idgbooks.com/database/
                                        book/isbn/generic-book.tmpl?
                                        query=1-56884-641-X;   http:
                                        //www.idgbooks.com/database/
                                        book_result.msql?uid=uidg323&
                                        isbn=1-56884-641-X.

Walsh:1996:PJ

[Wal96p] Aaron E. Walsh. Programacion en Java. Anaya Multimedia, Madrid, Spain, 1996. ISBN 84-415-0132-7. ???? pp. LCCN ???? 5495 Ptas. URL http://

                                        www.anaya.es/Catalogo/CGA?
                                        act=L&art=2318005&ed=2300.

Walsh:1996:CV

[Wal96q] Brian Walsh. Corporate view. Network Computing, 7(13):39– ??, September 1, 1996. CODEN NCOMEV. ISSN 1046-4468.

Waldo:1997:III

[Wal97a] J. Waldo. Interface and implementation. II. UNIX review, 15 (2):71–74, February 1997. CO-DEN UNRED5. ISSN 0742-3136.

Walnum:1997:PDG

[Wal97b] Clayton Walnum. Professional Developer’s Guide Java. Que Corporation, Indianapolis, IN, USA, 1997. ISBN 0-7897-1168-0. ???? pp. LCCN ???? 99,00 DM. URL http:/

                                        /www.mut.ch/profit.htx?&x_
                                        pk=kd84mf7ms9dm.

Walnum:1997:SEU

[Wal97c] Clayton Walnum. Special Edition Using Java Development Tools. Que Corporation, Indianapolis, IN, USA, 1997. ISBN 0-7897-1317-9. ???? pp. LCCN ???? 99,00 DM; US$50. URL

http://www.mut.ch/profit.
htx?&x_pk=kd84mf7ms9dm.

Creek:1997:JDT

[Wal97d] Walnut Creek. Jolt! Developer’s Toolkit (for Java). Walnut Creek, ????, October 1997. ISBN 1-57176-211-6. ???? pp. LCCN ???? US$40.

URL http://www.cdrom.com/
titles/prog/jolt_dev.htm.

Creek:1997:JGJ

[Wal97e] Walnut Creek. Jolt! Games (for Java). Walnut Creek, ????, February 1997. ISBN 1-57176-197-7. ???? pp. LCCN ???? US$40. URL http://www.

cdrom.com/titles/games/jolt_
gam.htm.

Walsh:1997:JB

[Wal97f] Aaron Walsh. Java Bible. I D G Books Worldwide, Indianapolis, IN, USA, 1997. ISBN 0-7645-8030-2. 960 pp. LCCN ???? US$49.99.

URL http://www.idgbooks.
com/database/book_result.
msql?uid=uidg323&isbn=0-7645-
8030-2.

Walsh:1997:JDb

[Wal97g] Aaron Walsh. Java für Dummies. International Thomson Computer Press, 20 Park Plaza Suite 1001, Boston, MA 02116 USA, 1997. ISBN 3-8266-2716-4 (??invalid ISBN??). 382 pp. LCCN ???? 50 DM. URL http:

//www.itp.de/. Includes CD-

ROM.

Walsh:1997:JDa

[Wal97h] Aaron E. Walsh. Java for Dummies. IDG Books, San Mateo, CA, USA, second edition, May 1997. ISBN 0-7645-0140-2 (softcover). xx + 373 pp. LCCN QA76.73.J38W36 1997. US$22.49. URL http://

                                        www.idgbooks.com/database/
                                        book_result.msql?uid=uidg323&
                                        isbn=0-7645-0140-2.

Walworth:1997:JGT

[Wal97i] Alan Walworth. Java GUI testing. Dr. Dobb’s Journal of Software Tools, 22(2):30, 32, 34, February 1997. CODEN DDJOEB. ISSN 1044-789X.

Waldo:1998:SRP

[Wal98a] Jim Waldo. Spotlight: Remote procedure calls and Java Remote Method Invocation. IEEE Concurrency, 6(3):5–7, July/September 1998. CO-DEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic).

                                        URL http://dlib.computer.
                                        org/pd/books/pd1998/pdf/p3005.
                                        pdf.

Walnum:1998:ABB

[Wal98b] Clayton Walnum. AFC Black Book. Coriolis Group Books, Scottsdale, AZ, USA, April 1998. ISBN 1-57610-235-1. xvii + 638 pp. LCCN QA76.73.J38 W345 1998. US$40.

Walsh:1998:JDa

[Wal98c] Aaron Walsh. Java for Dummies. IDG Books, San Mateo, CA, USA, third edition, 1998. ISBN 0-7645-0417-7. LCCN QA76.73.J38 W36 1998. US$24.99.

Walsh:1998:JDb

[Wal98d] Aaron Walsh. Java for Dummies. Apogeo srl, Viale Papiniano 38, 20123 Milano, Italy, April 1998. ISBN 88-7303-423-3. LCCN ????

39,500 L. URL http://www.
apogeonline.com/catalogo/
375.html.

Waldo:1999:JAN

[Wal99] Jim Waldo. The Jini architecture for network-centric computing. Communications of the ACM, 42(7):76–82, July 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.
acm.org:80/pubs/citations/
journals/cacm/1999-42-7/p76-
waldo/.

Walmsley:19xx:PJ

[Walxxa] N. Walmsley. Programming in Java. Bernard Babani, ????, 19xx. ISBN 0-85934-436-3. ???? pp. LCCN ???? ????

Walnum:19xx:JEE

[Walxxb] Clayton Walnum. Java em Exemplos: a Maneira Mais Fácil de Aprender a Programar para a Web. Makron Books do Brasil Editora Ltda., Rua Tabapuã 1348, São Paulo SP - 04533-004 - Brasil, 19xx. ISBN 85-346-0680-3. 644 pp. LCCN ???? R$ 79.00. URL http:/

/safe.tesla.com.br/makron/
options/livro.asp?id=25;

                                        http://www.makron.com/. Por-

tuguese translation of [Wal96k].

Walsh:19xx:FJP

[Walxxc]  Aaron Walsh.  Foundations

of Java Programming for the World Wide Web. Dialectica Computer Pub., ????, 19xx. ISBN ???? ???? pp. LCCN ????

                                        URL http://www.dcp.kiev.
                                        ua/koi8/books/foundation/
                                        f_javaww.html.

Walsh:19xx:JH

[Walxxd] Aaron Walsh. Het Java Handboek. Addison-Wesley Nederland, ????, 19xx. ISBN 90-6789-733-7. ???? pp. LCCN ???? fl.

80. Includes CD-ROM.

Walsh:19xx:JD

[Walxxe] Aaron Walsh. Java für Dummies. International Thomson Computer Press, 20 Park Plaza Suite 1001, Boston, MA 02116 USA, 19xx. ISBN 3-8266-2716-4 (??invalid ISBN??). LCCN ????

                                        50 DM. URL http://www.itp.
                                        de/dummies/2716/2716.html.

Walsh:19xx:JVDa

[Walxxf] Aaron Walsh. Java voor dummies. Addison-Wesley Nederland, ????, 19xx. ISBN 90-6789-747-7. ???? pp. LCCN ???? fl.

70. Includes CD-ROM.

Walsh:19xx:JVDb

[Walxxg] Aaron Walsh. Java voor dummies. Addison-Wesley Nederland, ????, second edition, 19xx. ISBN 90-6789-897-X. ???? pp. LCCN ???? fl. 70.

Warford:1999:BNO

[War99] J. Stanley Warford. BlackBox:

a new object-oriented framework for CS1/CS2. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(1):271–275, March 1999. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Watt:1995:NMU

[Wat95a] Peggy Watt. News: Microsoft unfolds Internet road map. Network World, page ??, December 11, 1995. ISSN 0887-7661 (print), 1944-7655 (elec-

tronic). URL http://www.
iquest.net/index.html.

Watt:1995:NNS

[Wat95b] Peggy Watt. News: Netscape, Sun announce JavaScript. Network World, page ??, December 01, 1995. ISSN 0887-7661 (print), 1944-7655 (elec-

tronic). URL http://www.
iquest.net/index.html.

Watt:1995:NSV

[Wat95c] Peggy Watt. News: Several vendors announce plans for VRML. Network World, page ??, December 11, 1995. ISSN 0887-7661 (print), 1944-7655 (elec-

tronic). URL http://www.
iquest.net/index.html.

Watt:1995:NSA

[Wat95d] Peggy Watt. News: Sun– Netscape–GI announce their Web browser. Network World, page ??, December 11, 1995. ISSN 0887-7661 (print), 1944-7655 (electronic). URL http://

www.iquest.net/index.html.

Watt:1995:NTM

[Wat95e] Peggy Watt. News: Travel made easy via the Web. Network World, page ??, December 18, 1995. ISSN 0887-7661 (print), 1944-7655 (elec-

tronic). URL http://www.
iquest.net/index.html.

Wattawa:1995:API

[Wat95f] Scott Wattawa. Agent-based paradigm for integration of interactive cable television operations and business support systems. Proceedings of the SPIE — The International Society for Optical Engineering, 2609: 67–79, ???? 1995. CODEN PSISDG. ISBN 0-8194-1973-7. ISSN 0277-786X (print), 1996-756X (electronic).

Watson:1997:JCR

[Wat97a] Mark Watson. Creating JavaBeans?: Components for Distributed Applications. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 1997. ISBN 1-55860-476-6. xvii + 238 pp. LCCN QA76.73.J38W37 1998. US$30. URL http://

www.mkp.com/books_catalog/
1-55860-476-6.asp.

Watson:1997:IJA

[Wat97b] Mark Watson. Intelligent Java Applications for the Internet and Intranets. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, March 1997. ISBN 1-55860-420-0. xx + 377 pp. LCCN QA76.73.J38W38 1997. US$39.95. URL http://

www.mkp.com/books_catalog/
1-55860-420-0.asp.

Watson:1998:JPW

[Wat98]  Mark Watson.  Java Pro-

gramming for Windows: Using Microsoft AFC, WFC, and XML. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 1998. ISBN 1-55860-516-9. xxiii + 344 pp. LCCN QA76.73.J38W383 1999. US$39.95. URL http://

www.mkp.com/books_catalog/
catalog.asp?ISBN=1-55860-
516-9.

Watson:1999:JPW

[Wat99]  Mark Watson.  Java pro-

gramming for Windows: using Microsoft AFC, WFC, and XML. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 1999. ISBN 1-55860-516-9. xxiii + 344 pp. LCCN QA76.73.J38W383 1999.

Wayner:1996:IN

[Way96a] P. Wayner. Inside the NC.

BYTE Magazine, 21(11):105– 106, 108, 110, November 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wayner:1996:BJP

[Way96b] Peter Wayner. Better Java programming: Knowing how Java’s dynamic linking works can help you improve a program’s performance. BYTE Magazine, 21 (9):63–64, September 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wayner:1996:INN

[Way96c] Peter Wayner. Inside the NC: Are network computers just stripped-down terminals? No way. The official NC platform covers everything from a set-top box to a Cray. BYTE Magazine, 21(11):105, 106, 108, 110, November 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic). Discusses the importance of Java in a Network Computer.

Wayner:1996:NPM

[Way96d] Peter Wayner. Net programming for the masses. BYTE Magazine, 21(2):101–102, 104, February 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wayner:1996:SGJ

[Way96e] Peter Wayner. Sun gambles on Java chips: Are Java chips better than general purpose CPUs? or will new compilers make them obsolete? BYTE Magazine, 21(11):79, 80, 82, 84, 86, 88, November 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wayner:1996:XTB

[Way96f] Peter Wayner. X terminal + browser + Java = Web PC. BYTE Magazine, 21(10):50–??, October 1, 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wayner:1996:XBJ

[Way96g] Peter Wayner. XTerminal + browser + Java = Web PC — HDS’s (a) workStation is a lowcost box for Web access, Java applets, and Windows. BYTE Magazine, 21(10):50–??, October 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wayner:1997:JJP

[Way97a]  Peter Wayner.  Java and

JavaScript Programming. AP Professional, Boston, MA, USA, 1997. ISBN 0-12-738769-2. x + 335 pp. LCCN QA76.73.J38W39 1996. US$29.95. URL http://www.

apcatalog.com/cgi-bin/AP?
ISBN=0-12-738769-2&LOCATION=
US&FORM=FORM2;   http://www.
apnet.com/.

Wayner:1997:NCE

[Way97b] Peter Wayner. Novera composes Epic for Java. BYTE Magazine, 22(7):122–123, July 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wayner:1998:HSJ

[Way98a] Peter Wayner. How to soup up Java, part II: Nine recipes for east, easy Java: a maturing Java brings a wide variety of programming solutions to developers. BYTE Magazine, 23(5):76–80, May 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Wayner:1998:JRP

[Way98b]  Peter Wayner.  JavaBeans

for Real Programmers. AP Professional, Boston, MA, USA, January 1998. ISBN 0-12-738670-X. ???? pp. LCCN QA76.73.J38W393 1998.

US$40.  URL http://www.
apcatalog.com/cgi-bin/AP?
ISBN=012738670X&LOCATION=
US&FORM=FORM2.

Wayner:1998:MJD

[Way98c] Peter Wayner. Making Java development JSafe — RSA’s cryptographic toolkit targets Java as the platform of choice for secured Internet applications. BYTE Magazine, 23 (1):117, January 1998. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Weber:1996:SEU

[WB^{+}96a] Joseph Weber, David Baker, et al. Special Edition Using Java. Que Corporation, Indianapolis, IN, USA, second edition, 1996. ISBN 0-7897-0936-8. xxx + 1134 pp. LCCN QA76.73.J38W4 1996. US$49.99. URL http:

//www.mcp.com/cgi-bin/bag?

isbn=0-7897-0936-8&last=/
bookstore.

Wehling:1996:LNA

[WB^{+}96b] Jason Wehling, Vidya Bharat, et al. Late Night Advanced Java. Ziff-Davis Press, Emeryville, CA, USA, December 1996. ISBN 1-56276-407-1. xxvii + 588 pp. LCCN ???? US$35. URL http://

merchant.superlibrary.com:
8000/catalog/hg/PRODUCT/PAGE/
15627/bud/1562764071.html.

Wynkoop:1996:RPW

[WB^{+}96c] Stephen Wynkoop, John Burke, et al. Running a Perfect Web Site: Learn how to implement the hottest technologies like Java, JavaScript, VB-Script and ActiveX into your site. Que Corporation, Indianapolis, IN, USA, second edition, 1996. ISBN 0-7897-0944-9 (softcover). xxviii + 711 pp. LCCN TK5105.888.W96 1996. US$44.99. Includes CD-ROM.

Weitzel:1998:NJD

[WB98] Steven Weitzel and Lyle Bates.

Novell’s Java Developer’s Guide. I D G Books Worldwide, Indianapolis, IN, USA, May 1998. ISBN 0-7645-4541-8. 600 pp. LCCN ???? US$59.99.

URL http://www.novell.ru:
8081/programs/press/4541.

html. Includes CD-ROM.

Watt:1999:PLP

[WB99] David A. (David Anthony)

Watt and Deryck F. Brown.

Programming language processors in Java: compilers and interpreters. Prentice Hall, Harlow, England, 1999. ISBN 0-13-025786-9 (case). xvi + 436 pp. LCCN QA76.73.J38 W385 1999.

Wood:1996:WSC

[WBA^{+}96] Frank Wood, Daniel Brown, Robert A. Amidon, Jonathan Alferness, Brian Joseph, Richard E. Gillilan, and Carlos Faerman. WorkSpace and the study of Chagas’ disease. IEEE Computer Graphics and Applications, 16(4):72–78, July 1996. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). WorkSpace is a 3-D extension of the 2-D graphics model provided by the Java Abstract Window Toolkit.

Wutka:1997:HJP

[WBB^{+}97] Mark Wutka, David Baker, David Boswell, Ken Cartwright, and Da(??) Liebke. Hacking Java: the professional’s resource kit. Que Corporation, Indianapolis, IN, USA, 1997. ISBN 0-7897-0935-X. xviii + 852 pp. LCCN QA76.73.J38W88 1997. US$59.99; 119,00 DM. URL

http://www.mcp.com/cgi-bin/
bag?isbn=0-7897-0935-X&last=
/bookstore;   http://www.
mut.ch/profit.htx?&x_pk=kd84mf7ms9dm.

Includes CD-ROM.

Wharton:1996:EPS

[WBC^{+}96] John Wharton, John Banning, Brian Case, David S. Hardin, Martin Hopkins, John Novitsky, and Marc Tremblay. Evening panel session, Lagunita Court: Software or silicon: What’s the best route to Java? In IEEE [IEE96e], pages 145–146. ISBN ???? LCCN ????

Weber:1997:SEU

[WBC^{+}97] Joe Weber, David Baker, Joe Carpenter, Jamie Costa, Anil Hemrajani, and Joseph Weber. Special Edition Using Java 1.1. Que Corporation, Indianapolis, IN, USA, third edition, 1997. ISBN 0-7897-1094-3. xlix + 1234 pp. LCCN QA76.73.J38W4 1997. US$49.99. URL http://

merchant.superlibrary.com:
8000/catalog/hg/PRODUCT/PAGE/
07897/bud/0789710943.html.

Wallach:1997:ESA

[WBDF97] Dan S. Wallach, Dirk Balfanz, Drew Dean, and Edward W. Felten. Extensible security architectures for Java. Operating Systems Review, 31(5):116–128, December 1997. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wagner:1996:JU

[WC96] Richard Wagner and Paul Colton. JavaScript Unleashed. Sams.net, Indianapolis, IN 46268, USA, October 1, 1996. ISBN 1-57521-118-1. xxx + 869 pp. LCCN QA76.73.J39J38 1996. US$49.99. URL http://

merchant.superlibrary.com:
8000/catalog/hg/PRODUCT/PAGE/
15752/bud/1575211181.html;

                                         http://www.mcp.com/samsnet/

.

Walrath:1998:JST

[WC98] Kathy Walrath and Mary Campione. The JFC Swing Tutorial: a Guide to Constructing GUIs. Addison-Wesley, Reading, MA, USA, 1998. ISBN 0-201-43321-4. 752 pp. LCCN QA76.9.U83W36 1998. URL

                                        http://www2.awl.com/cseng/
                                        javaseries/swingtut.html.

Wilson:1999:PBG

[WC99] Gregory V. Wilson and Steve Chartley. Programmer’s bookshelf: The general, the particular, and the just plain odd. Dr. Dobb’s Journal of Software Tools, 24(4):117–118, April 1999. CODEN DDJOEB. ISSN 1044-789X. URL http:

                                        //www.ddj.com/1999/9904q/
                                        9904q.htm.

White:1998:EJP

[WCF98] Seth White, Rick Cattell, and Shel Finkelstein. Enterprise Java platform data access. SIG-MOD Record (ACM Special Interest Group on Management of Data), 27(2):504–505, June 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Warner:1997:JSP

[WCL97] Simeon Warner, Simon Catterall, and Edward Lipson. Java simulations for physics education. Concurrency: Practice and Experience, 9(6):477–484, June 1997. CODEN CPEXEI. ISSN 1040-3108. URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=13872;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=13872&PLACEBO=IE.pdf.

Wu:1999:JBD

[WCS99] X. Wu, Q. Chen, and X.-H.

Sun. A Java-based distributed debbuger supporting MPI and PVM. Parallel and Distributed Computing Practices, 2(4):??, ???? 1999. CODEN ???? ISSN 1097-2803. URL http://www.

cs.okstate.edu/~pdcp/vols/
vol02/vol02no4abs.html#wu.

Walnum:1996:UJW

[WD^{+}96] Clayton Walnum, Michael Desmond, et al. Using Java Workshop. Que Corporation, Indianapolis, IN, USA, August 16, 1996. ISBN 0-7897-0900-7. xx + 425 pp. LCCN QA76.73.J38W355 1996. US$34.99, CDN$49.95, UK£32.49. URL http:/

/www.mcp.com/cgi-bin/bag?
isbn=0-7897-0900-7&last=/
bookstore.

Weiss:1998:TJJ

[WDDF98] M. Weiss, F. De Ferriere, B. Detsart, and C. Fabre. Turbo J, a Java bytecode-to-native compiler. Lecture Notes in Computer Science, 1474:119– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weiss:1998:TJB

[WdFD^{+}98] Michael Weiss, François de Ferrière, Bertrand Delsart, Christian Fabre, Frederick Hirsch, E. Andrew Johnson, Vania Joloboff, Fred Roy, Fridtjof Siebert, and Xavier Spengler. TurboJ, a Java bytecode-to-native compiler. Lecture Notes in Computer Science, 1474:119– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/
                                        service/series/0558/bibs/
                                        1474/14740119.htm;   http:
                                        //link.springer-ny.com/link/
                                        service/series/0558/papers/
                                        1474/14740119.pdf.

Weakley:1997:IAW

[Wea97a]  William R. Weakley.  In-

teractive authoring on the Web: an HTML tutorial using JavaScript. Thesis (m.a.), Illinois State University, Springfield, IL, USA, 1997. iv + 93 pp.

Weatherford:1997:MVI

[Wea97b] M. Weatherford. Micro view: IBM bets on JavaBeans. IEEE Micro, 17(3):80–??, May/June 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //pascal.computer.org/mi/
                                        books/mi1997/pdf/m3080.pdf.

Webster:1996:HSG

[Web96] John S. Webster. HDS’ secondgeneration network computer.

SunExpert Magazine, 7(12):16, 18, December 1996. ISSN 1053-9239. Notes that HDS’ Network Computer contains a 66-MHz Intel i960 and supports up to 1600 × 1200 resolution. HDS NetOS includes its inhouse Java Virtual Machine, with about twice the performance of a 133MHz Intel Pentium PC. It is the first Network Computer (formerly, X terminal) to support a Java-enabled browser in the machine itself.

Webster:1997:IDW

[Web97a] John S. Webster. Insignia delivers Windows to Java desktops. SunExpert Magazine, 8(4): 21–22, April 1997. ISSN 1053-9239. Describes NTrigue Client for Java.

Webster:1997:JAJ

[Web97b] John S. Webster. Java 3D adds to Java’s CAD utility. Sun-Expert Magazine, 8(5):48, May 1997. ISSN 1053-9239.

Webster:1997:JGDa

[Web97c] John S. Webster. JavaSoft gives developers compatibility tests. SunExpert Magazine, 8(3):16, 18, 20, 22, March 1997. ISSN 1053-9239.

Webster:1997:JGDb

[Web97d] John S. Webster. JavaSoft gives developers compatibility tests. WebServer Magazine: For Managers of World Wide Web Sites, 2(2):11–12, March 1997. ISSN 1087-4232.

Weber:1998:LBS

[Web98] Stephan Weber. E-Lab: Berechnung und Simulation von Schaltungen; das virtuelle Elektronik-Labor. Franzis-Verlag, Feldkirchen, Germany, 1998. ISBN 3-7723-8764-0. 27 pp. LCCN ???? 40 DM. URL ????. Includes CD-ROM.

Weber:19xx:SEUb

[Webxx] Joseph Weber. Special Edition Using Java. BHV Russia, ????, 19xx. ISBN 5-7791-0051-9. LCCN ????

                                        ???? URL http://www.bhv.
                                        ru/Books/1997/TechnologyJavaPdl/
                                        TechnologyJavapdl.htm.

Wehling:1996:PMA

[Weh96] Jason Wehling. PC Magazine Advanced Java Programming. Macmillan Computer Publishing, Indianapolis, IN, USA, November 1996. ISBN 0-614-20298-1. ???? pp. LCCN ???? US$34.99.

Wehling:1998:ANC

[Weh98a] Jason Wehling. Aproveche las Noches con Java. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1998. ISBN 970-17-0034-1 (??invalid ISBN??). LCCN ???? ????

Wehrli:1998:JDR

[Weh98b] Rob Wehrli. JDBC developer’s resource. Linux Journal, 45: ??, January 1998. CODEN LI-JOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Weiss:1997:CIG

[Wei97a] Aaron Weiss. The Complete Idiot’s Guide to JavaScript. Que Corporation, Indianapolis, IN, USA, second edition, February 1997. ISBN 0-7897-1136-2. xx + 316 pp. LCCN QA76.73.J39W45 1997. US$19.99.

Weiss:1997:JF

[Wei97b] Aaron Weiss. JavaScript ¡Fácil!

Prentice Hall HispanoAmericana, Mexico, DF, Mexico / Nueva York, NY, USA, 1997. ISBN 968-880-975-6. 320 pp. LCCN ???? URL http:/

/prentice.com.mx/;   http:
//prentice.com.mx/libros/
javascfac.html.  Spanish

translation of [Wei97a].

Weiss:1997:ETD

[Wei97c] Mark Allen Weiss. Experiences teaching data structures with Java. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(1): 164–168, March 1997. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Weiss:1998:DSP

[Wei98] Mark Allen Weiss. Data Structures and Problem Solving Using Java. Addison-Wesley, Reading, MA, USA, 1998. ISBN 0-201-54991-3. xxxii + 780 pp. LCCN QA76.73.J38 W45 1998. US$50. URL http:

//cseng.aw.com/bookdetail.
qry?ISBN=0-201-54991-3.

Welch:1997:LEI

[Wel97] Peter Welch. Letters to the editor: Ignore the monitor methods! Computer, 30(7):7, 9, July 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). Comments on [Lew97b], and raises concerns about the inadequacy of the Java threads model, and points

                                        to http://www.cs.bris.ac.
                                        uk/~alan/javappp.html as a

source for Java classes implementing C. A. R. Hoare’s Communicating Sequential Processes. See also [Ols97].

Welsh:1999:SSH

[Wel99a] Matt Welsh. A system supporting high-performance communication and I/O in Java. M.S. in Computer Science, University of California, Berkeley, Dept. of Computer Science, Berkeley, CA, USA, 1999. 46 pp.

Welstead:1999:JOL

[Wel99b] Steve Welstead. A Java object list dialog. C/C++ Users Journal, 17(1):??, January 1999. CODEN CCUJEX. ISSN 1075-2838.

Welzel:1999:JS

[Wel99c] Doug Welzel. Java servlets. Linux Journal, 66:??, October 1999. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http://

                                        noframes.linuxjournal.com/
                                        lj-issues/issue66/3119.html.

Westmacott:1997:WJE

[Wes97a] Ian Westmacott. WebAdmin:

JavaScript: An example. Web-Server Magazine: For Managers of World Wide Web Sites, 2(2):14–16, 18, March 1997. ISSN 1087-4232.

Westphal:1997:LNS

[Wes97b] Mario Westphal. Lernen Sie in nur 16 Stunden Visual J++. te-wi, München, Germany, 1997. ISBN 3-89362-496-1. 288 pp. LCCN ???? 39

DM. URL http://www.tewi.
de/scripts/search/details_
ISBN.idc?ISBN=3-89362-496-
1.

Westley:1998:WJA

[Wes98] Terry J. Westley. Writing Java applets in Ada: a beginner’s guide. Ada User Journal, 19(1): 57–82, April 1998. CODEN AU-JOET. ISSN 0268-652X.

Winsor:1997:JJ

[WF97] Janice Winsor and Brian Freeman. Jumping Javascript. Sun-Soft Press Java series. Prentice-Hall and SunSoft Press, Englewood Cliffs, NJ 07632, USA and 2550 Garcia Avenue, Mountain View, CA 94043-1100, USA, 1997. ISBN 0-13-841941-8. xl + 1173 pp. LCCN QA76.73.J39 W56 1997. US$34.95. Includes CD-ROM.

Winsor:1999:MJJ

[WF99] Janice Winsor and Brian Freeman. More Jumping JavaScript. SunSoft Press Java series. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1999. ISBN 0-13-922832-2. xxviii + 512 pp. LCCN QA76.73.J39 W56

1999. URL http://www.phptr.
com/bookseri/java.html;
http://www.phptr.com/ptrbooks/
ptr_0139228322.html;   http:
//www.sun.com/books/catalog/
winsor8/index.html.

White:1999:JAT

[WFC^{+}99] Seth White, Maydene Fisher, Rick Cattell, Graham Hamilton, and Mark Hapner. JDBC API tutorial and reference: universal data access for the Java 2 platform. Addison-Wesley, Reading, MA, USA, 1999. ISBN 0-201-43328-1. xxv + 1059 pp. LCCN QA76.625.J43 1999. US$44.95.

Wall:1997:IJA

[WG97a] David A. Wall and Arthur Griffith. 101 Instant Java Applets. IDG Books, San Mateo, CA, USA, May 1997. ISBN 0-7645-3082-8 (softcover). ???? pp. LCCN ???? US$26.99.

Wall:1997:DVC

[WG97b] David A. Wall and Arthur Griffith. Discover Visual Café. IDG Books, San Mateo, CA, USA, 1997. ISBN 0-7645-3062-3. xxviii + 416 pp. LCCN QA76.73.J38W34 1997. US$25.

URL http://www.idgbooks.
com/database/book_result.
msql?isbn=0-7645-3062-3.

Wall:1999:GPJ

[WG99] David Wall and Arthur Griffith. Graphics Programming with JFC. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 1999. ISBN 0-471-28307-X. xv + 381 pp. LCCN T385.W352 1999. US$44.99. Includes CD-ROM.

Waldron:1999:AVM

[WH99] John Waldron and Owen Harrison. Analysis of virtual machine stack frame usage by Java methods. In Anonymous [Ano99h], pages 271–274. ISBN ???? LCCN ????

Wheeler:1997:ACC

[Whe97] David A. Wheeler. Ada, C, C++, and Java vs. The Steelman. ACM SIGADA Ada Letters, 17(4):88–112, July/August 1997. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

White:1996:JCA

[Whi96a] Greg White. Java commandline arguments. Dr. Dobb’s Journal of Software Tools, 21 (2):58–??, 60, 62, 96, 98– 99, February 1996. CODEN DDJOEB. ISSN 1044-789X.

White:1996:DDJ

[Whi96b] I. White. Designing and developing Java applications. In Anonymous [Ano96-142], pages 347–348. ISBN ???? LCCN ????

White:1996:DJC

[Whi96c] I. White. Designing with the Java class libraries. In Anonymous [Ano96-142], pages 277– 282. ISBN ???? LCCN ????

Whitehouse:1996:AWW

[Whi96d] K. Whitehouse. Applications: Weather without the weatherman [on the World Wide Web]. IEEE Computer Graphics and Applications, 16(2): 12–15, March 1996. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). Discusses use of Java for browsing of weather data on the Web, with samples avail-

                                        able at http://cirrus.sprl.
                                        umich.edu/javaweather/.

White:1997:CJ

[Whi97a] Conor White. From C to Java. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, April 1997. ISBN 1-55558-169-2 (softcover). ???? pp. LCCN ???? US$29.95.

White:1997:UA

[Whi97b] I. White. Understanding the AWT. Object Expert, 2(3):14– 17, March–April 1997. CODEN OBEXFJ. ISSN 1360-3426.

Whybrow:1996:EPJ

[Why96] M. Whybrow. Epidemic proportions [Java programming language]. Banking Technology, 13 (2):20–??, March 1996. CODEN BATEEM. ISSN 0266-0865.

Wiedling:1996:JG

[Wie96] Hans-Peter Wiedling. Java ohne Geheimnis. Verlag Heinz Heise, Helsdorfer Straße 7, D-30625, Hannover, Germany, 1996. ISBN 3-88229-077-3. 163 pp. LCCN ???? 39,80 DM. URL

http://www.emedia.de/bin/
bookshop?show=4862&id=;
http://www.heise.de/buch/.

Wiedling:1997:JG

[Wie97] H.-P. Wiedling. Java ohne Geheimnis. Verlag Heinz Heise, Helsdorfer Straße 7, D-30625, Hannover, Germany, 1997. ISBN ???? 164 pp. LCCN ???? 39,80 DM. URL

http://www.emedia.de/bin/
bookshop?show=4862&id=.

Wijela:19xx:JBP

[Wijxx] Michael R. Wijela. Java Bahasa Pemograman Internet. Dinas Tino, ????, 19xx. ISBN 979-552-300-7. ???? pp. LCCN ???? ????

Wilde:1996:CCS

[Wil96a] Candee Wilde. Computer careers: Steaming brew: Demand for Java skills is hotter than ever. it’s easy to learn, and it’s heating up the competition. Computerworld, 30(24):87, June 10, 1996. CODEN CMPWAB. ISSN 0010-4841.

Wilder:1996:JGC

[Wil96b] C. Wilder. Java in gear [CSX freight transporter]. Information Week, ????(592):14–16, ???? 12, 1996. CODEN IN-FWE4. ISSN 8750-6874.

Wilder:1996:PCI

[Wil96c] C. Wilder. Pouring cash into the Internet [electronic commerce]. Information Week, ????(560): 14–16, January 1, 1996. CO-DEN INFWE4. ISSN 8750-6874.

Wilder:1996:RW

[Wil96d] C. Wilder. Reaching for the Web. Information Week, ???? (594):47–48, 52, 58, ???? 26, 1996. CODEN INFWE4. ISSN 8750-6874.

Williams:1996:JPD

[Wil96e] T. Williams. Java perks developer interest from IS to embedded systems. Computer Design (International Edition), 35(6): 32–4, 37, May 1996. CODEN CDESEL. ISSN 0010-4566.

Williams:1996:JGW

[Wil96f] Tom Williams. Java goes to work controlling networked embedded systems. Computer Design, 35(9):36–??, August 1, 1996. CODEN CMPDAM. ISSN 0010-4566.

Wills:1996:J

[Wil96g] A. C. Wills. Java++. Object Expert, 1(4):48–50, 57, May–June 1996. CODEN OBEXFJ. ISSN 1360-3426.

Willston:1996:LEJ

[Wil96h] John B. Willston. Letter to the editor: Java dive. Dr. Dobb’s Journal of Software Tools, 21 (12):8, December 1996. CO-DEN DDJOEB. ISSN 1044-789X.

Wilkinson:1997:DJA

[Wil97a] Bill Wilkinson. Developing Java and ActiveX Business Applications: Using Supercede for Cutting Edge Web Development. Osborne/McGraw-Hill, Berkeley, CA, USA, June 1997. ISBN 0-07-882342-0 (softcover). ???? pp. LCCN ???? US$31.49. Publication cancelled.

Williamson:1997:SEU

[Wil97b] Alan Williamson. Special Edition Using Java Servlet API. Que Corporation, Indianapolis, IN, USA, 1997. ISBN 0-7897-1319-5. ???? pp. LCCN ???? 99,00 DM; US$50. URL http:/

/www.mut.ch/profit.htx?&x_
pk=kd84mf7ms9dm.  Includes

CD-ROM.

Wilson:1997:IGS

[Wil97c] A. Wilson. An Internet game server in Java. Web Techniques, 2(3):32–37, March 1997. CO-DEN WETEFA. ISSN 1086-556X.

Wilson:1997:JNM

[Wil97d] Andrew Wilson. The Java Native Method Interface and Windows. Dr. Dobb’s Journal of Software Tools, 22(8):46, 48– 50, 94, August 1997. CODEN DDJOEB. ISSN 1044-789X.

Anonymous:1997:BPC

[Wil97e] Gregory V. Wilson. Bookshelf:

First-year C++ text outdated by Java: Programming with Class: A C++ Introduction. the craft of software testing. operating systems. executive guide to preventing IT disasters. managing a programming projectprocesses and people. IEEE Software, 14(3):122–124, May/ June 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Wilson:1997:BFY

[Wil97f] Gregory V. Wilson. Bookshelf: First-year C++ text outdated by Java: Programming with Class: A C++ Introduction. the craft of software testing. operating systems. executive guide to preventing IT disasters. managing a programming projectprocesses and people. IEEE Software, 14(3):122–124, May/ June 1997. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Wilson:1997:PBM

[Wil97g] Gregory V. Wilson. Programmer’s bookshelf. Dr. Dobb’s Journal of Software Tools, 22 (3):117–??, March 1997. CO-DEN DDJOEB. ISSN 1044-789X.

Wilson:1997:PBMa

[Wil97h] Gregory V. Wilson. Programmer’s bookshelf. Dr. Dobb’s Journal of Software Tools, 22 (3):117–??, March 1997. CO-DEN DDJOEB. ISSN 1044-789X.

Jenkins:1997:AAH

[Wil97i]  Gregory V. Wilson.  Pro-

grammer’s bookshelf: From ActiveX to cargo-cult science. Dr. Dobb’s Journal of Software Tools, 22(9):117–119, September 1997. CODEN DDJOEB. ISSN 1044-789X. Contains reviews of Java books [App97, CM97c].

Wilson:1997:PBA

[Wil97j]  Gregory V. Wilson.  Pro-

grammer’s bookshelf: From ActiveX to cargo-cult science. Dr. Dobb’s Journal of Software Tools, 22(9):117–119, September 1997. CODEN DDJOEB. ISSN 1044-789X. Contains brief book reviews of [App97, CM97c].

Wilhelms:1998:JB

[Wil98a] Wilhelms. Das Java Buch. International Thomson Computer Press, 20 Park Plaza Suite 1001, Boston, MA 02116 USA, January 1998. ISBN 3-8266-0395-8 (??invalid ISBN??). LCCN ???? ????

Wilson:1998:EMJ

[Wil98b] Andrew Wilson. Examining Microsoft’s J/Direct: Making Java native code easy on Windows 95/NT. Dr. Dobb’s Journal of Software Tools, 23(1):90, 92–96, 104–105, January 1998. CO-DEN DDJOEB. ISSN 1044-789X.

Wilson:1998:NPV

[Wil98c] Andy Wilson. Network Programming with Visual J++. Microsoft Press, Bellevue, WA, USA, June 1998. ISBN 1-57231-855-4. xiv + 277 pp. LCCN A76.625 .W57 1999. US$40. Includes CD-ROM.

Williamson:1999:JSE

[Wil99a] Alan R. Williamson. Java Servlets By Example. Manning Publications, Greenwich, CT, USA, September 1999. ISBN 1-884777-66-X. LCCN QA76.73.J38 W552 1999. US$50.

URL http://www.manning.
com/Williamson/index.html.

Wilson:1999:JQH

[Wil99b] Andrew Wilson. Java Q and A: How do I implement Microsoft’s delegate in Pure Java? Dr. Dobb’s Journal of Software Tools, 24(3):107–110, March 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/1999/9902/9902toc.
htm;   http://www.ddj.com/
ftp/1999/1999_03/jqa399.txt;
 http://www.ddj.com/ftp/1999/
1999_03/jqa399.zip.

Wilson:1999:JQW

[Wil99c] Andy Wilson. Java Q&A: What is the Java VM Profiler interface? Dr. Dobb’s Journal of Software Tools, 24(9): 103–106, September 1999. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/1999/1999_09/
jqa999.txt;   http://www.

ddj.com/ftp/1999/1999_09/
jqa999.zip.

Wilson:1999:PBJ

[Wil99d] Gregory V. Wilson. Programmer’s bookshelf: a joke too far. Dr. Dobb’s Journal of Software Tools, 24(11):127–128, November 1999. CODEN DDJOEB. ISSN 1044-789X.

Wilson:1999:PBS

[Wil99e] Gregory V. Wilson. Programmer’s bookshelf: Small is beautiful — kind of. Dr. Dobb’s Journal of Software Tools, 24 (2):135, 137, February 1999. CODEN DDJOEB. ISSN 1044-

789X. URL http://www.ddj.
com/1999/9902/9902toc.htm;
http://www.ercb.com/ddj/
1999/ddj.9902.html.

Wingfield:1995:IMA

[Win95a]  Nick Wingfield.  Interview:

Marc Andreessen tackles Internet issues. InfoWorld, page ??, October 23, 1995. CODEN IN-WODU. ISSN 0199-6649. URL

http://www.infoworld.com/.

Wingfield:1995:ISE

[Win95b]  Nick Wingfield.  Interview:

Sun’s Eric Schmidt sizes up Internet rivals. InfoWorld, page ??, October 16, 1995. CODEN INWODU. ISSN 0199-6649. URL http://www.

infoworld.com/.

Wingfield:1995:NAI

[Win95c] Nick Wingfield. News: All-out Internet war begins as Microsoft enters fray. InfoWorld, page ??, December 03, 1995. CODEN INWODU. ISSN 0199-6649.

URL http://www.infoworld.
com/.

Wingfield:1995:NGM

[Win95d] Nick Wingfield. News: General Magic looks to ride Internet wave. InfoWorld, page ??, November 30, 1995. CODEN INWODU. ISSN 0199-6649.

URL http://www.infoworld.
com/.

Wingfield:1995:NJG

[Win95e] Nick Wingfield. News: Java gets up head of steam and IBM readies Notes support. InfoWorld, page ??, December 10, 1995. CODEN INWODU. ISSN 0199-6649. URL http://www.

infoworld.com/.

Wingfield:1995:NMS

[Win95f] Nick Wingfield. News: Microsoft storms the Web. InfoWorld, page ??, December 10, 1995. CODEN INWODU. ISSN 0199-6649. URL http://www.

infoworld.com/.

Wingfield:1995:NNI

[Win95g] Nick Wingfield. News: Netscape inks pact with Sun, Macromedia. InfoWorld, page ??, May 30, 1995. CODEN INWODU. ISSN 0199-6649. URL http://

www.infoworld.com/.

Wingfield:1995:NNS

[Win95h] Nick Wingfield. News: Netscape sets Navigator free. InfoWorld, page ??, September 21, 1995. CODEN INWODU. ISSN 0199-6649. URL http://www.

infoworld.com/.

Wingfield:1995:NNT

[Win95i] Nick Wingfield. News: Netscape takes stab at easing Web chores. InfoWorld, page ??, September 09, 1995. CODEN INWODU. ISSN 0199-6649. URL http://

www.infoworld.com/.

Wingfield:1995:NOR

[Win95j] Nick Wingfield. News: Oracle reveals strategy to cut Netscape’s lead on Web. InfoWorld, page ??, November 02, 1995. CODEN INWODU. ISSN 0199-6649. URL http://www.

infoworld.com/.

Wingfield:1995:NSP

[Win95k] Nick Wingfield. News: SGI, ParaGraph usher in 3-D Web sites. InfoWorld, page ??, December 07, 1995. CODEN IN-WODU. ISSN 0199-6649. URL

http://www.infoworld.com/.

Wingfield:1995:NSJ

[Win95l] Nick Wingfield. News: Smell of Java lures scores of vendors. InfoWorld, page ??, October 28, 1995. CODEN INWODU. ISSN 0199-6649. URL http://www.

infoworld.com/.

Wingfield:1995:NSF

[Win95m] Nick Wingfield. News: Sun focuses on eclipsing Windows. InfoWorld, page ??, November 14, 1995. CODEN INWODU. ISSN 0199-6649. URL http://www.

infoworld.com/.

Wingfield:1995:NWA

[Win95n] Nick Wingfield. News: Web apps promise to help harness Internet power. InfoWorld, page ??, August 26, 1995. CODEN INWODU. ISSN 0199-6649. URL http://www.

infoworld.com/.

Winston:1996:JBH

[Win96] Peter Winston. The Java bandwagon: How to time your jump. Computerworld, 30(11): 48, March 11, 1996. CO-DEN CMPWAB. ISSN 0010-

4841.  URL http://www.
computerworld.com/.

Winzeler:1998:PJM

[Win98] Loren Winzeler. Programming With the Java Media Framework. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, March 1998. ISBN 0-471-25169-0. LCCN ???? US$40. URL http:

//www.wiley.com/compbooks/
catalog/25169-0.htm.

Wipke:1996:RVT

[Wip96]  K. B. Wipke.  Reducing

VMTs through transit-on-demand with GPS and satellite communications. In IEEE [IEE96g], page ?? ISBN 0-7803-3277-6 (softbound), 0-7803-3278-4 (casebound). LCCN TK7868.D5 I34 1996. IEEE catalog number 96CH35928.

Wiseman:1997:HJ

[Wis97] Raymond Wiseman. HTML und Java. Sybex Verlag GmbH, Erkrather Straße 345-349, 40231 Düsseldorf, Germany, 1997. ISBN 3-8155-7509-5. 300 pp. LCCN ???? 20

DM. URL http://www.sybex.

de/. Includes CD-ROM.

Withers:1997:DJE

[Wit97] John Withers. Developing Java Entertainment Applets. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, March 1997. ISBN 0-471-16506-9. vii + 392 pp. LCCN 96-049129. US$29.95.

URL http://www.wiley.com/
compbooks/catalog/16506-9.
htm.

Weaver:1997:IJW

[WJ97] Lynn Weaver and Bob Jervis.

Inside Java Workshop. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1997. ISBN 0-13-858234-3. xxvi + 267 pp. LCCN QA76.73.J38W398 1997. US$39.95. URL http://www.

sun.com/smi/ssoftpress/books/
Weaver/Weaver.html.

Weaver:1998:IJW

[WJ98] Lynn Weaver and Bob Jervis.

Inside Java WorkShop 2.0. Sun-Soft Press Java series. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, second edition, 1998. ISBN 0-13-899584-2. xxix + 380 pp. LCCN QA76.73.J38W397 1998.

                                        US$40.  URL http://www.
                                        fdds.com/books/books/Weaver2/
                                        Weaver2.html;   http://www.
                                        prenhall.com/ptrbooks/ptr_
                                        0138995842.html.

Wright:1998:CJT

[WJUH98] A. Wright, S. Jagannathan, C. Ungureanu, and A. Hertzmann. Compiling Java to a typed lambda-calculus: a preliminary report. Lecture Notes in Computer Science, 1473:9– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wingfield:1995:NCS

[WK95] Nick Wingfield and Sari Kalin. News: Consortium to standardize Web objects. InfoWorld/ IDG International News Service, page ??, December 16, 1995.

Wolz:1999:SJPa

[WK99a] Ursula Wolz and Elliot Koffman. simpleIO: a Java package for novice interactive and graphics programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(3): 139–142, September 1999. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Wolz:1999:SJPb

[WK99b] Ursula Wolz and Elliot Koffman. simpleIO: a Java package for novice interactive and graphics programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31 (3):212, September 1999. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Weber:1998:PCAb

[WKE98] Andreas Weber, Wolfgang Küchlin, and Bernhard Eggers. Parallel computer algebra software as a Web component. Concurrency: Practice and Experience, 10(11–13): 1179–1188, September 1998. CODEN CPEXEI. ISSN 1040-3108. URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=10050416;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=10050416&PLACEBO=IE.pdf.

Special Issue: Java for Highperformance Network Computing.

Weber:1998:PCAa

[WKES98] Andreas Weber, Wolfgang Küchlin, Bernhard Eggers, and Volker Simonis. Parallel computer algebra software as a Web component. In ACM [ACM98a], page ?? CODEN CPEXEI. ISSN 1040-3108. LCCN ????

URL http://www.cs.ucsb.
edu/conferences/java98/papers/
algebra.pdf;   http://www.
cs.ucsb.edu/conferences/java98/
papers/algebra.ps. Also pub-

lished as Concurrency: Practice and Experience, 10(11– 13), September 1998, CODEN CPEXEI, ISSN 1040-3108.

Wingfield:1995:NJB

[WL95] Nick Wingfield and Martin La-Monica. News: Java brews trouble for Microsoft. InfoWorld, page ??, November 18, 1995. CODEN INWODU. ISSN 0199-6649. URL http://www.

                                        infoworld.com/.

Waddington:1996:JVM

[WL96] Simon Waddington and Stephen Li. Java: Virtual machine for virtually any platform. Embedded Systems Programming, 9 (6):26–28, 30, 32, 34, 36, 38, 40, 42, June 1, 1996. CODEN EYPRE4. ISSN 1040-3272.

Widom:1997:LES

[WL97a] Allan Widom and Steen Lehmann. Letters to the editor: Swaine on Java. Dr. Dobb’s Journal of Software Tools, 22(4):10, 12, April 1997. CODEN DDJOEB. ISSN 1044-789X.

Wigglesworth:1997:OJP

[WL97b] Joe Wigglesworth and Paula Lumby. Official Java Programmer Certification Training: Moving from C++ to Java. International Thomson Computer Press, 20 Park Plaza Suite 1001, Boston, MA 02116 USA, December 1997. ISBN 1-85032-922-2. LCCN ???? US$55.

Winters:1997:AVJ

[WL97c] Patrick Winters and Laura Lemay. Aprendiendo Visual J++ en 21 Dias. Prentice Hall HispanoAmericana, Mexico, DF, Mexico / Nueva York, NY, USA, 1997. ISBN 968-880-884-9. 584 pp. LCCN

????  ????  URL http:/

/prentice.com.mx/;   http:
//prentice.com.mx/libros/

visualj.html. Spanish translation of [W^{+}96].

Woytaszek:1997:LEQ

[WL97d] Tom Woytaszek and Ted Lewis.

Letter to the editor: Questions Java answers. Computer, 30(4): 7, April 1997. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Winters:1998:AVJ

[WL98] Patrick Winters and Laura Lemay. Aprendiendo Visual J++ en 21 d’as. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1998. ISBN 968-880-849-9. LCCN ????

???? URL http://www.mcl.
com.mx/sotano/prentice.fm$
Retrive?RecID=2673&html=HTMLRecord1.

Welter:1997:SSJ

[WM97a] Pete Welter and John Meier.

Server-side Java. SysAdmin, 6 (8):41, 43, 45–47, 49–51, August 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Williamson:1997:JDP

[WM97b] Alan (Alan R.) Williamson and Ceri Moran. Java database programming: servlets and JDBC. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1997. ISBN 0-13-737917-X (paperback). ???? pp. LCCN QA76.73.J38W55 1997. US$50. URL http://

                                        www.prenhall.com/ptrbooks/
                                        ptr_013737917x.html.

Wooldridge:1997:UJ

[WMDH97] Andrew Wooldridge, Mike Morgan, Rick Darnell, and Jerry Honeycutt. Using JavaScript. Que Corporation, Indianapolis, IN, USA, second edition, May 1, 1997. ISBN 0-7897-1138-9. xxx + 856 pp. LCCN QA76.73.J38 W664 1997. US$49.99.

Winston:1996:J

[WN96] Patrick Henry Winston and Sundar Narasimhan. On to Java. The Programming Languages Library. Addison-Wesley, Reading, MA, USA, July 1, 1996. ISBN 0-201-49826-X. xi + 329 pp. LCCN QA76.73.J38W56 1996. US$17.95. URL http://www.

                                        ascent.com/books;   http:/
                                        /www.aw.com/.

Whittington:1997:LEC

[WN97] Paul Whittington and Nancy Nicolaisen. Letter to the editor: Could Ada95 outmuscle Java? Object Magazine, 6(12): 6, 8, February 1997. CODEN OBMAFO. ISSN 1055-3614.

Winston:1998:J

[WN98] Patrick Henry Winston and Sundar Narasimhan. On to Java. Addison-Wesley, Reading, MA, USA, second edition, 1998. ISBN 0-201-38598-8. xi + 379 pp. LCCN QA76.73.J38W56 1998. US$18.95.

West:1997:SAW

[WO97] R. Webster West and R. Todd Ogden. Statistical analysis with Webstat, Java applet for the World Wide Web. Journal of Statistical Software, 2(3):1– 7, 1997. CODEN JSSOBK. ISSN 1548-7660. URL http://

www.jstatsoft.org/v02/i03;
 http://www.jstatsoft.org/
v02/i03/HTML/;   http://www.
jstatsoft.org/v02/i03/updates;
 http://www.jstatsoft.org/
v02/i03/webstat.pdf;   http:
//www.jstatsoft.org/v02/i03/
webstat.ps.

Welsh:1999:UNS

[WOC99] Matt Welsh, David Oppenheimer, and David Culler. U-Net/SLE: a Java-based usercustomizable virtual network interface. Scientific Programming, 7(2):147–156, ???? 1999. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:

//iospress.metapress.com/
app/home/contribution.asp%
3Fwasp=f277qlrwwjr5m4vxjywv%
26referrer=parent%26backto=
issue%2C6%2C8%3Bjournal%2C7%
2C9%3Blinkingpublicationresults%
2C1%2C1.

Wodaski:1996:CCW

[Wod96] Ron Wodaski. Creating Cool Web Pages with Netscape Gold and Javascript. I D G Books Worldwide, Indianapolis, IN, USA, October 1996. ISBN 0-7645-3021-6. xvi + 412 pp.

LCCN TK5105.883.N48W63 1996. US$29.99.

Woehr:1997:CWK

[Woe97] Jack Woehr. A conversation with William Kahan: How important is numerical accuracy? Dr. Dobb’s Journal of Software Tools, 22(11):18–20, 22, 24, 26, 30, 32, November 1997. CO-DEN DDJOEB. ISSN 1044-789X. Kahan, the father of the IEEE 754 floating-point standard, talks about floating-point arithmetic issues, and numerical deficiencies in Java.

Wolbrueck:1997:JCN

[Wol97] David Wolbrueck. Java compiler now available for HP 3000. HP Chronicle, 14(6):10, 15, May 1997. ISSN 0892-2829.

                                        URL http://jazz.external.
                                        hp.com/src/java.

Wolff:1998:EJI

[Wol98] Christian T. Wolff. Einführung in Java. Informatik & Praxis. Teubner, Stuttgart, Germany; Leipzig, Germany, 1998. ISBN 3-519-02993-6. 300 pp. LCCN ???? 45 DM.

Wolczko:1999:UTJ

[Wol99] M. Wolczko. Using a Tracing Java Virtual Machine to gather data on the behavior of Java programs. Technical report, Sun Microsystems, Inc., Menlo Park, CA, USA, 1999. ???? pp. URL http:

                                        //research.sun.com/people/
                                        mario/tracing-jvm/.

Wolmeringer:19xx:JTa

[Wolxxa] Gottfried Wolmeringer. Java Taschentabelle. Hofacker Verlag, ????, 19xx. ISBN 3-88963-376-5. LCCN ???? 10 DM.

URL http://www.hofacker.
de/TAB376.HTM.

Wolmeringer:19xx:JTb

[Wolxxb] Gottfried Wolmeringer. Java Taschentabelle. Hofacker Verlag, ????, 19xx. ISBN 3-88963-847-3. LCCN ???? 30 DM.

Winters:1996:PVJ

[WOLP96] Patrick Winters, David Olhasso, Laura Lemay, and Charles Perkins. Programmare in Visual J++. Apogeo srl, Viale Papiniano 38, 20123 Milano, Italy, April 1996. ISBN 88-7303-316-4. 544 pp. LCCN ???? 58000 Lire. URL http://www.

apogeonline.com/catalogo/
316.html;   http://www.urra.

it/316.html. Includes CD-ROM.

Wong:1997:MHJ

[Won97] Wylie Wong. Microsoft hones Java strategy: Virtual machine, class libraries are on top. Computerworld, 31(14):113, April 7, 1997. CODEN CMPWAB. ISSN 0010-4841.

Wong:1998:DMJ

[Won98] Simon K. Wong. Developing multilingual Java applications for Oracle8. In Unicode Consortium [Uni98b], page ?? ISBN ???? LCCN ???? URL http:

//www.unicode.org/unicode/
iuc13/program.html.

Wong:1999:JNE

[Won99] Hinkmond Wong. Jini and network-enabled devices. Dr. Dobb’s Journal of Software Tools, 24(7):21–22, 24–26, July 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://

                                        www.ddj.com/ftp/1999/1999_
                                        07/jini.txt;   http://www.
                                        ddj.com/ftp/1999/1999_07/
                                        jini.zip.

Wood:1996:VJ

[Woo96]  Charles A. Wood.  Vi-

sual J++. Prima Publishing, Rocklin, CA, USA, 1996. ISBN 0-7615-0814-7. 603 pp. LCCN QA76.73.J38W66 1996. US$35. URL http:

                                        //www.primapublishing.com/
                                        cpd/76150814.html.

Wood:1997:PVJ

[Woo97a] Charles A. Wood. Programmieren in Visual J++. Franzis-Verlag, Feldkirchen, Germany, 1997. ISBN 3-7723-7443-3. 511 pp. LCCN ???? 78 DM.

                                        URL http://www.franzis-
                                        buch.de/katalog/vjprogrammieren/

.

Wood:1997:VJM

[Woo97b] Charles A. Wood. Visual J++ Master’s Handbook. Prima Publishing, Rocklin, CA, USA, 1997. ISBN 0-7615-0819-8. ???? pp. LCCN ???? ????

Wood:1999:ACF

[Woo99a] Dave Wood. Ada: a commercial flop and proud of it! — or — how to deal with Java envy. ACM SIGADA Ada Letters, 19 (4):32–36, December 1999. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Woods:1999:DGJ

[Woo99b] Dan Woods. The Developer’s Guide to the Java Web Server: building effective and scalable server-side applications. Addison-Wesley, Reading, MA, USA, June 1999. ISBN 0-201-37949-X. xxi + 577 pp. LCCN QA76.73.J38W662 1999. US$39.95.

Wood:19xx:PVJa

[Wooxx] Charles A. Wood. Programmer en Visual J++. Sybex France, Paris, France, 19xx. ISBN 2-7361-2291-7. LCCN ????

229 FF. URL http://www-
db.sybex.fr/cgi-bin/Sybex/
Webdriver?MIval=Fiche_Book.
html&Reference=2291;   http:
//www.sybex.fr/.  Includes

CD-ROM.

WWW:1997:AHS

[Wor97] World Wide Web Consortium.

Advancing HTML: style and substance, volume 2(1) of World Wide Web journal. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, Winter 1997. ISBN 1-56592-264-6. ISSN 1085-2298. vii + 246 pp. LCCN ????

Wingfield:1995:NDM

[WP95] Nick Wingfield and Michael Parsons. News: Database, mainframe vendors pin hopes on network device. InfoWorld/ IDG International News Service, page ??, November 23, 1995.

Winters:1997:TYV

[WPL97] Patrick Winters, Charles Perkins, and Laura Lemay. Teach Yourself Visual J++ In 21 Days. Sams.net, Indianapolis, IN 46268, USA, second edition, 1997. ISBN 1-57521-244-7. ???? pp. LCCN ???? US$40.

Wong:1999:JMA

[WPM99] David Wong, Noemi Paciorek, and Dana Moore. Java-based mobile agents. Communications of the ACM, 42(3):92– 95, 97–102, March 1999. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.
acm.org:80/pubs/citations/
journals/cacm/1999-42-3/p92-
wong/.

Weaver:1998:JE

[WR98a] Lynn Weaver and Leslie Robertson. JavaStudio by Example. SunSoft Press Java series. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1998. ISBN 0-13-899519-2 (softcover). xxx + 392 pp. LCCN A76.73.J38 W399 1998. US$35.95. URL http:

//www.phptr.com/ptrbooks/
ptr_0138995192.html;   http:
//www.sun.com/books/catalog/
books_comingsoon.html;   http:
//www.sun.com/smi/ssoftpress/

catalog/books_comingsoon.
html.

Winder:1998:DJS

[WR98b] R. (Russel) Winder and Graham Roberts. Developing Java Software. Worldwide series in computer science. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 1998. ISBN 0-471-97655-5. xxi + 818 pp. LCCN QA76.73.J38 W557 1998. US$66. URL

http://www.dcs.kcl.ac.uk/
DevJavaSoft/.

Walker:1999:UJH

[WR99a] D. Walker and O. Rana. The use of Java in high performance computing: a data mining example. Lecture Notes in Computer Science, 1593:863– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Whaley:1999:CPE

[WR99b] John Whaley and Martin Rinard. Compositional pointer and escape analysis for Java programs. ACM SIGPLAN Notices, 34(10):187–206, October 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

URL http://www.acm.org/
pubs/citations/proceedings/
oops/320384/p187-whaley/.

Winder:19xx:DJS

[WRxx] Russel Winder and Graham Roberts. Developing Java Software. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 19xx. ISBN 0-471-19983-4. ???? pp. LCCN ???? US$53.

Wood:1997:GTJ

[WRB^{+}97] Kenneth R. Wood, Tristan Richardson, Frazer Bennett, Andy Harter, and Andy Hopper. Global teleporting with Java: Toward ubiquitous personalized computing. Computer, 30(2):53–59, February 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Weber:1998:SEU

[WRTM98] Joseph L. Weber, Tim Ryan, Jeff Taylor, and Ben Milstead. Special edition, using Java 1.2. Que Corporation, Indianapolis, IN, USA, fourth edition, 1998. ISBN 0-7897-1529-5. xxv + 1414 pp. LCCN QA76.73.J38 W4 1998. US$50. URL

http://iq-mcp3.iquest.net/
catalog/mcp/PRODUCT/PAGE/
07897/bud/0789715295.html.

Edited by Tim Ryan, Jeff Taylor, and Ben Milstead.

Wollrath:1996:DOMa

[WRW96a] A. Wollrath, R. Riggs, and J. Waldo. A distributed object model for the Java system. In Association [Ass96a], pages 261–??, 219–231. ISBN ???? LCCN ????

Wollrath:1996:DOMb

[WRW96b] Ann Wollrath, Roger Riggs, and Jim Waldo. A distributed object model for the Java system.

Computing Systems, 9(4):265– 290, Fall 1996. CODEN CM-SYE2. ISSN 0895-6340.

Wollrath:COMPSYS-9-4-265

[WRW96c] Ann Wollrath, Roger Riggs, and Jim Waldo. A distributed object model for the Java system. Computing Systems, 9(4):265– 290, Fall 1996. CODEN CM-SYE2. ISSN 0895-6340.

Ward:1998:JFP

[WS98] Robert Ward and Martin Smith. JavaScript as a first programming language for multimedia students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(3):249–253, September 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Welch:1999:DKE

[WS99] I. Welch and R. Stroud. From Dalang to Kava — the evolution of a reflective Java extension. Lecture Notes in Computer Science, 1616:2–??, 1999. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wester:1997:VPJ

[WSK97] Frank Wester, Marleen Sint, and Peter Kluit. Visual programming with Java; an alternative approach to introductory programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(3):57–58, September 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Wilder:1996:MNG

[WSL96] C. Wilder, J. Swenson, and R. Levin. Microsoft’s Net gains [Internet strategy]. Information Week, ????(571):14–16, March 18, 1996. CODEN INFWE4. ISSN 8750-6874.

Wu:1997:CSI

[Wu97]  Amy Wu.  Chip shots II.

Dr. Dobb’s Journal of Software Tools, 22(1):12, January 1997. CODEN DDJOEB. ISSN 1044-789X. Short note about Sun’s JavaChip silicon project, and the picoJava I chip.

Wu:1999:IOO

[Wu99] C. Thomas Wu. An Introduction to Object-Oriented Programming With Java. McGraw-Hill, New York, NY, USA, January 1999. ISBN 0-07-115612-7. LCCN ???? US$47.

Wu:19xx:IPO

[Wuxx] C. Thomas Wu. An Introduction to Programming: An Object-Oriented Approach with Java. Richard D. Irwin, Inc., ????, 19xx. ISBN 0-256-25462-1. ???? pp. LCCN ???? US$47.

                                        URL http://www.irwin.com/
                                        catalogs/999103.html.

Walter:1996:BMJ

[WW96a] Scott J. Walter and Aaron Weiss. Beginnen met JavaScript. Academic Service, Postbus 81, 2870 AB Schoonhoven, The Netherlands, 1996. ISBN 90-395-0507-1. 304 pp. LCCN ???? fl. 49,00. URL http:/

/www.acaserv.nl;   http://
www.acaserv.nl/5_0507.htm.

Dutch translation of [WW96b]. Includes CD-ROM.

Walter:1996:CIG

[WW96b] Scott J. Walter and Aaron Weiss. The Complete Idiot’s Guide to JavaScript. Que Corporation, Indianapolis, IN, USA, March 1, 1996. ISBN 0-7897-0798-5. xviii + 284 pp. LCCN QA76.73.J39W34 1996. US$19.99. URL http://www.

mcp.com/175067741767563/cgi-
bin/bag?isbn=0-7897-0798-
5&last=/bookstore.

Waldo:1997:DCJ

[WW97] Jim Waldo and Ann Wollrath. Distributed computing with Java remote method invocation. Tutorial session at [USE97b]., 1997.

URL http://dxsting.cern.
ch/sting/COOTS97.html.

Weinman:1998:CHD

[WW98] Lynda Weinman and William Weinman. Creative HTML design. New Riders Publishing, Carmel, IN, USA, 1998. ISBN 1-56205-704-9. xxx + 434 pp. LCCN QA76.76.H94W45 1998.

Wollrath:1997:JCD

[WWR97a] Ann Wollrath, Jim Waldo, and Roger Riggs. Java-centric distributed computing: Providing a homogeneous view of a heterogeneous group of machines.

IEEE Micro, 17(3):44–53, May/ June 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //pascal.computer.org/mi/
                                        books/mi1997/pdf/m3044.pdf.

Wollrath:1997:JDC

[WWR97b] Ann Wollrath, Jim Waldo, and Roger Riggs. Java-centric distributed computing: Providing a homogeneous view of a heterogeneous group of machines. IEEE Micro, 17(3):44–53, May/ June 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //pascal.computer.org/mi/
                                        books/mi1997/pdf/m3044.pdf.

Wang:19xx:JPL

[WYxx] Mingsheng Wang and Shih-Yang Yang. The Java Programming Language. Flag Publishing, ??, Taiwan, 19xx. ISBN ???? ca. 400

pp.  LCCN ????  400

                                        NT$. URL http://203.66.
                                        203.145/JavaBookEnglish.html;
                                         http://www.seed.net.tw/~

flaginfo/. Includes CD-ROM.

Wygant:19xx:HPM

[Wygxx] Daniel Wygant. How to Program Microsoft Visual J++. Ziff-Davis Press, Emeryville, CA, USA, 19xx. ISBN 1-56276-525-6. ???? pp. LCCN ???? US$40.

Xu:1997:CSU

[Xu97] Linghua Xu. A comparative study of the use of HTML, Java, perl and C in the implementation of World Wide Web applications. Thesis (m.s.), Major in Engineering/Computer Science, South Dakota State University, Brookings, SD, USA, 1997. ix + 86 pp.

Yalamov:1998:BRC

[Yal98] Plamen Y. Yalamov. Book review: Concurrent Programming: The Java Programming Language. IEEE Concurrency, 6(3):89, July/September 1998. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL

http://dlib.computer.org/
pd/books/pd1998/pdf/p3085.
pdf.

Yang:1996:IJT

[Yan96] Jin Yang. Integrating Java technology with an Oracle relational database. Thesis (m.s.), Marquette University, Milwaukee, WI, USA, 1996. iv + 59 pp.

Yang:19xx:IPD

[Yanxx] We Shing Yang. Internet Programming and Design (Java and HTML). Chuen Hwa Books & Technology Co., Ltd., ????, 19xx. ISBN ???? ???? pp. LCCN ???? 250 NT$.

Yasumura:1996:WIT

[Yas96] Y. Yasumura. WWW-OODB integration technology. NEC Technical Journal = NEC giho, 49(7):51–56, July 1996. CO-DEN NECGEZ. ISSN 0285-

Yates:1997:JWD

[Yat97] Andy Yates. Java: The wonder drug? UNIX Developer, 1(1): 16–17, January/February 1997. ISSN 1090-2279.

Yang:1999:UJS

[YB99] Andrew Yang and Yael Bachar. Using Java and the socket interface in teaching client/server programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(3):206, September 1999. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Yu:1997:JDP

[YC97] Weimin Yu and Alan Cox. Java/DSM: a platform for heterogeneous computing. Concurrency: Practice and Experience, 9(11):1213–1224, November 1997. CODEN CPEXEI. ISSN 1040-3108. URL http:

                                        //www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=13817;
                                         http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=13817&PLACEBO=IE.pdf. Spe-

cial Issue: Java for computational science and engineering — simulation and modeling II.

Yalamanchilli:1998:CPJa

[YC98a] Narendar Yalamanchilli and William Cohen. Communication performance of Java based Parallel Virtual Machines. In ACM [ACM98a], page ?? CODEN CPEXEI. ISSN 1040-3108. LCCN ????

URL http://www.cs.ucsb.
edu/conferences/java98/papers/
passing.pdf;   http://www.
cs.ucsb.edu/conferences/java98/
papers/passing.ps. Also pub-

lished as Concurrency: Practice and Experience, 10(11– 13), September 1998, CODEN CPEXEI, ISSN 1040-3108.

Yalamanchilli:1998:CPJb

[YC98b] Narendar Yalamanchilli and William Cohen. Communication performance of Javabased parallel virtual machines. Concurrency: Practice and Experience, 10(11–13): 1189–1196, September 1998. CODEN CPEXEI. ISSN 1040-3108. URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=10050407;
http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=10050407&PLACEBO=IE.pdf.

Special Issue: Java for Highperformance Network Computing.

Yee:19xx:JPL

[Yeexxa] Huang Chung Yee. Java Programming Language. Chuen Hwa Books & Technology Co., Ltd., ????, 19xx. ISBN 957-21-1533-2. ???? pp. LCCN ???? 300 NT$.

Yee:19xx:DJP

[Yeexxb] Wang Yee. Developing JavaScript Program Handbook. Unalis Corporation, ??, Taiwan, 19xx. ISBN 957-22-2261-9. ???? pp. LCCN ???? 430 NT$. URL

http://www.unalis.com.tw/

index.html. Includes CD-ROM.

Yellin:1995:LLS

[Yel95] F. Yellin. Low level security in Java. In O’Reilly and Associates and Web Consortium (W3C) [OW95], pages 369–380. ISBN 1-56592-169-0. ISSN 1085-2301. LCCN TK5105.888 .I68 1995. US$39.95. URL http://www.

ora.com/gnn/bus/ora/item/

wj1.html. The World Wide Web Journal is a quarterly publication that provides timely, indepth coverage of the issues, techniques, and research developments in the World Wide Web. The December issue contains the Conference Proceeding papers that were chosen for the 4th International World Wide Web conference in Boston, MA.

Yelland:1999:CAJ

[Yel99] Phillip M. Yelland. A compositional account of the Java virtual machine. In ACM [ACM99a], pages 57–69. ISBN 1-58113-095-3. LCCN

????  URL http://www.
acm.org:80/pubs/citations/
proceedings/plan/292540/p57-
yelland/.

Yetongnon:1996:PII

[YH96] K. Yetongnon and S. Hariri, editors. Proceedings of the ISCA International Conference. Parallel and Distributed Computing Systems, Dijon, France, 25–27 September, 1996. International Society of Computers and Their Applications (ISCA), Raleigh, NC, USA, 1996. ISBN ???? LCCN ???? Two volumes.

Yong:1998:JIM

[YHW98] Leong Kok Yong, Liu Hai, and Oliver P. Wu. Java Input Method Engine. Computer Networks and ISDN Systems, 30(1– 7):271–279, April 1, 1998. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

URL http://www.elsevier.
com/cas/tree/store/comnet/
sub/1998/30/1-7/1915.pdf.

Yerosheva:1999:PEM

[YK99] L. Yerosheva and P. M. Kogge.

Prototyping execution models for HTMT petaflop machine in Java. Lecture Notes in Computer Science, 1602:32–46, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yang:1999:LMJ

[YLP^{+}99] Byung-Sun Yang, Junpyo Lee, Jinpyo Park, Soo-Mook Moon, Kemal Ebcioğlu, and Erik Altman. Lightweight monitor for Java VM. ACM SIGARCH Computer Architecture News, 27(1):35–38, March 1999. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Yang:1998:DIW

[YLQ98] Andrew Yang, James Linn, and David Quadrato. Developing integrated Web and database applications using JAVA applets and JDBC drivers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30(1):302–306, March 1998. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Yagel:1996:PVS

[YN96] R. Yagel and G. M. Nielson, editors. Proceedings. Visualization ’96. San Francisco, CA, USA. 27 October–1 November 1996. ACM Press, New York, NY 10036, USA, 1996. ISBN 0-89791-864-9. LCCN T385 .V59 1996. IEEE catalog number 96CB36006. IEEE Computer Society Press order number RS00118.

Yoshida:1997:MDW

[Yos97] Akitoshi Yoshida. MOWS: Distributed Web and cache server in Java. Computer Networks and ISDN Systems, 29(8–13): 965–975, September 30, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/
browse/browse.cgi?year=1997&
volume=29&issue=08-13&aid=
1700.

Yourdon:1996:DAI

[You96a] E. Yourdon. Developing applications for the Internet: advice for the Java generation. American Programmer, 9(12):36–41, December 1996. CODEN AM-PRFD. ISSN 1048-5600.

Yourdon:1996:JWS

[You96b] Edward Yourdon. Java, the Web, and software development. Computer, 29(8):25–30, August 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Young:1997:NDG

[You97a] Douglas A. Young. Netscape Developer’s Guide to Plug-Ins. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1997. ISBN 0-13-270992-9. 512 pp. LCCN QA76.625.Y68 1997. US$44.95. Includes some discussion of Java.

Young:1997:MMV

[You97b] Ted M. Young. Moving from Microsoft Visual Basic to Java. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, July 1997. ISBN 0-13-839069-X. ???? pp. LCCN ???? US$39.95.

Young:1997:PBJ

[You97c] Warren Young. Programmer’s bookshelf: Java pro. Dr. Dobb’s Journal of Software Tools, 22 (10):133–134, October 1997. CODEN DDJOEB. ISSN 1044-789X. Contains book reviews of [CH97c, WBB^{+}97].

Yourdon:1997:JJT

[You97d] E. Yourdon. Java and Java tools. American Programmer, 10(1):2–12, January 1997. CO-DEN AMPRFD. ISSN 1048-

Young:1998:AFC

[You98a] Michael J. Young. Application Foundation Class (AFC) Developer’s Workshop. Microsoft Press, Bellevue, WA, USA, March 1998. ISBN 1-57231-697-7. xiii + 573 pp. LCCN QA76.625 .Y69 1998. US$45.

Yourst:1998:IJC

[You98b] Matt T. Yourst. Inside Java class files: The key to Java’s binary portability. Dr. Dobb’s Journal of Software Tools, 23 (1):46, 49, 51–52, 102–103, January 1998. CODEN DDJOEB. ISSN 1044-789X.

Yurkoski:1998:UIE

[YRE98] C. F. Yurkoski, L. R. Rau, and B. K. Ellis. Using Inferno to execute Java on small devices. Lecture Notes in Computer Science, 1474:108–??, 1998. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yelick:1998:THPa

[YSP^{+}98a] Kathy Yelick, Luigi Semenzato, Geoff Pike, Carleton Miyamoto, Ben Liblit, Arvind Krishnamurthy, Paul Hilfinger, Susan Graham, David Gay, Phil Colella, and Alex Aiken. Titanium: a high-performance Java dialect. In ACM [ACM98a], page ?? CODEN CPEXEI. ISSN 1040-3108. LCCN ????

URL http://www.cs.ucsb.
edu/conferences/java98/papers/
titanium.pdf;   http://www.
cs.ucsb.edu/conferences/java98/

papers/titanium.ps.  Also

published as Concurrency: Practice and Experience, 10(11– 13), September 1998, CODEN CPEXEI, ISSN 1040-3108.

Yelick:1998:THPb

[YSP^{+}98b] Kathy Yelick, Luigi Semenzato, Geoff Pike, Carleton Miyamoto, Ben Liblit, Arvind Krishnamurthy, Paul Hilfinger, Susan Graham, David Gay, Phil Colella, and Alex Aiken. Titanium: a high-performance Java dialect. Concurrency: Practice and Experience, 10 (11–13):825–836, September 1998. CODEN CPEXEI. ISSN 1040-3108. URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=10050392;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=10050392&PLACEBO=IE.pdf.

Special Issue: Java for Highperformance Network Computing.

Zachary:1997:GSP

[Zac97]  Joseph L. Zachary.  The

gestalt of scientific programming: problem, model, method, implementation, assessment. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29 (1):238–242, March 1997. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Zannoli:1996:OMJ

[Zan96] S. Zannoli. An optimised microprocessor for Java. Elettronica Oggi, ????(226):91–92, 94, ???? 15, 1996. CODEN ELOGDA. ISSN 0391-6391.

Zeigler:1997:OSP

[Zei97] Bernard P. Zeigler. Objects and Systems: Principled Design with Implementations in C++ and Java. Undergraduate texts in computer science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. ISBN 0-387-94781-7. xxi + 221 pp. LCCN QA76.73.C153Z45 1996. US$39.95.

Zhang:1997:SCD

[Zha97] X. Nick Zhang. Secure code distribution. Computer, 30 (6):76–79, June 1997. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). Discusses distribution security problem for Java, Safe Tcl, and Safe Perl.

Zhou:1997:CCJ

[Zho97] Ling Zhou. A comparison of C++, Java and Python. Thesis (m.s.), Department of Computer Science, Texas A&M University, College Station, TX, USA, 1997. vi + 44 pp.

Zie:19xx:JPL

[Ziexx] Shin She Zie. Java Programming Language. GOTOP Information Inc., 5F, No.7, Lane 50, Sec.3 Nan Kang Road Taipei, Taiwan; Unit 1905,Metro Plaza Tower 2, No.223 Hing Fong Road, Kwai Chung, N.T., Hong Kong, 19xx. ISBN 957-641-842-9. ???? pp. LCCN ???? 420

NT$. URL http://www.gotop.
com.tw.

Ziff-Davis:1996:HPJ

[Zif96] Ziff-Davis Press Staff. How to Program Java. Ziff-Davis Press, Emeryville, CA, USA, 1996. ISBN 1-56276-478-0. 376 pp. LCCN ???? US$39.99.

URL http://www.mcp.com/
cgi-bin/bag?isbn=1-56276-
478-0&last=/bookstore.

Ziff-Davis:1997:LNJ

[Zif97] Ziff-Davis Development Group.

Late Night Java Beans Programming. Ziff-Davis Press, Emeryville, CA, USA, April 1997. ISBN 1-56276-531-0 (softcover). ???? pp. LCCN ???? US$44.99. Publication cancelled.

Zuckerman:1996:JLP

[ZJ96] A. E. Zuckerman and L. B.

Jacques. A Java language program for writing clinical cases using custom extensions to HTML. In Cimino [Cim96], pages 906–?? ISBN 1-56053-208-4. LCCN W 26.5 S978 1996.

Zadroiny:1998:IFQ

[ZK98a] S. Zadroiny and J. Kacprzyk.

Implementing fuzzy querying via the Internet/WWW: Java applets. ActiveX controls and cookies. Lecture Notes in Computer Science, 1495:382– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zadrozny:1998:IFQ

[ZK98b] Sławomir Zadrozny and Janusz Kacprzyk. Implementing fuzzy querying via the Internet/ WWW: Java applets, ActiveX controls and cookies. Lecture Notes in Computer Science, 1495:382–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/
service/series/0558/bibs/
1495/14950382.htm;   http:
//link.springer-ny.com/link/
service/series/0558/papers/
1495/14950382.pdf.

Zirintsis:1999:HPJ

[ZKM99] Evangelos Zirintsis, Graham N. C. Kirby, and Ronald Morrison. Hyper-programming in Java. In Atkinson et al. [AOV^{+}99], pages 734–737. ISBN 1-55860-615-7. LCCN QA76.9.D3 I559 1999. URL

http://www.vldb.org/dblp/
db/conf/vldb/ZirintsisKM99.

html. Also known as VLDB’99.

Zaki:1999:TSP

[ZLGS99] Omer Zaki, Ewing Lusk, William Gropp, and Deborah Swider. Toward scalable performance visualization with Jumpshot. The International Journal of High Performance Computing Applications, 13(3):277–288, Fall 1999. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic).

Zingirian:1998:ESS

[ZMN98] N. Zingirian, M. Maresca, and S. Nalin. Efficiency of standard software architectures for Java-based access to remote databases. Lecture Notes in Computer Science, 1401:479– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zingirian:1999:ESS

[ZMN99] N. Zingirian, M. Maresca, and S. Nalin. Efficiency of standard software architectures for Java-based access to remote databases. Future Generation Computer Systems, 15(3):417– ??, ???? 1999. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Ziv:19xx:MLJ

[ZOxx] Abraham Ziv and Moshe Olshansky. Math library for Java. World-Wide Web document, 19xx. URL http:

//www.alphaworks.ibm.com/
tec/mathlibrary4java.

Zoltan:1996:SGI

[Zol96] E. Zoltan. “slap a GUI on it”.

your M future will be easier if it’s object oriented. M Computing, 4(5):16–18, December 1996. CODEN MCPUEF. ISSN 1060-7684.

Zhu:1998:SUPa

[ZTLV98a] Jinsong Zhu, Maria Toeroe, Victor Leung, and Son Vuong. Supporting universal personal computing on Internet with Java and CORBA. In ACM [ACM98a], page ?? CODEN CPEXEI. ISSN 1040-3108. LCCN ????

URL http://www.cs.ucsb.
edu/conferences/java98/papers/
upc.pdf;   http://www.cs.
ucsb.edu/conferences/java98/

papers/upc.ps. Also published as Concurrency: Practice and Experience, 10(11–13), September 1998, CODEN CPEXEI, ISSN 1040-3108.

Zhu:1998:SUPb

[ZTLV98b] Jinsong Zhu, Maria Törö, Victor Leung, and Son Vuong. Supporting universal personal computing on the Internet with Java and CORBA. Concurrency: Practice and Experience, 10(11–13):1007–1013, September 1998. CODEN CPEXEI. ISSN 1040-3108. URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=10050418;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=10050418&PLACEBO=IE.pdf.

Special Issue: Java for Highperformance Network Computing.

Zukowski:1997:BJN

[Zuk97a]  John Zukowski.  Borland’s

JBuilder: No Experience Required. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 1997. ISBN 0-7821-2135-

7. xviii + 502 pp. LCCN QA76.73.J38Z849 1997. US$30.

URL http://www.sybex.com/

cgi-bin/bookpg.pl?2135back.
html.

Zukowski:1997:JAR

[Zuk97b] John Zukowski. Java AWT Reference. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1997. ISBN 1-56592-240-9. xxv + 1045 pp. LCCN QA76.73.J38 Z841 1997. US$39.95. URL http:/

/www.oreilly.com/catalog/
javawt.

Zukowski:1998:MJ

[Zuk98]  John Zukowski.  Mastering

Java 1.2. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 1998. ISBN 0-7821-2180-2. xxxiii + 1244 pp. LCCN QA76.73.J38 Z854 1998. US$50. Includes CD-ROM.

Zukowski:1999:JZD

[Zuk99] John Zukowski. John Zukowski’s Definitive Guide to Swing for Java 2. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. ISBN 1-893115-02-X. xvi + 861 pp. LCCN QA76.73.J38 Z8493 1999. US$49.95.